A Complete Bibliography of Publications in Applied Mathematics and Computation: 2005–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

25 March 2025 Version 1.35

Title word cross-reference

(1 + 1) [ZW09a, CW05b, Waz08l]. (1 + 3) [CIS09]. (2 + 1) [DW09a, FL09b, HC09, LZ09a, MWQ09, TB09a, Waz09c, ZWT09c, Zha09f, CW05a, lH08a, MD08a, SC08a, Waz08h, Waz08l, Waz08t, XZ05, ZL07b, ZLZ08]. (2, 2) [CL05d]. (3 + 1) [MZ09, Waz09a, Waz08i, lYpZqL08, ZLYL08]. (Ax = λBx) [NS08]. (Ax = λx) [NS08]. (AXB, CXD) = (E, F) [SC07a]. (\frac{G′}{G}) [AÖ09a, Asl09a, Asl09b, AÖ09b, GZ09a, LXML09, ZG09, ZWT09c]. (G′/G) [GZ09b]. (γ, r) [HN07b]. (G′G) [WZL08b]. (I + S(α)) [ZHLG07]. (k, n) + 1 [LCZ05d]. (Kη) [DW09c]. (m) [AV09, AZ08]. (m + 1) [WS09b]. (M, N) [PHZ08]. (N + 1) [GKK09, Wan06e, GHL07]. (P Q)

[ZZH09, ILZ09]. (S_{+}) [WCO08]. (t, n) [CHY05a]. −1, 0, 0, , 0 [Rim07e]. −1, 0, 0, , 0, 1 [Rim07d, Rim07f]. 0 [KR07d, MD05]. 1 [Ahm05a, AKM06, DZ09, Dön06b, KS05a, KR07d, LLFM08, MD05, Moh07a, Moh09, Ram07d, Ram07e]. 1, 0, 0, , 0, 1 [Rim07g, Rim07h]. 1, 1, 1, , 1 [Rim07d, Rim07e, Rim07g, Rim07h, Rim07f]. 1.839 [SG05b]. 1/7 [Cha05h]. 2 [AM08d, Bra05, CJL05a, CGS08, HBE09, LX06a, LX06b, Moh07b, QZ09, SDL09, TD07a, WCZ09, YW09, ZW05g, ZH05]. 2.732 [Ren08b]. 2μ [RLZ08]. 2n [XZ09b]. 2r [hCxC05]. 2s [Fra09a]. 3 [AB05a, AS07b, ALSL08, DMA07b, Hlo05, Hsu06, KK07c, KSM09, KM07, Lee05c, LM08a, Lin05c, Ma06b, Ohm08, SP09, WZZ06a, XL06b]. 3 × 3 [lZXL09]. 3x + 1 [XX07]. 5/9 [Kah05c]. = [dCD03]. [0, 1] [EFI06]. [0, ∞) [DMJE06b].

∗ [YL08b]. ^{+} [WS07e]. ^{1} [Zhu06a]. ^{2} [KSP08, NKE05, QZ09]. ^{3} [NKE05]. ^{∞} [AM08b, AM08c, AM08d]. ^{p} [LWW09]. _{1} [HT07, hPyHZ06, PHZ06b]. _{2} [HT07, Liu07a, Ma08b, OSS09, hPyHZ06, PHZ06b]. _{m} [Aki05b]. A [AV09, Bil04, Bil07, yLcCsL09, PST06, SH06b]. A < 0 [SA06b]. A_{T,S}^{(2)}

[LYH09, WSL08]. A_{d,W} [QW05]. A_{T,S}^{(2,4)} [YLX08]. AA_{MN}^{†} [XSG08]. α [PR09a, Ste09d, HED06b, KRRGT07, LFW08, RC06, Ste09c, YL07]. αz + βx′(z) = x(az + bx′(z)) [Liu09c]. AX = B [ZZH09, AT07a, QZL07a, Ben07a]. AX = B, XC = D [QW08b]. AXB + CX^{T} D = E [WCW07a]. AXB + CY D = D [fLyHZ09]. AXB = C [PHZ05]. B [AOS09, SY09, AS07b]. BB_{MN}^{†} [XSG08]. β [HED06b]. BXA^{T} = F [QZL07b]. C[a, b] [Kat08]. C^{0} [Ram05a]. C^{1} [Ram05a]. C^{1} · N_{p} [Lal09]. C^{2} [DWT05a]. C^{∞} [Boy05, BO06]. H_{2} [Mah09b]. H_{∞} [PJWL08]. SL^{∗} [Sok09]. cos [MAY09]. coth [MAY09, GS08a, Waz07h, Waz07i]. d [LO08]. δ [JLA07]. DGDM(1, 1, 1) [Tie05]. d ≥ 2 [Kar09b]. E [KK05a]. ℓ [MMS^{+}08]. ℓ_{2}^{∞}(p) [GÇ05]. ϵ [Mav09, KP06a, KPS06a, MA05f, PS06b]. F [MBS07, SC08a, ZX06b, ZX07a]. F_{C,2} [DSY08]. F_{S,2} [DSY08]. * [OML09]. g [GWJ07, JH09, MBS07]. G_{x_{0}}^{ϵ} [mHsLzC06]. γ [Den08c]. γ < λ^{2}/4 [KAF08]. γ = λ^{2}/4 [VHFF08]. γ > λ^{2}/4 [VHFF08]. GMC(1, n) [WC05d]. H [AH07b, CKKS08, CKR09, CN09c, HL07a, HWF07, LS09c, LH06b, LH07c, yS06c, Taw07, XHL07, Yua05, ZHGG08, Bie08a, Bie08b]. H(·, ·) [ZH09b, ZH08d]. H^{1} [LL09, LL12]. H^{2} [EGFAE06]. H_{μ}^{∞} [Ste09c]. H_{∞} [HL08b, LYL^{+}08, MZW08, Muk08b, Muk08a, SMA08, SFFY08, SWX07b]. H_{J} [Pra07]. J [Sad09]. J_{ν} [LLG06]. K [GÜ07b, SLZ05a, SZL^{+}07, DOU08a, DOU08b, Dem08, FP09a, FA06, GKR06, Kah05a, Kan09a, KAT06, KB05c, Küç05a, NAUH09, ÖTA05, RSSS07, SA08c, Van09, WYL08b, YMCL06, lY08b, YK03, ZW05d]. K(d, 2) [AAG05]. K(m, n) [Waz05l]. K(n + 1, n + 1) [Waz07h]. K(n, −n, 2n) [TH08]. K(n, n) [Waz05o, Waz06f, Waz07h]. κ [PC07a]. KP (n, n) [Waz05o]. L [Inc07e, Inc07b, Sim07, WH09, You09, YK08, ACOS06]. L^{2} [TY07]. L^{2}(R^{s}) [LY09a]. L^{2}[a, b] [SP08]. L^{α}(Ω) [Pog07b]. L^{∞} [JZ06b]. L^{p} [CL09b]. L_{1} [KJ08, SMA08]. L_{2} [BMY08]. L_{∞} [XW09, Fab08]. L_{p} [Yoo05]. L_{p}(p ≥ 1) [EkES05]. LM [UP07a]. M [JZ07, qJ06, Li07a, tLlWpL08, LXSH06, PH06, ST05c, WZSL08, Xin06, BC09c, CZ07c, DZ08f, Ere09b, Kha07b, PDW06, SY07c, SWZL07, Su08c, SL06c, WP05, Wei06a, Wei06b, Xu08a, XZ09b, XYL10, YW08, ZW07b, ZYW08]. M/G/1 [WWP05]. M^{X}/G/1 [CD09, MAR05a]. C [Gra09]. C^{N} [SW09c, Ste09g, LZZC09]. R^{n} [KM09a]. R^{\} [Kat09]. min_{x∈R^{n}} \frac{||Ax−b||^{2}}{1+||x||^{2}} [Sal09]. μ [HM06, Li05a]. μ_{q} [Kim07a]. N [CLH06, Che07b, Nis07, Smi06a, Su09, WWP05, Waz08q, XMGW09, AR06, Ant05, Cas07b, Che05a, Che07c, DKL09, EOM03, EOM08, Gor06a, Kis08, KK08c, KB05c, Küç05a, LL05f, LGw07, Luo07, LG08, PDW06, PSP05, QD08, SWZL07, Su08c, TNL09, WC07b, WS09b, WZ06g, XCH05b, YW08, YL07, ZW05d, ZXL09, Zho08a]. n + 1 [GB07]. N_{0} [JAT09]. ν [NCÁHC^{+}06]. O(k^{2}h_{ℓ}^{−1} + kh_{ℓ} + h_{ℓ}^{3}) [Moh09]. O(mn) [CS06b]. O(τ^{2} + h^{4}) [DZ09]. Ω [LY09a]. P [Reb06, WCL08, ANM07a, AMN07a, ARS07a, Alt07b, Aou07, DLT08, DH09, DHG09, Fat06, GZ07c, HC08a, HS07b, Jan09c, JGY07, JG07b, Li08d, Li08e, LZ08d, Liu08b, LOA05, MA09, NMM09, Pin06, SM06c, SWX06a, Su08c, SLS08, SL06c, SG07c, SQG08, SMG08, WC08b, Xu07, ZLCZ06]. p(x) [Dai09, Den09c]. P_{∗}(κ) [AP09, CK06a]. P_{0} [TLM09, ZWG09].

P_{2} − P_{1} [BH07]. P_{4} [DSY08]. Φ [Kar09a]. ψ [Ma08a]. P X = sXP [QZL07a]. Q [SB08b, WC06b, BK08c, CAS08, CAS09, FB08, Gup08b, GF09, JK08, JV07, KG08, Kim07a, Ost08, Rad09, RMS05, Sim07]. q(t) < 0 [SWC05]. Q_{1} − P_{0} [DC08]. R [HR07a, HR09a, HY07b, HY07c, JOB08, IOS07, PL05, PG08b]. R^{2} [JH09, DC06a, DC06c]. R^{3} [Sid05, Sid06]. R_{1}^{3} [AB05a]. R^{N} [Fig08]. S [WC06b, KBÖP07]. S((d + 1)/2, d) [AAG05]. S = 2 [Oma06]. S_{m} [Sid05, Sid06]. σ = 1 [Oma06]. sin [MAY09]. ∑ f(n)n^{k} [Van09]. T [BMJED06, Ma08a]. tanh [MAY09, GS08a, Waz05o, Waz07h, Waz07i]. T · C_{1} [MS08c]. θ [WQL09, HBBA07, LGY07, Ran08b, TG08a, WL07c, WWL08b, WWY08, JRW08]. × [HM09a, BZZ08b]. U [SPX07]. u′(t) = au(t) + a_{0}u([t]) + a_{1}u([t − 1]) [YLS05]. UV [SPX07]. V [Reb06]. W [LWW05, LX07b, MD08b, QW05, WG05, DwWmW05, IR08a]. X + A^{2} + ^{2^{n}} \sqrt{X^{−1}A} = I [RES06]. X − A^{∗}X^{−n}A = I [HI05]. x^{−p}e^{−1/x} [DMJE06b]. x′(t) + ax(t) + a_{1}x([t − 1]) = 0 [LGY07]. X^{s} + A^{∗}X^{−t}A = Q [Yue07]. X^{s} + A^{T} X^{t}A = Q [JW09b]. x_{n+1} = (α + βx_{n−k+1})/(A + Bx_{n−k+1} + Cx_{n−2k+1}) [DMS06c]. x_{n+1} = (α + βx_{n−k+1} + γx_{n−2k+1})/(Bx_{n−k+1} + Cx_{n−2k+1}) [DD05]. x_{n+1} = (x_{n} + αx_{n−1})/(β + x_{n}) [CW06a]. x_{n+1} = px_{n−s} + x_{n−t}qx_{n−s} + x_{n−t} [YY08a]. x_{n+1} = x_{n−2k+1}/(x_{n−2k+1} + αx_{n−2t}) [DMS07]. x_{n} = (α + βx_{n−k+1})/(A + Bx_{n−2k+1}) [DD06b]. x_{n} = (ax_{n−1} + bx_{n−2})/(c + dx_{n−1}x_{n−2}) [YSC^{+}05]. x_{n} = (p + x_{n−1})/(qx_{n−1} + x_{n−1}) [YYL07]. x_{n} = (x_{n−1}x_{n−2} +x_{n−3} +a)/(x_{n−1} +x_{n−2}x_{n−3} +a) [Yan05d]. x_{n} = A + 1/y_{n−p} [ZYME06]. x_{n} = A/x_{n−k} + B/x_{n−3k} [DDR06c]. x_{n} = a/y_{n−3} [Özb07a]. x_{n} = a/y_{n−p} [YLB05]. XC = D [QZL07a]. XF − AX = C [WDY06, WDY06]. XH = HX [RES07]. X ± A^{T} X^{−2}A = I [REDES05]. y′ [Moh06]. y′ = f(x, y, y′) [Moh06]. y_{n+1} = (α + β exp(y_{n}))/(γ + y_{n−1}) [OBO06]. y_{n+1} = (α + y_{n−1})/(β + y_{n}) + y_{n−1}/y_{n} [OOB07]. y_{n+1} = (p + qy_{n} + ry_{n−k})/(1 + y_{n}) [MSD06b]. y_{n+1} = A + y_{n−k}/y_{n} [SA05a]. y_{n+1} = A + y_{n}/y_{n−k} [SA06c]. y_{n+1} = A + y_{n}/yn − k [SA06b]. y_{n} = A + y_{n−1}/(x_{n−r}y_{n−s}) [ZYME06]. y_{n} = by_{n−3}/(x_{n−q}y_{n−q}) [Özb07a]. y_{n} = by_{n−p}/(x_{n−q}y_{n−q}) [YLB05]. Z [HCC06, Yun07c]. Z^{2} [Aki05a]. ζ [AA09, Yan07d]. ZK(n, n) [WF07c, Waz06f, Waz05k].

-accretive [ZH08d, ZH09b]. -actions [Aki05a]. -Analytic [KRRGT07]. -approximation [Yoo05]. -B-spline [Lin05c]. -band [LXSH06]. -based [KK05a]. -Bernoulli [CAS09]. -Bernstein [Ost08]. -beta [FB08]. -Bloch [Ste09c, Ste09d]. -body [Smi06a]. -calculus [Rad09]. -centrum [PL05, PC07a]. -Chlodowsky [KG08]. -cocoercive [HN07b]. -constrained [Muk08b]. -constraint [Mav09]. -convergence [Bil07, Bil04]. -convex [KBÖP07, YL07]. -convolutions [SY09]. -cycles [Cas07b, Zho08a]. -cyclic [HS07b, Li08e]. -D [AS07b, AM08d, AKM06, DMA07b, Dön06b, HBE09, Hlo05, Lee05c, Moh07a, Ram07d, Ram07e, ZW05g, ZH05]. -d’Alembert [Prá08a]. -decomposition [YL08b]. -differentiability [Taw07]. -digit [SA08c]. -dimension [SC08a, ZLZZ06]. -dimensional [CW05a, CW05b, CZ08b, CIS09, DW09a, DKL09, FHL07, FL09b, GHL07, Gor06a, HC09, Inc06c, KK08c, LZ09a, LGw07, LH08d, Luo07, LG08, Ma07c, MWQ09, MD08a, MZ09, zPY06, SZ06b, SZ07d, SZ07e, TB09a, WSLZ07, WLZ05, WCZ06, WZ07c, WSZ07b, Waz08h, Waz08l, Waz08i, Waz08t, Waz09a, Waz09c, WS09b, XZ05, lYpZqL08, Yu08b, ZW09a, ZX06a, ZX06c, Zha07c, Zha07d, Zha08d, ZWT09c, Zha09f, ZL07b, ZLZ08, ZLYL08, Kis08, Su09, Wan06e]. -dissimilarity [BC09c]. -distance [IR08a]. -distribution [BMJED06]. -Durrmeyer [Gup08b, GF09]. -error [JZ06b]. -Euler [CAS09]. -exceptional [Li05a]. -expansion [AÖ09a, Asl09a, Asl09b, AÖ09b, GZ09a, GZ09b, LXML09, SC08a, WZL08b, ZG09, ZX06b, ZX07a, ZWT09c]. -extensions [CAS08]. -Fibonacci [FP09a]. -fold [GKR06]. -Fractional [Nis07]. -function [Yan07d, JZ07, Sim07, SB08b]. -fuzzy [ACOS06]. -Galerkin [LL12, LL09]. -gamma [FB08]. -Gauss [MBS07]. -generalized [ÖTA05, lY08b]. -harmonic [GÜ07b]. -Hurwitz [CAS09, Sim07]. -indeterminate [ST05c]. -integral [BK08c, FB08]. -inverse [PH06, UP07a, YW09]. -inverses [GWJ07]. -isogenies [MMS^{+}08]. -Jacobi [MBS07]. -Lagrangian [SPX07]. -Laplace [ZLCZ06]. -Laplacian [ANM07a, AMN07a, Dai09, Den09c, DLT08, DH09, DHG09, GZ07c, HC08a, Jan09c, JGY07, JG07b, Li08d, LZ08d, Liu08b, LOA05, Pin06, SM06c, SWX06a, Su08c, SLS08, SL06c, SG07c, SQG08, SMG08, WC08b, Xu07, Kar09a]. -Lucas [Kah05a, DOU08a, DOU08b, Dem08, YK03]. -machine [Ere09b, XYL10]. -matrices [WH09, AH07b, CKKS08, CKR09, CN09c, HL07a, HCC06, HWF07, qJ06, LS09c, LH06b, LH07c, XHL07, Yun07c, YK08, ZHGG08]. -matrix [Li07a, tLlWpL08, yS06c, Yua05]. -maximal [AV09]. -median [Fat06]. -method [TG08a]. -methods [LGY07, WL07c, WWL08b, WQL09, WWY08, LX07b]. -moduli [ZXL09]. -monotone [DW09c, yLcCsL09]. -Navier [JH09]. -neighborhood [YMCL06]. -neural [JRW08]. -Newton [RMS05]. -NNAF [DwWmW05]. -norm [Ma08b]. -order [SWZL07, Su08c, TNL09, WS09b, Ant05, HR07a, HR09a]. -Orthogonality [LO08]. -out-of- [ZW05d]. -plane [AA09]. -player [AR06, PSP05]. -point [CZ07c, DZ08f, IOS07, Kha07b, PDW06, Su08c, SL06c, WC07b, WP05, Wei06a, Wei06b, Xu08a, XZ09b, YW08, ZW07b, ZYW08]. -preconditioner [qJ06]. -processes [Pog07b]. -property [OML09]. -Pseudomonotone [Sad09]. -quasiconvex [YL07]. -records [FA06]. -refinement [Bie08a, Bie08b]. -relaxed [AV09]. -Restricted [WYL08b]. -schemes [Ran08b]. -Separable [AOS09]. -series [JV07, PST06]. -solitary-wave [XMGW09]. -soliton [Waz08q]. -space [AB05a]. -species [Che05a, Che07c, LL05f, QD08, XCH05b, CLH06, Che07b]. -spline [NKE05]. -stability [RC06, AM08b, AM08c, AM08d]. -stable [Inc07e, Inc07b, You09]. -starlike [Kan09a]. -step [CL05d]. -symmetric [YL07, HY07b, HY07c, PHZ08]. -th [EOM08, Fra09a]. -transform [BMY08, DSY08]. -two-variable [Sim07]. -type [ZHLG07]. -uniformly [JK08, Kan09a, NAUH09, RSSS07, KP06a, KPS06a, PS06b]. -valent [ARS07a, Aou07, MA09, NMM09]. -valently [Alt07b]. -Volkenborn [Kim07a]. -wavelet [LY09a]. -weakly [JOB08]. -weighted [LWW05, MD08b, QW05, WG05].

0-1 [JHST05, JRR^{+}07].

1 [CK06b, CK07b, hLsT07, MX08b, XM08, Zho05]. 1-dimensional [Zhi09]. 1-Soliton [BR09b]. 1.839 [RW07]. 10th [SA07b]. 10th-order [SA07b]. 12th [SA08e]. 145 [dCD06a]. 147 [KAK05, KAK06, Kas06b, Sub07]. 152 [Pan08c]. 160 [Pan12]. 1684 [EM11]. 174 [Mar14]. 186 [Hai08]. 191 [Den08b]. 196 [WZ13]. 199 [SQ11].

2-cyclic [DH06b]. 2-isogenies [MMS^{+}06]. 2-variate [AH08a]. 201 [hMfWzLpZ09]. 212 [LL12]. 215 [Tua13]. 2DPCA [KSC^{+}08]. 2SGP [HTO06c].

3* [KH05a]. 3-node [BDDB08]. 3-space [KA06c]. 3-species [Wan08f]. 352 [Den08b]. 3R [BNdFPS06]. 3rd [WZS07].

4-dimensional [LM09]. 4-point [KLY07b]. 40m [dCDLN^{+}07].

5-stage [NBHYV09].

752 [HES05].

802.11e [ZWM08].

= [AV06, KHZ07, LpL07, yP05, hPyHZ06, PHZ06b, XL09b, dCD06a].

A. [Bra09]. Abaffian [AMAP07]. Abel [pLT06, You06b]. Abelian [ZL08e]. Abo [Pan08b]. Abo-Eldahab [Pan08b]. above [Sel05]. ABS [AMAP07]. ABS-type [AMAP07]. absolute [CZ07a, Cha05a, Cha05b, Cha05c, CDT07, CGSS09, TZX07, WL06g]. absolute-value [Cha05a, Cha05b]. Absorbing [WL08b, Ahm05a, Ahm07b, O’R09, OB09]. absorption [Ahm06, MTL08]. absorptivity [CYL07]. abstract [AAG05, KWX08]. abstraction [TCDV07]. Abundant [SQHY08, Waz06a, ZZB08, MLW09, Waz07a]. accelerate [ZW06d]. Accelerated [GSPG07, RJ07, AC07a, Chu07a, DHY06, EH06, GSDB07b, KAF08, KLW07e, SG07b, Wan06a]. accelerating [CZ05a, Thu07a, Thu07b, Thu07c, VFF08a, WZ08f]. Acceleration [And09, MCMA07, Ned05]. accelerometer [CYL^{+}08]. accept [Nwo06e]. access [CH05b, Che08c, LY05, WC05d]. accessible [GMS06]. according [MSM07a]. account [ESA05]. Accounting [Sae07]. accretion [Dön06a]. accretive [ĆRR^{+}08, ZH08d, ZH09b]. accumulation [TG06a]. accumulative [Oht08]. Accuracy [HM09a, AY05b, CR06, Cot08, KM06, KSM09, pLT06, MV09b, Moh07a, Moh09, RLZ08, ST06b, SYH07, Shi05b, Shi05c, Tie05, Wan07b, Yoo05]. Accurate [CL07c, DL09, Dan06, EF07, CCJR07, EFI06, Kar05a, KM05d, LJ07, MD09b, NLD08, SFH09, SdHT09, Vaj07b, Zha07h]. ACM [ZSNS05]. ACM-EC’03 [ZSNS05]. ACO [CK08, KF08b]. Acoustic [Ahm07a, ARM09, Ahm07b, Ate06, FHG07, HFL06, ZL08c]. acting [SiU09]. action [SANN07]. actions [Aki05a, LCJ08]. activations [GH09]. active [JjQjWmL06, LZ06a]. active-set [JjQjWmL06, LZ06a]. activity [CYL^{+}08, KL05a]. actor [DF08]. actuated [ZS08b]. actuator [Yu05]. actuators [LL06a]. acute [CL08a]. ad [HRM07, Zha08h]. Adams [GSNDB09, OISM07]. Adams-like [GSNDB09]. adaptable [GPR^{+}09]. adaptation [BDCW06, CGS08, KR07b, YYYM07, ZBH06]. adapted [Ami05, Ram05e, Van07]. Adaptive [Bie08a, Day05, HLY09, Jia07b, Pen09, Ql08, SW09b, XXTL09, ZWLX08, Zhu09a, lS08a, Bie08b, Che07h, CCZ08, CR05a, DAQ09, EH08a, FS05, FB07, GJ07a, GLLC08, GHZS07, Han06, HL08a, HH09a, HCX06, Li07c, LZX08, MX08b, PLTW07, SJ08a, Sal05, WW06, YZDW07, YZZP08, YZW07c, YZD08, ZS07c, dlS06c, HK06]. adaptive-homotopy [HH09a]. added [SW09b]. added-order [SW09b]. Addendum [dCD06a]. addition [LXH06]. additional [HH06a, SG08c, SDM08]. Additive [JZ07, Qia09, yWjCnG07, YY05a, Aki05a, Aki05b, Akı06, Bry07a, CS08a, Dao05, DG07b, JJ09b, gLXyZ06, PH09, RSZ08, WGL08, XRZ09, ZZ07c]. ADI [LZ08a, YGY09]. adiabatic [Fag05, Smi06b]. adjoint [AS09a, KA05a, KK07a, Pat05b, RK07b]. adjustable [WC05c, Wu06c]. Adleman [HC08b]. ADM [GK05b, KI05a, Kay05b, YCL09a]. ADM-Padé [YCL09a]. admissibility [ADM05]. admissible [O’R09, ZN05, ZLW06b]. admission [CD09]. admits [Hub09a]. Adomian [Abb05, AD05a, AD05b, Abb06a, Abb06c, Abb06d, All05a, All07a, AH07d, Att05b, BDDB08, BD05a, BGA05, BA05a, BE05, BIK05, Bia05, BIK06, Bia06, BB05, BKBK06, Bou06a, BB06d, BEF08, DGJ07, DGB08, DHS07, EG07a, EWAE06, EWEA06, EA06, GJB05, Gor07, GbL07, HNB06, HM07, Hos06b, Hos06a, HN06b, HN07a, IRSR05, JDG06a, JDG06b, JDG06c, JDG06d, Jan07a, JL05, Lay08, Luo05b, Mak07a, Mak07b, MO06, Pam05a, Pam06, Pou06, QS09, RN09, RB05a, RB05b, RB06, Ray08, Sya05a, Wan06e, Wan06f, Waz05b, Waz05a, Waz05c, Waz06k, Wu06b, ZWL06, ZLW06a, ZCW05]. adsorbate [RY05b]. adsorbate-induced [RY05b]. adsorption [Hua05a, Rem07]. Advanced [BUP07, Hua08c, Jan08b, Jan09c, Psi09]. advantages [Vaj07b]. advection [DIT06, Deh05b, DCS07, Li09f, LZA^{+}07, Ram09b]. advertising [Çet06, YX06b]. aeroassisted [ES05a]. affected [NCÁHC^{+}06]. affecting [Liu06a]. affine [AST08, Hos08, JZ08b, WZ09c, WZ09a, ZW08b, Zhu05a, Zhu06a, Zhu07]. after [Gal06b, MJ07e]. again [Dan08]. against [HW05a, Hsu05b, RYE07, RR06]. AGE [BÖ05, AAA08, AAA09, CH09, LGw07, Luo06, Luo07, LG08, PRM06, PLL08, RS08b, SMM08, mZzH05, ZH07a, ZWZ05]. age-dependent [AAA08, AAA09, LGw07, Luo06, Luo07, LG08, PRM06, PLL08, mZzH05, ZWZ05]. age-distributed [CH09]. age-structured [SMM08, ZH07a]. agent [Dam08, DvdHD08, GAS08]. agent-based [Dam08, DvdHD08]. agents [MTD05]. aggregate [BCMT06]. aggregation [DE07b]. agreement [CYY05, CY05b, CJL05b, DL05, KRY05, KHKL05, LL05b, LKY05b, LKY05c, LKY05d, Shi05a, SW05b, SC05c, YW05, YRY05c]. agricultural [DSMK08]. agriculture [RR06]. AGV [LFM06]. AHP [AGI08, GANG07]. aid [Lep08]. aided [ESAH10, LW06a, WC05b]. AIDS [DXH08, KRG06, Kov07, NTO06]. aimed [DS06b]. airflow [CLNCG07]. airfoils [CL08e]. Aitken [LMV09]. AKNS [SZ08c]. al [MK08b, MK10, ZW05c, d’O05a, Bul09, XFTY09]. Al-Oboudi [Bul09]. al. [BCW05, HH05, Hwa05, KJY05, PKH05, PK05a, SHSZ07, Xie05, YRY05b, YRY05d, YY05d, ZSNS05, ZAX05]. alcohol [DRS^{+}08]. Alexander [d’O05c]. algal [SMBB07, SMC08]. Algebra [Zot08, AV06, Alt06, DW09d, ILZ09, KH08, ML07, MH06b, MH06a, YL09c]. algebraic [Att06c, ANK05, BLW06b, Bog05d, CL09a, CW05a, CW05b, CMS05, CMR08, FGfFL06, Far05, GJ07a, GAV08, GB06a, GB06b, Hos05b, Hos06a, Ide08b, Ide08a, Iva08, Jan08a, Jan09b, LYX08, LXSH06, LS06e, LS07d, LN06, Luo05a, MG06, Muk07c, Muk07b, Muk08b, NP05, PVN06a, PVN06b, SC05b, SJ05, WCZ06, WZSZ06, WHL08, ZZL05, ZXDL05]. algebraically [Gin07]. algebric [GJ07d]. Algorithm [DF09, YGL05, Zha05d, Zha06c, AJAH06, AAES07, AZ08, ASAS06, ASAEE06, ARK07, ABK^{+}06, ANM06, AMN06, ANM07a, AMN07a, AMN07b, AMN07g, ANM07g, AB05b, ASK05, AT07a, AKB09a, AKB09b, Ali05, AV06, ACDR09, AH09, ACD^{+}05, Ars07b, Ars07d, AH08c, ADDdM07, AGTS07, AGA^{+}07, APS05, BK07b, BK08a, BK07c, BHCC09, BIK06, Bog08, BS05c, Cas07a, CK08, CS07d, CL06a, CCFC08, CC08c, CC09c, CWLH08, CKY05, CH06a, CYL^{+}08, CLW09a, CAA07, Chi07a, Chi07c, Chi08a, Chi08b, Chi06b, Chi06c, CK06a, CJK07, CTL09, CK06c, CYH^{+}07, CT09b, CS06b, CST07, CMS05, CL06d, Cui08, DHL05, DZZ07, DS05a, DN07b, DJM08, DD08a, DD08b, DSKM09, DH08, DT06, DE07b, DM08a, Din05, DX07b, DN07c, DS06a, DS06b, DZ08e, DZ06, EAEG07, EM05a, EM08, EMR08, EMR09, ESK07, EH09a, FY08, FE07]. algorithm [FG07, Fat06, FAM07, FMH07, FL06, FWJK09, GXY06, wG07a, GC06b, Gee08, GTA06, Geo07, GH07, GAFK07, GK06b, GWX06, GY07b, xGyZtZ^{+}07, GG05b, Gui05, GBG07, HEZ07, HEZ08, HE08a, Ham05b, Ham08, HWH09, Har06, HJB06, HN06a, Has06b, HCLPJL^{+}08, HRT07, HC05a, HZW06, HD07, HL07b, HH06b, HL05a, HZ06d, HTO06a, HTO06b, HW06, HJL08, IL08, Ji05, JS06, JZQ07, JWL^{+}07, JZ08b, Jia05, JZ07, JHHW07, JC08b, JRR^{+}07, JRA^{+}07, JM06c, KWdW08, KA09a, Kar06, Kas07, KK05b, KK07d, KMN09, KS05a, KEM08, KO08, KCK08, Kıy06, KA06d, KSP08, KY06, KT08b, LWH06a, LMT^{+}09, LpL07, LZC05, Li06d, LZ06b, Li07b, LD07b, LYKC07, LM08b, LLFM08, LL08c, LML08, LL08b, Li09d, LGJ06, LJG06, LHB06, LKC05, LC06b, Lin07b, Lin07d, LG07, LZX08, LH06a, LG06b, LL06c, pLT06, LLY06, LHC07, LXZ09, LLWC07, LP08c, Lu07, LXZ07, LC06d]. algorithm [LX06a, MFD07, MYSA05, MAT08, MBM06, MM07c, XSX08, MD05, MKSK07, MMS^{+}06, Moh07b, MK07b, MM05d, Mor05, MCGS05, MED08, Muk07c, Nab09, NS05, NN06a, NG06a, NG06b, NR07b, NS08, NM08a, NGLHCLP06, NV07, NA08, NK07b, NTZ08, NB07, Ohm08, Ony05a, OS09, OASM05c, OASM05a, OASM05b, OASM06, Oss05b, Pal09, PL05, PL06a, PHL06, Pan08a, PU08, PQJ08, PN06, PN08, PK08d, PL06b, hPyHZ06, hPY06, PLTW07, Pen09, PFA07, Pou06, Ql08, fQZgH08, QZZ07, QZJ07a, QZJ07b, RKS08, RM05, RM06, Ree07, Rig08, RS06, SBS08, Sae06a, SSB06, Sal07a, ST05a, SKT06, SS06a, SCR08c, Sha08a, SG05b, SPX07, Sha08d, SHK06, SM06b, SCS05a, She06, SG08c, SS05c, SLZZ06, SLWX06, SDR07, Sog08a, Sog08c, SM09a, SCI08, SQ07, SCS06, SW06b, SHSZ07, Sun07a, SC07f, Tak05, TCW08, TLB07, TMMVA06]. algorithm [TMSZ06, TK07, TS08b, TS06, TG08b, TA05a, THLC07, TAK07b, Tso08, Tso09, Tze07, UP07b, Vaj07c, Wan04, Wan05g, Wan05h, WL05e, Wan05d, WTW05, WL05d, WL05c, Wan06g, WL06c, WJ06, WZT06, WL06b, Wan06i, Wan06h, Wan06j, WZ06a, WHH07, WWWL07, WZ07b, WC08a, WD08, WYZL08, WS08a, WZ09a, WB09, WZ05, WCX08, Wu05a, Wu05e, WL06g, WLZ06, WS06b, Wu07f, WWHL08, Wu08d, WWC09, XLwW07, XZM^{+}09, XW07a, nXwCpC06, XCS07, XQ06b, XLSL06, XBZ06, XLH^{+}07, XL07, YB03, Yam05a, YCW05, YC06a, YDL07, YPZ05, YLY07, YC07a, YLK09, YE07, Ye06, YZZP08, YC09c, YSCC05, YQ09, YRY05d, You05, YEA07, YHL08b, YHL09a, YGY09, YK06a, (Ag06, (Ag07, ZGH06, ZGK06, ZL05a, ZMW05, ZZLL07, Zha07b, ZLH^{+}09, bZlDL09, ZW08b, ZW05g, ZL06b, Zho07, Zho08a, jZzFbL06, ZW06f, gZW08, Zho09a, Zhu05a, Zhu05f, ZCW05, ZYZ06, ZZ06b, Zhu06c, AG08, Lin07e]. Algorithm [PBG07]. algorithm-fuzzy [OASM05c]. Algorithmic [Vel06, CS07b, JLI06a, LWL07a, LH06b, Mou08, WLT07b, WLL07a, XSKC07]. Algorithmization [Zha05e]. Algorithms [Bog05b, Lin07c, Abd06, ANM07h, AK08a, AM08b, AM08c, AM08d, And09, AZ07, BNR06, Bog05a, Bog05c, CPY09, CXZL09, CH07f, CY06b, DL09, DT07a, DT07b, DLS06d, EG07c, Fad05, FYZ07, Ghi06, GHZS07, HH06a, HM09a, HBJ06, HGN06, HRM07, HCLP07b, HZ05, KM06, Kar07b, Kar07a, KW05b, KW05a, Kay05d, KB05a, KTÖA08, KLZ07, KX08, LZ08a, LZ07a, LY07c, LLX07b, Lin06b, Lin06a, LG06c, LWKU09, MK05b, cMpZfR06, MCAF08, MM05a, Mei06, Mis05, MK06, Mor08b, NKE05, NS05, NS06a, Nak08, Nwo06a, OHT05a, PB07b, PSW07, PBPF08, PE07b, Qia09, QCS09, RK07a, SMA06, Sal06b, SCR08b, SSJK07, Sog08d, Sog08b, SG05d, STV09, Sun08, TH07b, TTZ09, Thu06, Thu07b, TC05, Vaj07a, Ver07, VAMV07, WY05a, WDL07, WCW07a, WH08a, WXW09, XD07, Yan07c, YLM07, ZLW06b, ZY08c, Zhu05b]. alien [BD06b]. Allee [SS05f, SS05h]. allelopathy [ABB06e, MKG09]. Allen [TB09b]. Allocating [AK05d]. allocation [ATM06, Cai06a, ÇE06a, CK08, Che06l, GL06b, JLM05, Kun08, LG07, LPW08, OASM05a, Sad07, YW06a, YW06b, YYWW07]. alloy [dCJLNÁA08]. Almost [Che06a, LC09b, QD08, XL08a, Bai08, Che05c, CS06a, Che06d, CJP08, hLqSlZ09, LC08a, MXC07, TQC09, WL07e, WS07b, ZS07b, dAC09]. along [Mak06, SAM07]. also [AKR08]. alternate [BIK06, MMW09, PSS07]. Alternating [Gw05, Alt07a, BÖ05, BA05b, Bog05a, CY08b, LY07a, LY07b, NCGdSM07, PST06, TD07b, Thu07c, WT09, ZCH07, HL06a]. Alternating-direction [Gw05, LY07a, LY07b, NCGdSM07]. Alternative [QCDS09, BRRGT07, CCC07, Cor05b, ESK07, Li06d, LJE07a, PH06, WY05a, Sil06a, Sil06b]. alternatives [GANG07, MF09]. Altman [ZZZ09]. Amartya [Tan06b]. American [AS07a, CWY06, CK06d, DL09, KF08c, SS06e, YC08b, You09]. American-style [CWY06, DL09, YC08b]. among [HN06a, Ino08, MS07c, OHJ07, SML05a, SZL^{+}07]. amplifiers [Liu06g]. amplitude [BV06, CHH09, CDT07]. AMRM [AV09]. Analog [Myc06, Bou06c, HJL08, TTG07, WLLL07]. analogous [FBB08]. analogue [CZF09]. analogy [ACZ05]. analyses [WC05b]. Analysis [ASL07, DHT09, FX09, FA05b, HZYL08, Kad07a, Li05a, gLXyZ06, NTZ08, PDGR07, Sil05c, Sil05a, Sil05d, Sil05b, Sil06g, Sil06c, Sil06d, Sil06a, Sil06f, Sil06b, Sil06e, Sil06h, Sil07b, Sil07a, SJW09, TG06a, Toh05, TY07, cWmZ06, Wen05, Yan08e, ZSS08a, ZSS08b, ATL07, AEFFA08, AS09a, All07a, AP06, AJK05, Ami06, AKS06b, AN06, Ami07, AKA07, AC06, AGH05, ALN06, BD06a, BT05, BY05, BE06, BM05a, BK06b, Çak08, CZH07, ÇFÇ06b, Cel06a, CN06, CS07g, mCLB08, CL08e, CLA08, CXZL09, Che09b, CFJ^{+}05, CTGD08, CW08, CK06d, CG06, Cor05b, DY06, DD07, DBR07, DESG08, dCDLN^{+}07, dCDNPN09, DW09b, DK05, DLS06d, DW09c, DEFT05, DX06, DC06e, EH06, EMP08, ET05b, FK09, GM06a, GI06, GK09, GANG07, Gop07, GS06e, HBBA07, Hu07a, HH09b, HTO06a, HCZ06, Inc06a, JF05]. analysis [JMHS05, JLZ05b, JLS^{+}05b, JHS^{+}05, JPZ06, JSdM07, JLBR07, JK08, JL07, JW09b, JSW05, JY06, JC08a, JY09, dCJLNÁA08, Kan06a, KWdW08, KI05a, KKKM09, KL06a, KC06b, KC07b, Kho07, KWZ08, KP09b, LPP08, LJK07, LW09a, Li09f, LYJ09, LX09a, Lia05, Lie06, LC05c, Liu06a, LWL06, LC06c, LGY07, LMY07, LJE07b, LZWK09, Ma08a, MUZG09, MC07, Mai05, MDZ08, Mit06, MB08, NNN05, NPGW09, OF06, Omr07a, Ozb07b, Par06c, PKT06a, PKT06b, PG07b, PG07d, QCS09, Ram07b, RW09, REN^{+}05, SMM08, Sae05, SML05b, SML05a, SB08a, SH08a, SS08a, SMG07, SL09b, SP05, SMC08, SGB05, SMD06, SP06b, Sun05a, SC06c, SW06c, SHSZ07, SM08d, SM08e, TKA^{+}08, TG06b, TKJ08, WWP05, WLT05b, WB06, WC06a, WZ06a, WF07a, WZC08, WLZ08, WC09b, WLH09, Wan09a, WC05d, WYL06d, WS06a, WRB07, Wu07d, WLL07b, WCH08b, XLHK09, YAO09, hYG05, YZW07a]. analysis [YX06b, YLL07, YWL07b, YZFL08, YQYL07, Yüc08, YB05, ZLMW05, Zha05c, mZzH05, yZ07e, ZLZZ07, ZXL05, ZD06b, ZZW09b, Zho05, ZW06e, ZFT06, ZWZZ07, ZY08d, ZXL06, d’O05a]. Analytic [AÖ09a, CKS09, HAK07a, KS08f, LS07e, Mom05a, PB07a, SMB08, SOS07, TB09a, Waz06b, Waz08a, XMGW09, ZM09, AJMR07, Alt07b, Aou07, BA05a, BW09, CLW09b, Den09b, Eli07, GOPP07, HT06b, HL08d, Irm07, KKA09b, Kum08, LL05e, LS09d, LL05g, Liu07c, LN07, LL08f, Liu09c, NO07, NM08b, Noo08a, NB09, Ozd05, PS06c, PKT06a, PKT06b, QS09, SOT07, SS07b, Sok07b, Sok08a, SEŞ09, WL06d, WZ13, Waz06i, Waz06m, Waz09f, Yuj07, KRRGT07]. Analytical [Ben07b, EEAES06, ESS05a, MNG09, MSH08, Mom05b, MO06, QS06, RN09, RB05a, RCB06, RB06, Waz05c, ZPT^{+}06, Cai06b, DNHF05, Kum09a, Mar05c, ÖK05, Ram05h, Ram05i, RCB08, Sil07a, Soh06, SI06, TF07, YC09a, Yig08]. analyze [KJ09, YWJ07]. Analyzing [YJW06, AGT07]. anatomy [CDT07]. ancient [Wu05a]. and/or [Shi09a]. angiogenic [EP07]. Angled [MC06, McC05, McC06]. angular [SM08e]. anharmonic [SMG07]. animals [LW06b, WL05a]. anisotropic [Ang05, BC09a, CMM07, GC07c, HS09b, SL09b, SW09a]. annealing [AJAH06, Alr05, AD09a, GAFK07, GÜ07b, JJT06, LPA^{+}09, SK07a, TMSG06, WZ06e]. ANNs [TG06b]. annular [Rad05b, YC06c]. anomalous [Su09]. anomaly [TX08]. anonymous [YTWY05]. anorexia [HCW06]. ANOVA [AGS07]. ANP [Shy06]. Ant [FYZ07, SK08c, Tok06, YW06a, BHCC09, CK08, DT06, DS06b, HP07, KF08b, LZX08, SAK08, SSJK07, TS08b]. ante [KC09]. anterior [AS06i]. Anti [DXH07b, Ami07, CS06b, EMR08, EMR09, GG08e, HY07b, HYG07, MJ07e, Pen05, Rim08a, Rim08b, Rim08c, Rim09b, Rim09a, SW09b, WL06e, Yin08a, ZZH09]. anti-cent-dian [CS06b]. anti-diagonals [GG08e]. anti-efficient [Ami07]. anti-Hermitian [HY07b]. anti-ideal [WL06e]. anti-malicious [MJ07e]. anti-pentadiagonal [EMR09, Rim09a]. Anti-periodic [DXH07b]. anti-reflexive [Pen05, ZZH09]. anti-symmetric [HYG07]. anti-synchronization [SW09b]. anti-tridiagonal [EMR08, GG08e, Rim08a, Rim08b, Rim08c, Rim09b, Yin08a]. antiangiogenic [dG06]. antibiotic [MBC08]. anticancer [dG06]. anticipated [WC06b]. anticipation [AS08]. Anticontrol [GZ07b]. antiperiodic [ZS08a]. antiplane [LD06a, Shi09a]. antiviral [d’O05b]. AntPart [ABK^{+}06]. ants [ABK^{+}06]. ANU [KK05d, KK05c]. ANU-cluster [KK05d, KK05c]. any [WZSZ06, ZZ07a]. AOR [WH09, CK07c, DH06a, GH06, HS07b, LD07c, tLxLlW07, LY08a, YK08, Yun09]. apex [She05b]. Apostol [CAS08]. appearance [FLO07]. Appell [OSS09]. Appl [Bil07, EOM08, ESAH10, HLSS10, KAK05, KAK06, Kas06b, Kum09c, KP06b, LL12, hMfWzLpZ09, MK10, Pak11, Pan08b, Pre06b, Sal07d, Sol07, Tua13, UNS07, WL08a, Wan08b, WZ13, Waz08b, Wey08, WH09, Yin08b, YHL09b, YHL09a, dCD06a]. Appl. [Mar14]. applicable [Boy07b, ZFN07]. Application [ADK06a, BDE06, BI08b, CL07a, DN07c, Dog05, DD09, EGN07, EK06, FAM07, GKB07a, GJ07b, GJ07c, HR07a, KPS06b, Kuo05, MK06, ONG08, RCB08, SDL09, SP05, SZ08a, WZL08b, WLK08, XXL09, ZS09b, ZWZL06, Ahm05b, AN06, An05b, AGJ07, BL06, CW06c, DZZ07, DM07a, DFGV07, EN06, ESS05a, EA06, FLLH07a, FL09a, GZ09a, GZ09b, GK05a, GK09, GHZS07, GVMT07, GBG07, Hub09c, JZ06c, KI05a, KI05b, KKY07b, KOS08, KH07b, LL05a, lLWhL06, LZ07b, Lin05c, Liu07e, LZ07d, MBM06, MS07c, Men08, MKSK07, MO07, MD09b, Ony05b, Pam05a, PK05d, PSBM08, PG08c, RVN07, RNV07b, Ray08, Rey05a, SYT06, SZ07d, SZ07e, TT06a, WL05c, Wan06e, WGF^{+}06, WZSZ06, WSB06, WSL07, WCX08, WS08e, YM06, lYqLW07, YZZP08, YLL07, You09, YZD08, YQYL07, ZW09a, ZC06b, Zha07c, Zha07d, ZDL07, ZZZ09, ZWT09c, ZZZ07, ZZYZ06b, ZXZL07, ZY08d]. application [ZH08d, BDG08, Boy06b, GAV08, KRG06, MUZG09, OME08, VGMV^{+}07]. Applications [BR08b, DLL09, Dzi09, GPS09, HMNT09, KWS05, ZZ07d, AK06a, Abd07, ADZ05, AMJD08, Alt07a, ÁCTZ05, AA06d, Bar09, Bel05, BJ08, Che07g, CS07i, Cui05, Dan06, DSZ07, Den09c, DSY08, EFE08, EELZS05, FO09, FHZ08, FPP06, GGOS08, GGSNR06, HPD07, HHC08, Irm07, JV07, Kat08, KZ07, KK05d, Noo07a, OSS09, Özt05, PS08, zPY06, RJM07a, SWH07, uITA06, SM05c, Sjö07, Sok08c, SOS07, TF07, TS08c, WWL06, Wan06c, WL07a, Wan07e, WCN08, WYL08a, WSL08, WN08b, WN08c, Wu05e, Wu08c, XBZ06, XS06, YKL08, YLX08, YL08b, YYMZ09, ZG09, ZWT09b, ZWT09a, ZTWL09, ZY08c, Zha09h, Zha09g, ZWZ09, ZYD07, GGSN06, Kat09, RBvB08, Tou09]. Applied [Den08b, EM11, HV08, Hai08, HES05, KPL09a, Li05f, ORG^{+}09, Pan08c, Pan12, SQ11, Sil06h, Sub07, Yam05b, ZW06b, Alr05, Bah05, CCE06, Dem05d, DM07b, EKS08, ESMM06, GS08a, HP07, Hua08c, IR08b, Lee05c, LWH06b, LS07d, MM08b, ÖÖ06, RNRRMRCH09, SDM^{+}09, SY08b, WH06a, dSYY08]. Applying [ÁFVZ06, CW06b, HAS07b, LMSV07, Lee05b, ANA07]. Approach [Zha09a, ASL07, ASAEE06, AD06b, AuH07, AA07d, AM08b, AM08c, AM08d, AK05d, AJK05, AS06f, AKR06, Ang06, ALN06, APS05, BC09a, BS05a, BCEGMRMP07, BHN08, BIK05, BOB05, BBCP09, BB09, BKFM05, BKM06, CTT08, CF06a, ÇE06a, CN05, Cha05a, Cha05b, Cha05c, CZZ05, Che06l, Che06k, CHW07, Che07g, CAL08, CHY^{+}09, Dan05, DES07b, DZ08e, EARS07, EF07, EMC05, Ema06b, FZN07, FK09, Fil05, GBK06, GXS05, GTJ07, Ghi06, GG07, Gül06, GS07c, HK06, HL08a, HN06a, HYZ07, HYL09, HH06b, HZ06b, HCX06, HLK06, Hub09c, HS09c, Ibr08, ICM08, Isk07, IR06, IR07, Isl08, JLM05, JLS^{+}05b, Jin08, JL09, KKF07, SM08f, KL06a, KHL06, KSSA07, KSM09, KR07c, KM05c, Kop22, KS09a, KP06b, KP06c, KPL09b, LH05, LL05e, Li06e, LZL^{+}07, LLLX07, Li09e, LYL^{+}08, Liu06e, LJE07a]. approach [LMD^{+}08, MUZG09, MM06a, Mak07a, MSH08, MA05f, Mir06a, ME07, NCÁHCLP05b, NHCLP07, NNN05, NHKZ06, OB09, OT06, OASM05a, OASM05b, ÖA05, PS06a, Par05c, Par05a, PKL08b, PKL08a, PN08, PSS07, PG07b, PG07c, QZ09, RAES05, Ran08c, RMJG07, RZ07a, RGM07, Ray08, RR07, Saa05, Sae07, SD09, Sal06d, SAK08, SK07a, SG05a, SGA06, SK06b, SG08b, SJ06c, SC08a, SDM08, SLZ05b, SLZ05a, SZL^{+}07, uIKTT05, uIT06, uINTK06, uITiHK08, SSTD05a, SMD06, SdJA06, SW06c, SM07d, Tan06c, TK06c, Tse05, VRG06, VL09, WC05a, WZ09b, WS09b, WL08c, XZB05, XLHK09, XLKH09, YWZW07, YW06b, YEN08, Yu05, ZGH06, ZC05a, ZWZZ07, dEdSNL05]. approaches [DDD05, HCLP07a, KE05b, MJ07a, MPT08, Mou08, NMM08, yN06, Nwo06c, Nwo06d, PC07a, RLZ07, RAS08, ST06a, SD06, WYX05, WS07c, WLT07a, YC07b]. appropriate [JCL08]. approximant [GIT07]. approximants [Waz06h, Waz06k]. Approximate [ADH05, AKMA05, BG07b, Çel05a, Çel05b, ÇG05a, CM09, CM08b, DSK08, GSG06, GS07c, GS06e, HIST06, HMSZ07, KKT05, Mar05c, MM08c, OM06, QS09, SG06, Sun06a, Wan08a, Yan05a, Yig08, AKBZ08, AKM05, AO05c, AO06c, BJ08, Cha05a, CW06c, EL05a, ET05a, Fer09, GG08a, GEG07, GS05, Inc07a, Jbi06, JR07, KPS06b, Kia05, KT07b, KPG07, LS09b, MMO07, Mom05b, Mom05a, ÖA05, PSD07, QM05, Ram06e, Ram06f, RJF07, RB05b, RCB06, ST05a, Sal06d, SKT06, SPX07, WCLZ05, WZL08b, WLK08, WLH09, WLZ06, YS06a]. approximated [Boy07b]. Approximating [CLCK09, yMT08, WZDgH09, AKR08, HR07b, KK07c, KLY07b, Liu08a, Özb06, Sya05a, YLY07]. Approximation [BM06a, Beh06, KKS06, Lal09, Tia09, YKL08, AA06a, AA06b, AÖÖ06, ADY06, AH05b, ALL07b, BFR07, BS06, BA05a, BE05, BS07c, Boy06b, BW09, BX09, BK05b, CMGMD06, CY07a, CY08a, Chi07a, CE07b, CE07a, Dai07, DD08a, DMA07a, DH08, DNS08, DR06, DWT05b, EA07a, EDS06, Fag05, GC06b, GÖ05, Gup08a, GI08, Gup08b, GLMLP09, HA05, HL08a, He06, HL06b, HD07, HY07b, HY07c, Hua09, IDS06, IM07, Ism05, JKN09, KMB05, KBH06b, KG08, KKA09a, KKA05, Kha09, LLG06, LX07a, LHZ09, LLAT09, gLL07, LjZH08, MW09, MM06a, Mak07c, MC06, Oca07, Ouj06, PK08d, Pen05, PHZ05, PHZ06b, Rad09, Sal06e, SAS07, SA08c, Sha07c, SG09b, SP06a, SM09a, Wan05e, WYQ05, Xia06, Yoo05, Yu08a, YKR06, ZSM07, ZS07a, ZXF07, ZHZ08, Zho08a]. Approximations [Kha07a, Odi06a, AK05b, CCJR07, DNHF05, Deh05a, FA06, Fro06, mHsLzC06, IRSR05, KKT07, Kha07b, Kha08a, KB05b, LKU07, LKU09, LST09, Mel08, NRLT07, Omr06, RJW07, Sha08b, SFH09, Wan05b, WX09, XPS07, You08]. approximative [SC07b, Wan08g]. aquarium [FJ06]. Aqueous [AS06h, ORG^{+}09, AS06i, AS07c, BTI06, MH08a]. Arbitrarily [Dai09]. arbitrary [Ahm05a, EMESES05, HK09, HHK^{+}05, IMB09, IH09, LR07a, LR07b, Li08a, Oss05b, RNV07a, Rim05g, Rim05a, Rim05h, Rim05b, Rim05e, Rim05c, Rim05d, Rim05f, Rim06c, Rim06g, Rim06f, Rim06d, Rim06h, Rim06a, Rim06b, Rim06e, Rim07c, Rim07g, Rim07h, Rim07d, Rim07f, Rim07a, Rim07b, Rim07e, Rim08a, Rim08b, Rim08c, Rim08d, Rim08e, Rim09b, Rim09a, Van09, Yin08a, lS08b]. arc [PSRN06, WYZL08]. arc/circle [WYZL08]. arched [SC06c]. archetypal [dSF05]. Archimedean [Ali05, MA05f]. architecture [INR08, JY07a]. arcs [SP06d]. area [Chi07c, Chi07h, CR05b, HT06a, Imr06]. area-compaction [HT06a]. Argument [Aou08a, BUÇ07, BKS^{+}06, SG05a, SdHT09, YS06a, ZM08c]. arguments [Ada07, Che06f, Dyk09, Jan08b, Jan09a, Jan09c, LMY07, LYL07, MX07a, MX07b, PRA^{+}09]. ARIMA [OHT05a]. arising [AV07a, AVM^{+}09, BLW06b, Cao06, CD08, DMA07a, GK09, HES05, HAM04, KPS06a, KB06a, MR05a, MGMC05, MAD07, NDD05, O’R08, RJW07, RMJ07, SP08, TY07, Yan08b]. arithmetic [CK05, FP09a, KP07, MK06, Moh06, WLLC07, ZXL09, Gal06a]. arm [LWL07b]. ARMA [Nwo06d]. armed [KX08]. armlets [CWZ07]. Arnoldi [NG06a, NG06b, TGSNB07, Ye06]. arrangement [RH08, TTH08]. array [KCW07, Mis05, Sil06d, Sil06b, Sil06e]. arrays [Sil05b]. Arrhythmia [PG07b]. arrival [Cho07c, KL06b, KHL06]. arrivals [Cha09, WY09, Zho05]. Arrow [Tan06a, Tan06c, Tan06b, Tan07a]. arterial [SVMG09]. artery [BA07, IAMA08, MCMA07]. Article [Che08c, YSZG05]. artifact [Ma05b]. artifact-free [Ma05b]. Artificial [ANA07, Imr06, PSBM08, AN07, AAES07, AGTS07, AGA^{+}07, CCFC08, EAG08, EEAES06, ESS05a, ESS05b, ESAR05, EA08b, Fil05, Gop07, KBTS07, PG07d, QM05, Ram08b, Ser09, SS05e, SLX^{+}05b, YY07c, KA09a]. ARX [WDL07]. ARX-like [WDL07]. Ashraf [Pan08c]. Asian [CL07c, DHL05, DL09, HL07b, YC08b]. aspect [Fad06]. aspects [AH06b, GGSNR06, GGSN06, Hos08, MPG07, NKR^{+}08]. assembly [AK06i, KO06, TMRV06]. assess [PB07a, SMB08]. Assessing [YF08]. assessment [AGA^{+}07, AGM^{+}07, HWC07, JLBR07, MPT08, WL06e, WL05g, Whi06, Wu06a, WLL07a]. asset [CWY06, GL06b, Kun08, PS06a]. assets [Zha07e]. assignment [ÇE06a, Dan05, DT06, GH07, JA08, Kum06, LL06c, SAK08, SAB08, SP06c, XHSL09, YLLL08]. assisted [YY09]. associated [AMJD06b, Aou08a, CC08a, CK07a, CS05c, Ery06, GQS08, HY07b, Kos07, LTW07, LCS09, LN07, LP08b, Mar05b, Neh07a, QBH06a, RES07, Sad09, SS07b, Sof09, SSR05, SYX08, Sri09, SEŞ09, TNL09, lZXL09]. association [SM05c, Sjö07, TRLD08]. associative [JSL06b, Oht08, PK08c, PPKL08, TLC05, XCL09, YLL^{+}05]. ASSOR [YJ07]. assumption [LS06d]. assurance [Ali05]. astrophysical [DK06]. asymmetric [AH07c, EEEE06, Rey05b, SG09c, WS08d, YY06a]. asymptotes [WN08a]. Asymptotic [BK08c, BS07d, DXH08, DC06a, DB07d, mHzC05, mH07a, HHW09, Hua08b, HT09a, HTM09, HT09b, JW07, Kar09c, KWZ08, MeT05, Ram07b, RM08, Ser09, SR06, SL09a, TG08a, WZ06b, XZ08b, BR07a, BR08a, BR09a, Che05a, CHH09, DDR06b, FBB07, Geu07, Guo09, HK09, Has06d, mHsLzC06, KS08d, LZdC06, LFC08, Liu09a, MSHeT07, Par05a, Par06a, Pen06, RB08a, RB08b, SSA09, SL05b, SS05h, WGL08, Wan08f, WWYL08, Wu05d, YLZ06, Yan05d, ZZ07b, ZZ08b]. Asymptotical [GL06a, WZL08a]. Asymptotically [MT09, CLCK09, DL08, FK06a, GG09b, JK06, LjZH08, NS06b, NS07b, OS09, QCS09, SZ07a, Son08b, WW05b, XCL09, YC06b, YN07]. asynchronous [HWH09, YGL05, Yua05]. ATC [yLLjP06]. ATL [SL06a]. ATM [Par08, SS05e]. atmosphere [NSS06b]. atom [KWS05]. Attack [ZSNS05, HW05a]. attacks [Hsu05b, Shi05a]. attainable [IMB08]. attention [SY07b, YLL07]. Attenuation [RLdLSF07, SA06e]. attitude [TK06b, TK06c]. attracting [ABP05]. attraction [LDC09, SK09b]. attractive [AA08, XT06]. attractivity [AOS08, BR09a, CS06a, CW06a, CXC06, Che06e, JWS^{+}09, LC08a, Rös07, She07a, WZ08b, XZ07b, YSZG05, ZZ05a]. attractor [JY07b, JH09, Yan08d, ZZL08]. attractors [DTC08, DC06b, FLO07, HHYW07, LZ07b, LWW09]. attributes [Lin07f]. auction [LJL06]. Augmented [KPL09b, MM08b, CCWH09, DH06c, DZG06, HS07a, HW06, LK08, Lin06d, UP07a, WLL09, dCdSJM08, dSLV^{+}08]. authenticated [CYY05, CJL05b, Hsu05a, Hwa05, HL05c, HL05d, KRY05, LL05b, LKY05c, LKY05d, LWK05a, SW05b, SC05c, TLH05, YW05, YRY05c, ZAX05, ZW05b, ZL05b]. Authentication [CC05b, CY05b, Hsu05b, KJY05, LKY05a, LC05e, Sha05c, Shi05a, SL05a, SCS05b, SC05d, SCS05c, SZS05, TWL05, WLT05b, YWC05, YTWY05, YRY05a, YRY05d, YY05c, ZK05]. Author [Ano05d, Ano05h, Ano05e, Ano05f, Ano05a, Ano05g, Ano05b, Ano05c, Ano05i, Ano06a, Ano06b, Ano07a]. authorities [Chi05]. authority [CCH05]. autoadaptative [ZG08b]. autocatalytic [AS06d]. autocorrelated [IM08]. automata [Aki05a, Aki05b, Akı06, AVSB08, JY07a, JY07b, Rom09a, WdRS07, dRMS05]. Automated [WLLL07, GG05b]. Automatic [KB05c, OISM07, TTC08, Asa06b, Asa07b, BP09, Su07c, TCMPMC07, TCMP07, WF07d, PGdSS08]. automation [HJL08]. automaton [MB05a]. automorphic [dAC09]. autonomous [Che05a, Che06a, Che06h, LWW09, Meh06, MN05, Pod06, Rig08, TG08a, TQC09, XZ08b, YEM05, Yan08d, Zen09]. autoregressive [MAD08]. Auxiliary [KK07d, ARM^{+}05, DCL07, FqZ08, KC05a, Lay09a, LL06b, LXZ09, MLW09, RGA06, Wu06c, Zha07c, Zha07d]. av [KG06]. availability [AGT07, KC07b, RKS08, WC06a, WDK06, WC09b]. available [ARM^{+}05, RGA05, RGA06]. Average [Pe¸ 09, ADK07, CS07a, GK06a, MK06, Moh06, WML07, WXW09]. averaged [Son08b]. averages [ST07]. averaging [OS09, YKR06]. avian [EH08a]. AVK [BKA07]. Avrachenkov [KES09]. AX [KHZ07, XL09b, AV06]. AXB [LpL07, yP05]. axial [CYL^{+}08]. axially [CZ05a, CZ05b, CZZ05, ZC06a]. axis [LR07a, Ram06a, Rim06a, Rim06b, Rim06e, Rim07c, Rim07a, Rim07b]. axis- [Rim06a]. axis-I [LR07a]. axis-II [Rim07c]. Axisymmetric [ZO05, ASH09, Kim05b]. Ayala [BW05a, Che06h, LZC07a]. Aziz [Pan08b].

b [AV06, AÖ06b, Aks06a, ASA07a, BB06c, CCE06, CC06a, CC08b, DSB05, DIS05, FFKV08, GC06a, GK05b, JS07a, KA05a, KA05b, KK07b, KY08, KYK08, KK08a, KNM05, KHZ07, Lin05c, LA07a, ML06, NR07a, NRA07, NHKZ06, ÖEK05, PLZ05, RK07b, Ras05a, SJ08b, SD09, Sha06a, VFF08a, XGTS07, XL09b, ZW09c]. B-Spline [KY08, AÖ06b, Aks06a, CCE06, CC06a, CC08b, DSB05, GK05b, JS07a, KA05a, KA05b, KK07b, KYK08, KK08a, LA07a, ML06, NHKZ06, ÖEK05, PLZ05, RK07b, Ras05a, SD09, ZW09c]. B-splines [ASA07a, BB06c, DIS05, GC06a, KNM05]. B.V.P. [Att06d]. Babajee [Ren08a]. back [LWL07a, SL07b, SS05e, jW06a, ZZLL07]. back-order [LWL07a]. back-propagation [ZZLL07]. backbone [SCR08a]. background [WHD^{+}08, ZL08c]. backhauls [TMSZ06]. backlogging [JTMRS06]. Backlund [AB05a, Sha07a]. backorder [Lee05a, LWH06a]. Backpropagation [HT06b]. backstepping [CCZ08]. backtracking [WZ09c, Zhu05b]. Backward [Li09a, DSMK08, IOS07, Kar07a, Kıl08a, KT07c, gLxW05, OISM07, QFS07, SZ05, TT09, Tua13, XFQ06]. bacteria [HES05, HAM04, HMSZ07]. bacterial [GCSS05]. bacteriocin [ABB06e]. Bad [HJL08]. Bagley [RB05a]. balance [CGJ08]. balanced [Che07d, SH06b, XHX07, YH09]. Balancing [KOP07a]. Balazs [Gup08a]. balking [Oma06]. ball [LS09d, LS09h, PL06a, RW07, RW09, Ste08, SW09c, SiU09, Ste09e, Ste09h, WR06, WRB07, ZTWL09]. Ballistic [HC07b]. BAM [HLT09, Li09c, SSA09, YZW07a, ZS07b, ZLL06b]. Banach [AOS09, ASK05, sCLC09, CKQ08, DF09, GS09, ILZ09, JK08, KB05a, KKKM09, Kum08, LL07a, LYH09, yMT08, NS07a, Raf06, Son07, SLW08, WCO08, WWL08a, WWL08b, WCO09, Wan08g, WL05f, WYL06b, WYL06c, WWYL08, WZ06h, WZ06i, WZ06j, XZ09b, XF08b, ZSM07, Zha08e, ZLW08, bZlDL09, ZW08c, ZWC06, ZH08d, ZH09b]. Banach-valued [AOS09]. Banachiewicz [Den09a]. band [LXSH06, XQ06b, YXSX05]. banded [KLR07]. bandit [KX08]. bank [WYL06d]. bankruptcy [Nwo06a, Nwo07a, zWxY07]. bankruptcy/recovery [Nwo07a]. banks [GZL08]. bar [BL06, KTÖA08, WC05b]. Barcode [Che08c]. Barnes [Eli07]. barrel [LYCW09]. barrier [AZ07, Cho05, GY08, ND06, SYL07a, dCdSJM08]. barriers [ÇFÇ06a, PSRN06, PSS07]. base [AVKA09, AS07b, Ali05, dCRDSSN^{+}08, WLCL06]. base- [AS07b]. Based [Ohm08, AV09, AÖÖ06, AS06c, AS07b, AKB09b, ACJ09, AP06, AP09, AK05d, AJK05, AS06f, AKR06, AK06h, AHIP07, BDDB08, BK08a, BBMN07, BP09, Baw05, BKFM05, Boy07b, BKS^{+}06, Cai09, CL05a, CLS05b, Cha05g, CT05b, CY05b, CLH05, Che05g, CDP06, Che06l, CL08c, CZ08a, Che09a, lCS06, Chi05, CK06a, CJL05b, CYH^{+}07, CH07f, CHY^{+}09, CFVZ06, DY06, Dam08, DN07a, DN07b, DD08a, DvdHD08, DCL07, DDZ08, fDxZfWyY09, DWZ07, DjT05a, DjT05b, DWT05a, Dua06, EFE08, EARS07, EH08b, FWL08, FTG07, FAQ08, Fuk06a, GPR^{+}09, GY06, GAV08, GFH08, GANG07, GJ07g, GJ07h, GW07, Gui05, GT05a, GT05b, mHsLzC06, HW07a, HCLPJL^{+}08, HIPA08, Hor07, HL06b, HHC08, HC08b, HYL09, Hsu05a, HZYL08, HC08c, HLS05, HH05, HL05d, ID08, Ide06, IZIO09, IM08, JY07a, JY07b, JWL^{+}07, JZCW05, JY06, JCX^{+}07]. based [JB08, JZHW08, JL09, KA06a, KS08a, Kad07a, KF08a, KF08b, KK07c, KK05a, Kat09, KMN09, KS05a, Kho07, KSC^{+}08, KCW07, KPG07, KLS09, KM05e, KHKL05, LKY05a, LKY05b, LHY05, Lee06a, LM08a, LS08c, Li05a, LC05a, LTH05, LXL06, LSW06b, LZL^{+}07, LLX07a, LLLX07, LD07b, LML08, LT09, Li09e, LL06b, Lin05a, LWZH05, Lin05b, Lin06b, Lin07d, Lin07e, LG07, yLLjP06, LHC07, LYGL07b, LY09c, LLC09b, LCC05c, LM07a, LSN09, LCZ05d, LCZ05b, LLC06a, LLC06b, LXZ07, LMD^{+}08, LHWW07, cMpZfR06, Ma06b, MW07, MW09, MAR05a, MUZG09, MJFAS07, MPEL09, Meh06, MM05c, MY07b, MZ07, MA05f, MG06, Mit06, MSM08, NCÁHCLP05b, NHCLP07, NK05b, PW05, PLTW07, PVN06b, PSBM08, PG07a, PG07b, PG07d, PG08c, PLJ05a, PLJ05b, QCB05a, Qiu05, QC05, Rao06a, RJM07a, RZX07, RGA05, RGA06, RV08, Sae05, Sal06a, SM06a, SB06b, SSW09, Sev07, Sha08d, SC05a, Sha05c, Sha05d, SCS05a, She06]. based [SY07b, SP06a, SW05b, Sin07b, Sin07c, uITA06, uIK06, Sog08d, SZ07d, SZ07e, SCL05b, SLC05, SD06, SCS05c, SCS06, SW06c, SC07e, Sun07a, TD07b, TKJ08, TTG07, Tsa05, WTW05, WYX05, WZT06, Wan06a, WHH07, WF07a, WXW09, WCX08, Wel08, WLHH05, WL05g, WS07g, WW05c, WY05d, WWT05, WLC05, WS06b, WLC07b, Wu08d, WLL09, WWC09, XYZ09, XCW08, XX07, XBZ06, XZB06, XLKH09, XHSL09, XTCC08, YZY08, YY06a, YWZW07, Yin07, YY07c, Yin09, YRY05d, You08, YY06c, YK06a, ZSNS05, ZCZ05, ZD06a, ZHLG07, ZHZ07a, ZZDL07, ZSCW09, ZHX06, ZGH07, ZXL09, ZCL05, ZZYT08, ZCZX07, ZWC09, dRMS05, lS07b]. bases [IL08, Kar09b, Lin05c, ML06, MY06b, PMRAS05b, ZWT09a, ZTWL09]. basic [AM08b, BK05a, FBB07]. basins [LDC09, SK09b]. basis [AE07a, Boy07b, BW09, CMM07, DZ08e, DjT06, Dua06, FWJK09, GS06b, GS06c, GS07a, Hu07a, HZ05, KWS05, LKC05, MW09, MMA07, MJ07d, Pan08a, RLY06, RAN07, RAN08a, Yoo05, Zha06a, Zha06b, Zha07g, ZWT09b, ZWT09a]. Baskakov [GAG09, WG08]. Basto [Pet07]. batch [BLH06, Cho07c, KL06b, KHL06, LHH05a, WFX07]. batches [GMS06]. bath [SH08a]. bathtub [WWT05]. bathtub-shaped [WWT05]. Bayes [CL06b]. Bayesian [Dan05, JZW08, NK07b, WZ06g]. Bazilevic [KS06a, Kim09]. BBM [AKO06, Omr06, Waz05f, Waz05h]. BBS [HZYL08]. BC [HT05]. BDF [AHIP07, HIPA08, Psi07, VAMV07]. be [Bou06c, Boy07b, CS07b, Sok07a, Via08]. beach [KTÖA08]. beam [Gal08a, Gal08b, JSL05, LeT05, Li07e, Lin05c, MRN09, XFTY09, Yao08b]. beams [BTMB09, WCH08b]. bearing [YB05]. bearings [BK06b, BK06a]. beat [GPR^{+}09]. beaver [BKS^{+}06]. bed [DSMK08, QS09]. beddington [WZ06b, DK05, LH07a, LC09b, XW07b]. beddington-type [WZ06b]. Bee [KA09a, FAM07]. bees [KBTS07]. behaved [XSX08]. Behavior [OMST07, AB06, AAA08, BS07d, CD05, CLH06, DXH08, DDR05a, DDR05b, DDR06b, HZYL08, HT09a, HTM09, HT09b, JZ08a, JW07, JSLL07, KK08c, gLXyZ06, MeT05, MSHeT07, RK08, SL09a, SS05h, WZ06b, WGL08, XCL09, ZLT09, ZM07, ZSZS09, lS08b]. behavioral [TCDV07]. behaviors [WL07e, Zha09f]. behaviour [CCS05, GY07a, Kar09c, Qui09, Sey05]. being [PR06]. Bell [WN08c]. Bellman [ZZ07e]. belonging [Lal09]. below [MR09, SDS05, SSS06, SM08d]. BEM [DS06a, GK09]. Bending [Shu05, BTMB09, KNJ05, Li07e, Lin05c, ZL07a]. Bending-torsion [Shu05]. benefit [AEFFA08, WC06a]. Benjamin [AKO06, AKMA05, Omr06, ZXLS06, ZX07b]. Berge [AK05a, AK06a]. Bergman [Ste09e, Ste09h]. Berlekamp [IL08]. Bernoulli [AF07, AM05, CAS08, CAS09, Cho07c, CD09, JR07, LeT05, PR07, Wen05, WLY06, XM08, ZFL^{+}09]. Bernstein [BP05, BB06b, Bha08, Boy07b, Boy08b, CMGMD06, DNS08, GLMLP09, MB07, Ost08, RLY06, RAN07, RAN08a, Yu08a]. Bernstein-type [CMGMD06]. Bertrand [SK08b]. Besicovitch [LS08e]. Besov [AS09c]. Bessel [AES08, FBB07, KS08d, LLG06, NK07a, SY07a, SY09, SdHT09, WLT07a]. Best [Cai09, SA08c, Sim09a, Cai06b, CGSS09, GANG07, GIT07, HN06a, HMJD06, KMB05, XSX08, OM08, SLZ05a, SZL^{+}07]. beta [FB08, GN06b]. better [AKR08, HT06a, Kar09b]. between [Ahu08, AR08, AS09b, BS09a, BN07, BB09, CC08d, ESES06, ESD07b, Fuk06b, Gor06b, HBN06, IRSR05, JyGcL^{+}09, KAS06a, KAP06, KLX06, LS08e, Lia05, LYH09, MNM07, PVS^{+}08, RAG^{+}07, REN^{+}05, SANN07, ST06c, Sla06, SZXL06, SW09c, SiU09, Ste09g, TS05, VFF08b, WL05a, WDK06, WC09b, Waz06d, Yan09c, ZYH08]. Beverton [DAQ09, lSAQ08]. beyond [Myc06]. Bézier [Gup08a, GI08, Lin09b, Rab05, RLY06, ZWT09a, ZTWL09]. Bezout [ADTGV09, TLB07]. BFGS [KLS09, XW07a]. Bi [ASA06a, BS07c, DM07a, Ema06b, xGyZtZ^{+}07, JRR^{+}07, JRA^{+}07, KH05a, LD06a, MJ07a, MFF09, RSA07, RAS08, XXTL09, MMC06, Vog07]. bi-3* [KH05a]. Bi-CG [MMC06]. Bi-CGSTAB [Vog07]. bi-conjugate [xGyZtZ^{+}07]. bi-criteria [JRA^{+}07]. bi-directionally [XXTL09]. bi-harmonic [BS07c]. bi-level [DM07a, Ema06b, RSA07, RAS08]. bi-material [LD06a]. Bi-objective [ASA06a, JRR^{+}07, MFF09]. bi-obstacle [MJ07a]. bi-variable [Yan08a]. biasing [AÖ05b]. BIBO [LmZ08b, LmZ08c, LZ09b]. BIC [Fuk06a]. BIC-based [Fuk06a]. BiCG [Vog07]. Bicriteria [Aza06, EG06b, EG06c, XY11]. bidding [LJL06]. bidiagonal [Yam08]. bidiagonalization [SB06b]. bidirectional [JSL06b, KSC^{+}08, PK08c, PPKL08, XCL09, YLL^{+}05]. bidomain [NRLT07]. Bieberbach [HTM09]. Bifurcation [FLLH07b, GZC08, lH08a, JY09, LW09a, LM09, Lou07, MB08, SXJ05, SMG07, hYG05, ZLC06, ZZW09a, ZYC09, ZL07b, ZLZ08, AGR09, Chr08, DM05, Dil05, GCT08, HRCL08, HH09b, LL08a, LLG09, LX09a, LW09c, MD08a, SP06b, SHG07, TMS08, YL06a, Yan07b, Yan09d, ZXF07, ZLY08, ZW06a, ZSS08b, ZH07b, ZHZ07b]. Bifurcations [FHL07, LSL06, MV09a, SLX05a, TL06, TH07a, TH08, ZZH06, ZXLS06, ZWX08, ZXZZ07, ANK05, FX09, FF06, MG09, QBH06b, SXL09b, WFL06, WCZ09]. biharmonic [AMK06, AE07a, CPW09, DM06, Fad06, KSM09, LL05a, LAL07, Ony05b]. biharmonical [gLXyZ06]. bilaterally [Lee06b]. Bilevel [SP06c, AHS06, JS06, LHWW07, MD06c, SZL05, SLZ05b, SLZ05a, SLZZ06, SZL^{+}07, WWWL07]. bilinear [Far05, LC05a, LC05e, Men08, Par08, Waz07e, Waz08c, jZzFbL06]. billing [OHJ07]. Bin [SS09b]. Bin-Packing [SS09b]. binarization [CT09b]. binary [ÁCTZ05, CL07a, Cha07a, DN07b, FZN07, LLX07b, MZL07, Pak11, Sal08d, SCS06, Tan08a, WLC07a, YGL05, YLM07, ZS07d]. binding [dSYGM^{+}05, TSS09]. binomial [hB07, GM05b, MTM05, SYL07a, SYL07b, WWC06]. bio [AS06d, FJ06, HHL08]. bio-chemical [AS06d, FJ06]. bio-inspired [HHL08]. biochemical [ZYC09]. bioconversion [WFX07]. bioheat [ZZKY05]. Biological [EFE08, YLLL08, JC06, JCL08, MR05a, RBvB08, dCRDSSN^{+}08, VGMV^{+}07]. Biology [GPS09, AE06b, BSN^{+}08, EH08a]. Biomagnetic [MS09b]. biomass [MBM07, NSS06a]. Biomathematical [d’O05a]. biomedical [MUZG09]. Biomedicine [GPS09]. Biorthogonal [Yan06a, AM08c, hCxC05, CWZ07, JV07]. biorthogonalization [WSB06]. biosensor [DRS^{+}08, TSS09]. bipancyclicity [XHZ06]. bipartite [KH05b]. Birkhoff [NBHYV09]. birth [JY09]. Bisection [Wu05d, XXX06]. bispherical [DT08]. bit [JY07b]. Bivariate [CMV05, DZT06, GN06c, GN07, KFM08, Nad05, Nad06b, NK07a]. BKM [ZT05]. BKP [MWQ09]. black [Dön06a, CKK05, CJSSP05, Wu07a]. blackbox [Zho07]. BLAGE [Moh07b]. Blasius [Has06b, Bat08c, Bat08b, Cor05c, Faz09, Lee06b, Wan04, Waz07k, YC09a]. blind [CLS05b, CHYC05, FTMA07, LHY05, LCZ05c, Qiu05, WHH05, WW05c]. Bloch [LS08d, LS09f, LS09e, LS09g, Ste08, Ste09a, Ste09f, Ste09c, Ste09d, Yan09c, Zhu09b]. Bloch-type [LS09f, LS09e, LS09g, Ste08, Ste09a, Ste09f, Yan09c]. Block [AAAA06, LB06, ZW07c, ÁFVZ06, AKVH05, BK09, BZZ08b, BLZZ09, CGDR06, CL06d, ES05b, bGgLgW09, HS07b, IOS07, KT06, Kat05, KT07b, KLR07, LCH09, LXZ07, MS07a, MQ07, MSCG09, NS06a, OISM07, PW08b, RH06, SNH07a, ST05a, Tia09, Wan07f, WL09c, YHTR05, Yan07e, YGY09, KV05a, MSK05, YJ07]. block-by-block [SNH07a]. block-pulse [AKVH05, Wan07f, WL09c]. block-tridiagonal [CL06d, LXZ07, MQ07]. blockability [Ish09, KI08]. blocked [AA07d]. Blocking [ACDR09]. blocks [TAK07b]. blood [BA07, DRS^{+}08, GM05a, IAMA08, NMS07, SH07c, YSXL08]. bloom [SMBB07, SMC08]. Blow [CT05a, CG07b, PY08a, Vul07, WM07, CC06b, DX06, GR06a, KV05b, LCM08, LZZ06b, LZ07c, SS08d, WY08, YXW07, ZY08b, ZLW09a, ZQ06, ZQ08, ZL07c]. Blow-up [CT05a, PY08a, WM07, CC06b, DX06, GR06a, KV05b, LCM08, LZZ06b, LZ07c, SS08d, WY08, YXW07, ZY08b, ZLW09a, ZQ06, ZQ08, ZL07c]. blowflies [Lin09a, Sak05]. Blowup [WMX07, HT09a, HT09b]. BLUE [WLC06]. BMAP [BGP06]. BMOA [AS09c]. board [Ano05j, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p]. Boas [CL07b]. bodies [Att06b, GKK09]. body [BS09a, CQL08, GPT09, KSKK06, Kös06a, MCMA07, MB05b, SA06e, Smi06a, WsZ07a]. Bogoyavlenskii [Waz08j]. Boltzmann [CLSZ06, CLZ^{+}07, Che09a, DLJ08, WYQ05, YS06d, ZYD09]. Bolza [BGRS07]. Bona [AKO06, AKMA05, Omr06, ZXLS06, ZX07b]. bond [SS06e]. bonded [Li05f, Li05g, SLX05a]. bonds [FV05, GL06b]. Bonhoeffer [ZXF07]. Boolean [GM06b, MG06, NO09]. boosted [HJB06]. Bootstrap [CK06b]. border [DM05]. borrowing [AS06c, WZ06k]. both [DDD05, LJL06, Sae06d, Sae06e, SYL07b, Wu05d]. bound [CKL06, FYZ07, wG07a, HC05a, JZ08b, JLA07, LHL06, LZZ06b, LH07b, LC06d, MYSA05, MAM06, MD05, MZL07, Pak11, SLZZ06, TMMVA06, XW07a, YMCL06, YPZ05, YCYL05, ZSHB06]. bound-and-stopped [wG07a]. bound-constrained [JZ08b]. boundaries [Deh05a, Li05f, Li05g, Udw05b, Wu07a]. Boundary [CMC06, Dyk09, GR06a, KB09, Udw05b, WsZ07a, WC07a, ASL07, Afr05, AMN07d, AN05a, AOLL05, AOS09, AS06a, AS06b, AS05, ASN07, AAK07, And06, Ang05, AO05d, ADK05, ADK06a, ADK06b, AA06e, Att05b, AL06, AK05i, BS05a, BZ09, Bat08b, Baw05, Bay05a, BŞ07a, BUÇ07, BLLM08, BBB08, BS07c, BR09c, BD09b, BAM07, BT08, CCE06, CC06a, CC07a, CA05, ÇFÇ06b, ÇFÇ06a, Cen06, CN05, CN06, CC06b, Cha05h, CTG06, Che09b, CSC08, CMS05, CZ07c, CG08b, CC08d, DM06, DMH05, DXH07b, DB07d, DZ07a, Du07, DZ08f, DLT08, DHG09, EPER07, EG07a, EG07b, Elb05, EESF04, EGX05, EM07, Fad05, FSG07, Fan06, Fil05, Fra09a, GB07, GC07b, GL08a, Gen09a, Gen09b, GJ07c, Gül05, GHX08, HK09, HM08a, mHzC05, HH08a, HJ08, HAU05, Hsu06, HZS07, HLW08, HT09a, HT09b, Hus09]. boundary [IMS09, JDG06a, Jah07, Jan08c, Jan08b, Jan09b, Jan09c, JS07a, JGY07, JG07b, JGZ07, JFG08b, JZ08a, JZ06a, JZ08c, JC09, JLW09, JV05, KA05a, KA05b, KK07b, KK07a, KY08, KK08b, KS08a, KW08a, KW08b, KWX08, KR05, KB06a, Kan07, Kar09a, KE05a, KKA06, Kha07a, Kha07b, Kha08a, Kha08b, KV05b, KB05c, Küç05a, KS09a, Küç05b, Kum03, KA06e, Kum07, Kum09c, LeT05, LJ08, LS06b, LS06c, Li08a, Li08d, LCJ08, Li08b, LL08e, LXJ09, LS07c, LPS08, LZ08e, LDL08, Liu05a, Liu06h, LZW07, LLW07, LLW08, LLW09, LA07b, LA07a, LST09, Ma05c, Ma07b, Mai05, MM06b, MD09a, Meš07, MJ05, MK06, MA06e, MS06b, MA06f, MD09b, MMN06b, MAO06, MN07, Mor08a, MD08c, ND07, NMD07, OM06, Odi09, Ola06, OB09, OIL09, PDW06, PFG08, PB09, Pan06, Pan08c, Pao07, PK08a, PC08]. boundary [PR06, QM05, Qua07, RAN08a, RLZ07, RLZ08, Ram05i, Rao06b, RK07b, RG05, RMJG07, RJM07a, RMG07b, RMJ07, RMG07a, RGM07, RJF07, RSZ08, SF07, SH07b, SS07a, SS09c, SH06a, SR06, SM06c, SL07b, SA06f, SA06g, SAM07, SA07a, SAN07, SA07b, SA07d, SA07c, SA07e, SAE08, SA08e, uIKTT05, uITA06, uIK06, uIT06, uINTK06, uITiHK08, SFH09, SWX06a, SWZL07, Su08c, SL06c, SG07c, SQG08, SMG08, SCO09, SA05b, SA06h, TWZ07, TS09, TJG08, TiHuI08, Vul07, WJ05, WMX07, WZS07, WS07d, WC07b, WLK08, WTW08, WCH08a, WLW08, WH08b, WLW09, WSL09, Waz06h, WP05, Wei06a, Wei06b, WZ08e, Wey06, Wey08, WL08b, fW07b, WY08, XCM07, XL08b, Xu07, Xu08a, XZ09b, YC06c, YXW07, YS07, Yan08b, YW08, YWL09, YZ09a, YZ09b, YC07a, YC09b, Yi05, YC08b, YJO08, ZLCZ06, ZW07b, ZL07a]. boundary [Zha08f, Zha08e, ZS08a, ZYW08, ZS09a, Zha07h, ZC08b, Zha09h, Zha09g, Zhe08, ZL08c, ZM07, ZLC07, ZHZ07b, ZJO08, LWG09]. boundary-layer [And06, Elb05]. boundary-value [CG08b, EG07a, HJ08, KKA06, Kum07, ND07, Ram05i, RMJG07, RJM07a, RGM07, RJF07, RSZ08, uIKTT05, uITA06, uIK06, uIT06, uINTK06, uITiHK08, TiHuI08, WLK08]. Bounded [DJJL06, HRLC08, Att07, AS09c, AH05c, CSV07, Den08a, DZ08d, Djo06, Fro06, GR06a, GZ07d, GBG08, Jah07, Kar09c, Lee06b, LLZ06, LKU09, LYH09, NM08b, Sok08d, SS07d, VL09, WZ13, YY06b, YY07b, Zha07e, Zhe08]. Bounded-rational-prisoner [DJJL06]. Boundedness [CY08d, FO09, JSL06a, AO08, ABM09, FeT05, Sta06a]. Bounding [NG09, Pal09, PLTW07]. Bounds [Cui05, Jia07a, ARS07a, BDE06, DL09, DC06c, Fro06, GBC09, Hua05b, Liu06c, MR05a, SG08c, Sim09a, UE07, VCRCG09, WCX05, WW09, Wu08a, Yam08, YZLD08, YHL08a, YHL09b, Zhu06a, ZL07c]. Boussinesq [Waz08b, Che05e, CL08b, Dem05d, FHL07, HAK07a, Inc07c, Kay05c, LSL06, MGG^{+}09, Sha06a, WM07, WSKZ07, WX08a, Waz05i, Waz06c, Waz07d, Waz08o]. Boussinesq-like [Sha06a]. Boussinesq-type [FHL07]. Boussinseq [FLLH07a]. box [HMW07, JmBdXgXm05, LT08a, MK05a, WSZ07c, Wan09a, YHL08b, YHL09a, Zhu05e, LS08e]. box-constrained [LT08a]. BP [YWZW07, YZD08]. brain [IZIO09, SSTD05b, SLX^{+}05b, XCW08]. branch [CKL06, Fen06, GXS05, HC05a, LC06d, MD05, MZL07, Pak11, SJ06c, SLZZ06, SG05d, TMMVA06, YPZ05, YC09a]. branch-and-bound [CKL06, TMMVA06, YPZ05]. branch-and-prune [SG05d]. branch-and-pruning [SJ06c]. branch-and-reduce [GXS05]. branched [IIS05]. branches [WYL06d]. Bratu [MSM08, SH06c, Waz05a]. Bratu-type [Waz05a]. break [Fuk06b]. breakdowns [GI06, WWP05]. breaking [XZ05, lYpZqL08, Zha07d, Zha08d]. breaks [Jön05]. breast [LT07]. breath [DRS^{+}08]. Bretherton [Rom09b]. Brinkman [MSH08]. Broadcast [HRM07]. broadcasting [She07c]. Broer [ESK05, JZ06c, SXJ05, ZS09b, LH08d, SZ06b, WSLZ07, WZL08b, ZX06a, ZX07a, ZWT09c, ZLZ08]. brokerage [Nwo07b]. Brouwer [Tan06a]. Brownian [GY08]. Broyden [GKK09, JY06, MCW08, YJ07, ZG08b]. Broyden-like [MCW08]. BSP [CPTZ05]. Buck [CL07b]. Bucy [RLdLSF07]. budget [Lin07e]. budworm [BKP06]. buffers [AKH06, HYL09]. build [dCD09]. building [TKJ08]. bulk [GMS06]. bulk-service [GMS06]. Bulkley [VSB05]. bullwhip [MM07d]. bundle [KMA08]. Bünyamin [Sub07]. Burger [EHAR05, HAA07, Jav06a, Jav06b, Jav06d, KS08b, SM05c]. Burgers’ [DLJ08, KAF08, KS09a, Lia08, PVV09, SG09a, WLH09, You08, ZW09c, AD05a, AD05b, Aks05, AÖÖ06, Aks06a, AKMA05, BS05a, Bha08, CW06b, CA08a, CIS09, CG07c, DSB05, DIS05, DJ06b, DHS07, Dem05c, Den08d, Gor06b, GÖ05, Gül06, HSH05, HXH08, Inc05, Jer08, KSA05, KA06a, KY05a, KS09b, Kud09, ÖEK05, ÖA05, zPY06, She08, SW09d, TSD07, VHFF08, Wan06f, WZ07c, WSZ07b, Waz05p, Waz05q, Waz07c, Waz08a, Waz08g, Waz08h, Waz08i, ZLYL08]. Burmeister [RZX07]. Burr [WY05c]. bursting [WS09a]. business [MG09]. businesses [HWC07]. busy [BKS^{+}06]. BVP [ANM06, DH09, Wan07e]. BVPs [And06, Bie08a, Bie08b, CD08, KYK08, Liu08b]. bypassing [She06]. BYY [LM08b].

c [NKE05, HC05a, LpL07, Moh09, yP05, hPyHZ06, PHZ06b, AA07d, AC06]. C-SPLAGE [Moh09]. CA [WTG07]. cache [ACDR09]. CAD [XL06b]. Cahn [BN09, DN06, TB09b, YC05a, YLH09]. calculate [PRA^{+}09, Wan06b]. calculating [Cas07a, CC08c, CC09c, ESK07, Pou06, TGSNB07, WLK08, ZCW05]. Calculation [GM05b, JSL05, MTM05, yZsZ06, BH06, BB06c, Çel05a, Hub08a, KM06, KB05c, LZ09e, SWC05, WGCH05, YW09]. calculations [Dem05b, NK07b]. calculus [CGW06, Fel05, KZ07, KSKK06, Kös06a, LS08e, LTW07, MBS07, Nis07, Rad09, SOT07, SOS07, ST09, WLT07a, WLS08]. Calheiros [Pet07]. calibration [Ewa06, WWXX07]. call [CK06d]. callable [FV05]. Calogero [Waz08j]. Camassa [HRCL08, TXW09, Waz05d, Waz05e, Waz05j, Waz06l, Waz07f]. campaign [Çet06]. Can [Cha07a, Bou06c, CS07b, NR09, SMBB07]. canal [AS06h]. Canard [BKP06]. cancellation [ZC06a]. cancelled [CL07a]. cancer [dPCdBYV08, Çet07, LT07]. candidates [HN06a]. Canonical [ESA05, LXJL09, AM08a]. cantilever [XFTY09, Yao08b]. cantilever-beam [XFTY09]. cap [Che07a]. capabilities [LT09]. capacitance [Fan06]. capacitated [TMSG06]. Capacity [CHY^{+}09, AEFFA08, BDPP08, Chi07c, Chi07i, DAQ09, Lin06c, Lin07d, SP06d, Wan05f, Yu07]. capillaries [LLLW06]. capillary [BMD08, EP07, PE07a, SM07b]. capillary-tissue [SM07b]. capital [Cai06a, DM07a, ZC05a]. capturing [KK08d, LWL07b]. Caputo [Odi06a]. car [CW08]. car-following [CW08]. Carbon [WLX05, WL05g, WWLX06]. carcinoma [YSXL08]. card [She05c, SS05e]. cardinal [BW09, LFW08, ML06, Sae06d, Sae06e, cWxChY07]. cardiovascular [NMS07]. cardiovascular/respiratory [NMS07]. cards [Hsu05b, LKY05a]. care [EMH07, RS08b]. Carlitz [CAS09]. Carlo [CWY06, EFE08, FE07, FE08, JLRB05, JLBR07, JLBR08, PQJ08, RAG^{+}07, VV06, Vaj07d, WMC08, ZZX05]. Carlson [NMM09]. Carreau [AH07c]. Carrier [TNL09]. carrying [BDPP08, DAQ09]. CAS [DX07a, YB06]. cascade [CAL08]. case [AKH06, AVSB08, AH06d, AM08b, AM08d, CS07a, Dem05d, HRE05, HCS05, Isk07, KAF08, KC07c, LY09c, SML05b, SWC05, SAM07, Sta06c, TA05b, WC05d, WY08, YSCC05, ZS07d]. case-based [LY09c]. cases [AM08c, BIK06, DAQ09, GGSN06, VHFF08]. cash [CJC05, NN06a]. Casson [SM07b, SVMG09]. Catalan [SSMI09]. catalytic [Bay05b]. catch [MBM07]. categorical [GG05a]. categorizing [CS07c]. cathode [Vyn07]. Cauchy [AM09a, CFF09, GC07c, GC09, KPS06b, KA06d, Kou07a, Kou07d, Kum08, LC07, MA06c, Mar05b, Ost08, QW09, QW08a, WX09, Wu08a, XL06a, XF06, Zha05e, Zha05d, Zha06c, ZL09b, dAC09]. Caudrey [Waz08k, XZY08]. caused [SMC08].

Cavalcante [Hub09a]. cavitations [Ouj06]. cavities [GD06]. cavity [Dem05a, Gür05, XZB06]. Cayley [ZW05g]. CCA [SC07e]. CD4 [WS07e]. CDF [GL08b]. CDG [Waz08k, XZY08, Waz08q]. CDT [yN06]. cell [Ahm06, BMD08, LBZY08, MJFAS07, SANN07, Sal07b, TMASA05, Vyn07, dG06]. cells [PE07a, WS07e, YSXL08]. Cellular [QSS08, Aki05a, Aki05b, Akı06, AS06c, AVSB08, CS05a, EP07, JY07a, JY07b, KP08a, LH08c, MJFAS07, MB05a, Par06a, Par06b, Par06c, Par07, TMJJS07, WL07b, WdRS07, XCL09, ZWX05, ZH08b, dRMS05]. cement [Liu06a]. CEMOP [OASM06]. censored [Bak05, HWC07, WWT05, WLC05, WY05c, WLC06, WLC07b]. censoring [WWC06]. cent [CS06b]. Center [QBH06a, Sil07c, LM09, QBH06b, Sil07d]. centers [HZ05, SG05d, ZL08e]. Central [Tou09, XFL06, XF06, CFX05, LZR07, Sil07f, Tan05a, Wan09a, XCFX05]. central-upwind [CFX05, XCFX05]. centrohermitian [LCC05b, LTT06]. centrosymmetric [CD08, LHZ06, LHZ09, LCC05b, LCC06, TG07]. centrum [PL05, PC07a]. Certain [Aou08b, CS05c, Chu07a, CN09a, Den09b, GZ07a, LP08b, MA09, NO07, Aki05b, Alt07b, Aou07, Aou08a, AA05, Att06c, BKZ06, CPW09, CK07a, EA08a, EDS06, Fra07, GAPT08, GS08c, GI08, GF09, GG08b, HL08d, KAPM07, KS06a, KP08c, LTW07, LCS09, Liu07c, LN07, Liu09b, MAN06, MS08b, MX07b, Mor05, Mus07, Noo07a, Noo08a, Par05a, PC07b, RB07, ST06b, SK08a, SOT07, Sok07a, Sok07b, SFG07, SEŞ09, TS08c, Van09, WS08d, YHL08c, ZM08b]. certified [BCL05, CCH05, HW05b, LWK05a, Sha05b, TLH05, Tsa05]. Cesàro [LS09d, ABM09, AS09b, SC07b, Wan08g]. Cesàro-type [LS09d]. CFD [KK05c]. CG [FHSS08, MMC06, Sal06b]. CG-type [Sal06b]. CGM [KK05c]. CGSTAB [Vog07]. CH [Den08c]. CH- [Den08c]. chain [Che07c, Chi07a, ESAH10, HL05b, LL05f, LW06a, LZC07b, LGw07, Luo07, LG08, MPS08, Mel08, MQW05, MB08, PMS09, RSA07, SA08b, She07a, She07b, SS06g, WLCL06, XCH05a, XCH05b, XW07b, Zen09, ZGL08]. chains [Att06b, BSP05, Çet07, FLO07, ZC06b]. chamber [AS06i, LJ08]. Chan [EDS06]. chance [Hua06]. chance-constrained [Hua06]. Chandrasekhar [Nak07]. Chandrasekhar-type [Nak07]. Chang [Sun07a]. change [Liu06a, Pal08, Yu07]. changes [CGS08, ES05a, VG07, ZYW08]. changing [ANM08, AOLL05, JFG08b, LOA05, MM08a, OMST07, SLS08, XL08b, YY05d]. channel [AH07c, AH05c, BAK07, EEEE06, HAA07, IIS05, LY05, LCZ05c, Mak09, Oma06, Ram07b, Rao05, Rao06b, SG09c, VSB05, YTWY05]. channels [ZY09]. Chaos [AB05c, GL05b, HH08b, LXL06, UBKC08, YY08b, Yil06, DD06a, GZ07b, GLLC08, HLY09, RSD07, TH07b, Upa09, XC06a, YLC09, YEN08, YY09, ZXF07]. Chaotic [DD07, WS09a, Che07d, CCZ08, DTC08, GW07, HHYW07, HLY09, gLXyZ06, hL08c, LYGL07b, PKL08b, PJWL08, PNL05, PCCHLG08, SW09b, SH07e, TH07b, XLwW07, XXTL09, YY08b]. character [DD06b]. Characteristic [DZ08c, BL06, EH09c, Fra09b, HE08b, JLZ05a, LY09d, SWC05, SW09a, Su07b, ZZL05]. characteristics [BE05, DMS06a, ESES06, Kim06a, qQcM05, Ram06b]. Characterization [AK05h, CG07a, ETJ06, Ahm05b, AJRLRGOG09, BE06, Gao07, KI08, LH06b, Ma05b, MS08b, SDM^{+}09, VR08]. characterizations [XW08]. Characterizing [AS06f]. charged [OHJ07, TS08b]. charges [Li06c]. Charpit [Gor07]. chart [IM08]. charts [ANA07, HIST06]. Chaum [WHH05].

Chebychev [LS09b]. Chebyshev

[Pan12, AMJD08, AA07a, AD06b, AHR08, BFR07, BF07a, BF07b, Baz08, Boy06a, BV06, Boy07b, Boy08a, BX09, Çel05b, ÇG05a, Çel06b, CL05d, CM08b, Chu07a, Chu07j, Chu07l, DMJE05b, EF05, Elb05, EEEE05, EO06, EH08b, EDMJ05c, EDMJ05a, FFRF07, GJ07g, GSDB09, HEDMJ06, HMJED06, Hos06b, JYX06, KA05b, KV05a, KT07a, KLW06d, Kou07c, KL07d, KL07e, KL07f, KW07, KL07g, MSR07, Mao09, MJHED05, MJHED06, PMRAS05b, RLY06, SF07, VS07, WL09c, WZ06h, WZ07d, Xia06, Xia08, YHL08c, ZXF07]. Chebyshev-type [JYX06, Mao09]. check [HRT07]. checkerboard [WWXX07]. checks [Che05h]. Chemical [GAS08, AS06d, CGJ08, EM11, FJ06, JC06, JCL08, KR09, MR05a, XHSL06]. chemostat [MKG09, ZGL08]. chemotaxis [SS05a]. chemotherapy [MKA07]. Chen [LW05c, PNL05, Yan05b]. Chien [YRY05c]. China [ZLMW05]. Chinese [CL05a, Su07a, Sun07a, Wu05a]. Chlodowsky [İBİ09, KG08]. chlorinated [Slo07]. Choi [LL05g]. choice [BA09, EGE06, Raj08, Tan06d, Tan07c, Tan07b, Tan07d, Tan08a, WZ06k]. choices [RS06]. Cholesky [LZ06c]. choosing [Fad06, GANG07, JP09]. Choquard [CG07b]. Chord [RH08]. Chree [Waz08w]. chromosomes [CCFC08]. chronic [CL08a, LH05]. Chua [TCDV07, TCMPMC07, TCMP07, WC06b]. Church [Gal06b, Smi06a, Zie09]. Chyan [EDS06]. ciphers [Kat05, LMSV07]. circle [KSGM06, WYZL08]. circles [CH07f]. circuit [jF09, Li09e, MSM07a, TCDV07, TCMPMC07, TCMP07, WC06b, WLLL07]. circuits [Fen05]. circulant [ES05b, GG08c, GG08d, MM05d, Rim05b, Rim05c, Rim06c, Rim06d, Sol05a, Sol07, TAK07b, Tzo08, UNS07]. circular [ASL07, BL06, CC07b, CLW09b, CB06, KS08d, KS08c, YAO09, Zhu05c, Zhu05d]. circumvent [LMH06]. city [NSS06b]. claims [BCMT06]. Class [SC07e, ANM06, AAN05, ACP07, AGR09, AC07b, AZ07, Att07, BAF08, BK08c, BG07b, Bou06a, Bou05, BZZ08b, CZH07, CCWJ08, CZH08, Cen06, CS09b, CG07b, CH09, CAA07, CKOS07, CAS09, CT08b, CD08, CY07c, DNS08, DF09, DDR06b, DZG06, DZ08f, EJ07a, EGE06, GR06a, GWX06, GQS08, GGOS08, HA05, HW08b, HC05a, HH08b, Hub08a, Hub08b, IS09b, JSL05, JZ08c, KK07b, Kah07, Kan09a, KR05, KB06a, Kar09c, KMÖS09, KuI08, KA06e, KK08d, Lay09b, LW05a, LH08c, LTW07, LCS09, LL05g, LS07e, LW09c, MKKS05a, ML09, MSD06c, MJ05, Moh06, MHA06, Odi08, Par05b, QBH06a, QBH06b, QG08, Rad09, RB07, RSSS07, RJM07a, RMG07a, RAP05, Ras05c, Ren07, RW08, RWB09, RGA06, Sai07, ST06b, SA09, SJ06b, SS06c, SDG05, SZ06a, Sid05, Sid06, uITA06, Śmi07, Sok08a, Sok09, SLW08, SHD07, Sun08]. class [TL06, TL08, TZW09, TZH09a, TXW09, TZH09b, eT09, WZ06a, WR07a, WL07d, WWWL07, WL07e, WS07b, WLP07, Wan08a, WZ08c, WL08a, WCH08a, WJZG08, Wan09c, Waz05e, WG07b, WYL06c, WWY08, WZY07, WY08, WC09c, Xia06, XT06, XQ06a, Xin06, XB06, XL07, XZ09b, Yan07a, YLL^{+}05, YLL06b, YC06a, YZ07b, YL08a, YH09, YWL09, YC07a, YQ09, You08, YY08c, YHL08b, Yua08, YHL09a, ZZ08a, ZW09b, Zha05a, Zha05c, ZLC06, ZHLG07, ZZ07a, ZT07a, ZLZZ07, ZL08a, ZXLS06, ZXWL06, ZX07b, ZWX08, Zha09g, ZXZL07, ZYZ07, ZX08, Zhu05e, ZXL06, d’O05a, dlS06b, lS07a, Lal09]. Classes [Sri09, AMN07a, ANM07e, AMN07e, ANM08, AM06, CK07a, Djo06, EM06, KKA09b, MB07, NO07, Noo07b, NMM09, Özk07, Raf07a, SEŞ09, ST05c, WCO09]. classic [ÇE06b]. classical [Ang06, BT06b, Cha05h, CG07a, Cor05c, Deh05a, GRZ06, GJ07g, Has06b, JLL08, NO09, REN^{+}05, Wan04, WLS09, Yi05]. Classification [PG07a, SGG^{+}09, WHD^{+}08, XL09a, ABK^{+}06, AG08, CL06b, Dan05, DCS07, Hu07b, HZYL08, JLS^{+}05b, KM09b, PBKM09, PG08c, YSCC05]. classifications [JLS^{+}05c, LLL09]. classified [BHCK08]. classifier [CYL^{+}08, GPR^{+}09, PG07a]. clearing [ALN06]. Clenshaw [KKY07b]. clique [SLWX06, XSKC07]. clique-transversal [XSKC07]. close [Noo08a, NAUH09, OP09, PC07b, PVY09, RAES05]. close-to-convex [Noo08a, NAUH09, OP09, PVY09]. close-to-convexity [PC07b]. Closed [Cvi08, Kol06, WFHD09, Att06b, Cvi09a, DMJE05a, ETJ06, EDMJ05c, LS06a, Muk07a, Sim09b]. closed-chains [Att06b]. Closed-form [Cvi08, WFHD09, Cvi09a]. closed-loop [Muk07a]. closure [LL08a]. cloth [XL05]. cloud [TTC08]. clouds [ME07]. cluster [SCL^{+}05a, KK05c, KK05d]. Clustered [AT07a, MAT08]. Clustering [XTCC08, AGA^{+}07, CC07b, FAM07, GÜ07b, JCSR07, MCAF08, PLZ07, SCL^{+}05a]. Clustering-based [XTCC08]. clusters [LWL06]. CMAC [ZS08b]. CNNs [WL09a]. Co [MED08, Bel07b, zHWH07, LY08d, OASM06, ZY08c, Liu07a, XFTY09]. co-coercivity [ZY08c]. Co-evolutionary [MED08, zHWH07, OASM06]. co-evolving [LY08d]. co-state [Bel07b]. coagulation [ZZX05]. coalitions [II07]. coating [O’R08, SH08a]. coaxial [Rad05a]. cobweb [DW09b]. cocoercive [HG09, HN07b, yLcCsL09]. code [ACZ05, CVMM06, Che08c, Dön06a, DK06, SS05e, Tso08]. codec [Che08b]. coded [DT07a, DT07b, DSKM09, KSSA07, MBM06]. codes [AS07b]. codesign [KBTS07]. codewords [WY05a]. coding [CK05, Che07h, KH07b, Sha08d]. Coefficient [ARS07a, LW09d, Cen05, CG07c, DH06a, DH06b, DKA06a, DSZ08, FSG07, FCG06, Fuk06b, HED06c, mH07a, Hon09, LLL09, MGG^{+}09, MTL08, SML05a, SZ07e, TM08, TW09, XMGW09, ZDY08]. coefficients [ADH05, AES08, AM08a, Alt07b, Aou07, BW07, BRT07, Bou06b, Boy07a, Bra05, Bra06, ÇFÇ06a, CL07b, Den07, EBP09, EDMJ05b, Fat05a, Fat05b, FK06b, Fra07, GS05, GM05b, HMJD06, KK05a, KS08d, KRRGT07, Kıy06, LZ08a, Liu09d, MSTY06, zMsCxL08, MCS08, MS08c, Moh05a, Moh05b, Mom05b, NLD08, OMST07, OIL09, Ozb07b, Par05a, SAD06, SW06a, Sil05a, Sil06c, Sil06h, SM08a, WL07d, WL07e, WL08a, XC06b, XC07, XCL09, lYqL06, lYqLW07, Yan08a, ZS07b, ZCMQ07, ZHX06]. coercivity [ZY08c]. coevolution [CZ08a]. Cohen [CZF09, JSL06a, LYJ09, Li09c, MGH08, WZ08a, Wan05c, WRL06, YX06a, YZW07a, ZD06b, ZW06a]. Coiflets [AM08c]. Coifman [MLR07]. coins [QL06]. cold [AKKS06, JW09a]. cold-standby [AKKS06, JW09a]. Cole [GS08a, Gor06b, Waz09b]. coli [PG07c]. collage [KH07b]. collapsible [SI06]. collected [DRS^{+}08]. Collinear [Shi09a, LF08]. collision [DM05]. Collocation [Çel06b, Ras05a, SH05b, AV07a, ADK05, ADK06a, Çel05a, Çel05b, ÇG05a, CM08b, CJP08, DN06, DjT05a, DjT06, EHZES05, Fad06, GJ07g, GSG06, IMB08, JS07a, Jav06b, Jav06d, Jia07b, KA05a, KY08, KYK08, KK08a, KKY07b, KAM06, MD06a, MAM06, MD06b, MD07a, RK07b, Raw06a, Raw06b, SG06, SGT06, SG07a, STU09, WZ08c, Wan08c, Wel08, WCH08b, YS06a, YHL08c, RZ05]. collusion [HW05a]. colony [BHCC09, CK08, DT06, DS06b, FYZ07, HP07, KF08b, SAK08, SK08c, SSJK07, TS08b, Tok06, YW06a, KA09a]. Color [Che07h, Che08b, MM05c, WLL07b]. colored [Alt06, XLKH09]. coloring [AASAA06]. Column [BT09a, GJD^{+}07, UP07a]. combination [Dzi07, GK06a, Kan06a, pLT06, MD06b, MD07a, LY09c]. combinations [BR08b, Ben07b, ÖSÖG09, SÖ08, Yu08a]. Combinatorial [EMC05, HGN06, Iri06, JCSR07, AE06b, Chi07a, JDSR08, SSMI09]. Combined [PD09, Bie08a, BGR05, DZ08e, HW08a, KI05b, KH09, Li09g, NJG08, OASM05c, Pet08a, WH06a, WW08b, WSL09, WDLM10, Waz06o, Waz06q, Waz08a, Waz08q, WC05d, XLwW07, YEA07, YHW06, ZT05, ZZ05b, ZT06, ZZ07d, dCdSJM08]. Combining [CGS08, GZL08, Jia05, PG07c]. combustion [AD06c, CLZ^{+}07, ESS05c, SG05c]. comets [SBS08]. comfort [XL06b]. Comment [ÇT07a, CX08b, Den08b, EM11, HV08, HES05, Kah05a, Kah06a, Pan06, Pan08b, Pan08c, Pan12, RJM07a, SCS05b, WW05c, WY05d, FG09, Ma11, gZW08, Sub07]. Commentary [WW08a]. Comments [And06, Chi05, Has06b, Pet07, Pre06a, Pre06b, Sal06c, SQ11, WDLM10, WZ13, XYL10, XY11, ZW06b, ZZW09b, BFR05, XF08b]. commerce [AH07a]. commercial [LCC05c]. commodity [Lin06c, Lin07f]. Common [AR09, ACOS06, IR08a, LL06a, AKH06, Ber09, sCLC09, ĆC09d, GR07a, JLM05, JMLR05, JOB08, KA06c, KAO08, KA06f, OS09, Son07, SQ07, Wan06d, WCN08, WCL08, YVU06, YLM07, ZS07a]. common-multiplicand [YLM07]. Communicability [EH09d]. communication [OHJ07, Zho07]. communications [SL05a, VV06, YTWY05]. community [EH09d]. commutativity [Den08a]. commuting [JOB08, ÖSÖG09]. comonotonic [OML09]. comonotonic- [OML09]. Compact [Odi08, Waz05f, ZC06c, dAC09, DZ09, Djo06, DM08c, HGN07, Inc07a, Inc07c, Lia08, MMW09, MH06b, MH06a, O’R09, SG09a, Sha06b, Sht07, WZZ06a, WZZ06b, Waz05d, Waz05e, Waz05h, Waz05j, Waz05k, Waz05m, Waz06f, ZZ05b, ZZC06, Zha07h]. compaction [HT06a]. compactly [LFW08, yXhWZ06, YH09]. compacton [WF07c]. Compactons [Waz05g, Waz06c, RV08, RV09, Waz05i, Waz06l, Waz07h]. Comparability [YSXL08]. Comparative [KYK08, KC07b, REN^{+}05, WC09b, DESG08, KA09a, KuI08]. compare [II07]. compared [CCE06]. Comparing [DES07a, ESES06, EM07, FCG06, GAFK07, Has06c, KTW08]. Comparison [Cao06, DMA07a, EG07a, GK05b, HPD07, IRSR05, IT06b, JLL08, Lia05, MHM09, NMM08, Pin06, Ser05, SH06a, TH07b, WDK06, Abb06a, Abb06c, AH07d, ACD^{+}05, BE05, BIK05, BN07, Bou06b, Çak08, EGX05, Gor06b, Hor07, HW09, HT06b, Mal06, MN07, NA06b, SLLL07, SH08b, Sol05c, yS06c, Utk06, WYX05, Wan06l, Waz06d, WLY06, WLC06, WLC07b, WLC07c, WLC07d, ZS07c, ZW05j, Zot07d, Kan06a]. comparisons [JF05, Su08a]. compartmental [ZLMW05]. compass [TK06b, TK06c]. compatibility [Li06a]. compatible [Ber09]. compensation [JCX^{+}07]. competence [Lin06b]. competing [Luo06]. competition [Che06h, Che06e, CF06b, Chu05, Day05, Hua08a, JW07, Li07d, LZC07a, LZ09c, MKG09, PW08a, WW05b, YC06b, ZY07]. Competitive [Lis06, Che05a, Che06a, Che06c, GLV08, GL06a, LKW06, LFC08, QD08, QC08, SM07d, WS07b, ZT07a]. competitor [Yan09d]. complement [CI09, MSCG09, SC08b, WLC07a, ZW07c]. complementarity [AP09, BNKZ09, CR05b, Hua05b, HW06, Jia05, bJyKxC09, JZ07, Li05a, LD07c, LM07a, LMN07, Ma05a, MTC07, MJ07a, MCW08, MT08, qMpXjH09, yNyF05, yN05, PL06b, TLM09, Taw07, XL07, Yan07a, YM07b, YHW06, ZMW05, Zha08a, ZWG09, ZZ06b]. complementary [BK07c, BDY06, HGH06]. complements [CKKS08, CN09c, YLM07]. Complete [CC06b, Cha07e, DK05, AA07b, JMLR05, WC06b, XXTL09, Zha07a]. completely [GQS08, QG08]. completeness [HAS07b, KI08]. completion [AASAA06, Ere09a, WNCL06, WSJW06]. completions [JAT09]. Complex [Dem05c, NR08a, YX06b, AM08a, AC08, Att07, BDCW06, Cot08, Deh08, Den09b, EH08a, EH09d, GB07, GS07c, GG08c, GG08d, Ism08b, KNM05, LZC05, Ma05b, MWQ09, MP08b, Men08, MD09c, NBO09, NAUH09, Özk07, Ree07, SB06b, SG06, SGT06, SdHT09, dSYGM^{+}05, SEŞ09, Sum07, WL06d, WLLC07, XQ06a, Xin06, XS09]. complexities [AFH05]. complexiton [AC08, LZ08c, SZ07b, WSZ07b]. Complexitons [WZF08]. Complexity [AP06, AP09, Cha07a, CS07a, CS07b, DESK06, DES07b, ESD07a, JSdM07, KC07c, PRGH05, XSKC07, Zie09, EG07c]. compliant [SK09a]. complicated [WHD^{+}08]. component [AK05e, AK05g, AN06, AKKS06, CL05b, EGS05, Gop07, GBG07, KA05c, LM08a, qLJ05, PG07b, PG07d, RS06, Waz09e, YX08]. components [AGT07, SANN07, WDK06, WLC07b]. componentwise [WCX05]. composed [KLW07b]. composing [XRZ09]. composite [AR08, BTMB09, Chu07c, DD06a, FK09, Jun09b, KLW07a, OAHR09, RVN07, RNV07b, RNV07c, Sha05e, Tha07, WN08c]. composites [MA06d]. composition [GK06b, KG06, KGM06, KLL^{+}08, LS08d, SW09c, Ste09c, Ste09d, Ste09g, Ste09h, Ste09i]. compound [hB07, SY07a, She08]. compounded [GY08]. comprehensive [TS06]. Compressed [Ma08c, PBKM09]. compressibility [EA08b]. compressible [EESESD05, Ere06, Gar07, GS08b, PPS09, Rad05b, Ras05c]. compression [MJ05, Tse05]. compromise [AR08, BTPGAPRU06, BTPGAl06]. Comput [Bil07, EOM08, ESAH10, HLSS10, KAK05, KAK06, Kas06b, Kum09c, KP06b, LL12, hMfWzLpZ09, MK10, Pak11, Pan08b, Pre06b, Sal07d, Sol07, Tua13, UNS07, WL08a, Wan08b, WZ13, Waz08b, Wey08, WH09, Yin08b, YHL09b, YHL09a, dCD06a]. Comput. [Mar14]. Computability [LCM07, Coo09, Tan07d, Tan08a]. computability-theoretic [Coo09]. computable [MN05, Vel06]. Computation [AAG05, BUÇ07, CLSZ06, CK06d, DEH05c, Den08b, EM11, HV08, Hai08, HES05, LRMS06, Li05f, LCZT08, Nav08a, ORG^{+}09, Pan08c, Pan12, PSRN06, SQ11, Sil06h, Sub07, Van09, Wan09b, Yam05b, ZW06b, AH09, Ari07, Att06d, ALN06, BF07a, BT06a, Çel05b, CZ05a, CW05a, CW05b, Chi07c, CK07b, CH07f, CT09b, DwWmW05, DM07b, EFI06, EHL08, GVMT07, GWX06, GM06b, Hos07, Hua08c, Inc07d, JZH07, KYA05, KM05b, LMSV07, LZ08c, LKC05, LCC06, LHC07, LL05h, MG06, Muk08b, Muk08a, Myc06, NO09, PBPF08, PGdSS08, PVM07, Şen07, SdHT09, SZ07b, SZ07c, Sor06, TSP07, TS08a, Vaj08, WLZ05, WZZ06b, WZSZ06, WZ07c, WYD07, WLHH05, WY05d, XZ05, YLLL08, ZC06b, ZX06c, ZX07a, ZZB08, Zha09b, ZZYZ06a, ZS09b, KPL09a]. Computational [ANM08, AC06, Att06b, AFH05, DHG06, EB05c, GBC09, HWC07, Kat07, Kat08, KP07, LWH06a, LWL07a, Liu06c, MYN07, NKE05, NS06a, Pet08a, RMAM07, SS09a, Sey05, SdJA06, SHSZ07, WL06g, WLY06, WLC06, WLT07b, WLC07b, WLC07c, WLC07d, WLL07a, ANM06, AMN06, AA07c, AA07d, Amm05, ACD^{+}05, Baw05, BM05b, BD09a, CG07c, CG08b, Dam08, DJ06a, DDD05, DESK06, DES07b, EMC05, EM05a, ESK07, ESD07a, FSG07, GG06, GSDB06, GGSN06, HGH06, Hos08, Hub09c, IIS05, JSdM07, JCM05, Kar06, KEM08, KM05c, Liu07d, MY06a, MLR07, Mir06a, NZ08, ND07, NKR^{+}08, NA08, Öza05, QW09, RMA07a, Ras05b, SY07b, Wu07f, ZZ06a, ZLC07, Zie09, dSYY08, GGSNR06]. Computationally [ADK07, ADK06b, Ars07b, JA08, SH07a, Sha08a]. Computations [BT09b, Fen06, Sha06c, WL05b, XCFX05, AR06, BQYW07, Du06, IR05, KK05c, SZF05, TC05, VV06]. compute [ANK05, BSC06, MMS^{+}06, Sal06d, SC09, SCS06]. Computer [Alt06, BAK07, ESAH10, FA05a, LW06a, WC05b, YY09, ACD^{+}05, AGI08, CPTZ05, DB06, GD08, MYSA05, ML07, MH06b, MH06a, MJ07e, MS07d, MS07e, PA09, SA08c, SGP07, Zot08]. Computer-aided [WC05b]. Computer-assisted [YY09]. Computerized [RMA07b, WW06]. computers [CH06a, ND06, TTZ09]. Computing [AR08, Boy06a, Bry07a, Cha07b, CT05b, CDP06, Eng06, FS09, GS06f, LBS05, LTT06, LZR07, MMS^{+}08, Odi09, SG08a, Sch08, SG09b, ST05b, SP06e, Abd06, ACJ09, Att05b, Att06c, AL06, Baz08, BE06, BIK06, CS07c, CL05a, DN07a, Del09, Djo07a, Du05, Ghi06, xGyZtZ^{+}07, Hua07, Hua08c, HHL08, KKT05, Kat09, KAT06, KSP08, KWP07, KC05b, LR07a, LR07b, Li05d, Li06d, Li07b, Li07c, Li09d, LHB06, LSBL07, LCC05b, LYH09, LHL08, MYSA05, MA05f, Nač09, NS06a, NG06a, NG06b, NR07a, NRA07, NS08, NM08a, NBO09, Oca07, PU08, PH06, QBH06b, Rim05g, Rim05a, Rim05h, Rim05b, Rim05e, Rim05c, Rim05d, Rim05f, Rim06c, Rim06g, Rim06f, Rim06d, Rim06h, Rim06a, Rim06b, Rim06e, Rim07c, Rim07g, Rim07h, Rim07d, Rim07f, Rim07a, Rim07b, Rim07e, Rim08a, Rim08b, Rim08c, Rim08d, Rim08e, Rim09b]. computing [Rim09a, SS09b, SCI08, UP07b, VRG06, WTW05, WL07f, WS07h, WN09a, YYWW07, Yin08a, YT06, YHL08a, YHL09b, ZZL05, ZCW06, ZYD07, Zie09, Ant05]. comrade [Kar07b, Sog08d]. concave [LC05b, MA08d, XXF06]. concentration [DRS^{+}08, NAM^{+}07]. concept [Lee05d, MFF09, NN08, Oma06, PDGR07, Sal07c, SGUK08, Sla06]. concepts [Cha05g, Ino08]. conceptual [HAS07b]. concern [JC06, JCL08]. concerned [NO07]. concerning [LCZ05a, ML09, MC08, Wu07b]. concrete [BOB05, dCDLN^{+}07, WGCH05]. concurrent [Kos05, TTZ09]. condensation [Hu05]. Condensed [GW06a]. condensing [ZZZ09]. Condition [CX08b, DQW06, LWW05, MD08b, MD09c, WG05, WWW03, AMN07d, AAK07, Bat08b, BŞ07a, Bou06a, BT06b, Cai06b, CA05, CZ07a, CC06c, Cui05, DZ07a, DjT06, Has06c, Hsu06, HWS07, KE05a, Küç05b, LMH06, LXJ09, LZZZ06, LHWW07, Mai05, OIL09, PK08a, Sid09, Sok07a, Sok08b, Son08a, WMX07, fWZpXzL09, Wei06a, fW07b, XL09c, Yan05c, YM06, Zhe08, Zhu05c, Zhu05d]. conditional [Li06a]. conditioned [KB05b, TK06c, Wu06e, Wu07d]. conditions [AP06, BZ09, BM06b, CZH08, Cha07c, DM06, DMR07, DB07d, EPER07, Fil05, GH09, HK09, HM08a, HGH06, Hus09, ID08, IMS09, JOB08, JLW09, JV05, JM06b, KWX08, Kar09a, KV05b, KTV07, Lee05c, LL08e, LZ08e, hLqSlZ09, LA07b, MV09a, MD08c, NL05, OB09, PC07b, Pe¸ 09, QM05, RB08b, Ren06, RAS08, SS07a, SS09c, SGT06, TS09, TNL09, VCRCG09, Via08, Vul07, WJ05, Wan07c, WTW08, WL08b, Woo05, Wu06f, Wu07b, WX08b, Yam08, YCL07, Yüc08, ZGK06, ZS08a, ZW08c, ZL08c, d’O05b]. conducting [EB05c, ESMM06, You06c]. conduction [ACZ05, Hsu06, KBLT08, Li09a, LLW05, MM05a, MC06, OIL09, QFX06, QFS07, SK05, SZ05, SP06a, Waz05p, XFL06, XFQ06, YM06]. conductive [BS09a, TA06]. conductivity [KJ09, KE05c, PVS^{+}08]. cone [AH05a, AZ09, Che07i, Che09b, Ece06, EM06, FHH09, FJCW05, Kas09, Sol05b, SCO09, WB09, ZMW05]. cone-invex [Kas09]. cones [Mor05]. conference [RIY05]. Confidence [CK06b]. Configurations [KB06b]. confirmer [GT05a, ZSNS05]. conflict [XLHK09]. conflicts [HKH06]. confluent [HRS07, XL06a]. conformal [ASL07]. conforming [AAN05, GMCM06, LAL07]. congested [SM06b]. congestion [Chi08c]. conic [JjQjWmL06, Köp07b, Köp07d, LCC05c, LCZ05b, LN08b, WZT06]. conjugacy [Shi09b]. Conjugate [AJE05, And09, BMP06, CS07i, qDyC08, xGyZtZ^{+}07, GG05b, HWS07, MSM07b, PVN06a, SWH07, mTxWyL07, yWjL06, jW06b, WYL06a, WLQ06a, WFHD09, YZW07b, YGW09, Zha09b, Zhu07, HL06a]. conjugation [JV05]. connectedness [KH05a, KLL^{+}08]. connection [AES08, JR07]. Connections [Ahu08, AR08, BS09c, GD06, KAP06]. Connectivity [ME05, CT07b, LCX09, PK08d, WYL08b]. conquer [MCGS05]. Consequences [dCD03, dCD06a]. Conservation [Bri08, Bay05b, Che05e, KK06, KK08d, Mel08, NMM08, Sea06, SM05c, Sjö07, SMHQ06, TmChL06, Wan09a, XL09a, YL06b, Zah09, ZCZ08]. Conservative [WZC07, HRE05, OAAK08, Ozb07b, cWmZ06, WZ08c, Zha05a]. consideration [IR06, LMT^{+}09, RAG^{+}07]. Considerations [MS06a, BWB05, PBG07]. considering [ONG08, Oma06, dCRDSSN^{+}08, TMMVA05, Uje09]. consistency [DJ06a, WYX05]. consistent [tLxLlW07, WWL06, WSL07, YL09b, Yu09, Zhe08]. CONSLAW [YL06b]. constant [AMN07c, AK05f, BRT07, Geu07, GG08e, GG08f, HS09a, IS09a, Kil08b, Mak07a, Mak07b, WS07h, YKL08, ZzLZ08]. constant-diagonals [Kil08b, WS07h]. constantly [VFF08a]. constants [Cho07a, Gin07]. constellations [ESETA^{+}06]. Constitutive [NKR^{+}08]. Constrained [HY07b, TK06b, ASA06a, AL09, AKB09a, AKB09b, BGRS09, BS05c, Che06m, DD08b, DQW06, DWT05a, EJ07a, FYZ07, FB07, Ham05b, HH06a, HWH09, HW07a, HCS05, Hua06, zHWH07, HMW07, JDSR08, JJT06, JZ08b, JZHT05, JZTH06, bJyKxC09, JZX07b, JZX07a, Jin08, KMA08, KKY07a, LT08a, LZ06e, MK05a, MO05, MP08b, MSD06d, Mor08a, Muk08b, OB09, PSW07, PN08, hPY06, PL06c, PFL09, Ran08c, SAB08, SI06, Tso09, Udw05c, WSZ07c, WZ09c, WLY08, WLL09, XW07a, Yan06b, Yu08c, YK06a, YKJ07, Zhu05e, Zhu06c, Zhu07]. constrains [JjQjWmL06, XW08]. constraint [CWLH08, GS06e, IR06, IR07, LWH06a, fLyHZ09, LD07d, Mav09, QZL07a, Sal08d, SS06a, SJ06a, SGB05, YC08a, YE07, ZLZ06, lZ06d]. constraints [AL09, ALD08, CCL06, CCC07, DZG06, GK06b, HH06a, JXH06, JGS06, KG06, KGM06, Lin07b, Lin07d, Lin07e, MD06c, OASM06, PS06a, QZL07b, QW08b, QZZ07, RAN08a, SG08c, SDM08, SLC07, WS07c, WZ08f, WX08b, Yan07a, YM08a, YES07, YHL08b, Yua08, YHL09a, ZCH07, Zhu05a, ZZ06b]. construct [CW05a, CW05b, CW06d, FZN07, Kin05, LZ08c, Nar07b, YL06b, ZZYZ06a]. constructed [Aks05, EDS06, Tsa05]. Constructing [ABP05, CCJR07, YZ07c, Yu08b, Fil05, JZCW05, Li06e, ST05c, YY06c, Zha08c, ZL05b, ZZ06a]. Construction [Ars07c, Chu07b, LY09a, LZ07c, LZ09e, PP07b, XY07, YH09, ZDY08, AVKA09, CC09b, CHL06, CHT06, CYL^{+}08, Chu07f, hCxC05, FA05a, HHZ09, JH08a, KT08b, LXSH06, SS09d, WD06, XZ05]. Constructive [JC09, CT05c, LP08c]. consumption [AGS07, AGTS07, ESS05c, GL06b, LCZ05a, OMR07b, SLC07, WZ06k]. contact [BS09a, LYCW09, WCX08]. contained [BŞ07a]. container [Cho07b]. containing [Arg09, DW09d, Li05b, RCB06, ST09, Wan06k]. contaminant [RM05, RM06]. context [ABB06e, JA08]. contiguous [GG05a]. contingency [RSDC08]. Continuation [Att06c, CS09c, DM05, DH09, KLW06a, RX09c, Wu05a, Wu05c, Wu06c, Kad07a]. continuations [Eli07]. continued [Kıl08a]. continuity [AR09, Sht07, Tia09, ZW08c]. Continuous [DCR07, Zha09a, Ahm05b, AS09a, AGM^{+}07, BAB05, BS05c, CPY09, CLH06, DXH07a, DS06b, GFWX05, Ham08, HMW07, JPCL08, LZ07b, LMY07, LYL07, MV09a, Pam05b, PLTW07, Pha06, RYW06, RA09, Sak05, Sal08c, SSJK07, TmChL06, TS08b, VL09, WCO09, WZ05, XYZ09, YFLX09, YJO08, ZZYT08, Zhu05e, dlS06b]. continuously [Li09c, MGH08, Shi09b]. continuum [BDCW06]. contour [Boy07a, TM07, YZY08]. contracting [Nwo07c]. contraction [OS08, PHL06, SD06]. contractive [Ber09, ĆRR^{+}08, HBN06, JOB08, O’R09, WZDgH09]. contractivity [TG08a]. contrasting [SD06]. contribution [ZYD07]. Control [DAQ09, KM05a, LL08a, ZŠ05, AD06a, AJOGRLRG08, AJRLRGOG09, AK08b, AK06i, BK07a, BY05, BS05b, Bel07a, Bel07b, Bel08a, Bel08b, BS09b, BS07c, BS08, CCS05, CZH07, CZ07a, CZH08, CPSS08a, CPSS08b, CMMC09, Che05a, CH05b, Che05d, Che06b, CF06b, Che07d, CH09, Chi07c, Chi07h, CTGD08, CST07, Cot08, DS05a, DT09, DF08, DWT05a, EJER06, EB05e, EG05, FDL08, GBK06, Gal08b, GLLC08, GM08, GZ07d, HC05b, HCX06, HLY09, IM08, JPYW09, JC06, JCL08, JLL08, KKF07, KRG06, KO08, Kla08, Kos06b, Kos07, KS09a, KB07, KP06b, KP06c, LY05, LCC05a, LP07b, LYX08, LL06a, LYL^{+}08, LCZ05a, LWL07b, LLZ08, Luo06, LG08, LHW07, MZW08, MI08, Mah09b, MP08b, Mor08a, MB05b, Muk08b, Muk08a, NAM^{+}07, PK05b, RBvB08, RY05b, SMA08, SJ08b, Sal06a, Sal05, SYT06, SMBB07, SFFY08, SH07f, SWX07b]. control [TZX07, TCMPMC07, TCMP07, Udw05b, UBKC08, Wan07f, WS07a, Wan08d, WJZG08, zWxY07, WGCH05, WF07d, XW09, XS09, YZ07b, YF08, YCL07, YCL09b, YWZW07, Yu05, ZZ08a, ZW09b, ZLMW05, ZC06a, ZLZZ07, ZWY07, ZS08b, ZWLX08, dlS06c, lS08a, lSAQ08]. Controllability [PNL05, Kla08, RES07, Wan08a, Yan06b]. controllable [Lee05a, LWH06a, LWL07a, Wan06d]. Controlled [Haj06, BCR09, Chi07g, Chi07i, GZC08, WXW09]. controller [GZG09, LZX08, OASM05c, Par05c, PKL08b, Par09b, WLP07, WS09a]. Controlling [PRGH05, WY09, Yan05b, LXL06, SKSJ06, Wan08d]. controls [CLH06, Che06g, Che06c, Che07b, CYCX09, GM07, QD08, WS07b, Wan09b, Yan06b, dlS06b]. controversy [Rös07]. convection [AES05, AB05c, AH05a, AL08, Baz08, Bi06, BB09, Bog05a, Bog08, Bra05, Cen05, Che06n, Che07i, Che09b, DG07a, Ece06, EB05c, EG06a, EWAE06, EESF04, FM05, GY06, Hu05, HTM09, JK07, KY08, KS05b, Kuo05, LC05b, Lin08, Lin05f, LLZ09, Mur06, PM09, Pan08c, qQcM05, Rem07, Sal06e, SW09a, SDG05, SFA08, SM09b, SAS05, SCB07, STU09, TM08, Zha08h]. convection-diffusion [Cen05, GY06]. convection-dominated [Bra05, EG06a, SW09a, SM09b]. convective [Bat08b, EM11, KR09, KJ09, Mak09]. conventional [YW05, YRY05c]. Convergence [AS07a, Cai07, sCLC09, CL05c, CS07g, CXZL09, CAA07, DLS06d, EGE06, GH06, GG09b, HZ06a, KB05a, xLyW06, Li09f, LX07b, MMN06a, MGH08, NS09, PLL08, Pog07a, QCS09, RG05, RMJG07, RW07, RW09, RHY06, RY06, RZ07c, RMWk09, SS05b, SS06b, SCX07, SL09b, Wan05b, WB06, WDL07, sWX07a, Wei09, WR06, WRB07, WH09, XZ09a, XCL09, YN07, YKKY09, Yun07a, YK08, Yun09, Zha05a, Zha06a, ZH07a, ZHGG08, ZZC06, ZW08b, ZWZZ07, ZH08c, AKO06, Ahm05b, AHR08, Ami05, AM09a, AH08b, BD06a, BDDB08, BRW08b, Bil04, Bil07, BM06b, BCC^{+}09, CPY09, CS07e, Che06j, Chu07a, Chu07f, Chu07i, CH08b, CN09a, CR05b, CK07c, DJ06a, DH06a, DHY06, DH06b, qDyC08, EHR09, FGfFL06, GKR06, GSDB07a, GSDB07b, GK06c, GBG08, Guy09, HWH09, HR07b, Hor07, HN06b, HBN06, Inc05, Inc06a, Ism05, İBİ09, JK06, JC05, Jun09b]. convergence [KSM05, KOR08, KI05a, Kha07a, Kha07b, KLW06e, KLW06b, Kou07c, KL07a, KL07d, KL07e, KL07g, KLW07e, LZ07a, LLAT09, LZA^{+}07, LL07a, MT08, MNY07, MV08, NS07b, NNMDS06, NN07a, NN07e, Omr06, Pet08b, PKT06a, PKT06b, QSZ08, RN09, RYW06, Ren06, Ren07, RA09, RWB09, SA08a, SZ07a, SG07b, SGS09, SC05b, SAM07, SC06b, Son08a, Son08b, SL06b, SQ07, SC07c, TK07, Tha07, Thu07a, Thu07b, Thu07c, Vaj07c, WG08, yWjL06, Wan07c, WKL08, WLW09, WYL06a, WLY08, Wu05d, Wu05c, WZ06h, WH06b, WZ06i, WZ06j, WZ07d, WR07b, Wu08b, WMC08, WLL09, XCD06, XL09c, YNCL08, YLK09, rYzS07, YZW07c, Yua05, ZG08a, ZSCW09, ZWC06, ZZYT08]. convergent [Ash06, BP06a, Bra05, CL06c, CS07h, Chi07e, Chu07g, DHL05, HFL06, HL07b, JZTH06, JXH06, bJyKxC09, KP06a, KPS06a, Kah07, KS08b, KT08a, Liu08a, MTC07, Ma08b, MCW08, MKKS05a, NA06a, PS06b, WWWL07, WN09b, YGW09, Zhu05c, Zhu05d, Zhu05g, ZZ06b]. Conversion [EM06, Boy07b]. convertible [LWK05a, ZW05b]. Converting [Qiu05, Boy08b]. convex [AK06e, ĆC09d, Dzi07, EB05a, EGN07, Fab08, FY08, GK06a, Ham05a, Ham05b, HC08a, Kan09a, KKY07a, KBÖP07, Lin08, MA08d, MM08b, NO07, Noo08a, Noo08b, NAUH09, OP09, Özk07, PSW07, PL06c, PFL09, PVY09, QCDS09, QW06, RSSS07, SG08c, WL07f, WZDgH09, WZ08f, YC08a, YLY07, YM08a, You05, YL07, Zha07b, ZZZ09, Zha09h, gZW08]. convex-power [ZZZ09]. convexity [AH06a, AKA07, PC07b]. convolution [Bry07a, BO06, Irm07, LP06a, MA05d, MS08a, Noo06a, RFFFCC08a, SEŞ09, ZW05f]. convolutions [BH06, Djo07b, SY09]. cooling [Hlo05, Mak09]. Cooperation [Chu05, CLH06, Che07b, DJJL06]. cooperative [FK06a, JZHW08, LKW06, LL08d, LTZ09, NZHW07]. coordinate [KA06f]. Coordinated [Rig08]. coordinates [AFH05, Bri08, DT08]. coordination [yLLjP06, WS07g]. core [MFF09]. corner [Boy06b, Lin09b]. cornered [Soh05]. corners [WWXX07]. corporate [Nwo06a, Nwo07a, Nwo07c, WZ06k]. Corput [QCN08]. correct [MV09b]. Corrected [GS06a, LW06c, LD06b, Ran08b, UE07, ZCMQ07]. correction [BN07, Çel06b, EHL08, HL06b, JB08, LZ06c, LZ07a, LY07c, Wan06b, WLK08]. corrections [ESA05]. corrector [Noo06b, NN07b, SMA06, Sal07a, SS07c, Wu06d, Wu06e, Wu07d, XD07]. correlation [CL07a, HCLPJL^{+}08, SML05a]. correspondences [Tan07c]. Corrigendum [Bil07, Hai08, KPL09a, LL12, Tua13]. cosh [Waz06q]. cosh-Gordon [Waz06q]. cosine [KGB06, NK05b, RV09, Sil06e, TB09a, Waz05m, Waz06m, Waz08w, YC09c, KH09, Waz06o]. cosine-Gordon [KH09, Waz06o]. cosine/sine [RV09]. cosmic [Sha06c]. Cost [AEFFA08, WC06a, AKR06, APS05, DM07a, HGN06, IR06, JSdM07, JCM05, KMN09, Kos07, LWL07a, Lin07f, Liu06c, Liu06d, MR06c, MH08b, PK05b, Par09b, QW06, RKS08, RS06, SD06, SM09d, TT06a, TT06b, TT07, WLT07b, XXF06, YX06b, Yu05]. Cost-benefit [AEFFA08]. costs [AK05d, CJK07, ER06b, GBG07, Hu07a, JM06a, SYL07a, SYL07b]. Cot [All05b]. cotangent [Cvi09b]. Cotes [Ade07, DMJE05a, DMJE05c, DMJE06a, EDMJ05c, HEDMJ06, HMJED06, Ser05, Sim09b]. coth [SS08c, Waz07e, Waz08c, Waz08d, Waz08u, Waz08w]. COTS [Shy06]. Couette [Att06a]. count [AN07, CS07d]. countable [NS09, YLK09]. countably [LZ08d]. Counterexamples [YSZG05, Smi06b]. counterterrorism [Bel05]. counting [DFGV07, DFG07, WLC07a, Zho08a]. couple [Sel06, SA06e]. coupled [AKARA08, An05b, Bab07, BUP07, Bob08, BB06d, CTT08, CY06a, CLZ^{+}07, CA08b, CA08a, DG07a, DHS07, DH08, Den08b, FqZ08, Inc06b, Ism08a, Iva08, JZ08a, KY05a, KSM09, Kum09a, LMT^{+}09, Liu06g, LHF07, Muk07b, Muk08b, Muk08a, ND07, RKOK09, RN09, Ray06, SMY08, SXX05, SMH07, WsZ07a, WSZ07b, WZ08c, WZC08, Waz07c, WP06, XCM07, XXTL09, ZSHB06, ZS06, ZQ08, ROK09]. coupling [Ahm07b, GRH07, He06, Lin07b, MX08a, SMG07, YZ07c, YZ07d, YL09c]. couplings [SZ08c, YX08, YZ08, YL09b, Yu09]. coupon [SS06e]. Cournot [DHT09]. course [Mir06a]. Covariance [AR05, NCÁHCLP05a, NGLHCLP06, NHCLP07]. covariances [HCLPJL^{+}08]. cover [Cas07a, dCDLN^{+}07]. coverage [WC06a]. covolume [HJK08, YJ09]. Cox [SLO08]. CPSO [JCSR07]. crack [Li05f, Li05g]. cracked [LD06a, MRN09]. cracks [ASL07, CLW09b, LF08, Shi09a]. Cramer [GW06a, WX05]. Crank [GR06b, HXH08, KA06a, Sid09]. crashing [WLT07b]. create [Zha07a]. credibility [ZL06b]. credit [HY07a, HTO06c]. creeping [AuH07]. Creutz [MB05a]. criteria [ASA06a, AH07b, CL09a, CT05c, CLWP09, EHP08, GZ07d, HBN06, JRA^{+}07, KMÖS09, KY06, KWP07, KPL08c, Li07a, LS09c, LM08c, LX09b, LW09e, MMAHN08, MSTY06, MD05, MX07a, MX07b, Oht08, QCSF09, dCRDSSN^{+}08, Sin09, Sit06, SM08b, TMRV06, WZM06, XZ08a, XM06, XM07a, XM07b, YE07, YX09, YT06, ZY08a, ZYT09, ZM08c, lS07b]. criterion [HL07a, Kök06, KC09, KP08c, KP08b, KPL08b, LH07c, Nam08, Nam09, Par05b, PK05c, Par06a, PK08c, PPKL08, RC06, SZ06a, WZ07a, WZL08a, WS07f, ZZW09b]. criterions [XS09]. critic [DF08]. Critical [Dem05b, ML09, XCM07, AJKM09, ALN06, DJJL06, DLT08, Efi09, ESA05, ESAR05, GD06, Kan09b, SZ06a, SP06d, WCZ09, WY08, Yan08d, ZLW09a]. critique [Nwo06c, Nwo06d, Nwo07a]. cross [Muk07b, Muk08b, PWY08, WYL06d, Yin07]. cross-coupled [Muk07b, Muk08b]. cross-diffusion [PWY08]. cross-region [WYL06d]. crossover [DT07a, KTW08]. crosswise [JSL05]. Crouzeix [SP09]. Cryptanalysis [BLH06, CYY05, CKY05, Hsu05a, JmBdXgXm05, KRY05, KHKL05, LL05b, LW05c, PKH05, QCB05a, TLH05, YW05, YRY05a, YY05b, ZK05, ZF05, LMSV07, SCS05b]. cryptographic [ÁCTZ05]. cryptography [DwWmW05]. cryptosystem [Che05g, CFVZ06, SCS05a, SLC05]. cryptosystems [PLJ05b, Sha05c, TC05, Tsa05]. crystal [Prá08a, Prá08b]. crystallization [Ram07h]. CS [BR08b]. CSII [NAM^{+}07]. CSPs [SB06a]. CTL [ZSZS09]. cubature [SV06]. cube [Sil06g, Sil06f]. Cubes [Sil06h, CCWH09, HS07a, Sil05c, Sil05a, Sil05d, Sil06c, Sil06a, YMCL06]. Cubic [ASNR06, KR05, KB06a, Kan07, RMG07b, AV07a, Beh06, Boy06a, CS09c, Chu07f, DIS05, DN06, DB07b, DMR07, DWT05b, FGfFL06, GK06c, KLW06e, MY06b, NNMDS06, NN07e, RMJG07, RW09, SDL09, SA07a, SA07c, Sil05a, Sil06h, SMQ06, ZZH06, ZL06a, ZXLZ07, ZX08, ZW09c, Sil05b]. cubic-spline [RMJG07]. cultivation [MM07c]. cultivations [LBZY08]. cultures [GFWX05, WFX07, YFLX09]. cumulants [NO09]. cumulative [Nwo06c]. current [KK05a, Ma06b, Ma08a, SdJA06]. currents [GS08b]. Curtailing [SG08b]. Curtis [KKY07b]. Curvature [LM08a, Sil06c, WC05b, Yu08b]. curve [BK07c, Che05g, CFVZ06, DwWmW05, Hsu05a, HLS05, HL05d, Jia07a, JS05, KNM05, Lin09b, SK08b, SS05c, SS05d, SCL05b, SLC05, TC05, WZ06a, ZL05b, ŽMO09]. curved [CLW09b, dCDNPN09]. curvelet [Ma08c]. curves [CJL05a, DWT05a, HH05, KYA05, LWZH05, MMS^{+}06, MMS^{+}08, Nar07b, RLY06, Sha07c, UP07c, XCM07]. curvilinear [AA09, Bri08, Özt05, SFH09, WZ09c]. Customer [EG07c, LT09, Oss05a]. customers [Cha09, MC07]. cut [Ghi06, HMW07, TT06a]. cut-peak [HMW07]. cuts [Lin07f]. cutting [CLW09a, Kas07]. cutting-plane [Kas07]. CWENO [CFX05]. CWENO-type [CFX05]. cybermall [MC07]. cybernetic [LBZY08]. Cycle [HS07a, Li05e, AKH06, AS07b, CY08b, CHW07, CL08e, KLL^{+}08, MG09, SD06]. cycle-embedding [CY08b]. cycles [AM06, AGR09, Bay05b, Cas07b, LM09, Ram08a, ZZH06, ZL06a, ZZW09a, Zho08a, ZXZZ07, ZXLZ07, ZX08]. cyclic [BK09, DH06b, GL06a, HS07b, Li08e, OS09]. cylinder [Lee05c, LWH06b, LLW05, MP08a, SG05c, YC06c, ZM06a]. cylinders [KS08d, KS08c, Rad05a]. cylindrical [CKS09, FSÇ06, KS08f, Mak07c, ÖŞ06].

D [AS07b, AM08d, ALSL08, AKM06, Bra05, CGS08, DMA07b, DZ09, Dön06b, HBE09, Hlo05, HC05a, Hsu06, KK07c, KS05a, KSM09, KM07, Lee05c, LM08a, LLFM08, LX06a, LX06b, Ma06b, Moh07a, Moh07b, Moh09, Ohm08, QZ09, Ram07d, Ram07e, SDL09, SP09, TD07a, WZZ06a, WCZ09, XL06b, ZW05g, ZH05]. D-C [HC05a]. D-integral [KK07c]. D. [Ren08a]. DAE [SJ06d, Wan06c]. DAEs [Bar05b, Hos05d, Wan05g]. d’Alembert [Prá08a]. dam [AABD07, WGCH05]. dam-detouring [WGCH05]. damage [NKR^{+}08, QSS08, Yan09b]. damped [EHP08, GM08, LM07b, RN09, WC05a, WC09c]. dampers [EB05e]. Damping [Ram06b, AKBZ08, ÇT07a, HM08b, HZ09b, KL05b, LM08c, MM07e, SQ11, SM05d, SM06e, Sun06b, WF07b, XZ08a, ZMM06, Zha06c, ZM06b]. dampness [JH09]. dams [HBBA07, WGCH05]. Daphnia [PW09]. Darboux [Bel07b, QLG08, WZF08]. Darcian [Pan08b, AEA05a]. Darcy [Che06n, JK07, MSH08]. dark [Cai06b]. Darvishi [DJ06b, DKK06, Jav06a, Jav06c]. Data [Ami06, AKA07, HJL08, KR07c, Li06a, LT07, SMD06, AR05, AH05b, AK05g, AJK05, AKS06b, Bak05, BBCP09, BP06a, CL05b, CHY^{+}09, DX07b, DRS^{+}08, EH06, ET05b, Fan06, GÜ07b, HK05, Har06, HH07b, JMHS05, JLZ05b, JLI06a, JLS06, JLI06b, JPZ06, JLBR07, JLBR08, JWL^{+}07, Jön05, Kho07, KR07d, LJK07, MK05b, qMpXjH09, NK07b, PQJ08, PG08c, RLdLSF07, RGA05, Sae05, SML05b, SML05a, Sae06d, Sae06e, SCL^{+}05a, SL07b, Sil06d, Sil07b, TARm08, Wan06a, WL06f, WLX05, WL05g, WWLX06, WL08c, WYL06d, WLL07b]. databases [CTL09]. datasets [XCW08]. date [Wan06d, ZLZ06]. dates [Ere09a]. Datum [Men08, Sil06d]. Daubechies [AM08b, LKC05, MD06a, MYN07]. Dauhoo [Ren08a]. Davey [GC09]. Davies [EMH07]. Dawson [Boy08a]. DCT [SLLL07]. DDEs [KPS06a, LWG09, Sha08b]. DDM [ZT05]. de-lighting [JCX^{+}07]. De-Novo [CH06b]. DEA [HV08, AK05d, Ali05, AA07b, AA07c, AK05e, AK05f, AS06f, AKR06, AS06g, Ami07, AGI08, FG09, HK05, JVFM04, JLM05, JSdRm05, JLZ05a, JMLR05, JLS^{+}05b, JLS^{+}05c, JHS^{+}05, JLS^{+}05d, MA05f, SM05a, SMD06, WL06e, Wu06a]. dea-based [AK05d, AS06f, AKR06]. DEA-like [AK05e]. DeAngelis [DK05, LH07a, LC09b, XW07b]. deblurring [FL06, SCX07, WM09]. decagonal [XFTY09]. Decay [FFKV08, Gor06a, KM09a, PK08a]. Decaying [WC08b]. dechlorination [Slo07]. deciding [WW06]. Decision [Nwo06a, Nwo07a, AJK05, AS06g, AGA^{+}07, Bog05d, CL06b, EMP08, HKH06, II07, JLI06a, JLI06b, JPZ06, Lin06b, LG07, MMAHN08, PDGR07, PG07a, RR06, SML05b, Sae06b, Sae06c, Sae06d, WL06e, WF07a, YT06]. decision-based [Lin06b, LG07]. Decision-making [Nwo06a, Nwo07a, JLI06a, JLI06b, RR06]. decisions [Nwo06a, SA08b]. decomposable [Dje09, LLX07b]. decomposing [ASAS06]. Decomposition [Ham05a, HT07, MMN06b, MS06c, Abb05, AD05b, Abb06a, AEJ06, Abb06c, Abb06d, AE07a, AKM05, All05a, All07a, AH07d, AZ07, AA07f, Att05b, BDDB08, BD05a, BVC05, BR08b, BA05a, BE05, Bia05, BTI06, Bia06, BS07b, BB05, BKBK06, Bog08, Bou06b, BEF08, lCS06, CY06b, DGJ07, DGB08, DG07b, DH06e, Du05, Dua06, EG07a, EWAE06, EWEA06, EA06, GK05a, GMCM06, Gor06b, GW07, Ham05c, HNB06, HM07, Hos06b, Hos06a, HN06b, HN07a, Inc05, Inc06b, Inc06c, Inc06a, IRSR05, JDG06a, JDG06b, JDG06c, JDG06d, Jan07a, JL08, JL05, JM06b, JM06c, Jun09a, KY05a, KI05b, KES05, KW05b, Lay08, Lee06a, LMZ08a, Les06, LZ06b, Luo05b, MAD08, MCAH05, Mak07a, Mak07b, Meš07, MY07b, MGMC05, MMC06, Mom05b, MAK05, MO06, NPGW09, Odi06b, Odi06c, Pam05a, Pam05b, Pam06]. decomposition [QS09, RN09, RB05a, RB05b, RCB06, RB06, Ray06, RCB08, Ray08, SB06b, SPX07, SC05b, SB05b, Sun06a, Sya05a, TD07b, Wan06e, Wan06f, Waz05b, Waz05a, Waz05c, Waz06d, Waz06i, Waz06h, Waz06k, Yan05a, Yan05c, YL08b, Yu07, (Ag06, ZWL06, ZLW06a, ZZYY08, ZW05i]. decompositions [WS08c]. deconvolution [LS09b, Mal06, MM05d]. Deconvolving [DRS^{+}08]. decoupled [ANK05, DG07a, Ma08a]. decreasing [MeT05]. decrementing [MX08b]. defect [EHL08]. defect-correction [EHL08]. defective [CL06b, IMB09, QL06, yZsZ06, yZ07e]. defects [AN07]. defence [JMC07]. defensive [SSEA06]. deficient [CC06c, DKA06b, Du05, Pan08a, ZW05h, ZW06d, ZW06b]. deficient-basis [Pan08a]. Definability [Coo06]. defined [AJMR07, Aou08b, Boy05, BO06, GOPP07, Noo06a, OP09, RSSS07, Sav07, Sok07a, YL08a]. definite [AJE05, AD07, BI08b, GY07b, Hai07, Hai08, Li07a, LCS09, MHA06, REDES05, YK06b, ZH08c]. definiteness [HZ09a]. definition [AHS06, Chi06b, SZL05, dCD03]. deflections [YWJ07]. deformable [KM07]. deformation [Lee05b, SA06d, Sel07]. deformed [HZ06a, WZ06h, WZ07d]. Degasperis [HRLC08, MYG08, SGC09, SX06, Waz07f]. degenerate [Bar09, BDLS05, CC06b, CY08d, GD06, LCM08, MMN06a, MTL08, PK08a, PY08a, TM08, Wan08a, WG07b, XCM07, ZY08b, ZLW09b]. degradation [EH06, JC07]. degraded [GCSS05]. Degree [WCO08, CC08b, CO09, DJ06a, KSP08, LA07a, MCS08, RLY06, SA06g, SA07e, WCO09, XB06]. dehydratation [BM05a]. Delaminating [LWWS09]. delamination [FK09]. Delaunay [CGS08, MCGS05]. Delay [GZG08, KP09a, LWG09, LH07a, LLZ08, MZW08, RB08a, SWX06b, SWX07b, TZX07, ZWY07, AO08, AOS08, Ahm09b, AA08, AST08, BB06a, BC09b, BI08a, BI08b, CZ07a, CZH08, Che05c, CO06, Che06b, CC07b, CZF09, CLWP09, DXH07a, DHG09, FT07, FK07, GZ07d, GZG09, HHP09, HHW09, Hua08b, HL05b, HL07d, IM07, IMB08, JSL06a, JSL06b, JZW08, KR07a, KR07b, KK08a, KS08a, KMÖS09, KeTL08, Kov07, KPL08a, KP08c, KP08b, KPL08c, LH05, LRMS06, LL05d, LX09a, LZdC06, Lin07a, Lin09a, LC05c, LW06d, LKU07, LL08f, LMC08, LKU09, LW09c, LWKU09, LC08a, MM07a, hMzLfW^{+}08, hMfWzLpZ09, MG09, MI08, MPS08, Mat09, zMsCxL08, MS07e, Nam09, Par05a, PK05c, PK05d, PK08c, PK09, Par09b, PS06b, PCCHLG08, QD08, QCSF09, RX09a, RX09b, RHY06, RZ07c, Rös07, Sak05, SL09a, SLC09, SFFY08, SS05g, SC07d, SM09c, TFX07, TG08a, TQC09, TTG07, Wan06k]. delay [sWX07a, WL07b, WZ07a, Wan07f, WS07a, WS07d, sWMX07, WLZ08, WWL08a, WC09b, WQL09, Wei09, WL05f, WWY08, WWY09, WS08b, WS09c, WZY07, WRL06, WZM06, XwWL07, XW09, XXTL09, Xu06, XM06, XZ07c, XM07b, XZ08b, XS09, YSZG05, Yan09d, YX09, YCC08, YS06d, YWL07a, YWL07b, YYMZ09, ZZ08a, ZW07a, ZCZ07, ZY08a, ZG08a, ZL08a, Zha08g, ZQJ09, ZYT09, ZXL05, ZXDL05, ZZW09b, ZSS08a, ZSS08b, ZXY06, dlS07a, lS08a]. Delay-dependent [GZG08, LWG09, LLZ08, MZW08, RB08a, SWX06b, SWX07b, TZX07, CZ07a, CZH08, CLWP09, KPL08a, KP08c, KP08b, PK08c, Par09b, ZZW09b]. delay-difference [LZdC06]. delay-differential [DXH07a, KPL08c, PK05c, PK05d]. delay-integro-differential [ZQJ09, ZXL05]. Delay-range-dependent [KP09a]. delayed [hBxY07, Che06h, CXC06, CS07f, FO09, FC06, GCT08, GH09, HCLP07b, JPYW09, JWS^{+}09, JPCL08, LmZ08c, LTC07, LFC08, LTCC08, NMP08, PW09, Par06c, PKL08a, PJWL08, RB08a, SB08a, Sin08b, Sin09, SHG07, XW07b, XCL09, YL06a, Yan07b, hYG05, YCZ06, ZWX05, ZLY08, ZW08d, ZL08d]. delaying [KL06b]. delays [Bai08, BR08a, CL09a, CZH07, CZ07a, Che05a, CLH06, Che06c, Che07c, Che07b, Che08d, CLWP09, GZG08, GTX08, GS07b, Guo09, HH09b, Hua08a, Ibr08, Kla08, KP06b, KP06c, KPL08a, KP08a, KP08b, KPL08b, KPL08c, KPL09b, KP09a, KP09b, Li07d, LmZ08b, LL08d, LZ09b, LYJ09, Li09c, LZC07b, LTZ09, LS07e, LLZ08, LZWK09, Mah09b, MXC07, MCC07, MGH08, NCÁHCLP05a, Nam08, Par05c, Par05b, Par06a, Par06b, Par07, PK08c, PKL08c, QC08, RB08a, RB08b, RB08c, RC06, SSA09, SA09, SP06b, SJW09, SWX06b, SWX07b, SM07d, TZX07, WZ08a, Wan05c, WML07, WS07b, WZL08a, Wan08f, WL09a, XCH05a, XCD05b, XCH05b, Xu08b, YLL^{+}05, YX06a, Yan06c, YZW07a, YZFL08, ZWX05, ZC05b, ZS07b, ZT07a, ZLT09, ZD06b, ZZW09b, ZY07, ZH08b, ZL09a, ZH07b, Zhu08, dlS06a, dlS06b, dlS07a, dlS07b, lS07b]. delegation [JY07b]. delivery [PN08]. delta [XPS07]. demand [GBG07, HZ06b, IR06, JTMRS06, Lee05a, LWH06a, LWL07a, Liu06c, PBG07, WS05a, WLT07b, YVU06]. demands [Chi08c, JS06, SA08b, TMJJS07]. Dengue [DB06, MMC^{+}08, YF08]. deniable [LC05e, YRY05d]. Denoising [Sha08d, LG06c, SCX07]. dense [dAC09]. densities [Efi09, Tak07a]. density [ALN06, Cai06b, CL08b, DM07b, HS09a]. dental [dCJLNÁA08]. departure [ETJ06]. depend [Dyk09, Ram07b]. dependence [Che07a, Fig08, YT06]. dependent [Ahm05a, AB05c, AAA08, AAA09, ANK05, Bad05, BOB05, CZ07a, CZH08, CS07f, Che08d, CLWP09, CG05b, Deh05a, DKL09, EM11, EESF04, EG06c, Ery06, GZG08, GBG07, HL05b, IR06, Jav06c, JTMRS06, KSA05, KR09, Kis08, KK08c, KKS06, KJ09, KPL08a, KP08c, KP08b, KP09a, LP07b, Li07d, LWG09, LCZ05a, LW06d, LLZ08, LL08f, LGw07, LC06d, Luo06, Luo07, LG08, MM07a, Ma06a, MW07, MZW08, Mak07b, Mon07, MBM07, PM09, PRM06, PLL08, Pan08c, PK08c, Par09b, PR06, Psi07, RB08a, RK08, SMM08, SMY08, SB08a, SCR08a, SH09a, SC06a, SP09, SL07b, SWX06b, SWX07b, SAS05, SSK06, TZX07, yWjL06, Waz05c, Xu08b, ZGH06, ZZ08a, Zen09, mZzH05, ZT07a, ZWZ05, ZWY07, ZZW09b, ZY07, ZL09a, ZJO08, dlS06a]. dependented [SFA08]. depending [DM08b, LS07e, MD06c]. deposit [CJC05, WZ06k]. depth [HKS09b]. Derivation [All07a, AA06d, Bay05b, CL08b, NGLHCLP06, COG08, VGMV^{+}07, YSdS^{+}06]. Derivative [Cvi09a, BP05, CA08b, Chu07j, Chu07l, CH08b, DM08b, Gee08, HR07a, HR09a, HMT09a, HMT09b, Ide06, Kan09a, KLW06d, KLW06c, KLW07d, KL07f, Kou07d, LLC09a, Lin09b, LL08f, LM07a, MC06, Moh07c, MMR09, yNyF05, PL06b, Pog07b, Ram08c, RCB06, Ren08b, RA09, Sok08b, Wel08, WX06, Xia08, XPS07, YM06, ZWY07, ZZYT08, dEdSNL05]. derivative-free [LM07a, yNyF05, PL06b, Ram08c, WX06]. derivatives [CLA08, CM08a, CA08a, FH07, GBG08, HED06d, HR09b, HY07a, Iri06, JYX06, Li05b, LD07a, LKC05, McC05, McC06, MA06e, NG07, NKH07, Noo07e, NNuH07, Odi06a, Odi08, Odi09, QFXW06, QFF06, Raf07a, RMG07b, SS09c, TF07, WL06f, WN08c, WR06, Yoo05, yZsZ06, Zha08f]. derived [Çak08, Fel05, Gon07, GSG09, Kwo07a]. descendant [HH07a]. Descent [CHLS08, GH07, AJ06a, AJ06b, AE07c, BNK07, Chi07b, HWS07, Jia05, SS05b, SS05c, YNCL08, YZW07b, Zho09a]. descent-projection [BNK07]. describe [CC08d, LBZY08]. describing [Efi09, KM06, Liu06g]. description [HA07, Wey06, Wey08, ZZX05]. Design [BNdFPS06, CLS05a, GZG09, NCÁHCLP05b, NHCLP07, PK08b, TMJJS07, CLS05b, Chi07e, Chi07d, Chi07g, DD09, GRZ06, Gee08, GHZS07, GCJ08, HBJ06, Kök06, LFM06, LY05, LCC05a, LML08, Li09e, MFE05, Mor08a, MH08b, Par05c, Par09b, fQZgH08, Sad07, SCR08c, SCQ06, SM09a, WLLL07, WLP07, YCC08, ZLZG06, Zot07b]. Designated [LC05d, GT05a]. designed [nXwCpC06]. Designing [LFM06, MJFAS07, MMAHN08, TCDV07]. designs [MWW05]. desirability [II07, PN06]. desirable [KA05c]. desired [CW08]. desired-speed [CW08]. desorption [Hua05a]. Despeckling [GZL08]. Detecting [XCW08, CH07f]. Detection [DM05, PG07b, WML05, ZXZL07, CYL^{+}08, Fuk06a, Gop07, IZIO09, KFM08, LLX07a, SD06, TX08, Yan09b]. detector [GAS08]. deteriorating [CO06, JTMRS06, WNCL06, lZyT05]. determinant [HE08b, Kop22, LHL08, RR07, Sog07, Sog08a, Sog08c]. Determinantal [LYW09, Thu06]. determinants [EM08, Kar05b, KEM08, Kwo07b, Li07b, Li09d, SG09b, WS07h, ZWS08]. Determination [DKA06a, HED06a, KB08, Liu05a, MM07c, SCL^{+}05a, ADH05, AEFFA08, Dan08, DS05a, ESS05a, Fat05a, Fat05b, GJD^{+}07, JL09, QC07, Ram06a, Ram08b, Sha08a, TCMPMC07, TK06b, TK06c]. determine [BOB05, BP06b]. determined [LZ09e]. Determining [Çet06, CFF09, HZ05, SML05b, Bar05a, Hu07a, HW05c, SSB06, TL07, WL06g, XFL06]. Deterministic [RSD07, CP05, JZQ07, JC08b, MA05e, NDD05, Tie05, WL05e]. detouring [WGCH05]. developed [AH05c]. Developing [FEOC08, Sae05, ZGK06, CHW07]. Development [GTJ07, Gui05, SLX^{+}05b, SH07f, Thu07a, Kos05, KK05c, SJ08a, SGA06, Thu07b, yZ07e]. deviated [Dyk09]. deviating [Che06f, Jan09a, MX07a, MX07b, ZM08c]. deviation [Cha05c, WL06g, Zha07e]. deviations [CGSS09]. devices [Ouj06]. devil’s [Sum07]. devil’s-staircase-like [Sum07]. DFT [Ant05, Che08b, YW09]. DGMRES [ZW06e]. diabetes [CMMC09, NAM^{+}07]. diagnosis [FTG07, Mea05]. diagnostics [VRG06]. diagonal [BK09, CN09c, DZ06, EM05a, HL05a, Kar06, KLR07, LL08g, Mat09, NA08, Rim08b, Rim09b, WD06, WYL06b]. diagonal-Schur [CN09c]. diagonalization [IMB09]. diagonally [AKIS07, DH06a, HZ06d, LCH09, WWL08a, YY06c]. diagonals [GG08e, GG08f, Kil08b, Rim07g, Rim07h, Rim07d, Rim07f, Rim07e, Rim08b, Rim09b, WS07h]. diagram [WZSL08]. diagrams [FS09]. diameter [Cha05g, dCDLN^{+}07]. diamond [SY09]. dian [CS06b]. dielectric [KS08d, KS08c, Sel05]. difference [AKO06, Alo06a, Alo06b, Alo09, ACJ09, Ami05, AKD06, AŞ06e, AR07a, AV07b, AST08, AO06e, AO06d, AY05b, AA06e, Asl09a, AR07b, Bah05, BS05a, Bie08b, BG07b, Bra05, CCE06, CA05, Cen05, CZZ05, CLSZ06, CW06a, mCLB08, CL08d, lCX06, DSMK08, DDD05, Deh05b, DMS06a, DMS06b, DMA07b, Dem05a, DZ09, Djo06, DM08c, DDR05b, EOM03, EOM08, Elb05, EEEE05, EO06, EI09, Ere06, FZ09, GC07a, GS05, GSG06, Gül06, HED06a, HSH05, HY05, HH08a, HL07c, HXH08, IS09b, IS09a, IMS09, JKN09, Jer08, JOA05, JV05, KM06, KSA05, KA06b, KP06a, KYK08, KK08a, KS08a, KRA07, KV05b, Kum03, KA06e, Kum09c, Lei09, LS05, Li05b, Li08a, LZdC06, Lia08, LLAT09, LZA^{+}07, LKU09, MTL07, MM08a, MMW09, MMO07, MA08d, MSD06a, MSD06c, MY07b, MV09b, Moh05a]. difference [Moh05b, Moh07c, Mou08, MP08c, NDD05, OAAK08, Özb07a, Ozb07b, OBO06, PVV09, PFG08, Pan12, PS06b, Pen06, RK08, Ram06b, Ram06c, Ram06d, Ram07d, SF07, SS05a, Sal07b, SA05a, SA06b, SAB06, SA06c, Sal06e, SG09a, SG05a, SAD06, SG07a, SS08b, SYH07, SP05, SM07a, SFH09, SLL08, Ste09b, SL05b, Sun05b, SS05g, SS07e, SM08c, SW09d, TD06, TD07a, Wan05a, WZZ06a, WTZ06, WL07e, WTW08, WS06a, XFL06, XF06, XZ08b, YS06a, YLZ06, Yan05d, YEM05, YLB05, YLL06a, YEM06, YYL07, YY08a, rYzS07, YJW06, YWJ07, Zha05a, Zha05b, ZZ05a, ZZ05b, ZYME06, Zha09e, ZZKY05, ZZC06, ZC06c, ZZ06a, ZZ08b, ZXY06, ZYZ06, lS08b]. difference-differential [ZZ06a]. difference-integral [Jer08]. Differences [SW09c, Gal06a, Gül05, RAP05, RYW06]. Different [HCLP07a, Vaj07a, AH06b, AuH07, AS09b, BKBK06, BKM06, Dön06b, ESES06, HPD07, HN07b, Kay05a, KAP06, NMM08, QL06, ST06c, TS08a, TH07b, YY08b, Zha07f, Zhu09a]. differentiability [GM06a, Taw07]. Differentiable [Noo08b, Nar07a, Shi09b, Arg09]. Differential [LS07d, SS07b, ST08, Abb06a, Abb06b, AC05, Ade07, AJKM09, AMN06, AO08, AO05a, AD06a, AS08, Ahm09b, AVM^{+}09, AD06b, AGR09, AKB09b, AŞ06e, AAK07, Ang06, AE07b, Aou08b, AGPJV09, AHIP07, AO05d, AO06e, AO06d, AKM06, AA07f, Asl09a, Att06c, ADDdM07, ANK05, AKR05, BVC05, BF07b, BR07a, BM06a, BUP07, BQ09, BSKK06, BK08b, BC09b, BSN^{+}08, BDG08, BD05b, BM08, Bia05, BG07a, BB05, BKBK06, BRRGT07, BRT07, BGF09, BT08, BSHZ07, Boy08a, Bri08, BSKS07, Bul09, CH07a, CC08b, CL09a, CD05, CC08c, CC09c, CT05a, CS05b, CCJR07, CW07, Che08b, CE07b, CE07a, DGJ07, DX07a, DE07a, DSZ07, DMK07, Deh05a, DBC^{+}07, DXH07a, DT08, DMH05, DXH07b, DB07d, DZ07b, DB09, DFGV07, Dyk09, ERAS07, ER06a, EBsENMA05, EB08, EHZES05, EMESES05, ET05a]. differential [ETETH05, ETETH07, EWEA06, EMH07, FO09, Fel05, Fra09b, Fro06, GJ07a, GKB07a, GkB07b, GKD07, GF06, GJ07b, GS07a, GCJ08, GBC09, GAPT08, GbL07, GS07b, GG07, Gül08, GS07c, GR06b, GB06a, GB06b, HW08b, HES05, HAM04, HBVM07, HIPA08, Hos05b, Hos05c, Hos06a, Hos06c, HN07a, HRS07, HW07b, HLW08, HHW09, zHWH07, Hua08b, Hus09, Inc07e, Inc07b, IM07, IMB08, JRW08, JDG06c, JKN09, Jan08a, Jan08c, Jan08b, Jan09b, Jav06c, JWPS08, Jia07b, JSLL07, Jia09, JZX07b, JZX07a, Jin08, JJ09b, JR07, KR07a, KR07b, KK08a, KS08a, KGB06, KMÖS09, KKA09b, KRRGT07, Kim07b, KM05c, Kıy06, KSKK06, Kös06a, KL08, Kum09a, Kuo05, KP08c, KPL08c, Lak05, Lar05, Lep06, Lep08, Li05b, LLZ06, LS06c, LCZT08, LYX08, LZ08c, Li08b, Li08c, LC09a, LS07c, LZ08e, LS06e, LW06d, Liu06h, LS07e, LKU07]. differential [LLW07, LMY07, LLY07, LM08c, LL08f, Liu09c, LW09c, LLW09, LWKU09, LP06b, LCM07, LZ09e, LW09e, LM07b, LYL07, MM07a, Ma07a, Ma07b, MCAH05, MPEL09, MS07a, MB06, MA06c, MSTY06, MRN09, Mat09, MX07a, MX07b, MM07e, MSHeT07, MSR08, MA05g, Moh07a, Moh07b, MAD07, MAK05, MS06c, MN06, Mom06, MD08c, Mus07, MT09, NMM08, NL05, OP09, OME08, OAHR09, Ozb07b, PS08, PPF09, Par05a, PK05c, PK05d, PS06b, Pod06, PCS07, PG07b, PMRAS05a, PMRAS05b, QBH06a, QBH06b, Ql08, RAS07, RJW07, RK08, Ram05d, Ram05f, Ram05g, Ram06d, Ram07d, Ram07e, Ram07g, Ram08c, Ram08f, Ram08d, Ram08e, Ram09a, RZ07b, RB08c, Raw06a, RB05b, RX09a, RX09b, RHY06, RZ07c, SKHZ06, Sai07, SA06a, Sal08c, SG06, SGT06, SAD06, Sha05a, SR06, SK06b, SS08b, SL07a, SL09a, SLC09, SI06, Sol05b, Sol05c]. differential [Sol05d, SHG07, SLW08, SLL08, Sta06a, SM08a, SM05d, SJ05, SS06f, SM06e, Sun06b, SHD07, SM07c, SM08b, SM09c, TL07, TFX07, TG08a, TJG08, Wan05b, WJ05, dWqL05, Wan06g, WL06b, WZSZ06, WL06a, WC07a, sWX07a, WL07a, WZ07b, WML07, WL07c, WLT07a, sWMX07, WLS08, WWL08a, WWL08b, WH08b, WHC08, WQL09, Waz06i, WZ08e, WZ08d, Wei09, WL05f, WYL06b, WYL06c, WWY08, WWYL08, WWY09, WS08b, WS09c, WZY07, WL08c, WCW07b, Wu07c, Wu09, XZ08a, qXbL05, XM06, XM07a, XM07b, Xu08b, XZ09b, Xu09, YAO09, YSS09, YC06a, YLB06, YS07, YZ07b, YX09, YBA05, YJW06, YWJ07, YL06c, YC06d, You06a, YWL07a, YWL07b, (Ag07, ZG09, ZZ08a, ZSHB06, ZZL05, ZT05, ZLC06, ZT06, Zha06b, ZLW07, ZFY07, ZCMQ07, ZZDL07, Zha07g, ZZYY08, ZG08a, ZLW08, ZM09, Zha09e, ZQJ09, ZXL05, ZXDL05, ZC06c, ZZYZ06a, ZC08b, ZZ06a, ZWZZ07, ZLC07, Zho08b, Zho09c, ZL09b]. differential [ZW05j, ZJO08]. differential-algebraic [ANK05, CL09a, GJ07a, GB06b, Hos05b, Jan08a, Jan09b, LS06e, ZXDL05]. differential-difference [Asl09a, SS08b]. differentiated [MTD05]. Differentiating [Fuk06b]. differentiation [Asa06b, Asa07b, CL08c, IOS07, JJ09a, OISM07, Rao06a, Ste09c, Ste09i, Yan08c]. differintegral [LP08b]. difficulties [AA07c]. Diffie [YRY05d]. Diffraction [Ahm07b, Ahm07a, ARM09, KS08c, SSR05]. diffuse [Kim05b, PSD07]. diffuse-interface [Kim05b]. diffuser [CLNCG07]. diffusion [ANM07c, Ahm06, AKM05, AL08, Ang05, AEA07, Bab07, BW05a, BG06, BP09, Baz08, BHN08, Bi06, Bog05a, BP06a, Bog08, Bra06, Cen05, Çet06, CX06, CS07f, CSC08, CG05b, DGB08, DIT06, DG07a, DCS07, DKL09, EA07a, EA07b, EG06a, ESS05c, EWAE06, EESF04, FM05, FK06a, GD08, GY06, GGSNR06, GGSN06, Has06c, HZS07, HY07a, JDG06d, JW07, JCM05, KM06, KY08, KS05a, Kis08, KV05b, Lay09b, LS06a, LS08c, LZ07b, Li09f, LXJ09, LWW09, LLAT09, Lin05f, LG06c, LZA^{+}07, LLZ09, LGw07, Luo07, MPG07, MR05a, MV08, MMR09, ND07, Odi06b, Pan08c, PWY08, PD09, qQcM05, Ram05b, Ram06b, Ram06e, Ram06f, Ram07c, Ram09b, RB06, RCB08, Ray08, Rem07, Sal06e, SDG05, SL07b, SFA08, Sid09, SM07b, STU09, TD06, TD07a, Vyn07, WZ08a, WiN06, Wan08f, Wan09b, XO07, XCD06, Xu06]. diffusion [Yan07b, ZH07a, Zha08h, ZW08a, ZWZ05, ZQ06, ZS06, ZSS08a, Zhu08]. diffusion-based [LS08c]. diffusion-thermo [EESF04, Pan08c]. diffusion-wave [AKM05, Odi06b]. diffusive [ANM07b, AMN07c, BW07, Che07i, Che09b, DFGV07, KS05b, Sak05, SCB07, ZW07a]. digit [LLWC07, SA08c, YGL05]. Digital [CCH05, Gop07, BLH06, BBMN07, Che07h, GRZ06, HHC05, LLC06b, PC05, PLJ05a, QCB05b, Sha05d, SCL05b, TTG07, XC05]. digitalized [Imr06, UP07c]. digraphs [CLZ05]. Dijkstra [XLH^{+}07]. dilation [Kar09b, LFW08, SH06b, SY07c]. dilemma [AHE05, DJJL06, JLS06]. dimension [Bad05, CG06, GRH07, Imr06, Imr09, KA06b, LS08e, Lin05e, LZ09d, MMW09, NM09, PU08, SC08a, UP07b, UP07c, Zha05a, ZLZZ06]. dimension-splitting [Lin05e]. dimensional [Akm09, AST08, ASA07a, Ari07, Att06b, BAF08, BS05a, BE06, BB05, Bra06, Cel06a, CC08c, CC09c, CW05a, CW05b, CZ08b, CG08a, CFJ^{+}05, CIS09, CG07c, CCZS08, CC08d, DW09a, Dao05, DS05a, DG07b, Deh05b, DM06, DMA07a, DCS07, DC06b, DHG09, DKL09, EFE08, EB05d, EMY05, Eng06, Fen06, FHL07, FL09b, FHG07, GC07a, GHL07, GKD07, GJD^{+}07, Gor06a, GPT09, Gül05, HFL06, Has06c, mH07a, lH08a, HC09, HM06, Inc05, Inc06c, Ism06, IMS09, Jan07a, Jan09c, JZ08c, JZ06c, Jun09a, Kar05a, KY05a, Kay05b, KK08c, Kut05, LC06a, Li07c, LF08, LWWS09, LZ09a, LZ08d, LH08c, Lia08, LLC09b, LM09, LST09, LGw07, LH08d, Luo07, LG08, Ma07a, MJ07a, Ma07c, MWQ09, MD08a, ML08b, Mar05b, MFF09, MZ09, MB05a, MM08c, MV09b, Moh05a, Moh05b, MS06b, Moh07c, NKR^{+}08, OB09, OIL09, zPY06, QS09, Ram05c]. dimensional [Ram05b, Ram06b, Ram06f, Ram07c, RM05, RM06, Ras05c, RNV07a, Rem07, RM08, Sal07d, SK08b, SS08a, Shi09a, SLC09, SZ05, SFPE07, Sid09, Sil05c, Sil05a, Sil05d, Sil05b, Sil06g, Sil06c, Sil06d, Sil06a, Sil06f, Sil06b, Sil06e, Sil06h, Sil07b, Sil07a, SZ06b, SZ07d, SZ07e, Su09, SQG08, TWZ07, TB09a, TH07b, TT09, Tua13, TD05, WSLZ07, WLZ05, Wan06e, WCZ06, WsZ07a, WZ07c, WSZ07b, WCO09, Waz08h, Waz08l, Waz08i, Waz08r, Waz08t, Waz09a, Waz09c, Waz09d, WL06f, WS09b, Whi06, XZ05, XL05, Xu07, lYpZqL08, rYzS07, Yig08, Yu08b, YGY09, Yuj07, ZW09a, ZX06a, ZX06c, Zha07c, Zha07d, Zha08d, ZWT09c, Zha05e, ZZKY05, ZZL08, Zha09f, Zhi09, ZL07b, ZY07, ZLZ08, ZCZ08, ZLYL08, lS07b, Kis08]. Dimensionally [ASA07c]. dimensions [Eli07, MB05a, NLD08, PL06a, SY08a, Yan08e]. diminishing [KK06]. diode [DS05b]. Diophantine [Tze07]. Dirac [DW09d, Köp06, YZ08, YL09b]. Direct [Deo08, GS08c, LKC05, MC09, MP08b, SFG07, YC06d, BMS07, ES05b, LL05a, LWL06, MMW09, Waz08d, Waz08e, lYqLW07, YC09c, ZM06a, Chi06b, Chi06c]. directed [DES07b, Pud05, SP06d]. Direction [HL06a, Ars07d, GH07, Gw05, JZHT05, JZTH06, JXH06, LY07a, LY07b, NCGdSM07, SY07c, YHS09, Yan06a, ZCH07]. directional [Aki05a, GZL08]. directionally [XXTL09]. directions [bJyKxC09]. directly [JZCW05]. Dirichlet [ANM06, AK06c, AK06b, ANM07a, AMN07b, ANM07h, ANM07j, Cai09, DM06, DjT05a, KE05a, KS06b, Li09a, Li09b, Liu05a, Utk06, ZJO08]. Dirichlet-to-Neumann [Liu05a]. Dirichlet-type [Li09b]. disappearance [FA05b]. disaster [Cha09]. disc [AS09c]. discharge [SMC08]. discipline [BGP06, Dav06]. discontinuity [Cha07c]. Discontinuous [An09, KBH06a, BR07a, BUÇ07, Bel07b, BP06a, Boy06b, Bra05, Bra06, CH07a, Cen05, GH09, KBH06b, KK05a, Kim06b, Ozb07b, RAP05, Ras05c, Sea06, SY08b, SM09b, YKR06]. discount [CCL06, Liu07e, Liu09d, MBM06]. Discounted [JM06a, hB06, hBxY07, LS09a, NN06a, WSJW06]. discounts [AK08b, JS06, LWL07a]. discovery [Zho07]. discrepancy [SK06b, Su08b]. Discrete [Asl09a, BST06, CGW06, CZF09, Chr08, GMS06, IT06a, SMM08, XM08, ZWZ09, ACJ09, AM09b, BZ09, BR08a, BB06a, BEF08, CLH05, Che06e, CF06b, Che07c, Che07b, CC07b, CT09a, CK06c, FLO07, GBK06, GZG08, Gee08, GTJ07, HC08a, Hsu05a, HL05b, HL05d, ID08, Ibr08, Ide06, Iva08, JPYW09, JZ06a, JZ08c, KP08a, LHY05, LPP08, LP08a, LTH05, LL05f, hLsT07, Li07d, LZC07b, LZC07a, LOZ07, LTZ09, LZWK09, MZW08, MI08, Mah09b, MŘ07b, Mor08b, MH08b, NCÁHCLP05b, Nak07, NHCLP07, Nak08, Omr06, Par06a, PN08, PLJ05a, QCB05a, RK07a, RSD07, Sak05, SLLL07, SH09a, Sha05c, Sha05d, SP08, SFA08, SP06b, SFFY08, SCL05b, SLC05, SS05f, SS05h, SCS05c, SWX06b, SJ06d, SS06g, SWX07b, TT06b, TT07, Toh05, Wan05b, Wan09a, WW05c, XCH05a, XW07b, XT08, YB03, YB04, Yam05a, Yam05b, YZ09a, YM08a, Yu05, YZFL08, YJO08]. discrete [ZW05f, ZW06a, lZXL09, Zho05, ZL06c, ZLL06b, ZH07b, ZMV05]. discrete-ordinate [ZMV05]. Discrete-time [GMS06, ID08, Ibr08, Iva08, JPYW09, LP08a, hLsT07, LZWK09, MI08, Mah09b, NCÁHCLP05b, Nak07, NHCLP07, SWX06b, SWX07b, Toh05, YB03, YB04, Yam05a, Yam05b, YM08a, Yu05, ZW06a, Zho05, ZL06c, ZLL06b]. Discrete-velocity [BST06]. discretionary [LJE07b]. discretisation [DM06, SB05a]. discretisations [AAN05, GJ05]. discretization [AD05b, Aks05, AD06c, BC09a, BDCW06, Che05e, HIST06, LLC09b, cMpZfR06, MK06, Moh06, MS06b, MA06f, MD09b, MAD07, Raj08, Reb06, Sil07e, WZZ06a, Wan07a]. discretizations [Wan08c, Wan09a]. Discriminant [JLBR07, DY06, KR07c]. Discriminating [SH09b]. discrimination [SSB06]. Discussion [MNY07, Mon07]. Disease [MA05e, BK05a, GCSS05, MCC07, PG07d]. diseases [RYE07]. disintermediation [Nwo07b, Nwo07d]. Disjunctive [EKM08]. disk [Att05a, Dön06a, GOPP07, LL05a]. disks [KAS06a]. disparity [HLCH08]. dispersal [CH07c, LL08d, XCD05a, ZT07a]. Dispersion [Pat05a, BM09, FEOC08, KBMZ08, MXC07, Sha06a, Sha06b]. dispersion-managed [KBMZ08]. dispersive [CW05a, CW05b, Dem05d, Fen06, GL08c, Inc06c, Inc07c, Odi08, WCZ06, Waz05k, Waz06m, WF07c, YX08, ZW09a, ZLZZ06, ZL07b]. Displacement [QW05, LH07b, SY08a]. displayed [GBG07]. disretizations [CL08b]. dissimilar [Li05f, Li05g, TS05]. dissimilarity [BC09c]. dissipation [CB06, EO06, LeT05, Shu05]. Dissipative [Sha08b, DTC08, SA06e]. Dissipativity [WWY08, WWY09, TFX07, TG08a]. dissociation [TSS09]. distance [AJK05, IR08a, JH08c, PK08d, Sha06c, Su07a, PG08c]. distance-based [AJK05]. Distances [Rab05]. distilling [MQW05]. distinct [MR07c, Waz08r]. distinguished [Chi05]. distortion [Dzi09]. Distributed [KO08, SB06a, BR08a, BE06, Che06b, CH09, GZG08, xGyZtZ^{+}07, Guo09, KNJ05, KP08a, Li07d, Li09c, yLLjP06, LGHS06, MPG07, MGH08, NGLHCLP06, ONG08, Par06a, RB08b, Rig08, SP06b, Wan05c, WiN06, YLL^{+}05, Yan06b, YYWW07, YZFL08, ZLT09, ZM08c, ZH08b, ZH07b]. Distribution [LW06b, PLTW07, WCW07b, AJAH06, AÖ05b, AH06d, AA07d, BMJED06, Bak05, BCMT06, CL06b, Cha07e, Dam08, GN06c, GN06b, HZW06, Kim07a, KA06f, Lee05a, MK05b, Ma11, MS06a, MA06d, Nad05, Nad06b, Nad07b, NK05a, ONG08, PSW07, RIY05, SK06a, SMB08, SNK06, WL05a, Wan06a, WZY07, WWT05, WLC05, WY05c, WLC06, WWC06, WLC07b, WLC07c, WLC07d, WLL07a, XZ07a]. Distribution-based [PLTW07]. distribution-like [MK05b]. distributional [AL07]. distributions [AL07, GS06f, GN07, GGOS08, Haa09, HGN07, IT06a, LWH06a, MS08b, MR05b, Nad06a, NK07a, SA08d, SSR05, ST05c]. Divergence [BX09, OY05, ZH05]. diversity [Pal07]. divide [MCGS05]. divide-and-conquer [MCGS05]. divided [RYW06]. dividend [hB07, GY08]. divider [UP07c]. division [CL06a, Che08c, SGP07]. DL [PLJ05b]. DL-based [PLJ05b]. DLP [Qiu05]. DLP-based [Qiu05]. DMUs [AK06f, AA07b, AA07c, AK06h, JLZ05a, JLRB05]. DNA [DN07a, DN07b, HHC08, HC08b, LXH06, LD07b, LT07, QSS08, SS09b, WY05b, WXLH06, ZXL09]. DNA-based [HHC08, HC08b]. documents [MCAF08]. Dodd [Waz08k, XZY08]. Does [DD06a, Hub09a]. Dogleg [jW06a]. Domain [JM06b, ZZYY08, ZW05i, AE07a, Bog08, BD09a, Boy05, BO06, CSV07, lCS06, CGS08, DG07b, Dua06, GR06a, JL08, JM06c, Jun09a, Lee06a, MY07b, MGMC05, MMC06, PLZ07, RM05, RM06, SG06, Soh05, SLW08, SW09d, TD07b, Toh05, Vis08, VL09, Waz06h, Waz07k, YM06, ZL08c, dAC09]. domains [BHN08, DC06b, EHR09, FM08, Küç05b, Ma07a, MR07a, QM05, SGT06, ZLW08]. domestic [Chu05]. dominance [Fra09b, HL05a]. dominant [BK08a, CSY07, DH06a, HZ06d, LCH09, WSJW06, Wan08b, YY06c]. dominated [Bra05, EG06a, SW09a, SM09b]. dominating [AK08b]. dot [WLC07a]. Douady [PRA^{+}09]. Double [GÇ05, KS05b, Sjö07, SCB07, BQYW07, Che07i, Che09b, DMS09, Den09c, GLLC08, GHL07, Gra09, HT07, JSW05, dCJLNÁA08, Lee05b, Rad05a, Sav07, SC08a, TH07a, WC08a, Waz06o, Waz06q, WLCL06, XFTY09, YC06c]. double-base [WLCL06]. Double-diffusive [SCB07, Che09b]. double-layer [YC06c]. double-threaded [dCJLNÁA08]. double-trapdoor [JSW05]. doublecast [YTLH06]. doubly [CW05a, CY06a, Inc06b, KS05b, Nis07, WLA09, Yan05c]. doubly-periodic [CW05a]. downside [SLC07]. drag [ESS05a, KF07]. dragon [Mer07]. drainage [HM07]. draining [O’R08]. Drazin [CX08b, CGDR06, CGVC07, CX08a, CI09, Den08a, Den09a, LWW05, MSCG09, MD08b, PH09, QW05, SC09, VCRCG09, WG05, WWW03, WDN06, ZC06b]. Drinfel’d [Inc06b, WLS09, Waz09b]. driven [Dem05a, KAS06a, Wei09]. driving [SJ08a]. Droop [JB05]. droplet [ALD08]. drug [Kwo07a, TG06a]. drug-resistant [Kwo07a]. drugs [dG06]. drying [DSMK08, GRF07]. DSA [LHH05a]. Dual [ZWT09b, ZWT09a, ARK07, Ang05, Bis08, CF06a, DLS06d, GG09a, Ham05b, Ham05c, JM07, KAV05, KSKK06, Kös06a, LY09a, Lin07d, LWL07b, MY05, RAN07, RAN08a, WB09, ZTWL09]. dual-arm [LWL07b]. dual-power [Bis08]. dual-rate [DLS06d]. dual-reciprocity [Ang05]. dual-type [Lin07d]. Duality [AH06a, MAN06, Amm06, GM06b, Kas06c, Kas09, SG08c]. duality-bounds [SG08c]. Dubrovsky [FL09b, lH08a, Hon08b, SZ07c, ZX06b, Zhi09]. ducts [Özt05]. due [Ahm05a, Ahm07b, Ben07b, Cel06a, dCDNPN09, KR08, MP08a, PR09b, PR06, SK06b, VFF08a, Wan06d, ZLZ06]. Duffing [jF09, GLLC08, gLXyZ06, Ram07f, XXC05, XXC06, (Ag06]. Duggan [Tan07c]. Dullin [XXL09]. dunes [KMK07]. Dunkl [CG07a]. Dunkl-classical [CG07a]. duopoly [SM09d, YX06b]. duplicate [WY05a]. Durand [Zhu05c]. duration [Çet06, CL07a, KL05a]. during [SANN07]. Durrmeyer [GBG08, Gup08b, GF09, GLMLP09]. dust [MR09, SDS05, SSK06, SSS06]. dusty [Att06a, HS09a]. Dykstra [Mor05]. Dynamic [AS06c, MC07, SAB08, SH07e, XL06b, ZLT09, ZD06b, ASAS06, AH07a, Ami06, ATM06, BL06, Bar09, BS09a, BDCW06, BE06, BBCP09, Bog05c, BS07d, Buk05, Buk09, ÇFÇ06a, CH05b, CLH06, CHY^{+}09, Cor05b, DMDB08, DS06b, ER06a, EJER06, ET05b, EHP08, EBP09, EM09, EMH07, FG09, GZ07c, GTJ07, GL06b, GO08, HHP09, HTO06a, Jan09a, JL09, Kar09c, KO08, KMK07, KOS08, Kun08, KPL08b, KP09a, LHGT06, LCC05c, MPT08, MNM07, Nwo06a, PS06a, Par05c, PKL08b, Par08, PK08d, PLZ05, RCB06, Sit06, SLS08, SZ08b, SJ05, SJ06d, TMASA05, TS08b, Tie05, WD08, Whi06, WZM06, XXC08, XCJ09, YYYM07, YZD08, YQYL07, Yua09, ZW09b, ZD06a, ZZDL07, ZWD07, ZLR07, ZFT06, ZL08d, dlS07b, lS07a, BC09a]. Dynamical [BM06b, LYJ09, SB08a, WSCL07, ZSZS09, AH06c, BWB05, CLZ05, DLLN09, DX07b, Ewa06, Far05, GFWX05, GJL06, JV05, KV05b, LCC05a, LH08b, LDC09, MV09a, PNL05, SH07e, TTZ09, Vul07, WL07e, Woo05, YFLX09, YS06d, ZDL07]. Dynamics [Alo06a, Alo06b, CS05a, DDD05, DMS06b, DMS07, HLT09, Kov07, LS05, LCS08, LZ09c, MSD06a, MB05b, SAB06, Upa09, WL07d, WZSL08, WL08a, Zen09, ANM07c, AHE05, ABBP08, AAA09, Att06b, BR07b, BP06b, CTGD08, Chu05, CQL08, DZ08a, DHT09, DKL09, ESS05b, HHZ09, KWdW08, KV06, LMT^{+}09, LH05, LCZ05a, LWL07b, LBZY08, MR05a, MGH08, zMsCxL08, MC08, MD07b, NR08a, PS08, PR09a, PMS09, RS08b, SG08b, Sta06b, Sum07, TML07, VR08, WLP07, Wan08d, XFTY09, XCW08, XC06a, XX07, XMGW09, YX06b, ZYH08, ZSZS09, dSF05, lSAQ08, lS08b]. dysphagia [SGG^{+}09]. Dziok [AJMR07, Dzi07, HL08d, MA09, PS06c, RSSS07, Sok08c]. Dzurina [SM06d].

e-cash [CJC05]. e-commerce [AH07a]. E. [PG07c, Zim07]. E.F [Pan12]. each [Sal07c]. Earliness [IGM07, JRA^{+}07, OT06, TMMVA05, TMMVA06]. earliness/tardiness [OT06]. early [KH07a, LMT^{+}09]. Earth [ESS05a, ESS05b]. earthquakes [EMA06a]. EAS [Psi09]. EAS2 [Psi09]. EC’03 [ZSNS05]. ECA [EdR07]. ECC [Tsa05]. ECC-based [Tsa05]. ECG [GPR^{+}09, PG07b]. echelon [Rey05b]. eco [UBKC08, ZLY08]. eco-epidemiological [UBKC08, ZLY08]. ecological [Che05b, CS06a, CG06, HCX06, KWdW08, MB08, Upa09, WL08c]. ecology [DAQ09, YKJ07]. econometric [HN06a]. Economic [TG06b, AKH06, Cai06a, IR07, JTMRS06, MPT08, Pal08, PVM07, PFA07, Vel09, ZY08d, ZYD07]. Economics [HMNT09, SYT06]. economies [KOS08, MNM07]. economizing [Day05]. economy [Dam07, DvdHD08, MTD05]. ecosystem [WWLX06]. ecosystems [FK07]. ED50 [Men08]. EDA [XYZ09]. eddy [BST06, Che09a, KK05a, Ma06b, Ma08a]. eddy-current [Ma08a]. Edge [Su07a, XHZ06, XHX07, ZXL06, Li05e, LCX09, Su08a, Upa09, WYL08b]. Edge-bipancyclicity [XHZ06]. edge-connectivity [LCX09]. edge-fault [XHZ06]. Edge-pancyclicity [XHX07]. edges [Lin06c]. Editor [Cas07c, HM09b, Nad07a]. Editorial [Ano05j, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p]. education [ZYD07]. EEG [PG07a, PG08c, XC06a, XCJ09]. Effect [BA07, EB05e, ESMM06, JK07, MPS08, MCMA07, MR09, Mur06, Sin08a, SDS05, SSK06, SSS06, AEG05, ADK07, CSY07, Coo06, DV06, dCDNPN09, Ere09a, Ere09b, Far08, GCSS05, Kov07, LC06a, MM07d, MA06d, MB05a, NSS06a, NAM^{+}07, RM08, SML05a, Sel06, SJC09, dSYGM^{+}05, SS05f, SS05h, SAS05, TNS07, Wan08b, XZ07c, XYL10]. effected [KOS08]. Effective [Mav09, CK08, HYL09, HH06b, zHWH07, Hua09, JA08, Kol06, lLWhL06, Li07b, Li09d, MMC06, NBO09, OASM05a, OASM05b, Wu06d, ZLH^{+}09]. effectiveness [ÇFÇ06a]. Effects [BAB08, HHZ09, LN06, MH08a, ORG^{+}09, Bat08c, Bat08b, BK06a, Cha07d, CTGD08, EB05c, EEAES06, ESS05b, ESAR05, EESF04, GTX08, IHB06, Pan08c, RYE07, SM06c, SV07, TS09, XL08a, Yan07b, YX06a, YF08, ZPT^{+}06]. efficiencies [KTW08]. Efficiency [JF05, WYL06d, AK06f, AK05f, AK05g, AKR06, AS06g, AN06, Ami07, ET05b, GKK09, GSDB06, GSDB07a, JLZ05a, JMLR05, JLS^{+}05c, JSdM07, Kho07, KJ09, LJK07, LJL06, NTZ08, Pet08a, SML05b, SML05a, SHSZ07, TARm08, VMK05, WL06e, Wu06a]. Efficient [AK05f, AJRLRGOG09, CLS06, Che05h, CJL05b, HW05a, KK05b, KLW06a, LKY05a, LKY05b, Muk07a, TC05, You09, AAES07, ÁCTZ05, Ami07, ADK06b, ADK07, AFS06, AL06, ANK05, BDCW06, CY05b, CHC05, Chi07c, Chi08a, CTL09, DHL05, DL09, Dan08, DwWmW05, FMH07, wG07a, GOMA06, HA05, HK05, HK09, HH07a, HLS05, JF05, JLZ05a, JLS^{+}05a, JHST05, JLRB05, Jun09a, KC09, LFM06, LJ07, LJK07, Lin07b, Lin07d, LJE07b, LLWC07, MA05f, MK06, Ony05a, PL05, PL06a, hPyHZ06, Pet08b, QZ09, QZZ07, RAZ07, Raj08, RY05a, SSB06, SH07a, SRMRGRL08, Sha08a, Sha05b, SC05a, SH09b, SH06c, WYZL08, XC05, YTWY05, YC05b, (Ag07, Zah09, ZN05, ZLW06b, ZBH06, Zho07, Zho08a, ZLC07, Ami07]. efficiently [DES07a]. effort [Mak07b]. Egoist [JLS06]. Ehrhardt [Bra09]. EHTA [DLL09]. eigen [Sun06a]. eigen-decomposition [Sun06a]. eigenelements [Att05b, AL06, Odi09]. eigenenergies [NHKZ06]. eigenfunctions [ACJ09, BUÇ07]. eigenpairs [LHZ06, LHZ09, WD06]. eigenproblem [HY07b, Kay05e]. eigenproblems [BMP06, NS08]. eigensensitivity [yZ07e]. Eigenvalue [YE07, Zha08f, jB06, BW05b, CC07b, DY07, Gha08, LMZ08a, LZC05, LB06, gLxW05, MH09, NR07b, Nat09, Pen05, PHZ06a, TMS08, WJ06, WSB06, Ye06, Yin08b, YHL08a, YHL09b, Yua09, yZsZ06, ZLW06c, ZLZ07, ZL08b]. eigenvalue-clustering [CC07b]. Eigenvalues [Cha07c, ZS08a, AN05a, BUÇ07, Çel05a, Çel05b, Cha07b, Cha05e, JSL05, LR07a, LHL06, gLL07, NG06a, NG06b, NR07a, NRA07, NM08a, Pin06, Rim06a, Rim06b, Rim06e, Rim07c, Rim07a, Rim07b, TGSNB07, TM07]. eigenvector [tLlWpL08, WC06d, yZsZ06]. eigenvectors [HE08b, Lin06d, TGSNB07]. eighth [AS06a, BWR09, GJ07c, Meš07]. eighth-order [BWR09, GJ07c, Meš07]. Eikonal [CR05a]. El-Saka [Pan08c]. elastic [Akm09, BL06, CB06, Gal08a, Gal08b, Gar07, HAU05, HS09c, Li07e, LZ08b, Sel06, SA06d, ST06c, SK07b, TS05, Vin09]. elasticity [CCK06]. elastomeric [EB05e]. elastoplastic [Sey05]. elastoplasticity [NKR^{+}08]. Elbarbary [Pan12]. Eldabe [Pan08c, Pan12]. Eldahab [Pan08b]. election [FJT05]. electoral [MS07c, Oca07]. electric [ESMM06]. electrical [AGS07, AGTS07, KE05c, OMR07b, SANN07]. electro [ACZ05, KK07c, MUZG09, ZY09]. electro-oculogram [MUZG09]. electro-osmotic [ZY09]. electro-static [KK07c]. electro-thermal [ACZ05]. electrocardiogram [YZZP08]. electrode [QS09]. electroencephalogram [MUZG09]. electrolyte [Vyn07]. Electromagnetic [Sel05, ASA07a, FTT^{+}09, KS08c, Liu05b, Ma06b, SC06c, TT07, TA05b]. electromagnetism [Köp07b, TGSNB07]. electronic [BV06, Che05h, KKS06, LJL06, WLC07b, YSCC05]. element [AAN05, AABD07, Aks05, AÖ06b, Aks06a, ALL07b, An09, Ang05, BR07a, BH07, BS05a, BCCK05, Bi06, BS07c, CCE06, CR06, ÇFÇ06b, ÇFÇ06a, CN06, Che08a, CL09b, CS09c, CFJ^{+}05, DIT06, Dai07, DC08, DXH07a, dCDLN^{+}07, DJ07, Eid05, EGX05, FM05, FJCW05, GJ05, GY06, GFH08, GK05a, GG08a, Gw05, GG07, mHsLzC06, mH06c, He06, mH07b, HL06b, Ide06, JZ06b, KYK08, KBH06a, KBH06b, KBLT08, KNJ05, KE05b, KE05c, Li06b, LMH06, LAL07, LL08e, Li09f, Lin05a, LWL06, LL09, LL12, cMpZfR06, Ma07a, Ma08a, Mai05, Nat09, OY05, Ony05a, Ony05b, PD09, QBH06b, qQcM05, Sak06, SH09a, SL09b, SW09a, SP09, SRH09, SFA08, TmChL06, TD09, TY07, WC05a, Wan05e, WHL06, WsZ07a, WYQ05, XO07, XC06b, XC07, YM07a, Yan08e, ZO05, Zha05c, ZOZZ08, ZC08a, ZWZL06]. element-continuation [CS09c]. element-free [ZOZZ08]. elementary [ADDdM07]. elements [ADK05, ADK06a, KYL05, KE05d, MKR06, Mar08, ÖEK05, Rim07g, Rim07h, Rim07d, Rim07f, Rim07e, Rim08b, Rim09b]. eleventh [SA07e]. Elgazery [Pan12]. Eliaz [Tan06d]. elimination [ACDR09]. ellipses [CH07f]. ellipsoid [Men08]. Elliptic [BK07c, WL07a, Afr05, AMN06, AK06c, AK06d, AK06e, AMN07f, ANM07c, AMN07d, AMN07g, ANM07g, An05b, AV07b, AA06e, BZ09, Bay05a, Bou05, BSKS07, CS05b, Che05g, CW06c, CG05b, CFVZ06, CMS05, CY07c, DMA07a, DMA07b, DwWmW05, Dod08, EHZES05, GS06a, GR06a, mH07b, Hon09, Hsu05a, HL06c, HTM09, Hub09d, HLS05, HH05, HL05d, HJK08, Inc06c, JZ06b, KS06b, LWZH05, LCS09, LY07d, LY08c, MM07b, MTM05, MY07a, MY08, MMS^{+}06, MMS^{+}08, MSR08, MA05g, MS06b, Moh07b, MAD07, OY05, Özt05, Pao07, SS05a, Sal06f, SGT06, SCL05b, SLC05, TC05, WLZ05, WHL06, WR07a, WLW09, WLCL06, fW07b, WY08, XC06b, XZB05, Yan05a, Yan06d, YXW07, YCL07, YY06b, YY07b, YY08c, ZZYY08, ZL05b, ZZC06, ZYZ07, SL07b]. elliptic-type [Sal06f]. elliptically [Nad06a]. Elsayed [Pan12]. Elshehawey [Pan12]. ELSP [JGTK07]. Emad [Pan08b]. emanating [CLW09b]. Embedded [AKIS07, AEA05a, Boy05, BO06, JZW08, Pan08b]. Embedding [BT06a, CY08b, HS07a, Li05e, MLP07, MFE05, Wan05f, Yan09a]. EMD [ZL06c]. Emden [HM08a, KM09b, LX09b, Waz05b, Waz05c, You06a]. Emergence [Coo09, MA05e]. Emergent [GPS09]. Emission [Pal08]. emissions [Liu07a]. emphasis [Hua05b]. Empirical [CL06b, CS07c, ZZW08]. Employee [Nwo06b]. Employee-stock-options [Nwo06b]. emulation [BK07b]. enclosing [PL06a, PG06]. enclosure [KS05b]. enclosures [Sch08]. encoding [HHC08, LLX07b]. encryption [HHYW07, Hsu05a, Hwa05, HL05c, HL05d, JXW05, JSW05, JZCW05, LWK05a, TLH05, ZAX05, ZW05b, ZL05b]. end [Pud05, Shu05]. end-directed [Pud05]. endemics [BAB08]. endoscope [HAAS06]. endothelial [PE07a, dG06]. ends [Ray08]. energetic [MUZG09, dCRDSSN^{+}08]. Energy [SKHZ06, TmChL06, dSLV^{+}08, Ang05, AGTS07, BSHZ07, CB06, Gor06a, Ide06, Kla08, LP08c, MD08a, OMR07b, Shu05]. engine [SG05c]. engineering [BDCW06, HPD07, RJM07a, uITA06, WCX08, ZLH^{+}09]. English [LLX07a]. enhance [MSM08]. Enhanced [WHH07, DT09, JSdRm05, LHXL05, TWL05, WHH05]. enhancement [DFGV07, FAQ08, YL06c]. Enhancing [GCJ08, Thu07d, Mir06a]. ENO [Rao05]. enrichment [GFH08]. ensemble [CH06a, WS07f, WS07g]. enterprises [ZLZ06]. Entire [MY07a, MY08, YY06b, YY07b]. entrant [KO06]. entries [Kar05b, Kwo07b, Li07b]. Entropy [SM06a, Aki05a, Akı06, BM06b, BP06b, Deh08, DM07b, GG05a, JM07, SM08f, KL06b, LH08c, Pud05, Sal07c, SDM^{+}09, WWP05, YLC09, Yin07, Prá08b]. entropy-optimization [JM07]. Entropy-regular-solutions [Prá08b]. enumeration [FA05a]. envelope [SMQ06]. envelopes [WL07f]. envelopment [AJK05, Ami06, AKS06b, AKA07, ET05b, JMHS05, JLZ05b, JPZ06, Kho07, LJK07, Sae05, SML05b, SML05a, SMD06, WYL06d]. environment [DAQ09, DCR07, KO06, LZ09c, PWL06, SB08a, ZŠ05, ZT07a]. environmental [HHZ09, Kum09b, MPT08, MB05b]. environmentally [GCSS05]. environments [KNJ05, LGHS06, LL05h, MMC06, Sch05]. enzymatic [dSYGM^{+}05]. enzyme [dSYGM^{+}05, YSdS^{+}06]. enzymes [BY05]. EP [ZW05e]. epidemic [Bia06, BK05a, CL08a, CLG09, CP05, JL07, JY09, JWS^{+}09, KMVL09, LW09a, LLG09, Mak07a, MMC^{+}08, MCC07, MC08, MS07e, NTO06, RGD07, RDG07, SJC09, SJW09, SXL09b, TQC09, dWqL05, WTS09, YC08c, ZW07a, ZzLZ08, ZLT09, d’O05a]. epidemics [WdRS07]. epidemiological [PA09, UBKC08, ZLY08]. epileptiform [PG07a, PG08c]. epistasis [BAB05]. EPQ [IR06, PN08]. epsilon [AK06f, Ali05]. equal [Dog05, Ese05b, EDMJ05b, HED06c, Ras05a, Sak06]. equalities [BT09a, WSL08]. Equality [Yam08, DZG06, JjQjWmL06, JXH06, JZX07a, LZ06e, MSD06d, ZCH07, Zhu05a]. equality-constrained [JZX07a, MSD06d]. equation [AD05a, AD05b, Abb06a, Abb06b, AESD05, AMK06, AL09, AKO06, AC05, AE07a, AK06d, AMN07f, ANM07d, ANM07f, AMN07d, ANM07b, AMN07c, AMN07g, ANM07g, ANM07i, AN05a, AKBZ08, AA07a, Aks05, AÖÖ06, AÖ06b, Aks06a, AV07a, AVM^{+}09, AKM05, AK05c, AA06c, ACZ05, AK06g, Alo06a, Alo06b, Alo09, AD09b, AH07d, AC08, AEA07, AA07f, Bad05, Bah05, BS05a, BUP07, BQ09, BZ09, BSKK06, BK08b, BR09a, BN09, BLW06a, Bar05a, BSN^{+}08, BNH07, Bay05a, BŞ07a, Baz08, BHN08, BB06b, Bha08, BM08, BA05a, BG08, BÖ05, BA05b, BB05, BKBK06, BQYW07, BS07c, BZZ08a, BR09b, BD09b, BGF09, Boy06b, Boy08a, BT06b, Buk05, Buk09, CC08a, CTT08, CSV07, Che05e, Che05f, CW05a, CW05b, CW06b, CLSZ06, CW06a, CW06c, CXC06, CG07b, CW07, CL08b, mCLB08, CLW09b, CLDH09]. equation [lCS06, CFF09, CKS09, CSC08, CIS09, CE07a, CJSSP05, CR05a, CG07c, CD07, DGJ07, DSB05, DIS05, DIT06, Dai09, DLLN09, DN06, DG07a, DJ06b, DKK06, DEB06, DDD05, Deh05a, Deh05b, DMS06a, DMS06b, DM06, DMA07a, DAQ09, DCS07, Dem06, Dem05d, DXH07a, Den08c, Den08d, DT08, DSK08, DSZ08, DZ09, DEFT05, DM08b, Dog05, DS06a, DC07, DDR05b, DX06, DC06e, DB07d, DLJ08, EA07a, EG06a, EHAR05, ESK06, ETETH05, ETETH07, EA06, EI09, Eng06, Ese05b, EK06, Fat05a, Fat05b, FCG06, FH06a, FLLH07b, FHL07, FLLH07a, FL09b, FL09a, FXLZ05, FHG07, Gal08a, Gal08b, GA08, Gao06, GB07, GC07a, GL08b, GL08c, GK05a, GK05b, GKD07, GK06b, GG06, G07, Gor06a, GÖ05, Gül06, GS06d, GS07c, GHX08, HAK07a, HFL06, HW08a, HKS09a, HM08a, HES05, HI05, HDG07, HED06a, Has06b]. equation [HNB06, HSH05, HAM04, mH07a, HRLC08, HRCL08, HM07, Hol07, Hon09, HZS07, HL07c, HXH08, HC09, HTM09, Hub08a, Hub08b, Hub09a, Hus09, HS09b, Inc05, Inc06b, Inc06c, Inc06a, Inc07a, Inc07d, IS09a, IDS06, IRSR05, Ism06, Ism08a, Ism08b, JRW08, Jav06a, Jav06b, Jav06d, Jer08, JFG08a, Ji08, JSL05, JW09b, JC08a, JZ06c, JL05, JZX07b, JZX07a, Jin08, JR07, KA06a, KKT05, KGB06, KPS06b, KS08b, Kay05c, Kay05b, KI05b, KES05, KE05a, KM09b, KuI08, KG06, KGM06, Kim08, Kin06, KL05b, KB05b, KKS06, KV05b, KWZ08, KHZ07, Köp06, KS09a, KL08, KM05d, Kum09a, KS08f, KH09, LeT05, LL05a, Lay09a, Laz08, Lee06b, LpL07, Lei09, LS05, LL05d, Li06a, LG06a, LZ08b, Li08c, Li09a, LXJ09, Li09g, LZ09a, fLyHZ09, Lia08, Lin07a, LLAT09, Lin09a, LXML09, LZZZ06, LZZ06b, LS07e, LZ07c, LGY07, LZA^{+}07, LD08, LL08f, Liu09c, Liu09a]. equation [LX09b, LWKU09, LHSB09, Lou07, LN06, Ma07c, MYG08, MMW09, MC09, MWQ09, MLW09, MMO07, Mah05, MMN06a, MM06b, MAM06, MR06b, MA06c, MY06b, MMA07, MSD06a, qMqZ05, MSH08, MSHeT07, MY07b, MAY09, Moh05a, Moh05b, Mom05c, MO06, MMR09, MD08c, MTL08, NG09, NA06a, NKH07, NNuH07, NH07c, OM06, OY05, OB09, Omr06, Omr07a, OAAK08, Ony05b, ÖK05, ÖEK05, ÖA05, ÖÖ06, OBO06, Pam06, PVV09, Par05a, PK08a, Par09a, Pat05a, yP05, PHZ05, hPyHZ06, PY08a, Pen08, Pod06, PCS07, PKT06a, PKT06b, QS06, QW09, QN08, QW08a, Qui09, RMAM07, RJW07, RES06, RES07, Ram07f, RZ07b, Ras05b, Ras05a, RB05a, RB05b, RB06, Ray06, RCB08, Ray08, Rom09b, RZ07c, SKHZ06, Sak06, SD09, Sal08b, Sal08a, SGC09, Sal07b, SA05a, SA06b, SAB06, SA06c, Sal06e, SG09a, SY07a, SM05b, Sha07a, SQHY08]. equation [Sha06c, SS08b, SP08, SXJ05, SX06, SC08a, She08, SL09a, SYH07, SWC05, SDL09, SLC09, SDG05, SZ06a, SS08c, Sil06e, ST05b, Sim09b, SBA09, SLO08, Soh06, SI06, SH05b, Sol05c, SZ07c, SZ07d, SZ07e, SLL08, Ste09b, SL05b, Su09, Sun05a, SMHQ06, SW09d, Sya05a, TSD07, TWZ07, Tan05b, TL06, TH07a, TZW09, TZH09a, TXW09, TZH09b, TB09b, TD06, TD07a, TM08, TW09, TNL09, Tze07, Van05, Wan04, WW05a, WLZ05, Wan06g, cWmZ06, WCZ06, WCX06, WZZ06a, WL06b, Wan06i, Wan06h, Wan06j, Wan06f, WMX07, WM07, WL07a, WL07e, Wan07d, WCW07a, WZ07c, WSZ07b, WSKZ07, Wan07a, WLT07a, sWMX07, WGL08, WX08a, WLH09, WC09a, WDLM10, Waz05b, Waz05e, Waz05f, Waz05h, Waz05j, Waz05k, Waz05l, Waz05m, Waz06a, Waz06c, Waz06g, Waz06h, Waz06k, Waz07c, Waz07e, Waz07d, Waz07g, Waz07k, Waz08b, Waz08c, Waz08d, Waz08g, Waz08k, Waz08n]. equation [Waz08m, Waz08o, Waz08t, Waz08q, Waz08v, Waz08u, Waz09b, Waz09a, Waz09c, WZ08e, WLS09, WYQ05, WL08b, WL08c, WZF08, WC08b, WFHD09, XXL09, XZ05, XF06, XC07, XZB05, XZY08, XMGW09, XL09a, Yan05a, Yan05c, lYqL06, lYqLW07, Yan08a, lYpZqL08, hYG05, Yan05d, YEM05, YLS05, YLL06a, YYL07, YY08a, Yan08b, Yan08d, Yao08a, YC09a, YC09b, Yax07, YC05a, YX08, You06b, You08, YZ07c, YZ07d, Yu08b, Yüc08, Yue07, (Ag06, YB07, ZW09a, Zha05b, ZZL05, ZZ05a, ZX06a, ZX06c, Zha07c, Zha07d, ZZ07a, ZZB08, ZDY08, Zha08b, Zha08c, Zha08d, ZM09, Zha09e, ZZH09, ZC09, ZnCxX09, ZYD09, ZLW09a, Zha09c, Zha05e, Zha05d, ZZKY05, Zha06c, ZXLS06, ZZYZ06a, ZZTJ06, ZZZ07, ZM07, ZMY08, Zhi09, ZLZZ06, ZL07b, ZWZZ07, ZLZ08, ZLYL08, ZXY06, ZL07c, Zhu08, ZW09c, ZWC09, ZMV05, lSAQ08, DLL09, Hub09d]. equation-based [SZ07d, SZ07e]. equations [AJE05, Abb05, AEJ06, Abb06c, AE06a, Abb06d, AJ06a, AJ06b, ATL07, AAS06, AC05, Ade07, AJKM09, AMN06, AMN07a, AO08, AO05a, AD06a, Ahm09b, ADY06, AD06b, AKMA05, AKARA08, AM08a, AGR09, AL08, AD07, All05a, All05d, AAAA06, AB06, AŞ06e, AAK07, An05b, AC09, AE07b, AH08b, Arg09, AHIP07, AO05d, AO06e, AO06d, AKM06, AKVH05, AY05b, AA06e, AÖ09a, Asl09a, Asl09b, AÖ09b, Att06c, ADDdM07, AKR05, BD05a, BGA05, BVC05, BDE06, BFR07, BF07a, BF07b, BLP07, BAF08, BMS08, BR07a, BM06a, BS06, BLW06b, BC09b, BP09, BSC06, BK09, BS05b, Bel06, Bel07a, Bel08a, Bel08b, BS09b, BDG08, BRW08a, BWR09, BE05, Bia05, BG07a, Bog05d, BS07d, BKZ06, BRRGT07, BRT07, BKM06, BG07b, BM06b, Bou06a, BB06d, BJ08]. equations [BEF08, BSKS07, BK05b, CC08b, CFX05, CL09a, CD05, Cao06, CHLS08, CM09, Cha07e, CT05a, CS05b, CW06c, CW06d, CCC07, Che07f, CS07i, CY08c, CjC08, CA08b, CM08a, CA08a, CL08d, CZ08b, CS09c, Chu06, Chu07c, Chu07d, Chu07e, Chu07h, CH07d, Chu07k, Chu07m, Chu07n, Chu08, CGJ08, CJP08, CE07b, CT08b, CC05d, CY05c, CD06, CY07c, DGB08, DC08, DX07a, DE07a, DB07a, DB07b, DB07c, Dar06, DSZ07, DMK07, DHS07, DMA07b, DH08, DBC^{+}07, Dem05c, Den09c, DMH05, DXH07b, DLD08, DZ08b, fDxZfWyY09, DC06c, Dön06b, DZ07b, DB09, Dyk09, ERAS07, EN05, ER06a, EAEG07, EBsENMA05, EB08, EGFAE06, EHZES05, EMESES05, EOM03, EOM08, ES05b, ESK05, ESE05a, ET05a, EWAE06, EWEA06, EWAH08, EkES05, Elb05, EELZS05, EHP08, EBP09, Fad06, Fad05, FHYZ09, FGfFL06, FZ09, FO09, FE08, Fel05]. equations [Fen06, FH07, FLLH08, Fig08, Fra09b, Fro06, Fus08, Gal06a, GZ05, GJ05, GJ07a, GL09, GZ09a, GZ09b, GFH08, GHL07, GZ07c, GKB07a, GKB07c, GkB07b, GK06a, GF06, GJB05, GS06b, GS06c, GJ07b, GJ07d, GJ07e, GJ07f, GS07a, GJ07h, GR06a, Gor06b, Gor07, GY07a, GBC09, GAPT08, GO08, GSPG07, Gw05, GW06a, GbL07, Gu07a, GS07b, GG07, GS05, GSG06, GR06b, GR07b, GC07c, GC09, GB06a, GB06b, HA05, HAK07b, HC07a, HX07, HW08b, HBE09, HY05, lH08a, HZ08, HBVM07, HIPA08, Hon08b, Hos05b, Hos05c, Hos06a, Hos06c, HN07a, HRS07, HR09b, HW07b, HLW08, HHW09, Hua08b, HZ09b, Hub05, Hub09b, Hub09c, HMT08, Ibr07, Ide08b, Ide08a, Inc07c, Inc07e, Inc07b, IS09b, IM07, IMB08, IMS09, Iva08, JDG06b, JDG06c, JDG06d, Jai07, Jan08c, Jan08b, Jan09a, Jav06c, Jav07, JG07a]. equations [Jbi06, JWPS08, JL08, JZ08a, JOA05, Jia07b, JH08a, JH08b, Jia09, JH09, JV05, JJ09b, KTK06, KKT07, KM06, KSA05, KR07a, KR07b, KK08a, KS08a, KAV05, KV05a, KK07c, KK05a, Kan06b, KSGM06, Kar09c, KMÖS09, KY05a, Kha09, KSM09, KRRGT07, Kim06a, Kim06b, KM05c, Kıy06, KSS07, Kos07, KLW06a, KLW07a, Kou07c, KLW07c, Kou07a, KL07f, Kou07e, KW07, Kou07f, KL07h, KL08, KS09b, KD09, Kum08, KP08c, LY07a, Lak05, Lar05, Lay09b, Lee06a, Lep06, Lep07, Lep08, Lep09, Li05b, LSL06, LLZ06, LDZ06, Li06b, LSW06b, LSW06a, LS06c, LMH06, Li07e, LZ07b, zLmLL08, LZ08c, fLsLdXlF08, Li08b, LC09a, LWW09, LLX05, LS07c, LZ08e, Lin05d, Lin06d, LTW07, LYM09, LCL05, qLJ05, Liu06g, pLT06, LS06e, LW06d, Liu06h, LS07d, LKU07, LMY07, LLY07, Liu08b, LW09d, LKU09, LL09, Liu09a, LLL09, LW09c, LLW09, LLC09b, hLqSlZ09, LL12, LP06b]. equations [LP06a, LCM07, LW09e, LSN09, LOA05, LM07b, Luo05a, LYL07, MM07a, cMpZfR06, Ma06a, Ma07a, MW07, Ma07b, MTL07, MJ07b, MM08a, MX08a, MM07b, MD08a, Mah09a, MCAH05, MPG07, MPEL09, MS07a, MMT07, MB06, ML05, MKA05, MSK05, MK05c, MA05d, MM05b, MK05d, MD06a, MY06a, ML06, MD06b, MAR06, MKR06, MNY07, MLR07, MLM07, MS07b, MD07a, MSR07, MS08a, MKKS05b, MB07, MSTY06, MQ07, Mar05b, MJD06, Mat09, MSD06c, MC06, Mei06, MV08, MX07a, MX07b, MO07, MY07a, MZ07, MR07c, MSR08, Moh07a, Moh07b, Moh07c, Moh09, MAD07, Mom05b, Mom05a, MAK05, MS06c, MN06, Mom06, Mor08b, MED08, Muk07c, Muk07b, Muk08b, Mus07, MT09, NDD05, NMM08, Ned05, Net08b, NL05, yN06, NRLT07, NBO09, NNMDS06, NNKA06, NA06b, NN06b, NAJ06, NN06c, NN07c]. equations [NN07d, NN07a, NNM07, NG07, NN07e, Noo07b, Noo07c, Noo07d, NN07b, Noo07e, O’R08, Odi06b, Odi08, OME08, OAHR09, Ord06b, Özb07a, Ozb07b, PS08, PPF09, Pam05a, PPS09, PRM06, PLL08, PFG08, PG08a, PG06, PG08b, PS06b, Pen06, PHZ06b, zPY06, PP07a, Pet08a, PMRAS05a, PMRAS05b, QBH06a, QBH06b, QZL07a, QW08b, RMA07a, ROK09, RKOK09, Raf06, Raf07a, Raf07b, Raf08, RAS07, RN09, RMS05, REDES05, RK08, Ram05d, Ram05b, Ram05f, Ram05g, Ram06b, Ram06c, Ram06d, Ram06e, Ram06f, Ram07d, Ram07e, Ram07g, Ram08c, Ram08f, Ram08d, Ram08e, Ram09a, Ram09b, Ran08b, RZ05, RZ07a, RJK07, RAP05, RB08c, Raw06a, Raw06b, RX09a, RX09b, RMA07b, RHY06, RMWk09, RV08, RV09, SNH07a, SNH07b, SNH07c, SA08a, Sai07, SG08a, SA06a, Sal08c, Sal06b, ST06a, SB09, SK08a, SS08a, SPV08]. equations [SG05a, SG06, SGT06, SAD06, SG07a, Sha05a, SH09a, Sha06a, Sha06b, SH07d, SHY08, SR06, Sha05e, SK06b, Sha07c, SL07a, SS09d, SC07a, SXL09a, SP09, Shi09b, SRH09, Sil05b, Sil06b, Sil07c, SM05c, Śmi07, Sol05c, SZ06b, SZ07b, SHG07, SZ08a, SLW08, ST08, SB05b, Sta06a, SLS08, SZ08b, SM08a, Sun05b, SS05g, SJ05, SS06f, Sun06b, SMQ06, SHD07, SS07e, SM07c, SM08b, SM08c, SY08b, SM09c, SM09b, Sya05b, SH06c, TF08, TWZ07, TL07, TH08, TL08, TB09a, TS09, Thu08, TFX07, TG08a, TJG08, TG08b, Tou09, UHuI09, Uje06a, UEL07, Uje09, Vin09, Vul07, WJ05, WCLZ05, WC05a, Wan05e, Wan05a, WX05, WWL06, WZZ06b, WZSZ06, WiN06, WL06a, WTZ06, WC07a, sWX07a, WZC07, WL07a, WSL07, WML07, WSZ07b, WSKZ07, WL07c, WZL08b, WZ08c, WCN08, WYL08a, WLS08, WTW08, WWL08a, WWL08b, WW08b, WH08b, WCL08, WQL09, WL09b]. equations [Waz05a, Waz05c, Waz05d, Waz05g, Waz05i, Waz05o, Waz05n, Waz05p, Waz05q, Waz06b, Waz06d, Waz06f, Waz06e, Waz06i, Waz06j, Waz06l, Waz06m, Waz06n, Waz06o, Waz06p, Waz06q, Waz07b, Waz07a, Waz07c, Waz07h, Waz07f, Waz07i, Waz08a, Waz08e, Waz08f, Waz08h, Waz08p, Waz08l, Waz08j, Waz08i, Waz08r, Waz08s, Waz08w, Waz09d, Waz09e, Wei09, WL05f, WYL06b, WYL06c, WWY08, WWYL08, WWY09, WS08b, WZY07, WN08b, Wu05d, Wu06d, Wu06b, Wu06e, WZM06, WDY06, Wu06c, WCW07b, Wu07d, Wu07c, WMBY07, WF07c, WF07b, Wu07g, WF07d, WHL08, Wu09, Xia06, XCM07, XWZ06, XCFX05, XZ08a, qXbL05, XBZ06, XZB06, XHH06, XM06, XM07a, XM07b, XGTS07, XZ08b, Xu08b, XZ09b, Xu09, YSdS^{+}06, YS06a, YSS09, YST09, YLZ06, YHTR05, YLB05, Yan06d, YC06a, YLB06, Yan06b, YEM06, YXW07, YS07, Yan08e, YCL09a, YL06b, Yao08b, YX09, YBA05, YL06c, YY06b, YLH09, YMY05, YC07b]. equations [You06a, YB06, YWL07a, YWL07b, Yu08b, YY08c, Yun08a, Yun08b, (Ag07, ZG09, ZZ08a, ZSHB06, Zha05a, Zha05c, mZzH05, ZT05, ZX06a, ZX06b, ZLCZ06, ZT06, Zha06b, ZYME06, ZLW07, Zha07c, ZX07a, Zha07d, ZZ07a, ZFY07, ZCMQ07, ZZDL07, Zha07g, ZM08b, ZZYY08, ZS08a, Zha08d, ZYH08, ZM08a, ZG08a, ZLW08, ZWT09c, ZQJ09, ZXDL05, ZZC06, ZC06c, ZZYZ06a, ZXWL06, ZZ07b, ZX07b, ZWX08, ZC08b, ZM08c, ZZ06a, ZQ08, ZZ07d, ZLR07, ZSL05, ZH05, ZLC07, ZYZ07, ZZ07e, Zho08b, ZZ08b, Zho09c, Zho09b, ZLDW09, ZL09b, Zhu05b, ZW05j, ZG08b, ZJO08, Zuo09, lS08b]. equator [ZLC06, ZZW09a]. equatorial [Boy06b]. equidistant [EFI06]. equidistribution [SSW09]. equilibration [XLSL06]. equilibria [AK06a, Dah07, PVM07]. Equilibrium [XT06, AK05a, Bar09, CY08a, Din05, gDH07, GJD^{+}07, Ino08, Jun09b, LOZ07, PP08, Vel06, Woo05, WSXL09, XD07, YFLX09, ZH09a]. equilibrium-like [Din05]. equispaced [BX09]. equitable [AS06f]. equity [Che07a]. equity-linked [Che07a]. Equivalence [Ram05a, Tan07a, ZY08a, FA05a, HBN06, Tan06a, Tan06b, Tan07c, Tan08b, XZ07a]. Equivalent [SC06c, Wan07c]. era [SA08c]. ergodic [Aki05b]. Erickson [GS08b]. erosion [MDZ08]. Erratum [EOM08, ESAH10, KAK05, KAK06, Kas06b, Kum09c, KP06b, Li05f, hMfWzLpZ09, MK10, Pak11, Pre06b, Sal07d, Sil06h, Sol07, UNS07, WL08a, Wan08b, Waz08b, Wey08, WH09, Yam05a, Yam05b, Yin08b, YHL09b, YHL09a]. Error [EkES05, GG07, mH07b, Kin06, Mor08b, OF06, VCRCG09, WLH09, Wan09a, Wu07d, Bou05, BDY06, Cai07, CL09b, lCX06, GJ05, Geu07, HZ06a, HED06b, Hos05d, Hua05b, IT06b, JZ06b, Kök06, KY06, LAL07, pLT06, LH07b, MAM06, Omr07a, RW09, Sha05a, SY08b, UE07, Vaj07d, WS07f, WRB07, Xia06, YZLD08, ZZ05b]. Error-free [Mor08b]. Errors [MR05b, AA06d, Bog05b, HBN06, HN07c, JK06, gLxW05, MNN05, NS07a, Vaj08, lS08b]. ERT [WZ09b]. escape [WZSL08]. escape-line [WZSL08]. escrow [LCC05c]. essential [LX07a, Ste09c]. ESSs [PSP05]. establishment [SL05a]. estate [Nwo07b]. Estimate [ZL08e, lCX06, DRS^{+}08, HZ06a, JZ06b, LH08c, YLC09, Zho09c, ZYD07]. estimated [WWC06]. Estimates [Zho06, Aou08a, Bou05, Cai07, Cai09, CL09b, Dzi09, FA05b, Gor06a, GG07, mH07b, LAL07, Moh06, SY08b, TK07, ZY08b]. Estimating [AJAH06, CYL07, Hsu06, LS06a, LS08c, PY07, Sal07c, HKS09b, JSdRm05, Kos05, Yan07e, Yuj07]. Estimation [Bog05e, CL05b, CKK05, Kim05a, KTÖA08, LYCW09, Slo07, WWLX06, AR05, ARM^{+}05, Bak05, Bog05b, DJM08, DLS06d, EkES05, FSG07, FSÇ06, HV08, Hos05d, JVFM04, lLWhL06, LLLX07, Lin07b, pLT06, LJE07a, MAM06, Mug05, NCÁHC^{+}06, Nak07, PKL08c, PK09, PG07c, RAGA^{+}05, Sha05a, SS06d, Sil06d, Sil07b, WTW05, WGF^{+}06, WDL07, WHH07, WWT05, WZ06g, Xia06, YC06c, YZLD08, ZL05a]. Estimations [EGS05, Wan06a]. estimator [AÖ05b, Gon07, KC05a, PK08b, WLY06]. estimators [BCEGMRMP07, KC06a, NCÁHCLP05a, RGA06, WLC06]. ETD [CP08]. ethanol [DRS^{+}08]. ETRS89 [Men08]. Euclidean [Gao07, PL05, PC07a, RNV07a]. Euler [ABBP08, CAS08, CAS09, Djo06, FPP06, GS07b, KAPM07, LeT05, LYL07, PLL08, PPV08, PMP08, Ram06a, RB08c, SS08a, sWMX07, WWYL08, Wu07c]. Euler-like [PMP08]. Eulerian [ZL08c]. EURACE [DvdHD08]. European [AS07a, Che08c, DHL05, DvdHD08, HL07b, Wu07a]. European-style [HL07b]. European/American [AS07a]. evaluate [KP07]. evaluated [KF08b]. Evaluating [Boy08a, BCMT06, EM08, JC07, OB09]. Evaluation [BP05, Boy07a, PSD07, ZSHB06, AKKS06, ÇFÇ06b, GK09, KWP07, MM05c, Pra07, Shy06]. Evaluations [KAPM07, Wu05e]. evaporation [ALD08]. even [Bou09b, LLZ06, MX07a, MX07b, Rim05a, Rim05h, Rim05e, Rim06c, Rim06g, Rim06d, Rim06a, Rim06b, Rim07a, Rim07b, Rim08a, Rim08b, Rim08d, Rim09a, Yin08a]. even-order [LLZ06]. event [Ben07b]. events [AA07d, LMT^{+}09]. eventual [MS09a]. Eventually [IS09a, Zho08b]. evidence [Fuk06a, Mit06]. Evolution [LL06b, AAES07, AKB09b, AÖ09a, Asl09b, BKZ06, CW06d, EWAH08, GCJ08, HAK07b, zHWH07, Hub05, Hub08a, Hub08b, Hub09b, Hub09c, JL05, KCW07, KD09, Lep07, Liu09a, hLqSlZ09, Mel08, Ql08, RV08, SZ07b, SZ08a, TB09a, WHC08, Waz06b, Waz06n, Waz08r, Waz09a, YSL05, YL06b, YC07b, YHL08a, YHL09b, Zha05c, ZWL06, ZZ07a, ZZ07d]. Evolutionarily [SSEA06]. Evolutionary [RSDC08, YK06a, ZGH07, ASPA09, Bar09, Bog08, zHWH07, INR08, JZH07, KX08, LMSV07, LYKC07, Li09e, LLX07b, MED08, OASM06, PGdSS08, TTG07, WZ07b, WD08, ZLH^{+}09]. evolutionary-programming-based [TTG07]. evolving [LY08d]. EVT [Nwo06d]. EW [Ras05a, Ram06c]. ex [KC09]. ex-ante [KC09]. Exact [Asl09b, CY08c, CJSSP05, Den08c, Efi09, EB08, ESK05, FXY08, GA08, GHL07, GL08b, GL08c, HAAS06, HRCL08, Hos07, Jan07a, JRR^{+}07, KAF08, KS09b, LF08, LZ08b, qLJ05, LD08, MAY09, MM05d, Oua05, Pen08, QM05, Rom09b, Sal08b, Waz05h, Waz06e, Zha09g, ZY09, ZLZZ06, AM08a, AuH07, Asl09a, BQYW07, CW06d, CjC08, CY05c, CD06, DC06e, ESK06, Fab08, Fel05, FLLH08, FL09b, Fil05, GAR09, HEZ08, Inc07d, Ism05, Ji08, JC08a, Kay05c, KM05b, Kıy06, Küç05b, Lay09b, Lay09a, LG06a, LZ08c, Li09g, LZ09a, LLL09, LH08d, MYG08, MC09, MWQ09, MLW09, Pal09, RS06, SG08a, SPV08, Sha06a, Sha07a, SQHY08, SHY08, SZ06b, SZ07c, WSLZ07, WCLZ05, WCZ06, WZ07c, WSKZ07, WDLM10, Waz05n, WLS09, XGTS07, Yan06b, YMY05, YC07b, ZX06a, ZX06b, ZX06c, Zha08c, ZZ06a, Yan08a]. exactly [LZ07c, WC05c, ZX07a, Zha08b]. examination [Mir06b, WW06, Azi05]. example [dPCdBYV08, Liu06a, Rey05b, ZCZ05]. examples [DJ06a]. Exception [TRLD08]. exceptional [Li05a]. exchange [DD07, SH07a, WWLX06, WL08c]. excitable [KM05a]. excitation [HBBA07, gLXyZ06]. excitations [XXC05, XXC06]. excited [BSHZ07, ETETH07, YXSX05]. exciters [BFR05]. exclusion [LL06b, Lis06, Sya05b]. exclusion-based [LL06b]. Exercise [YC08b]. exhaustive [JY07b]. exhibiting [SS05f, SS05h]. Existence [AS08, BQ09, BC09b, Bar09, BBB08, BT08, CH07a, DEB06, DZ08f, DB09, FZ09, Fus08, GL09, GF06, GHX08, HL05b, HL07d, JZ06a, KKA09a, Li09c, LTZ09, LKU07, LY07d, LY08c, LKU09, LC08a, Ma05c, MM08a, MH09, PK08a, PVM07, RX09a, SRMRGRL08, Sor06, SS05f, SC07d, SHD07, SQG08, WZS07, WWD09, WG07b, WY08, WC09c, XZ07b, Yan06d, YCZ06, YXW07, YWL09, YZ09a, YFLX09, YY08c, YJO08, ZS07b, ZLW07, ZW07b, Zha08e, bZlDL09, Zha09h, ZLL06b, ZYZ07, ZJO08, AC07b, AA08, DZ07b, DZ07a, DH09, ESE05a, HY05, Hlo05, JGY07, JFG08a, JOA05, JC09, Jia09, KeTL08, LLZ06, Li07e, LCM08, LZ08d, LDL08, LZZZ06, LZW07, MN05, MSHeT07, MY08, RX09b, Sae06d, Sal08c, SH05a, SZ06a, SWX06a, Wei06a, WRL06, XL08b, YCL09b, YY06b, YY07b, ZW09b, ZY08b, DJ06a]. exotic [dCD03, dCD06a]. Exp [zLmLL08, Zha08b, Zha08c, GA08, Sal08b, XZY08, MAY09, NR09, SDL09, Waz08u, XXL09, ZS09b]. Exp-function [zLmLL08, Zha08b, Zha08c, GA08, XZY08, MAY09, NR09, SDL09, Waz08u, XXL09, ZS09b]. expand [TF07]. expanded [CL09b, SY08a]. expanding [Yu08c]. Expansion [FBB07, ML06, AES08, AÖ09a, Asl09a, Asl09b, AÖ09b, CW06c, CW06d, CY06a, CL08c, GZ09a, GZ09b, Has06d, mHzC05, Ji08, LZ08c, Lin06b, LXML09, ML05, MA05d, MA06c, MAR06, MTM05, Pog07a, RMA07a, SC08a, SZ07b, SZ07c, Su07d, WSZ07b, WSKZ07, WZL08b, WC09a, Yan07d, ZG09, ZW09a, ZX06b, yZ07e, ZX07a, ZHZ07a, ZZ07a, ZWT09c, ZZYZ06a, ZZZ07]. expansions [BK08c, BS09c, Chi07c, Has07a, mHsLzC06, mH07a, JV07]. expansive [LL07a, SC07b, Son08b, SL06b]. expected [hB06, Ben07b, CF06a, CH06c, LS09a, WWC06]. experience [Kay05a]. experiment [GJD^{+}07, XCJ09]. Experimental [MA06d, Kay05c, ZS07c]. Experimentation [ZWL06]. Experiments [Bie08b, KBLT08, LZ08a, MH06b, MH06a, PRGH05, WL05g]. expert [dPCdBYV08, PG07b]. Explicit [CC09b, FL09b, JWPS08, Kıl08a, LTW07, LCL05, MCS08, OS09, Psi09, Ram06c, Vin09, WLS08, Waz06f, XHSL06, AES08, AS09a, AM08b, AM08c, AM08d, Asl09b, BÖ05, BA05b, CW06d, CCC07, CP05, DG07b, Deh05b, ESK06, FW07a, FLLH08, GL08b, HN07b, Ism08a, KK08b, Kay05c, KMVL09, KPG07, LW05a, LLAT09, MW07, MZ09, Moh07c, Ram07a, SPV08, Sha07a, SQHY08, TD06, TD07b, TD07a, Van05, Van07, WFX07, WWL08a, WYL06b, WF07b, XGTS07, Zha08c]. explicit-magnetic-field [MW07]. Explicitly [TM05]. exploitation [EJER06, Mak07c]. Exploiting [Boy08b, ZL05a, ACDR09, MFF09, Sch08]. exploration [AKB09b]. exploration-based [AKB09b]. Exploring [AP09]. explosive [LY08c, Yan06d]. exponent [CKY05, FX09, LL08a, MM07d, SJ06b, SZ06a, Yan08d]. Exponential [Bah05, EB05d, Gor06a, Isk06, KP08a, KP09b, Nam09, PRM06, QC08, Ram05b, Ram06d, RHQ06, VAMV07, WZ08a, WL07b, ZZ07b, ZZW09b, ZW08d, ADM05, Bak05, BDY06, CL06c, Deo08, FO09, GN06c, Hub09b, JSL06a, Kah07, LWH06a, Li09c, Liu06f, MK05b, McC06, Nam08, NL05, NMP08, Par06b, Par07, Ram06f, eT09, WS05a, Wan06a, WL09a, WZ09b, WZ08d, WLC05, WLT07b, WLC07c, WLL07a, XQ06a, XLL06, YLL^{+}05, YZFL08, ZWX05, ZS07b, ZSCW09, ZLL06b, ZH08b, ZL09a, ZXY06, Zhu08]. exponential-type [Deo08]. Exponentially [Ram05e, AR07b, Ram05c, Ram05d]. Exponentially-fitted [Ram05e, Ram05c, Ram05d]. exponentiation [ÁFVZ06, LLWC07, SCS06, SHSZ07, WLLC07, YLM07]. exponents [CDP06, MB05a, WC06c]. exposed [BM05a]. Expression [CI09, CDP06, PGdSS08, WWL06, WYL08a, YSXL08]. expressions [PE07b]. extend [JLI06a]. Extended [Abb05, AMAP07, AS09b, HCLP07b, KL08, Occ09, Abd07, ASA07a, Bie08a, BAK07, CB06, EELZS05, Gre06, Inc07b, Jan07b, JZ06c, JZH07, gLXyZ06, LXL06, Liu07a, LP08b, NV07, Prá08a, Prá08b, Psi07, Rad09, RAG^{+}07, Sal06a, SHY08, SLZ05b, SLZ05a, SLZZ06, SZ07c, SZ08a, WSKZ07, Waz07b, Waz07a, Waz08p, ZW09a, ZX06a, Zha08d, ZCH07, LZ09a, MAY09, ZSHB06]. extending [Boy05]. Extension [JLI06b, KKY07a, KES09, PR09b, Boy05, BO06, EMH07, GA07, RZ08, Sam06, d’O05a]. Extensions [ASA05, JV07, STV09, CAS08, GKAS07, QG08, Waz09c]. exterior [He06]. external [Far08, Lin08, PRA^{+}09, dlS07a, lS08a]. Extinction [LC06a, CX06, ZCZ07]. Exton [PR09b]. extra [EP07, Wu05e]. extra-cellular [EP07]. Extraction [GRH07, NP05, Su07a, Su07c, Su08a, YZY08, YZZP08]. extragradient [CY07a]. extragradient-like [CY07a]. extrapolated [CT05c, Wan07c]. extrapolation [Haj06, mH06c, Mal06, MO07, MP08c]. Extrema [KFM08]. Extremal [GO08, WWL06, WW09, WX08b]. Extreme [WSL07, WYL08a, Dzi09, JF05, MV09a]. eye [AS06i].

F [OSS09]. fabric [LX06a, LX06b]. Face [xGZZ08, KSC^{+}08, Su07b, BT06b, DY06, JCX^{+}07, KF08a, KF08b, QZ09, ZCZ05, WLX05, WL05g]. Facet [TKJ08]. Facet-based [TKJ08]. facial [LM08a]. facility [TMJJS07]. factor [BHCK08, Den08a, DX07b, EP07, HTO06a, Kar09b, LFW08, SH06b, SY07c, TZW09, YLX08]. Factored [KT07b, Bor09]. factoring [CLH05, LTH05, LCZ05d, PLJ05a, QCB05a, Sha05d, ZCL05]. Factorization [HRS07, CH06a, Ram06e, Ram06f, XTCC08]. Factorizations [XL06a]. factorized [TD05]. factors [AKS06b, AKA07, HV08, JVFM04, KA05c, Liu06a, Liu07a, Sae06a, Sae06c, ST05a, Sal06d, SKT06, VMK05, XZ07a]. factory [OMR07b]. facultative [LTCC08]. Faddeev [STV09]. Fail [CLS05b]. Fail-stop [CLS05b]. failure [GL05c, HLK06, WY05c, Yu05]. failure-censored [WY05c]. failures [WDK06, WC09b]. fair [JLM05]. Falkner [Yan08b, YC09b, ZC09, ZWC09]. falling [Mak06]. falsi [Che07f, CS07h, PG06]. Famelis [Bra09]. families [CT07b, CS05c, Chu07k, Cvi08, Ji08, JH08b, KT07a, OT06, PST06, WLZ05, WL09b, ZX06c]. family [ADM05, AK09, AM08a, BRW08a, BWR09, Chu07c, Chu07g, Chu07h, CH07d, Chu07m, Chu07n, CWZ07, CS07j, EG07c, GJ07d, GSDB07a, GSDB07b, HW08a, KSM05, Kan06b, KH05b, KAO08, KLW06d, KLW07b, KLW07c, KL07a, KL07b, Li05a, MA06e, MKS09, NP05, NS09, Noo07c, OHT05a, PR05, PP07c, Raf07b, Raf08, RW09, SZ07a, SG07b, Sha07c, Sha07d, SC06b, SFG07, UHuI09, UEL07, Waz06l, WZ06h, WH06b, WZ07d, YLK09, ZS07a, ZXWL06, ZW08c, ZWC06, ZLDW09]. fan [Ran08c, FLLH07a, FL09a, YRY05d, ZX06a, Zha08d]. Fang [HT05]. far [Ate06, Mai05, NR09, TA06]. far-boundary [Mai05]. farm [BDPP08]. fashion [GHZS07]. Fast [BK07d, CT09b, LL05a, LP07a, Liu06g, LL05h, ML08a, MK07a, fQZgH08, SdHT09, WLCL06, WLC07a, WLLC07, BSKS07, CL05a, Chi08b, CGJ08, EM08, FDL08, FL06, HE08a, Hos07, Lan07, PBPF08, PK08d, PVN06a, Pet08b, PG07a, Sog08a, Sog08c, SHSZ07, WL06c, WJ06, WYZL08, WM09, XG08, YGL05, YZZP08]. fatigue [FK09, KTV07]. Fault [CY08b, FTG07, KLL^{+}08, YLL06b, CCWH09, CHT06, KH05b, XHZ06, ZXL06]. Fault-tolerant [CY08b, KLL^{+}08, YLL06b, CCWH09, CHT06]. faults [Li05e]. faulty [MLP07, YLL06b]. FCFS [Wen05]. FCM [GAFK07]. FDE [ZFYH08]. FDH [SH09b]. FDM [KK07a]. FDSD [Zha08h]. feasibility [FG07, HW07a, MWW05, YK06a]. feasibility-based [HW07a, YK06a]. Feasible [PHL06, Ars07d, JMHS05, Jia05, JZHT05, JZTH06, JXH06, bJyKxC09, hPY06, Zhu05f, Zhu05g, Zhu06c]. Feature [LLC06a, Mor09, DES07a, GRH07, KF08b, PG07c, Alb06]. Feature-sensitive [Mor09]. features [BNdFPS06, DESG08, KF08a, Kos07, WLL07b]. fed [WFX07]. fed-batch [WFX07]. Feedback [LYX08, Che05a, Che05d, CLH06, Che06b, Che06g, Che06c, CF06b, Che07d, Che07b, CYCX09, Day05, GZ07d, JPYW09, KO08, LP07b, MAR05a, Muk08a, Par05c, PK05b, QD08, WS07b, WS07a, YZ07b, YC08a, YT06, ZZ08a, ZC05b]. feedbacks [SJ08b]. feeder [ONG08]. feedforward [fQZgH08, SM09a, ZZLL07, Zha09d]. feeding [MD07b]. Feistel [LMSV07]. FEM [DS06a, ETETH07, GK05b, dCJLNÁA08]. FENE [Ola06]. FENE-P [Ola06]. Fermi [Yao08a]. ferroconvection [SSK06]. ferrofluid [SM08d, SM08e]. ferromagnetic [MR09, SDS05, SAS05, SSS06, SCB07]. fetal [YZZP08]. fever [DB06]. few [MS07c, NR07a, NRA07]. fiber [Liu06g]. fibers [BTMB09, KBMZ08]. Fibonacci [Kah05a, Kah06a, Sol07, Sub07, UNS07, DOU08b, DM08a, FP09a, FZ08, Kar05b, Kwo07b, ÖTA05, Sol05a, SYY04, lY08b, YK03]. fibroadenoma [LT07]. field [Ate06, BA07, dCDNPN09, Ece06, ESMM06, GKAS07, KLX06, Lin08, LC08a, MW07, Mur06, OSS09, RAES05, Sel05, SAS05, SSK06, TA06, Wan06e, Wan08d, WS09b, WS05b, CJ06]. fields [AH06d, BZZ08b, CJL05a, GMCM06, MMS^{+}06, MMS^{+}08, MGMC05, MMC06, WHH07]. Fifth [CC08b, NN07c, CC07a, Chu07d, CT07c, EG07b, G07, HC07a, KLW07b, KL07d, KL07g, LD08, RJF07, SG08a, Sal08b, SA06f, SAM07, SAE08, SS08c, uIK06, SA05b, WLK08, WFS08, Waz06b, Waz06a, Waz07b, Waz08k, XZY08]. Fifth-degree [CC08b]. Fifth-order [NN07c, Chu07d, CT07c, EG07b, G07, HC07a, KLW07b, KL07d, KL07g, LD08, RJF07, SAM07, SS08c, uIK06, WLK08, WFS08, Waz06a, Waz07b]. files [Vis08]. fill [HBBA07]. filled [GW06b, LYM09, lSgPpJ07, WZ06f, WSZ07c, WMBY07, YS06b, Zhu05e]. film [BK06a, Hu05, Li07c, Mak06, Soh06]. filter [CC09b, DD09, GZL08, JC07, KOR08, KLS09, LL08c, LM07a, LMN07, Nak08, yNyF05, yN05, yNfM06, PL06b, PFL09, Ran08c, RLdLSF07, SC07c, Wu08d, YDL07, YZY08]. filter-and-fan [Ran08c]. filter-based [Wu08d]. filter-variable-metric [PFL09]. filtering [HL08b, HCLP07a, HCLP07b, Jwo07, Mit06, WS07g]. filters [LML08]. filtration [Ras05c]. final [LLFM08]. Finance [HMNT09]. financial [Day05, MNM07, Nwo07c, Sev07]. find [AK06c, ADDdM07, HN06a]. finder [GSNDB09, GSG09]. finders [HCS05, PP07c]. Finding [JLZ05a, JLZ05b, PSS07, lSgPpJ07, SS07c, ANM06, AK06b, ANM07a, AMN07a, AMN07b, AMN07f, ANM07d, ANM07c, ANM07f, AMN07d, ANM07b, AMN07c, AMN07g, ANM07g, ANM07h, ANM07j, ANM07i, BDE06, CH07e, DSZ08, GSDB06, GSPG07, GSDB09, GDB06, HK06, Hu07b, HHK^{+}05, JMLR05, KY05a, Kop22, MR07c, Muk07c, PB07b, PG08b, PR05, Pet07, PP07b, PY08b, PHP09, RR07, SW05a, SXL09a, SLWX06, Sya05b, Tok06, WR06, YST09, YC07b, Zhu05d]. Finger [Su07c, Su07d, Su08a]. Fingerprint [HHYW07, WHH07]. fingers [Su08a]. Finite [ALL07b, CK05, mCLB08, Dai07, FJCW05, mHsLzC06, JKN09, KE05b, KE05c, Ma07a, MMO07, MV09b, RAP05, SA08d, WZ09a, YM07a, AAN05, AABD07, AKH06, Aks05, AÖ06b, Aks06a, Ami05, An09, AR07a, AV07b, AD06c, AGPJV09, ADK05, ADK06a, AR07b, BR07a, BH07, Bah05, BS05a, BCCK05, Bi06, Bie08b, BS07c, Bra05, CCE06, CA05, CR06, Cha07c, CLSZ06, Che08a, CL09b, CHH09, lCX06, CS09c, CFJ^{+}05, CJL05a, CS07j, Cvi08, Cvi09b, DIT06, DSMK08, Deh05b, Dem05a, DXH07a, dCDLN^{+}07, dCDNPN09, DX07b, DC07, Eid05, ESAH10, Elb05, EEEE05, EO06, EGX05, Ere06, FM05, FA05a, FK07, Gal06a, GJ05, GXY06, GY06, GRK07, GFH08, GK05a, GG08a, Gw05, GG07, Gül05, Gül06, HSH05, He06, mH07b, HZX07, HH08a, HL06b, HXH08, Ide06, IGM07, JGTK07, JZ06b, KP06a, KYK08, KS08a, KBH06a, KBH06b]. finite [KBLT08, KRA07, KYL05, KR07c, KV05b, KNJ05, Kum03, KA06e, Kum09c, KE05d, Li05b, LW06a, Li06b, LMH06, LAL07, Li08a, LL08e, LYJ09, Lia08, Lin05a, LLAT09, LWL06, LL09, LL12, cMpZfR06, Ma08a, MA08d, Mai05, Mar08, MB05a, MY07b, MMS^{+}06, MMS^{+}08, Moa06, Moh05b, MP08c, Nat09, OY05, OAAK08, Ozb07b, ÖEK05, PVV09, PFG08, Pan12, PS06b, PD09, qQcM05, Ram06b, Ram06c, Ram07c, Ram07d, Rom09a, SF07, SS05a, Sak06, Sal07a, Sal06e, SG09a, SZ07a, SH09a, SC07a, SW09a, SP09, SRH09, SP05, SFA08, SM07a, SFH09, SC06b, SS07d, TmChL06, TD06, TD07a, TD09, Tou09, TY07, TD05, WC05a, Wan05e, Wan05a, WHL06, WL07f, WCO09, WYQ05, WS06a, XO07, XC06b, XC07, YLL^{+}05, Yan08e, YJW06, YWJ07, YZLD08, ZS07a, ZO05, Zha05b, Zha05c, ZW08a, Zha09e, ZZKY05, ZZC06, ZC06c, ZC08a, ZWC06, ZWZL06, ZYZ06]. finite-amplitude [CHH09]. finite-data-window [DX07b]. Finite-difference [MV09b, Bah05, Bie08b, Deh05b, HH08a, LLAT09, MA08d]. finite-element [qQcM05]. Finite-precision [CK05]. finite-size [MB05a]. fins [KJ09]. FIR [DDZ08]. fire [AVSB08, BM05a, DM05]. firing [SS08b]. firmly [KT08a]. First [SFH09, AD06a, ADDdM07, BLP07, BG07b, BT08, BSHZ07, CC06b, CT09a, CC05d, DBR07, Den08c, DMH05, DXH07b, DC06e, EDMJ05c, EDMJ05a, GS07b, HEDMJ06, IOS07, JC08a, pLT06, MMN06a, MAM06, MLM07, MS07b, MMA07, MJHED05, MA06e, Moh07c, Mus07, PS08, RMG07b, SH07d, TB09b, VHFF08, WTW08, YS07, ZW09b]. first-integral [Den08c]. First-order [SFH09, AD06a, BT08, CT09a, GS07b, IOS07, Moh07c, WTW08, ZW09b]. fiscal [SYT06]. Fischer [RZX07]. Fischer-Burmeister [RZX07]. fish [Mon07, MBM07]. Fisher [DM08b, EA07a, HKS09a, Jav06d, Nad06a, Waz05p, Waz08a]. Fisher-like [DM08b]. fisheries [MD07b]. fit [CT08a]. Fitted [AR07a, KA05a, KK08a, AR07b, BFKT07, FW07a, KP06a, KPS06a, MKS09, MP08c, Pat05b, Ram05c, Ram05d, Ram05e, SK06b, Sim09b, ZW08a, KYK08]. Fitted-mesh [KYK08]. Fitting [MK05b, Cai06b, CK05, JS05, KNM05, Sam06, VAMV07, WZ06a, MBDD08]. Fitzhugh [LG06a]. Five [Shi05b, MCS08, Sil05c, Sil05d, Sil06a, WD06, Sil06c]. five- [Sil05c, Sil06a]. five-diagonal [WD06]. five-point [Sil05d, Sil06c]. five-term [MCS08]. fix [CMGMD06]. Fixed [DSMK08, MJ07e, ZZZ09, AK05a, AKR08, AR09, ACOS06, AJKM09, AO05a, Ahm05b, AK05d, AHIP07, Ber09, BA06, BT06b, CY07a, CY08a, sCLC09, ĆC09d, GR07a, GTA06, GG09b, HCLPJL^{+}08, IR08a, JOB08, Jun09b, KT08a, yMT08, NCÁHCLP05b, NGLHCLP06, NHCLP07, Nak08, NS06b, OZ07, O’R09, OS09, PP08, QSZ08, SCX07, Son07, SQ07, SYL07b, SCO09, Tan06a, WZDgH09, XZ08b, YKKY09, ZS07a, Zha09h, ZH09a]. Fixed-bed [DSMK08]. fixed-interval [NGLHCLP06]. fixed-lag [NHCLP07]. fixed-point [AK05a, HCLPJL^{+}08, NCÁHCLP05b, Nak08]. Fixing [YRY05b]. flat [AH05c, AK05i, Cor05c, Pan06]. flat-plate [Cor05c]. flavonoid [RNRRMRCH09]. flaws [SL05a]. FLC [OASM05c]. FLDA [xGZZ08]. fleet [GTA06]. Fletcher [yWjL06]. flexibility [IR06, PBG07, SM05a]. Flexible [FSMC05, Vog07, AKH06, IGM07, JGTK07, Kia05, LWL07b, NK07b, RO07]. Flierl [Waz06k]. flight [Mea05]. floating [EF07, MB05b, Tak05]. floating-point [EF07, Tak05]. floor [AVKA09]. Floquet [KS08c]. Flow [AEA05a, KR08, ÖŞ06, Pan08b, PR06, VFF08a, VFF08b, AES05, ADH05, AKH06, AV07a, AVM^{+}09, AH06d, And06, Ari07, AGH05, Att05a, Att06a, AS06i, AH05c, AK05i, BA07, Bat08c, BAK07, CN05, Cha05g, Che09a, CSY07, Chi07a, Cor05a, Cor07, Dem05a, DT08, DJ07, DMZ08, Ece06, Eid05, EESF04, EEEE05, EO06, EHL08, FXY08, FG07, FSÇ06, GH07, Ghi06, GA07, GM05a, Gür05, GD06, HS09a, HGN06, HAAS06, HA07, He06, HZX07, HYL09, HS09c, IHB06, IIS05, IAMA08, JK07, JGTK07, JL08, KAS06a, KAMR05, Kim05b, KBKV09, LWH06b, LC05b, Lin06c, Lin07f, Lin08, LY09d, MNG09, MC07, MCMA07, MS09b, NMS06, Ola06, Ony05a, Oua05, Özt05, Pan06, Pan08c, Pan12, Reb06, SH07b, SH07c, SVMG09, SMH07, SI06, SK09a, Su09, VC06, Vyn07, WNCL06, WY06, WSJW06, Wan08b]. flow [XHSL09, YZ06, ZGH06, ZO05, ZY09, Zhe08, ZL08c, ZS07d, dCdSJM08]. flows [AAN05, BCCK05, Bat08b, Bat08a, CLSZ06, Chi07b, Chi07f, Chi08b, DEH05c, DC06a, DC06b, Ere06, EA08b, KAF08, LL05a, MM06a, NN06a, NCGdSM07, O’R08, Ony05b, Ram07b, RM05, Rao05, RM06, Rao06b, RM08, TY07, XBZ06, XZB06, Yu08c, ZCFX07]. flowshop [CCFC08, EG06b, IGM07, LGJ06, XY11]. flu [EH08a]. fluctuating [SB08a]. fluctuations [Kum09b, WSCL07]. Fluid [AH06d, VC06, AHM09a, AEA05a, AES05, AEG05, Ahm07b, ASH09, AB05c, AV07a, AVM^{+}09, AuH07, AH07c, AGH05, Att05a, Att06a, Bat08a, BBCP09, Buk09, CLA08, Cor05a, DC06b, EB05c, EB05d, EESF04, EEEE05, EO06, EHL08, FXY08, FFKV08, Gar07, GS08b, HAA07, HH09a, HS09c, ICM08, JK07, KAS06a, KAF08, Küç05b, LWL06, MNG09, Mak07c, Mak09, MCMA07, MP08a, MS09b, MM08c, MR09, NMS06, NMM08, Ola06, ÖŞ06, PPS09, Pan08b, Pan08c, Pan12, PR06, Rad05b, Ras05c, RM08, SH07b, SH08a, SM07b, SVMG09, SG09c, SDS05, SAS05, SSS06, SCB07, VSB05, VFF08a, VFF08b, WsZ07a, Wey06, Wey08, XBZ06, XMGW09, YM07a, YZ06, YB05, ZLL06a, ZZL08]. Fluid-particle [AH06d]. fluid-saturated [HH09a, ZLL06a]. fluids [CC08d, IHB06, Mur06, VHFF08, ZY09]. flutter [CL08e]. flux [BT06b, CFF09, KS08f, LYCW09, Sea06, XCM07, XFL06, XF08a]. fluxes [YC06c]. fMRI [XCW08]. foam [HM07]. focal [BW05b]. Fock [Ste09g]. Fock-type [Ste09g]. focus [SH05a]. focusing [Fen06]. Fokker [EA06, Su09]. fold [GKR06, ZHZ07b, QBH06a]. Fold-Hopf [QBH06a]. folding [LLWC07]. followed [Ste09c]. followers [SZL^{+}07]. following [ARK07, BG06, CW08, FY08, FAQ08].

FOM [Lin05d, NR07a, NRA07]. FOM-inverse [NR07a, NRA07]. food [BSP05, Che07c, FLO07, HL05b, LL05f, LZC07b, LGw07, Luo07, LG08, MPS08, MB08, PMS09, She07a, She07b, SS06g, VKVHN05, XCH05a, XCH05b, XW07b, Zen09, ZGL08]. food-chain [Che07c, LL05f, LZC07b, MB08, She07a, She07b, SS06g, XCH05a, XCH05b, XW07b]. food-chains [FLO07]. force [SJ08b, d’O05c]. Forced [SS06f, SS07e, EESF04, GL09, LX09b, PM09, Pan08c, SM07c, WL08b, XZ08a, ZSL05]. forcing [Kut05, PRGH05, SG08a, Zha08b, ZFYH08]. forecast [WH08a, WC05d]. Forecasting [AGS07, DZ08a, Hon08a, HZ06b, JZW08, Sin07b, Sin07c, WML05]. foreign [DD07]. forest [AVSB08, DM05, ZZW08]. forgery [Gop07]. forgetting [DX07b]. fork [ETJ06]. form [AA06a, AD06b, AE07b, BP05, Boy06a, Boy07b, Cvi08, Cvi09a, DG07a, DMA07b, EH08a, GS07b, HRLC08, KW05a, Kay05d, Li09g, LXJL09, cMpZfR06, MYG08, MD09b, OY05, RLY06, Sha05a, SC05b, Van09, WDLM10, Waz06c, WFHD09, jZzFbL06]. formal [KM05b, Yax07]. formation [EP07, LC05b, NN08, PE07a, TMASA05]. formed [NSS06a]. forms [Bri08, Den09a, G07, KES09, KC05b, Ord06a, QBH06a, Sof09, Waz05p, Waz05q, Waz06a, Waz07b, Waz07f, Waz07k, Waz08a]. formula [BK09, BŞ07a, CS07c, EH09c, IOS07, Kıl08a, Li05d, Lin09b, Moh05b, OS07, OISM07, Pod06, QLG08, SK08b, SV06, Wu05a, Wu06b, Wu07a, Yin09, ZPH06, ZT07b, Zha05e]. formulae [AES08, AA06d, CL06c, Chu07g, Cvi09a, Cvi09b, DB07a, FPP06, Iri06, Kah07, MKKS05a, PPV08, PH06, Sim09b, Thu06, Wel08]. Formulas [CGDR06, CL07c, CT07c, HT07, Mar05c, MCS08, OSS09, PB07b, PCS07, RNV07c, Sri09, WLQ06a].

Formulating [RR06]. formulation [Cha05f, CC08d, EGFAE06, GMCM06, MPEL09, Mar08, Rao06b, dCD06a]. formulations [ZS07d]. Fornberg [CLDH09]. Forward [AS09a, CHYC05, GWJ07, CHY05b, HLS05, KT07c, WHC08]. Forward-secure [CHYC05]. Fouad [Pan08c]. foundation [BS09a]. foundations [ÇFÇ06b, ÇFÇ06a, Vel06, Zie09]. Four [FC06, Waz09c, AA07f, BFR05, Boy06b, Deh05a, DW09d, HH09b, Jan08c, JGY07, Kha08b, Li08d, Lou07, MS07a, PB09, PC08, REN^{+}05, Shi05c, Sil05d, Sil06g, SMG08, WDK06, WC05b, ZXZZ07]. four- [Sil05d]. four-bar [WC05b]. four-mode [Boy06b]. four-point [Jan08c, JGY07, Kha08b, Li08d, MS07a, PB09, PC08, Sil06g, SMG08]. Fourier [AKVH05, Bor09, Boy05, Boy06b, BO06, Che05e, Che07h, mCLB08, CDT07, DEH05c, GVMT07, KS08d, Lal09, LPP08, LL06b, Mal06, MSM07b, MS08c, PG07a, QFXW06, Wan07a, YC05a, YKR06, ZCMQ07]. Fourth [HSH05, KLW07d, Kou07a, LX06a, NA06a, YL06c, AMJD08, ASNR06, An05b, ALL07b, An09, AR07a, AM09b, AL06, AKR05, BR07a, Bas08, BM09, CC08b, Chu07a, Chu07c, CH07d, Chu07m, Chu08, CH08a, CH08b, CN09a, DB07a, DMA07a, DMA07b, EM07, FZ09, GL08a, Gen09a, GC07c, HY05, HH08a, Hub08a, Hub08b, HMT09a, Ism08a, KWX08, KLW07a, KLW07c, KL07a, KLW07e, KNJ05, Lia08, LA07b, Ma05c, Mah09a, MR07c, MS06b, MMN06b, MN06, MN07, Nat09, NN07a, PB09, PR05, RG05, RJM07a, RWB09, SR06, SGS09, SXL09a, SM06c, SA07a, SA07d, SA07c, uITA06, SI06, TXW09, WZZ06a, Waz05e, WZ08e, YZ09b, YJO08, ZW07b, ZLZ07, ZL07a]. Fourth-order [HSH05, KLW07d, Kou07a, LX06a, NA06a, YL06c, ASNR06, ALL07b, An09, AR07a, AM09b, AL06, AKR05, CC08b, Chu07a, Chu07c, CH07d, Chu07m, Chu08, CH08a, CH08b, CN09a, DB07a, EM07, FZ09, Gen09a, GC07c, HY05, HH08a, Hub08b, Ism08a, KWX08, KLW07a, KLW07c, KL07a, KLW07e, KNJ05, Lia08, LA07b, Ma05c, MR07c, MMN06b, MN06, MN07, NN07a, PB09, PR05, RG05, RJM07a, RWB09, SR06, SXL09a, SM06c, SA07d, uITA06, SI06, WZZ06a, WZ08e, YZ09b, YJO08, ZW07b, ZLZ07, ZL07a]. Fowler [LX09b, Waz05b, Waz05c]. Fox [BI08a]. FR [qDyC08]. Fractal [IZIO09, Imr09, TSS09, UP07c, Imr06, KVŽŽ08, Nar07b, NM09, PU08, UP07b, WC06c, XX07, ZWZL06]. Fractal-based [IZIO09]. Fractals [GPS09, Zha09f]. Fractional [AC05, BRRGT07, Cha05b, GKAS07, Nar07a, ST09, AH06a, ACP07, AKM05, Amm06, BQ09, BRT07, Cha05a, mCLB08, CLA08, CA08a, DGJ07, DGB08, DEB06, DMDB08, EM06, EB08, EMESES05, EMA06a, EWEA06, Fel05, GD08, GZ07b, GHX08, GS06e, GITV06, IPPT05, JDG06a, JDG06c, JDG06d, JGS06, KRRGT07, KZ07, KL05b, LeT05, Lep09, LD07a, LYX08, LXJ09, LS08e, LTW07, LLAT09, LZA^{+}07, LP08b, MPG07, MBDD08, MSHeT07, MBS07, Mom05b, Mom05a, MAK05, MO06, MS06c, MN06, Mom06, Odi06a, OM06, Odi06b, Odi08, OME08, Odi09, Raw06a, RB05b, RCB06, RB06, RCB08, Ray08, Sal08c, SK08a, SOT07, SCM09, SGB05, SOS07, Sta06b, Su09, TF07, VFF08b, WZT06, Wan06f, Wan07d, WLT07a, WLS08, WS08a, Wu09, YZ06, YZ07d, Yür07, Zha08g, Zha09e, dEdSNL05, Nis07]. fractional-calculus [WLT07a]. fractional-order [ACP07]. fractionary [WN08c]. fractions [Kıl08a, WS08c]. fracture [KTV07]. fragile [LYGL07b]. frame [AHM09a]. frames [LY09a, YZ07a]. framework [Ami06, AGM^{+}07, Chi07b, LLX07a, LY08d, PN06]. frameworks [AV09]. framing [Nwo06e]. Fréchet [Arg09, RA09, ZZYT08]. Fréchet-differentiable [Arg09]. Fredholm [Ord06b, AD06b, AD07, BGA05, BLP07, BMS08, Bad05, BG08, BKM06, CM09, CD06, DC06e, FE08, GKB07c, GG06, Hol07, JRW08, JG07a, LSN09, MK05c, MY06a, MMN06a, MAM06, MR06b, MA06c, MAR06, MLR07, MLM07, MS07b, PMRAS05a, QS06, RMAM07, RJW07, SG05a, SG07a, Sha05a, SH07d, Wan06g, Wan06j, WN08b, XWZ06, YS06a, YSS09, YB06]. Free [Ece06, YAO09, AEA05a, AES05, BCCK05, Boy07b, Bry07a, CN05, CN06, Che06m, CM08a, Chu07j, Chu07l, CH08b, EB05c, EESF04, HMT09a, HMT09b, JK07, JYX06, KLW06d, KLW06c, KLW07d, KL07f, Kou07d, Kuo05, LLC09a, LM07a, LG08, Ma05b, MMC^{+}08, Mor08b, NO09, yNyF05, NG07, Noo07e, NNuH07, Pan08b, Pan08c, PL06b, Raf07a, Ram08c, Ren08b, SS07c, UHuI09, WZ06i, WX06, Xia08, YJW06, Yi05, ZOZZ08, ZH05, Zhu05g, d’O05b]. free-convection [AES05]. free-forced [EESF04, Pan08c]. free-stream [AES05]. freight [XHSL09]. frequencies [GG05a, MRN09]. Frequency [TCMPMC07, DBR07, FTMA07, KV06, ML08a, Vis08]. frequent [JWL^{+}07]. fresh [BOB05]. Fresnel [KVŽŽ08]. Freund [GN06c]. frictional [LG06b]. friendly [Hsu05b, SZS05, WLT05b]. Frobenius [BM06b]. frogging [Geu07]. front [Waz07c, Waz08g]. frontier [AK06h, Ami07, JLZ05a, JLZ05b]. fronts [PW09, Ram07h, Rao06b]. frustum [Che07i]. FS LSSVM [PG07c]. FSS [WLY06]. fuel [ALD08, DN07c, ER06b, ESS05c, Vyn07]. fueling [WS07c]. full [hL08c, Ram06e, WB09]. full-state [hL08c]. Fully [AH05c, DHG06, DH06e, MS07a, Omr06, SH09a, Sha06a, Sha06b, SW09b]. Function [Cha05e, AVKA09, Ada07, AOS09, AMJD06a, ABBP08, AP06, AP09, Ant05, Aou08a, ALN06, BMJED06, hBxY07, Boy07a, Boy05, BO06, CC08a, Cha06, CAL08, CK07a, COG08, CS05c, CAS08, Dam08, DD08a, DMJE06b, DZ08e, DWT05a, EGFAE06, Eli07, EGX05, EELZS05, Fab08, FqZ08, FWJK09, GA08, wG07a, GK06b, GS06c, GS07a, GW06b, GQS08, GL05c, Ham08, HL08a, Has06a, HED06b, mHsLzC06, Hu07a, HCZ06, HMW07, Hus09, Inc06c, IT06a, Isk06, JLZ05b, JZ07, KWS05, KKA05, KG06, KT08b, Kum07, LS09a, LP07b, zLmLL08, LYX08, LL06b, LYM09, Liu07b, LP09, LHW07, LHWW07, cMpZfR06, MH09, MAN06, MJ07d, Men08, MBM07, Mug05, NR09, Neh07a, NHKZ06, OSS09, Öza05, PWL06, PDW06, PN06, QW06, Ree07, RZX07, RAS08, Sal08b, Sea06, SAB08, SGUK08, lSgPpJ07, Sha07b, SCM09, SDL09, Sim07, Sla06, Sol05b, SOS07]. function [Sri07, ST09, SB08b, TLM09, Via08, VL09, WLZ05, WZ06f, WSZ07c, WSL09, Waz08u, Wu05e, Wu06c, WZ06j, cWxChY07, WMBY07, WLL09, XXL09, XY07, XX07, XXX06, XPS07, XZY08, Yan05a, YW05, Yan06a, YS06b, Yan07d, YH09, Yoo05, YRY05c, YZ07c, Zha08b, Zha08c, ZHX06, ZZ07c, ZS09b, ZZYZ06b, ZZ07d, ZXZL07, dSLV^{+}08, MAY09]. functional [AS08, BAF08, BC09b, CH07a, CS07f, CE07b, DK05, DMH05, DXH07b, Fra09b, GAPT08, GO08, GM06b, HH06a, HW08b, HW07b, JOA05, Jia09, JJ09b, Kim08, KPL09b, LC09a, LX09a, LS07c, LH07a, LC09b, LL08f, Liu09c, MM07a, Ma07b, MCGS05, NL05, PS08, PMS09, QBH06a, QBH06b, RX09a, RX09b, She07a, SL07a, Sta06a, Sun06b, SHD07, Thu07b, TG08a, Wan05b, WZ06b, WCX06, WFL06, WC07a, WL07e, WL07c, WTW08, WW05b, WYL06b, WYL06c, WWYL08, WCW07b, XCD05b, XW07b, YCZ06, YC06b, YS07, YBA05, ZZ08a, ZFY07, ZCZ07, ZM09, Zho09c]. functional-type [Thu07b]. functionally [MRN09]. functionals [Deh08, HZ09a, ML09, RAN07, WL07b]. functions [AA06b, AA09, Abd09, AE07a, ANM08, AJMR07, Ahu08, AK09, ARS07a, AD07, All05b, All05c, Alt07b, Aou07, Aou08a, Aou08b, ADK07, AKVH05, Att07, BMS07, BMS08, BD05b, Bil04, Bil07, BS09c, Bog05b, Boy07b, Boy08a, BW09, BR07b, Bry07b, Cao09, CMM07, ÇFÇ06b, Cen06, Cha05a, Cha05b, CC08c, CC09c, CHL06, CK06a, CKOS07, CK07a, CAS09, CMR08, CT06, CCZS08, DNHF05, Den07, Den09b, DjT06, Dua06, EA08a, FBB07, FB08, Fra07, GOPP07, GZ07a, GS06b, Gre06, GQS08, GM06b, HT07, Hon09, HR09b, HL08d, Hub09d, Iri06, Irm07, Jah07, JV07, KV05a, Kan09a, KK07c, KFM08, Kas09, KAP06, KH08, KKA09b, KS06a, KR07c, Kim07b, KS08e, Kim09, KBÖP07, KS06b, Lal09, LLG06, LS09d, LS09h, LS08e, LKC05, LL05g, Liu07c, LN07, LP08b, Liu09b, LA07a, MW09, MSK05, MYN07, MJ07d, MBS07]. functions [MD09b, MA09, MT09, Nis07, NO07, Noo06a, NM08b, Noo08a, Noo08b, NB09, NMM09, NAUH09, Nwo06b, OP09, Ord06b, OSY07, Özk07, PP07b, PY08b, PS06c, PVY09, Qad07, QLG08, QG08, RAN08a, RZ08, RB07, RSSS07, RJM07a, RAP05, Ras05c, RNV07a, RFFFCC07, SOT07, SHY08, SS07b, Sha07b, SW05a, SY07c, uITA06, uIK06, SdHT09, Sok07a, Sok07b, Sok08a, Sok08b, Sok08d, SSR05, SYX08, SEŞ09, Ste09a, SW09c, Sum07, Tan07d, Taw06, TF07, VGMV^{+}07, WL05d, Wan06k, Wan07f, WL09c, WX09, WZ13, WZ05, WW09, WN08c, WS08c, WLY06, nXwCpC06, XB06, YL08a, Yu08a, YL07, YKR06, ZW05a, Zha06a, Zha06b, Zha07f, Zha07g, ZWT09b, ZTWL09, Zhu05e, dlS06a]. functions-II [PVY09]. fundamental [AL09, GS06a, KB05c, Küç05a, TYCL06]. fundamentalists [NR08b]. Fundamentals [Ish09]. Further [CKKS08, Dod08, Liu07c, LHC07, Nwo06d, Par06a, Par07, PK09, Thu07b, yWyCqD06, ZX06a, CFC06c, DX06, Nwo06c, WS08e, Yan07d, YRY05a, Zha08d]. fused [SJ08b]. fusion [JLCH08, KSC^{+}08, WLL07b]. FUSS [CCZS08]. future [WLC05, WLC07b, WLC07c, WLC07d]. Fuzzy [AAA05, BTPGAPRU06, dPCdBYV08, CO06, CH06b, GL05a, GW07, HGH06, Hua06, IR07, KL06a, Liu06d, MR06a, Moa06, MNN05, NNN05, SJ06a, TA05a, YY06a, ZL06b, AJE05, AEJ06, AA06a, AA06b, AE06a, ABA06, AO06a, AJ06a, AJ06b, ASAS06, ASAEE06, ACOS06, AB05b, All05a, AO05c, All05b, AH05b, All05c, All05d, AAAA06, AO06c, AK06g, AY05a, AGA^{+}07, BGA05, BTPGAl06, CS09a, Cha05b, Che06k, Cho07b, CK06c, DHG06, DH06d, DH06e, DCR07, Ema06b, FFRF07, FFRFCC08, FFRFCC09, GK06b, GK06a, GG06, Hu07b, HC05b, HCX06, HLK06, HAS07b, Isk06, Isk07, IR06, JLI06b, JS06, KAV05, KHL06, KP07, KG06, Lak05, LCC05a, LM08a, Lin06b, LL06c, LLY06, Liu08a, MAN06, MMAHN08, NJG08, OASM05c, PPF09, RFFFCC07, RFFFCC08a, RFFFCC08b, Saa05, SM05a]. fuzzy [SGA06, SM06b, Sin07b, Sin07c, SW06c, Tse05, WC06d, WZ06c, WZ06d, WF07a, WYL06d, WLZ06, Wu07a, YCC08, ZW06c].

G [BGP06, Bra09, CK06b, CK07b, GS06f, MX08b, Pan08c, XM08, Zho05, CK07b, HT07, LMT^{+}09]. G-protein-coupled [LMT^{+}09]. GA [OASM05c, EARS07, KF08a, KMN09, KSSA07, LLLX07, Sad07, ZCZX07]. GA-based [KF08a, KMN09, LLLX07, ZCZX07]. Galerkin [CM08b, LL12, AÖ06b, An09, Cai07, Cai09, DSB05, DXH07a, Dog05, DjT05b, Dua06, EG06a, EG07a, Ese05b, EK06, GkB07b, GR06b, GR07b, HBE09, HL06b, Ism08b, KTK06, KBH06a, KBH06b, Kim06b, KM05d, KM05f, LY07a, LY07b, LL09, LXZ07, Mah05, MK05c, ML06, MA06c, MKR06, MLM07, Mar08, MK07a, NKR^{+}08, NHKZ06, Omr06, Ram08b, SD09, SY08b, SM09b, XWZ06, Zha06a, ZOZZ08]. gambler [KCK08]. game [DHT09, gDH07, EMH07, Nwo06b, Nwo07c, RR06, Rey05a, Tan07a, Tan07c]. games [AH06c, AR06, Ish09, PSP05, SMY08]. gamma [FB08, GQS08, GL05c, Nad05, Öza05, SSR05, XZ07a]. gamma-type [GL05c, SSR05]. GAOR [Yua05]. gap [BAK07, GANG07, Kim05a]. gaps [LS09h, Ste09a]. GARCH [Nwo06d]. GARCH/ARMA/VAR/EVT [Nwo06d]. Gardner [Waz08v, XMGW09, XL09a, ZDY08]. Gardner-KP [XL09a]. garment [XL06b]. Gas [VRG06, BP06b, CLSZ06, HS09a, PVS^{+}08, Vyn07, ZYH08]. gated [XM08]. Gauss [Pre06b, RVNI07, Wan08c, AF07, AT07a, AAAA06, AA07e, BJE05, BDE06, CL05c, Dan08, DMJE05b, EF05, EDMJ05a, EMJB05, Has06a, HED06b, KS05a, KLS09, LC07, ML05, MJED05, Men08, MBS07, Pre06a, RVN07, RNV07b, RVNI07, RNV07c, yS06c, TD07b, WLS08, yXhWZ06, YEA07, Yun07b, ZHL07]. Gaussian [AO05c, AO06c, BH06, BW09, BO06, HC08c, JS05, KK07c, KBMZ08, LM08b, Ser05, SM09a, ZC05a, ZL05a]. Gautschi [QG08]. GCS [Par05c]. Gear [DJ06a]. Gegenbauer [JP09]. gel [DV06]. Gel’Fand [Kin06]. GEM [ASPA09]. gene [PGdSS08, PG07c, QSS08, YSXL08]. General [AV09, FPP06, yLcCsL09, MX08b, QN08, SQ07, ZW06c, AD06b, AE07b, Ars07c, Ars07d, AAA05, AM05, BN07, BN08, CZ07b, CW05a, CW05b, EMR08, EMR09, ESMM06, GSG06, Haj06, HR07b, HYG07, KP06a, KPS06a, Kar05b, Kar07a, Kar09b, KW05a, KKA09a, KKKM09, KV06, LT08b, LML08, LZC07a, MA05g, MF09, Mor08b, Muk07c, yNfM06, NB07, Noo08b, Ouj06, PRA^{+}09, QSZ08, Rad09, RGA06, SWK^{+}06, Sal08b, SS06a, SAD06, SG07a, Sha05a, Sog07, TK06a, Vel06, WJ05, WWP05, WZZ06a, WZSZ06, WL07d, WLT07a, WL08a, Wan08e, WC09b, WL09c, XF08a, XS09, YS06a, YCZ06, YC06b, YWL07b, bZlDL09, ZH08a, IRSR05]. generalised [ICM08, RR06, dEdSNL05]. Generalization [ADZ05, TM07, WCX06, dSYY08, Att07, Bul09, CJ08, HH05, Kad07b, KAO08, LN08a, Qad07, QCN08, She05c, UE07, ZZ09, d’O05c]. generalizations [SNK06, WS07i, lSAQ08, Sin08b]. Generalized [HG09, HAU05, Jah07, Kha07b, Kha08a, Kha09, KKY07b, KVŽŽ08, Kos07, LV05, LD07c, LZ08c, Li09b, LZ06e, Mah09b, OS07, OME08, Sad09, Sal06f, Siv07, UP07a, WSLZ07, WCZ06, WL06d, WC06b, WCO09, Waz05i, You06c, ZWS08, ZLW09b, ZZZ07, ZYCI08, AR09, AA09, AMJD06a, AB05b, ASK05, AH06a, AKMA05, Aou08a, AH08b, Att06d, AS09c, BK08b, BLW06a, BNH07, BE06, BS09c, BJ08, BSHZ07, Bul09, CS05a, CL05a, CT05b, CY06a, CLA08, Chi07d, CKK05, CK07a, CKQ08, CAS09, CI09, DH06a, DHY06, DH08, Dem06, DL08, Den08d, Din05, DLD08, DSZ08, DF09, DW09c, DM08b, Djo07a, DW09d, DZ08e, EKM08, EHAR05, ESD07b, FTG07, FC06, FqZ08, FFKV08, FS09, GRK07, Gha08, Gre06, GC09, GM05b, HAK07a, HFL06, HX07, HM08a, Has06a, HDG07, HNB06, HRLC08, HRCL08, HZ08, HW09, HZ06d]. generalized [Hub09b, ILZ09, Jav06a, Jav06b, Jav06d, JK08, Jia05, JSL06a, JGS06, Kas09, Kay05c, Kay05b, KB05a, KK07d, KM09b, KAP06, KAT06, KD09, Lal09, Lay09a, zLmLL08, LHZ09, LPS08, Lin06d, LC06b, LCL05, LCC05b, LYW09, LYH09, Lou07, LH08d, MLW09, MTM05, MSCG09, MV08, MAY09, NMS06, Nač09, NR07b, NK05a, NCGdSM07, Neh07a, NM08b, NB09, ÖTA05, PR09a, Pog07a, PDA06, RJ07, Rom09b, SNH07a, Sal08a, SGC09, SB09, SPV08, SZ07a, She05a, SXX05, SJ06c, SX06, She07c, SC08b, Shi05a, SS08c, Sin05, SKS06, Sin07a, SZ07b, ST09, SP06e, Sun07a, SB08b, Tan05b, TL06, TH07a, TZW09, TZH09a, TXW09, TZH09b, TS08a, Ver07, VFF08a, WL05e, WC06c, yWjCnG07, WM07, Wan07e, WZSL08, WX08a, WSL08, WZDgH09, WCZ09, WT09, WZ13, Waz05e, Waz05p, Waz05q, Waz06c, Waz06g, Waz08l, Waz08u, WZ08d, WP06, Wu08c, XD07]. generalized [XG08, XQ06a, Xin06, XX07, XRZ09, lYqL06, YMCL06, lY08b, YLX08, Yu08b, ZW09a, Zha05b, ZMW05, ZX06b, Zha07c, ZX07a, Zha07d, ZT07b, ZZ07a, ZDY08, ZWT09c, ZZH09, Zha09c, ZTWL09, ZXLS06, ZX07b, ZLR07, ZW07c, ZL08d, ZZ09]. generalized-Zakharov [zLmLL08]. generate [Wu07a]. generated [Aki05a]. Generating [KH08, Cao09, Gre06, HH07a, IT06a, JF05, JHST05, WC06d, Wan06l, ZWD07]. Generation [ADH06, Elg09, PB07b, BB06e, CMS05, Özt05, RMA07b, WYX05, WTG07, ZWZL06]. generator [ÁCTZ05, KJ08, Shi05b, YY09]. generators [CT08a, KT07c, Liu07b, ONG08, PJ10]. Genesio [PKL08b]. Genetic [BS05c, CCFC08, JRA^{+}07, PE07b, AH08c, AGTS07, APS05, BAB08, DT07a, DT07b, DD08a, DD08b, DSKM09, DN07c, FZN07, Fat06, GC06b, GAFK07, GHZS07, GG05b, GBG07, Ham08, Har06, HJB06, HBJ06, HN06a, HH06b, HTO06c, KTW08, KY06, KTÖA08, KB07, LMT^{+}09, LML08, LLX07b, Lin06b, Lin06a, LG07, LL06c, LLY06, MK05b, MBM06, MM07c, MKSK07, NN06a, NJG08, OHT05a, OASM05c, OASM05a, OASM05b, PN06, PN08, PLTW07, PFA07, RK07a, RKS08, SSB06, SCR08b, SCR08c, SM06b, SCS05a, Tso08, Tso09, Wan05d, WTW05, WLLL07, WH08a, WZ05, nXwCpC06, XCS07, YHL08b, YHL09a, PBG07]. genetic-conjugate [GG05b]. GENLS [MED08]. genus [CJL05a, KB05c, Küç05a].

Geo [hLsT07, XM08]. Geodesic [BBMN07, CCWH09, KYA05, KA06c, KAO08]. Geodesic-pancyclicity [CCWH09]. geography [PVM07]. Geometric [JM07, KK07a, KH07a, Sai07, GY08, GS09, Hos07, IR06, IR07, Isl08, Liu06e, Liu07e, Liu09d, QZJ07b, SM06a, Sha07b, She05a, SJ06c, dCRDSSN^{+}08, Sok08b, WL05e, WF07a, YCYL05]. geometrical [VCRCG09]. geometries [SFH09]. geometry [Che05e, KSKK06, Kös06a]. Gerber [hBxY07]. GGM [CHL06]. GGM-construction [CHL06]. Ghodousian [MK10, MK07b, MK08b, Zim07]. GI [hLsT07, Zho05]. Gibbard [Tan07b, Tan07d]. Gibbon [XZY08, Waz08k]. Gibbs [AH08a]. Gilpin [BW05a, Che06h, LZC07a]. Ginzburg [GB07, LZ09d, SDL09, Tan05b, WGL08]. given [KT08a, MFE05]. giving [VGMV^{+}07]. GKN [ZC06d]. Gl-LSQR [TK06a]. gliding [HES05, HAM04]. Global [Alo09, AM09a, AAA08, Bai08, BR08a, CLG09, CZH07, Che05a, CS06a, CW06a, Che06b, Che07d, DDR06a, DC06b, qDyC08, FK06a, GY07b, HL07c, Hua05b, JSL06b, JWS^{+}09, JGS06, KOR08, Laz08, LCM08, LTC07, MSD06b, MSHeT07, OM08, Par06b, Pen06, Pha06, SSA09, SW06a, SL06a, Ste09b, SL05b, TK06a, dWqL05, yWjL06, WZ07b, WS07e, WZ08b, Wan08f, WX08a, WL09a, XCD05a, XCD06, XLL06, YLZ06, YLL^{+}05, YZW07a, Yi05, ZY08b, ZzLZ08, ZSCW09, ZH08b, ZXY06, Zhu08, AOS08, AK08a, AKB09a, AKB09b, BV06, BS05c, CL08a, CCWJ08, CS07e, CXC06, Che06e, Che06f, CY08d, DDR05a, EHR09, FO09, GCT08, GHX08, GH09, Guo09, HK06, HR07b, HMW07, HL05b, JZQ07, JSL06a, JH09, JS07b, JC08b, KR07c, KL05b, LT08a, LKW06, Li09c, LZC07a, Lin05d, LZZZ06, LC08a, LZ07d, Par07, PK08c, QZJ07a, QZJ07b, RB08b, Rös07, SA09, lSgPpJ07, SHK06, She05a]. global [SW05a, She07a, SZ06a, Sim09a, Sin08b, Sin09, SXL09b, SG05d, SC07c, Tok06, THLC07, Tso08, WL05e, WZ06f, WSZ07c, WS08a, WZ05, Wu05d, WZ08f, WLL09, nXwCpC06, XZ07b, YL06a, YSZG05, Yan05d, YS06b, YDL07, YZW07c, YHL08b, YHL09a, ZWX05, ZLL06b, Zhu05e, d’O05b]. Global-best [OM08]. Globally [LFC08, WL07e, YGW09, AA08, Ash06, Chi07e, KH05a, MTC07, Ma08b, MCW08, QZZ07, WWWL07, YPZ05, Zhu05g]. Glucose [CMMC09, FA05b, NAM^{+}07]. glycerol [WFX07]. GMRES [HIPA08, Lin05d, NZ08]. GMRES-based [HIPA08]. GN [SC05c]. GN-authenticated [SC05c]. GNG3D [NTZ08]. GNLS [MAY09]. go [NR09]. Goal [JA08, Çak08, Cha05b, Cha05c, Cha06, HGH06, Isk06, Isk07, RR06, RAS08, Saa05, SGA06, Wan06l]. Gödelian [BKS^{+}06]. Goertler [LC05b]. goes [BCC^{+}09]. gold [CSBS07]. Goldstein [DSY08]. Gompertz [GGSNR06, MR05b, WWC06]. Gomportz [LMC08]. good [Che05e, CL08b, SB05a]. goodness [CT08a]. goodness-of-fit [CT08a]. Gordon [WZ08c, Wan08d, Zha05a, BQYW07, BZZ08a, CY08c, FLLH07b, GZ05, GHL07, HC09, HZ09b, Hub09d, Kay05b, KE05a, KLX06, KH09, qLJ05, Ma07c, MD08a, Ray06, SB09, SMG07, Sun05a, TH07a, WW05a, WC05a, Wan05e, Wan06e, WZ07c, Waz05n, Waz05m, Waz06o, Waz06p, Waz06q, Yüc08]. Gordon-Equation [Hub09d]. Gottwald [XXL09]. Goursat [AA09, Bel07b, Waz07j]. governance [Nwo06a, Nwo07a]. governed [BS07c, Gal08b, MD08a, ZW09b]. governing [FTT^{+}09]. GPS [KAM06, TK06b, TK06c]. grade [ASH09, AuH07, AGH05, CLA08, FXY08, Mak07c, Mak09, MP08a, ÖŞ06]. graded [MRN09]. gradient [AJE05, And09, Ars07d, BMP06, BS07c, CS07i, qDyC08, Ewa06, Fig08, GG05b, HYG08, HWS07, HT09a, HT09b, KA09b, KF08c, xLyW06, LYKC07, LLFM08, LY07d, LXZ09, PVN06a, Rig08, SWH07, SS05d, SS06c, mTxWyL07, yWjL06, jW06b, WHH07, WXW09, WYL06a, WLQ06a, YZW07b, YZW07c, YGW09, ZYH08, ZSCW09, Zha09b, Zha09d, Zhu07, dCdSJM08, HL06a]. gradient-based [WHH07, ZSCW09]. gradient-type [BS07c, xLyW06]. gradients [KR08]. granular [HS09a]. granularity [HZW06, JWL^{+}07]. graph [AASAA06, Cas07a, Cas07b, EH09d, HKH06, HHC08, HC08b, Lin07d, PK08d, RMA07b, SB06a, XLHK09, Zho08a, Deh08]. graph-theoretic [RMA07b]. graphical [ZWD07]. graphics [WYD07]. graphs [CY08b, CHT06, DESK06, DES07a, DES07b, DESG08, ESD07a, ESD07b, KH05b, Li05e, LCX09, ME05, RH08, TTH08, UP07b, XHZ06, Yan09a]. grassland [LW06b, WL05a]. grating [KS08d, KS08c]. gravitational [BK07b, Rad05a, SBS08]. Gravity [Köp07b, Bat08a, Köp07d, MGG^{+}09, Vin09]. gray [Hos07, Hos08, Su07a, AS07b]. gray-scale [Hos08]. grazing [CRC08, LW06b, WL05a]. Green [CC08a, Cen06, mHsLzC06, Hus09, Ony05a, Ony05b, PDW06, SV06, WSL09]. Grey [Hu07a, SCS05a, Tie05, WH08a, WC05d, YY07c, ZFL^{+}09]. GRH [Dön06a]. grid [AC07a, Bie08a, Bie08b, BX09, CS05b, CL09b, CMS05, CMR08, HL06b, KK07b, KR07b, KK05a, cMpZfR06, MH06a, Özt05, qQcM05, Rao05, SSW09, Sil07c, TD05, WHL06]. gridding [SM07a]. grids [BB06e, Sil05c, Sil05a, Sil05d, Sil05b, Sil06g, Sil06c, Sil06d, Sil06a, Sil06f, Sil06b, Sil06e, Sil06h, Sil07b, Sil07a, WZZ06b, YJ09]. Griewank [COG08]. Gröbner [IL08]. Gromov [JLA07]. Gromov-hyperbolic [JLA07]. Gronwall [ADZ05, Wan07e]. Gronwall-like [ADZ05, Wan07e]. gross [Chu05]. gross-domestic [Chu05]. Grossberg [Li09c, CZF09, JSL06a, LYJ09, MGH08, WZ08a, Wan05c, WRL06, YX06a, YZW07a, ZD06b, ZW06a]. Grossberg-type [Li09c]. Group [AC07a, BZZ08b, BLZZ09, CHY05b, EM11, KR09, WF07a, ZD06a, BÖ05, BA05b, CY08b, CHC05, DL05, GT05b, II07, KM09b, KB05c, Küç05a, LZ06c, LZ07a, LY07c, LLL09, LWK05b, Ma07c, SS08a, SC09, SLO08, TD06, TD07b, TD07a, XL09a, YSCC05, YY05b, YC05b, YC08c, ZW05c]. group-invariant [SLO08]. group-oriented [CHC05, GT05b]. grouping [YK06a]. groups [AAG05, LDZ06, Lin07c, PPS09, Sht07]. growing [LMD^{+}08]. growth [AK05b, BT05, BT08, Cai06a, Dam08, DK05, Far08, Kim09, Liu06f, MKA07, Mon07, SKSJ06, ZY08d, ZYD07]. GSAOR [LCH09]. GSOR [LK08]. GSOR-like [LK08]. GSPN [GI06]. GSWW [Waz08u]. GTST [XY07, YH09]. guaranteed [Kos07, PK05b, Par09b, Yu05]. guessing [Shi05a]. Gumbel [Nad06b]. gun [LYCW09]. Gurtin [Kim06a]. gypsum [BM05a]. gyroscope [GL05b].

H [Kop22, KS08c, LCC06]. H-matrices [LCC06]. H-polarized [KS08c]. Haar [HKS09b, HKS09a, HW07b, Lep06, Lep07, Lep08, Lep09, MPEL09, MM05b, Ord06b]. Hadamard [Aou07, CS09a, Irm07, KBÖP07, LS09h, Sok07b, Ste09a, TS08c]. Hadamard-type [KBÖP07]. HADM [Jan07b]. half [Ahm05a, Ahm07a, Ahm07b, EB05d, GRH07, KW08a, KW08b, LZ08d, SKS06, ST06c, Sin08a, SM08a, SM08c, TS05, WLW08, Waz07k, XZ08a, ZLW07]. half-infinite [Waz07k]. half-line [KW08a, KW08b, LZ08d, WLW08, ZLW07]. half-linear [SM08a, SM08c, XZ08a]. half-plane [EB05d]. half-space [SKS06, Sin08a]. half-spaces [ST06c, TS05].

Hall [GS08b]. Halley [Chu07a, Chu07j, Chu07l, CH08b, GSDB09, KT07a, KLW06c, Kou07c, KL07a, KL07d, KL07e, KW07, KL07g, KLW07e, LLC09a, NKH07, NN07b, PM05, RAZ07, WY07, Xia08, XL09c, ZW08c]. Halley-like [PM05]. haloes [Cai06b]. Halpern [Son08a]. halting [Tan08a]. halving [WLCL06]. Hamel [MM06a]. Hamilton [BS06, ZW05g, ZZ07e]. Hamiltonian [AS07b, CHT06, CT09a, FF06, KH05b, KLL^{+}08, LSW06b, LSW06a, LYX08, LMZ09, ML08b, ML09, NM09, Ser09, TmChL06, Wan06c, Wan08c, Wan09a, XHX07, XT08, YX08, YZ08, ZC06d, ZXZZ07, ZXZL07, ZXLZ07, ZX08]. Hamiltonians [ZL08e]. Hamiltonicity [TTH08, YLL06b]. Hammerstein [BFR07, MKA05, MD06a, MKR06, MD07a, Ord06b, RZ07a, YZLD08]. Han [ZSNS05]. hand [GY07b, KT06, TK06a]. Handling [Çet07]. Hankel [Boy07a, GRK07, WLZN08, Wu08a]. Hansen [SGS09]. haplotype [QYYL08, WWC09]. hardware [JY07a]. Hardy [AS09b, FBB08, KS06a, LS09h, RFFFCC08b, Ste09e]. Hardy-type [RFFFCC08b]. harmful [SMBB07]. harmonic [Ahu08, AK09, Ate06, BS07c, GÜ07b, JCX^{+}07, LZZZ06, LZZ06b, LZ07c, NM08a, Sof09, TWZ07, XXC05, XXC06, ZLW09a]. harmony [Gee08, MFD07, OM08]. harvesting [AMN07c, AAA09, BW05a, BW07, BI08a, JPCL08, LGw07, Luo06, Luo07, Mak07b, MBM07, MD07b, SMM08, ZWZ05]. Hash [ZW05a]. Haskell [ML07]. Hassell [PMS09]. Hausdorff [Jia07a]. having [HK09, KJ09]. hazard [Mug05]. hazards [KOS08]. health [Whi06]. healthy [FAQ08, SDM^{+}09]. heartbeat [DBR07]. Heat [ASH09, NPR09, PVS^{+}08, AEA05a, AEG05, AH05a, ACZ05, Ang05, Asa06b, Att05a, Att06a, BA05a, BT06b, Che06n, CFF09, CKS09, CSC08, Cor07, DLLN09, DSK08, ESMM06, EEEE05, EO06, HBE09, Hsu06, HS09c, JZ08a, KBLT08, Kha09, KV05b, KK05d, KS08f, LYCW09, Li06a, Li09a, LLW05, LX06a, LX06b, MM05a, MC06, MY07b, MHM09, Mom05b, OIL09, Özt05, Pam05a, Pan08b, Pan12, QS06, QFX06, QFS07, Qui09, SK05, SZ05, SP06a, SM05c, SK09a, SG09c, TT09, Tua13, WY06, Waz05p, XFL06, XFQ06, XF08a, XL05, YM06, YC06c, ZQ08, ZMY08]. heat-conduction [Li09a]. heat-diffusion [KV05b]. heat-like [Mom05b]. heated [CLA08, Mak06, MR09, SDS05, SSS06, SM08d]. heating [EO06, MM07f]. Heatlet [BHN08]. Hedging [Hor08]. height [MMS^{+}08]. helium [KWS05]. helix [KY05b]. Hellman [YRY05d]. Helmholtz [BÖ05, DEH05c, mH07a, Mar05b, QW08a]. Helmholtz-type [Mar05b]. hematopoiesis [Rös07, WL07d, WL08a, YSZG05]. hemi [GG09b, WCO09]. hemi-continuous [WCO09]. hemi-pseudocontractive [GG09b]. hemiequilibrium [Noo06b]. hemivariational [CS09b]. hepatitis [PG07d]. hereditary [SV07, XL08a]. hereditary-type [SV07]. Hermite [AA06d, CS09a, CS09c, DLBT08, Gre06, Has06a, KH08, LO08, gLxW05, LL07b, MM06a, Mak07c, MY06b, NBHYV09, PC09, SL09b, WZ08d]. Hermite-type [SL09b]. Hermitian [HY07b, Pen05, XL09b, ZHZ08, ZH08c]. Herschel [VSB05]. Hessenberg [EH09a, EH09c, Hos05a, KW05a]. Hessian [ZL05a, Zhu05a]. Hestenes [BGR05]. heteroclinics [FT07]. heterogemous [WGCH05]. heterogeneity [HHZ09]. Heterogeneous [NR08b, OKEL05, DHT09, GTA06]. Heun [HRS07, HR09b]. Heuristic [Oss05a, WS07c, GÜ07b, HK06, HT06a, KT07c, MCAF08, Oss05b, Sal08d, SAK08, TGSNB07, XLSL06, ZGK06]. heuristics [Azi05, JGTK07]. HEX [Tan07a, Tan07c]. hexagonal [LF08, Shi09a, YTLH06]. Heys [ORG^{+}09]. Hidden [TX08, Yuj07]. hiding [CHY^{+}09, Vis08]. hierarchical [Che07h, DES07b, DLD08, KBLT08, LKW06, LYGL07a]. Hierarchy [WY05a, CH05b, DW09d, GHZS07, HT06b, KS09b, Kud09, MX08a, Ozd05, PB07a, SMB08, YX08, YZ07c, YZ07d, YZ08, Yu08c, YL09b, Yu09, YL09c]. Hierarchy-oriented [WY05a]. High [AV07b, CG05b, FH07, GKR06, GG08a, Li05b, pLT06, Pat05b, Pat07, PMP08, RLZ08, SH06b, Sim09b, WW05a, WZZ06b, WL06f, AN07, AS06c, CC09b, Cha05g, CZF09, CSC08, CTL09, Cot08, CCZS08, DM06, Eng06, EHR09, GS05, GSG06, HM06, HLT09, JS07a, KK06, KR08, KH07a, KSM09, KK08d, LW06a, LS07d, MO05, MB06, MH06b, MH06a, Mar08, Moh07a, ÖA05, PL06a, PLZ05, PSD07, QFXW06, QFF06, Sal07b, TD05, Wan06g, WH08a, Waz08r, WRL06, XCFX05, XB06, XLL06, YSS09, YLB05, lY08b, You08, ZPH06, Zha05e, ZL06c, ZCZ08, ZL09a, ZYZ06, LJ07]. high-dimensional [Waz08r]. High-order [AV07b, GKR06, Li05b, SH06b, WW05a, CZF09, GS05, GSG06, HLT09, KH07a, LW06a, Wan06g, WRL06, XCFX05, YSS09, lY08b, ZL09a, ZYZ06]. high-pass [CC09b]. high-performance [MO05]. high-resolution [ZCZ08]. high-speed [ZL06c]. high-yields [AN07]. Higher [BŞ07a, Eid05, Kas09, Kum07, AÖ09b, CD05, Cvi09a, DMS06a, DMS06b, DDR06a, DBC^{+}07, DDR05a, DDR06c, DDR06b, ESAH10, Elg09, Eli07, Fen06, FS09, GZ09a, GZ09b, GRH07, HHK^{+}05, Hub09c, JZ08c, Kar09c, KY05a, Li05b, LXML09, LCL05, Liu06h, LKU07, Liu09a, LWKU09, NMD07, PDA06, SAB06, SS07a, SS09c, SH06a, SZ08b, SS06f, SS07e, SM07c, Waz06n, Waz09d, WZ06k, WH06b, XHH06, YLZ06, ZZ05a, ZFY07, Zha08f, ZYW08, Zha09c, ZWZZ09, ZY07, Zho08b, ZZ08b, ZX08, Zho09b]. higher-dimensional [Fen06, KY05a, Waz09d, ZY07]. Higher-order [Kas09, AÖ09b, DMS06a, DMS06b, ESAH10, FS09, GZ09a, GZ09b, GRH07, Kar09c, LXML09, Liu09a, SS09c, SZ08b, SS06f, SM07c, Waz06n, WH06b, XHH06, Zha08f, ZYW08, Zha09c, ZWZZ09, ZX08, Zho09b]. Highly [Zha07h, LWWS09, MK08a, Pet08b]. Hilbert [Gen09a, Gen09b, HG09, Li09b, MA05g, QN08, QSZ08]. Hill [CC08a, SM05b, SWC05, ST05b]. Hilliard [BN09, DN06, YC05a, YLH09]. Hindmarsh [WC06b, WS09a]. Hirota [CY06a, FqZ08, Ji08, SXX05, Waz07e, Waz08c, Waz08d, Waz08e, Waz08u, Waz09b, WP06, jZzFbL06]. histogram [XYZ09]. histogram-based [XYZ09]. history [BM06a]. history-valued [BM06a]. HIV [KRG06, Kov07, Kwo07a, NDD05, TNS07, WS07e, ZSS08b]. HIV/AIDS [KRG06, Kov07]. HLS [WDL07]. HMM [LLX07a]. hoc [HRM07]. Hölder [BLLM08, BDLS05, RZ08, RYW06, RA09, Wu08c, ZW08c, ZZYT08]. hole [ASL07, AA09, Cha05g, CLW09b, CB06, Dön06a]. Holling [Che06i, CS07f, HCZ06, JMC07, LX09a, PWY08, She07a, SS06g, WW05b, XCD05b]. Holling-type [XCD05b]. hollow [Lee05c, LWH06b]. Holm [Waz05d, HRCL08, TXW09, Waz05e, Waz05j, Waz06l, Waz07f, XXL09]. holomorphic [SW09c]. Holt [DAQ09, lSAQ08]. Homoclinic [FF06, FZ09, LZ09a, SH05a]. homogeneity [Sae07]. homogeneous [Jan07a, Jan07b, Kıy06, LBS05, Liu08b, MD09b, SML05b, Sae06b, Sey05, SM05b, TWZ07, TLB07, YL09a, ZMV05]. homomorphisms [JR07, Sht07]. Homotopy [Abb06a, CL08e, CA08b, LXJ09, Wan07d, Yüc08, Abb06b, Abb06c, ATL07, All07a, AC08, AGH05, CO09, CX09, EL05b, FH06b, FHG07, GKB07a, GKB07c, GkB07b, GKD07, GJ07b, GJ07c, GJ07h, HDG07, HH09a, JG07a, Li05a, LBS05, Lia05, MHM09, Odi07, RGD07, RDGP07, Ram08f, SZ08a, Sun05a, WSL09, Wu05a, Wu05c, Wu06b, Wu06c, YHW06, Abb06d]. honey [FAM07]. honey-bee [FAM07]. honeycomb [YTLH06]. Hopf [Chr08, Dil05, GCT08, GS08a, Gor06b, HH09b, LL08a, LX09a, LW09c, MG09, NH07c, QBH06a, QBH06b, QN08, Waz09b, YL06a, Yan07b, Yan09d, ZXF07, ZLY08, ZSS08b]. Hopfield [Bai08, HH08b, KB06b, WiN06, Wan09b, XLL06, ZW08d]. Hopfield-type [KB06b]. horizon [AKH06, AK08b, CPSS08b, IGM07, JGTK07, KOS08, LG08, LPW08, MMM06]. horizons [BS08]. horizontal [AK05i, Pan06]. Horseshoe [jF09, LH08c, YLC09]. Horvitz [BCEGMRMP07]. host [AOS08, MA05e, TML07]. Householder [CM08a, NNM07, NG07, SB06b]. Houston [EMH07]. hp [EGX05, KBH06b]. hp-finite [KBH06b]. HPBILc [XYZ09]. HPM [BG08, GJ07e]. Hsu [HL05c]. HTA [DLL09]. HTLV [SL06a]. HTLV-I [SL06a]. Huard [AT07a]. Hubbard [PRA^{+}09]. Hubbell [GM05b]. hull [Abd06]. Hulthen [Dem05b]. human [CYL^{+}08, NMS07, TML07, XCW08, YSXL08]. Hummel [GSDB07a]. Humor [ORG^{+}09, AS06i, AS07c, MH08a]. Hunter [Boy06b]. Hurwitz [Ada07, Boy07a, CS05c, CAS08, CAS09, GGOS08, Li07a, Sim07, lS07b]. Huxley [BNH07, Den08d, HDG07, HNB06, Jav06a, Jav06b, KS08b, Waz05q, Waz08a]. Hwang [SCS05b, ZK05, Hsu05a, HH05, HL05d, LW05c, QCB05a]. HWWM [LKY05c]. HWWM-authenticated [LKY05c]. Hybrid [Azi05, CY08a, Fad05, HP07, KR07a, Lee05c, LWH06b, LPA^{+}09, Liu06f, Sit06, WTW05, AS08, Ars07d, AKVH05, BTMB09, Bou06c, Cen05, ĆC09d, DD08a, DZ08e, FW07a, FM05, FWJK09, GTJ07, GG05b, HW07a, HZ06b, Ide08a, JM06a, JGTK07, lLWhL06, hL08c, Li09d, LHB06, Lin06b, Lin06a, LG07, LLY06, MB06, MR06c, OHT05b, PGdSS08, PG07a, TKA^{+}08, TGSNB07, TLB07, TMJV^{+}05, TMSG06, Van07, Wan05d, WLT05a, Wan06k, WH06a, Wan07f, WL09c, WZ05, XZ09a, YNCL08, YM08a, Yax07, YYWW07, YHL08b, YHL09a, ZGH06, ZZLL07, ZL06b, ZCFX07, lS07a, KV05a]. hybridized [SSJK07]. hydrocephalus [SSTD05a]. hydrodynamics [DK06, Dön06b]. hydrofoil [GK09]. hydrogen [SLX05a]. hydrogen-bonded [SLX05a]. Hydromagnetic [NMS06, Rad05a, EO06, HA07]. hydrostatic [SKS06, Sin08a]. hydrothermal [BGRS07, BGRS09]. hyper [KH05a, WS05a]. hyper-exponential [WS05a]. Hyper2d [CVMM06]. Hyperbolic [LLX05, SS08b, ACJ09, AY05b, BE05, Bou06a, Bou09a, DSK08, DZ08b, DZ09, GC07a, Hsu06, Ism05, JLA07, KK06, Kim06b, KL05b, Köp06, KM05f, KK08d, LY07a, LY07b, LLW05, LL09, LLC09b, LL12, MC06, Moh05a, Moh05b, Moh07c, QM05, Ram07d, Ram07e, RJK07, SY07a, SHY08, Tou09, Waz06l, Yan08d, Zah09, ZZYZ06b, ZCZ08]. hyperchaos [Yan05b]. hyperchaotic [CAL08, Yan05b, ZWLX08, Zhu09a]. hypercomputation [Dav06, DC06d, Ord06a, Smi06b, Sta06c, dCD06b]. hypercomputational [BD06b, Coo06]. hypercomputer [dCD09]. hypercomputing [BKS^{+}06]. hypercube [KK05b]. hypercubes [MLP07, XHX07]. hyperelliptic [CJL05a]. Hypergeometric [Sha07b, Ahu08, Aou08a, CK07a, EMC05, HT07, HR09b, Neh07a, Noo06a, OSS09, Sri09]. hyperharmonic [DM08a]. hyperinterpolation [CVMM06]. Hypernumbers [Köp07c]. hyperparameters [ZL05a]. hyperparasitism [Kum09b]. hyperplane [wG07a]. hyperprismatic [TD05]. hyperrectangles [Chi06b]. hypersingular [AA07e, Fad05]. hypersurfaces [Öza05]. Hypotheses [Zie09]. hypothesis [WZ06e]. hysteresis [Bel05, DEB06, Dar06].

IAs [Che06l]. IBVP [Ism05]. ICA [YZZP08]. ICS [ZZL05]. ID [Chi05, LC05a, SW05b, SCL05b, ZSNS05]. ID-based [Chi05, LC05a, SW05b, SCL05b, ZSNS05]. Ideal [VL09, BFR05, CFX05, CCC07, Gar07, WL06e, Wu06a]. idempotency [ÖSÖG09]. idempotent [SÖ08, ZW07c]. idempotents [CGVC07]. identical [OT06, WZ06g]. Identification [FWJK09, LT09, Bog05c, DCL07, DLD08, DDZ08, DMDB08, FDL08, GFWX05, GL05b, JLS^{+}05a, JLCH08, KSSA07, LL08b, LW09d, MBDD08, OHT05a, RY05a, Su07c, Su08a, WFX07]. identifying [JLZ05a, SDR07]. Identities [BDY06, BMY08, Yür07, DSY08, EMC05, HZ09a, Sri07, Sri09, SSMI09]. Identity [GT05a, GT05b, QC05, CJL05b, CI09, LWZH05, SC08b]. identity-based [CJL05b, LWZH05]. idle [CSY07, TMMVA05, TMMVA06, Wan08b]. IEEE [ZWM08]. if [Hu07b]. ignorable [PQJ08]. II [ZL08b, AM08c, HWC07, Kay05e, KSKK06, MH06a, PVY09, Prá08b, Ram07e, Rim05c, Rim06g, Rim06f, Rim06d, Rim06h, Rim06b, Rim07c, Rim07h, Rim07f, Rim07b, Sid06, SS06g, VHFF08, WWT05, WLC05, WLC06, WLC07b, Zhu05d]. II. [YSCC05]. III [AM08d, HCZ06, WW05b]. Ill [TK06c, AC09, CHLS08, CC05d, KB05b, LZ06a, LH08b, Qia08, Raj08, SP05, Tad08, WS06a, Wu06e, Wu07d]. Ill-conditioned [TK06c, KB05b, Wu06e, Wu07d]. ill-posed [AC09, CHLS08, CC05d, LZ06a, LH08b, Qia08, Raj08, SP05, Tad08, WS06a]. illustration [RR06]. Image [NK05b, SLLL07, BHCK08, Che07h, Che08b, DZZ07, DFG07, xGZZ08, Gop07, HC08c, Lan07, LM08a, LC05c, LG06c, LC06c, LYGL07b, LLC06b, LMD^{+}08, MM05c, PLZ07, QWX08, SC06a, Su07c, Tse05, Wan05f, WWXX07, WM09, YL06c, YW09, ZCZ05]. images [GZL08, HHYW07, HYZ07, Hos07, Hos08, Imr06, Sha08d, TTC08, Vic09, WHD^{+}08]. imaginary [LR07a, Rim06a, Rim06b, Rim06e, Rim07c, Rim07a, Rim07b]. imaging [LP07a, PLZ05]. IMGS [Yun07b]. imitative [NR08b]. immersed [LJ08]. immiscible [Kim05b]. immune [HYL09, PG07d, RYE07, THLC07, YSXL08, ZGH06, ZSZS09]. immunity [BAB08, MJ07e, TML07]. Impact [LWL07b, BL06, FX09, She05b, WML05]. impatient [Cha09]. impedance [ÇFÇ06b]. imperfect [MR06c, WC06a]. Imperfection [CH07b]. impersonation [Hsu05b]. implant [dCJLNÁA08]. implement [MCGS05]. Implementation [Dam07, DFGV07, MS07a, ACD^{+}05, BD05a, CP08, HH07a, Kay05b, KES05, KNJ05, LP09, LX06b, Mav09, Nav08b, Pou06, RO07, SSW09, SL07b, jZzFbL06]. implementations [DLLN09]. Implementing [ML07]. implications [WL05g]. Implicit [AVM^{+}09, HL06a, IOS07, AV09, AKIS07, ACJ09, AD06c, AFS06, CPY09, ĆRR^{+}08, Deh05b, Din05, DF09, KSA05, Lia08, MMW09, MS07a, Moh07a, NS05, NCGdSM07, PLL08, QCS09, Ram06f, Ran08b, RB08c, SZ07a, TSD07, Taw07, Tha07, sWMX07, WWL08a, WYL06b, WWYL08, XD07, XHH06, bZlDL09]. Implicitly [Lin05d]. implied [Li05d]. important [Sae07]. impossibility [Tan06a, Tan06c, Tan06b, Tan07a]. imprecise [AK05g, Utk06]. improper [SE07]. improve [CKL06, CR05b, ZZW08]. Improved [AK08a, AKB09a, CLWP09, HH06a, HYZ07, LDZ06, LHH05a, MV08, NMP08, NN07d, PM05, Sin08b, SC05d, THLC07, Wu05d, XZM^{+}09, Yoo05, ZWG09, ASPA09, AH09, CYY05, FLLH07a, FL09a, GS08a, GSDB07a, xGyZtZ^{+}07, HZ06d, HZ06c, JJT06, JHHW07, JL09, KF08b, Kho07, KP08c, MFD07, XSX08, Nam08, OT06, Pal09, PBKM09, PG06, PG08b, PK08c, Pre06a, Pre06b, Psi07, QSS08, QCB05a, Ren07, SSJK07, WZ07b, XLwW07, XLH^{+}07, YW05, YRY05a, ZW05a, ZX06a, ZHZ07a, Zha08d, Zha09b]. Improvement [BCW05, HH05, Hwa05, LKY05c, LKY05d, LTH05, Sha05b, TK07, WWHL08, Xie05, YRY05c, YRY05d, ZAX05, AA07c, AKR06, AGM^{+}07, AGI08, BJE05, BLH06, BIK06, BA06, CHY05a, CHL06, CGS08, DMJE05a, DMJE05b, DMJE05c, DMJE06a, EDMJ05b, EDMJ05c, EDMJ05a, EMJB05, GDB06, HEDMJ06, HMJED06, Hsu05a, JSW05, JmBdXgXm05, Kah06a, KJY05, KL07c, LHL06, LW05c, MJHED05, MJED05, MJHED06, RAG^{+}07, SAK08, SYY04, Sub07, SZS05, WLT05b, Wu06a, Wu07g, YWC05, YRY05a, YHL08c, ZLW06a, ZQJ09, SKHZ06]. improvements [Chu07a, Chu07i, CN09a, KL07d, Kou07b, WX06, WS07i]. Improving [CY05a, FAQ08, Ham08, HWF07, tLxLlW07, Mir06b, Zot07a, BHCC09, BL08, GJ05, KM06, yS06c, Yun07b]. impulse [BS09b, Cha07d, Hos05c, LCJ08]. impulses [BS05b, BGF09, HLT09, LW06d, LX09b, SSA09, SL09a, ZFY07, ZFYH08, ZZ07b]. Impulsive [WS09c, YX06a, AO05a, AS08, Ahm09b, AA08, BS09b, BBB08, DMH05, DXH07b, GM07, GF06, GO08, HLW08, HL07d, Hus09, Jan08c, Jan08b, Jan09b, JMC07, JPCL08, LS06c, Li08b, Li09c, LS07c, LCZ05a, Liu06h, LLY07, LMC08, MS09a, zMsCxL08, SM06c, SJC09, Sol05b, SLW08, Sta06a, SC07d, TJG08, WJ05, WL06a, WC07a, WL07b, WFX07, Wu07c, YZW07a, YS07, YZLD08, ZL08a, ZLW08, ZC08b, ZYC09, ZXY06, Zhu08, ZJO08]. impulsively [GZC08]. imputation [RGA05]. in-cylinder [SG05c]. inactivation [dSYGM^{+}05]. incidence [CLG09, JWS^{+}09, KS08d, LW09a, LLG09, MCC07, SJW09, dWqL05, WTS09]. incident [KS08c]. inclination [ESA05, ESAR05]. inclined [Mak06]. include [Dod08]. Inclusion [CK07a, SEŞ09, AV09, yLcCsL09, PM05, PMP08, WX08b, XW08, ZPH06]. inclusions [AB05b, ASK05, BT08, DF09, DW09c, JK08, KB05a, KKA09a, NH07a, Shi09a, Wan05b, ZH08d, ZH09b]. income [hB06, hBxY07, LS09a]. incoming [MM08c]. incommensurate [dlS07a]. incompatible [OT06]. incomplete [AR05, HT06b, LCS09, Zha07a]. incompressible [DJ07, EA08b, HZX07, Küç05b, LL05a, LY09d, NCGdSM07, Ran08b, TY07, YM07a, ZZL08]. inconsistency [Ozd05]. Inconsistent [WZ06c, GG07, LpL07, Wan06l, ZW05e]. Incorporate [HCX06]. Incorporating [ARM^{+}05, PLZ07, HH06a, HL08a, HCZ06]. incorporation [ZBH06]. increase [Jön05, MB05a]. increased [MV09b]. incremental [CT09b, LM08b]. indefinite [BK08b, CC08a, Che07g, CS07i, MH09, ML09, Muk07c, Muk07b]. indentation [Sey05]. independence [Cas07a]. independent [AK06f, Bel08a, CS07b, CCC07, EGS05, Gin07, Gop07, NCÁHC^{+}06, NM09, QLG08, TMSG06, WZ06g, XS06, ZM08b, ZM08a]. indeterminate [ST05c]. Index [Ano05d, Ano05h, Ano05e, Ano05f, Ano05a, Ano05g, Ano05b, Ano05c, Ano07a, AK05f, Ano05i, Ano06a, Ano06b, Ano07b, Ano08, Att06c, BTPGAl06, CK05, DD06a, GL06b, GB06a, GB06b, HWC07, Hos05a, Kos06b, LS06e, LS07d, O’R09, Su08a, SJ05, SJ06d, WLL07a, Yu07, ZW06f]. index-1 [Att06c]. index-2 [GB06a, SJ05, SJ06d]. index-3 [GB06b]. index-splitting [CK05]. indexed [Imr09]. indexes [FP09a]. indexing [ALSL08, LY09c]. Indian [Mit06]. indicators [Mit06]. indicatrix [KY05b]. Indirect [RGA05]. individual [Tan06c]. individuals [LS06a]. induced [ANK05, LH07a, Pat05a, RY05b, XHLL05]. induction [KJ08]. industrial [WH08a]. industry [KO06, Liu06a, Nwo06f]. Inefficiency [YLM07, AK06f]. Inequalities [SB08b, Bar09, BN06, BNR06, BNK07, BN07, BN08, Bou05, CLCK09, CS09b, CKQ08, FB08, FFRFCC09, FB07, HG09, HN07b, HN07c, IR08b, JB08, Kim07b, KBÖP07, Ma08b, MS07c, NB07, NY07, NH07b, NH07c, Noo08b, OP09, Pin06, QG08, QN08, Sad09, SS06a, XS06, YHS09, YNCL08, YM07b, YYMZ09, ZM08b, ZM08a, ZY08c, ZWZ09]. inequality [AL09, ADZ05, ACP07, CS09a, CY07a, FBB08, FFRF07, FFRFCC08, Han06, HT05, JjQjWmL06, JZHT05, JZTH06, JXH06, JZX07b, Jin08, KMA08, KK07d, KKKM09, KP06b, KP06c, Lin07b, LN08a, LOA05, MK05a, Mao09, hPY06, QCN08, QSZ08, RFFFCC08a, RFFFCC08b, Sim09a, Wan07e, WZ09a, WS07i, WS08e, Wu08c, Yua08, lZ06d, bZlDL09, ZW08b, ZW05i, ZCH07, Zhu05a, Zhu06c]. inertia [NS08]. Inexact [jB06, RX09c, YHTR05, BGR05, CJ08, Wu08b, ZZYT08, ZZ09, Zhu05b]. inf [LMH06]. infeasible [ARK07]. infectiology [KRG06]. infection [GN06a, HCW06, NDD05, SL06a, TNS07, WS07e, ZSS08b, d’O05c]. infectious [RYE07, ZzLZ08]. infectives [TNS07]. infectivity [dWqL05]. inference [GGSN06, WY05c]. Infiltration [AKMH06]. Infinite [KOS08, ASL07, AA09, AST08, AK08b, BC09b, BS08, BW09, Che05a, EM05b, FK07, IR05, KS08d, KS08c, KeTL08, Küç05b, LF08, LYJ09, LCM07, LPW08, Mai05, ML08b, ML09, Nis07, RX09a, RX09b, Tan07b, Tan07d, Tan08a, Van09, Waz07k, YLK09, YQ09, YMY05, Zha08e, ZXZZ07]. infinite-dimensional [ML08b]. Infinite-horizon [KOS08, AK08b]. Infinitely [YZ09b, BT06a, CY08a, CB06, LZ08d, ZH09a]. infinity [BCC^{+}09, ZL06a]. inflation [JTMRS06]. inflection [Sch08]. inflow [Far08]. Influence [PM09, BCEGMRMP07, BD09a, II07, SK09a, Vin09]. Information [Deh08, ESD07a, GN07, Nad06b, Vis08, ARM^{+}05, BHCK08, Çet06, EMP08, HL08a, INR08, Nad06a, NCÁHCLP05a, NGLHCLP06, NHCLP07, Rey05b, RGA06, SH07a, Sal07c, Thu07d, WW05d, YC08c, ZCZX07]. infrared [ESS05b]. Ingersoll [SLO08]. inhibit [dG06]. inhibition [dSYGM^{+}05]. inhibitory [WL09a]. Inhomogeneous [GD08, GB07, TNL09]. initial [AGPJV09, CC06b, CCJR07, DSZ07, EHN07, HEZ08, Has06d, JC05, KK08b, KM05c, Les06, Mam09, NKE05, PCS07, Psi09, Ram05f, Ram05g, Ram06g, Ram07a, Rao06a, SA06a, Sel06, SA06e, SKS06, Sin08a, SLL08, Vin09, WB06, Wan06b, Yan05c, YEA07, Yüc08]. initial-boundary [CC06b]. initial-value [DSZ07, JC05, Mam09, Ram05f, Ram05g, Ram06g, Ram07a]. initialization [MV09b, XTCC08, ZBH06]. initially [CFC06c, CH07b, SA06d, Sel07]. injection [KAS06a]. injury [LH05]. inner [AAAA06]. innovation [DCL07, LXZ09]. innovations [NCÁHCLP05b, NHCLP07]. inpainting [BBMN07]. input [CZH07, DDZ08, EKM08, KP06b, KP06c, LML08, LZ09c, LXZ09, SML05a, WLP07, YY08b, Yu07, ZC05b, ZHX06, ZFT06, lS08a]. Inputs [HV08, JVFM04, JLS^{+}05d, LJE07b]. Inputs/outputs [HV08, JVFM04]. insert [TMMVA05, TMMVA06]. insertion [KC07c]. insight [yWyCqD06]. inspired [HHL08, YKJ07]. instabilities [Dil05, ESMM06]. Instability [CC05c, HM08a, HZ09b, Rad05b, GS08b, Lin08, Udw05b]. Instantaneous [TM08]. instrument [Sil07a]. insulated [Ray08]. insulating [KS08d, KS08c]. insulin [FA05b]. insurance [Kat07, Kat08, Kat09]. Integer [Zha09a, AKA07, CKL06, Cha05d, DSKM09, ElS06, Ema06b, wG07a, GW06b, GG08b, GG08c, GG08d, HH06b, Kar09b, LL08a, LR07a, LR07b, LD07b, MBM06, MCS08, Mir06a, Mir06b, Rim05g, Rim05a, Rim05h, Rim05b, Rim05e, Rim05c, Rim05d, Rim05f, Rim06c, Rim06g, Rim06f, Rim06d, Rim06h, Rim06a, Rim06b, Rim06e, Rim07c, Rim07g, Rim07h, Rim07d, Rim07f, Rim07a, Rim07b, Rim07e, Rim08a, Rim08b, Rim08c, Rim08d, Rim08e, Rim09b, Rim09a, Saa05, WY05b, ZS07d, KR07d]. integer-valued [AKA07]. integrability [AGR09, ZLW09b]. integrable [EBP09, GS08a, MX08a, SZ08c, Waz06b, Waz08f, Waz09b, Waz09c, Waz09e, YX08, YZ07c, YZ07d, YZ08, Yu08c, YL09b, Yu09, YL09c, ZnCxX09, lZXL09, Zha09f]. Integral [LS09e, Noo06a, SOT07, Sof09, SiU09, WML07, ZM08a, Zhu09b, Abb06d, AESD05, ADY06, AD07, AD09b, AAK07, AKVH05, BD05a, BGA05, BFR07, BF07a, BLP07, BAF08, BMS08, Bad05, BR09a, Bel07a, Bel08a, Bel08b, BK08c, Bia05, BG08, BKM06, Bou06a, BDY06, Buk05, Buk09, CM09, CLW09b, CKOS07, CJP08, CD06, CC08d, DEB06, Den08c, DEFT05, DC06e, ESE05a, EkES05, EGX05, Fad05, FE08, FBB08, FB08, FFRFCC08, GKB07c, GJB05, GG06, GS06b, GS06c, Gor07, HA05, Has06c, HRCL08, Hol07, Hu07b, Ism06, JG07a, Jer08, Jia07b, JC08a, JLW09, KV05a, KK07c, KWX08, KPS06b, KES09, LM08a, Lep08, Lep09, LS09f, LS09g, LL05g, pLT06, LN07, LSN09, Mah05, ML05, MSK05, MK05c, MA05d, MM05b, MK05d, MY06a, MMN06a, ML06, MM06b, MAM06, MR06b]. integral [MD06b, MY06b, MAR06, MKR06, MNY07, MLR07, MLM07, MS07b, MD07a, MSR07, MMA07, MS08a, MB07, MJD06, MO07, MD09b, Noo07a, O’R08, Ord06b, OML09, PK05b, PPV08, Pin06, Pra07, QS06, QLG08, RMAM07, RMA07a, RZ05, RZ07a, RFFFCC07, SNH07a, SNH07b, SNH07c, SH07d, Sid09, SBA09, SFG07, ST09, TF08, TWZ07, TL08, TZW09, TB09b, WL05d, WL05c, Wan06i, Wan06h, Wan06j, Waz06d, WN08b, WS07i, XWZ06, XS06, YDL07, Yan09c, You06b, YB06, YYMZ09, Yür07, ZM08b, ZWZ09, Boy08a]. Integral-type [LS09e, SiU09, Zhu09b, LS09f, LS09g, Yan09c]. integrals [AO05c, AO06c, AA07e, BB06c, Boy07a, CS09a, CX09, FFRF07, FFRFCC09, GK09, GM05b, KAPM07, KVŽŽ08, LC07, LWWS09, LCS09, MYN07, MTM05, Mao09, MK08a, Ngô09, Odi06a, Pra07, RFFFCC08a, RFFFCC08b, Sal06f, SAS07, SE07, TM07, WX09, XG08, yXhWZ06, Yür07, ZL08e]. integrands [Sid05, Sid06]. Integrated [Bog05d, Cot08, AGA^{+}07, AGM^{+}07, CCFC08, GAG09, MA05f, PBG07, PFA07]. Integrating [KAM06, RAS08, LY05, ZLH^{+}09]. Integration [AGTS07, AGI08, CPSS08a, AMJD06a, All05b, All05c, AGS07, BMJED06, BF07a, BF07b, CTG06, Cor05b, DX07a, Elg09, FW07a, GAR09, HEZ08, Hai07, Hai08, HED06a, HED06d, Has07a, HBES09, Inc07b, KM05b, Mus08, Neh07b, PR09a, PCS07, Ram07a, RVN07, RNV07b, RNV07a, RNV07c, RGAFVA07, Sid05, Sid06, Van07, XXC05, XXC06, XXC08, ZDL07, ZMV05]. integrations [IH09]. integrator [Che05f]. integrators [Che08a, LSW06b]. integro [Ahm09b, AD06b, AŞ06e, Ang06, AO05d, Beh06, BM08, DX07a, DE07a, DMK07, ERAS07, GKB07a, GkB07b, GJ07b, GS07a, GR06b, JRW08, JKN09, Jia07b, KGB06, Lep06, LS06c, Li08b, LP06b, Ma07a, MA06c, MN06, Pao07, PMRAS05a, PMRAS05b, RAS07, RJW07, Ram07g, Ram09a, RZ07b, Raw06a, SG05a, Sha05a, SI06, SLW08, Wan06g, WL06a, WWY09, XZ09b, YS06a, YSS09, YC06a, YWL07a, (Ag07, ZLW08, ZQJ09, ZXL05, ZC06c, ZL09b]. integro-difference [SG05a, YS06a]. integro-differential [Ahm09b, AD06b, AŞ06e, Ang06, AO05d, BM08, DX07a, DE07a, DMK07, ERAS07, GKB07a, GkB07b, GJ07b, GS07a, GR06b, JRW08, JKN09, Jia07b, KGB06, Lep06, LS06c, Li08b, LP06b, Ma07a, MA06c, MN06, PMRAS05a, PMRAS05b, RAS07, RJW07, Ram07g, Ram09a, RZ07b, Raw06a, Sha05a, SI06, SLW08, Wan06g, WL06a, WWY09, XZ09b, YSS09, YC06a, YWL07a, (Ag07, ZLW08, ZC06c, ZL09b]. integro-elliptic [Pao07]. integrodifferential [CY07b, Kum08, LLX05, SG07a, WL09c]. integrodifferential-difference [SG07a]. Intellectual [LT09]. intelligence [HCX06, TKJ08]. Intelligent [ZWM08, BE06, Hua07, Hua08c, JLL08, nXwCpC06, ZL06b]. intensities [Bog05e]. intensity [FLLH07b, KC06b]. interacting [Efi09, Ere06, Pam05b, ZWZ05]. interaction [Bis08, CFJ^{+}05, JB05, KLX06, Ram07h, SI06]. Interactions [ZYH08, DW09a, DW09b, JyGcL^{+}09, KRA07, ME05, RNRRMRCH09]. Interactive [ElS06, GHZS07, Kas07, ASAS06, ASAEE06, AK06i, WXW09]. interconnection [AS07b, CT07b, WYL08b, ZXL06]. interconnections [Par05c]. interdependent [AS06g]. interest [SS06e]. interface [DG07b, Kim05b, ST06c, SK07b, SZXL06, TS05, YC06c]. interfaces [AV07b, CC08d, Rad05a]. interfacial [Li05f, Li05g, Wan07b]. interfacially [LD06a]. interfacially-cracked [LD06a]. interference [CL07a, Che06k, LC09b]. interference-cancelled [CL07a]. interior [AP06, AP09, BL08, CK06a, EM06, JZ08b, KMA08, KKY07a, NV07, PHL06, RX09c, SR06, SY08b, WZ09c, WB09, XFL06, YHW06, Zhu05f, Zhu06a]. interior-point [AP06, AP09, WB09, Zhu05f]. interleaved [LH07c]. intermediate [AKH06, NSS06a]. intermittent [HZ06b]. internal [SG05c, SM08e, WWXX07, XHSL06, dlS07a, dlS07b]. Internalization [LMT^{+}09]. Internet [TWL05, WC05d]. interpolant [DWT05a]. interpolated [XC06b, XC07]. interpolating [Guy09, Sil07a]. interpolation [AA06d, BX09, CCE06, CMV05, CW06b, CW07, DWT05a, DWT05b, DZT06, DLBT08, Kah05b, LMV09, LZZC09, Lin05e, LL07b, ML05, MKA05, MD06b, MD07a, MJ07d, Occ09, SP06e, Yoo05, ZW09c, Sil06d]. interpolatory [BCR09, HMJD06, KLY07a]. Interpretation [Har06, BC09c]. interpretations [BD09a]. interrelationship [LT09]. interrelationships [Ino08]. interruption [hLsT07]. intersecting [AV07b]. Intersection [AK06a]. Intertwined [LDC09]. Interval [CCWJ08, LM08c, SM08b, WYX05, WC09a, ZCW06, ACP07, Ali05, AKR06, Boy07b, Çak08, Cha07c, Dam08, GBG07, HK05, HGN06, JMHS05, JLI06a, JLS06, JLBR07, JLBR08, KPL08a, KP09a, KP09b, LT08b, Liu06c, LW07, Liu09d, qMpXjH09, Moh07c, NGLHCLP06, NG09, PKL08c, Sin08b, Sin09, SMD06, TARm08, WC06d, Wan06l, Wu05d, YB03, YB04, Yam05a, Yam05b, YHL08a, YHL09b, ZYT09, ZZW09b]. intervals [CK06b, LRMS06, WLC05, WLC07c, WLC07d, Zha08e]. intervalwise [OISM07]. intracellular [TG06a]. intratrophic [JB05]. Introduction [CC09a, DC06d, GPS09, Thu06, Thu08, Hub09c, MPT08]. intuitional [GTA06]. invalidates [PJ10]. invariance [ZLR07]. Invariant [BKZ06, ESETA^{+}06, ACP07, CS07b, HM08a, MM07f, SLO08, Su07a, Su07b, Su08b, WYD07, XL09a, dlS07a, lS07b, lS08a]. invariants [AKBZ08, ADDdM07, Hos08, ST08]. Inventory [Lee05a, CO06, DCR07, GBG07, Isl08, LWH06a, LWL07a, Liu06c, MBM06, MMM06, Rey05b, SH07a, SA08b, WS05a, WLT07b, YVU06]. Inverse [EdR07, LD07d, OIL09, TD09, ZMM06, jB06, BZZ08b, BLZZ09, CGDR06, CGVC07, CYL07, CX08b, CX08a, CI09, Den08a, DQW06, EFE08, EPER07, EH09b, FSG07, FCG06, FH06b, FHG07, Gha08, GG08a, GY07a, GEG07, HE08a, HH09a, Hsu06, HZ06c, HY07b, HY07c, ILZ09, JLS^{+}05c, KKT05, KB09, KS05a, Kıl08a, KT07b, KPG07, KH07b, LMV09, LWW05, Li06a, LHZ06, LHZ09, Liu05b, LYH09, LHL08, Ma08c, MSCG09, MD08b, MD09c, NS06a, NR07a, NRA07, Pen05, PHZ06a, PH06, QFX06, QW05, SS09a, ST05a, Sal06d, SKT06, Sal07d, Sal05, Sev07, SC09, SK05, SZ05, SP06a, SFPE07, SDR07, SP06e, TSP07, TA06, UP07a, Vaj07c, Vaj08, VCRCG09, WG05, Wan06k, WSL08, fWZpXzL09, WWW03, WCX05, WDN06, WZ06i, Wu07f, XFL06, XL09b, YC06c, rYzS07, YW09, Yua09, ZM06a, ZC06b, ZCW06, ZT07b, ZH08a]. inverse-free [WZ06i]. inverses [CT05b, Den09a, Djo07a, GWJ07, LYW09, Nač09, PH09, RH06, SC08b, Siv07, SCI08, TS08a, Tia09, WS07h, YLX08, ZYCI08]. inversion [AGJ07, EF06, HFL06, HPD07, LLFM08, MMR09, TAK07b, Vaj07a, Vaj07b, Vaj07d, WLX05, WL08c, ZLL06a, ZZDL07, ZLH^{+}09, ZSCW09]. invert [Ye06]. inverter [LML08]. invertible [BLZZ09]. inverting [EMR08, EMR09, EH09a, WZ09b]. investigate [BWB05]. Investigation [BH07, ZM06a, AH08a, CSC08, GKK09, MA06d, QS06]. Investigations [Li06b]. Investment [WW05d, AK05h, KC09, NN06a, Wan09c, ZL08d]. invex [AJOGRLRG08, Kas09]. inviscid [EA08b, TY07]. involutive [SÖ08, ZZL05]. involved [AA06d]. involving [Ahu08, AC07b, Bab07, BB06c, Dai09, DSY08, DW09c, HC08a, HL08d, Kas09, KKA09b, Lee05a, LWL07a, MYN07, MA09, NMM09, PQJ08, WLT07b, ZH09b]. IO [ZHX06]. iodine [Har06]. ion [QSS08]. IPIC [JM06c]. IR [QSS08]. Iran [SYT06]. IROST [SML05b]. irrational [GKR06, SA08c]. irreducible [DZ06, WH09, YK08]. irregular [Boy05, BO06, CS05b, Nar07b, PCCHLG08, Sel07, UP07b, UP07c, YZ07a]. irrotational [DT08]. IS-LM [ZL08d]. ISD [ZS07d]. Ishikawa [HBN06, JK06]. Ising [MB05a]. Island [AVSB08]. isogenies [MMS^{+}06, MMS^{+}08]. isolation [CP05]. isomorphism [HHC08, HC08b]. isoparametric [Eid05]. isospectral [SZ08c, Yu09]. isothermal [PM09, ZO05]. isotropic [Hlo05]. issue [DC06d, GPS09, Hua07, Hua08c, HHL08, Wan08d, Wan09b]. Issues [Nwo07b, DAQ09, Nwo07a, Nwo07c, Nwo07d]. IT-CEMOP [OASM06]. item [CCL06, CTL09, IR07, MBM06, MR06c]. items [AN07, CL06b, CO06, DM07a, JTMRS06]. itemsets [CTL09]. Iterated [Abb06b, Pal07, LS09f, LS09g, WN08b, Ye06, Zhu09b]. iterates [Geo09, Nar07a, WN08c]. Iteration [AV06, Ahm05b, Akı06, BNH07, BG07a, BV06, CT05c, CAA07, Chu07g, DjT05a, FGfFL06, GL08a, HZ06c, HN07c, Ide08a, JK06, KB06c, KHZ07, LDZ06, Luo05a, MKKS05a, MAO06, NR07a, NRA07, NS06b, NMD07, PHZ05, RDG07, Ram08e, SZ07a, Sha07d, SCX07, Son08b, SQ07, SJ05, SG07c, SQG08, SMG08, Tha07, Waz07k, Waz07j, Waz09f, XB06, YKKY09, YEA07, ZPH06, Zhu05c, Zhu05d]. iterations [AAAA06, CS08a, CT05c, ĆRR^{+}08, Dan08, HBN06, KLW07b, yLcCsL09, NS07a, NS07b, NY07, Raf06, SKT06, Son08a, SJ06d, Wan08g, XF08b, YN07]. Iterative [lCS06, DH06d, DL08, Din05, DLD08, Djo07a, ER06a, GLV08, HN05, Jav07, Nač09, NN07a, Qua07, RAS07, REDES05, Ram06e, Ram09a, Son07, WCW07a, AAS06, AB05b, AS08, AH07b, AL08, All05d, ABP05, ABBP08, AHR08, AC09, BDDB08, BLW06b, BSC06, BDG08, BWR09, BA06, Bie08a, BA05b, BN06, BNR06, Cao06, CPY09, CT05b, CL05d, CL06c, Che06j, CM08a, Chi07e, Chi07d, Chi07f, Chu06, Chu07f, Chu07c, Chu07d, Chu07e, CH07d, Chu07k, Chu08, CT08b, Cui08, Dai07, DC08, DB07b, DH08, fDxZfWyY09, Du05, EJER06, EGE06, FH06a, FH07, GS06b, GJ07d, GJ07e, GJ07f, GSPG07, GSDB07a, Gw05, GK06c, HC07a, HW08a, HL07a, HR07a, HR09a, HW09, HL05a, HZ06d, HCC06, HMT08, HMT09a, HMT09b, JH08a, Jun09b, Kah07, KAV05, KB05a, KKA09a, KSS07, KLW07a, KLW07d, LHL06, LY08a, fLsLdXlF08, tLlWpL08, LH06a, LKU07]. iterative [LH07c, LKU09, LYH09, LWKU09, Lu07, qMpXjH09, Mah09a, Mam09, MKKS05b, MQ07, MZ07, MR07c, MK06, MA06f, MD09b, MGMC05, Ned05, NP05, NK05b, NBO09, NA06a, NNMDS06, NNKA06, NA06b, NN06b, NAJ06, NN06c, NN07c, NB07, NN07d, NNM07, NG07, NN07e, Noo07b, Noo07c, Noo07d, Noo07e, NH07a, NNuH07, OASM06, ÖÖ06, Pao07, yP05, PHZ06b, PP08, QSZ08, QCS09, QZL07b, RHAA07, Raf07a, Raf07b, Ram08c, Ram08e, Ram09b, RW09, Saa05, Sad09, SA08a, ST06b, Sam06, SG05b, SH08b, SC07a, SC09, SXL09a, SM06c, Son05, SC06b, SL06b, Tad08, TG07, TNL09, Uje06b, WL07f, WTW08, WT09, WH06b, WX06, WR07b, WF07d, XHL07, YCW05, YLY07, YY07a, Yao08b, YLK09, YBA05, Yue07, Yun08a, Yun08b, ZCW06, ZM09, ZC09, ZH09a, ZWC06]. iteratively [KS05a]. Ito [LZ09a, Waz08d, Waz08l, WL08c]. IV [He06, LX09a, She07a]. IVGTT [FAQ08]. IVPs [AKIS07, LW05a, Li05c, PDA06].

J [WC06c]. Jacket [LMZ08a]. Jacobi [ZZ07e, AAAA06, BS06, BS07b, Gha08, HED06b, Hon09, HWF07, IMB09, Inc06c, KPS06b, MBS07, PHZ06a, RVNI07, WLZ05, WLS08, Yan05a, YY06c]. Jacobi-like [IMB09]. Jacobian [Hub09d, MK05a, Wu05e, dSLV^{+}08]. Jacobsthal [Djo07b]. Jameson [EA08b]. Jan [RIY05, YRY05c]. Jarratt [Chu07i, KL07c, WKL08, WZ06i]. Java [GEG07]. Jeffery [MM06a]. Jeng [QCB05b]. Jensen [Sim09a]. jet [DEH05c, DK06, Ere06]. jets [ESMM06]. Jimbo [WSKZ07, Waz08j]. JIT [TMRV06]. job [AASAA06, HP07, LJG06, NJG08, OT06, SK07a, SK08c, TMJV^{+}05, THLC07]. job-shop [HP07, LJG06, THLC07]. jobs [JRA^{+}07, WNCL06, lZyT05]. join [ETJ06]. joint [IMB09, PFA07, YVU06, ZS08b, Zho06]. joint-ordering [YVU06]. jointly [GH07]. joints [BK06b]. Jones [Alt06, Alt07a]. Jordan [WS08e]. judge [ZLZG06]. Julia [SS07d, WL06d, yWjCnG07, WY07, XQ06a, XW06, XB06]. jump [HL08b, HY07a, RZ07c, sWX07a, Wei09, WMC08, ZW08a]. jump-diffusion [HY07a]. jumps [RHY06, sWMX07]. just [KOP07a]. just-in-time [KOP07a].

Kaczmarz [CHLS08]. Kadomstev [ZX06c]. Kadomtsev [Waz08c, Waz08g, Waz08m, BR09b, HZ08, Waz08s]. Kahya [Sub07]. Kalman [JC07, Jwo07, RLdLSF07, WS07g]. Kamenev [MX07a, XM07a, Xu09]. Kamenev-type [MX07a, XM07a, Xu09]. Kansei [BHCK08]. Kantorovich [Gup08a, AD09b, RMS05, WZ06i, WZ07d, ZW08c]. Kapteyn [Pog07a]. Kármán [CS09c]. Karmarkar [NV07]. Karush [BGR05]. Kassem [EM11]. Kau [ZLZ08]. Kaup [WSLZ07, ZX06a, ESK05, Inc06a, JZ06c, LH08d, SXJ05, SZ06b, WZL08b, ZX07a, ZWT09c, ZS09b]. Kawahara [CSV07, PKT06a, PKT06b, Waz06j, Zuo09]. KdV [WDLM10, Waz08b, AC08, BB06b, Bha08, CY06a, Dem05c, DSZ08, FqZ08, GA08, GK05a, GK05b, G07, Inc07a, IRSR05, JZ06c, KI05b, KuI08, Li09g, LCL05, qMqZ05, Par09a, SG08a, Sal08b, SPV08, Sha06b, SXX05, SS08c, TL06, TZW09, TZH09a, TZH09b, Wan06f, Wan07d, Waz05g, Waz05q, Waz06a, Waz06e, Waz06g, Waz07b, Waz08o, Waz08t, Waz08q, Waz09c, WL08b, WP06, WZF08, Yan05a, Yan05c, Yan08a, Zha07c, ZZB08, Zha08b, Zuo09]. KdV-like [Sha06b, Waz06e]. KdV6 [GS08a, Waz08f, ZnCxX09]. KdVB [SD09]. Keban [AABD07]. Kelvin [Boy06b]. Kepler [Boy06b]. Kernel [MW09, AP06, CjC08, CK06a, DC06e, GC07b, Gen09a, Gen09b, JC08a, KPS06b, LS06d, LZ09e, MA05d, MMN06a, MA06c, MeT05, MD08c, Ost08, SY09, SM09a, ST09, YC06a, YW09, ZHZ07a, ZL09b, AP09]. kernels [BO06, MR06b, TL08, eT09, WN08b, BH06]. Kerner [Zhu05c]. key [CYY05, CY05b, CH05b, CJL05b, DL05, JXW05, JZCW05, KRY05, KHKL05, LL05b, LKY05b, LKY05c, LKY05d, LT09, Liu07a, LCC05c, PLJ05b, RIY05, Sha05c, Shi05a, SL05a, SW05b, SC05c, SLC05, SCS05b, SC05d, SCS05c, Tsa05, YW05, YRY05a, YRY05c, YY05c, ZK05]. keying [Che08b]. keys [BCL05, CCH05, LWK05a, Sha05b, TLH05]. Khorram [Zim07, MK07b, MK08b]. Kieu [Smi06b]. Killingbeck [BSHZ07]. kind [AO08, AMJD08, AKVH05, BD05a, BGA05, BLP07, BMS08, BKM06, CM09, CC05d, CY05c, DMJE05b, DM06, DC06e, ESE05a, EDMJ05c, EDMJ05a, FE08, GG06, GS06b, GS06c, HM08a, HEDMJ06, HMJED06, Hol07, JC08a, JL05, KV05a, LY07a, LY07b, Li08a, pLT06, LSN09, MSK05, MA05d, MK05d, MMN06a, MAM06, MR06b, MY06b, MAR06, MLR07, MLM07, MS07b, MSR07, MMA07, MJHED05, MJHED06, MO07, PHZ06a, RMA07a, SNH07b, SNH07c, SH07d, She07b, SL07a, SHG07, SZ08a, TF08, WX05, Wan06h, Wan06j, XWZ06, YZ09b, ZL08e]. kinds [Li06b, WZ07c]. Kinematic [KSKK06, Kös06a, BT09b]. kinematics [BT06a, WHC08]. kinetic [SK08a, dSYGM^{+}05]. kinetics [TSS09]. King [Ma11, Chu07m, HT05, Nad07b]. kink [Li09g, WDLM10, Waz07i, Waz08h]. Kinks [Waz06g, Waz06l, Waz07g, Waz08w]. Kirchhoff [DEH05c, Gor06a, Mar05c, TS09, TNL09]. Klein [BZZ08a, FLLH07b, GZ05, HC09, HZ09b, Kay05b, KE05a, KLX06, qLJ05, MD08a, SB09, SMG07, Sun05a, WW05a, WC05a, WZ07c, WZ08c, Wan08d, Waz05m, Zha05a]. knapsack [BK08a, DN07a, DN07b, GTJ07, JRR^{+}07, MFF09, SLC05]. knots [AAG05]. Knowledge [JZHW08, PLZ07]. Knowledge-based [JZHW08]. known [HW05a, KB08, RAGA^{+}05, WLC06, WLC07d, Xie05]. Koch [Jia07a]. Koiter [Fus08]. Kolmogorov [Che06c, Pe¸ 09]. Konopelchenko [FL09b, lH08a, Hon08b, SZ07c, ZX06b, Zhi09]. Korteweg [ROK09, She08, AÖ06b, Bah05, CSV07, DW09a, DKK06, HX07, IDS06, Ism08b, KM05d, ÖK05, ÖÖ06, RKOK09, SZ07d, SZ07e, Sya05a]. Kotera [Waz08c, Waz08d, LD08, Sal08a, ZZB08]. KP [Waz08m, Waz05d, Waz05g, Waz05h, Waz07e, Waz08v, XL09a]. Kronecker [BUP07, KZ07, ZLDW09]. Krylov [AL08, ASA07c, BLW06a, HJ06, LB06]. KT [AJOGRLRG08, AJRLRGOG09]. KT-invex [AJOGRLRG08]. Kuhn [BGR05, LHWW07, RAS08, SLZ05b]. Kupershmidt [Inc06a, WSLZ07, ZX06a, ZLZ08]. Kuramoto [UHuI09, Waz06j]. Kutta [AKIS07, LSW06b, AS09a, AFS06, CLS06, Chr08, LL05d, LSW06a, LMY07, LLY07, MKS09, PPF09, Pod06, SJ05, Van05, WWL08a, WYL06b, WWY09, Wu06f, XT06, YLS05, YWL07a, ZDL07, ZXWL06].

Kutta-like [Wu06f]. Kuznetsov [Pen08, Waz05k, lYqL06, Zha09c, ZZTJ06].

L [Has06b, AM08b, AM08c, AM08d, Ma08b, lCS06, LWW09]. L-shape [lCS06]. L.P. [GANG07]. label [SC07e]. label-based [SC07e]. labor [Cai06a]. laceability [XHX07]. Lacunary [Bil04, Bil07]. lag [NHCLP07, Qui09, SA06a, WC06b, WS09a]. lagged [HHZ09]. lagging [DG07b]. Lagrange [CR06, Lan07, MD06b, MD07a, Mar08, Occ09]. Lagrangian [CL08b, CHH09, MM08b, QZJ07b, RZX07, SPX07, Taw07, WLL09, dCdSJM08]. Lagrangians [AKBZ08, DZG06]. lags [SMBB07]. Laguerre [BR07b, KWS05]. Laguerre-type [BR07b]. Lai [Sun07a]. lake [SMC08, MD07b]. lakes [BFI06]. Lambda [GM06b, MG06]. Lambda-Boolean [GM06b, MG06]. Lame [Liu09a, Iri06]. Laminar [AK05i, Mak06, Pan06]. Lamport [CY05a]. Lanczos [JZ08b, NR07b]. Landau [CPW09, GB07, LZ09d, SDL09, Tan05b, WGL08]. landslides [Val07]. Lane [HM08a, KM09b, You06a]. language [MCGS05]. LANs [ZWM08]. Laplace [AA07a, AK05c, CFF09, Dem06, HPD07, MMR09, QW09, SNK06, Tak07a, XF06, (Ag06, ZLCZ06, ZZDL07, ZYD09]. Laplacian [Abd09, ANM07a, AMN07a, ANM08, Dai09, Den09c, DLT08, DH09, DHG09, EGX05, GZ07c, HC08a, Jan09c, JGY07, JG07b, Kar09a, Li08d, LZ08d, Liu08b, LOA05, Pin06, SM06c, SWX06a, Su08c, SLS08, SL06c, SG07c, SQG08, SMG08, WC08b, Xu07]. Lapwood [MSH08]. large [ASA05, ASAEE06, AP09, ASA07c, BST06, CKY05, Che09a, CK06a, CTL09, DES07a, DC06a, Ham05c, HWH09, HJ06, HH06b, HT09a, HTM09, HT09b, Jbi06, KMA08, Lan07, LZ06b, LZ06c, LY07c, LB06, LW06c, Lin07d, LH06a, LGHS06, LY07d, LY08c, Lu07, MAD07, Muk07b, Muk08a, Neh07b, Par05c, RX09c, SMY08, ST06a, SdHT09, SCS06, eT09, WN08a, WLQ06a, Wu06d, Xia06, XW07a, YWJ07, YZW07b, YY08c, ZŠ05]. large-eddy-based [Che09a]. large-scale [ASA05, HWH09, Lan07, LZ06b, LY07c, LB06, Muk07b, Muk08a, Par05c, RX09c, SMY08, WN08a, WLQ06a, XW07a, YZW07b, ZŠ05]. large-update [CK06a]. largest [Ars07c, NR07a, NRA07]. laser [CYL07, LZL^{+}07]. laser-based [LZL^{+}07]. Lasserre [KES09]. late [Cha05g]. lateness [Ere09b, XYL10]. latent [DJM08]. Lateral [FSÇ06]. Lates [MD07b]. Lattice [DLJ08, BW09, CLZ^{+}07, Che09a, DHL05, DL09, EG05, Hor07, HL07b, Kat08, Kat09, MX08a, SLC09, YCL09a, Yax07, YS06d, Yu09, Zha08c, ZYD09, ZZYZ06b]. lattice-subspaces [Kat08, Kat09]. launch [MWW05]. law [AES05, Bat08a, Bis08, BM09, BR09b, Cha05h, Cor05a, Dem06, GWJ07, IAMA08, KK06, LHSB09, MNG09, Mel08, YB05]. laws [Bri08, Che05e, KK08d, MS06a, NMM08, Sea06, SM05c, Sjö07, Wan09a, XL09a, YL06b, Zah09, ZCZ08]. Lax [Waz08m, GA08, Par09a, Sha07a, Waz08n, Waz08m, Waz08q, Zhi09]. layer [Ami05, And06, AK05i, Che09b, DF08, Elb05, EESF04, FK09, HT06b, Lin05f, LP07c, MR09, Ola06, Pan06, Pan08c, RK08, Ram05e, SH07b, Sel05, SR06, STU09, Vyn07, WLLL07, Waz06h, Wey06, Wey08, YC06c, Yan08b, Zha09d]. layer-adapted [Ami05, Ram05e]. layer-resolving [STU09]. layered [LLW05]. layers [SMH07]. layout [TMJJS07]. LC [Zhu06a]. LCPs [CK06a]. leaching [HMSZ07]. lead [AK06i, CCL06, Kos05, Lee05a, LWH06a, LWL07a, WLT07b]. lead-time [CCL06]. leading [SH07a]. Leaf [DWZ07, WHD^{+}08]. leap [Geu07]. leap-frogging [Geu07]. leapfrog [Sha09]. leapfrog/midpoint [Sha09]. learned [PE07b]. learning [CSY07, DZ08e, DF08, Ere09a, Ere09b, FWJK09, GAFK07, GB08, HH06a, Hon08a, Hu07a, HCS05, KX08, LM08b, Sch05, Wan08b, XYL10, YQYL07, ZS07c]. Least [DIT06, GR07b, JS05, NCÁHC^{+}06, QW08b, AÖÖ06, AHR08, Arg09, BX09, Cha07a, CC06c, Chu07f, CGSS09, DX07b, DCL07, DDZ08, DJ07, GR06b, HR07a, HZ05, HW05c, KT06, KKY07b, LTT06, Ma08b, MNN05, Mug05, yP05, hPyHZ06, PG07b, PG07c, QZL07b, RJ07, Sam06, SG05b, TK06a, TK06b, TY07, WF07a, fWZpXzL09, WL06g, Yin09, ZW05h]. least-square [LTT06, TK06b]. Least-squares [DIT06, GR07b, NCÁHC^{+}06, BX09, DJ07, GR06b, KKY07b, hPyHZ06, TY07]. Lebesgue [Kim07a, MJ07d]. LEDE [AKB09a]. Lee [CGW06, KRY05, KHKL05, LKY05d, SCS05b, ZK05]. Leffler [ST09]. Left [LHZ06, LHZ09, SCS06, YLX08]. leg [TG08a, WL07c, WQL09]. Legendre [Pre06b, RVNI07, BJE05, CM08b, EMC05, LTW07, MNY07, MS07b, PMRAS05b, Pre06a, RVN07, RNV07b, RNV07c, SH07d, SP06a, Wan07f, Wan08c, Wan08e, YSS09, You06a, You06b]. Leggett [AO05a, OZ07]. length [AC06, GS06f, PSRN06, TMSG06, WTG07, YGY09]. lengths [Ram07b]. Leontief [EKM08]. Lerch [Boy07a, CS05c, GGOS08]. lesion [BP09]. Lesser [Tan08b]. lessons [EH08a]. Letter [HM09b, Nad07a]. Level [YZY08, You05, gZW08, AA05, BSP05, DM07a, Deh05b, DMZ08, Ema06b, GBG07, Ham05c, Hos07, LWH06a, Li06b, MD06c, RSA07, RAS08, SH07a, Su07a, SS06g, TCDV07, XW07b, YCYL05, ZZKY05]. levels [BSHZ07, GK09, MM05c]. Levenberg [FP09b, MTC07, MJ07b, Ma08b, MT08, lZZ06a]. Leverrier [ZW05g]. Levin [Thu06]. Levin-type [Thu06]. Levitan [Kin06]. lexicographic [Çak08, Wan06l]. LFSR [JZCW05]. LHL [YRY05a]. LHL-key [YRY05a]. Li [SZS05, QCB05a, SCS05b, ZK05]. liability [PS06a]. Liapunov [Gin07, SK08b, Sol05b]. Libera [AS09c, Sok08d]. liberal [Tan06b]. library [SCR08a]. Lidstone [AM09b]. Lie [Hon08b, KM09b, DW09d, HMNT09, KH08, LDZ06, LLL09, SS08a, Sht07, YL09c]. Lienard [KES05, LXL06, ZSL05]. Liénard-type [ZSL05]. life [CL05b, CHW07, GL05c, HHL08, Wan06a]. life-testing [Wan06a]. LifeRobot [SLX^{+}05b]. lifetime [ESS05a, HWC07, WWT05, WLC07b, WLL07a]. lifetimes [CL05b, CJK07]. lighting [JCX^{+}07]. like [ADZ05, ASK05, AK05e, AH08b, CY07a, Che05i, Din05, DM08b, GJ07f, GSNDB09, HRE05, Hon08b, IMB09, KSM05, KAK06, KB05a, KK07d, KKKM09, KLW06a, KL07b, KLR07, Kut05, LK08, LHSB09, MK05b, MCW08, qMqZ05, Mom05b, yN06, PM05, PMP08, PY08b, Ren07, Ren08b, Sha06a, Sha06b, Śmi07, Sum07, SMG08, WDL07, Wan07e, Waz06e, Wu07e, XZ05, Yan08a, ZWZZ09, ZLYL08, Zhu05d, JGY07, Wu06f, YWL09]. Limit [Ram08a, AM06, AGR09, CL08e, DMS09, DS05b, LM09, Myc06, ZZH06, ZL06a, ZZW09a, ZXZZ07, ZXLZ07, ZX08]. Limited [KMA08, AKH06, BGP06, Tan07b, Tan08b, XW07a, ZG08b]. limiting [Sol05c]. limits [LCM07]. Lin [CHY05a, YRY05b, YY05c]. Lindelof [Ram08d]. Lindstedt [Nav08b]. line [Che05h, EH08b, Gon07, HWS07, KW08a, KW08b, KO06, xLyW06, LZ08d, LM07a, RSDC08, SS05b, SS06b, SS06d, Shi09a, Shi05a, TMRV06, jW06b, WZSL08, WLW08, YGW09, ZLW07, Zho09a]. Linear [GPTU05, KG06, MD05, MD06c, ZHZ08, AJE05, AEJ06, AJ06b, AZ08, ARK07, AJMR07, AKBZ08, ACP07, AD06b, AK06f, All05a, All05d, AAAA06, AK06g, AM08b, AM08c, AM08d, AP06, AP09, Ant05, ADK06a, AKM06, Ars07c, Ars07b, Ars07d, AAA05, AHS06, BD05a, BGA05, BDE06, BBES07, BSKK06, BK08b, BK09, Baw05, Ben07a, BR08b, BL08, BG07a, BG08, BA09, Bog05b, BRT07, BT08, Bou09a, Boy06a, BK05b, Cao06, CS08a, CF06a, ÇE06b, CKL06, Cha06, CS05b, CLZ05, CC06c, CTG06, Che07d, CS07i, CC07b, CX08b, CjC08, CCK06, CPTZ05, CL06d, CR05b, DGB08, Dan05, DH06a, DKA06b, DJM08, DHG06, DH06d, DH06e, DMA07a, DQW06, dCDLN^{+}07, DK05, DR06, DZ08d, fDxZfWyY09, DZ09, ER06a, EN06, EL05a, EL05b, EM05a, ES05b, ESE05a, Ema06b, EBP09, EDS06]. linear [FX09, FA05b, FAQ08, GXY06, wG07a, GC07a, GZG08, GAR09, GAV08, GOMA06, GK06b, GK06a, GG06, GJ07f, GS07a, GWX06, GY07b, Gu07a, GS05, GSG06, GS07c, GS06e, HA05, Han06, Han07, HK09, HES05, Has06d, HAM04, HHP09, Hor08, Hos05a, Hos05b, HW09, HHW09, HW06, HWF07, HT09a, HT09b, Hua09, Hub08a, Hub08b, Hub09b, Hub09c, ID08, Ide08b, JDG06d, JLZ05b, JHST05, Jai07, JS07a, JJT06, JCX^{+}07, JH09, JGS06, dCJLNÁA08, KK08b, Kad07a, KGB06, KMB05, KB06a, Kan07, KSM05, Kar06, Kar07b, Kar07a, KT06, KSM09, KGM06, KRRGT07, Kim06b, Kıy06, KLW06a, KLW07a, Kou07c, KLW07c, Kou07a, KL07f, Kou07e, KW07, Kou07f, KV06, KM05e, KB07, LY07b, LZ06a, Lar05, Lei09, LD07c, tLxLlW07, Li08a, LY08a, fLsLdXlF08, LK08, LZ09b, LMZ09, LL06a, Liu06b, LGHS06, LC06c, LLY07, LYH09, LW09c, LZ09e, LST09, LXZ07]. linear [LHWW07, LL08g, MJ07a, qMpXjH09, MA08d, MAN06, MI08, MBM06, MY05, MA07d, MK05c, MM05a, MM05b, ML06, MQ07, Mat09, MSD06a, Meh06, MK07a, MX07b, MM07e, MB05a, MO07, MZ07, MR07c, MNN05, Moh05a, Moh05b, MK06, MS06b, MA06f, Moh07a, MD09b, MGMC05, MAD07, Mor08b, MT09, NCÁHCLP05b, Nak07, NHCLP07, Nak08, Nam08, NNN05, NDD05, Neh07b, NL05, NA08, NPGW09, yNfM06, Odi07, OMR07b, ÖSÖG09, Pam05a, Pan08a, PFG08, PK08a, PQJ08, yP05, PVN06b, Pha06, QZL07b, Rad09, RAS07, RSSS07, Ram09a, RS08a, RZ05, RMJ07, RB08c, RNV07a, RC06, RGAFVA07, RSA07, RAS08, RH08, Saa05, SNH07a, SNH07b, SNH07c, SANN07, Sai07, Sal08c, ST06b, Sam06, SH07c, SÖ08, SAD06, SG07a, Sha05a, SH07d, SS07b, SHK06, SW06a, SC07a, SZL05, SLZ05b, SLZ05a]. linear [SLZZ06, SZL^{+}07, SAM07, SAN07, SAE08, SGB05, Sog08d, Sog08b, Son05, ST08, SB05b, SYX08, SM08a, Sun05b, SM08c, SM09b, TF08, TCW08, TLC05, TK06a, Uje06b, VG07, VR08, Vog07, WG05, WY05b, WL06c, WZ06c, WZ06d, Wan07b, Wan07c, WSL07, WLP07, WYL08a, WS08a, WZ09c, fWZpXzL09, WLH09, WL09b, Waz07j, Waz09f, WWW03, WCX05, Wei06a, Wei06b, WZ08e, WL05f, WYQ05, Wu06d, WLZ06, Wu06e, Wu07d, WF07d, XCM07, XLwW07, XXC08, XZ08a, XQ06b, XM06, XM07b, XL07, XZ08b, Xu09, YS06a, YSS09, YXSX05, YLB06, Yan06b, YL08a, lY08b, YPZ05, YMY05, YHW06, YWL07b, Yua08, YCYL05, Yun08a, ZG09, ZW09b, ZW05e, ZMW05, ZHLG07, ZHL07, ZCMQ07, ZHGG08, ZZYY08, ZY08a, Zha08g, ZXL05, Zha06c, ZHX06, ZW05h, ZW06d, ZW06b, ZW06c, ZC06d, ZW06e, ZW06f, ZCH07, ZH08c, Zhu05a, ZZ06b, Zhu07, dlS06a, dlS06b, dlS06c, dlS07a, lS07a]. linear [lS07b, lS08a, lS08b, Kah05b, KR07d]. linear-plus-linear [GS06e, SGB05]. linearities [EB05e]. Linearity [ŽMO09]. Linearizability [DMR07]. Linearization [CL07b, Ram05f, Ram05g, She05a, SJ06b, DBC^{+}07, EA07b, JC05, Liu05b, QZJ07a, Sev07, VR08]. linearization-based [Sev07]. Linearized [MDZ08, Ram08b, MA08d, Ram06g, Ram07d]. Linearly [Ram06f, HWH09, KKY07a, LZ06e]. Linearly-implicit [Ram06f]. lines [CH07f, EP07, JGTK07, KOP07a, PE07a, RSZ08, SB05a]. linewise [WGCH05]. link [Chi07c, Chi07h, MC07, WC05c]. linkage [WC05b]. linkages [Shi05b, Shi05c, ZAX05]. linked [Che07a, Chi07b]. Linking [Mel08]. links [Alt07a]. Linstedt [Ram07f]. Linux [LWL06, MAT08]. Liouville [AN05a, Att05b, AL06, BŞ07a, BR09c, BD09b, ÇG05a, Cha07b, Cha07c, Cha07d, CM08b, JGY07, JLW09, KB08, KB09, MBS07, SMG08, YWL09, YHL08c, lZXL09]. Liouville-like [JGY07, YWL09, SMG08]. Lipschitz [Lal09, MS09a, Sol05c, XL09c]. Lipschitzian [sCLC09, Hua05b, ID08, NS07a, NS09]. Lipton [HC08b]. liquid [ALD08, ESMM06, EESESD05, Mak06]. lists [GPTU05]. Liu [AÖ05b, HWS07, mTxWyL07, Zha09b]. living [Upa09]. LLL [CKY05]. LLPO [Tan08b]. LM [ZL08d]. LMI [GH09, ID08, Ibr08, KJ08, LYL^{+}08, Par05c, Par05a, PKL08b, PKL08a, RB08b, SM07d, Yu05]. LMIs [SMA08]. LNA [Li09e]. load [FSÇ06, GG07, WML05]. loading [KO08, KTV07, Lee05c, LD06a]. loads [Cel06a, MB05b]. Lobatto [EF05, EMJB05, MJHED05]. Local [AGR09, AH08b, BCEGMRMP07, Bor09, LL08f, Liu09c, cMpZfR06, Ola06, SXL09b, WCZ09, Alb06, AKB09b, BDDB08, BM06a, BR09a, Bie08a, Bie08b, CC07a, Che05e, CKK05, DBC^{+}07, Den09c, Dod08, Fil05, Ham08, Has07a, IMS09, JC05, Liu07b, MD08c, ME05, OKEL05, SM05c, SS08d, WG07b, Wu08b, Yan06b, ZY08b]. localized [CY08d, DX06, KWZ08, LCM08]. Locally [CMR08, DAQ09, HEZ08, HZX07, Ram05i]. locally-analytical [Ram05i]. Locating [BEvdB07, Su08b, GKR06, Geu07, VL09]. location [PL05, PC07a, RSDC08, SM06b, TKA^{+}08, TS06, WWXX07]. locations [Fro06]. locomotion [Che05i]. LOD [Kar05a]. Log [Cho05, KAT06]. Log-barrier [Cho05]. logarithm [Hsu05a, HL05d, LMV09, LHY05, SCL05b, SLC05, WW05c]. logarithmic [AA08, BNKZ09, Cha06, LS09f, LS09g, Ste08, yXhWZ06, ZM08d, Zhu09b, Che05c, Che06d, WF07a, WS07a, WWD09]. logarithmic-quadratic [BNKZ09]. logarithmically [GQS08, QG08, QCDS09]. logarithms [CLH05, LTH05, PLJ05a, QCB05a, Sha05c, Sha05d, SCS05c]. logharmonic [Abd09]. Logic [LW09b, HAS07b, OASM05c]. logistic [ANM07d, ANM07f, ANM07b, AMN07c, ANM07i, AA06c, AAA08, AAA09, CXC06, HZ06b, PQJ08, hYG05, JSLL07, SP06b, SC07d]. Logistics [Rey05a, Chi08b, Mir08]. Long [GY07a, Kim06a, WZC07, CW05a, CW05b, CW06c, DD06a, EK06, Inc06c, Liu07d, MGG^{+}09, Ras05b, WCLZ05, WCZ06, WZL08b, YX08, ZW09a, Zha05b, ZLZZ06, ZL07b]. long-run [Liu07d]. Long-time [GY07a, Kim06a]. long-wave [Zha05b]. longest [Ery06]. longitudinal [MP08a, TS05]. Looking [YEN08]. loop [Muk07a, SLWX06]. Lorenz [KH07b]. loss [BWB05]. losses [Nwo06e]. lot [AKH06, JTMRS06, MR06c, PFA07]. lot-size [PFA07]. Lotka [Che05a, Che05d, Che06e, Che06f, CF06b, FK06a, GLV08, LC06a, Li07d, LL08d, LZC07b, Lis06, LZ09c, MXC07, PW08a, QD08, QC08, SC08c, SM07d, WS07b, Wan08f, XCH05a, XCH05b, ZY07]. lottery [ZLZG06]. love [WSCL07]. Low [Jbi06, SP09, CLZ^{+}07, DESK06, DES07b, DMDB08, Efi09, ESS05a, ESD07a, Erb06, JCM05, NCGdSM07, SW09a, WLHH05, WY05d, XBZ06]. low-rank [Erb06]. Lower [Fro06, MJ07a, DC06c, LZZ06b, MS08b, MD06c, PB09, Wan06b, WLK08, Wu08a, Yam08, YMCL06, ZL07c]. Lower-dimensional [MJ07a]. LP [ACD^{+}05, AGJ07, YST09]. LPO [Tan07b]. LPV [LP07b]. LQG [Muk08b]. LQP [FB07]. LRT [ADM05]. LSD [Dam07]. LSMS [HHL08]. LSQR [TK06a]. Lu [QCB05b, AEJ06, KW05b, KL07h, PK05a].

LU-SGS [KL07h]. lubrication [BK06b, Ouj06]. Lucas [Kah05a, Sol07, UNS07, DOU08a, DOU08b, Dem08, FZ08, Kwo07b, LMV09, ÖTA05, Sol05a, YK03]. lumped [Ese05b, EK06]. lumping [GG07]. Luo [Wu06a]. Lur’e [CLWP09, GZG09, LP08a]. Lurie [CZ07a, TZX07]. LUT [CH07f]. LUT-based [CH07f]. LV [WCZ09]. Lyapunov [ACP07, BD05b, CDP06, FX09, KPL09b, LP07b, LCZT08, MA08e, MM07d, MT09, Sta06a, TG08b, WC06c, WL06a, WL07b, YCC08, dlS06a].

M [AA07d, AC06, Bra09, EM11, Pan08b, Ren08a, DM08c, Sha06a, Sha06b, XGTS07, AA07d, CK06b, MX08b, WC06c]. M. [EM11]. M.E [Pan12]. Ma [Nad07b]. MAC [ZWM08]. MacCamy [Kim06a]. MacCormack [LJ07]. Mach [Ahm05a, CLZ^{+}07]. Mach- [Ahm05a]. machine [Che06k, Del09, EARS07, EG07c, Ere09a, Ere09b, Hon08a, JRA^{+}07, KTW08, LC06d, MG06, PG07b, PG07c, TMMVA05, TMMVA06, WNCL06, Wan06d, XYL10, YES07, ZS07c, lZyT05]. machines [CSY07, GB08, HZ06b, JGTK07, LY09b, OT06, WSJW06, Wan08b, YY06a]. machining [Tie05]. Maclaurin [LYL07]. macroeconomic [Mit06]. macromolecular [LZ07d]. macroparasite [AOS08]. macroscopic [Mel08]. magic [Dem06]. magna [PW09]. magnetic [BA07, CC08d, Ece06, Lin08, MW07, Mur06, NCGdSM07, PLZ05, SAS05, SSK06]. magnetically [DS05b]. magnetized [Rad05b, SM08d, SM08e]. Magneto [SZXL06, You06c]. magneto-thermoelasticity [You06c]. Magneto-thermoviscoelastic [SZXL06]. magnetohydrodynamic [Cor05a, MNG09, NCGdSM07]. Magnus [VR08]. Mahoney [ZX07b]. Mahony [AKMA05, AKO06, Omr06, ZXLS06]. main [Rim08b, Rim09b, SANN07]. maintenance [Kia05]. MAJ [JPZ06, SH09b]. majorants [AD09b]. majority [PDGR07]. majorizing [WZ06j]. make [Dan08]. maker [HKH06]. makespan [ASA06a, LLLX07, LGJ06, LJG06]. making [AJK05, AS06g, AGA^{+}07, EMP08, JLI06a, JLI06b, JPZ06, MMAHN08, Nwo06a, Nwo07a, PDGR07, RR06, SML05b, WL06e, YT06]. malaria [CTGD08, Kum09b, TML07]. malicious [MJ07e, MS07e]. Mall [LT09]. Malliavin [Ewa06]. Malmquist [Yu07]. mammary [LT07]. man [FAQ08]. managed [KBMZ08]. Management [GPTU05, AEA05b, CPSS08a, CPSS08b, CH05b, FSMC05, GZC08, JC06, JCL08, KOS08, PS06a, RAG^{+}07, SC08c]. Mandarin [LCL06]. Mandelbort [XQ06a]. Mandelbrot [yWjCnG07, XRZ09]. maneuvers [ES05a]. Manhattan [PK08d]. manifold [zPY06, QBH06b]. manifolds [QBH06a]. manipulation [SJ08a, SJ08b]. manipulator [SJ08b]. manipulators [BNdFPS06, WHC08]. Mann [CAA07, ĆRR^{+}08, Geo09, HBN06]. manpower [DT09]. manufacturing [AGM^{+}07, Kia05, MJFAS07, MQW05, RO07, TMJJS07, YY07c]. Many [WZ07c, AR06, CY08a, Dai09, LZ08d, LKU09, ML09, Ord06a, Waz07b, YZ09b]. many-valued [AR06]. MAOR [CR05b]. map [AF07, HLCH08, Liu05a, PC05, XLwW07, XQ06a, AC06, GS06f]. MAPLE [Bay05b, YZ07c, NS05, QBH06b, Shi05b, Shi05c, YL06b, jZzFbL06, Zot07d]. Maple-V [Shi05b, Shi05c]. Mapping [Kim07b, ASL07, Ahm05b, AS07b, Boy06b, KSP08, WL06d, Wu05b, Xin06]. mappings [AR09, AB05b, Ahu08, Ber09, CY08a, CPY09, CLCK09, CO09, CPW09, CAA07, ĆRR^{+}08, ĆC09d, DL08, DW09c, ESD07b, GR07a, GG09b, HN07b, HN07c, JK06, JOB08, KT08a, LL07a, LjZH08, yMT08, NS07a, NS06b, NS07b, NS09, NH07a, NY07, NH07b, NH07c, OZ07, PP08, QN08, QCS09, SC06b, SC07b, Son08b, Sor06, SL06b, SQ07, WCO08, WZDgH09, WCO09, Wan08g, YN07, YY07a, YLK09, ZSM07, ZS07a, ZH09a, ZWC06]. maps [BC09c, IR08a, Nar07a, O’R09, OS09, SZ07a, SY07b, TH07b, ZBH06]. Marangoni [Mur06]. Marcum [SB08b]. margins [DES07a]. marine [Cho07b]. mark [LS06a]. marked [LS06a]. market [DW09b, IR08b, LG06b, Rey05b, Sal07c, SYL07a, SYL07b, zWxY07, Zha07a]. marketing [GBG07, Isl08, PBG07]. markets [Mar08, Sev07, ZC05a]. Markov [ESAH10, Bog05e, Çet07, DR06, Ery06, FFRFCC09, HL08b, LW06a, Mel08, TX08, Yuj07, ZC06b]. Markovian [Cha09, CMC06, RB08c, RY06, RHQ06, RZ07c, RMWk09, sWX07a]. Marquardt [FP09b, MTC07, MJ07b, Ma08b, MT08, lZZ06a]. mask [KLY07a]. masked [CL05b]. mass [Che06n, ESMM06, EESF04, EO06, HZS07, Hua05a, JMC07, Pan08c, SK09a, XL05]. mass-spring [Hua05a]. masses [RAES05]. massive [RAES05]. massively [DvdHD08]. Master [Ano07b, Ano08, Eng06, GZG09]. match [LLC06a]. matching [HLCH08, KK05a, LLY06, Ohm08, Su07c, ZWD07]. material [EB05e, LD06a, Sin07a, You06c, dEdSNL05]. materials [Sey05, Sin05]. maternal [RS08b]. Math [Bil07, EOM08, ESAH10, HLSS10, KAK05, KAK06, Kas06b, Kum09c, KP06b, LL12, hMfWzLpZ09, Mar14, MK10, Pak11, Pan08b, Pre06b, Sal07d, Sol07, Tua13, UNS07, WL08a, Wan08b, WZ13, Waz08b, Wey08, WH09, Yin08b, YHL09b, YHL09a, dCD06a]. Mathematica [Zot07d]. Mathematical [BM05a, BTMB09, DV06, FJ06, FTMA07, HMNT09, Mav09, MS07d, MKA07, RBvB08, SVMG09, WC05c, ALD08, BMD08, Che06k, CZ08b, CTGD08, Cho07a, DK05, ET05b, EP07, GN06a, HA07, HAU05, Kos05, MR06a, MJFAS07, NN08, OF06, PE07a, QS09, RJW07, Rey05b, RO07, Sae07, SH07c, SKSJ06, SS08b, SG08b, TML07, WW06, Wey06, Wey08, Yan07a, ZZ06b, dSYY08, DB06, SJ08a]. Mathematics [Den08b, EM11, HV08, Hai08, HES05, KPL09a, Li05f, Mar07, ORG^{+}09, Pan08c, Pan12, SQ11, Sil06h, Sub07, Yam05b, ZW06b, Hua08c, KK07e, Zot07b, Zot07c]. Mathieu [DNHF05, PST06]. mating [BAB05, FAM07]. Matlab [Sha08c, LLLW06, Kat09, SSW09]. Matlab-based [Kat09]. matrices [ADTGV09, AH07b, ÁFVZ06, BZZ08b, BLZZ09, Çak08, CGDR06, CGVC07, CT05b, CT05c, CFVZ06, CD08, CKKS08, CKR09, CN09c, DH06a, DH06b, DKA06a, DZ08c, DZ06, EMR08, EMR09, Elg09, EH09a, EH09c, FZ08, Gha08, GWJ07, GN07, GG08b, GG08c, GG08d, GG08e, GG08f, HM09a, HL07a, HT06b, HZ06d, HCC06, HY07b, HYG07, HWF07, HY07c, HS07b, IMB09, qJ06, JAT09, Kar05b, KZ07, Kop22, LpL07, LR07a, LR07b, LHL06, LHZ06, Li08e, LHZ09, LS09c, LHB06, LSBL07, LCC05b, LCC06, LH06b, LZR07, LH07c, LD07d, LCH09, MS08a, MSCG09, NS06a, ÖSÖG09, Pen05, RH06, RR07, Rim05g, Rim05a, Rim05h, Rim05b, Rim05e, Rim05c, Rim05d, Rim05f, Rim06c, Rim06g, Rim06f, Rim06d, Rim06h, Rim06a, Rim06b, Rim06e, Rim07c, Rim07g, Rim07h, Rim07d, Rim07f, Rim07a, Rim07b]. matrices [Rim07e, Rim08a, Rim08b, Rim08c, Rim08d, Rim08e, Rim09b, Rim09a, SÖ08, Sol05a, Sol07, SP06e, STV09, TG07, Tia09, TAK07b, Tzo08, UNS07, WYX05, WJ06, Wan06l, WLZN08, WS07h, WS08d, Wu08a, WH09, XL06a, XHL07, YB03, YB04, Yam05a, Yam05b, YL08b, Yin08a, YHL08a, YHL09b, Yun07c, YK08, ZHLG07, ZHGG08, ZHZ08, ZW07c]. matricies [KT07b]. Matrix [ABM09, Cha05f, CLZ05, DM08c, HJ06, HY07c, QZL07b, YLX08, Abd06, AR05, AH09, AGJ07, BF07a, BF07b, BUP07, BSKK06, BK08b, BS07b, Bog05b, Bog05e, BJ08, Cha05e, CL08c, CX08a, CJSSP05, Cui05, CI09, DX07a, DH08, Dem06, DY07, DLD08, Du05, Du06, EF06, EM03, ESAH10, EH09b, EP07, FZ07, FHZ08, Gao06, Gon07, GW06a, GS06d, bGgLgW09, GS06e, HE08b, HE08a, HC07b, HI05, HL05a, HYG07, Jbi06, JW09b, KKT07, KW05a, Kay05d, Kil08b, Kıl08a, KHZ07, KP06b, KP06c, LPP08, LpL07, LZC05, LW06a, LDZ06, Li07a, tLlWpL08, fLyHZ09, Lin05d, LCC05b, LTT06, LM08c, LHL08, MR07a, MA06d, MM07e, MD09c, Nad06b, NS06a, NG06a, NG06b, NM08a, NPGW09, Oss05b, yP05, PHZ05, hPyHZ06, PHZ06b, PHZ06a, PH06, QZL07a, RLY06, REDES05, RES06, RES07, Sal06c, Sal06b, Sal06d, SKT06]. matrix [SB06b, SZF05, SG09b, SC07a, SC08b, SC05b, SGB05, SS07c, Sog07, Sog08a, Sog08c, SCI08, SSMI09, Su07b, SM05d, yS06c, SM06e, TGSNB07, Tze07, UP07a, Vaj07a, Vaj07c, Vaj07b, Vaj07d, Vaj08, WX05, WD06, WWL06, WSL07, WCW07a, WCN08, WYL08a, WCL08, WCX05, WDN06, Wu05b, Wu05e, WDY06, Wu07f, WFHD09, XL09b, XLHK09, XLKH09, Xu09, XTCC08, Yam08, Yan07e, Ye06, YW09, YL09c, Yua05, Yue07, YK06b, ZL05a, yZsZ06, yZ07e, ZSCW09, ZZH09, ZHX06, ZH08a, lZXL09, ZZ07c, ZYCI08, ZLDW09]. matrix-based [XLKH09]. matrix-free [SS07c]. matter [Cai06b]. max [ABA06, AV06, CCWJ08, EI09, GK06b, GK06a, Isk07, KG06, KGM06, LP08c, Sha08a, Pal09, XXX06]. Max-2-SAT [Pal09]. max-average [GK06a]. Max-Bisection [XXX06]. max-ordering [Sha08a]. max-prod [GK06b]. max-star [KGM06]. max-type [EI09]. MaxEnt [SA08d]. Maxentropic [GITV06]. maximal [AV09, AC06, CO09, Mou08, SLWX06, Vaj07c]. maximization [CHY^{+}09, Liu06e, Liu07d, Liu07e, Liu09d, MD05]. Maximum [An05b, DJ06a, GG05a, SM08f, KL06b, WWP05, Amm05, BM06b, DM07b, Ere09b, GOPP07, GPT09, HK05, JRA^{+}07, LLY06, Pal07, Ste09b, TMMVA05, TMMVA06, WTG07, XYL10, YLL06a]. Maxwell [Ma06a, MW07, SP09, VFF08b, YL09a, YZ06]. MC [Lin07e]. MC-based [Lin07e]. MCDM [Cho07b]. McKibben [ZS08b]. MCPSO [NZHW07]. MCTDHF [KKS06]. MCUSUM [ANA07, IM08]. Mean [CK07b, Guo09, KY06, RB08c, XXF06, ZG08a, CL05b, Cha05c, CK06b, DC07, GPT09, GL05c, HYZ07, XSX08, Ngô09, QCDS09, RAGA^{+}05, SML05a, Wan08g, WMC08, Zha07e, ZWS08, Gal06a, Kök06]. mean-absolute [Cha05c]. mean-reverting [WMC08]. means [BB06c, EDS06, EA08b, FA06, GM07, GSG06, GÜ07b, JLZ05a, LTW07, RZ07b, SOT07, SG07a, SC07b, Sun05a, WLS08, YL08a, dSLV^{+}08]. measure [AK06h, BKFM05, CS07b, DESK06, EHN07, GBK06, HAS07b, Jia07a, Sae05, WFX07, Wei09, ZL06b, ŽMO09]. measure-theoretical [GBK06]. measured [DRS^{+}08]. measurement [AK05e, AK05g, KA05c, LC06c, SdJA06, VMK05, dEdSNL05]. measurements [AR05, NCÁHCLP05a, SK05]. measures [Ahm09b, Bry07a, CZ07b, CS07a, ESD07a, GS09, Lee05d, Liu06d, SDM^{+}09, Sil07f, WL06a, WTZ06, WYD07, ZLMW05]. Measuring [AS06g, TARm08, BL06, Pud05]. Mechanical [Wan06h, CCC07, Smi06b, WL06b, Wan06i, Wan06j]. mechanics [CGW06, Mea05, NMM08]. mechanism [BWB05, BK06b, CD09, SJ08a, WY05a, WC05b, WC05c]. mechanisms [Kay05a]. Mechanization [LW05b, HEZ07, Wan05g, Wan05h, WL05b, WL05d, Wan06g, Zha05e, Zha05d, Zha06c]. media [AH05a, AH06d, ADK06b, BM09, ÇE06a, Che06n, Che07i, HS09a, HH09a, HMSZ07, KR08, KM05a, LY09d, MD08a, SZXL06, Su09, Vin09, ZLL06a, ZMV05]. median [Fat06]. mediated [NN08]. medical [PLZ07, WGF^{+}06]. Medium [Pin07, AEA05a, AEG05, Bat08a, Che09b, CY08d, DX06, EEEE06, GS08b, JK07, MR09, Pam06, Pan08b, Ras05c, Sel06, SA06e, SA06d, Sel07, SG09c, SDS05, SM08e, WMX07, You06c]. meets [Dan08, Smi06a]. Mehrotra [SMA06, Sal07a]. Mehrotra-type [SMA06, Sal07a]. Melnikov [LXL06]. MEM [Kan06a]. membership [Isk06]. memetic [TMSZ06]. memories [ACDR09]. memory [AD06a, CWY06, FM08, JPYW09, JSL06b, KMA08, Kay05a, KW05b, KW05a, Kay05d, Kay05e, KNJ05, PK08a, PK08c, PPKL08, SS05d, TLC05, TS09, WYZL08, XW07a, XCL09, YLL^{+}05, YZW07c, ZG08b]. memory-efficient [WYZL08]. Meng [Mar14]. Menten [Slo07]. MEPAR [AG08]. MEPAR-miner [AG08]. Merchant [ORG^{+}09]. meromorphic [CKOS07, CK07a, Irm07]. meromorphically [Aou08a, EA08a, Liu09b, SYX08, YL08a]. mesh [Ami05, AKM06, BDCW06, CGS08, CR05a, KA05a, KP06a, KK07a, KR07b, KYK08, KK08a, KS08b, KA06e, MJ05, MK06, Moh06, MA06e, Moh07a, MD09b, Moh09, NTZ08, OY05, SK06b, She07c, SS08d, UHuI09, YTLH06, ZWZL06]. mesh-free [UHuI09]. meshes [HL06c, Mor09, Ram05e, SL09b, WFS08, XCFX05, YLL06b]. meshfree [XXC08]. Meshless [DjT05a, DjT05b, Dua06, SV06, Cai07, Cai09, CG08a, DjT06, FTT^{+}09, Mar05b, PSD07, Zha06a, ZT06, Zha07g]. meshsize [WZZ06a]. meshwork [AS07c]. message [CCH05, HW05b, LC05d, YGY09, ZF05, ZAX05]. Meta [PLJ05a, JGTK07, MCAF08, SAK08, Sha05d, TGSNB07, QCB05b]. Meta-He [PLJ05a, Sha05d, QCB05b]. meta-heuristic [MCAF08, SAK08, TGSNB07]. meta-heuristics [JGTK07]. Metabolic [BY05]. metaheuristic [AKH06]. metaheuristics [TMASA05]. metal [MA06d]. metallic [dCDNPN09]. metals [SV07]. metamathematics [dCDB07]. metapopulation [HHZ09]. Method [Lee06b, RSZ08, ST05c, WN09b, ZL05b, AJE05, Abb05, AD05a, AD05b, Abb06a, Abb06b, AEJ06, Abb06c, AE06a, Abb06d, AJ06a, AJ06b, ATL07, Abd07, AZ08, Ade07, AK06c, AK06b, ANM07d, ANM07c, AMN07d, ANM07b, AMN07c, ANM07e, AMN07e, ANM08, Ahm06, AS07a, ASPA09, AABD07, Aks05, AÖ06b, Aks06a, AV07a, ADY06, AKM05, AKARA08, ASNR06, All05a, All05d, AAAA06, All07a, ACJ09, AB06, AHR08, AD09b, AH07d, Ami05, AKD06, AŞ06e, AK06h, AC08, An09, AR07a, Ang05, AE07b, AE07c, Arg09, AHIP07, AO05d, AO06e, AO06d, AKM06, ACD^{+}05, AGJ07, Asa06b, AGH05, AÖ09a, Asl09a, Asl09b, AÖ09b, Att05b, Att06c, AFS06, AL06, AR07b, ANK05, AK06i, BD06a, BDDB08, BD05a, BGA05, BVC05, BBES07, BF07a, BMS07, BLP07, BAF08, BR07a]. method [BKA07, Bah05, BN09, BLW06b, Bar05b, BSN^{+}08, BSC06, Bas08, BNH07, Baz08, Bel07a, Bel08b, BL08, BDG08, Bi06, BRW08b, BA05a, BE05, BM05b, BIK05, Bia05, BA06, Bia06, BG07a, Bie08a, BS07b, BA05b, BB05, BKBK06, BS07c, BN06, BNK07, BN08, BNKZ09, BP06a, Bog05d, BGR05, BM06b, Bou06a, BB06d, BSHZ07, Boy06b, BEF08, BP06b, BCMT06, BSKS07, CA05, CR06, CCWJ08, ÇFÇ06b, ÇFÇ06a, Çel05a, Çel05b, ÇG05a, Çel06b, CY07a, CN06, CPR07, CKL06, Cha07b, CL05a, CS07e, CL05c, CZ05a, CLZ05, CW05a, CW05b, CZ05b, CL05d, CW06c, CC06c, Che06m, CW06d, CY06a, Che07e, Che07f, CS07i, CS07h, CM08b, CL08e, CA08b, CM08a, Che08a, CX09, lCX06, lCS06, CG08a, CS09c, Chi06c, Cho05, Cho07b, CCK06, Chu06, Chu07b, Chu07i, CH08a, CH08b, CN09a, CN09b, CPTZ05, CT06, hCxC05, CG07c]. method [CG08b, CGSS09, CR05b, CK07c, DGB08, DIT06, Dai07, DC08, DLLN09, DN06, Dao05, DG07b, DE07a, DH06a, DHY06, DH06b, DKA06a, DJ06b, DKK06, DH06c, DB07a, DB07c, DMA07b, DBC^{+}07, DOU08b, Dem08, Dem05d, DXH07a, Den08c, dCDLN^{+}07, dCDNPN09, DwWmW05, DM07b, DSZ08, Dog05, DC07, qDyC08, DjT05a, DjT05b, DjT06, Dua06, DMZ08, DFGV07, ERAS07, EFE08, EN05, EM06, EHN07, EG06a, ES05b, EWAE06, EWEA06, Elb05, EEEE05, EO06, EA06, EGX05, EELZS05, EP07, Ere06, EHL08, Ese05b, EK06, EA08b, Ewa06, FGfFL06, FW07a, FHH09, FM05, FE07, FE08, FDL08, FLLH07a, FqZ08, FHSS08, FL09a, FJT05, Fra09a, FS05, FH06b, FB07, FHG07, Gal06a, GA08, GH06, Gao06, GY06, GJ07a, GZ09a, GZ09b, GAR09, GFH08, GL08a, Gen09a, Gen09b, GKR06, Geu07, GK05a, GKB07a, GKB07c, GkB07b, GKD07]. method [GJB05, GJ07b, GJ07c, GJ07g, GJ07h, GS08a, GCJ08, Gor06b, GSDB07b, GDB06, GW06b, GbL07, Gu07a, GSG06, GS06d, GK06c, HBBA07, Haj06, HC07a, HZ06a, Han06, HYG08, HK09, HKS09a, HDG07, HNB06, HBE09, Has06d, HSH05, mHzC05, mHsLzC06, He06, mH07b, HZX07, HRCL08, HH09a, HM07, HBVM07, HH08a, HIPA08, HJ08, Hos05a, Hos05d, Hos06b, Hos06a, HN06b, Hos06c, HN07a, HL06b, HW09, HZS07, HCC06, HN07b, HWS07, HMW07, HS07b, HC08c, Hua09, Hub08a, Hub09b, HJK08, Ide06, Ide08a, IMB09, Inc05, Inc06b, Inc06c, Inc06a, Inc07e, Inc07b, II07, IM07, IIS05, IRSR05, Ism08b, Iva08, JDG06a, JDG06b, JDG06c, JHST05, JLRB05, JLI06a, JLI06b, JLBR07, JA08, JLBR08, Jan07a, JS07a, Jav06a, Jav06b, Jav06c, Jav06d, JG07a, JL08, Jer08, JjQjWmL06, JZHT05, JZTH06, JXH06, bJyKxC09, JY06]. method [JB08, JZ06c, JL05, JYX06, JCM05, Joy06, JM06b, Jun09a, KKT07, KA05a, KA06a, KK07b, KR07a, KY08, KYK08, KK08a, KS08a, Kad07a, Kad07b, Kah05a, Kah05c, Kah05b, Kah06a, KC07a, KAV05, KB09, KF08b, KK07c, KBH06a, KBLT08, KK05a, KSM05, KSGM06, KT06, KMA08, Kat09, KS08b, KY05a, KI05b, KES05, KKY07a, KP07, Kha09, KSM09, Kia05, KYL05, Kim06a, KKY07b, KSC^{+}08, Kin05, KB06c, KSS07, KLX06, KHZ07, KLW06a, KLW06d, KLW06c, KLW06b, KLW07a, Kou07a, KL07c, Kou07b, KL07f, Kou07d, Kou07f, KL07h, KLW07e, KLR07, KLS09, KL08, Kum03, KM05d, Kum06, KA06e, Kum07, Kum09c, KH07b, Kuo05, KWP07, KAM06, KF08c, LY07b, LV05, Lal09, LZ06a, Lan07, LFM06, Lar05, Lay08, Lay09a, Lee05b, Lee05c, LWH06b, Lee06a, LS08c, LJ08, Lep06, Lep07, Lep09, Les06, LW05b, Li05c, LMH06, tLxLlW07, LH08b]. method [zLmLL08, LYX08, LZ08c, LY08a, LT08b, Li08e, LL08e, Li09f, LWWS09, LXJ09, LPA^{+}09, LLC09a, LSBL07, Lia05, Lin05a, LD06b, LYM09, LXML09, Liu06c, LGHS06, Liu07d, LS07d, LH07b, Liu07a, LW07, LCH09, LLC09b, LM07a, LMN07, LY09d, LZ07d, LN08b, LP09, Luo05a, Luo05b, LHW07, LHWW07, MK05a, Ma05a, Ma06a, MTC07, MW07, MJ07b, Ma08b, MCW08, MT08, MMW09, qMpXjH09, MLW09, MA08d, Mah09a, MCAH05, Mah05, MPEL09, MS07a, Mak07b, MB06, MKA05, MK05c, MA05d, MK05d, MD06a, MY06a, ML06, MM06b, MAM06, MA06c, MD06b, MAR06, MLR07, MLM07, MD07a, Mam09, MQ07, Mar05b, MRN09, Mav09, McC05, McC06, MFE05, Meh06, MK07a, MDZ08, MV08, MKSK07, MO07, Meš07, MSD06d, MZW05, Moh05a, MA06f, MD09b, Moh09, MGMC05, MMC06, Mom05b, MAK05, MO06, MMN06b, MS06c, MAO06, MHA06, MZL07, Muk07b]. method [NG06a, NG06b, NR07a, NR07b, NRA07, NZ08, NM08a, Nar07b, Nat09, ND07, NKR^{+}08, NR09, Net08b, yNyF05, yN05, yNfM06, NBO09, NA06a, NNMDS06, NNM07, NG07, NKH07, NN07b, Odi06b, Odi06c, Odi07, OME08, OF06, Ony05b, OISM07, Ouj06, Öza05, ÖK05, Pak11, Pam05a, Pam05b, Pam06, PE07a, PLL08, PB09, Pan12, PG08a, PG06, PG08b, PR09a, Pat05b, Pat07, PC09, PPV08, PC08, yP05, PHZ05, PHZ06b, zPY06, PL06c, PFL09, Pet07, PMP08, PHP09, PD09, PP08, Pod06, PCS07, PKT06a, PKT06b, PG08c, PMRAS05a, PMRAS05b, PSD07, QFXW06, QFF06, QFX06, QFS07, Qia08, QW09, QSZ08, QW08a, QS09, QWX08, Qua07, RMA07a, RGD07, RDGP07, RDG07, RAZ07, RZ08, RAS07, RN09, RMS05, Ram05c, Ram05d, Ram05h, Ram07c, Ram08c, Ram08d, Ram08e, Ram09b, RK07b, RZ05, RJ07, RZ07b]. method [RMG07a, RJMG07, Ras05b, Ras05a, RAP05, RB08c, Raw06a, Raw06b, RB05a, RB05b, RCB06, RB06, Ray06, RCB08, Ray08, RYW06, Ren06, RW07, Ren08b, RA09, RAG^{+}07, RX09c, SBS08, SF07, SNH07a, SNH07b, SNH07c, SA08a, SKHZ06, Sak06, Sal08b, SM06a, SSW09, SG06, SGT06, SG07a, Sha05a, SAB08, SH06a, Sha09, SH09a, SHY08, SR06, SB05a, Sha05e, SGS09, SXJ05, She05a, SJ06b, SS09d, SC07a, SC09, SXL09a, SWH07, SS05b, SS05d, SS06b, SDL09, SW09a, SP05, SFA08, SAM07, SA07a, SM07a, SS08c, SS05e, Sin07b, Sin07c, SBA09, uIK06, Śmi07, SS08d, SC06b, SZ07b, SZ07c, SZ07d, SZ07e, SZ08a, SY08a, SB05b, Sta06a, SL06b, SC07c, Su08b, SYY04, Sub07, SH07f, Sun05a, yS06c, SC06c, SW06c, SY08b, SM09b, STU09, SK09b, Sya05a, SH06c, SA06h, TKA^{+}08, Tad08, TL07, mTxWyL07, TZW09]. method [TLM09, TB09a, TB09b, TSP07, TMJV^{+}05, TD06, TD07b, TD07a, Thu07c, Thu08, TG08a, TK06a, TW09, TYCL06, TY07, TT09, Tua13, UHuI09, Uje06a, Uje06b, UP07c, Van07, VR08, WSLZ07, WJ05, WC05a, Wan05e, Wan05a, Wan06e, WB06, WC06d, yWyCqD06, WZT06, WCZ06, Wan06l, WZSZ06, Wan06f, WH06a, WL06a, WHL06, Wan06b, Wan07b, WsZ07a, Wan07c, WSZ07c, Wan07d, WML07, WY07, WZ07c, WSZ07b, WSKZ07, sWMX07, WLK08, WTW08, WN08a, WCH08a, WW08b, WFS08, WKL08, WZ09c, WY09, WM09, WC09a, WT09, WSL09, Waz05b, Waz05a, Waz05c, Waz05o, Waz05n, Waz05p, Waz06b, Waz06a, Waz06d, Waz06i, Waz06h, Waz06k, Waz06o, Waz06p, Waz07b, Waz07a, Waz07h, Waz07e, Waz07i, Waz07k, Waz07j, Waz08c, Waz08d, Waz08e, Waz08u, Waz09f, WLY08, WWYL08, WGCH05, Wu05d, WC05d, Wu05c, WL06g, WR06, WZ06h, WS06a, WZ06i, Wu06c, WZ06j, WRB07, WMBY07]. method [WLL07b, Wu07g, WZ07d, Wu07e, Wu07b, WZ08f, WHL08, WCH08b, WH09, Wu09, XWZ06, XXL09, XHL07, XXC08, XW06, XFL06, XF06, XF08a, XXX06, XHH06, XZY08, XL09c, YAO09, YS06a, Yan05a, Yan05c, lYqLW07, YHS09, YS06b, Yan07e, YJ07, YZY08, Yan08c, YJ09, YY07a, Yao08b, YC05a, YBA05, YJW06, YWJ07, YCC08, YKKY09, YK03, You06a, You09, YHW06, YT06, YZW07b, YZW07c, Yua05, YY06c, YQYL07, YHL08a, YHL08c, YHL09b, Yüc08, YCYL05, Yue07, YKR06, Yun07b, Yun07c, YK08, Yun08a, Yun08b, Yun09, YB05, ZFN07, ZG09, ZW09a, ZSHB06, ZCZ05, Zha06a, ZWL06, ZX06a, ZX06b, ZBH06, ZLW06a, lZ06d, lZZ06a, ZT06, ZC06a, ZLL06a, yZ07e, Zha07c, ZX07a, Zha07d, ZHZ07a, ZZ07a, Zha08b, Zha08c, Zha08d, ZOZZ08, ZW08a, Zha09e, ZWT09c, ZC09, Zha09b, ZYT09, Zha09d, ZZX05, ZZC06, ZC06c, ZZYZ06a, ZHX06, ZC08a, ZW08c, ZH09a, ZZ06a, ZZZ07, ZS09b, ZWZZ09]. method [ZZYZ06b, ZZ07d, ZW05i, ZWC06, ZXZL07, ZCH07, ZLC07, ZZ09, Zho09b, Zhu05c, Zhu05d, Zhu05g, Zhu06a, Zhu07, ZCZX07, ZG08b, dCdSJM08, CT07c, KBH06b, SL07b]. methodological [Nwo07a]. methodology [AGT07, BDCW06, CYL07, GAS08, Hua07, Hua08c, dSYY08]. Methods [HR07b, AAS06, AE07a, AMN07f, ANM07f, ANM07j, ANM07i, AMJD06a, AKH06, AC07a, All05b, AH05b, ABP05, AH07d, AP06, AMAP07, AP09, AC09, AH08b, AKR05, BDDB08, BMJED06, BCCK05, jB06, BST06, BS06, BLW06a, BRW08a, BWR09, BÖ05, BN07, Boy05, BFKT07, BO06, Buk09, CCE06, Cai07, Cai09, CP08, Cao06, CHLS08, CMC06, CLS06, CWY06, CT05b, CT05c, Che06j, CS07g, mCLB08, CJ08, CL09b, lCS06, Chr08, Chu07a, Chu07f, Chu07c, Chu07d, Chu07h, CH07d, Chu07j, CH07e, Chu07k, Chu07l, Chu07m, Chu07n, Chu08, CG05b, CJP08, CT08b, CD08, DSB05, DJ06a, DKA06b, DB07b, DSZ07, DMK07, DHG06, DCL07, DZ08b, fDxZfWyY09, Djo07a, Dön06b, Du05, EBsENMA05, EG07a, EGE06, EHAR05, EHZES05, EMESES05, EM07, EHR09, FCG06, Faz09, Fen05, FH06a, FH07, Fer09, FXLZ05, GAV08, GG06, GJ07d]. methods [GJ07e, GJ07f, GKK09, GSDB06, GSPG07, GSDB07a, GSNDB09, GSDB09, GSG09, Gw05, HL06a, HAK07b, Ham05a, HW08a, Has06c, HPD07, HRE05, HR09a, HJ06, HN05, HHW09, HWF07, Hua08b, HMT08, HMT09a, HMT09b, Ide08b, IMB08, Jai07, Jav07, JWPS08, JZCW05, Jia07b, JH08a, JH08b, Jun09b, KTK06, KP06a, KPS06a, KR07b, Kah06b, Kan06a, Kan06b, KT07a, KOR08, Kat07, Kat08, Kim06b, KLW06e, Kou07c, KLW07b, KLW07c, KL07a, KL07b, KLW07d, KL07d, KL07e, Kou07e, KW07, KL07g, LY07a, LP07a, LMSV07, LMV09, LW05a, LL05d, xLyW06, Li06b, LSW06b, LSW06a, LD07c, Li08a, fLsLdXlF08, LK08, LBS05, LB06, LW06c, Lin06d, Liu06g, LZ06e, Liu06f, LX07b, LGY07, LMY07, LZA^{+}07, LLY07, LL09, LYH09, LW09c, LL12, LSN09, LjZH08, LP07c, LYL07, Ma07a, ML05, Mal06, MYN07, MKKS05b, MH06b, MH06a, MP08b, MZ07, MR07c, MM08c]. methods [MV09b, MAY09, MJ05, MA06e, MSM08, MGMC05, MN06, MN07, MKS09, MP08c, Nač09, NP05, NNKA06, NA06b, Noo06b, NAJ06, NN06c, NN07c, NN07d, NN07a, NN07e, Noo07b, Noo07c, Noo07e, NH07a, NNuH07, NH07b, OB09, OAHR09, Omr07a, Özb06, PPF09, Pao07, PS06b, PW08b, PM05, PR05, PP07b, PP07a, Pet08a, PY08b, Psi09, RMAM07, RHAA07, Raf07a, Raf07b, Raf08, RK08, Ram05e, Ram05a, Ram05b, Ram06b, Ram06c, Ram06d, Ram06e, Ram06f, Ram06g, Ram07d, Ram07e, Ram08f, Ram09a, Rao06a, RJM07a, RJM07b, RJK07, Ren07, RW08, RW09, RWB09, RGAFVA07, RMA07b, REN^{+}05, RGA05, RAGA^{+}05, RV08, SS09a, Sal06a, SA06a, ST06b, Sea06, Ser05, SH06a, SG07b, Sha07c, Sha07d, SH08b, SS06c, SS06d, SP09, Sil06a, uITA06, Sit06, SS07c, Son05, Son07, SC07b, SJ05, TSD07, TGSNB07, TmChL06, TG07]. methods [TFX07, UEL07, Van05, jW06a, yWjL06, jW06b, WZ06d, WF07a, WL07c, WWL08a, WWL08b, WZWW08, WQL09, WL09b, Wan08g, Waz05m, Waz06d, Waz06m, Waz06q, Waz08w, WYL06a, WLQ06b, WL05f, WYL06b, WWY08, WWY09, WLY06, WH06b, Wu06b, Wu06f, WX06, WR07b, Wu08b, WLL09, XT06, XZ09a, Xia08, XFQ06, Yan07a, YM07a, YY05a, YHTR05, YLS05, Yan06b, Yan08e, YNCL08, YWL07a, YWL07b, YGW09, YK06b, Yun07a, ZSM07, ZS07a, ZD06a, ZCW06, Zha07g, ZHGG08, ZZYY08, ZG08a, ZWG09, ZXL05, ZXDL05, ZXWL06, Zha07h, ZW05h, ZW06d, ZW06b, ZZYT08, ZH08c, LWG09]. metric [AR09, ACOS06, Ber09, ĆC09d, Geo09, IR08a, yMT08, PL06c, PFL09, WZDgH09]. MEW [Inc07d]. MHD [Pan12, AES05, Att06a, CFX05, EEEE05, HS09c, KBKV09, MM06a, Oua05, PM09, Prá08b, SBA09, SK09a]. MHD-PDE [Prá08b]. Michaelis [Slo07]. Micro [HM06, Mur06, ZY09]. micro-channels [ZY09]. Micro-particle [HM06]. micro-polar [Mur06]. microarray [SCL^{+}05a]. microbial [GFWX05, YFLX09]. microeconomics [MP08b]. microfluidic [LD07b]. micropolar [AEA05a, AEG05, DC06b, EB05d, EEEE05, EO06, IHB06, KAS06a, MR09, Pan08b, Pan12, RM08, SK07b, SZXL06, SCB07]. Microscopic [CW08, Mel08]. microstructure [BL06]. microwave [MM07f]. middle [Su08a]. midpoint [RK08, Sha09, UE07]. midpoint-trapezoid [UE07]. migrating [DD08b]. migration [LZX08, Mar05c]. migratory [CRC08]. mild [hLqSlZ09]. Miller [LHC07]. milling [HIST06]. MILP [AEFFA08]. Milstein [OAHR09]. MIMO [CZH08, FWJK09, JLL08]. min [ABA06, AV06, CCWJ08, EI09, GK06a, Isk07, LP08c, HRM07]. Min-Power [HRM07]. min-type [EI09]. mind [JZH07]. Mindlin [WH06a]. minds [BKS^{+}06]. miner [AG08]. minima [COG08, VL09, nXwCpC06]. Minimal [Lin06d, ZW07a, FAQ08, KTK06, LpL07, Lin07f, Vaj07d]. minimax [AH06a, CWLH08, FHSS08, GM06a, Mor08a, WSXL09, YQ09]. minimization [CG06, ER06b, Fag05, HWH09, KMA08, KF07, KLS09, KJ09, Ma05b, Nwo06e, PSW07, RS06, WZT06, YLY07, Yua08, Zha07b, Zhu06a]. minimize [AASAA06, LGJ06, LJG06, TMMVA05, TMMVA06, WSJW06]. minimizer [lSgPpJ07]. minimizers [SW05a]. Minimizing [Ere09a, HED06b, OT06, WNCL06, Ere09b, JRA^{+}07, LC06d, QZZ07, SP06c, Taw07, XYL10]. Minimum [RH08, GL05a, Ghi06, HK05, HK06, HGN06, HYG07, Kla08, Liu08a, NG06a, TT06a, Tok06, VV06, Vaj08, Yin07, Zhu05a]. Mining [AG08, CTL09, LT07, TRLD08, WML05]. Minkowski [AB05a, KA06c]. minus [XSKC07]. MIP [OT06]. Mirakyan [GBG08, GS08c]. miscible [SY08a]. misclassification [MR05b]. misclassified [NK07b]. mismatch [XXTL09]. mismatched [WJZG08]. MISO [DCL07]. missed [SS05e]. missile [CQL08, HC07b, HJB06, HBJ06]. missing [PQJ08, RGA05, Sil06d, Sil07b, SMD06]. Mittag [ST09]. Mittag-Leffler [ST09]. Miwa [WSKZ07, Waz08j]. Mixed [Djo07b, Hu05, YJ09, ZC08a, ZL09a, ZS07d, AB05b, AASAA06, ACJ09, AS09c, BS05a, Bay05a, BS07c, BNR06, BN08, Cha05d, CL09b, DSKM09, Din05, DW09c, DCR07, EESF04, FM05, Gw05, HA05, HG09, HAU05, HH06b, Ibr07, JK07, JK08, JCM05, KOP07a, KNJ05, LS06c, LmZ08b, Li09f, LC09a, Liu05a, LLZ08, Liu08b, LL09, LX09b, LL12, MMO07, MA08d, MBM06, MMR09, Mor08a, Nat09, Pan08c, Par05b, PSP05, Sad09, SSA09, SG05a, SGT06, SRH09, SY08a, Ste09i, SM08b, SM09c, TLM09, TMRV06, TNL09, XD07, YS06a, YM06, Zha07h, KR07d, MD05]. mixed-boundary-value [HAU05]. mixed-hybrid [FM05]. mixed-model [KOP07a, TMRV06]. mixed-norm [AS09c, Ste09i]. mixture [CZ07b, Haa09, HC08c, LWH06a, LYGL07a, LM08b, NK05a, SK06a, ZC05a]. mixtures [Lee05a, SA08d]. Miyaji [ZW05c]. MKdV [KI05b, Waz06p, Hon09, SS09d, TW09]. MKdV-sine-Gordon [Waz06p]. MKdV-sinh-Gordon [Waz06p]. mKdV-type [SS09d]. ML [WLY06]. MLE [WWC06]. mobile [OHJ07, SL05a, SH07f]. Mocanu [WZ13, NM08b]. Mock [BX09]. Modal [CL08c, yZ07e]. mode [Boy06b, JDSR08, Shi09a, SWX06b, SWX07b, TD09]. mode-dependent [SWX06b, SWX07b]. Model [AA07c, JS06, OHT05a, WL05g, ABB06e, ADM05, AOS08, AH07a, ALD08, AVSB08, AH06d, Alt07a, AA08, AAA08, ASA07c, ANK05, BW05a, BL06, BW07, BMD08, BB06a, hB07, BT05, Bia06, BTPGAl06, BK05a, Bog05e, BAB05, BKP06, Cai06a, CL08a, CLG09, CS05a, Cha05c, CH05a, CWLH08, Che05b, Che05c, Che06a, CO06, CS06a, Che06b, Che06h, Che06d, CS07f, CLZ^{+}07, CH07c, CY07b, CLW09a, Che09a, CYCX09, CJ06, Chi07h, CKK05, CP05, Dam07, DvdHD08, DM05, DW09b, DT09, DCL07, DXH08, DMDB08, gDH07, DLJ08, DCR07, EKM08, EH06, EAG08, EMA06a, ET05b, EP07, FEOC08, FDL08, Fen05, FAQ08, GN06a, GTX08, GZC08, GL06b, GJD^{+}07, GL05c, GBG07, GGSN06, GM06b, HN06a, HJ06, HWC07, HAU05, HZW06, HC08b, HY07a, HHK^{+}05, HC05b, HCZ06, HTO06c, HC08c, Hua08a, HL05b, HAS07b, IR06, IR07]. model [Isl08, IAMA08, JLS^{+}05b, JPZ06, JPYW09, JW07, JL07, JY09, JWS^{+}09, JC06, JMC07, JCL08, JPCL08, JZH07, JTMRS06, KOP07a, KRG06, Kho07, Kim05b, Kos05, Kov07, Kum09b, Kwo07a, LS09a, LL08a, Lee05a, LWH06a, LP07b, LJ08, LKW06, LYGL07a, LJK07, LL08b, LW09a, LT09, LLG09, LZC07b, LOZ07, LC09b, LCZ05a, LC05c, Liu06c, LG06c, LS06d, LBZY08, LMC08, LTCC08, LXZ09, LZ09d, LGw07, LN08b, Luo07, LG08, LPW08, Ma06b, MG09, MAD08, MR06a, MJFAS07, MMAHN08, MBM06, MMM06, MPS08, Mak07a, MR06c, MKG09, MJD07, MBC08, MDZ08, MCC07, zMsCxL08, MC08, MB05a, Mir08, MS07e, MQW05, MM07f, MF09, MD07b, MB08, NDD05, OKEL05, OHT05b, OMR07b, PE07a, PW09, PN08, PMS09, PWY08, PLZ05, PA09, PFA07, QS09, RGD07, RDG07, Rey05b, RO07, Rös07, RY05b, SMM08, Sae05, Sae06b, Sae06c, Sae06d, SWK^{+}06, SB08a]. model [Sch05, SKSJ06, SM06b, SLX05a, SMG07, SC06a, SY07b, SJC09, Shu05, SGP07, SSTD05b, SI06, SL06a, SP06b, SJW09, SXL09b, SQ07, SS05f, SS05h, SS06g, SC07d, TX08, TG06a, TMRV06, TD09, Tie05, TML07, TD05, Upa09, Utk06, VG07, VKVHN05, VFF08b, dWqL05, WH06a, WL07d, WS07e, WS07a, WL08a, WZ08b, WT08, WWD09, WTS09, Waz08e, Waz08p, WCX08, Whi06, WC05c, WC05d, Wu06a, XL08a, XC06a, XZ07b, XCD05a, XCH05a, XCD05b, XCH05b, XL06b, XCD06, Xu06, XW07b, XLHK09, YSZG05, YX06b, YKKY09, YZ06, YY07c, YLL07, YS06d, YZD08, YC08c, YY09, Yua09, Zen09, ZLMW05, ZW07a, ZzLZ08, ZLT09, ZYD09, ZHX06, ZWD07, ZYC09, ZL06b, ZGL08, ZS09b, ZL06c, ZFT06, ZSS08a, ZSS08b, ZY08d, ZL08d, ZSZS09, ZFL^{+}09, ZH07b, ZWM08, dSF05, YML07]. Model-based [WL05g]. Modeling [AKS06b, BFI06, CRC08, EKS08, EH08a, KTV07, LMT^{+}09, NSS06a, NSS06b, SMC08, TNS07, WdRS07, WL08c, XWSL08, BHCK08, Che07e, Chu05, DV06, DX07b, FJ06, GD08, HHL08, KAMR05, KO06, LH05, ML08a, MS06a, Mea05, MKA07, NKR^{+}08, Nwo06g, PSBM08, QSS08, RBvB08, RM05, Rao05, RM06, Rao06b, SJ08a, Sey05, SS08b, SG08b, ZPT^{+}06]. modelled [SM07b]. Modelling [AS06i, AS07c, AT07b, Che05i, GCSS05, GRF07, KM07, NTO06, NAM^{+}07, NMS07, RYE07, RS08b, Cel06a, Dem05a, DB06, DS05b, GGSN06, JRW08, PDGR07, SS05a, SVMG09, SSTD05a, Val07, Vyn07]. Models [Ji05, Abd07, AK05b, Ali05, AA07b, AK05e, BSN^{+}08, BI08a, Bog05c, BD09a, BRRGT07, CG06, DJM08, DK05, DDZ08, EN06, EG05, ESK07, Ewa06, FT07, FA05b, Fer09, GB08, GG09a, HKH06, Hon08a, Hor07, HMSZ07, JMLR05, JLS^{+}05c, JLS^{+}05d, JSdM07, JB05, KMN09, KMVL09, LS09b, LW06b, LL06c, Liu06f, LLY06, MA05e, Mel08, MA05f, MS07d, MNN05, MK07b, ME05, Nwo06a, Nwo06d, Nwo07a, OF06, Pam05b, PQJ08, Sak05, SH07c, SC08c, TYCL06, WL05a, WLT05a, WW06, Wan08f, WXW09, WSCL07, WWLX06, YZ07b, YZLD08, Yuj07, ZGK06, ZWL06, Zha07e, d’O05a, ZLW06b]. models/algorithms [Nwo06a]. modification [BBES07, CN09b, JL05, Kos06b, KLW06b, Kou07f, MK08b, MK10, NG06a, Odi06c, Odi07, yWyCqD06]. Modifications [Tso08, Chu07b, CH08a, Lay08, WX06, Zho09b]. Modified [Abb06c, BNK07, EELZS05, HN07a, HCC06, JPZ06, Kah06b, KT07a, KLW06c, KL07e, KL07f, NNM07, NG07, NN07e, Raf06, RAZ07, Sin09, Van05, Xia08, Abb05, AD05a, AHR08, BLW06b, BB06b, Bha08, Bou06b, BK05b, CPY09, Cha05c, Cha06, CS07e, CZZ05, Che07f, Che08b, CM08a, CX08a, DH06b, Dem05d, DR06, qDyC08, DLJ08, jF09, GZ07b, GK05a, GJ07e, GJ07h, Gon07, GbL07, HRLC08, Hos06c, HBN06, Ide08b, Inc07a, Ism08b, Jön05, KC07a, KBH06b, KLW06e, KL07a, Kou07b, Kum06, LL05b, MYG08, MLW09, MD08a, MYSA05, Meš07, MK07b, NS07a, NKH07, PK08d, PA09, PKT06a, PKT06b, Psi07, QFF06, QFX06, QFS07, QW08a, RCB06, Ray06, RCB08, RW07, RW08, Ren08b, RA09, RV08, SNH07c, SGC09, Sha08d, SG07b, Shy06, SM07b, SC07c, Wan06e, WC06b, Waz05j, Waz06d, Waz06i, Waz06h, Waz07f, Waz08s, Wu05a, WR06]. modified [WZ06j, WLLC07, WZF08, Yan05a, Yan08a, YHS09, YYYM07, YY07a, YGW09, Yun07b, Yun07c, ZZTJ06, ZW08c, ZZYZ06b]. modular [CL05a, SCS06, SHSZ07, Sun07a, WLLC07, YCW05]. modulation [Che07h]. moduli [CL05a, LHL06, ZXL09]. modulo [Wan05f]. modulus [Bil04, Bil07, GOPP07, Sav07]. Moghadas [d’O05c]. Mohamed [Pan08b]. moisture [LX06a, LX06b, YC06c]. Molecular [DN07a, YSCC05]. molecule [Har06]. MOLP [JF05]. moment [BLP07, GITV06, Hos08, IPPT05, IT06a, SGUK08]. moments [Bry07a, Eng06, Hos07, IPPT05, PBPF08, WLC06, PBKM09]. momentum [SM08e, Zha09d]. monetary [LXJL09, MTD05, SYT06]. monitor [LZL^{+}07]. Monitoring [AN07, DAQ09]. monodomain [NRLT07]. monopoly [NR08a]. monosyllables [LCL06]. Monotone [Pao07, SM06c, WTW08, ZS09a, AS08, Bog08, CO09, DW09c, JB08, yLcCsL09, xLyW06, Mou08, PP08, RFFFCC07, RX09c, SH06a, Taw06, WZ09a, YY07a, ZW08b, ZH09a]. monotonic [GQS08, QG08, QCDS09, SCM09, Shi09b, WFL06]. Monotonically [Yao08b]. Monotonicity [GL05c, AV09, HH07b, WH08b]. Monte [CWY06, EFE08, FE07, FE08, JLRB05, JLBR07, JLBR08, PQJ08, RAG^{+}07, VV06, Vaj07d, WMC08, ZZX05]. Montgomery [JY07a, LLWC07]. Moore [HZ06c, PH06, TSP07, Tia09, fWZpXzL09, ZCW06]. MORAP [CK08]. MORAPs [OASM05a]. MORPs [OASM05b]. Morrison [BK09, Yin09]. Mortar [Bi06, KYL05, CR06, WYQ05]. mosquito [TML07]. most [AD06b, Ghi06, SG07a]. motion [AL09, AH07c, BMD08, Çet07, CCC07, EB05d, GY08, GC06b, HES05, HAM04, HAA07, KSKK06, Kös06a, MP08a, RMA07b, Rig08, SBS08, Shi05b, Udw05c, Wan07b, WsZ07a]. motions [CC08d, dSF05]. motivated [AE06b]. motor [HJB06]. mountain [Afr05, Den09c, YZ09b, Fig08]. movable [Fro06]. movement [MC07]. Moving [ST07, SS08d, ADH06, Ahm07b, Cel06a, CZ05b, CZZ05, Gül05, LXJ09, Ram07h, Rao06b, SSW09, WN08a, WXW09, ZC06a, ZC08a]. MP [CI09, DD09]. MP-inverse [CI09]. MPEG [Par08, YZDW07]. MR [LP07a]. MRA [Kan06a]. MRI [IZIO09, Su08b]. MRLW [KRA07]. MSE [Kök06]. MTL [VR08]. much [Bou06c]. mud [FSÇ06]. mud-flow [FSÇ06]. Muhammad [WZ13]. Muller [Wu05d, BRW08b, WRB07]. Multi [AK05g, DCL07, Isl08, Kun08, Lin05a, LXZ09, LST09, MR06c, MD05, PN06, ROK09, TMRV06, zWxY07, YML07, YYWW07, ZZX05, ASA05, AH06c, AD09a, AK05e, AN06, AA05, Ash06, AK06i, ATM06, AKKS06, BCW05, BA09, CK08, CC05a, CWY06, CHY05a, CCL06, CCFC08, Che05b, CH06b, CS06a, Che06g, Che06c, CZ08a, Chi05, CT08b, DGJ07, DGB08, Dao05, DS05a, DG07b, DDZ08, DHG09, EMESES05, GHZS07, GSDB06, GT05a, HHYW07, HWH05, HT06b, HBN06, IR07, JHST05, JDSR08, JWPS08, JGZ07, JyGcL^{+}09, KW08a, Kök08, KLZ07, KA05c, KX08, LC05a, LL05d, LSW06a, LYKC07, LY08a, LWWS09, LW09b, LBS05, LZ08e, Lin05b, LDL08, qLJ05, Liu08b, LSN09, LCZ05c, LW05c, MMW09, MR06a, MBM06, MFF09, MD09a, Moh07c, Mom06, NNN05, NZHW07, Noo06b, PW05, PN08, PCS07, Ram06f, RSA07, SA08b, SH07d, SC05a, SJ06a]. multi [Sit06, SY08a, SWZL07, SHSZ07, TLB07, TWL05, TTC08, WL05b, WW05a, WGF^{+}06, WS07d, WD08, WZC08, WLLC07, nXwCpC06, YX08, YT06, YZ07c, ZL07a, Zha08f, Zha08e, ZOZZ08, ZS09a, Zha09h, ZLYL08, Zhu05c, ZHZ07b, Mav09]. multi-armed [KX08]. multi-asset [CWY06]. multi-choice [BA09]. Multi-component [AK05g, AK05e, AN06, AKKS06, KA05c, qLJ05, YX08]. multi-confirmer [GT05a]. Multi-criteria [MD05, TMRV06, Sit06, YT06]. multi-dimension [MMW09]. multi-dimensional [Dao05, DG07b, LWWS09, MFF09, Moh07c, Ram06f]. multi-exponentiation [SHSZ07, WLLC07]. multi-function [YZ07c]. multi-homogeneous [LBS05, TLB07]. Multi-innovation [DCL07, LXZ09]. multi-input [DDZ08]. Multi-item [MR06c, IR07, MBM06]. multi-layer [HT06b]. multi-minima [nXwCpC06]. multi-mode [JDSR08]. Multi-Monte [ZZX05]. Multi-objective [Isl08, YYWW07, ASA05, AD09a, Ash06, AK06i, ATM06, CK08, CCFC08, JHST05, LYKC07, LW09b, MR06a, NNN05, RSA07, WD08, Mav09]. multi-order [DGJ07, Mom06]. multi-pantograph [LL05d]. multi-parameters [LY08a]. Multi-period [Kun08, zWxY07, YML07]. Multi-point [LST09, CT08b, DHG09, JGZ07, KW08a, LZ08e, LDL08, Liu08b, MD09a, SWZL07, WS07d, ZL07a, Zha08f, Zha08e, ZS09a, Zha09h, ZHZ07b]. multi-population [GHZS07]. multi-precision [GSDB06]. multi-product [SJ06a]. multi-products [PN08]. multi-projection [LSN09]. multi-proxy [BCW05, LC05a, LW05c]. multi-resolution [Lin05b]. Multi-response [PN06, Kök08]. multi-resultant [WL05b]. multi-scale [ZOZZ08]. multi-scroll [HHYW07]. multi-secret [CC05a, CHY05a, PW05, SC05a]. multi-server [AA05, TWL05]. multi-signature [BCW05, Chi05, HWH05, LC05a, LCZ05c, LW05c]. multi-solitary [ZLYL08]. multi-solitons [JyGcL^{+}09]. Multi-solution [Lin05a]. multi-species [Che05b, CS06a, Che06g, Che06c, CZ08a]. multi-stage [CH06b, SA08b, WGF^{+}06]. multi-step [HBN06, JWPS08, PCS07, SY08a, Zhu05c]. multi-supplier [CCL06]. multi-swarm [NZHW07]. multi-symplectic [KLZ07, LSW06a, WW05a, WZC08]. multi-team [AH06c]. multi-temporal [TTC08]. multi-term [DGB08, EMESES05]. multi-valued [Noo06b]. multi-wavelets [SH07d]. multiagent [ZLH^{+}09]. multibody [PR09a]. multicommodity [GH07, MH08b]. multicomponent [YZ08, YL09c]. Multicriteria [EMP08, Lin06a]. multidigraph [XLKH09]. multidimensional [AD06c, DESG08, MM05d, Ram06e]. multiestimation [dlS06c]. multiexponentiation [YGL05]. multifollower [SZL^{+}07]. Multigrid [DM06, Mei06, BN09, BSKS07, GAV08, KH07a, MSM08, PVN06a, PVN06b, SC05b, TK07]. multigrid-based [MSM08]. multigroup [Mei06]. multihop [Alr05]. multilayer [KM05f, LYCW09, ZMM06]. multilayered [Lee05c, LWH06b]. Multilevel [Yin07, AE07a]. multilinear [KES09]. multimedia [AS06c]. multimodal [GG05b, IZIO09, WZ05]. Multiobjective [GPTU05, Lin06b, LG07, ASAS06, AEA05b, AY05a, AJRLRGOG09, CXZL09, ElS06, GG09a, Kas07, OASM05a, OASM05b, OASM06, Ql08, Saa05, YSL05]. multiparameter [AMN07b, ANM07h, Muk07c]. Multiparametric [SGB05, GS06e]. Multiple [Che08c, CT09a, GZ07c, HW08b, JG07b, JFG08b, JZ08c, KW08a, KW08b, LS06b, Li08d, MD05, Mea05, PS08, PFG08, SL06c, TJG08, WFL06, Waz07c, Waz07e, Waz07d, Waz08b, Waz08g, Waz08h, Waz08k, Waz08n, Waz08m, Waz08p, Waz08l, Waz08o, Waz08j, Waz08i, Waz09d, WLL07b, fW07b, Xu08a, XZ09b, XT08, ZZ08a, ASAEE06, AMN07f, ANM07f, ANM07j, ANM07i, AO06c, BK08a, BR08a, BLH06, BEvdB07, CL09a, CZ07a, CX06, CT08a, CLW09b, CAS08, CN09b, FLO07, GY07b, GS06f, HKH06, JY07b, JWL^{+}07, JOA05, Jia09, KOP07a, KT06, KS08d, Kla08, KT07c, LL05e, LmZ08b, LZ08c, LLC09a, LXZ09, MX08b, MMAHN08, MY08, MR07c, MF09, Nam08, Net08a, PG08b, RM05, Rao05, RM06, RB08c, RC06, RLdLSF07, RS06, SML05a, SAS07, Sam06, Sha08a, SXL09a, SW05b, Tak05, TWZ07, TZX07, TKJ08, TTG07, WZL08a, Waz08c, Waz08d, Waz08e, Waz08f]. multiple [Waz08t, Waz09b, Waz09a, Waz09c, Waz09e, WLZ06, WX06, XW06, XXX06, YW05, YRY05c, ZnCxX09, ZFT06, Zhu05d, ZJO08, dlS07a, Waz08f, Waz09c, Waz09e, ZnCxX09]. multiple-attractor [JY07b]. multiple-decision-maker [HKH06]. Multiple-front [Waz07c, Waz08g]. multiple-input [LXZ09]. multiple-key [SW05b, YW05, YRY05c]. multiple-precision [Tak05]. multiple-singular [Waz09d]. Multiple-soliton [Waz07e, Waz07d, Waz08b, Waz08k, Waz08n, Waz08m, Waz08p, Waz08l, Waz08o, Waz08j, Waz09d, Waz08c, Waz08d, Waz08e, Waz08t, Waz09c]. Multiple-watermarking [Che08c]. multiplicand [YLM07]. multiplication [BK07c, BK07d, DwWmW05, Tak05, WLCL06, WLC07a, YCW05, YLM07]. multiplicative [CS08a, GXY06, SJ06b, SG08c, SDM08, XRZ09, You05, ZZ07c, gZW08]. multiplicities [Gra09]. Multiplicity [BR09c, LCJ08, ZC08b, AO05a, hCxC05, DLT08, MH09, PG08b]. Multiplier [AK09, BGR05, Ham05a, SWC05]. multipliers [CR06]. multiply [WLC05, WLC06, WLC07b]. Multipoint [GK06c, JFG08b, Kan06b, SQG08, Xu07]. multiprocessor [KW05b, KW05a, Kay05d, Kay05e, ZŠ05]. multiprocessors [Kay05a]. Multiprogramming [XCS07]. multiquadric [CW06b, CW07, Lin05e]. multiresolution [AM08b, AM08c, AM08d]. Multiresponse [Kök06, KY06]. Multiscale [QWX08, BC09a, FHG07, GFH08, mHsLzC06, Reb06, ZLL06a, Zha07f]. multisignature [CHYC05, LWK05b, YY05b]. multispecies [Che05c, Che06e, Che06d, WS07a, WW05b, YC06b]. multisplitting [SH08b, Son05, Yua05, YK06b, Yun07a, Yun09, ZHGG08]. multisplittings [Yun09]. Multistability [VKVHN05]. multistage [AK06i, Lin06b, Lin06a, LG07]. Multistart [VL09]. multistate [Lin06c]. Multistep [LY09d, HHW09, Hua08b, Li08a, LW09c, Psi09, Rao06a, RGAFVA07, SA06a, TFX07, VR08, WL05f, XT06, XZ09a]. Multisymplectic [Che05e, Che05f, Che07e, CL08b, KLX06, LSW06b, Wan07a]. multisymplecticity [Wan08c]. multithreading [GEG07]. multivalent [Aou08a, EA08a, Fra07, HL08d, Liu07c, LP08b, Liu09b, PS06c, Sok08a, SYX08, YL08a]. multivalently [Irm07, SEŞ09]. multivalued [ASK05, OZ07]. multivariable [WDL07]. Multivariate [HYG08, Lin05e, WN08c, CX09, Li06d, Li07b, MJ07d, SAS07, VS07]. multiwavelet [SCQ06]. multiwavelets [AM08c, CC09b, hCxC05, MK05c, SH06b]. muscles [ZS08b]. musical [HT05]. mutants [Kwo07a]. mutation [DT07b, ZGH07]. mutual [LC09b]. mutualism [CY07b, CYCX09, LTCC08]. MV [CLW09a]. myeloma [LBZY08].

N [Pan08c, Akı06, CAS08, EM08, Fen06, GJ07b, HJ08, HL05c, KEM08, Kıy06, LZ08e, LM09, Luo06, Neh07b, PW05, Prá08a, Sha06a, Sha06b, SC05a, Wei06b, XGTS07, ZLW08, BGP06, GS06f]. N. [CI05]. N.T [Pan12]. Nagumo [LG06a]. Nano [SJ08a, SJ08b]. Nano-robotic [SJ08a, SJ08b]. nanomanipulator [SJ08a, SJ08b]. nanoparticles [MKA07]. nanotechnology [Hua05a]. narrow [BMD08, BAK07, YXSX05]. narrow-band [YXSX05]. narrowing [YST09]. Nash [gDH07, Ino08, SMY08]. Nasser [Pan12]. national [KAM06]. nations [Chu05]. Natural [MRN09, AH05a, Che06n, HYZ07, Kol06, KOS08, KS05b, LC05b, Lin08, Sha08d]. nature [RR06]. Navier [JH09, AM08a, Cao06, DC06c, EGFAE06, FHYZ09, GJ05, GFH08, Li06b, cMpZfR06, MO06, Ran08b, SH09a, SRH09, ZH05]. NCP [RZX07, WLL09, lZZ06a]. NCPs [Hua05b]. NDE [Joy06]. near [Ahm05a, Che07i, Dil05, FZN07, Gür05, GD06, Kıy06]. near-optimal [FZN07]. nearest [AA06a]. nearly [LL08g, NA08]. nearness [HYG07]. necessarily [MeT05]. Necessary [Via08, Cai06b, CC06c, DZ07a, Wei06a]. neck [PB07a]. Necmettin [Sub07]. needs [LT09]. neg [Fat06]. negative [Alt07b, Aou07, CTL09, Den07, Fra07, LWH06a, MSTY06, STKB05, SD06, WLT07b, XTCC08, ZQ08, Zho05]. neighborhood [ASPA09, SEŞ09, YMCL06]. Neighborhoods [Alt07b, Fra07]. neighbouring [Rim07g, Rim07h, Rim07d, Rim07f, Rim07e, Rim08b, Rim09b]. Nesterov [WB09]. net [CHW07, Far08, KT08a, WWLX06, WL08c]. NetLogo [Dam08]. Network [YC08c, AO06a, AN07, AS07b, Alr05, AGA^{+}07, Bar09, CS05a, Chi07a, Chi07b, Chi07e, Chi07d, Chi07f, Chi07g, Chi07i, Chi08a, Chi08b, Chi08c, CK06c, DF08, EB05b, EB05a, EN06, EGN07, EJ07a, ETJ06, FEOC08, FG07, GOMA06, Gop07, Hu07a, HHK^{+}05, JLS^{+}05a, JZW08, LFM06, LHGT06, Lin06c, Lin07e, Lin07f, MA07d, MMC^{+}08, Mel08, MS07e, MH08b, Par08, PFA07, fQZgH08, QWX08, SMB08, SS05e, SP06d, SM09a, TLC05, Wan05d, WiN06, Wan09b, WS07g, WRL06, XLSL06, XLH^{+}07, XHSL09, YC08a, YM08a, YWZW07, ZGK06, ZZLL07, ZSCW09, ZW06a, Zho07, ZH07b, AGS07, MB06]. networked [CZH08, DF08, FDL08, YWZW07]. networks [AS06c, ANA07, AGTS07, APS05, Aza06, ATM06, Bai08, BSKK06, BR08a, BK08b, CZH07, CMC06, CT07b, CZF09, Deh08, DZ08e, EJ07b, EAG08, EH09d, FWJK09, GPR^{+}09, GS06c, GS07a, GH09, Guo09, HH06a, HRM07, HZ05, HH08b, HLT09, JRW08, JSL06a, JSL06b, JL09, JH08c, KH05a, KH05b, KB06b, KLL^{+}08, KAM06, KP08a, KP08b, KP09b, LX07a, LYJ09, Li09c, LH08c, LL07b, LP08c, LZWK09, MYSA05, MO05, MY05, MGH08, OKEL05, Par06a, Par06b, Par06c, Par07, PK08b, PK08c, PKL08a, PPKL08, Par08, PKL08c, PK09, PSRN06, PK08d, PSBM08, PCCHLG08, QSS08, RB08a, RB08b, Rey05a, Sal05, SSA09, SGP07, Sin08b, Sin09, WZ08a, Wan05c, WL07b, WZL08a, WYL08b, XLL06, XCL09, XS09, YLL^{+}05, YX06a, YZW07a, YM07b, YY07c, YZD08, YZFL08, ZWX05, ZS07b, ZYT09, Zha09d, ZD06b, ZZW09b, ZL06c, ZLL06b, ZW08d, ZH08b, ZL09a]. networks [ZXL06]. Neumann [AMN07d, BD09b, Dai07, DjT05a, GB07, JM06b, LCJ08, Li09a, Liu05a, SCO09, YZ09b]. Neural [EN06, Sal05, AO06a, AN07, ANA07, AGTS07, AGA^{+}07, Bai08, BSKK06, BR08a, BK08b, CZH07, CZF09, DZ08e, DF08, EB05b, EB05a, EGN07, EJ07a, EJ07b, EAG08, GPR^{+}09, GOMA06, Gop07, GH09, Guo09, HH06a, HCS05, HZ05, HHK^{+}05, HH08b, HLT09, JRW08, JSL06a, JSL06b, JZW08, JL09, KB06b, KP08a, KP08b, KP09b, LX07a, LYJ09, Li09c, LH08c, LHGT06, LL07b, LC08a, LZWK09, MO05, MY05, MB06, MA07d, MGH08, Par06a, Par06b, Par06c, Par07, PK08b, PK08c, PKL08a, PPKL08, Par08, PKL08c, PK09, PSBM08, PCCHLG08, fQZgH08, QWX08, RB08a, RB08b, SSA09, SA09, SS05e, Sin08b, Sin09, SM09a, TLC05, WZ08a, Wan05d, Wan05c, WiN06, WL07b, WLP07, WZL08a, Wan09b, WRL06, XLL06, XCL09, YX06a, YZW07a, YC08a, YM07b, YM08a, YWZW07, YY07c, YZD08, YZFL08, ZWX05, ZS07b, ZZLL07, ZSCW09, ZYT09, Zha09d, ZD06b]. neural [ZW06a, ZZW09b, ZLL06b, ZW08d, ZH08b, ZL09a, AGS07]. neuron [GH09, MSM07a, ZH07b]. neuronal [SS08b]. neurons [HH09b]. neuroscience [OF06]. neurosciences [RJW07]. Neutral [JJ09b, Bai08, BC09b, BGF09, CD05, Che05d, Che06d, CE07b, CE07a, Fra09b, GS07b, HL08b, HHW09, HL07d, HS09b, Kar09c, KMÖS09, KPL08a, KP08c, KPL08c, KPL09b, LLZ06, LKU07, LLZ08, LKU09, LWKU09, LYL07, MM07a, MX07a, MX07b, NL05, NPGW09, PK05b, Par05a, Par05b, PK05c, PK05d, PK08b, PKL08a, PPKL08, PKL08c, PK09, Par09b, RB08b, RX09a, RX09b, SA09, Sha08b, SL07a, SS05g, TTG07, Wan05b, WZ07a, WL07c, WLZ08, WWL08a, WWL08b, WWD09, WQL09, WWY08, WWY09, WZY07, WZM06, XM06, XM07a, XM07b, YZ07b, YX09, YWL07a, YWL07b, ZLZZ07, ZG08a, ZL08a, ZQJ09, ZXL05, ZXDL05, ZXWL06, ZWY07, ZM08c, ZY07, Zho08b, ZZ08b]. neutral-type [PK08b, PKL08a, PPKL08, PKL08c, PK09, RB08b, SA09]. neutron [Mei06]. Neville [ACDR09, LMV09]. newborns [Far08]. newsboy [DM07a, JS06, SJ06a]. Newton [ATL07, EDMJ05c, HEDMJ06, HMJED06, Abb05, AE06a, Ade07, All05b, AH08b, Arg09, BD06a, jB06, Bas08, BGR05, BV06, CL05c, Chu07b, CH08a, CN09a, CN09b, CT06, CT07c, Dan08, DB07c, DMJE05a, DMJE05c, DMJE06a, EHR09, FHH09, FHSS08, Gao06, GKR06, Geu07, GA07, GJ07f, GJ07h, GKK09, GSDB07b, HZ06a, HZ06c, Ide08b, KT07a, KS05a, KLW06a, KLW06e, KLW06b, KLW07b, KL07b, Kou07b, Kou07f, KLR07, KLS09, LV05, LDZ06, tLlWpL08, LN08b, Ma05a, Muk07b, PHP09, RMS05, RA09, Ser05, Sha05e, Sim09b, Śmi07, SK09b, TLM09, Thu08, yWyCqD06, WZT06, WY07, WY09, WLQ06b, WLY08, Wu05c, WZ06i, WS06b, Wu06c, WZ06j, Wu07g, WZ07d, Wu07e, Wu07b, Wu08b, WHL08, XW06, YHTR05, ZMW05, ZWG09, ZW08c, ZZYT08, Zho09b, Zhu05d, Zhu05b]. Newton-based [KLS09]. Newton-CG [FHSS08]. Newton-homotopy [ATL07, Wu05c, Wu06c]. Newton-like [AH08b, GJ07f, KLW06a, KL07b, KLR07, Śmi07, Wu07e, Zhu05d]. Newton-Quasi [WY09]. Newton-type [DB07c, EHR09, FHH09, Thu08, ZMW05]. Newtonian [Pan08c, AES05, AAN05, AB05c, AV07a, AuH07, And06, Att05a, BT06a, BT09b, DC06a, ESAR05, EESF04, GM05a, ICM08, JK07, MCMA07, NMS06, SG09c, ZZL08]. newtonization [Gra09]. Ni [XFTY09]. niche [HHZ09, WZ05]. Nicholson [Lin09a, Sak05]. Nicolson [GR06b, HXH08, KA06a, Sid09]. Nikodym [Kim07a]. niloticus} [MD07b, MD07b]. nilpotent [AGR09]. nine [Sil05c, Sil05b, Sil06a, Sil06b, Sil06e, Sil07a]. nine-point [Sil05c, Sil05b, Sil06a, Sil06b, Sil06e, Sil07a]. ninth [Waz06e, Waz08b, Waz08o]. ninth-order [Waz06e, Waz08b, Waz08o]. Nizhnik [WLZ05]. NLS [ESK06]. NMNF [Lin07d]. NMR [CH06a]. NN [LLLX07]. NNAF [DwWmW05]. no [CC05b, CSY07, Dav06, Wan08b]. no-idle [CSY07, Wan08b]. Nodal [MM07b, CMV05]. node [BDDB08, Lin07e, She05b]. nodes [EF05, EFI06, HED06a, Has07a, Lin06c, MJ07e, TD05]. Noise [LHXL05, CH05a, RSZ08, SMY08, WGL08, XXTL09, XRZ09, XHLL05, YXSX05]. noise-perturbed [XXTL09]. noisy [JJ09a, LL05h, PWL06, Sch05, WL06f]. nominal [lS08b]. Non [AD06a, AAN05, BS06, CZ08b, Che06n, DAQ09, dCDLN^{+}07, EH06, dCJLNÁA08, LP07c, RJM07b, SH07b, SH07c, STKB05, Shi09b, uIKTT05, uITiHK08, Soh06, TiHuI08, ZQ06, ZQ08, ZM08d, AEA05a, AES05, AKBZ08, AS06b, AV07a, AH07b, AuH07, Ali05, And06, Ang06, ADK06a, AKM06, Att05a, BM06a, BFR05, BGRS09, BG07a, BG08, BT08, BT06b, Che05a, Che06a, Che06h, CCC07, Che07g, CRC08, ĆC09d, DGB08, DLLN09, DJM08, Deh05a, fDxZfWyY09, Dod08, DC06a, EB05e, EGS05, EL05a, EL05b, ESE05a, EESF04, Ema06b, FX09, GOPP07, GR07a, GJ07f, GJ07g, GM05a, GMS06, HBBA07, HES05, HAM04, HD07, HL06c, HT09a, HT09b, Hub08a, Hub08b, Hub09b, Hub09c, Ide08b, IMB09, ICM08, IMS09, Jai07, Jan07a, Jan07b, JK07, KK08b, Kad07a, KK05a, KB06a, Kan07, KWS05, KS08b, KSM09, KLW06a, KLW07a]. non [Kou07c, KLW07c, Kou07a, KL07f, Kou07e, KW07, Kou07f, KX08, LL08a, Lep08, gLXyZ06, xLyW06, LMH06, fLsLdXlF08, LWW09, Liu06b, LL07a, Liu08b, LJE07b, Ma07c, MA08d, MI08, MBM06, MK05c, MH06a, MCMA07, MSD06a, Meh06, MK07a, MN05, MZ07, MR07c, MA05f, MK06, Moh06, MA06e, MS06b, MA06f, Moh07a, MD09b, MD08c, NS06a, NCÁHC^{+}06, NS07a, NDD05, yNfM06, NHKZ06, Noo08b, OMR07b, PM09, Pan08b, Pan08c, PK08a, PQJ08, PS06b, Psi09, QFX06, RAS07, Ram06e, Ram07a, Ram08a, Ram08c, Ram09a, RJM07a, RMJ07, RX09c, SBS08, SML05b, Sae06b, Sae07, Sal08c, Sal08d, SB09, SM05b, SH06a, SX06, SG08c, SA07a, SA07b, SA07d, SA08e, SM07a, uITA06, uINTK06, SM05c, SC07b, Son08b, SL06b, SC05d, Sun05b, SZ08c, TGSNB07, TWZ07, TG08a, TQC09, VR08, WL06d, WCO08, WFS08, WLH09, WL09b, WP05, WN08b]. non [WZ06g, WZ08f, XCM07, XXC08, XFL06, XZ08b, XTCC08, YEM05, YXSX05, Yan08d, YL09a, Yu09, Yun08a, ZG09, Zen09, ZO05, ZX06c, ZCMQ07, ZDY08, ZY08b, ZHX06, ZGH07, ZZL08, ZMV05, dAC09]. non- [AH07b]. non-accessible [GMS06]. Non-adaptive [DAQ09]. non-analytic [GOPP07, WL06d]. non-Archimedean [Ali05, MA05f]. non-autonomous [Che05a, Che06a, Che06h, LWW09, Meh06, MN05, TG08a, TQC09, XZ08b, YEM05, Yan08d, Zen09]. non-binary [Sal08d]. non-classical [Ang06, BT06b, Deh05a, GJ07g].

Non-conforming [AAN05]. non-convex [Noo08b, SG08c, WZ08f]. non-Darcian [Pan08b, AEA05a]. Non-Darcy [Che06n, JK07]. non-defective [IMB09]. non-dense [dAC09]. non-discretionary [LJE07b]. non-expansive [LL07a, SC07b, Son08b, SL06b]. non-gravitational [SBS08]. non-homogeneity [Sae07]. non-homogeneous [Jan07a, Jan07b, Liu08b, SML05b, Sae06b, SM05b, TWZ07, YL09a, ZMV05]. non-ideal [BFR05, CCC07]. non-identical [WZ06g]. non-ignorable [PQJ08]. non-independent [EGS05, NCÁHC^{+}06, WZ06g]. non-integer [LL08a]. non-interior [RX09c]. non-isospectral [SZ08c, Yu09]. non-isothermal [PM09, ZO05]. non-iterative [Ram06e, Ram08c, Ram09a, Yun08a]. Non-linear [dCDLN^{+}07, dCJLNÁA08, SH07c, AKBZ08, ADK06a, AKM06, BG07a, BG08, DGB08, DJM08, fDxZfWyY09, EL05a, EL05b, ESE05a, Ema06b, FX09, GJ07f, HES05, HAM04, HT09a, HT09b, Hub08a, Hub08b, Hub09b, Hub09c, Jai07, KK08b, Kad07a, KB06a, Kan07, KSM09, KLW06a, KLW07a, Kou07c, KLW07c, Kou07a, KL07f, Kou07e, KW07, Kou07f, fLsLdXlF08, Liu06b, MA08d, MI08, MBM06, MK05c, MSD06a, MK07a, MZ07, MR07c, MK06, MS06b, MA06f, Moh07a, MD09b, NDD05, yNfM06, OMR07b, PK08a, RAS07, Ram09a, RMJ07, Sal08c, VR08, WLH09, WL09b, XCM07, XXC08, YXSX05, Yun08a, ZG09, ZHX06]. non-linearities [EB05e]. non-Lipschitzian [NS07a]. non-local [BM06a, Dod08, IMS09, MD08c, SM05c, ZY08b]. non-matching [KK05a]. non-monotone [xLyW06, SH06a]. Non-monotonic [Shi09b]. Non-negative [STKB05, XTCC08]. non-Newtonian [Pan08c, AES05, AAN05, AV07a, AuH07, And06, Att05a, DC06a, EESF04, GM05a, ICM08, JK07, MCMA07, ZZL08]. non-orthogonal [KWS05]. Non-oscillation [AD06a, Sun05b]. Non-oscillatory [BS06]. Non-parametric [EH06]. non-periodic [BT08]. Non-perturbative [Soh06]. Non-polynomial [RJM07b, uITiHK08, TiHuI08, AS06b, RJM07a, SA07a, SA07b, SA07d, SA08e, uITA06, uINTK06]. non-propagating [Ma07c]. non-reflexive [WCO08]. non-repudiation [SC05d]. non-resonance [WP05]. non-self [ĆC09d]. non-self-mappings [GR07a]. Non-similar [SH07b]. Non-simultaneous [ZQ06, ZQ08, ZM08d]. non-singular [ZCMQ07]. non-smooth [BGRS09, FX09, HD07, Ram08a, SX06]. non-square [NS06a]. Non-standard [LP07c, PS06b, QFX06, Ram07a, XFL06]. non-static [SM07a]. non-stationary [DLLN09, HBBA07, KX08, LMH06]. non-stiff [Psi09]. non-symmetric [Che07g, TGSNB07, WN08b]. non-symmetry [gLXyZ06]. non-topological [SB09]. non-toxic [CRC08]. Non-travelling [CZ08b, ZX06c, ZDY08]. non-uniform [HL06c, KS08b, Lep08, MH06a, Moh06, MA06e, WFS08, ZGH07]. non-variational [NHKZ06]. nonautonomous [Che06f, Che06g, CS07f, CH07c, FK06a, Lar05, LC06a, LZ07b, MXC07, WZ06b, ZLT09, ZZL08, ZSS08a]. nonce [CY05b, LKY05a]. nonce-based [CY05b, LKY05a]. noncommuting [AR09]. noncompact [Inc07c, Waz05d, Waz05e, Waz05f, Waz05h, Waz05j, Waz05k, Waz05m, Waz06f]. nonconfirming [GFH08]. nonconforming [AN07, HJK08, KYL05, SW09a, SP09, SRH09, WHL06]. nonconvex [AH08c, GXS05, QZZ07, SDM08, WLL09, Yua08, Zhu05f]. noncylindrical [FM08]. nondifferentiable [AH06a, Ren06]. nondiscretionary [Sae05, Sae06a, Sae06c]. nondominated [JLS^{+}05a]. Nonexistence [AM06, KL05b]. nonexpansive [CY08a, CLCK09, DL08, HN07c, JK06, KT08a, LjZH08, yMT08, NS06b, NS07b, NH07a, NY07, NH07b, NH07c, PP08, QN08, QCS09, SZ07a, SC06b, SQ07, Wan08g, YN07, YY07a, YLK09, ZS07a, ZH09a, ZWC06]. nonholonomic [GC06b]. nonhomogeneous [GSG06, TT09, Tua13, WFHD09]. Nonic [AS06a]. noninsulated [DS05b]. Nonlinear [BAM07, CN05, EJ07b, HBBA07, JSLL07, JZW08, KR07d, OS08, SS09d, Sta06b, Sya05b, WWL08a, WWL08b, Waz05k, WL05f, WWYL08, WF07c, XCJ09, XPS07, Xu08b, YS07, YWL07a, Zha09a, dSF05, Abb05, Abb06c, AE06a, AJ06a, ATL07, AESD05, Abd07, ASA05, ADH06, AAS06, AK06e, ANM07c, AB05b, ASK05, AV07a, ADY06, AKARA08, ALD08, AD07, AD09b, Amm06, ALL07b, AE07c, AGPJV09, AH08c, Asa06b, AÖ09a, Asl09a, Asl09b, AÖ09b, BFR07, BK07a, BKA07, BCCK05, BZ09, Bar09, BP09, BSN^{+}08, BSC06, Bay05a, BRW08a, BWR09, BB05, BZZ08a, BNKZ09, Bog05c, BKZ06, BRRGT07, BKM06, BG07b, Bou05, BEF08, BK05b, CLG09, ÇT07a, CS05a, CZH08, CHLS08, CC08c, CC09c, Cha05h, CWLH08, CT05a, CZ05a, CC05c, Che05f, CFC06c, CS06a, CX06, Che06m, CW06d, CG07b, CCJR07, CH07b, Che07f]. nonlinear [CA08b, CM08a, CL08d, CZ08b, Che08d, CH09, CKQ08, CSC08, Chu06, Chu07c, Chu07d, Chu07e, Chu07h, CH07d, Chu07k, Chu07m, Chu07n, Chu08, CFVZ06, CG06, CE07a, CT08b, CY05c, CD06, CY08d, DB07a, DB07b, DB07c, Den08b, DF09, DDR05b, DDR06b, DZG06, DZ07b, Du07, ERAS07, EN05, EB05b, EB05a, ER06a, EHN07, EJ07a, EA07a, EA07b, EBsENMA05, EB08, EWEA06, EWAH08, EkES05, ElS06, EELZS05, EHP08, EM09, FGfFL06, FTG07, Fen05, Fen06, FH06a, FLLH07b, FH07, Fer09, Fro06, FH06b, Gal08a, Gal08b, Gal06a, GZ05, GFWX05, GY06, GJ07a, GZ09a, GZ09b, GC07b, Gen09a, GKB07a, GKB07c, GkB07b, GKD07, GTJ07, GJB05, GJ07d, GJ07e, GJ07h, GAPT08, GSPG07, GW06b, GbL07, GHX08, HAK07b, HC07a, HW08a, HW08b, Has06d, HCLP07a, Hos06a, HLW08, HHW09, HC09, HLK06, HN07b, Hua08b, HM08b, Hua08a, HZ09b]. nonlinear [Hub05, HMT08, HMT09a, HMT09b, ID08, Inc05, IS09b, Ism08a, JDG06a, JDG06b, JDG06d, JKN09, Jan09b, Jav07, JK08, JOB08, JjQjWmL06, JFG08a, JZ08b, JXH06, JZ06b, JZ06a, JZ07, JH08a, JH08b, JC09, JLW09, JGS06, JS07b, KAV05, KMB05, KKF07, KBLT08, Kan06b, KSGM06, KOR08, Kas07, KL06a, Kha07a, Kha08a, KCW07, KSS07, Kök08, KL07h, KLR07, KL08, KD09, KC05b, KPL08b, LY07a, yLcCsL09, Lay09b, Laz08, LP08a, Lep06, LL05e, LXL06, LDZ06, LmZ08b, LmZ08c, LH08b, LZ08b, LYX08, LZ08c, tLlWpL08, LL08e, Li08c, LL08b, LW09a, LZ09b, LLG09, LYL^{+}08, LD06b, LC06b, LDL08, LYM09, LLAT09, LXML09, qLJ05, LG06c, LZZZ06, Liu06h, LZZ06b, LZ07c, LKU07, LZW07, LLW07, LKU09, Liu09a, LLZ09, LWKU09, LM07a, LMN07, LW09e, LHSB09, LHF07, LC08b, LXJL09, Ma05a, MTC07, MTL07, MJ07b, Ma08b, MCW08, MX08a, MT08, MC09, MH09, MM07b, Mah09a, Mah05]. nonlinear [MM06b, MSR07, MKKS05b, MŘ07b, Mea05, MCC07, MKSK07, MAY09, Moh09, MHM09, MAO06, MHA06, MED08, Nam09, Ned05, Net08a, Net08b, NMP08, yNyF05, yN05, yN06, NBO09, NA06a, NNMDS06, NNKA06, NA06b, NN06b, NAJ06, NN06c, NN07c, NN07d, NN07a, NNM07, NG07, NN07e, Noo07b, Noo07c, Noo07d, NKH07, NN07b, Noo07e, NNuH07, Odi07, Omr07a, Ord06b, OASM05c, OASM06, PS08, Pam05a, Pao07, PG08a, PG06, PG08b, Par05b, PL06b, Pen06, PP07a, qQcM05, Qua07, Raf06, Raf07a, Raf07b, Raf08, REDES05, Ram07d, Ram07e, Ram08b, Ram08c, Ram08f, Ram08d, Ram08e, Ram09b, RB05b, RCB06, RZX07, RV08, SF07, SA08a, Sal07d, SS07a, SS09c, SG05b, Sha06a, Sha06b, SQ11, Sha05e, Sha07c, SMG07, She07b, SC08a, SL09a, SXL09a, SDG05, SFPE07, SZ06a, SZ07b, SHG07, SZ08a, SLL08, SB05b, Sta06a, SL05b, SWX06a, SS05f, SS05h]. nonlinear [SS05g, SS06f, SMHQ06, SS07e, SM07c, SY08b, SM08d, SM08e, SV07, SA06h, TMS08, TXW09, TB09a, Taw07, Thu08, TM08, TNL09, Uje06a, UEL07, Uje09, Ver07, VJD09, Wan05h, dWqL05, WW05a, WC05a, Wan05a, Wan06g, cWmZ06, WZSZ06, WL07a, WZS07, WWWL07, WL07e, Wan07a, WL07c, WC07b, WLP07, WFX07, WTW08, WW08b, WJZG08, WC09a, WLA09, WTS09, WQL09, Waz05e, Waz05l, Waz05p, Waz05m, Waz06b, Waz06d, Waz06m, Waz06n, Waz07a, Waz07i, Waz07j, Waz08r, Waz09a, Waz09f, WLQ06a, WYL06b, WWY08, WS09c, Wu05d, Wu06b, WZM06, Wu06c, WMBY07, fW07b, WF07b, Wu07g, WHL08, XW09, XCW08, qXbL05, XHH06, XM07a, Xu08a, YST09, YLZ06, YHTR05, YC06a, YC06b, YCL09a, YX06b, YL06b, Yao08b, YM08a, YFLX09, YW06a, YW06b, Yin08b, YC06d, YC07b, YZW07b, YWL07b, Yu08b, YZLD08, YJO08, Yue07, Yun08b, ZPT^{+}06, Zha05c, ZC05b, ZZ05a, ZD06a, ZLW06c, lZ06d, ZZ07a]. nonlinear [ZFY07, ZDL07, ZZDL07, ZL08b, ZFYH08, ZLW08, ZQJ09, ZWG09, Zha09b, ZLW09a, Zha09c, ZZYZ06a, ZM06b, ZZ07b, ZWX08, ZWZ09, ZZ07d, ZM07, ZLZZ06, ZL07b, ZWZZ07, ZLC07, ZSS08a, Zho08b, ZCZ08, ZZ08b, Zho09b, ZFL^{+}09, Zhu05a, Zhu05f, Zhu05g, Zhu05b, ZL07c, ZW05j, ZG08b, d’O05c]. nonlinearities [AOLL05, CZ07a, CLWP09, LP08a, LC09a, LOA05, SM08b, SM09c, WR07a, Waz06g, XL08b, ZJO08]. nonlinearity [Bis08, BR09b, CSC08, DM08b, ESK07, GZ07d, JFG08b, LCL05, LHSB09, SLS08, YY08b, ZYW08]. nonlinearly [Cor07, GL08c, Inc07c, Liu06g, Odi08, VC06]. Nonlocal [Bab07, Kum08, AA06e, CA05, DB09, Kar09a, LL08d, hLqSlZ09, PY08a, WMX07, Xu06, Zha09g, ZS06, ZMY08]. Nonmonotone [FS05, Zhu05b, Che06m, MZW05, SS06b, Wu08d, YZW07c]. nonmonotonic [YQ09]. Nonnegative [WLA09, DZ06, GZ05, HWF07, LZ06a, Yan07e]. nonnegativity [HRT07]. Nonoscillation [SM08c, EBP09, SM08a]. Nonoscillatory [CL08d, MŘ07b, LKU07, LKU09]. nonoverlapping [DG07b]. Nonparametric [KC06b]. Nonpolynomial [SAM07, uIT06, RLZ07, RLZ08, SA06f]. nonrectangular [LY07a]. nonrepudiable [BCW05, HW05a, Xie05]. nonrepudiation [HW05b]. nonresonance [GO08]. Nonresonant [AOLL05]. nonself [DL08, Wan08g]. nonself-mappings [Wan08g]. nonseparable [LY09a]. Nonsimilar [Che07i, Che09b]. nonsingular [Cao06, Pen09]. nonsingularity [CD08]. nonsmooth [Amm06, KMA08, PL06c, PFL09, SRMRGRL08, SW05a, Taw06, Zha07b]. nonstandard [Jer08]. nonstationary [KSS07, ML08a, SK05]. nonsymmetric [BLW06b, KT06, tLlWpL08, Lin05d, LW06c, LD07d, LN06, NG06a, NG06b, NM08a, Vog07]. Nontrivial [BZ09, CZ07c, HZ08, Yin08b, ZLW06c, ZL08b, FZ09, GZC08, JZ06a, JZ08c, LW06d, ZYC09]. nonuniform [WZZ06b]. Noor [WZ13, LN07, NS07a, Raf06, XF08b]. Nordsieck [DJ06a]. Nörlund [MSM07b]. Norm [Kan09a, Ste09c, Ste09d, AS09c, JZHT05, JZTH06, JXH06, bJyKxC09, LZ07b, Ma08b, OZ07, Occ09, Ste09i]. norm-relaxed [JZHT05, JZTH06, JXH06, bJyKxC09]. norm-to-weak [LZ07b]. norm-type [OZ07]. normal [Gar07, HW05c, Li05f, Li05g, LT07, QBH06a, Qiu05, SM07b, SGG^{+}09, ZYCI08]. normality [WL06g]. normalized [KNM05, LM08a, WC06d]. normalizing [CST07]. Norms [Ste09e, Sol05a, Sol07, UNS07, Wu08a]. Note [Cas07c, CX09, Lay09b, LLZ09, SM06d, Wu06e, Wu07g, AY05b, AA06e, AA07f, AHS06, BDDB08, hB07, BJ08, Bra09, CL08a, CM08a, CSY07, Chu07d, Chu07e, Cor05a, CK07c, Dah07, Den09a, fDxZfWyY09, Du06, EM03, EM05b, Ere09b, FP09b, FFRFCC08, Gar07, GC09, HED06c, HNB06, HYL09, Hub08b, HMT08, HW05c, IL08, JMLR05, JS07b, JC08b, JC07, Kim09, Kop22, gLxW05, LCC05b, LS06d, gLL07, Liu07b, Lu07, Luo05a, LHL08, LL08g, Mao09, Mus08, Neh07a, PC05, Par07, Par09a, PP07a, Raf07a, Raf07b, Raf08, Ren08a, RX09b, Sal09, Sal06c, SWH07, Sog08c, SJ06d, SW06b, Sun07a, WG08, Wan07a, Wan08b, WLZN08, WCX05, Wu06a, WS06a, Wu07f, WR07b, WSXL09, XYL10, YZ07a, Yun07b, Yun07c, Yun08b, ZW05e, Zha09a, ZnCxX09, Zho09b, Zim07]. Notes [SS05a, Uje09]. notice [Ano07c, FV05]. Novel [CZH08, Gee08, ID08, MCAF08, Par05b, QCSF09, RC06, ZLYL08, ASPA09, CH05b, CJC05, CLZ^{+}07, CH07f, DW09a, HTO06a, HTO06b, Hub08a, Hub09c, HJL08, KWdW08, LS08c, LLX07b, Mai05, OHT05b, PG07c, PG08c, SCS05a, SC08a, TCW08, Tso09, WSZ07b, WZ07a, XwWL07, YSL05, YLL07, ZLH^{+}09, ZWLX08, Zhu09a]. Novikov [WLZ05]. Novo [CH06b]. Noyes [CJ06]. nozzle [MFE05]. NP [dCD06a, dCD03, dCDB07]. NS [ZWT09b]. NS-power [ZWT09b]. NSFOT [XZM^{+}09]. nuclear [YY09]. nucleus [Wan08d]. null [AH09]. null-space [AH09]. Number [Che08c, Ben07b, Cas07b, CLZ^{+}07, CX08b, CT08a, COG08, Cui05, DjT06, EdR07, Gin07, HK05, HS09a, HW05c, JRA^{+}07, KR07c, KT07c, Liu07b, MAN06, MR06c, MD08b, MD09c, NCGdSM07, ÖA05, PJ10, SCL^{+}05a, SS05e, TLB07, WG05, WW06, WWW03, WTG07, WL06g, ZXLZ07, ZL08e, Zho09c]. numbers [AY05a, CL06a, CX08b, CAS09, DOU08a, DOU08b, Dem08, DQW06, DM08a, Ere06, FP09a, FZ08, Kah05a, Kar05b, KAT06, LWW05, LJ07, ÖTA05, SA08c, Sof09, Sol05a, Sol07, SW06c, UNS07, fWZpXzL09, WWW03, lY08b, YK03]. Numeric [RGAFVA07]. Numerical [AA06b, ABA06, AO06a, Abb06d, AE07a, AK06d, AMN07f, ANM07f, AÖ06b, AH05a, AK05b, AK05c, ASN07, AS06d, ACZ05, AD07, AH05b, AC08, Ang06, Asa07b, Att05a, BD05a, BGA05, BF07a, BF07b, BK07a, Bad05, BCCK05, Bar05a, BNH07, Bay05a, BE06, BB06b, BM08, BG08, BBCP09, BN07, BN08, Bra06, Buk09, Cen06, CLA08, CA08a, CSC08, CLNCG07, Cor05b, Cor05c, DLLN09, DX07a, Deh05a, Dem05a, dCDNPN09, DEFT05, Dön06a, ERAS07, EBsENMA05, EHAR05, EMESES05, Ere06, Fag05, FV05, Fat05a, Fat05b, Faz09, FK09, GL08a, GKB07c, GkB07b, GKD07, GK09, GS06c, GÖ05, GB06a, HX07, HDG07, HED06d, Has07a, HBES09, Hos05b, Hos05c, HW07b, IR05, IH09, Ism08b, JRW08, JJ09a, KA05b, KR07b, KGB06, KAS06a, KBLT08, KMK07, KLX06, Kut05, KE05d]. Numerical [LZ08a, Lep07, LT08b, LC05c, LLW05, Liu06b, LW09c, LA07b, LA07a, MO05, MB06, MA07d, ML05, MKA05, MSK05, MK05c, MM05a, MA05d, MM06b, MAM06, MR06b, MA06c, MD06b, MY06b, MAR06, MLR07, MLM07, MS07b, MD07a, MSR07, MMA07, MS08a, MB07, MSR08, MK08a, MAK05, MN06, MN07, MMR09, Muk08b, Muk08a, NA06b, Omr07a, ÖEK05, Özt05, PW08a, PBPF08, PMRAS05a, PMRAS05b, PDA06, RMA07a, Ram07d, Ram07e, Rao05, Rao06b, RZ05, Ras05c, RNV07a, Raw06a, Raw06b, Rem07, RV08, SMY08, Sal06a, ST06b, Sha05a, SH07d, SG05c, SK05, SP06a, SA07a, Sid05, Sid06, dCRDSSN^{+}08, Sog08d, SA05b, TMS08, TF08, Tan05b, TFX07, Val07, WC05a, Wan05e, Wan05a, Wan06k, WiN06, Wan06f, Wan06c, Wan07f, WZC08, WT08, Wan08e, WL09c, XXC08, XL05, XZB06, Yan05c]. Numerical [You06b, YB06, YB05, (Ag06, mZzH05, ZL08c, ZCFX07, ZW09c, ZWC09, ZMV05, AD05a, AD05b, Abd06, ARK07, AK06c, AK06b, AK06e, ANM07a, AMN07a, AMN07b, ANM07d, ANM07c, AMN07d, ANM07b, AMN07c, ANM07e, AMN07e, ANM07h, ANM07j, ANM07i, AMJD06a, Ahm06, Aks05, AÖÖ06, Aks06a, AH07d, AAK07, AE07b, AD06c, AEA07, BJE05, BMJED06, BAF08, BFR05, BT05, BSN^{+}08, Baz08, BM05a, Ben07b, BÖ05, BOB05, BB09, Bog05d, BJ08, CVMM06, CL09a, CSV07, CPR07, CS05b, CZ05b, CS09c, CJSSP05, CD07, DSB05, DIS05, Dan06, Dao05, DE07a, DJ06b, DKK06, DMJE05a, DMJE05b, DMJE05c, DMJE06a, Dön06b, DS05b, EFE08, EPER07, EHN07, EAEG07, EG07b, ESS05c, ESK05, ESK06, EP07, EM07, Ese05b, EA08b, EDMJ05b, EDMJ05c, EDMJ05a, EMJB05, FW07a, FE07, FCG06, FJ06]. numerical [Fen05, FHZ08, GS06a, GJ07g, GS07a, GKK09, Gu07a, Gül05, bGgLgW09, GB05, GB06b, HA05, HE08a, HAK07a, HED06a, HEDMJ06, HMJED06, Has06c, Has06d, HPD07, HBVM07, HJ08, Hos05d, Hos06c, Inc05, Inc06b, Inc06c, Inc06a, Inc07e, Ism06, JP09, Jav06a, Jav06b, JG07a, JCM05, Joy06, dCJLNÁA08, KA06a, KPS06a, KR07a, KS08a, KK07c, KYA05, KS08b, KI05b, KES05, KAMR05, KE05a, KuI08, KSM09, Kia05, Kim06a, KS09a, Lar05, Lee05c, LWH06b, Liu06f, LG06c, LWL06, LW07, LGY07, LLY07, LHSB09, LX06b, MA08e, Mah05, Mal06, MMN06a, MNY07, Mar05b, MJHED05, MJED05, MJHED06, MP08a, MA06d, Meh06, Mom05c, Mom06, Muk07a, Muk07c, NLD08, OF06, OB09, OAHR09, ÖK05, PE07a, PRM06, Pao07, PR09a, Pat05b, Pat07, PKT06a, PKT06b, QS06, QFXW06, QFF06, RAS07, RLZ07]. numerical [Rao06a, RG05, RZ07a, RZ07b, RMJ07, RVN07, RNV07b, RNV07c, Ray06, RHY06, RY06, RHQ06, RZ07c, RMWk09, SD09, Sal07d, SG09a, Sea06, SGT06, SR06, SS08b, SDG05, SZ05, SFPE07, SDR07, Sim09b, uIK06, SMH07, Sog08a, Sog08c, Sun06a, Sun08, TSD07, TS06, TG08b, THLC07, UHuI09, Van05, Van07, WTW05, WY06, Wan07b, sWX07a, WL07f, WL06f, Wei09, XO07, XFQ06, XZB05, XL06b, XBZ06, YML07, YLS05, YLB06, Yan06b, Yan08c, Yan09b, Yig08, You08, YK06a, YKJ07, ZPT^{+}06, Zha05c, ZH07a, ZZDL07, ZQJ09, ZXL05, ZH05, Zot07a, ZHZ07b, GMCM06]. numerically [Bha08, Pre06a, Pre06b, SS06e, Wu09]. Numerov [Bie08a, BFKT07, Van07]. Numerov-type [Van07]. nutrient [JB05]. nutrients [SMC08]. NYSE [DD06a]. Nystrom [AKIS07, Fra09a, LSW06b].

O [Kop22, Cha07a]. Object

[Zot07b, Zot07c, ALSL08, HLCH08, LWL07b]. Object-oriented [Zot07b, Zot07c]. Objective [MD05, ASA06a, ASA05, ASAEE06, AD09a, Ash06, AK06i, ATM06, CK08, CCFC08, wG07a, Isl08, JHST05, JRR^{+}07, KG06, LYKC07, LW09b, MR06a, MFF09, NNN05, RSA07, Sha08a, WD08, WLZ06, YYWW07, ZFT06, Mav09]. objectives [KOP07a]. objects [DESG08, MS07e]. oblate [RAES05]. Oblique [JyGcL^{+}09, KS08d, SC09, XSG08]. obliquely [KS08c]. oblivious [ZW05d]. Oboudi [Bul09]. Obreshkov [GSG09]. observable [KC05b]. observation [GLV08, WLC05, WLC07c, WLC07d, d’O05c]. Observations [WFS08, Bog05b, GG05a, HCLP07a, HCLP07b, HCLPJL^{+}08, LLFM08, NCÁHC^{+}06, XFL06]. Observer [KK07e, TTG07, FTG07, ID08, KC05b]. observer-based [ID08]. obstacle [ASNR06, An09, BN09, CS09b, lCX06, GC06a, MJ07a, MMN06b, RJM07b, RJMG07, WC08a, YCL07, YCL09b]. obtain [ANM07e, AMN07e, ANM08, ML06, Vaj07a, Vaj07c, Vaj07b, Vaj07d]. obtained [EG07a, GK05b, Imr06, MB05a]. obtaining [KA06f, SA06h]. occlusion [BG06]. occupational [PB07a]. occurrences [Ben07b]. occurring [SSR05]. occurs [SK06b]. octahedron [Sil06f]. octonions [Köp06]. oculogram [MUZG09]. odd [LR07a, LR07b, Rim05b, Rim05c, Rim05d, Rim05f, Rim06f, Rim06h, Rim06e, Rim07c, Rim08c, Rim08e, Rim09b]. ODE [ÇT07a, SQ11, HM08b, NBHYV09, Ram05a, TW09, Waz06o, Waz06p, Waz06q, ZM06b]. ODEs [BBES07, Bar05b, Bie08a, Bie08b, IOS07, LC08b, Nav08a, Neh07b, OISM07, STKB05, ST07, TM05, Waz09f]. ODEs/DAEs [Bar05b]. Off [Shi05a, APS05, KMN09, RK07a, TT06b]. Off-line [Shi05a]. offsets [Sam06]. Ohm [Dem06]. Ohmic [EO06]. oil [AKMH06, EJER06, GKAS07, YZD08]. Oldroyd [FFKV08, VFF08a]. Oldroyd-B [FFKV08, VFF08a]. Olver [SQHY08, Waz07g]. omission [AS06f]. omniscience [Tan07b, Tan08b]. On-line [Gon07, Che05h]. One [GJD^{+}07, NBHYV09, PLJ05b, TLM09, Wu05b, AE07c, ASA07a, Bad05, BS05a, BT06b, CY05a, CC08c, CG07b, CL08c, Chu07g, Chu07h, CH07d, CG07c, Deh05b, DHG09, GC07a, GRH07, mH07a, HCLP07b, Inc05, JLS^{+}05b, Jan07a, Jan09c, JH08b, KA06b, Kay05b, KTW08, Kut05, LR07a, LR07b, LF08, LZ08d, LOZ07, Lia08, LLC09b, LST09, MM08c, MV09b, Moh05b, NKR^{+}08, NR09, OIL09, PP07c, QCDS09, Ram05c, Ram05b, Ram06b, Ram07c, RM05, Ras05c, Rim05g, Rim05a, Rim05h, Rim05b, Rim05e, Rim05c, Rim05d, Rim05f, Rim06c, Rim06g, Rim06f, Rim06d, Rim06h, Rim06a, Rim06b, Rim06e, Rim07c, Rim07a, Rim07b, Rim08a, Rim08c, Rim08d, Rim08e, Rim09a, Sal06a, Sal07d, lSgPpJ07, Sha07d, Shi09a, SZ05, SFPE07, SW05b, SQG08, TH07b, TG08a, WSZ07c, WL07f, WL07c, WQL09, WL06f, WLC05, XB06, Xu07, YW05, Yan09d, rYzS07, YRY05c, Zha05a]. one [ZW05a, ZG08a, ZZKY05, ZW06f, lS07b, dSYY08]. One-dimensional [GJD^{+}07, ASA07a, BS05a, CC08c, CG07c, Deh05b, DHG09, mH07a, Inc05, Jan09c, Kay05b, Kut05, LZ08d, Lia08, LST09, MM08c, MV09b, NKR^{+}08, OIL09, Ram05c, Ram05b, Ram06b, Ram07c, RM05, Sal07d, Shi09a, SZ05, SFPE07, SQG08, TH07b, WL06f, Xu07, rYzS07, lS07b]. one-leg [TG08a, WL07c, WQL09]. one-model [JLS^{+}05b]. one-parameter [Chu07g, Chu07h, CH07d, JH08b, QCDS09, lSgPpJ07, Sha07d, WLC05, XB06]. one-phase [BT06b]. one-predator [LOZ07]. one-prey [Yan09d]. one-space-dimensional [GC07a, LLC09b]. One-step [NBHYV09, TLM09, HCLP07b, Sal06a, ZG08a]. one-time [CY05a]. One-to-one [Wu05b]. one-way [CL08c, YW05, YRY05c, ZW05a]. Online [CYL^{+}08, LJL06, ZSCW09, Zha09d]. onto [AS07b, Men08, Mor05]. Open [Pod06, DMJE05c, DMJE06a, dCDNPN09, HEDMJ06, HMJED06, MJHED06, Rao05, Rao06b, VKVHN05, WW08a]. operation [ZXL09]. Operational [DSZ07, Sil07d, Sil07e, Sil07f, BF07a, BF07b, DX07a, DMK07, KZ07, MS08a, Mir08, Wu09]. operations [CS07c, Cha05f, KP07]. operator [AJMR07, Aou08b, AEA07, Ate06, BM06b, Bul09, CjC08, CC05d, CY05c, DT07a, DT07b, Dzi07, ETETH07, Gül08, HL08d, Ibr07, JGY07, yLcCsL09, Li09b, LS09f, LS09g, LL05g, LN07, LP08b, LYH09, Mar05c, MA05g, Moh05a, MA09, MP08c, NMM09, Pat05b, PS06c, RSSS07, Rem07, SS07b, SK06b, Sil07a, Sok08c, SYX08, ST09, Ste08, Su08c, TA06, Wan05f, cWmZ06, WZ08d, XHH06, Xu07, YL08a, Yan09c, ZLH^{+}09, ZZZ09, ZH08d, ZH09b]. operators [AC07b, AM09a, Arg09, AS09b, Bel05, CMGMD06, CL08c, CKOS07, Den08a, DNS08, Deo08, DR06, DZ08d, Djo06, DM08c, Dod08, EDS06, Fro06, GS08c, Gup08a, GBG08, GI08, Gup08b, GF09, GAG09, GLMLP09, HR07a, HR09a, HBN06, KG08, KTW08, yLcCsL09, LS08d, LS09d, LS09e, LL06b, LLX07b, LZ09e, MA08e, Mou08, Nač09, NJG08, Noo06a, Noo07a, Odi07, PDGR07, Rad09, Ren06, Siv07, SFG07, SW09c, SiU09, Ste09e, Ste09d, Ste09g, Ste09h, Ste09i, SCS06, WG08, WLS08, WN08b, Yu08a, Zha09h, Zhu09b]. opinion [HZYL08]. Optical [BM09, WN09a, KS05a, KBMZ08, LHSB09]. optima [BEvdB07]. Optimal [AAA09, BS05b, Bel07b, Bel08a, BS09b, BS08, CZ07b, CS08b, CH09, Chi07g, DMDB08, EG05, GC06b, HL06a, HC05b, JOA05, KNM05, KRG06, KB07, Kwo07a, LML08, LK08, LY09b, LGw07, Luo06, Luo07, LG08, Mis05, MBM07, MH08b, NRLT07, OY05, PS06a, RK07b, RY05b, SH07a, SLC07, SG05d, TMMVA05, WLT05a, WWT05, YTLH06, Yua08, ZWZ05, AMN07g, AEFFA08, AK05h, AK08b, BK07a, BW05a, BW07, Bel07a, Bel08b, BS07c, BKFM05, Cai06a, CPSS08a, Çet06, CMMC09, CC06c, CHW07, Chi06a, Chi06b, CJK07, CP05, CH06c, Dah07, DKA06a, DH08, EJER06, EGE06, FZN07, GBK06, Gal08b, GL06b, HY07c, KK07b, KKF07, KF08a, KH05b, KJ08, Kia05, KT08b, KS09a, LWL07a, LAL07, Li08e, LHZ09, LBS05, Lin06a, MP08b, PBG07, PHZ05, PHZ06b, RSDC08, RO07, SSB06, Sal06a, SYT06, Sha08a, SLO08, SQ07, TTG07, WZ06a]. optimal [Wan07f, WZ09c, WGCH05, WZ06k, WLT07b, XW08, XXF06, YCL07, YCL09b, YGY09, ZW09b, dCdSJM08]. optimality [Çak08, JC07, Kim05a, LW09b, Sla06, WX08b, YCL07]. optimisation [PLTW07, PE07b]. Optimization [BDPP08, CZ08a, Chi08c, DOU08a, DT06, ES05a, KMN09, QW06, WS05b, WS07f, AEA05b, AE06b, AH06b, ASPA09, AK08a, AKB09a, AKB09b, AD09a, AP06, And09, Ash06, AGI08, BHCC09, BGRS07, BNdFPS06, BMP06, BS07c, BEvdB07, BP06b, BS05c, CK08, Cha05c, Che06m, CXZL09, Chi07h, Chi08a, CMR08, CCZS08, Dan06, DD08a, DD08b, DSKM09, DN07c, DS06b, DMZ08, EJ07a, EKS08, Fab08, FAM07, FYZ07, FS05, GRZ06, GH07, GG05b, Ham05c, Ham08, HYG08, HL08a, HW07a, HP07, HZW06, HM06, zHWH07, HMW07, JM07, JCSR07, JDSR08, JjQjWmL06, JJT06, JZQ07, Jia05, JZHT05, JZTH06, JXH06, bJyKxC09, JHHW07, JGS06, JS07b, JC08b, JZHW08, JZX07b, JZX07a, Jin08, KM06, Kah05c, Kah05b, Kah06a, Kah06b, Kah07, KC07a, KF08b, KWdW08, Kas06c, Kas09, KG06, KGM06, KR07c, KY06, KPG07, LT08a, Lee06b, Lee06a, LMT^{+}09, LKW06, lLWhL06]. optimization [LZ06c, LY07c, LW09b, Li09e, LGJ06, LJG06, LZ06e, LZ07d, LN08b, LY08d, MR06a, MFD07, MBDD08, XSX08, MBC08, MP08b, MPT08, MSD06d, MQW05, MZW05, NJG08, Nwo06a, ONG08, OASM06, Oss05a, OMR07b, PWL06, PC07a, Par05c, Par05a, PKL08b, PKL08a, PN06, hPY06, PL06c, PFL09, Ql08, QYYL08, QZJ07a, QZJ07b, RSDC08, RW08, Ren08b, Sad07, SKHZ06, SRMRGRL08, SK08c, lSgPpJ07, SHK06, SSJK07, She05a, SW06a, SG05d, SYY04, Sub07, Sya05b, TLC05, TH07b, TS08b, Tok06, THLC07, Tso08, Tso09, Wan05b, WL05e, Wan05d, WTW05, jW06a, WZ06f, WZ06e, WSZ07c, WZ07b, WD08, WS08a, WN08a, WZWW08, WZ09c, WB09, WZ05, WLQ06a, WLQ06b, zWxY07, WLY08, WS07g, WS06b, WZ08f, Wu08d, WLL09, WWC09, XLwW07, XW07a, XYZ09, nXwCpC06, XCS07, YSL05, YS06b, YDL07, YC08a, YE07, YW06a, YW06b, Yin07, YYWW07, YZW07b, YHL08b, YHL09a, YK06a, YKJ07, ZLZG06]. optimization [lZ06d, ZZLL07, ZLH^{+}09, Zha09b, Zho09a, ZFL^{+}09, Zhu05a, Zhu05e, Zhu06c, Zhu06b, Zhu07, ZCZX07, MB06]. optimization-based [Lee06a]. optimizations [JZH07]. optimize [PN08, SMB08]. Optimized [VGMV^{+}07, LS09b, SM07a]. optimizer [HM06, KCW07, NZHW07, XwWL07, YYYM07]. Optimizing [HJB06, JW09a, SA08b, Alr05, GK06b]. optimum [BTPGAl06, LKP06]. Option [ZW05f, Hor07, IR08b, Mar08, SS06e, SYL07a, SYL07b, YC08b]. optional [MAR05a]. options [AS07a, CWY06, CL07c, CK06d, DHL05, DL09, Hor08, HL07b, KF08c, Nwo06b, TYCL06, Wu07a, You09, ZW08a]. orbit [AA05, Dan08, ES05a, QC07, Wen05]. orbital [AAES07, FW07a]. orbits [ABP05, CJ06, ESS05a, ESETA^{+}06, FZ09, SH05a]. order [AJKM09, AO08, AO05a, AD06a, ACP07, AS06a, AS06b, AKIS07, AS05, ASNR06, ASN07, Alo06a, ABP05, AHR08, An05b, ALL07b, An09, AR07a, AM09b, AV07b, AE07b, Arg09, AY05b, AA07f, AÖ09b, AFS06, AL06, Att07, ADDdM07, AZ09, AR07b, AKR05, BD06a, BR07a, BCEGMRMP07, Bas08, BŞ07a, BBB08, BRW08a, BWR09, BTI06, Bie08a, Bie08b, BM09, BS07d, BGF09, BG07b, BAM07, BT08, CC07a, CC08b, ÇT07a, CD05, CPR07, Che06j, CS07h, CL08d, CZF09, CT09a, CAS08, Chu07a, Chu07b, Chu07c, Chu07d, Chu07h, CH07d, Chu07i, CH07e, Chu07k, Chu07m, Chu07n, Chu08, CH08a, CH08b, CN09a, CN09b, CG05b, CT07c, CG08b, DGJ07, DGB08, DB07a, DB07c, DMS06a, DMS06b, DDR06a, DM06, DMA07a, DMA07b, DBC^{+}07, Den09b, DMH05, DXH07b, DMDB08, DM08c, DDR05a, DDR06c, DDR06b, DB07d, Du07]. order [EM06, Eid05, EG07b, EM08, EOM03, EOM08, ESS05a, ESETA^{+}06, ESAH10, Elg09, EBP09, EG07c, EM07, EHR09, FZ09, FHH09, Fen05, FH06a, FH07, FS09, Fra09a, GA08, GC06a, GZ09a, GZ09b, GZ07b, GC07b, GL08a, Gen09a, Gen09b, GKR06, GF06, GJ07b, GJ07c, GJ07h, G07, GAPT08, GSDB07b, GSDB09, GS07b, Gül08, GS05, GSG06, GRH07, GWJ07, GC07c, HC07a, HSH05, HY05, mH06c, mH07a, HC08a, HH08a, HR07a, HR07b, HR09a, HJ08, HN07a, HLW08, HHK^{+}05, HM08b, Hub08a, Hub08b, Hub09c, HMT09a, HMT09b, HLT09, IOS07, IS09b, IMB08, Ism08a, Jan08c, Jan08b, JS07a, JWPS08, JGZ07, JZCW05, JH08b, JLW09, KW08b, KWX08, Kan06b, Kar09c, KMÖS09, Kas09, KH07a, KAT06, KEM08, KM05c, Kıy06, KSS07, KLW06b, KLW07a, Kou07c, KLW07b, KLW07c, KL07a, KLW07d, Kou07a, KL07d, KL07e, Kou07e, KW07, KL07g]. order [Kou07f, KLW07e, KNJ05, Kum03, Kum07, Kum09c, LY07b, LWL07a, Lei09, LR07a, LR07b, Li05b, LW05b, LW06a, LLZ06, LS06b, LX07a, Li08a, LCJ08, Li08b, LC09a, LS07c, LPS08, LZ08e, Lia08, LB06, LDL08, LLAT09, LXML09, LCL05, Liu06h, LKU07, LD08, LL08f, LLW08, LKU09, Liu09a, LLW09, LWKU09, LA07b, LA07a, LW09e, LM07b, LX06a, Ma05c, MTL07, MM08a, Mah09a, MPG07, MB06, MKKS05b, MH06b, MH06a, MSTY06, Mar08, MŘ07b, MN05, MX07a, MX07b, MM07e, MD09a, Meš07, MR07c, MA05g, Moh05b, MK06, MS06b, MA06f, Moh07c, MD09b, MMN06b, MN06, Mom06, MN07, Mus07, NKE05, NCÁHC^{+}06, Nat09, Nav08a, Neh07b, Net08a, NBHYV09, NLD08, NA06a, NN07c, NN07a, NNuH07, NMD07, NAUH09, OME08, Özk07, PS08, PDW06, PB09, PG08a, Par09a, Pat05b, Pat07, PC09, PC08, PR05, PMP08, PSD07, PDA06, QFXW06, QFF06]. order [RLZ07, RLZ08, Ram07d, Ram07e, Ram08d, RG05, RJM07a, RJF07, RW07, Ren08b, RWB09, Rim05a, Rim05h, Rim05b, Rim05e, Rim05c, Rim05d, Rim05f, Rim06c, Rim06g, Rim06f, Rim06d, Rim06h, Rim06a, Rim06b, Rim06e, Rim07c, Rim07a, Rim07b, Rim08a, Rim08b, Rim08c, Rim08d, Rim08e, Rim09b, Rim09a, SF07, SH07b, SMA06, SA06a, Sal07b, SAB06, SS07a, SS09c, SG09a, SH06a, SG05b, SQ11, SR06, Sha05e, SG07b, Sha07c, Sha07d, SGS09, SL07a, SYH07, SXL09a, SM06c, SW09b, SW09a, SP09, SRH09, SH06b, SA06f, SA06g, SAM07, SA07a, SAN07, SA07b, SA07d, SA07c, SA07e, SAE08, SA08e, SS08c, Sim09b, uIKTT05, uITA06, uIK06, uIT06, uINTK06, uITiHK08, SFH09, SdHT09, SI06, SLW08, SEŞ09, SL05b, SWZL07, Su08c, SZ08b, Sun05b, SS05g, SS06f, Sun06b, SS07e, SM07c, SG07c, SM08b, SM08c]. order [SM09c, SCO09, SA05b, TXW09, TJG08, TNL09, WW05a, Wan06g, WZZ06a, WZZ06b, WZSZ06, WSB06, WHL06, WC07a, WZS07, WML07, WL07f, WC07b, WLT07a, WLK08, WTW08, WFS08, WKL08, WB09, WL09b, Waz05e, Waz06b, Waz06a, Waz06e, Waz06n, Waz07b, Waz08b, Waz08d, Waz08k, Waz08n, Waz08o, Waz09b, WL06f, Wei06b, WZ08e, WZ08d, WS09b, WS08b, WS09c, WRL06, WH06b, WZM06, Wu06f, WZ06j, WR07b, WX08b, XCFX05, XZ08a, XLL06, XHH06, XM06, XM07a, XM07b, Xu08b, XZY08, XZ09b, XT08, YSS09, YLZ06, YLB05, YLB06, YS07, Yan08e, lY08b, YW08, YZ09b, YX09, YL06c, Yin08a, Yin08b, You08, YZ07d, YJO08, ZGK06, ZW09b, ZZ05a, ZPH06, ZLW06c, ZLW07, ZW07b, ZZ07a, ZFY07, ZLZ07, ZL07a, Zha08f, ZS08a, ZL08b, ZFYH08, ZYW08, Zha08g, ZLW08, ZM09, ZS09a, Zha09c, ZZC06, ZM06b, ZC08b, ZM08c, Zha09g, ZWZZ09, ZLC07, Zho08b, ZZ08b, ZX08, Zho09c, ZL09a]. order [Zho09b, ZYZ06, ZW05j, ZJO08, lS08b, Ant05, WLW09]. order- [KAT06]. ordering [LWL07a, Sal08d, Sha08a, YVU06]. orders [EMESES05, LS08e, Oss05a, PN08, Tan06c]. ordinal [Sae06d, Sae06e]. ordinary [AES08, ADDdM07, BR07a, BG07a, DBC^{+}07, ETETH05, ETETH07, HBVM07, Hos05c, Hos06c, HN07a, Inc07e, Inc07b, JDG06c, KM05c, Kıy06, Lar05, LZ08e, MCAH05, MS07a, Mus07, Pod06, PCS07, Ram05d, Ram05f, Ram05g, Ram06d, Ram08c, Ram08f, Sai07, SR06, WL06b, ZCMQ07]. ordinate [ZMV05]. ordination [Ser09]. Oreochromis [MD07b]. organ [Çet07]. organized [Bog05c]. organizing [DD08b, ZBH06]. orientation [Su07a, Su08b, WHH07]. oriented [Alt07a, CHC05, GT05b, GBG07, KB05c, Küç05a, LWZH05, QC05, WY05a, Zot07b, Zot07c]. originating [ASL07]. oropharyngeal [SGG^{+}09]. Orthogonal [AV07a, DN06, Lin05c, AMJD06b, ADK05, ADK06a, ADK07, BMS07, BMS08, BT09a, BR08b, BB06e, Boy08a, Boy08b, CC09b, CG07a, CMS05, GCJ08, HT06a, HZ05, KWS05, KCW07, Lin05b, LFW08, MJ07c, MCS08, SLLL07, Tia09, WZ08c, cWxChY07]. orthogonal-array-based [KCW07]. Orthogonality [LO08]. Orthomin [AZ08]. orthonormal [LXSH06]. Orthotropic [Sel06, Li05f, Li05g, LD06a, SA06d, YWJ07]. Oscillating [YZ06, Ahm07a, ARM09, mH07a, KS08f, Ram06g]. Oscillation [AOS08, BGF09, ÇT07a, CD05, DDR06b, EHP08, EBP09, HM08b, KMÖS09, LC09a, LGY07, LX09b, LM07b, MSTY06, MX07b, MM07e, Sak05, SQ11, Sun05b, SM05d, SS05g, SM06e, Sun06b, SM07c, SM09c, WS08b, WZM06, XM07b, YSZG05, YX09, ZM06b, ZM08c, AD06a, EB05e, EM09, Gar07, GAPT08, LHXL05, Lei09, LLZ06, LMZ09, LH07a, LM08c, LW09e, MX07a, SS06f, SS07e, SM08b, SM08c, WML07, XZ08a, XM06, XM07a, Xu08b, ZZ05b].

Oscillations [ZFY07, ZFYH08, CH07b, MP08a]. oscillator [BV06, KV06, XXC05, XXC06, XHSL06, YXSX05]. oscillators [Bob08, Ram08a, Ram08b, Van07]. Oscillatory [SS05h, ZZ08b, BS06, CX09, Dai09, DD06b, GD08, GAR09, LWWS09, LC09a, LP07c, MK08a, Mus07, Mus08, SE07, WX09, XG08, lS08b]. oscillatory/layer [LP07c]. osculating [KSGM06]. Oseen [JL08]. Oseen-viscoelastic [JL08]. osimotic [WGCH05]. osmotic [ZY09]. OSPA [YRY05b]. Ostrovsky [Boy06b, YB07]. Ostrowski [CH07e, FPP06, GDB06, LN08a, SG07b]. OSV [YKKY09]. other [II07, Ram08e]. out-of-orbit [ES05a]. outage [LL06a]. outbreak [ZLMW05]. outcome [GXY06]. outcome-space [GXY06]. outer [DQW06, ILZ09, MD09c, SCI08]. Outflow [ORG^{+}09, AS06h, AS07c, MH08a]. outflux [ZM07]. outliers [HW05c, WL06g]. Output [LP07b, ZC05b, DDZ08, gDH07, EKM08, LXZ09, Muk08a, Par05c, WH08a, ZHX06, ZFT06, lS08a]. outputs [HV08, JVFM04, JLS^{+}05d, Kho07]. over-relaxation [Haj06]. overall [Ali05, TARm08]. Overholt [Thu07a]. overland [KAMR05]. overlap [JLBR07]. overlapped [CPTZ05]. Overlapping [DG07a, DS06a, IAMA08, MY07b]. overrelaxation [DHY06, DH06b, DKA06b, PW08b, RJ07, WZ06d, ZW05h, ZW06d, ZW06b, AC07a]. overrelaxation-type [PW08b]. overview [Kos07, Nab09]. OWA [PDGR07]. oxygen [Ahm06, AT07b, BG06]. Ozone [BTI06].

P [dCD06a, Aou08b, DJ06a, EGX05, Ola06, dCD03, dCDB07]. p-valent [Aou08b]. p/hp [EGX05]. P53 [QSS08]. P5P [TCW08]. package [qXbL05, YL06b, ZZL05]. packed [QS09]. packed-bed [QS09]. Packing [SS09b]. Padé [YCL09a, DHS07, Ism05, IRSR05, KB06c, MM06a, Mak07c, SG09b, Waz06h, Waz06k, XHH06, You08]. PageRank [SW06b]. pain [PB07a, XWSL08]. Painlevé [qXbL05]. Pair [GR07a, AKIS07, NR07a, NRA07, SC07a, SMQ06, WCL08, WZDgH09, XSG08, Zhi09]. pairing [LHC07, PC05]. pairings [CLS05b, CJL05b, DL05, LC05a, LWZH05, LC05e, SW05b]. pairs [LYGL07b, Sha07a, dCRDSSN^{+}08]. pairwise [HT06b, Utk06]. Palmprint [JLCH08]. panconnectivity [CCWH09, TTH08]. pancyclicity [CCWH09, XHX07]. panel [Jön05]. Panpositionable [TTH08]. pantograph [LL05d, RMWk09, ZXWL06]. paper [Bra09, ÇT07a, HLSS10, JS07b, Kop22, Mar14, Mir06b, RJM07a, Ren08a, SQ11, SM06d, Sun07b, Uje09, Zim07]. papers [KPL09a, MK08b, MK10]. parabolic [AV07a, Ang06, AKM06, AY05b, AKR05, Bar05a, BDLS05, BM09, BS08, Bou09a, CC08b, CC06b, CG08a, CY08d, Dao05, DS05a, DG07b, EFE08, EPER07, EJER06, EB08, FE07, Fat05a, Fat05b, FCG06, FXLZ05, GJ07g, GY07a, Gu07a, Gui05, GR06b, Ism05, IMS09, Jan07a, Jan07b, JV05, JM06b, JM06c, Jun09a, KA06b, KBH06a, KBH06b, Kar05a, KWZ08, KM05f, Lee06a, LCM08, LW09d, MMW09, Moh07a, Moh09, Mor08a, MD08c, PY08a, Qia09, Ram05c, Ram05d, Ram05h, SS05a, Sal07d, SYH07, SFPE07, SB05b, Tad08, Vul07, Waz07i, WG07b, Xia06, XCM07, XC07, XF08a, YY05a, Yan07c, Yan08e, YJ09, rYzS07, You09, ZY08b, Zha05d, ZM07, ZM08d, ZW05i]. parabolic-elliptic [SS05a]. paradox [Tan06b]. Parallel [AC09, BMP06, KW05b, KW05a, Kay05d, KK05d, MAT08, MY07b, PGdSS08, VV06, YK06b, ZXL09, AH06d, BT06a, BB09, CL06d, DG07b, DvdHD08, DJ07, GEG07, GWX06, xGyZtZ^{+}07, HWH09, HL07a, JGTK07, KK05b, KNJ05, KK05c, LPA^{+}09, LZX08, LWL06, LGHS06, LCH09, LXZ07, cMpZfR06, MQ07, MSR08, MMC06, OT06, PSW07, SKT06, SH08b, SYH07, SMH07, Son05, TD07b, TD07a, WHC08, XZ09a, XQ06b, Yua05, YGY09, ZO05, Zhu05c, Zhu05d, ZYZ06]. parallel-plate [ZO05]. parallelizable [WYZL08]. Parameter [DDZ08, Fab08, KM06, KSSA07, YZLD08, ZFL^{+}09, dlS06a, AST08, ANK05, Bak05, Bar05a, BŞ07a, BP06b, CC06c, CG08a, Chu07g, Chu07h, CH07d, Chu07n, Dah07, DS05a, DKA06a, Dem05b, EGE06, EHL08, FP09b, GL05b, GJD^{+}07, Gon07, GGSNR06, HBVM07, JH08b, KSA05, KA06b, KY08, LP07b, lLWhL06, LML08, Li08e, LX07b, LN06, Muk07a, NGLHCLP06, Pat07, PP07c, QBH06a, QCDS09, Raj08, Ram08b, RSDC08, lSgPpJ07, Sha07d, WTW05, WSZ07c, WFX07, WWT05, WLC05, WY05c, WLY06, WLC06, WWC06, WLC07c, WLC07d, WLL07a, XB06, YZ09a, ZBH06]. parameter-dependent [ANK05, LP07b]. parameter-uniform [KSA05]. parameterization [Mor09]. parameterized [AKD06, CJ08, Fro06, JSLL07, ZZ09]. Parameters [DZ08b, GFWX05, AJAH06, ADTGV09, ASAEE06, AÖ05b, AM09b, BKA07, Bog05c, BKZ06, Che06k, CAL08, CA08b, CCZ08, DW09d, EGS05, ESK07, FAQ08, HED06b, HLK06, JP09, JL09, KHL06, KAP06, KTÖA08, KA06f, LY08a, LK08, MR06c, Mug05, PGdSS08, SW09b, Slo07, WLA09, WWLX06, WZ06g, WLZ06, YZD08, ZW07b, ZWLX08, Zhu09a]. Parametric [FH06a, HLK06, KHL06, RMG07a, AA06a, ASAS06, CE07a, DW09c, EH06, EJ07a, KKKM09, KKA05, KV06, ML08a, PK05d, PLZ05, QZJ07a, XXTL09, YE07]. parametrical [lS08b]. parametrically [YXSX05]. parametrization [LZ07d]. Parametrized [DOU08b, Śmi07]. parenchyma [SSTD05b]. parent [MR05b, YKJ07]. Parental [EMH07]. Pareto [GN07, HWC07, Sla06, WLC06, WLC07b, WLC07d]. parity [Boy08b]. Park [Sun07b]. Parseval [BMY08, DSY08]. Part [AM08b, AM08c, AM08d, He06, Kay05e, KSKK06, Kös06a, Lie06, MH06b, MH06a, SJ08a, SJ08b, Sid05, Sid06]. Partial [WS08c, AMN06, AE07b, AKM06, AKR05, BDG08, BKBK06, BSKS07, CC08b, CS05b, CW07, Deh05a, DT08, EB08, EHZES05, Jav06c, JTMRS06, JAT09, Kum09a, LZ08c, MM07a, Ma07a, MSR08, Moh07a, Moh07b, MAD07, NMM08, NMS07, Ram07d, Ram07e, SS08b, SZL^{+}07, WL07a, WML07, WN08c, Wu09, yXhWZ06, qXbL05, YLB06, YL06c, YC06d, ZG09, ZZL05, ZT05, ZT06, Zha06b, Zha07g, ZZYY08, Zha09e]. partially [KB08, PBG07, WLHH05, WY05d]. partials [Dan06]. participation [CPSS08a]. Particle [CCZS08, NJG08, PWL06, QYYL08, SSJK07, AK08a, AH06d, BEvdB07, HL08a, HW07a, HM06, JCSR07, JDSR08, JHHW07, JZHW08, KWdW08, KCW07, Lee06b, LYKC07, LGJ06, LJG06, LY08d, MBDD08, XSX08, MA06d, NZHW07, Ohm08, ONG08, SKHZ06, WWC09, XLwW07, XwWL07, YYYM07, YE07, YW06b, Yin07, YYWW07, ZZLL07, ZZX05, ZFL^{+}09]. particles [HS09b, MR09, SDS05, SSK06, SSS06]. particular [AL07]. particulate [MA06d]. parties [LKY05b, OHJ07]. partition [Ali05, Cha07e, GY06, LBS05, SC08b, YGY09]. partitional [JCSR07]. partitioned [CI09, DjT05b, LSW06a, MKS09]. Partitioning [LZ06c, Che07h, HHK^{+}05, KBTS07, LZ07a, LY07c, OISM07, SB06a, TSP07]. partly [Zhu09a]. Partner [ZLZ06]. parts [HZ06b]. Pascal [WL06c, WJ06]. pass [Afr05, CC09b, Den09c, YZ09b, Fig08]. Passenger [GTA06]. password [KJY05, KHKL05, Shi05a, YWC05, YY05d]. password-based [KHKL05]. password-guessing [Shi05a]. past [AEA05a, Ari07, EEEE05, EO06, KBKV09, Pan08b, Pan12]. patchy [WZ06b, ZT07a]. patchy-environment [ZT07a]. patent [CS08b]. Path [MLP07, XXC05, Yan09a, ARK07, Aza06, BBMN07, BKFM05, Che07a, CK06c, EJ07b, FY08, Ji05, JZ08b, Li05a, LSBL07, Lin06a, Moa06, Oht08, PSS07, RKS08, She06, WXLH06, WZ09c, XXC06, XXC08, XLH^{+}07, YTLH06, ZFN07, Zhu07]. path-following [ARK07]. pathogen [MA05e]. pathological [SDM^{+}09]. paths [Imr09, JLS^{+}05a, Li06c, LP08c, jW06a, XLKH09]. patient [GPR^{+}09]. Patrick [SGS09]. Pattern [PBKM09, FYZ07, LZ06e, LYGL07b, RNRRMRCH09, TH07b, WWXX07, WS06b, WF07c, Wu08d]. patterns [Dil05, GL06a, HH07a, JY07b, JWL^{+}07, MD07b, PWY08, Sha06a, Waz06l, ZLMW05]. payment [CO06]. payoff [KT08b]. pays [CS07d]. PC [LWL06]. PCA [EARS07]. PCALDA [QZ09]. PCNN [LMD^{+}08]. PCR [GWX06, SWK^{+}06]. PD [GZG09]. PDE [Gui05, MZ09, Prá08a, Prá08b, Sun06a, jZzFbL06]. PDEPtest [qXbL05]. PDEs [EJER06, FTT^{+}09, GS06a, HK09, MA08e, MAT08, NLD08, Psi07, Sjö07, Wan08c, Wan09a]. peak [HMW07]. peaked [Den08c]. Peakons [Waz06l]. pear [GRF07]. Pearson [Nad06a]. Peclet [LJ07]. peer [JWPS08]. Peinado [YRY05a]. Pekar [CG07b]. Pell [KAT06]. Penalty [LL08e, WGCH05, hB06, hBxY07, Che06m, Dai07, DC08, Fab08, HZS07, LS09a, LL06c, LHW07, LHWW07, MSD06d, RAS08, SAB08, SY08b, XXX06]. penalty-free-type [Che06m]. pencil [AN05a]. penetration [Cha05g, FJCW05]. Pennes [ZZKY05]. Penrose [HZ06c, PH06, TSP07, Tia09, fWZpXzL09, ZCW06]. penta [BK09, Kar06, LL08g, NA08]. penta-diagonal [BK09, Kar06, LL08g, NA08]. pentadiagonal [EM08, EMR09, HE08b, HE08a, KK05b, KEM08, KT07b, LHL08, Rim08d, Rim08e, Rim09a, Sog08a, Sog08b, Sog08c, ZH08a]. peptides [SCR08a, SCR08b]. perceptron [HT06b]. perfect [LXSH06]. perfectly [EB05c, Li05f, Li05g]. perfectly-bonded [Li05f, Li05g]. Performance [Bar05b, CHL06, JY06, LWL06, Zho05, Zot07d, AK05e, AGA^{+}07, BHCC09, Cao06, EARS07, GG08a, GCJ08, GEG07, Ham08, HWC07, JC07, KA05c, Kos06b, Lin07f, MO05, SdJA06, WLL07a]. performing [WLC07a, YLM07]. perimeter [Imr06]. period [GHX08, Kun08, MJ07e, zWxY07, YML07, YWZW07, ZzLZ08]. Periodic [BI08a, Che05c, Che06d, DMH05, HC08a, LS06c, Li08b, Li08c, LS07c, Lin07a, MTL07, ML08b, SL07a, TQC09, WS07a, XCH05a, XCD05b, XCH05b, YZ07b, ZSL05, ZY07, AOS08, AA06c, AA08, ABP05, AAA08, AAA09, BW07, Bai08, BZ09, BBB08, BT08, Boy05, CCS05, ÇG05a, Che05f, CW05a, Che05b, Che05d, Che06a, CS06a, CXC06, Che06i, Che06c, CF06b, CY06a, CZF09, CT09a, CJ06, DXH07b, EM05a, EMR09, EOM03, EOM08, EWAH08, Fen06, FC06, FK06a, GL09, GC07b, GZC08, G07, HW08b, mHzC05, HL05b, HL07d, Hus09, Inc06b, JW07, JOA05, JSL06b, Jia09, JC06, Kar06, KY05a, KeTL08, Lee05c, LW05a, LL05f, Li05c, Li07d, Li09c, Li09g, LC09b, Lis06, Liu06h, LTC07, LFC08, LTCC08, Liu09a, hLqSlZ09, LC08a, MM08a, ML09, MN05, qMqZ05, MXC07, Nav08a, PS08, QD08, RKOK09, Ram08b, SA06a, SQHY08, She07b, Shi09a, Sog08b, SL06a, SM08a]. periodic [SS05f, SS05h, SS06g, SC07d, SHD07, Tan05b, TJG08, WC06c, WL07d, WL07e, WS07e, WS07b, WL08a, WZ08b, WWD09, WLA09, WDLM10, Waz05g, Waz05i, Waz06n, Waz08w, WW05b, WRL06, WC09c, XWZ06, XL08a, XZ07b, XCL09, XT08, Yan05c, YL06a, YCZ06, YC06b, YLH09, ZZ08a, ZS07b, ZT07a, ZCZ07, ZS08a, ZLY08, ZYC09, ZLL06b, Zuo09, BO06]. periodic-like [qMqZ05]. periodical [MQ07]. Periodically [d’O05b, Toh05]. periodicities [VG07]. Periodicity [Che08d, ZZ05a, ZL08a, YLL^{+}05, ZH08b]. periods [WCZ09]. peristalsis [NMS06]. Peristaltic [AH07c, EESESD05, EEEE06, HAA07, SG09c, VSB05, HAAS06, HA07, ICM08, SK09a]. Permanence [Che06h, CX06, CXC06, Che06e, Che06f, Che06g, Che06i, Che07c, Che07b, CY07b, FK06b, Hua08a, KeTL08, LZC07b, LOZ07, hMzLfW^{+}08, hMfWzLpZ09, She07a, XW07b, ZT07a, ZCZ07, JWS^{+}09, LZC07a, XZ07c]. permanents [LHB06, LSBL07]. Permeability [ORG^{+}09, MH08a]. permeable [AK05i, KBKV09, ÖŞ06, Pan06]. permissible [CO06]. permittivity [Fan06]. Permutation [WSJW06, WTG07, LGJ06, XCS07]. perpendicular [Sam06, VFF08b]. Perron [BM06b, DZ06]. Persistence [FK07, LHF07, WW05b, Yan06c, YC06b, Den08b]. person [xGZZ08, Su08a, ZCZ05]. persymmetric [GG08e]. PERT [APS05, ATM06]. Perturbation [KBMZ08, XSG08, Abb06a, Abb06b, Abb06c, Abb06d, AKD06, AC08, AR07a, AR07b, BZZ08a, CZH08, CPR07, CA08b, CCZ08, CX09, Dem05d, DC06a, EL05b, GKB07a, GKB07c, GkB07b, GKD07, GJ07b, GJ07c, GJ07h, HEZ07, HEZ08, HDG07, Hub09d, JG07a, KP06a, KK07a, LWW05, LmZ08c, LXJ09, Lia05, LX07b, LP06a, LP09, MDZ08, Muk07a, Odi07, PY07, PB07b, Pat05b, Pat07, QW05, RGD07, RDGP07, SZ08a, VCRCG09, Wan05h, Wan07d, Wan08d, WSL09, WCX05, XZ09a, Yan08c, ZCZ05, ZXLZ07]. perturbations [AST08, JMC07, JJ09b, KPL08b, xLyW06, LmZ08b, tLlWpL08, LZ09b, LLX05, LMC08, LP06b, qMpXjH09, NMP08, Par05b, ZZH06, ZYC09, ZXZL07]. perturbative [Soh06]. Perturbed [AGJ07, Ahm09b, Ami05, AŞ06e, AAK07, BR07a, Baw05, Bie08a, Bie08b, BP06a, Bra06, CA05, CGVC07, Cen05, CPY09, CG05b, CG08b, Dai09, Den09c, DZ08c, DZ08d, Du07, EG06a, GY08, GAR09, HBVM07, HH08a, JZ06c, KA05a, KA06b, KPS06a, KR07a, KR07b, KY08, KYK08, KK08a, KK08b, KS08a, KS08b, KKA06, yLcCsL09, LP07c, MJ05, MA06e, MS06b, Moh07b, Muk07c, ND07, Nav08a, PS06b, PD09, PNL05, Rad05a, RJW07, Ram05c, Ram05d, Ram05g, Ram05h, Ram05i, Ram06d, Ram07g, RK07b, RMG07b, RGM07, RGAFVA07, SS06a, SH06a, SR06, SK06b, Sol05d, Sor06, SA05b, TiHuI08, Van07, WJ05, yWjCnG07, XO07, XXTL09, XRZ09, ZXZZ07, ZX08]. perturbing [Sol05b]. Pest [GM07, GZC08, HC05b, JC06, JCL08, RBvB08, SC08c]. petroleum [Mir08]. Petrov [Ism08b, MK05c, MKR06, Zha06a]. Petviashivili [Waz06k]. Petviashvili [BR09b, HZ08, Waz08c, Waz08g, Waz08m, Waz08s, ZX06c]. pexiderized [Kim08]. pharmacoeconomics [PSBM08]. Phase [BFKT07, BT06b, Che07h, Che08b, Cho07c, Kim05b, LL05h, Pud05, Qui09, Ras05c, RNRRMRCH09, RY05b, SA06a, Su07c, Su09, Vyn07, ZCFX07]. Phase-fitted [BFKT07]. phase-lag [SA06a]. phase-shift [Che08b]. phase-space [Pud05]. phased [RO07]. phases [CD09]. phenomena [DJJL06, SM07b]. Phenomenon [BX09, AH08a, Coo09, ICM08, WS08d, XC06a, ZMY08]. phi [Lou07]. phi-four [Lou07]. phylogenetic [HH07a]. physical [Abd07, Att06a, Hub09a, Hub09b, Waz05f, Waz05l, Waz08r, Zie09]. Physically [Zie09]. Physically-relativized [Zie09]. physics [CZ08b, XMGW09, Zie09]. physiology [KB06a, RMJ07]. phytoplankton [CRC08, GN06a, SMBB07]. Picard [Ahm05b, Ram08d, Ram09b, YEA07]. pickup [GTA06]. Picone [HZ09a].

Piecewise [AGPJV09, Ram06g, Ram07g, Ram08f, Cha06, DR06, DLBT08, JLZ05b, LZ09b, LMY07, LYL07, Nar07a, Ram05h, SK08b, VG07, XLwW07, ZYT09]. piecewise-analytical [Ram05h]. Piecewise-linearized [Ram06g]. Piecewise-quasilinearization [Ram07g]. piezoelectric [BS09a, Sin05]. pigs [SKSJ06]. piles [FSÇ06]. pin [SS05e]. pinning [XS09]. pipe [Mak07c, YZ06]. pitch [LC06c]. pitfalls [SdJA06]. pivotal [Ars07d]. pivoting [LSBL07]. pixel [LYGL07b]. pixel-pairs [LYGL07b]. pixellated [Vic09]. Piyavskii [RZ08]. placement [GAS08, dlS06c]. plan [Kia05, Smi06b, WY05c]. planar [Ahu08, HAA07, LCZT08, WWXX07, WC05b]. Planck [EA06, Su09]. Plane [SA06d, AA09, Ahm05a, Ahm07a, ALL07b, CLW09a, CGSS09, EB05d, ES05a, Kas07, KS08c, ST06c, Soh05, Sum07, TS05]. planes [Ahm07b]. plankton [JB05, LL08a]. planning [BKFM05, CPSS08a, DT09, GC06b, IGM07, Isl08, Lin06a, yLLjP06, MMM06, Mir08, RSA07, SGA06, ZGK06]. plant [DWZ07, NSS06a, WHD^{+}08]. plasma [JyGcL^{+}09, XMGW09]. plasterboards [BM05a]. plastic [FSÇ06]. plate [ASL07, AA09, AK05i, Cha05g, CB06, Cor05c, FXY08, KA09b, KNJ05, Lin05c, Mak06, NPR09, Pan06, VFF08a, VFF08b, WH06a, YJW06, YWJ07, ZO05]. plates [AH05c, BB09, CC05c, CFC06c, CH07b, KBLT08, TS09, YAO09]. platform [CH07f]. player [AR06, PSP05]. players [DHT09]. plot [Thu07d]. plus [GS06e, SGB05]. PN [ESA05]. PN-corrections [ESA05]. Pochhammer [Waz08w]. POD [Joy06]. Poincaré [Gin07, Nav08b, Ram07f, SK08b, Thu07d]. point [AK05a, AR09, AAS06, ACOS06, AO05a, ASK05, Ahm05b, AP06, AP09, AHIP07, ADK05, ADK06b, Att05b, Ber09, BA06, BW05b, BAM07, CCE06, Cao06, Cen06, CY07a, CY08a, Cha05g, CTG06, CJ08, CK06a, ĆC09d, CT08b, CZ07c, Cui08, DB07d, DZ07a, DZ08f, DHG09, EM06, EF07, Gen09b, Ham05b, mHzC05, HCLPJL^{+}08, Hon08b, HLW08, IOS07, JF05, Jan08c, Jan08b, Jan09b, Jan09c, JOB08, JGY07, JGZ07, JZ06a, JC09, Jun09b, KK07b, KYK08, KK08b, KW08a, KW08b, KR05, Kan07, KMA08, KKY07a, Kha07a, Kha07b, Kha08a, Kha08b, Kıy06, KLY07b, KT08a, Kum03, KA06e, Kum09c, KA06f, LW06b, LS06b, Li08d, LPS08, LZ08e, LD06b, LW06c, LC06b, LDL08, LH06a, LZW07, Liu08b, LLW08, LST09, Lu07, MNG09, MS07a, MD09a, MJ05, MK06, Moh06, MA06e, MA06f, MD09b, ME07, MHA06, NCÁHCLP05b, Nak08, NV07, NS06b]. point [OZ07, O’R09, PHL06, PNB06, PDW06, PB09, PC08, PW08b, PP08, Psi09, QSZ08, RLZ08, Ram05i, RMG07a, SS07a, SS09c, Sch08, SCX07, SAM07, Sil05c, Sil05d, Sil05b, Sil06g, Sil06c, Sil06a, Sil06b, Sil06e, Sil07a, Son07, SWZL07, Su08c, SL06c, SG07c, SMG08, SCO09, SA05b, SA06h, Tak05, Tan06a, Ver07, WJ05, WS07d, WC07b, WCH08a, WB09, WSL09, WP05, Wei06a, Wei06b, WLCL06, Wu05d, Wu06a, Xu08a, XZ09b, YW08, YKKY09, YHW06, YC08c, ZS07a, ZLCZ06, ZW07b, ZLZ07, ZL07a, Zha08f, Zha08e, ZYW08, ZZZ09, ZLW09b, ZS09a, ZWD07, Zha07h, ZC08b, Zha09h, ZH09a, ZZ09, Zhu05f, ZHZ07b, dlS06a, dlS06b, dlS07a, dlS07b, lS07b, lS08a, lSAQ08]. point-based [AHIP07]. point-sets [ZWD07]. point-to-group [YC08c]. points [AKR08, AJKM09, AA07f, Att06c, Att06d, sCLC09, CGSS09, DMS09, DLT08, Dzi09, Fad06, GR07a, GG09b, GD06, IR08a, LZ07c, LN08a, ML09, yMT08, OS09, Ram05d, Shi05b, Shi05c, Sil07c, Sil07d, SS07c, SQ07, TMS08, TS06, WL05a, WZDgH09, YFLX09, YL07, Zha09h, ZXZZ07]. Pointwise [Bou05, Wan09b]. poisoning [HC05b]. poisoning-pest [HC05b]. Poisson [AM09a, BA05b, lCS06, LMH06, OB09, Pal08, RZ07c, Sal07b, WZZ06a, sWX07a, Wei09, WYQ05]. Pol [jF09, GZ07b, GLLC08, gLXyZ06, Lin07a, ZXF07]. Polak [YGW09]. polar [DT08, Du05, Mur06]. polarized [KS08c, TA05b]. pole [Kos06b, dlS06c]. pole-placement [dlS06c]. policies [CH05a, CHW07, KHL06, SYT06, d’O05c]. policy [BW05a, BW07, CP05, JW09a, LWL07a, MAR05a, MBM06, MBM07, PBG07, Pal08, WWP05, WLT07b, YVU06, ZN05, ZL06c]. pollutants [KM06, NSS06b]. polluted [LZ09c]. pollution [BFI06, FEOC08, MMT07, Pat05a]. Polyak [YGW09]. polydisk [Ste09e]. polygon [GKK09]. polygonal [Mor09]. polyhedral [Mor05, ZMW05]. polymer [Ram07h, Vyn07]. polymerization [DV06]. Polynomial [RLZ07, SMA06, SG07a, Abd06, AN05a, AS06b, AD06b, Alt07a, AH09, BK08a, BP05, Bou09b, Boy06a, Boy07b, BX09, CC07a, CMV05, CL07b, CCJR07, Chi08b, DJ06a, DZ08a, EH09c, FY08, GBC09, HE08b, HR09b, JW09b, KA05b, Kan07, Kıy06, LLG06, Li07b, LCZT08, Liu07b, MNY07, NCÁHC^{+}06, NP05, PM05, PR05, Pet08a, PMP08, PHP09, PE07b, RJM07a, RJM07b, SS09b, SG05a, SAD06, SA07a, SA07b, SA07d, SA08e, Sil06b, uIKTT05, uITA06, uIK06, uINTK06, uITiHK08, SP06e, STV09, SS07d, TiHuI08, Tze07, Wu05e, YSS09, YW09, ZLC06, ZL06a, ZPH06, ZXF07, Zhu05d]. polynomial-time [SS09b]. Polynomially [LSN09]. polynomials [AA06b, AO06a, ADTGV09, AMJD06b, AMJD08, AES08, Alt06, ADK07, BB06b, Bha08, BIK06, BV06, Boy07b, Boy08b, Cao09, CMGMD06, Cas07a, CG07a, CAS08, CAS09, Cvi09a, DZ08c, EMC05, EH08b, EDS06, Gre06, GS05, bGgLgW09, Guy09, Hos06b, HHK^{+}05, İBİ09, KPS06b, LO08, Li06d, MSR07, MB07, MJ07c, MCS08, Ost08, PY07, PR07, Pou06, SLLL07, SP06a, Sum07, VS07, Wan07f, WL09c, WZ08d, WN08c, XB06, ZXLZ07, ZLDW09, ZCW05]. polytopic [dlS06b]. Pon [QCB05b]. poorly [Bog05b]. Popovski [Net08b]. Population [BSP05, ANM07c, AK05b, AA08, AAA08, AAA09, BCC^{+}09, BR07b, BKP06, Che05c, Che06d, CH09, DFGV07, DFG07, FT07, GHZS07, KC06a, KWdW08, LL08a, LS06a, LS08c, Mon07, MBM07, PS08, Pam05b, PW09, PRM06, PLL08, RBvB08, RS08b, SMM08, VJD09, WS07a, WWD09, WC05d, XL08a, YKJ07, mZzH05, ZH07a, lSAQ08]. populations [Efi09, Far08, MR05b]. pore [EESESD05]. poroelastic [BK06b, BK06a]. porosity [ZLL06a]. porous [AEA05a, AEG05, AH05a, AH06d, AKMH06, ADK06b, Att05a, AH05c, Bat08a, Che06n, Che07i, Che09b, CY08d, DX06, EEEE06, FXY08, GS08b, HS09a, HH09a, HMSZ07, JK07, KAS06a, KR08, KS05b, LY09d, MR09, Oua05, Pam06, Pan08b, PVS^{+}08, Ras05c, ST06c, SMH07, SK09a, SG09c, Su09, SDS05, SM08e, TS05, WMX07, ZLL06a]. port [Cho07b]. Portfolio [CF06a, AEA05b, AH06b, BTPGAPRU06, BTPGAl06, Cha05c, CLW09a, CJK07, Hor08, Hua06, HTO06b, Kat07, Kat08, Kat09, LHGT06, LG06b, OHT05b, SLC07, zWxY07, WZ06k, WSXL09, XXF06, YML07, ZN05, ZLW06b, Zha07e]. portrait [RNRRMRCH09]. pos [Fat06]. pos/neg [Fat06]. posed [AC09, CHLS08, CC05d, LZ06a, LH08b, Qia08, Raj08, SP05, Tad08, WS06a]. posedness [AA06e, GC07c, Lie06]. position [Dan08, XSX08]. Positive [Che05d, CY07c, DHG09, FT07, GG08b, GG08c, GG08d, HLW08, JDG06a, Jan08c, Jan08b, Jan09c, JGZ07, KWX08, Kan09b, Kha08b, LL05f, Li07d, tLlWpL08, LPS08, LZ08e, LLW07, Liu08b, LLW08, LLW09, LOA05, Ma07b, She07b, SWZL07, Su08c, SLS08, SS06g, SCO09, WS07b, WS07d, WC07b, WLW08, WP05, Wei06b, WZ08e, Xu07, YW08, ZLZ07, ZL07a, ZYW08, AJE05, Afr05, ANM06, AK06c, AK06b, AK06d, AMN07a, ANM07d, ANM07c, AMN07d, ANM07b, AMN07c, ANM07e, AMN07e, AMN07g, ANM07g, ANM08, AD07, AC07b, BQ09, CCS05, DZ07b, DZ07a, DZ08f, EDS06, Ghi06, GY07b, HW08b, HY05, HL07d, JGY07, JG07b, JFG08a, JFG08b, JOA05, Jia09, KW08a, KeTL08, LR07a, LR07b, LLZ06, LS06b, Li07e, Li07a, Li08d, LCJ08, LZ08d, LZW07, Ma05c, MH09, MSTY06, MY07a, MY08, PDW06, PFG08, Rad09, REDES05, Rim05g, Rim05a, Rim05h, Rim05b]. positive [Rim05e, Rim05c, Rim05d, Rim05f, Rim06c, Rim06g, Rim06f, Rim06d, Rim06h, Rim06a, Rim06b, Rim06e, Rim07c, Rim07g, Rim07h, Rim07d, Rim07f, Rim07a, Rim07b, Rim07e, Rim08c, Rim08d, Rim08e, Rim09b, Rim09a, SWX06a, SS05f, SS05h, SL06c, SC07d, SHD07, SG07c, SQG08, SMG08, SA06h, TJG08, WZ08b, WWD09, Wei06a, fW07b, Xu08a, XZ09b, Yan06d, YXW07, YWL09, YY06b, YY07b, Yin08a, YJO08, YK06b, ZZ08a, ZLW07, ZW07b, Zha08e, ZLW08, ZS09a, ZQ08, ZYZ07, ZH08c, Zho08b]. positive-definite [Li07a]. positivity [dlS07a, lS07a]. possessing [LP07c]. Possibilistic [Zha07e, GPTU05]. possible [Sim09a, UBKC08]. Post [Çak08, ESAR05]. post-Newtonian [ESAR05]. Post-optimality [Çak08]. postcritical [SS07d]. posteriori [pLT06, Raj08]. potential [BN09, BSHZ07, CC08a, Dem05b, EM06, FFKV08, LZZZ06, LZZ06b, LZ07c, MA05g, SANN07, SCR08a, SCR08b, SZ07e, ZLW09a, ZL07c]. potential-YTSF [SZ07e]. potentially [Chi06b]. potentials [Bou09b, GZ05, KB08, Kin05, LC09a, Ma06b, YLH09]. Powell [HWS07]. power [AES05, AEFFA08, ANK05, ALN06, Bat08a, Bis08, BR09b, Cha05h, Cor05a, EARS07, GA07, Hub09d, IAMA08, Jön05, KM05b, Kıy06, MNG09, Mug05, Nis07, SBS08, WML05, Waz06g, WN08c, YLC09, ZWT09b, ZZZ09, dCdSJM08, HRM07, YB05]. power-law [AES05, Bat08a, Cor05a, IAMA08, MNG09, YB05]. powered [HJB06]. Powers [GG08e, GG08f, EH09b, GG08b, GG08c, GG08d, LR07a, LR07b, Rim05g, Rim05a, Rim05h, Rim05b, Rim05e, Rim05c, Rim05d, Rim05f, Rim06c, Rim06g, Rim06f, Rim06d, Rim06h, Rim06a, Rim06b, Rim06e, Rim07c, Rim07g, Rim07h, Rim07d, Rim07f, Rim07a, Rim07b, Rim07e, Rim08a, Rim08b, Rim08c, Rim08d, Rim08e, Rim09b, Rim09a, WN08c, WS08d, Yin08a]. Practical [Li07a, LS09c, LWK05a, SLL08, JC07, MSD06d, Sol05d, WWC09]. practice [Cha07a]. Pragmatical [GLLC08]. Pre [Che07g, Kan09a, SA08c]. Pre-computer [SA08c]. pre-Schwarzian [Kan09a]. Pre-symmetric [Che07g]. Precise [CTG06, ZDL07]. precision [CK05, FAQ08, GSDB06, LC06c, Tak05]. precondition [MLM07]. preconditional [Zhu07]. Preconditioned [KTK06, MAD07, AL08, BMP06, Fad05, LY08a, LH07b, PVN06a, WH09, YK08]. Preconditioner [Yin09, BH07, qJ06, KKT05, PVN06b]. Preconditioners [Ben07a, HL06c, JY06, KK05d, KM05e, MR06b, PNB06, Sun06a, YJ07, YY06c, ZHLG07, HL06a]. preconditioning [DJ06b, DMA07a, EA08b, GG08a, GMCM06, GEG07, Jav06a, Jav06c, KPG07, ZHLG07, ZHL07]. preconditionings [DKK06]. predation [JB05]. predator [AM06, BB06a, BM05b, BIK05, Bou06b, Che06e, Che06f, Che06g, Che06i, CS07f, CH07c, Che08d, DK05, DKL09, EG05, FC06, FK06b, GCT08, HCZ06, HL07d, JMC07, JPCL08, Kis08, KK08c, LCS08, LX09a, LZC07a, LOZ07, LH07a, LTC07, LMC08, hMzLfW^{+}08, hMfWzLpZ09, Mak07b, MXC07, zMsCxL08, MB08, PWY08, RDGP07, RDG07, SB08a, SC08c, TQC09, WZ06b, WFL06, WW05b, XZ07b, XCD05a, XCD05b, XCD06, Xu06, XZ07c, YL06a, Yan07b, YCZ06, Yan06c, YC06b, ZCZ07, ZSS08a, HCW06]. predator-prey [CH07c, LCS08, hMfWzLpZ09, TQC09, XCD05a, ZSS08a]. predators [JMC07]. predict [AGTS07]. Predicting [SCR08a, GM05a, SLWX06]. Prediction [Par08, PG07d, SS05e, WLC05, AVSB08, BN07, DG07b, GW07, HC07b, JB08, LYGL07a, MAD08, Nwo07a, OKEL05, Tie05, WLX05, WC05d, WLC07b, WLC07c, WLC07d, YWZW07, YZD08]. prediction-based [YWZW07]. predictive [JPYW09, LP07b]. Predictor [Noo06b, NN07b, XD07, SMA06, Sal07a, SS07c, Wu06d, Wu06e, Wu07d]. Preface [BCN06, SP07, Sim09c, Udw05a, YM08b]. preference [HKH06, WF07a]. preferences [Tan06c]. preferential [FJT05]. prefilters [SCQ06]. prefixed [HR07b]. prejudiced [Sch05]. Preliminary [Bak05, QC07, XWSL08]. preponderant [LHXL05]. prescribed [CWZ07, YW09]. presence [AA07c, ESMM06, GS08b, Jön05, xLyW06, LL06a, LJE07b, PSRN06, PSS07, Sae06a, Sae06c, Sae06e, SSK06, SSS06, ZMM06, lS08b]. present [CHW07]. presentation [SSR05]. presentations [BS09c]. preserving [CMGMD06, CKOS07, HH07b, Ide06, Sun08, ZD06a, ZZ07c]. Pressure [GJ05, Ran08b, Ahm07b, dCDNPN09, EEAES06, ESS05b, KR08, KA09b, LMH06, MFE05, PR06, Ram07b, SG05c, SSTD05b, ZYH08]. pressure-gradient [ZYH08]. pressure-Poisson [LMH06]. pressures [NMS07]. pretreatment [MZL07, Pak11]. prey [AM06, BB06a, BM05b, BIK05, Bou06b, Che06e, Che06f, Che06g, Che06i, CS07f, CH07c, Che08d, DK05, DKL09, EG05, FC06, FK06b, GCT08, HCZ06, Hua08a, HL07d, JMC07, JPCL08, Kis08, KK08c, LCS08, LX09a, LZC07a, LOZ07, LH07a, LCZ05a, LTC07, LMC08, hMzLfW^{+}08, hMfWzLpZ09, Mak07b, MXC07, zMsCxL08, PWY08, RDGP07, RDG07, SB08a, SC08c, TQC09, WZ06b, WFL06, WW05b, XZ07b, XCD05a, XCD05b, XCD06, Xu06, XZ07c, YL06a, Yan07b, YCZ06, Yan06c, YC06b, Yan09d, ZCZ07, ZSS08a, HCW06]. prey-competition [Hua08a]. prey-dependent [LCZ05a]. price [CCL06, LHW07, MBM07, SSB06, WSXL09]. price-quantity [CCL06]. prices [AK08b]. Pricing [CWY06, Che07a, SS06e, ZW08a, AS07a, CL07c, CS08b, Chi06a, Chi08c, CK06d, FV05, Hor07, HL07b, Mar08, SLO08, SYL07a, SYL07b, TYCL06, Wu07a, You09]. Primal [MY05, ARK07, Ham05b, Ham05c, Pan08a, WB09]. primal-dual [ARK07, Ham05b, WB09]. primary [NSS06b, RAES05]. prime [CL05a]. primitives [JZCW05]. principal [Che07a, LC07, LCC06, PG07b, PG07d, Rim07g, Rim07h, Rim07d, Rim07f, Rim07e, WX09]. principal-protected [Che07a]. principle [DLD08, DDZ08, GOPP07, SSW09, Tan07b, Tan08b, ZLR07]. principles [An05b, BK08a, Zot07b]. prion [TSS09]. priori [HL08a, MR05a, SY08b]. Prioritization [AK06h]. priority [Çak08]. prism [GY06, Sil07a]. prismatic [Sil05b, Sil06d, Sil06b, Sil06e]. prisoner [AHE05, DJJL06]. privacy [LL05c]. private [PLJ05b]. private-key [PLJ05b]. probabilistic [FK09, Gu07a, HZ05, JOB08, JZCW05, Kad07b, LKW06, MJD07, OS08, QWX08, RSA07, SWK^{+}06, YES07]. probability [ALN06, CK05, ESAH10, GG05a, LS06a, LW06a, MR05b, SSR05, ST05c, Tak07a]. probability-fitting [CK05]. probability/entropy [GG05a]. probe [DLLN09]. Problem [Lie06, AHM09a, ASAS06, ABK^{+}06, Afr05, AK06b, AK06e, ANM07j, AP09, AKD06, AK05h, Amm05, AY05a, ALL07b, An09, AM09b, Ang06, AJOGRLRG08, AGPJV09, AA06e, AK06i, ATM06, Azi05, BC09a, BH07, BK07b, BK08a, BS09a, BGRS07, BGRS09, Bay05a, BUÇ07, BDLS05, BM05b, BIK05, BS07c, BA09, Bog05a, BP06a, Bog08, BD09b, Bra05, Bra06, BKS^{+}06, BT06b, Buk05, Buk09, Cai09, CA05, Cen05, CK08, CN05, CN06, CKL06, CS09b, Cha05h, CH06b, Che06k, CC06c, CLA08, lCX06, CG08a, CFF09, Chi06a, Chi07a, Chi07c, Chi07d, CJK07, CK06c, CS06b, Cor05c, CI05, Dai07, Dao05, DS05a, DG07b, DN07a, DN07b, DM07a, DT06, DE07b, DY07, Du07, DHG09, DS05b, EFE08, EPER07, EJER06, EJ07b, EB05d, EMY05, ESA05, ESAR05, EkES05, Ema06b, EG06b, Ere09b, FSG07]. problem [FE07, FG07, Fat06, FCG06, Faz09, FMH07, FHSS08, FK09, FHG07, FeT05, Gal08b, GL05a, GFWX05, GM06a, GY06, GB07, GH07, GK06a, GPTU05, GA07, GJ07g, GKK09, GW06b, Gül05, GC07c, GHX08, GC09, GS06e, HWH09, HGN06, HRM07, Has06d, mHzC05, HH09a, HP07, Hlo05, HC08b, Hsu05a, Hsu06, Hua05a, HW06, HY07b, HYG07, HY07c, IHB06, IPPT05, IIS05, IR08b, JHST05, JA08, JM07, JCSR07, Ji05, JS06, JGY07, JZ06b, JZ07, JRR^{+}07, JRA^{+}07, JR07, KA06b, Kas06c, Kas09, KL05a, KMN09, KTW08, KSSA07, KS05a, KJ08, KO08, KA06d, Kök08, Kum08, KH07b, Kuo05, KS06b, KE05b, KE05c, KE05d, LFM06, LMV09, Lee05b, LHY05, LL05e, LZC05, Li06a, LHZ06, LS06b, LD07c, LD07b, LAL07, LL08e, LHZ09, fLyHZ09, LS07c, LPS08, Lin05c, Lin06b, LG07, LDL08, Lin05f, Liu05a, Liu05b, LL06c, Liu06d, LZW07]. problem [LY08c, LLW08, LLW09, LZ07d, Luo06, Luo07, LHW07, Ma05c, MTC07, MJ07a, MCW08, MT08, MH09, MSM07a, MM05a, Mar05b, MP08b, MFF09, MeT05, MD09a, Moa06, MSM08, MN07, MH08b, Muk08b, NP08, NN06a, NR07b, NJG08, Oca07, OMST07, OASM06, Pal07, Pal09, PL05, PL06a, PC07a, PDW06, PB09, Pen05, PL06b, PHZ08, PCS07, QS06, QFX06, QFS07, Qia08, QYYL08, QW09, QW08a, QZL07b, QW06, QZZ07, Qua07, RAES05, RDGP07, RDG07, RES07, RK07a, RKS08, Ran08c, Rao06a, RG05, RJF07, Rey05a, RSA07, Sal07d, SM06a, SS07a, SS09c, SAK08, SK08c, Sev07, SAB08, SJ06a, She05c, SW06a, SCX07, SW09a, SP05, SK05, SP06a, SFPE07, SDR07, SLC07, SGB05, Smi06a, Soh05, SY08a, SP06c, SCL05b, SQG08, STU09, SA05b, Tan08a, TCW08, TLM09, TT06a, TT06b, TT07, TMASA05, TMSZ06, TMRV06]. problem [TMJJS07, TA05b, TA06, TY07, TT09, Tua13, UBKC08, VHFF08, WXLH06, Wan06b, WC07a, WZS07, WWWL07, WS07c, WS07d, WC07b, fWZpXzL09, WZ09a, Waz07j, WYL06c, WW05c, WZ06k, WS06a, WX08b, Wu08d, XFL06, XF06, XC06b, XFQ06, XSKC07, XW08, Xu08a, XL09b, XZ09b, XYL10, XY11, XXF06, YSL05, YLLL08, YW08, YZ09b, YLY07, rYzS07, YWJ07, Yi05, Yig08, YW06a, YW06b, YHW06, Yua09, ZMM06, ZMW05, ZW07b, ZOZZ08, ZYW08, Zha05e, Zha05d, Zha06c, ZW08b, ZHZ08, lZXL09, ZW05h, ZHZ07b, dCdSJM08, SS09b]. problems [ASA05, ASAEE06, AK06c, ANM07a, AMN07b, ANM07e, ANM07h, AN05a, AOLL05, AV09, AOS09, Akm09, AS06a, AS06b, AS05, ASNR06, ASN07, ACJ09, AR07a, AV07b, AE07c, ASA07a, AJRLRGOG09, AO05d, ADK05, ADK06a, ADK06b, AH08c, Asa07b, AK08b, Att05b, AL06, AR07b, Bab07, BMS07, BK07a, BKA07, jB06, Bar09, BDCW06, Baw05, BLLM08, BBB08, Bi06, BNKZ09, BW05b, BR09c, BKFM05, BAM07, BT08, CCE06, CC06a, CC07a, CP08, CCWJ08, ÇG05a, Cen06, CY07a, CY08a, CPR07, CC06b, Cha07b, Cha07c, Cha07d, Cha05d, CCFC08, CWLH08, Che06l, CTG06, CCJR07, Che07g, CM08b, CJ08, CXZL09, CSY07, Cho07b, CG05b, Cor05b, CZ07c, CG08b, Cui08, CR05b, DSZ07, DSKM09, Din05, DMH05, DXH07b, Dön06b, DJ07, DZG06, DZ07a, DZ08f, DLT08, Dyk09, EB05b, EB05a, EN06, EGN07, EHN07, EJ07a]. problems [EA07b, EG07a, EG07b, EL05a, EL05b, ElS06, EGX05, EM07, Fab08, FW07a, FHH09, FM05, Fra09a, FS05, FH06b, GBK06, GKAS07, GXS05, GC06a, Gar07, GTA06, GC07b, GL08a, Gen09a, Gen09b, Gha08, GOMA06, GTJ07, GJ07c, GY07a, GG05b, GITV06, HEZ07, HEZ08, Ham05c, Han06, mH07b, HH08a, HRE05, HMNT09, HJ08, HJ06, HN05, HAU05, HHC08, HL06c, HLW08, HH06b, Hua05b, HM06, HT09a, HT09b, Hua09, HL05d, HJK08, JDG06a, JLI06a, JLI06b, Jan07a, Jan07b, Jan08a, Jan08c, Jan08b, Jan09b, Jan09c, JDSR08, JS07a, JJT06, JG07b, JGZ07, JFG08b, Jia05, JZ06a, JZ08c, JC09, JLW09, JC05, Jin08, JCM05, JH08c, JM06b, JM06c, Jun09a, Jun09b, KA05a, KA05b, KK06, KP06a, KPS06a, KK07b, KK07a, KY08, KK08b, KS08a, KM09a, KB09, KKF07, KBH06a, KBH06b]. problems [KW08a, KW08b, KWX08, Kan09b, KR05, KB06a, Kan07, Kar05a, Kas07, KK07d, KKKM09, KH07a, KKA06, Kha07a, Kha07b, Kha08a, Kha08b, KGM06, KM05c, KBTS07, KX08, KT08b, KNJ05, Kum03, KM05f, Kum06, KA06e, Kum07, Kum09c, Kut05, LY07b, yLcCsL09, LZ06a, Lan07, Les06, Li05a, LZ06b, LS06c, LYKC07, LH08b, Li08a, Li08d, LCJ08, Li08b, Li09a, Li09f, LB06, LD06b, LW06c, LC06b, Lin07b, Lin07d, LH06a, gLxW05, LLW05, LTT06, LZ06e, Liu06h, LX07b, LY07d, LD07d, LLW07, Liu08a, LA07b, LA07a, LM07a, LMN07, LST09, LP08c, Lu07, LGw07, LP07c, LP09, LG08, MK05a, Ma05a, Ma07b, Ma08a, qMpXjH09, MA08d, MFD07, MO05, MY05, MA07d, Mam09, MM08b, Mav09, MK07a, MKSK07, Meš07, MSD06d, MV09b, MJ05, MK06, MA06e, MS06b, MA06f, MD09b, MMN06b, MAO06, MHA06, NKE05, NPR09, Nat09, ND07, NKR^{+}08, yNyF05]. problems [yN05, Noo06b, NMD07, OM06, Odi09, Oht08, OSS09, OASM05a, OASM05b, PNB06, PSW07, PFG08, Pao07, Pat05b, Pat07, PC08, PHZ06a, PLTW07, PW08b, PD09, PP08, Psi09, Ql08, Qia09, QM05, qQcM05, QSZ08, Raj08, RLZ07, RLZ08, Ram05c, Ram05d, Ram05f, Ram05g, Ram05h, Ram05i, Ram06g, Ram07c, Ram07d, Ram07e, Ram07a, RK07b, RJ07, RMJG07, RJM07a, RMG07b, RJM07b, RMJ07, RMG07a, RGM07, RJMG07, Reb06, Rem07, RAS08, RSZ08, Saa05, SF07, SS09a, Sal06a, SA06a, Sal08d, SK07a, SH06a, SG05b, SR06, SM06c, SZ05, SFA08, SA06f, SA06g, SAM07, SA07a, SAN07, SA07b, SA07d, SA07c, SA07e, SAE08, SA08e, uIKTT05, uITA06, uIK06, uIT06, uINTK06, uITiHK08, SS08d, Sok09, Son07, Su08c, Sun06a, SL06c, SG07c, SMG08, SCO09, SA06h, TMS08, TMJV^{+}05]. problems [TMSG06, Taw07, TS06, TJG08, TD09, TiHuI08, THLC07, Tso09, Ver07, Wan05h, Wan05b, WB06, WSB06, WZ06a, WHL06, WR07a, WLK08, WC08a, Wan08b, WCH08a, WLW08, WH08b, WLW09, WSL09, WP05, WLQ06a, WLQ06b, Wei06a, Wei06b, WZ08e, WLZ06, WC09c, XO07, XD07, XZ09a, XL08b, XCS07, XXX06, Xu07, XL07, YY05a, Yan06b, YS07, Yan07c, YC08a, YWL09, YZ09a, YJ09, YC07a, YM07b, YM08a, YE07, Ye06, YCL07, YCL09b, Yin09, Yin08b, YRY05b, You05, YEA07, You09, YJO08, YHL08b, YHL08c, Yua08, YHL09a, YK06a, YKJ07, YCYL05, ZFN07, ZLW06c, ZLCZ06, ZLZ07, ZL07a, Zha08h, Zha08f, Zha08e, ZL08b, Zha08a, ZWG09, ZS09a, ZWZ05, lZyT05, Zha07h, ZC08b, Zha09h, Zha09g, ZH09a, ZWZZ07, ZCH07, ZLC07, gZW08, ZZ09, Zhu06b, ZS07d, ZJO08, dAC09]. procedure [BL08, CKL06, DH06e, FSG07, Fat05a, Fat05b, HWC07, HW05c, JSdRm05, LWL07a, Pal09, Sam06, TT07, WXLH06, WSB06, WLT07b, WLL07a]. procedures [Bar05a, Deh05b, Fag05, GR06b, GR07b, Muk07a, Utk06, WS05b]. Procesi [HRLC08, MYG08, SGC09, SX06, Waz07f]. process [hB06, hBxY07, Bog05e, CH05a, EM11, ETJ06, GY08, GGSNR06, HT06b, IM08, JK06, KR09, KY06, LZ07b, LZZC09, LY09c, MQW05, Ozd05, PB07a, Sad09, SWK^{+}06, SZ07a, SMB08, Tha07, Wu06d, Wu06e, Wu07d, WMC08, ZL05a]. processes [AN07, ANA07, Cha05h, HIST06, HR07a, NR08b, Ned05, Pog07b, WS05a, WL05g, ZW08a, ZL08d]. processing [CYL07, Deh08, DFG07, Lin05a, LC05c, NJG08, Wan06d, HJL08]. Procrustes [PHZ08]. prod [GK06b]. producing [GN06a]. product [BUP07, CHW07, CMR08, GWJ07, LCX09, Mis05, NSS06a, SJ06a, Sok07b, TS08c, YVU06]. production [ABB06e, BI08a, BDPP08, IGM07, IR06, IR07, JLZ05b, JTMRS06, Kia05, yLLjP06, MR06c, MTD05, Nwo06b, PBG07, QW06, TMRV06, WS05a, ZGK06, ZHX06]. production-inventory [WS05a]. production/service [Nwo06b]. productivity [EARS07, Yu07]. Products [Nad05, AL07, Aou07, CS08b, Chu05, DSMK08, GN06c, IR05, LCC05b, Mir08, Nis07, PN08, WLL07a]. profile [Cha05h, Fan06, HKS09b, Oss05b]. profiles [Cai06b]. Profit [Liu07e, Liu06a, Liu06c, Liu06e, Liu07d, Liu09d, TARm08]. Profitless [MXC07, MCC07]. program [AK06f, QBH06b, YSdS^{+}06]. Programmable [JY07a]. Programming [Zha09a, ASA05, ASAS06, ASAEE06, ARK07, AH06a, AR08, AE07c, AZ07, AH08c, AHS06, AZ09, BK07a, BL08, BTPGAPRU06, BTPGAl06, BA09, BI08b, Çak08, CN05, CKL06, Cha05a, Cha05b, Cha05c, Cha05d, Cha06, CH06b, Che06k, Cho05, CHY^{+}09, DZG06, EKM08, EB05b, EB05a, EM06, ER06a, EN06, EJER06, EGN07, ElS06, Ema06b, Fab08, FY08, FHH09, FA05a, GXS05, GXY06, GOMA06, GTJ07, GK06a, GW06b, GG09a, GS06e, HGH06, HC05a, Hor08, HD07, HH06b, HLK06, HTO06c, Hua09, Isk06, Isk07, IR06, IR07, Isl08, JHST05, JA08, JM07, JGS06, KOR08, Kas07, KL06a, KHL06, KKY07a, Kök08, KOS08, KB07, Kun08, LZ06b, LD07b, LT08b, LLLW06, Liu06e, LW07, Liu07e, Liu09d, LHWW07, MR06a, MAN06, MBM06, MY05, MA07d, Meh06, MD06c, Mir06a, Mir06b, MHA06, MZL07, NNN05]. programming [yNfM06, OASM05c, OMR07b, Pak11, Pan08a, PS06a, PGdSS08, QZJ07b, RSA07, RAS08, Saa05, SM06a, Sha08a, SGA06, She05a, SJ06b, SJ06c, SZL05, SLZ05b, SLZ05a, SLZZ06, SZL^{+}07, SGB05, Sit06, TTG07, WL05e, WY05b, WZT06, Wan06l, WLLL07, WWWL07, WS08a, WLZ06, XW08, YQ09, You05, YCYL05, ZGH07, ZWZZ07, gZW08, Zhu05f, Zhu05g, ZS07d, Zot07c, GPTU05, KR07d, MD05, Mav09]. programs [Amm06, Ars07c, Ars07b, Ars07d, wG07a, Ham05a, Ham05b, HW06, SG08c, SDM08, SCM09, Yan07a, ZZ06b]. progression [SL06a]. progressive [WWC06]. Project [KL05a, ASA06a, JDSR08, RK07a, RKS08, Ran08c, TT06a, WZ06k]. projecting [Mor05]. projection [BNK07, EGN07, Han06, HW09, HN07b, Jia05, JB08, KF08c, LSN09, Men08, NM08a, RMAM07, SS06a, SC09, YHS09]. projection-based [JB08]. projections [yMT08]. projective [BL08, KKY07a, hL08c, Zhu05a]. projectors [BT09a, BR08b, XSG08]. projects [TG06b, ZS07d]. proliferation [dG06]. PROLOG [MG06]. promoter [PG07c]. proof [Bri08, CCC07, HWS07, JC09, PH06, Tan06d, Tan07c]. proofs [Cao09, QCDS09]. propagating [DV06, Ma07c, ZM06a]. Propagation [Ram07h, SA06e, ADH06, CB06, Kin06, SS05e, Sin07a, SZXL06, SC06c, YC08c, ZZLL07]. propagations [ZL08c]. Properties [Ate06, HL08d, SZF05, TA06, AMK06, Aki05b, Att06a, BD06a, Bor09, CMM07, CHY05b, CK07a, CWZ07, Dod08, Du06, DZT06, EA08a, Gup08b, Kar09b, KG08, Kim07b, KS08e, KVŽŽ08, Liu07c, LP08b, Liu09b, MTL08, QCDS09, RB07, Sai07, SEŞ09, TmChL06, Taw06, VAMV07, yWjL06, WYL06a, XT06, You06c, YL07, ZZ08b, dEdSNL05]. property [CO09, xLyW06, OML09, XCS07, ZZ05b, LT09]. proportional [EH06, IM07, IMB08, SH07f, SYL07b, WQL09]. proposed [ETETH05, MK07b, d’O05a]. prospect [Nwo06c]. Prospective [DT09]. Prospects [GRZ06]. prostate [dPCdBYV08]. protected [Che07a, CS08b, PKH05, PK05a, YY05d, ZCL05]. protection [INR08, LL05c, RYE07]. protein [LMT^{+}09, RNRRMRCH09, SCR08c, SLWX06]. proteins [TSS09]. protocol [CYY05, CJC05, CJL05b, DL05, HHC05, LKY05b, LKY05c, LJL06, LC05e, Shi05a, SW05b, WLT05b, YW05, YTWY05, YRY05b, YRY05c, YRY05d]. protocols [BCC^{+}09, SL05a]. Provably [ZCL05, ZW05d]. proved [Bou06c]. provided [LT09]. providers [GCSS05]. provisioning [AS06c]. Proximal [ASK05, Ver07, BNKZ09, Ham05b, Ham05c, MHA06]. proximal-point [MHA06]. Proxy [BCL05, LCZ05c, BCW05, CJC05, HW05a, HWH05, HW05b, HH05, LL05c, LC05a, LC05d, LHH05b, LCZ05d, LW05c, PKH05, PK05a, Sha05b, Xie05, ZCL05]. proxy-protected [PKH05, PK05a, ZCL05]. PRP [SS06b]. prune [SG05d]. pruning [fQZgH08, SW05a, SJ06c]. Pseudo [RNRRMRCH09, BBES07, ĆRR^{+}08, HBN06, JY07b, Li06d, LZX08, LL09, hLqSlZ09, LL12, SG07c, Yan08e, GKR06]. pseudo-contractive [ĆRR^{+}08, HBN06]. pseudo-exhaustive [JY07b]. pseudo-hyperbolic [LL09, LL12]. pseudo-Newton [GKR06]. pseudo-parabolic [Yan08e]. Pseudo-phase [RNRRMRCH09]. pseudo-remainder [Li06d]. pseudo-spectral [BBES07]. pseudo-symmetric [SG07c]. pseudoantiharmonics [PRA^{+}09]. pseudocontraction [sCLC09]. pseudocontractive [CPY09, CAA07, GG09b, OS09, ZSM07]. pseudoharmonics [PRA^{+}09]. pseudohyperbolic [GR07b]. pseudoinverse [Gon07]. pseudoinvex [AJRLRGOG09]. Pseudomonotone [Sad09]. pseudorandom [CHL06, WTG07]. pseudospectra [Du06, SZF05]. Pseudospectral [Hos05d, Jav06c, Baz08, Boy06b, Che05e, Che07e, DJ06b, DKK06, Elg09, Hos06c, Jav06a, Wan07a, Xia06]. PSO [HM06]. PSOSA [lLWhL06]. PSPICE [ACZ05]. PTC [KE05b]. public [BCL05, CCH05, JXW05, JZCW05, LKY05b, LWK05a, Sha05b, SLC05, SC05d, TLH05, Tsa05]. public-key [JXW05, JZCW05, SLC05]. published [KPL09a]. Pull [ACD^{+}05]. Pullback [LZ07b, LWW09]. pulsatile [MCMA07, SVMG09, SSTD05b]. Pulse [WTS09, AKVH05, GTX08, HFL06, JY09, JWS^{+}09, LZ09c, MCC07, MC08, SJW09, Wan07f, WL09c, KV05a, MSK05]. pulse-spectrum [HFL06]. pulses [LCS08]. pumping [LJ08, VSB05]. pure [Dai07, Day05]. purpose [LML08]. Push [ACD^{+}05]. Push-and-Pull [ACD^{+}05]. pyramid [JH08c, She05b, She06]. PZMI [KF08a].

QMRCGSTAB [LGHS06]. QoS [AS06c]. QP [Zhu05g]. QP-free [Zhu05g]. QR [NS05]. Quadratic [Aks06a, DD08a, uINTK06, dlS07b, Abb06a, Abb06b, AZ07, BSKK06, BK08b, BR09a, BZZ08a, BNKZ09, BSHZ07, Cho05, DT06, EB05b, EN06, Fab08, Far05, GL05a, GXS05, Gao06, GOMA06, HZ09a, HL07b, LZ06b, LDZ06, LT08b, LB06, LL06c, LW07, MT08, MO05, MA07d, MM08b, MZL07, ÖEK05, Pak11, PR09b, QC05, QZZ07, Sha07c, SG08c, SDM08, SCM09, SL09b, TF07, WSB06, WLY06, WZ08f, XC07, YLLL08, Ye06, SAK08]. quadratic-time [HL07b]. quadratical [CWLH08]. Quadratically [CWLH08, CL06c, Chu07g, Kah07, MKKS05a]. quadrature [AA06d, AA07e, BDDB08, BJE05, BDE06, CT07c, DB07a, DMJE05a, DMJE05b, DMJE05c, DMJE06a, DMJE06b, fDxZfWyY09, EDMJ05b, EDMJ05c, EDMJ05a, EMJB05, Has06a, HEDMJ06, HED06b, HED06c, HMJD06, HMJED06, KKY07b, LC07, LWWS09, ML05, MJHED05, MJED05, MJHED06, MJ07c, MJD07, MRN09, MO07, MZ07, Pre06a, Pre06b, RVN07, RNV07b, RVNI07, RNV07c, SNH07b, SNH07c, SKHZ06, Ser05, SE07, Wel08, XG08, yXhWZ06, ZSHB06]. quadratures [AO05c, AO06c, FPP06, GVMT07]. quadrilateral [IH09]. Qualitative [JL07, KK08c, DDR05b]. quality [Kök08, KWP07, TT06b, TT07]. quantiles [RAGA^{+}05]. quantitative [ZYD07]. quantities [Moa06]. quantity [AA06a, CCL06, IR07, Liu07e, Liu09d, ZLW09b]. quantization [Tse05, WY05a]. Quantum [CH06a, Ade07, AR06, GRZ06, MSM07a, XSX08, Smi06b, WTW05, YSCC05, ZSHB06]. quantum-behaved [XSX08]. quantum-classical [GRZ06]. Quartic [AS05, SD09, SAE08, GC06a, LLC09b, Ras05a, SA08a, ZL08e, EHZES05, NKE05]. Quasi [Deh05b, ASK05, Aou07, Ber09, BN08, Bou05, CW06b, CW07, DL08, DF09, DW09c, Gu07a, JK08, KK05b, KKKM09, LY07b, Lin05e, Liu06h, LN08b, Ma05a, MX07b, NS06a, PFG08, SZ07a, ST08, WZT06, WLQ06b, WLY08, WS06b, XD07, XM06, XM07b, Yi05, Zhu05b, ZW09c, Ree07, WY09]. quasi-equilibrium [XD07]. quasi-Hadamard [Aou07]. Quasi-implicit [Deh05b]. quasi-interpolation [CW06b, CW07, Lin05e, ZW09c]. quasi-inverse [NS06a]. quasi-linear [Gu07a, LY07b, MX07b, PFG08, ST08, XM06, XM07b]. quasi-Newton [LN08b, Ma05a, WZT06, WLQ06b, WS06b, Zhu05b]. quasi-Newton-SQP [WLY08]. quasi-nonexpansive [DL08, SZ07a]. quasi-periodic [Liu06h]. Quasi-shrinking [Ree07]. quasi-stationary [Yi05]. quasi-variational [BN08, Bou05, DF09, DW09c, JK08]. quasi-variational-like [ASK05, KKKM09]. quasiconvex [YL07]. quasicrystal [LF08, Shi09a]. quasicrystals [Akm09, XFTY09]. Quasilinear [Fig08, AC07b, CY07c, JGZ07, Kan09b, LY07d, LY08c, MY07a, MY08, Vul07, WG07b, WY08, Yan06d, YXW07, YCL07, YX09, YY06b, YY07b, YY08c, ZY08b, ZS06, ZYZ07]. quasilinearization [HS09c, LV05, MV08, PC08, Ram07g, ZWC09]. Quaternion [SB06b, BS07b, WWL06, WSL07, WCN08, WYL08a, WCL08]. Qubit [MSM07a]. Qudits [AR06]. Quenching [JZ08a, ZM07, ZM08d, ZMY08]. query [ZCZX07]. question [WW08a, dCDB07]. questions [WW06]. queue [AA07d, AC06, AM05, BGP06, Cha09, CD09, GS06f, KL06b, hLsT07, MX08b, MAR05a, Wen05, XM08, Zho05]. queueing [CMC06, Cho07c, CK06b, CK07b, KL06a, KC06b, WWP05, WY09, ZL06c]. queues [AA05, Aza06, GMS06, KHL06, Oma06]. queuing [SM06b]. quick [ÁFVZ06]. quiet [Udw05b]. Quintic [SAN07, Ram07f, SDL09, ZLC06, ZXZZ07, ZXZL07, ZXLZ07, NKE05]. quo [XLHK09]. quota [QW06]. quotient [BMP06, CI09, gLL07, SC08b].

R [Ren08a, KSP08]. Radar [WS07g, WS07f]. Radau [MJED05]. Rademacher [HTM09]. Radial [GS07a, WCH08b, AE07a, BW09, CMM07, CLNCG07, DZ08e, DjT06, Dua06, FWJK09, GS06b, GS06c, Hu07a, HZ05, MW09, MJ07d, Ony05a, Su09, Yoo05, Zha06a, Zha06b, Zha07g]. Radially [CB06, DV06]. Radiation [AEG05, Bat08c, Bat08b, AH05a, EEAES06, ESS05b, Kha09, LH05, LLW05, OSS09, Oua05, PM09, QS06, QSS08]. radiative [ZMV05]. radii [AST08]. radio [Alr05]. radiotherapy [Çet07]. Radius [Sok09, Yan07e, Zho06]. Radon [Kim07a, WZ09b]. railway [AGI08]. rain [NSS06b]. Rainfall [Hon08a]. ramjet [HJB06]. ramp [KE05c]. Random [LSBL07, Sum07, AB05b, hB06, hBxY07, CT08a, CJK07, CD07, DC07, ETETH05, ETETH07, FZ07, HK06, KC06a, KOS08, KT07c, Liu07b, Liu08a, MMM06, PJ10, RSD07, Utk06, Wei09, Wu07c, YXSX05, ZZ07b]. randomization [WHH05]. randomization-enhanced [WHH05]. Randomized [BK07b, hB07, Cha07a, CH07f, TT06a]. randomly [ETETH07, HCLP07b]. randomness [Kat05, KCK08, Liu07b]. range [Abd06, DL09, bGgLgW09, KP09a]. Rank [WSL08, AMAP07, CC06c, DKA06b, DE07b, Du05, Erb06, FZ07, FHZ08, Jbi06, WL07f, ZW05h, ZW06d, ZW06b]. rank-deficient [Du05]. rank-one [WL07f]. ranked [Dan05, SAS07]. Ranking [AJK05, Utk06, AA07b, EARS07, JMLR05, JLRB05, JPZ06, JLBR08, KC09, LJK07, MAN06, Sae06a, Sae06e, SW06c, TA05a]. ranks [WWL06, WSL07, WYL08a]. Raphson [Dan08, GA07]. rapid [Kat09, Kha07a, Kha07b, MV08, XL07]. rapidly [mH07a]. rarefaction [ZYH08]. Rashed [EM11]. Rate [Gup08a, GBG08, GI08, İBİ09, JC05, AS06c, DLS06d, Far08, GL05c, GBG07, Hor07, Lee05a, LWH06a, LZZ06b, Mon07, MBM07, SS06e, SQ07, TK07, Vaj07c, dWqL05, WTS09, WZ06k, ZLW09a, ZL07c]. rate-based [AS06c]. rates [AAA08, DD07, Thu07d]. rather [CS07b, SK06b]. Ratio [KC06a, RAGA^{+}05, BCEGMRMP07, CS07f, DKL09, HL05b, Kis08, KK08c, Mak07b, SB08a, Zen09]. ratio-dependent [CS07f, DKL09, HL05b, Kis08, KK08c, Mak07b, SB08a, Zen09]. Rational [IM07, KB05b, Yax07, Ada07, Alo06a, Alo06b, Alo09, ADDdM07, Boy08a, CW06c, CW06d, CY06a, DJJL06, DDD05, DMS06a, DMS06b, DDR06a, DDR06c, DWT05a, DWT05b, DZT06, DLBT08, GA08, HL07c, Ino08, IS09a, JH08a, JV07, Lay09a, LS05, LZ08c, Lin09b, MSD06c, Özb07a, Rab05, RB07, SA05a, SAB06, SA06c, SA08c, SZ07b, SZ07c, WSKZ07, WS08c, YLZ06, YLB05, YEM06, ZW09a, ZYME06, ZZ07a, ZZZ07]. Rationality [ZLZG06]. rationalized [MM05b, Ord06b]. ratios [GN06c, GN06b, JS07b, Nad05, QZZ07, SW06a, SCM09, YPZ05]. Raviart [SP09]. ray [LLLW06]. Rayleigh [AHM09a, BMP06, CLA08, GL09, GGSN06, gLL07, Vin09, XXC05, XXC06]. Rayleigh-type [GL09]. RBF [Cai09, WLP07]. RBF-based [Cai09]. RBFs [ZT06, Zha07f]. RCGA [MBM06]. Re [PJ10, CN06, KO06, SM07a, XBZ06]. re-analysis [CN06]. re-entrant [KO06]. re-gridding [SM07a]. Re-seeding [PJ10]. reach [AA05]. reaction [Bab07, BP06a, Bra06, mCLB08, CG05b, CGJ08, DG07a, ESS05c, EWAE06, GD08, JW07, LZ07b, LWW09, MR05a, MV08, ND07, PD09, Ram05b, Ram06b, Ram06e, Ram06f, Ram07c, Ram09b, dSYGM^{+}05, Sta06b, WZ08a, XO07, XCD06, Xu06, YS06d, ZQ06, ZS06, Zhu08]. reaction-diffusion [Bab07, Bra06, ZQ06, ZS06]. reactive [Mak07c, Mak09]. reactor [DN07c]. Real [HLCH08, MTD05, Boy06a, DT07a, DT07b, DSKM09, Erb06, GKR06, Geu07, GL06b, GY07b, GG08e, GGSN06, KM07, LL07a, MBM06, MNM07, NBO09, Nwo07b, SB06b, Tso08, Van09, WD06, WLZN08, WZ06k, YZDW07, KSSA07]. real-coded [MBM06, KSSA07]. Real-time [HLCH08, Erb06, YZDW07]. real-valued [Van09]. realistic [MB05b]. reals [LCM07]. reasoning [LY09c]. reassignment [DFGV07]. reboot [WC09b]. Rebuild [HT05]. recapture [LS06a]. Receding [CPSS08b]. recently [JWL^{+}07]. receptors [LMT^{+}09]. reciprocity [Ang05]. recoding [BK07c, LLWC07, WLC07a, YGL05]. Recognition [WWXX07, ALSL08, DY06, DWZ07, xGZZ08, KF08a, KF08b, KSC^{+}08, LM08a, LCL06, PG07d, QZ09, RNRRMRCH09, Su07a, Su07b, Su07d, WWHL08, ZCZ05, ZS07c]. Recognizing [AK06f]. reconfiguration [ONG08]. reconsideration [NS07b]. reconsidered [Haa09]. Reconstruct [Zha07f]. Reconstruction [Fan06, JP09, LXSH06, ME07, OIL09, Pog07b, QYYL08, SLLL07, WWC09, YM06]. record [MS08b]. records [FA06]. recovery [Bou09b, CCH05, HW05b, LC05d, Nwo07a, ZF05]. recruitment [VKVHN05]. rectangle [Boy05, BO06, GXS05, Ree07, SJ06c]. rectangles [CVMM06]. Rectangular [Odi06b, DKA06a, GM05b, LD06a, Odi06c, Pra07, SG06, XC06b, YL08b, YWJ07, ZY09]. rectifiable [UP07b]. recurrence [EM03, MCS08, Sal06c, Sog07, lY08b]. Recurrences [AES08]. Recurrent [LHGT06, BR08a, Guo09, MA07d, Par08, RB08a, SBS08, YZFL08]. Recursion [OSS09, BB06c, EH09c]. recursions [Wu06f]. Recursive [HCLPJL^{+}08, NCÁHCLP05a, Wu05e, CT08a, DD05, DDR06a, DD06b, DMS06c, DDR05a, DDR06c, DDR06b, EMR08, EMR09, EH09a, HZ05, KT07c, LL08b, NCÁHCLP05b, Nak07, OOB07, PH06, TS08a, TAK07b, YSC^{+}05]. recursively [CHT06, HHK^{+}05]. red [BMD08]. reduce [GXS05]. Reduced [dCdSJM08, AMAP07, ASA07c, Whi06, Zhu05a, Zhu07, AG08]. reducibility [EM05b]. Reducing [SM05a, LW05b, Liu07a]. reduction [ASA07c, Bel08b, CWY06, EM06, Fen05, HJ06, Hos05a, HC09, KW05a, Kay05d, LWL07a, LY09b, MG06, Oss05b, PG08c, RLY06, Sjö07, Sri09, WLT05a, XL09a]. reductions [AKBZ08, CIS09, MMT07, MZ09, lYpZqL08, YC06d, Zhi09]. reductive [Dem05d, Slo07]. redundancy [Che06l]. redundant [AKKS06, KC07b]. reentrant [DEH05c]. Reeves [yWjL06]. reference [BK06b]. refinable [SY07c, XY07, Yan06a, YH09]. refinement [Bie08a, Bie08b, CGS08, DJ06a, Gui05, LHC07, SS08d, WF07d, WS08e, ZWZL06]. Refinements [QG08]. reflected [AS07b, SL07b]. Reflection [SKS06, ST06c, ADH05, RLdLSF07, SK07b]. reflectionless [Kin05]. reflexive [DH08, LL07a, Pen05, hPyHZ06, WCO08, XL09b, ZZH09, ZHZ08, ZWC06]. Reformulation [Cha06]. refuge [HCZ06]. refuting [Smi06b]. regarding [GK06b, KG06]. regime [LS06a]. Region [PVY09, Ars07c, Boy05, BO06, Che06m, CC07b, gDH07, FHSS08, FS05, GCSS05, IH09, JjQjWmL06, KS06b, LL08c, LMD^{+}08, LN08b, MSD06d, MZW05, MHA06, yNfM06, PHL06, hPY06, RVN07, RVNI07, jW06a, WZT06, WZWW08, WYL06d, YQ09, ZL05a, lZ06d, Zha07b, Zhu06a]. regional [SGA06]. regions [DjT05b, LY07a, ST05b]. registration [Su07c]. regression [CC07a, ÇE06b, ESK07, HN06a, HZ06b, IM08, JWL^{+}07, LC06c, MNN05, NNN05, Sam06, YY06a, ZHZ07a]. regression-based [JWL^{+}07]. regression/econometric [HN06a]. regressors [ZZW08]. Regret [Nwo06e]. regula [Che07f, CS07h, PG06]. regula-falsi [PG06]. Regular [Waz08s, BŞ07a, FA05a, GKK09, Prá08b, TD05, dlS07a]. regularity [Bar09, DB09, Geo09, Kar09b]. Regularization [WZ09b, AC09, CXZL09, FXLZ05, JZW08, KB06c, LZ06a, LLFM08, LC05c, LS06e, ME07, NK05b, QW08a, Sad09, TK06c, XF06, XF08a, Mou08]. regularizations [KH07a]. Regularized [CS07i, PR09a, BŞ07a, Cha07b, EK06, Gül08, KS05a, Ras05b, TT09, Tua13, Yin09, Zha05b, WZC07]. regularizing [CMM07, CHLS08]. regulation [GM07]. regulatory [QSS08]. reinforced [dCDLN^{+}07]. Reinforcement [KX08]. reinsurance [CZ07b]. Reissner [WH06a]. rejuvenation [KPG07]. related [ABM09, CGVC07, CL05d, FZ08, KK06, Kim07b, KVŽŽ08, MD09c, Muk08b, Noo08a, NB09, Nwo06c, Nwo06d, SSMI09, WG05, WSL08]. relating [Dah07, Wu07d]. relation [GK06b, GK06a, II07, KG06, MJ07c]. Relational [Ino08, Hu07a, WC05d]. relations [BB06c, BMY08, FA05a, Ish09, KH08, KI08, MB05a, SANN07, SV07, WF07a, lY08b]. relationship [Cha05g, Cha05h, Imr06, LS08e, Sla06]. relationships [Bay05b, MNM07, RAG^{+}07]. Relative [SDM^{+}09, ESETA^{+}06, SML05b, SGUK08, WS07f]. Relativistic [ND06, BT09b, DK06, Dön06b, EEAES06]. relativity [KK07e, Köp07c]. relativized [Zie09]. relaxation [All05d, BST06, BS06, DG07a, DKA06a, EGE06, EMY05, Haj06, QZJ07a, QZJ07b, SHK06, SD06, WB06, WHL08, ZZ07e, ZH08c]. relaxation-based [SD06]. Relaxed [AZ09, AV09, HG09, JZHT05, JZTH06, JXH06, bJyKxC09, yLcCsL09, YNCL08, YCYL05, YK06b, ZHGG08]. release [Ahm07b, CH05a, Ere09a]. relevance [Hub09b]. Reliability [AKKS06, XZ07a, XHSL09, AÖ09b, Che06l, ESES06, GGOS08, IR06, LYGL07a, MR06a, PBG07, RS06, WDK06]. reliabilty [MYSA05]. reliable [EM08, HAK07b, KL06b, KSP08, KA06f, LYL^{+}08, Odi06c, Pou06, Waz05b, Waz05a, Waz05o, Waz05l, Waz07j, XW09, Yu05, Waz06m]. relinking [RKS08]. reload [DN07c]. remainder [CL05a, Li06d, QLG08, Sun07a]. Remark [Lay08, SC05b, Sun07b]. Remarks [CLH05, DC06c, HLSS10, Jwo07, Lei09, LD07a, Mar14, MKSK07, SCS05c, ZLW09a, jW06b, Wu07e]. remote [Hsu05b, LKY05a, SZS05, WLT05b]. removal [NSS06b, TTC08]. removals [WWC06]. Removing [Gra09]. renewal [hBxY07, WS05a]. repair [JW09a, WC09b]. repairable [AGT07, HLK06, JM06a, KC07b]. repairs [GI06]. repeated [AR05, Cha05e, SNH07c]. Repel [LKP06]. Replacement [CSBS07, JW09a]. replenishment [CHW07]. Reply [Sub07]. repository [RV09]. Representation [CD06, DC06e, YMY05, KS08c, KH08, LYW09, ÖTA05, TD05]. Representations [CX08a, AR06, HT06a, Sal06f]. representing [DT08]. represents [DD06a]. reproducing [CjC08, DC06e, GC07b, Gen09a, Gen09b, LZ09e, MD08c, YC06a]. repudiation [SC05d]. required [Wan05f]. Research [WH06a, WC06c, ZHX06, DN07c, JXW05, MJ07b, Tie05, XCJ09]. Researches [XC06a]. resembling [DESG08]. Reserve [Chi07i, AEFFA08]. reservoir [Mai05, YZD08]. reset [Rom09a, SYL07b]. residual [Çel06b, Dje09, xGyZtZ^{+}07, GL05c, HYG07, IM08, KTK06, Lee05b, LpL07, Lin06d, SA06h, Wan06b, WLK08, WXW09]. residuals [ANA07, Çel05a]. Residue [KJ09, CL06a]. residues [QC05]. Resilient [MI08]. resistance [BA07, BAB08, GPT09, LYCW09, YC06c]. resistant [Kwo07a]. resolution [AH08a, Ben07a, KK06, KK08d, Lin05b, PLZ05, XLHK09, ZCZ08]. resolvent [Kum08, NH07a]. resolving [STU09]. resonance [DZZ07, DH09, Li08c, LM09, MD09a, PLZ05, WP05, XHSL06]. resonant [ZLW09b]. resource [ASA06a, ATM06, CK08, CCL06, Dam08, FSMC05, JDSR08, KPG07, LW06b, LG07, LPW08, MKG09, NN06a, OASM05a, RK07a, RKS08, Ran08c, WL05a, YW06a, YW06b]. resource-constrained [ASA06a, JDSR08, Ran08c]. resources [AS06f, CPSS08a, CPSS08b, Yil06, ZGL08]. respect [Bil04, Bil07, HED06b, Kim07a, KI08, YL07]. respiratory [NMS07]. responding [QSS08]. Response [ORG^{+}09, AABD07, CS07f, CK06b, CK07b, DK05, HCW06, HCZ06, Kök08, LX09a, LH07a, LC09b, PN06, She07a, WZ06b, WFL06, WW05b, XCD05b, XW07b, YCZ06, YC06b, YZLD08, ZCZ07, ZSZS09, dG06]. Responses [YXSX05, PMS09]. restart [ZW06e]. Restarted [GJB05, CN06, Lin05d, TGSNB07]. restarting [NG06a, NG06b, NR07b, NM08a]. resting [XCW08]. restoration [Lan07, NK05b, SC06a]. Restricted [CT07b, CLWP09, GW06a, JLCH08, LP08a, WX05, WW05d, WYL08b]. restriction [ZHZ08]. restrictions [AA07b, BA05a, LJE07a, ZLW06a]. restrictive [GÖ05, Ism05, IRSR05]. Restudy [FWL08]. result [AO05a, BBB08, Deo08, DLT08, Par06a]. resultant [WL05b]. resulting [LHSB09]. results [AK05a, AR09, AK06d, AK06e, AS08, BIK05, BR09c, Bou06b, CH07a, CSV07, CKKS08, Dön06b, gDH07, EM09, Ery06, GZ07c, GMCM06, GS08c, HW09, Kay05c, KV06, LMZ09, LDL08, LX07b, Liu08a, MN05, MM07e, Özk07, PK09, PH09, RB08b, SA09, SH08b, SFG07, SM05d, SM06e, WT09, Yan08b, YCL09b, YY06b, YY07b, Zha08g, ZQJ09]. retailing [Nwo06f]. retardation [AS08]. retarded [ADZ05, BM06a, BUÇ07, DB09, GF06, LMY07, PK05b, QBH06a, QBH06b, Wan07e, XS06, ZM08b, ZM08a]. retina [AT07b]. RETRACTED [EHZES05]. Retraction [Ano07c]. retrial [AA07d, AA05, AC06, AM05, Cho07c, CD09, GI06, Wen05]. retrieval [WS05b, ZCZX07]. returns [AK05f, JSdRm05, LJE07a]. returns-to-scale [JSdRm05]. revenue [JLM05]. reversible [CHY^{+}09, FF06]. reverting [WMC08]. Review [Fen05, DCR07]. Revised [JDG06b, JDG06c, EHN07, HMW07, KT07c]. revisions [LJ07]. revisited [CSBS07, CS07d, SK09b]. revocation [LHH05b]. Reynolds [Ere06, NCGdSM07, ÖA05]. Rezaee [Kop22]. Rezaifar [Kop22]. RFIMiner [JWL^{+}07]. Ribière [YGW09]. Riccati [Abb06a, Abb06b, AHIP07, BUP07, BSKK06, BK08b, BLW06b, GS06d, HIPA08, Iva08, Kos07, LZ08c, LN06, MS06c, Muk07c, Muk07b, Muk08b, SZ07c, TL07, ZZ07a, ZZYZ06a, ZZZ07]. Rich [PMS09]. Richardson [MP08c]. Riemann [MBS07, Yan07d]. Riesz [SY09]. right [BW05b, GY07b, HWC07, KT06, LHZ06, LHZ09, SCS06, TK06a, WWT05]. right-hand [GY07b, KT06, TK06a]. right-to-left [SCS06]. righthand [mH07b]. rigid [Att06b, CFJ^{+}05, KSKK06, Kös06a, Shi09a]. Rijndael [JmBdXgXm05]. rill [MDZ08]. ring [HH09b, Su07b, Su08a]. Risk [Haa09, hB06, hBxY07, CZ07b, CG06, GY08, LS09a, Liu08a, Nwo06a, Nwo06g, Nwo07a, PB07a, RAG^{+}07]. risky [TG06b]. river [MMT07]. rivers [KM06]. Rivlin [GS08b]. RKMK [ZD06a]. RLS [DZ08e, Nak08]. RLW [Ras05b, FLLH08, Ram06c, SH05b].

RNS [YCW05]. road [Chi07g, Chi07i, Chi08a, Chi08c]. Robbins [HN05]. Robertson [NB09]. Robin [GM06a]. robot [LWL07b, SH07f, TKJ08, ZS08b]. robotic [SJ08a, SJ08b]. Robs [Geo07]. Robust [ACP07, CC07b, JPYW09, Kos06b, KP06b, KP06c, LCC05a, LP08a, LYL^{+}08, LLC06b, LZWK09, Par09b, SFFY08, XW09, dlS06b, dlS06c, AZ09, BI08b, ÇE06b, CS07a, Dan08, EH08b, GH09, Ibr08, Kök06, Kök08, KPL08a, KP08b, KPL08b, LLZ08, LCZ05d, LY08d, Par05b, PK05d, QCSF09, SJ08a, SDM08, Sin07b, Sin08b, Sin09, SWX06b, SWX07b, STU09, WLP07, YY05c, lZ06d, ZWY07, ZZW09b, dlS06a, ND07, SJ08b]. Robustness [YCC08, ASA06a]. rock [HBBA07]. rock-fill [HBBA07]. rocket [HJB06]. rod [LZ08b, TD09]. Rogers [Cao09]. Role [GL06b, JY07b, AN06, EP07, KA06f, PE07a]. Romberg [All05c, MO07]. roofs [dCDNPN09]. root [CH07e, DZ06, Fuk06a, GDB06, HCS05, Jön05, LHSB09, Nak08, PB07b, Pet07, PP07c, Pet08b, PY08b, WCX06, WMC08, XW06]. root-finders [PP07c]. root-finding [CH07e, GDB06]. root-solver [Pet08b]. Rootfinding [BV06]. roots [Bas08, Boy06a, CN09b, Fra09b, HR07b, HHK^{+}05, LLC09a, LCC06, LZR07, NP05, Net08a, NBO09, Özb06, PY07, PG08b, SXL09a]. Rose [WC06b, WS09a]. Rosenau [HXH08, OAAK08, Zuo09]. Rosenbrock [ZXDL05]. Ross [SLO08]. Rossby [SMQ06]. rotamer [SCR08a]. Rotated [AL08, AC07a]. Rotating [AGH05, AHM09a, Att05a, Dön06a, Gar07, LC05b, Lin08, SM08d]. rotation [GS08b, Ram06a, SK07a, SK08c, Su07b, SSK06, SSS06, VG07, Vic09, ZZW08]. rotational [TY07]. rotations [YY06c]. Rothe [Ram07e, ZZZ09]. Rouche [DD09]. rough [BK06b, LY09c]. roughness [BK06a]. roulette [ZLH^{+}09]. round [SW05b]. route [TMSG06, WS07c]. Routh [lS07b]. routine [SE07]. routing [Dah07, GTA06, OASM05b, She05b, She06, TMSG06, TMSZ06, ZL06b]. roving [LY09c]. Row [CY06b, FSÇ06, Rim05a, Rim05d, Rim06f, Tia09]. Row-wise [CY06b]. rows [LR07a, LR07b]. RRP [SJ08a, SJ08b]. RSA [BLH06, CKY05, WLHH05, WY05d, XC05]. RSA-based [WLHH05, WY05d]. RTS [JLZ05a]. ruin [hB06, KCK08]. rule [AA07e, BDDB08, BDE06, CL06b, DMJE05c, EdR07, GW06a, HED06b, HMJED06, HW07a, IT06b, KKY07b, MJHED05, NR08a, SS05c, SS05d, TRLD08, UE07, WX05, WSXL09]. Ruled [AAABH04, KAK05, KAK06, Kas06b]. rules [AM08b, AG08, BJE05, DMJE05a, DMJE05b, DMJE06a, DMJE06b, EdR07, EDMJ05b, EDMJ05c, EDMJ05a, EMJB05, Has06a, HED06a, HEDMJ06, HED06c, HMJD06, Hu07b, LT08a, LC07, ML05, MJED05, MJHED06, MJ07c, MJD07, RVN07, RNV07b, RVNI07, Tan08a, TRLD08, XG08, yXhWZ06]. rulings [AAABH04, KAK05, KAK06, Kas06b]. run [Ery06, Liu07d, MJ07e, PVM07]. Runge [BX09, AKIS07, AS09a, AFS06, CLS06, Chr08, LL05d, LSW06b, LSW06a, LMY07, LLY07, MKS09, PPF09, Pod06, SJ05, Van05, WWL08a, WYL06b, WWY09, Wu06f, XT06, YLS05, YWL07a, ZDL07, ZXWL06]. rupture [CY07c]. rural [SGA06]. RVT [ETETH07, HS09b]. RWS [CCZS08].

S [EM11, Pan12, JmBdXgXm05]. S-box [JmBdXgXm05]. SA [Sad07]. Saddle [GD06, Cao06, CJ08, Cui08, LD06b, LW06c, LC06b, LH06a, Lu07, PNB06, PW08b, SH05a, ZZ09]. saddle-focus [SH05a]. Saigo [LL05g]. Saka [Pan08c]. Sakaguchi [OSY07]. Sakiadis [Bat08b]. salesman [BK07b]. saliency [SY07b]. Salton [UBKC08]. sample [HWC07, HW05c, NK07b, SC07e, WGF^{+}06, WWT05, WLC06]. samples [WLC05, WLC07b]. sampling [Cha07b, Chi06c, KC06a, Kim05a, Pog07b, SAS07, Thu07d, WGF^{+}06, WL05g, WTG07, WY05c, cWxChY07, YWZW07]. sand [KMK07]. sandwich [SS07b]. SAR [GZL08, QWX08]. SARS [ZLMW05]. SAT [Pal09]. satellite [AAES07, EEAES06, ESS05b, ESAR05, ESETA^{+}06, QC07, Ser09]. satellites [ESS05a]. satisfaction [Sal08d]. satisfying [RAN08a, Sok08b]. Satsuma [Waz09b, CY06a, FqZ08, SXX05, WP06, d’O05a]. Satterthwaite [Tan07b, Tan07d]. saturated [Che06n, HH09a, JK07, ZLL06a]. saturating [GS08b, MR09, SDS05, SM08e, dlS06b]. saturation [JWS^{+}09, SJW09]. Saulyev [LJ07]. Sawada [ZZB08, LD08, Sal08a, Waz08c, Waz08d]. scalar [BK07c, BK07d, DwWmW05, FT07, Kim06b, WLCL06]. Scalarization [Sla06]. Scale [MMC^{+}08, ASA05, ASAEE06, AK05f, ASA07c, BDPP08, DD06a, Ham05c, HWH09, HJ06, Hos08, HH06b, JSdRm05, Lan07, LZ06b, LZ06c, LY07c, LM08b, LB06, LJE07a, Ma08c, Muk07b, Muk08a, Par05c, RX09c, SMY08, WN08a, WLQ06a, WLC06, Wu06d, XW07a, YZW07b, ZŠ05, ZOZZ08]. Scale-free [MMC^{+}08]. scale-incremental [LM08b]. scaled [JLA07]. scales [BW05b, Che08d, EHP08, EBP09, EM09, GZ07c, GO08, HT05, HZ09a, LN08a, MS09a, Mea05, Ngô09, PB09, SS07a, SZ08b, SL06c, ZW09b, ZS08a, ZLR07]. scaling [AMN07g, ANM07g, AKBZ08, HL05a, JZ08b, MB05a, TCMPMC07, WZ09c, cWxChY07, Zhu05a, Zhu06a, Zhu07]. scanning [WC06c, WS07f, WS07g]. scatter [TT07]. scattered [WL06f]. scattering [HS09b, KS08d, Liu05b, SS09a, TA05b, TA06, WM07, ZMM06]. SCD [GY07b]. schedule [AM05, Cho07c, LC06d, MM07c, OT06]. scheduling [ASA06a, AKH06, AASAA06, CCFC08, CSY07, EG07c, EG06b, EG06c, Ere09a, Ere09b, HP07, HYL09, IGM07, JDSR08, JRA^{+}07, KL05a, KO08, KBTS07, LGJ06, LJG06, NJG08, RK07a, RKS08, Ran08c, SK07a, SK08c, TMJV^{+}05, TMMVA05, THLC07, WSJW06, Wan06d, WS07c, Wan08b, XYL10, XY11, ZGH06, lZyT05, ZWM08, ZS07d]. schema [She07c]. scheme [AS06c, ACJ09, BCW05, BCL05, BP09, BGR05, Bra05, CFX05, Cen05, CC05a, CLS05b, CY05a, CHY05a, CY05b, CHC05, CH05b, Che05g, CHT06, Che07h, Che08b, Che08c, Chi06a, Chi07e, Chi07d, Chi07f, CD07, DSMK08, DTC08, DL08, DZ09, EA07a, FWL08, GLV08, GT05b, HHC08, Hsu05a, HW05a, HWH05, Hsu05b, HXH08, HLS05, Hwa05, HL05c, HL05d, JSW05, KSA05, KK06, KA06a, KA06b, KK08b, Kar05a, KRA07, KRY05, KLY07b, KHKL05, LL05b, LKY05a, LKY05d, LL05c, LTH05, Lia08, LYGL07b, LCC05c, LC05d, LHH05b, LCZ05c, LCZ05d, LCZ05b, LWK05b, Ma06b, MW07, Ma08a, Moh07a, Mom05c, Mom06, NCGdSM07, NS06b, OAAK08, ÖÖ06, PVV09, PW05, PC05, PK05a, Psi07, Psi09, qQcM05, QC05, Raj08, Ram07a, Rao05, Rem07, RY05a, SG09a, SC05a, Sha05c, Sha05d, SRH09, SDG05, Sid09, SCS05b, SC05d, SCS05c, SZS05, SW09d, TLH05, TWL05, WZZ06a, WL07f]. scheme [WZC08, WHH05, Wu07c, Xie05, XC05, rYzS07, YRY05a, YY05b, YY05c, YY05d, YC05b, Zah09, Zha05a, ZK05, Zha05b, ZW05a, ZAX05, ZW05c, ZZ05b, ZLZG06, ZL05b, ZZKY05, ZCFX07, ZZ07e, ZCZ08, dRMS05]. schemes [AKO06, AS09a, AL08, AMAP07, AV07b, AY05b, CY08a, CC05b, CHY05b, CLH05, CL09b, CHYC05, Chu07e, GC07a, Gül05, HGH06, HW05b, HN07c, HH05, Ide06, Ism08a, IMS09, JXW05, JC05, JV05, KM06, KJY05, KLY07a, KM05f, KK08d, Kut05, Li05b, LC05a, LJ07, LWG09, LWZH05, Lin05e, LWK05a, LW05c, MY07b, Moh07c, NN06b, Noo07d, Ozb07b, PKH05, QCB05a, QCB05b, Ran08b, SS05a, Sal07b, Sha05b, SYH07, SMH07, SFH09, SCL05b, SMHQ06, Tan05a, Tou09, TNL09, Tsa05, WW05a, cWmZ06, WZC07, WZ08c, Waz06l, XCFX05, YWC05, You08, ZF05, ZW05d, ZD06a, ZCL05]. Schiff [Waz08j]. Schlemm [AS06h]. Scholes [CKK05, CJSSP05, Wu07a]. Schröder [WY07, XB06]. Schrödinger [AC05, BEF08, Che05f, CG07b, GL08c, GC07c, GHX08, Ism08a, KKS06, KLX06, LXML09, qLJ05, LZZZ06, LZZ06b, LZ07c, LHSB09, MC09, MAY09, SKHZ06, SP08, SZ06a, Sim09b, SMHQ06, SMQ06, Van05, Wan05a, cWmZ06, Wan07a, WZ08c, WZC08, Wan08d, XHH06, lYqLW07, YL09a, Zha05a, ZLW09a, ZL07c]. Schroedinger [CTT08]. Schur [CI09, CKKS08, CN09c, Den09a, MSCG09, SC08b, ZW07c]. Schwartz [Tan07c]. Schwarz [Bog05a, DG07a, JZ07, Qia09, YY05a, Yan07c]. Schwarzian [WDLM10, Kan09a, Li09g]. Schwarzschild [Dön06a]. score [JLZ05a]. scoring [HTO06c, NP08]. Scramjet [HBJ06]. screening [Dem05b, TNS07]. Screw [LC06c]. scroll [HHYW07]. SDCP [RX09c]. SDDEs [RY06, RHQ06]. Sea [UBKC08]. seal [Su07a]. search [ASPA09, Alb06, BK07b, Bog05c, Cha07a, Chi08a, DOU08b, EH08b, FZN07, FYZ07, Gee08, GIT07, HK06, Ham08, HWS07, Kah05a, Kah05c, Kah05b, Kah06a, LFM06, xLyW06, LZ06e, LM07a, MC09, MFD07, MKSK07, OM08, Pal07, SH07a, SS05b, SS05c, SS05d, SS06b, SS06d, SYY04, Sub07, TT07, TH07b, jW06b, WS06b, Wu08d, YK03, YGW09, Zho09a]. searching [CGSS09, HL05a, WY05a, WLT05a, XLKH09]. secant [CS07g, CS07j, RW07, WR06, HRE05, Kah06b, KC07a, KSM05, RYW06, Ren06, Ren07, RW08, Ren08b]. Secant-like [HRE05, KSM05, Ren07, Ren08b]. Secant-type [Kah06b]. sech [GA08]. Second [JLW09, Kou07d, WZ08d, AO05a, ASH09, AKIS07, AS05, AuH07, An05b, AE07b, AKVH05, AY05b, AFS06, ADDdM07, AZ09, AR07b, BD05a, BGA05, BMS08, BCEGMRMP07, BTI06, Bie08a, Bie08b, BS07d, BGF09, BKM06, ÇT07a, CM09, CLA08, CM08a, CL08d, Chu07j, Chu07l, CH08b, DBR07, DMJE05b, DM06, DB07d, EM06, ESS05a, ESETA^{+}06, ESE05a, EBP09, FHH09, FE08, FH06a, GC07b, Gen09b, GF06, GG06, GS06b, GS06c, Gül08, HM08a, HMJED06, mH06c, mH07a, HC08a, HR07a, HR09a, Hol07, HN07a, HLW08, HM08b, HMT09a, HMT09b, IHB06, Jan08c, Jan08b, JWPS08, JGZ07, JYX06, KV05a, KW08b, KMÖS09, KM05c, KSS07, KLW06d, KLW06c, KLW07d, KL07f, Kum03, Kum09c, LY07b, Lei09, LS06b, LCJ08, Li08b, LC09a, LLC09a, LS07c, LPS08, LB06, LL08f, LLW08, LKU09, LLW09, LW09e, LSN09, LM07b, MTL07, MM08a, MSK05]. second [MA05d, MK05d, MR06b, MY06b, MAR06, MLR07, MSR07, MSTY06, MJHED06, MP08a, MŘ07b, MM07e, MD09a, MO07, MA05g, Moh05b, Moh07c, NKE05, Nav08a, NLD08, NG07, NKH07, Noo07e, NNuH07, ÖŞ06, PC08, RMA07a, Raf07a, RLZ07, Ram07d, Ram07e, Ram08d, Ren08b, SF07, SNH07b, SNH07c, SMA06, SA06a, SQ11, Sha07d, SL07a, SYH07, SRH09, uIT06, uINTK06, SLW08, SL05b, Sun05b, SS05g, Sun06b, SG07c, SM08b, SM08c, SM09c, SCO09, TF08, TJG08, Wan06h, Wan06j, WSB06, WHL06, WC07a, WML07, WC07b, WL09c, WB09, WZ08e, WS08b, WS09c, WZM06, WX08b, XWZ06, Xia08, XZ08a, XM06, XM07a, XM07b, Xu08b, XT08, YLB06, Yan08e, YX09, ZLW07, ZS08a, ZFYH08, ZM09, ZS09a, ZZC06, ZM06b, ZC08b, ZM08c, Zha09g, ZLC07, Zho09c, ZW05j, ZJO08]. Second-derivative-free [Kou07d, Chu07j, CH08b, KLW06d, LLC09a, Ren08b]. second-grade [AuH07]. second-kind [HM08a]. Second-order [JLW09, AKIS07, AR07b, Bie08a, Bie08b, BS07d, BGF09, ÇT07a, DB07d, EM06, ESS05a, FHH09, FH06a, GC07b, mH07a, HLW08, HM08b, JWPS08, JGZ07, KW08b, KMÖS09, KM05c, KSS07, LY07b, Lei09, LS06b, Li08b, LC09a, LS07c, LB06, LL08f, LLW08, LLW09, LW09e, LM07b, MTL07, MŘ07b, MD09a, Moh07c, NKE05, Nav08a, PC08, Ram07d, Ram07e, Ram08d, SF07, SQ11, Sha07d, uIT06, uINTK06, SL05b, Sun05b, SS05g, SG07c, SM08b, SM09c, WSB06, WC07a, WC07b, WB09, WZ08e, WS09c, WZM06, XT08, YLB06, ZLW07, ZS08a, ZS09a, ZZC06, ZM06b, ZC08b, ZM08c, Zha09g, ZJO08]. second-order-cone [AZ09]. secondary [NSS06b]. secrecy [HLS05]. secret [CC05a, CHY05a, PW05, SC05a, Wan05f, Zha07a, dRMS05]. sector [Cai06a, CLWP09, EARS07, GZ07d, LP08a, Nwo07b]. sector-bounded [GZ07d]. sectors [MNM07]. secular [Vin09]. secure [CHY05b, CHYC05, LCZ05c, YTWY05, ZW05d, ZCL05]. securities [Che07a]. securitization [Nwo07d]. Security [HL05c, HL05d, JSW05, LWK05b, PK05a, QCB05b, RIY05, Sha05d, SL05a, WLT05b, CLS05a, LL05c, LJL06, RY05a, Tsa05, YY05c, ZW05b, ZW05c]. sedenions [Köp07b, Köp07d]. sediments [AKMH06]. Seebeck [GSDB07a]. seeding [PJ10]. segment [Sal07c]. segmentation [BP09, DZZ07, HLCH08, HYZ07, LG06c, LMD^{+}08, MM05c, QWX08, WYZL08]. Segmented [CE07b, CE07a]. segments [ŽMO09]. segregation [BOB05]. SEI [CL08a]. Seidel [yS06c, TD07b, YEA07, Yun07b]. SEIR [GTX08, WTS09]. SEIRS [MCC07, MS07e]. seismic [AABD07, Che07e, RLdLSF07, SA06e]. select [YEN08]. Selecting [BTPGAl06, QL06, KWP07, Sae06b, Sae06c]. selection [BTPGAPRU06, CF06a, ÇE06b, CLW09a, CJK07, Cho07b, EAG08, FJT05, HMJD06, Hua06, HTO06b, KF08a, KF08b, LL06b, Lin06a, LHGT06, yLLjP06, LG06b, Nwo06f, OHT05b, PG07c, Sae06d, Sae07, SLC07, Son07, TKA^{+}08, TA05a, WSXL09, YML07, Yin07, YKJ07, ZLZ06, ZN05, ZLW06b, Zha07e]. Selective [BAB05, AKA07]. Self [HW05b, Kud09, BFR05, BCL05, Ber09, CCH05, ĆC09d, DD08b, dCDNPN09, FB07, GR07a, Han06, KA05a, KK07a, Kas06c, LWK05a, PC05, Pat05b, RK07b, Sha05b, TLH05, Tsa05, YL09b, Yu09, ZBH06]. self-adaptive [FB07, Han06]. self-adjoint [KA05a, KK07a, Pat05b, RK07b]. Self-certified [HW05b, BCL05, CCH05, LWK05a, Sha05b, TLH05, Tsa05]. self-consistent [YL09b, Yu09]. self-organizing [DD08b, ZBH06]. self-pairing [PC05]. Self-similar [Kud09]. self-synchronization [BFR05]. self-weighted [dCDNPN09]. Semi [SI06, XHH06, ADDdM07, BI08b, CY06a, CLW09b, DMJE05c, HIST06, HEDMJ06, KM07, LF08, LLC09b, MJHED06, MCS08, MHA06, ÖA05, PLL08, RB08c, SFA08, Soh06, ST08, TX08, Taw06, sWMX07, Wu08b, YML07, YQ09]. semi-algorithm [ADDdM07]. semi-analytic [CLW09b]. Semi-analytical [SI06, Soh06]. semi-approximate [ÖA05]. semi-definite [BI08b, MHA06]. semi-deformable [KM07]. semi-discrete [SFA08]. semi-discretization [HIST06, LLC09b]. Semi-implicit [XHH06, PLL08, RB08c, sWMX07]. semi-infinite [LF08, YQ09]. semi-integer [MCS08]. semi-linear [ST08]. semi-local [Wu08b]. semi-Markov [TX08]. semi-monotone [Taw06]. semi-open [DMJE05c, HEDMJ06, MJHED06]. semi-rational [CY06a]. semi-variance [YML07]. Semiao [Pet07]. semiconductor [KO06, LT09]. Semiconductor-Intellectual-Property [LT09]. Semiconvergence [CT05c, HS07b, CS08a, CC06c]. semiconvergent [CT05c]. semidefinite [Hua05b, Zha08a]. semidifferential [Raw06b]. semigroups [sCLC09, SS07d]. semilinear [Afr05, AK06c, AMN07f, BP06a, Bog08, CP08, CC06b, KV05b, hLqSlZ09, WR07a, WLW09, fW07b, Xia06, XC06b, XC07, Yan06b, ZZYY08, ZM08d, ZMY08, dAC09]. Semilocal [XL09c, BRW08b, EHR09, RYW06]. semipositone [AMN07b, ANM07e, AMN07e, ANM07h, LZW07, ZLW07, Zha08f, ZJO08]. semisimple [Sht07]. semismooth [AB06, Śmi07]. Sen [Tan06b]. sensing [Ma08c]. sensitive [Mor09]. sensitivities [SH07e]. Sensitivity [DW09c, JLS^{+}05c, JHS^{+}05, JK08, KKKM09, LJE07b, AS09a, Ars07c, ALN06, CH07b, CK06d, FA05b, FK09, GM06a, GS06e, JLS^{+}05b, JW09b, SGB05]. sensor [SJ08b]. sentence [LLX07a]. Separable [Dem06, Hua09, Via08, WCO09, AOS09]. separated [LA07b, Waz06o, Waz06p, Waz06q]. Separation [MA05g, CQL08, FTMA07, GJ05, KR07c, KR07d, SS09d]. septic [SH05b]. Sequence [Sad07, AMJD06b, ABM09, ÁCTZ05, Bil04, Bil07, CC09b, DD05, DDR06a, DD06b, DMS06c, DDR06c, EG06c, Ery06, GÇ05, LC06d, MS08c, Oss05a, OOB07, Sav07, SCR08c, YSC^{+}05, ZGH06]. sequence-dependent [EG06c, ZGH06]. sequences [Ben07b, CL07a, DMS09, Djo06, DM08c, DDR05a, DDR06b, Hu07a, JZCW05, MS07c, PG07c, Thu07a, Thu07b]. sequencing [KOP07a, TMMVA06, TMRV06]. Sequential [GJL06, CLZ05, YQYL07]. sequestration [WLX05, WWLX06]. Series [FA06, Pam06, YC09b, AES08, AGPJV09, AGS07, Bar05b, Bel06, BS09c, Bor09, Boy07a, Boy06a, Boy07b, Bry07b, BAK07, Çel06b, CW06d, CS05c, DJM08, DEH05c, EGS05, EM05b, GW07, JZW08, JV07, KMVL09, KM05b, Kıy06, Lal09, LL06b, MLW09, MAD08, MA05d, MA06c, MAR06, MSM07b, OKEL05, PST06, RS06, SBS08, SH07b, SP08, Sin07b, Sin07c, Sri09, Thu07c, Van09, Wan05e, WN08c, WN09b, WZ06g, XGTS07, Yao08a, ZCMQ07]. server [AA05, AM05, KL06a, KL06b, LKY05b, MAR05a, TWL05, WWP05, WY09, Wen05]. servers [GI06]. service [BGP06, CD09, GCSS05, GMS06, KWP07, LWH06a, MX08b, Nwo06b, TS06, XM08]. services [LT09, TWL05]. services/capabilities [LT09]. servoing [SJ08b]. Set [BD05b, Lak05, Che07h, DMZ08, JMLR05, JjQjWmL06, KB05a, yLcCsL09, LZ06a, Lin06b, LY09c, SS07d, WZSL08, XW06, XRZ09, XW08, YZY08, You05, ZZL05, ZXL09, gZW08]. set-valued [KB05a, yLcCsL09]. sets [CMV05, CL07b, Cha07e, DESG08, FA05a, SS06a, SGUK08, SS07d, Utk06, WL06d, WC06c, yWjCnG07, WY07, Wu07a, XQ06a, Xin06, XB06, ZWD07]. setting [AK05d, RSDC08]. settings [Chi07e, Chi07f]. setup [EG07c, EG06b, EG06c, LC06d, XY11, ZGH06]. seven [ZL06a]. seventh [BRW08a, CPR07, GA08, Par09a, Waz06e, Waz08d, Waz08n]. seventh-order [BRW08a, GA08, Par09a, Waz06e, Waz08d, Waz08n]. Several [SS07a, SS09c, Tsa05, CT06, RZ08, TK06a, WC09a, Waz06a, YPZ05, ZY08c]. sex [RS08b, XZ07b]. sex-structure [XZ07b]. Sextic [KKA06, RJMG07, SAM07, uIK06]. SFEM [ETETH05]. SGS [KL07h]. Shadow [AK08b]. Shaffer [NMM09]. shallow [ADH06, CW06c, RAP05, Sha07a, Tou09, WCLZ05, Waz08e, Waz08p, Waz08u, Waz09d, XCFX05, ZS09b]. Shamanskii [yWyCqD06]. Shamir [PW05]. Shannon [ML06]. Shape [BCR09, CMGMD06, DMZ08, GM06a, lCS06, DWZ07, Su07d, Su08a, WY05c, WLC06, WLC07d, YM06]. Shape-topology [DMZ08]. shaped [NPR09, WWT05]. shared [AS06f, BCW05, HL05c, Kay05a, KW05b, KW05a, Kay05d, Kay05e, SZL^{+}07]. shares [Sal07c]. sharing [CC05a, CHY05a, PW05, SC05a, dRMS05]. Sharma [SQHY08, Waz07g]. Sharp [LZZZ06, SZ06a, ZL07c]. Sharpe [BTPGAl06]. sharpened [Wu08c]. Sheaf [Şen07]. shear [LD06a]. sheet [Ari07, Cor05a, Cor07, dCDLN^{+}07, EEEE05, KBKV09, MS09b, Oua05, Pan12, SH07b, VC06, YC09a]. shell [Gar07, Lee05b]. shellfish [BDPP08]. shells [Fus08]. Sherman [BK09, Yin09]. Shieh [YWC05]. shift [AF07, AD09b, Che08b, HYZ07, NS05, Ye06]. shift-and-invert [Ye06]. shifted [GAV08, PR07]. Shifting [NS05, Kos06b, Sil07a, WGCH05]. ships [dSF05]. Shishkin [KR07b]. Shiu [hBxY07]. Shock [She08, CGJ08, EMY05, KK08d]. shocks [EGS05]. Shooting [Yan06b, Asa06b]. shop [AASAA06, CSY07, HP07, HYL09, LJG06, NJG08, TMJV^{+}05, THLC07, WNCL06, WSJW06, Wan08b, ZS07d]. shops [AKH06, ZGH06]. short [BTMB09, Bri08, Liu07d, PVM07, Rom09a, Sal09]. short-run [Liu07d, PVM07]. shortest [Aza06, CK06c, EJ07b, Ji05, Moa06, Oht08, PSS07, WXLH06, XLH^{+}07, ZFN07]. shot [CH05a]. shot-noise [CH05a]. shoulder [PB07a]. shrinkage [AÖ05b]. shrinking [Ree07, TM08, YC09a]. shunting [WL09a]. SI [JC06, JCL08]. sic [All05b]. side [mH07b, Lee05b, SAB08, VFF08b]. sided [gLL07]. Sidel [AAAA06]. sides [GY07b, KT06, TK06a]. sideways [FXLZ05, XF08a]. Sigmoid [PY08b]. Sigmoid-like [PY08b]. sign [ANM08, AOLL05, CC08a, Hus09, JFG08b, LOA05, MM08a, Muk07c, Muk07b, NN08, SLS08, Tzo08, XL08b, ZYW08]. sign-changing [ANM08, MM08a]. sign-indefinite [Muk07c, Muk07b]. sign-mediated [NN08]. signal [AD06a, Chi07b, Chi07e, Chi07f, Chi07g, Chi07i, DFGV07, FAQ08, GH07, JJ09a, LMT^{+}09, Lin05a, LL05h, XHSL06]. signal-controlled [Chi07g, Chi07i]. signaling [AH06c]. signals [ML08a, MUZG09, NCÁHCLP05a, NCÁHC^{+}06, PG08c, Sel05, XCJ09]. Signature [Köp07d, BCW05, BCL05, CCH05, CLS05b, CY05a, CHC05, CHY05b, CLH05, Che05g, Chi05, CHYC05, GT05a, GT05b, HW05a, HWH05, HW05b, HH05, LHY05, LL05c, LC05a, LTH05, LWZH05, LC05d, LHH05b, LCZ05c, LCZ05d, LCZ05b, LW05c, PKH05, PC05, PK05a, QCB05a, QCB05b, QC05, Sha05b, Sha05d, SCL05b, WLHH05, WHH05, WW05c, WY05d, Xie05, XC05, YC05b, ZSNS05, ZF05, ZW05a, ZW05c, ZCL05]. signatures [BLH06, PLJ05a, Qiu05]. signcryption [HLS05]. signed [LLWC07, YGL05]. signed-digit [YGL05]. signed-digit-recoding [LLWC07]. signer [LL05c]. signers [HW05a, Xie05]. significance [Hub09a]. significant [MA06e, Moh07c, RMG07b]. signing [Chi05]. signomial [Cha05d, QZJ07b]. sike [Ahm06]. similar [Che08c, Kud09, LGJ06, LJG06, SH07b]. Similarity [CIS09, MMT07, SS08a, DESK06, DES07b, DCS07, Hub08b, Ola06, ÖÖ06, PPS09, lYqL06, lYpZqL08]. similarity-transformation-iterative [ÖÖ06]. Simple [LJ07, Oss05b, BLW06b, Bas08, CS07c, DMA07b, GKR06, Geu07, Ish09, KC06a, LL08a, LCL06, LC06c, OHJ07, ÖÖ06, PG06, RLY06, SH06a, Sin07c, YLC09, YC09c, YSCC05, Zhu06c]. Simpler [AKR08]. simplest [KL08]. simplex [ACD^{+}05, AGJ07, Nab09, Pan08a, YC09c]. simplexes [Lin07c]. simplices [KES09]. simplification [Hol07, NTZ08]. Simplified [YZFL08, GOMA06, LMD^{+}08, Ony05b]. simulated [AJAH06, Alr05, AD09a, Cai06b, GAFK07, GÜ07b, JJT06, KC06b, LPA^{+}09, SAS07, SK07a, TMSG06, WZ06e, CK07b]. simulating [CZ05b, WsZ07a]. Simulation [CN06, DK06, SMQ06, TCMP07, ZS08b, AD09a, AD06c, AGI08, BFR05, BST06, BG06, BNH07, ÇE06b, CLZ^{+}07, Che09a, CK06b, CLNCG07, DSMK08, Dam07, DB06, Dön06a, EKS08, FEOC08, Fen05, FJCW05, GRF07, HDG07, Jwo07, KMK07, KLX06, Li07c, Li09e, LLLW06, Liu06b, LWL06, LX06a, PN06, RAG^{+}07, dCRDSSN^{+}08, SW06c, TG06a, TCMPMC07, Wan05d, WC05c, XL05, XL06b, XZB06, XM08, Zhe08, ZCFX07, Fuk06a]. simulation-based [Li09e, Fuk06a]. Simulations [CZZ05, BBCP09, Dem05a, HAK07a, PRGH05, Tan05b, WMC08, ZL08c]. Simultaneous [YC06c, BK07d, Fat05a, Fat05b, GI08, GLMLP09, LX07a, LH07b, NP05, PR05, PHP09, ZPH06, ZQ06, ZQ08, ZM08d]. simultaneously [Zhu05d]. Sinc [EG07b, RZ05, RZ07b, Wan08c]. sine [BQYW07, CY08c, GHL07, KGB06, Ma07c, Ray06, RV09, TB09a, Wan05e, Waz05n, Waz05m, Waz06m, Waz06p, Waz08w, Yüc08, Hub09d, Sil06e, KH09, Wan06e, Waz06o]. sine- [GHL07]. Sine-and-cosine [Sil06e]. sine-Gordon [BQYW07, Ma07c, Ray06, Wan05e, Waz05n, Yüc08, Wan06e]. sine-Gordon-type [CY08c]. Single [PCS07, Sal07b, Wan06d, Waz08t, XHLL05, lZyT05, AM05, BW07, BTPGAl06, CCL06, Che06b, CX06, DDZ08, EG07c, Ere09a, FT07, xGZZ08, GS06f, JOA05, JRA^{+}07, KK08b, LS06a, LXZ09, LC06d, MAR05a, MPG07, MM07c, Mon07, Pam05b, RSDC08, TMMVA05, TMMVA06, WS09a, Wen05, XL08a, YGY09, ZCZ05, lS08a]. single-index [BTPGAl06]. single-input [lS08a]. single-output [DDZ08, LXZ09, lS08a]. single-server [AM05, Wen05]. single-species [BW07, MM07c]. Singular [LP06b, LP06a, O’R08, AOLL05, AC07b, AD09b, AKD06, AM09a, AR07a, Att06c, AR07b, BQ09, BSKK06, BK08b, BM08, BS07b, Buk09, CC06a, Cao06, CS08a, Cen06, CPR07, Cha07b, CC07b, CX08b, CLW09b, Cui05, CZ07c, CY07c, DQW06, DEFT05, DM08b, DZ07b, DZ07a, DZ08f, EGX05, GC07b, Gen09b, GK09, GWX06, GW07, HEZ07, HEZ08, mH07b, HRT07, HLW08, HS07b, Ism06, JC08a, JLW09, KA05b, KP06a, KK07b, KK07a, KW08b, KWX08, KR05, KB06a, Kan07, KPS06b, Kha08b, KM05c, Kum03, KA06e, Kum07, KB07, Kum09c, LCM08, LLX05, LX07b, LZW07, LLW07, LLW08, LLW09, LOA05, LC08b, LP09, MZW08, MA06c, MS08a, MN05, MJ05, MA06e, MA06f, Moh07a, Muk07a, Pat05b, Pat07, zPY06, Ram05f, RMJ07, RMG07a, Sal08c, SB06b, SBA09, Son05, SWX06a, SWZL07, Su08c, SLS08, SCO09, TL08, Wan05h]. singular [WG05, WJ05, WR07a, WCH08a, WLW08, Waz08f, Waz08h, Waz08i, Waz08s, Waz08v, Waz09a, Waz09c, Waz09d, Waz09e, WWW03, WP05, WDN06, Wei06a, Wei06b, WZ08e, WC08b, yXhWZ06, XZ09a, XL08b, Xu08a, XZ09b, Yam08, YC07a, YJO08, ZLZ07, ZCMQ07, ZYW08, ZLW08, ZLW09b, ZnCxX09, ZW05f, ZW06e, ZW06f, ZXZZ07, ZYZ07, ZJO08, dlS07a, lS08a, Sid06]. singularities [EGX05, Fro06, JZ08a, LZ08d, Yu08a, ZM08d]. singularity [ANK05, BBES07, QBH06a, QBH06b, ZMY08]. Singularly [Du07, Ami05, AŞ06e, AAK07, BR07a, Baw05, Bie08a, Bie08b, BP06a, Bra06, CA05, Cen05, CG05b, CG08b, EG06a, HBVM07, HH08a, KA05a, KA06b, KPS06a, KR07a, KR07b, KY08, KYK08, KK08a, KK08b, KS08a, KS08b, KKA06, LP07c, MJ05, MA06e, MS06b, Moh07b, Muk07c, ND07, PS06b, PD09, Ram05c, Ram05d, Ram05g, Ram05h, Ram05i, Ram06d, Ram07g, RK07b, RMG07b, RGM07, SH06a, SR06, SK06b, SA05b, TiHuI08, XO07]. singularly-perturbed [Ram05d, Ram05g, RGM07]. sinh [TH07a, Waz05n, Waz06p, GHL07, Waz06q]. sinh-Gordon [TH07a, Waz05n, GHL07]. SIR [JL07, JY09, KMVL09, LLG09, Mak07a, MC08, SJC09, ZzLZ08]. SIRS [SXL09b, ZLT09]. SIS [KMVL09]. Sisko [AHM09a, AVM^{+}09, SH08a]. Site [Nwo06f]. Sivashinsky [Mom05c, Omr07a, UHuI09, Waz06j]. six [MB05a, WCN08]. six-dimensional [MB05a]. Sixth [KW07, AS06b, AHR08, Chu07i, CH07e, GSDB09, Kou07c, KL07e, Kou07e, LA07a, MA06f, PG08a, SG09a, SG07b, SAN07, uITiHK08, WKL08, WL09b, Waz08b, Waz08o, Waz09b]. Sixth-order [KW07, Chu07i, CH07e, Kou07c, KL07e, Kou07e, SG09a, SAN07, uITiHK08, WKL08, WL09b, Waz08b, Waz08o, Waz09b]. size [HEZ08, JTMRS06, KA06d, LS06a, LS08c, MR06c, MA06d, MB05a, MBM07, NK07b, PFA07, SS06d, TCMPMC07, TCMP07, WGF^{+}06, WF07d, YMCL06]. sizing [AKH06]. SK [FL09a]. Skan [Yan08b, YC09b, ZC09, ZWC09]. sketch [ZS07c]. skew [BZZ08b, CS07i, GG08c, LHZ06, Rim06a, Rim06b, Rim06e, Rim07c, Rim08b, Rim09b, WCL08]. skew-centrosymmetric [LHZ06]. skew-symmetric [CS07i, GG08c, Rim06a, Rim06b, Rim06e, Rim07c]. skin [BP09, XWSL08]. slack [AK06h]. slack-based [AK06h]. slackness [HGH06]. slacks [Sae05]. slacks-based [Sae05]. SLAE [VV06]. slant [KY05b]. slanting [KOR08]. slave [GZG09]. slider [YB05]. sliding [Shi09a]. slightly [SML05b, Sae06b]. slip [LL08e, Ram07b, WY06]. slope [Boy06b, CLWP09, LP08a, Sch08]. Slot [MFE05]. slow [CGJ08, dSYGM^{+}05]. slow-binding [dSYGM^{+}05]. slowly [WN09b]. small [Bah05, BKZ06, Ere06, mHzC05, IDS06, KK08a, Muk07a, PS06b, RAES05]. smallest [NR07a, NRA07, PL06a, Yam08]. Smart [LY09c, Hsu05b, LKY05a, Moh07b]. smart-BLAGE [Moh07b]. smoking [SG08b]. Smooth [SX06, AM09a, BGRS09, FX09, GHX08, Ham05a, HD07, JK08, KKKM09, Ram05i, Ram08a, Sid05, Sid06, Zhu05b, Sid05]. smoother [MSM08, NHCLP07, Nak08]. smoothers [NCÁHCLP05b]. Smoothing [HW06, Sid09, FHH09, FHSS08, HCLPJL^{+}08, KNM05, MK05a, Ma05a, MCW08, MT08, NGLHCLP06, RM08, TLM09, WZ09a, Yan07a, Yan07e, lZZ06a, ZWG09, ZW08b]. Smoothing-type [HW06, WZ09a, ZW08b]. smoothness [AM09a]. snake [Che05i]. snake-like [Che05i]. SNP [QYYL08]. Sobolev [GC07c, MW09, SY08b, SM09b]. social [Sla06, Tan06d, Tan07c, Tan07b, Tan07d, Tan08a]. society [LWZH05, Tan07b, Tan07d, Tan08a]. Soft [ZYD07, KWP07, VRG06, YT06]. Software [CH05a, KPG07, KC07c, LYGL07a, MH06b, MH06a, MJ07e, VGMV^{+}07, Zot07a]. soil [Cel06a, CFJ^{+}05, GJD^{+}07, HKS09b]. soil-column [GJD^{+}07]. Sokolov [Inc06b, Waz09b, WLS09]. solar [EEAES06]. solid [CLSZ06, HJB06, Hlo05, LLW05, Liu06d, SKS06, ST06c, Sin08a, TS05]. solidification [MM08c]. Solitary [Hub05, Waz06n, Waz08u, YL09a, DW09a, HZ08, Hub08b, Inc07a, IRSR05, KY05a, KRA07, Sha06a, Sha06b, SQHY08, Waz06c, Waz06j, Waz06l, Waz07a, Waz07f, WF07c, XMGW09, ZLYL08]. Soliton [BZZ08a, Bis08, BR09b, EWAH08, Fen06, G07, Hon08b, Inc06a, LD08, LHSB09, qMqZ05, ROK09, RKOK09, SC08a, TW09, Waz07e, Waz07d, Waz08b, Waz08c, Waz08d, Waz08e, Waz08f, Waz08k, Waz08n, Waz08m, Waz08p, Waz08l, Waz08o, Waz08j, Waz08i, Waz08s, Waz08t, Waz08q, Waz09b, Waz09a, Waz09c, Waz09d, Waz09e, XZ05, lYpZqL08, YZ07d, Yu08b, YL09b, Zha07d, Zha08d, ZnCxX09, ZLYL08]. soliton-like [Hon08b, LHSB09, qMqZ05, XZ05, ZLYL08]. Solitonic [MGG^{+}09]. Solitons [Waz08v, Zuo09, BM09, HX07, JyGcL^{+}09, KBMZ08, Li09g, Ma07c, SB09, SMQ06, WDLM10, Waz05g, Waz05i, Waz06a, Waz06g, Waz07b, Waz07h, Waz07g, Waz07i, Waz08w]. solute [LLFM08]. soluted [SDS05, SSS06]. Solution [AS06b, AK06g, AO05d, AO06e, AO06d, ADK05, BSKK06, BK08b, Bia05, Bia06, BKBK06, CC07a, DH06e, EA07b, HS09b, KA09b, KS06b, Ord06b, RGD07, RDGP07, SA06f, SA07b, SA07d, SA07c, SB05b, TL08, ZZDL07, AD05a, AD05b, ABA06, AO06a, AESD05, AE07a, AK06c, AMN07b, ANM07d, ANM07c, AMN07d, ANM07b, AMN07c, ANM07e, AMN07e, AMN07g, ANM07g, ANM08, AA07a, Aks05, AÖÖ06, AÖ06b, Aks06a, ADY06, AKM05, AS06d, AT07a, ACZ05, AL08, AD07, AC07b, AA08, AH07d, Amm05, Ang06, AGPJV09, Ars07b, Ars07d, Asa07b, AEA07, AHS06, AKR05, BGA05, BDE06, BF07b, BK07a, Bad05, BH07, BQ09, BHCC09, Bay05a, Baz08, BTMB09, BMP06, BB06b, BM08, BA05a, BE05, BM05b, BIK05, BG08, Bie08a, Bie08b, BÖ05, BA05b, BB05, BR09b, Bog08]. solution [BGR05, Bra06, BAK07, BK05b, CC06a, CC08b, ÇG05a, Cha05h, CL05c, CS05b, CW05b, Che06a, CDP06, CF06b, CM08b, CjC08, CHH09, CLW09b, CJSSP05, CD06, CD07, DIS05, DX07a, Dao05, DG07a, DE07a, DN07a, DJ06b, DKK06, DH06d, DM06, DHS07, Den08c, DT08, DZ08b, DC07, DC06e, Dyk09, ERAS07, EFE08, EPER07, Eid05, EAEG07, EB05d, EG07b, EHAR05, EL05a, EM11, ESA05, ESS05c, ESK06, ESE05a, Elb05, EP07, Ese05b, FK06a, Fus08, GL08a, GKD07, GS06a, GS06b, GS06c, GJ07g, GS07a, GWX06, Gu07a, GG07, Gül05, GS05, GÖ05, GSG06, Gül06, GS06d, GS07c, GHX08, GB05, GB06a, GB06b, GITV06, HK09, HES05, HAM04, Hlo05, Hol07, HAU05, Hos05b, Hos05c, Hos05d, Hos06a, Hos06c, HW07b, HC08b, HT09a, HTM09, HT09b, HL07d, Inc06a, Inc07a, IRSR05, Ism06, Ism08b]. solution [JRW08, JA08, Jav06a, Jav06b, Jav06d, JG07a, JZ06b, JSL06b, JC08a, KM06, KA05b, KPS06a, KR07a, KKT05, KGB06, KAS06a, KK07c, KPS06b, KR09, KS08d, KE05a, KKA06, KMVL09, KSM09, KJ08, Kök08, Kol06, KeTL08, Kum09a, KE05c, Lee05b, Lee06b, Lep07, LLZ06, LT08b, tLlWpL08, Li08c, Li09c, Lin05a, LC09b, LWL06, LW06d, LW07, LGY07, LLY07, LL08f, LA07b, LA07a, MJ07a, Ma08b, MA08d, MNG09, Mah05, MY05, MB06, MA07d, MKA05, MSK05, MK05c, MM05a, MA05d, Mal06, MMN06a, MAM06, MR06b, MA06c, MD06b, MY06b, MAR06, MNY07, MLM07, MS07b, MD07a, MSR07, MMA07, MS08a, MB07, Mar05b, MD06c, MY07a, MM08c, MSR08, MA06f, Mom05b, Mom05c, MO06, Mom06, MMR09, MH08b, MD08c, MTL08, Muk07c, OMST07, OAHR09, Oma06, Ony05b, Oua05]. solution [ÖEK05, Pam06, PE07a, PB09, yP05, PHZ05, hPyHZ06, PHZ06b, PKT06a, PKT06b, PMRAS05a, PMRAS05b, Psi09, PDA06, QD08, qQcM05, RMA07a, ROK09, RAS07, RLZ07, RG05, RZ05, RMJG07, RJM07a, RMG07b, RZ07a, RJM07b, RZ07b, RMJ07, RGM07, RJF07, RJMG07, Ras05c, Raw06a, Raw06b, RB05a, RB05b, RCB06, RB06, Ray06, RCB08, Ray08, Rem07, RX09b, RZ07c, SMY08, SKHZ06, SH07a, SH07b, SD09, Sal07d, SS09b, SK07a, Sev07, SG06, SGT06, SG07a, Sha05a, SH07d, SK06b, SS08b, SXJ05, She07b, SDM08, SZL05, SK05, SP06a, SFPE07, SDR07, SAM07, SA07a, SAN07, SAE08, Sim09b, uIKTT05, uITA06, uIK06, uIT06, uINTK06, uITiHK08, SI06, SH05b, SL06a, SC07d, Sya05a, SA05b, TSD07, TMS08, Tad08, TF08, TL07, TT07, TD06, TD07a, TM08, TQC09]. solution [TiHuI08, TK06b, Tze07, UHuI09, Van05, WY05b, WC05a, Wan05e, Wan06k, WM07, WSL07, WS07e, WS07a, WCN08, WZ08b, WX08a, Wan08e, Wan09c, Waz05c, WRL06, WF07d, WC08b, XO07, XZ07b, XL05, XW08, XMGW09, YS06a, Yan05c, YML07, YLS05, YLB06, YLLL08, Yao08a, YC09a, YC09b, Yi05, Yin08b, YMY05, YY05d, You06b, YB06, You08, Yuj07, (Ag06, YB07, ZO05, ZLW06c, ZS07b, ZL08b, ZYW08, ZLW08, ZZYZ06a, ZH05, ZLL06b, ZLZZ06, ZW09c, ZWC09, ZW05j]. Solutions [DMK07, Jan07b, SA06g, SA07e, SA08e, SZ08b, ZCMQ07, ZLDW09, Abb06d, AMK06, Afr05, ANM06, AK06b, AK06d, ANM07a, AMN07a, AMN07f, ANM07f, ANM07h, ANM07j, ANM07i, AO08, AS06a, AKMA05, AK05c, AKARA08, AM08a, AS05, ASN07, AuH07, AK05h, Amm06, AC08, AR08, AJRLRGOG09, Asl09a, Asl09b, BM06a, BCCK05, Bai08, BZ09, BR09a, BC09b, Bar09, Bel06, Ben07b, BTI06, BQYW07, BKZ06, BG07b, BT08, Bou09a, BJ08, BKP06, CL09a, CCWJ08, CM09, CMC06, CC06b, Cha07e, CLCK09, CS09b, CT05a, CW05a, Che05d, Che05c, Che06d, Che06c, CW06d, CY06a, CG07b, CY08c, CA08a, CL08d, CZ08b, CT09a, CLDH09, Che07i, CKS09, CS09c, CGJ08, Cor05c, CC05d, CY05c, CZ07c, CY07c, CY08d, DSB05, Dai09, DH08, DCS07, Dem05c, Den08d, DSK08, Din05, DLD08, DSZ08, DC06c, DDR05b, DB07d, DZ07b, DZ07a, DZ08f]. solutions [DH09, DHG09, DB09, Efi09, EB08, EG07a, EGFAE06, EOM03, EOM08, ESK05, ET05a, EWAH08, Eng06, EBP09, FXY08, FHYZ09, Fel05, Fen06, FC06, FLLH07b, FHL07, FLLH08, FL09b, GD08, GA08, GL09, GHL07, GL08b, GL08c, GZC08, GK05b, GKB07c, GkB07b, G07, GR06a, Gor06a, GY07a, GO08, GSPG07, HW08b, Han07, Has06c, HY05, lH08a, HRLC08, HRCL08, HC08a, Hon09, Hon08b, HLW08, HYG07, Hub05, Hub08a, Hub08b, HL05b, Inc05, Inc06b, Inc06c, Inc07d, Inc07e, IS09a, IDS06, Ism05, JDG06a, JHST05, Jan07a, Jan08c, Jan08b, Jan09c, Jbi06, JGY07, JG07b, JGZ07, JFG08a, JFG08b, Ji08, JZ08a, JW07, JOA05, JZ06a, JZ08c, Jia09, KW08a, KW08b, KWX08, Kan09b, Kar09c, KY05a, Kay05c, KKA09a, KL06b, Kha07a, Kha07b, KAF08, Kha08b, KuI08, KRRGT07, Kim06a, KL05b, Kıy06, KV05b, Kos07, KT08b, KS09b, Kud09]. solutions [KD09, KM05d, KS08f, Kuo05, KH09, KE05d, Lay09b, Lay09a, Laz08, LL05e, LL05f, LSL06, LS06b, LG06a, Li07e, Li07d, LF08, LZ08b, LZ08c, Li08d, LCJ08, Li09a, Li09g, LZ09a, LZ08d, LPS08, LZ08e, LTW07, Lin07a, LY09b, LXML09, LCL05, qLJ05, LLW05, LTT06, LS07e, LZ07c, LKU07, LY07d, LZW07, LLW07, LY08c, LD08, Liu08b, LLW08, LTCC08, LKU09, Liu09c, Liu09a, LLL09, LLW09, hLqSlZ09, Lou07, LC08a, LOA05, LH08d, LP07c, Ma05c, Ma07b, MTL07, MM08a, MYG08, MC09, MH09, MWQ09, MLW09, MMO07, MM07b, MA08e, MAT08, MO05, ML08b, ML09, MŘ07b, MN05, qMqZ05, MZ09, MGG^{+}09, MY08, MAY09, MM07f, MHM09, MK08a, Mom05a, MAK05, Mus07, MT09, Nav08a, OM06, Ola06, PS08, PPS09, PRM06, PDW06, PFG08, Pao07, PW08a, Par09a, PHZ05, PHZ06b, Pen08, Prá08b, QS09, QZL07b, QW08b, Qui09, RKOK09, REDES05, RLZ08, Ram06g, Ram08b]. solutions [RJK07, RM08, RX09a, Rom09b, RHY06, RY06, RHQ06, RMWk09, Sai07, SG08a, Sal08b, Sal08a, SGC09, Sal08c, SRMRGRL08, SG09a, SY07a, SK08a, SS08a, SPV08, SAD06, Sha08a, STKB05, ST07, Sha06a, Sha06b, Sha07a, SQHY08, SHY08, SLX05a, SXX05, SX06, SMG07, SL07a, SC08a, She08, SL09a, SS09d, Shi09b, SLO08, Soh06, SZ06b, SZ07b, SZ07c, Sta06a, SWX06a, SWZL07, Su08c, SLS08, Su09, SS05f, SS05h, SL06c, SS06g, SHD07, SG07c, SQG08, SMG08, SCO09, Sya05b, SA06h, TL06, TH07a, TH08, TZH09a, TXW09, TZH09b, TB09a, TB09b, TJG08, TW09, TYCL06, WSLZ07, WCLZ05, WLZ05, WZ06b, WCZ06, WiN06, Wan06f, Wan06c, Wan06b, WMX07, sWX07a, WL07a, WZS07, WZ07c, WSZ07b, WSKZ07, Wan07f, WS07b, WS07d, WC07b, WLT07a, WZL08b, WLK08, WLS08, WW08b, WLW08, WCL08, WLH09, WWD09, WLA09, WL09c, WDLM10, Waz05d, Waz05e, Waz05g, Waz05h]. solutions [Waz05i, Waz05j, Waz05n, Waz05m, Waz05q, Waz06a, Waz06c, Waz06f, Waz06e, Waz06g, Waz06j, Waz06l, Waz06n, Waz06o, Waz06p, Waz07b, Waz07a, Waz07c, Waz07h, Waz07e, Waz07d, Waz07g, Waz07f, Waz07i, Waz08b, Waz08c, Waz08d, Waz08e, Waz08f, Waz08g, Waz08h, Waz08k, Waz08n, Waz08m, Waz08p, Waz08l, Waz08o, Waz08j, Waz08i, Waz08r, Waz08s, Waz08t, Waz08q, Waz08u, Waz08w, Waz09b, Waz09a, Waz09c, Waz09d, Waz09e, WP05, Wei06a, Wei06b, WG07b, WZ08e, Wei09, WLS09, WZY07, WDY06, WP06, WCW07b, fW07b, WF07c, WF07b, WY08, WFHD09, WC09c, XZ05, XL08a, XCH05a, XCD05b, XCH05b, Xu07, XGTS07, Xu08a, XL09a, XL09b, XZ09b, XT08, YSS09, YST09, Yan05a, YL06a, lYqL06, Yan08a, lYpZqL08, Yan06d, YCZ06, Yan06b, YXW07, YZ07b, YW08, YWL09, YZ09a, YZ09b, Yax07, Yig08, YY06b, YY07b, YC07b, YY08c, YJO08, Yua08, ZM06a, ZZ08a, ZX06a, ZX06b]. solutions [ZX06c, ZH07a, ZLW07, ZW07b, ZLZ07, ZL07a, ZZB08, ZDY08, Zha08c, Zha08e, ZY08b, ZLY08, ZM09, ZZH09, ZS09a, ZnCxX09, Zha09c, ZXLS06, ZZTJ06, ZX07b, ZWX08, ZC08b, ZYC09, Zha09g, ZY09, ZZ06a, ZSL05, ZL07b, ZYZ07, ZY07, ZLZ08, Zho08b, ZLYL08, Zho09c, ZJO08, Zuo09, dAC09, dSLV^{+}08]. Solvability [BS09a, Bou06b, Dar06, Ibr07, JW09b, Kar09a, LWKU09, MD09a, ZLCZ06, Zha08a, SS07a, SS09c]. solvation [SCR08a, SCR08b]. Solve [ZT06, Zha06b, BKA07, BK09, Bha08, BSKS07, CW06b, Chu07h, DB07a, DB07b, DB07c, FGfFL06, FE07, Gui05, Ham05c, Iva08, KBTS07, LL05e, Mor08b, RAS08, Sal06b, Sha07c, XZB05, YPZ05, ZWL06, ZBH06, ZG08b]. solvents [Slo07]. solver [DJ07, FTT^{+}09, KK05c, LL05a, MMC06, NBHYV09, NLD08, PVN06a, Pet08b, RS08a, YGY09]. solvers [Net08a, Sog08d]. solves [Boy06a]. Solving [AKVH05, BVC05, BKFM05, CW07, CR05a, DGJ07, DGB08, Dön06b, EJER06, EGX05, GC06a, GC07b, Gen09b, GK06a, Han06, HBE09, HRE05, JDG06d, KV05a, KGM06, Lep08, Lep09, LC08b, Mak07b, MFF09, Psi07, RK07a, RKS08, Ran08c, SNH07c, SCM09, Sun05a, TMASA05, Tso09, WX05, WSL09, Wu06c, Wu08d, XWZ06, XZY08, YM07b, ZT05, Zha07g, ZL09b, AEJ06, AE06a, AJ06a, Abd07, AZ08, AAS06, AKIS07, ASNR06, AV06, AC09, AE07b, AE07c, Arg09, AHIP07, ADK06a, ADK06b, AH08c, AR07b, BFR07, BBES07, BF07a, BMS07, BLP07, BAF08, BMS08, BWR09, BG07a, BNK07, BKM06, BB06d, Boy08a, Buk09, CMC06, CC06c, Cho07b, Chu06, Chu07c, Chu07d, Chu07k, Chu07n, Chu08, CPTZ05, CG07c, CG08b, DN07b, DSKM09, Deh05a, Deh05b, DHG06, DMA07a, DBC^{+}07, DF09, EN05, EB05b]. solving [EB05a, EM06, ER06a, EN06, EGN07, EHN07, EJ07a, EJ07b, EM05a, ES05b, ESAH10, ETETH05, ETETH07, EWAE06, EWEA06, EM07, EA08b, Fad06, FE08, FG07, FH06a, FMH07, FH07, FqZ08, Fra09a, GBK06, Gal06a, GXS05, GXY06, GTA06, Gen09a, GOMA06, GTJ07, GG06, GJ07c, GJ07d, GJ07f, Gor06b, Gor07, GW06a, HEZ07, HAK07b, HC07a, HW08a, HKS09a, Has06b, HSH05, HM07, HIPA08, HJ08, HC05a, HD07, HW09, HM06, HW06, Hub09b, IOS07, Ide08b, Ide08a, IIS05, JDG06b, JDG06c, Jai07, JDSR08, Jav06c, JG07a, JJT06, JZ08b, JH08a, JH08b, JL05, JCM05, KTK06, KKT07, KA05a, KSA05, KAV05, KKF07, Kan06b, KSGM06, Kar06, Kar07b, Kar07a, KT06, KSS07, KLW06a, KLW07a, Kou07c, KLW07c, Kou07a, Kou07e, KW07, Kum06, Lar05, LW05b, Li05c, LW06a, LZ06b, LD07b, LYKC07, LH08b]. solving [zLmLL08, LK08, Lie06, LD06b, LW06c, Lin06d, Liu06g, pLT06, LLC09b, LM07a, LMN07, Luo05a, LX06a, LHW07, LHWW07, LL08g, Ma06a, MTC07, Ma08a, MCW08, MMW09, MFD07, Mah09a, MPEL09, MS07a, ML05, MM05b, MK05d, MKR06, MLR07, MKKS05b, MQ07, MP08b, Meh06, Mei06, MO07, MGMC05, MK07b, MAD07, MMN06b, MS06c, MAO06, MN07, MHA06, Muk07a, Muk07b, NKE05, Ned05, NA08, yNyF05, yN05, NA06b, NN07c, NN07d, NN07e, NMD07, OISM07, ÖK05, ÖA05, PP07a, Pod06, PCS07, Qia08, Qua07, RMAM07, RDG07, Raf08, Raj08, RMS05, Rao06a, Rey05a, RX09c, Saa05, SNH07a, SNH07b, SA08a, Sal06a, ST06b, ST06a, SG05a, SG05b, Sha05e, SK06b, SC07a, SDG05, SBA09, Sog08d, Sog08b, Sun06a, SH06c, TZW09, TLM09, TM05, TMJV^{+}05, TS06, Thu08, TK06a, TYCL06, Uje06a, Uje06b, UEL07]. solving [VV06, Wan04, Wan05g, Wan05h, WJ05, Wan06g, WXLH06, WL06c, WL06b, Wan06i, Wan06h, Wan06j, WS07c, WCW07a, WL07c, WC08a, WZ08c, WCH08a, WL09b, Waz06d, Waz06k, Waz06q, Waz07k, Wu05d, WLZ06, Wu06b, WHL08, Wu09, XD07, XFQ06, XF08a, XXX06, YHS09, YHTR05, YC08a, YM08a, YE07, You05, You06a, YHL08c, Yue07, Yun08a, (Ag07, Zah09, ZFN07, ZW05e, ZHLG07, ZX07a, ZDL07, Zha08b, ZC09, ZY08c, ZW05h, ZW06e, ZW06f, ZWZZ07, gZW08, Zho09b, Zhu05b, ZH08d]. SOM [Ohm08]. SOM-Based [Ohm08]. Some [AK06e, BFR05, BS09c, BSHZ07, Che06j, Cho07a, CAS08, Chu07i, Chu07j, CH07e, Chu07k, Chu07l, Chu07m, Chu08, CH08a, CH08b, DSY08, EA08a, Eli07, EM09, Ery06, FB08, GS09, GSDB09, Gup08b, HN07c, Ide06, Irm07, JH08b, Kan06a, KG08, KAP06, KKA09b, KS08e, Kou07e, KL07g, KV06, LAL07, LMZ09, LLC09a, Liu09b, MJ07b, MPG07, MH06b, MH06a, MN05, MZ07, MM07f, NK07a, Ngô09, NL05, NN06b, Noo07e, NH07a, Nwo07c, Nwo07d, Özk07, PH09, PST06, Raf07a, RB07, Sal08a, SGS09, SC08b, SNK06, Sri07, SYX08, TSD07, WW06, jW06b, WT09, WW09, WS08d, WX06, WS07i, XM06, XF08b, YCL07, YY06b, YL07, YYMZ09, ZG09, Zha08g, Zha09f, dCD06b, AK05a, AMK06, AN05a, AHE05, ABM09, AP06, Asl09a, Asl09b, Bay05a, BKFM05, BB06d, Bry07b, Buk09, CSV07, CLH05, Chu07e, DZ08c, fDxZfWyY09]. some [Djo06, DM08c, DC06b, EM06, EBsENMA05, EB08, ET05a, FK07, Fer09, Fro06, GAR09, Gre06, GG09a, GN07, HV08, HK09, Hon08b, HRS07, Hub05, HMT08, JVFM04, JF05, JMLR05, JG07b, JFG08a, KRRGT07, KZ07, Li08d, LS09d, MM07b, MA08d, MS06a, MKKS05b, MB07, MJD06, MHA06, NMM08, Nis07, Noo07a, NMM09, OSS09, Ouj06, PB07b, PP07a, Pet08a, QZL07b, QW08b, RNV07a, SS09a, SG08a, Sav07, SS07b, SS09d, Smi06b, Sok08c, SdJA06, Ste09e, SMG08, Tak07a, VCRCG09, Vul07, cWmZ06, Wan08c, WW08b, WYL08b, WP05, WYL06a, WS07h, WS08c, WN09b, Wu06f, Wu07e, WR07b, YZ09a, ZF05, ZY08a, Zha09h, Zhu06b, lSAQ08, Zho09b]. SOR [DKA06a, DH06c, HCC06, Li08e, Yun07c]. SOR- [Li08e]. SOR-type [HCC06]. sort [CSBS07, HD07, KC07c]. sorting [SC07f]. sound [Ahm05a]. sounds [DBR07]. source [Ahm05a, BR07a, Bou05, CG08a, CKS09, Far08, FTMA07, GJD^{+}07, GM05b, JZ06b, KWZ08, Li06a, LLFM08, MAY09, Pam06, PY08a, Pra07, Soh06, Zha06c]. source-type [Soh06]. sources [CY08d, EGS05, LCM08, YLH09, YL09b, Yu09, ZS06, ZQ08]. space [AB05a, ASAS06, AH09, Ant05, AEA07, BC09a, BHCK08, BT09a, CA08a, Djo06, DJ07, DC06e, EkES05, FTMA07, GXY06, GC07a, GC07b, Gen09a, Gen09b, GÇ05, Gui05, GC07c, HL06b, IR08a, IR06, IR07, KA06b, KA06c, KS06a, KL05b, Li09b, LWL07b, LZA^{+}07, LLC09b, Ma08c, yMT08, Moh05a, Moh05b, Moh07c, Mom05a, MD08c, PN08, Pud05, RNV07a, RCB08, Ray08, SP08, SZ05, SGP07, SKS06, Sin08a, Son07, SLW08, SK09a, Ste08, Ste09e, SM09b, WWL08a, WWL08b, WYL06c, WZ06h, WZ06i, WZ06j, YC06a, Zha05a, bZlDL09, ZW08c]. space- [Mom05a]. space-division [SGP07]. space-fractional [CA08a]. spaces [AR09, ACOS06, AOS09, ASK05, ABM09, AST08, AS09b, AS09c, BLLM08, Ber09, BDLS05, sCLC09, CjC08, CKQ08, ĆC09d, Dar06, DF09, Fen06, Gao07, GS09, HG09, JK08, JOB08, KWX08, KB05a, KKKM09, Kum08, LS08d, LS09d, LS09f, LS09e, LS09g, LL07a, LYH09, MW09, MA05g, NS07a, OS08, QN08, QSZ08, Raf06, Sav07, ST06c, SZ06b, SW09c, SiU09, Ste09f, Ste09c, Ste09d, Ste09g, Ste09h, Ste09i, TS05, WCO08, WZDgH09, WCO09, Wan08g, WL05f, WYL06b, WWYL08, XZ09b, XF08b, Yan09c, ZSM07, Zha08e, ZLW08, ZWC06, Zhu09b, ZH08d, ZH09b]. spacing [BTMB09]. spanning [GL05a, MKSK07]. spare [HZ06b]. sparse [BMP06, KT07b, KK05c, KM05e, LPP08, LZ06c, LY07c, LHB06, LH06a, Lu07, MAD07, Oss05b, PVN06b, Sal06d, SKT06, Sha08d, XLH^{+}07]. Spatial [Qui09, SM09d, Bab07, BW05a, GL06a, KAO08, KSKK06, Kös06a, Ma07a, SB05a, Yan08e]. spatially [Boy05, MTD05]. spatio [Lin09a]. spatio-temporal [Lin09a]. spatiotemporal [XCW08]. Special [CN09c, G07, Hua07, Hua08c, HHL08, Rao06a, AKIS07, AA07c, AR07b, BIK06, BG08, BR07b, Bry07b, CGDR06, CD08, DK06, Dön06b, DC06d, EM05b, GPS09, HRE05, JWPS08, KEM08, LD07b, hL08c, MN07, Mor08b, MHA06, Qad07, RJM07a, Sha06a, SDG05, SAM07, SS08c, uITA06, uIK06, SS07c, TF07, Waz05k, WS07h]. specialized [OF06]. specially [DY07]. species [BW07, Che05a, Che05b, Che06a, CLH06, CS06a, Che06b, CX06, Che06g, Che06c, Che07c, Che07b, CZ08a, DWZ07, JW07, LL05f, LTCC08, LZ09c, Luo06, MM07c, Mon07, MBM07, Pam05b, QD08, Wan08f, XCH05b, ZWZ05]. specific [HN07a]. specification [Ang05, Deh05a, HZS07]. specified [CC07b, LWK05b, MFE05, YY05b]. specimen [XFTY09]. spectra [SKHZ06]. Spectral [CCK06, Jav06d, BBES07, BŞ07a, CL08c, CD08, Dod08, DJ07, qDyC08, EHAR05, FXLZ05, Gar07, GJ07g, HYG08, Jav06b, Kan06a, REN^{+}05, Toh05, WLZN08, XF08a, Yan07e, YL08b, YC05a, lZXL09, Zho06]. spectral/spline [EHAR05]. spectrogram [DFG07]. spectroscopic [Har06, LP07a]. spectrum [CDT07, DZ08d, GRH07, HFL06]. speech [LCL06]. Speed [CH07f, AS06c, CW08, ZW07a, ZL06c]. speeding [CT09b, TC05]. Spetses [AVSB08]. sphere [Hu05, Men08]. spheres [GS09]. Spherical [JCX^{+}07, Ahm07b, KY05b]. Spider [KH05b, KH05a]. SPIHT [Che08b]. spin [YY09]. spinning [Ece06]. spinor [LZ09d]. spiral [KM05a, WCX06, XHLL05]. spirallike [KS08e]. spirals [KVŽŽ08]. SPLAGE [Moh09]. splicing [LLX07b]. splicing/decomposable [LLX07b]. Spline [AKR05, Baw05, KY08, RGM07, RJF07, RJK07, AS06a, AS06b, AÖ06b, Aks06a, AV07a, CCE06, CC06a, CC08b, DSB05, DN06, DZ08b, EHAR05, EHZES05, GK05b, IDS06, IT06b, JS07a, KA05a, KA05b, KK07b, KYK08, KK08a, KFM08, KR05, KB06a, Kan07, KKA05, KKA06, Kum03, Kum07, Kum09c, KE05d, Lin05b, Lin05c, LA07a, ML06, MD07a, MJ05, MA06e, NKE05, NHKZ06, ÖEK05, PLZ05, RLZ07, RLZ08, RK07b, RMJG07, RJM07a, RMG07b, RJM07b, RMG07a, RJMG07, Ras05a, SD09, SA06f, SA06g, SAM07, SA07a, SAN07, SA07b, SA07d, SA07c, SA07e, SAE08, SA08e, uIKTT05, uITA06, uIK06, uIT06, uINTK06, SH05b, STU09, TiHuI08, WZ08c, WCH08b, ZW09c, MD06b]. spline-collocation [MD07a, MD06b]. splines [AS05, ASNR06, ASA07a, Beh06, BB06c, DIS05, GC06a, KNM05, Kol06, LLC09b, MY06b, uITiHK08]. split [Sin08b, Wan05a]. split-step [Wan05a]. Splitting [Fil05, CS08a, CK05, Dao05, DS05a, DG07b, KK05a, Lin05e, Moh05a, Rem07, XHH06]. splittings [CT05b, Nač09, Yun07a]. SPOT [TTC08]. SPP [LW05b]. spread [AVSB08, GCSS05, Mir06b, NTO06, TNS07, YE07]. spring [Hua05a]. spruce [BKP06]. SQP [CS07e, Jia05, bJyKxC09, MHA06, SC07c, WCX08, WLY08, ZZ06b, Zhu06c]. SQP-filter [SC07c]. Square [Nak08, CMV05, CL05a, Cha05e, Guo09, HR07b, HZ05, Kök06, KY06, LTT06, LCC06, LZR07, LHSB09, MNN05, NS06a, Özb06, PP07c, PG07b, PG07c, RB08c, Sun07a, TK06b, Vic09, WCX06, Wu07f, WMC08, ZG08a]. square-pixellated [Vic09]. Square-root [Nak08, LHSB09]. squares [AÖÖ06, BX09, CC06c, DIT06, DX07b, DCL07, DDZ08, DJ07, GR06b, GR07b, HW05c, JS05, KT06, KKY07b, Mug05, NCÁHC^{+}06, yP05, hPyHZ06, QZL07b, QW08b, RJ07, Sam06, SG05b, TK06a, TY07, WF07a, fWZpXzL09, Yin09, ZW05h]. squaring [SCI08]. squeeze [BK06a]. SRH [Dön06b]. Srivastava [AJMR07, Dzi07, EDS06, HL08d, LL05g, MA09, PS06c, RSSS07, Sok08c]. SSOR [CC06c, YJ07, Yun09]. ST [CY06b]. stabilisation [Val07]. stabilities [ZL08c]. Stability [Ahm09b, AST08, CL09a, DW09b, Geo09, HHP09, HH09b, HCZ06, Ibr08, IMS09, JV05, LLG09, LLAT09, LW06d, LMY07, LZA^{+}07, LLY07, LYL07, MG09, Mat09, Moh07c, Sha09, SLC09, Sht07, SP06b, SHG07, Wan05c, WL06a, WTZ06, WQL09, Wan09c, WYL06b, Yan07b, YLS05, YWL07b, ZLZZ07, ZXL05, ZXDL05, ZW06a, ZXWL06, ZH07b, lS07b, AO08, ACP07, Alo09, AH07d, ALN06, Bai08, BR08a, BR09a, BWB05, BI08b, Bob08, CL08a, CLG09, CS05a, CZH07, CZ07a, CZH08, Che05a, Che06b, Che06f, CW08, CLWP09, DJ06a, DDR06a, DXH08, DC06a, DC06e, ElS06, FO09, FK06a, FeT05, GL06a, GZG08, GCT08, GF06, GH09, Guo09, HIST06, Hlo05, HL07c, HHW09, Hua08b, HL05b, JP09, JHS^{+}05, JSL06a, JSL06b, JC08a, Kas07, KB05a, Kim06a, Kim08, KPL08a, KP08a, KP08c, KP08b, KPL08b, KPL08c, KPL09b, KP09b, Lee05d, LRMS06, Li05c, Li06a, LWG09]. stability [Li09c, LZdC06, LZC07a, LOZ07, LTC07, LLZ08, LFC08, LZWK09, MZW08, MS09a, Mak07c, Mak09, MSD06b, Nam08, Nam09, NL05, NMP08, NPGW09, PRM06, PBPF08, Par05a, Par05b, PK05c, Par06a, Par06b, Par06c, Par07, PK08c, PKL08a, PPKL08, Pen06, PNL05, Prá08a, Prá08b, QCSF09, QC08, Rad05a, RB08a, RB08b, RB08c, RC06, RX09a, RY06, RHQ06, SMM08, SSA09, SA09, SYH07, ST05b, Sin08b, Sin09, Sol05c, Sol05d, SL06a, SLL08, Ste09b, SL05b, SWX06b, SWX07b, SM07d, SM08d, SM08e, eT09, TZX07, TG08a, Udw05b, WZ08a, dWqL05, WL07b, WS07e, WZ07a, WZL08a, Wan08f, WLZ08, WWL08a, WWL08b, WL09a, WL05f, WWYL08, WRL06, XCD05a, XLL06, XZ08b, YB03, YB04, Yam05a, Yam05b, YLZ06, Yan05d, YX06a, YZW07a, rYzS07, YFLX09, YWL07a, YZFL08, ZWX05, ZS07b, ZY08a, ZzLZ08, ZLY08, ZSCW09, ZQJ09, ZYT09, ZWY07, ZZ07b, ZZW09b, ZLL06b]. stability [ZFT06, ZSS08b, ZW08d, ZH08b, ZL09a, ZXY06, Zhu08, d’O05b, dlS06a, dlS07b, AM08b, AM08c, AM08d]. Stabilization [CCZ08, AD06a, HHP09, ID08, Ibr08, KP06b, KP06c, KP09a, LP08a, LmZ08b, LmZ08c, LZ09b, PK05d, PKL08b, Pha06, SWX06b, WS09c, ZC05b, dlS06b, dlS07b]. Stabilized [Kim06b, GFH08, He06, HZX07, Li06b, LMH06, LD06b, SH09a]. stabilizers [LL06a]. stabilizing [Far08]. Stable [CF06b, Sea06, AKM06, Ars07b, DZ09, gDH07, GC07a, Inc07e, Inc07b, Jin08, LTCC08, Moh05a, Moh05b, SSEA06, SDR07, TD06, TD07a, You09, dlS06c]. Stackelberg [Muk07a]. stacks [FEOC08]. Stacy [GN06b]. stage [AAAA06, BB06a, CLG09, Che06a, CH06b, Che06i, Cho05, FK06b, FK07, GCT08, HTO06c, JL07, JMC07, JPCL08, KeTL08, LMC08, hMzLfW^{+}08, hMfWzLpZ09, NBHYV09, SA08b, WB06, WGF^{+}06, XCD05a, XCD06, Xu06, XZ07c, Yun07a, Yun09, ZCZ07]. stage-structure [XZ07c]. stage-structured [BB06a, CLG09, JMC07, JPCL08, KeTL08, XCD05a]. Staged [SC08c]. Staged-structured [SC08c]. stages [CL08a]. stagnation [Cha05g, MNG09]. stagnation-point [MNG09]. staircase [Sum07]. stamping [HHC05]. Standard [Mar08, SMH07, CSBS07, LP07c, PS06b, QFX06, Ram07a, WY07, XFL06, Zha07e]. standby [AKKS06, JW09a, WC06a, WDK06]. Standing [GZ05, CHH09, HZ09b]. star [KGM06, Li05e, XHZ06, Yan09a]. Starlike [OP09, Aou08b, Att07, GZ07a, Kan09a, NPR09, NO07, Özk07, Sok07a]. Starlikeness [Sok07b, Sok08d]. startup [WWP05]. State [PKL08c, Alt07a, Ant05, BWB05, Bel07b, CC07b, Che08d, CFF09, DJM08, EHL08, HC07b, HCLP07a, JPYW09, Li07d, hL08c, Lin07b, LW06d, LS07e, LL08f, MM07a, MI08, MP08b, NPGW09, PK05b, PK08b, PK09, SMY08, XCW08, Xu08b, YSdS^{+}06, ZZ08a, ZY07, d’O05b]. state-constrained [MP08b]. state-delay [MI08]. state-dependent [Che08d, LW06d, MM07a, SMY08, Xu08b, ZZ08a, ZY07]. states [KB06b, YSCC05, ZSHB06]. Static [Sel07, HRM07, KK07c, SM07a, WCH08b]. Stationary [PWY08, DLLN09, DM07b, GZL08, GFH08, Gür05, HBBA07, KTV07, KX08, Li06b, LMH06, SRH09, WL08b, Yi05, ZH05]. statistic [Ery06]. Statistical [EDS06, WY05c, CS07a, DMS09, KC07c, LCL06, MJD06, MM05c, Rad09, Val07, GGSN06]. statistically [CS07d]. status [XLHK09]. Stavroulakis [SM06d]. steady [AVM^{+}09, BWB05, BB09, CFF09, EHL08, KAS06a, KB06b, SI06, YSdS^{+}06]. steady-state [EHL08, YSdS^{+}06]. Steepest [AJ06a, AJ06b, AE07c, YNCL08, CHLS08]. steepest-descent [YNCL08]. Steepest-Descent-Kaczmarz [CHLS08]. Stefan [Asa07b, BT06b, Kut05, MV09b, OMST07, Yi05, Yig08]. Stefan-like [Kut05]. Steffensen [AB06, Jai07, PC09, RWB09, Sha05e, WS07i, ZWZZ09]. Steffensen-like [ZWZZ09]. Steganography [Wan05f, Che07h, Che08b]. stenosed [IAMA08, MCMA07]. stenosis [SVMG09]. stenotic [BA07, SM07b]. step [CP08, CL05d, CJP08, fDxZfWyY09, HEZ08, HBE09, HCLP07b, HBN06, HMT08, Inc07b, JWPS08, KK08b, KA06d, LW05a, Li05c, LX07b, Luo05b, MZ07, NBHYV09, NS06b, NS07b, NAJ06, NN06c, NY07, NH07b, PP07c, PCS07, Psi09, RWB09, SNH07b, Sal06a, SS06d, SS08d, SY08a, TLM09, TCMPMC07, TCMP07, Wan05a, Wan06a, WB09, WF07d, YN07, YNCL08, Yun08b, ZWL06, ZG08a, ZW06f, Zhu05c]. step-point [Psi09]. step-size [SS06d, TCMPMC07, TCMP07, WF07d]. step-stress [Wan06a]. stepping [Ang05, You08]. steps [BNR06, Neh07b]. stereo [HLCH08]. Stewartson [GC09]. stickers [HC08b]. Stieltjes [AMJD06a, AMJD06b, HGN07, IPPT05, ST05c]. stiff [GAR09, GB05, IOS07, MCAH05, Psi09, SJ06d, WL05f, WWYL08]. stiffness [ÇFÇ06a]. Still [Gal06b]. stirred [ZGL08]. Stochastic [AABD07, Alb06, Bob08, Hua05a, Kla08, KO06, Oht08, RJW07, Vaj08, WZ06e, ZXF07, Ade07, AÖ05b, AZ07, BR08a, BK08b, BC09b, Cha05h, Cho05, DLS06d, DXH08, DCR07, EBsENMA05, EG05, ET05a, Ewa06, FX09, Guo09, GGSNR06, HBBA07, HS09b, Isk06, Isk07, Ji05, JJ09b, KL05a, KMN09, KTV07, KT08b, Kun08, LS09a, LT08a, Lin07f, LXZ09, LZWK09, MM05a, MA05e, MDZ08, NCÁHCLP05b, NCÁHCLP05a, Nak07, NHCLP07, Nak08, OAHR09, PRM06, PLL08, Pog07b, RB08a, RB08c, RX09a, RX09b, RHY06, RZ07c, RMWk09, RSZ08, SMY08, SA08b, SGP07, TMJV^{+}05, TMJJS07, Vaj07a, VL09, WZ08a, sWX07a, sWMX07, WGL08, WXW09, zWxY07, WL08c, XXC05, XXC06, XHSL06, XHSL09, You09, mZzH05, ZH07a, ZZYY08, ZG08a, ZD06b, ZW08d, ZL09a, GGSN06, Nwo06d]. stochastic-flow [Lin07f, XHSL09]. Stochastic-Volatility [Nwo06d]. stock [GBG07, JTMRS06, Nwo06b, TA05a, WW05d]. stock-dependent [JTMRS06]. stocking [JPCL08]. stocks [Hor08]. Stokes [AM08a, BC09a, BH07, Cao06, CLA08, DC08, DS06a, DC06c, DMZ08, Eid05, EGFAE06, FHYZ09, GJ05, GFH08, Gw05, Gür05, GD06, He06, IHB06, JH09, Li06b, LMH06, LL08e, cMpZfR06, MO06, Ony05b, Ran08b, SH09a, SRH09, Soh05, VHFF08, ZOZZ08, ZH05]. Stolarsky [FFRFCC08]. stop [CLS05b]. stopped [wG07a]. Stopping [LT08a]. storage [MBM06, MMM06]. Storey [mTxWyL07]. straight [GY06, KJ09]. strain [SA06d, SV07]. strata [GKAS07]. strategies [CTGD08, KC09, Kwo07a, LL06b, MKA07, Mor05, Muk07a, SSEA06, WS07f, Zot08]. strategy [hB07, EAG08, GY08, Hor08, HH07a, JP09, JjQjWmL06, KR07b, lLWhL06, LCZ05a, Mak07a, MC08, Tan07c, Wan05d, WLT05a, WL05g, YST09, YSL05, YHL08a, YHL09b, YKJ07]. strategy-proof [Tan07c]. stratified [KC06a, Kim05a, KS05b]. Stratonovich [OAHR09]. stream [AEA05a, AES05, EGFAE06, SM08f, cMpZfR06, Pan08b]. streamfunction [AFH05, WZZ06b, Zhe08]. streaming [Rad05a]. Streamline [Gür05, SFA08]. streams [JWL^{+}07]. Strength [HKH06]. stress [KO06, LLX07a, SKS06, SV07, Vin09, Wan06a]. stressed [CFC06c, CH07b, SA06d, Sel07]. stresses [LYCW09, Sel06, SA06e, Sin08a]. stretch [Wan08e]. stretched [PR06]. stretching [AEA05a, AES05, Ari07, Cor05a, Cor07, EEEE05, EO06, KA09b, KBKV09, MNG09, MS09b, Oua05, Pan08b, Pan12, SH07b, VC06, YC09b]. strict [Jia05, bJyKxC09]. strictly [Fab08, OS09, Taw06]. string [CZ05b, CZZ05, ZC06a]. strings [CZ05a]. strip [ARM09, LF08]. strips [Li05f, Li05g]. Strong [CPY09, Jun09b, LL07a, QSZ08, SZ07a, SC06b, Son08b, SL06b, WMC08, YNCL08, YLK09, YZW07c, ZWC06, lS08b, ABM09, Bil04, Bil07, HWS07, JK06, JH08c, LL05c, MSM07b, Son08a, Tha07, YC06b]. Strongly [KT08a, ĆRR^{+}08, Dem05d, HBN06, JZHT05, JZTH06, JXH06, bJyKxC09, LL05e, ML09, Muk08a, Noo08a, WZC08, YXSX05, ZMV05]. Structural [DES07b, CG06, Jön05, LT09, PE07b, Whi06, Yan09b, Yua09]. Structure [SS07d, Bab07, Che06a, Che06i, Far05, FK06b, GCT08, mHzC05, HLY09, JL07, JL09, LZ07d, hMzLfW^{+}08, hMfWzLpZ09, MC07, PVS^{+}08, QW05, SAK08, SCR08a, SCR08b, She05b, SLWX06, SI06, SEŞ09, WC06c, WJZG08, Waz05l, XZ07b, XCD06, Xu06, XZ07c, YZ08, ZCZ07]. structure-based [JL09]. Structured [Far08, BB06a, CLG09, FK07, Ham05a, JB08, JMC07, JPCL08, KeTL08, LWW05, LMC08, RS08b, SMM08, SC08c, XCD05a, ZH07a]. structures [DW09a, EH09d, Inc07c, LHSB09, Odi08, She06, She07c, Waz05f, Waz05h, Waz05k, Waz06f, Waz08r, YX08]. studies [FLLH07b, HAK07a, HP07, Wan05a, ZHZ07b]. Study [AY05a, BLLM08, LC05b, MNM07, NDD05, WH08b, XFTY09, XXC06, ARK07, AH05a, AÖ09a, Att05a, Cen06, ÇE06b, CS07a, CFC06c, Dam08, Ere06, HCS05, IRSR05, Joy06, KYK08, KA09a, KYA05, KuI08, KLY07a, KC07c, KH09, Liu06h, LHSB09, MP08a, MSD06c, Özt05, REN^{+}05, SML05b, SCR08b, She05c, SG05c, SM07b, WY06, WT08, Waz06b, Waz06d, Waz08a, WC05d, Wu05c, XO07, XWSL08, ZPT^{+}06, ZZW08, ZS08b, ZS07d]. studying [KNM05]. Sturm [AN05a, Att05b, AL06, BŞ07a, BR09c, BD09b, ÇG05a, Cha07b, Cha07c, Cha07d, CM08b, JGY07, JLW09, KB08, KB09, SMG08, YWL09, YHL08c, ZW05j]. STWS [PDA06]. style [CWY06, DL09, HL07b, YC08b]. Sub [Ahm05a, TW09, AS06g, Cel06a, FLLH07a, FL09a, JZHT05, JZTH06, JXH06, bJyKxC09, LY08c, MMO07, Sil07c, SZ07d, SZ07e, WSZ07b, Wei06b, WZ08e, YXW07, ZX06a, Zha08d]. sub-equation [FLLH07a, FL09a, SZ07d, SZ07e, ZX06a, Zha08d]. sub-equations [WSZ07b]. sub-feasible [JZHT05, JZTH06, JXH06, bJyKxC09]. sub-grid [Sil07c]. sub-linear [Wei06b, WZ08e]. Sub-ODE [TW09]. sub-soil [Cel06a]. sub-superlinear [MMO07]. sub-supersolutions [LY08c]. sub-units [AS06g]. subblock [BLZZ09]. subclass [CKR09, Den09b, GZ07a, KS06a, Liu07c, Sok07a]. subclasses [Ahu08, Aou07, Aou08b, EA08a, HL08d, MA09, SS07b, SYX08, WW09]. subdiffusion [mCLB08]. subdivision [BCR09, Boy06a, KLY07a, KLY07b]. subdomains [DjT05b, ZT06, Zha07f]. subject [ALD08, Ang05, HZS07, gLXyZ06, Lin08, WWL06, WYL08a, XXC05, XXC06, ZLCZ06, Zhu06a, dlS07a]. subjected [EGS05, LD06a, YY08b]. sublattices [Kat09]. subliminal [LY05]. sublinear [AMN06, ANM07e, AMN07e, DZ07a]. submatrix [fLyHZ09, LD07d]. Suboptimal [Mor08a]. subordination [CKOS07, KS08e, Özk07, Sok07a]. Subordinations [AJMR07, KKA09b, LN07]. subregion [MFE05]. Subresultant [ADTGV09]. subresultants [Li06e]. subsampling [LLC06b]. subscales [BC09a]. subset [BX09]. subsidy [SM09d]. subsistence [GL06b]. Subspace [Erb06, ASA07c, BLW06a, HJ06, LB06, WM09, XW07a, YQYL07]. subspaces [Kat08, Kat09, cWxChY07]. substance [OMST07]. substances [LC06a]. substantial [WX06]. substation [HJL08]. substitutable [DM07a, YVU06]. substrate [dSYGM^{+}05]. substructure [DjT05a]. subterranean [LS08c]. subtree [HH07a]. successful [AA07d]. Successive [All05d, SCI08, SG07c, SMG08, DH06b, DKA06b, KKT07, RJ07, WZ06d, ZW05h, ZW06b]. successively [HBN06]. such [Dav06]. suction [KA09b]. Sufficient [Woo05, CC06c, DZ07a, HWS07, NL05, PC07b, Ren06, SS07a, SS09c, Sok07a, Son08a, Via08, Wei06a]. Sugeno [Hu07b, Mao09, OML09]. suitability [SGUK08]. Suitable [MGMC05, xGyZtZ^{+}07]. sum [JS07b, Neh07a, OT06, QZZ07, SW06a, SCM09, TMMVA05, TMMVA06, YPZ05]. summability [ABM09, Lal09, MSM07b, MS08c]. Summation [CS07j, Cvi09b, Cvi08, EF07, SFG07, WN09b]. Sums [GN06c, GN06b, CS07j, Cvi08, Cvi09b, MK05b, Nis07, Sof09]. Sun [Mar14]. sup [LMH06]. Super [DB07b, BT08, Bou06c, CHT06, CH08b, IMB08, KSM05, Kho07, KL07a, KLW07e, LJK07, LCX09, RAZ07, SS05d, Udw05b, Wei06a, Xu09, YL09c]. super- [Bou06c]. super-attainable [IMB08]. super-efficiency [Kho07, LJK07]. super-Halley [CH08b, KL07a, KLW07e, RAZ07]. super-instability [Udw05b]. super-integrable [YL09c]. super-linear [BT08, KSM05, Wei06a, Xu09]. super-memory [SS05d]. super-stability [Udw05b]. supercompact [AM08c]. Superconvergence [Att06d, XC06b, XC07, mH06c, SL09b]. superiority [CL05d]. superlinear [ANM06, AK06b, AK06d, ANM07j, CZ07c, LM07b, MMO07, WLY08, Wu05d]. superlinearly [JZTH06, bJyKxC09, MCW08, Zhu05g, ZZ06b]. supermemory [SS06c]. superordination [CKOS07]. Superposition [LZZC09]. superquadratic [CT09a, XT08]. Supersingular [CJL05a]. supersolutions [LY08c, YXW07]. Superstability [Bal05]. Supplements [GQS08]. supplier [CCL06, yLLjP06]. suppliers [Sae06a, Sae06c, Sae07]. supply [Chi07a, MQW05, PFA07, RSA07, SA08b]. Support [IM08, MTL08, CGVC07, HGN07, HZ06b, Inc07a, LY09b, PG07b, PG07c, Sha06b, yXhWZ06, YZY08, YY06a, YES07]. supported [LFW08, yXhWZ06, YH09]. suppression [HLY09]. Surface [BK06a, AEA05a, AES05, BCCK05, Bat08b, Cel06a, Che06n, CFF09, DN07b, DLBT08, EO06, JK07, KB05c, Küç05a, LC05b, Lin08, LLW05, Ma05b, MNG09, ME07, Pan08b, RNV07b, RNV07c, XFL06, XF08a]. surface-based [DN07b]. Surfaces [KA06c, AAABH04, AK05f, FS09, KAK05, KAK06, Kas06b, KAO08, KB05c, Küç05a, Rab05, Sid05, Sid06, TSS09]. surgical [GPTU05]. surrounded [BMD08]. survey [AMJD08, Gha08, KKA05, KLZ07, SANN07, WGF^{+}06, WLT07a]. survivability [AGT07]. SVD [CYH^{+}07, FWL08, HM09a, ZCZ05]. SVD-based [CYH^{+}07, FWL08]. SVEIR [JWS^{+}09]. SVEIRS [SJW09]. swallowing [SGG^{+}09]. swarm [AK08a, BEvdB07, CCZS08, HL08a, HW07a, HM06, JCSR07, JDSR08, JHHW07, JZHW08, KWdW08, KCW07, Lee06b, LKP06, LGJ06, LJG06, LY08d, MBDD08, XSX08, NZHW07, NJG08, ONG08, PWL06, QYYL08, SKHZ06, SSJK07, WWC09, XLwW07, XwWL07, YYYM07, YE07, YW06b, Yin07, YYWW07, ZZLL07, ZFL^{+}09]. swept [AVSB08]. switched [HHP09, Ibr08, LZ09b, LLZ08, MZW08, Mah09b, SFFY08, SWX06b, SWX07b, XW09, ZLZZ07, dlS07b]. switches [Mat09, SGP07]. switching [DTC08, GZG08, HLK06, MB08, RB08c, RY06, RHQ06, RZ07c, RMWk09, SSEA06, WDK06, sWX07a, WC09b]. Sylvester [BLW06a, BJ08, DH08, DLD08, Jbi06, KTK06, KKT07, KKT05, Lin05d, Lin06d, ST06a]. symbiotic [CZ08a]. Symbolic [KC05b, TSP07, TS08a, WLZ05, XZ05, ZX06c, BQYW07, CW05a, CW05b, Inc07d, KM05b, KM05c, Kıy06, Li07b, LZ08c, Li09d, MG06, Pou06, SZ07b, SZ07c, WL05c, WZSZ06, WZ07c, ZX07a, ZZB08, ZZYZ06a, ZZ06a, ZS09b]. Symlets [AM08c]. Symmetric [DH06c, DKA06b, Kas06c, RNV07c, SA06a, VS07, WZ06d, ZW05h, ZW06b, AJE05, AMJD06b, Alo09, BK07b, CG07a, Che07g, CS07i, CB06, DH06b, DKA06a, DM08a, FBB08, GG09a, GG08c, GG08d, HY07b, HYG07, HY07c, JGY07, JFG08a, Kas09, Kay05d, Kay05e, KLY07a, LpL07, LWG09, Lin07c, Nad06a, yP05, PHZ05, PHZ06b, PHZ08, RS08a, RNV07b, Rim05b, Rim05e, Rim05c, Rim05f, Rim06c, Rim06g, Rim06d, Rim06h, Rim06a, Rim06b, Rim06e, Rim07c, Rim08a, Rim08c, Rim08d, Rim08e, Rim09a, Sal06b, SG07c, SMG08, TGSNB07, Tzo08, WD06, WZC07, WCL08, WN08b, YH09, YWL09, YL07, Yua09, YK06b, ZS08a]. symmetrical [hL08c]. symmetries [DSK08, SM05c, Sjö07]. symmetrized [Hai07, Hai08]. Symmetry [CTT08, FHYZ09, Ma07c, MZ09, PPS09, lYqL06, lYpZqL08, Zhi09, Hon08b, HC09, gLXyZ06, LZR07, MHM09, lYqLW07, YLX08, ZZ07d]. symplectic [CLS06, KLZ07, LSW06a, MKS09, TmChL06, WW05a, WZC08]. synchronisation [Kay05a]. Synchronization [GZ07d, PCCHLG08, XS09, BFR05, Che07d, CAL08, ETJ06, GL05b, GZG09, hL08c, PJWL08, SW09b, WC06b, WS09a, XXTL09, YY08b, ZWLX08, Zhu09a]. Synchronizing [Rom09a]. synchronous [TTZ09]. synovial [BK06b]. Synthesis [Shi05c, AD06c, LLZ08, RSD07, Shi05b, WC05c, ZLZZ07]. synthesizing [DTC08]. System [GB08, Lin07f, AJE05, Abb05, AEJ06, AE06a, AJ06b, AEFFA08, ACP07, AS05, ASN07, AS06d, All05a, All05d, AAAA06, AK06g, Ami05, AGT07, An05b, ANK05, ALN06, AGI08, BSKK06, BK08b, BY05, BWB05, BFI06, Bog05c, Bou06b, Bou09a, dPCdBYV08, CTT08, CL07a, CS07b, CCH05, CLS05a, CL05b, CCL06, CMMC09, Che05d, CDP06, CLH06, CX06, Che06f, Che06g, Che06i, Che06c, CF06b, CY06a, CG07b, Che07c, Che07b, CS07i, CjC08, Che08d, CCZ08, Chi07b, Cho07c, CK06b, CK07b, Cot08, CC05d, hCxC05, CY08d, Dah07, DW09a, DG07a, Den08b, DQW06, DW09c, DMR07, DF08, DKL09, EN05, EGS05, EM09, FTG07, FM08, FDL08, FC06, FqZ08, FLLH08, FX09, FK06a, FK06b, GN06a, GLV08, GFWX05, GL06a, GCT08, GL05b, GLLC08, GL08a, GJB05, GJ07d, GJ07f, GS08a, GSPG07, GbL07, GM08]. system [Gui05, GZ07d, HG09, HBVM07, HCX06, HLK06, HCW06, HN07b, HHL08, HL07d, INR08, JDG06b, JKN09, Jav06c, JG07a, JK08, JW07, JW09a, JLL08, Kad07a, KF08b, KK07d, KKA09a, KKKM09, KC06b, KC07b, Kia05, Kos05, KeTL08, KLR07, KB07, KH07b, Lar05, Lay09a, gLXyZ06, LXL06, LLX07a, LmZ08b, LmZ08c, LCS08, LH08b, LCM08, LL08d, LX09a, LZC07a, Lin05c, LH07a, Lis06, LZW07, LTC07, LZ09c, LLZ09, LHF07, LH08d, LC08b, Luo06, LXJL09, hMzLfW^{+}08, hMfWzLpZ09, qMpXjH09, MJFAS07, Mak07b, ML07, MSK05, MY06a, ML06, MAR06, MS08a, MBDD08, ML08b, Mat09, MŘ07b, MXC07, MGG^{+}09, MY07a, MGMC05, MMN06b, MB05b, MED08, NDD05, ND07, NPGW09, yN06, NMS07, OHJ07, Özb07a, PPS09, PW08a, PKL08b, PHZ06b, PNL05, PG07a, PG07b, PG07d, PMRAS05a, QD08, QC08, RMA07a, ROK09, RKOK09, RN09, RS08a, RMJG07, RJM07b, RJMG07]. system [RCB06, RGAFVA07, RIY05, SF07, SMA08, SS05a, SH07a, Sal08c, ST06b, SMBB07, SCS05a, SXX05, She07a, She07b, SA07a, SA07d, SA07c, SLO08, uIKTT05, uIT06, uINTK06, SFFY08, SWX06a, SWZL07, SM07d, SZ08c, TF08, TMRV06, TQC09, WSLZ07, WWP05, WZ06b, WFL06, WSL07, WZS07, WZ07a, WS07b, WLP07, WFX07, WZC08, WLZ08, WY09, WS09a, Wan09b, WCZ09, WLA09, Wan09c, WGCH05, Wu05e, WZ06g, Wu06d, Wu06e, WP06, Wu07d, Wu07g, WF07d, WSXL09, Xin06, XZ07c, XM08, YVU06, Yan05b, YL06a, Yan07b, YLB05, YCZ06, Yan06c, YC06b, Yan09d, YLC09, YL09a, YFLX09, Yil06, YMY05, YC06d, YZ07c, YZ07d, Yu08c, YL09c, (Ag07, ZZH06, ZSHB06, ZLC06, ZL06a, ZD06a, ZYME06, ZC06a, ZH07a, ZT07a, ZCZ07, ZXF07, ZY08b, ZLY08, ZL08a, Zha08g, ZXL05, ZWY07, Zha09f, ZW05g, ZQ06, ZS06, ZM08d, jZzFbL06, ZXZZ07, ZXLZ07, ZWLX08, ZX08]. system [ZH09b]. systematic [YEN08]. Systems [SL07b, ABA06, AZ08, ANM07c, AMN07e, ANM08, AM06, AS08, AGR09, AC07b, AST08, AC09, Ant05, AAA05, AFS06, ASA07c, AGM^{+}07, AK06i, AKKS06, Bab07, BDE06, BUP07, Bal05, BK09, BT09b, Bel07b, BE06, Ben07a, BD05b, Bia05, BG07a, BI08b, BS08, BGR05, BRT07, BT08, BB06d, Bou06c, CH07a, CCS05, Cao06, CZ07a, CZH08, CS08a, CHLS08, Che05a, CLZ05, CJC05, Che06e, CCC07, Che07d, CC07b, CX08b, CAL08, CT09a, CH09, CKQ08, CLWP09, CPTZ05, CST07, CL06d, DTC08, DH06a, DH06c, DKA06b, DB07a, DB07b, DB07c, DHG06, DH06d, DH06e, DMA07a, DLS06d, DCL07, DDZ08, DMDB08, Du07, Efi09, EF05, EFI06, EM05a, ESES06, ES05b, EH08a, Erb06, Far05, FSMC05, FF06, FWJK09, GD08, Gal06a, GZG08, GAR09, GJL06, GAV08, GC06b, GZ07b, GI06, GWX06, GY07b, GZG09]. systems [GB05, Han07, HC08a, HL08b, HCLP07a, HRT07, HHP09, Hos05a, Hos06c, HW09, Hua05a, HWF07, HMT09a, HMT09b, HLY09, HJL08, ID08, Ibr08, IGM07, JDG06c, JFG08a, JPYW09, JZ08b, JSLL07, JZX07b, JZX07a, KMB05, KNM05, Kar06, Kar07b, Kar07a, KT06, KK05b, KL06a, KB05b, Kis08, KK08c, Kla08, KLW06a, Kou07f, KL07h, KM05e, KP06b, KP06c, KPL08a, KPL08b, KPL08c, KPL09b, KP09a, LCC05a, LP07b, LP08a, LRMS06, LL05f, LC06a, tLxLlW07, LD07b, Li07d, LCZT08, LYX08, LY08a, LK08, hL08c, LZ09b, LDC09, LMZ09, LL06a, LZdC06, LTZ09, LYL^{+}08, LGHS06, LLW07, LLZ08, LFC08, LM08c, LXZ09, LLW09, LXZ07, LZWK09, LL08g, MZW08, MX08a, MS09a, MCAH05, MI08, Mah09b, ML09, MS07c, MV09a, MPT08, MR05a, Meh06, MN05, MM07e, MY08, MSR08, MHM09, MAD07, MAK05, MF09, Mor08a, Mor08b, Muk07c, Muk07b, Muk08a, NCÁHCLP05b, NGLHCLP06]. systems [Nak07, NHCLP07, Nak08, Nam08, Nam09, NM09, NA08, NMP08, NBO09, Par05c, PK05b, Par05b, PK05c, PK05d, PJWL08, Par09b, PVN06b, Pe¸ 09, Pha06, Pud05, PDA06, QCSF09, RBvB08, RMS05, RC06, RO07, RS06, SMY08, Sal05, SA09, SM06b, SLX05a, SW09b, SH07e, Sog08d, Sog08b, Sol05b, Sol05c, Sol05d, Son05, SM05d, SWX06b, SM06e, SWX07b, Tad08, TmChL06, TMJJS07, TTZ09, TZX07, Toh05, Tou09, TK06a, TTG07, Uje06b, VGMV^{+}07, Vog07, WG05, WS05a, WC06a, WDK06, WWL06, WL06c, WC06b, WZ06c, Wan06k, WZ06d, WDL07, Wan07c, Wan07f, Wan08a, Wan08e, WJZG08, WC09b, WL09c, Waz09e, WWW03, WCX05, WW05b, WG07b, WZ08e, WS09c, Woo05, fW07b, WY08, XO07, XW09, XXTL09, XXC08, XQ06b, XT08, YSdS^{+}06, YY08b, YCC08, YWZW07, YYWW07, YY07b, Yu05, YS06d, ZŠ05, ZW09b, ZZL05, ZW05e, ZC05b, ZHLG07, ZHL07]. systems [ZDL07, ZLZZ07, ZHGG08, ZY08a, lZXL09, ZW06d, ZW06b, ZW06c, ZC06d, ZW06e, ZW06f, ZXZL07, ZY07, ZH08c, Zhu09a, ZG08b, dlS06a, dlS06b, dlS06c, dlS07a, dlS07b, lS07a, lS07b, lS08a, Zot08]. systems-based [LD07b]. systolic [Mis05]. Szasz [GS08c, GBG08]. Szego [Cao09, GVMT07].

T [Pan08c, HL05c, PW05, SC05a, WS07e]. table [CC05b]. tabu [LFM06, MKSK07, Pal07]. TAGE [MK06, MA06f, MD09b]. Taiwan [Liu06a]. taking [ESA05]. tandem [LFM06, dSF05]. Tangent [HL06b, Cvi08]. tanh [Abd07, AKARA08, EELZS05, HM07, SS08c, Waz05n, Waz05p, Waz05m, Waz06b, Waz06a, Waz06m, Waz06q, Waz07b, Waz07a, Waz07e, Waz08c, Waz08d, Waz08u, Waz08w, ZZ07d]. tanh-coth [SS08c]. tanh-function [EELZS05]. Tanner [PWY08]. tapered [EESESD05, IAMA08, SH07c]. tardiness [IGM07, KTW08, LC06d, OT06, TMMVA05, TMMVA06]. tardy [JRA^{+}07]. target [AK05d, ÇE06a, LL05h]. task [YYWW07]. tasks [Smi06b]. Tasso [SQHY08, Waz07g]. Tau [AA07a, CE07b, CE07a, ERAS07, PMRAS05a, PMRAS05b, RAS07, Sha05a]. taxation [AH07a, MTD05]. Taylor [NBHYV09, AA07f, Bar05b, BK05b, GS05, GÖ05, GSG06, GS06d, Has07a, IRSR05, İBİ09, KM05f, LST09, MA05d, MA06c, MAR06, MK07a, OS07, SG06, SGT06, SAD06, SG07a, YS06a, ZHZ07a]. Taylor-series [MA05d, MA06c, MAR06]. TCSC [RSDC08]. TE [TA05b]. team [AH06c, DHT09]. team-game [DHT09]. technique [AS08, AA07a, AS06b, AA07c, AD09a, ADK06b, ADK07, Baw05, CN06, Che08c, CCZ08, DHS07, EL05b, ETETH05, ETETH07, EH08b, GJ05, Gw05, Ham05c, HK06, HFL06, HS09b, KK08b, Kad07a, KBH06b, KJ09, Li07c, LZ09a, Lin07d, Mak07c, Moh05a, Nak07, Nav08b, NMD07, NH07c, PDA06, QN08, Ram05i, Sal07d, SAK08, SZ05, SFPE07, SA06f, SA07b, SA07d, SA08e, jW06a, WZ09c, Wu05a, YCL09a, Yan09b, YEA07, dSYY08]. techniques [Che08b, CGS08, DMA07a, FV05, Fig08, GSNDB09, GSG09, GEG07, HW09, LZ06c, LLWC07, MAD07, Ram05f, Ram05g, Ram07f, Ram07e, Ram07g, Ram08b, Ram08e, RSDC08, Sla06, SdJA06, WC06c, ZHL07, Zot08]. technological [Hon08a, Pal08]. Technologies [Sae06e, Sae06b]. technology [Sae06a, Sae06c, Sae06d, WH08a]. telecommunication [Lin07e]. telegraph [EAEG07, Mom05a, WLA09]. Teleonomic [Pud05]. teleoperation [SMA08]. temperature [AB05c, EESF04, GKAS07, HKS09b, KJ09, PM09, Pan08c, SV07]. temperature-dependent [KJ09, PM09]. tempering [LPA^{+}09]. template [LLC06a]. temporal [Lin09a, TTC08, ZW07a]. temporary [MJ07e, RYE07, TML07]. tendency [Sil07f]. Tenenblat [Hub09a]. tension [MA06e]. tensor [CMR08, Mis05]. tenth [SA07e]. tenth-order [SA07e]. term [AAA08, AAA09, BR07a, DGB08, DXH07a, Dod08, EMESES05, EM03, HZ09b, JZ06b, KT07c, Kut05, LLFM08, LY07d, MCS08, MMR09, Pam06, SG08a, Sal06c, Sog07, WF07b, ZL07a, Zha08b, ZFYH08]. termination [Sal07a, WZ09a]. termites [LS08c]. terms [Ahm09b, Bou05, Dai09, FBB07, GS05, HT09a, HTM09, HT09b, Hub09d, KS06b, Lee05d, Lin07f, Moh07c, Öza05, QBH06b, TF07, WZSZ06, WL06a, WTZ06, WRL06, ZZ07a, Zha09c, Zha06c, ZX08]. ternary [BCR09, KLY07b, LXH06]. terrestrial [ESS05b, WL05g]. Tesi [PKL08b]. test [ACP07, Bak05, Boy07b, CE07b, Dön06b, Kat05, LZ09a, SW05a, WYX05, WZ06e, WYL06c, WWC06, qXbL05]. testing [Boy06a, KO06, Wan06a, WLL07a]. Tests [KCK08, CT08a, Jön05, PJ10, Sya05b, WW06, YEN08, lS07b]. tethered [QC07]. tetrahedra [RNV07a]. tetrahedral [RVN07, RVNI07]. texture [MM05c, WLL07b, YQYL07]. Th [Bra09, Akı06, Alo06a, Cha05h, DM08c, EM08, EOM03, EOM08, Fra09a, GJ07b, HJ08, KEM08, Kıy06, LZ08e, MS08b, NCÁHC^{+}06, Neh07b, PDW06, RLZ08, WLW09, Wei06b, XZ09b, YW08, ZLW08, SLZ05a, SZL^{+}07]. th-best [SLZ05a, SZL^{+}07]. th-order [DM08c, GJ07b, HJ08, KEM08, LZ08e, NCÁHC^{+}06, Wei06b, XZ09b, ZLW08]. their [AK06a, AMJD08, AA06d, CL09a, CX08b, CKKS08, CN09c, DSY08, EDMJ05b, GD08, GHZS07, GGOS08, LS08e, LKU09, MR05b, Sal06f, SOS07, VAMV07, WL07a, WWW03, WN08c, XS06, YL08b, YX08, YYMZ09, ZWT09b, ZWT09a, ZTWL09, ZWZ09]. theorem [AB05a, AO05a, Ber09, BRW08b, CL05a, sCLC09, CPW09, Den09c, DZ08d, DD09, DLT08, DH09, JOB08, KES09, Kim07a, Kim09, KT08a, MTM05, RYW06, Ren07, Sht07, SM08a, yS06c, Sun07a, SCO09, Tan06a, Tan06c, Tan06b, Tan06d, Tan07a, Tan07c, Tan07b, Tan07d, Tan08b, WZ06h, WZ06i, WZ06j, cWxChY07, WZ07d, Wu08b, XL08b, bZlDL09, ZW08c, ZW05g, ZZYT08, ZW05j, Bri08]. theorems [AK05a, AK06a, ACOS06, ĆC09d, DZ07b, GG09b, LL07a, LM07b, ML09, Ngô09, NS09, OZ07, OS08, SS07b, SC06b, SL06b, WC09c, Xu09, ZZZ09, ZWC06]. theoretic [Coo09, ESD07a, RMA07b]. Theoretical [Wan08d, DDD05, GBK06, SdJA06, WZ06a, YSCC05]. theories [Dje09, Pin07]. Theory [HMNT09, BLW06b, BZZ08a, BKFM05, EHN07, ESS05a, ESAR05, ESETA^{+}06, Gar07, HR09a, Hua07, Hua08c, JOA05, Kah06a, KK07c, LXL06, Lin05a, Lin05b, LOA05, LXZ07, NN08, Nwo06b, Nwo06c, Nwo07c, PY07, PB07b, Rey05a, SYT06, Sha07b, SCS05a, Sin05, SSR05, SOS07, SYY04, Sub07, Tan06d, Tan07b, TTZ09, Vel06, Vel09, WCO08, WH08a, Wu07a, Yan08b, YY07c, ZHX06, ZC06d, ZWZL06, Zie09, ZJO08, lSAQ08]. therapies [d’O05b]. there [Dav06]. Thermal [EESF04, IHB06, Lin08, Mak07c, Pan08c, AB05c, AH05a, ACZ05, KJ09, LYCW09, Liu06b, Mak09, OMST07, Oua05, PM09, PVS^{+}08, TS09, XL06b, XWSL08]. Thermal-diffusion [EESF04, Pan08c]. thermistor [KE05b, KE05c, KE05d]. thermo [CB06, EESF04, Pan08c]. thermo-elastic [CB06]. Thermoconvective [SM08e]. thermoelastic [FM08, SKS06, Sin07a, Sin08a]. thermoelasticity [Bou09a, EMY05, Sin05, You06c]. thermomechanical [EMY05]. Thermosolutal [GS08b, SAS05]. thermoviscoelastic [SZXL06]. these [Gal06b, Waz08a]. thesis [Gal06b, Smi06a]. theta [Sri07]. theta-function [Sri07]. thin [CB06, Li07c, Soh06]. thin-film [Li07c, Soh06]. thinning [Pen09]. Third [AJKM09, HMT09a, HMT09b, MR07c, MD09b, NNuH07, WZ06j, AO08, AMJD08, ASN07, ABP05, AGH05, AA07f, BD06a, BR07a, Bas08, BBB08, BAM07, CS07h, Chu07b, Chu07h, Chu07k, Chu07n, CN09b, CG08b, DB07c, Du07, FXY08, GC06a, GJ07h, GAPT08, GSDB07b, HR09a, IS09b, JZCW05, JC08a, JH08b, Kan06b, KLW06b, KLW07b, Kou07f, LDL08, Mak07c, Mak09, MKKS05b, MN05, MK06, Net08a, SH07b, Sha05e, Sha07c, SGS09, uIKTT05, WR07b, Yin08b, ZLW06c, ZL08b]. third-grade [Mak09]. Third-order [MR07c, NNuH07, WZ06j, ASN07, ABP05, AA07f, BAM07, CS07h, Chu07b, Chu07h, Chu07k, Chu07n, CN09b, CG08b, DB07c, Du07, GC06a, GJ07h, GSDB07b, IS09b, JZCW05, JH08b, Kan06b, KLW06b, KLW07b, Kou07f, LDL08, MKKS05b, SH07b, Sha07c, uIKTT05, WR07b, Yin08b, ZLW06c, ZL08b]. Thirteen [ZX08, SA06g]. Thomas [Yao08a]. Thome [CI05]. Thompson [BCEGMRMP07]. thoughts [dCD06b]. threaded [dCJLNÁA08]. Three [Akm09, BNR06, Cel06a, CS09b, CFJ^{+}05, CGSS09, NN06c, NY07, NH07b, Smi06b, ABP05, Ari07, Att06b, BSP05, BAM07, CTT08, Che06j, CT07b, CC08d, DM06, DMA07a, fDxZfWyY09, DB07d, DZ07a, EGS05, EM03, Gen09b, GMCM06, GRH07, GBG07, HR07a, HMT08, Jan08b, Jan09c, JW07, Jun09a, KW08b, Kha07a, Kha08a, LKY05b, LS06b, LPS08, LH08c, LLW08, LZ09d, Mar05b, MM05c, MZ07, Moh05a, Moh06, MGMC05, MMC06, NLD08, NS06b, NS07b, PW08a, PC09, Qui09, Ras05c, RNV07a, Sal07b, Sal06c, SS07a, SS09c, SS08a, Sil05c, Sil05a, Sil05d, Sil05b, Sil06g, Sil06c, Sil06d, Sil06a, Sil06f, Sil06b, Sil06e, Sil06h, Sil07b, Sil07a, SY08a, SS06g, SG07c, WD06, WsZ07a, WC09b, Waz08e, Whi06, WLC07d, XW07b, Yan08e, YN07, YNCL08, YGY09, Yun08b, ZLCZ06, ZLZ07, ZCZ08]. three-competition [PW08a]. three-component [GBG07]. Three-dimensional [Akm09, Cel06a, CFJ^{+}05, Ari07, Att06b, CC08d, DM06, DMA07a, Jun09a, LH08c, Mar05b, RNV07a, Sil05c, Sil05a, Sil05d, Sil05b, Sil06g, Sil06c, Sil06d, Sil06a, Sil06f, Sil06b, Sil06e, Sil06h, Sil07b, Sil07a, WsZ07a, Whi06, YGY09, ZCZ08]. three-level [SS06g, XW07b]. three-order [Che06j]. three-parameter [WLC07d]. three-phase-lag [Qui09]. three-point [BAM07, DB07d, DZ07a, Gen09b, Jan08b, Jan09c, KW08b, Kha08a, LS06b, LPS08, LLW08, SS07a, SS09c, SG07c, ZLCZ06]. three-species [JW07]. Three-step [NN06c, NY07, NH07b, fDxZfWyY09, HMT08, MZ07, NS06b, NS07b, YN07, YNCL08, Yun08b]. Three-steps [BNR06]. three-term [EM03, Sal06c]. Threshold [LCZ05b, BCW05, CC05a, CHY05a, Che05g, GGSNR06, HW05a, HW05b, LCC05c, LCZ05d, SCL05b, Xie05, XC05, Yin07, YC05b]. thresholding [HC08c, Ma08c]. throughflow [Mur06]. thumb [NR08a]. TiC [AD06c]. Tie [SZS05]. Tikhonov [CXZL09, QW08a]. Tikhonov-type [CXZL09]. Time [DG07b, DBR07, KM05d, SMBB07, AD05b, Ahm05a, ACP07, Aks05, AASAA06, AN06, Ang05, AA05, Ate06, ALN06, AGS07, APS05, AK06i, BC09a, Bad05, Bai08, Bak05, BK08a, BB06a, BW05b, Bog05b, BS08, CZ07a, CS07b, CY05a, CCL06, Cha05g, CLH06, Che06b, CTG06, Che07c, Che07b, CA08a, Che08d, Chi08b, CLWP09, CK06b, CK07b, CG05b, DD06a, DJM08, Deh05a, DJ07, EM11, Erb06, EHP08, EBP09, EM09, Ere09a, FTMA07, GZG08, GTX08, GZ07c, GMS06, GY07a, GO08, GW07, GZ07d, GZG09, HLCH08, HHZ09, HK09, HBE09, HHP09, HZ09a, HW07b, HL07b, HHC05, ID08, Ibr08, Imr09, IDS06, Iva08, Jav06c, JPYW09, JWL^{+}07, JSL06b, JZW08, KSA05, KOP07a, KAK06, KR09, KMN09, KAT06, Kim06a, KL05b, KM07, KCW07, KKS06, Kos05, KV06, Kov07, KPL08a, KP08a, KP08b, KPL08b, KPL08c, KPL09b, KP09a, KP09b, Lee05a]. time [LWH06a, LWL07a, LP08a, LH05, LRMS06, hLsT07, LmZ08c, LXJ09, LX09a, LZC07b, Lin07a, LC05c, LZA^{+}07, LN08a, LMC08, LC08a, LZWK09, Ma06a, MW07, MZW08, ML08a, hMzLfW^{+}08, hMfWzLpZ09, MS09a, MAD08, MI08, Mah09b, MPS08, Mea05, Mom05a, MO06, Mon07, NCÁHCLP05b, Nak07, NHCLP07, Nam08, Nam09, NM09, Ngô09, NMP08, NJG08, OM06, Odi08, OKEL05, PB09, Par05a, Par06a, PK08c, PKL08c, PJWL08, PK09, Pha06, Psi07, QCSF09, RB08a, RB08b, RK08, RK07a, RC06, SMA06, SA09, SS09b, SS07a, SH09a, SC06a, SP09, SL07b, SFA08, Sin07b, Sin07c, SS08d, SLL08, SFFY08, SJW09, SP06c, SZ08b, SWX06b, SL06c, SWX07b, SM07d, SM09b, TT06a, TT06b, TT07, TZX07, TQC09, Toh05, TTG07, WNCL06, Wan06k, WSJW06, WL07b, WS07c, WZ07a, Wan07f, WZL08a, Wan08f, WLZ08, Wan08e, WSCL07, Waz05c, WRL06, WWC06, WLT07b, Xia06, XwWL07, XW09, XXTL09, XCH05a]. time [XCD05b, XCH05b, XZ07c, XM08, XS09, YB03, YB04, Yam05a, Yam05b, YSZG05, Yan09d, YM08a, YZDW07, YCC08, You08, Yu05, YS06d, YZFL08, ZW09b, ZWX05, ZT07a, ZS08a, ZY08a, Zha08g, ZYT09, ZD06b, ZW06a, ZWY07, ZL08c, ZZW09b, ZLR07, Zho05, ZL06c, ZLL06b, ZSS08a, ZSS08b, ZL08d, ZL09a, dlS06a, dlS06b, dlS07a, lS07b, lS08a, lS08b, Ram07d]. time- [CA08a]. Time-accurate [KM05d]. time-cost [APS05, KMN09]. time-delay [GZ07d, GZG09, JZW08, LH05, LRMS06, LC05c, MZW08, MPS08, Nam09, TQC09, TTG07, XwWL07, XXTL09, YCC08, ZY08a, Zha08g, dlS07a]. time-delayed [LmZ08c, PJWL08]. time-delays [CZ07a, Ibr08, TZX07, ZL09a]. time-dependent [Deh05a, KSA05, KKS06, Ma06a, MW07, RK08, SH09a, SP09, Waz05c, ZT07a]. time-dependented [SFA08]. time-derivative [ZWY07]. time-domain [ZL08c]. time-fractional [LXJ09, Mom05a, MO06, OM06, Odi08]. Time-frequency [DBR07]. time-independent [NM09]. time-indexed [Imr09]. time-invariant [ACP07, dlS07a, lS07b, lS08a]. time-lagged [HHZ09]. Time-linearized [Ram07d]. time-scales [Mea05]. time-stamping [HHC05]. time-stepping [Ang05]. time-varying [Bai08, CLWP09, HK09, HW07b, KCW07, KV06, KPL08a, KP08a, KP08b, KPL08b, KPL08c, KPL09b, KP09a, KP09b, LZWK09, NMP08, Par05a, Par06a, PK08c, PKL08c, PK09, Pha06, QCSF09, SWX06b, SM07d, Toh05, WZ07a, WZL08a, WLZ08, Wan08e, WSCL07, YZFL08, ZWX05, ZYT09, ZZW09b, dlS06a, dlS06b, lS08b]. time/cost [TT06a]. time/resource [RK07a]. timelike [AAABH04, KAK05, Kas06b]. times [AM05, Bah05, BS09b, EMY05, EG07c, EG06b, EG06c, JW09a, KL05a, LC06d, WWP05, Wan06d, WC09b, XY11, ZGH06]. timetables [Mir06b]. timetabling [Mir06a, Azi05]. timing [KT08b]. Timoshenko [GM08]. Timoshenko-type [GM08]. tip [SJ08b]. tissue [BMD08, LT07, SKSJ06, SM07b]. titanium [dCJLNÁA08]. titled [Kah05a, Kah06a]. Toda [ZZYZ06b]. Todd [WB09]. Toeplitz [AH09, Dem06, KEM08, LHL08, MR06b, WDN06, Wu08a]. tolerance [ZXL06]. tolerant [CCWH09, CY08b, CHT06, KH05b, KLL^{+}08, XHZ06, YLL06b]. toll [Chi06a, Chi07e]. tolls [Chi07h]. tomography [KS05a, SL07b]. tool [KNM05]. tools [BE06, Şen07, Waz06m]. top [FX09]. topography [ADH05, ADH06, Ma05b]. Topological [ESD07b, SB09, Akı06, CO09, KSP08, Lee05d, LH08c, Tan06c, Tan06b, Tan06d, YLC09]. topologies [FA05a, Gür05]. topology [DMZ08, Zho07]. TOPSIS [ASA05, ASAEE06, JLI06a, JLI06b, MMAHN08, Shy06]. torsion [Shu05]. torsional [EB05e, MP08a, ZM06a]. torus [AS07b, AAG05]. Torvik [RB05a]. total [AASAA06, CGW06, Ere09a, KK06, KTW08, Liu06d, LC06d, Ma05b, TK06b, WNCL06, WSJW06]. Tournament [NP08]. toxic [CRC08, LC06a, NSS06a]. toxicant [LZ09c, NSS06a]. toxin [GN06a]. trabecular [AS07c]. trace [BŞ07a, Gül08]. traceability [HW05b, WLHH05, WY05d]. traceable [CLS05a]. track [Cel06a, WWHL08]. tracker [TTG07]. tracking [Erb06, Ohm08, jW06a]. trade [APS05, KMN09, RK07a, TT06b]. trade-off [APS05, KMN09, RK07a, TT06b]. tradeoff [TT06a, TT07]. traditional [Waz06d]. traffic [BBCP09, Chi07c, Chi07h, Dah07, GH07, SM08f, Par08, SAB08, ZWM08]. train [Cel06a, ER06b]. train-track [Cel06a]. training [xGZZ08, Mel08, ZZLL07]. trajectory [HC07b]. transaction [CJK07, SYL07a, SYL07b, XXF06]. transactions [Nwo07c]. transcendental [Boy07b]. transdermal [DRS^{+}08]. transduction [LMT^{+}09]. transfer [AEA05a, AEG05, ASH09, AH05a, Ant05, Asa06b, Att05a, Att06a, Che06n, Cor07, ESMM06, EESF04, EEEE05, EO06, HS09c, KK05d, LX06a, LX06b, MHM09, Özt05, Pan08b, Pan08c, Pan12, PVS^{+}08, SK09a, SG09c, VGMV^{+}07, WY06, XL05, ZW05d, ZMV05]. transform [AO05d, AO06e, AO06d, AS09c, BKBK06, BMY08, CC08c, CC09c, Che07h, CDT07, DSY08, GZL08, GRK07, GVMT07, HGN07, Hub08a, Hub09b, LMZ08a, LPP08, LS07d, MMR09, NK05b, OME08, PG07a, Sok08d, Su07a, Su07b, Su07d, WS07g, YLB06, ZT07b, ZZDL07]. transform-based [NK05b]. Transformation [BA09, AM08a, DZ06, FHYZ09, Faz09, GS08a, Gor06b, Kad07b, Kuo05, KA06f, Men08, ÖÖ06, PR09b, Sha07a, Thu07a, Waz09b, WX06, WZF08, YAO09, YJW06, YWJ07]. transformations [ABM09, DM08c, ET05a, HR09b, RLY06, SB06b, Sid05, Sid06, ZZ07c]. transformed [XBZ06, XZB06]. transforms [BK08c, BDY06, FBB08, Tak07a, WZ09b, Yür07]. transient [AGT07, ALN06, BC09a, BWB05, CN05, CMC06, CKS09, KS08f, LX06a, Oma06, OIL09, RM05, Sel05]. transients [FTT^{+}09]. transistor [Li07c, TCDV07]. transit [WS07c]. transition [ESAH10, HC07b, LW06a, RY05b]. translating [GG05a]. Transmission [TS05, ADH05, BLLM08, CTGD08, DXH08, MS07e, NTO06, ST06c, YF08, ZLMW05]. transmitted [SL07b]. transmural [AT07b]. transport [AT07b, BLW06b, CD07, DC07, EA07b, EESESD05, EEEE06, HS09b, ICM08, KM06, Mei06, RM05, RM06, SW09a, SG09c]. transportation [AY05a, ER06b, LLFM08, Liu06d, SM06a]. transshipment [Cho07b, Rey05a]. transversal [VR08, XSKC07]. transversality [MV09a]. transverse [CZ05b, CZZ05, ST06c]. trapdoor [JSW05]. trapezoid [SNH07c, UE07]. trapezoidal [IT06b]. Traveling [AKARA08, KD09, LL08d, PW09, BK07b, Bou09a, DSZ08, ESK05, FT07, GL08b, lH08a, Ji08, LTZ09, MYG08, MWQ09, SXJ05, SLC09, WCZ06, ZXLS06, ZWX08, ZY08d]. Travelling [CLDH09, Den08d, Lin09a, SXX05, TZH09a, TXW09, TZH09b, TB09b, Waz05q, Waz06o, Waz06p, WP06, Zha09c, ZZTJ06, ZX07b, Boy06b, BKP06, CW05b, CZ08b, Dem05c, Den08b, FLLH07b, FHL07, GA08, GHL07, HRLC08, HRCL08, LSL06, LZ08b, LXML09, Lou07, LHF07, MLW09, MAY09, Par09a, Pen08, Rom09b, SLX05a, SX06, SMG07, Tan05b, TL06, TH07a, TH08, WCLZ05, WZL08b, WW08b, Waz05d, Waz06f, Waz06e, XGTS07, Yan08a, YB07, ZX06c, ZZB08, ZDY08, ZZYZ06a, ZL07b, ZLZ08]. travelling-wave [Par09a]. treatment [AAK07, CMMC09, EB05c, EEAES06, EL05a, Kwo07a, LLG09, RN09, Waz05b, Waz05a, Waz05o, Waz05l, Waz06i, Waz06m, Waz07j, Waz09f]. tree [AS07a, GL05a, MKSK07, PG07a]. trees [Che07h, ESD07b, FZN07, HH07a]. tri [BK09, CYL^{+}08, EM05a, ZH07b]. tri-axial [CYL^{+}08]. tri-diagonal [BK09, EM05a]. tri-neuron [ZH07b]. trials [Ery06]. triangle [KS06b]. triangles [MR07a]. triangular [ÁFVZ06, BMS07, BMS08, GY06, HM09a, mH06c, Rab05, RNV07b, RNV07c, SL09b, XCFX05, XC07, YJ09]. triangulated [FS09]. triangulation [MCGS05, Pen09]. Tricomi [KH08]. tridiagonal [AR07a, CPTZ05, CL06d, EM03, EMR08, ES05b, EH09b, GG08b, GG08e, GG08f, HE08a, Han07, Kar07a, Kay05d, Kay05e, Kıl08a, KT07b, LR07a, LR07b, LXZ07, MQ07, RH06, Rim05g, Rim05a, Rim05h, Rim05e, Rim05d, Rim05f, Rim06g, Rim06f, Rim06h, Rim06a, Rim06b, Rim06e, Rim07c, Rim07g, Rim07h, Rim07d, Rim07f, Rim07a, Rim07b, Rim07e, Rim08a, Rim08b, Rim08c, Rim09b, Sal06c, Siv07, Sog08d, Sog08b, WS07h, WS08d, Yin08a]. Trigonometric [WLP07, Kim08, MCS08, WL05d]. trigonometrically [FW07a, MKS09, Sim09b]. trigonometrically-fitted [Sim09b]. trinomial [AS07a]. trio [BY05]. tripartite [SW05b]. triple [WYL08a]. tripotency [ÖSÖG09]. tripotent [ÖSÖG09, SÖ08]. Trisection [Dem08]. trivial [LS06d]. Troesch [FMH07]. trophic [BSP05]. trophic-level [BSP05]. tropical [BC09c]. truck [KO08]. Truncated [LY07c, Che09b, Lan07]. Truncation [ZZ05b, Küç05b, QFXW06]. trust [Che06m, FHSS08, FS05, JjQjWmL06, LL08c, LN08b, MSD06d, MZW05, MHA06, yNfM06, hPY06, jW06a, WZT06, WZWW08, YQ09, ZL05a, lZ06d, Zha07b, Zhu06a]. trust-region [Che06m, FHSS08, FS05, JjQjWmL06, MSD06d, hPY06, WZWW08, ZL05a, Zhu06a]. trustworthy [CCH05]. Tseng [Hwa05, RIY05, ZAX05]. TSP [ZBH06]. tsunami [ZPT^{+}06]. tsunamis [ADH06]. tube [HA07, NMS06, ÖŞ06, SI06]. tubes [HAAS06, KM07, SH07c]. Tucker [BGR05, LHWW07, RAS08, SLZ05b]. tumor [BT05, EP07, IZIO09, MKA07, WT08]. tumors [TG06a]. tuning [CH07f, PGdSS08]. tuning- [CH07f]. tunnel [SC06c]. turbine [KJ08, VRG06]. turbine-induction [KJ08]. turbulence [LS09b, LWL06]. turbulent [Cha05h, Che09a, CLNCG07, Hu05]. Turing [Zie09, Bou06c, Dil05, Gal06b, Myc06, ND06]. turning [Att06d, Jah07, Ram05d, Sok08d]. twelfth [SA06g]. Twersky [KS08c]. twin [WLA09]. twin-parameters [WLA09]. twisted [Sim07]. Two [AAS06, ANM07h, ANM07j, ANM07i, AKH06, BW05b, CQL08, CJP08, GG06, GPT09, HAK07b, Ham05c, HRM07, HC09, HTO06c, JGTK07, JZCW05, JZX07b, JZX07a, Kar07b, Kar07a, Kwo07b, Li05f, Li05g, MMM06, MCC07, NAJ06, PC07a, PSW07, qQcM05, Ram06a, SY07c, TWZ07, WHL06, WL09b, Waz09e, Wu07b, XFQ06, ZW05d, AM06, Ahm09b, AAAA06, AMAP07, Arg09, ADK05, ADK06b, AÖ09a, Att05b, BAF08, Bak05, Bel08a, BR08b, BB05, BK05a, BB09, BB06d, Bra06, CCE06, Cai06a, Cen06, CC09c, Che06a, CTG06, CZ08b, CL09b, CG08a, Cho05, Cho07c, CD09, Chu07n, CPTZ05, Cvi08, Dan08, DM07a, Deh05b, DCS07, DC06b, DF08, EFE08, EB05d, EMY05, ESES06, ESK07, Eli07, EHL08, FK09, FHG07, GTX08, GKD07, Gül05, GWJ07, GM06b, HFL06, HM09a, Has06c, mHzC05, HL06b, HLW08, Inc07b, Ism06, IMS09, Jan09b]. two [JK06, JZ06a, JC09, JZ06c, JV07, KK07b, KY08, KYK08, KK08b, KC05a, KAS06a, KR05, Kan07, Kar05a, Kim05b, Kum03, KA06e, Kum09c, KT07c, LJ08, Lee05d, LR07a, LR07b, LW05a, Li05c, LC06a, Li06b, Li07c, LCZT08, LF08, LOZ07, Lin06c, Lin07f, LX07b, LZW07, gLL07, LFC08, LTCC08, LZ09c, LLL09, LC08b, Luo05b, cMpZfR06, Ma07a, MX08a, MBM06, MJ05, MK06, MA06e, MS06b, MA06f, MD09b, MK08b, MK10, MBM07, MB05b, Mou08, OB09, ÖSÖG09, PP07c, QL06, QLG08, QS09, RAES05, REDES05, RLZ08, Ram05i, RM06, RMG07a, RVN07, RNV07b, Rem07, RM08, RWB09, Rey05b, SH07a, SÖ08, SMBB07, SK08b, SLC09, Shu05, SAM07, Sid09, Sil07b, Sim07, ST06c, Sog07, SZXL06, SJW09, SL06b, Su09, SA05b, SA06h, Tan07d, TS05, TT09, Tua13, VFF08b, Vyn07, WL05a, WJ05, WB06, yWjL06, WNCL06, WZSZ06, WL06a]. two [WTZ06, WLLL07, WL07a, Wan07e, WCH08a, WSL09, Waz05j, Waz06g, Waz07k, Waz08p, Waz09d, WWT05, WWC06, WLC07c, WLL07a, XXTL09, XL05, XS06, YVU06, Yan06a, YEM06, Yan09d, Yig08, YC07b, YCYL05, Yuj07, Yun07a, Yun09, ZWL06, Zha06b, ZM08b, ZYH08, ZM08a, Zha09d, ZZKY05, Zha07h, ZC08b, ZZL08, ZGL08, ZWZ09, ZW06f, ZCFX07, ZL08e, Zhu09a]. Two-axis [Ram06a]. Two-body [CQL08, MB05b]. two-chamber [LJ08]. two-commodity [Lin06c, Lin07f]. two-competitor [Yan09d]. two-component [Waz09e]. Two-dimensional [GPT09, HC09, TWZ07, Bra06, CC09c, CG08a, DCS07, DC06b, EFE08, EB05d, EMY05, FHG07, GKD07, Gül05, HFL06, Has06c, Ism06, IMS09, JZ06c, Kar05a, Li07c, Ma07a, OB09, QS09, Rem07, RM08, SK08b, SLC09, Sid09, XL05, Yig08, Yuj07, ZZL08]. Two-direction [SY07c, Yan06a]. two-disease [BK05a]. two-echelon [Rey05b]. two-end [Shu05]. Two-grid [qQcM05, WHL06, CL09b, HL06b, cMpZfR06]. two-layer [DF08, WLLL07, Zha09d]. Two-level [Ham05c, Deh05b, Li06b, SH07a, YCYL05]. two-machine [WNCL06]. two-parameter [Chu07n, EHL08, KY08, LX07b, WWT05, WWC06, WLC07c, WLL07a]. two-phase [Kim05b, Su09, Vyn07, ZCFX07]. Two-point [AAS06, BW05b, CCE06, Cen06, CTG06, HLW08, Jan09b, JC09, KK07b, KYK08, KK08b, KR05, Kan07, Kum03, KA06e, Kum09c, LZW07, MK06, MA06e, MD09b, Ram05i, RMG07a, SAM07, SA06h, WJ05, WCH08a, WSL09, ZC08b]. two-prey [LOZ07]. two-product [YVU06]. two-sector [Cai06a]. two-sided [gLL07]. two-species [Che06a, LTCC08, LZ09c, MBM07]. Two-stage [HTO06c, AAAA06, Cho05, WB06, Yun07a, Yun09]. Two-step [CJP08, NAJ06, Inc07b, LW05a, Li05c, LX07b, Luo05b, PP07c, RWB09, ZWL06, ZW06f]. two-storage [MBM06]. two-term [KT07c, Sog07]. two-variable [GM06b]. two-way [CPTZ05]. Type [Tan07d, AÖ05b, AB06, AMAP07, AM09a, BFR07, BUP07, BKBK06, BS07c, BKZ06, BFKT07, BR07b, BMY08, CFX05, CMGMD06, CL05d, Che06m, Che06i, Che06c, CG07b, CS07g, CS07f, CY08c, CA08b, CXZL09, DB07c, DNS08, Deo08, DSY08, Djo07b, DHG09, EI09, EDMJ05c, EHR09, FHH09, FM08, FHL07, FFRF07, FFRFCC08, FFRFCC09, GZG08, GL09, GAR09, GJ07h, Gor06a, GM08, GT05a, GL05c, GF09, GAG09, HK09, HEDMJ06, HMJED06, HZ09a, HWC07, HCC06, HCZ06, HW06, HTM09, Ibr07, Jai07, JYX06, JS05, Kah06b, KB06b, Kay05c, KAPM07, KBÖP07, KT08a, KPL08c, LO08, LR07a, LR07b, xLyW06, LLZ06, LS06c, LS09d, Li09c, Li09b, LS09f, LS09e, LS09g, LX09a, LZC07b, Lin07d, Liu08b, LX09b, LP06a, MA08e, MMA07, MS08a, MTM05, Mao09, Mar05b, MJHED06, MJ07c, MX07a, Mug05, Nak07, NAM^{+}07, OZ07]. type [OISM07, OSY07, PK05c, PK08b, PKL08a, PPKL08, PKL08c, PK09, PMS09, PW08b, PHP09, Pog07b, QW08a, RB08b, RWB09, Rim05g, Rim05a, Rim05h, Rim05b, Rim05e, Rim05c, Rim05d, Rim05f, Rim06c, Rim06g, Rim06f, Rim06d, Rim06h, Rim06a, Rim06b, Rim06e, Rim07c, Rim07a, Rim07b, Rim08a, Rim08c, Rim08d, Rim08e, Rim09a, RFFFCC08a, RFFFCC08b, SMA06, Sal07a, Sal08a, Sal06b, Sal06f, SA09, She07a, SS09d, SL09b, SP09, Soh06, Son08a, ST08, SSR05, SFG07, Ste08, Ste09a, SW09c, SiU09, Ste09f, Ste09d, Ste09g, Ste09h, Ste09i, SS06g, SV07, Thu06, Thu07b, Thu08, Van07, Vul07, WG08, WZ06b, WL06c, WS07b, WZDgH09, WZ09a, Waz05a, Waz05c, Waz05i, Waz05k, WW05b, WS07h, WWT05, WLC05, WY05c, WLC06, WLC07b, WZ07d, WS08e, XCH05a, XCD05b, XCH05b, XLL06, XM07a, Xu09, YZW07a, Yan07c, Yan09c, YLH09, You06a, ZMW05, ZHLG07]. type [ZHL07, ZTWL09, ZXWL06, ZW08b, ZM08d, Zhu09b, ZSL05]. types [Ism05, Lin06c, SS08c, TS08a, Yun07a, ZZZ09]. typical [OMR07b]. Tzeng [QCB05a, BCW05, Hsu05a, HL05d, Xie05].

U [KOP07a]. U-lines [KOP07a]. UDV [YL08b]. UHG [ESD07a]. Ulam [JR07]. ultra [SY07a]. ultra-hyperbolic [SY07a]. ultralow [PVS^{+}08]. umbrella [dCDLN^{+}07]. Un-desirable [KA05c]. un-reliable [KL06b]. un-stirred [ZGL08]. unanimity [KI08]. unaware [TNS07]. unbalanced [Kum06]. Unbiased [HL08b, Wan06a]. unbounded [AD06a, BR08a, BHN08, DC06b, GS07b, HR07a, HR09a, Küç05b, LMY07, Ma07a, QM05, SLW08, SW09d, Waz06h, ZLW08]. Unboundedness [LeT05]. uncertain [BKA07, Bog05b, CC07b, CAL08, CLWP09, HCLP07a, HCLPJL^{+}08, HTO06b, ID08, JPYW09, KP06b, KP06c, KPL08a, KP08a, KP08b, KPL08c, KPL09b, KP09a, KP09b, LYL^{+}08, LLZ08, MZW08, Par07, Par09b, QCSF09, SWX07b, WLZ08, XW09, ZZW09b, ZWLX08, ZL09a, Zhu09a]. uncertainties [HHP09, Nam09, PK05d, RAG^{+}07, WJZG08]. uncertainty [AK05h, EMP08, HCLPJL^{+}08, Mir08, NCÁHC^{+}06, RR06, ZGK06]. unclassified [CL05b]. uncommensurate [dlS07b]. Uncomputability [Vel09]. uncomputable [Smi06b]. Unconditional [SYH07]. Unconditionally [GC07a, DZ09, Moh05a, Moh05b]. unconstrained [And09, FYZ07, FS05, HYG08, Kah06b, Kah07, KC07a, KLS09, LZ06c, LY07c, LN08b, MZW05, PC07a, RW08, Ren08b, lSgPpJ07, jW06a, WZT06, WZ06f, WN08a, WZWW08, WLQ06a, WLQ06b, WS06b, Wu08d, YS06b, YDL07, YZW07b, Zho09a]. uncountably [LKU09]. Undecidability [Tan08b, Vel09]. undeniable [CHC05, GT05b, LCZ05b, ZSNS05]. Undesirable [JLS^{+}05d, VMK05, AKS06b, HV08, JVFM04]. unearthed [CSBS07]. unequal [HW05c]. ungauged [Ma06b]. uni [Yan08a]. uni/bi [Yan08a]. uni/bi-variable [Yan08a]. unidimensional [Kah05c, Kah05b]. unified [BB09, CW06d, HR09a, HN07c, HLY09, Lee05d, RSSS07, SK06a, SS09d, SSR05, Tan06d]. Uniform [AKD06, AŞ06e, FeT05, MS09a, WX09, WW05b, Yan06c, Yan08d, ZZL08, AEA05a, AES05, AH06d, AM09a, Guy09, HBVM07, HL06c, KSA05, KA06b, KM09a, KNM05, KS08b, KA06e, Lep08, MH06a, Moh06, MA06e, NMS06, Occ09, Pan08b, PK08a, Pat07, WFS08, ZGH07]. Uniformly [Bra05, YC06b, BP06a, Dod08, JK08, Kan09a, KS08b, KKKM09, NO07, NAUH09, RSSS07, KP06a, KPS06a, PS06b]. Unifying [TTZ09]. union [Pal08]. uniparametric [JYX06, KL07h]. unique [ESE05a, Sch08, ZLW08]. Uniqueness [LDL08, BC09b, CL05c, DZ08f, Fus08, GL09, Hlo05, Liu05b, PK08a, RX09a, RX09b, SLC09, WWD09, WG07b, YJO08]. unit [AS09c, Fuk06a, GOPP07, IR06, Jön05, KJ08, LS09d, LS09h, Liu06c, fQZgH08, Ste08, SW09c, SiU09, Ste09e, Ste09h]. unit-root [Fuk06a]. Unitary [Vic09]. units [AJK05, AS06g, AGA^{+}07, HK05, JPZ06, JM06a, LJK07, LJE07b, RGA05, RGA06, SML05b, SH09b, WC06a, WL06e]. univalence [Abd09]. univalent [AK09, Den07, Kim07b, WW09]. univariate [RAN07, SG05d]. universal [DES07b, Del09]. Universality [NM09]. unknown [Bar05a, Dyk09, EPER07, FSG07, Fat05a, Fat05b, Mah09b, SW09b, XXTL09]. unlinkability [WHH05]. unmodeled [WLP07, lS08b]. unrelated [JGTK07]. Unreliable [WS05a, CD09, KL06a, WY09]. unscented [HCLP07b]. Unsolvability [Zhu06b]. unstable [LLZ06, dSLV^{+}08]. unstaggered [Tou09]. Unsteady [Att06a, ASH09, AM08a, Bat08a, Ere06, GM05a, IHB06, ICM08, KS08b, MCMA07, Tou09]. unsupervised [ABK^{+}06]. unsymmetric [PW08b]. untraceability [LHY05, WW05c]. update [AMAP07, AP09, CK06a, ZHX06]. updated [DY07]. updating [Yua09]. upon [MY07b]. Upper [JLA07, LZZ06b, PB09, ÁFVZ06, DC06c, KW05a, LHL06, MYSA05, MS08b, MD06c, Wan06b, WLK08, WCO09, ZL07c]. upper-bound [MYSA05]. uptake [NSS06a]. Upwind [RK08, Tan05a, Bi06, CFX05, GY06, Li09f, SH06a, SY08a, XCFX05]. upwind-mixed [SY08a]. Urysohn [Dar06]. USAOR [ZHGG08]. Use [Bha08, SCR08b, Wan06b, BK09, FTT^{+}09, MAN06, MBC08, Rig08, SM06a, TF07, WHC08]. used [WZ07c]. user [CH05b, Hsu05b, HZYL08, LKY05a, RY05a, SZS05, TWL05, WLT05b]. user-friendly [SZS05, WLT05b]. Using [AL09, AASAA06, BMS07, BMS08, BKA07, BWB05, ETETH07, GANG07, Isk07, JLRB05, JLBR08, Jön05, KBTS07, Liu09d, MM05b, MK05d, MKR06, RAG^{+}07, SCR08c, WH08a, Wu07a, YWJ07, ZH05, AJAH06, ASL07, AE07a, ABK^{+}06, ACP07, AS06b, AKM05, AKARA08, AM08a, AR05, AA07b, All07a, Ant05, Aou08b, AO05d, AO06e, AO06d, ADK05, Asa06b, Asa07b, AÖ09a, AÖ09b, AK06i, BF07a, BF07b, BK07a, BSKK06, BK07c, BK08b, BCL05, Bay05b, BE06, BNdFPS06, BB06b, BTPGAl06, BQYW07, BKFM05, Boy07a, Boy08a, Bri08, BEvdB07, CC07a, ÇFÇ06b, ÇFÇ06a, Çel06b, Cen06, Çet06, CN06, Cha07b, CCH05, CL05b, CKY05, Che07h, CW07, Che08b, Che08c, CYL^{+}08, CLW09b, CH07f, CT07c, DGJ07, DIS05, Dai07, DC08, DSMK08, Dam07, Dam08, DX07a, DKK06, DEH05c, DDD05, DH06e, DHS07]. using [Dem05a, DT06, DLD08, Dön06a, DK06, Dön06b, DF08, DjT06, Dua06, ES05a, ET05a, ET05b, Ere06, Fab08, FAQ08, Fro06, FWJK09, GZL08, xGZZ08, GC06b, GL08a, Gen09b, GKB07c, Ghi06, GS06b, GS06c, GCJ08, Gop07, GEG07, GRH07, GM05b, HEZ08, HLCH08, HC07b, Har06, HJB06, HBJ06, HIST06, HED06a, HED06d, Has07a, HBES09, HBE09, HRCL08, HCLP07b, HW09, HC08b, HYL09, HW06, HS09b, Imr06, Inc07d, IDS06, JDG06a, JMLR05, JLBR07, JG07a, JLCH08, KK07b, KC05a, KGB06, KBLT08, KNM05, Kay05a, KS05a, KH08, Kim05a, KSC^{+}08, KTÖA08, KSKK06, Kös06a, KM05d, Kum07, KB07, LFM06, LKY05a, LS06a, Lee06b, LP07b, LM08a, LS08c, LJ08, LAL07, Li07c, Lin06a, LC06c, LXZ09, LLWC07, LOA05, LLC06a, LWK05a, Ma05b, MLW09, MC07, Mak07b, MO05, MY05, MB06, MA07d, MK05c, MM05a, MA05d, MR06b]. using [MA06c, MD06b, MY06b, MAR06, MLR07, MS07b, MD07a, MSR07, MS08a, MTM05, MH06b, MH06a, MB07, MBDD08, MRN09, MFE05, Mir06b, MKA07, MA06f, Mom05b, NCÁHCLP05a, NGLHCLP06, NHCLP07, Neh07b, OHT05a, OISM07, PY07, PBKM09, PC05, Par08, PRGH05, PN06, PK08d, PG07a, PG07b, PFA07, PDA06, QS09, QZJ07a, RLY06, RMA07a, Ran08c, Rao05, Rao06b, RZ05, Ras05a, Ray06, Rem07, RMA07b, RLdLSF07, Sad07, SJ08a, SJ08b, SKHZ06, Sal07c, SAS07, SB06b, SLLL07, SS08a, SH07d, Sha05b, SHK06, SK06b, Shi05b, Shi05c, Shy06, SA06f, SA06g, SA07a, SA07b, SA07d, SA07c, SA07e, SA08e, SS07c, SZ07b, SZ07c, SGG^{+}09, Su07b, Su08a, SH09b, TT07, TB09a, TB09b, TSP07, TMASA05, TTZ09, TA05a, TLH05, TK06c, Tsa05, Tso09, Utk06, VRG06, WY05a, WC05a]. using [Wan05e, WLT05a, Wan05f, WL06e, WZ06e, WLLL07, WSL09, Waz06b, Waz06a, Waz06l, Waz06p, WdRS07, WS07f, WL08c, WLCL06, WC05d, WYL06d, Wu06a, WZ06j, WLLC07, XZY08, YW05, YM07b, YJW06, YCC08, YRY05c, You06b, YB06, YK06a, Yun07a, Yun09, ZGK06, ZW05a, Zha06a, ZBH06, Zha07f, ZCMQ07, Zha07g, ZZW08, ZW05f, ZS09b, ZCFX07, ZFL^{+}09, ZLDW09]. utility [CF06a, CTL09, Nwo06g]. Utilization [CDT07, MJFAS07]. utilized [WC06c]. utilizing [Asl09b, BS07c]. Uzawa [CJ08, LD06b, LW06c, LC06b, Tan08b, ZZ09].

V [AJRLRGOG09, Shi05b, Shi05c]. V-KT-pseudoinvex [AJRLRGOG09]. vacation [Cho07c, GS06f, KHL06, hLsT07, MAR05a, Wen05]. vacations [BGP06, KL06b, hLsT07, MX08b, MAR05a]. Vaccination [d’O05c, GTX08, JY09, JWS^{+}09, Mak07a, MCC07, MC08, RYE07, SJC09, WTS09]. Vakhnenko [MLW09]. valent [ARS07a, Aou07, Aou08b, MA09, NMM09]. valently [Alt07b]. valid [Gal06b]. validation [ZZX05]. Validity [Ozd05, AÖ09b]. Valuation [KF08c, ZW05f]. Value [Haa09, LWG09, Afr05, AN05a, AOLL05, AOS09, AS06a, AS06b, AS05, ASN07, AGPJV09, AO05d, ADK05, ADK06a, ADK06b, AA06e, Att05b, AL06, Baw05, Bay05a, BUÇ07, BLLM08, BBB08, BS07b, BR09c, BD09b, BAM07, BT08, CCE06, CC06a, CC07a, Cen06, CMC06, CC06b, Cha05a, Cha05b, Cha05h, CTG06, CCJR07, CHW07, CX09, CH06c, CZ07c, CG08b, DKA06a, DSZ07, Dem05b, DMH05, DXH07b, DZ07a, Du07, DZ08f, DLT08, DHG09, Dyk09, EHN07, EG07a, EG07b, EM07, FSG07, Fra09a, GB07, GC07b, GL08a, Gen09a, Gen09b, GJ07c, GW07, GHX08, HEZ08, Has06d, mHzC05, HH08a, HJ08, HAU05, HLW08, JDG06a, Jan08c, Jan08b, Jan09b, Jan09c, JS07a, JGY07, JG07b, JGZ07, JFG08b, JZ06a, JZ08c, JC09, JC05, KA05a, KA05b, KK07b, KK07a, KY08, KK08b, KS08a, KB09, KW08a, KW08b, KWX08, KR05]. value [KB06a, Kan07, KKA06, Kha07a, Kha07b, Kha08a, Kha08b, KM05c, Kum03, KA06e, Kum07, Kum09c, Les06, LS06b, LS06c, LC07, Li08a, Li08d, LCJ08, Li08b, LS07c, LPS08, LDL08, Liu06h, LZW07, LLW07, LLW08, LLW09, LA07a, LST09, Ma05c, Ma07b, Mam09, MD09a, Meš07, MJ05, MK06, MA06e, MS06b, MA06f, MD09b, MMN06b, MAO06, MN07, NKE05, ND07, Ngô09, NMD07, OM06, Odi09, PDW06, PFG08, PB09, PC08, PCS07, Psi09, Qua07, RLZ07, RLZ08, Ram05f, Ram05g, Ram05i, Ram06g, Ram07a, Rao06a, RK07b, RG05, RMJG07, RJM07a, RMG07b, RMJ07, RMG07a, RGM07, RJF07, RSZ08, SF07, SA06a, SS07a, SS09c, SB06b, SH06a, SR06, SM06c, SA06f, SA06g, SAM07, SA07a, SAN07, SA07b, SA07d, SA07c, SA07e, SAE08, SA08e, uIKTT05, uITA06, uIK06, uIT06, uINTK06, uITiHK08]. value [SWX06a, SWZL07, Su08c, SL06c, SG07c, SQG08, SMG08, SCO09, SA05b, SA06h, TJG08, TiHuI08, WB06, Wan06b, WC07a, WZS07, WS07d, WC07b, WLK08, WTW08, WCH08a, WLW08, WH08b, WLW09, WSL09, WX09, WP05, Wei06a, Wei06b, WZ08e, WW05d, XL08b, Xu07, Xu08a, XZ09b, Yam08, YS07, YZY08, YW08, YWL09, YZ09a, YZ09b, YC07a, YEA07, YJO08, ZLCZ06, ZW07b, ZL07a, Zha08f, Zha08e, ZYW08, ZS09a, Zha07h, ZC08b, Zha09h, Zha09g, ZLC07, ZHZ07b, ZJO08]. Value-at-Risk [Haa09]. valued [AOS09, AR06, AKA07, BM06a, GBG07, KB05a, yLcCsL09, LFW08, Noo06b, Van09]. values [CTL09, DWT05a, LCZT08, MS08b, QCDS09, SMD06, ZWS08]. Valuing [HY07a]. valveless [LJ08]. Vandermonde [EF05, EF06, EFI06, XL06a, ZWS08]. vapor [LWH06b]. VAR [Nwo06d, ZC05a]. variability [Ben07b, PVY09]. Variable [CP08, HEZ08, ORG^{+}09, WJZG08, AH05a, AKM06, AEA07, Att06a, BS09b, BTMB09, Bou06b, ÇE06b, CCL06, Che09b, Chi08c, CST07, CG07c, DXH07a, DSZ08, FK06b, GS05, GM06b, HBE09, Hon09, HLY09, JK07, JSL06a, KRRGT07, KA06d, Lee05a, LLAT09, LLL09, Mak06, Men08, zMsCxL08, MGG^{+}09, MH08a, MJ05, Moh05a, Moh05b, MK06, Moh07a, MD09b, Moh09, Mom05b, MTL08, NLD08, Oma06, PBG07, Par06b, Par07, PL06c, PFL09, Rao06b, RGAFVA07, SNH07b, SAD06, SS09d, Sid05, Sid06, Sim07, SZ07e, TM08, TW09, Waz06o, Waz06p, Waz06q, Wei09, WZM06, WLT07b, XMGW09, lYqL06, lYqLW07, Yan08a, YX06a, YWZW07, You06c, Yu07, ZS07b, ZW07b, ZDY08, ZH08b, Zhu08]. variable-coefficient [CG07c, Hon09, LLL09, MGG^{+}09, TW09, XMGW09, ZDY08]. variable-order [LLAT09]. variable-period [YWZW07]. variables [Bel08a, BS05c, CT06, Gee08, KC05a, LBS05, QLG08, RZ08, Sal07b, SZL^{+}07, Wan07e, WC09a, XS06, ZM08b, ZM08a, ZWZ09, Zhu06a]. variance [ARM^{+}05, CD07, KC06a, Vaj08, WS07g, XXF06, YML07]. variances [HW05c]. variant [AZ08, Bas08, FLLH08, GSDB07b, GI08, HGN07, KLW07e, LHH05a, MGG^{+}09, NV07, RAZ07, TXW09, WKL08, WZWW08, Waz05e]. Variants [CT06, CT07c, GAV08, AHE05, BD06a, Chu07j, CH07e, Chu07l, Chu07m, CH08b, Faz09, GSDB09, KLW06d, Kou07a, Kou07d, KW07, KL07g, LLC09a, SGS09, Waz05g, Waz05o, Waz05j, Waz06f, Waz06k, ZWZZ09]. variate [AH08a]. variation [CGW06, GBG08, KK06, Ma05b, NM08b, WL06a, WZ13]. Variational [Che08a, IR08b, Mam09, MAO06, NMD07, RDG07, AV09, AB05b, ASK05, AKARA08, BMS07, Bar09, BNH07, BG07a, BN06, BNR06, BNK07, BN07, BN08, Bou05, CY07a, CLCK09, CS09b, CDP06, CKQ08, DF09, DW09c, DMZ08, FB07, GL08a, Han06, HG09, HN07b, HN07c, Ide06, JK08, JB08, KB05a, KK07d, KKA09a, KKKM09, MK05a, NHKZ06, NB07, NH07a, NY07, NH07b, NH07c, Noo08b, QN08, QSZ08, Ram08e, RMA07b, Sad09, SS06a, Ver07, WC05a, Wan05e, WZ09a, Waz07k, Waz07j, Waz09f, WS09b, YHS09, YNCL08, YM07b, bZlDL09, ZW08b, ZY08c, ZW05i, ZCH07, ZH08d, ZH09b]. variational-like [KB05a, KK07d]. various [Ahu08, GK09, LZ09e, SLLL07, Yun07a, Yür07]. Varley [PMS09]. varying [Bai08, BAK07, CZH07, CLWP09, HK09, HW07b, JSL06b, KCW07, KV06, KPL08a, KP08a, KP08b, KPL08b, KPL08c, KPL09b, KP09a, KP09b, Lin06c, LZWK09, NMP08, Par05a, Par06a, PK08c, PKL08c, PK09, Pha06, QCSF09, RB08a, SWX06b, SM07d, Toh05, WZ07a, WZL08a, WLZ08, Wan08e, WSCL07, YZFL08, ZWX05, ZCMQ07, ZYT09, ZZW09b, d’O05b, dlS06a, dlS06b, lS08b]. varying-time [JSL06b]. vascular [MKA07]. VC [CG06]. VC-dimension [CG06]. Vector [MBDD08, WFX07, Amm05, BD05b, Cha05f, CX09, HZ06b, IM08, Kas06c, Kas09, Kat09, KSKK06, Kös06a, LY09b, LCC05b, LFW08, MD05, NR07a, NRA07, PG07b, PG07c, SRMRGRL08, Tse05, TML07, UP07a, WY05a, WN08c, XY07, YF08, YH09, YY06a, YES07, YGY09]. vector-value [CX09]. vector-valued [LFW08]. vectorial [TM05]. vectors [Çak08, Dan08, DES07a, SML05a]. vehicle [Kos05, LZL^{+}07, MWW05, TMSG06, TMSZ06, WS07c, ZL06b]. vehicles [Rig08]. velocimetry [Ohm08]. velocity [AH06d, BST06, Cha05g, Cha05h, GJ05, TY07]. venerable [Cha05h]. ventricular [SANN07, SSTD05b]. verifiable [CHY05a, SC05a]. verification [BCW05, CC05b, Has06d, HL05c, Jwo07, LHH05a, YY09]. Verified [Neh07b]. verifier [LKY05b, LC05d]. verifier-based [LKY05b]. verifiers [LWK05b, YY05b]. verifying [BLH06]. version [EGX05, KKT07, LGHS06, ST05a, mTxWyL07, Wu08c]. versions [AKB09a, PSD07]. versus [LV05, Lak05, SA08c, SC07e, ZQ08]. vertex [Cas07a, SD06]. vertex-cover [Cas07a]. vertical [Che06n, Che09b, Ece06, JK07, NTO06, SG09c]. very [Pet08b, XBZ06]. Veselov [WLZ05]. via [AO05a, Asl09a, BMS07, BT09b, BBCP09, BO06, Çet07, CK05, Che07d, CCZ08, DLT08, DMZ08, EHAR05, EMC05, FE07, FX09, Fig08, FB08, GM06a, Haa09, HHYW07, HW07b, HL06c, IPPT05, JjQjWmL06, JZ08b, KA05b, KYK08, KK08b, KR07c, KR07d, Kos06b, LMT^{+}09, LYL^{+}08, LY08c, MM07c, MKA05, MG06, MT09, NPGW09, Ord06b, PB07b, PBG07, PKL08b, Psi07, QZJ07b, SMA08, SKT06, SCM09, Sid05, Sid06, Sin08b, Sol05c, SCO09, Thu07d, WL06a, WL07b, WW08b, WS09a, XCM07, XY07, Yan05a, Yan05c, YXW07, ZSHB06, ZXF07, ZS06]. viability [BSP05]. vibrating [TD09]. Vibration [ZC06a, CZ05a, CFC06c, CH07b, JSL05, Shu05, WC09c, YAO09]. vibrations [ÇFÇ06b, Cel06a, CZ05b, CZZ05, YJW06]. vibro [FX09]. vibro-impact [FX09]. Victoria [MD07b]. video [Par08, YLL07]. videos [YZDW07]. Viéta [Wu05e]. view [FTMA07, PRA^{+}09, lSAQ08]. viewpoint [Ish09]. Vinokur [Bri08]. viral [GN06a, HCW06]. virotherapy [WT08]. virtual [YY07c, ZLZ06]. virus [YC08c, ZSZS09, d’O05b]. virus-epidemic [YC08c]. viruses [LT07, MS07d, PA09]. visco [FSÇ06, HS09c]. visco-elastic [HS09c]. visco-plastic [FSÇ06]. viscoelastic [CZ05a, CC05c, Dem05a, EB05c, EHL08, FeT05, Fus08, GM08, JL08, KM09a, MeT05, MS09b, SK07b, SSTD05a, SSTD05b, SZXL06, VFF08b, YZ06, ZY09, dEdSNL05]. viscoelasticity [eT09]. Viscosity [LjZH08, SC07b, Wan08g, ZSM07, ZS07a, AB05c, AH05a, BOB05, CY08a, Che09b, EESF04, JK07, Mak06, PM09, Pan08c, PR06, SAS05, SSK06]. viscosity-dependent [BOB05]. Viscous [Cor07, EO06, EESESD05, EA08b, Jer08, KAS06a, KS09a, NCGdSM07, PPS09, YM07a]. visibility [HP07]. vision [KH07a]. visual [SJ08b, SY07b]. Visualization [HH07b]. Visualizing [XX07]. vital [AAA08]. VMOA [AS09c]. VMSS [SC05a]. voice [Alr05, SDM^{+}09]. voice-multihop [Alr05]. voids [Sin07a]. Voigt [BS09c, KAP06]. volatilities [CKK05]. Volatility [Nwo06g, Fuk06b, Li05d, Oca07, You09, Nwo06d]. volcanoes [MMS^{+}06, MMS^{+}08]. Volkenborn [Kim07a]. Vols [Ano07b, Ano08]. Volterra [AD06b, AŞ06e, BD05a, Bad05, BR09a, BS05b, Bel06, Bel07a, Bel08a, Bel08b, BS09b, Che05a, Che05d, Che06e, Che06f, CF06b, CJP08, CD06, Dar06, ERAS07, FK06a, GLV08, GKB07c, GJB05, Jia07b, LC06a, Li07d, LL08d, LZC07b, LC09b, Lis06, LZ09c, LP06a, Ma07a, MA05d, MY06a, MSR07, MMA07, MS08a, MXC07, MO07, Ord06b, PW08a, PMRAS05b, QD08, QC08, RMA07a, RJW07, Ram07g, Ram09a, SNH07a, SNH07b, SG07a, Sha05a, SP08, SC08c, SM07d, TF08, TL08, Wan06g, Wan06i, Wan06h, WS07b, WZ08b, Wan08f, WYL06b, WYL06c, WWYL08, XCH05a, XCH05b, YS06a, Yu09, ZM08a, ZXL05, ZZ07b, ZY07]. Volume [Ano05d, Ano05e, Ano05f, Ano05a, Ano05g, Ano05b, Ano05c, Ano07a, AD06c, Bi06, Bra05, CCE06, CR06, dCDNPN09, DC07, GY06, HZX07, PD09, Ram07c, Sun08, Tou09, Yan08e, ZW08a]. volume-preserving [Sun08]. Volumes [Ano05h, Ano06a, Ano06b, Ano05i]. volumetric [AZ07]. Voronoi [FS09]. vortex [FFKV08, Zhe08]. vortices [LC05b]. vorticity [EGFAE06, SFH09, TY07, WZZ06b, Zhe08]. voting [CH07f, MF09]. Vries [CSV07, AÖ06b, Bah05, DW09a, DKK06, HX07, IDS06, Ism08b, KM05d, ÖK05, ÖÖ06, ROK09, RKOK09, She08, SZ07d, SZ07e, Sya05a]. vs [AF07, dCDB07].

Waerden [Sht07]. waiting

[GPTU05, JW09a]. wall [Gür05, Zhe08].

wall-bounded [Zhe08]. walls [Mak09, SK09a, VFF08b]. Wang [Has06b, SZS05, Wu06a, ZWT09a, ZTWL09]. war [Mer07]. warm [WC06a, WDK06]. warranty [JM06a]. water [ADH05, ADH06, CPSS08a, CPSS08b, FSMC05, RAP05, Sha07a, Tou09, WZL08b, Waz08e, Waz08p, Waz08u, Waz09d, XCFX05, Yil06, ZS09b]. water-wave [ADH05]. watermark [LYGL07b, LLC06a]. watermarking [Che08c, CYH^{+}07, FWL08, HHC05, LLC06a, LLC06b, YZDW07]. Wave [Sin07a, SK07b, ADH05, Ahm07b, AKBZ08, AKM05, AKMA05, AKARA08, Bou09a, ÇFÇ06a, CW05a, CW05b, CW06c, CZ08b, CLDH09, CB06, DGB08, DW09a, DG07a, Dem05c, Den08b, Den08c, Den08d, DSZ08, Dog05, ESK05, Ese05b, EK06, FLLH07b, FHL07, FLLH07a, FHG07, GHL07, GL08b, Gor06a, HFL06, lH08a, HRLC08, HRCL08, Inc06b, Inc06c, Inc07a, IRSR05, JDG06d, KRA07, KD09, Laz08, LSL06, LSW06b, LSW06a, LZ08b, LXML09, Liu06b, Lou07, LHF07, MYG08, MWQ09, MLW09, MK07a, MAY09, Mom05b, NHKZ06, OM06, Odi06b, OB09, PW09, Par09a, Pat05a, Pen08, RKOK09, RN09, Ras05b, Rom09b, Sha06b, SQHY08, SP08, SXJ05, SLX05a, SXX05, SX06, SMG07, SC08a, She08, SZXL06, ST08, TL06, TH07a, TH08, TZH09a, TXW09, TZH09b, TB09b, TNL09, Vis08, WSLZ07, WCLZ05, cWmZ06, WCZ06, WZL08b, WW08b, Waz05d, Waz05q, Waz06c, Waz06f, Waz06e]. wave [Waz06j, Waz06n, Waz06o, Waz06p, Waz07a, Waz07f, Waz08u, Waz09d, WP06, WF07b, XHLL05, XGTS07, XMGW09, Yan08a, Yan06b, YX08, YB07, ZW09a, Zha05b, ZX06c, ZW07a, ZZB08, ZDY08, Zha09c, Zha05e, Zha06c, ZXLS06, ZZYZ06a, ZZTJ06, ZX07b, ZWX08, ZL08c, ZLZZ06, ZL07b, ZY08d, ZLZ08, ZLYL08, WZC07, ZC06a]. wave-like [Mom05b, Yan08a]. wave-type [ST08]. waveform [WB06, WHL08, ZH08c]. waveforms [PU08]. wavefronts [LL08d, LTZ09, SLC09]. wavefunction [Fag05]. Wavelet [BLP07, DY06, FXLZ05, GMCM06, Kar09b, KM05e, KM05f, MAD08, Mah05, NKR^{+}08, SM07a, BF07a, BF07b, BSKS07, DX07a, EG07a, FHG07, GJ07a, GZL08, GAV08, GC06b, HKS09b, HKS09a, HBE09, HH09a, KM05d, Lep06, Lep07, Lep09, LY09a, MUZG09, MPEL09, MM05b, ML06, MY06b, MLR07, MMA07, MK07a, Mit06, PBKM09, PVN06b, Qia08, SBA09, SGG^{+}09, Su07d, cWxChY07, Wu09, XWZ06, Yan06a, YZ07a, ZLL06a, EG06a, GkB07b]. wavelet-based [GAV08, MUZG09, PVN06b]. Wavelet-Galerkin [NKR^{+}08, EG07a, HBE09, EG06a, GkB07b]. wavelet-multigrid [BSKS07]. wavelet-Taylor [MK07a]. Wavelets [GPS09, AVKA09, AM08c, HBES09, HW07b, KGB06, Lep08, LC07, LXSH06, LFW08, Ma05b, MD06a, MYN07, MLM07, MS07b, MNM07, SH07d, SY07c, Wan08e, yXhWZ06, You06a, You06b, YB06, ZH05]. wavelets-Galerkin [MLM07]. waves [Ate06, BL06, Boy06b, BKP06, CHH09, DV06, FT07, GZ05, HZ08, HZ09b, KS08c, KM05a, Lin09a, LHF07, MGG^{+}09, SA06e, Sha07a, SKS06, ST06c, Sin08a, SC06c, Tan05b, TS05, TA05b, Vin09, VJD09, Waz08e, Waz08p, YL09a, ZM06a, ZYH08]. wavy [Che06n, Che07i, KS05b]. way [CL08c, CPTZ05, Vaj07d, YW05, YRY05c, ZW05a]. Weak [JK06, MPEL09, Tha07, ZY08c, AK06f, LZ07b, Lin05f, LHW07, SY07a, SR06, Tan06c, XL09c, ZM08d, ZMY08, lS08b]. weakly [JOB08, Muk07b, SMY08, SRMRGRL08, TL08]. Weakness [SW05b, SZS05, YY05d]. wealth [Dam08]. weapon [ÇE06a]. wearout [EH06]. web [KH05a, KH05b, MCAF08, VKVHN05, ZCZX07, WYD07]. Weber [ZT07b]. Website [SAK08]. wedge [AH05a, PM09]. Wei [CI05, HWS07, Zha09b]. Weibull [AJAH06, CL06b]. Weierstrass’ [PHP09, CY06a, PP07b, Hub08a]. weigh [CS07d]. weight [ANM08, AMJD06a, ADK07, BMJED06, DMJE06b, Has06a, Lin06c, LJE07a, MH09, SM05a, WYX05]. Weighted [ASA07a, DMJE06b, LS08d, LS09h, hLqSlZ09, NG06b, NRA07, RAN08a, Ste09g, Ste09h, Ste09i, SA06h, YL08b, Yu08a, CGVC07, Çel05a, CX08a, Dar06, dCDNPN09, DWT05b, Ere09a, GSDB07b, HZ06c, Isk07, Kan09b, KTW08, KWZ08, Lee05b, LWW05, LLY06, LYW09, XSX08, MJ07c, MA05g, MD08b, Occ09, PPV08, QW05, RAN07, SW09c, SiU09, Ste09d, TSP07, WG05, WSJW06, fWZpXzL09, WLC06, XLKH09, ZTWL09]. Weighted-Hardy [LS09h]. weighted-type [SW09c, SiU09, Ste09d, Ste09h, Ste09i]. weighting [TA05a]. weights [Fat06, Isk07, JMHS05, JMLR05, WC06d, Wan06l]. welfare [Sla06]. Well [GC07c, AA06e, Lie06]. Well-posedness [GC07c, AA06e]. Wendel [QG08]. WENO [Tan05a, WFS08, Zah09]. WGM [MLM07]. wheel [CFJ^{+}05]. which [CCE06, CMGMD06, Dyk09]. Whitham [CLDH09, ESK05, JZ06c, SXJ05, ZS09b]. Whittaker [Pog07b]. Whittaker-type [Pog07b]. widely [HFL06]. width [Dog05, Ese05b, Ras05a, Sak06]. Wiener [NCÁHCLP05b, Nak07, Nak08, NH07c, QN08]. Wigert [AMJD06b]. wildfire [AVSB08]. Wilkinson [WF07d]. Williams [AO05a, OZ07]. willingness [Nwo06e]. willingness-to-accept-losses [Nwo06e]. Wilson [HBBA07, Inc06b, WLS09]. Wilson- [HBBA07]. wind [dCDNPN09, KJ08, WS05b]. window [DX07b]. Windows [YSdS^{+}06]. wine [Mar07]. Wing [KF07]. Winograd [CS07d]. Wire [SH08a]. wireless [AS06c, YTWY05, ZWM08]. wise [CY06b]. withdrawal [MM07c, SH08a]. within [GD06, PN06, SB08a]. without [AR09, CCH05, CB06, FH07, HLT09, Jia05, bJyKxC09, LKY05b, LCZ05c, NKH07, SS05b, jW06b, Wu05e, YW05, YRY05c, ZW05a, ZL08c, Zho09a]. Wolf [DFG07, DFGV07]. Wolfe [HWS07, YGW09]. wood [CDT07]. Woodbury [BK09]. worded [Bog05b]. words [GJL06, Rom09a]. working [hLsT07]. workspace [BNdFPS06]. workstations [MAT08]. world [Dam07]. worlds [Pin07]. worsted [LY09c, YY07c]. WPAN [CL07a]. WPG [MK05d]. Wu [CHY05a, HL05c, SHSZ07, YY05c]. WYL [SWH07]. WZ [LZC05].

X [LLLW06]. X-ray [LLLW06]. XB [hPyHZ06, PHZ06b]. Xiao [YY05b]. XII [WY05c]. XOR [MSM07a].

Y. [CI05]. Yakubovich [WFHD09]. Yakubovich-conjugate [WFHD09]. Yang [KJY05, YWC05, YY05d, YL09c].

Yang-Shieh [YWC05]. Yao [Zha09b]. yarn [YY07c]. YCH [SC05a]. years [Gal06b]. yield [AMN07c]. yielding [eT09]. yields [AN07]. Yildirim [Sub07]. Yildiz [Sub07]. Yoo [KHKL05]. YTSF [DLL09, SZ07e, Waz08j].

Z [Ren08a, Aki05b]. Zakharov [Lay09a, zLmLL08, Pen08, SHY08, Waz05k, lYqL06, Zha09c, ZZTJ06]. Zernike [PBKM09, PBPF08]. Zero [GSG09, AE07c, CL07a, GKR06, Geu07, GSDB06, GSNDB09, GSDB09, LR07a, LR07b, PP07b, Rim05a, Rim05d, Rim06f, SA06a, SS06e, WR06, Yu08b]. zero-correlation-duration [CL07a]. zero-coupon [SS06e]. Zero-finder [GSG09, GSNDB09]. zero-finding [GSDB06, GSDB09, PP07b]. zero-one [AE07c]. zeros [Boy07b, MR07c, NG09, PG06, PM05, PR05, PMP08, PHP09, PR07, Ree07, Rim08b, Rim09b, Sor06, WZY07, WX06, WCW07b, ZPH06, ZL08e, Zho09c, Zhu05d]. zeta [Boy07a, CAS08, CAS09, Eli07, Sim07, CS05c, GGOS08]. Zhang [YY05b]. Zhou [PKH05]. ZK [Inc07d, Waz05f]. ZK-MEW [Inc07d]. Zolotarev [LZ08a]. zooplankton [CRC08, GN06a, SMBB07]. Zygmund [LS08d, LS09e, Zhu09b]. Zygmund-type [LS09e, Zhu09b].

References

Apaolaza:2005:TRC

[AA05] N. M. Apaolaza and J. R. Artalejo. On the time to reach a certain orbit level in multi-server retrial queues. Applied Mathematics and Computation, 168(1):686–703, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006320.

Abbasbandy:2006:NAFa

[AA06a] S. Abbasbandy and M. Amirfakhrian. The nearest approximation of a fuzzy quantity in parametric form. Applied Mathematics and Computation, 172(1):624–632, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002274.

Abbasbandy:2006:NAFb

[AA06b] S. Abbasbandy and M. Amirfakhrian. Numerical approximation of fuzzy functions by fuzzy polynomials. Applied Mathematics and Computation, 174(2):1001–1006, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005278.

AlSharawi:2006:PLE

[AA06c] Ziyad AlSharawi and James Angelos. On the periodic logistic equation. Applied Mathematics and Computation, 180(1):342–352, September 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030501129X.

Ashour:2006:DEI

[AA06d] S. A. Ashour and H. M. Ahmed. Derivation of the errors involved in Hermite interpolation and their applications to quadrature formulae. Applied Mathematics and Computation, 181(2):1482–1489, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306002773.

Ashyralyev:2006:NWP

[AA06e] A. Ashyralyev and N. Altay. A note on the wellposedness of the nonlocal boundary value problem for elliptic difference equations. Applied Mathematics and Computation, 175(1):49–60, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005904.

Ahmadi:2007:CTT

[AA07a] M. R. Ahmadi and H. Adibi. The Chebyshev Tau technique for the solution of Laplace’s equation. Applied Mathematics and Computation, 184(2):895–900, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007788.

Alirezaee:2007:CRD

[AA07b] M. R. Alirezaee and M. Afsharian. A complete ranking of DMUs using restrictions in DEA models. Applied Mathematics and Computation, 189(2):1550–1559, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017413.

Alirezaee:2007:MIC

[AA07c] M. R. Alirezaee and M. Afsharian. Model improvement for computational difficulties of DEA technique in the presence of special DMUs. Applied Mathematics and Computation, 186(2):1600–1611, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010964.

Amador:2007:DSB

[AA07d] J. Amador and J. R.

Artalejo. On the distribution of the successful and blocked events in the M / M / c retrial queue: a computational approach. Applied Mathematics and Computation, 190(2):1612–1626, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002068.

Ashour:2007:GQR

[AA07e] S. A. Ashour and H. M. Ahmed. A Gauss quadrature rule for hypersingular integrals. Applied Mathematics and Computation, 186(2):1671–1682, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011027.

Ashyralyev:2007:NTD

[AA07f] Allaberen Ashyralyev and Doghonay Arjmand. A note on the Taylor’s decomposition on four points for a third-order differential equation. Applied Mathematics and Computation, 188(2):1483–1490, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015013.

Alzabut:2008:EGA

[AA08] Jehad O. Alzabut and Thabet Abdeljawad. On existence of a globally attractive periodic solution of impulsive delay logarithmic population model. Applied Mathematics and Computation, 198(1):463–469, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700882X.

Abdou:2009:GFI

[AA09] M. A. Abdou and S. A.

Aseeri. Goursat functions for an infinite plate with a generalized curvilinear hole in ζ-plane. Applied Mathematics and Computation, 212(1):23– 36, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001143.

Asady:2005:FGL

[AAA05] B. Asady, S. Abbasbandy, and M. Alavi. Fuzzy general linear systems. Applied Mathematics and Computation, 169(1):34–40, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007453.

Anita:2008:GBA

[AAA08] Laura-Iulia Aniţa, Sebastian Aniţa, and Viorel Arnăutu. Global behavior for an age-dependent population model with logistic term and periodic vital rates. Applied Mathematics and Computation, 206(1):368–379, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006851.

Anita:2009:OHP

[AAA09] Laura-Iulia Aniţa, Sebastian Aniţa, and Viorel Arnăutu. Optimal harvesting for periodic agedependent population dynamics with logistic term. Applied Mathematics and Computation, 215(7):2701– 2715, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030900808X.

Allahviranloo:2006:BJT

[AAAA06] T. Allahviranloo, E. Ahmady, N. Ahmady, and Kh. Shams Alketaby. Block Jacobi two-stage method with Gauss–Sidel inner iterations for fuzzy system of linear equations. Applied Mathematics and Computation, 175(2):1217–1228, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006715.

Abdel-All:2004:RST

[AAABH04] Nassar H. Abdel-All, Rashad A.

Abdel-Baky, and Fathi M. Hamdoon. Ruled surfaces with timelike rulings. Applied Mathematics and Computation, 147(1):241– 253, January 5, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See errata [KAK05, KAK06, Kas06b].

Akkose:2007:SSR

[AABD07] Mehmet Akköse, Süleyman Adanur, Alemdar Bayraktar, and A. Aydın Dumanoğlu. Stochastic seismic response of Keban dam by the finite element method. Applied Mathematics and Computation, 184(2):704–714, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007466.

Abdel-Aziz:2007:EAO

[AAES07] Yehia A. Abdel-Aziz and F. A. Abd El-Salam. An efficient algorithm for orbital evolution of artificial satellite. Applied Mathematics and Computation, 191(2):415–428, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700269X.

Ankaralioglu:2005:CAG

[AAG05]  Nurullah  Ankaralıoğlu,

Hüseyin Aydın, and Inci Gültekin. Computation of abstract groups S((d + 1)/2, d) and torus knots K(d, 2). Applied Mathematics and Computation, 167(1):695–710, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005247.

Amiraliyev:2007:NTS

[AAK07] G. M. Amiraliyev, I. G. Amiraliyeva, and Mustafa Kudu. A numerical treatment for singularly perturbed differential equations with integral boundary condition. Applied Mathematics and Computation, 185(1):574–582, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008927.

Agouzal:2005:NCF

[AAN05] A. Agouzal, K. Allali, and K. Najib. Non-conforming finite element discretisations of a class of non-Newtonian flows. Applied Mathematics and Computation, 168(2):1004–1019, September 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304006587.

Abu-Alshaikh:2006:TPI

[AAS06] Ibrahim Abu-Alshaikh and Ali Sahin. Two-point iterative methods for solving nonlinear equations. Applied Mathematics and Computation, 182(1):871– 878, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306004048.

Al-Anzi:2006:UMG

[AASAA06] Fawaz S. Al-Anzi, Yuri N.

Sotskov, Ali Allahverdi, and George V. Andreev. Using mixed graph coloring to minimize total completion time in job shop scheduling. Applied Mathematics and Computation, 182(2):1137–1148, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004280.

Abdel-Baky:2005:BTM

[AB05a] Rashad A. Abdel-Baky. The Backlund’s theorem in Minkowski 3space R_{1}^{3}. Applied Mathematics and Computation, 160(1):41–50, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303007161.

Ahmad:2005:IAR

[AB05b] Rais Ahmad and Fabián Flores Bazán. An iterative algorithm for random generalized nonlinear mixed variational inclusions for random fuzzy mappings. Applied Mathematics and Computation, 167(2):1400–1411, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304005752.

Akyildiz:2005:CTC

[AB05c] F. Talay Akyildiz and Hamid Bellout. Chaos in the thermal convection of a Newtonian fluid with a temperature dependent viscosity. Applied Mathematics and Computation, 162(3):1103–1118, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001523.

Amat:2006:STM

[AB06] Sergio Amat and Sonia Busquier. On a Steffensen’s type method and its behavior for semismooth equations. Applied Mathematics and Computation, 177(2):819–823, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009628.

Abbasbandy:2006:NSFa

[ABA06] S. Abbasbandy, E. Babolian, and M. Allame. Numerical solution of fuzzy max–min systems. Applied Mathematics and Computation, 174(2):1321–1328, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005527.

Abbasbandy:2005:ENM

[Abb05]  S. Abbasbandy.  Ex-

tended Newton’s method for a system of nonlinear equations by modified Adomian decomposition method. Applied Mathematics and Computation, 170(1):648–656, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000068.

Abbasbandy:2006:HPM

[Abb06a] S. Abbasbandy. Homotopy perturbation method for quadratic Riccati differential equation and comparison with Adomian’s decomposition method. Applied Mathematics and Computation, 172(1):485– 490, January 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305002171.

Abbasbandy:2006:IHH

[Abb06b]  S. Abbasbandy.  It-

erated He’s homotopy perturbation method for quadratic Riccati differential equation. Applied Mathematics and Computation, 175(1):581–589, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006314.

Abbasbandy:2006:MHP

[Abb06c] S. Abbasbandy. Modified homotopy perturbation method for nonlinear equations and comparison with Adomian decomposition method. Applied Mathematics and Computation, 172(1):431–438, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002134.

Abbasbandy:2006:NSI

[Abb06d] S. Abbasbandy. Numerical solutions of the integral equations: Homotopy perturbation method and Adomian’s decomposition method. Applied Mathematics and Computation, 173(1):493–500, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003693.

Abell:2006:MAC

[ABB06e] Martha Abell, James Braselton, and Lorraine Braselton. A model of allelopathy in the context of bacteriocin production. Applied Mathematics and Computation, 183(2):916–931, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006254.

Amat:2008:DEI

[ABBP08] S. Amat, C. Bermúdez, S. Busquier, and S. Plaza. On the dynamics of the Euler iterative function. Applied Mathematics and Computation, 197(2):725– 732, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008314.

Abdollahi:2006:PNH

[Abd06] A. Abdollahi. The polynomial numerical hull of a matrix and algorithms for computing the numerical range. Applied Mathematics and Computation, 180(2):635–640, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600052X.

Abdou:2007:ETM

[Abd07] M. A. Abdou. The extended tanh method and its applications for solving nonlinear physical models. Applied Mathematics and Computation, 190(1):988– 996, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000756.

Abdulhadi:2009:UFL

[Abd09]  Zayid Abdulhadi.  On

the univalence of functions with logharmonic Laplacian. Applied Mathematics and Computation, 215(5):1900–1907, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006924.

Admane:2006:AAU

[ABK^{+}06] Lotfi Admane, Karima Benatchba, Mouloud Koudil, Lamri Siad, and Said Maziz. AntPart: an algorithm for the unsupervised classification problem using ants. Applied Mathematics and Computation, 180(1): 16–28, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305010933.

Altay:2009:MTS

[ABM09] Bilâl Altay, Feyzi Başar, and Eberhard Malkowsky. Matrix transformations on some sequence spaces related to strong Cesàro summability and boundedness. Applied Mathematics and Computation, 211(2):255–264, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000812.

Amat:2005:CAP

[ABP05] Sergio Amat, Sonia Busquier, and Sergio Plaza. Constructing attracting periodic orbits of three third-order iterative methods. Applied Mathematics and Computation, 161(2):679–690, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000207.

Adda:2005:FDE

[AC05] Fayçal Ben Adda and Jacky Cresson. Fractional differential equations and the Schrödinger equation. Applied Mathematics and Computation, 161(1):323–345, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012748.

Artalejo:2006:CAM

[AC06] Jesus R. Artalejo and Srinivas R. Chakravarthy. Computational analysis of the maximal queue length in the MAP / M / c retrial queue. Applied Mathematics and Computation, 183(2):1399–1409, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006886.

Ali:2007:GAO

[AC07a] Norhashidah Hj. Mohd.

Ali and Lee Siaw Chong. Group accelerated Over-Relaxation methods on rotated grid. Applied Mathematics and Computation, 191(2):533–542, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002810.

Alves:2007:EPS

[AC07b] Claudianor O. Alves and Francisco J. S. A. Corrêa. On the existence of positive solution for a class of singular systems involving quasilinear operators. Applied Mathematics and Computation, 185(1):727–736, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009088.

An:2008:NCS

[AC08] Hong-Li An and Yong Chen. Numerical complexiton solutions for the complex KdV equation by the homotopy perturbation method. Applied Mathematics and Computation, 203(1):125–133, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002142.

Anh:2009:PIR

[AC09] Pham Ky Anh and Cao Van Chung. Parallel iterative regularization methods for solving systems of illposed equations. Applied Mathematics and Computation, 212(2):542–550, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001866.

Arsham:2005:CIP

[ACD^{+}05] H. Arsham, G. Cimperman, N. Damij, T. Damij, and J. Grad. A computer implementation of the Push-and-Pull algorithm and its computational comparison with LP simplex method. Applied Mathematics and Computation, 170(1):36–63, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008690.

Alonso:2009:BNE

[ACDR09] P. Alonso, R. Cortina, I. Díaz, and J. Ranilla. Blocking Neville elimination algorithm for exploiting cache memories. Applied Mathematics and Computation, 209(1):2–9, March 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004682.

Aloy:2009:DEM

[ACJ09] R. Aloy, M.-C. Casabán, and L. Jódar. A discrete eigenfunctions method for computing mixed hyperbolic problems based on an implicit difference scheme. Applied Mathematics and Computation, 215(1):333– 343, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309004743.

Adibi:2006:CFP

[ACOS06] H. Adibi, Y. J. Cho, D. O‘Regan, and R. Saadati. Common fixed point theorems in L-fuzzy metric spaces. Applied Mathematics and Computation, 182(1):820–828, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004000.

Ahn:2007:RST

[ACP07] Hyo-Sung Ahn, YangQuan Chen, and Igor Podlubny. Robust stability test of a class of linear time-invariant interval fractional-order system using Lyapunov inequality. Applied Mathematics and Computation, 187(1):27– 34, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011337.

Alvarez:2005:EBS

[ÁCTZ05] Rafael Álvarez, Joan-Josep Climent, Leandro Tortosa, and Antonio Zamora. An efficient binary sequence generator with cryptographic applications. Applied Mathematics and Computation, 167(1):16–27, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004709.

Alhama:2005:NSH

[ACZ05] Francisco Alhama, Antonio Campo, and Joaquín Zueco. Numerical solution of the heat conduction equation with the electrothermal analogy and the code PSPICE. Applied Mathematics and Computation, 162(1):103–113, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000748.

Abbasbandy:2005:NSBa

[AD05a] S. Abbasbandy and M. T.

Darvishi. A numerical solution of Burgers’ equation by modified Adomian method. Applied Mathematics and Computation, 163(3):1265–1272, April 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003017.

Abbasbandy:2005:NSBb

[AD05b] S. Abbasbandy and M. T.

Darvishi. A numerical solution of Burgers’ equation by time discretization of Adomian’s decomposition method. Applied Mathematics and Computation, 170(1):95–102, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008732.

Agarwal:2006:NOF

[AD06a] Ravi P. Agarwal and Alexander Domoshnitsky. Non-oscillation of the firstorder differential equations with unbounded memory for stabilization by control signal. Applied Mathematics and Computation, 173(1):177–195, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500319X.

Akyuz-Dascioglu:2006:CPA

[AD06b] Ayşegül Akyüz-Daşcıoğlu. A Chebyshev polynomial approach for linear Fredholm–Volterra integrodifferential equations in the most general form. Applied Mathematics and Computation, 181(1):103–112, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000816.

Aoufi:2006:IFV

[AD06c] A. Aoufi and G. Damamme.

An implicit finite volume discretization for the multidimensional numerical simulation of TiC combustion synthesis. Applied Mathematics and Computation, 176(1):99–116, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007757.

Alipanah:2007:NSN

[AD07] Amjad Alipanah and Mehdi Dehghan. Numerical solution of the nonlinear Fredholm integral equations by positive definite functions. Applied Mathematics and Computation, 190(2):1754–1761, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002238.

Alrefaei:2009:SAT

[AD09a] Mahmoud H. Alrefaei and Ali H. Diabat. A simulated annealing technique for multi-objective simulation optimization. Applied Mathematics and Computation, 215(8):3029–3035, December 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309008807.

Amer:2009:MKM

[AD09b] S. M. Amer and S. Dardery. The method of Kantorovich majorants to nonlinear singular integral equation with shift. Applied Mathematics and Computation, 215(8):2799– 2805, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030900887X.

Adamchik:2007:HFR

[Ada07] V. S. Adamchik. On the Hurwitz function for rational arguments. Applied Mathematics and Computation, 187(1):3–12, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011301.

Avellar:2007:SAF

[ADDdM07] J. Avellar, L. G. S. Duarte, S. E. S. Duarte, and L. A. C. P. da Mota. A semialgorithm to find elementary first order invariants of rational second order ordinary differential equations. Applied Mathematics and Computation, 184(1):2–11, January 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004796.

Adeyeye:2007:NCM

[Ade07]  John O. Adeyeye.  A

Newton–Cotes method for quantum stochastic differential equations. Applied Mathematics and Computation, 190(1):402–409, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000707.

Abou-Dina:2005:ADT

[ADH05] M. S. Abou-Dina and F. M. Hassan. Approximate determination of the transmission and reflection coefficients for waterwave flow over a topography. Applied Mathematics and Computation, 168(1):283–304, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006046.

Abou-Dina:2006:GPN

[ADH06] M. S. Abou-Dina and F. M. Hassan. Generation and propagation of nonlinear tsunamis in shallow water by a moving topography. Applied Mathematics and Computation, 177(2):785–806, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009604.

Arora:2005:STP

[ADK05] Shelly Arora, S. S. Dhaliwal, and V. K. Kukreja. Solution of two point boundary value problems using orthogonal collocation on finite elements. Applied Mathematics and Computation, 171(1):358–370, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001013.

Arora:2006:AOC

[ADK06a] Shelly Arora, S. S. Dhaliwal, and V. K. Kukreja. Application of orthogonal collocation on finite elements for solving non-linear boundary value problems. Applied Mathematics and Computation, 180(2):516–523, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000129.

Arora:2006:CET

[ADK06b] Shelly Arora, S. S. Dhaliwal, and V. K. Kukreja. A computationally efficient technique for solving two point boundary value problems in porous media. Applied Mathematics and Computation, 183(2):1170–1180, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006618.

Arora:2007:CET

[ADK07] Shelly Arora, S. S. Dhaliwal, and V. K. Kukreja. Computationally efficient technique for weight functions and effect of orthogonal polynomials on the average. Applied Mathematics and Computation, 186(1):623–631, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009738.

Abu-Dayyeh:2005:ALE

[ADM05] Walid Abu-Dayyeh and Kailash C. Madan. The admissibility of the LRT for the exponential family model. Applied Mathematics and Computation, 160(2):303–308, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303008427.

Abdeljaoued:2009:BMS

[ADTGV09] Jounaidi Abdeljaoued, Gema M. Diaz-Toca, and Laureano Gonzalez-Vega. Bezout matrices, Subresultant polynomials and parameters. Applied Mathematics and Computation, 214(2):588–594, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309003580.

Akyuz-Dascioglu:2006:AMS

[ADY06] A. Akyüz-Daşcıoğlu and H. Çerdik Yaslan. An approximation method for the solution of nonlinear integral equations. Applied Mathematics and Computation, 174(1):619–629, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305004893.

Agarwal:2005:GRG

[ADZ05] Ravi P. Agarwal, Shengfu Deng, and Weinian Zhang.

Generalization of a retarded Gronwall-like inequality and its applications. Applied Mathematics and Computation, 165(3):599–612, June 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004059.

Abbasbandy:2006:NMS

[AE06a] S. Abbasbandy and R. Ezzati. Newton’s method for solving a system of fuzzy nonlinear equations. Applied Mathematics and Computation, 175(2):1189–1199, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006697.

Ahmed:2006:COM

[AE06b] E. Ahmed and M. F.

Elettreby. On combinatorial optimization motivated by biology. Applied Mathematics and Computation, 172(1):40–48, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001840.

Adibi:2007:NSB

[AE07a] H. Adibi and Jaafar Es’haghi. Numerical solution for biharmonic equation using multilevel radial basis functions and domain decomposition methods. Applied Mathematics and Computation, 186(1):246–255, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009349.

Anjidani:2007:NNM

[AE07b] M. Anjidani and S. Effati. A new numerical method for solving the general form of the second order partial differential equations. Applied Mathematics and Computation, 188(2):1087–1093, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012276.

Anjidani:2007:SDM

[AE07c] M. Anjidani and S. Effati. Steepest descent method for solving zeroone nonlinear programming problems. Applied Mathematics and Computation, 193(1):197–202, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003864.

Abo-Eldahab:2005:FHT

[AEA05a] Emad M. Abo-Eldahab and Mohamed A. El Aziz. Flow and heat transfer in a micropolar fluid past a stretching surface embedded in a non-Darcian porous medium with uniform free stream. Applied Mathematics and Computation, 162(2):881–899, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001365.

See comments [Pan08b].

Ahmed:2005:MOP

[AEA05b] E. Ahmed and M. El-Alem. On multiobjective optimization in portfolio management. Applied Mathematics and Computation, 167(1):616–621, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005314.

Ashyralyev:2007:NSD

[AEA07] Allaberen Ashyralyev, Abdullah Said Erdogan, and Nurullah Arslan. On the numerical solution of the diffusion equation with variable space operator. Applied Mathematics and Computation, 189(1):682– 689, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016444.

Afshar:2008:CBA

[AEFFA08] K. Afshar, M. Ehsan, M. Fotuhi-Firuzabad, and N. Amjady. Cost-benefit analysis and MILP for optimal reserve capacity determination in power system. Applied Mathematics and Computation, 196(2): 752–761, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307007278.

Abo-Eldahab:2005:REH

[AEG05] Emad M. Abo-Eldahab and Ahmed F. Ghonaim. Radiation effect on heat transfer of a micropolar fluid through a porous medium. Applied Mathematics and Computation, 169(1):500–510, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007891.

Abbasbandy:2006:LDM

[AEJ06] S. Abbasbandy, R. Ezzati, and A. Jafarian. LU decomposition method for solving fuzzy system of linear equations. Applied Mathematics and Computation, 172(1):633–643, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002286.

Abo-Eldahab:2005:MFC

[AES05] Emad M. Abo-Eldahab and Ahmed M. Salem. MHD free-convection flow of a non-Newtonian powerlaw fluid at a stretching surface with a uniform free-stream. Applied Mathematics and Computation, 169(2):806–818, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008124.

Ahmed:2008:REF

[AES08] H. M. Ahmed and S. I. El-Soubhy. Recurrences and explicit formulae for the expansion and connection coefficients in series of ordinary Bessel polynomials. Applied Mathematics and Computation, 199(2):482– 493, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010144.

Abdou:2005:SNI

[AESD05] M. A. Abdou, W. G. El-Sayed, and E. I. Deebs. A solution of a nonlinear integral equation. Applied Mathematics and Computation, 160(1):1–14, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303006131.

Al-Fadhel:2007:GMV

[AF07] Tariq A. Al-Fadhel. Gauss map vs Bernoulli shift. Applied Mathematics and Computation, 194(2):520– 526, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307005048.

Awartani:2005:CCS

[AFH05] M. M. Awartani, R. A. Ford, and M. H. Hamdan. Computational complexities and streamfunction coordinates. Applied Mathematics and Computation, 169(2):758–777, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008094.

Afrouzi:2005:PMP

[Afr05] G. A. Afrouzi. On positive mountain pass solutions for a semilinear elliptic boundary value problem. Applied Mathematics and Computation, 167(1):76–80, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004746.

Attili:2006:EIR

[AFS06] Basem S. Attili, Khalid Furati, and Muhammed I. Syam. An efficient implicit Runge–Kutta method for second order systems. Applied Mathematics and Computation, 178(2):229– 238, July 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305009690.

Alvarez:2006:AQE

[ÁFVZ06] Rafael Álvarez, Francisco Ferrández, José-Francisco Vicent, and Antonio Zamora. Applying quick exponentiation for block upper triangular matrices. Applied Mathematics and Computation, 183(2):729–737, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006072.

Yusufoglu:2006:NSD

[(Ag06] Elçin Yusufoğlu (Agadjanov). Numerical solution of Duffing equation by the Laplace decomposition algorithm. Applied Mathematics and Computation, 177(2):572–580, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009379.

Yusufoglu:2007:EAS

[(Ag07] Elcin Yusufoğlu (Agadjanov). An efficient algorithm for solving integrodifferential equations system. Applied Mathematics and Computation, 192(1): 51–55, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307002937.

Aydogan:2008:MCR

[AG08] Emel Kizilkaya Aydogan and Cevriye Gencer. Mining classification rules with Reduced MEPAR-miner Algorithm. Applied Mathematics and Computation, 195(2):786–798, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006182.

Azadeh:2007:IANb

[AGA^{+}07] A. Azadeh, S. F. Ghaderi, M. Anvari, M. Saberi, and H. Izadbakhsh. An integrated artificial neural network and fuzzy clustering algorithm for performance assessment of decision making units. Applied Mathematics and Computation, 187(2):584–599, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011143.

Asghar:2005:RFT

[AGH05] S. Asghar, M. Mudassar Gulzar, and T. Hayat. Rotating flow of a third grade fluid by homotopy analysis method. Applied Mathematics and Computation, 165(1):213–221, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003753.

Azadeh:2008:IDA

[AGI08] A. Azadeh, S. F. Ghaderi, and H. Izadbakhsh. Integration of DEA and AHP with computer simulation for railway system improvement and optimization. Applied Mathematics and Computation, 195(2):775–785, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006170.

Arsham:2007:PMI

[AGJ07] Hossein Arsham, Janez Grad, and Gašper Jaklič. Perturbed matrix inversion with application to LP simplex method. Applied Mathematics and Computation, 188(1):801–807, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014184.

Azadeh:2007:IFC

[AGM^{+}07] A. Azadeh, S. F. Ghaderi, Y. Partovi Miran, V. Ebrahimipour, and K. Suzuki. An integrated framework for continuous assessment and improvement of manufacturing systems. Applied Mathematics and Computation, 186(2):1216–1233, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010472.

Arenas:2009:PFS

[AGPJV09] Abraham J. Arenas, Gilberto González-Parra, Lucas Jódar, and Rafael-J. Villanueva. Piecewise finite series solution of nonlinear initial value differential problem. Applied Mathematics and Computation, 212(1):209–215, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001350.

Algaba:2009:LBL

[AGR09] A. Algaba, C. García, and M. Reyes. Local bifurcation of limit cycles and integrability of a class of nilpotent systems of differential equations. Applied Mathematics and Computation, 215(1):314–323, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004718.

Azadeh:2007:FEC

[AGS07] A. Azadeh, S. F. Ghaderi, and S. Sohrabkhani. Forecasting electrical consumption by integration of Neural Network, time series and ANOVA. Applied Mathematics and Computation, 186(2):1753–1761, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011106.

Amiri:2007:MAT

[AGT07] Maghsoud Amiri and Farhad Ghassemi-Tari. A methodology for analyzing the transient availability and survivability of a system with repairable components. Applied Mathematics and Computation, 184(2):300–307, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007053.

Azadeh:2007:IANa

[AGTS07] A. Azadeh, S. F. Ghaderi, S. Tarverdian, and M. Saberi. Integration of artificial neural networks and genetic algorithm to predict electrical energy consumption. Applied Mathematics and Computation, 186(2):1731–1741, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011088.

Al-Harbi:2005:NSN

[AH05a] Saleh M. Al-Harbi. Numerical study of natural convection heat transfer with variable viscosity and thermal radiation from a cone and wedge in porous media. Applied Mathematics and Computation, 170(1):64–75, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008707.

Allahviranloo:2005:NMA

[AH05b] T. Allahviranloo and T. Hajari. Numerical methods for approximation of fuzzy data. Applied Mathematics and Computation, 169(1):16–33, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007441.

Awartani:2005:FDF

[AH05c] M. M. Awartani and M. H.

Hamdan. Fully developed flow through a porous channel bounded by flat plates. Applied Mathematics and Computation, 169(2):749–757, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008082.

Ahmad:2006:DNM

[AH06a] I. Ahmad and Z. Husain. Duality in nondifferentiable minimax fractional programming with generalized convexity. Applied Mathematics and Computation, 176(2):545–551, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008544.

Ahmed:2006:DAP

[AH06b] E. Ahmed and A. S. Hegazi. On different aspects of portfolio optimization. Applied Mathematics and Computation, 175 (1):590–596, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305006326.

Ahmed:2006:DMT

[AH06c] E. Ahmed and A. S. Hegazi. On dynamical multi-team and signaling games. Applied Mathematics and Computation, 172(1):524–530, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002201.

Allan:2006:FPM

[AH06d] F. M. Allan and M. H.

Hamdan. Fluid-particle model of flow through porous media: The case of uniform particle distribution and parallel velocity fields. Applied Mathematics and Computation, 183(2):1208–1213, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006655.

Ahmed:2007:DMC

[AH07a] E. Ahmed and A. S.

Hegazi. A dynamic model for e-commerce taxation. Applied Mathematics and Computation, 187(2):965–967, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601232X.

Alanelli:2007:ICN

[AH07b] M. Alanelli and A. Hadjidimos. On iterative criteria for H- and non-H-matrices. Applied Mathematics and Computation, 188(1):19–30, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306013154.

Ali:2007:PMC

[AH07c] Nasir Ali and Tasawar Hayat. Peristaltic motion of a Carreau fluid in an asymmetric channel. Applied Mathematics and Computation, 193(2):535–552, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004377.

Aminataei:2007:CSA

[AH07d] A. Aminataei and S. S. Hosseini. The comparison of the stability of Adomian decomposition method with numerical methods of equation solution. Applied Mathematics and Computation, 186(1):665–669, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009787.

Abdi:2008:IRV

[AH08a] Ali Abdi and Seyed Mohammad Hosseini. An investigation of resolution of 2-variate Gibbs phenomenon. Applied Mathematics and Computation, 203(2):714–732, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003962.

Argyros:2008:LCN

[AH08b] Ioannis K. Argyros and Saïd Hilout. Local convergence of Newtonlike methods for generalized equations. Applied Mathematics and Computation, 197(2):507–514, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007965.

Aryanezhad:2008:NGA

[AH08c] M. B. Aryanezhad and Mohammad Hemati. A new genetic algorithm for solving nonconvex nonlinear programming problems. Applied Mathematics and Computation, 199(1):186–194, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009861.

Anaya:2009:ITA

[AH09] J. C. Zúñiga Anaya and D. Henrion. An improved Toeplitz algorithm for polynomial matrix nullspace computation. Applied Mathematics and Computation, 207(1):256– 272, January 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300308008114.

Ahmed:2005:SVP

[AHE05] E. Ahmed, A. S. Hegazi, and A. S. Elgazzar. On some variants of prisoner’s dilemma dynamics. Applied Mathematics and Computation, 163(1):163– 168, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002176.

Arias:2007:FPB

[AHIP07] Enrique Arias, Vicente Hernández, Jacinto Javier Ibáñez, and Jesús Peinado. A fixed point-based BDF method for solving differential Riccati equations. Applied Mathematics and Computation, 188(2):1319–1333, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601486X.

Ahmad:2005:SMS

[Ahm05a] Bashir Ahmad. Sub Mach-1 sound due to an arbitrary time dependent source near an absorbing half plane. Applied Mathematics and Computation, 163(1):39– 50, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002061.

Ahmed:2005:CCP

[Ahm05b] M. A. Ahmed. A characterization of the convergence of Picard iteration to a fixed point for a continuous mapping and an application. Applied Mathematics and Computation, 169(2):1298–1304, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008525.

Ahmed:2006:NMO

[Ahm06] S. G. Ahmed. A numerical method for oxygen diffusion and absorption in a sike cell. Applied Mathematics and Computation, 173(1):668–682, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305003887.

Ahmad:2007:ADO

[Ahm07a] Bashir Ahmad. Acoustic diffraction from an oscillating half plane. Applied Mathematics and Computation, 188(2):2029–2033, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014718.

Ahmad:2007:DSA

[Ahm07b] Bashir Ahmad. Diffraction of a spherical acoustic wave due to the coupling of pressure release and absorbing half planes in a moving fluid. Applied Mathematics and Computation, 188(2):1897–1907, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030601558X.

Abelman:2009:RPS

[AHM09a] S. Abelman, T. Hayat, and E. Momoniat. On the Rayleigh problem for a Sisko fluid in a rotating frame. Applied Mathematics and Computation, 215(7):2515–2520, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007875.

Ahmad:2009:STT

[Ahm09b] Bashir Ahmad. Stability in terms of two measures for perturbed impulsive delay integro-differential equations. Applied Mathematics and Computation, 214(1):83–89, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002811.

Amat:2008:MCI

[AHR08] S. Amat, M. A. Hernández, and N. Romero. A modified Chebyshev’s iterative method with at least sixth order of convergence. Applied Mathematics and Computation, 206(1):164–174, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006449.

Audet:2006:NDL

[AHS06] Charles Audet, Jean Haddad, and Gilles Savard. A note on the definition of a linear bilevel programming solution. Applied Mathematics and Computation, 181(1):351–355, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001032.

Ahuja:2008:CBV

[Ahu08] Om P. Ahuja. Connections between various subclasses of planar harmonic mappings involving hypergeometric functions. Applied Mathematics and Computation, 198(1):305–316, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008909.

Abbasbandy:2006:SDMa

[AJ06a] S. Abbasbandy and A. Jafarian. Steepest descent method for solving fuzzy nonlinear equations. Applied Mathematics and Computation, 174(1):669– 675, March 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305004935.

Abbasbandy:2006:SDMb

[AJ06b] S. Abbasbandy and A. Jafarian. Steepest descent method for system of fuzzy linear equations. Applied Mathematics and Computation, 175(1):823–833, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006478.

Abbasi:2006:EPW

[AJAH06] Babak Abbasi, Abdol Hamid Eshragh Jahromi, Jamal Arkat, and Mehdi Hosseinkouchack. Estimating the parameters of Weibull distribution using simulated annealing algorithm. Applied Mathematics and Computation, 183(1):85–93, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005406.

Abbasbandy:2005:CGM

[AJE05] S. Abbasbandy, A. Jafarian, and R. Ezzati. Conjugate gradient method for fuzzy symmetric positive definite system of linear equations. Applied Mathematics and Computation, 171(2):1184–1191, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Amirteimoori:2005:RDM

[AJK05] Alireza Amirteimoori, Gholamreza Jahanshahloo, and Sohrab Kordrostami. Ranking of decision making units in data envelopment analysis: a distance-based approach. Applied Mathematics and Computation, 171(1):122–135, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000858.

Adjabi:2009:TOD

[AJKM09] Yasin Adjabi, Fahd Jrad, Arezki Kessi, and Uḡurhan Muḡan. Third order differential equations with fixed critical points. Applied Mathematics and Computation, 208(1):238–248, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008965.

Aghalary:2007:SAF

[AJMR07] R. Aghalary, S. B. Joshi, R. N. Mohapatra, and V. Ravichandran. Subordinations for analytic functions defined by the Dziok– Srivastava linear operator. Applied Mathematics and Computation, 187(1):13– 19, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011313.

Arana-Jimenez:2008:KIC

[AJOGRLRG08] M. Arana-Jiménez, R. Osuna-Gómez, A. Rufián-Lizana, and G. Ruiz-Garzón. KTinvex control problem. Applied Mathematics and Computation, 197(2):489– 496, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007989.

Arana-Jimenez:2009:ESV

[AJRLRGOG09] Manuel Arana-Jiménez, Antonio Rufián-Lizana, Gabriel Ruiz-Garzón, and Rafaela Osuna-Gómez. Efficient solutions in V-KT-pseudoinvex multiobjective control problems: a characterization. Applied Mathematics and Computation, 215(2):441–448, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005566.

Abalo:2005:BES

[AK05a] Kokou Abalo and Michael Kostreva. Berge equilibrium: some recent results from fixed-point theorems. Applied Mathematics and Computation, 169(1):624–638, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007994.

Al-Khaled:2005:NAP

[AK05b] Kamel Al-Khaled. Numerical approximations for population growth models. Applied Mathematics and Computation, 160(3):865–873, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303011998.

Al-Khaled:2005:NSL

[AK05c] Kamel Al-Khaled. Numerical solutions of the Laplace’s equation. Applied Mathematics and Computation, 170(2):1271–1283, November 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305000627.

Amirteimoori:2005:AFC

[AK05d] Alireza Amirteimoori and Sohrab Kordrostami. Allocating fixed costs and target setting: a deabased approach. Applied Mathematics and Computation, 171(1):136–151, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500086X.

Amirteimoori:2005:DLM

[AK05e] Alireza Amirteimoori and Sohrab Kordrostami. DEAlike models for multicomponent performance measurement. Applied Mathematics and Computation, 163(2):735–743, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002553.

Amirteimoori:2005:ESE

[AK05f] Alireza Amirteimoori and Sohrab Kordrostami. Efficient surfaces and an efficiency index in DEA: a constant returns to scale. Applied Mathematics and Computation, 163(2):683–691, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002528.

Amirteimoori:2005:MCE

[AK05g] Alireza Amirteimoori and Sohrab Kordrostami. Multicomponent efficiency measurement with imprecise data. Applied Mathematics and Computation, 162(3):1265–1277, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001869.

Ammar:2005:COS

[AK05h] E. Ammar and H. A. Khalifa. Characterization of optimal solutions of uncertainty investment problem. Applied Mathematics and Computation, 160(1):111–124, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303010828.

Aydin:2005:LBL

[AK05i] Orhan Aydın and Ahmet Kaya. Laminar boundary layer flow over a horizontal permeable flat plate. Applied Mathematics and Computation, 161(1):229–240, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012232.

See comment [Pan06].

Abalo:2006:ITT

[AK06a] Kokou Y. Abalo and Michael M. Kostreva. Intersection theorems and their applications to Berge equilibria. Applied Mathematics and Computation, 182(2):1840–1848, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005339.

Afrouzi:2006:NMFb

[AK06b] G. A. Afrouzi and S. Khademloo. A numerical method for finding positive solutions of a superlinear Dirichlet problem. Applied Mathematics and Computation, 181(2):1513–1518, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002815.

Afrouzi:2006:NMFa

[AK06c] G. A. Afrouzi and S. Khademloo. A numerical method to find positive solution of semilinear elliptic Dirichlet problems. Applied Mathematics and Computation, 174(2):1408–1415, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030500559X.

Afrouzi:2006:NRP

[AK06d] G. A. Afrouzi and S. Khademloo. Numerical results for positive solutions of a superlinear elliptic equation. Applied Mathematics and Computation, 180(2):599–604, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306000488.

Afrouzi:2006:SNR

[AK06e] G. A. Afrouzi and S. Khademloo. Some numerical results on a convex nonlinear elliptic problem. Applied Mathematics and Computation, 175(1):465–471, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006211.

Alirezaee:2006:REW

[AK06f] Mohammad R. Alirezaee and Masoud Khalili. Recognizing the efficiency, weak efficiency and inefficiency of DMUs with an epsilon independent linear program. Applied Mathematics and Computation, 183(2):1323–1327, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006795.

Allahviranloo:2006:SFS

[AK06g] T. Allahviranloo and M. Afshar Kermani. Solution of a fuzzy system of linear equation. Applied Mathematics and Computation, 175(1):519–531, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006260.

Amirteimoori:2006:PMF

[AK06h] Alireza Amirteimoori and Sohrab Kordrostami. Prioritization method for frontier DMUs: a slackbased measure. Applied Mathematics and Computation, 174(1):409–418, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004674.

Azaron:2006:MOL

[AK06i] Amir Azaron and Farhad Kianfar. A multi-objective lead time control problem in multistage assembly systems using an interactive method. Applied Mathematics and Computation, 176(2):609–620, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008593.

Ali:2008:IPS

[AK08a] M. M. Ali and P. Kaelo. Improved particle swarm algorithms for global optimization. Applied Mathematics and Computation, 196(2):578–593, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307007126.

Aseev:2008:SPI

[AK08b] S. M. Aseev and A. V. Kryazhimskiy. Shadow prices in infinite-horizon optimal control problems with dominating discounts. Applied Mathematics and Computation, 204(2):519–531, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002993.

Al-Khal:2009:MFH

[AK09]  R. A. Al-Khal.  Mul-

tiplier family of harmonic univalent functions. Applied Mathematics and Computation, 215(6):2238–2242, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900736X.

Amirteimoori:2007:DEAb

[AKA07] Alireza Amirteimoori, Sohrab Kordrostami, and Omme Habibe Azmayande. Data envelopment analysis with selective convexity and integervalued factors. Applied Mathematics and Computation, 188(1):734–738, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014111.

Al-Khaled:2008:TWS

[AKARA08] Kamel Al-Khaled, Mohammed Al-Refai, and Ameen Alawneh. Traveling wave solutions using the variational method and the tanh method for nonlinear coupled equations. Applied Mathematics and Computation, 202(1):233–242, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800088X.

Ali:2009:IVL

[AKB09a] M. M. Ali and Z. Kajee-Bagdadi. Improved versions of the LEDE algorithm for constrained global optimization. Applied Mathematics and Computation, 215(2):431– 440, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309005050.

Ali:2009:LEB

[AKB09b] M. M. Ali and Z. Kajee-Bagdadi. A local explorationbased differential evolution algorithm for constrained global optimization. Applied Mathematics and Computation, 208(1):31–48, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008862.

Ahmad:2008:ALI

[AKBZ08] Ferhana Ahmad, A. H. Kara, A. H. Bokhari, and F. D. Zaman. On approximate Lagrangians and invariants for scaling reductions of a nonlinear wave equation with damping. Applied Mathematics and Computation, 206(1):16–20, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006243.

Amiraliyev:2006:UDMa

[AKD06] G. M. Amiraliyev, Mustafa Kudu, and Hakki Duru. Uniform difference method for a parameterized singular perturbation problem. Applied Mathematics and Computation, 175(1):89– 100, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005941.

Akrami:2006:TMM

[AKH06] B. Akrami, B. Karimi, and S. M. Moattar Hosseini. Two metaheuristic methods for the common cycle economic lot sizing and scheduling in flexible flow shops with limited intermediate buffers: The finite horizon case. Applied Mathematics and Computation, 183(1):634–645, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005996.

Akin:2005:DEA

[Aki05a] Hasan Akin. On the directional entropy of Z^{2}- actions generated by additive cellular automata.

Applied Mathematics and Computation, 170(1):339– 346, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304009294.

Akin:2005:EPC

[Aki05b] Hasan Akin. On the ergodic properties of certain additive cellular automata over Z _{m}. Applied Mathematics and Computation, 168(1):192–197, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304005983.

Akin:2006:TET

[Akı06] Hasan Akın. The topological entropy of n th iteration of an additive cellular automata. Applied Mathematics and Computation, 174(2):1427–1437, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005618.

Al-Khasawneh:2007:EDI

[AKIS07] Raed Ali Al-Khasawneh, Fudziah Ismail, and Mohamed Suleiman. Embedded diagonally implicit Runge–Kutta–Nystrom 4(3) pair for solving special second-order IVPs. Applied Mathematics and Computation, 190(2):1803–1814, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002305.

Azaron:2006:REM

[AKKS06] Amir Azaron, Hideki Katagiri, Kosuke Kato, and Masatoshi Sakawa. Reliability evaluation of multicomponent cold-standby redundant systems. Applied Mathematics and Computation, 173(1):137– 149, February 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305003152.

Al-Khaled:2005:ASF

[AKM05] Kamel Al-Khaled and Shaher Momani. An approximate solution for a fractional diffusion-wave equation using the decomposition method. Applied Mathematics and Computation, 165(2):473–483, June 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003960.

Arora:2006:NSV

[AKM06] Urvashi Arora, Samir Karaa, and R. K. Mohanty. A new stable variable mesh method for 1-D non-linear parabolic partial differential equations. Applied Mathematics and Computation, 181(2):1423–1430, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600261X.

Akmaz:2009:TDE

[Akm09] Hakan K. Akmaz. Threedimensional elastic problems of three-dimensional quasicrystals. Applied Mathematics and Computation, 207(2):327–332, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800814X.

Al-Khaled:2005:AWS

[AKMA05] Kamel Al-Khaled, Shaher Momani, and Ahmed Alawneh. Approximate wave solutions for generalized Benjamin–Bona– Mahony–Burgers equations. Applied Mathematics and Computation, 171(1):281–292, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000949.

Allan:2006:IOP

[AKMH06] F. M. Allan, M. T. Kamel, T. A. Mughrabi, and M. H. Hamdan. Infiltration of oil into porous sediments. Applied Mathematics and Computation, 177(2):659–664, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009471.

Achouri:2006:CDS

[AKO06] T. Achouri, N. Khiari, and K. Omrani. On the convergence of difference schemes for the Benjamin–Bona– Mahony (BBM) equation. Applied Mathematics and Computation, 182(2):999–1005, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004140.

Aziz:2005:SMS

[AKR05] Tariq Aziz, Arshad Khan, and Jalil Rashidinia. Spline methods for the solution of fourth-order parabolic partial differential equations. Applied Mathematics and Computation, 167(1):153–166, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004813.

Amirteimoori:2006:ICE

[AKR06] Alireza Amirteimoori, Sohrab Kordrostami, and Aliakbar Rezaitabar. An improvement to the cost efficiency interval: a DEAbased approach. Applied Mathematics and Computation, 181(1):775–781, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001585.

Abbas:2008:SBA

[AKR08] M. Abbas, Safeer Hussain Khan, and B. E. Rhoades. Simpler is also better in approximating fixed points. Applied Mathematics and Computation, 205(1):428–431, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006103.

Aksan:2005:NSB

[Aks05] E. N. Aksan. A numerical solution of Burgers’ equation by finite element method constructed on the method of discretization in time. Applied Mathematics and Computation, 170(2):895–904, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000317.

Aksan:2006:QBS

[Aks06a] E. N. Aksan. Quadratic B-spline finite element method for numerical solution of the Burgers’ equation. Applied Mathematics and Computation, 174(2):884–896, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005187.

Amirteimoori:2006:MUF

[AKS06b] Alireza Amirteimoori, Sohrab Kordrostami, and Maryam Sarparast. Modeling undesirable factors in data envelopment analysis. Applied Mathematics and Computation, 180(2):444–452, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000075.

Asady:2005:SSK

[AKVH05] B. Asady, M. Tavassoli Kajani, A. Hadi Vencheh, and A. Heydari. Solving second kind integral equations with hybrid Fourier and blockpulse functions. Applied Mathematics and Computation, 160(2):517–522, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303011780.

Attili:2006:EMC

[AL06] B. S. Attili and D. Lesnic. An efficient method for computing eigenelements of Sturm–Liouville fourthorder boundary value problems. Applied Mathematics and Computation, 182(2):1247–1254, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004395.

Aguirre:2007:DPP

[AL07] M. A. Aguirre and C. K. Li. The distributional products of particular distributions. Applied Mathematics and Computation, 187(1):20–26, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011325.

Ali:2008:RKP

[AL08] Norhashidah Hj. Mohd Ali and Sam Teek Ling. Rotated Krylov preconditioned iterative schemes in the solution of convection– diffusion equations. Applied Mathematics and Computation, 206(1):425– 437, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308006930.

Abramova:2009:UFE

[AL09] Inna Abramova and Simon Latyshev. Using the fundamental equation of constrained motion with inequality constraints. Applied Mathematics and Computation, 215(8):2858–2876, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008273.

Albrecht:2006:SLS

[Alb06] Andreas A. Albrecht. Stochastic local search for the Feature. Applied Mathematics and Computation, 183(2):1148–1164, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006503.

Alexandre:2008:MME

[ALD08] R. Alexandre, Nguyen Thanh Long, and A. Pham Ngoc Dinh. A mathematical model for the evaporation of a liquid fuel droplet, subject to nonlinear constraints. Applied Mathematics and Computation, 199(1):139–154, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009812.

Alirezaee:2005:OAI

[Ali05] Mohammad Reza Alirezaee. The overall assurance interval for the non-Archimedean Epsilon in DEA models; a partition base algorithm. Applied Mathematics and Computation, 164(3):667–674, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003145.

Allahviranloo:2005:ADM

[All05a] Tofigh Allahviranloo. The Adomian decomposition method for fuzzy system of linear equations. Applied Mathematics and Computation, 163(2):553–563, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002425.

Allahviranloo:2005:NCM

[All05b] Tofigh Allahviranloo. Newton Cot’s [sic] methods for integration of fuzzy functions. Applied Mathematics and Computation, 166(2):339–348, July 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004412.

From the text, the author clearly means Cotes, not Cot’s, but the mistake is endemic. Allahviranloo:2005:RIF

[All05c] Tofigh Allahviranloo. Romberg integration for fuzzy functions. Applied Mathematics and Computation, 168(2):866–876, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006459.

Allahviranloo:2005:SRI

[All05d] Tofigh Allahviranloo. Successive over relaxation iterative method for fuzzy system of linear equations. Applied Mathematics and Computation, 162(1):189– 196, March 4, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304000797.

Allan:2007:DAD

[All07a] Fathi M. Allan. Derivation of the Adomian decomposition method using the homotopy analysis method. Applied Mathematics and Computation, 190(1):6–14, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000227.

An:2007:FEA

[ALL07b] Rong An, Yuan Li, and Kaitai Li. Finite element approximation for fourthorder nonlinear problem in the plane. Applied Mathematics and Computation, 194(1):143–155, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004535.

Ayasun:2006:SAC

[ALN06] Saffet Ayasun, Yiqiao Liang, and Chika O.

Nwankpa. A sensitivity approach for computation of the probability density function of critical clearing time and probability of stability in power system transient stability analysis. Applied Mathematics and Computation, 176(2):563–576, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008568.

Aloqeili:2006:DKT

[Alo06a]  Marwan Aloqeili.  Dy-

namics of a k th order rational difference equation. Applied Mathematics and Computation, 181(2):1328–1335, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002517.

Aloqeili:2006:DRD

[Alo06b] Marwan Aloqeili. Dynamics of a rational difference equation. Applied Mathematics and Computation, 176(2):768–774, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008775.

Aloqeili:2009:GSR

[Alo09] Marwan Aloqeili. Global stability of a rational symmetric difference equation. Applied Mathematics and Computation, 215(3):950–953, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309005815.

Alrefaei:2005:SAA

[Alr05] Mahmoud H. Alrefaei. A simulated annealing applied for optimizing a voice-multihop radio network. Applied Mathematics and Computation, 167(1):496–507, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304005065.

Aouat:2008:OIR

[ALSL08] Saliha Aouat, Nacéra Laiche, Feryel Souami, and Slimane Larabi. 3D object indexing and recognition. Applied Mathematics and Computation, 196(1):318–332, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006741.

Altintas:2006:CAC

[Alt06]  Ismet Altintas.  Com-

puter algebra and colored Jones polynomials. Applied Mathematics and Computation, 182(1):804–808, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003985.

Altintas:2007:OSM

[Alt07a] Ismet Altintas. An oriented state model for the Jones polynomial and its applications alternating links. Applied Mathematics and Computation, 194(1):168–178, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004559.

Altintas:2007:NCV

[Alt07b] Osman Altıntaş. Neighborhoods of certain pvalently analytic functions with negative coefficients. Applied Mathematics and Computation, 187(1):47– 53, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011350.

Atencia:2005:SSR

[AM05] I. Atencia and P. Moreno. A single-server retrial queue with general retrial times and Bernoulli schedule. Applied Mathematics and Computation, 162(2):855–880, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001353.

Aghajani:2006:NLC

[AM06] A. Aghajani and A. Moradifam. Nonexistence of limit cycles in two classes of predator–prey systems. Applied Mathematics and Computation, 175(1):356– 365, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006132.

Al-Mdallal:2008:NFE

[AM08a] Qasem M. Al-Mdallal. A new family of exact solutions to the unsteady Navier–Stokes equations using canonical transformation with complex coefficients. Applied Mathematics and Computation, 196(1):303–308, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006728.

Amat:2008:SLMa

[AM08b] Sergio Amat and María Moncayo. l ^{∞}-Stability for linear multiresolution algorithms: a new explicit approach. Part I: The basic rules and the Daubechies case. Applied Mathematics and Computation, 206(1):74–91, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006309.

Amat:2008:SLMb

[AM08c] Sergio Amat and María Moncayo. l ^{∞}-Stability for linear multiresolution algorithms: a new explicit approach. Part II: The cases of Symlets, Coiflets, biorthogonal wavelets and supercompact multiwavelets. Applied Mathematics and Computation, 206(1):92–103, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006310.

Amat:2008:SLMc

[AM08d] Sergio Amat and María Moncayo. l ^{∞}-Stability for linear multiresolution algorithms: a new explicit approach. Part III: The 2-D case. Applied Mathematics and Computation, 206(1):104–112, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006322.

Anastassiou:2009:GSU

[AM09a] George A. Anastassiou and Razvan A. Mezei. Global smoothness and uniform convergence of smooth Poisson–Cauchy type singular operators. Applied Mathematics and Computation, 215(5):1718–1731, November 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309006729.

Anderson:2009:DFO

[AM09b] Douglas R. Anderson and Feliz Minhós. A discrete fourth-order Lidstone problem with parameters. Applied Mathematics and Computation, 214(2):523–533, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003518.

Amini:2007:ERR

[AMAP07] K. Amini, N. Mahdavi-Amiri, and M. R. Peyghami. Extended reduced rank two Abaffian update schemes in the ABS-type methods. Applied Mathematics and Computation, 185(1):255–265, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008575.

Amiraliyev:2005:CFD

[Ami05] G. M. Amiraliyev. The convergence of a finite difference method on layeradapted mesh for a singularly perturbed system. Applied Mathematics and Computation, 162(3):1023–1034, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001468.

Amirteimoori:2006:DEA

[Ami06] Alireza Amirteimoori. Data envelopment analysis in dynamic framework. Applied Mathematics and Computation, 181(1):21–28, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000622.

Amirteimoori:2007:DEAa

[Ami07] Alireza Amirteimoori. DEA efficiency analysis: Efficient and anti-efficient frontier. Applied Mathematics and Computation, 186(1):10–16, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306007934.

Aghigh:2006:NIM

[AMJD06a] Kamal Aghigh, M. Masjed-Jamei, and Mehdi Dehghan. On numerical integration methods with the generalized Stieltjes weight function. Applied Mathematics and Computation, 182(2):1184–1190, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004322.

Aghigh:2006:SSO

[AMJD06b] Kamal Aghigh, M. Masjed-Jamei, and Mehdi Dehghan. A symmetric sequence of orthogonal polynomials associated with the Stieltjes–Wigert polynomials. Applied Mathematics and Computation, 182(1):194–199, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003250.

Aghigh:2008:STF

[AMJD08] Kamal Aghigh, M. Masjed-Jamei, and Mehdi Dehghan. A survey on third and fourth kind of Chebyshev polynomials and their applications. Applied Mathematics and Computation, 199(1):2–12, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009605.

AbdulHadi:2006:SPS

[AMK06] Z. AbdulHadi, Y. Abu Muhanna, and S. Khuri. On some properties of solutions of the biharmonic equation. Applied Mathematics and Computation, 177(1):346–351, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500915X.

Ammar:2005:CSV

[Amm05]  E. E. Ammar.  On

computational solution of vector maximum problem. Applied Mathematics and Computation, 167(1):167–178, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004825.

Ammar:2006:SDN

[Amm06] E. E. Ammar. On solutions and duality of nonlinear nonsmooth fractional programs. Applied Mathematics and Computation, 172(2):865–875, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305002420.

Afrouzi:2006:CAS

[AMN06] G. A. Afrouzi, S. Mahdavi, and Z. Naghizadeh. A computational algorithm for sublinear elliptic partial differential equations. Applied Mathematics and Computation, 183(1):610–616, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005960.

Afrouzi:2007:NAFb

[AMN07a] G. A. Afrouzi, S. Mahdavi, and Z. Naghizadeh. A numerical algorithm for finding positive solutions for classes of p-Laplacian equations. Applied Mathematics and Computation, 187(2):1126–1130, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012495.

Afrouzi:2007:NAFc

[AMN07b] G. A. Afrouzi, S. Mahdavi, and Z. Naghizadeh. A numerical algorithm for finding solution of multiparameter semipositone Dirichlet problems. Applied Mathematics and Computation, 190(1):287–291, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000483.

Afrouzi:2007:NMFg

[AMN07c] G. A. Afrouzi, S. Mahdavi, and Z. Naghizadeh. A numerical method for finding positive solution of diffusive logistic equation with constant yield harvesting. Applied Mathematics and Computation, 191(1):234–238, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002500.

Afrouzi:2007:NMFe

[AMN07d] G. A. Afrouzi, S. Mahdavi, and Z. Naghizadeh. A numerical method for finding positive solution of elliptic equation with Neumann boundary condition. Applied Mathematics and Computation, 189(1):23– 26, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015736.

Afrouzi:2007:NMOb

[AMN07e] G. A. Afrouzi, S. Mahdavi, and Z. Naghizadeh. A numerical method to obtain positive solution for classes of sublinear semipositone systems. Applied Mathematics and Computation, 186(2):1113–1119, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030601037X.

Afrouzi:2007:NMFa

[AMN07f] G. A. Afrouzi, S. Mahdavi, and Z. Naghizadeh. Numerical methods for finding multiple solutions of a semilinear elliptic equation. Applied Mathematics and Computation, 186(1):801–805, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010010.

Afrouzi:2007:OSA

[AMN07g] G. A. Afrouzi, S. Mahdavi, and Z. Naghizadeh. On optimal scaling algorithm for finding positive solution of elliptic equation. Applied Mathematics and Computation, 189(2):1255–1259, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017103.

Agamaliyev:2005:ESB

[AN05a] Agamali Agamaliyev and Anar Nabiyev. On eigenvalues of some boundary value problems for a polynomial pencil of Sturm– Liouville equation. Applied Mathematics and Computation, 165(3):503–515, June 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003984.

An:2005:MPC

[An05b] Yukun An. Maximum principles for a coupled system of second and fourth order elliptic equations and an application. Applied Mathematics and Computation, 161(1):121–127, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303012141.

Amirteimoori:2006:RTM

[AN06] Alireza Amirteimoori and Reza A. Nashtaei. The role of time in multicomponent efficiency analysis: an application. Applied Mathematics and Computation, 177(1):11– 17, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008829.

Abbasi:2007:MHY

[AN07] Babak Abbasi and Seyed Taghi Akhavan Niaki. Monitoring high-yields processes with defects count in nonconforming items by artificial neural network. Applied Mathematics and Computation, 188(1):262– 270, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306013488.

An:2009:DGF

[An09]  Rong An.  Discontin-

uous Galerkin finite element method for the fourth-order obstacle problem. Applied Mathematics and Computation, 209(2):351–355, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009715.

Arkat:2007:ANN

[ANA07] Jamal Arkat, Seyed Taghi Akhavan Niaki, and Babak Abbasi. Artificial neural networks in applying MCUSUM residuals charts for AR(1) processes. Applied Mathematics and Computation, 189(2):1889–1901, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601784X.

Andersson:2006:CBN

[And06] Helge I. Andersson. Comments on BVPs in non-Newtonian boundary-layer flow. Applied Mathematics and Computation, 182(2):1714–1716, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005200.

Andrei:2009:ACG

[And09]  Neculai Andrei.  Ac-

celeration of conjugate gradient algorithms for unconstrained optimization. Applied Mathematics and Computation, 213(2):361–369, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002380.

Ang:2005:TSD

[Ang05] Whye-Teong Ang. A timestepping dual-reciprocity boundary element method for anisotropic heat diffusion subject to specification of energy. Applied Mathematics and Computation, 162(2):661–678, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001213.

Ang:2006:NSN

[Ang06] Whye-Teong Ang. Numerical solution of a nonclassical parabolic problem: an integro-differential approach. Applied Mathematics and Computation, 175(2):969–979, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006557.

Ayasun:2005:EMC

[ANK05] Saffet Ayasun, Chika O.

Nwankpa, and Harry G. Kwatny. An efficient method to compute singularity induced bifurcations of decoupled parameterdependent differential-algebraic power system model. Applied Mathematics and Computation, 167(1):435– 453, August 5, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030400503X.

Afrouzi:2006:CAF

[ANM06] G. A. Afrouzi, Z. Naghizadeh, and S. Mahdavi. A computational algorithm for finding positive solutions for a class of superlinear Dirichlet BVP. Applied Mathematics and Computation, 183(2):1381–1385, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006862.

Afrouzi:2007:NAFa

[ANM07a] G. A. Afrouzi, Z. Naghizadeh, and S. Mahdavi. A numerical algorithm for finding solutions of p-Laplacian Dirichlet problems. Applied Mathematics and Computation, 185(1):213–217, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008526.

Afrouzi:2007:NMFf

[ANM07b] G. A. Afrouzi, Z. Naghizadeh, and S. Mahdavi. A numerical method for finding positive solution of diffusive logistic equation. Applied Mathematics and Computation, 190(2):1730–1734, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002196.

Afrouzi:2007:NMFc

[ANM07c] G. A. Afrouzi, Z. Naghizadeh, and S. Mahdavi. A numerical method for finding positive solution of elliptic systems with nonlinear diffusion in population dynamics. Applied Mathematics and Computation, 187(2):957–961, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012306.

Afrouzi:2007:NMFb

[ANM07d] G. A. Afrouzi, Z. Naghizadeh, and S. Mahdavi. A numerical method for finding positive solution of logistic equation. Applied Mathematics and Computation, 186(2):1497–1501, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010757.

Afrouzi:2007:NMOa

[ANM07e] G. A. Afrouzi, Z. Naghizadeh, and S. Mahdavi. A numerical method to obtain positive solution for classes of sublinear semipositone problems. Applied Mathematics and Computation, 184(2):445–450, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007193.

Afrouzi:2007:NMFd

[ANM07f] G. A. Afrouzi, Z. Naghizadeh, and S. Mahdavi. Numerical methods for finding multiple solutions of a logistic equation. Applied Mathematics and Computation, 188(1):314–321, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306013543.

Afrouzi:2007:SAF

[ANM07g] G. A. Afrouzi, Z. Naghizadeh, and S. Mahdavi. On scaling algorithm for finding positive solution of elliptic equation. Applied Mathematics and Computation, 189(1):298–301, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016031.

Afrouzi:2007:TNA

[ANM07h] G. A. Afrouzi, Z. Naghizadeh, and S. Mahdavi. Two numerical algorithms for finding solutions of multiparameter semipositone Dirichlet problems. Applied Mathematics and Computation, 189(1):201–207, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030601592X.

Afrouzi:2007:TNMb

[ANM07i] G. A. Afrouzi, Z. Naghizadeh, and S. Mahdavi. Two numerical methods for finding multiple solutions of a logistic equation. Applied Mathematics and Computation, 193(1):203–210, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003906.

Afrouzi:2007:TNMa

[ANM07j] G. A. Afrouzi, Z. Naghizadeh, and S. Mahdavi. Two numerical methods for finding multiple solutions of a superlinear Dirichlet problem. Applied Mathematics and Computation, 188(1):981–988, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014469.

Afrouzi:2008:CMO

[ANM08] G. A. Afrouzi, Z. Naghizadeh, and S. Mahdavi. Computational method to obtain positive solution for classes of Laplacian systems with sign-changing weight functions. Applied Mathematics and Computation, 195(2):460–465, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005681.

Anonymous:2005:AIVe

[Ano05a] Anonymous. Author index to volume. Applied Mathematics and Computation, 163(3):1273–1278, April 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305001475.

Anonymous:2005:AIVg

[Ano05b] Anonymous. Author index to volume. Applied Mathematics and Computation, 165(3):759–763, June 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305003346.

Anonymous:2005:AIVh

[Ano05c] Anonymous. Author index to volume. Applied Mathematics and Computation, 166(3):??, July 26, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005035.

Anonymous:2005:AIVa

[Ano05d] Anonymous. Author index to volume 160. Applied Mathematics and Computation, 160(3):955–959, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007179.

Anonymous:2005:AIVc

[Ano05e] Anonymous. Author index to volume 161. Applied Mathematics and Computation, 161(3):1041–1046, February 25, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304007611.

Anonymous:2005:AIVd

[Ano05f] Anonymous. Author index to volume 162. Applied Mathematics and Computation, 162(3):??, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009361.

Anonymous:2005:AIVf

[Ano05g] Anonymous. Author index to volume 164. Applied Mathematics and Computation, 164(3):935–939, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002663.

Anonymous:2005:AIVb

[Ano05h]  Anonymous.  Author

index to volumes 151– 160. Applied Mathematics and Computation, 160(3):961–999, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007222.

Anonymous:2005:AIVi

[Ano05i] Anonymous. Author index to Volumes 161–

170. Applied Mathematics and Computation, 170(2):1469–1476, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008453.

Anonymous:2005:EBa

[Ano05j]  Anonymous.  Editorial

board. Applied Mathematics and Computation, 160 (1):ii, January 5, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304006897.

Anonymous:2005:EBb

[Ano05k]  Anonymous.  Editorial

board. Applied Mathematics and Computation, 161(1):ii, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400726X.

Anonymous:2005:EBc

[Ano05l]  Anonymous.  Editorial

board. Applied Mathematics and Computation, 162 (1):ii, March 4, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304008847.

Anonymous:2005:EBd

[Ano05m]  Anonymous.  Editorial

board. Applied Mathematics and Computation, 163 (1):ii, April 5, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304009439.

Anonymous:2005:EBe

[Ano05n]  Anonymous.  Editorial

board. Applied Mathematics and Computation, 164 (1):ii, May 5, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030500233X.

Anonymous:2005:EBf

[Ano05o]  Anonymous.  Editorial

board. Applied Mathematics and Computation, 165 (1):ii, June 6, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305003048.

Anonymous:2005:EBg

[Ano05p]  Anonymous.  Editorial

board. Applied Mathematics and Computation, 166 (1):ii, July 6, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305003474.

Anonymous:2006:AIVa

[Ano06a]  Anonymous.  Author

index to volumes 171–

180. Applied Mathematics and Computation, 180(2):722–725, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013099.

Anonymous:2006:AIVb

[Ano06b]  Anonymous.  Author

index to volumes 171–

180. Applied Mathematics and Computation, 180(2):726–763, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013142.

Anonymous:2007:AIV

[Ano07a] Anonymous. Author index to volume. Applied Mathematics and Computation, 192(2):592–595, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009411.

Anonymous:2007:MIV

[Ano07b] Anonymous. Master index: Vols 181–190. Applied Mathematics and Computation, 190(2):i–ix, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2007:RN

[Ano07c] Anonymous. Retraction notice. Applied Mathematics and Computation, 189 (1):1010, June 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306013415. See

[EHZES05].

Anonymous:2008:MIV

[Ano08] Anonymous. Master index: Vols 191–200. Applied Mathematics and Computation, 200(2):i–iv, July 1, 2008. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

Antoniou:2005:OLS

[Ant05] George E. Antoniou. n-Order linear state space systems: Computing the transfer function using the DFT. Applied Mathematics and Computation, 170(2):1077–1084, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000482.

Agarwal:2005:MRS

[AO05a] Ravi P. Agarwal and Donal O’Regan. A multiplicity result for second order impulsive differential equations via the Leggett Williams fixed point theorem. Applied Mathematics and Computation, 161(2):433–439, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000037.

Akdeniz:2005:DSS

[AÖ05b] Fikri Akdeniz and Fikri Öztürk. The distribution of stochastic shrinkage biasing parameters of the Liu type estimator. Applied Mathematics and Computation, 163(1):29– 38, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400205X.

Allahviranloo:2005:GQA

[AO05c] T. Allahviranloo and M. Otadi.

Gaussian quadratures for approximate of fuzzy integrals. Applied Mathematics and Computation, 170(2):874–885, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000299.

Arikoglu:2005:SBV

[AO05d] Aytac Arikoglu and Ibrahim Ozkol. Solution of boundary value problems for integro-differential equations by using differential transform method. Applied Mathematics and Computation, 168(2):1145–1158, September 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304006691.

Abbasbandy:2006:NSFb

[AO06a] S. Abbasbandy and M. Otadi.

Numerical solution of fuzzy polynomials by fuzzy neural network. Applied Mathematics and Computation, 181(2):1084–1089, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001925.

Aksan:2006:NSK

[AÖ06b] E. N. Aksan and A. Özdeş. Numerical solution of Korteweg–de Vries equation by Galerkin B-spline finite element method. Applied Mathematics and Computation, 175(2):1256– 1265, April 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305006752.

Allahviranloo:2006:GQA

[AO06c] T. Allahviranloo and M. Otadi. Gaussian quadratures for approximate of fuzzy multiple integrals. Applied Mathematics and Computation, 172(1):175–187, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305001967.

Arikoglu:2006:SDE

[AO06d] Aytac Arikoglu and Ibrahim Ozkol. Solution of difference equations by using differential transform method. Applied Mathematics and Computation, 174(2):1216–1228, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005436.

Arikoglu:2006:SDD

[AO06e] Aytac Arikoglu and Ibrahim Ozkol. Solution of differential– difference equations by using differential transform method. Applied Mathematics and Computation, 181(1):153–162, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000853.

Afuwape:2008:SBS

[AO08] A. U. Afuwape and M. O.

Omeike. On the stability and boundedness of solutions of a kind of third order delay differential equations. Applied Mathematics and Computation, 200(1):444–451, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011514.

Aslan:2009:AST

[AÖ09a] İsmail Aslan and Turgut Öziş. Analytic study on two nonlinear evolution equations by using the (\frac{G′}{G})-expansion method. Applied Mathematics and Computation, 209(2):425–429, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800979X.

Aslan:2009:VRE

[AÖ09b] İsmail Aslan and Turgut Öziş. On the validity and reliability of the (\frac{G′}{G})- expansion method by using higher-order nonlinear equations. Applied Mathematics and Computation, 211(2):531–536, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900112X.

Agarwal:2005:NSB

[AOLL05] Ravi P. Agarwal, Donal O’Regan, V. Lakshmikantham, and S. Leela. Nonresonant singular boundary value problems with sign changing nonlinearities. Applied Mathematics and Computation, 167(2):1236–1248, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005648.

Aksan:2006:NSB

[AÖÖ06] E. N. Aksan, A. Özdeş, and T. Öziş. A numerical solution of Burgers’ equation based on least squares approximation. Applied Mathematics and Computation, 176(1):270–279, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007915.

Agarwal:2008:OGA

[AOS08] R. P. Agarwal, Donal O’Regan, and S. H. Saker. Oscillation and global attractivity in a delay periodic host macroparasite model. Applied Mathematics and Computation, 196(1):340–352, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006765.

Agarwal:2009:SBV

[AOS09] Ravi P. Agarwal, Donal O’Regan, and Veli B. Shakhmurov. B-Separable boundary value problems in Banach-valued function spaces. Applied Mathematics and Computation, 210(1):48–63, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008175.

Aouf:2007:QHP

[Aou07] M. K. Aouf. The quasi-Hadamard products of certain subclasses of analytic p-valent functions with negative coefficients. Applied Mathematics and Computation, 187(1):54– 61, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011362.

Aouf:2008:AEC

[Aou08a]  M. K. Aouf.  Argu-

ment estimates of certain meromorphically multivalent functions associated with generalized hypergeometric function. Applied Mathematics and Computation, 206(2):772–780, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308007066.

Aouf:2008:CSP

[Aou08b] M. K. Aouf. Certain subclasses of p-valent starlike functions defined by using a differential operator. Applied Mathematics and Computation, 206(2):867–875, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800708X.

Amini:2006:CAI

[AP06] K. Amini and M. R.

Peyghami. Complexity analysis of interiorpoint methods for linear optimization based on some conditions on kernel function. Applied Mathematics and Computation, 176(1):194–207, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007848.

Amini:2009:ECL

[AP09] Keyvan Amini and M. Reza Peyghami. Exploring complexity of large update interior-point methods for P_{∗}(κ) linear complementarity problem based on Kernel function. Applied Mathematics and Computation, 207(2):501–513, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008370.

Azaron:2005:GAA

[APS05] Amir Azaron, Cahit Perkgoz, and Masatoshi Sakawa. A genetic algorithm approach for the time-cost trade-off in PERT networks. Applied Mathematics and Computation, 168(2):1317–1339, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304007040.

Al-Rawwash:2005:CME

[AR05] Mohammad Y. Al-Rawwash. Covariance matrix estimation using repeated measurements when data are incomplete. Applied Mathematics and Computation, 167(1):299–315, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030400493X.

Al-Rabadi:2006:QRC

[AR06] Anas N. Al-Rabadi. Qudits representations and computations of n-player many-valued quantum games. Applied Mathematics and Computation, 175(1):691– 714, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006399.

Andargie:2007:FFO

[AR07a] Awoke Andargie and Y. N.

Reddy. Fitted fourthorder tridiagonal finite difference method for singular perturbation problems. Applied Mathematics and Computation, 192(1):90–100, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002974.

Awoke:2007:EFS

[AR07b] Andargie Awoke and Y. N.

Reddy. An exponentially fitted special second-order finite difference method for solving singular perturbation problems. Applied Mathematics and Computation, 190(2):1767–1782, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002251.

Andre:2008:CCS

[AR08] F. J. André and C. Romero.

Computing compromise solutions: On the connections between compromise programming and composite programming. Applied Mathematics and Computation, 195(1):1–10, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005115.

Abbas:2009:CFP

[AR09] M. Abbas and B. E. Rhoades. Common fixed point results for noncommuting mappings without continuity in generalized metric spaces. Applied Mathematics and Computation, 215(1):262–269, September 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309004561.

Argyros:2009:NMS

[Arg09] Ioannis K. Argyros. On Newton’s method for solving equations containing Fréchet-differentiable operators of order at least two. Applied Mathematics and Computation, 215(4):1553–1560, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006584.

Ariel:2007:CTD

[Ari07]  P. Donald Ariel.  On

computation of the threedimensional flow past a stretching sheet. Applied Mathematics and Computation, 188(2):1244–1250, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014688.

Achache:2007:NSI

[ARK07] M. Achache, H. Roumili, and A. Keraghel. A numerical study of an infeasible primal-dual pathfollowing algorithm for linear programming. Applied Mathematics and Computation, 186(2):1472–1479, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010721.

Arcos:2005:IAI

[ARM^{+}05] A. Arcos, M. Rueda, M. D.

Martínez, S. González, and Y. Román. Incorporating the auxiliary information available in variance estimation. Applied Mathematics and Computation, 160(2):387–399, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303011640.

Ayub:2009:ADO

[ARM09] M. Ayub, M. Ramzan, and A. B. Mann. Acoustic diffraction by an oscillating strip. Applied Mathematics and Computation, 214(1):201–209, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030900294X.

Ali:2007:CBV

[ARS07a] Rosihan M. Ali, V. Ravichandran, and N. Seenivasagan. Coefficient bounds for pvalent functions. Applied Mathematics and Computation, 187(1):35– 46, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011349.

Arsham:2007:CSS

[Ars07b]  H. Arsham.  A com-

putationally stable solution algorithm for linear programs. Applied Mathematics and Computation, 188(2):1549–1561, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015074.

Arsham:2007:CLS

[Ars07c] H. Arsham. Construction of the largest sensitivity region for general linear programs. Applied Mathematics and Computation, 189(2):1435–1447, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017309.

Arsham:2007:HGF

[Ars07d] Hossein Arsham. A hybrid gradient and feasible direction pivotal solution algorithm for general linear programs. Applied Mathematics and Computation, 188(1):596–611, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013981.

Al-Said:2005:QSS

[AS05] Eisa A. Al-Said. Quartic splines solutions for a system of second order boundary value problems. Applied Mathematics and Computation, 166(2):254–264, July 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004345.

Akram:2006:NSS

[AS06a] Ghazala Akram and Shahid S. Siddiqi. Nonic spline solutions of eighth order boundary value problems. Applied Mathematics and Computation, 182(1):829–845, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004012.

Akram:2006:SSO

[AS06b] Ghazala Akram and Shahid S. Siddiqi. Solution of sixth order boundary value problems using non-polynomial spline technique. Applied Mathematics and Computation, 181(1):708–720, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001500.

Al-Sharaeh:2006:DRB

[AS06c] Saleh H. Al-Sharaeh. Dynamic rate-based borrowing scheme for QoS provisioning in high speed multimedia wireless cellular networks. Applied Mathematics and Computation, 179(2):714–724, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010799.

Al-Showaikh:2006:NSA

[AS06d] F. N. M. Al-Showaikh. Numerical solution of an autocatalytic bio-chemical system. Applied Mathematics and Computation, 176 (1):177–193, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305007836.

Amiraliyev:2006:UDMb

[AŞ06e] G. M. Amiraliyev and Sebaheddin Şevgin. Uniform difference method for singularly perturbed Volterra integro-differential equations. Applied Mathematics and Computation, 179(2):731–741, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010817.

Amirteimoori:2006:CEO

[AS06f] Alireza Amirteimoori and Mahmoud Shafiei. Characterizing an equitable omission of shared resources: a DEA-based approach.

Applied Mathematics and Computation, 177(1):18– 23, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008830.

Amirteimoori:2006:MEI

[AS06g] Alireza Amirteimoori and Mahmoud Shafiei. Measuring the efficiency of interdependent decision making sub-units in DEA. Applied Mathematics and Computation, 173(2):847–855, February 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305003978.

Avtar:2006:AOS

[AS06h] Ram Avtar and Rashmi Srivastava. Aqueous outflow in Schlemm’s canal. Applied Mathematics and Computation, 174(1):316– 328, March 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305004583.

Avtar:2006:MFA

[AS06i] Ram Avtar and Rashmi Srivastava. Modelling the flow of aqueous humor in anterior chamber of the eye. Applied Mathematics and Computation, 181(2):1336–1348, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002529.

Ahn:2007:CTT

[AS07a] Jaemin Ahn and Minsu Song. Convergence of the trinomial tree method for pricing European/ American options. Applied Mathematics and Computation, 189(1):575–582, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016316.

Al-Sharaeh:2007:HCM

[AS07b] Saleh H. Al-Sharaeh. On the Hamiltonian cycle mapping onto 3-D torus interconnection network based on base-b reflected Gray codes. Applied Mathematics and Computation, 186(2):1311–1321, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010551.

Avtar:2007:MAH

[AS07c] Ram Avtar and Rashmi Srivastava. Modelling aqueous humor outflow through trabecular meshwork. Applied Mathematics and Computation, 189 (1):734–745, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306016511.

Ahmad:2008:ERM

[AS08] Bashir Ahmad and S. Sivasundaram. Existence results and monotone iterative technique for impulsive hybrid functional differential systems with anticipation and retardation. Applied Mathematics and Computation, 197(2):515– 524, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007977.

Alexe:2009:FAS

[AS09a] Mihai Alexe and Adrian Sandu. Forward and adjoint sensitivity analysis with continuous explicit Runge–Kutta schemes. Applied Mathematics and Computation, 208(2):328– 346, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300308008576.

Avetisyan:2009:ECO

[AS09b] Karen Avetisyan and Stevo Stević. Extended Cesàro operators between different Hardy spaces. Applied Mathematics and Computation, 207(2):346–350, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008187.

Avetisyan:2009:GLT

[AS09c] Karen Avetisyan and Stevo Stević. The generalized Libera transform is bounded on the Besov mixed-norm, BMOA and VMOA spaces on the unit disc. Applied Mathematics and Computation, 213(2):304–311, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900232X.

Abo-Sinna:2005:ETM

[ASA05] Mahmoud A. Abo-Sinna and Azza H. Amer. Extensions of TOPSIS for multiobjective large-scale nonlinear programming problems. Applied Mathematics and Computation, 162(1): 243–256, March 4, 2005.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304000852.

Abbasi:2006:BOR

[ASA06a] Babak Abbasi, Shahram Shadrokh, and Jamal Arkat. Bi-objective resourceconstrained project scheduling with robustness and makespan criteria. Applied Mathematics and Computation, 180(1):146–152, September 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305011057.

Asaithambi:2006:SMN

[Asa06b]  Asai Asaithambi.  A

shooting method for nonlinear heat transfer using automatic differentiation. Applied Mathematics and Computation, 180(1):264–269, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305011215.

Apaydin:2007:WES

[ASA07a] Gökhan Apaydin, Selim Seker, and Niyazi Ari. Weighted extended Bsplines for one-dimensional electromagnetic problems. Applied Mathematics and Computation, 190(2):1125– 1135, July 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307001324.

Asaithambi:2007:NSS

[Asa07b]  Asai Asaithambi.  Nu-

merical solution of Stefan problems using automatic differentiation. Applied Mathematics and Computation, 189(1):943–948, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016730.

Awais:2007:DRK

[ASA07c] M. M. Awais, Shafay Shamail, and Nisar Ahmed. Dimensionally reduced Krylov subspace model reduction for large scale systems. Applied Mathematics and Computation, 191(1):21–30, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012641.

Abo-Sinna:2006:IAL

[ASAEE06] Mahmoud A. Abo-Sinna and Tarek H. M. Abou-El-Enien. An interactive algorithm for large scale multiple objective programming problems with fuzzy parameters through TOP-SIS approach. Applied Mathematics and Computation, 177(2):515–527, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500932X.

Abo-Sinna:2006:IAD

[ASAS06] Mahmoud A. Abo-Sinna, A. H. Amer, and Hend H. El Sayed. An interactive algorithm for decomposing the parametric space in fuzzy multiobjective dynamic programming problem. Applied Mathematics and Computation, 174(1):684–699, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004959.

Ashry:2006:GCM

[Ash06] G. A. Ashry. On globally convergent multi-objective optimization. Applied Mathematics and Computation, 183(1):209–216, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005534.

Ahmad:2009:HTU

[ASH09] I. Ahmad, M. Sajid, and T. Hayat. Heat transfer in unsteady axisymmetric second grade fluid. Applied Mathematics and Computation, 215(5):1685–1695, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006687.

Ahmad:2005:PPA

[ASK05] Rais Ahmad, A. H. Siddiqi, and Zubair Khan. Proximal point algorithm for generalized multivalued nonlinear quasivariational-like inclusions in Banach spaces. Applied Mathematics and Computation, 163(1):295–308, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002267.

Abdelmoula:2007:ACO

[ASL07] R. Abdelmoula, K. Semani, and J. Li. Analysis of cracks originating at the boundary of a circular hole in an infinite plate by using a new conformal mapping approach. Applied Mathematics and Computation, 188(2):1891–1896, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015578.

Aslan:2009:DES

[Asl09a] İsmail Aslan. Discrete exact solutions to some nonlinear differential-difference equations via the (\frac{G′}{G})- expansion method. Applied Mathematics and Computation, 215(8):3140–3147, December 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300309008856.

Aslan:2009:EES

[Asl09b]  İsmail Aslan.  Exact

and explicit solutions to some nonlinear evolution equations by utilizing the (\frac{G′}{G})-expansion method. Applied Mathematics and Computation, 215(2):857–863, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309005426.

Al-Said:2007:NST

[ASN07] Eisa A. Al-Said and Muhammad Aslam Noor. Numerical solutions of third-order system of boundary value problems. Applied Mathematics and Computation, 190(1):332– 338, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000537.

Al-Said:2006:CSM

[ASNR06] Eisa A. Al-Said, Muhammad Aslam Noor, and Themistocles M. Rassias. Cubic splines method for solving fourth-order obstacle problems. Applied Mathematics and Computation, 174(1):180–187, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004492.

Ahrari:2009:GNE

[ASPA09] A. Ahrari, M. Shariat-Panahi, and A. A. Atai. GEM: a novel evolutionary optimization method with improved neighborhood search. Applied Mathematics and Computation, 210(2):376–386, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000058.

Anh:2008:SRD

[AST08] Bui The Anh, Nguyen Khoa Son, and Duong Dang Xuan Thanh. Stability radii of delay difference systems under affine parameter perturbations in infinite dimensional spaces. Applied Mathematics and Computation, 202(2):562–570, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308001410.

Al-Towaiq:2007:CGH

[AT07a] Mohammad Al-Towaiq. Clustered Gauss–Huard algorithm for the solution of Ax = b. Applied Mathematics and Computation, 184(2):485–495, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007247.

Avtar:2007:MTT

[AT07b] Ram Avtar and Deepti Tandon. Modelling the transmural transport of oxygen to the retina. Applied Mathematics and Computation, 186(1):540– 547, March 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306009659.

Ates:2006:PFF

[Ate06] Ülkü Dinlemez Ateş. Properties of the far field operator for time harmonic acoustic waves. Applied Mathematics and Computation, 174(2):1206–1215, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005424.

Abbasbandy:2007:NHA

[ATL07] S. Abbasbandy, Y. Tan, and S. J. Liao. Newtonhomotopy analysis method for nonlinear equations. Applied Mathematics and Computation, 188(2):1794– 1800, May 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306015475.

Azaron:2006:MOR

[ATM06] Amir Azaron and Reza Tavakkoli-Moghaddam. A multi-objective resource allocation problem in dynamic PERT networks. Applied Mathematics and Computation, 181(1):163– 174, October 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306000865.

Attia:2005:NSF

[Att05a] Hazem Ali Attia. Numerical study of flow and heat transfer of a non-Newtonian fluid on a rotating porous disk. Applied Mathematics and Computation, 163(1):327–342, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002280.

Attili:2005:ADM

[Att05b]  Basem S. Attili.  The

Adomian decomposition method for computing eigenelements of Sturm– Liouville two point boundary value problems. Applied Mathematics and Computation, 168(2):1306– 1316, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304007039.

Attia:2006:UMC

[Att06a] Hazem A. Attia. Unsteady MHD Couette flow and heat transfer of dusty fluid with variable physical properties. Applied Mathematics and Computation, 177(1):308–318, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009124.

Attia:2006:CDT

[Att06b] Hazem Ali Attia. Computational dynamics of threedimensional closed-chains of rigid bodies. Applied Mathematics and Computation, 172(1):286–304, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002043.

Attili:2006:CMC

[Att06c] Basem S. Attili. Continuation method for computing certain singular points for index-1 differential algebraic equations. Applied Mathematics and Computation, 175(2):1026–1038, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006594.

Attili:2006:SCG

[Att06d] Basem S. Attili. Superconvergence and the computation of generalized turning points for B.V.P.’s. Applied Mathematics and Computation, 183(2):885–889, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006229.

Attiya:2007:GCB

[Att07]  A. A. Attiya.  On a

generalization class of bounded starlike functions of complex order. Applied Mathematics and Computation, 187(1):62– 67, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011374.

Ali:2007:DAE

[AuH07] Muhammad Ali and Mahmood ul Hassan. A different approach to exact solutions in non-Newtonian second-grade of creeping fluid. Applied Mathematics and Computation, 191(2):484–489, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002767.

Allame:2006:IAS

[AV06] Masoud Allame and Behnaz Vatankhahan. Iteration algorithm for solving Ax = b in max–min algebra. Applied Mathematics and Computation, 175(1):269– 276, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006065.

Akyildiz:2007:OCS

[AV07a] F. Talay Akyildiz and K. Vajravelu. Orthogonal cubic spline collocation method for the nonlinear parabolic equation arising in non-Newtonian fluid flow. Applied Mathematics and Computation, 189 (1):462–471, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306016195.

Angelova:2007:HOF

[AV07b] Ivanka Tr. Angelova and Lubin G. Vulkov. Highorder finite difference schemes for elliptic problems with intersecting interfaces. Applied Mathematics and Computation, 187(2):824– 843, April 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030601215X.

Agarwal:2009:GIV

[AV09] Ravi P. Agarwal and Ram U. Verma. General implicit variational inclusion problems based on A-maximal (m)-relaxed monotonicity (AMRM) frameworks. Applied Mathematics and Computation, 215(1):367–379, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309004780.

Abbas:2009:NCW

[AVKA09] Zulkifly Abbas, S. Vahdati, M. Tavassoli Kajani, and K. A. Atan. New construction of wavelets base on floor function. Applied Mathematics and Computation, 210(2):473–478, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000162.

Akyildiz:2009:IDE

[AVM^{+}09] F. Talay Akyildiz, K. Vajravelu, R. N. Mohapatra, Erik Sweet, and Robert A. Van Gorder. Implicit differential equation arising in the steady flow of a Sisko fluid. Applied Mathematics and Computation, 210(1):189–196, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009922.

Alexandridis:2008:CAM

[AVSB08] A. Alexandridis, D. Vakalis, C. I. Siettos, and G. V. Bafas. A cellular automata model for forest fire spread prediction: The case of the wildfire that swept through Spetses Island in 1990. Applied Mathematics and Computation, 204(1):191–201, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004943.

Ammar:2005:SMT

[AY05a] E. E. Ammar and E. A.

Youness. Study on multiobjective transportation problem with fuzzy numbers. Applied Mathematics and Computation, 166(2):241–253, July 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004333.

Ashyralyev:2005:NSO

[AY05b] A. Ashyralyev and H. A.

Yurtsever. A note on the second order of accuracy difference schemes for hyperbolic–parabolic equations. Applied Mathematics and Computation, 165(3):517–537, June 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003996.

Ariyawansa:2007:CVB

[AZ07] K. A. Ariyawansa and Yuntao Zhu. A class of volumetric barrier decomposition algorithms for stochastic quadratic programming. Applied Mathematics and Computation, 186(2):1683–1693, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011039.

Abe:2008:VAO

[AZ08] Kuniyoshi Abe and Shao-Liang Zhang. A variant algorithm of the Orthomin(m) method for solving linear systems. Applied Mathematics and Computation, 206(1):42–49, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006267.

Averbakh:2009:RRS

[AZ09] I. Averbakh and Y. B.

Zhao. Relaxed robust second-order-cone programming. Applied Mathematics and Computation, 210(2):387–397, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900006X.

Azaron:2006:BSP

[Aza06] Amir Azaron. Bicriteria shortest path in networks of queues. Applied Mathematics and Computation, 182(1):434–442, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600350X.

Azimi:2005:HHE

[Azi05] Zahra Naji Azimi. Hybrid heuristics for Examination Timetabling problem. Applied Mathematics and Computation, 163(2):705–733, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002541.

Biazar:2005:AAS

[BA05a] J. Biazar and A. R. Amirtaimoori. An analytic approximation to the solution of heat equation by Adomian decomposition method and restrictions of the method. Applied Mathematics and Computation, 171(2):738–745, December 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bildik:2005:SPE

[BA05b] Necdet Bildik and Gülhan Avci. On the solution of Poisson equation by alternating group explicit iterative method. Applied Mathematics and Computation, 166(2):411–425, July 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400445X.

Biazar:2006:IFP

[BA06] Jafar Biazar and Alireza Amirteimoori. An improvement to the fixed point iterative method. Applied Mathematics and Computation, 182(1):567–571, November 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300306003651.

Bali:2007:EMF

[BA07] Rekha Bali and Usha Awasthi. Effect of a magnetic field on the resistance to blood flow through stenotic artery. Applied Mathematics and Computation, 188(2):1635–1641, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015153.

Biswal:2009:TMC

[BA09] M. P. Biswal and S. Acharya.

Transformation of a multichoice linear programming problem. Applied Mathematics and Computation, 210(1):182–188, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009910.

Braselton:2005:SMC

[BAB05] James Braselton, Martha Abell, and Lorraine Braselton. Selective mating in a continuous model of epistasis. Applied Mathematics and Computation, 171(1):225–241, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000913.

Babak:2007:NPI

[Bab07] Petro Babak. Nonlocal problems involving spatial structure for coupled reaction-diffusion systems. Applied Mathematics and Computation, 185(1):449–463, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008794.

Braselton:2008:EGI

[BAB08] James Braselton, Martha Abell, and Lorraine Braselton. Effects of genetic immunity and resistance on endemics. Applied Mathematics and Computation, 197(2):733–744, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008326.

Badr:2005:NSF

[Bad05] Abdalah A. Badr. Numerical solution of Fredholm– Volterra integral equation in one dimension with time dependent. Applied Mathematics and Computation, 167(2):1156–1161, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005570.

Babolian:2008:NMS

[BAF08] E. Babolian, S. Abbasbandy, and F. Fattahzadeh. A numerical method for solving a class of functional and two dimensional integral equations. Applied Mathematics and Computation, 198(1):35– 43, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008648.

Bahadir:2005:EFD

[Bah05] A. Refik Bahadır. Exponential finite-difference method applied to Korteweg– de Vries equation for small times. Applied Mathematics and Computation, 160(3):675–682, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303011871.

Bai:2008:GSA

[Bai08]  Chuanzhi Bai.  Global

stability of almost periodic solutions of Hopfield neural networks with neutral time-varying delays. Applied Mathematics and Computation, 203(1):72– 79, September 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308002075.

Baklizi:2005:PTE

[Bak05] Ayman Baklizi. Preliminary test estimation in the two parameter exponential distribution with time censored data. Applied Mathematics and Computation, 163(2):639–643, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002498.

Bujurke:2007:CES

[BAK07] N. M. Bujurke, V. B. Awati, and N. N. Katagi. Computer extended series solution for flow in a narrow channel of varying gap. Applied Mathematics and Computation, 186(1):54– 69, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009167.

Balakrishnan:2005:SS

[Bal05] A. V. Balakrishnan. Superstability of systems.

Applied Mathematics and Computation, 164(2):321– 326, May 16, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304003194.

Boucherif:2007:NTP

[BAM07] Abdelkader Boucherif and Nawal Al-Malki. Nonlinear three-point thirdorder boundary value problems. Applied Mathematics and Computation, 190(2):1168–1177, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001476.

Baran:2005:NPD

[Bar05a] Emine Can Baran. Numerical procedures for determining of an unknown parameter in parabolic equation. Applied Mathematics and Computation, 162(3):1219–1226, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001821.

Barrio:2005:PTS

[Bar05b] Roberto Barrio. Performance of the Taylor series method for ODEs/ DAEs. Applied Mathematics and Computation, 163(2):525–545, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002401.

Barbagallo:2009:ERS

[Bar09] Annamaria Barbagallo. Existence and regularity of solutions to nonlinear degenerate evolutionary variational inequalities with applications to dynamic network equilibrium problems. Applied Mathematics and Computation, 208(1):1–13, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308007923.

Basu:2008:TFO

[Bas08] Dhiman Basu. From third to fourth order variant of Newton’s method for simple roots. Applied Mathematics and Computation, 202(2):886–892, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308001161.

Bataller:2008:UGF

[Bat08a] Rafael Cortell Bataller. On unsteady gravity flows of a power-law fluid through a porous medium. Applied Mathematics and Computation, 196(1):356–362, February 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307006881.

Bataller:2008:REBb

[Bat08b] Rafael Cortell Bataller.

Radiation effects for the Blasius and Sakiadis flows with a convective surface boundary condition. Applied Mathematics and Computation, 206(2):832– 840, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308007339.

Bataller:2008:REBa

[Bat08c] Rafael Cortell Bataller.

Radiation effects in the Blasius flow. Applied Mathematics and Computation, 198(1):333–338, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008934.

Bawa:2005:SBC

[Baw05] Rajesh K. Bawa. Spline based computational technique for linear singularly perturbed boundary value problems. Applied Mathematics and Computation, 167(1):225–236, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004874.

Bayrak:2005:NSM

[Bay05a] Aylin Bayrak. Numerical solution of mixed boundary value problem for some nonlinear elliptic equation. Applied Mathematics and Computation, 166(1):204– 212, July 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030400431X.

Bayram:2005:DCR

[Bay05b] Mustafa Bayram. Derivation of conservation relationships for catalytic cycles using MAPLE. Applied Mathematics and Computation, 160(1):189–195, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303011238.

Bazan:2008:CPM

[Baz08] F. S. V. Bazán. Chebyshev pseudospectral method for computing numerical solution of convection– diffusion equation. Applied Mathematics and Computation, 200(2):537–546, July 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011332.

Bildik:2005:STD

[BB05] Necdet Bildik and Hatice Bayramoglu. The solution of two dimensional nonlinear differential equation by the Adomian decomposition method. Applied Mathematics and Computation, 163(2):519–524, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002395.

Bandyopadhyay:2006:SSP

[BB06a] M. Bandyopadhyay and Sandip Banerjee. A stagestructured prey–predator model with discrete time delay. Applied Mathematics and Computation, 182(2):1385–1398, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004553.

Bhatta:2006:NSK

[BB06b] Dambaru D. Bhatta and Muhammad I. Bhatti. Numerical solution of KdV equation using modified Bernstein polynomials. Applied Mathematics and Computation, 174(2):1255–1268, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005461.

Bhatti:2006:CII

[BB06c] M. Bhatti and P. Bracken. The calculation of integrals involving B-splines by means of recursion relations. Applied Mathematics and Computation, 172(1):91–100, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030500189X.

Bougoffa:2006:AMS

[BB06d] Lazhar Bougoffa and Smail Bougouffa. Adomian method for solving some coupled systems of two equations. Applied Mathematics and Computation, 177(2):553–560, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009355.

Bourchtein:2006:GOG

[BB06e] Andrei Bourchtein and Ludmila Bourchtein. On generation of orthogonal grids. Applied Mathematics and Computation, 173(2):767–781, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003917.

Boaca:2009:UNA

[BB09] Tudor Boaca and Ioana Boaca. An unified numerical approach of steady convection between two parallel plates. Applied Mathematics and Computation, 215(7):2673–2685, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008054.

Benbouziane:2008:ERI

[BBB08] Zehour Benbouziane, Abdelkader Boucherif, and Sidi Mohamed Bouguima. Existence result for impulsive third order periodic boundary value problems. Applied Mathematics and Computation, 206(2):728–737, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006978.

Blandin:2009:NST

[BBCP09] S. Blandin, G. Bretti, A. Cutolo, and B. Piccoli. Numerical simulations of traffic data via fluid dynamic approach. Applied Mathematics and Computation, 210(2):441–454, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000137.

Babolian:2007:MPS

[BBES07] E. Babolian, M. Bromilow, R. England, and M. Saravi. A modification of pseudo-spectral method for solving a linear ODEs with singularity. Applied Mathematics and Computation, 188(2):1260–1266, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030601472X.

Barcelos:2007:GPB

[BBMN07] Celia A. Zorzo Barcelos, Marcos Aurélio Batista, Adriana M. Martins, and Antonio Carlos Nogueira. Geodesic path based digital inpainting. Applied Mathematics and Computation, 189(2):1662–1674, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017589.

Badia:2009:MAT

[BC09a] Santiago Badia and Ramon Codina. On a multiscale approach to the transient Stokes problem: Dynamic subscales and anisotropic space–time discretization. Applied Mathematics and Computation, 207(2):415–433, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008291.

Bao:2009:EUS

[BC09b] Haibo Bao and Jinde Cao.

Existence and uniqueness of solutions to neutral stochastic functional differential equations with infinite delay. Applied Mathematics and Computation, 215(5):1732–1743, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006730.

Bocci:2009:TIM

[BC09c] Cristiano Bocci and Filip Cools. A tropical interpretation of m-dissimilarity maps. Applied Mathematics and Computation, 212(2):349–356, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001507.

Bournez:2009:CPP

[BCC^{+}09] Olivier Bournez, Philippe Chassaing, Johanne Cohen, Lucas Gerin, and Xavier Koegler. On the convergence of population protocols when population goes to infinity. Applied Mathematics and Computation, 215(4):1340–1350, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030900410X.

Bai:2005:NSN

[BCCK05] K. J. Bai, S. M. Choo, S. K. Chung, and D. Y. Kim. Numerical solutions for nonlinear free surface flows by finite element methods. Applied Mathematics and Computation, 163(2):941–959, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002723.

Barranco-Chamorro:2007:LIR

[BCEGMRMP07] I. Barranco-Chamorro, A. Enguix-González, J. L. Moreno-Rebollo, and J. M. Muñoz-Pichardo. Local influence on the ratio and Horvitz–Thompson estimators through the second order approach. Applied Mathematics and Computation, 190(1):851–865, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001178.

Bao:2005:PSS

[BCL05] Haiyong Bao, Zhenfu Cao, and Rongxing Lu. Proxy signature scheme using self-certified public keys. Applied Mathematics and Computation, 169(2):1380–1389, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008586.

Bruno:2006:NME

[BCMT06] Maria Giuseppina Bruno, Emanuela Camerini, Angelo Manna, and Alvaro Tomassetti. A new method for evaluating the distribution of aggregate claims. Applied Mathematics and Computation, 176(2):488–505, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500826X.

Bai:2006:P

[BCN06] Zhong-Zhi Bai, Wai-Ki Ching, and Michael K. Ng. Preface. Applied Mathematics and Computation, 172(2):679–681, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009014.

Beccari:2009:SCI

[BCR09] C. Beccari, G. Casciola, and L. Romani. Shape controlled interpolatory ternary subdivision. Applied Mathematics and Computation, 215(3):916–927, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005785.

Bao:2005:ITA

[BCW05] Haiyong Bao, Zhenfu Cao, and Shengbao Wang. Improvement on Tzeng et al.’s nonrepudiable threshold multi-proxy multisignature scheme with shared verification. Applied Mathematics and Computation, 169(2):1419–1430, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008628.

Babolian:2005:NIA

[BD05a] E. Babolian and A. Davari.

Numerical implementation of Adomian decomposition method for linear Volterra integral equations of the second kind. Applied Mathematics and Computation, 165(1):223–227, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003765.

Bhaskar:2005:SDS

[BD05b] T. Gnana Bhaskar and J. Vasundhara Devi. Set differential systems and vector Lyapunov functions. Applied Mathematics and Computation, 165(3):539–548, June 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030400400X.

Babajee:2006:APV

[BD06a] D. K. R. Babajee and M. Z. Dauhoo. An analysis of the properties of the variants of Newton’s method with third order convergence. Applied Mathematics and Computation, 183(1):659–684, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006011.

See note [Ren08a].

Boker:2006:HA

[BD06b] Udi Boker and Nachum Dershowitz. A hypercomputational alien. Applied Mathematics and Computation, 178(1):44– 57, July 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008325.

Boker:2009:IDI

[BD09a] Udi Boker and Nachum Dershowitz. The influence of domain interpretations on computational models. Applied Mathematics and Computation, 215(4):1323–1339, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004196.

Bonanno:2009:NBV

[BD09b] Gabriele Bonanno and Giuseppina D’Aguì. A Neumann boundary value problem for the Sturm– Liouville equation. Applied Mathematics and Computation, 208(2):318–327, February 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308009296.

Basu:2006:TME

[BDCW06] S. Basu, D. DasGupta, S. Chakraborty, and M. Walker. Towards a more efficient dynamic mesh adaptation methodology for continuum discretization in complex engineering problems. Applied Mathematics and Computation, 180(2):469–487, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000099.

Babajee:2008:NLC

[BDDB08] D. K. R. Babajee, M. Z.

Dauhoo, M. T. Darvishi, and A. Barati. A note on the local convergence of iterative methods based on Adomian decomposition method and 3-node quadrature rule. Applied Mathematics and Computation, 200(1):452–458, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011198.

Babolian:2006:AGQ

[BDE06] E. Babolian, Mehdi Dehghan, and M. R. Eslahchi. Application of Gauss quadrature rule in finding bounds for solution of linear systems of equations. Applied Mathematics and Computation, 179(2):707–713, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010787.

Bhalekar:2008:NIM

[BDG08] Sachin Bhalekar and Varsha Daftardar-Gejji. New iterative method: Application to partial differential equations. Applied Mathematics and Computation, 203(2):778–783, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004037.

Berroug:2005:DPP

[BDLS05] Tarik Berroug, Hua Ding, Rabah Labbas, and Boubaker-Khaled Sadallah. On a degenerate parabolic problem in Hölder spaces. Applied Mathematics and Computation, 162(2):811–833, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400133X.

Brigolin:2008:OSP

[BDPP08] D. Brigolin, A. Davydov, R. Pastres, and I. Petrenko. Optimization of shellfish production carrying capacity at a farm scale. Applied Mathematics and Computation, 204(2):532–540, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003007.

Brown:2006:IEI

[BDY06] David Brown, Neşe Dernek, and Osman Yürekli. Identities for the exponential integral and the complementary error transforms. Applied Mathematics and Computation, 182(2):1377–1384, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004541.

Biazar:2005:ASH

[BE05] J. Biazar and H. Ebrahimi. An approximation to the solution of hyperbolic equations by Adomian decomposition method and comparison with characteristics method. Applied Mathematics and Computation, 163(2):633–638, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002486.

Bellamine:2006:NCD

[BE06] F. H. Bellamine and A. Elkamel. Numerical characterization of distributed dynamic systems using tools of intelligent computing and generalized dimensional analysis. Applied Mathematics and Computation, 182(2):1021–1039, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004176.

Bratsos:2008:DAD

[BEF08] Athanassios Bratsos, Matthias Ehrhardt, and Ioannis Th. Famelis. A discrete Adomian decomposition method for discrete nonlinear Schrödinger equations. Applied Mathematics and Computation, 197(1):190–205, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007692.

Behforooz:2006:AIC

[Beh06] Hossein Behforooz. Approximation by integro cubic splines. Applied Mathematics and Computation, 175(1):8–15, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305005874.

Belbas:2005:NHO

[Bel05] S. A. Belbas. New hysteresis operators with applications to counterterrorism. Applied Mathematics and Computation, 170(1):425–439, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009592.

Belbas:2006:SSV

[Bel06] S. A. Belbas. On series solutions of Volterra equations. Applied Mathematics and Computation, 181(2):1287–1304, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306002487.

Belbas:2007:NMO

[Bel07a] S. A. Belbas. A new method for optimal control of Volterra integral equations. Applied Mathematics and Computation, 189(2):1902–1915, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017851.

Belbas:2007:OCG

[Bel07b] S. A. Belbas. Optimal control of Goursat–Darboux systems with discontinuous co-state. Applied Mathematics and Computation, 186(1):101–116, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009210.

Belbas:2008:OCV

[Bel08a] S. A. Belbas. Optimal control of Volterra integral equations in two independent variables. Applied Mathematics and Computation, 202(2):647–665, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001537.

Belbas:2008:RMO

[Bel08b] S. A. Belbas. A reduction method for optimal control of Volterra integral equations. Applied Mathematics and Computation, 197(2):880–890, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008582.

Benhamadou:2007:PRL

[Ben07a] Mongi Benhamadou. Preconditioners for the resolution of the linear systems Ax = b. Applied Mathematics and Computation, 189(1):927–942, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016729.

Benke:2007:ANS

[Ben07b] Roland R. Benke. Analytical and numerical solutions of the expected number of occurrences for combinations of event sequences due to variability. Applied Mathematics and Computation, 190(2):1824–1838, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002329.

Berinde:2009:CFP

[Ber09] Vasile Berinde. A common fixed point theorem for compatible quasi contractive self mappings in metric spaces. Applied Mathematics and Computation, 213(2):348–354, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002367.

Brits:2007:LMO

[BEvdB07] R. Brits, A. P. Engelbrecht, and F. van den Bergh. Locating multiple optima using particle swarm optimization. Applied Mathematics and Computation, 189(2):1859–1883, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017826.

Babolian:2007:NCM

[BF07a] E. Babolian and F. Fattahzadeh. Numerical computation method in solving integral equations by using Chebyshev wavelet operational matrix of integration. Applied Mathematics and Computation, 188(1):1016–1022, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014494.

Babolian:2007:NSD

[BF07b] E. Babolian and F. Fattahzadeh. Numerical solution of differential equations by using Chebyshev wavelet operational matrix of integration. Applied Mathematics and Computation, 188(1):417–426, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013786.

Biazar:2006:MPS

[BFI06] J. Biazar, L. Farrokhi, and M. R. Islam. Modeling the pollution of a system of lakes. Applied Mathematics and Computation, 178(2):423–430, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009872.

Bratsos:2007:PFN

[BFKT07] A. G. Bratsos, I. Th. Famelis, A. Kollias, and Ch. Tsitouras. Phasefitted Numerov type methods. Applied Mathematics and Computation, 184(1):23–29, January 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004814.

Balthazar:2005:SCN

[BFR05] José M. Balthazar, Jorge L. Palacios Felix, and Reyolando MLRF Brasil. Some comments on the numerical simulation of self-synchronization of four non-ideal exciters. Applied Mathematics and Computation, 164(2):615–625, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003418.

Babolian:2007:CAS

[BFR07] E. Babolian, F. Fattahzadeh, and E. Golpar Raboky. A Chebyshev approximation for solving nonlinear integral equations of Hammerstein type. Applied Mathematics and Computation, 189(1):641– 646, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016377.

Bang:2006:SOD

[BG06] Keumseong Bang and JaeGwi Go. The simulation of oxygen diffusion following occlusion. Applied Mathematics and Computation, 183(2):1301–1309, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006771.

Biazar:2007:HVI

[BG07a] J. Biazar and H. Ghazvini.

He’s variational iteration method for solving linear and non-linear systems of ordinary differential equations. Applied Mathematics and Computation, 191(1):287–297, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002561.

Borzabadi:2007:ASC

[BG07b] Akbar H. Borzabadi and Mortaza Gachpazan. Approximate solutions for a class of first order nonlinear difference equations. Applied Mathematics and Computation, 190(2):1108–1115, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001300.

Biazar:2008:NSS

[BG08] J. Biazar and H. Ghazvini. Numerical solution for special non-linear Fredholm integral equation by HPM. Applied Mathematics and Computation, 195(2):681–687, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005978.

Babolian:2005:NSL

[BGA05] E. Babolian, H. Sadeghi Goghary, and S. Abbasbandy. Numerical solution of linear Fredholm fuzzy integral equations of the second kind by Adomian method. Applied Mathematics and Computation, 161(3):733–744, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000256.

Bonotto:2009:OSO

[BGF09] E. M. Bonotto, L. P.

Gimenes, and M. Federson. Oscillation for a second-order neutral differential equation with impulses. Applied Mathematics and Computation, 215(1):1–15, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003877.

Guo:2009:BNR

[bGgLgW09] Hong bo Guo, Xin guo Liu, and Wei guo Wang. The block numerical range of matrix polynomials. Applied Mathematics and Computation, 211(2):265– 283, May 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309001027.

Banik:2006:BGQ

[BGP06] A. D. Banik, U. C. Gupta, and S. S. Pathak. BMAP / G / 1 / N queue with vacations and limited service discipline. Applied Mathematics and Computation, 180(2):707–721, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306000592.

Bonettini:2005:INM

[BGR05] Silvia Bonettini, Emanuele Galligani, and Valeria Ruggiero. An inexact Newton method combined with Hestenes multipliers’ scheme for the solution of Karush–Kuhn–Tucker systems. Applied Mathematics and Computation, 168(1):651–676, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304006307.

Bayon:2007:BPH

[BGRS07] L. Bayón, J. M. Grau, M. M. Ruiz, and P. M. Suárez. A Bolza problem in hydrothermal optimization. Applied Mathematics and Computation, 184(1):12–22, January 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004802.

Bayon:2009:CNS

[BGRS09] L. Bayón, J. M. Grau, M. M. Ruiz, and P. M. Suárez. A constrained and non-smooth hydrothermal problem. Applied Mathematics and Computation, 209(1):10–18, March 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004694.

Bailon:2006:CCG

[BH06] Martin L. Bailon and David J. Horntrop. On the calculation of convolutions with Gaussian Kernels. Applied Mathematics and Computation, 176(1):383– 387, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008131.

Bagheroskoui:2007:IPF

[BH07] M. Bagheroskoui and S. Mohammad Hosseini. Investigation of a preconditioner for the P_{2} − P_{1} finite element solution of Stokes problem. Applied Mathematics and Computation, 186(1):490–495, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009581.

Bhatta:2008:UMB

[Bha08] Dambaru Bhatta. Use of modified Bernstein polynomials to solve KdV– Burgers equation numerically. Applied Mathematics and Computation, 206(1):457–464, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800698X.

Baskan:2009:NSA

[BHCC09] Ozgur Baskan, Soner Haldenbilen, Huseyin Ceylan, and Halim Ceylan. A new solution algorithm for improving performance of ant colony optimization. Applied Mathematics and Computation, 211(1):75– 84, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030900040X.

Baek:2008:KFS

[BHCK08] Sunkyoung Baek, Myunggwon Hwang, Hyunsook Chung, and Pankoo Kim. Kansei factor space classified by information for Kansei image modeling. Applied Mathematics and Computation, 205(2):874– 882, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308003664.

Bhaskar:2008:HAD

[BHN08] T. Gnana Bhaskar, S. Hariharan, and Neela Nataraj. Heatlet approach to diffusion equation on unbounded domains. Applied Mathematics and Computation, 197(2):891–903, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008399.

Bi:2006:MUF

[Bi06] Chunjia Bi. Mortar upwind finite volume element method for convection diffusion problems. Applied Mathematics and Computation, 183(2):831–841, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006175.

Berezansky:2008:PFP

[BI08a] L. Berezansky and L. Idels.

Periodic Fox production harvesting models with delay. Applied Mathematics and Computation, 195(1):142–153, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005267.

Bliman:2008:ASD

[BI08b] Pierre-Alexandre Bliman and Tetsuya Iwasaki. Application of semi-definite programming to robust stability of delay systems. Applied Mathematics and Computation, 200(2):517– 528, July 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011319.

Biazar:2005:SSI

[Bia05]  J. Biazar.  Solution

of systems of integral– differential equations by Adomian decomposition method. Applied Mathematics and Computation, 168(2):1232–1238, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304006976.

Biazar:2006:SEM

[Bia06] J. Biazar. Solution of the epidemic model by Adomian decomposition method. Applied Mathematics and Computation, 173(2):1101–1106, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004145.

Bieniasz:2008:ASB

[Bie08a] L. K. Bieniasz. Adaptive solution of BVPs in singularly perturbed secondorder ODEs, by the extended Numerov method combined with an iterative local grid h-refinement. Applied Mathematics and Computation, 198(2):665– 682, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009344.

Bieniasz:2008:ELA

[Bie08b] L. K. Bieniasz. Experiments with a local adaptive grid h-refinement for the finite-difference solution of BVPs in singularly perturbed secondorder ODEs. Applied Mathematics and Computation, 195(1):196–219, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005310.

Biazar:2005:NAS

[BIK05] J. Biazar, M. Ilie, and A. Khoshkenar. A new approach to the solution of the prey and predator problem and comparison of the results with the Adomian method. Applied Mathematics and Computation, 171(1):486–491, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001104.

Biazar:2006:IAA

[BIK06] J. Biazar, M. Ilie, and A. Khoshkenar. An improvement to an alternate algorithm for computing Adomian polynomials in special cases. Applied Mathematics and Computation, 173(1):582–592, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003772.

Bilgin:2004:LSC

[Bil04] Tunay Bilgin. Lacunary strong A-convergence with respect to a sequence of modulus functions. Applied Mathematics and Computation, 151(3):595–600, April 15, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See corrigendum [Bil07].

Bilgin:2007:CLS

[Bil07] Tunay Bilgin. Corrigendum to “Lacunary strong A-convergence with respect to a sequence of modulus functions” [Appl. Math. Comput. 151 (3) (2004) 595–600]. Applied Mathematics and Computation, 185(1):755, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009015.

See [Bil04].

Biswas:2008:SSI

[Bis08] Anjan Biswas. Soliton– soliton interaction with dual-power law nonlinearity. Applied Mathematics and Computation, 198 (2):605–612, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307009265.

Bouhamidi:2008:NNA

[BJ08] A. Bouhamidi and K. Jbilou.

A note on the numerical approximate solutions for generalized Sylvester matrix equations with applications. Applied Mathematics and Computation, 206(2):687–694, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006693.

Babolian:2005:NIG

[BJE05] E. Babolian, M. Masjed Jamei, and M. R. Eslahchi. On numerical improvement of Gauss–Legendre quadrature rules. Applied Mathematics and Computation, 160(3):779–789, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303011949.

See comment [Pre06a].

Jian:2009:SCN

[bJyKxC09] Jin bao Jian, Xiao yan Ke, and Wei xin Cheng. A superlinearly convergent norm-relaxed SQP method of strongly sub-feasible directions for constrained optimization without strict complementarity. Applied Mathematics and Computation, 214(2):632–644, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309003634.

Blyuss:2005:BMT

[BK05a] Konstantin B. Blyuss and Yuliya N. Kyrychko. On a basic model of a twodisease epidemic. Applied Mathematics and Computation, 160(1):177–187, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303011184.

Bulos:2005:MTA

[BK05b] B. Bulos and Suheil A.

Khuri. On the modified Taylor’s approximation for the solution of linear and nonlinear equations. Applied Mathematics and Computation, 160(3):939–953, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012049.

Bujurke:2006:SRE

[BK06a] N. M. Bujurke and R. B.

Kudenatti. Surface roughness effects on squeeze film poroelastic bearings. Applied Mathematics and Computation, 174(2):1181– 1195, March 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305005400.

Bujurke:2006:ARP

[BK06b] N. M. Bujurke and Ramesh B.

Kudenatti. An analysis of rough poroelastic bearings with reference to lubrication mechanism of synovial joints. Applied Mathematics and Computation, 178(2):309–320, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009768.

Badakhshan:2007:NSN

[BK07a] K. P. Badakhshan and A. V. Kamyad. Numerical solution of nonlinear optimal control problems using nonlinear programming. Applied Mathematics and Computation, 187(2):1511–1519, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012902.

Balachandar:2007:RGE

[BK07b] S. Raja Balachandar and K. Kannan. Randomized gravitational emulation search algorithm for symmetric traveling salesman problem. Applied Mathematics and Computation, 192(2):413–421, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700344X.

Balasubramaniam:2007:ECS

[BK07c] P. Balasubramaniam and E. Karthikeyan. Elliptic curve scalar multiplication algorithm using complementary recoding. Applied Mathematics and Computation, 190(1):51– 56, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000276.

Balasubramaniam:2007:FSS

[BK07d] P. Balasubramaniam and E. Karthikeyan. Fast simultaneous scalar multiplication. Applied Mathematics and Computation, 192(2):399–404, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003426.

Balachandar:2008:NPT

[BK08a] S. Raja Balachandar and K. Kannan. A new polynomial time algorithm for 0–1 multiple knapsack problem based on dominant principles. Applied Mathematics and Computation, 202(1):71–77, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010831.

Balasubramaniam:2008:SGM

[BK08b] P. Balasubramaniam and N. Kumaresan. Solution of generalized matrix Riccati differential equation for indefinite stochastic linear quadratic singular system using neural networks. Applied Mathematics and Computation, 204(2):671–679, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002294.

Bettaibi:2008:AEC

[BK08c] Néji Bettaibi and Jamel El Kamel. Asymptotic expansions for a class of qintegral transforms. Applied Mathematics and Computation, 198(1):433– 444, April 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307009125.

Batista:2009:USM

[BK09] Milan Batista and Abdel Rahman A. Ibrahim Karawia. The use of the Sherman–Morrison– Woodbury formula to solve cyclic block tri-diagonal and cyclic block pentadiagonal linear systems of equations. Applied Mathematics and Computation, 210(2):558–563, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000125.

Badakhshan:2007:UAM

[BKA07] K. P. Badakhshan, A. V.

Kamyad, and A. Azemi. Using AVK method to solve nonlinear problems with uncertain parameters. Applied Mathematics and Computation, 189(1):27– 34, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015748.

Bildik:2006:SDT

[BKBK06] Necdet Bildik, Ali Konuralp, Funda Orakçı Bek, and Semih Küçükarslan. Solution of different type of the partial differential equation by differential transform method and Adomian’s decomposition method. Applied Mathematics and Computation, 172(1):551–567, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305002237.

Borzabadi:2005:SSO

[BKFM05] A. H. Borzabadi, A. V. Kamyad, M. H. Farahi, and H. H. Mehne. Solving some optimal path planning problems using an approach based on measure theory. Applied Mathematics and Computation, 170(2):1418–1435, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000743.

Borzabadi:2006:DAS

[BKM06] Akbar Hashemi Borzabadi, Ali Vahidian Kamyad, and Hamed Hashemi Mehne. A different approach for solving the nonlinear Fredholm integral equations of the second kind. Applied Mathematics and Computation, 173(2):724–735, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305003863.

Buric:2006:CST

[BKP06] Lubor Buřič, Alois Klíč, and Lucie Purmová. Canard solutions and travelling waves in the spruce budworm population model. Applied Mathematics and Computation, 183(2):1039– 1051, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306006394.

Bringsjord:2006:NGA

[BKS^{+}06] Selmer Bringsjord, Owen Kellett, Andrew Shilliday, Joshua Taylor, Bram van Heuveln, Yingrui Yang, Jeffrey Baumes, and Kyle Ross. A new Gödelian argument for hypercomputing minds based on the busy beaver problem. Applied Mathematics and Computation, 176(2):516–530, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008349.

Bokhari:2006:ISC

[BKZ06] A. H. Bokhari, A. H. Kara, and F. D. Zaman. Invariant solutions of certain nonlinear evolution type equations with small parameters. Applied Mathematics and Computation, 182(2):1075–1082, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600422X.

Bai:2006:MIW

[BL06] Yan Ping Bai and Ying Jie Lei. The model of impact waves in a circular elastic bar and application in measuring dynamic characteristic of microstructure. Applied Mathematics and Computation, 173(2):1350–1356, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305004364.

Benterki:2008:IPI

[BL08] D. Benterki and A. Leulmi. An improving procedure of the interior projective method for linear programming. Applied Mathematics and Computation, 199(2):811–819, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010375.

Bao:2006:CIB

[BLH06] Feng Bao, Cheng-Chi Lee, and Min-Shiang Hwang. Cryptanalysis and improvement on batch verifying multiple RSA digital signatures. Applied Mathematics and Computation, 172(2):1195–1200, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002869.

Belhamiti:2008:SBV

[BLLM08] Omar Belhamiti, Rabah Labbas, Keddour Lemrabet, and Ahmed Medeghri. Study of boundary value and transmission problems in the Hölder spaces. Applied Mathematics and Computation, 202(2):608– 619, August 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308001501.

Babolian:2007:WMM

[BLP07] E. Babolian, T. Lotfi, and M. Paripour. Wavelet moment method for solving Fredholm integral equations of the first kind. Applied Mathematics and Computation, 186(2):1467–1471, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601071X.

Bao:2006:KSM

[BLW06a] Liang Bao, Yiqin Lin, and Yimin Wei. Krylov subspace methods for the generalized Sylvester equation. Applied Mathematics and Computation, 175(1):557– 573, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006296.

Bao:2006:MSI

[BLW06b] Liang Bao, Yiqin Lin, and Yimin Wei. A modified simple iterative method for nonsymmetric algebraic Riccati equations arising in transport theory. Applied Mathematics and Computation, 181(2):1499–1504, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002797.

Bu:2009:GIB

[BLZZ09] Changjiang Bu, Min Li, Kuize Zhang, and Lan Zheng. Group inverse for the block matrices with an invertible subblock. Applied Mathematics and Computation, 215(1):132–139, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004111.

Belmiloudi:2005:MNA

[BM05a] A. Belmiloudi and G. Le Meur. Mathematical and numerical analysis of dehydratation of gypsum plasterboards exposed to fire. Applied Mathematics and Computation, 163(3):1023–1041, April 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002760.

Biazar:2005:CMS

[BM05b] J. Biazar and R. Montazeri.

A computational method for solution of the prey and predator problem. Applied Mathematics and Computation, 163(2):841–847, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002644.

Bahuguna:2006:ASN

[BM06a] D. Bahuguna and M. Muslim. Approximation of solutions to non-local history-valued retarded differential equations. Applied Mathematics and Computation, 174(1):165–179, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305004480.

Bose:2006:DCC

[BM06b] Christopher J. Bose and Rua Murray. Dynamical conditions for convergence of a maximum entropy method for Frobenius– Perron operator equations. Applied Mathematics and Computation, 182(1):210–212, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003298.

Bhattacharya:2008:NSS

[BM08] Subhra Bhattacharya and B. N. Mandal. Numerical solution of a singular integro-differential equation. Applied Mathematics and Computation, 195(1):346–350, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005516.

Biswas:2009:OSP

[BM09] Anjan Biswas and Daniela Milovic. Optical solitons in a parabolic law media with fourth order dispersion. Applied Mathematics and Computation, 208(1):299–302, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009041.

Bali:2008:MMR

[BMD08] Rekha Bali, Swati Mishra, and Shraddhya Dubey. A mathematical model for red cell motion in narrow capillary surrounded by tissue. Applied Mathematics and Computation, 196(1):193–199, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006583.

Babolian:2006:NIM

[BMJED06] E. Babolian, M. Masjed-Jamei, M. R. Eslahchi, and Mehdi Dehghan. On numerical integration methods with T -distribution weight function. Applied Mathematics and Computation, 174(2):1314–1320, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005515.

Bergamaschi:2006:PPC

[BMP06] Luca Bergamaschi, Ángeles Martínez, and Giorgio Pini. Parallel preconditioned conjugate gradient optimization of the Rayleigh quotient for the solution of sparse eigenproblems. Applied Mathematics and Computation, 175(2):1694–1715, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007320.

Babolian:2007:UDM

[BMS07] E. Babolian, R. Mokhtari, and M. Salmani. Using direct method for solving variational problems via triangular orthogonal functions. Applied Mathematics and Computation, 191(1):206–217, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002470.

Babolian:2008:UTO

[BMS08] E. Babolian, H. R. Marzban, and M. Salmani. Using triangular orthogonal functions for solving Fredholm integral equations of the second kind. Applied Mathematics and Computation, 201(1–2):452–464, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307012210.

Brown:2008:IPT

[BMY08] David Brown, John Maceli, and Osman Yürekli. Identities and Parseval type relations for the L_{∈}- transform. Applied Mathematics and Computation, 196(1):426–432, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006960.

Bnouhachem:2006:NIM

[BN06] Abdellah Bnouhachem and Muhammad Aslam Noor. A new iterative method for variational inequalities. Applied Mathematics and Computation, 182(2):1673–1682, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005169.

Bnouhachem:2007:NCB

[BN07] Abdellah Bnouhachem and Muhammad Aslam Noor. Numerical comparison between prediction–correction methods for general variational inequalities. Applied Mathematics and Computation, 186(1):496–505, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009593.

Bnouhachem:2008:NMG

[BN08] Abdellah Bnouhachem and Muhammad Aslam Noor. Numerical method for general mixed quasi-variational inequalities. Applied Mathematics and Computation, 204(1):27–36, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004396.

Banas:2009:MMC

[BN09] L’ubomír Baňas and Robert Nürnberg. A multigrid method for the Cahn– Hilliard equation with obstacle potential. Applied Mathematics and Computation, 213(2):290–303, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002318.

Bergamaschi:2006:DOM

[BNdFPS06] Paulo Roberto Bergamaschi, Antônio Carlos Nogueira, and Sezimária de Fátima Pereira Saramago. Design and optimization of 3R manipulators using the workspace features. Applied Mathematics and Computation, 172(1):439–463, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002146.

Batiha:2007:NSG

[BNH07] B. Batiha, M. S. M.

Noorani, and I. Hashim. Numerical simulation of the generalized Huxley equation by He’s variational iteration method. Applied Mathematics and Computation, 186(2):1322– 1325, March 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306010563.

Bnouhachem:2007:MDP

[BNK07] Abdellah Bnouhachem, Muhammad Aslam Noor, and Mohamed Khalfaoui. Modified descent-projection method for solving variational inequalities. Applied Mathematics and Computation, 190(2):1691–1700, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002147.

Bnouhachem:2009:NLQ

[BNKZ09] Abdellah Bnouhachem, Muhammad Aslam Noor, Mohamed Khalfaoui, and Sheng Zhaohan. A new logarithmic-quadratic proximal method for nonlinear complementarity problems. Applied Mathematics and Computation, 215(2):695–706, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309005499.

Bnouhachem:2006:TSI

[BNR06] Abdellah Bnouhachem, Muhammad Aslam Noor, and Th.M. Rassias. Threesteps iterative algorithms for mixed variational inequalities. Applied Mathematics and Computation, 183(1):436–446, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005765.

Bildik:2005:NSH

[BÖ05] Necdet Bildik and Sevil Özlü. On the numerical solution of Helmholtz equation by alternating group explicit (AGE) methods. Applied Mathematics and Computation, 163(1):505– 518, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002383.

Bueno-Orovio:2006:FED

[BO06] Alfonso Bueno-Orovio. Fourier embedded domain methods: Periodic and C^{∞} extension of a function defined on an irregular region to a rectangle via convolution with Gaussian kernels. Applied Mathematics and Computation, 183(2):813–818, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006151.

Bilgil:2005:NAD

[BOB05] Ahmet Bilgil, Baki Ozturk, and Halis Bilgil. A numerical approach to determine viscositydependent segregation in fresh concrete. Applied Mathematics and Computation, 162(1):225–241, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000827.

Bobryk:2008:SSC

[Bob08] Roman V. Bobryk. Stochastic stability of coupled oscillators. Applied Mathematics and Computation, 198(2):544–550, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307009204.

Boglaev:2005:SAA

[Bog05a] Igor Boglaev. Schwarz alternating algorithms for a convection–diffusion problem. Applied Mathematics and Computation, 165(3):647–668, June 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004096.

Boguslavskiy:2005:ALE

[Bog05b] J. A. Boguslavskiy. Algorithms of a linear estimation of a poorly worded matrix and errors of observations — uncertain functions of time. Applied Mathematics and Computation, 164(1):25– 44, May 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002942.

Boguslavskiy:2005:AOS

[Bog05c] J. A. Boguslavskiy. The algorithms of the organized search for identification of parameters of models of nonlinear dynamic system. Applied Mathematics and Computation, 163(1):357– 372, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002309.

Boguslavskiy:2005:IMN

[Bog05d] J. A. Boguslavskiy. Integrated method of the numerical decision of the algebraic equations. Applied Mathematics and Computation, 166(2):324–338, July 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004400.

Boguslavskiy:2005:EMI

[Bog05e] Josef A. Boguslavskiy. Estimation of a matrix of intensities for model of Markov process. Applied Mathematics and Computation, 166(2):282–290, July 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004369.

Boglaev:2008:SSE

[Bog08] Igor Boglaev. The solution of a semilinear evolutionary convection–diffusion problem by a monotone domain decomposition algorithm. Applied Mathematics and Computation, 197(2):536– 547, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008004.

Bor:2009:LPF

[Bor09]  Hüseyin Bor.  Local

properties of factored Fourier series. Applied Mathematics and Computation, 212(1):82–85, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001209.

Boulbrachene:2005:PEE

[Bou05] Messaoud Boulbrachene. Pointwise error estimates for a class of elliptic quasi-variational inequalities with nonlinear source terms. Applied Mathematics and Computation, 161(1):129–138, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303012153.

Bougoffa:2006:AMC

[Bou06a] Lazhar Bougoffa. Adomian method for a class of hyperbolic equations with an integral condition. Applied Mathematics and Computation, 177(2):545–552, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009343.

Bougoffa:2006:SPP

[Bou06b] Lazhar Bougoffa. Solvability of the predator and prey system with variable coefficients and comparison of the results with modified decomposition. Applied Mathematics and Computation, 182(1):383–387, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003456.

Bournez:2006:HMC

[Bou06c]  Olivier Bournez.  How

much can analog and hybrid systems be proved (super-)Turing. Applied Mathematics and Computation, 178(1):58–71, July 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008337.

Bougoffa:2009:TWS

[Bou09a] Lazhar Bougoffa. On the traveling wave solutions for a parabolic–hyperbolic system in linear thermoelasticity. Applied Mathematics and Computation, 215(2):850–853, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309005128.

Boumenir:2009:REP

[Bou09b] Amin Boumenir. The recovery of even polynomial potentials. Applied Mathematics and Computation, 215(8):2914–2926, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309008625.

Boyd:2005:FED

[Boy05] John P. Boyd. Fourier embedded domain methods: extending a function defined on an irregular region to a rectangle so that the extension is spatially periodic and C^{∞}. Applied Mathematics and Computation, 161(2):591–597, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400013X.

Boyd:2006:CRR

[Boy06a] John P. Boyd. Computing real roots of a polynomial in Chebyshev series form through subdivision with linear testing and cubic solves. Applied Mathematics and Computation, 174(2):1642–1658, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005850.

Boyd:2006:FPM

[Boy06b] John P. Boyd. Fourier pseudospectral method with Kepler mapping for travelling waves with discontinuous slope: Application to corner waves of the Ostrovsky–Hunter equation and equatorial Kelvin waves in the four-mode approximation. Applied Mathematics and Computation, 177(1):289–299, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009100.

Boyadzhiev:2007:ESH

[Boy07a] Khristo N. Boyadzhiev.

Evaluation of series with Hurwitz and Lerch zeta function coefficients by using Hankel contour integrals. Applied Mathematics and Computation, 186(2):1559–1571, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010836.

Boyd:2007:TBC

[Boy07b] John P. Boyd. A test, based on conversion to the Bernstein polynomial basis, for an interval to be free of zeros applicable to polynomials in Chebyshev form and to transcendental functions approximated by Chebyshev series. Applied Mathematics and Computation, 188(2):1780–1789, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015451.

Boyd:2008:EDI

[Boy08a] John P. Boyd. Evaluating of Dawson’s Integral by solving its differential equation using orthogonal rational Chebyshev functions. Applied Mathematics and Computation, 204(2):914–919, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800578X.

Boyd:2008:EPC

[Boy08b] John P. Boyd. Exploiting parity in converting to and from Bernstein polynomials and orthogonal polynomials. Applied Mathematics and Computation, 198(2):925–929, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009642.

Barrio:2005:EDP

[BP05] R. Barrio and J. M. Peña.

Evaluation of the derivative of a polynomial in Bernstein form. Applied Mathematics and Computation, 167(1):125–142, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004795.

Boglaev:2006:UCM

[BP06a] Igor Boglaev and Sophie Pack. A uniformly convergent method for a singularly perturbed semilinear reaction–diffusion problem with discontinuous data. Applied Mathematics and Computation, 182(1):244–257, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600333X.

Brizaut:2006:OMD

[BP06b] Jean-Sebastien Le Brizaut and Marc Pogu. An optimization method to determine an entropy parameter in gas dynamics. Applied Mathematics and Computation, 177(2):629–634, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009434.

Barcelos:2009:ABN

[BP09] C. A. Z. Barcelos and V. B. Pires. An automatic based nonlinear diffusion equations scheme for skin lesion segmentation. Applied Mathematics and Computation, 215(1):251–261, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309004536.

Bai:2009:EPS

[BQ09] Zhanbing Bai and Tingting Qiu. Existence of positive solution for singular fractional differential equation. Applied Mathematics and Computation, 215(7):2761–2767, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008157.

Bin:2007:NES

[BQYW07] He Bin, Meng Qing, Long Yao, and Rui Weiguo. New exact solutions of the double sine-Gordon equation using symbolic computations. Applied Mathematics and Computation, 186(2):1334–1346, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010587.

Babu:2007:AFE

[BR07a] A. Ramesh Babu and N. Ramanujam. An asymptotic finite element method for singularly perturbed third and fourth order ordinary differential equations with discontinuous source term. Applied Mathematics and Computation, 191(2):372–380, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002640.

Bretti:2007:LTS

[BR07b] Gabriella Bretti and Paolo E. Ricci. Laguerre-type special functions and population dynamics. Applied Mathematics and Computation, 187(1):89–100, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011404.

Balasubramaniam:2008:GAS

[BR08a] P. Balasubramaniam and R. Rakkiyappan. Global asymptotic stability of stochastic recurrent neural networks with multiple discrete delays and unbounded distributed delays. Applied Mathematics and Computation, 204(2):680–686, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002798.

Benitez:2008:ACD

[BR08b] Julio Benítez and Vladimir Rakočević. Applications of CS decomposition in linear combinations of two orthogonal projectors. Applied Mathematics and Computation, 203(2):761–769, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004013.

Banas:2009:LAA

[BR09a] Józef Banaś and Beata Rzepka. On local attractivity and asymptotic stability of solutions of a quadratic Volterra integral equation. Applied Mathematics and Computation, 213(1):102–111, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002070.

Biswas:2009:SSK

[BR09b] Anjan Biswas and Arjuna Ranasinghe. 1-soliton solution of Kadomtsev– Petviashvili equation with power law nonlinearity. Applied Mathematics and Computation, 214(2):645– 647, August 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030900318X.

Bonanno:2009:MRS

[BR09c] Gabriele Bonanno and Giuseppa Riccobono. Multiplicity results for Sturm– Liouville boundary value problems. Applied Mathematics and Computation, 210(2):294–297, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009909.

Brayanov:2005:UCF

[Bra05] Iliya A. Brayanov. Uniformly convergent finite volume difference scheme for 2D convection-dominated problem with discontinuous coefficients. Applied Mathematics and Computation, 163(2):645–665, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002504.

Brayanov:2006:NST

[Bra06] Iliya A. Brayanov. Numerical solution of a twodimensional singularly perturbed reaction-diffusion problem with discontinuous coefficients. Applied Mathematics and Computation, 182(1):631–643, November 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300306003730.

Bratsos:2009:NPG

[Bra09] A. G. Bratsos. A note on a paper by A. G. Bratsos, M. Ehrhardt and I. Th. Famelis. Applied Mathematics and Computation, 211(1):242–245, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000393.

Bridges:2008:CLC

[Bri08] Thomas J. Bridges. Conservation laws in curvilinear coordinates: a short proof of Vinokur’s Theorem using differential forms. Applied Mathematics and Computation, 202(2):882–885, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001112.

Bonilla:2007:FDE

[BRRGT07] B. Bonilla, M. Rivero, L. Rodríguez-Germá, and J. J. Trujillo. Fractional differential equations as alternative models to nonlinear differential equations. Applied Mathematics and Computation, 187(1):79– 88, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011398.

Bonilla:2007:SLF

[BRT07] B. Bonilla, M. Rivero, and J. J. Trujillo. On systems of linear fractional differential equations with constant coefficients. Applied Mathematics and Computation, 187(1):68– 78, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011386.

Bi:2008:NFS

[BRW08a] Weihong Bi, Hongmin Ren, and Qingbiao Wu. New family of seventh-order methods for nonlinear equations. Applied Mathematics and Computation, 203(1):408–412, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002658.

Bi:2008:NSC

[BRW08b] Weihong Bi, Hongmin Ren, and Qingbiao Wu. A new semilocal convergence theorem of Müller’s method. Applied Mathematics and Computation, 199(1):375–384, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010065.

Bryc:2007:CMF

[Bry07a] Włodzimierz Bryc. Computing moments of free additive convolution of measures. Applied Mathematics and Computation, 194(2):561–567, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004699.

Brychkov:2007:SNS

[Bry07b]  Yu.A. Brychkov.  On

some new series of special functions. Applied Mathematics and Computation, 187(1):101–104, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011416.

Bahadir:2005:MFD

[BS05a] A. Refik Bahadır and Mustafa Sağlam. A mixed finite difference and boundary element approach to one-dimensional Burgers’ equation. Applied Mathematics and Computation, 160(3):663–673, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030301186X.

Belbas:2005:OCV

[BS05b] S. A. Belbas and W. H. Schmidt. Optimal control of Volterra equations with impulses. Applied Mathematics and Computation, 166(3):696–723, July 26, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004679.

Bunnag:2005:GAC

[BS05c] Dhiranuch Bunnag and Min Sun. Genetic algorithm for constrained global optimization in continuous variables. Applied Mathematics and Computation, 171(1):604–636, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305001207.

Banda:2006:NOM

[BS06] Mapundi Banda and Mohammed Seaı¨ d. Nonoscillatory methods for relaxation approximation of Hamilton–Jacobi equations. Applied Mathematics and Computation, 183(1):170–183, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005509.

Bayramoglu:2007:HOR

[BŞ07a] Mehmet Bayramoğlu and Hülya Şahintürk. Higher order regularized trace formula for the regular Sturm–Liouville equation contained spectral parameter in the boundary condition. Applied Mathematics and Computation, 186(2):1591–1599, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010952.

Bihan:2007:JMQ

[BS07b] Nicolas Le Bihan and Stephen J. Sangwine. Jacobi method for quaternion matrix singular value decomposition. Applied Mathematics and Computation, 187(2):1265–1271, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601263X.

Bingjie:2007:GTO

[BS07c] Li Bingjie and Liu Sanyang. On gradient-type optimization method utilizing mixed finite element approximation for optimal boundary control problem governed by bi-harmonic equation. Applied Mathematics and Computation, 186(2):1429–1440, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010691.

Bohner:2007:ABS

[BS07d] Martin Bohner and Stevo Stević. Asymptotic behavior of second-order dynamic equations. Applied Mathematics and Computation, 188(2):1503–1512, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015037.

Bokhari:2008:OCP

[BS08] M. A. Bokhari and I. S. Sadek. Optimal control of parabolic systems with infinite time horizons. Applied Mathematics and Computation, 206(2):678–684, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003913.

Barboteu:2009:SDC

[BS09a] Mikael Barboteu and Mircea Sofonea. Solvability of a dynamic contact problem between a piezoelectric body and a conductive foundation. Applied Mathematics and Computation, 215(8):2978–2991, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008704.

Belbas:2009:OCI

[BS09b] S. A. Belbas and W. H.

Schmidt. Optimal control of impulsive Volterra equations with variable impulse times. Applied Mathematics and Computation, 214(2):353–369, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003208.

Bin-Saad:2009:SPC

[BS09c] Maged G. Bin-Saad. Some presentations, connections and series expansions for the generalized Voigt functions. Applied Mathematics and Computation, 207(2):559–568, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008448.

Basto:2006:NIM

[BSC06] Mário Basto, Viriato Semiao, and Francisco L. Calheiros. A new iterative method to compute nonlinear equations. Applied Mathematics and Computation, 173(1):468–483, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030500367X.

See comments [Pet07].

Boumedjane:2007:SFE

[BSHZ07] Y. Boumedjane, H. Saidi, S. Hassouni, and A. Zerarka. Some first excited energy levels for the generalized Killingbeck potential with the differential quadratic method. Applied Mathematics and Computation, 194(1):243–249, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004638.

Balasubramaniam:2006:SMR

[BSKK06] P. Balasubramaniam, J. Abdul Samath, N. Kumaresan, and A. Vincent Antony Kumar. Solution of matrix Riccati differential equation for the linear quadratic singular system using neural networks. Applied Mathematics and Computation, 182(2):1832–1839, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005327.

Bujurke:2007:FWM

[BSKS07] N. M. Bujurke, C. S.

Salimath, R. B. Kudenatti, and S. C. Shiralashetti. A fast waveletmultigrid method to solve elliptic partial differential equations. Applied Mathematics and Computation, 185(1):667–680, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009027.

Barro:2008:NMS

[BSN^{+}08] Geneviève Barro, Ousséni So, Jean Marie Ntaganda, Benjamin Mampassi, and Blaise Some. A numerical method for some nonlinear differential equation models in biology. Applied Mathematics and Computation, 200(1):28– 33, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030701079X.

Bonneuil:2005:PVT

[BSP05] Noël Bonneuil and Patrick Saint-Pierre. Population viability in three trophiclevel food chains. Applied Mathematics and Computation, 169(2):1086–1105, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008392.

Banda:2006:DVR

[BST06] Mapundi Banda, Mohammed Seaïd, and Ioan Teleaga. Discrete-velocity relaxation methods for large eddy simulation. Applied Mathematics and Computation, 182(1):739– 753, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030600381X.

Barrea:2005:NAM

[BT05] Andrés Barrea and Cristina Turner. A numerical analysis of a model of growth tumor. Applied Mathematics and Computation, 167(1):345–354, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004965.

Beggs:2006:EIP

[BT06a] E. J. Beggs and J. V.

Tucker. Embedding infinitely parallel computation in Newtonian kinematics. Applied Mathematics and Computation, 178(1): 25–43, July 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305008313.

Briozzo:2006:OPS

[BT06b] Adriana C. Briozzo and Domingo A. Tarzia. A one-phase Stefan problem for a non-classical heat equation with a heat flux condition on the fixed face. Applied Mathematics and Computation, 182(1):809–819, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003997.

Boucherif:2008:EPN

[BT08] Abdelkader Boucherif and Christopher C. Tisdell. Existence of periodic and nonperiodic solutions to systems of boundary value problems for first-order differential inclusions with super-linear growth. Applied Mathematics and Computation, 204(1):441– 449, October 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308005328.

Baksalary:2009:CSE

[BT09a] Oskar Maria Baksalary and Götz Trenkler. Column space equalities for orthogonal projectors. Applied Mathematics and Computation, 212(2):519–529, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001842.

Beggs:2009:CNR

[BT09b] E. J. Beggs and J. V. Tucker. Computations via Newtonian and relativistic kinematic systems. Applied Mathematics and Computation, 215(4):1311–1322, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004214.

Biazar:2006:ODS

[BTI06] J. Biazar, M. Tango, and R. Islam. Ozone decomposition of the second order in aqueous solutions. Applied Mathematics and Computation, 177(1):220– 225, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009033.

Benatta:2009:MSB

[BTMB09] M. A. Benatta, A. Tounsi, I. Mechab, and M. Bachir Bouiadjra. Mathematical solution for bending of short hybrid composite beams with variable fibers spacing. Applied Mathematics and Computation, 212(2):337–348, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001490.

Bilbao-Terol:2006:SOP

[BTPGAl06] A. Bilbao-Terol, B. Pérez-Gladish, and J. Antomillbias. Selecting the optimum portfolio using fuzzy compromise programming and Sharpe’s single-index model. Applied Mathematics and Computation, 182(1):644–664, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003742.

Bilbao-Terol:2006:FCP

[BTPGAPRU06] Amelia Bilbao-Terol, Blanca Pérez-Gladish, Mar Arenas-Parra, and Maria Victoria Rodríguez-Uría. Fuzzy compromise programming for portfolio selection. Applied Mathematics and Computation, 173(1):251– 264, February 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305003413.

Bayramov:2007:CEE

[BUÇ07] A. Bayramov, S. Öztürk Uslu, and S. Kizilbudak Çalişkan. Computation of eigenvalues and eigenfunctions of a discontinuous boundary value problem with retarded argument. Applied Mathematics and Computation, 191(2):592–600, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002871.

Bukhari:2005:DPI

[Buk05] Abdul-Fattah K. Bukhari.

On dynamic problem and integral equation. Applied Mathematics and Computation, 160(1):51–63, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303007264.

Bukhari:2009:NMS

[Buk09] Abdul-Fattah K. A. Bukhari.

Numerical methods of singular integral equation for solving some problem in fluid dynamic. Applied Mathematics and Computation, 213(1):39– 46, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001787.

Bulut:2009:GGA

[Bul09] Serap Bulut. The generalization of the generalized Al-Oboudi differential operator. Applied Mathematics and Computation, 215(4):1448–1455, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005724.

Bahuguna:2007:ATC

[BUP07] D. Bahuguna, A. Ujlayan, and D. N. Pandey. Advanced type coupled matrix Riccati differential equation systems with Kronecker product. Applied Mathematics and Computation, 194(1):46–53, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004419.

Boyd:2006:RTG

[BV06] John P. Boyd and Robert M. Visser. Rootfinding through global Newton iteration and Chebyshev polynomials for the amplitude of an electronic oscillator. Applied Mathematics and Computation, 182(1):166–174, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003237.

Babolian:2005:SDE

[BVC05] E. Babolian, A. R. Vahidi, and Gh. Asadi Cordshooli. Solving differential equations by decomposition method. Applied Mathematics and Computation, 167(2):1150–1155, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005569.

Bai:2005:GAM

[BW05a] Ling Bai and Ke Wang.

Gilpin–Ayala model with spatial diffusion and its optimal harvesting policy. Applied Mathematics and Computation, 171(1):531–546, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500113X.

Boey:2005:TPR

[BW05b] K. L. Boey and Patricia J. Y. Wong. Twopoint right focal eigenvalue problems on time scales. Applied Mathematics and Computation, 167(2):1281–1303, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005673.

Bai:2007:DSS

[BW07] Ling Bai and Ke Wang.

A diffusive single-species model with periodic coefficients and its optimal harvesting policy. Applied Mathematics and Computation, 187(2):873–882, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012197.

Boyd:2009:AAC

[BW09] John P. Boyd and Lei Wang. An analytic approximation to the cardinal functions of Gaussian radial basis functions on an infinite lattice. Applied Mathematics and Computation, 215(6):2215–2223, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309007322.

Belato:2005:UTS

[BWB05] Débora Belato, Hans I. Weber, and José M. Balthazar. Using transient and steady state considerations to investigate the mechanism of loss of stability of a dynamical system. Applied Mathematics and Computation, 164(2):605–613, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003406.

Bi:2009:NFE

[BWR09] Weihong Bi, Qingbiao Wu, and Hongmin Ren. A new family of eighthorder iterative methods for solving nonlinear equations. Applied Mathematics and Computation, 214(1):236–245, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002987.

Boyd:2009:DRP

[BX09] John P. Boyd and Fei Xu.

Divergence (Runge Phenomenon) for least-squares polynomial approximation on an equispaced grid and Mock–Chebyshev subset interpolation. Applied Mathematics and Computation, 210(1):158–168, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009867.

Bayram:2005:MCA

[BY05] Mustafa Bayram and Necmettin Yildirim. Metabolic control analysis of trio enzymes system. Applied Mathematics and Computation, 170(2):948–957, November 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305000366.

Bai:2009:NSN

[BZ09] Liang Bai and Guang Zhang. Nontrivial solutions for a nonlinear discrete elliptic equation with periodic boundary conditions. Applied Mathematics and Computation, 210(2):321–333, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308009351.

Zhang:2009:ETA

[bZlDL09] Qing bang Zhang, Ru liang Deng, and Liu Liu. Existence theorem and algorithm for a general implicit variational inequality in Banach space. Applied Mathematics and Computation, 215(2):636–643, September 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300309005438.

Biswas:2008:SPT

[BZZ08a] Anjan Biswas, Chenwi Zony, and Essaid Zerrad. Soliton perturbation theory for the quadratic nonlinear Klein–Gordon equation. Applied Mathematics and Computation, 203(1):153–156, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002191.

Bu:2008:GIC

[BZZ08b] Changjiang Bu, Jiemei Zhao, and Jinshan Zheng. Group inverse for a class 2×2 block matrices over skew fields. Applied Mathematics and Computation, 204(1):45–49, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004529.

Cakir:2005:FDM

[CA05] M. Cakir and G. M. Amiraliyev. A finite difference method for the singularly perturbed problem with nonlocal boundary condition. Applied Mathematics and Computation, 160(2):539–549, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303011809.

Chen:2008:NSC

[CA08a] Yong Chen and Hong-Li An. Numerical solutions of coupled Burgers equations with time- and spacefractional derivatives. Applied Mathematics and Computation, 200(1):87– 95, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010892.

Chen:2008:HPM

[CA08b] Yong Chen and Hongli An. Homotopy perturbation method for a type of nonlinear coupled equations with parameters derivative. Applied Mathematics and Computation, 204(2):764–772, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005511.

Chidume:2007:CMI

[CAA07] C. E. Chidume, M. Abbas, and Bashir Ali. Convergence of the Mann iteration algorithm for a class of pseudocontractive mappings. Applied Mathematics and Computation, 194(1):1–6, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004341.

Cai:2006:TSE

[Cai06a] Donghan Cai. A twosector economic growth model with optimal labor and capital allocation. Applied Mathematics and Computation, 183(2):1359–1377, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006837.

Caimmi:2006:NCB

[Cai06b] R. Caimmi. A necessary condition for best fitting analytical to simulated density profiles in dark matter haloes. Applied Mathematics and Computation, 174(1):447–482, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004789.

Cai:2007:CEE

[Cai07] Zhijie Cai. Convergence and error estimates for meshless Galerkin methods. Applied Mathematics and Computation, 184(2):908–916, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007806.

Cai:2009:BER

[Cai09]  Zhijie Cai.  Best es-

timates of RBF-based meshless Galerkin methods for Dirichlet problem. Applied Mathematics and Computation, 215(6):2149–2153, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309007231.

Cakir:2008:POA

[Çak08] Ozan Çakır. Post-optimality analysis of priority vectors derived from interval comparison matrices by lexicographic goal programming. Applied Mathematics and Computation, 204(1):261–268, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005031.

Chen:2008:FCS

[CAL08] Yong Chen, Hongli An, and Zhibin Li. The function cascade synchronization approach with uncertain parameters or not for hyperchaotic systems. Applied Mathematics and Computation, 197(1):96–110, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007588.

Cao:2006:CPI

[Cao06]  Zhi-Hao Cao.  Com-

parison of performance of iterative methods for singular and nonsingular saddle point linear systems arising from Navier– Stokes equations. Applied Mathematics and Computation, 174(1):630–642, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500490X.

Cao:2009:NPG

[Cao09] Jian Cao. New proofs of generating functions for Rogers–Szegő polynomials. Applied Mathematics and Computation, 207(2):486–492, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008357.

Cash:2007:ACI

[Cas07a] Gordon G. Cash. An algorithm for calculating the independence and vertex-cover polynomials of a graph. Applied Mathematics and Computation, 190(2):1487–1491, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700183X.

Cash:2007:NCG

[Cas07b] Gordon G. Cash. The number of n-cycles in a graph. Applied Mathematics and Computation, 184(2):1080–1083, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007971.

Casti:2007:NE

[Cas07c] John Casti. Note from the Editor. Applied Mathematics and Computation, 191(2):299–301, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001749.

Choi:2008:SEA

[CAS08] Junesang Choi, P. J. Anderson, and H. M. Srivastava. Some q-extensions of the Apostol–Bernoulli and the Apostol–Euler polynomials of order n and the multiple Hurwitz zeta function. Applied Mathematics and Computation, 199(2):723–737, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010594.

Choi:2009:CBE

[CAS09] Junesang Choi, P. J. Anderson, and H. M. Srivastava. Carlitz’s q-Bernoulli and q-Euler numbers and polynomials and a class of generalized q-Hurwitz zeta functions. Applied Mathematics and Computation, 215(3):1185–1208, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900633X.

Choudhuri:2006:RST

[CB06] S. K. Roy Choudhuri and Nupur Bandyopadhyay. Radially symmetric thermo-elastic wave propagation without energy dissipation in an infinitely extended thin plate with a circular hole. Applied Mathematics and Computation, 179(1):267–281, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010350.

Chan:2005:STM

[CC05a] Chao-Wen Chan and Chin-Chen Chang. A scheme for threshold multi-secret sharing. Applied Mathematics and Computation, 166(1):1–14, July 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304004163.

Chang:2005:ASN

[CC05b] Ya-Fen Chang and Chin-Chen Chang. Authentication schemes with no verification table. Applied Mathematics and Computation, 167(2):820–832, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304005296.

Chen:2005:INV

[CC05c] Li-Qun Chen and Chang-Jun Cheng. Instability of nonlinear viscoelastic plates. Applied Mathematics and Computation, 162(3):1453–1463, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001997.

Cui:2005:ASS

[CC05d] Minggen Cui and Zhong Chen. All solutions of system of ill-posed operator equations of the first kind. Applied Mathematics and Computation, 165(3):749–758, June 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004151.

Caglar:2006:BSS

[CC06a] Nazan Caglar and Hikmet Caglar. B-spline solution of singular boundary value problems. Applied Mathematics and Computation, 182(2):1509–1513, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004784.

Chan:2006:CBS

[CC06b] C. Y. Chan and W. Y.

Chan. Complete blowup of solutions for degenerate semilinear parabolic first initial-boundary value problems. Applied Mathematics and Computation, 177(2):777–784, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009598.

Chen:2006:NSC

[CC06c] Xin Chen and Yong-Lin Chen. A necessary and sufficient condition for semiconvergence and optimal parameter of the SSOR method for solving the rank deficient linear least squares problem. Applied Mathematics and Computation, 182(2):1108–1126, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004267.

Caglar:2007:SFO

[CC07a] Hikmet Caglar and Nazan Caglar. Solution of fifth order boundary value problems by using local polynomial regression. Applied Mathematics and Computation, 186(2):952–956, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010204.

Chen:2007:REC

[CC07b] Shinn-Horng Chen and Jyh-Horng Chou. Robust eigenvalue-clustering in a specified circular region for linear uncertain discrete singular systems with state delay. Applied Mathematics and Computation, 186(2):1660–1670, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011015.

Cabada:2008:SGF

[CC08a] Alberto Cabada and José Ángel Cid. On the sign of the Green’s function associated to Hill’s equation with an indefinite potential. Applied Mathematics and Computation, 205(1):303–308, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005882.

Caglar:2008:FDB

[CC08b] Hikmet Caglar and Nazan Caglar. Fifth-degree Bspline solution for a fourthorder parabolic partial differential equations. Applied Mathematics and Computation, 201(1–2):597–603, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000052.

Chang:2008:NAC

[CC08c] Shih-Hsiang Chang and I-Ling Chang. A new algorithm for calculating one-dimensional differential transform of nonlinear functions. Applied Mathematics and Computation, 195(2):799–808, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006194.

Cunha:2008:NBI

[CC08d] F. R. Cunha and H. L. G. Couto. A new boundary integral formulation to describe three-dimensional motions of interfaces between magnetic fluids. Applied Mathematics and Computation, 199(1):70– 83, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009708.

Calude:2009:I

[CC09a] C. S. Calude and J. F. Costa. Introduction. Applied Mathematics and Computation, 215(4):1309– 1310, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0096300309004202.

Cen:2009:ECH

[CC09b] Yi-Gang Cen and Li-Hui Cen. Explicit construction of high-pass filter sequence for orthogonal multiwavelets. Applied Mathematics and Computation, 213(1):235–242, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002239.

Chang:2009:NAC

[CC09c] Shih-Hsiang Chang and I-Ling Chang. A new algorithm for calculating two-dimensional differential transform of nonlinear functions. Applied Mathematics and Computation, 215(7):2486–2494, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900784X.

Ciric:2009:CFP

[ĆC09d] Ljubomir Ćirić and Nenad Cakić. On common fixed point theorems for non-self hybrid mappings in convex metric spaces. Applied Mathematics and Computation, 208(1):90–97, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008503.

Chen:2007:APE

[CCC07] Po-Chih Chen, Chia-Ou Chang, and W. T. Chang Chien. An alternative proof for the explicit equations of motion for mechanical systems with independent non-ideal constraints. Applied Mathematics and Computation, 190(2):1445–1449, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001774.

Caglar:2006:BSI

[CCE06] Hikmet Caglar, Nazan Caglar, and Khaled Elfaituri. B-spline interpolation compared with finite difference, finite element and finite volume methods which applied to two-point boundary value problems. Applied Mathematics and Computation, 175(1):72– 79, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005928.

Chang:2008:GAI

[CCFC08] Pei-Chann Chang, Shih-Hsin Chen, Chin-Yuan Fan, and Chien-Lung Chan. Genetic algorithm integrated with artificial chromosomes for multiobjective flowshop scheduling problems. Applied Mathematics and Computation, 205(2):550–561, November 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308003330.

Chang:2005:DSM

[CCH05] Ya-Fen Chang, Chin-Chen Chang, and Hui-Feng Huang. Digital signature with message recovery using self-certified public keys without trustworthy system authority. Applied Mathematics and Computation, 161(1):211–227, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012220.

Chen:2007:CAP

[CCJR07] B. Chen, R. Company, L. Jódar, and M. D. Roselló. Constructing accurate polynomial approximations for nonlinear differential initial value problems. Applied Mathematics and Computation, 193(2):523–534, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004365.

Chu:2006:SML

[CCK06] Jeong Ho Chu, Hi Jun Choe, and Man-Hoe Kim. Spectral method for linear elasticity. Applied Mathematics and Computation, 182(1):269–282, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003365.

Chang:2006:SIM

[CCL06] Ching-Ter Chang, Chen-Lung Chin, and Mei-Feng Lin. On the single item multi-supplier system with variable leadtime, price-quantity discount, and resource constraints. Applied Mathematics and Computation, 182(1):89–97, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600316X.

Canto:2005:PBP

[CCS05] Begoña Cantó, Carmen Coll, and Elena Sánchez. On positive behaviour of periodic control systems. Applied Mathematics and Computation, 161(3):779–786, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000293.

Chan:2009:GPF

[CCWH09] Hung-Chang Chan, Jou-Ming Chang, Yue-Li Wang, and Shi-Jinn Horng. Geodesicpancyclicity and faulttolerant panconnectivity of augmented cubes. Applied Mathematics and Computation, 207(2):333–339, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008151.

Cao:2008:IMG

[CCWJ08] Dexin Cao, Meirong Chen, Haijun Wang, and Ji Jianxiong. Interval method for global solutions of a class of min–max– min problems. Applied Mathematics and Computation, 196(2):594–602, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307007138.

Chen:2008:SPP

[CCZ08] Fengxiang Chen, Lin Chen, and Weidong Zhang. Stabilization of parameters perturbation chaotic system via adaptive backstepping technique. Applied Mathematics and Computation, 200(1):101–109, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010910.

Cui:2008:PSO

[CCZS08] Zhihua Cui, Xingjuan Cai, Jianchao Zeng, and Guoji Sun. Particle swarm optimization with FUSS and RWS for high dimensional functions. Applied Mathematics and Computation, 205(1):98–108, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003214.

Candan:2005:OBH

[CD05] T. Candan and R. S. Dahiya. Oscillation behavior of higher order neutral differential equations. Applied Mathematics and Computation, 167(2):1267–1280, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005661.

Cui:2006:RES

[CD06] Minggen Cui and Hong Du.

Representation of exact solution for the nonlinear Volterra–Fredholm integral equations. Applied Mathematics and Computation, 182(2):1795–1802, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005261.

Cunha:2007:NSV

[CD07] Maria Cristina C. Cunha and Fábio Antonio Dorini. A numerical scheme for the variance of the solution of the random transport equation. Applied Mathematics and Computation, 190(1):362–369, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000574.

Costabile:2008:NSC

[CD08] F. A. Costabile and F. Dell’Accio.

On the nonsingularity of a special class of centrosymmetric matrices arising in spectral methods in BVPs. Applied Mathematics and Computation, 206(2):991–993, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006462.

Choudhury:2009:URQ

[CD09] Gautam Choudhury and Kandarpa Deka. An M^{X}/G/1 unreliable retrial queue with two phases of service and Bernoulli admission mechanism. Applied Mathematics and Computation, 215(3):936–949, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005803.

Chen:2006:CLE

[CDP06] Z.-M. Chen, K. Djidjeli, and W. G. Price. Computing Lyapunov exponents based on the solution expression of the variational system. Applied Mathematics and Computation, 174(2):982–996, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005254.

Csoka:2007:UFT

[CDT07] Levente Csoka, Ferenc Divos, and Katsuhiko Takata. Utilization of Fourier transform of the absolute amplitude spectrum in wood anatomy. Applied Mathematics and Computation, 193(2):385– 388, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307004110.

Cetin:2006:WTA

[ÇE06a] Eyüp Çetin and Seda Tolun Esen. A weapon–target assignment approach to media allocation. Applied Mathematics and Computation, 175(2):1266–1275, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006764.

Cetin:2006:SSC

[ÇE06b] Meral Çetin and Aydın Erar. A simulation study on classic and robust variable selection in linear regression. Applied Mathematics and Computation, 175(2):1629–1643, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305007204.

Cordero:2007:STAb

[CE07a] Luis F. Cordero and René Escalante. Segmented Tau approximation for a parametric nonlinear neutral differential equation. Applied Mathematics and Computation, 190(1):866– 881, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700118X.

Cordero:2007:STAa

[CE07b] Luis F. Cordero and René Escalante. Segmented Tau approximation for test neutral functional differential equations. Applied Mathematics and Computation, 187(2):725–740, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011271.

Celik:2005:ACEa

[Çel05a] İbrahim Çelik. Approximate calculation of eigenvalues with the method of weighted residuals– collocation method. Applied Mathematics and Computation, 160(2):401– 410, January 14, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300303011688.

Celik:2005:ACEb

[Çel05b] İbrahim Çelik. Approximate computation of eigenvalues with Chebyshev collocation method. Applied Mathematics and Computation, 168(1):125–134, September 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304005910.

Celebi:2006:TDM

[Cel06a] E. Celebi. Three-dimensional modelling of train-track and sub-soil analysis for surface vibrations due to moving loads. Applied Mathematics and Computation, 179(1):209–230, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010313.

Celik:2006:CMR

[Çel06b] İbrahim Çelik. Collocation method and residual correction using Chebyshev series. Applied Mathematics and Computation, 174(2):910–920, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005205.

Cen:2005:HDS

[Cen05] Zhongdi Cen. A hybrid difference scheme for a singularly perturbed convectiondiffusion problem with discontinuous convection coefficient. Applied Mathematics and Computation, 169(1):689–699, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008033.

Cen:2006:NSC

[Cen06] Zhongdi Cen. Numerical study for a class of singular two-point boundary value problems using Green’s functions. Applied Mathematics and Computation, 183(1):10–16, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003341.

Cetin:2006:DOD

[Çet06] Eyüp Çetin. Determining the optimal duration of an advertising campaign using diffusion of information. Applied Mathematics and Computation, 173(1):430–442, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003632.

Cetin:2007:HOM

[Çet07] Eyüp Çetin. Handling organ motion in radiotherapy of cancer via Markov chains. Applied Mathematics and Computation, 184(2):149–155, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006916.

Cenci:2006:PSL

[CF06a] M. Cenci and F. Filippini. Portfolio selection: a linear approach with dual expected utility. Applied Mathematics and Computation, 179(2):523–534, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030501057X.

Chen:2006:SPS

[CF06b] Xiaoxing Chen and Chen Fengde. Stable periodic solution of a discrete periodic Lotka–Volterra competition system with a feedback control. Applied Mathematics and Computation, 181(2):1446–1454, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002736.

Celebi:2006:EWB

[ÇFÇ06a] E. Çelebi, S. Fırat, and İ. Çankaya. The effectiveness of wave barriers on the dynamic stiffness coefficients of foundations using boundary element method. Applied Mathematics and Computation, 180(2):683–699, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000579.

Celebi:2006:EIF

[ÇFÇ06b] E. Çelebi, S. Fırat, and İ. Çankaya. The evaluation of impedance functions in the analysis of foundations vibrations using boundary element method. Applied Mathematics and Computation, 173(1):636–667, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500384X.

Chen:2006:FSN

[CFC06c] Chun-Sheng Chen, Chin-Ping Fung, and Rean-Der Chien. A further study on nonlinear vibration of initially stressed plates. Applied Mathematics and Computation, 172(1):349–367, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002079.

Cheng:2009:DSH

[CFF09] Hao Cheng, Chu-Li Fu, and Xiao-Li Feng. Determining surface heat flux in the steady state for the Cauchy problem for the Laplace equation. Applied Mathematics and Computation, 211(2):374–382, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900085X.

Chiroux:2005:TDF

[CFJ^{+}05] R. C. Chiroux, W. A.

Foster, Jr., C. E. Johnson, S. A. Shoop, and R. L. Raper. Threedimensional finite element analysis of soil interaction with a rigid wheel. Applied Mathematics and Computation, 162(2):707–722, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001262.

Climent:2006:NEC

[CFVZ06] Joan-Josep Climent, Francisco Ferrández, José-Francisco Vicent, and Antonio Zamora. A nonlinear elliptic curve cryptosystem based on matrices. Applied Mathematics and Computation, 174(1):150–164, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305004479.

Cai:2005:CTC

[CFX05] Li Cai, Jian-Hu Feng, and Wen-Xian Xie. A CWENO-type central-upwind scheme for ideal MHD equations. Applied Mathematics and Computation, 168(1):600–612, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304006265.

Celik:2005:ASP

[ÇG05a] İbrahim Çelik and Guzin Gokmen. Approximate solution of periodic Sturm– Liouville problems with Chebyshev collocation method. Applied Mathematics and Computation, 170(1):285– 295, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304009269.

Clavero:2005:HOM

[CG05b] C. Clavero and J. L. Gracia. High order methods for elliptic and time dependent reaction–diffusion singularly perturbed problems. Applied Mathematics and Computation, 168(2):1109–1127, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006678.

Corani:2006:VDS

[CG06] Giorgio Corani and Marino Gatto. VC-dimension and structural risk minimization for the analysis of nonlinear ecological models. Applied Mathematics and Computation, 176(1):166– 176, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007824.

Cheikh:2007:CDC

[CG07a] Y. Ben Cheikh and M. Gaied. Characterization of the Dunkl-classical symmetric orthogonal polynomials. Applied Mathematics and Computation, 187(1):105– 114, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011428.

Chen:2007:BSO

[CG07b] Jianqing Chen and Boling Guo. Blow up solutions for one class of system of Pekar–Choquard type nonlinear Schrödinger equation. Applied Mathematics and Computation, 186 (1):83–92, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306009192.

Cui:2007:CMS

[CG07c] Minggen Cui and Fazhan Geng. A computational method for solving one-dimensional variablecoefficient Burgers equation. Applied Mathematics and Computation, 188(2):1389–1401, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014913.

Cheng:2008:MMT

[CG08a] Rongjun Cheng and Hongxia Ge. The meshless method for a two-dimensional parabolic problem with a source parameter. Applied Mathematics and Computation, 202(2):730–737, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001719.

Cui:2008:CMS

[CG08b] M. Cui and F. Geng. A computational method for solving third-order singularly perturbed boundaryvalue problems. Applied Mathematics and Computation, 198(2):896–903, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700968X.

Castro-Gonzalez:2006:FDI

[CGDR06] N. Castro-González, E. Dopazo, and J. Robles. Formulas for the Drazin inverse of special block matrices. Applied Mathematics and Computation, 174(1):252–270, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305004546.

Conforto:2008:SSB

[CGJ08] F. Conforto, M. Groppi, and A. Jannelli. On shock solutions to balance equations for slow and fast chemical reaction. Applied Mathematics and Computation, 206(2):892–905, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308007509.

Coll:2008:CIR

[CGS08] Narcís Coll, Marité Guerrieri, and J. Antoni Sellarès. Combining improvement and refinement techniques: 2D Delaunay mesh adaptation under domain changes. Applied Mathematics and Computation, 201(1–2):527–546, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307012283.

Cupec:2009:TPM

[CGSS09] Robert Cupec, Ratko Grbić, Kristian Sabo, and Rudolf Scitovski. Three points method for searching the best least absolute deviations plane. Applied Mathematics and Computation, 215(3):983–994, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005876.

Castro-Gonzalez:2007:WDI

[CGVC07] N. Castro-González and J. Y. Vélez-Cerrada. The weighted Drazin inverse of perturbed matrices with related support idempotents. Applied Mathematics and Computation, 187(2):756–764, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011817.

Chen:2006:DTV

[CGW06] Jing-Bo Chen, Han-Ying Guo, and Ke Wu. Discrete total variation calculus and Lee’s discrete mechanics. Applied Mathematics and Computation, 177(1):226– 234, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009045.

Chang:2005:SRP

[CH05a] Yen-Chang Chang and Wen-Liang Hung. Software release policies on a shot-noise process model. Applied Mathematics and Computation, 171(2):746– 759, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2005:NKM

[CH05b] Tzer-Shyong Chen and Jen-Yan Huang. A novel key management scheme for dynamic access control in a user hierarchy. Applied Mathematics and Computation, 162(1):339–351, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304000992.

Chen:2006:QFA

[CH06a] Chien-Yuan Chen and Chih-Cheng Hsueh. Quantum factorization algorithm by NMR ensemble computers. Applied Mathematics and Computation, 174(2):1363–1369, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005552.

Chen:2006:FMS

[CH06b] Yuh-Wen Chen and Hung-En Hsieh. Fuzzy multistage De-Novo programming problem. Applied Mathematics and Computation, 181(2):1139–1147, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001986.

Croicu:2006:EOV

[CH06c] Ana-Maria Croicu and M. Yousuff Hussaini. On the expected optimal value and the optimal expected value. Applied Mathematics and Computation, 180(1):330–341, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305011288.

Cabada:2007:ERD

[CH07a] A. Cabada and S. Heikkilä.

Existence results for discontinuous functional differential systems. Applied Mathematics and Computation, 188(2):1251–1259, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014706.

Chen:2007:ISN

[CH07b] Chun-Sheng Chen and Chun-Yao Hsu. Imperfection sensitivity in the nonlinear vibration oscillations of initially stressed plates. Applied Mathematics and Computation, 190(1):465– 475, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000781.

Chen:2007:NPP

[CH07c] Fengde Chen and Aimei Huang. On a nonautonomous predator-prey model with prey dispersal. Applied Mathematics and Computation, 184(2):809–822, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007697.

Chun:2007:OPFc

[CH07d] Changbum Chun and YoonMee Ham. A oneparameter fourth-order family of iterative methods for nonlinear equations. Applied Mathematics and Computation, 189(1):610– 614, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016353.

Chun:2007:SSO

[CH07e] Changbum Chun and YoonMee Ham. Some sixth-order variants of Ostrowski root-finding methods. Applied Mathematics and Computation, 193(2):389–394, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004122.

Chung:2007:SCR

[CH07f] Kuo-Liang Chung and Yong-Huai Huang. Speed up the computation of randomized algorithms for detecting lines, circles, and ellipses using novel tuning- and LUT-based voting platform. Applied Mathematics and Computation, 190(1):132–149, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000367.

Chun:2008:SFOb

[CH08a] Changbum Chun and YoonMee Ham. Some fourth-order modifications of Newton’s method. Applied Mathematics and Computation, 197(2):654– 658, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008223.

Chun:2008:SSD

[CH08b] Changbum Chun and YoonMee Ham. Some second-derivative-free variants of super-Halley method with fourth-order convergence. Applied Mathematics and Computation, 195(2):537–541, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005838.

Chen:2009:OCC

[CH09] Jun-Jie Chen and Ze-Rong He. Optimal control for a class of nonlinear age-distributed population systems. Applied Mathematics and Computation, 214(2):574–580, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309003567.

Chang:2005:AAF

[Cha05a]  Ching-Ter Chang.  An

approximate approach for fractional programming with absolute-value functions. Applied Mathematics and Computation, 161(1):171–179, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012189.

Chang:2005:FPA

[Cha05b] Ching-Ter Chang. Fractional programming with absolute-value functions: a fuzzy goal programming approach. Applied Mathematics and Computation, 167(1):508–515, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005077.

Chang:2005:MGP

[Cha05c] Ching-Ter Chang. A modified goal programming approach for the meanabsolute deviation portfolio optimization model. Applied Mathematics and Computation, 171(1):567– 572, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305001165.

Chang:2005:MIS

[Cha05d]  Ching-Ter Chang.  On

the mixed integer signomial programming problems. Applied Mathematics and Computation, 170(2):1436–1451, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000755.

Chang:2005:FSM

[Cha05e] Feng Cheng Chang. Function of a square matrix with repeated eigenvalues. Applied Mathematics and Computation, 160(3):819–849, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303011974.

Chang:2005:MFV

[Cha05f] Feng Cheng Chang. Matrix formulation of vector operations. Applied Mathematics and Computation, 170(2):1135–1165, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000548.

Chant:2005:HVP

[Cha05g] Lawrence J. De Chant. A high velocity plate penetration hole diameter relationship based on late time stagnation point flow concepts. Applied Mathematics and Computation, 170(1):410–424, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009580.

Chant:2005:VPL

[Cha05h] Lawrence J. De Chant.

The venerable 1/7th power law turbulent velocity profile: a classical nonlinear boundary value problem solution and its relationship to stochastic processes. Applied Mathematics and Computation, 161(2):463–474, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000050.

Chang:2006:RMG

[Cha06] Ching-Ter Chang. Reformulation of the modified goal programming for logarithmic piecewise linear function. Applied Mathematics and Computation, 174(1):13–23, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004388.

Chakraborty:2007:CRB

[Cha07a] Soubhik Chakraborty. Can a randomized binary search have an O(1) complexity at least in practice? Applied Mathematics and Computation, 189(2):1884–1888, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017838.

Chanane:2007:CES

[Cha07b] B. Chanane. Computing the eigenvalues of singular Sturm–Liouville problems using the regularized sampling method. Applied Mathematics and Computation, 184(2):972–978, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007879.

Chanane:2007:ESL

[Cha07c] B. Chanane. Eigenvalues of Sturm–Liouville problems with discontinuity conditions inside a finite interval. Applied Mathematics and Computation, 188(2):1725–1732, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015268.

Chanane:2007:SLP

[Cha07d]  B. Chanane.  Sturm–

Liouville problems with impulse effects. Applied Mathematics and Computation, 190(1):610–626, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000951.

Chang:2007:CSS

[Cha07e] Feng Cheng Chang. Complete sets of solutions to partition–distribution equations. Applied Mathematics and Computation, 190(2):1094–1107, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001294.

Chakravarthy:2009:DQM

[Cha09] Srinivas R. Chakravarthy.

A disaster queue with Markovian arrivals and impatient customers. Applied Mathematics and Computation, 214(1):48–59, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900277X.

Chen:2005:EUG

[CHC05] Tzer-Shyong Chen, Jen-Yan Huang, and Tzer-Long Chen. An efficient undeniable group-oriented signature scheme. Applied Mathematics and Computation, 165(1):95–102, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003649.

Chen:2005:GAS

[Che05a]  Fengde Chen.  Global

asymptotic stability in nspecies non-autonomous Lotka–Volterra competitive systems with infinite delays and feedback control. Applied Mathematics and Computation, 170(2):1452–1468, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000767.

Chen:2005:PMS

[Che05b] Fengde Chen. On a periodic multi-species ecological model. Applied Mathematics and Computation, 171(1):492–510, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305001116.

Chen:2005:PSA

[Che05c] Fengde Chen. Periodic solutions and almost periodic solutions for a delay multispecies Logarithmic population model. Applied Mathematics and Computation, 171(2):760–770, December 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2005:PPS

[Che05d] Fengde Chen. Positive periodic solutions of neutral Lotka–Volterra system with feedback control. Applied Mathematics and Computation, 162(3):1279–1302, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001882.

Chen:2005:MGL

[Che05e] Jing-Bo Chen. Multisymplectic geometry, local conservation laws and Fourier pseudospectral discretization for the “good” Boussinesq equation. Applied Mathematics and Computation, 161(1):55–67, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012098.

Chen:2005:MIP

[Che05f] Jing-Bo Chen. A multisymplectic integrator for the periodic nonlinear Schrödinger equation. Applied Mathematics and Computation, 170(2):1394– 1417, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305000731.

Chen:2005:TSS

[Che05g]  Tzer-Shyong Chen.  A

threshold signature scheme based on the elliptic curve cryptosystem. Applied Mathematics and Computation, 162(3):1119–1134, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001535.

Chen:2005:ELE

[Che05h]  Wei-Kuei Chen.  Effi-

cient on-line electronic checks. Applied Mathematics and Computation, 162(3):1259–1263, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001857.

Chernousko:2005:MSL

[Che05i] F. L. Chernousko. Modelling of snake-like locomotion. Applied Mathematics and Computation, 164(2):415–434, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003273.

Chen:2006:APS

[Che06a] Fengde Chen. Almost periodic solution of the nonautonomous two-species competitive model with stage structure. Applied Mathematics and Computation, 181(1):685–693, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001482.

Chen:2006:GSS

[Che06b] Fengde Chen. Global stability of a single species model with feedback control and distributed time delay. Applied Mathematics and Computation, 178(2):474–479, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009938.

Chen:2006:PSP

[Che06c] Fengde Chen. On the periodic solutions of periodic multi-species Kolmogorov type competitive system with delays and feedback controls. Applied Mathematics and Computation, 180(1):366–373, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305011318.

Chen:2006:PSA

[Che06d] Fengde Chen. Periodic solutions and almost periodic solutions of a neutral multispecies Logarithmic population model. Applied Mathematics and Computation, 176(2):431–441, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008192.

Chen:2006:PGAb

[Che06e]  Fengde Chen.  Perma-

nence and global attractivity of a discrete multispecies Lotka–Volterra competition predator–prey systems. Applied Mathematics and Computation, 182(1):3–12, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003055.

Chen:2006:PGS

[Che06f]  Fengde Chen.  Perma-

nence and global stability of nonautonomous Lotka– Volterra system with predator– prey and deviating arguments. Applied Mathematics and Computation, 173(2):1082–1100, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004133.

Chen:2006:PNM

[Che06g] Fengde Chen. Permanence in nonautonomous multispecies predator–prey system with feedback controls. Applied Mathematics and Computation, 173(2):694–709, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003826.

Chen:2006:PDN

[Che06h]  Fengde Chen.  Perma-

nence of a delayed nonautonomous Gilpin–Ayala competition model. Applied Mathematics and Computation, 179(1):55– 66, August 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305010064.

Chen:2006:PPH

[Che06i]  Fengde Chen.  Perma-

nence of periodic Holling type predator–prey system with stage structure for prey. Applied Mathematics and Computation, 182(2):1849–1860, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005340.

Chen:2006:SNI

[Che06j]  Jinhai Chen.  Some

new iterative methods with three-order convergence. Applied Mathematics and Computation, 181(2):1519–1522, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002839.

Chen:2006:MPA

[Che06k] Shih-Pin Chen. A mathematical programming approach to the machine interference problem with fuzzy parameters. Applied Mathematics and Computation, 174(1):374–387, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004650.

Chen:2006:IBA

[Che06l] Ta-Cheng Chen. IAs based approach for reliability redundancy allocation problems. Applied Mathematics and Computation, 182(2):1556–1567, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005042.

Chen:2006:PFT

[Che06m] Z. W. Chen. A penaltyfree-type nonmonotone trust-region method for nonlinear constrained optimization. Applied Mathematics and Computation, 173(2):1014–1046, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004091.

Cheng:2006:NDN

[Che06n] Ching-Yang Cheng. Non-Darcy natural convection heat and mass transfer from a vertical wavy surface in saturated porous media. Applied Mathematics and Computation, 182(2):1488–1500, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004760.

Chen:2007:PEL

[Che07a]  Fen-Ying Chen.  Pric-

ing the equity-linked and principal-protected securities with cap and path dependence. Applied Mathematics and Computation, 194(2):502–513, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005024.

Chen:2007:PDSb

[Che07b]  Fengde Chen.  Perma-

nence of a discrete Nspecies cooperation system with time delays and feedback controls. Applied Mathematics and Computation, 186(1):23– 29, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600912X.

Chen:2007:PDSa

[Che07c]  Fengde Chen.  Perma-

nence of a discrete nspecies food-chain system with time delays. Applied Mathematics and Computation, 185(1):719–726, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009076.

Chen:2007:GSC

[Che07d] Heng-Hui Chen. Global synchronization of chaotic systems via linear balanced feedback control. Applied Mathematics and Computation, 186(1):923–931, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010150.

Chen:2007:MPM

[Che07e] Jing-Bo Chen. A multisymplectic pseudospectral method for seismic modeling. Applied Mathematics and Computation, 186(2):1612–1616, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010976.

Chen:2007:NMR

[Che07f] Jinhai Chen. New modified regula falsi method for nonlinear equations. Applied Mathematics and Computation, 184(2):965–971, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007867.

Chen:2007:PSA

[Che07g] Jinhai Chen. Pre-symmetric approach and applications to indefinite nonsymmetric problems. Applied Mathematics and Computation, 192(2):457– 465, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307003499.

Chen:2007:CIS

[Che07h] Wen-Yuan Chen. Color image steganography scheme using set partitioning in hierarchical trees coding, digital Fourier transform and adaptive phase modulation. Applied Mathematics and Computation, 185(1):432–448, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008782.

Cheng:2007:NSD

[Che07i] Ching-Yang Cheng. Nonsimilar solutions for double diffusive convection near a frustum of a wavy cone in porous media. Applied Mathematics and Computation, 194(1):156–167, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004547.

Chen:2008:VIF

[Che08a] Jing-Bo Chen. Variational integrators and the finite element method. Applied Mathematics and Computation, 196(2):941–958, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007473.

Chen:2008:CIS

[Che08b] Wen-Yuan Chen. Color image steganography scheme using DFT, SPIHT codec, and modified differential phase-shift keying techniques. Applied Mathematics and Computation, 196(1):40–54, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700625X.

Chen:2008:MWS

[Che08c] Wen-Yuan Chen. Multiplewatermarking scheme of the European Article Number Barcode using similar code division multiple access technique. Applied Mathematics and Computation, 197(1):243–261, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307007734.

Chen:2008:PNP

[Che08d] Xiaoxing Chen. Periodicity in a nonlinear predator– prey system on time scales with state-dependent delays. Applied Mathematics and Computation, 196(1):118–128, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006509.

Chen:2009:LEB

[Che09a] Sheng Chen. A large-eddy-based lattice Boltzmann model for turbulent flow simulation. Applied Mathematics and Computation, 215(2):591–598, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309005335.

Cheng:2009:NBL

[Che09b] Ching-Yang Cheng. Nonsimilar boundary layer analysis of double-diffusive convection from a vertical truncated cone in a porous medium with variable viscosity. Applied Mathematics and Computation, 212(1):185–193, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001337.

Chen:2009:LAS

[CHH09] Yang-Yih Chen, Hung-Chu Hsu, and Hwung-Hweng Hwung. A Lagrangian asymptotic solution for finite-amplitude standing waves. Applied Mathematics and Computation, 215(8):2891–2900, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008443.

Chien:2005:CIB

[Chi05] Hung-Yu Chien. Comments on ID-based multisignature with distinguished signing authorities. Applied Mathematics and Computation, 170(2):1284–1289, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000639.

Chiou:2006:OST

[Chi06a] Suh-Wen Chiou. An optimal scheme for toll pricing problem. Applied Mathematics and Computation, 182(2):1127–1136, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004279.

Chiter:2006:DAN

[Chi06b] Lakhdar Chiter. DIRECT algorithm: a new definition of potentially optimal hyperrectangles. Applied Mathematics and Computation, 179(2):742–749, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010829.

Chiter:2006:NSM

[Chi06c] Lakhdar Chiter. A new sampling method in the DI-RECT algorithm. Applied Mathematics and Computation, 175(1):297–306, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006089.

Chiou:2007:CAA

[Chi07a]  Suh-Wen Chiou.  A

combinatorial approximation algorithm for supply chain network flow problem. Applied Mathematics and Computation, 186(2):1526–1536, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010794.

Chiou:2007:DFL

[Chi07b] Suh-Wen Chiou. A descent framework for linked signal system with network flows. Applied Mathematics and Computation, 194(2):441–452, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004936.

Chiou:2007:ECA

[Chi07c] Suh-Wen Chiou. An efficient computation algorithm for area traffic control problem with link capacity expansions. Applied Mathematics and Computation, 188(2):1094–1102, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013208.

Chiou:2007:GIS

[Chi07d] Suh-Wen Chiou. A generalized iterative scheme for network design problem. Applied Mathematics and Computation, 188(2):1115–1123, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306013798.

Chiou:2007:GCI

[Chi07e]  Suh-Wen Chiou.  A

globally convergent iterative scheme for toll design network with signal settings. Applied Mathematics and Computation, 187(2):1086–1099, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012458.

Chiou:2007:ISS

[Chi07f] Suh-Wen Chiou. An iterative scheme for signal settings and network flows. Applied Mathematics and Computation, 189(2):1808–1815, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017760.

Chiou:2007:ODS

[Chi07g]  Suh-Wen Chiou.  Op-

timal design of signalcontrolled road network. Applied Mathematics and Computation, 189(1):1–8, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015190.

Chiou:2007:OMA

[Chi07h] Suh-Wen Chiou. An optimization model for area traffic control with link tolls. Applied Mathematics and Computation, 192(2):520–532, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003554.

Chiou:2007:RCS

[Chi07i]  Suh-Wen Chiou.  Re-

serve capacity of signalcontrolled road network. Applied Mathematics and Computation, 190(2):1602– 1611, July 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307002056.

Chiou:2008:ESA

[Chi08a] Suh-Wen Chiou. An efficient search algorithm for road network optimization. Applied Mathematics and Computation, 201(1–2):128–137, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011770.

Chiou:2008:FPT

[Chi08b]  Suh-Wen Chiou.  A

fast polynomial time algorithm for logistics network flows. Applied Mathematics and Computation, 199(1):162–170, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009836.

Chiou:2008:OCP

[Chi08c] Suh-Wen Chiou. Optimization of congestion pricing road network with variable demands. Applied Mathematics and Computation, 195(2):382–391, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005577.

Chen:2006:PIG

[CHL06] Yu-Sheng Chen, Gwoboa Horng, and Chao-Liang Liu. Performance improvement for the GGMconstruction of pseudorandom functions. Applied Mathematics and Computation, 183(1):486–490, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005820.

Cezaro:2008:SDK

[CHLS08] A. De Cezaro, M. Haltmeier, A. Leitão, and O. Scherzer. On Steepest-Descent-Kaczmarz methods for regularizing systems of nonlinear ill-posed equations. Applied Mathematics and Computation, 202(2):596–607, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001495.

Cho:2005:LBM

[Cho05]  Gyeong-Mi Cho.  Log-

barrier method for twostage quadratic stochastic programming. Applied Mathematics and Computation, 164(1):45– 69, May 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002954.

Choi:2007:SMC

[Cho07a]  Junesang Choi.  Some

mathematical constants. Applied Mathematics and Computation, 187(1):122– 140, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030601143X.

Chou:2007:FMM

[Cho07b] Chien-Chang Chou. A fuzzy MCDM method for solving marine transshipment container port selection problems. Applied Mathematics and Computation, 186(1):435–444, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009520.

Choudhury:2007:TPB

[Cho07c] Gautam Choudhury. A two phase batch arrival retrial queueing system with Bernoulli vacation schedule. Applied Mathematics and Computation, 188(2):1455–1466, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014986.

Christodoulou:2008:DHB

[Chr08] Nikolaos S. Christodoulou.

Discrete Hopf bifurcation for Runge–Kutta methods. Applied Mathematics and Computation, 206(1):346–356, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800670X.

Chen:2006:RCS

[CHT06] Y-Chuang Chen, Lih-Hsing Hsu, and Jimmy J. M. Tan. A recursively construction scheme for super fault-tolerant Hamiltonian graphs. Applied Mathematics and Computation, 177(2):465–481, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009252.

Chukwu:2005:CCM

[Chu05] E. N. Chukwu. Cooperation and competition in modeling the dynamics of gross-domestic products of nations. Applied Mathematics and Computation, 163(2):991–1021, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001304.

Chun:2006:NIM

[Chu06]  Changbum Chun.  A

new iterative method for solving nonlinear equations. Applied Mathematics and Computation, 178(2):415–422, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009860.

Chun:2007:CIC

[Chu07a] Changbum Chun. Certain improvements of Chebyshev– Halley methods with accelerated fourth-order convergence. Applied Mathematics and Computation, 189 (1):597–601, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030601633X.

Chun:2007:CTO

[Chu07b] Changbum Chun. Construction of third-order modifications of Newton’s method. Applied Mathematics and Computation, 189(1):662–668, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016419.

Chun:2007:FCF

[Chu07c] Changbum Chun. A family of composite fourthorder iterative methods for solving nonlinear equations. Applied Mathematics and Computation, 187(2):951–956, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601229X.

Chun:2007:NFO

[Chu07d]  Changbum Chun.  A

note on fifth-order iterative methods for solving nonlinear equations. Applied Mathematics and Computation, 189(2):1805–1807, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017759.

Chun:2007:NSI

[Chu07e] Changbum Chun. A note on some iterative schemes for nonlinear equations. Applied Mathematics and Computation, 188(1):271– 274, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030601349X.

Chun:2007:CIM

[Chu07f] Changbum Chun. On the construction of iterative methods with at least cubic convergence. Applied Mathematics and Computation, 189(2):1384–1392, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017243.

Chun:2007:OPFa

[Chu07g]  Changbum Chun.  A

one-parameter family of quadratically convergent iteration formulae. Applied Mathematics and Computation, 189(1):55– 58, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015785.

Chun:2007:OPFb

[Chu07h]  Changbum Chun.  A

one-parameter family of third-order methods to solve nonlinear equations. Applied Mathematics and Computation, 189(1):126– 130, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015876.

Chun:2007:SIJ

[Chu07i] Changbum Chun. Some improvements of Jarratt’s method with sixth-order convergence. Applied Mathematics and Computation, 190(2):1432–1437, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001750.

Chun:2007:SSD

[Chu07j] Changbum Chun. Some second-derivative-free variants of Chebyshev–Halley methods. Applied Mathematics and Computation, 191(2):410–414, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002688.

Chun:2007:STO

[Chu07k] Changbum Chun. Some third-order families of iterative methods for solving nonlinear equations. Applied Mathematics and Computation, 188(1):924– 933, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014305.

Chun:2007:SVC

[Chu07l] Changbum Chun. Some variants of Chebyshev– Halley methods free from second derivative. Applied Mathematics and Computation, 191(1):193–198, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002457.

Chun:2007:SVK

[Chu07m] Changbum Chun. Some variants of King’s fourthorder family of methods for nonlinear equations. Applied Mathematics and Computation, 190(1):57– 62, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700029X.

Chun:2007:TPT

[Chu07n]  Changbum Chun.  A

two-parameter third-order family of methods for solving nonlinear equations. Applied Mathematics and Computation, 189(2):1822–1827, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017784.

Chun:2008:SFOa

[Chu08] Changbum Chun. Some fourth-order iterative methods for solving nonlinear equations. Applied Mathematics and Computation, 195(2):454–459, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700567X.

Chen:2007:NPV

[CHW07] Cheng-Kang Chen, Ta-Wei Hung, and Tzu-Chun Weng. A net present value approach in developing optimal replenishment policies for a product life cycle. Applied Mathematics and Computation, 184(2):360–373, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007132.

Chang:2005:ILW

[CHY05a] Ting-Yi Chang, Min-Shiang Hwang, and Wei-Pang Yang. An improvement on the Lin–Wu (t, n) threshold verifiable multisecret sharing scheme. Applied Mathematics and Computation, 163(1):169– 178, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002188.

Chen:2005:GSS

[CHY05b] Lin Chen, Xiaoqin Huang, and Jinyuan You. Group signature schemes with forward secure properties. Applied Mathematics and Computation, 170(2):841–849, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000275.

Chung:2009:CMR

[CHY^{+}09] Kuo-Liang Chung, Yong-Huai Huang, Wei-Ning Yang, Yu-Chiao Hsu, and Chyou-Hwa Chen. Capacity maximization for reversible data hiding based on dynamic programming approach. Applied Mathematics and Computation, 208(1):284–292, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008217.

Chow:2005:FSM

[CHYC05] Sherman S. M. Chow, Lucas C. K. Hui, S. M.

Yiu, and K. P. Chow. Forward-secure multisignature and blind signature schemes. Applied Mathematics and Computation, 168(2):895–908, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006472.

Cvetkovic-Ilic:2005:PTW

[CI05] Dragana S. Cvetković-Ilić.

On a problem of N. Thome and Y. Wei. Applied Mathematics and Computation, 166(1):233–236, July 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002602.

Cvetkovic-Ilic:2009:EDM

[CI09] Dragana S. Cvetković-Ilić.

Expression of the Drazin and MP-inverse of partitioned matrix and quotient identity of generalized Schur complement. Applied Mathematics and Computation, 213(1):18– 24, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002136.

Christou:2009:SRD

[CIS09] M. A. Christou, N. M. Ivanova, and C. Sophocleous. Similarity reductions of the (1 + 3)- dimensional Burgers equation. Applied Mathematics and Computation, 210 (1):87–99, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300308008850.

Cheng:2006:FNM

[CJ06] Yongkuan Cheng and Jifa Jiang. The Field–Noyes model with periodic orbits. Applied Mathematics and Computation, 182(1):36–40, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003109.

Chen:2008:GIP

[CJ08] Fang Chen and Yao-Lin Jiang. A generalization of the inexact parameterized Uzawa methods for saddle point problems. Applied Mathematics and Computation, 206(2):765–771, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007054.

Chen:2005:NPD

[CJC05] Yu-Yi Chen, Jinn-Ke Jan, and Chin-Ling Chen. A novel proxy deposit protocol for e-cash systems. Applied Mathematics and Computation, 163(2):869–877, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002668.

Chen:2008:ESSb

[CjC08] Zhong Chen and Zhi jie Chen. The exact solution of system of linear operator equations in reproducing kernel spaces. Applied Mathematics and Computation, 203(1):56– 61, September 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308001938.

Choi:2007:AOP

[CJK07] U Jin Choi, Bong-Gyu Jang, and Hyeng-Keun Koo. An algorithm for optimal portfolio selection problem with transaction costs and random lifetimes. Applied Mathematics and Computation, 191(1):239–252, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002512.

Choie:2005:SHC

[CJL05a] Young Ju Choie, Eun Kyung Jeong, and Eun Jeong Lee. Supersingular hyperelliptic curves of genus 2 over finite fields. Applied Mathematics and Computation, 163(2):565–576, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002437.

Choie:2005:EIB

[CJL05b] Young Ju Choie, Eunkyung Jeong, and Eunjeong Lee. Efficient identity-based authenticated key agreement protocol from pairings. Applied Mathematics and Computation, 162(1):179– 188, March 4, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304000785.

Conte:2008:TSA

[CJP08] D. Conte, Z. Jackiewicz, and B. Paternoster. Twostep almost collocation methods for Volterra integral equations. Applied Mathematics and Computation, 204(2):839–853, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005699.

Cortes:2005:ENS

[CJSSP05] J. C. Cortés, L. Jódar, R. Sala, and P. Sevilla-Peris. Exact and numerical solution of Black–Scholes matrix equation. Applied Mathematics and Computation, 160(3):607–613, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303011779.

Chan:2005:FPA

[CK05] Din-Yuen Chan and Cheng-Yuan Ku. Finite-precision arithmetic coding via index-splitting and probabilityfitting. Applied Mathematics and Computation, 167(2):976–995, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005430.

Cho:2006:NLU

[CK06a] Gyeong-Mi Cho and Min-Kyung Kim. A new largeupdate interior point algorithm for P_{∗}(κ) LCPs based on kernel functions. Applied Mathematics and Computation, 182(2):1169–1183, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004310.

Chu:2006:CIM

[CK06b] Yunn-Kuang Chu and Jau-Chuan Ke. Confidence intervals of mean response time for an M / G / 1 queueing system: Bootstrap simulation. Applied Mathematics and Computation, 180(1):255–263, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305011203.

Chuang:2006:NAD

[CK06c] Tzung-Nan Chuang and Jung-Yuan Kung. A new algorithm for the discrete fuzzy shortest path problem in a network. Applied Mathematics and Computation, 174(1):660–668, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004923.

Constales:2006:CSA

[CK06d] D. Constales and J. Kačur.

Computation and sensitivity analysis of the pricing of American call options. Applied Mathematics and Computation, 176(1):302– 307, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007940.

Cho:2007:IPC

[CK07a] Nak Eun Cho and In Hwa Kim. Inclusion properties of certain classes of meromorphic functions associated with the generalized hypergeometric function. Applied Mathematics and Computation, 187(1):115– 121, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011441.

Chu:2007:MRT

[CK07b] Yunn-Kuang Chu and Jau-Chuan Ke. Mean response time for a G / G / 1 queueing system: Simulated computation. Applied Mathematics and Computation, 186(1):772–779, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009970.

Cvetkovic:2007:NCA

[CK07c] Ljiljana Cvetković and Vladimir Kostić. A note on the convergence of the AOR method. Applied Mathematics and Computation, 194(2):394–399, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004808.

Chaharsooghi:2008:EAC

[CK08] S. K. Chaharsooghi and Amir H. Meimand Kermani. An effective ant colony optimization algorithm (ACO) for multiobjective resource allocation problem (MORAP). Applied Mathematics and Computation, 200(1):167– 177, June 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307010971.

Cho:2005:ELV

[CKK05] Chung-Ki Cho, Taekkeun Kim, and YongHoon Kwon. Estimation of local volatilities in a generalized Black– Scholes model. Applied Mathematics and Computation, 162(3):1135–1149, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001742.

Cvetkovic:2008:FRM

[CKKS08] Ljiljana Cvetković, Vladimir Kostić, Maja Kovačević, and Tomasz Szulc. Further results on H-matrices and their Schur complements. Applied Mathematics and Computation, 198(2):506– 510, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009162.

Chan:2006:MII

[CKL06] Din-Yuen Chan, Cheng-Yuan Ku, and Ming-Chai Li. A method to improve integer linear programming problem with branch-and-bound procedure. Applied Mathematics and Computation, 179(2):484–493, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010544.

Cho:2007:CIO

[CKOS07] Nak Eun Cho, Oh Sang Kwon, Shigeyoshi Owa, and H. M. Srivastava. A class of integral operators preserving subordination and superordination for meromorphic functions. Applied Mathematics and Computation, 193(2):463–474, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004298.

Cho:2008:SGN

[CKQ08] Yeol Je Cho, Shin Min Kang, and Xiaolong Qin. On systems of generalized nonlinear variational inequalities in Banach spaces. Applied Mathematics and Computation, 206(1):214–220, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006668.

Cvetkovic:2009:NSM

[CKR09] Ljiljana Cvetković, Vladimir Kostić, and Sonja Rauški. A new subclass of Hmatrices. Applied Mathematics and Computation, 208(1):206–210, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008904.

Chiang:2009:ATS

[CKS09] Ko-Ta Chiang, Kuang Yuan Kung, and H. M. Srivastava. Analytic transient solutions of a cylindrical heat equation with a heat source. Applied Mathematics and Computation, 215(8):2877–2885, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008285.

Chen:2005:CLR

[CKY05] Chien-Yuan Chen, Cheng-Yuan Ku, and David C. Yen. Cryptanalysis of large RSA exponent by using the LLL algorithm. Applied Mathematics and Computation, 169(1):516–525, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400791X.

Chang:2005:FMS

[CL05a] Chin-Chen Chang and Yeu-Pong Lai. A fast modular square computing method based on the generalized Chinese remainder theorem for prime moduli. Applied Mathematics and Computation, 161(1):181–194, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303012190.

Chang:2005:ECM

[CL05b] Shui-Ching Chang and Tze Fen Li. Estimation of component mean lifetimes of a masked system using unclassified system life data. Applied Mathematics and Computation, 169(2):797–805, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008112.

Chen:2005:CGN

[CL05c] Jinhai Chen and Weiguo Li. Convergence of Gauss– Newton’s method and uniqueness of the solution. Applied Mathematics and Computation, 170(1):686–705, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030500010X.

Chen:2005:SNT

[CL05d] Mei-Qin Chen and Xiezhang Li. The superiority of a new type (2, 2)- step iterative method over the related Chebyshev method. Applied Mathematics and Computation, 162(2):605–625, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001183.

Chang:2006:DAR

[CL06a] Chin-Chen Chang and Yeu-Pong Lai. A division algorithm for residue numbers. Applied Mathematics and Computation, 172(1):368–378, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002080.

Chang:2006:EBD

[CL06b] Shui-Ching Chang and Tze Fen Li. Empirical Bayes decision rule for classification on defective items in Weibull distribution. Applied Mathematics and Computation, 182(1):425–433, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003493.

Chen:2006:NEQ

[CL06c] Jinhai Chen and Weiguo Li. On new exponential quadratically convergent iterative formulae. Applied Mathematics and Computation, 180(1):242–246, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305011185.

Cui:2006:PAB

[CL06d] Xining Cui and Quanyi Lü. A parallel algorithm for block-tridiagonal linear systems. Applied Mathematics and Computation, 173(2):1107–1114, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305004157.

Cha:2007:ABZ

[CL07a] Jaesang Cha and Chonghyun Lee. Application of binary zero-correlation-duration sequences to interferencecancelled WPAN system. Applied Mathematics and Computation, 184(1):30–36, January 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004826.

Chaggara:2007:LCB

[CL07b] H. Chaggara and I. Lamiri.

Linearization coefficients for Boas–Buck polynomial sets. Applied Mathematics and Computation, 189(2):1533–1549, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017401.

Chen:2007:APF

[CL07c] Kuan-Wen Chen and Yuh-Dauh Lyuu. Accurate pricing formulas for Asian options. Applied Mathematics and Computation, 188(2):1711–1724, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015256.

Cai:2008:NGS

[CL08a] Liming Cai and Xuezhi Li. A note on global stability of an SEI epidemic model with acute and chronic stages. Applied Mathematics and Computation, 196(2):923–930, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700745X.

Chen:2008:DLD

[CL08b] Jing-Bo Chen and Hong Liu. Derivation of Lagrangian density for the “good” Boussinesq equation and multisymplectic disretizations. Applied Mathematics and Computation, 204(1):58–62, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004542.

Chen:2008:MEO

[CL08c] Jing-Bo Chen and Hong Liu. Modal expansion of one-way operators based on spectral differentiation matrix. Applied Mathematics and Computation, 206(1):193–197, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006541.

Chen:2008:NSS

[CL08d] Shuming Chen and Chenshun Li. Nonoscillatory solutions of second order nonlinear difference equations. Applied Mathematics and Computation, 205(1):478–481, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006139.

Chen:2008:HAM

[CL08e] Y. M. Chen and J. K. Liu.

Homotopy analysis method for limit cycle flutter of airfoils. Applied Mathematics and Computation, 203(2):854–863, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004141.

Campbell:2009:SCD

[CL09a] Stephen L. Campbell and Vu Hoang Linh. Stability criteria for differentialalgebraic equations with multiple delays and their numerical solutions. Applied Mathematics and Computation, 208(2):397– 415, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308009089.

Chen:2009:EET

[CL09b] Yanping Chen and Li Li.

L^{p} error estimates of two-grid schemes of expanded mixed finite element methods. Applied Mathematics and Computation, 209(2):197–205, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800934X.

Chen:2008:NAR

[CLA08] Chang-Ming Chen, F. Liu, and V. Anh. Numerical analysis of the Rayleigh– Stokes problem for a heated generalized second grade fluid with fractional derivatives. Applied Mathematics and Computation, 204(1):340–351, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005122.

Chang:2009:ASV

[CLCK09] S. S. Chang, H. W. J. Lee, Chi Kin Chan, and J. K. Kim. Approximating solutions of variational inequalities for asymptotically nonexpansive mappings. Applied Mathematics and Computation, 212(1):51– 59, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001167.

Chen:2009:TWS

[CLDH09] Aiyong Chen, Jibin Li, Xijun Deng, and Wentao Huang. Travelling wave solutions of the Fornberg–Whitham equation. Applied Mathematics and Computation, 215(8):3068–3075, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008868.

Cai:2009:GSS

[CLG09] Li-Ming Cai, Xue-Zhi Li, and Mini Ghosh. Global stability of a stagestructured epidemic model with a nonlinear incidence. Applied Mathematics and Computation, 214(1):73–82, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900280X.

Chen:2005:RSS

[CLH05] Tzung-Her Chen, Wei-Bin Lee, and Gwoboa Horng. Remarks on some signature schemes based on factoring and discrete logarithms. Applied Mathematics and Computation, 169(2):1070–1075, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008379.

Chen:2006:DBS

[CLH06] Fengde Chen, Xinyuan Liao, and Zhenkun Huang. The dynamic behavior of N-species cooperation system with continuous time delays and feedback controls. Applied Mathematics and Computation, 181(2):803–815, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001603.

Colaciti:2007:NSR

[CLNCG07] Alysson Kennerly Colaciti, Luis Miguel Valdés López, Hélio Aparecido Navarro, and Luben Cabezas-Gómez. Numerical simulation of a radial diffuser turbulent airflow. Applied Mathematics and Computation, 189(2):1491–1504, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017371.

Chang:2005:DTS

[CLS05a] Henry Ker-Chang Chang, Erl-Huei Lu, and Pin-Chang Su. Design of traceable security system. Applied Mathematics and Computation, 168(2):933–944, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006514.

Chang:2005:FSB

[CLS05b] Henry Ker-Chang Chang, Erl-Huei Lu, and Pin-Chang Su. Fail-stop blind signature scheme design based on pairings. Applied Mathematics and Computation, 169(2):1324–1331, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008549.

Chan:2006:ESR

[CLS06] R. P. K. Chan, Hongyu Liu, and Geng Sun. Efficient symplectic Runge– Kutta methods. Applied Mathematics and Computation, 172(2):908–924, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002468.

Chen:2006:CGS

[CLSZ06] Sheng Chen, Zhaohui Liu, Baochang Shi, and Chuguang Zheng. Computation of gas–solid flows by finite difference Boltzmann equation. Applied Mathematics and Computation, 173(1):33–49, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003000.

Chen:2009:CPA

[CLW09a] Guohua Chen, Xiaolian Liao, and Shouyang Wang. A cutting plane algorithm for MV portfolio selection model. Applied Mathematics and Computation, 215(4):1456–1462, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006146.

Chen:2009:SAS

[CLW09b] Y. Z. Chen, X. Y. Lin, and Z. X. Wang. A semi-analytic solution for multiple curved cracks emanating from circular hole using singular integral equation. Applied Mathematics and Computation, 213(2):389–404, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002410.

Choi:2009:IDD

[CLWP09] S. J. Choi, S. M. Lee, S. C. Won, and Ju H. Park. Improved delaydependent stability criteria for uncertain Lur’e systems with sector and slope restricted nonlinearities and time-varying delays. Applied Mathematics and Computation, 208(2):520–530, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009259.

Chen:2005:MML

[CLZ05] William Y. C. Chen, Xueliang Li, and Jie Zheng. Matrix method for linear sequential dynamical systems on digraphs. Applied Mathematics and Computation, 160(1):197–212, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303011317.

Chen:2007:NCL

[CLZ^{+}07] Sheng Chen, Zhaohui Liu, Chao Zhang, Zhu He, Zhiwei Tian, Baochang Shi, and Chuguang Zheng. A novel coupled lattice Boltzmann model for low Mach number combustion simulation. Applied Mathematics and Computation, 193(1):266–284, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003967.

Chen:2008:NMH

[CM08a] Bilian Chen and Changfeng Ma. A note on modified Householder iterative method free from second derivatives for nonlinear equations. Applied Mathematics and Computation, 203(2):913–915, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002178.

Chen:2008:ASS

[CM08b] Lan Chen and He-Ping Ma. Approximate solution of the Sturm–Liouville problems with Legendre– Galerkin–Chebyshev collocation method. Applied Mathematics and Computation, 206(2):748–754, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007030.

Chakrabarti:2009:ASF

[CM09] A. Chakrabarti and S. C.

Martha. Approximate solutions of Fredholm integral equations of the second kind. Applied Mathematics and Computation, 211(2):459–466, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001039.

Chan:2006:BVM

[CMC06] Raymond H. Chan, Ka-Chun Ma, and Wai-Ki Ching. Boundary value methods for solving transient solutions of Markovian queueing networks. Applied Mathematics and Computation, 172(2):690–700, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009038.

Cardenas-Morales:2006:SPA

[CMGMD06] D. Cárdenas-Morales, P. Garrancho, and F. J. Muñoz-Delgado. Shape preserving approximation by Bernstein-type operators which fix polynomials. Applied Mathematics and Computation, 182(2):1615– 1622, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030600508X.

Casciola:2007:RPA

[CMM07] G. Casciola, L. B. Montefusco, and S. Morigi. The regularizing properties of anisotropic radial basis functions. Applied Mathematics and Computation, 190(2):1050–1062, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016432.

Chavez:2009:GOC

[CMMC09] Irma Y. Sánchez Chávez, Rubén Morales-Menéndez, and Sergio O. Martínez Chapa. Glucose optimal control system in diabetes treatment. Applied Mathematics and Computation, 209(1):19– 30, March 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004700.

Ma:2006:LPF

[cMpZfR06] Yi chen Ma, Zhi peng Zhang, and Chen feng Ren. Local and parallel finite element algorithms based on two-grid discretization for the stream function form of Navier–Stokes equations. Applied Mathematics and Computation, 175(1):786– 813, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006454.

Conti:2008:LTP

[CMR08] Costanza Conti, Rossana Morandi, and Christophe Rabut. Locally tensor product functions in algebraic grid optimization. Applied Mathematics and Computation, 197(2):745– 754, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008338.

Conti:2005:AEA

[CMS05] Costanza Conti, Rossana Morandi, and Rosa Maria Spitaleri. An algebraic– elliptic algorithm for boundary orthogonal grid generation. Applied Mathematics and Computation, 162 (1):15–27, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304000529.

Caliari:2005:BPI

[CMV05] Marco Caliari, Stefano De Marchi, and Marco Vianello. Bivariate polynomial interpolation on the square at new nodal sets. Applied Mathematics and Computation, 165(2):261–274, June 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003790.

Chakib:2005:NPA

[CN05] A. Chakib and A. Nachaoui. Nonlinear programming approach for a transient free boundary flow problem. Applied Mathematics and Computation, 160(2):317–328, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303009536.

Chakib:2006:SFB

[CN06] A. Chakib and A. Nachaoui. Simulation of a free boundary problem using boundary element method and restarted re-analysis technique. Applied Mathematics and Computation, 177(2):824–838, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500963X.

Chun:2009:CIN

[CN09a] Changbum Chun and Beny Neta. Certain improvements of Newton’s method with fourth-order convergence. Applied Mathematics and Computation, 215(2):821–828, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005700.

Chun:2009:TOM

[CN09b] Changbum Chun and Beny Neta. A third-order modification of Newton’s method for multiple roots. Applied Mathematics and Computation, 211(2):474–479, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001052.

Cvetkovic:2009:SMT

[CN09c] Ljiljana Cvetković and Maja Nedović. Special H-matrices and their Schur and diagonal-Schur complements. Applied Mathematics and Computation, 208(1):225–230, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800893X.

Chen:2006:FIM

[CO06] Liang-Ho Chen and Liang-Yuh Ouyang. Fuzzy inventory model for deteriorating items with permissible delay in payment. Applied Mathematics and Computation, 182(1):711–726, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003791.

Chen:2009:HPT

[CO09] Yuqing Chen and Donal O’Regan. On the homotopy property of topological degree for maximal monotone mappings. Applied Mathematics and Computation, 208(2):373– 377, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300308009053.

Cho:2008:DNM

[COG08] Huidae Cho, Francisco Olivera, and Seth D. Guikema. A derivation of the number of minima of the Griewank function. Applied Mathematics and Computation, 204(2):694–701, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005419.

Cooper:2006:DHE

[Coo06] S. Barry Cooper. Definability as hypercomputational effect. Applied Mathematics and Computation, 178(1):72– 82, July 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008350.

Cooper:2009:ECT

[Coo09] S. Barry Cooper. Emergence as a computabilitytheoretic phenomenon. Applied Mathematics and Computation, 215(4):1351– 1360, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309004159.

Cortell:2005:NMF

[Cor05a] Rafael Cortell. A note on magnetohydrodynamic flow of a power-law fluid over a stretching sheet. Applied Mathematics and Computation, 168(1):557– 566, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304006241.

Cortell:2005:NAD

[Cor05b] Rafael Cortell. Numerical analysis of dynamic problems: an alternative integration. Applied Mathematics and Computation, 170(1):158–171, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008793.

Cortell:2005:NSC

[Cor05c] Rafael Cortell. Numerical solutions of the classical Blasius flat-plate problem. Applied Mathematics and Computation, 170(1):706–710, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000111.

Cortell:2007:VFH

[Cor07] Rafael Cortell. Viscous flow and heat transfer over a nonlinearly stretching sheet. Applied Mathematics and Computation, 184(2):864–873, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007740.

Cotsaftis:2008:ICH

[Cot08] Michel Cotsaftis. Integrated control for high accuracy complex system. Applied Mathematics and Computation, 206(2):555–560, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003780.

Clancy:2005:EOI

[CP05] Damian Clancy and Alexei B.

Piunovskiy. An explicit optimal isolation policy for a deterministic epidemic model. Applied Mathematics and Computation, 163(3):1109–1121, April 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002838.

Calvo:2008:VSI

[CP08] M. P. Calvo and A. M. Portillo. Variable step implementation of ETD methods for semilinear problems. Applied Mathematics and Computation, 196(2): 627–637, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307007163.

Chakravarthy:2007:SON

[CPR07] P. Pramod Chakravarthy, K. Phaneendra, and Y. N. Reddy. A seventh order numerical method for singular perturbation problems. Applied Mathematics and Computation, 186(1):860–871, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010083.

Castelletti:2008:IPO

[CPSS08a] Andrea Castelletti, Francesca Pianosi, and Rodolfo Soncini-Sessa. Integration, participation and optimal control in water resources planning and management. Applied Mathematics and Computation, 206(1):21–33, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700999X.

Castelletti:2008:RHC

[CPSS08b] Andrea Castelletti, Francesca Pianosi, and Rodolfo Soncini-Sessa. Receding horizon control for water resources management. Applied Mathematics and Computation, 204(2):621–631, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800310X.

Climent:2005:OTW

[CPTZ05] Joan-Josep Climent, Carmen Perea, Leandro Tortosa, and Antonio Zamora. An overlapped two-way method for solving tridiagonal linear systems in a BSP computer. Applied Mathematics and Computation, 161(2):475–500, February 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304000062.

Chen:2009:LTC

[CPW09] Sh. Chen, S. Ponnusamy, and X. Wang. Landau’s theorem for certain biharmonic mappings. Applied Mathematics and Computation, 208(2):427–433, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009120.

Ceng:2009:SCM

[CPY09] Lu-Chuan Ceng, Adrian Petruşel, and Jen-Chih Yao. Strong convergence of modified implicit iterative algorithms with perturbed mappings for continuous pseudocontractive mappings. Applied Mathematics and Computation, 209(2):162–176, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007704.

Cicci:2008:TBM

[CQL08] D. A. Cicci, C. Qualls, and G. Landingham. Twobody missile separation dynamics. Applied Mathematics and Computation, 198(1):44–58, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700865X.

Covello:2005:SEE

[CR05a] Paul Covello and Garry Rodrigue. Solving the Eikonal equation on an adaptive mesh. Applied Mathematics and Computation, 166(3):678–695, July 26, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004667.

Cvetkovic:2005:HIM

[CR05b] Ljiljana Cvetković and Sanja Rapajiç. How to improve MAOR method convergence area for linear complementarity problems. Applied Mathematics and Computation, 162(2):577–584, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400116X.

Cao:2006:AMF

[CR06] Yufei Cao and Hongxing Rui. On the accuracy of the mortar finite volume element method with Lagrange multipliers. Applied Mathematics and Computation, 183(2):1386–1398, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006874.

Chowdhury:2008:MMG

[CRC08] Tanmay Chowdhury, Shovonlal Roy, and J. Chattopadhyay. Modeling migratory grazing of zooplankton on toxic and non-toxic phytoplankton. Applied Mathematics and Computation, 197(2):659–671, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008235.

Ciric:2008:MII

[ĆRR^{+}08] Ljubomir Ćirić, Arif Rafiq, Stojan Radenović, Miloje Rajović, and Jeong Sheok Ume. On Mann implicit iterations for strongly accretive and strongly pseudo-contractive mappings. Applied Mathematics and Computation, 198(1):128–137, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008739.

Cancelliere:2005:DSG

[CS05a] Rossella Cancelliere and Angela Slavova. Dynamics and stability of generalized cellular nonlinear network model. Applied Mathematics and Computation, 165(1):127–136, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003662.

Chen:2005:IGN

[CS05b] Yurong Chen and Jiachang Sun. An irregular grid for the numerical solution of linear elliptic partial differential equations. Applied Mathematics and Computation, 166(1):84– 94, July 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004230.

Choi:2005:CFS

[CS05c] Junesang Choi and H. M.

Srivastava. Certain families of series associated with the Hurwitz–Lerch Zeta function. Applied Mathematics and Computation, 170(1):399–409, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009579.

Chen:2006:GAA

[CS06a] Fengde Chen and Chunling Shi. Global attractivity in an almost periodic multispecies nonlinear ecological model. Applied Mathematics and Computation, 180(1):376–392, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305011331.

Colebrook:2006:AAC

[CS06b] M. Colebrook and J. Sicilia. An O(mn) algorithm for the anti-cent-dian problem. Applied Mathematics and Computation, 183(1):350–364, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005674.

Chakraborty:2007:HRA

[CS07a] Soubhik Chakraborty and Suman Kumar Sourabh. How robust are average complexity measures? A statistical case study. Applied Mathematics and Computation, 189(2):1787– 1797, June 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306017735.

Chakraborty:2007:WAT

[CS07b] Soubhik Chakraborty and Suman Kumar Sourabh. On why an algorithmic time complexity measure can be system invariant rather than system independent. Applied Mathematics and Computation, 190(1):195–204, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000409.

Chakraborty:2007:SEF

[CS07c] Soubhik Chakraborty and Kiran Kumar Sundararajan. A simple empirical formula for categorizing computing operations. Applied Mathematics and Computation, 189(1):326– 340, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016067.

Chakraborty:2007:WAS

[CS07d] Soubhik Chakraborty and Kiran Kumar Sundararajan. Winograd’s algorithm statistically revisited: It pays to weigh than to count! Applied Mathematics and Computation, 190(1):15– 20, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000239.

Che:2007:MSM

[CS07e] Jianren Che and Ke Su.

A modified SQP method and its global convergence. Applied Mathematics and Computation, 186(2):945–951, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010174.

Chen:2007:DNR

[CS07f] Fengde Chen and Jinlin Shi. On a delayed nonautonomous ratio-dependent predator–prey model with Holling type functional response and diffusion. Applied Mathematics and Computation, 192(2):358– 369, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307003372.

Chen:2007:CAS

[CS07g] Jinhai Chen and Zuhe Shen. Convergence analysis of the secant type methods. Applied Mathematics and Computation, 188(1):514– 524, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306013919.

Chen:2007:TOC

[CS07h] Jinhai Chen and Zuhe Shen. On third-order convergent regula falsi method. Applied Mathematics and Computation, 188(2):1592–1596, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015128.

Chen:2007:RCG

[CS07i] Jinhai Chen and Zuhe Shen. Regularized conjugate gradient method for skew-symmetric indefinite system of linear equations and applications. Applied Mathematics and Computation, 187(2):1484–1494, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012872.

Cvijovic:2007:SFF

[CS07j] Djurdje Cvijović and H. M.

Srivastava. Summation of a family of finite secant sums. Applied Mathematics and Computation, 190(1):590– 598, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000938.

Cao:2008:SAM

[CS08a] Guangxi Cao and Yongzhong Song. On the semiconvergence of additive and multiplicative splitting iterations for singular linear systems. Applied Mathematics and Computation, 204(2):794–801, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005626.

Chen:2008:OPP

[CS08b] Miao-Sheng Chen and Yu-Ti Shih. Optimal pricing for patent protected new products. Applied Mathematics and Computation, 197(2):525–535, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307007990.

Caballero:2009:HHI

[CS09a] J. Caballero and K. Sadarangani. Hermite–Hadamard inequality for fuzzy integrals. Applied Mathematics and Computation, 215(6):2134–2138, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309007218.

Chang:2009:TSO

[CS09b] Gao Chang and Zifei Shen. Three solutions for an obstacle problem for a class of variational– hemivariational inequalities. Applied Mathematics and Computation, 215(6):2063–2069, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007152.

Chien:2009:CHF

[CS09c] C.-S. Chien and Y.-T.

Shih. A cubic Hermite finite element-continuation method for numerical solutions of the von Kármán equations. Applied Mathematics and Computation, 209(2):356–368, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009727.

Chakraborty:2007:RSR

[CSBS07] Soubhik Chakraborty, Suman Kumar Sourabh, Mausumi Bose, and Kumar Sushant. Replacement sort revisited: The “gold standard” unearthed! Applied Mathematics and Computation, 189(1):384–394, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016134.

Christou:2008:NIN

[CSC08] M. A. Christou, C. Sophocleous, and C. I. Christov. Numerical investigation of the nonlinear heat diffusion equation with high nonlinearity on the boundary. Applied Mathematics and Computation, 201(1–2):729–738, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000404.

Coll:2007:ANV

[CST07] Carmen Coll, Elena Sánchez, and Néstor Thome. An algorithm for normalizing variable control systems. Applied Mathematics and Computation, 192(2):439–445, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003475.

Ceballos:2007:KVK

[CSV07] Juan Carlos Ceballos, Mauricio Sepúlveda, and Octavio Paulo Vera Villagrán. The Korteweg–de Vries–Kawahara equation in a bounded domain and some numerical results. Applied Mathematics and Computation, 190(1):912– 936, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001221.

Cheng:2007:NFS

[CSY07] MingBao Cheng, ShiJie Sun, and Ying Yu. A note on flow shop scheduling problems with a learning effect on no-idle dominant machines. Applied Mathematics and Computation, 184(2):945–949, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007843.

See erratum [Wan08b].

Chen:2005:BSN

[CT05a] Y. C. Chen and L. Y.

Tsai. Blow-up solutions of nonlinear differential equations. Applied Mathematics and Computation, 169(1):366–387, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007805.

Chen:2005:CGI

[CT05b] Yong-Lin Chen and Xue-Yuan Tan. Computing generalized inverses of matrices by iterative methods based on splittings of matrices. Applied Mathematics and Computation, 163(1):309–325, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002279.

Chen:2005:SCI

[CT05c] Yong-Lin Chen and Xue-Yuan Tan. Semiconvergence criteria of iterations and extrapolated iterations and constructive methods of semiconvergent iteration matrices. Applied Mathematics and Computation, 167(2):930–956, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304005405.

Cordero:2006:VNM

[CT06] A. Cordero and Juan R. Torregrosa. Variants of Newton’s method for functions of several variables. Applied Mathematics and Computation, 183(1):199–208, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005522.

Cakmak:2007:CPO

[ÇT07a] Devrim Çakmak and Aydın Tiryaki. Comment on the paper “Oscillation of second-order nonlinear ODE with damping”. Applied Mathematics and Computation, 191(1):298, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002275.

See [ZM06b].

Chen:2007:RCT

[CT07b] Y-Chuang Chen and Jimmy J. M. Tan. Restricted connectivity for three families of interconnection networks. Applied Mathematics and Computation, 188(2):1848–1855, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015529.

Cordero:2007:VNM

[CT07c] A. Cordero and Juan R.

Torregrosa. Variants of Newton’s Method using fifth-order quadrature formulas. Applied Mathematics and Computation, 190(1):686–698, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700104X.

Chen:2008:GFT

[CT08a] Chi-Chi Chen and Hui-Chin Tang. On goodnessof-fit tests for multiple recursive random number generators. Applied Mathematics and Computation, 200(1):70–79, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010867.

Cordero:2008:CMP

[CT08b] A. Cordero and Juan R. Torregrosa. A class of multi-point iterative methods for nonlinear equations. Applied Mathematics and Computation, 197(1):337–344, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307007813.

Chen:2009:MPS

[CT09a] Shang-Jie Chen and Chun-Lei Tang. Multiple periodic solutions for superquadratic first-order discrete Hamiltonian systems. Applied Mathematics and Computation, 208(2):495–500, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009223.

Chung:2009:FIA

[CT09b] Kuo-Liang Chung and Chia-Lun Tsai. Fast incremental algorithm for speeding up the computation of binarization. Applied Mathematics and Computation, 212(2):396– 408, June 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309001714.

Chen:2006:PTI

[CTG06] Biaosong Chen, Liyong Tong, and Yuanxian Gu. Precise time integration for linear two-point boundary value problems. Applied Mathematics and Computation, 175(1):182–211, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006004.

Chiyaka:2008:MAE

[CTGD08] C. Chiyaka, J. M. Tchuenche, W. Garira, and S. Dube. A mathematical analysis of the effects of control strategies on the transmission dynamics of malaria. Applied Mathematics and Computation, 195(2):641–662, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005930.

Chu:2009:EAM

[CTL09] Chun-Jung Chu, Vincent S. Tseng, and Tyne Liang. An efficient algorithm for mining high utility itemsets with negative item values in large databases. Applied Mathematics and Computation, 215(2):767–778, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030900561X.

Catania:2008:SAT

[CTT08] A. Catania, M. Torrisi, and R. Tracinà. Symmetry approach for a three coupled Schroedinger equation system. Applied Mathematics and Computation, 204(1):408–415, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005195.

Cui:2005:BCN

[Cui05]  Xiaoke Cui.  Bounds

on condition number of a singular matrix and its applications. Applied Mathematics and Computation, 162(1):81– 93, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000724.

Cui:2008:IAS

[Cui08] Mingrong Cui. On the iterative algorithm for saddle point problems. Applied Mathematics and Computation, 204(1):10–13, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002774.

Cvijovic:2008:CFS

[Cvi08] Djurdje Cvijović. Closedform summation of two families of finite tangent sums. Applied Mathematics and Computation, 196(2):661–665, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007199.

Cvijovic:2009:DPC

[Cvi09a] Djurdje Cvijović. Derivative polynomials and closedform higher derivative formulae. Applied Mathematics and Computation, 215(8):3002–3006, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309008728.

Cvijovic:2009:SFF

[Cvi09b] Djurdje Cvijović. Summation formulae for finite cotangent sums. Applied Mathematics and Computation, 215(3):1135–1140, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006250.

Caliari:2006:HNC

[CVMM06] Marco Caliari, Marco Vianello, Stefano De Marchi, and Roberto Montagna. Hyper2d: a numerical code for hyperinterpolation on rectangles. Applied Mathematics and Computation, 183(2):1138–1147, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006485.

Chen:2005:NGAa

[CW05a] Yong Chen and Qi Wang. A new general algebraic method with symbolic computation to construct new doubly-periodic solutions of the (2 + 1)- dimensional dispersive long wave equation. Applied Mathematics and Computation, 167(2):919–929, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005399.

Chen:2005:NGAbb

[CW05b] Yong Chen and Qi Wang.

A new general algebraic method with symbolic computation to construct new travelling wave solution for the (1 + 1)- dimensional dispersive long wave equation. Applied Mathematics and Computation, 168(2):1189–1204, September 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304006848.

Chen:2006:GAD

[CW06a] Haibo Chen and Haihua Wang. Global attractivity of the difference equation x_{n+1} = (x_{n} + αx_{n−1})/(β + x_{n}). Applied Mathematics and Computation, 181(2):1431–1438, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306002621.

Chen:2006:AMQ

[CW06b] Ronghua Chen and Zongmin Wu. Applying multiquadric quasi-interpolation to solve Burgers’ equation. Applied Mathematics and Computation, 172(1):472–484, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030500216X.

Chen:2006:NEE

[CW06c] Yong Chen and Qi Wang. A new elliptic equation rational expansion method and its application to the shallow long wave approximate equations. Applied Mathematics and Computation, 173(2):1163–1182, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305004212.

Chen:2006:URE

[CW06d] Yong Chen and Qi Wang. A unified rational expansion method to construct a series of explicit exact solutions to nonlinear evolution equations. Applied Mathematics and Computation, 177(1):396–409, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009203.

Chen:2007:SPD

[CW07] Ronghua Chen and Zongmin Wu. Solving partial differential equation by using multiquadric quasiinterpolation. Applied Mathematics and Computation, 186(2):1502–1510, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010769.

Cho:2008:MAD

[CW08] Hsun-Jung Cho and Yuh-Ting Wu. Microscopic analysis of desired-speed car-following stability. Applied Mathematics and Computation, 196(2):638– 645, March 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307007175.

Chao:2008:QCQ

[CWLH08] Mian-Tao Chao, Zhong-Xing Wang, Yu-Mei Liang, and Qing-Jie Hu. Quadratically constraint quadratical algorithm model for nonlinear minimax problems. Applied Mathematics and Computation, 205(1):247–262, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006334.

Wang:2006:ASN

[cWmZ06] Ting chun Wang and Lu ming Zhang. Analysis of some new conservative schemes for nonlinear Schrödinger equation with wave operator. Applied Mathematics and Computation, 182(2):1780–1794, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600525X.

Wu:2007:COS

[cWxChY07] Guo chang Wu, Zheng xing Cheng, and Xiao hui Yang. The cardinal orthogonal scaling function and sampling theorem in the wavelet subspaces. Applied Mathematics and Computation, 194(1):199–214, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004596.

Chan:2006:PMA

[CWY06] Raymond H. Chan, Chi-Yan Wong, and Kit-Ming Yeung. Pricing multi-asset American-style options by memory reduction Monte Carlo methods. Applied Mathematics and Computation, 179(2):535–544, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010581.

Cui:2007:FBA

[CWZ07] Lihong Cui, Weixing Wang, and Jinming Zhang. On a family of biorthogonal armlets with prescribed properties. Applied Mathematics and Computation, 193(1):36–65, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003736.

Chen:2006:PEN

[CX06] Fengde Chen and Xiangdong Xie. Permanence and extinction in nonlinear single and multiple species system with diffusion. Applied Mathematics and Computation, 177(1):410– 426, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009215.

Chen:2008:RWD

[CX08a] Jianbiao Chen and Zhaoliang Xu. Representations for the weighted Drazin inverse of a modified matrix. Applied Mathematics and Computation, 203(1):202–209, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002300.

Chen:2008:CCN

[CX08b] Jueping Chen and Zhaoliang Xu. Comment on “Condition number of Drazin inverse and their condition numbers of singular linear systems”. Applied Mathematics and Computation, 199(2):512– 526, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030701017X.

See [WWW03].

Chen:2009:NHP

[CX09] Ruyun Chen and Shuhuang Xiang. Note on the homotopy perturbation method for multivariate vectorvalue oscillatory integrals. Applied Mathematics and Computation, 215(1):78– 84, September 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309003853.

Chen:2006:PGAa

[CXC06] Fengde Chen, Xiangdong Xie, and Xiaoxing Chen. Permanence and global attractivity of a delayed periodic logistic equation. Applied Mathematics and Computation, 177(1):118– 127, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008945.

Chen:2009:CAT

[CXZL09] Zhe Chen, Changhe Xiang, Kequan Zhao, and Xuewen Liu. Convergence analysis of Tikhonov-type regularization algorithms for multiobjective optimization problems. Applied Mathematics and Computation, 211(1):167–172, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000666.

Chang:2005:ILO

[CY05a] Ming-Hsin Chang and Yi-Shiung Yeh. Improving Lamport one-time signature scheme. Applied Mathematics and Computation, 167(1):118–124, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004783.

Chen:2005:ENB

[CY05b] Yen-Cheng Chen and Lo-Yao Yeh. An efficient nonce-based authentication scheme with key agreement. Applied Mathematics and Computation, 169(2):982–994, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008288.

Cui:2005:ESK

[CY05c] Ming Gen Cui and Li Hong Yang. The exact solutions of a kind of nonlinear operator equations. Applied Mathematics and Computation, 167(1):607–615, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030400517X.

Chen:2006:WSR

[CY06a] Yong Chen and Zhenya Yan. Weierstrass semirational expansion method and new doubly periodic solutions of the generalized Hirota–Satsuma coupled KdV system. Applied Mathematics and Computation, 177(1):85– 91, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008891.

Cordeiro:2006:RWA

[CY06b] Celso José Cordeiro and Jin Yun Yuan. Rowwise algorithms of ST decomposition. Applied Mathematics and Computation, 175(1):748–761, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006429.

Ceng:2007:ELA

[CY07a] Lu-Chuan Ceng and Jen-Chih Yao. An extragradientlike approximation method for variational inequality problems and fixed point problems. Applied Mathematics and Computation, 190(1):205–215, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000410.

Chen:2007:PIM

[CY07b] Fengde Chen and Minsheng You. Permanence for an integrodifferential model of mutualism. Applied Mathematics and Computation, 186(1):30– 34, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009131.

Cui:2007:PSR

[CY07c] Zhoujin Cui and Zuodong Yang. Positive singular rupture solutions to a class for quasilinear elliptic equations. Applied Mathematics and Computation, 188(1):399–405, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306013737.

Ceng:2008:HVA

[CY08a] Lu-Chuan Ceng and Jen-Chih Yao. Hybrid viscosity approximation schemes for equilibrium problems and fixed point problems of infinitely many nonexpansive mappings. Applied Mathematics and Computation, 198(2):729–741, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009496.

Chang:2008:FTC

[CY08b] Jou-Ming Chang and Jinn-Shyong Yang. Faulttolerant cycle-embedding in alternating group graphs. Applied Mathematics and Computation, 197(2):760– 767, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008351.

Chen:2008:ESSa

[CY08c] Yuzhen Chen and Xuelin Yong. Exact solutions to the sine-Gordon-type equations. Applied Mathematics and Computation, 202(2):539–543, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001380.

Cui:2008:BGS

[CY08d] Zhoujin Cui and Zuodong Yang. Boundedness of global solutions for a nonlinear degenerate parabolic (porous medium) system with localized sources. Applied Mathematics and Computation, 198(2):882– 895, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009630.

Chen:2009:MMF

[CYCX09] Fengde Chen, Jinghui Yang, Lijuan Chen, and Xiangdong Xie. On a mutualism model with feedback controls. Applied Mathematics and Computation, 214(2):581–587, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309003579.

Chung:2007:SBW

[CYH^{+}07] Kuo-Liang Chung, Wei-Ning Yang, Yong-Huai Huang, Shih-Tung Wu, and Yu-Chiao Hsu. On SVD-based watermarking algorithm. Applied Mathematics and Computation, 188(1):54–57, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030601318X.

Chen:2007:EAL

[CYL07] Wen-Lih Chen, Yu-Ching Yang, and Haw-Long Lee. Estimating the absorptivity in laser processing by inverse methodology. Applied Mathematics and Computation, 190(1):712– 721, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001087.

Chen:2008:OCC

[CYL^{+}08] Yen-Ping Chen, Jhun-Ying Yang, Shun-Nan Liou, Gwo-Yun Lee, and Jeen-Shing Wang. Online classifier construction algorithm for human activity detection using a tri-axial accelerometer. Applied Mathematics and Computation, 205(2):849–860, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003640.

Chang:2005:CIA

[CYY05] Ting-Yi Chang, Chou-Chan Yang, and Ya-Wen Yang. Cryptanalysis of the improved authenticated key agreement protocol. Applied Mathematics and Computation, 171 (2):771–774, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:2005:CMN

[CZ05a] Li-Qun Chen and Wei-Jia Zhao. A computation method for nonlinear vibration of axially accelerating viscoelastic strings. Applied Mathematics and Computation, 162(1):305–310, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000955.

Chen:2005:NMS

[CZ05b] Li-Qun Chen and Wei-Jia Zhao. A numerical method for simulating transverse vibrations of an axially moving string. Applied Mathematics and Computation, 160(2):411–422, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030301169X.

Cao:2007:NDD

[CZ07a] Jiuwen Cao and Shouming Zhong. New delaydependent condition for absolute stability of Lurie control systems with multiple time-delays and nonlinearities. Applied Mathematics and Computation, 194(1):250–258, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700464X.

Cao:2007:ORU

[CZ07b] Yusong Cao and Yi Zhang.

Optimal reinsurance under the general mixture risk measures. Applied Mathematics and Computation, 185(1):229–239, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600854X.

Cui:2007:NSS

[CZ07c] Yujun Cui and Yumei Zou. Nontrivial solutions of singular superlinear m-point boundary value problems. Applied Mathematics and Computation, 187(2):1256–1264, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012628.

Chen:2008:OBS

[CZ08a] Hanning Chen and Yunlong Zhu. Optimization based on symbiotic multi-species coevolution. Applied Mathematics and Computation, 205(1):47–60, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003172.

Chen:2008:NTW

[CZ08b] Lanxin Chen and Junxian Zhang. Non-travelling wave solutions of two (2 + 1)-dimensional nonlinear mathematical physics equations. Applied Mathematics and Computation, 198(1):184–197, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008788.

Chen:2009:DAH

[CZF09] Zhang Chen, Donghua Zhao, and Xilin Fu. Discrete analogue of highorder periodic Cohen– Grossberg neural networks with delay. Applied Mathematics and Computation, 214(1):210–217, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002951.

Cao:2007:GSA

[CZH07] Jiuwen Cao, Shouming Zhong, and Yuanyuan Hu. Global stability analysis for a class of neural networks with varying delays and control input. Applied Mathematics and Computation, 189(2):1480–1490, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601736X.

Cao:2008:NDD

[CZH08] Jiuwen Cao, Shouming Zhong, and Yuanyuan Hu. Novel delay-dependent stability conditions for a class of MIMO networked control systems with nonlinear perturbation. Applied Mathematics and Computation, 197(2):797–809, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008405.

Chen:2005:STV

[CZZ05] Li-Qun Chen, Wei-Jia Zhao, and Jean W. Zu. Simulations of transverse vibrations of an axially moving string: a modified difference approach. Applied Mathematics and Computation, 166(3):596–607, July 26, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004606.

deAndrade:2009:CAA

[dAC09] Bruno de Andrade and Claudio Cuevas. Compact almost automorphic solutions to semilinear Cauchy problems with non-dense domain. Applied Mathematics and Computation, 215(8):2843–2849, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008236.

Dahl:2007:NPR

[Dah07] Geir Dahl. A note on a parameter relating traffic equilibria and system optimal routing. Applied Mathematics and Computation, 191(2):445–450, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700272X.

Dai:2007:FEA

[Dai07] Xiaoxia Dai. Finite element approximation of the pure Neumann problem using the iterative penalty method. Applied Mathematics and Computation, 186(2):1367–1373, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010617.

Dai:2009:AMS

[Dai09] Guowei Dai. Arbitrarily many solutions for a perturbed p(x)-Laplacian equation involving oscillatory terms. Applied Mathematics and Computation, 215(7):2544–2556, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007929.

Damaceanu:2007:ISM

[Dam07] Romulus-Catalin Damaceanu.

Implementation of simulation model of world economy using LSD. Applied Mathematics and Computation, 189(2):1011–1033, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016092.

Damaceanu:2008:ABC

[Dam08] Romulus-Catalin Damaceanu.

An agent-based computational study of wealth distribution in function of resource growth interval using NetLogo. Applied Mathematics and Computation, 201(1–2):371–377, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307012143.

Danchick:2005:RLA

[Dan05] Roy Danchick. A ranked linear assignment approach to Bayesian classification. Applied Mathematics and Computation, 162(1):265– 281, March 4, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304000876.

Danchick:2006:ANP

[Dan06] Roy Danchick. Accurate numerical partials with applications to optimization. Applied Mathematics and Computation, 183(1):551–558, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600590X.

Danchick:2008:GMN

[Dan08]  Roy Danchick.  Gauss

meets Newton again: How to make Gauss orbit determination from two position vectors more efficient and robust with Newton–Raphson iterations. Applied Mathematics and Computation, 195(2):364–375, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003360.

Daoud:2005:NSM

[Dao05] Daoud S. Daoud. On the numerical solution of multi-dimensional parabolic problem by the additive splitting up method. Applied Mathematics and Computation, 162(1):197– 210, March 4, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304000803.

DelaSen:2009:CIB

[DAQ09] M. De la Sen and S. Alonso-Quesada. Control issues for the Beverton–Holt equation in ecology by locally monitoring the environment carrying capacity: Non-adaptive and adaptive cases. Applied Mathematics and Computation, 215(7):2616–2633, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007991.

Darwish:2006:SUU

[Dar06] Mohamed Abdalla Darwish. Solvability of Urysohn and Urysohn– Volterra equations with hysteresis in weighted spaces. Applied Mathematics and Computation, 176(2):722–726, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008726.

Davis:2006:WTN

[Dav06] Martin Davis. Why there is no such discipline as hypercomputation. Applied Mathematics and Computation, 178(1):4–7, July 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008295.

Day:2005:AEF

[Day05] Richard H. Day. Adaptive economizing and financial feedback in pure competition. Applied Mathematics and Computation, 164(2):435–451, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003285.

Derouich:2006:DFM

[DB06] M. Derouich and A. Boutayeb. Dengue fever: Mathematical modelling and computer simulation. Applied Mathematics and Computation, 177(2):528–544, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009331.

Darvishi:2007:FOM

[DB07a] M. T. Darvishi and A. Barati.

A fourth-order method from quadrature formulae to solve systems of nonlinear equations. Applied Mathematics and Computation, 188(1):257–261, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013476.

Darvishi:2007:SCI

[DB07b] M. T. Darvishi and A. Barati.

Super cubic iterative methods to solve systems of nonlinear equations. Applied Mathematics and Computation, 188(2):1678–1685, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015220.

Darvishi:2007:TON

[DB07c] M. T. Darvishi and A. Barati.

A third-order Newtontype method to solve systems of nonlinear equations. Applied Mathematics and Computation, 187(2):630–635, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011180.

Du:2007:ASS

[DB07d] Zengji Du and Zhanbing Bai. Asymptotic solutions for a secondorder differential equation with three-point boundary conditions. Applied Mathematics and Computation, 186(1):469–473, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009556.

Dubey:2009:ERS

[DB09] Shruti A. Dubey and Dhirendra Bahuguna. Existence and regularity of solutions to nonlocal retarded differential equations. Applied Mathematics and Computation, 215(7):2413–2424, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007668.

DelaCruz:2007:HOL

[DBC^{+}07] H. De la Cruz, R. J. Biscay, F. Carbonell, T. Ozaki, and J. C. Jimenez. A higher order local linearization method for solving ordinary differential equations. Applied Mathematics and Computation, 185(1):197–212, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008514.

Debbal:2007:TFA

[DBR07] S. M. Debbal and F. Bereksi-Reguig. Time-frequency analysis of the first and the second heartbeat sounds. Applied Mathematics and Computation, 184(2):1041– 1052, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306007922.

Dong:2006:ASN

[DC06a] Bo-Qing Dong and Zhi-Min Chen. Asymptotic stability of non-Newtonian flows with large perturbation in R^{2}. Applied Mathematics and Computation, 173(1):243–250, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305003401.

Dong:2006:GAT

[DC06b] Bo-Qing Dong and Zhi-Min Chen. Global attractors of two-dimensional micropolar fluid flows in some unbounded domains. Applied Mathematics and Computation, 182(1):610–620, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003705.

Dong:2006:RUL

[DC06c] Bo-Qing Dong and Zhi-Min Chen. Remarks on upper and lower bounds of solutions to the Navier– Stokes equations in R^{2}. Applied Mathematics and Computation, 182(1):553– 558, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306003638.

Doria:2006:ISI

[DC06d] Francisco Antonio Doria and José Félix Costa. Introduction to the special issue on hypercomputation. Applied Mathematics and Computation, 178(1):1–3, July 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008283.

Du:2006:RES

[DC06e] Hong Du and Minggen Cui. Representation of the exact solution and a stability analysis on the Fredholm integral equation of the first kind in reproducing kernel space. Applied Mathematics and Computation, 182(2):1608–1614, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005078.

Dorini:2007:FVM

[DC07] Fábio A. Dorini and M. Cristina C. Cunha. A finite volume method for the mean of the solution of the random transport equation. Applied Mathematics and Computation, 187(2):912–921, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012239.

Dai:2008:IPM

[DC08] Xiaoxia Dai and Xiaoliang Cheng. The iterative penalty method for Stokes equations using Q_{1} − P_{0} element. Applied Mathematics and Computation, 201(1–2):805–810, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011587.

daCosta:2003:CED

[dCD03] N. C. A. da Costa and F. A. Doria. Consequences of an exotic definition for P = NP. Applied Mathematics and Computation, 145(2–3):655–665, December 25, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See addendum [dCD06a].

daCosta:2006:ACE

[dCD06a] N. C. A. da Costa and F. A. Doria. Addendum to “Consequences of an exotic formulation for P = NP” [Appl. Math. Comput. 145 (2–3) (2003) 655–665]. Applied Mathematics and Computation, 172(2):1364–1367, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002717.

See [dCD03].

daCosta:2006:STH

[dCD06b] N. C. A. da Costa and F. A. Doria. Some thoughts on hypercomputation. Applied Mathematics and Computation, 178(1):83– 92, July 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008362.

daCosta:2009:HBH

[dCD09] N. C. A. da Costa and F. A. Doria. How to build a hypercomputer. Applied Mathematics and Computation, 215(4):1361–1367, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004135.

daCosta:2007:MPV

[dCDB07] N. C. A. da Costa, F. A.

Doria, and E. Bir. On the metamathematics of the P vs. NP question. Applied Mathematics and Computation, 189(2):1223–1240, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017073.

Diaz:2007:NLA

[dCDLN^{+}07] J. J. del Coz Díaz, L. Peñalver Lamarca, P. J. García Nieto, Alfonso Lozano Martínez-Luengas, and J. A. Vilán Vilán. Non-linear analysis of a reinforced concrete sheet cover (umbrella) of 40m diameter by the finite element method. Applied Mathematics and Computation, 184(1):37–51, January 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004838.

Diaz:2009:NAP

[dCDNPN09] J. J. del Coz Díaz, P. J. García Nieto, J. A. González Pérez, and A. Martín Navarro. Numerical analysis of the pressure field on curved and open selfweighted metallic roofs due to the wind effect by the finite volume method. Applied Mathematics and Computation, 209(1):31– 41, March 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004712.

deCarvalho:2008:RGM

[dCdSJM08] Esdras Penêdo de Carvalho, Anésio dos Santos Júnior, and To Fu Ma. Reduced gradient method combined with augmented Lagrangian and barrier for the optimal power flow problem. Applied Mathematics and Computation, 200(2):529–536, July 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011320.

Juez:2008:NLN

[dCJLNÁA08] F. J. de Cos Juez, F. Sánchez Lasheras, P. J. García Nieto, and A. Álvarez-Arenal. Non-linear numerical analysis of a double-threaded titanium alloy dental implant by FEM. Applied Mathematics and Computation, 206(2):952–967, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007674.

Ding:2007:MIL

[DCL07] Feng Ding, Huibo Chen, and Ming Li. Multiinnovation least squares identification methods based on the auxiliary model for MISO systems. Applied Mathematics and Computation, 187(2):658–668, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011210.

Dutta:2007:CRI

[DCR07] Pankaj Dutta, Debjani Chakraborty, and A. R. Roy. Continuous review inventory model in mixed fuzzy and stochastic environment. Applied Mathematics and Computation, 188(1):970–980, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014445.

Silva:2008:NSB

[dCRDSSN^{+}08] Sani de C. R. Da Silva, Dimitrios Samios, P. A. Netz, Dagoberto A. R. Justo, and A. L. De Bortoli. Numerical simulation of biological base pairs considering geometric and energetic criteria. Applied Mathematics and Computation, 200(2):602– 609, July 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030701140X.

Demetriou:2007:CSS

[DCS07] E. Demetriou, M. A. Christou, and C. Sophocleous. On the classification of similarity solutions of a two-dimensional diffusion–advection equation. Applied Mathematics and Computation, 187(2):1333–1350, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012720.

Dehghan:2005:RSI

[DD05] Mehdi Dehghan and Majid Jaberi Douraki. On the recursive sequence x_{n+1} = (α + βx_{n−k+1} + γx_{n−2k+1})/(Bx_{n−k+1} + Cx_{n−2k+1}). Applied Mathematics and Computation, 170(2):1045–1066, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000469.

Das:2006:DCI

[DD06a] Atin Das and Pritha Das. Does composite index of NYSE represents chaos in the long time scale? Applied Mathematics and Computation, 174(1):483–489, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004790.

Dehghan:2006:OCR

[DD06b] Mehdi Dehghan and Majid Jaberi Douraki. The oscillatory character of the recursive sequence x_{n} = (α + βx_{n−k+1})/(A + Bx_{n−2k+1}). Applied Mathematics and Computation, 175(1):38–48, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005898.

Das:2007:CAF

[DD07] Atin Das and Pritha Das. Chaotic analysis of the foreign exchange rates. Applied Mathematics and Computation, 185(1):388–396, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008721.

Deep:2008:QAB

[DD08a] Kusum Deep and Kedar Nath Das. Quadratic approximation based hybrid genetic algorithm for function optimization. Applied Mathematics and Computation, 203(1):86– 98, September 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308002099.

Deep:2008:SOM

[DD08b] Kusum Deep and Dipti.

A self-organizing migrating genetic algorithm for constrained optimization. Applied Mathematics and Computation, 198(1):237– 250, April 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307008843.

Dolecek:2009:ART

[DD09] Gordana Jovanovic Dolecek and Vlatko Dolecek. Application of Rouche’s theorem for MP filter design. Applied Mathematics and Computation, 211(2):329–335, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000782.

Dehghan:2005:DRD

[DDD05] Mehdi Dehghan, Majid Jaberi Douraki, and Marjan Jaberi Douraki. Dynamics of a rational difference equation using both theoretical and computational approaches. Applied Mathematics and Computation, 168(2):756–775, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304006368.

Douraki:2005:GBH

[DDR05a] M. Jaberi Douraki, M. Dehghan, and A. Razavi. On the global behavior of higher order recursive sequences. Applied Mathematics and Computation, 169(2):819–831, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008148.

Douraki:2005:QBS

[DDR05b] Majid Jaberi Douraki, Mehdi Dehghan, and Mohsen Razzaghi. The qualitative behavior of solutions of a nonlinear difference equation. Applied Mathematics and Computation, 170(1):485–502, November 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304009634.

Dehghan:2006:GSH

[DDR06a] Mehdi Dehghan, Majid Jaberi Douraki, and Mohsen Razzaghi. Global stability of a higher order rational recursive sequence. Applied Mathematics and Computation, 179(1):161–174, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010180.

Douraki:2006:OAB

[DDR06b] M. Jaberi Douraki, M. Dehghan, and M. Razzaghi. Oscillation and asymptotic behavior of a class of higher order nonlinear recursive sequences. Applied Mathematics and Computation, 179(1):175–189, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010192.

Douraki:2006:HOR

[DDR06c] Majid Jaberi Douraki, Mehdi Dehghan, and Mohsen Razzaghi. On the higher order rational recursive sequence x_{n} = A/x_{n−k} + B/x_{n−3k}. Applied Mathematics and Computation, 173(2):710– 723, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305003851.

Ding:2008:PIM

[DDZ08] Jie Ding, Feng Ding, and Shi Zhang. Parameter identification of multiinput, single-output systems based on FIR models and least squares principle. Applied Mathematics and Computation, 197(1):297–305, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007771.

Darania:2007:MNS

[DE07a] P. Darania and Ali Ebadian. A method for the numerical solution of the integro-differential equations. Applied Mathematics and Computation, 188 (1):657–668, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306014044.

Didehvar:2007:ARA

[DE07b] Farzad Didehvar and Changiz Eslahchi. An algorithm for rank aggregation problem. Applied Mathematics and Computation, 189(2):1847–1858, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017814.

Darwish:2006:EFI

[DEB06] Mohamed Abdalla Darwish and A. A. El-Bary. Existence of fractional integral equation with hysteresis. Applied Mathematics and Computation, 176(2):684–687, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008672.

deEspindola:2005:GFD

[dEdSNL05] J. J. de Espıñdola, João M.

da Silva Neto, and Eduardo M. O. Lopes. A generalised fractional derivative approach to viscoelastic material properties measurement. Applied Mathematics and Computation, 164(2):493–506, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003315.

Diogo:2005:NAS

[DEFT05] Teresa Diogo, John T. Edwards, Neville J. Ford, and Sophy M. Thomas. Numerical analysis of a singular integral equation. Applied Mathematics and Computation, 167(1):372–382, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304005004.

Dehghan:2005:NAS

[Deh05a]  Mehdi Dehghan.  Nu-

merical approximations for solving a time-dependent partial differential equation with non-classical specification on four boundaries. Applied Mathematics and Computation, 167(1):28–45, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004710.

Dehghan:2005:QIT

[Deh05b] Mehdi Dehghan. Quasiimplicit and two-level explicit finite-difference procedures for solving the one-dimensional advection equation. Applied Mathematics and Computation, 167(1):46–67, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004722.

DeLillo:2005:CHK

[DEH05c] Thomas K. DeLillo, Alan R. Elcrat, and Chenglie Hu. Computation of the Helmholtz– Kirchhoff and reentrant jet flows using Fourier series.

Applied Mathematics and Computation, 163(1):397– 422, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002334.

Dehmer:2008:IPC

[Deh08]  Matthias Dehmer.  In-

formation processing in complex networks: Graph entropy and information functionals. Applied Mathematics and Computation, 201(1–2):82–94, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011733.

Delvenne:2009:WUC

[Del09] Jean-Charles Delvenne.

What is a universal computing machine? Applied Mathematics and Computation, 215(4):1368–1374, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004147.

Demir:2005:NMV

[Dem05a] H. Demir. Numerical modelling of viscoelastic cavity driven flow using finite difference simulations. Applied Mathematics and Computation, 166(1):64– 83, July 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004229.

Demiralp:2005:CVC

[Dem05b] Metin Demiralp. Critical value calculations for the screening parameter of Hulthen potential. Applied Mathematics and Computation, 168 (2):1380–1399, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304007362.

Demiray:2005:CTW

[Dem05c] Hilmi Demiray. Complex travelling wave solutions to the KdV and Burgers equations. Applied Mathematics and Computation, 162(2):925–930, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001390.

Demiray:2005:MRP

[Dem05d] Hilmi Demiray. The modified reductive perturbation method as applied to Boussinesq equation: strongly dispersive case.

Applied Mathematics and Computation, 164(1):1–9, May 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400284X.

Demidenko:2006:SLE

[Dem06] Eugene Demidenko. Separable Laplace equation, magic Toeplitz matrix, and generalized Ohm’s law. Applied Mathematics and Computation, 181(2):1313–1327, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002505.

Demir:2008:TML

[Dem08]  Ali Demir.  Trisection

method by k-Lucas numbers. Applied Mathematics and Computation, 198(1):339–345, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008946.

Deng:2007:UFN

[Den07] Qin Deng. On univalent functions with negative coefficients. Applied Mathematics and Computation, 189(2):1675–1682, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017590.

Deng:2008:DIB

[Den08a] Chun Yuan Deng. The Drazin inverse of bounded operators with commutativity up to a factor. Applied Mathematics and Computation, 206(2):695–703, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006942.

Deng:2008:CPT

[Den08b] Xijun Deng. Comment on: Persistence of travelling wave for a coupled nonlinear wave system, Applied Mathematics and Computation 191 (2) (2007) 347–352. Applied Mathematics and Computation, 206(2):989–990, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006358.

See [LHF07].

Deng:2008:EPW

[Den08c] Xijun Deng. Exact peaked wave solution of CHγ equation by the firstintegral method. Applied Mathematics and Computation, 206(2):806–809, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800711X.

Deng:2008:TWS

[Den08d]  Xijun Deng.  Travel-

ling wave solutions for the generalized Burgers– Huxley equation. Applied Mathematics and Computation, 204(2):733–737, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800547X.

Deng:2009:NDI

[Den09a] Chun Yuan Deng. A note on the Drazin inverses with Banachiewicz–Schur forms. Applied Mathematics and Computation, 213(1):230– 234, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002227.

Deng:2009:CSA

[Den09b] Qin Deng. Certain subclass of analytic functions with complex order. Applied Mathematics and Computation, 208(2):359–362, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308009016.

Deng:2009:LMP

[Den09c] Shao-Gao Deng. A local mountain pass theorem and applications to a double perturbed p(x)- Laplacian equations. Applied Mathematics and Computation, 211(1):234– 241, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000757.

Deo:2008:DRE

[Deo08] Naokant Deo. Direct result on exponential-type operators. Applied Mathematics and Computation, 204(1):109–115, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004633.

Dehmer:2007:CLG

[DES07a] Matthias Dehmer and Frank Emmert-Streib. Comparing large graphs efficiently by margins of feature vectors. Applied Mathematics and Computation, 188(2):1699–1710, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015244.

Dehmer:2007:SSD

[DES07b] Matthias Dehmer and Frank Emmert-Streib. Structural similarity of directed universal hierarchical graphs: a low computational complexity approach. Applied Mathematics and Computation, 194(1):7–20, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004389.

Dehmer:2008:CAM

[DESG08] Matthias Dehmer, Frank Emmert-Streib, and Tanja Gesell. A comparative analysis of multidimensional features of objects resembling sets of graphs. Applied Mathematics and Computation, 196(1):221–235, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006625.

Dehmer:2006:SMG

[DESK06] Matthias Dehmer, Frank Emmert-Streib, and Jürgen Kilian. A similarity measure for graphs with low computational complexity. Applied Mathematics and Computation, 182(1):447–459, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003523.

Du:2008:TLN

[DF08] Dajun Du and Minrui Fei. A two-layer networked learning control system using actor–critic neural network. Applied Mathematics and Computation, 205(1):26–36, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003159.

Ding:2009:ASN

[DF09] Xie Ping Ding and Hai Rong Feng. Algorithm for solving a new class of generalized nonlinear implicit quasi-variational inclusions in Banach spaces. Applied Mathematics and Computation, 208(2):547–555, February 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308009284.

Dugnol:2007:WPC

[DFG07] B. Dugnol, C. Fernández, and G. Galiano. Wolf population counting by spectrogram image processing. Applied Mathematics and Computation, 186(1):820– 830, March 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306010046.

Dugnol:2007:IDD

[DFGV07] B. Dugnol, C. Fernández, G. Galiano, and J. Velasco. Implementation of a diffusive differential reassignment method for signal enhancement: an application to wolf population counting. Applied Mathematics and Computation, 193(2):374–384, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004109.

dOnofrio:2006:RAA

[dG06] Alberto d’Onofrio and Alberto Gandolfi. The response to antiangiogenic anticancer drugs that inhibit endothelial cell proliferation. Applied Mathematics and Computation, 181(2):1155–1162, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002001.

Daoud:2007:OSW

[DG07a] Daoud S. Daoud and Jürgen Geiser. Overlapping Schwarz wave form relaxation for the solution of coupled and decoupled system of convection diffusion reaction equation. Applied Mathematics and Computation, 190(1):946– 964, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001245.

Daoud:2007:TLE

[DG07b] Daoud S. Daoud and N. Gurbuz. Time lagging and explicit interface prediction for nonoverlapping domain decomposition with parallel additive splitting method for multidimensional parabolic problem. Applied Mathematics and Computation, 186(2):1094–1103, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010356.

Daftardar-Gejji:2008:SMT

[DGB08]  Varsha  Daftardar-Gejji

and Sachin Bhalekar. Solving multi-term linear and non-linear diffusion–wave equations of fractional order by Adomian decomposition method. Applied Mathematics and Computation, 202(1):113–120, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000647.

Daftardar-Gejji:2007:SMO

[DGJ07]  Varsha  Daftardar-Gejji

and Hossein Jafari. Solving a multi-order fractional differential equation using Adomian decomposition. Applied Mathematics and Computation, 189(1):541– 548, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016274.

Darvishi:2006:CGAa

[DH06a] M. T. Darvishi and P. Hessari. On convergence of the generalized AOR method for linear systems with diagonally dominant coefficient matrices. Applied Mathematics and Computation, 176(1):128–133, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305007782.

Darvishi:2006:CSM

[DH06b] M. T. Darvishi and P. Hessari. On convergence of the symmetric modified successive overrelaxation method for 2-cyclic coefficient matrices. Applied Mathematics and Computation, 183(2):953–960, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006291.

Darvishi:2006:SSM

[DH06c] M. T. Darvishi and P. Hessari. Symmetric SOR method for augmented systems. Applied Mathematics and Computation, 183(1):409–415, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005728.

Dehghan:2006:ISF

[DH06d] Mehdi Dehghan and Behnam Hashemi. Iterative solution of fuzzy linear systems. Applied Mathematics and Computation, 175(1):645– 674, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006375.

Dehghan:2006:SFF

[DH06e] Mehdi Dehghan and Behnam Hashemi. Solution of the fully fuzzy linear systems using the decomposition procedure. Applied Mathematics and Computation, 182(2):1568–1580, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005054.

Dehghan:2008:IAR

[DH08] Mehdi Dehghan and Masoud Hajarian. An iterative algorithm for the reflexive solutions of the generalized coupled Sylvester matrix equations and its optimal approximation. Applied Mathematics and Computation, 202(2):571–588, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001422.

Du:2009:NCT

[DH09] Bo Du and Xueping Hu. A new continuation theorem for the existence of solutions to p-Laplacian BVP at resonance. Applied Mathematics and Computation, 208(1):172–176, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008783.

Dehghan:2006:CMS

[DHG06] Mehdi Dehghan, Behnam Hashemi, and Mehdi Ghatee. Computational methods for solving fully fuzzy linear systems. Applied Mathematics and Computation, 179(1):328–343, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010404.

Du:2009:PST

[DHG09] Bo Du, Xueping Hu, and Weigao Ge. Positive solutions to a type of multi-point boundary value problem with delay and one-dimensional p-Laplacian. Applied Mathematics and Computation, 208(2):501–510, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009235.

Dai:2005:ECL

[DHL05] Tian-Shyr Dai, Guan-Shieng Huang, and Yuh-Dauh Lyuu. An efficient convergent lattice algorithm for European Asian options. Applied Mathematics and Computation, 169(2):1458–1471, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008665.

Dehghan:2007:SCB

[DHS07] Mehdi Dehghan, Asgar Hamidi, and Mohammad Shakourifar. The solution of coupled Burgers’ equations using Adomian– Padé technique. Applied Mathematics and Computation, 189(2):1034–1047, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016808.

Ding:2009:ADC

[DHT09] Zhanwen Ding, Qinglan Hang, and Lixin Tian. Analysis of the dynamics of Cournot team-game with heterogeneous players. Applied Mathematics and Computation, 215(3):1098–1105, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006171.

Darvishi:2006:CGAb

[DHY06] M. T. Darvishi, P. Hessari, and J. Y. Yuan. On convergence of the generalized accelerated overrelaxation method. Applied Mathematics and Computation, 181(1):468–477, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001160.

Dilao:2005:TIP

[Dil05] Rui Dilão. Turing instabilities and patterns near a Hopf bifurcation. Applied Mathematics and Computation, 164(2):391–414, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003261.

Ding:2005:IAS

[Din05]  Xie Ping Ding.  Iter-

ative algorithm of solutions for generalized mixed implicit equilibriumlike problems. Applied Mathematics and Computation, 162(2):799–809, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001328.

Dag:2005:NSB

[DIS05] İdris Dağ, Dursun Irk, and Bülent Saka. A numerical solution of the Burgers’ equation using cubic B-splines. Applied Mathematics and Computation, 163(1):199–211, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002218.

Dag:2006:LSF

[DIT06] İdris Dağ, Dursun Irk, and Mustafa Tombul. Leastsquares finite element method for the advection– diffusion equation. Applied Mathematics and Computation, 173(1):554–565, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003759.

Danchick:2006:MPD

[DJ06a] Roy Danchick and Mario L.

Juncosa. Maximum polynomial degree Nordsieck– Gear ( k p ) methods: Existence, stability, consistency, refinement, convergence and computational examples. Applied Mathematics and Computation, 182(1):907–933, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004085.

Darvishi:2006:NSB

[DJ06b] M. T. Darvishi and M. Javidi. A numerical solution of Burgers’ equation by pseudospectral method and Darvishi’s preconditioning. Applied Mathematics and Computation, 173(1):421–429, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003620.

Dorao:2007:PTS

[DJ07] C. A. Dorao and H. A. Jakobsen. A parallel time– space least-squares spectral element solver for incompressible flow problems. Applied Mathematics and Computation, 185(1):45–58, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008022.

Djelloul:2009:DRT

[Dje09]  Khalil Djelloul.  From

decomposable to residual theories. Applied Mathematics and Computation, 215(5):1761–1779, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006766.

Dalmagro:2006:BRP

[DJJL06] Fermín Dalmagro, Juan Jiménez, Raúl Jiménez, and Haydée Lugo. Boundedrational-prisoners’ dilemma: On critical phenomena of cooperation. Applied Mathematics and Computation, 176(2):462–469, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008234.

Date:2008:NAL

[DJM08] Paresh Date, Luka Jalen, and Rogemar Mamon. A new algorithm for latent state estimation in nonlinear time series models. Applied Mathematics and Computation, 203(1):224–232, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002336.

Djolovic:2006:SBE

[Djo06] Ivana Djolović. On the space of bounded Euler difference sequences and some classes of compact operators. Applied Mathematics and Computation, 182(2):1803–1811, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005273.

Djordjevic:2007:IMC

[Djo07a] Dragan S. Djordjević. Iterative methods for computing generalized inverses. Applied Mathematics and Computation, 189(1):101– 104, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015840.

Djordjevic:2007:MCJ

[Djo07b] Gospava B. Djordjević. Mixed convolutions of the Jacobsthal type. Applied Mathematics and Computation, 186(1):646–651, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009763.

Duan:2005:MCM

[DjT05a] Yong Duan and Yong ji Tan. Meshless collocation method based on Dirichlet–Neumann substructure iteration. Applied Mathematics and Computation, 166(2):373–384, July 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004436.

Duan:2005:MGM

[DjT05b] Yong Duan and Yong ji Tan. Meshless Galerkin method based on regions partitioned into subdomains. Applied Mathematics and Computation, 162(1):317–327, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000979.

Duan:2006:CNM

[DjT06] Yong Duan and Yong ji Tan. On condition number of meshless collocation method using radial basis functions. Applied Mathematics and Computation, 172(1):141–147, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305001931.

Dimitrov:2005:CMA

[DK05] Dobromir T. Dimitrov and Hristo V. Kojouharov. Complete mathematical analysis of predator–prey models with linear prey growth and Beddington– DeAngelis functional response. Applied Mathematics and Computation, 162(2):523–538, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001122.

Donmez:2006:SAJ

[DK06] Orhan Dönmez and Refik Kayali. Simulation of astrophysical jet using the special relativistic hydrodynamics code. Applied Mathematics and Computation, 182(2):1296–1308, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004449.

Darvishi:2006:DOV

[DKA06a] M. T. Darvishi and R. Khosro-Aghdam. Determination of the optimal value of relaxation parameter in symmetric SOR method for rectangular coefficient matrices. Applied Mathematics and Computation, 181(2):1018–1025, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001858.

Darvishi:2006:SSO

[DKA06b] M. T. Darvishi and R. Khosro-Aghdam. Symmetric successive overrelaxation methods for rank deficient linear systems. Applied Mathematics and Computation, 173(1):404–420, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003619.

See comments [ZW06b].

Darvishi:2006:NSK

[DKK06] M. T. Darvishi, S. Kheybari, and F. Khani. A numerical solution of the Korteweg–de Vries equation by pseudospectral method using Darvishi’s preconditionings. Applied Mathematics and Computation, 182(1):98–105, November 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306003171.

Duque:2009:DDR

[DKL09] Cosme Duque, Krisztina Kiss, and Marcos Lizana. On the dynamics of an ndimensional ratio-dependent predator–prey system with diffusion. Applied Mathematics and Computation, 208(1):98–105, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008515.

Duursma:2005:GKA

[DL05] Iwan Duursma and Hyang-Sook Lee. A group key agreement protocol from pairings. Applied Mathematics and Computation, 167(2):1451–1456, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030400579X.

Deng:2008:ISN

[DL08] Lei Deng and Qifei Liu. Iterative scheme for nonself generalized asymptotically quasi-nonexpansive mappings. Applied Mathematics and Computation, 205(1):317–324, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005900.

Dai:2009:AEL

[DL09] Tian-Shyr Dai and Yuh-Dauh Lyuu. Accurate and efficient lattice algorithms for Americanstyle Asian options with range bounds. Applied Mathematics and Computation, 209(2):238–253, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009417.

Duan:2008:HIP

[DLBT08] Qi Duan, Shilong Li, Fangxun Bao, and E. H. Twizell. Hermite interpolation by piecewise rational surface. Applied Mathematics and Computation, 198(1):59–72, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008661.

Ding:2008:ISG

[DLD08] Feng Ding, Peter X. Liu, and Jie Ding. Iterative solutions of the generalized Sylvester matrix equations by using the hierarchical identification principle. Applied Mathematics and Computation, 197(1):41–50, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307007515.

Duan:2008:LBM

[DLJ08] Yali Duan, Ruxun Liu, and Yanqun Jiang. Lattice Boltzmann model for the modified Burgers’ equation. Applied Mathematics and Computation, 202(2):489–497, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308000593.

Dai:2009:AHE

[DLL09] Zhengde Dai, Jun Liu, and Donglong Li. Applications of HTA and EHTA to YTSF Equation. Applied Mathematics and Computation, 207(2):360–364, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008205.

Daido:2009:NID

[DLLN09] Y. Daido, Y. Lei, J. J. Liu, and G. Nakamura.

Numerical implementations of dynamical probe method for non-stationary heat equation. Applied Mathematics and Computation, 211(2):510–521, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001106.

delaSen:2006:PDL

[dlS06a] M. de la Sen. Parameter dependent Lyapunov functions for robust stability of time-varying linear systems under point delays. Applied Mathematics and Computation, 179(2):612–621, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010672.

delaSen:2006:RSC

[dlS06b] M. de la Sen. Robust stabilization of a class of polytopic linear time-varying continuous systems under point delays and saturating controls. Applied Mathematics and Computation, 181(1):73–83, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000695.

delaSen:2006:RSP

[dlS06c] M. de la Sen. Robust stable pole-placement adaptive control of linear systems with multiestimation. Applied Mathematics and Computation, 172(2):1145–1174, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305002833.

Ding:2006:CAE

[DLS06d] Feng Ding, Peter X. Liu, and Yang Shi. Convergence analysis of estimation algorithms for dualrate stochastic systems. Applied Mathematics and Computation, 176(1):245– 261, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305007897.

delaSen:2007:PSR

[dlS07a] M. de la Sen. On positivity of singular regular linear time-delay timeinvariant systems subject to multiple internal and external incommensurate point delays. Applied Mathematics and Computation, 190(1):382–401, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000690.

delaSen:2007:QSS

[dlS07b] M. de la Sen. Quadratic stability and stabilization of switched dynamic systems with uncommensurate internal point delays. Applied Mathematics and Computation, 185(1):508–526, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008861.

Du:2008:MRP

[DLT08] Zengji Du, Xiaojie Lin, and Christopher C. Tisdell. A multiplicity result for p-Laplacian boundary value problems via critical points theorem. Applied Mathematics and Computation, 205(1):231–237, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005407.

Dercole:2005:DCB

[DM05] Fabio Dercole and Stefano Maggi. Detection and continuation of a border collision bifurcation in a forest fire model. Applied Mathematics and Computation, 168(1):623–635, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304006289.

Dehghan:2006:MSH

[DM06] Mehdi Dehghan and Akbar Mohebbi. Multigrid solution of high order discretisation for threedimensional biharmonic equation with Dirichlet boundary conditions of second kind. Applied Mathematics and Computation, 180(2):575–593, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306000464.

Das:2007:ABL

[DM07a] Barun Das and Manoranjan Maiti. An application of bi-level newsboy problem in two substitutable items under capital cost. Applied Mathematics and Computation, 190(1):410– 422, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000719.

Ding:2007:MEM

[DM07b] Jiu Ding and Lawrence R.

Mead. The maximum entropy method applied to stationary density computation. Applied Mathematics and Computation, 185(1):658–666, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009003.

Dil:2008:SAH

[DM08a] Ayhan Dil and István Mező. A symmetric algorithm for hyperharmonic and Fibonacci numbers. Applied Mathematics and Computation, 206(2):942–951, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007662.

Djebali:2008:SGF

[DM08b] Smaïl Djebali and Karima Mebarki. On the singular generalized Fisherlike equation with derivative depending nonlinearity. Applied Mathematics and Computation, 205(1):336–351, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005924.

Djolovic:2008:MTC

[DM08c] Ivana Djolović and Eberhard Malkowsky. Matrix transformations and compact operators on some new m th-order difference sequences. Applied Mathematics and Computation, 198(2):700–714, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009460.

Dehghan:2007:CPT

[DMA07a] Mehdi Dehghan and Seyedeh Mahboubeh Molavi-Arabshahi. Comparison of preconditioning techniques for solving linear systems arising from the fourth order approximation of the three-dimensional elliptic equation. Applied Mathematics and Computation, 184(2):156–172, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006928.

Dehghan:2007:SFF

[DMA07b] Mehdi Dehghan and Seyedeh Mahboubeh Molavi-Arabshahi. A simple form for the fourth order difference method for 3-D elliptic equations. Applied Mathematics and Computation, 184(2):589–598, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007326.

Djamah:2008:OLO

[DMDB08] T. Djamah, R. Mansouri, S. Djennoune, and M. Bettayeb. Optimal low order model identification of fractional dynamic systems. Applied Mathematics and Computation, 206(2):543–554, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003779.

Ding:2005:PBV

[DMH05] Wei Ding, Junrong Mi, and Maoan Han. Periodic boundary value problems for the first order impulsive functional differential equations. Applied Mathematics and Computation, 165(2):433–446, June 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003923.

Dehghan:2005:NIC

[DMJE05a] M. Dehghan, M. Masjed-Jamei, and M. R. Eslahchi.

On numerical improvement of closed Newton–Cotes quadrature rules. Applied Mathematics and Computation, 165(2):251–260, June 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003741.

Dehghan:2005:NIS

[DMJE05b] Mehdi Dehghan, M. Masjed-Jamei, and M. R. Eslahchi. On numerical improvement of the second kind of Gauss–Chebyshev quadrature rules. Applied Mathematics and Computation, 168(1):431–446, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304006162.

Dehghan:2005:SON

[DMJE05c] Mehdi Dehghan, M. Masjed-Jamei, and M. R. Eslahchi. The semi-open Newton– Cotes quadrature rule and its numerical improvement. Applied Mathematics and Computation, 171(2):1129– 1140, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dehghan:2006:NIO

[DMJE06a] Mehdi Dehghan, M. Masjed-Jamei, and M. R. Eslahchi. On numerical improvement of open Newton– Cotes quadrature rules. Applied Mathematics and Computation, 175(1):618– 627, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006351.

Dehghan:2006:WQR

[DMJE06b] Mehdi Dehghan, M. Masjed-Jamei, and M. R. Eslahchi. Weighted quadrature rules with weight function x^{−p}e^{−1/x} on [0, ∞). Applied Mathematics and Computation, 180(1):1–6, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010891.

Dattoli:2007:SID

[DMK07] G. Dattoli, M. Migliorati, and Subuhi Khan. Solutions of integro-differential equations and operational methods. Applied Mathematics and Computation, 186(1):302–308, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306009416.

Dolicanin:2007:LCC

[DMR07] Diana Dolićanin, Gradimir V. Milovanović, and Valery G. Romanovski. Linearizability conditions for a cubic system. Applied Mathematics and Computation, 190(1):937–945, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001233.

Dehghan:2006:CHO

[DMS06a] Mehdi Dehghan and Reza Mazrooei-Sebdani. The characteristics of a higherorder rational difference equation. Applied Mathematics and Computation, 182(1):521–528, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003596.

Dehghan:2006:DHO

[DMS06b] Mehdi Dehghan and Reza Mazrooei-Sebdani. Dynamics of a higher-order rational difference equation. Applied Mathematics and Computation, 178(2):345–354, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500980X.

Dehghan:2006:RSI

[DMS06c] Mehdi Dehghan and Reza Mazrooei-Sebdani. On the recursive sequence x_{n+1} = (α+βx_{n−k+1})/(A+Bx_{n−k+1}+ Cx_{n−2k+1}). Applied Mathematics and Computation, 178(2):273–286, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009732.

Dehghan:2007:DI

[DMS07] Mehdi Dehghan and Reza Mazrooei-Sebdani. Dynamics of x_{n+1} = x_{n−2k+1}/(x_{n−2k+1}+ αx_{n−2t}). Applied Mathematics and Computation, 185(1):464–472, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008800.

Das:2009:SLP

[DMS09] Pratulananda Das, Prasanta Malik, and Ekrem Savaş. On statistical limit points of double sequences. Applied Mathematics and Computation, 215(3):1030–1034, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309005967.

Duan:2008:STO

[DMZ08] Xian-Bao Duan, Yi-Chen Ma, and Rui Zhang. Shapetopology optimization of Stokes flow via variational level set method. Applied Mathematics and Computation, 202(1):200–209, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308000830.

Danumjaya:2006:OCS

[DN06] P. Danumjaya and A. K. Nandakumaran. Orthogonal cubic spline collocation method for the Cahn– Hilliard equation. Applied Mathematics and Computation, 182(2):1316–1329, November 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306004462.

Darehmiraki:2007:MSK

[DN07a] Majid Darehmiraki and Hasan Mishmast Nehi. Molecular solution to the 0–1 knapsack problem based on DNA computing. Applied Mathematics and Computation, 187(2):1033–1037, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012409.

Darehmiraki:2007:SBD

[DN07b] Majid Darehmiraki and Hasan Mishmast Nehi. A surface-based DNA algorithm for the solving binary knapsack problem. Applied Mathematics and Computation, 188(2):1991–1994, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015682.

Do:2007:AGA

[DN07c] Binh Quang Do and Lan Phuoc Nguyen. Application of a genetic algorithm to the fuel reload optimization for a research reactor. Applied Mathematics and Computation, 187(2):977–988, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012343.

Dartora:2005:NAA

[DNHF05] C. A. Dartora, K. Z.

Nobrega, and H. E. Hernandez-Figueroa. New analytical approximations for the Mathieu functions. Applied Mathematics and Computation, 165(2):447–458, June 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003935.

Deo:2008:ACN

[DNS08] Naokant Deo, Muhammad Aslam Noor, and M. A. Siddiqui. On approximation by a class of new Bernstein type operators. Applied Mathematics and Computation, 201(1–2):604–612, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000064.

dOnofrio:2005:BAE

[d’O05a] Alberto d’Onofrio. Biomathematical analysis and extension of the new class of epidemic models proposed by Satsuma et al. (2004). Applied Mathematics and Computation, 170(1):125–134, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008768.

dOnofrio:2005:PVA

[d’O05b] Alberto d’Onofrio. Periodically varying antiviral therapies: conditions for global stability of the virus free state. Applied Mathematics and Computation, 168(2):945–953, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006526.

dOnofrio:2005:VPN

[d’O05c] Alberto d’Onofrio. Vaccination policies and nonlinear force of infection: generalization of an observation by Alexander and Moghadas (2004). Applied Mathematics and Computation, 168(1):613–622, September 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304006277.

Dodds:2008:FSP

[Dod08]  Niall Dodds.  Further

spectral properties of uniformly elliptic operators that include a non-local term. Applied Mathematics and Computation, 197(1):317–327, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307007795.

Dogan:2005:AGM

[Dog05] Abdulkadir Dogan. Application of Galerkin’s method to equal width wave equation. Applied Mathematics and Computation, 160(1):65–76, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303009287.

Donmez:2006:NSR

[Dön06a] Orhan Dönmez. Numerical simulation of rotating accretion disk around the Schwarzschild black hole using GRH code. Applied Mathematics and Computation, 175(2):902–922, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006521.

Donmez:2006:SDS

[Dön06b] Orhan Dönmez. Solving 1-D special relativistic hydrodynamics (SRH) equations using different numerical methods and results from different test problems. Applied Mathematics and Computation, 181(1):256–270, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000956.

Demir:2008:OLN

[DOU08a] Ali Demir, Nese Omur, and Yucel Turker Ulutas. Optimization by k-Lucas numbers. Applied Mathematics and Computation, 197(1):366–371, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007849.

Demir:2008:PFS

[DOU08b] Ali Demir, Nese Omur, and Yucel Turker Ulutas. Parametrized Fibonacci search method with k-Lucas numbers. Applied Mathematics and Computation, 198(1):355–360, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700896X.

Castanho:2008:FES

[dPCdBYV08] Maria José de Paula Castanho, Laécio Carvalho de Barros, Akebo Yamakami, and Laércio Luis Vendite. Fuzzy expert system: an example in prostate cancer. Applied Mathematics and Computation, 202(1):78–85, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011411.

Diao:2006:CNO

[DQW06] Huaian Diao, Mei Qin, and Yimin Wei. Condition numbers for the outer inverse and constrained singular linear system. Applied Mathematics and Computation, 174(1): 588–612, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030500487X.

Ding:2006:MPL

[DR06] J. Ding and N. H. Rhee. A modified piecewise linear Markov approximation of Markov operators. Applied Mathematics and Computation, 174(1):236–251, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305004534.

delRey:2005:SSS

[dRMS05] A. Martín del Rey, J. Pereira Mateus, and G. Rodríguez Sánchez. A secret sharing scheme based on cellular automata. Applied Mathematics and Computation, 170(2):1356–1364, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000718.

Dumett:2008:DEB

[DRS^{+}08] M. A. Dumett, I. G. Rosen, J. Sabat, A. Shaman, L. Tempelman, C. Wang, and R. M. Swift. Deconvolving an estimate of breath measured blood alcohol concentration from biosensor collected transdermal ethanol data. Applied Mathematics and Computation, 196(2):724– 743, March 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307007254.

Daoud:2005:SAD

[DS05a] Daoud S. Daoud and D. Subasi. A splitting up algorithm for the determination of the control parameter in multi dimensional parabolic problem. Applied Mathematics and Computation, 166(3):584–595, July 26, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030400459X.

Dulov:2005:NML

[DS05b] Eugene V. Dulov and Alexander V. Sinitsyn. A numerical modelling of the limit problem for the magnetically noninsulated diode. Applied Mathematics and Computation, 162(1):115–154, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030400075X.

Dominguez:2006:BFO

[DS06a] Víctor Domínguez and Francisco-Javier Sayas. A BEM–FEM overlapping algorithm for the Stokes equation. Applied Mathematics and Computation, 182(1):691–710, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600378X.

Dreo:2006:ACA

[DS06b] J. Dréo and P. Siarry. An ant colony algorithm aimed at dynamic continuous optimization. Applied Mathematics and Computation, 181(1):457–467, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001159.

Dag:2005:BSG

[DSB05] İdris Dağ, Bülent Saka, and Ahmet Boz. B-spline Galerkin methods for numerical solutions of the Burgers’ equation. Applied Mathematics and Computation, 166(3):506–522, July 26, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004540.

deSouza:2005:NDA

[dSF05] Jessé Rebello de Souza, Jr.

and Claudio Gomes Fernandes. Nonlinear dynamics of an archetypal model of ships motions in tandem. Applied Mathematics and Computation, 164(2):649–665, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009166.

Diatta:2008:ASS

[DSK08] B. Diatta, C. Wafo Soh, and C. M. Khalique. Approximate symmetries and solutions of the hyperbolic heat equation. Applied Mathematics and Computation, 205(1):263–272, November 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308004876.

Deep:2009:RCG

[DSKM09] Kusum Deep, Krishna Pratap Singh, M. L. Kansal, and C. Mohan. A real coded genetic algorithm for solving integer and mixed integer optimization problems. Applied Mathematics and Computation, 212(2):505–518, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001830.

deSouza:2008:EFU

[dSLV^{+}08] A. C. Zambroni de Souza, Rafael C. Leme, Luiz Frederico B. Vasconcelos, B. Isaias Lima Lopes, and Yuri C. da Silva Ribeiro. Energy function and unstable solutions by the means of an augmented Jacobian. Applied Mathematics and Computation, 206(1):154–163, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006437.

Dalpasquale:2008:FBD

[DSMK08] Valdecir Antoninho Dalpasquale, Décio Sperandio, Luiz Henry Monken e Silva, and Evandro Kolling. Fixed-bed drying simulation of agricultural products using a new backward finite difference scheme. Applied Mathematics and Computation, 200(2):590– 595, July 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011381.

Dernek:2008:SPG

[DSY08] Neşe Dernek, H. M. Srivastava, and Osman Yürekli. Some Parseval–Goldstein type identities involving the F_{S,∈}-transform, the F_{C,∈}-transform and the P_{△}-transform and their applications. Applied Mathematics and Computation, 202(1):327–337, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800115X.

Solo:2005:KEI

[dSYGM^{+}05] C. Garrido del Solo, J. M.

Yago, M. García-Moreno, B. H. Havsteen, F. García-Cánovas, and R. Varón. The kinetic effect of the inactivation of the enzyme– substrate complex in an enzymatic reaction with slowbinding inhibition. Applied Mathematics and Computation, 171(1):450–463, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001086.

deSampaio:2008:GOT

[dSYY08] Raimundo J. B. de Sampaio, S. Yuan, and J. Y. Yuan. Generalization: One technique of computational and applied mathematical methodology. Applied Mathematics and Computation, 200(2):574–582, July 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011368.

Dattoli:2007:OMD

[DSZ07] G. Dattoli, H. M. Srivastava, and K. Zhukovsky. Operational methods and differential equations with applications to initial-value problems. Applied Mathematics and Computation, 184(2):979–1001, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007880.

Ding:2008:NMF

[DSZ08] Shuangshuang Ding, Weijun Sun, and Danchen Zhu. A new method for finding traveling wave solutions of the generalized KdV equation with variable coefficient. Applied Mathematics and Computation, 199(1):268–274, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009940.

Demirel:2006:OQA

[DT06] Nihan Çetin Demirel and M. Duran Toksarı. Optimization of the quadratic assignment problem using an ant colony algorithm. Applied Mathematics and Computation, 183(1):427–435, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005753.

Deep:2007:NCO

[DT07a] Kusum Deep and Manoj Thakur. A new crossover operator for real coded genetic algorithms. Applied Mathematics and Computation, 188(1):895–911, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014287.

Deep:2007:NMO

[DT07b] Kusum Deep and Manoj Thakur. A new mutation operator for real coded genetic algorithms. Applied Mathematics and Computation, 193(1):211–230, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307003918.

Deo:2008:SPD

[DT08] Satya Deo and Ashish Tiwari. On the solution of a partial differential equation representing irrotational flow in bispherical polar coordinates. Applied Mathematics and Computation, 205(1):475–477, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006127.

Dimitriou:2009:PCE

[DT09] V. A. Dimitriou and N. Tsantas. Prospective control in an enhanced manpower planning model. Applied Mathematics and Computation, 215(3):995–1014, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005906.

Danca:2008:SSS

[DTC08] Marius-F. Danca, Wallace K. S. Tang, and Guanrong Chen. A switching scheme for synthesizing attractors of dissipative chaotic systems. Applied Mathematics and Computation, 201(1–2):650–667, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800012X.

Du:2005:IMC

[Du05] Kui Du. The iterative methods for computing the polar decomposition of rank-deficient matrix. Applied Mathematics and Computation, 162(1):95– 102, March 4, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304000736.

Du:2006:NPC

[Du06]  Kui Du.  A note on

properties and computations of matrix pseudospectra. Applied Mathematics and Computation, 174(2):1007–1019, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030500528X.

Du:2007:SPT

[Du07] Zengji Du. Singularly perturbed third-order boundary value problem for nonlinear systems. Applied Mathematics and Computation, 189(1):869–877, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016651.

Duan:2006:MGM

[Dua06]  Yong Duan.  Mesh-

less Galerkin method using radial basis functions based on domain decomposition. Applied Mathematics and Computation, 179(2):750–762, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010854.

Devadoss:2006:MMR

[DV06] D. E. Devadoss and V. A. Volpert. Mathematical modeling of radially propagating polymerization waves with the gel effect. Applied Mathematics and Computation, 172(2):1036–1053, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002766.

Deissenberg:2008:EMP

[DvdHD08] Christophe Deissenberg, Sander van der Hoog, and Herbert Dawid. EURACE: a massively parallel agentbased model of the European economy. Applied Mathematics and Computation, 204(2):541–552, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003019.

Dai:2009:NSW

[DW09a] Chao-Qing Dai and Yue-Yue Wang. The novel solitary wave structures and interactions in the (2 + 1)-dimensional Korteweg– de Vries system. Applied Mathematics and Computation, 208(2):453–461, February 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S009630030800917X.

Dieci:2009:SAC

[DW09b] Roberto Dieci and Frank Westerhoff. Stability analysis of a cobweb model with market interactions. Applied Mathematics and Computation, 215(6):2011–2023, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309007097.

Ding:2009:SAS

[DW09c] Xie Ping Ding and Zhong Bao Wang. Sensitivity analysis for a system of parametric generalized mixed quasi-variational inclusions involving (Kη)-monotone mappings. Applied Mathematics and Computation, 214(1):318–327, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309003075.

Dong:2009:LAC

[DW09d] Huan-He Dong and Xin-Zeng Wang. A Lie algebra containing four parameters for the generalized Dirac hierarchy. Applied Mathematics and Computation, 215(2):459–463, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900486X.

Wang:2005:GSE

[dWqL05] La di Wang and Jian quan Li. Global stability of an epidemic model with nonlinear incidence rate and differential infectivity. Applied Mathematics and Computation, 161(3):769–778, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000281.

Duan:2005:NRI

[DWT05a] Qi Duan, Liqiu Wang, and E. H. Twizell. A new C^{2} rational interpolation based on function values and constrained control of the interpolant curves. Applied Mathematics and Computation, 161(1):311–322, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012530.

Duan:2005:NWR

[DWT05b] Qi Duan, Liqiu Wang, and E. H. Twizell. A new weighted rational cubic interpolation and its approximation. Applied Mathematics and Computation, 168(2):990–1003, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006575.

Ding:2005:NME

[DwWmW05] Yong Ding, Kwok wo Wong, and Yu min Wang. A w-NNAF method for the efficient computation of scalar multiplication in elliptic curve cryptography. Applied Mathematics and Computation, 167(1):81–93, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004758.

Du:2007:LSB

[DWZ07] Ji-Xiang Du, Xiao-Feng Wang, and Guo-Jun Zhang. Leaf shape based plant species recognition. Applied Mathematics and Computation, 185(2):883– 893, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030600806X.

Du:2006:FBA

[DX06] Lili Du and Zhaoyin Xiang.

A further blow-up analysis for a localized porous medium equation. Applied Mathematics and Computation, 179(1):200–208, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010301.

Danfu:2007:NSI

[DX07a] Han Danfu and Shang Xufeng. Numerical solution of integro-differential equations by using CAS wavelet operational matrix of integration. Applied Mathematics and Computation, 194(2):460–466, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700495X.

Ding:2007:FDW

[DX07b] Feng Ding and Yongsong Xiao. A finite-data-window least squares algorithm with a forgetting factor for dynamical modeling. Applied Mathematics and Computation, 186(1):184–192, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009283.

Deng:2007:GCF

[DXH07a] Kang Deng, Zhiguang Xiong, and Yunqing Huang. The Galerkin continuous finite element method for delay-differential equation with a variable term. Applied Mathematics and Computation, 186(2):1488– 1496, March 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306010745.

Ding:2007:APB

[DXH07b] Wei Ding, Yepeng Xing, and Maoan Han. Antiperiodic boundary value problems for first order impulsive functional differential equations. Applied Mathematics and Computation, 186(1):45– 53, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009155.

Ding:2008:ABS

[DXH08] Yongsheng Ding, Min Xu, and Liangjian Hu. Asymptotic behavior and stability of a stochastic model for AIDS transmission. Applied Mathematics and Computation, 204(1):99–108, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004621.

Dai:2006:WBD

[DY06] Dao-Qing Dai and P. C.

Yuen. Wavelet based discriminant analysis for face recognition. Applied Mathematics and Computation, 175(1):307–318, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006090.

Ding:2007:EPS

[DY07] Jiu Ding and Guangming Yao. The eigenvalue problem of a specially updated matrix. Applied Mathematics and Computation, 185(1):415–420, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008769.

Dyki:2009:BVP

[Dyk09] Agnieszka Dyki. Boundary value problems for differential equations with deviated arguments which depend on the unknown solution. Applied Mathematics and Computation, 215(5):1895–1899, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006900.

Duan:2006:ADT

[DZ06] Fujian Duan and Kecun Zhang. An algorithm of diagonal transformation for Perron root of nonnegative irreducible matrices. Applied Mathematics and Computation, 175(1):762– 772, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006430.

Du:2007:NSC

[DZ07a] Xinsheng Du and Zengqin Zhao. A necessary and sufficient condition for the existence of positive solutions to singular sublinear three-point boundary value problems. Applied Mathematics and Computation, 186(1):404–413, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009490.

Du:2007:ETP

[DZ07b] Xinsheng Du and Zengqin Zhao. On existence theorems of positive solutions to nonlinear singular differential equations. Applied Mathematics and Computation, 190(1):542– 552, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000872.

Dimitrova:2008:FPD

[DZ08a] Elena S. Dimitrova and Andrew Zardecki. Forecasting polynomial dynamics. Applied Mathematics and Computation, 204(1):233–239, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004992.

Ding:2008:PSM

[DZ08b] Hengfei Ding and Yuxin Zhang. Parameters spline methods for the solution of hyperbolic equations. Applied Mathematics and Computation, 204(2):938–941, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005833.

Ding:2008:CPS

[DZ08c] Jiu Ding and Aihui Zhou.

Characteristic polynomials of some perturbed matrices. Applied Mathematics and Computation, 199 (2):631–636, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307010429.

Ding:2008:STP

[DZ08d] Jiu Ding and Aihui Zhou. A spectrum theorem for perturbed bounded linear operators. Applied Mathematics and Computation, 201(1–2):723–728, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000398.

Du:2008:HLA

[DZ08e] Ji-Xiang Du and Chuan-Min Zhai. A hybrid learning algorithm combined with generalized RLS approach for radial basis function neural networks. Applied Mathematics and Computation, 205(2):908–915, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003706.

Du:2008:EUP

[DZ08f] Xinsheng Du and Zengqin Zhao. Existence and uniqueness of positive solutions to a class of singular m-point boundary value problems. Applied Mathematics and Computation, 198(2):487–493, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009149.

Ding:2009:NUS

[DZ09] Hengfei Ding and Yuxin Zhang. A new unconditionally stable compact difference scheme of O(τ^{2} + h^{4}) for the 1D linear hyperbolic equation. Applied Mathematics and Computation, 207(1):236–241, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007868.

Du:2006:CAL

[DZG06] Xuewu Du, Liansheng Zhang, and Yuelin Gao. A class of augmented Lagrangians for equality constraints in nonlinear programming problems. Applied Mathematics and Computation, 172(1):644–663, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305002298.

Dziok:2007:CCD

[Dzi07] J. Dziok. On the convex combination of the Dziok–Srivastava operator. Applied Mathematics and Computation, 188(2):1214–1220, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014652.

Dziok:2009:AEP

[Dzi09] J. Dziok. Applications of extreme points to distortion estimates. Applied Mathematics and Computation, 215(1):71– 77, September 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300309003841.

Duan:2006:BRI

[DZT06] Qi Duan, Yunfeng Zhang, and E. H. Twizell. A bivariate rational interpolation and the properties. Applied Mathematics and Computation, 179(1):190–199, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010295.

Dai:2007:ARA

[DZZ07] Fengzhi Dai, Jinbin Zhao, and Huailin Zhao. The application of resonance algorithm for image segmentation. Applied Mathematics and Computation, 194(2):453–459, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004948.

Elhanbaly:2006:NAA

[EA06] A. Elhanbaly and M. A.

Abdou. New application of Adomian decomposition method on Fokker–Planck equation. Applied Mathematics and Computation, 182(1):301–312, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003389.

El-Azab:2007:ASN

[EA07a] M. S. El-Azab. An approximation scheme for a nonlinear diffusion Fisher’s equation. Applied Mathematics and Computation, 186(1): 579–588, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306009696.

El-Azab:2007:SNT

[EA07b] M. S. El-Azab. Solution of nonlinear transport– diffusion problems by linearization. Applied Mathematics and Computation, 192(1):205–215, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307003232.

El-Ashwah:2008:SPC

[EA08a] R. M. El-Ashwah. Some properties of certain subclasses of meromorphically multivalent functions. Applied Mathematics and Computation, 204(2):824–832, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005675.

Esfahanian:2008:JNM

[EA08b] Vahid Esfahanian and Pooria Akbarzadeh. The Jameson’s numerical method for solving the incompressible viscous and inviscid flows by means of artificial compressibility and preconditioning method. Applied Mathematics and Computation, 206(2):651–661, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003883.

El-Azab:2007:NAS

[EAEG07] M. S. El-Azab and Mohamed El-Gamel. A numerical algorithm for the solution of telegraph equations. Applied Mathematics and Computation, 190 (1):757–764, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307001105.

Egrioglu:2008:NMS

[EAG08] Erol Eğrioğlu, Çağdaş Hakan Aladağ, and Süleyman Günay. A new model selection strategy in artificial neural networks. Applied Mathematics and Computation, 195(2):591–597, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005887.

Ebrahimipour:2007:GPA

[EARS07] V. Ebrahimipour, A. Azadeh, K. Rezaie, and K. Suzuki. A GA–PCA approach for power sector performance ranking based on machine productivity. Applied Mathematics and Computation, 186(2):1205–1215, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010460.

Effati:2005:NNNb

[EB05a] S. Effati and M. Baymani. A new nonlinear neural network for solving convex nonlinear programming problems. Applied Mathematics and Computation, 168(2):1370–1379, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304007350.

Effati:2005:NNNa

[EB05b] S. Effati and M. Baymani. A new nonlinear neural network for solving quadratic programming problems. Applied Mathematics and Computation, 165(3):719–729, June 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004138.

El-Bary:2005:CTF

[EB05c] A. A. El-Bary. Computational treatment of free convection effects on perfectly conducting viscoelastic fluid. Applied Mathematics and Computation, 170(2):801–820, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000251.

El-Bary:2005:ESP

[EB05d] A. A. El-Bary. Exponential solution of a problem of two-dimensional motion of micropolar fluid in a half-plane. Applied Mathematics and Computation, 165(1):81– 93, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003637.

El-Bassiouny:2005:ENL

[EB05e] A. F. El-Bassiouny. Effect of non-linearities in elastomeric material dampers on torsional oscillation control. Applied Mathematics and Computation, 162(2):835–854, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001341.

El-Borai:2008:ESS

[EB08] Mahmoud M. El-Borai. Exact solutions for some nonlinear fractional parabolic partial differential equations. Applied Mathematics and Computation, 206(1):150–153, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006425.

Erbe:2009:ONS

[EBP09] Lynn Erbe, Jia Baoguo, and Allan Peterson. Oscillation and nonoscillation of solutions of second order linear dynamic equations with integrable coefficients on time scales. Applied Mathematics and Computation, 215(5):1868–1885, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006882.

El-Borai:2005:NMS

[EBsENMA05] Mahmoud M. El-Borai, Khairia El said El-Nadi, Osama L. Mostafa, and Hamdy M. Ahmed. Numerical methods for some nonlinear stochastic differential equations. Applied Mathematics and Computation, 168(1):65– 75, September 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304005867.

Ece:2006:FCF

[Ece06] Mehmet Cem Ece. Free convection flow about a vertical spinning cone under a magnetic field. Applied Mathematics and Computation, 179(1):231– 242, August 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305010325.

Eslahchi:2005:NIF

[EDMJ05a] M. R. Eslahchi, M. Dehghan, and M. Masjed-Jamei. On numerical improvement of the first kind Gauss–Chebyshev quadrature rules. Applied Mathematics and Computation, 165(1):5–21, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400356X.

Eslahchi:2005:ECQ

[EDMJ05b] M. R. Eslahchi, Mehdi Dehghan, and M. Masjed-Jamei. The equal coefficients quadrature rules and their numerical improvement. Applied Mathematics and Computation, 171(2): 1331–1351, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Eslahchi:2005:FKC

[EDMJ05c] M. R. Eslahchi, Mehdi Dehghan, and M. Masjed-Jamei. The first kind Chebyshev–Newton–Cotes quadrature rules (closed type) and its numerical improvement. Applied Mathematics and Computation, 168(1):479–495, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304006198.

Encinas:2007:IRE

[EdR07] L. Hernández Encinas and A. Martín del Rey. Inverse rules of ECA with rule number 150. Applied Mathematics and Computation, 189(2):1782–1786, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017723.

Erkus:2006:SAC

[EDS06] Esra Erkuş, Oktay Duman, and H. M. Srivastava.

Statistical approximation of certain positive linear operators constructed by means of the Chan– Chyan–Srivastava polynomials. Applied Mathematics and Computation, 182(1):213–222, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003304.

El-Enna:2006:ATR

[EEAES06] Ahmed A. El-Enna, M. K. M.

Ahmed, and F. A. Abd El-Salam. Analytical treatment of the relativistic and solar radiation pressure effects on an artificial satellite. Applied Mathematics and Computation, 175(2):1525–1542, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007113.

Eldabe:2005:CFD

[EEEE05] N. T. Eldabe, E. F. Elshehawey, Elsayed M. E. Elbarbary, and Nasser S. Elgazery. Chebyshev finite difference method for MHD flow of a micropolar fluid past a stretching sheet with heat transfer. Applied Mathematics and Computation, 160(2):437–450, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030301172X.

See comment [Pan12].

Elshehawey:2006:PTA

[EEEE06] E. F. Elshehawey, N. T. Eldabe, E. M. Elghazy, and A. Ebaid. Peristaltic transport in an asymmetric channel through a porous medium. Applied Mathematics and Computation, 182(1):140–150, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003201.

Elwakil:2005:MET

[EELZS05] S. A. Elwakil, S. K. El-Labany, M. A. Zahran, and R. Sabry. Modified extended tanh-function method and its applications to nonlinear equations. Applied Mathematics and Computation, 161(2):403–412, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303012943.

Elshehawey:2005:PTC

[EESESD05] E. F. Elshehawey, Abd El-Rahman El-Saman, Mostafa El-Shahed, and Mohamed Dagher. Peristaltic transport of a compressible viscous liquid through a tapered pore. Applied Mathematics and Computation, 169(1):526–543, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007921.

Eldabe:2004:TDD

[EESF04] N. T. Eldabe, A. G. El-Saka, and Ashraf Fouad. Thermal-diffusion and diffusionthermo effects on mixed free-forced convection and mass transfer boundary layer flow for non-Newtonian fluid with temperature dependent viscosity. Applied Mathematics and Computation, 152(3):867–883, May 13, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See comment [Pan08c].

Eisinberg:2005:VSG

[EF05] A. Eisinberg and G. Fedele.

Vandermonde systems on Gauss–Lobatto Chebyshev nodes. Applied Mathematics and Computation, 170(1):633–647, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000056.

Eisinberg:2006:IVM

[EF06] A. Eisinberg and G. Fedele. On the inversion of the Vandermonde matrix. Applied Mathematics and Computation, 174(2):1384– 1397, March 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305005576.

Eisinberg:2007:AFP

[EF07] A. Eisinberg and G. Fedele. Accurate floating-point summation: a new approach. Applied Mathematics and Computation, 189 (1):410–424, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306016158.

Ebrahimi:2008:BAN

[EFE08] Morteza Ebrahimi, Rahman Farnoosh, and Somayeh Ebrahimi. Biological applications and numerical solution based on Monte Carlo method for a two-dimensional parabolic inverse problem. Applied Mathematics and Computation, 204(1):1–9, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010879.

Eisinberg:2006:VSE

[EFI06] A. Eisinberg, G. Fedele, and C. Imbrogno. Vandermonde systems on equidistant nodes in [0, 1]: accurate computation. Applied Mathematics and Computation, 172(2):971–984, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002523.

Efimova:2009:ESS

[Efi09] Olga Yu. Efimova. Exact solutions of systems describing interacting populations with low critical densities. Applied Mathematics and Computation, 208(1):134–143, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008710.

El-Gohary:2005:OCS

[EG05]  A. El-Gohary.  Opti-

mal control of stochastic lattice of prey–predator models. Applied Mathematics and Computation, 160(1):15–28, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303006684.

El-Gamel:2006:WGM

[EG06a] Mohamed El-Gamel. A Wavelet-Galerkin method for a singularly perturbed convection-dominated diffusion equation. Applied Mathematics and Computation, 181(2):1635–1644, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306002955.

Eren:2006:BFS

[EG06b] Tamer Eren and Ertan Güner. A bicriteria flowshop scheduling problem with setup times. Applied Mathematics and Computation, 183(2):1292–1300, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600676X.

See comments [XY11].

Eren:2006:BSS

[EG06c] Tamer Eren and Ertan Güner. A bicriteria scheduling with sequence-dependent setup times. Applied Mathematics and Computation, 179(1):378–385, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010453.

El-Gamel:2007:CSO

[EG07a] Mohamed El-Gamel. Comparison of the solutions obtained by Adomian decomposition and wavelet-Galerkin methods of boundaryvalue problems. Applied Mathematics and Computation, 186(1):652–664, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009775.

El-Gamel:2007:SNS

[EG07b] Mohamed El-Gamel. Sinc and the numerical solution of fifth-order boundary value problems. Applied Mathematics and Computation, 187(2):1417–1433, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012793.

Erel:2007:COS

[EG07c] Erdal Erel and Jay B.

Ghosh. Customer order scheduling on a single machine with family setup times: Complexity and algorithms. Applied Mathematics and Computation, 185(1):11–18, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007983.

El-Gebeily:2006:COC

[EGE06] M. A. El-Gebeily and M. B. M. Elgindi. Convergence and the optimal choice of the relaxation parameter for a class of iterative methods. Applied Mathematics and Computation, 174(1):353–364, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004637.

El-Gebeily:2006:SSF

[EGFAE06] Mohamed A. El-Gebeily, Faisal Fairag, Rajai Alassar, and M. B. M. Elgindi. H^{2} solutions for the stream function and vorticity formulation of the Navier– Stokes equations. Applied Mathematics and Computation, 177(2):686–696, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009495.

Effati:2007:APN

[EGN07] S. Effati, A. Ghomashi, and A. R. Nazemi. Application of projection neural network in solving convex programming problems. Applied Mathematics and Computation, 188(2):1103–1114, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013774.

El-Gohary:2005:EPT

[EGS05] Awad I. El-Gohary and Ammar M. Sarhan. Estimations of the parameters in a three nonindependent component series system subjected to sources of shocks. Applied Mathematics and Computation, 160(1):29–40, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303006891.

Elliotis:2005:SLP

[EGX05] Miltiades Elliotis, Georgios Georgiou, and Christos Xenophontos. Solving Laplacian problems with boundary singularities: a comparison of a singular function boundary integral method with the p/ hp version of the finite element method. Applied Mathematics and Computation, 169(1):485–499, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400788X.

Ebrahem:2006:NPA

[EH06] Mohammed Al-Haj Ebrahem and James J. Higgins. Non-parametric analysis of a proportional wearout model for accelerated degradation data. Applied Mathematics and Computation, 174(1):365–373, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004649.

Elgazzar:2008:MAF

[EH08a] E. Ahmed A. S. Elgazzar and A. S. Hegazi. Modeling the avian flu, lessons form complex adaptive systems in biology. Applied Mathematics and Computation, 195(1):351–354, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005140.

Elgindy:2008:NRL

[EH08b] K. T. Elgindy and Abdel-Rahman Hedar. A new robust line search technique based on Chebyshev polynomials. Applied Mathematics and Computation, 206(2):853–866, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005985.

Elouafi:2009:NRA

[EH09a] Mohamed Elouafi and Ahmed Driss Aiat Hadj. A new recursive algorithm for inverting Hessenberg matrices. Applied Mathematics and Computation, 214(2):497–499, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003452.

Elouafi:2009:PIT

[EH09b] Mohamed Elouafi and Ahmed Driss Aiat Hadj. On the powers and the inverse of a tridiagonal matrix. Applied Mathematics and Computation, 211(1):137–141, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000502.

Elouafi:2009:RFC

[EH09c] Mohamed Elouafi and Ahmed Driss Aiat Hadj. A recursion formula for the characteristic polynomial of Hessenberg matrices. Applied Mathematics and Computation, 208(1):177–179, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008813.

Estrada:2009:CGC

[EH09d] Ernesto Estrada and Naomichi Hatano. Communicability graph and community structures in complex networks. Applied Mathematics and Computation, 214(2):500–511, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309003464.

El-Hawary:2005:NSG

[EHAR05] H. M. El-Hawary and E. O. Abdel-Rahman. Numerical solution of the generalized Burger’s equation via spectral/spline methods. Applied Mathematics and Computation, 170(1):267–279, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009245.

Ervin:2008:TPD

[EHL08] Vincent J. Ervin, Jason S. Howell, and Hyesuk Lee. A two-parameter defect-correction method for computation of steadystate viscoelastic fluid flow. Applied Mathematics and Computation, 196(2):818– 834, March 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307007357.

Effati:2007:NNM

[EHN07] S. Effati, M. Hoseini, and A. R. Nazemi. A new numerical method by revised measure theory for solving the nonlinear initial value problems. Applied Mathematics and Computation, 186(1):780–788, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009982.

Erbe:2008:OCN

[EHP08] Lynn Erbe, Taher S. Hassan, and Allan Peterson. Oscillation criteria for nonlinear damped dynamic equations on time scales. Applied Mathematics and Computation, 203(1):343–357, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002580.

Ezquerro:2009:NTM

[EHR09] J. A. Ezquerro, M. A. Hernández, and N. Romero. Newton-type methods of high order and domains of semilocal and global convergence. Applied Mathematics and Computation, 214(1):142–154, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002884.

El-Hawary:2005:RQS

[EHZES05] H. M. El-Hawary, E. A. Zanaty, and E. El-Sanousy. RETRACTED: Quartic spline collocation methods for elliptic partial differential equations. Applied Mathematics and Computation, 168(1):198–221, September 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/

S0096300304005995. See

retraction notice [Ano07c].

Elsayed:2009:MTM

[EI09] E. M. Elsayed and Bratislav D.

Iričanin. On a max-type and a min-type difference equation. Applied Mathematics and Computation, 215(2):608–614, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005384.

Eid:2005:HOI

[Eid05]  R. Eid.  Higher order

isoparametric finite element solution of Stokes flow. Applied Mathematics and Computation, 162(3):1083–1101, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001511.

Effati:2007:NNNa

[EJ07a] S. Effati and M. Jafarzadeh. A new nonlinear neural network for solving a class of constrained parametric optimization problems. Applied Mathematics and Computation, 186(1):814–819, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010034.

Effati:2007:NNNb

[EJ07b] S. Effati and M. Jafarzadeh. Nonlinear neural networks for solving the shortest path problem. Applied Mathematics and Computation, 189(1):567– 574, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016304.

Effati:2006:SOC

[EJER06] S. Effati, M. Janfada, M. Esmaeili, and H. Roohparvar. Solving of optimal control problem of parabolic PDEs in exploitation of oil by iterative dynamic programming. Applied Mathematics and Computation, 181(2):1505–1512, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306002803.

Esen:2006:ALG

[EK06] A. Esen and S. Kutluay. Application of a lumped Galerkin method to the regularized long wave equation. Applied Mathematics and Computation, 174(2):833–845, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005126.

El-kady:2005:EEP

[EkES05] M. El-kady and A. Abd El-Shafy. Error estimation problem of nonlinear integral equations in the space L_{p}(p ≥ 1). Applied Mathematics and Computation, 168(2):1251–1257, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400699X.

Ebiefung:2008:DPG

[EKM08] Aniekan Ebiefung, Michael Kostreva, and Ishita Majumdar. Disjunctive programming and the generalized Leontief input– output model. Applied Mathematics and Computation, 198(2):551–558, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009216.

El-Keib:2008:MSA

[EKS08] Abdurrahim El-Keib and Ibrahim Sadek. Modeling, simulation and applied optimization. Applied Mathematics and Computation, 206(2):509, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004608.

El-Latif:2005:AST

[EL05a] G. M. Abd El-Latif. An approximate solution treatment for non-linear problems. Applied Mathematics and Computation, 169(1):563–575, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304007945.

El-Latif:2005:HTP

[EL05b] G. M. Abd El-Latif. A homotopy technique and a perturbation technique for non-linear problems. Applied Mathematics and Computation, 169(1):576–588, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304007957.

Elbarbary:2005:CFD

[Elb05] Elsayed M. E. Elbarbary. Chebyshev finite difference method for the solution of boundary-layer equations. Applied Mathematics and Computation, 160(2):487–498, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303011755.

Elgindy:2009:GHO

[Elg09]  K. T. Elgindy.  Gen-

eration of higher order pseudospectral integration matrices. Applied Mathematics and Computation, 209(2):153–161, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006000.

Elizalde:2007:SAC

[Eli07] E. Elizalde. Some analytic continuations of the Barnes zeta function in two and higher dimensions. Applied Mathematics and Computation, 187(1):141– 152, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011453.

ElShafei:2006:ISM

[ElS06] Mervat Mohamed Kamel ElShafei. Interactive stability of multiobjective integer nonlinear programming problems. Applied Mathematics and Computation, 176(1):230–236, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305007873.

El-Mikkawy:2003:NTT

[EM03] Moawwad El-Mikkawy. A note on a three-term recurrence for a tridiagonal matrix. Applied Mathematics and Computation, 139(2– 3):503–511, July 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See comments [Sal06c].

El-Mikkawy:2005:NCA

[EM05a] M. E. A. El-Mikkawy. A new computational algorithm for solving periodic tri-diagonal linear systems. Applied Mathematics and Computation, 161(2):691–696, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304000220.

El-Mikkawy:2005:NRS

[EM05b] Moawwad El-Mikkawy. A note on the reducibility of special infinite series. Applied Mathematics and Computation, 162(1):311– 315, March 4, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304000967.

Effati:2006:CSC

[EM06] S. Effati and M. M.

Moghadam. Conversion of some classes of fractional programming to secondorder cone programming and solving it by potential reduction interior point method. Applied Mathematics and Computation, 181(1):563–578, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001287.

Erturk:2007:CNM

[EM07] Vedat Suat Ertürk and Shaher Momani. Comparing numerical methods for solving fourthorder boundary value problems. Applied Mathematics and Computation, 188(2):1963–1968, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015645.

El-Mikkawy:2008:FRA

[EM08] Moawwad E. A. El-Mikkawy. A fast and reliable algorithm for evaluating n th order pentadiagonal determinants. Applied Mathematics and Computation, 202(1):210–215, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308000842.

Erbe:2009:SNO

[EM09] Lynn Erbe and Raziye Mert. Some new oscillation results for a nonlinear dynamic system on time scales. Applied Mathematics and Computation, 215(7):2405–2412, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309007656.

El-Mistikawy:2011:CGS

[EM11] Tarek M. A. El-Mistikawy. Comment on “Group solution of a time dependent chemical convective process” by M. M. Kassem and A. S. Rashed, Applied Mathematics and Computation, 215 (2009) 1671– 1684. Applied Mathematics and Computation, 217(12): 5999–6003, February 15, 2011. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). See [KR09].

El-Misiery:2006:FME

[EMA06a] A. E. M. El-Misiery and E. Ahmed. On a fractional model for earthquakes. Applied Mathematics and Computation, 178(2):207–211, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008647.

Emam:2006:FAB

[Ema06b] O. E. Emam. A fuzzy approach for bi-level integer non-linear programming problem. Applied Mathematics and Computation, 172(1):62–71, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001864.

El-Mikkawy:2005:CHI

[EMC05] M. E. A. El-Mikkawy and G.-S. Cheon. Combinatorial and hypergeometric identities via the Legendre polynomials — A computational approach. Applied Mathematics and Computation, 166(1):181– 195, July 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004291.

El-Mesiry:2005:NMM

[EMESES05] A. E. M. El-Mesiry, A. M. A. El-Sayed, and H. A. A. El-Saka. Numerical methods for multi-term fractional (arbitrary) orders differential equations. Applied Mathematics and Computation, 160(3):683–699, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303011883.

Ewald:2007:PCD

[EMH07] Christian-Oliver Ewald, John McNamara, and Alasdair Houston. Parental care as a differential game: a dynamic extension of the Houston– Davies game. Applied Mathematics and Computation, 190(2):1450–1465, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001786.

Eslahchi:2005:NIG

[EMJB05] M. R. Eslahchi, M. Masjed-Jamei, and E. Babolian. On numerical improvement of Gauss–Lobatto quadrature rules. Applied Mathematics and Computation, 164(3):707–717, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003200.

Ekel:2008:MAD

[EMP08] P. Ya. Ekel, J. S. C. Martini, and R. M. Palhares. Multicriteria analysis in decision making under information uncertainty. Applied Mathematics and Computation, 200(2):501– 516, July 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011307.

El-Mikkawy:2008:NRA

[EMR08] Moawwad El-Mikkawy and El-Desouky Rahmo. A new recursive algorithm for inverting general tridiagonal and anti-tridiagonal matrices. Applied Mathematics and Computation, 204(1):368–372, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005158.

El-Mikkawy:2009:NRA

[EMR09] Moawwad El-Mikkawy and El-Desouky Rahmo. A new recursive algorithm for inverting general periodic pentadiagonal and anti-pentadiagonal matrices. Applied Mathematics and Computation, 207(1):164–170, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007753.

El-Maghraby:2005:TDT

[EMY05] Nasser M. El-Maghraby and Hamdy M. Yossef. A two-dimensional thermoelasticity problem for thermomechanical shock with two relaxation times. Applied Mathematics and Computation, 170(1):172– 184, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304008999.

Effati:2005:NMS

[EN05] S. Effati and A. R. Nazemi. A new method for solving a system of the nonlinear equations. Applied Mathematics and Computation, 168(2):877–894, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006460.

Effati:2006:NNM

[EN06] S. Effati and A. R.

Nazemi. Neural network models and its application for solving linear and quadratic programming problems. Applied Mathematics and Computation, 172(1):305–331, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002055.

Engblom:2006:CMH

[Eng06] Stefan Engblom. Computing the moments of high dimensional solutions of the master equation. Applied Mathematics and Computation, 180(2):498–515, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000117.

Eldabe:2006:CFD

[EO06] N. T. Eldabe and Mahmoud E. M. Ouaf. Chebyshev finite difference method for heat and mass transfer in a hydromagnetic flow of a micropolar fluid past a stretching surface with Ohmic heating and viscous dissipation. Applied Mathematics and Computation, 177(2):561–571, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009367.

El-Owaidy:2003:PSO

[EOM03] H. El-Owaidy and H. Y. Mohamed. On the periodic solutions for an nth order difference equations. Applied Mathematics and Computation, 135 (2–3):383–390, March 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See erratum [EOM08].

El-Owaidy:2008:EPS

[EOM08] H. El-Owaidy and H. Y. Mohamed. Erratum to “On the periodic solutions for an n-th order difference equations” [Appl. Math. Comput. 135 (2003) 383–390]. Applied Mathematics and Computation, 205(1):507, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004189.

See [EOM03].

Erdem:2007:MLN

[EP07] Arzu Erdem and Serdal Pamuk. The method of lines for the numerical solution of a mathematical model for capillary formation: The role of tumor angiogenic factor in the extracellular matrix. Applied Mathematics and Computation, 186(1):891–897, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010113.

Ebrahimian:2007:NSI

[EPER07] M. Ebrahimian, R. Pourgholi, M. Emamjome, and P. Reihani. A numerical solution of an inverse parabolic problem with unknown boundary conditions. Applied Mathematics and Computation, 189(1):228– 234, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015955.

Effati:2006:IDP

[ER06a] S. Effati and H. Roohparvar. Iterative dynamic programming for solving linear and nonlinear differential equations. Applied Mathematics and Computation, 175(1):247–257, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006041.

Effati:2006:MFC

[ER06b] S. Effati and H. Roohparvar. The minimization of the fuel costs in the train transportation. Applied Mathematics and Computation, 175(2):1415–1431, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006879.

Ebadi:2007:NSN

[ERAS07] G. Ebadi, M. Y. Rahimi-Ardabili, and S. Shahmorad. Numerical solution of the nonlinear Volterra integro-differential equations by the Tau method. Applied Mathematics and Computation, 188(2):1580– 1586, May 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306015104.

Erbay:2006:STL

[Erb06] Hasan Erbay. Subspace tracking in low-rank realtime systems. Applied Mathematics and Computation, 173(2):1300–1309, February 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305004327.

Eren:2006:NSU

[Ere06] Haydar Eren. Numerical study of unsteady interacting compressible jet flows using finite difference method for small Reynolds numbers. Applied Mathematics and Computation, 172(2):876–891, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002432.

Eren:2009:MTW

[Ere09a] Tamer Eren. Minimizing the total weighted completion time on a single machine scheduling with release dates and a learning effect. Applied Mathematics and Computation, 208(2):355–358, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009004.

Eren:2009:NMM

[Ere09b]  Tamer Eren.  A note

on minimizing maximum lateness in an m-machine scheduling problem with a learning effect. Applied Mathematics and Computation, 209(2):186–190, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800903X.

See comments [XYL10].

Eryilmaz:2006:SRA

[Ery06] Serkan Eryilmaz. Some results associated with the longest run statistic in a sequence of Markov dependent trials. Applied Mathematics and Computation, 175(1):119–130, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005965.

El-Salam:2005:OOP

[ES05a] F. A. Abd El-Salam. Optimization out-of-orbit plane changes using aeroassisted maneuvers. Applied Mathematics and Computation, 170(2):1303–1313, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000652.

El-Sayed:2005:DMS

[ES05b] Salah M. El-Sayed. A direct method for solving circulant tridiagonal block systems of linear equations. Applied Mathematics and Computation, 165(1):23– 30, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003571.

El-Salam:2005:CSC

[ESA05] F. A. Abd El-Salam and M. K. Ahmed. Canonical solution of the critical inclination problem taking into account PNcorrections. Applied Mathematics and Computation, 161(3):825–841, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000347.

El-Shehawey:2010:ECA

[ESAH10] M. A. El-Shehawey and A. S. Al-Henawy. Erratum to “Computer aided solving the higher-order transition probability matrix of the finite Markov chain” [Appl. Math. Comput. 172 (2006) 267–285]. Applied Mathematics and Computation, 217(8):4311–4312, December 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [LW06a].

El-Salam:2005:PNE

[ESAR05] F. A. Abd El-Salam, M. K. Ahmed, and M. Radwan. The post-Newtonian effects in the critical inclination problem in artificial satellite theory. Applied Mathematics and Computation, 161(3):813–823, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304000335.

Emmert-Streib:2007:ITM

[ESD07a] Frank Emmert-Streib and Matthias Dehmer. Information theoretic measures of UHG graphs with low computational complexity. Applied Mathematics and Computation, 190(2):1783–1794, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002263.

Emmert-Streib:2007:TMB

[ESD07b] Frank Emmert-Streib and Matthias Dehmer. Topological mappings between graphs, trees and generalized trees. Applied Mathematics and Computation, 186(2):1326–1333, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010575.

El-Shafy:2005:EUS

[ESE05a] A. Abd El-Shafy and Mamdouh Elkady. On existence of unique solution to nonlinear integral equations of the second kind. Applied Mathematics and Computation, 163(2):801–809, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002619.

Esen:2005:NSE

[Ese05b] A. Esen. A numerical solution of the equal width wave equation by a lumped Galerkin method. Applied Mathematics and Computation, 168(1):270–282, September 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304006034.

El-Said:2006:CRC

[ESES06] Khaled M. El-Said and M. Salah El-Sherbeny. Comparing of reliability characteristics between two different systems. Applied Mathematics and Computation, 173(2):1183–1199, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305004224.

El-Salam:2006:IRO

[ESETA^{+}06] F. A. Abd El-Salam, I. A. El-Tohamy, M. K. Ahmed, W. A. Rahoma, and M. A. Rassem. Invariant relative orbits for satellite constellations: a second order theory. Applied Mathematics and Computation, 181(1):6–20, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306000610.

El-Sayed:2005:ENT

[ESK05] Salah M. El-Sayed and Doğan Kaya. Exact and numerical traveling wave solutions of Whitham– Broer–Kaup equations. Applied Mathematics and Computation, 167(2):1339– 1349, August 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304005715.

El-Sayed:2006:NSE

[ESK06] Salah M. El-Sayed and Doğan Kaya. A numerical solution and an exact explicit solution of the NLS equation. Applied Mathematics and Computation, 172(2):1315–1322, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002948.

El-Shehawy:2007:ACA

[ESK07] S. A. El-Shehawy and A. A.

Karawia. An alternative computational algorithm for calculating the nonlinearity of regression models with two parameters. Applied Mathematics and Computation, 188(1):686– 692, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601407X.

El-Sayed:2006:EGA

[ESMM06] M. F. El-Sayed, A. A. Mohamed, and T. M. N. Metwaly. Effect of general applied electric field on conducting liquid jets instabilities in the presence of heat and mass transfer. Applied Mathematics and Computation, 172(2):1078–1102, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002791.

El-Salam:2005:ASO

[ESS05a] F. A. Abd El-Salam and L. Sehnal. Analytical second-order drag theory with application for lifetime determination of artificial satellites in low Earth orbits. Applied Mathematics and Computation, 171 (2):948–971, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Salam:2005:ETI

[ESS05b] F. A. Abd El-Salam and L. Sehnal. The effects of the terrestrial infrared radiation pressure on the Earth’s artificial satellite dynamics. Applied Mathematics and Computation, 162(3):1431–1451, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001985.

El-Salam:2005:NSR

[ESS05c] M. R. Abd El-Salam and M. H. Shehata. The numerical solution for reaction– diffusion combustion with fuel consumption. Applied Mathematics and Computation, 160(2):423–435, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303011706.

El-Tawil:2005:ASS

[ET05a] Magdy A. El-Tawil. The approximate solutions of some stochastic differential equations using transformations. Applied Mathematics and Computation, 164(1):167–178, May 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003030.

Emrouznejad:2005:MMD

[ET05b] Ali Emrouznejad and Emmanuel Thanassoulis. A mathematical model for dynamic efficiency using data envelopment analysis. Applied Mathematics and Computation, 160(2):363–378, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303011548.

Tatar:2009:LCK

[eT09] Nasser eddine Tatar. On a large class of kernels yielding exponential stability in viscoelasticity. Applied Mathematics and Computation, 215(6):2298–2306, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309007541.

El-Tawil:2005:PTS

[ETETH05] M. El-Tawil, W. El-Tahan, and A. Hussein. A proposed technique of SFEM on solving ordinary random differential equation. Applied Mathematics and Computation, 161(1):35–47, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303012074.

El-Tawil:2007:UFR

[ETETH07] M. El-Tawil, W. El-Tahan, and A. Hussein. Using FEM–RVT technique for solving a randomly excited ordinary differential equation with a random operator. Applied Mathematics and Computation, 187(2):856–867, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012173.

El-Taha:2006:CDP

[ETJ06] Muhammad El-Taha and Musa J. Jafar. Characterization of the departure process in a closed fork–join synchronization network. Applied Mathematics and Computation, 181(1):214–219, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000920.

Ewald:2006:MGM

[Ewa06] Christian-Oliver Ewald.

The Malliavin gradient method for the calibration of stochastic dynamical models. Applied Mathematics and Computation, 175(2):1332–1352, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006818.

El-Wakil:2006:ADMa

[EWAE06] S. A. El-Wakil, M. A.

Abdou, and A. Elhanbaly. Adomian decomposition method for solving the diffusion–convection– reaction equations. Applied Mathematics and Computation, 177(2):729–736, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009537.

El-Wakil:2008:NPS

[EWAH08] S. A. El-Wakil, M. A. Abdou, and A. Hendi. New periodic and soliton solutions of nonlinear evolution equations. Applied Mathematics and Computation, 197(2):497–506, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008168.

El-Wakil:2006:ADMb

[EWEA06] S. A. El-Wakil, A. Elhanbaly, and M. A. Abdou. Adomian decomposition method for solving fractional nonlinear differential equations. Applied Mathematics and Computation, 182(1):313–324, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003390.

Farrag:2005:CPC

[FA05a] A. S. Farrag and S. E. Abbas. Computer programming for construction and enumeration of all regular topologies and equivalence relations on finite sets. Applied Mathematics and Computation, 165(1):177– 184, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003716.

Fernandez:2005:AIS

[FA05b] M. Fernández and D. P.

Atherton. Analysis of insulin sensitivity estimates from linear models of glucose disappearance. Applied Mathematics and Computation, 167(1):528–538, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005119.

Fashandi:2006:SAM

[FA06] M. Fashandi and J. Ahmadi. Series approximations for the means of k-records. Applied Mathematics and Computation, 174(2):1290–1301, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005497.

Fabien:2008:POU

[Fab08] Brian C. Fabien. Parameter optimization using the L_{∞} exact penalty function and strictly convex quadratic programming problems. Applied Mathematics and Computation, 198(2):833–848, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009563.

Fadrony:2005:HPA

[Fad05] D. Rostamy Varnos Fadrony. Hybrid preconditioned algorithms for boundary hypersingular integral equations. Applied Mathematics and Computation, 166(3):531–550, July 26, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004564.

Fadrani:2006:NAC

[Fad06] D. Rostamy Varnos Fadrani. A new aspect for choosing collocation points for solving biharmonic equations. Applied Mathematics and Computation, 181(2):1112–1119, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001950.

Faghihi:2005:NMP

[Fag05]  F. Faghihi.  Numerical

minimization procedures of the adiabatic approximation wavefunction. Applied Mathematics and Computation, 162(3):1013–1021, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001456.

Fathian:2007:AHB

[FAM07] Mohammad Fathian, Babak Amiri, and Ali Maroosi. Application of honey-bee mating optimization algorithm on clustering. Applied Mathematics and Computation, 190(2):1502– 1513, July 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307001853.

Fang:2006:RPP

[Fan06] Weifu Fang. Reconstruction of permittivity profile from boundary capacitance data. Applied Mathematics and Computation, 177(1):178–188, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008994.

Fernandez:2008:IPM

[FAQ08] Margarita Fernandez, Dionisio Acosta, and Adolfo Quiroz. Improving the precision of model parameters using model based signal enhancement and the linear minimal model following an IVGTT in the healthy man. Applied Mathematics and Computation, 196(1):185–192, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006571.

Faruqi:2005:ASQ

[Far05] Farhan A. Faruqi. On the algebraic structure of quadratic and bilinear dynamical systems. Applied Mathematics and Computation, 162(2):751–797, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001316.

Farkas:2008:SPS

[Far08] Jozsef Z. Farkas. Structured populations: The stabilizing effect of the inflow of newborns from an external source and the net growth rate. Applied Mathematics and Computation, 199(2):547–558, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010302.

Fatullayev:2005:NPSa

[Fat05a] Afet Golayoglu Fatullayev. Numerical procedure for the simultaneous determination of unknown coefficients in a parabolic equation. Applied Mathematics and Computation, 162(3):1367–1375, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001948.

Fatullayev:2005:NPSb

[Fat05b] Afet Golayoglu Fatullayev. Numerical procedure for the simultaneous determination of unknown coefficients in a parabolic equation. Applied Mathematics and Computation, 164(3):697–705, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003182.

Fathali:2006:GAM

[Fat06]  Jafar Fathali.  A ge-

netic algorithm for the pmedian problem with pos/ neg weights. Applied Mathematics and Computation, 183(2):1071–1083, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006424.

Fazio:2009:NTM

[Faz09] Riccardo Fazio. Numerical transformation methods: Blasius problem and its variants. Applied Mathematics and Computation, 215(4):1513–1521, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006511.

Fu:2007:SAL

[FB07] Xiao-Ling Fu and Abdellah Bnouhachem. An selfadaptive LQP method for constrained variational inequalities. Applied Mathematics and Computation, 189(2):1586–1600, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017449.

Fitouhi:2008:SIB

[FB08] Ahmed Fitouhi and Kamel Brahim. Some inequalities for the q-beta and the q-gamma functions via some q-integral inequalities. Applied Mathematics and Computation, 204(1):385–394, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005171.

Fitouhi:2007:EAT

[FBB07] Ahmed Fitouhi, Fethi Bouzeffour, and Wafa Binous. Expansion and asymptotic in terms of basic Bessel functions. Applied Mathematics and Computation, 188(2):2034– 2044, May 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306015530.

Fitouhi:2008:HIS

[FBB08] Ahmed Fitouhi, Néji Bettaibi, and Rym H. Bettaieb. On Hardy’s inequality for symmetric integral transforms and analogous. Applied Mathematics and Computation, 198(1):346–354, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008958.

Feng:2006:FPS

[FC06] Jianwen Feng and Shihua Chen. Four periodic solutions of a generalized delayed predator– prey system. Applied Mathematics and Computation, 181(2):932–939, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001731.

Fatullayev:2006:CNM

[FCG06] Afet Golayoglu Fatullayev, Emine Can, and Nizami Gasilov. Comparing numerical methods for inverse coefficient problem in parabolic equation. Applied Mathematics and Computation, 179(2):567–571, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010623.

Fei:2008:FMI

[FDL08] Minrui Fei, Dajun Du, and Kang Li. A fast model identification method for networked control system. Applied Mathematics and Computation, 205(2):658–667, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003445.

Ding:2009:NSQ

[fDxZfWyY09] Heng fei Ding, Yu xin Zhang, San fu Wang, and Xiao ya Yang. A note on some quadrature based three-step iterative methods for non-linear equations. Applied Mathematics and Computation, 215(1): 53–57, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309003658.

Farnoosh:2007:MCM

[FE07] R. Farnoosh and M. Ebrahimi.

Monte Carlo method via a numerical algorithm to solve a parabolic problem. Applied Mathematics and Computation, 190(2):1593–1601, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002044.

Farnoosh:2008:MCM

[FE08] R. Farnoosh and M. Ebrahimi.

Monte Carlo method for solving Fredholm integral equations of the second kind. Applied Mathematics and Computation, 195(1):309–315, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005541.

Felber:2005:NES

[Fel05] F. S. Felber. New exact solutions of differential equations derived by fractional calculus. Applied Mathematics and Computation, 170(2):1261–1270, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000615.

Feng:2005:RMO

[Fen05]  Lihong Feng.  Review

of model order reduction methods for numerical simulation of nonlinear circuits. Applied Mathematics and Computation, 167(1):576–591, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304005156.

Feng:2006:CSS

[Fen06] Zhaosheng Feng. Computations of soliton solutions and periodic solutions for the focusing branch of the nonlinear dispersive K ( n n ) equations in higher-dimensional spaces. Applied Mathematics and Computation, 182(1):781–790, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003870.

Fatehifar:2008:DNM

[FEOC08] E. Fatehifar, A. Elkamel, A. Alizadeh Osalu, and A. Charchi. Developing a new model for simulation of pollution dispersion from a network of stacks. Applied Mathematics and Computation, 206(2):662–668, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003895.

Fernandez:2009:SAM

[Fer09] Francisco M. Fernández.

On some approximate methods for nonlinear models. Applied Mathematics and Computation, 215(1):168–174, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004251.

Furati:2005:UBS

[FeT05] Khaled M. Furati and Nasser eddine Tatar. Uniform boundedness and stability for a viscoelastic problem. Applied Mathematics and Computation, 167(2):1211–1220, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304005624.

Francisco:2006:HBR

[FF06] Gerson Francisco and André Fonseca. Homoclinic bifurcations in reversible Hamiltonian systems. Applied Mathematics and Computation, 176(2):654–661, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008635.

Fetecau:2008:DPV

[FFKV08] Corina Fetecau, C. Fetecau, M. Khan, and D. Vieru. Decay of a potential vortex in a generalized Oldroyd-B fluid. Applied Mathematics and Computation, 205(1):497–506, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006012.

Flores-Franulic:2007:CTI

[FFRF07] A. Flores-Franulič and H. Román-Flores. A Chebyshev type inequality for fuzzy integrals. Applied Mathematics and Computation, 190(2):1178–1184, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001488.

Flores-Franulic:2008:NFI

[FFRFCC08] A. Flores-Franulič, H. Román-Flores, and Y. Chalco-Cano. A note on fuzzy integral inequality of Stolarsky type. Applied Mathematics and Computation, 196(1):55–59, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006261.

Flores-Franulic:2009:MTI

[FFRFCC09] A. Flores-Franulič, H. Román-Flores, and Y. Chalco-Cano. Markov type inequalities for fuzzy integrals. Applied Mathematics and Computation, 207(1):242–247, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800790X.

Fathabadi:2007:NAS

[FG07] Hassan Salehi Fathabadi and Mehdi Ghiyasvand. A new algorithm for solving the feasibility problem of a network flow.

Applied Mathematics and Computation, 192(2):429– 438, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307003463.

Fare:2009:CDD

[FG09] Rolf Färe and Shawna Grosskopf. A comment on dynamic DEA. Applied Mathematics and Computation, 213(1):275–276, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900229X.

Fang:2006:NIM

[FGfFL06] Tiegang Fang, Fang Guo, and Chia fon F. Lee. A new iteration method with cubic convergence to solve nonlinear algebraic equations. Applied Mathematics and Computation, 175(2):1147–1155, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006673.

Feng:2006:PIM

[FH06a] Xinlong Feng and Yinnian He. Parametric iterative methods of secondorder for solving nonlinear equation. Applied Mathematics and Computation, 173(2):1060–1067, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500411X.

Fu:2006:HMN

[FH06b] H. S. Fu and B. Han.

A homotopy method for nonlinear inverse problems. Applied Mathematics and Computation, 183(2):1270–1279, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006746.

Feng:2007:HOI

[FH07] Xinlong Feng and Yinnian He. High order iterative methods without derivatives for solving nonlinear equations. Applied Mathematics and Computation, 186(2):1617–1623, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010988.

Fu:2007:WMH

[FHG07] H. S. Fu, B. Han, and G. Q.

Gai. A wavelet multiscale– homotopy method for the inverse problem of two-dimensional acoustic wave equation. Applied Mathematics and Computation, 190(1):576–582, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000914.

Fang:2009:NSN

[FHH09] Liang Fang, Guoping He, and Yunhong Hu. A new smoothing Newtontype method for secondorder cone programming problems. Applied Mathematics and Computation, 215(3):1020–1029, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005955.

Feng:2007:BTW

[FHL07] Dahe Feng, Tianlan He, and Junliang Lü. Bifurcations of travelling wave solutions for (2 + 1)- dimensional Boussinesqtype equation. Applied Mathematics and Computation, 185(1):402–414, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008757.

Feng:2008:STR

[FHSS08] Ye Feng, Liu Hongwei, Zhou Shuisheng, and Liu Sanyang. A smoothing trust-region Newton-CG method for minimax problem. Applied Mathematics and Computation, 199(2):581–589, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010338.

Fakhar:2009:STS

[FHYZ09] K. Fakhar, T. Hayat, Cheng Yi, and Kun Zhao. Symmetry transformation of solutions for the Navier– Stokes equations. Applied Mathematics and Computation, 207(1):213–224, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007819.

Feng:2008:NRM

[FHZ08] Xinlong Feng, Yinnian He, and Zhinan Zhang. The numerical rank of a matrix and its applications. Applied Mathematics and Computation, 196(1):416–421, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006947.

Figueiredo:2008:QED

[Fig08] Giovany M. Figueiredo. Quasilinear equations with dependence on the gradient via Mountain Pass techniques in R^{N}. Applied Mathematics and Computation, 203(1):14– 18, September 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030800218X.

Filatov:2005:SAC

[Fil05] Denis M. Filatov. Splitting as an approach to constructing local exact artificial boundary conditions. Applied Mathematics and Computation, 170(2):1209–1242, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000597.

Fazio:2006:MNM

[FJ06] Riccardo Fazio and Alessandra Jannelli. Mathematical and numerical modeling for a bio-chemical aquarium. Applied Mathematics and Computation, 174(2):1370–1383, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005564.

Foster:2005:FES

[FJCW05] Winfred A. Foster, Jr., Clarence E. Johnson, Robert C. Chiroux, and Tom R. Way. Finite element simulation of cone penetration. Applied Mathematics and Computation, 162(2):735–749, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001286.

Foroughi:2005:SMP

[FJT05] A. A. Foroughi, D. F.

Jones, and M. Tamiz. A selection method for a preferential election. Applied Mathematics and Computation, 163(1):107–116, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002139.

Fengying:2006:GSA

[FK06a] Wei Fengying and Wang Ke. Global stability and asymptotically periodic solution for nonautonomous cooperative Lotka–Volterra diffusion system. Applied Mathematics and Computation, 182(1):161–165, November 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306003225.

Fengying:2006:PVC

[FK06b] Wei Fengying and Wang Ke. Permanence of variable coefficients predator– prey system with stage structure. Applied Mathematics and Computation, 180(2):594–598, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000476.

Fengying:2007:PSS

[FK07] Wei Fengying and Wang Ke. Persistence of some stage structured ecosystems with finite and infinite delay. Applied Mathematics and Computation, 189(1):902–909, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016699.

Figiel:2009:NPA

[FK09] Łukasz Figiel and Marcin Kamiński. Numerical probabilistic approach to sensitivity analysis in a fatigue delamination problem of a two layer composite. Applied Mathematics and Computation, 209(1):75– 90, March 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004761.

Firsov:2006:FDA

[FL06] D. Firsov and S. H. Lui.

A fast deblurring algorithm. Applied Mathematics and Computation, 183(1):285–291, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005601.

Feng:2009:IFS

[FL09a] Dahe Feng and Guoxiang Luo. The improved Fan sub-equation method and its application to the SK equation. Applied Mathematics and Computation, 215(5):1949–1967, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007061.

Feng:2009:EES

[FL09b] Wei-Gui Feng and Chang Lin. Explicit exact solutions for the (2 + 1)- dimensional Konopelchenko– Dubrovsky equation. Applied Mathematics and Computation, 210(2):298– 302, April 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308009971.

Feng:2007:IFS

[FLLH07a] Dahe Feng, Jibin Li, Junliang Lü, and Tianlan He. The improved Fan subequation method and its application to the Boussinseq wave equation. Applied Mathematics and Computation, 194(2):309–320, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004717.

Feng:2007:BST

[FLLH07b] Dahe Feng, Junliang Lü, Jibin Li, and Tianlan He. Bifurcation studies on travelling wave solutions for nonlinear intensity Klein– Gordon equation. Applied Mathematics and Computation, 189(1):271–284, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016006.

Feng:2008:NEE

[FLLH08] Dahe Feng, Jibin Li, Junliang Lü, and Tianlan He. New explicit and exact solutions for a system of variant RLW equations. Applied Mathematics and Computation, 198(2):715– 720, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009472.

Fundinger:2007:AMA

[FLO07] Danny Fundinger, Torsten Lindström, and George Osipenko. On the appearance of multiple attractors in discrete foodchains. Applied Mathematics and Computation, 184(2):429–444, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007181.

Li:2008:NIM

[fLsLdXlF08] Tai fang Li, De sheng Li, Zhao di Xu, and Ya ling Fang. New iterative methods for nonlinear equations. Applied Mathematics and Computation, 197(2):755–759, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700834X.

Li:2009:SCP

[fLyHZ09] Jiao fen Li, Xi yan Hu, and Lei Zhang. The submatrix constraint problem of matrix equation AXB + CY D = D. Applied Mathematics and Computation, 215(7):2578–2590, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309007966.

Farhloul:2005:MHF

[FM05] M. Farhloul and A. Serghini Mounim. A mixedhybrid finite element method for convection–diffusion problems. Applied Mathematics and Computation, 171(2):1037–1047, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fatori:2008:TSM

[FM08] Luci Harue Fatori and To Fu Ma. A thermoelastic system of memory type in noncylindrical domains. Applied Mathematics and Computation, 200(2):583– 589, July 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030701137X.

Feng:2007:EAS

[FMH07] Xinlong Feng, Liquan Mei, and Guoliang He. An efficient algorithm for solving Troesch’s problem. Applied Mathematics and Computation, 189(1):500–507, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016237.

Faria:2009:BGE

[FO09] Teresa Faria and José J.

Oliveira. Boundedness and global exponential stability for delayed differential equations with applications. Applied Mathematics and Computation, 214(2):487–496, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003440.

Falcon:2009:FNA

[FP09a] Sergio Falcon and Angel Plaza. On k-Fibonacci numbers of arithmetic indexes. Applied Mathematics and Computation, 208(1):180–185, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008837.

Fan:2009:NLM

[FP09b] Jinyan Fan and Jianyu Pan. A note on the Levenberg–Marquardt parameter. Applied Mathematics and Computation, 207(2):351–359, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008199.

Franjic:2006:GEO

[FPP06] I. Franjić, I. Perić, and J. Pečarić. General Euler– Ostrowski formulae and applications to quadratures. Applied Mathematics and Computation, 177(1):92– 98, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008908.

Feng:2008:NAF

[FqZ08] Yang Feng and Hong qing Zhang. A new auxiliary function method for solving the generalized coupled Hirota–Satsuma KdV system. Applied Mathematics and Computation, 200(1):283–288, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011071.

Qiao:2008:FUP

[fQZgH08] Jun fei Qiao, Ying Zhang, and Hong gui Han. Fast unit pruning algorithm for feedforward neural network design. Applied Mathematics and Computation, 205(2):622–627, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003408.

Frasin:2007:NCM

[Fra07] B. A. Frasin. Neighborhoods of certain multivalent functions with negative coefficients. Applied Mathematics and Computation, 193(1):1–6, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700358X.

Frammartino:2009:NMS

[Fra09a] Carmelina Frammartino. A Nyström method for solving 2s-th order boundary value problems. Applied Mathematics and Computation, 215(1):100–109, September 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309003889.

Frasson:2009:DRC

[Fra09b] Miguel V. S. Frasson. On the dominance of roots of characteristic equations for neutral functional differential equations. Applied Mathematics and Computation, 214(1):66–72, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002793.

From:2006:LBA

[Fro06] Steven G. From. Lower bounds and approximations of the locations of movable singularities of some nonlinear differential equations using parameterized bounded operators. Applied Mathematics and Computation, 175(1):16– 37, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005886.

Fu:2005:NAT

[FS05] Jinhua Fu and Wenyu Sun. Nonmonotone adaptive trust-region method for unconstrained optimization problems. Applied Mathematics and Computation, 163(1):489–504, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002371.

Fort:2009:CGH

[FS09] Marta Fort and J. Antoni Sellarès. Computing generalized higher-order Voronoi diagrams on triangulated surfaces. Applied Mathematics and Computation, 215(1):235–250, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004524.

Firat:2006:LLE

[FSÇ06] Seyhan Fırat, Mehmet Sarıbıyık, and Erkan Çelebi. Lateral load estimation from visco-plastic mudflow around cylindrical row of piles. Applied Mathematics and Computation, 173(2):803–821, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003942.

Fakhraie:2007:CPE

[FSG07] M. Fakhraie, A. Shidfar, and M. Garshasbi. A computational procedure for estimation of an unknown coefficient in an inverse boundary value problem. Applied Mathematics and Computation, 187(2):1120–1125, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012483.

Faye:2005:FMW

[FSMC05] R. M. Faye, S. Sawadogo, and F. Mora-Camino. Flexible management of water resource systems. Applied Mathematics and Computation, 167(1):516–527, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304005107.

Faria:2007:PHT

[FT07] Teresa Faria and Sergei Trofimchuk. Positive heteroclinics and traveling waves for scalar population models with a single delay. Applied Mathematics and Computation, 185(1):594–603, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008940.

Fang:2007:FDN

[FTG07] Ming Fang, Yantao Tian, and Li Guo. Fault diagnosis of nonlinear system based on generalized observer. Applied Mathematics and Computation, 185(2):1131–1137, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600837X.

Fujita:2007:MVB

[FTMA07] Keiko Fujita, Yoshitsugu Takei, Akira Morimoto, and Ryuichi Ashino. Mathematical view of a blind source separation on a time frequency space. Applied Mathematics and Computation, 187(1):153–162, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011465.

Francomano:2009:UMS

[FTT^{+}09] E. Francomano, A. Tortorici, E. Toscano, G. Ala, and F. Viola. On the use of a meshless solver for PDEs governing electromagnetic transients. Applied Mathematics and Computation, 209(1):42– 51, March 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004724.

Fukuda:2006:BBU

[Fuk06a]  Kosei Fukuda.  BIC-

based unit-root detection: Simulation-based evidence. Applied Mathematics and Computation, 183(1):518–521, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005868.

Fukuda:2006:DBC

[Fuk06b] Kosei Fukuda. Differentiating between coefficient break and volatility break. Applied Mathematics and Computation, 176(1):262– 269, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305007903.

Fushan:2008:EUS

[Fus08]  Li Fushan.  Existence

and uniqueness of the solution to the viscoelastic equations for Koiter shells. Applied Mathematics and Computation, 200(1):407–412, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011459.

Farto:2005:NTP

[FV05] J. Farto and C. Vázquez.

Numerical techniques for pricing callable bonds with notice. Applied Mathematics and Computation, 161(3):989–1013, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000487.

Fang:2007:TFE

[FW07a] Yonglei Fang and Xinyuan Wu. A trigonometrically fitted explicit hybrid method for the numerical integration of orbital problems. Applied Mathematics and Computation, 189(1):178–185, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015906.

Wu:2007:MPS

[fW07b] Tsung fang Wu. Multiple positive solutions for semilinear elliptic systems with nonlinear boundary condition. Applied Mathematics and Computation, 189(2):1712–1722, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017632.

Fu:2009:IMS

[FWJK09] Yu-Yi Fu, Chia-Ju Wu, Jin-Tsong Jeng, and Chia-Nan Ko. Identification of MIMO systems using radial basis function networks with hybrid learning algorithm. Applied Mathematics and Computation, 213(1):184–196, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002173.

Fan:2008:RSB

[FWL08] Ming-Quan Fan, Hong-Xia Wang, and Sheng-Kun Li. Restudy on SVD-based watermarking scheme. Applied Mathematics and Computation, 203(2):926–930, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002865.

Wang:2009:CNW

[fWZpXzL09] Shu fan Wang, Bing Zheng, Zhi ping Xiong, and Zi zhen Li. The condition numbers for weighted Moore–Penrose inverse and weighted linear least squares problem. Applied Mathematics and Computation, 215(1):197–205, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004287.

Feng:2009:ABN

[FX09] Jinqian Feng and Wei Xu.

Analysis of bifurcations for non-linear stochastic non-smooth vibroimpact system via top Lyapunov exponent. Applied Mathematics and Computation, 213(2):577–586, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002707.

Fu:2005:WSR

[FXLZ05] Chu-Li Fu, Xiang-Tuan Xiong, Hong-Fang Li, and You-Bin Zhu. Wavelet and spectral regularization methods for a sideways parabolic equation. Applied Mathematics and Computation, 160(3):881–908, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303012013.

Fakhar:2008:EST

[FXY08] K. Fakhar, Zhenli Xu, and Cheng Yi. Exact solutions of a third grade fluid flow on a porous plate. Applied Mathematics and Computation, 202(1):376–382, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308001227.

Fan:2008:PPF

[FY08] Xiaona Fan and Bo Yu. A polynomial path following algorithm for convex programming. Applied Mathematics and Computation, 196(2):866–878, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307007400.

Feng:2007:ACP

[FYZ07] Y. J. Feng, L. Yu, and G. L. Zhang. Ant colony pattern search algorithms for unconstrained and bound constrained optimization. Applied Mathematics and Computation, 191(1):42–56, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013221.

Feng:2007:RRM

[FZ07] Xinlong Feng and Zhinan Zhang. The rank of a random matrix. Applied Mathematics and Computation, 185(1):689–694, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009040.

Fu:2008:MRF

[FZ08] Xudan Fu and Xia Zhou.

On matrices related with Fibonacci and Lucas numbers. Applied Mathematics and Computation, 200(1):96–100, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010909.

Fang:2009:ENH

[FZ09] Hui Fang and Dapeng Zhao. Existence of nontrivial homoclinic orbits for fourth-order difference equations. Applied Mathematics and Computation, 214(1):163–170, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002902.

Fatemi:2007:NGA

[FZN07] Afsaneh Fatemi, Kamran Zamanifar, and Naser NematBakhsh. A new genetic approach to construct near-optimal binary search trees. Applied Mathematics and Computation, 190(2):1514–1525, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001877.

GomezS:2007:SFF

[G07] Cesar A. Gómez S. Special forms of the fifthorder KdV equation with new periodic and soliton solutions. Applied Mathematics and Computation, 189(2):1066–1077, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016833.

Glavic:2007:ENR

[GA07] Mevludin Glavic and Fernando L. Alvarado. An extension of Newton– Raphson power flow problem. Applied Mathematics and Computation, 186(2):1192–1204, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010459.

Ganji:2008:ETS

[GA08] D. D. Ganji and M. Abdollahzadeh. Exact travelling solutions for the Lax’s seventh-order KdV equation by sech method and rational exp-function method. Applied Mathematics and Computation, 206(1):438–444, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006954.

Ghazanfari:2007:CSA

[GAFK07] M. Ghazanfari, S. Alizadeh, M. Fathian, and D. E. Koulouriotis. Comparing simulated annealing and genetic algorithm in learning FCM. Applied Mathematics and Computation, 192(1):56– 68, September 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307002949.

Gupta:2009:IBT

[GAG09] Vijay Gupta, P. N. Agrawal, and Asha Ram Gairola. On the integrated Baskakov type operators. Applied Mathematics and Computation, 213(2):419–425, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002525.

Galligani:2006:AMM

[Gal06a] Emanuele Galligani. The Arithmetic Mean method for solving systems of nonlinear equations in finite differences. Applied Mathematics and Computation, 181(1):579–597, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001299.

Galton:2006:CTT

[Gal06b]  Antony Galton.  The

Church–Turing thesis: Still valid after all these years? Applied Mathematics and Computation, 178(1):93– 102, July 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008374.

Galewski:2008:NEB

[Gal08a]  Marek Galewski.  On

the nonlinear elastic beam equation. Applied Mathematics and Computation, 202(1):427–434, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800091X.

Galewski:2008:OCP

[Gal08b] Marek Galewski. On the optimal control problem governed by the nonlinear elastic beam equation. Applied Mathematics and Computation, 203(2):916–920, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002233.

Ghazinoory:2007:UAP

[GANG07] S. Ghazinoory, A. Aliahmadi, S. Namdarzangeneh, and S. H. Ghodsypour. Using AHP and L.P. for choosing the best alternatives based the gap analysis. Applied Mathematics and Computation, 184(2):316–321, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007089.

Gao:2006:NMQ

[Gao06]  Yong-Hua Gao.  New-

ton’s method for the quadratic matrix equation. Applied Mathematics and Computation, 182(2):1772–1779, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005248.

Gao:2007:CES

[Gao07] Ji Gao. On characterization of Euclidean spaces. Applied Mathematics and Computation, 187(1):170– 175, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011489.

Grace:2008:OCT

[GAPT08] Said R. Grace, Ravi P. Agarwal, Raffaella Pavani, and E. Thandapani. On the oscillation of certain third order nonlinear functional differential equations. Applied Mathematics and Computation, 202(1):102–112, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308000623.

Garadzhaev:2007:NST

[Gar07] A. Garadzhaev. A note on the spectral theory of problems on normal oscillation of an ideal compressible fluid in rotating elastic shell. Applied Mathematics and Computation, 191(2):302–307, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002573.

Garcia-Alonso:2009:NME

[GAR09] F. García-Alonso and J. A.

Reyes. A new method for exact integration of some perturbed stiff linear systems of oscillatory type. Applied Mathematics and Computation, 215(7):2649–2662, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008030.

Gencer:2008:CAD

[GAS08] Cevriye Gencer, Emel Kızılkaya Aydogan, and Abdullah Soydemır. Chemical agent detector placement methodology. Applied Mathematics and Computation, 195(2):542–557, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700584X.

Garcia:2008:VAW

[GAV08] V. M. Garcia, L. Acevedo, and A. M. Vidal. Variants of algebraic waveletbased multigrid methods: Application to shifted linear systems. Applied Mathematics and Computation, 202(1):287–299, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308001094.

Guzel:2005:NSS

[GB05] Nuran Guzel and Mustafa Bayram. On the numerical solution of stiff systems. Applied Mathematics and Computation, 170(1):230–236, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030400921X.

Guzel:2006:NSDa

[GB06a] Nuran Guzel and Mustafa Bayram. Numerical solution of differential–algebraic equations with index-2. Applied Mathematics and Computation, 174(2):1279– 1289, March 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305005485.

Guzel:2006:NSDb

[GB06b] Nuran Guzel and Mustafa Bayram. On the numerical solution of differentialalgebraic equations with index-3. Applied Mathematics and Computation, 175(2):1320–1331, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006806.

Gao:2007:NIB

[GB07] Hongjun Gao and Charles Bu. Neumann inhomogeneous boundary value problem for the n + 1 complex Ginzburg–Landau equation. Applied Mathematics and Computation, 188(1):394–398, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306013634.

Guergachi:2008:SML

[GB08] Aziz Guergachi and Goran Boskovic. System models or learning machines? Applied Mathematics and Computation, 204(2):553– 567, October 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308003020.

Graca:2009:CBP

[GBC09] Daniel S. Graça, Jorge Buescu, and Manuel L. Campagnolo. Computational bounds on polynomial differential equations. Applied Mathematics and Computation, 215(4):1375–1385, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004093.

Gupta:2007:AGA

[GBG07] R. K. Gupta, A. K. Bhunia, and S. K. Goyal. An application of genetic algorithm in a marketing oriented inventory model with interval valued inventory costs and threecomponent demand rate dependent on displayed stock level. Applied Mathematics and Computation, 192(2):466–478, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003505.

Gupta:2008:RCS

[GBG08] M. K. Gupta, Man Singh Beniwal, and Prachi Goel. Rate of convergence for Szasz–Mirakyan–Durrmeyer operators with derivatives of bounded variation. Applied Mathematics and Computation, 199(2):828– 832, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010624.

Gachpazan:2006:MTA

[GBK06] Mortaza Gachpazan, Akbar Hashemi Borzabadi, and Ali Vahidian Kamyad. A measure-theoretical approach for solving discrete optimal control problems. Applied Mathematics and Computation, 173(2):736–752, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003875.

Gu:2007:MAM

[GbL07] Huifeng Gu and Zhi bin Li. A modified Adomian method for system of nonlinear differential equations. Applied Mathematics and Computation, 187(2):748–755, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011805.

Gokhan:2005:DSS

[GÇ05] A. Gökhan and R. Çolak. Double sequence space ℓ_{2}^{∞}(p). Applied Mathematics and Computation, 160(1):147–153, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303010981.

Gao:2006:STO

[GC06a] Feng Gao and Chun-Mei Chi. Solving third-order obstacle problems with quartic B-splines. Applied Mathematics and Computation, 180(1):270–274, September 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305011227.

Ge:2006:OMP

[GC06b] Xin-Sheng Ge and Li-Qun Chen. Optimal motion planning for nonholonomic systems using genetic algorithm with wavelet approximation. Applied Mathematics and Computation, 180(1):76–85, September 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305010994.

Gao:2007:USD

[GC07a] Feng Gao and Chunmei Chi. Unconditionally stable difference schemes for a one-space-dimensional linear hyperbolic equation. Applied Mathematics and Computation, 187(2):1272–1276, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012653.

Geng:2007:SSN

[GC07b] Fazhan Geng and Minggen Cui. Solving singular nonlinear second-order periodic boundary value problems in the reproducing kernel space. Applied Mathematics and Computation, 192(2):389–398, September 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307003414.

Guo:2007:WPC

[GC07c] Cuihua Guo and Shangbin Cui. Well-posedness of Cauchy problem of fourth-order Schrödinger equations in anisotropic Sobolev space. Applied Mathematics and Computation, 190(1):583–589, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000926.

Guo:2009:NCP

[GC09] Cuihua Guo and Shangbin Cui. A note on the Cauchy problem of the generalized Davey–Stewartson equations. Applied Mathematics and Computation, 215(6):2262–2268, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007395.

Gong:2008:EPD

[GCJ08] Wenyin Gong, Zhihua Cai, and Liangxiao Jiang. Enhancing the performance of differential evolution using orthogonal design method. Applied Mathematics and Computation, 206(1):56–69, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006280.

Ghosh:2005:MSB

[GCSS05] Mini Ghosh, Peeyush Chandra, Prawal Sinha, and J. B. Shukla. Modelling the spread of bacterial disease: effect of service providers from an environmentally degraded region. Applied Mathematics and Computation, 160(3):615–647, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303011846.

Gao:2008:HBG

[GCT08] Shujing Gao, Lansun Chen, and Zhidong Teng. Hopf bifurcation and global stability for a delayed predator–prey system with stage structure for predator. Applied Mathematics and Computation, 202(2):721–729, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001707.

Gurcan:2006:SCN

[GD06] F. Gürcan and A. Delıceoğlu.

Saddle connections near degenerate critical points in Stokes flow within cavities. Applied Mathematics and Computation, 172(2):1133–1144, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305002821.

Gafiychuk:2008:IOS

[GD08] V. Gafiychuk and B. Datsko. Inhomogeneous oscillatory solutions in fractional reaction–diffusion systems and their computer modeling. Applied Mathematics and Computation, 198(1):251–260, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008855.

Grau:2006:IOR

[GDB06] Miquel Grau and José Luis Díaz-Barrero. An improvement to Ostrowski rootfinding method. Applied Mathematics and Computation, 173(1):450–456, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305003656.

Du:2007:NRS

[gDH07] Jian guo Du and Tingwen Huang. New results on stable region of Nash equilibrium of output game model. Applied Mathematics and Computation, 192(1):12–19, September 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307002883.

Geem:2008:NDH

[Gee08] Zong Woo Geem. Novel derivative of harmony search algorithm for discrete design variables. Applied Mathematics and Computation, 199(1):223– 230, May 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307009897.

Gravvanis:2007:PPA

[GEG07] George A. Gravvanis, Victor N. Epitropou, and Konstantinos M. Giannoutakis. On the performance of parallel approximate inverse preconditioning using Java multithreading techniques. Applied Mathematics and Computation, 190(1):255– 270, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700046X.

Geng:2009:NRK

[Gen09a] Fazhan Geng. A new reproducing kernel Hilbert space method for solving nonlinear fourth-order boundary value problems. Applied Mathematics and Computation, 213(1):163– 169, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002148.

Geng:2009:SSS

[Gen09b]  Fazhan Geng.  Solv-

ing singular second order three-point boundary value problems using reproducing kernel Hilbert space method. Applied Mathematics and Computation, 215(6):2095–2102, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309007188.

George:2007:RA

[Geo07] Robins George. Robs algorithm. Applied Mathematics and Computation, 189 (1):314–325, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306016055.

Geoffroy:2009:SMI

[Geo09] Michel H. Geoffroy. Stability of Mann’s iterates under metric regularity. Applied Mathematics and Computation, 215(2):686–694, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005487.

Geum:2007:AEC

[Geu07] Young Hee Geum. The asymptotic error constant of leap-frogging Newton’s method locating a simple real zero. Applied Mathematics and Computation, 189(1):963–969, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016754.

Gimenes:2006:EIS

[GF06] L. P. Gimenes and M. Federson. Existence and impulsive stability for second order retarded differential equations. Applied Mathematics and Computation, 177(1):44– 62, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008866.

Gupta:2009:CDT

[GF09] Vijay Gupta and Zoltán Finta. On certain q-Durrmeyer type operators. Applied Mathematics and Computation, 209(2):415–420, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308009843.

Ge:2008:SNF

[GFH08] Zhihao Ge, Minfu Feng, and Yinnian He. A stabilized nonconfirming finite element method based on multiscale enrichment for the stationary Navier– Stokes equations. Applied Mathematics and Computation, 202(2):700–707, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308001598.

Gao:2005:PIP

[GFWX05] Caixia Gao, Enmin Feng, Zongtao Wang, and Zhilong Xiu. Parameters identification problem of the nonlinear dynamical system in microbial continuous cultures. Applied Mathematics and Computation, 169(1):476–484, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007878.

Grendar:2005:MPE

[GG05a] M. Grendar, Jr. and M. Grendar. Maximum probability/entropy translating of contiguous categorical observations into frequencies. Applied Mathematics and Computation, 161(2):347–351, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030301244X.

Gudla:2005:AHG

[GG05b] Pradeep Kumar Gudla and Ranjan Ganguli. An automated hybrid geneticconjugate gradient algorithm for multimodal optimization problems. Applied Mathematics and Computation, 167(2):1457–1474, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005818.

Goghary:2006:TCM

[GG06] Hossein Sadeghi Goghary and Motahare Sadeghi Goghary. Two computational methods for solving linear Fredholm fuzzy integral equation of the second kind. Applied Mathematics and Computation, 182(1):791–796, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003882.

Gudla:2007:EEI

[GG07] Pradeep Kumar Gudla and Ranjan Ganguli. Error estimates for inconsistent load lumping approach in finite element solution of differential equations. Applied Mathematics and Computation, 194(1):21–37, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004390.

Giannoutakis:2008:HPF

[GG08a] Konstantinos M. Giannoutakis and George A. Gravvanis. High performance finite element approximate inverse preconditioning. Applied Mathematics and Computation, 201(1–2):293–304, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307012052.

Gutierrez-Gutierrez:2008:PIPa

[GG08b] Jesús Gutiérrez-Gutiérrez.

Positive integer powers of certain tridiagonal matrices. Applied Mathematics and Computation, 202(1):133–140, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000660.

Gutierrez-Gutierrez:2008:PIPb

[GG08c] Jesús Gutiérrez-Gutiérrez.

Positive integer powers of complex skew-symmetric circulant matrices. Applied Mathematics and Computation, 202(2):798–802, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001835.

Gutierrez-Gutierrez:2008:PIPc

[GG08d] Jesús Gutiérrez-Gutiérrez.

Positive integer powers of complex symmetric circulant matrices. Applied Mathematics and Computation, 202(2):877–881, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000970.

Gutierrez-Gutierrez:2008:PRP

[GG08e] Jesús Gutiérrez-Gutiérrez. Powers of real persymmetric anti-tridiagonal matrices with constant antidiagonals. Applied Mathematics and Computation, 206(2):919–924, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007534.

Gutierrez-Gutierrez:2008:PTM

[GG08f] Jesús Gutiérrez-Gutiérrez. Powers of tridiagonal matrices with constant diagonals. Applied Mathematics and Computation, 206(2):885–891, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007480.

Gulati:2009:SSD

[GG09a] T. R. Gulati and Geeta. On some symmetric dual models in multiobjective programming. Applied Mathematics and Computation, 215(1):380–383, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004809.

Guo:2009:CTF

[GG09b] Weiping Guo and Qi Guo.

Convergence theorems of fixed points for asymptotically hemi-pseudocontractive mappings. Applied Mathematics and Computation, 215(1):16–21, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030900349X.

Gupta:2008:CHL

[GGOS08] Pushpa L. Gupta, Ramesh C.

Gupta, Seng-Huat Ong, and H. M. Srivastava. A class of Hurwitz–Lerch Zeta distributions and their applications in reliability. Applied Mathematics and Computation, 196(2):521– 531, March 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307007047.

Gutierrez:2006:SRD

[GGSN06] R. Gutiérrez, R. Gutiérrez-Sánchez, and A. Nafidi. The Stochastic Rayleigh diffusion model: Statistical inference and computational aspects. Applications to modelling of real cases. Applied Mathematics and Computation, 175(1):628–644, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006363.

Gutierrez:2006:NSG

[GGSNR06] R. Gutiérrez, R. Gutiérrez-Sánchez, A. Nafidi, and E. Ramos. A new stochastic Gompertz diffusion process with threshold parameter: Computational aspects and applications. Applied Mathematics and Computation, 183(2):738–747, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006084.

Gao:2006:CAM

[GH06] Zhong-Xi Gao and Ting-Zhu Huang. Convergence of AOR method. Applied Mathematics and Computation, 176(1):134–140, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305007794.

Ghatee:2007:DDA

[GH07] Mehdi Ghatee and S. Mehdi Hashemi. Descent direction algorithm with multicommodity flow problem for signal optimization and traffic assignment jointly. Applied Mathematics and Computation, 188(1):555– 566, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013956.

Guo:2009:LCG

[GH09] Zhenyuan Guo and Lihong Huang. LMI conditions for global robust stability of delayed neural networks with discontinuous neuron activations. Applied Mathematics and Computation, 215(3):889–900, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900575X.

Ghanbari:2008:SIG

[Gha08]  K. Ghanbari.  A sur-

vey on inverse and generalized inverse eigenvalue problems for Jacobi matrices. Applied Mathematics and Computation, 195(2):355–363, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005954.

Ghiyasvand:2006:NAC

[Ghi06] Mehdi Ghiyasvand. A new approach for computing a most positive cut using the minimum flow algorithms. Applied Mathematics and Computation, 176(1):27– 36, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305007678.

Geng:2007:ETW

[GHL07] Yixiang Geng, Tianlan He, and Jibin Li. Exact travelling wave solutions for the (n + 1)-dimensional double sine- and sinh-Gordon equations. Applied Mathematics and Computation, 188(2):1513–1534, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015049.

Guo:2008:EGS

[GHX08] Boling Guo, Yongqian Han, and Jie Xin. Existence of the global smooth solution to the period boundary value problem of fractional nonlinear Schrödinger equation. Applied Mathematics and Computation, 204(1):468–477, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005341.

Gong:2007:IGA

[GHZS07] Dun-Wei Gong, Guo-Sheng Hao, Yong Zhou, and Xiao-Yan Sun. Interactive genetic algorithms with multi-population adaptive hierarchy and their application in fashion design. Applied Mathematics and Computation, 185(2):1098–1108, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008344.

Gharbi:2006:GAR

[GI06] Nawel Gharbi and Malika Ioualalen. GSPN analysis of retrial systems with servers breakdowns and repairs. Applied Mathematics and Computation, 174(2):1151–1168, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005382.

Gupta:2008:RSA

[GI08] Vijay Gupta and Mircea Ivan. Rate of simultaneous approximation for the Bézier variant of certain operators. Applied Mathematics and Computation, 199(1):392–395, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009769.

Gine:2007:NAI

[Gin07] Jaume Giné. On the number of algebraically independent Poincaré–Liapunov constants. Applied Mathematics and Computation, 188(2):1870–1877, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015554.

Gzyl:2007:SBA

[GIT07] H. Gzyl, P. L. Novi Inverardi, and A. Tagliani. In search of the best approximant. Applied Mathematics and Computation, 189(1):652–661, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016407.

Gzyl:2006:MSF

[GITV06] Henryk Gzyl, Pier Luigi Novi Inverardi, Aldo Tagliani, and Minaya Villasana.

Maxentropic solution of fractional moment problems. Applied Mathematics and Computation, 173(1):109–125, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003139.

Ganesan:2005:PST

[GJ05] Sashikumaar Ganesan and Volker John. Pressure separation — a technique for improving the velocity error in finite element discretisations of the Navier– Stokes equations. Applied Mathematics and Computation, 165(2):275–290, June 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003819.

Gao:2007:AWM

[GJ07a] Jing Gao and Yao-Lin Jiang. An adaptive wavelet method for nonlinear differential-algebraic equations. Applied Mathematics and Computation, 189(1):208–220, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306015931.

Golbabai:2007:AHH

[GJ07b] A. Golbabai and M. Javidi. Application of He’s homotopy perturbation method for n thorder integro-differential equations. Applied Mathematics and Computation, 190(2):1409–1416, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001725.

Golbabai:2007:AHP

[GJ07c] A. Golbabai and M. Javidi. Application of homotopy perturbation method for solving eighthorder boundary value problems. Applied Mathematics and Computation, 191(2):334–346, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002615.

Golbabai:2007:NFI

[GJ07d] A. Golbabai and M. Javidi. A new family of iterative methods for solving system of nonlinear algebric equations. Applied Mathematics and Computation, 190(2):1717–1722, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002172.

Golbabai:2007:NIM

[GJ07e] A. Golbabai and M. Javidi. New iterative methods for nonlinear equations by modified HPM. Applied Mathematics and Computation, 191(1):122–127, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002342.

Golbabai:2007:NLI

[GJ07f] A. Golbabai and M. Javidi. Newton-like iterative methods for solving system of non-linear equations. Applied Mathematics and Computation, 192(2):546–551, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003578.

Golbabai:2007:NSN

[GJ07g] A. Golbabai and M. Javidi. A numerical solution for non-classical parabolic problem based on Chebyshev spectral collocation method. Applied Mathematics and Computation, 190(1):179–185, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000380.

Golbabai:2007:TON

[GJ07h] A. Golbabai and M. Javidi. A third-order Newton type method for nonlinear equations based on modified homotopy perturbation method. Applied Mathematics and Computation, 191(1):199–205, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002469.

Goghary:2005:RAM

[GJB05] H. Sadeghi Goghary, Sh. Javadi, and E. Babolian. Restarted Adomian method for system of nonlinear Volterra integral equations. Applied Mathematics and Computation, 161(3):745–751, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304000268.

Gongsheng:2007:ODE

[GJD^{+}07] Li Gongsheng, Cheng Jin, Yao De, Liu Hongliang, and Liu Jijun. Onedimensional equilibrium model and source parameter determination for soilcolumn experiment. Applied Mathematics and Computation, 190(2):1365– 1374, July 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307001671.

Garcia:2006:SDS

[GJL06] Luis David Garcia, Abdul Salam Jarrah, and Reinhard Laubenbacher. Sequential dynamical systems over words. Applied Mathematics and Computation, 174(1):500–510, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004819.

Geyikli:2005:AMK

[GK05a] Turabi Geyikli and Doğan Kaya. An application for a modified KdV equation by the decomposition method and finite element method. Applied Mathematics and Computation, 169(2):971–981, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008276.

Geyikli:2005:CSO

[GK05b] Turabi Geyikli and Doğan Kaya. Comparison of the solutions obtained by Bspline FEM and ADM of KdV equation. Applied Mathematics and Computation, 169(1):146–156, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304007660.

Ghodousian:2006:SLP

[GK06a] A. Ghodousian and E. Khorram. Solving a linear programming problem with the convex combination of the max–min and the max-average fuzzy relation equations. Applied Mathematics and Computation, 180(1):411–418, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306000038.

Ghodousian:2006:AOL

[GK06b] Amin Ghodousian and Esmaile Khorram. An algorithm for optimizing the linear function with fuzzy relation equation constraints regarding maxprod composition. Applied Mathematics and Computation, 178(2):502–509, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009975.

Gupta:2006:MIM

[GK06c] K. C. Gupta and Vinay Kanwar. Multipoint iterative method with cubic convergence. Applied Mathematics and Computation, 179(2):606–611, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010660.

Ghassemi:2009:NEV

[GK09] Hassan Ghassemi and Ahmad Reza Kohansal. Numerical evaluation of various levels of singular integrals, arising in BEM and its application in hydrofoil analysis. Applied Mathematics and Computation, 213(2):277–289, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002306.

Galue:2007:FET

[GKAS07] Leda Galue, S. L. Kalla, and B. N. Al-Saqabi. Fractional extensions of the temperature field problems in oil strata. Applied Mathematics and Computation, 186(1):35– 44, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009143.

Ghasemi:2007:AHH

[GKB07a] M. Ghasemi, M. Tavassoli Kajani, and E. Babolian. Application of He’s homotopy perturbation method to nonlinear integro-differential equations. Applied Mathematics and Computation, 188 (1):538–548, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306013932.

Ghasemi:2007:NSNb

[GkB07b] M. Ghasemi, M. Tavassoli kajani, and E. Babolian. Numerical solutions of the nonlinear integro-differential equations: Wavelet-Galerkin method and homotopy perturbation method. Applied Mathematics and Computation, 188(1):450–455, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013841.

Ghasemi:2007:NSNa

[GKB07c] M. Ghasemi, M. Tavassoli Kajani, and E. Babolian. Numerical solutions of the nonlinear Volterra–Fredholm integral equations by using homotopy perturbation method. Applied Mathematics and Computation, 188(1):446– 449, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601383X.

Ghasemi:2007:NST

[GKD07] M. Ghasemi, M. Tavassoli Kajani, and A. Davari. Numerical solution of twodimensional nonlinear differential equation by homotopy perturbation method. Applied Mathematics and Computation, 189(1):341– 345, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016079.

Gousidou-Koutita:2009:ENB

[GKK09] M. Gousidou-Koutita and T. J. Kalvouridis. On the efficiency of Newton and Broyden numerical methods in the investigation of the regular polygon problem of (N + 1) bodies. Applied Mathematics and Computation, 212(1):100– 112, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001234.

Geum:2006:HOC

[GKR06] Young Hee Geum, Young Ik Kim, and Min Surp Rhee. High-order convergence of the k-fold pseudo-Newton’s irrational method locating a simple real zero. Applied Mathematics and Computation, 182(1):492–497, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003560.

Gao:2005:FQM

[GL05a] Jinwu Gao and Mei Lu. Fuzzy quadratic minimum spanning tree problem. Applied Mathematics and Computation, 164(3):773–788, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003443.

Ge:2005:CSP

[GL05b] Z.-M. Ge and J.-K. Lee. Chaos synchronization and parameter identification for gyroscope system. Applied Mathematics and Computation, 163(2):667–682, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002516.

Gupta:2005:MFR

[GL05c] Ramesh C. Gupta and Sergey Lvin. Monotonicity of failure rate and mean residual life function of a gammatype model. Applied Mathematics and Computation, 165(3):623–633, June 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004072.

Gao:2006:ASS

[GL06a] Meng Gao and Zizhen Li. Asymptotical stability and spatial patterns of a spatial cyclic competitive system. Applied Mathematics and Computation, 183(2):1165–1169, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006515.

Gong:2006:RIB

[GL06b] Ning Gong and Tao Li. Role of index bonds in an optimal dynamic asset allocation model with real subsistence consumption. Applied Mathematics and Computation, 174(1):710–731, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004972.

Geng:2008:NSS

[GL08a] Fazhan Geng and Yingzhen Lin. Numerical solution of a system of fourth order boundary value problems using variational iteration method. Applied Mathematics and Computation, 200(1):231–241, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011022.

Geng:2008:EET

[GL08b] Yixiang Geng and Jibin Li. Exact explicit traveling wave solutions for the CDF equation. Applied Mathematics and Computation, 203(2):536–562, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002828.

Geng:2008:ESN

[GL08c] Yixiang Geng and Jibin Li. Exact solutions to a nonlinearly dispersive Schrödinger equation. Applied Mathematics and Computation, 195(2):420–439, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005632.

Gao:2009:EUP

[GL09] Hong Gao and Bingwen Liu. Existence and uniqueness of periodic solutions for forced Rayleightype equations. Applied Mathematics and Computation, 211(1):148–154, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000642.

Liu:2007:NAT

[gLL07] Xin guo Liu and Ma Li.

A note on the approximation of two-sided Rayleigh quotient to eigenvalues. Applied Mathematics and Computation, 184(2):523– 535, January 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300306007284.

Ge:2008:PAC

[GLLC08] Zheng-Ming Ge, Shih-Chung Li, Shih-Yu Li, and Ching-Ming Chang. Pragmatical adaptive chaos control from a new double van der Pol system to a new double Duffing system. Applied Mathematics and Computation, 203(2):513–522, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002804.

Gupta:2009:SAB

[GLMLP09] Vijay Gupta, Antonio-Jesús López-Moreno, and José-Manuel Latorre-Palacios. On simultaneous approximation of the Bernstein Durrmeyer operators. Applied Mathematics and Computation, 213(1):112– 120, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002082.

Gamez:2008:ISO

[GLV08] M. Gámez, I. López, and Z. Varga. Iterative scheme for the observation of a competitive Lotka– Volterra system. Applied Mathematics and Computation, 201(1–2):811–818, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011605.

Liu:2005:NBE

[gLxW05] Xin guo Liu and Ze xi Wang.

A note on the backward errors for Hermite eigenvalue problems. Applied Mathematics and Computation, 165(2):405–417, June 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400390X.

Li:2006:ACB

[gLXyZ06] Zhan guo Li, Wei Xu, and Xiao yan Zhang. Analysis of chaotic behavior in the extended Duffing–van der Pol system subject to additive non-symmetry biharmonical excitation. Applied Mathematics and Computation, 183(2):858–871, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006199.

Gonzalez:2005:PUN

[GM05a] Hernán A. González and Nelson O. Moraga. On predicting unsteady non-Newtonian blood flow. Applied Mathematics and Computation, 170(2):909– 923, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305000330.

Guseinov:2005:CGH

[GM05b] I. I. Guseinov and B. A. Mamedov. Calculation of the generalized Hubbell rectangular source integrals using binomial coefficients. Applied Mathematics and Computation, 161(1):285–292, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012505.

Gao:2006:SSA

[GM06a] Zhiming Gao and Yichen Ma. Shape sensitivity analysis for a Robin problem via minimax differentiability. Applied Mathematics and Computation, 181(2):1090–1105, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001937.

Guyer:2006:FCM

[GM06b] Tolga Güyer and Şeref Mirasyedioğlu. A functional computation model for the duality of twovariable Lambda-Boolean functions. Applied Mathematics and Computation, 182(1):879–887, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600405X.

Georgescu:2007:PRM

[GM07] Paul Georgescu and Gheorghe Moroşanu. Pest regulation by means of impulsive controls. Applied Mathematics and Computation, 190(1):790–803, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001130.

Guesmia:2008:CVD

[GM08] Aissa Guesmia and Salim A.

Messaoudi. On the control of a viscoelastic damped Timoshenko-type system. Applied Mathematics and Computation, 206(2):589–597, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003822.

Golbabaee:2006:WPT

[GMCM06] A. Golbabaee, M. R. Mokhtarzadeh, N. G. Chegini, and R. Mokhtary. Wavelet preconditioning for the three fields formulation: Numerical results in conforming decomposition. Applied Mathematics and Computation, 174(1):545–565, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004844.

Goswami:2006:DTB

[GMS06] V. Goswami, J. R. Mohanty, and S. K. Samanta. Discrete-time bulk-service queues with accessible and non-accessible batches. Applied Mathematics and Computation, 182(1):898– 906, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306004073.

Gakkhar:2006:MMV

[GN06a] Sunita Gakkhar and Kuldeep Negi. A mathematical model for viral infection in toxin producing phytoplankton and zooplankton system. Applied Mathematics and Computation, 179(1):301–313, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010386.

Gupta:2006:SRB

[GN06b] Arjun K. Gupta and Saralees Nadarajah. Sums and ratios for beta Stacy distribution. Applied Mathematics and Computation, 173(2):1310–1322, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004339.

Gupta:2006:SPR

[GN06c] Arjun K. Gupta and Saralees Nadarajah. Sums, products, and ratios for Freund’s bivariate exponential distribution. Applied Mathematics and Computation, 173(2):1334– 1349, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305004352.

Gupta:2007:IMS

[GN07] Arjun K. Gupta and Saralees Nadarajah. Information matrices for some bivariate Pareto distributions. Applied Mathematics and Computation, 184(2):1069–1079, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600796X.

Gulsu:2005:NSB

[GÖ05] Mustafa Gülsu and Turgut Öziş. Numerical solution of Burgers’ equation with restrictive Taylor approximation. Applied Mathematics and Computation, 171(2): 1192–1200, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Graef:2008:ESN

[GO08] John R. Graef and Abdelghani Ouahab. Extremal solutions for nonresonance impulsive functional dynamic equations on time scales. Applied Mathematics and Computation, 196(1):333–339, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006753.

Ghasabi-Oskoei:2006:ESN

[GOMA06] Hasan Ghasabi-Oskoei and Nezam Mahdavi-Amiri. An efficient simplified neural network for solving linear and quadratic programming problems. Applied Mathematics and Computation, 175(1):452–464, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500620X.

Gonzalez:2007:LPE

[Gon07] Graciela Adriana González.

On-line parameter estimator derived from modified matrix pseudoinverse. Applied Mathematics and Computation, 189(2):1139–1145, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016985.

Gopi:2007:DIF

[Gop07] E. S. Gopi. Digital image forgery detection using artificial neural network and independent component analysis. Applied Mathematics and Computation, 194(2):540–543, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005073.

Gageonea:2007:MMP

[GOPP07] Maria E. Gageonea, Shigeyoshi Owa, Radu N. Pascu, and Mihai N. Pascu. A maximum modulus principle for non-analytic functions defined in the unit disk. Applied Mathematics and Computation, 187(1):163– 169, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011477.

Gorain:2006:EED

[Gor06a] Ganesh C. Gorain. Exponential energy decay estimates for the solutions of n-dimensional Kirchhoff type wave equation. Applied Mathematics and Computation, 177(1):235– 242, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009057.

Gorguis:2006:CBC

[Gor06b] Alice Gorguis. A comparison between Cole– Hopf transformation and the decomposition method for solving Burgers’ equations. Applied Mathematics and Computation, 173(1):126–136, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003140.

Gorguis:2007:CAS

[Gor07] Alice Gorguis. Charpit and Adomian for solving integral equations. Applied Mathematics and Computation, 193(2):446–454, November 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307004183.

Gaetano:2009:PAE

[GPR^{+}09] A. De Gaetano, S. Panunzi, F. Rinaldi, A. Risi, and M. Sciandrone. A patient adaptable ECG beat classifier based on neural networks. Applied Mathematics and Computation, 213(1):243–249, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002252.

Guido:2009:ISI

[GPS09] Rodrigo Capobianco Guido, José Carlos Pereira, and Jan Frans Willem Slaets. Introduction to the special issue on Emergent Applications of Fractals and Wavelets in Biology and Biomedicine. Applied Mathematics and Computation, 207(1):3–4, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000155.

Gouveia:2009:TDB

[GPT09] Paulo D. F. Gouveia, Alexander Plakhov, and Delfim F. M. Torres. Twodimensional body of maximum mean resistance. Applied Mathematics and Computation, 215(1):37– 52, September 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309003646.

Gladish:2005:MSW

[GPTU05] Blanca Pérez Gladish, Mar Arenas Parra, Amelia Bilbao Terol, and M. Victoria Rodríuez Uría. Management of surgical waiting lists through a Possibilistic Linear Multiobjective Programming problem. Applied Mathematics and Computation, 167(1):477–495, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005053.

Guo:2008:SCL

[GQS08] Senlin Guo, Feng Qi, and H. M. Srivastava. Supplements to a class of logarithmically completely monotonic functions associated with the gamma function. Applied Mathematics and Computation, 197(2):768– 774, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008363.

Goncalves:2006:BBS

[GR06a] J. V. Goncalves and Angelo Roncalli. Boundary blow-up solutions for a class of elliptic equations on a bounded domain. Applied Mathematics and Computation, 182(1):13–23, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003079.

Guo:2006:CNL

[GR06b] Hui Guo and Hongxing Rui. Crank–Nicolson leastsquares Galerkin procedures for parabolic integrodifferential equations. Applied Mathematics and Computation, 180(2):622– 634, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306000518.

Gajic:2007:PNS

[GR07a] Ljiljana Gajić and Vladimir Rakočević. Pair of nonself-mappings and common fixed points. Applied Mathematics and Computation, 187(2):999–1006, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012367.

Guo:2007:LSG

[GR07b] Hui Guo and Hongxing Rui. Least-squares Galerkin procedures for pseudohyperbolic equations. Applied Mathematics and Computation, 189 (1):425–439, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030601616X.

Graca:2009:RMD

[Gra09] Mário M. Graça. Removing multiplicities in C by double newtonization. Applied Mathematics and Computation, 215(2):562–572, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900513X.

Greubel:2006:SGF

[Gre06] G. C. Greubel. On some generating functions for the extended generalized Hermite polynomials. Applied Mathematics and Computation, 173(1):547–553, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003747.

Guine:2007:MSP

[GRF07] R. P. F. Guiné, A. E.

Rodrigues, and M. M. Figueiredo. Modelling and simulation of pear drying. Applied Mathematics and Computation, 192(1):69– 77, September 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307002950.

Guo:2007:EHO

[GRH07] Yecai Guo, Wei Rao, and Yinge Han. Extraction of higher-order coupling feature using three and one half dimension spectrum. Applied Mathematics and Computation, 185(2):798–809, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007624.

Garg:2007:GFH

[GRK07] Mridula Garg, Alka Rao, and S. L. Kalla. On a generalized finite Hankel transform. Applied Mathematics and Computation, 190(1):705–711, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001075.

Gastaldo:2006:PQC

[GRZ06] Paolo Gastaldo, Sandro Ridella, and Rodolfo Zunino. Prospects of quantum-classical optimization for digital design. Applied Mathematics and Computation, 179(2):581–595, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010647.

Gulsu:2005:ASH

[GS05] Mustafa Gülsu and Mehmet Sezer. The approximate solution of high-order linear difference equations with variable coefficients in terms of Taylor polynomials. Applied Mathematics and Computation, 168(1): 76–88, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304005879.

Ghorbani:2006:CFS

[GS06a] Mehrzad Ghorbani and Ali Reza Soheili. Corrected fundamental solution for numerical solution of elliptic PDEs. Applied Mathematics and Computation, 181(1):175–184, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000877.

Golbabai:2006:ISS

[GS06b] A. Golbabai and S. Seifollahi. An iterative solution for the second kind integral equations using radial basis functions. Applied Mathematics and Computation, 181(2):903–907, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001706.

Golbabai:2006:NSS

[GS06c] A. Golbabai and S. Seifollahi. Numerical solution of the second kind integral equations using radial basis function networks. Applied Mathematics and Computation, 174(2):877–883, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005175.

Gulsu:2006:SRE

[GS06d] Mustafa Gülsu and Mehmet Sezer. On the solution of the Riccati equation by the Taylor matrix method. Applied Mathematics and Computation, 176(2):414–421, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008179.

Gupta:2006:AMS

[GS06e] Pankaj Gupta and Sanjeet Singh. Approximate multiparametric sensitivity analysis of the constraint matrix in linear-plus-linear fractional programming problem. Applied Mathematics and Computation, 179(2):662–671, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010714.

Gupta:2006:CQL

[GS06f] U. C. Gupta and Karabi Sikdar. Computing queue length distributions in MAP / G / 1 / N queue under single and multiple vacation. Applied Mathematics and Computation, 174(2):1498–1525, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500576X.

Golbabai:2007:RBF

[GS07a] A. Golbabai and S. Seifollahi. Radial basis function networks in the numerical solution of linear integrodifferential equations. Applied Mathematics and Computation, 188(1):427– 432, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013804.

Guan:2007:FON

[GS07b] Kaizhong Guan and Jianhua Shen. On first-order neutral differential equations of Euler form with unbounded delays. Applied Mathematics and Computation, 189(2):1419–1427, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017280.

Gulsu:2007:ASL

[GS07c] Mustafa Gülsu and Mehmet Sezer. Approximate solution to linear complex differential equation by a new approximate approach. Applied Mathematics and Computation, 185(1):636–645, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008988.

GomezS:2008:CHT

[GS08a] Cesar A. Gómez S. and Alvaro H. Salas. The Cole–Hopf transformation and improved tanh– coth method applied to new integrable system (KdV6). Applied Mathematics and Computation, 204(2):957–962, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005869.

Gupta:2008:TIC

[GS08b] Urvashi Gupta and Gaurav Sharma. Thermosolutal instability of a compressible Rivlin–Erickson fluid in the presence of rotation and Hall currents saturating a porous medium. Applied Mathematics and Computation, 196(1):158–173, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006558.

Gupta:2008:DRC

[GS08c] Vijay Gupta and Jyoti Sinha. Direct results on certain Szasz–Mirakyan operators. Applied Mathematics and Computation, 195(1):230–239, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005334.

Gao:2009:SGM

[GS09] Ji Gao and Satit Saejung. Some geometric measures of spheres in Banach spaces. Applied Mathematics and Computation, 214(1):102–107, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002859.

Grau-Sanchez:2006:CEM

[GSDB06] Miquel Grau-Sánchez and José Luis Díaz-Barrero. On computational efficiency for multi-precision zerofinding methods. Applied Mathematics and Computation, 181(1):402–412, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600110X.

Grau-Sanchez:2007:HSF

[GSDB07a] Miquel Grau-Sánchez and José-Luis Díaz-Barrero. A Hummel & Seebeck family of iterative methods with improved convergence and efficiency. Applied Mathematics and Computation, 191(1):263–271, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002536.

Grau-Sanchez:2007:WVF

[GSDB07b] Miquel Grau-Sánchez and José Luis Díaz-Barrero. A weighted variant family of Newton’s method with accelerated thirdorder convergence. Applied Mathematics and Computation, 186(2):1005–1009, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010265.

Grau-Sanchez:2009:SSO

[GSDB09] Miquel Grau-Sánchez and José Luis Díaz-Barrero. Some sixth order zerofinding variants of Chebyshev– Halley methods. Applied Mathematics and Computation, 211(1):108–110, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000459.

Gulsu:2006:ASG

[GSG06] Mustafa Gülsu, Mehmet Sezer, and Zekeriya Güney. Approximate solution of general high-order linear nonhomogeneous difference equations by means of Taylor collocation method. Applied Mathematics and Computation, 173(2):683– 693, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305003814.

Grau-Sanchez:2009:ZFM

[GSG09] Miquel Grau-Sánchez and José M. Gutiérrez. Zerofinder methods derived from Obreshkov’s techniques. Applied Mathematics and Computation, 215(8):2992–3001, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309008716.

Grau-Sanchez:2009:ALT

[GSNDB09] Miquel Grau-Sánchez, Miquel Noguera, and José Luis Díaz-Barrero. Adamslike techniques for zerofinder methods. Applied Mathematics and Computation, 211(1):130–136, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000496.

Grau-Sanchez:2007:AIM

[GSPG07] Miquel Grau-Sánchez, Josep M. Peris, and José M. Gutiérrez. Accelerated iterative methods for finding solutions of a system of nonlinear equations. Applied Mathematics and Computation, 190(2):1815–1823, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002317.

Guo:2005:IBD

[GT05a] Xing-Yang Guo and Chao-Jing Tang. Identity based designated multi-confirmer signature: a new type signature. Applied Mathematics and Computation, 169(2):1076–1085, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008380.

Guo:2005:IBG

[GT05b] Xing-Yang Guo and Chao-Jing Tang. Identity based group-oriented undeniable signature scheme. Applied Mathematics and Computation, 169(2):1448–1457, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008653.

Gencer:2006:NIA

[GTA06] Cevriye Gencer, İsmail Top, and Emel Kizilkaya Aydogan. A new intuitional algorithm for solving heterogeneous fixed fleet routing problems: Passenger pickup algorithm. Applied Mathematics and Computation, 181(2):1552–1567, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002864.

Ghassemi-Tari:2007:DHD

[GTJ07] Farhad Ghassemi-Tari and Eshagh Jahangiri. Development of a hybrid dynamic programming approach for solving discrete nonlinear knapsack problems. Applied Mathematics and Computation, 188(1):1023–1030, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014500.

Gao:2008:EPV

[GTX08] Shujing Gao, Zhidong Teng, and Dehui Xie. The effects of pulse vaccination on SEIR model with two time delays. Applied Mathematics and Computation, 201(1–2):282–292, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011964.

Gu:2007:PMN

[Gu07a] Wei Gu. A probabilistic method for numerical solution of quasilinear parabolic equations. Applied Mathematics and Computation, 189(2):1982–1992, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000173.

Gungor:2007:HMD

[GÜ07b] Zülal Güngör and Alper Ünler. K-harmonic means data clustering with simulated annealing heuristic. Applied Mathematics and Computation, 184(2):199–209, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006953.

Guine:2005:DAB

[Gui05] Raquel P. F. Guiné. Development of an algorithm based on space refinement to solve a system of parabolic PDE’s. Applied Mathematics and Computation, 161(3):963– 972, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304000463.

Gulkac:2005:FDS

[Gül05] Vildan Gülkaç. On the finite differences schemes for the numerical solution of two-dimensional moving boundary problem. Applied Mathematics and Computation, 168(1):549–556, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030400623X.

Gulsu:2006:FDA

[Gül06]  Mustafa Gülsu.  A fi-

nite difference approach for solution of Burgers’ equation. Applied Mathematics and Computation, 175(2):1245–1255, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006740.

Gul:2008:RTS

[Gül08] Erdal Gül. On the regularized trace of a second order differential operator. Applied Mathematics and Computation, 198(2):471– 480, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008831.

Guo:2009:MSG

[Guo09] Yingxin Guo. Mean square global asymptotic stability of stochastic recurrent neural networks with distributed delays. Applied Mathematics and Computation, 215(2):791–795, September 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309005669.

Gupta:2008:RAB

[Gup08a] Vijay Gupta. Rate of approximation for the Balazs–Kantorovich–Bézier operators. Applied Mathematics and Computation, 199(2):823–827, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307010600.

Gupta:2008:SAP

[Gup08b]  Vijay Gupta.  Some

approximation properties of q-Durrmeyer operators. Applied Mathematics and Computation, 197(1):172–178, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007667.

Gurcan:2005:STN

[Gür05] Fuat Gürcan. Streamline topologies near a stationary wall of Stokes flow in a cavity. Applied Mathematics and Computation, 165(2):329–345, June 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003856.

Guyker:2009:UCI

[Guy09] James Guyker. On the uniform convergence of interpolating polynomials. Applied Mathematics and Computation, 210(1):11– 17, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307007539.

Gonzalez-Vera:2007:ASQ

[GVMT07] P. González-Vera, H. Martínez, and J. J. Trujillo. An application of Szegő quadratures to the computation of the Fourier transform. Applied Mathematics and Computation, 187(1):183– 194, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011507.

Gu:2005:ADI

[Gw05] Haiming Gu and Xiaonan wu. Alternatingdirection iterative technique for mixed finite element methods for Stokes equations. Applied Mathematics and Computation, 162(3):1035–1047, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400147X.

Gu:2006:CCR

[GW06a] Chao Gu and Guorong Wang. Condensed Cramer rule for solving restricted matrix equations. Applied Mathematics and Computation, 183(1):301–306, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005625.

Gu:2006:NFF

[GW06b] Y. H. Gu and Z. Y.

Wu. A new filled function method for nonlinear integer programming problem. Applied Mathematics and Computation, 173(2):938–950, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004030.

Gu:2007:FPC

[GW07] Hong Gu and Hongwei Wang. Fuzzy prediction of chaotic time series based on singular value decomposition. Applied Mathematics and Computation, 185(2):1171–1185, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008411.

Guo:2007:FOL

[GWJ07] Wenbin Guo, Musheng Wei, and Zhao Jianli. Forward order law for ginverses of the product of two matrices. Applied Mathematics and Computation, 189(2):1749–1754, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017668.

Gu:2006:PAP

[GWX06] Chao Gu, Guorong Wang, and Zhaoliang Xu. PCR algorithm for the parallel computation of the solution of a class of singular linear systems. Applied Mathematics and Computation, 176(1):237–244, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305007885.

Gao:2005:NRB

[GXS05] Yuelin Gao, Honggang Xue, and Peiping Shen. A new rectangle branch-and-reduce approach for solving nonconvex quadratic programming problems. Applied Mathematics and Computation, 168(2):1409– 1418, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304007398.

Gao:2006:OSF

[GXY06] Yuelin Gao, Chengxian Xu, and Yongjian Yang. An outcome-space finite algorithm for solving linear multiplicative programming. Applied Mathematics and Computation, 179(2):494–505, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010556.

Gao:2006:UFV

[GY06] Fuzheng Gao and Yirang Yuan. The upwind finite volume element method based on straight triangular prism partition for nonlinear convection-diffusion problem. Applied Mathematics and Computation, 181(2):1229–1242, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306002426.

Gozukizil:2007:LTB

[GY07a] Ömer Faruk Gözükızıl and Metin Yaman. Longtime behaviour of the solutions to inverse problems for parabolic equations. Applied Mathematics and Computation, 184(2):669–673, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007417.

Gu:2007:GSA

[GY07b] Chuanqing Gu and Zhenxin Yang. Global SCD algorithm for real positive definite linear systems with multiple right-hand sides. Applied Mathematics and Computation, 189(1):59– 67, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015797.

Gao:2008:PRP

[GY08] Heli Gao and Chuancun Yin. A perturbed risk process compounded by a geometric Brownian motion with a dividend barrier strategy. Applied Mathematics and Computation, 205(1):454–464, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006061.

Gan:2005:SWN

[GZ05] Zaihui Gan and Jian Zhang. Standing waves for nonlinear Klein–Gordon equations with nonnegative potentials. Applied Mathematics and Computation, 166(3):551–570, July 26, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004576.

Gao:2007:CSS

[GZ07a] Chun-Yi Gao and Shi-Qiong Zhou. Certain subclass of starlike functions. Applied Mathematics and Computation, 187(1):176– 182, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011490.

Ge:2007:ACF

[GZ07b] Zheng-Ming Ge and An-Ray Zhang. Anticontrol of chaos of the fractional order modified van der Pol systems. Applied Mathematics and Computation, 187(2):1161–1172, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012537.

Geng:2007:MRL

[GZ07c] Fengjie Geng and Deming Zhu. Multiple results of p-Laplacian dynamic equations on time scales. Applied Mathematics and Computation, 193(2):311–320, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307003992.

Guo:2007:SCT

[GZ07d] Hongmei Guo and Shouming Zhong. Synchronization criteria of time-delay feedback control system with sector-bounded nonlinearity. Applied Mathematics and Computation, 191(2):550–559, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002834.

Gao:2009:NAE

[GZ09a] Hua Gao and Rong-Xia Zhao. New application of the (\frac{G′}{G})-expansion method to higher-order nonlinear equations. Applied Mathematics and Computation, 215(7):2781–2786, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007693.

Gao:2009:NAI

[GZ09b] Hua Gao and Rong-Xia Zhao. New application of the (G′/G)-expansion method to higher-order nonlinear equations. Applied Mathematics and Computation, 215(7):2781– 2786, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309007693.

Georgescu:2008:BNP

[GZC08] Paul Georgescu, Hong Zhang, and Lansun Chen. Bifurcation of nontrivial periodic solutions for an impulsively controlled pest management model. Applied Mathematics and Computation, 202(2):675– 687, August 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308001562.

Gao:2008:DDS

[GZG08] Fengyin Gao, Shouming Zhong, and Xiuzhi Gao. Delay-dependent stability of a type of linear switching systems with discrete and distributed time delays. Applied Mathematics and Computation, 196(1):24–39, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006248.

Guo:2009:DPC

[GZG09] Hongmei Guo, Shouming Zhong, and Fengyin Gao. Design of PD controller for master–slave synchronization of Lur’e systems with time-delay. Applied Mathematics and Computation, 212(1):86– 93, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001210.

Gao:2008:DSI

[GZL08] Qingwei Gao, Yanfei Zhao, and Yixiang Lu. Despeckling SAR images using stationary wavelet transform combining with directional filter banks. Applied Mathematics and Computation, 205(2):517–524, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003305.

Zhou:2008:CLS

[gZW08] Xue gang Zhou and Kun Wu. A comment on “Level set algorithm for solving convex multiplicative programming problems”. Applied Mathematics and Computation, 204(1):496–498, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004578.

See [You05].

Hadizadeh:2005:ENA

[HA05] M. Hadizadeh and M. Asgary. An efficient numerical approximation for the linear class of mixed integral equations. Applied Mathematics and Computation, 167(2):1090–1100, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400551X.

Hayat:2007:MDP

[HA07] T. Hayat and N. Ali.

A mathematical description of peristaltic hydromagnetic flow in a tube. Applied Mathematics and Computation, 188(2):1491–1502, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015025.

Hayat:2007:PMB

[HAA07] T. Hayat, N. Ali, and S. Asghar. Peristaltic motion of a Burger’s fluid in a planar channel. Applied Mathematics and Computation, 186(1):309–329, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009428.

Haas:2009:VRM

[Haa09] Markus Haas. Value-at-risk via mixture distributions reconsidered. Applied Mathematics and Computation, 215(6):2103–2119, November 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030900719X.

Hayat:2006:EPF

[HAAS06] T. Hayat, N. Ali, S. Asghar, and A. M. Siddiqui. Exact peristaltic flow in tubes with an endoscope. Applied Mathematics and Computation, 182(1):359–368, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003432.

Haidar:2007:SDI

[Hai07] Nassar H. S. Haidar.

On a symmetrized definite integration. Applied Mathematics and Computation, 186(1):70– 73, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009179.

See corrigendum [Hai08].

Haidar:2008:CSD

[Hai08] Nassar H. S. Haidar. Corrigendum to “On a symmetrized definite integration” [Applied Mathematics and Computation 186 (2007) 70–73]. Applied Mathematics and Computation, 204(1):513, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004372.

See [Hai07].

Hajjafar:2006:CRM

[Haj06] Ali Hajjafar. Controlled over-relaxation method and the general extrapolation method. Applied Mathematics and Computation, 174(1):188–198, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004509.

Hajji:2007:ASN

[HAK07a] Mohamed Ali Hajji and Kamel Al-Khaled. Analytic studies and numerical simulations of the generalized Boussinesq equation. Applied Mathematics and Computation, 191(2):320–333, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002603.

Hajji:2007:TRM

[HAK07b] Mohamed Ali Hajji and Kamel Al-Khaled. Two reliable methods for solving nonlinear evolution equations. Applied Mathematics and Computation, 186(2):1151–1162, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010411.

Hayat:2004:SNL

[HAM04] T. Hayat, Naseer Ahmad, and F. M. Mahomed. On solution of non-linear differential equation arising in gliding motion of bacteria. Applied Mathematics and Computation, 148(3): 743–752, January 30, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See comment [HES05].

Hamdi:2005:DSC

[Ham05a] Abdelouahed Hamdi. Decomposition for structured convex programs with smooth multiplier methods. Applied Mathematics and Computation, 169(1):218–241, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007714.

Hamdi:2005:PDP

[Ham05b] Abdelouahed Hamdi. A primal-dual proximal point algorithm for constrained convex programs. Applied Mathematics and Computation, 162(1):293–303, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400089X.

Hamdi:2005:TLP

[Ham05c] Abdelouahed Hamdi. Twolevel primal–dual proximal decomposition technique to solve large scale optimization problems. Applied Mathematics and Computation, 160(3):921–938, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303012037.

Hamzacebi:2008:IGA

[Ham08] Coşkun Hamzaçebi. Improving genetic algorithms’ performance by local search for continuous function optimization. Applied Mathematics and Computation, 196(1):309–317, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700673X.

Han:2006:SLV

[Han06] Deren Han. Solving linear variational inequality problems by a self-adaptive projection method. Applied Mathematics and Computation, 182(2):1765–1771, November 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0096300306005236.

Hanna:2007:STL

[Han07] Yousry S. Hanna. On the solutions of tridiagonal linear systems. Applied Mathematics and Computation, 189(2):2011–2016, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000203.

Hartfield:2006:ISD

[Har06] Roy J. Hartfield, Jr. Interpretation of spectroscopic data from the iodine molecule using a genetic algorithm. Applied Mathematics and Computation, 177(2):597–605, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009409.

Hascelik:2006:GQR

[Has06a] A. Ihsan Hascelik. Gauss quadrature rules for a generalized Hermite weight function. Applied Mathematics and Computation, 180(1):86–96, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305011008.

Hashim:2006:CNA

[Has06b]  Ishak Hashim.  Com-

ments on “A new algorithm for solving classical Blasius equation” by L. Wang. Applied Mathematics and Computation, 176(2):700–703, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008696.

See [Wan04].

Hashim:2006:CNM

[Has06c]  Ishak Hashim.  Com-

paring numerical methods for the solutions of two-dimensional diffusion with an integral condition. Applied Mathematics and Computation, 181(2):880–885, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001676.

Hassan:2006:AEN

[Has06d] I. H. Abdel-Halim Hassan. The asymptotic expansion and numerical verification method for linear and nonlinear initial value problem. Applied Mathematics and Computation, 180(1): 29–37, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305010945.

Hashemiparast:2007:NIU

[Has07a] S. M. Hashemiparast. Numerical integration using local Taylor expansions in nodes. Applied Mathematics and Computation, 192(2):332–336, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003335.

Hussain:2007:AFL

[HAS07b] Tauqeer Hussain, Mian M.

Awais, and Shafay Shamail. Applying fuzzy logic to measure completeness of a conceptual model. Applied Mathematics and Computation, 185(2):1078–1086, February 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306008319.

Hossain:2005:GMM

[HAU05] M. Zubaer Hossain, S. Reaz Ahmed, and M. Wahhaj Uddin. Generalized mathematical model for the solution of mixed-boundary-value elastic problems. Applied Mathematics and Computation, 169(2):1247– 1275, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304008495.

Bao:2006:EDP

[hB06] Zhen hua Bao. The expected discounted penalty at ruin in the risk process with random income. Applied Mathematics and Computation, 179(2):559–566, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010611.

Bao:2007:NCB

[hB07] Zhen hua Bao. A note on the compound binomial model with randomized dividend strategy. Applied Mathematics and Computation, 194(1):276–286, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004675.

Haciefendioglu:2007:NAR

[HBBA07] Kemal Hacıefendioğlu, Hasan Basri Başağa, Alemdar Bayraktar, and Şevket Ateş. Nonlinear analysis of rock-fill dams to non-stationary excitation by the stochastic Wilson-θ method. Applied Mathematics and Computation, 194(2):333–345, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004742.

Hashish:2009:SDH

[HBE09] H. Hashish, S. H. Behiry, and A. Elsaid. Solving the 2-D heat equations using wavelet-Galerkin method with variable time step. Applied Mathematics and Computation, 213(1):209– 215, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002203.

Hashish:2009:NIU

[HBES09] H. Hashish, S. H. Behiry, and N. A. El-Shamy. Numerical integration using wavelets. Applied Mathematics and Computation, 211(2):480–487, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001064.

Hartfield:2006:SMD

[HBJ06] R. J. Hartfield, J. E.

Burkhalter, and R. M. Jenkins. Scramjet missile design using genetic algorithms. Applied Mathematics and Computation, 174(2):1539–1563, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005783.

Huang:2006:ECC

[HBN06] Zhenyu Huang, Fanwei Bu, and Muhammad Aslam Noor. On the equivalence of the convergence criteria between modified Mann–Ishikawa and multistep iterations with errors for successively strongly pseudo-contractive operators. Applied Mathematics and Computation, 180(2):641–647, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000531.

Hemavathi:2007:PUN

[HBVM07] S. Hemavathi, T. Bhuvaneswari, S. Valarmathi, and J. J. H. Miller. A parameter uniform numerical method for a system of singularly perturbed ordinary differential equations. Applied Mathematics and Computation, 191(1):1–11, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006989.

Bao:2007:GSD

[hBxY07] Zhen hua Bao and Zhong xing Ye. The Gerber– Shiu discounted penalty function in the delayed renewal risk process with random income. Applied Mathematics and Computation, 184(2):857–863, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007739.

Honggang:2005:BBA

[HC05a] Xue Honggang and Xu Chengxian. A branch and bound algorithm for solving a class of D-C programming. Applied Mathematics and Computation, 165(2):291–302, June 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003820.

Huang:2005:OFC

[HC05b] Zhenkun Huang and Shuili Chen. Optimal fuzzy control of a poisoning-pest model. Applied Mathematics and Computation, 171 (2):730–737, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ham:2007:FOI

[HC07a] YoonMee Ham and Changbum Chun. A fifthorder iterative method for solving nonlinear equations. Applied Mathematics and Computation, 194(1):287–290, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004420.

Harlin:2007:BMT

[HC07b] W. J. Harlin and D. A. Cicci. Ballistic missile trajectory prediction using a state transition matrix. Applied Mathematics and Computation, 188(2):1832–1847, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015517.

He:2008:PSS

[HC08a] Tieshan He and Wenge Chen. Periodic solutions of second order discrete convex systems involving the p-Laplacian. Applied Mathematics and Computation, 206(1):124–132, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006371.

Hsieh:2008:DBS

[HC08b] Sun-Yuan Hsieh and Ming-Yu Chen. A DNA-based solution to the graph isomorphism problem using Adleman–Lipton model with stickers. Applied Mathematics and Computation, 197(2):672–686, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008247.

Huang:2008:NIT

[HC08c] Zhi-Kai Huang and Kwok-Wing Chau. A new image thresholding method based on Gaussian mixture model. Applied Mathematics and Computation, 205(2):899–907, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800369X.

Hu:2009:TDS

[HC09] Xiao-Rui Hu and Yong Chen. Two-dimensional symmetry reduction of (2 + 1)-dimensional nonlinear Klein–Gordon equation. Applied Mathematics and Computation, 215(3):1141–1145, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006262.

Huang:2006:MST

[HCC06] Ting-Zhu Huang, Guang-Hui Cheng, and Xiao-Yu Cheng. Modified SOR-type iterative method for Z-matrices. Applied Mathematics and Computation, 175(1):258–268, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006053.

Hermoso-Carazo:2007:DAS

[HCLP07a] A. Hermoso-Carazo and J. Linares-Pérez. Different approaches for state filtering in nonlinear systems with uncertain observations. Applied Mathematics and Computation, 187(2):708–724, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601126X.

Hermoso-Carazo:2007:EUF

[HCLP07b] A. Hermoso-Carazo and J. Linares-Pérez. Extended and unscented filtering algorithms using one-step randomly delayed observations. Applied Mathematics and Computation, 190(2):1375–1393, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001695.

Hermoso-Carazo:2008:RFP

[HCLPJL^{+}08] A. Hermoso-Carazo, J. Linares-Pérez, J. D. Jiménez-López, R. Caballero-Águila, and S. Nakamori. Recursive fixed-point smoothing algorithm from covariances based on uncertain observations with correlation in the uncertainty. Applied Mathematics and Computation, 203(1):243–251, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002452.

Huang:2005:CSC

[HCS05] De-Shuang Huang, Zheru Chi, and Wan-Chi Siu. A case study for constrained learning neural root finders. Applied Mathematics and Computation, 165(3):699–718, June 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004126.

Huang:2006:PPS

[HCW06] Zhenkun Huang, Fengde Chen, and Xinghua Wang. A Predator–Prey system with viral infection and anorexia response. Applied Mathematics and Computation, 175(2):1455–1483, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006892.

Huang:2006:IIE

[HCX06] Zhenkun Huang, Shuili Chen, and Yonghui Xia. Incorporate intelligence into an ecological system: An adaptive fuzzy control approach. Applied Mathematics and Computation, 177(1):243–250, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009069.

Cui:2005:MCB

[hCxC05] Li hong Cui and Zheng xing Cheng. A method of construction for biorthogonal multiwavelets system with 2r multiplicity. Applied Mathematics and Computation, 167(2):901–918, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005387.

Huang:2006:SAP

[HCZ06] Yunjin Huang, Fengde Chen, and Li Zhong. Stability analysis of a prey–predator model with Holling type III response function incorporating a prey refuge. Applied Mathematics and Computation, 182(1):672–683, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003766.

Hou:2007:AAS

[HD07] Zai-En Hou and Fu-Jian Duan. The approximation algorithm for solving a sort of non-smooth programming. Applied Mathematics and Computation, 186(2):1511–1519, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010770.

Hashemi:2007:NSG

[HDG07] S. H. Hashemi, H. R. Mohammadi Daniali, and D. D. Ganji. Numerical simulation of the generalized Huxley equation by He’s homotopy perturbation method. Applied Mathematics and Computation, 192(1):157–161, September 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307003049.

He:2006:SCM

[He06] Yinnian He. Stokes coupling method for the exterior flow Part IV: stabilized finite element approximation. Applied Mathematics and Computation, 172(2):1225–1238, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305002894.

Hadj:2008:FNA

[HE08a] Ahmed Driss Aiat Hadj and Mohamed Elouafi. A fast numerical algorithm for the inverse of a tridiagonal and pentadiagonal matrix. Applied Mathematics and Computation, 202(2):441–445, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001239.

Hadj:2008:CPE

[HE08b] Ahmed Driss Aiat Hadj and Mohamed Elouafi. On the characteristic polynomial, eigenvectors and determinant of a pentadiagonal matrix. Applied Mathematics and Computation, 198(2):634–642, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009319.

Hashemiparast:2006:DNN

[HED06a] S. M. Hashemiparast, M. R. Eslahchi, and Mehdi Dehghan. Determination of nodes in numerical integration rules using difference equation. Applied Mathematics and Computation, 176(1):117–122, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007769.

Hashemiparast:2006:MEF

[HED06b] S. M. Hashemiparast, M. R. Eslahchi, and Mehdi Dehghan. Minimizing the error function of Gauss– Jacobi quadrature rule with respect to parameters α and β. Applied Mathematics and Computation, 176(1):58– 64, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007708.

Hashemiparast:2006:NEC

[HED06c] S. M. Hashemiparast, M. R. Eslahchi, and Mehdi Dehghan. A note on equal coefficient quadrature rules. Applied Mathematics and Computation, 180(1):153–159, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305011070.

Hashemiparast:2006:NIU

[HED06d] S. M. Hashemiparast, M. R. Eslahchi, and Mehdi Dehghan. Numerical integration using the derivatives. Applied Mathematics and Computation, 180(1):194–199, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305011136.

Hashemiparast:2006:FKC

[HEDMJ06] S. M. Hashemiparast, M. R. Eslahchi, Mehdi Dehghan, and M. Masjed-Jamei. The first kind Chebyshev–Newton–Cotes quadrature rules (semiopen type) and its numerical improvement. Applied Mathematics and Computation, 174(2):1020–1032, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005291.

Haroon:2005:CSN

[HES05] Mohammed Haroon and Moustafa El-Shahed. Comment on: “On solution of non-linear differential equation arising in gliding motion of bacteria”, Applied Mathematics and Computation 148 (2004) 743–752. Applied Mathematics and Computation, 171(2):675–676, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [HAM04].

Habib:2007:ASS

[HEZ07] H. M. Habib and E. R.

El-Zahar. An algorithm for solving singular perturbation problems with mechanization. Applied Mathematics and Computation, 188(1):286–302, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601352X.

Habib:2008:VSS

[HEZ08] H. M. Habib and E. R. El-Zahar. Variable step size initial value algorithm for singular perturbation problems using locally exact integration. Applied Mathematics and Computation, 200(1):330–340, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011216.

Han:2006:WCG

[HFL06] B. Han, G. F. Feng, and J. Q. Liu. A widely convergent generalized pulse-spectrum technique for the inversion of two-dimensional acoustic wave equation. Applied Mathematics and Computation, 172(1):406–420, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002110.

He:2009:GSR

[HG09] Zhenhua He and Feng Gu. Generalized system for relaxed cocoercive mixed variational inequalities in Hilbert spaces. Applied Mathematics and Computation, 214(1):26–30, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002744.

Hashemi:2006:FGP

[HGH06] S. Mehdi Hashemi, Mehdi Ghatee, and Behnam Hashemi. Fuzzy goal programming: Complementary slackness conditions and computational schemes. Applied Mathematics and Computation, 179(2):506–522, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010568.

Hashemi:2006:CAM

[HGN06] S. Mehdi Hashemi, Mehdi Ghatee, and Ebrahim Nasrabadi. Combinatorial algorithms for the minimum interval cost flow problem. Applied Mathematics and Computation, 175(2):1200–1216, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006703.

Hayek:2007:VST

[HGN07] N. Hayek, B. J. González, and E. R. Negrin. A variant of the Stieltjes transform on distributions of compact support. Applied Mathematics and Computation, 187(1):202–206, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011520.

Hwang:2005:IHA

[HH05] Shin-Jia Hwang and Chia-Wei Huang. Improvement on Hwang et al.’s generalization of proxy signature schemes based on elliptic curves. Applied Mathematics and Computation, 170(2):941–947, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000354.

Han:2006:ICL

[HH06a] Fei Han and De-Shuang Huang. Improved constrained learning algorithms by incorporating additional functional constraints into neural networks. Applied Mathematics and Computation, 174 (1):34–50, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305004406.

Hua:2006:EGA

[HH06b] Zhongsheng Hua and Feihua Huang. An effective genetic algorithm approach to large scale mixed integer programming problems. Applied Mathematics and Computation, 174(2):897–909, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005199.

Hsieh:2007:ESG

[HH07a] Sun-Yuan Hsieh and Chao-Wen Huang. An efficient strategy for generating all descendant subtree patterns from phylogenetic trees with its implementation. Applied Mathematics and Computation, 193(2):408–418, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004146.

Hussain:2007:VDP

[HH07b] Malik Zawwar Hussain and Maria Hussain. Visualization of data preserving monotonicity. Applied Mathematics and Computation, 190(2):1353–1364, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700166X.

Herceg:2008:FOF

[HH08a] Dragoslav Herceg and Djordje Herceg. On a fourth-order finite-difference method for singularly perturbed boundary value problems. Applied Mathematics and Computation, 203(2):828–837, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004104.

Huang:2008:CNC

[HH08b] Wen-Zhi Huang and Yan Huang. Chaos of a new class of Hopfield neural networks. Applied Mathematics and Computation, 206(1):1–11, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006395.

He:2009:WAH

[HH09a] Ying He and Bo Han. A wavelet adaptive-homotopy method for inverse problem in the fluid-saturated porous media. Applied Mathematics and Computation, 208(1):189–196, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008874.

Hu:2009:SHB

[HH09b] Haijun Hu and Lihong Huang. Stability and Hopf bifurcation analysis on a ring of four neurons with delays. Applied Mathematics and Computation, 213(2):587–599, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002719.

Hwang:2005:TSP

[HHC05] Min-Shiang Hwang, Kuo-Feng Hwang, and Chin-Chen Chang. A timestamping protocol for digital watermarking. Applied Mathematics and Computation, 169(2):1276–1284, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008501.

Hsieh:2008:DBG

[HHC08] Sun-Yuan Hsieh, Chao-Wen Huang, and Hsin-Hung Chou. A DNA-based graph encoding scheme with its applications to graph isomorphism problems. Applied Mathematics and Computation, 203(2):502–512, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800252X.

Huang:2005:NPN

[HHK^{+}05] De-Shuang Huang, H. S. Ip Horace, C. K. Law Ken, Zheru Chi, and H. S. Wong. A new partitioning neural network model for recursively finding arbitrary roots of higher order arbitrary polynomials. Applied Mathematics and Computation, 162(3):1183–1200, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001791.

Huang:2008:SIL

[HHL08] Sunan Huang, Haibo He, and Kang Li. Special issue on life system modeling and bio-inspired computing for LSMS 2007. Applied Mathematics and Computation, 205(1):3–5, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003123.

Hien:2009:SSS

[HHP09] L. V. Hien, Q. P. Ha, and V. N. Phat. Stability and stabilization of switched linear dynamic systems with time delay and uncertainties. Applied Mathematics and Computation, 210(1):223–231, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009995.

Hu:2009:ASL

[HHW09] Peng Hu, Chengming Huang, and Shulin Wu. Asymptotic stability of linear multistep methods for nonlinear neutral delay differential equations. Applied Mathematics and Computation, 211(1):95– 101, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000423.

Han:2007:FIE

[HHYW07] Fengling Han, Jiankun Hu, Xinghuo Yu, and Yi Wang. Fingerprint images encryption via multi-scroll chaotic attractors. Applied Mathematics and Computation, 185(2):931–939, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008101.

Han:2009:ETL

[HHZ09] Xiaozhuo Han, Cang Hui, and Yanyu Zhang. Effects of time-lagged niche construction on metapopulation dynamics and environmental heterogeneity. Applied Mathematics and Computation, 215(2):449–458, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309004858.

Hasanov:2005:ME

[HI05] Vejdi I. Hasanov and Ivan G. Ivanov. On the matrix equation X − A^{∗}X^{−n}A = I. Applied Mathematics and Computation, 168(2):1340–1356, September 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304007337.

Hernandez:2008:GBB

[HIPA08] Vicente Hernández, Jacinto Javier Ibáñez, Jesús Peinado, and Enrique Arias. A GMRESbased BDF method for solving differential Riccati equations. Applied Mathematics and Computation, 196(2):613–626, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007151.

Hartung:2006:ASC

[HIST06] Ferenc Hartung, Tamás Insperger, Gábor Stépán, and Janos Turi. Approximate stability charts for milling processes using semi-discretization. Applied Mathematics and Computation, 174(1):51– 73, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004418.

Heyouni:2006:MKS

[HJ06] M. Heyouni and K. Jbilou.

Matrix Krylov subspace methods for large scale model reduction problems. Applied Mathematics and Computation, 181(2):1215–1228, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306002414.

Hesaaraki:2008:NMS

[HJ08] M. Hesaaraki and Y. Jalilian. A numerical method for solving n th-order boundary-value problems. Applied Mathematics and Computation, 196(2):889– 897, March 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307007424.

Hartfield:2006:OSR

[HJB06] Roy J. Hartfield, Rhonald M. Jenkins, and John E. Burkhalter. Optimizing a solid rocket motor boosted ramjet powered missile using a genetic algorithm. Applied Mathematics and Computation, 181(2):1720–1736, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003031.

Hyon:2008:NCM

[HJK08] Yun K. Hyon, Ho J. Jang, and Do Y. Kwak.

A nonconforming covolume method for elliptic problems. Applied Mathematics and Computation, 196(1):60–66, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006443.

Hyun:2008:NBD

[HJL08] Seung Ho Hyun, Bogun Jin, and Seung Jae Lee. A novel Bad Data Processing algorithm for analog data in substation automation systems. Applied Mathematics and Computation, 205(2):824–831, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003615.

Haghighat:2005:MMN

[HK05] Mehdi Sohrabi Haghighat and Esmaile Khorram. The maximum and minimum number of efficient units in DEA with interval data. Applied Mathematics and Computation, 163(2):919–930, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400270X.

Hamzacebi:2006:HAF

[HK06] Coşkun Hamzaçebi and Fevzi Kutay. A heuristic approach for finding the global minimum: Adaptive random search technique. Applied Mathematics and Computation, 173(2):1323–1333, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305004340.

Hansen:2009:EMA

[HK09] Scott K. Hansen and Bernard H. Kueper. An efficient method for asymptotic solution to some linear PDEs having arbitrary time-varying type I boundary conditions. Applied Mathematics and Computation, 207(1):273–278, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308007510.

Hamouda:2006:SPG

[HKH06] Luai Hamouda, D. Marc Kilgour, and Keith W. Hipel. Strength of preference in graph models for multiple-decision-maker conflicts. Applied Mathematics and Computation, 179(1):314–327, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010398.

Hariharan:2009:HWM

[HKS09a] G. Hariharan, K. Kannan, and K. R. Sharma. Haar wavelet method for solving Fisher’s equation. Applied Mathematics and Computation, 211(2):284–292, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009880.

Hariharan:2009:HWE

[HKS09b] G. Hariharan, K. Kannan, and Kal Renganathan Sharma. Haar wavelet in estimating depth profile of soil temperature. Applied Mathematics and Computation, 210(1):119– 125, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009405.

Huang:2005:IAS

[HL05a] Ting-Zhu Huang and Cheng Li. An iterative algorithm for searching a scaling matrix for diagonal dominance. Applied Mathematics and Computation, 164(3):821–828, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003480.

Huo:2005:EGS

[HL05b] Hai-Feng Huo and Wan-Tong Li. Existence and global stability of periodic solutions of a discrete ratio-dependent food chain model with delay. Applied Mathematics and Computation, 162(3):1333–1349, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001924.

Hwang:2005:SHW

[HL05c] Shin-Jia Hwang and Hao-Chih Liao. Security of Hsu– Wu’s authenticated encryption scheme with ( t n ) shared verification. Applied Mathematics and Computation, 167(1):281–285, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004916.

Hwang:2005:STH

[HL05d] Shin-Jia Hwang and Hao-Chih Liao. Security of Tzeng–Hwang’s authenticated encryption scheme based on elliptic curve discrete logarithm problems. Applied Mathematics and Computation, 168(1):717–721, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006344.

Hadjidimos:2006:OAD

[HL06a] A. Hadjidimos and M. Lapidakis. Optimal Alternating Direction Implicit Preconditioners for Conjugate Gradient methods. Applied Mathematics and Computation, 183(1):559–574, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005911.

Hou:2006:TSC

[HL06b] Yanren Hou and Kaitai Li.

Tangent space correction method for the Galerkin approximation based on two-grid finite element. Applied Mathematics and Computation, 175(1):413– 429, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500617X.

Hu:2006:PEP

[HL06c] Xiaozhe Hu and Xiaofeng Ling. Preconditioners for elliptic problems via nonuniform meshes. Applied Mathematics and Computation, 181(2):1182–1198, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306002037.

He:2007:PIC

[HL07a] Anqi He and Jianzhou Liu. A parallel iterative criterion for H-matrices. Applied Mathematics and Computation, 190(1):1–5, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017346.

Hsu:2007:CQT

[HL07b] William Wei-Yuan Hsu and Yuh-Dauh Lyuu. A convergent quadratic-time lattice algorithm for pricing European-style Asian options. Applied Mathematics and Computation, 189(2):1099–1123, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016869.

Hu:2007:GSR

[HL07c] Lin-Xia Hu and Wan-Tong Li. Global stability of a rational difference equation. Applied Mathematics and Computation, 190(2):1322–1327, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001634.

Huo:2007:EPP

[HL07d] Hai-Feng Huo and Wan-Tong Li. Existence of positive periodic solution of a neutral impulsive delay predator–prey system. Applied Mathematics and Computation, 185(1):499–507, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600885X.

Han:2008:NAF

[HL08a] Fei Han and Qing-Hua Ling. A new approach for function approximation incorporating adaptive particle swarm optimization and a priori information. Applied Mathematics and Computation, 205(2):792–798, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003573.

He:2008:UHF

[HL08b] Shuping He and Fei Liu. Unbiased H_{∞} filtering for neutral Markov jump systems. Applied Mathematics and Computation, 206(1):175–185, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006498.

Li:2008:SFS

[hL08c] Rui hong Li. A special full-state hybrid projective synchronization in symmetrical chaotic systems. Applied Mathematics and Computation, 200(1):321–329, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011204.

Huang:2008:PCS

[HL08d] Ya-Yuan Huang and Ming-Sheng Liu. Properties of certain subclasses of multivalent analytic functions involving the Dziok–Srivastava operator. Applied Mathematics and Computation, 204(1):137–149, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004852.

Han:2008:RTO

[HLCH08] Dongil Han, Byoungmoo Lee, Jae Il Cho, and Dae-Hwan Hwang. Real-time object segmentation using disparity map of stereo matching. Applied Mathematics and Computation, 205(2):770–777, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800355X.

Huang:2006:PNP

[HLK06] Hsin-I Huang, Chuen-Horng Lin, and Jau-Chuan Ke. Parametric nonlinear programming approach for a repairable system with switching failure and fuzzy parameters. Applied Mathematics and Computation, 183(1):508–517, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005856.

Hlomuka:2005:EUS

[Hlo05] Joe Hlomuka. On the existence, uniqueness and the stability of a solution to a cooling problem, for an isotropic 3-D solid. Applied Mathematics and Computation, 163(2):693–703, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400253X.

Liu:2009:WPA

[hLqSlZ09] Jing huai Liu, Xiao qiu Song, and Ping li Zhang. Weighted pseudo almost periodic mild solutions of semilinear evolution equations with nonlocal conditions. Applied Mathematics and Computation, 215(5):1647–1652, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007292.

Hwang:2005:ESS

[HLS05] Ren-Junn Hwang, Chih-Hua Lai, and Feng-Fu Su. An efficient signcryption scheme with forward secrecy based on elliptic curve. Applied Mathematics and Computation, 167(2):870–881, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005351.

Han:2010:RPA

[HLSS10] Zhenlai Han, Tongxing Li, Shurong Sun, and Yibing Sun. Remarks on the paper [Appl. Math. Comput. 207 (2009) 388–396]. Applied Mathematics and Computation, 215(11):3998–4007, February 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [YX09].

Li:2007:DTG

[hLsT07] Ji hong Li and Nai shuo Tian. The discrete-time GI / Geo / 1 queue with working vacations and vacation interruption. Applied Mathematics and Computation, 185(1):1–10, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007958.

Huo:2009:DHO

[HLT09] Hai-Feng Huo, Wan-Tong Li, and Sanyi Tang. Dynamics of high-order BAM neural networks with and without impulses. Applied Mathematics and Computation, 215(6):2120–2133, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309007206.

Hu:2008:PSN

[HLW08] Lili Hu, Lishan Liu, and Yonghong Wu. Positive solutions of nonlinear singular two-point boundary value problems for secondorder impulsive differential equations. Applied Mathematics and Computation, 196(2):550–562, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307007072.

Hung:2009:AVS

[HLY09] Yung-Ching Hung, Teh-Lu Liao, and Jun-Juh Yan. Adaptive variable structure control for chaos suppression of unified chaotic systems. Applied Mathematics and Computation, 209(2):391–398, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308009752.

Huang:2006:MPS

[HM06] Tony Huang and Ananda Sanagavarapu Mohan. Microparticle swarm optimizer for solving high dimensional optimization problems (μPSO for high dimensional optimization problems). Applied Mathematics and Computation, 181(2):1148–1154, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001998.

Helal:2007:TMA

[HM07] M. A. Helal and M. S.

Mehanna. The tanh method and Adomian decomposition method for solving the foam drainage equation. Applied Mathematics and Computation, 190(1):599–609, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030700094X.

Harley:2008:IIB

[HM08a] C. Harley and E. Momoniat. Instability of invariant boundary conditions of a generalized Lane–Emden equation of the second-kind. Applied Mathematics and Computation, 198(2):621–633, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009290.

Huang:2008:OSO

[HM08b] Yan Huang and Fanwei Meng. Oscillation of second-order nonlinear ODE with damping. Applied Mathematics and Computation, 199(2):644– 652, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010521.

See comments [SQ11].

Hari:2009:ATS

[HM09a] V. Hari and J. Matejaš. Accuracy of two SVD algorithms for 2×2 triangular matrices. Applied Mathematics and Computation, 210(1):232–257, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030801000X.

Heys:2009:LE

[HM09b] Jeffrey J. Heys and Bradley M. Merchant. Letter to the Editor. Applied Mathematics and Computation, 211(1):253, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300308009600.

Ma:2009:EPP

[hMfWzLpZ09] Zhi hui Ma, Shu fan Wang, Zi zhen Li, and Feng pan Zhang. Erratum to “Permanence of a predator-prey system with stage structure and time delay” [Appl. Math. Comput. 201 (2008) 65–71]. Applied Mathematics and Computation, 215(6):2397, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008029.

See [hMzLfW^{+}08].

Hashemiparast:2006:SBC

[HMJD06] S. M. Hashemiparast, M. Masjed-Jamei, and Mehdi Dehghan. On selection of the best coefficients in interpolatory quadrature rules. Applied Mathematics and Computation, 182(2):1240–1246, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004371.

Hashemiparast:2006:SKC

[HMJED06] S. M. Hashemiparast, M. Masjed-Jamei, M. R. Eslahchi, and Mehdi Dehghan. The second kind Chebyshev–Newton–Cotes quadrature rule (open type) and its numerical improvement. Applied Mathematics and Computation, 180(2):605–613, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600049X.

Hernandez:2009:LTA

[HMNT09] Isabel Hernández, Consuelo Mateos, Juan Núñez, and Ángel F. Tenorio. Lie theory: Applications to problems in mathematical finance and economics. Applied Mathematics and Computation, 208(2):446–452, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009156.

Horgan:2007:AMB

[HMSZ07] Graham Horgan, Ilya Molchanov, Vadim Shcherbakov, and Sergei Zuyev. Approximate models for bacteria leaching through porous media. Applied Mathematics and Computation, 188(2):1941–1954, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015621.

Hueso:2008:NST

[HMT08] José L. Hueso, Eulalia Martínez, and Juan R. Torregrosa. A note on some three-step iterative methods for nonlinear equations. Applied Mathematics and Computation, 202(1):252–255, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000933.

Hueso:2009:TFO

[HMT09a] José L. Hueso, Eulalia Martínez, and Juan R. Torregrosa. Third and fourth order iterative methods free from second derivative for nonlinear systems. Applied Mathematics and Computation, 211(1):190– 197, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000708.

Hueso:2009:TOI

[HMT09b] José L. Hueso, Eulalia Martínez, and Juan R. Torregrosa. Third order iterative methods free from second derivative for nonlinear systems. Applied Mathematics and Computation, 215(1):58– 65, September 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030900366X.

Huang:2007:RCP

[HMW07] Zheng-Hai Huang, Xin-He Miao, and Ping Wang. A revised cut-peak function method for box constrained continuous global optimization. Applied Mathematics and Computation, 194(1):224–233, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004614.

Ma:2008:PPP

[hMzLfW^{+}08] Zhi hui Ma, Zi zhen Li, Shu fan Wang, Ting Li, and Feng pan Zhang. Permanence of a predator– prey system with stage structure and time delay. Applied Mathematics and Computation, 201(1–2):65– 71, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011629.

See erratum [hMfWzLpZ09].

Ho:2005:IMR

[HN05] Andy C. Ho and Michael K. Ng. Iterative methods for Robbins problems. Applied Mathematics and Computation, 165(1):103–125, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003650.

Hasheminia:2006:GAA

[HN06a] Hamed Hasheminia and Seyed Taghi Akhavan Niaki. A genetic algorithm approach to find the best regression/econometric model among the candidates. Applied Mathematics and Computation, 183(1):337– 349, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306005662.

Hosseini:2006:CAD

[HN06b] M. M. Hosseini and H. Nasabzadeh.

On the convergence of Adomian decomposition method. Applied Mathematics and Computation, 182(1):536–543, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003614.

Hosseini:2007:MAD

[HN07a] M. M. Hosseini and H. Nasabzadeh.

Modified Adomian decomposition method for specific second order ordinary differential equations. Applied Mathematics and Computation, 186(1):117–123, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009222.

Huang:2007:EPM

[HN07b] Zhenyu Huang and Muhammad Aslam Noor. An explicit projection method for a system of nonlinear variational inequalities with different (γ, r)- cocoercive mappings. Applied Mathematics and Computation, 190(1):356– 361, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000562.

Huang:2007:SNU

[HN07c] Zhenyu Huang and Muhammad Aslam Noor. Some new unified iteration schemes with errors for nonexpansive mappings and variational inequalities. Applied Mathematics and Computation, 194(1):135–142, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004523.

Hashim:2006:NAD

[HNB06] I. Hashim, M. S. M.

Noorani, and B. Batiha. A note on the Adomian decomposition method for the generalized Huxley equation. Applied Mathematics and Computation, 181(2):1439–1445, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002633.

Hollands:2007:SFS

[Hol07] K. G. Terry Hollands.

A simplification to Fredholm’s solution to the Fredholm integral equation of the second kind. Applied Mathematics and Computation, 189(2):1078–1086, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016845.

Hong:2008:RFT

[Hon08a] Wei-Chiang Hong. Rainfall forecasting by technological machine learning models. Applied Mathematics and Computation, 200(1):41–57, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010843.

Hongyan:2008:LPS

[Hon08b] Zhi Hongyan. Lie point symmetry and some new soliton-like solutions of the Konopelchenko–Dubrovsky equations. Applied Mathematics and Computation, 203(2):931–936, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004074.

Hong:2009:NJE

[Hon09] Baojian Hong. New Jacobi elliptic functions solutions for the variablecoefficient mKdV equation. Applied Mathematics and Computation, 215(8):2908–2913, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309008601.

Horasanli:2007:CLB

[Hor07] Mehmet Horasanlı. A comparison of lattice based option pricing models on the rate of convergence. Applied Mathematics and Computation, 184(2):649–658, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007399.

Horasanli:2008:HSP

[Hor08] Mehmet Horasanlı. Hedging strategy for a portfolio of options and stocks with linear programming. Applied Mathematics and Computation, 199(2):804– 810, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010806.

Hosseini:2005:IRM

[Hos05a]  M. M. Hosseini.  An

index reduction method for linear Hessenberg systems. Applied Mathematics and Computation, 171(1):596–603, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001190.

Hosseini:2005:NSL

[Hos05b]  M. M. Hosseini.  Nu-

merical solution of linear differential-algebraic equations. Applied Mathematics and Computation, 162 (1):7–14, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304000438.

Hosseini:2005:NSO

[Hos05c] M. M. Hosseini. Numerical solution of ordinary differential equations with impulse solution. Applied Mathematics and Computation, 163(1):373–381, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002310.

Hosseini:2005:PMN

[Hos05d] M. M. Hosseini. Pseudospectral method for numerical solution of DAEs with an error estimation. Applied Mathematics and Computation, 170(1):115–124, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008756.

Hosseini:2006:ADMb

[Hos06a] M. M. Hosseini. Adomian decomposition method for solution of nonlinear differential algebraic equations. Applied Mathematics and Computation, 181(2):1737–1744, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003043.

Hosseini:2006:ADMa

[Hos06b] M. M. Hosseini. Adomian decomposition method with Chebyshev polynomials. Applied Mathematics and Computation, 175(2):1685– 1693, April 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305007241.

Hosseini:2006:MPM

[Hos06c] M. M. Hosseini. A modified pseudospectral method for numerical solution of ordinary differential equations systems. Applied Mathematics and Computation, 176(2):470–475, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008246.

Hosny:2007:EFC

[Hos07] Khalid M. Hosny. Exact and fast computation of geometric moments for gray level images. Applied Mathematics and Computation, 189(2):1214–1222, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017061.

Hosny:2008:CAA

[Hos08] Khalid M. Hosny. On the computational aspects of affine moment invariants for gray-scale images. Applied Mathematics and Computation, 195(2):762–771, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006157.

Heinonen:2007:HAC

[HP07] J. Heinonen and F. Pettersson. Hybrid ant colony optimization and visibility studies applied to a job-shop scheduling problem. Applied Mathematics and Computation, 187(2):989–998, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012355.

Hassanzadeh:2007:CDN

[HPD07] Hassan Hassanzadeh and Mehran Pooladi-Darvish. Comparison of different numerical Laplace inversion methods for engineering applications. Applied Mathematics and Computation, 189(2):1966–1981, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000161.

Peng:2006:FTR

[hPY06] Ye hui Peng and Shengbao Yao. A feasible trustregion algorithm for inequality constrained optimization. Applied Mathematics and Computation, 173(1):513–522, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003711.

Peng:2006:EAL

[hPyHZ06] Zhuo hua Peng, Xi yan Hu, and Lei Zhang. An efficient algorithm for the least-squares reflexive solution of the matrix equation A _{1} XB _{1} = C _{1}, A 2 XB _{2} = C _{2}. Applied Mathematics and Computation, 181(2):988–999, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001809.

Hernandez:2007:AIP

[HR07a] M. A. Hernández and N. Romero. Application of iterative processes of Rorder at least three to operators with unbounded second derivative. Applied Mathematics and Computation, 185(1):737–747, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600909X.

Hernandez:2007:MPO

[HR07b] M. A. Hernández and N. Romero. Methods with prefixed order for approximating square roots with global and general convergence. Applied Mathematics and Computation, 194(2):346–353, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004754.

Hernandez:2009:TUT

[HR09a] M. A. Hernández and N. Romero. Toward a unified theory for third R-order iterative methods for operators with unbounded second derivative. Applied Mathematics and Computation, 215(6):2248–2261, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007383.

Hounkonnou:2009:ADH

[HR09b] Mahouton Norbert Hounkonnou and André Ronveaux. About derivatives of Heun’s functions from polynomial transformations of hypergeometric equations. Applied Mathematics and Computation, 209(2):421–424, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800965X.

He:2008:ETW

[HRCL08] Bin He, Weiguo Rui, Can Chen, and Shaolin Li. Exact travelling wave solutions of a generalized Camassa–Holm equation using the integral bifurcation method. Applied Mathematics and Computation, 206(1):141–149, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006401.

Hernandez:2005:SSC

[HRE05] M. A. Hernández, M. J.

Rubio, and J. A. Ezquerro. Solving a special case of conservative problems by Secant-like methods. Applied Mathematics and Computation, 169(2):926–942, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008239.

He:2008:BTWb

[HRLC08] Bin He, Weiguo Rui, Shaolin Li, and Chan Chen. Bounded travelling wave solutions for a modified form of generalized Degasperis–Procesi equation. Applied Mathematics and Computation, 206(1):113–123, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800636X.

Hashemi:2007:TNA

[HRM07] S. Mehdi Hashemi, Mohsen Rezapour, and Ahmad Moradi. Two new algorithms for the Min-Power Broadcast problem in static ad hoc networks. Applied Mathematics and Computation, 190(2):1657–1668, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700210X.

Hounkonnou:2007:FSC

[HRS07] Mahouton Norbert Hounkonnou, André Ronveaux, and Komi Sodoga. Factorization of some confluent Heun’s differential equations. Applied Mathematics and Computation, 189(1):816–820, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016596.

Herrero:2007:ACN

[HRT07] Alicia Herrero, Antonio Ramírez, and Nestor Thome. An algorithm to check the nonnegativity of singular systems. Applied Mathematics and Computation, 189(1):355– 365, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016109.

Hsieh:2007:CEA

[HS07a] Sun-Yuan Hsieh and Jung-Yiau Shiu. Cycle embedding of augmented cubes. Applied Mathematics and Computation, 191(2):314–319, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002597.

Huang:2007:SBA

[HS07b] Yanli Huang and Yongzhong Song. Semiconvergence of block AOR method for singular p-cyclic matrices. Applied Mathematics and Computation, 189(2):1637–1647, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017486.

Hamdan:2009:FDG

[HS09a] M. H. Hamdan and H. I. Siyyam. On the flow of a dusty gas with constant number density through granular porous media. Applied Mathematics and Computation, 209(2):339–345, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308009697.

Hussein:2009:SST

[HS09b] A. Hussein and M. M. Selim. Solution of the stochastic transport equation of neutral particles with anisotropic scattering using RVT technique. Applied Mathematics and Computation, 213(1):250– 261, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002264.

Hymavathi:2009:QAH

[HS09c] T. Hymavathi and B. Shanker.

A quasilinearization approach to heat transfer in MHD visco-elastic fluid flow. Applied Mathematics and Computation, 215(6):2045–2054, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007127.

Hassanien:2005:FOF

[HSH05] I. A. Hassanien, A. A.

Salama, and H. A. Hosham. Fourth-order finite difference method for solving Burgers’ equation. Applied Mathematics and Computation, 170(2):781–800, November 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030500024X.

Hsu:2005:CIT

[Hsu05a] Chien-Lung Hsu. Cryptanalysis and improvement of the Tzeng–Hwang authenticated encryption scheme based on elliptic curve discrete logarithm problem. Applied Mathematics and Computation, 167(2):882–890, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005363.

Hsu:2005:UFR

[Hsu05b] Chien-Lung Hsu. A user friendly remote authentication scheme with smart cards against impersonation attacks. Applied Mathematics and Computation, 170(1):135–143, November 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030400877X.

Hsu:2006:EBC

[Hsu06] Pao-Tung Hsu. Estimating the boundary condition in a 3D inverse hyperbolic heat conduction problem. Applied Mathematics and Computation, 177(2):453–464, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009240.

He:2005:RKF

[HT05] Ji-Huan He and Jie Tang.

Rebuild of King Fang 40 BC musical scales by He’s inequality. Applied Mathematics and Computation, 168(2):909–914, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006484.

Hashemi:2006:BHA

[HT06a] S. Mehdi Hashemi and Maryam Tahmasbi. A better heuristic for areacompaction of orthogonal representations. Applied Mathematics and Computation, 172(2):1054–1066, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002778.

Hu:2006:BML

[HT06b] Yi-Chung Hu and Jung-Fa Tsai. Backpropagation multi-layer perceptron for incomplete pairwise comparison matrices in analytic hierarchy process. Applied Mathematics and Computation, 180(1): 53–62, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305010970.

Hasanov:2007:DFD

[HT07] Anvar Hasanov and Mamasali Turaev. Decomposition formulas for the double hypergeometric functions G _{1} and G _{2}. Applied Mathematics and Computation, 187(1):195–201, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011519.

Huang:2009:ABL

[HT09a] Shuibo Huang and Qiaoyu Tian. Asymptotic behavior of large solution for boundary blowup problems with non-linear gradient terms. Applied Mathematics and Computation, 215(8):3091–3097, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309008959.

Huang:2009:ABLb

[HT09b] Shuibo Huang and Qiaoyu Tian. Asymptotic behavior of large solution for boundary blowup problems with non-linear gradient terms. Applied Mathematics and Computation, 215(8):3091–3097, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008959.

Huang:2009:ABLa

[HTM09] Shuibo Huang, Qiaoyu Tian, and Chunlai Mu. Asymptotic behavior of large solution to elliptic equation of Bieberbach– Rademacher type with convection terms. Applied Mathematics and Computation, 210(2):284–293, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009764.

Huang:2006:NAD

[HTO06a] Jih-Jeng Huang, Gwo-Hshiung Tzeng, and Chorng-Shyong Ong. A novel algorithm for dynamic factor analysis. Applied Mathematics and Computation, 175(2):1288–1297, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006788.

Huang:2006:NAU

[HTO06b] Jih-Jeng Huang, Gwo-Hshiung Tzeng, and Chorng-Shyong Ong. A novel algorithm for uncertain portfolio selection. Applied Mathematics and Computation, 173(1):350–359, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003553.

Huang:2006:TSG

[HTO06c] Jih-Jeng Huang, Gwo-Hshiung Tzeng, and Chorng-Shyong Ong. Twostage genetic programming (2SGP) for the credit scoring model. Applied Mathematics and Computation, 174(2):1039–1053, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500531X.

Hu:2005:MCT

[Hu05] Hai-Ping Hu. Mixed convection turbulent film condensation on a sphere. Applied Mathematics and Computation, 170(2):1194– 1208, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305000585.

Hu:2007:GRA

[Hu07a] Yi-Chung Hu. Grey relational analysis and radial basis function network for determining costs in learning sequences. Applied Mathematics and Computation, 184(2):291–299, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007041.

Hu:2007:SFI

[Hu07b]  Yi-Chung Hu.  Sugeno

fuzzy integral for finding fuzzy if–then classification rules. Applied Mathematics and Computation, 185(1):72–83, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008046.

Huang:2005:SAD

[Hua05a] Shu-Shiang Huang. Stochastic adsorption–desorption problem for mass-spring systems in the nanotechnology. Applied Mathematics and Computation, 163(3):1043–1054, April 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002772.

Huang:2005:GLE

[Hua05b] Zheng-Hai Huang. Global Lipschitzian error bounds for semidefinite complementarity problems with emphasis on NCPs. Applied Mathematics and Computation, 162(3):1237–1258, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001845.

Huang:2006:FCC

[Hua06] Xiaoxia Huang. Fuzzy chance-constrained portfolio selection. Applied Mathematics and Computation, 177(2):500–507, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030500929X.

Huang:2007:SII

[Hua07] De-Shuang Huang. Special issue on intelligent computing theory and methodology. Applied Mathematics and Computation, 185(2):757–758, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007533.

Huang:2008:PNP

[Hua08a]  Aimei Huang.  Per-

manence of a nonlinear prey-competition model with delays. Applied Mathematics and Computation, 197(1):372–381, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007850.

Huang:2008:ASM

[Hua08b] Chengming Huang. Asymptotic stability of multistep methods for nonlinear delay differential equations. Applied Mathematics and Computation, 203(2):908–912, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002105.

Huang:2008:SIA

[Hua08c] De-Shuang Huang. Special issue on advanced intelligent computing theory and methodology in applied mathematics and computation. Applied Mathematics and Computation, 205(2):509–510, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006863.

Huang:2009:ELA

[Hua09] Chia-Hui Huang. An effective linear approximation method for separable programming problems. Applied Mathematics and Computation, 215(4):1496–1506, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006493.

Huber:2005:SSS

[Hub05]  Alfred Huber.  Soli-

tary solutions of some nonlinear evolution equations. Applied Mathematics and Computation, 166(2):464–474, July 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004515.

Huber:2008:CNC

[Hub08a] Alfred Huber. The calculation of novel class of solutions of a non-linear fourth order evolution equation by the Weierstraß transform method. Applied Mathematics and Computation, 201(1–2):668–677, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000131.

Huber:2008:NNS

[Hub08b] Alfred Huber. A note on new solitary and similarity class of solutions of a fourth-order non-linear evolution equation. Applied Mathematics and Computation, 202(2):787–797, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001823.

Huber:2009:CTE

[Hub09a]  Alfred Huber.  The

Cavalcante–Tenenblat equation — Does the equation admits physical significance? Applied Mathematics and Computation, 212(1):14–22, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001131.

Huber:2009:GET

[Hub09b] Alfred Huber. A generalized exponential transform method for solving non-linear evolution equations of physical relevance. Applied Mathematics and Computation, 215(1):344–352, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309004767.

Huber:2009:NLE

[Hub09c] Alfred Huber. On nonlinear evolution equations of higher order — the introduction and application of a novel computational approach. Applied Mathematics and Computation, 215(6):2337–2348, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309007589.

Huber:2009:SGE

[Hub09d] Alfred Huber. The Sine– Gordon-Equation with perturbation terms of power Jacobian elliptic functions. Applied Mathematics and Computation, 215(8):2850–2854, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309008248.

Huseynov:2009:SGF

[Hus09] Aydin Huseynov. On the sign of Green’s function for an impulsive differential equation with periodic boundary conditions. Applied Mathematics and Computation, 208(1):197–205, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008898.

Hadi-Vencheh:2008:CIO

[HV08] Abdollah Hadi-Vencheh.

Comment on “Inputs/ outputs estimation in DEA when some factors are undesirable” [Applied Mathematics and Computation 156 (2004) 19–32]. Applied Mathematics and Computation, 202(2):893–894, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009800.

See [JVFM04].

Hsu:2005:ENT

[HW05a] Chien-Lung Hsu and Tzong-Sun Wu. Efficient nonrepudiable threshold proxy signature scheme with known signers against the collusion attack. Applied Mathematics and Computation, 168(1):305–319, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304006058.

Hsu:2005:SCT

[HW05b] Chien-Lung Hsu and Tzong-Sun Wu. Self-certified threshold proxy signature schemes with message recovery, nonrepudiation, and traceability. Applied Mathematics and Computation, 164(1):201–225, May 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003066.

Hung:2005:NDN

[HW05c] Wen-Liang Hung and Jong-Wuu Wu. A note on determining the number of outliers in a normal sample with unequal variances by least squares procedure. Applied Mathematics and Computation, 162(3):1007–1012, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001420.

Huang:2006:STA

[HW06] Zheng-Hai Huang and Hui Wang. Smoothing-type algorithm for solving linear programs by using an augmented complementarity problem. Applied Mathematics and Computation, 177(1):330–345, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009148.

He:2007:HPS

[HW07a] Qie He and Ling Wang. A hybrid particle swarm optimization with a feasibilitybased rule for constrained optimization. Applied Mathematics and Computation, 186(2):1407–1422, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010678.

Hsiao:2007:NST

[HW07b] C. H. Hsiao and S. P.

Wu. Numerical solution of time-varying functional differential equations via Haar wavelets. Applied Mathematics and Computation, 188(1):1049–1058, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014548.

Han:2008:FCI

[HW08a] Danfu Han and Peng Wu. A family of combined iterative methods for solving nonlinear equation. Applied Mathematics and Computation, 195(2):448–453, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005668.

Han:2008:MPP

[HW08b] Fei Han and Quanyi Wang. Multiple positive periodic solutions for a class of nonlinear functional differential equations. Applied Mathematics and Computation, 205(1):383–390, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005961.

Hou:2009:GIM

[HW09] Guoliang Hou and Li Wang. A generalized iterative method and comparison results using projection techniques for solving linear systems. Applied Mathematics and Computation, 215(2):806–817, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005682.

Hwang:2005:ITA

[Hwa05]  Shin-Jia Hwang.  Im-

provement of Tseng et al.’s authenticated encryption scheme. Applied Mathematics and Computation, 165(1):1–4, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003170.

Hong:2007:CPP

[HWC07] Ching-Wen Hong, Jong-Wuu Wu, and Ching-Hsue Cheng. Computational procedure of performance assessment of lifetime index of businesses for the Pareto lifetime model with the right type II censored sample. Applied Mathematics and Computation, 184(2):336–350, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007119.

Huang:2007:IJM

[HWF07] Ting-Zhu Huang, Xue-Zhong Wang, and Ying-Ding Fu. Improving Jacobi methods for nonnegative H-matrices linear systems. Applied Mathematics and Computation, 186(2):1542–1550, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010812.

Hsu:2005:NPM

[HWH05] Chien-Lung Hsu, Tzong-Sun Wu, and Wei-Hua He. New proxy multisignature scheme. Applied Mathematics and Computation, 162(3):1201–1206, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001808.

Han:2009:CAP

[HWH09] Congying Han, Yongli Wang, and Guoping He. On the convergence of asynchronous parallel algorithm for large-scale linearly constrained minimization problem. Applied Mathematics and Computation, 211(2):434–441, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000915.

Huang:2007:PSD

[HWS07] Hai Huang, Zengxin Wei, and Yao Shengwei. The proof of the sufficient descent condition of the Wei– Yao–Liu conjugate gradient method under the strong Wolfe–Powell line search. Applied Mathematics and Computation, 189(2):1241–1245, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017085.

Han:2007:NSG

[HX07] Houde Han and Zhenli Xu. Numerical solitons of generalized Korteweg–de Vries equations. Applied Mathematics and Computation, 186(1):483–489, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600957X.

Hu:2008:CNF

[HXH08] Bing Hu, Youcai Xu, and Jinsong Hu. Crank– Nicolson finite difference scheme for the Rosenau– Burgers equation. Applied Mathematics and Computation, 204(1):311–316, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005080.

He:2005:EPS

[HY05] Zhimin He and Jianshe Yu. On the existence of positive solutions of fourth-order difference equations. Applied Mathematics and Computation, 161(1):139–148, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012165.

Hu:2007:VCD

[HY07a] Xinhua Hu and Zhongxing Ye. Valuing credit derivatives in a jumpdiffusion model. Applied Mathematics and Computation, 190(1):627–632, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000963.

Huang:2007:CIE

[HY07b] Guang-Xin Huang and Feng Yin. Constrained inverse eigenproblem and associated approximation problem for anti-Hermitian R-symmetric matrices. Applied Mathematics and Computation, 186(1):426– 434, March 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306009519.

Huang:2007:MIP

[HY07c] Guang-Xin Huang and Feng Yin. Matrix inverse problem and its optimal approximation problem for R-symmetric matrices. Applied Mathematics and Computation, 189(1):482– 489, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016213.

Yang:2005:BAD

[hYG05] Zhong hua Yang and Qian Guo. Bifurcation analysis of delayed logistic equation. Applied Mathematics and Computation, 167(1):454–476, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005041.

Huang:2007:GSM

[HYG07] Guang-Xin Huang, Feng Yin, and Ke Guo. The general solutions on the minimum residual problem and the matrix nearness problem for symmetric matrices or anti-symmetric matrices. Applied Mathematics and Computation, 194(1):85–91, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004468.

Han:2008:MSG

[HYG08] Le Han, Gaohang Yu, and Lutai Guan. Multivariate spectral gradient method for unconstrained optimization. Applied Mathematics and Computation, 201(1–2):621–630, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308000088.

Hsieh:2009:NUE

[HYL09] Y.-C. Hsieh, P.-S. You, and C.-D. Liou. A note of using effective immune based approach for the flow shop scheduling with buffers. Applied Mathematics and Computation, 215(5):1984–1989, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006791.

Hong:2007:IMS

[HYZ07] Yiping Hong, Jianqiang Yi, and Dongbin Zhao. Improved mean shift segmentation approach for natural images. Applied Mathematics and Computation, 185(2):940–952, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008113.

Huang:2005:DCR

[HZ05] De-Shuang Huang and Wen-Bo Zhao. Determining the centers of radial basis probabilistic neural networks by recursive orthogonal least square algorithms. Applied Mathematics and Computation, 162(1):461–473, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001092.

Han:2006:CEE

[HZ06a] Danfu Han and Jingfen Zhu. Convergence and error estimate of a deformed Newton method. Applied Mathematics and Computation, 173(2):1115–1123, February 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305004169.

Hua:2006:HSV

[HZ06b] Zhongsheng Hua and Bin Zhang. A hybrid support vector machines and logistic regression approach for forecasting intermittent demand of spare parts. Applied Mathematics and Computation, 181(2):1035–1048, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001871.

Huang:2006:INI

[HZ06c] Feng Huang and Xian Zhang. An improved Newton iteration for the weighted Moore–Penrose inverse. Applied Mathematics and Computation, 174(2):1460–1486, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005643.

Huang:2006:IIA

[HZ06d] Ting-Zhu Huang and Xiao-Qing Zhang. An improved iterative algorithm for generalized diagonally dominant matrices. Applied Mathematics and Computation, 181(1):742–746, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001548.

He:2008:NSW

[HZ08] Xiaoming He and Wenming Zou. Nontrivial solitary waves to the generalized Kadomtsev– Petviashvili equations. Applied Mathematics and Computation, 197(2):858– 863, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008557.

Hilscher:2009:PTI

[HZ09a] Roman Šimon Hilscher and Vera Zeidan. Picone type identities and definiteness of quadratic functionals on time scales. Applied Mathematics and Computation, 215(7):2425–2437, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007681.

Huang:2009:ISW

[HZ09b] Wenyi Huang and Jian Zhang. Instability of the standing waves for nonlinear Klein–Gordon equations with damping term. Applied Mathematics and Computation, 213(2):522–528, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002653.

Hu:2007:BPM

[HZS07] Xiaozhe Hu, Lingxiao Zhao, and Abdul Wasim Shaikh. The boundary penalty method for the diffusion equation subject to the specification of mass. Applied Mathematics and Computation, 186(1):735–748, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009945.

Hou:2006:MAO

[HZW06] Zai-En Hou, Ke-Cun Zhang, and Yan-Jun Wang. The model and algorithm for the optimization of the granularity distribution. Applied Mathematics and Computation, 179(2):763– 771, August 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305010866.

He:2007:LSF

[HZX07] Guoliang He, Yong Zhang, and Huang Xiaoqin. The locally stabilized finite volume method for incompressible flow. Applied Mathematics and Computation, 187(2):1399–1409, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601277X.

Huang:2008:AUB

[HZYL08] Weitong Huang, Yu Zhao, Shiqiang Yang, and Yuchang Lu. Analysis of the user behavior and opinion classification based on the BBS. Applied Mathematics and Computation, 205(2):668–676, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003457.

Ismail:2008:PLM

[IAMA08] Zuhaila Ismail, Ilyani Abdullah, Norzieha Mustapha, and Norsarahaida Amin. A power-law model of blood flow through a tapered overlapping stenosed artery. Applied Mathematics and Computation, 195(2):669–680, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005966.

Izgi:2009:RCC

[İBİ09] Aydin İzgi, İbrahim Büyükyazici, and Ertan İbikli. Rate of convergence by Chlodowsky– Taylor polynomials. Applied Mathematics and Computation, 213(2):426– 431, July 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300309002537.

Ibrahim:2007:SMT

[Ibr07] Ibrahim Abouelfarag Ibrahim. Solvability of mixed type operator equations. Applied Mathematics and Computation, 190(2):1344– 1352, July 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307001658.

Ibrir:2008:SRS

[Ibr08]  Salim Ibrir.  Stability

and robust stabilization of discrete-time switched systems with time-delays: LMI approach. Applied Mathematics and Computation, 206(2):570–578, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003809.

Ikbal:2008:UPT

[ICM08] Md. Asif Ikbal, Santabrata Chakravarty, and Prashanta Kumar Mandal. An unsteady peristaltic transport phenomenon of non-Newtonian fluid — A generalised approach. Applied Mathematics and Computation, 201(1–2):16–34, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011538.

Ibrir:2008:NLC

[ID08] Salim Ibrir and Sette Diopt. Novel LMI conditions for observer-based stabilization of Lipschitzian nonlinear systems and uncertain linear systems in discrete-time. Applied Mathematics and Computation, 206(2):579–588, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003810.

Ide:2006:SEP

[Ide06] Takanori Ide. Some energy preserving finite element schemes based on the discrete variational derivative method. Applied Mathematics and Computation, 175(1):277–296, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006077.

Ide:2008:NHI

[Ide08a] Nasr-Al-Din Ide. A new hybrid iteration method for solving algebraic equations. Applied Mathematics and Computation, 195(2):772–774, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006169.

Ide:2008:MNM

[Ide08b] Nasr-Al-Din Ide. On modified Newton methods for solving a non linear algebraic equations. Applied Mathematics and Computation, 198(1):138–142, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008740.

Irk:2006:STS

[IDS06] Dursun Irk, İdris Dağ, and Bülent Saka. A small time solutions for the Korteweg–de Vries equation using spline approximation. Applied Mathematics and Computation, 173(2):834–846, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003966.

Iranpoor:2007:ETP

[IGM07] M. Iranpoor, S. M. T. Fatemi Ghomi, and A. Mohamadinia. Earliness tardiness production planning and scheduling in flexible flowshop systems under finite planning horizon. Applied Mathematics and Computation, 184(2):950–964, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007855.

Islam:2009:NIA

[IH09] Md. Shafiqul Islam and M. Alamgir Hossain. Numerical integrations over an arbitrary quadrilateral region. Applied Mathematics and Computation, 210(2):515–524, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000319.

Ibrahem:2006:TES

[IHB06] F. S. Ibrahem, I. A. Hassanien, and A. A. Bakr. Thermal effects in Stokes’ second problem for unsteady micropolar fluids flow. Applied Mathematics and Computation, 173(2):916–937, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004029.

Ishikawa:2007:MCI

[II07] Keitarou Ishikawa and Takehiro Inohara. A method to compare influence of coalitions on group decision other than desirability relation. Applied Mathematics and Computation, 188(1):838–849, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014238.

Ismail:2005:CMS

[IIS05] I. A. Ismail, G. Ismail, and A. M. Sedqi. A computational method for solving the branched channel flow problem. Applied Mathematics and Computation, 161(1):19–34, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012050.

Insua:2008:NGB

[IL08] Manuel A. Insua and Manuel Ladra. A note on Gröbner bases and Berlekamp’s algorithm. Applied Mathematics and Computation, 196(1):77– 85, February 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307006467.

Ilic:2009:OGI

[ILZ09] Dragana Cvetkovic Ilic, Xiaoji Liu, and Jin Zhong. On the (pq) outer generalized inverse in Banach algebra. Applied Mathematics and Computation, 209(2):191–196, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009314.

Ishiwata:2007:RAM

[IM07] Emiko Ishiwata and Yoshiaki Muroya. Rational approximation method for delay differential equations with proportional delay. Applied Mathematics and Computation, 187(2):741–747, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011283.

Issam:2008:SVR

[IM08] Ben Khediri Issam and Limam Mohamed. Support vector regression based residual MCUSUM control chart for autocorrelated process. Applied Mathematics and Computation, 201(1–2):565–574, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307012350.

Ishiwata:2008:SAO

[IMB08] Emiko Ishiwata, Yoshiaki Muroya, and Hermann Brunner. A superattainable order in collocation methods for differential equations with proportional delay. Applied Mathematics and Computation, 198(1):227–236, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008818.

Iferroudjene:2009:NJL

[IMB09] R. Iferroudjene, K. Abed Meraim, and A. Belouchrani. A new Jacobilike method for joint diagonalization of arbitrary nondefective matrices. Applied Mathematics and Computation, 211(2):363–373, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000836.

Imre:2006:AFD

[Imr06] Attila R. Imre. Artificial fractal dimension obtained by using perimeter– area relationship on digitalized images. Applied Mathematics and Computation, 173(1):443–449, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003644.

Imre:2009:FDT

[Imr09] Attila R. Imre. Fractal dimension of timeindexed paths. Applied Mathematics and Computation, 207(1):90–94, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000222.

Ivanauskas:2009:SDS

[IMS09] F. Ivanauskas, T. Meškauskas, and M. Sapagovas. Stability of difference schemes for two-dimensional parabolic equations with non-local boundary conditions. Applied Mathematics and Computation, 215(7):2716– 2732, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309008108.

Inc:2005:NSO

[Inc05]  Mustafa Inc.  On nu-

merical solutions of onedimensional nonlinear Burgers’ equation and convergence of the decomposition method. Applied Mathematics and Computation, 170(1):76–85, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008719.

Inc:2006:NSS

[Inc06a]  M. Inc.  On numeri-

cal soliton solution of the Kaup–Kupershmidt equation and convergence analysis of the decomposition method. Applied Mathematics and Computation, 172(1):72–85, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001876.

Inc:2006:NDP

[Inc06b] Mustafa Inc. On numerical doubly periodic wave solutions of the coupled Drinfel’d–Sokolov–Wilson equation by the decomposition method. Applied Mathematics and Computation, 172(1):421–430, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002122.

Inc:2006:NJE

[Inc06c]  Mustafa Inc.  On nu-

merical Jacobi elliptic function solutions of the (1 + 1)-dimensional dispersive long wave equation by the decomposition method. Applied Mathematics and Computation, 173(1):372–382, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003577.

Inc:2007:ASW

[Inc07a] Mustafa Inc. An approximate solitary wave solution with compact support for the modified KdV equation. Applied Mathematics and Computation, 184(2):631–637, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007375.

Inc:2007:SET

[Inc07b]  Mustafa Inc.  An L-

stable extended two-step method for the integration of ordinary differential equations. Applied Mathematics and Computation, 186(2):1395–1401, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010654.

Inc:2007:NCN

[Inc07c] Mustafa Inc. New compact and noncompact structures of the nonlinearly dispersive Boussinesq equations. Applied Mathematics and Computation, 189(1):528– 540, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016262.

Inc:2007:NES

[Inc07d] Mustafa Inc. New exact solutions for the ZK-MEW equation by using symbolic computation. Applied Mathematics and Computation, 189(1):508–513, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016249.

Inc:2007:NSM

[Inc07e] Mustafa Inc. New L-stable method for numerical solutions of ordinary differential equations. Applied Mathematics and Computation, 188(1):779–785, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014160.

Inohara:2008:RNE

[Ino08] Takehiro Inohara. Relational Nash equilibrium and interrelationships among relational and rational equilibrium concepts. Applied Mathematics and Computation, 199(2):704– 715, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010570.

Isazadeh:2008:ESA

[INR08] Ayaz Isazadeh, Hamid Nouri, and Maral Rezvan. An evolutionary system architecture for information protection. Applied Mathematics and Computation, 197(2):687–691, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008259.

Ibrahim:2007:IPB

[IOS07] Zarina Bibi Ibrahim, Khairil Iskandar Othman, and Mohamed Suleiman. Implicit r-point block backward differentiation formula for solving firstorder stiff ODEs. Applied Mathematics and Computation, 186(1):558–565, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009672.

Inverardi:2005:SMP

[IPPT05] Pierluigi Novi Inverardi, Alberto Petri, Giorgio Pontuale, and Aldo Tagliani. Stieltjes moment problem via fractional moments. Applied Mathematics and Computation, 166(3):664–677, July 26, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004655.

Ibrahim:2005:NCI

[IR05] Adel K. Ibrahim and Medhat A. Rakha. Numerical computations of infinite products. Applied Mathematics and Computation, 161(1):271–283, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012499.

Islam:2006:FEM

[IR06] Sahidul Islam and Tapan Kumar Roy. A fuzzy EPQ model with flexibility and reliability consideration and demand dependent unit production cost under a space constraint: a fuzzy geometric programming approach. Applied Mathematics and Computation, 176(2):531–544, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008532.

Islam:2007:FMI

[IR07] Sahidul Islam and Tapan Kumar Roy. Fuzzy multiitem economic production quantity model under space constraint: a geometric programming approach. Applied Mathematics and Computation, 184(2):326–335, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007107.

Ilic:2008:CFP

[IR08a] Dejan Ilić and Vladimir Rakočević. Common fixed points for maps on metric space with w-distance. Applied Mathematics and Computation, 199(2):599– 610, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010351.

Israel:2008:VIA

[IR08b] V. P. Israel and M. A. Rincon. Variational inequalities applied to option market problem. Applied Mathematics and Computation, 201(1–2):384–397, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307012167.

Irimia:2006:CFD

[Iri06] Andrei Irimia. Combinatorial formulae for the derivatives of Lamé functions. Applied Mathematics and Computation, 182(2):1501–1508, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004772.

Irmak:2007:SAH

[Irm07]  Hüseyin Irmak.  Some

applications of Hadamard convolution to multivalently analytic and multivalently meromorphic functions. Applied Mathematics and Computation, 187 (1):207–214, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306011532.

Ismail:2005:CSB

[IRSR05] Hassan N. A. Ismail, Kamal R. Raslan, Ghada S. E. Salem, and Aziza A. Abd Rabboh. Comparison study between restrictive Taylor, restrictive Padé approximations and Adomian decomposition method for the solitary wave solution of the General KdV equation. Applied Mathematics and Computation, 167(2):849–869, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400534X.

Iricanin:2009:ECS

[IS09a] Bratislav Iričanin and Stevo Stević. Eventually constant solutions of a rational difference equation. Applied Mathematics and Computation, 215(2):854–856, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309005372.

Iricanin:2009:CTO

[IS09b] Bratislav Iričanin and Stevo Stević. On a class of third-order nonlinear difference equations. Applied Mathematics and Computation, 213(2):479–483, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002598.

Ishikawa:2009:FSG

[Ish09] Keitarou Ishikawa. Fundamentals of simple games from a viewpoint of blockability relations. Applied Mathematics and Computation, 211(1):45– 53, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308009806.

Iskander:2006:EMF

[Isk06] Maged George Iskander. Exponential membership function in stochastic fuzzy goal programming. Applied Mathematics and Computation, 173(2):782–791, February 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305003929.

Iskander:2007:UWM

[Isk07] Maged George Iskander.

Using the weighted max– min approach for stochastic fuzzy goal programming: a case of fuzzy weights. Applied Mathematics and Computation, 188(1):456– 461, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013853.

Islam:2008:MOM

[Isl08]  Sahidul Islam.  Multi-

objective marketing planning inventory model: A geometric programming approach. Applied Mathematics and Computation, 205(1):238–246, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005778.

Ismail:2005:CRP

[Ism05] Hassan N. A. Ismail. On the convergence of the restrictive Padé approximation to the exact solutions of IBVP of parabolic and hyperbolic types. Applied Mathematics and Computation, 162(3):1055–1064, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001493.

Ismail:2006:NST

[Ism06] A. S. Ismail. On the numerical solution of twodimensional singular integral equation. Applied Mathematics and Computation, 173(1):389–393, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305003590.

Ismail:2008:FOE

[Ism08a] M. S. Ismail. A fourthorder explicit schemes for the coupled nonlinear Schrödinger equation. Applied Mathematics and Computation, 196(1):273–284, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006686.

Ismail:2008:NSC

[Ism08b] M. S. Ismail. Numerical solution of complex modified Korteweg–de Vries equation by Petrov–Galerkin method. Applied Mathematics and Computation, 202(2):520–531, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001276.

Inverardi:2006:DDM

[IT06a] P. L. Novi Inverardi and A. Tagliani. Discrete distributions from moment generating function. Applied Mathematics and Computation, 182(1):200–209, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003262.

Ivaz:2006:CTR

[IT06b] K. Ivaz and H. Taghvafard. Comparison of the trapezoidal rule error with the spline rule error. Applied Mathematics and Computation, 183(2):1256–1261, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006710.

Ivanov:2008:MSD

[Iva08] Ivan Ganchev Ivanov. A method to solve the discrete-time coupled algebraic Riccati equations. Applied Mathematics and Computation, 206(1):34– 41, December 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308006255.

Iftekharuddin:2009:FBB

[IZIO09] Khan M. Iftekharuddin, Jing Zheng, Mohammad A. Islam, and Robert J. Ogg. Fractal-based brain tumor detection in multimodal MRI. Applied Mathematics and Computation, 207(1):23–41, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000179.

Jahanshahloo:2008:GPC

[JA08] G. R. Jahanshahloo and M. Afzalinejad. Goal programming in the context of the assignment problem and a computationally effective solution method. Applied Mathematics and Computation, 200(1):34– 40, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030701082X.

Jahangiri:2007:GBB

[Jah07] Jay M. Jahangiri. Generalized bounded boundary turning functions. Applied Mathematics and Computation, 187(1):215–218, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011544.

Jain:2007:STM

[Jai07] Pankaj Jain. Steffensen type methods for solving non-linear equations. Applied Mathematics and Computation, 194(2):527– 533, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030700505X.

Jang:2007:ESO

[Jan07a]  Bongsoo Jang.  Exact

solutions to one dimensional non-homogeneous parabolic problems by the homogeneous Adomian decomposition method. Applied Mathematics and Computation, 186(2):969– 979, March 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306010228.

Jang:2007:SNH

[Jan07b] Bongsoo Jang. Solutions to the non-homogeneous parabolic problems by the extended HADM. Applied Mathematics and Computation, 191(2):466–483, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002755.

Jankowski:2008:DAP

[Jan08a] Tadeusz Jankowski. On differential-algebraic problems. Applied Mathematics and Computation, 196(2): 542–549, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307007060.

Jankowski:2008:PST

[Jan08b] Tadeusz Jankowski. Positive solutions of threepoint boundary value problems for second order impulsive differential equations with advanced arguments. Applied Mathematics and Computation, 197(1):179–189, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007680.

Jankowski:2008:PSS

[Jan08c] Tadeusz Jankowski. Positive solutions to second order four-point boundary value problems for impulsive differential equations. Applied Mathematics and Computation, 202(2):550–561, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001409.

Jankowski:2009:DED

[Jan09a] Tadeusz Jankowski. On dynamic equations with deviating arguments. Applied Mathematics and Computation, 208(2):423–426, February 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308009107.

Jankowski:2009:NTP

[Jan09b] Tadeusz Jankowski. On nonlinear two-point boundary value problems for impulsive differential-algebraic problems. Applied Mathematics and Computation, 213(2):380–388, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002409.

Jankowski:2009:PST

[Jan09c] Tadeusz Jankowski. Positive solutions for threepoint one-dimensional p-Laplacian boundary value problems with advanced arguments. Applied Mathematics and Computation, 215(1):125–131, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309003919.

Jordan:2009:CPM

[JAT09] C. Jordán, C. Mendes Araújo, and Juan R. Torregrosa. N_{0} completions on partial matrices. Applied Mathematics and Computation, 211(2):303–312, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000769.

Javidi:2006:NSGa

[Jav06a] M. Javidi. A numerical solution of the generalized Burger’s–Huxley equation by pseudospectral method and Darvishi’s preconditioning. Applied Mathematics and Computation, 175(2):1619–1628, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007198.

Javidi:2006:NSGb

[Jav06b] M. Javidi. A numerical solution of the generalized Burger’s–Huxley equation by spectral collocation method. Applied Mathematics and Computation, 178(2):338–344, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009793.

Javidi:2006:PMD

[Jav06c] M. Javidi. Pseudospectral method and Darvishi’s preconditioning for solving system of time dependent partial differential equations. Applied Mathematics and Computation, 176(1):334–340, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008088.

Javidi:2006:SCM

[Jav06d] Mohammad Javidi. Spectral collocation method for the solution of the generalized Burger–Fisher equation. Applied Mathematics and Computation, 174(1): 345–352, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305004613.

Javidi:2007:IMN

[Jav07] M. Javidi. Iterative methods to nonlinear equations. Applied Mathematics and Computation, 193(2):360–365, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004055.

Jang:2005:DMN

[JB05] S. R.-J. Jang and J. Baglama. Droop models of nutrient– plankton interaction with intratrophic predation. Applied Mathematics and Computation, 169(2):1106– 1128, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304008409.

Bai:2006:INM

[jB06] Zheng jian Bai. Inexact Newton methods for inverse eigenvalue problems. Applied Mathematics and Computation, 172(2):682–689, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009026.

Jiang:2008:PBP

[JB08] Zhikai Jiang and Abdellah Bnouhachem. A projection-based prediction– correction method for structured monotone variational inequalities. Applied Mathematics and Computation, 202(2):747–759, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001756.

Jbilou:2006:LRA

[Jbi06]  K. Jbilou.  Low rank

approximate solutions to large Sylvester matrix equations. Applied Mathematics and Computation, 177(1):365–376, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305009173.

Jimenez:2005:RCL

[JC05] J. C. Jimenez and F. Carbonell. Rate of convergence of local linearization schemes for initialvalue problems. Applied Mathematics and Computation, 171(2):1282–1295, December 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jiao:2006:PMS

[JC06] Jianjun Jiao and Lansun Chen. A pest management SI model with periodic biological and chemical control concern. Applied Mathematics and Computation, 183(2):1018–1026, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006370.

Jwo:2007:PNE

[JC07] Dah-Jing Jwo and Ta-Shun Cho. A practical note on evaluating Kalman filter performance optimality and degradation. Applied Mathematics and Computation, 193(2):482–505, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004316.

Jiang:2008:ESS

[JC08a] Wei Jiang and Minggen Cui. The exact solution and stability analysis for integral equation of third or first kind with singular kernel. Applied Mathematics and Computation, 202(2):666–674, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001549.

Jiao:2008:NDG

[JC08b] Hongwei Jiao and Yongqiang Chen. A note on a deterministic global optimization algorithm. Applied Mathematics and Computation, 202(1):67–70, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800060X.

Jiang:2009:CPE

[JC09] Wei Jiang and Minggen Cui. Constructive proof for existence of nonlinear two-point boundary value problems. Applied Mathematics and Computation, 215(5):1937–1948, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900695X.

Jiao:2008:APM

[JCL08] Jianjun Jiao, Lansun Chen, and Guilie Luo. An appropriate pest management SI model with biological and chemical control concern. Applied Mathematics and Computation, 196(1):285–293, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006698.

Jodar:2005:LCC

[JCM05] L. Jódar and L. A. Caudillo-Mata. A low computational cost numerical method for solving mixed diffusion problems. Applied Mathematics and Computation, 170(1):673–685, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000093.

Jarboui:2007:CPS

[JCSR07] B. Jarboui, M. Cheikh, P. Siarry, and A. Rebai. Combinatorial particle swarm optimization (CPSO) for partitional clustering problem. Applied Mathematics and Computation, 192(2):337–345, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003347.

Jiang:2007:SHB

[JCX^{+}07] Xiaoyue Jiang, Yinglei Cheng, Rong Xiao, Ying Li, and Rongchun Zhao. Spherical harmonic based linear face de-lighting and compensation. Applied Mathematics and Computation, 185(2):857–868, February 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306007673.

Jafari:2006:PSN

[JDG06a] Hossein Jafari and Varsha Daftardar-Gejji. Positive solutions of nonlinear fractional boundary value problems using Adomian decomposition method. Applied Mathematics and Computation, 180(2):700–706, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000580.

Jafari:2006:RADa

[JDG06b] Hossein Jafari and Varsha Daftardar-Gejji. Revised Adomian decomposition method for solving a system of nonlinear equations. Applied Mathematics and Computation, 175(1): 1–7, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005862.

Jafari:2006:RADb

[JDG06c] Hossein Jafari and Varsha Daftardar-Gejji. Revised Adomian decomposition method for solving systems of ordinary and fractional differential equations. Applied Mathematics and Computation, 181(1):598–608, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001305.

Jafari:2006:SLN

[JDG06d] Hossein Jafari and Varsha Daftardar-Gejji. Solving linear and nonlinear fractional diffusion and wave equations by Adomian decomposition. Applied Mathematics and Computation, 180(2):488–497, September 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300306000105.

Jarboui:2008:CPS

[JDSR08] B. Jarboui, N. Damak, P. Siarry, and A. Rebai. A combinatorial particle swarm optimization for solving multimode resource-constrained project scheduling problems. Applied Mathematics and Computation, 195(1):299–308, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700553X.

Jerez:2008:NDI

[Jer08] Silvia Jerez. A nonstandard difference-integral method for the viscous Burgers’ equation. Applied Mathematics and Computation, 200(1):378–386, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011423.

Jahanshahloo:2005:EAG

[JF05] G. R. Jahanshahloo and A. A. Foroughi. Efficiency analysis, generating an efficient extreme point for an MOLP, and some comparisons. Applied Mathematics and Computation, 162(2):991–1005, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001444.

Fan:2009:HMV

[jF09] Qing ju Fan. Horseshoe in a modified Van der Pol– Duffing circuit. Applied Mathematics and Computation, 210(2):436–440, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000113.

Ji:2008:ESP

[JFG08a] Dehong Ji, Hanying Feng, and Weigao Ge. The existence of symmetric positive solutions for some nonlinear equation systems. Applied Mathematics and Computation, 197(1):51–59, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307007527.

Ji:2008:MPS

[JFG08b] Dehong Ji, Meiqiang Feng, and Weigao Ge. Multiple positive solutions for multipoint boundary value problems with sign changing nonlinearity. Applied Mathematics and Computation, 196(2):511–520, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007035.

Javidi:2007:NSS

[JG07a] M. Javidi and A. Golbabai.

A numerical solution for solving system of Fredholm integral equations by using homotopy perturbation method. Applied Mathematics and Computation, 189(2):1921–1928, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000112.

Ji:2007:MPS

[JG07b] Dehong Ji and Weigao Ge. Multiple positive solutions for some p-Laplacian boundary value problems. Applied Mathematics and Computation, 187(2):1315–1325, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012707.

Jiao:2006:GOG

[JGS06] Hongwei Jiao, Yunrui Guo, and Peiping Shen. Global optimization of generalized linear fractional programming with nonlinear constraints. Applied Mathematics and Computation, 183(2):717–728, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006060.

Jenabi:2007:THM

[JGTK07] M. Jenabi, S. M. T. Fatemi Ghomi, S. A. Torabi, and B. Karimi. Two hybrid meta-heuristics for the finite horizon ELSP in flexible flow lines with unrelated parallel machines. Applied Mathematics and Computation, 186(1):230– 245, March 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306009337.

Ji:2007:ESP

[JGY07] Dehong Ji, Weigao Ge, and Yitao Yang. The existence of symmetric positive solutions for Sturm– Liouville-like four-point boundary value problem with a p-Laplacian operator. Applied Mathematics and Computation, 189(2):1087–1098, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016857.

Ji:2007:PSS

[JGZ07] Yude Ji, Yanping Guo, and Jiehua Zhang. Positive solutions for secondorder quasilinear multipoint boundary value problems. Applied Mathematics and Computation, 189 (1):725–731, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306016493.

Jiang:2008:CRI

[JH08a] Dongdong Jiang and Danfu Han. The construction of rational iterative methods for solving nonlinear equations. Applied Mathematics and Computation, 204(2):802–808, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005638.

Jiang:2008:SOP

[JH08b] Dongdong Jiang and Danfu Han. Some one-parameter families of third-order methods for solving nonlinear equations. Applied Mathematics and Computation, 195(2):392–396, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005589.

Juan:2008:SDP

[JH08c] Justie Su-Tzu Juan and Chun-Ming Huang. On the strong distance problems of pyramid networks. Applied Mathematics and Computation, 195(1):154–161, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005279.

Jiang:2009:GAG

[JH09] Jinping Jiang and Yanren Hou. The global attractor of g-Navier–Stokes equations with linear dampness on R^{2}. Applied Mathematics and Computation, 215(3):1068–1076, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006109.

Jiang:2007:IPS

[JHHW07] Yan Jiang, Tiesong Hu, ChongChao Huang, and Xianing Wu. An improved particle swarm optimization algorithm. Applied Mathematics and Computation, 193(1):231–239, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700392X.

Jahanshahloo:2005:SSA

[JHS^{+}05] G. R. Jahanshahloo, F. Hosseinzadeh, N. Shoja, M. Sanei, and G. Tohidi. Sensitivity and stability analysis in DEA. Applied Mathematics and Computation, 169(2):897–904, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008203.

Jahanshahloo:2005:MGA

[JHST05] G. R. Jahanshahloo, F. Hosseinzadeh, N. Shoja, and G. Tohidi. A method for generating all efficient solutions of 0-1 multiobjective linear programming problem. Applied Mathematics and Computation, 169(2):874–886, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008185.

Ji:2005:MAS

[Ji05] Xiaoyu Ji. Models and algorithm for stochastic shortest path problem. Applied Mathematics and Computation, 170(1):503–514, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009646.

Ji:2008:NEN

[Ji08] Jie Ji. A new expansion and new families of exact traveling solutions to Hirota equation. Applied Mathematics and Computation, 204(2):881–883, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005730.

Jian:2005:NFD

[Jia05] Jin-Bao Jian. A new feasible descent algorithm combining SQP with generalized projection for optimization problems without strict complementarity. Applied Mathematics and Computation, 162(3):1065–1081, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400150X.

Jia:2007:BHM

[Jia07a] Baoguo Jia. Bounds of the Hausdorff measure of the Koch curve. Applied Mathematics and Computation, 190(1):559–565, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000896.

Jiang:2007:ACM

[Jia07b]  Yingjun Jiang.  Adap-

tive collocation methods for Volterra integral and integro-differential equations. Applied Mathematics and Computation, 191(1):67–78, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700015X.

Jiang:2009:EMP

[Jia09] Weihua Jiang. The existence of multiple positive periodic solutions for functional differential equations. Applied Mathematics and Computation, 208(1):165–171, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008771.

Jin:2008:SDE

[Jin08] Li Jin. A stable differential equation approach for inequality constrained optimization problems. Applied Mathematics and Computation, 206(1):186–192, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006528.

Jauberteau:2009:NDN

[JJ09a] F. Jauberteau and J. L. Jauberteau. Numerical differentiation with noisy signal. Applied Mathematics and Computation, 215(6):2283–2297, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030900753X.

Jovanovic:2009:NSF

[JJ09b] Miljana Jovanović and Svetlana Janković. Neutral stochastic functional differential equations with additive perturbations. Applied Mathematics and Computation, 213(2):370–379, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002392.

Ji:2006:CTR

[JjQjWmL06] Ying Ji, Shao jian Qu, Yan jun Wang, and Hui min Li. A conic trust-region method for optimization with nonlinear equality and inequality constrains via active-set strategy. Applied Mathematics and Computation, 183(1):217–231, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005546.

Ji:2006:ISA

[JJT06] Mingjun Ji, Zhihong Jin, and Huanwen Tang. An improved simulated annealing for solving the linear constrained optimization problems. Applied Mathematics and Computation, 183(1):251–259, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005571.

Jeong:2006:WSC

[JK06] Jae Ug Jeong and Soo Hwan Kim. Weak and strong convergence of the Ishikawa iteration process with errors for two asymptotically nonexpansive mappings. Applied Mathematics and Computation, 181(2):1394–1401, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600258X.

Jayanthi:2007:EVV

[JK07] S. Jayanthi and M. Kumari. Effect of variable viscosity on non-Darcy free or mixed convection flow on a vertical surface in a non-Newtonian fluid saturated porous medium. Applied Mathematics and Computation, 186(2):1643–1659, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011003.

Jeong:2008:SAS

[JK08] Jae Ug Jeong and Soo Hwan Kim. Sensitivity analysis for a system of generalized nonlinear mixed quasi-variational inclusions in q-uniformly smooth Banach spaces. Applied Mathematics and Computation, 202(1):178–187, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000726.

Jangveladze:2009:FDA

[JKN09] Temur Jangveladze, Zurab Kiguradze, and Beny Neta. Finite difference approximation of a nonlinear integro-differential system. Applied Mathematics and Computation, 215(2):615–628, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005402.

Jin:2005:NMA

[JL05] Chengri Jin and Mingzhu Liu. A new modification of Adomian decomposition method for solving a kind of evolution equation. Applied Mathematics and Computation, 169(2):953–962, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008252.

Jia:2007:QAS

[JL07] Jianwen Jia and Qiuying Li. Qualitative analysis of an SIR epidemic model with stage structure. Applied Mathematics and Computation, 193(1):106–115, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307003773.

Jenkins:2008:DDM

[JL08] Eleanor Jenkins and Hyesuk Lee. A domain decomposition method for the Oseen-viscoelastic flow equations. Applied Mathematics and Computation, 195(1):127–141, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005255.

Jin:2009:IDA

[JL09] Da-Wei Jin and Jun Lu. An improved determination approach to the structure and parameters of dynamic structure-based neural networks. Applied Mathematics and Computation, 215(7):2787–2797, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309008169.

Jonckheere:2007:UBS

[JLA07] E. A. Jonckheere, P. Lohsoonthorn, and F. Ariaei. Upper bound on scaled Gromov-hyperbolic δ. Applied Mathematics and Computation, 192(1):191– 204, September 1, 2007.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307003219.

Jahanshahloo:2007:DAI

[JLBR07] G. R. Jahanshahloo, F. Hosseinzadeh Lotfi, F. Rezai Balf, and H. Zhiani Rezai. Discriminant analysis of interval data using Monte Carlo method in assessment of overlap. Applied Mathematics and Computation, 191(2):521–532, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002809.

Jahanshahloo:2008:UMC

[JLBR08] G. R. Jahanshahloo, F. Hosseinzadeh Lotfi, F. Rezai Balf, and H. Zhiani Rezai. Using Monte Carlo method for ranking interval data. Applied Mathematics and Computation, 201(1–2): 613–620, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308000076.

Jia:2008:PIU

[JLCH08] Wei Jia, Bin Ling, Kwok-Wing Chau, and Laurent Heutte. Palmprint identification using restricted fusion. Applied Mathematics and Computation, 205(2):927–934, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800372X.

Jahanshahloo:2006:AME

[JLI06a] G. R. Jahanshahloo, F. Hosseinzadeh Lotfi, and M. Izadikhah. An algorithmic method to extend TOPSIS for decision-making problems with interval data. Applied Mathematics and Computation, 175(2):1375–1384, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006843.

Jahanshahloo:2006:ETM

[JLI06b] G. R. Jahanshahloo, F. Hosseinzadeh Lotfi, and M. Izadikhah. Extension of the TOP-SIS method for decisionmaking problems with fuzzy data. Applied Mathematics and Computation, 181(2):1544–1551, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002852.

Juang:2008:CCC

[JLL08] Jih-Gau Juang, Ren-Wei Lin, and Wen-Kai Liu. Comparison of classical control and intelligent control for a MIMO system. Applied Mathematics and Computation, 205(2):778–791, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003561.

Jahanshahloo:2005:DAF

[JLM05] G. R. Jahanshahloo, F. Hosseinzadeh Lotfi, and M. Moradi. A DEA approach for fair allocation of common revenue. Applied Mathematics and Computation, 160(3):719–724, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303011901.

Jahanshahloo:2005:UMC

[JLRB05] G. R. Jahanshahloo, F. Hosseinzadeh Lotfi, H. Zhiani Rezai, and F. Rezai Balf. Using Monte Carlo method for ranking efficient DMUs. Applied Mathematics and Computation, 162(1):371–379, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001018.

Jahanshahloo:2005:INE

[JLS^{+}05a] G. R. Jahanshahloo, F. Hosseinzadeh Lotfi, N. Shoja, G. Tohidi, and S. Razavyan. The identification of nondominated and efficient paths on a network. Applied Mathematics and Computation, 169(2):866–873, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008173.

Jahanshahloo:2005:OMA

[JLS^{+}05b] G. R. Jahanshahloo, F. Hosseinzadeh Lotfi, N. Shoja, G. Tohidi, and S. Razavyan. A one-model approach to classification and sensitivity analysis in DEA. Applied Mathematics and Computation, 169(2):887–896, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008197.

Jahanshahloo:2005:SEC

[JLS^{+}05c] G. R. Jahanshahloo, F. Hosseinzadeh Lotfi, N. Shoja, G. Tohidi, and S. Razavyan. Sensitivity of efficiency classifications in the inverse DEA models. Applied Mathematics and Computation, 169(2):905–916, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008215.

Jahanshahloo:2005:UIO

[JLS^{+}05d] G. R. Jahanshahloo, F. Hosseinzadeh Lotfi, N. Shoja, G. Tohidi, and S. Razavyan. Undesirable inputs and outputs in DEA models. Applied Mathematics and Computation, 169(2):917–925, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008227.

Jahanshahloo:2006:EDI

[JLS06] G. R. Jahanshahloo, F. Hosseinzadeh Lotfi, and S. Sohraiee. Egoist’s dilemma with interval data. Applied Mathematics and Computation, 183(1):94–105, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005418.

Jiang:2009:SON

[JLW09] Jiqiang Jiang, Lishan Liu, and Yonghong Wu. Second-order nonlinear singular Sturm–Liouville problems with integral boundary conditions. Applied Mathematics and Computation, 215(4):1573– 1582, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300309006602.

Jahanshahloo:2005:FES

[JLZ05a] G. R. Jahanshahloo, F. Hosseinzadeh Lotfi, and M. Zohrehbandian. Finding the efficiency score and RTS characteristic of DMUs by means of identifying the efficient frontier in DEA. Applied Mathematics and Computation, 170(2):985–993, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000391.

Jahanshahloo:2005:FPL

[JLZ05b] G. R. Jahanshahloo, F. Hosseinzadeh Lotfi, and M. Zohrehbandian. Finding the piecewise linear frontier production function in data envelopment analysis. Applied Mathematics and Computation, 163(1):483– 488, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400236X.

Jain:2006:DCR

[JM06a] Madhu Jain and Sandhya Maheshwari. Discounted costs for repairable units under hybrid warranty. Applied Mathematics and Computation, 173(2):887–901, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004005.

Jun:2006:DDM

[JM06b] Younbae Jun and Tsun-Zee Mai. Domain decomposition method for parabolic problems with Neumann conditions. Applied Mathematics and Computation, 182(2):1683–1695, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005170.

Jun:2006:IDD

[JM06c] Younbae Jun and Tsun-Zee Mai. IPIC domain decomposition algorithm for parabolic problems.

Applied Mathematics and Computation, 177(1):352– 364, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009161.

Jana:2007:GPD

[JM07] P. Jana and S. K. Mazumder. Geometric programming and dual entropy-optimization problem. Applied Mathematics and Computation, 193(2):340–345, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004031.

Jing-mei:2005:CRB

[JmBdXgXm05] Liu Jing-mei, Wei Baodian, Cheng Xiang-guo, and Wang Xin-mei. Cryptanalysis of Rijndael S-box and improvement. Applied Mathematics and Computation, 170(2):958–975, November 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305000378.

Jiao:2007:SSH

[JMC07] Jianjun Jiao, Xinzhu Meng, and Lansun Chen. A stagestructured Holling mass defence predator–prey model with impulsive perturbations on predators. Applied Mathematics and Computation, 189(2):1448–1458, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017310.

Jahanshahloo:2005:FIW

[JMHS05] G. R. Jahanshahloo, A. Memariani, F. Hosseinzadeh, and N. Shoja. A feasible interval for weights in data envelopment analysis. Applied Mathematics and Computation, 160(1):155–168, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303010993.

Jahanshahloo:2005:NSD

[JMLR05] G. R. Jahanshahloo, A. Memariani, F. Hosseinzadeh Lotfi, and H. Z. Rezai. A note on some of DEA models and finding efficiency and complete ranking using common set of weights. Applied Mathematics and Computation, 166(2):265–281, July 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004357.

Jiang:2005:OET

[JOA05] Daqing Jiang, Donal O’Regan, and R. P. Agarwal. Optimal existence theory for single and multiple positive periodic solutions to functional difference equations. Applied Mathematics and Computation, 161(2):441–462, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304000049.

Jesic:2008:CFP

[JOB08] Siniša N. Ješić, Donal O’Regan, and Nataša A. Babačev. A common fixed point theorem for R-weakly commuting mappings in probabilistic spaces with nonlinear contractive conditions. Applied Mathematics and Computation, 201(1–2):272–281, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011940.

Jonsson:2005:UPD

[Jön05] Kristian Jönsson. Using panel data to increase the power of modified unit root tests in the presence of structural breaks. Applied Mathematics and Computation, 171(2):832–842, December 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Joyner:2006:NSP

[Joy06] Michele L. Joyner. A numerical study of the POD method in NDE. Applied Mathematics and Computation, 174(1):732–754, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004984.

Jackiewicz:2009:SCG

[JP09] Z. Jackiewicz and R. Park.

A strategy for choosing Gegenbauer reconstruction parameters for numerical stability. Applied Mathematics and Computation, 212(2):418–434, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900174X.

Jiao:2008:DSS

[JPCL08] Jianjun Jiao, Guoping Pang, Lansun Chen, and Guilie Luo. A delayed stage-structured predator– prey model with impulsive stocking on prey and continuous harvesting on predator. Applied Mathematics and Computation, 195(1):316–325, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005553.

Ji:2009:RMS

[JPYW09] D. H. Ji, Ju H. Park, W. J. Yoo, and S. C. Won. Robust memory state feedback model predictive control for discrete-time uncertain state delayed systems. Applied Mathematics and Computation, 215(6):2035–2044, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309007115.

Jahanshahloo:2006:MMM

[JPZ06] G. R. Jahanshahloo, L. Pourkarimi, and M. Zarepisheh. Modified MAJ model for ranking decision making units in data envelopment analysis. Applied Mathematics and Computation, 174(2):1054–1059, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005321.

Jung:2007:UPA

[JR07] Soon-Mo Jung and Themistocles M. Rassias. Ulam’s problem for approximate homomorphisms in connection with Bernoulli’s differential equation. Applied Mathematics and Computation, 187(1):223–227, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011568.

Jolai:2007:GAB

[JRA^{+}07] Fariborz Jolai, M. Rabbani, S. Amalnick, A. Dabaghi, M. Dehghan, and M. Yazadn Parast. Genetic algorithm for bi-criteria single machine scheduling problem of minimizing maximum earliness and number of tardy jobs. Applied Mathematics and Computation, 194(2):552–560, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005103.

Jolai:2007:EAB

[JRR^{+}07] Fariborz Jolai, M. J.

Rezaee, M. Rabbani, J. Razmi, and Pariviz Fattahi. Exact algorithm for biobjective 0-1 knapsack problem. Applied Mathematics and Computation, 194(2):544–551, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005097.

Jackiewicz:2008:NSF

[JRW08] Z. Jackiewicz, M. Rahman, and B. D. Welfert. Numerical solution of a Fredholm integro-differential equation modelling θ̇-neural networks. Applied Mathematics and Computation, 195(2):523–536, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005826.

Jukic:2005:LSF

[JS05] Dragan Jukić and Rudolf Scitovski. Least squares fitting Gaussian type curve. Applied Mathematics and Computation, 167(1):286– 298, August 5, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304004928.

Ji:2006:MAB

[JS06] Xiaoyu Ji and Zhen Shao. Model and algorithm for bilevel newsboy problem with fuzzy demands and discounts. Applied Mathematics and Computation, 172(1):163–174, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001955.

Jator:2007:HOB

[JS07a] Samuel Jator and Zachariah Sinkala. A high order B-spline collocation method for linear boundary value problems. Applied Mathematics and Computation, 191(1):100–116, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001828.

Jiao:2007:NPG

[JS07b] Hongwei Jiao and Peiping Shen. A note on the paper global optimization of nonlinear sum of ratios. Applied Mathematics and Computation, 188(2):1812–1815, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015499.

Jahanshahloo:2007:CCC

[JSdM07] G. R. Jahanshahloo, M. Soleimanidamaneh, and A. Mostafaee. On the computational complexity of cost efficiency analysis models. Applied Mathematics and Computation, 188(1):638–640, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014019.

Jahanshahloo:2005:EPE

[JSdRm05] G. R. Jahanshahloo, M. Soleimanidamaneh, and M. Rostamymalkhalifeh. An enhanced procedure for estimating returns-to-scale in DEA. Applied Mathematics and Computation, 171(2):1226– 1238, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jia:2005:CEC

[JSL05] Gao Jia, Ying-Qing Song, and Qiong Li. Calculation of eigenvalues for a class of crosswise vibration equation of the beam. Applied Mathematics and Computation, 165(1):143– 154, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003686.

Jiang:2006:BGE

[JSL06a] Minghui Jiang, Yi Shen, and Xiaoxin Liao. Boundedness and global exponential stability for generalized Cohen–Grossberg neural networks with variable delay. Applied Mathematics and Computation, 172(1):379–393, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002092.

Jiang:2006:GSP

[JSL06b] Minghui Jiang, Yi Shen, and Xiaoxin Liao. Global stability of periodic solution for bidirectional associative memory neural networks with varying-time delay. Applied Mathematics and Computation, 182(1):509–520, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003584.

Jiang:2007:NBP

[JSLL07] Minghui Jiang, Yi Shen, Haigeng Luo, and Xiaoxin Liao. Nonlinear behavior of the parameterized Logistic differential systems. Applied Mathematics and Computation, 189(2):1694–1704, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017619.

Jiang:2005:SAI

[JSW05] Zhengtao Jiang, Xi Sun, and Yumin Wang. Security analysis and improvement of a double-trapdoor encryption scheme. Applied Mathematics and Computation, 169(1):41–50, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304007465.

Jolai:2006:EPL

[JTMRS06] Fariborz Jolai, R. Tavakkoli-Moghaddam, M. Rabbani, and M. R. Sadoughian. An economic production lot size model with deteriorating items, stock-dependent demand, inflation, and partial backlogging. Applied Mathematics and Computation, 181(1):380–389, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600107X.

Jun:2009:EDD

[Jun09a]  Younbae Jun.  An ef-

ficient domain decomposition method for threedimensional parabolic problems. Applied Mathematics and Computation, 215(8):2815–2825, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008200.

Jung:2009:SCC

[Jun09b] Jong Soo Jung. Strong convergence of composite iterative methods for equilibrium problems and fixed point problems. Applied Mathematics and Computation, 213(2):498–505, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002628.

Jovanovic:2005:SDS

[JV05] Bosko S. Jovanović and Lubin G. Vulkov. Stability of difference schemes for parabolic equations with dynamical boundary conditions and conditions on conjugation. Applied Mathematics and Computation, 163(2):849–868, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002656.

Joshi:2007:ETS

[JV07] C. M. Joshi and Yashoverdhan Vyas. Extensions of two q-series expansions with applications to biorthogonal rational functions. Applied Mathematics and Computation, 187(1):219–222, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011556.

Jahanshahloo:2004:IOE

[JVFM04] G. R. Jahanshahloo, A. Hadi Vencheh, A. A. Foroughi, and R. Kazemi Matin. Inputs/outputs estimation in DEA when some factors are undesirable. Applied Mathematics and Computation, 156(1):19–32, August 25, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See comment [HV08].

Wang:2006:DPT

[jW06a] Cheng jing Wang. Dogleg paths and trust region methods with back tracking technique for unconstrained optimization. Applied Mathematics and Computation, 177(1):159– 169, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008970.

Wang:2006:SRC

[jW06b] Cheng jing Wang. Some remarks on conjugate gradient methods without line search. Applied Mathematics and Computation, 181(1):370–379, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001068.

Jia:2007:ABS

[JW07] Yunfeng Jia and Jianhua Wu. Asymptotic behavior of solutions of a periodic reaction–diffusion system of a three-species of competition model. Applied Mathematics and Computation, 189(2):1683–1693, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017607.

Jia:2009:ORP

[JW09a] Jishen Jia and Shaomin Wu. Optimizing replacement policy for a coldstandby system with waiting repair times. Applied Mathematics and Computation, 214(1):133–141, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002872.

Jia:2009:SSA

[JW09b] Zhigang Jia and Musheng Wei. Solvability and sensitivity analysis of polynomial matrix equation X^{s} + A^{T} X^{t}A = Q. Applied Mathematics and Computation, 209(2):230–237, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009399.

Jia:2007:RRB

[JWL^{+}07] Lifeng Jia, Zhe Wang, Nan Lu, Xiujuan Xu, Dongbin Zhou, and Yan Wang. RFIMiner: a regressionbased algorithm for recently frequent patterns in multiple time granularity data streams. Applied Mathematics and Computation, 185(2):769–783, February 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306007600.

Jwo:2007:RKF

[Jwo07]  Dah-Jing Jwo.  Re-

marks on the Kalman filtering simulation and verification. Applied Mathematics and Computation, 186(1):159–174, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600926X.

Jebens:2008:EMS

[JWPS08] Stefan Jebens, Rüdiger Weiner, Helmut Podhaisky, and Bernhard A. Schmitt. Explicit multi-step peer methods for special secondorder differential equations. Applied Mathematics and Computation, 202(2):803–813, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001847.

Jiang:2009:GAP

[JWS^{+}09] Yu Jiang, Huiming Wei, Xinyu Song, Liquan Mei, Guanghui Su, and Suizheng Qiu. Global attractivity and permanence of a delayed SVEIR epidemic model with pulse vaccination and saturation incidence. Applied Mathematics and Computation, 213(2):312–321, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002331.

Jian:2006:SCN

[JXH06] Jin-Bao Jian, Qing-Juan Xu, and Dao-Lan Han. A strongly convergent normrelaxed method of strongly sub-feasible direction for optimization with nonlinear equality and inequality constraints. Applied Mathematics and Computation, 182(1):854–870, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004036.

Jiang:2005:RNP

[JXW05] Zhengtao Jiang, Mingsen Xiang, and Yumin Wang. A research on new publickey encryption schemes. Applied Mathematics and Computation, 169(1):51– 61, October 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304007519.

Jiang:2006:PAP

[JY06] Peng Jiang and Geng Yang. Performance analysis of preconditioners based on Broyden method. Applied Mathematics and Computation, 178(2):295–308, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009756.

Jeon:2007:PCA

[JY07a] Jun-Cheol Jeon and Kee-Young Yoo. Programmable cellular automata based Montgomery hardware architecture. Applied Mathematics and Computation, 186(1):915–922, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306010149.

Jeon:2007:RDM

[JY07b] Jun-Cheol Jeon and Kee Young Yoo. Role delegation on multiple-attractor cellular automata based on pseudo-exhaustive bit patterns. Applied Mathematics and Computation, 186(2):1423–1428, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601068X.

Jiang:2009:BAS

[JY09] Guirong Jiang and Qigui Yang. Bifurcation analysis in an SIR epidemic model with birth pulse and pulse vaccination. Applied Mathematics and Computation, 215(3):1035–1046, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309005979.

Jiang:2009:OIB

[JyGcL^{+}09] Xin Jiang, Xiu yun Gao, Sheng chang Li, Yu ren Shi, and Wen shan Duan. Oblique interactions between multi-solitons in plasma. Applied Mathematics and Computation, 214(1):60–65, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002781.

Jisheng:2006:UCT

[JYX06] Kou Jisheng, Li Yitian, and Wang Xiuhua. A uniparametric Chebyshevtype method free from second derivatives. Applied Mathematics and Computation, 179(1):296–300, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010374.

Jiang:2006:ENS

[JZ06a] Liqun Jiang and Zhan Zhou. Existence of nontrivial solutions for discrete nonlinear two point boundary value problems. Applied Mathematics and Computation, 180(1):318–329, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305011276.

Jiang:2006:EEF

[JZ06b] Ying-Jun Jiang and Jin-Ping Zeng. An L^{∞}- error estimate for finite element solution of nonlinear elliptic problem with a source term. Applied Mathematics and Computation, 174(1):134–149, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004467.

Jiao:2006:EMA

[JZ06c] Xiaoyu Jiao and Hong-Qing Zhang. An extended method and its application to Whitham– Broer–Kaup equation and two-dimensional perturbed KdV equation. Applied Mathematics and Computation, 172(1):664–677, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002511.

Jiang:2007:ASA

[JZ07] Ying-Jun Jiang and Jin-Ping Zeng. Additive Schwarz algorithm for the nonlinear complementarity problem with Mfunction. Applied Mathematics and Computation, 190(2):1007–1019, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014561.

Ji:2008:QBS

[JZ08a] Ruihong Ji and Sining Zheng. Quenching behavior of solutions to heat equations with coupled boundary singularities. Applied Mathematics and Computation, 206(1):403–412, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006899.

Jia:2008:ASI

[JZ08b] Chunxia Jia and Detong Zhu. An affine scaling interior algorithm via Lanczos path for solving bound-constrained nonlinear systems. Applied Mathematics and Computation, 195(2):558–575, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005851.

Jiang:2008:MNS

[JZ08c] Liqun Jiang and Zhan Zhou. Multiple nontrivial solutions for a class of higher dimensional discrete boundary value problems. Applied Mathematics and Computation, 203(1): 30–38, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307011873.

Jiang:2005:TMD

[JZCW05] Zhengtao Jiang, Yang Zhan, Dan Chen, and Yumin Wang. Two methods of directly constructing probabilistic public-key encryption primitives based on third-order LFSR sequences. Applied Mathematics and Computation, 171(2):900–911, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jie:2007:EME

[JZH07] Jing Jie, Jianchao Zeng, and Chongzhao Han. An extended mind evolutionary computation model for optimizations. Applied Mathematics and Computation, 185(2):1038–1049, February 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306008289.

Jian:2005:NNR

[JZHT05] Jinbao Jian, Haiyan Zheng, Qingjie Hu, and Chunming Tang. A new norm-relaxed method of strongly subfeasible direction for inequality constrained optimization. Applied Mathematics and Computation, 168(1):1–28, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005806.

Jie:2008:KBC

[JZHW08] Jing Jie, Jianchao Zeng, Chongzhao Han, and Qinghua Wang. Knowledgebased cooperative particle swarm optimization. Applied Mathematics and Computation, 205(2):861–873, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003652.

Ji:2007:DGO

[JZQ07] Ying Ji, Ke-Cun Zhang, and Shao-Jian Qu. A deterministic global optimization algorithm. Applied Mathematics and Computation, 185(1):382–387, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600871X.

Jian:2006:NSC

[JZTH06] Jin-Bao Jian, Hai-Yan Zheng, Chun-Ming Tang, and Qing-Jie Hu. A new superlinearly convergent norm-relaxed method of strongly sub-feasible direction for inequality constrained optimization. Applied Mathematics and Computation, 182(2):955– 976, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306004115.

Jiang:2008:NTS

[JZW08] Weijin Jiang, Lianmei Zhang, and Pu Wang. Nonlinear time series forecasting of time-delay neural network embedded with Bayesian regularization. Applied Mathematics and Computation, 205(1):123–132, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003238.

Jin:2007:TDEb

[JZX07a] Li Jin, Li-Wei Zhang, and Xian-Tao Xiao. Two differential equation systems for equality-constrained optimization. Applied Mathematics and Computation, 190(2):1030–1039, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015426.

Jin:2007:TDEa

[JZX07b] Li Jin, Li-Wei Zhang, and Xian-Tao Xiao. Two differential equation systems for inequality constrained optimization. Applied Mathematics and Computation, 188(2):1334–1343, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014871.

Zhou:2006:IAH

[jZzFbL06] Zhen jiang Zhou, Jing zhi Fu, and Zhi bin Li. An implementation for the algorithm of Hirota bilinear form of PDE in the Maple system. Applied Mathematics and Computation, 183(2):872–877, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006205.

Kadalbajoo:2005:FMB

[KA05a] Mohan K. Kadalbajoo and Vivek K. Aggarwal. Fitted mesh B-spline collocation method for solving self-adjoint singularly perturbed boundary value problems. Applied Mathematics and Computation, 161(3):973–987, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000475.

Kadalbajoo:2005:NSS

[KA05b] Mohan K. Kadalbajoo and Vivek K. Aggarwal. Numerical solution of singular boundary value problems via Chebyshev polynomial and B-spline. Applied Mathematics and Computation, 160(3):851–863, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303011986.

Kordrostami:2005:DFM

[KA05c] Sohrab Kordrostami and Alireza Amirteimoori. Undesirable factors in multicomponent performance measurement. Applied Mathematics and Computation, 171(2):721–729, December 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadalbajoo:2006:NMB

[KA06a] Mohan. K. Kadalbajoo and A. Awasthi. A numerical method based on Crank– Nicolson scheme for Burgers’ equation. Applied Mathematics and Computation, 182(2):1430–1442, November 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030600470X.

Kadalbajoo:2006:PUD

[KA06b] Mohan K. Kadalbajoo and A. Awasthi. A parameter uniform difference scheme for singularly perturbed parabolic problem in one space dimension. Applied Mathematics and Computation, 183(1):42–60, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600453X.

Kasap:2006:SCG

[KA06c] Emin Kasap and F. Talay Akyildiz. Surfaces with common geodesic in Minkowski 3-space. Applied Mathematics and Computation, 177(1):260– 270, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009082.

Kizilkan:2006:NVS

[KA06d] G. Çelik Kızılkan and K. Aydın. A new variable step size algorithm for Cauchy problem. Applied Mathematics and Computation, 183(2):878–884, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006217.

Kumar:2006:UMF

[KA06e] Manoj Kumar and Tariq Aziz. A uniform mesh finite difference method for a class of singular twopoint boundary value problems. Applied Mathematics and Computation, 180(1):173–177, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305011100.

Kutoglu:2006:RCP

[KA06f] Hakan S. Kutoglu and Tevfik Ayan. The role of common point distribution in obtaining reliable parameters for coordinate transformation. Applied Mathematics and Computation, 176(2):751–758, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008751.

Karaboga:2009:CSA

[KA09a] Dervis Karaboga and Bahriye Akay. A comparative study of Artificial Bee Colony algorithm. Applied Mathematics and Computation, 214(1):108–132, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002860.

Kudenatti:2009:SPG

[KA09b] Ramesh B. Kudenatti and Vishwanath B. Awati. Solution of pressure gradient stretching plate with suction. Applied Mathematics and Computation, 210(1):151–157, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009855.

Kadry:2007:ANL

[Kad07a] Seifedine Kadry. Analysis of a non-linear system by a new technique based on the Continuation method. Applied Mathematics and Computation, 187(2):1298–1307, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012689.

Kadry:2007:GPT

[Kad07b]  Seifedine Kadry.  On

the generalization of probabilistic transformation method. Applied Mathematics and Computation, 190(2):1284–1289, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001592.

Khan:2008:ESA

[KAF08] M. Khan, S. Hyder Ali, and C. Fetecau. Exact solutions of accelerated flows for a Burgers’ fluid. I. The case γ < λ^{2}/4. Applied Mathematics and Computation, 203(2):881–894, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004190.

Kahya:2005:CTF

[Kah05a] Emin Kahya. Comment on titled “On Fibonacci search method with k-Lucas numbers”. Applied Mathematics and Computation, 162(3):1321–1324, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001900.

See [YK03].

Kahya:2005:NUSb

[Kah05b] Emin Kahya. A new unidimensional search method for optimization: Linear interpolation method. Applied Mathematics and Computation, 171(2):912– 926, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kahya:2005:NUSa

[Kah05c] Emin Kahya. A new unidimensional search method for optimization: The 5/9 method. Applied Mathematics and Computation, 171(1):163–179, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000883.

Kahya:2006:CTI

[Kah06a] Emin Kahya. Comment on titled “An improvement on Fibonacci search method in optimization theory”. Applied Mathematics and Computation, 173(2):753–756, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003899.

See [SYY04].

Kahya:2006:MST

[Kah06b] Emin Kahya. Modified Secant-type methods for unconstrained optimization. Applied Mathematics and Computation, 181(2):1349–1356, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002530.

Kahya:2007:CEQ

[Kah07] Emin Kahya. A class of exponential quadratically convergent iterative formulae for unconstrained optimization. Applied Mathematics and Computation, 186(2):1010–1017, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010277.

Kasap:2005:ERS

[KAK05] Emin Kasap, İsmail Aydemir, and Nuri Kuruoğlu. Erratum to: “Ruled surfaces with timelike rulings” [Appl. Math. Comput. 147 (2004) 241– 248]. Applied Mathematics and Computation, 168(2):1461–1468, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006605.

See [AAABH04, KAK06, Kas06b].

Kasap:2006:ERSa

[KAK06] Emin Kasap, İsmail Aydemir, and Nuri Kuruoğlu. Erratum to: “Ruled surfaces with time like rulings” [Appl. Math. Comput. 147 (2004) 241– 253]. Applied Mathematics and Computation, 174(2):1660–1667, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005795.

See [AAABH04, KAK05, Kas06b].

Kutoglu:2006:IGN

[KAM06] Hakan S. Kutoglu, Tevfik Ayan, and Ç. Mekik. Integrating GPS with national networks by collocation method. Applied Mathematics and Computation, 177(2):508–514, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009318.

Kazezyilmaz-Alhan:2005:NMO

[KAMR05] Cevza Melek Kazezyılmaz-Alhan, Miguel A. Medina, Jr., and Prasada Rao. On numerical modeling of overland flow. Applied Mathematics and Computation, 166(3):724–740, July 26, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004680.

Kane:2006:SMS

[Kan06a] R. P. Kane. Some methods of spectral analysis: Comparison with MEM– MRA combination. Applied Mathematics and Computation, 181(2):949–957, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001755.

Kanwar:2006:FTO

[Kan06b] V. Kanwar. A family of third-order multipoint methods for solving nonlinear equations. Applied Mathematics and Computation, 176(2):409–413, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008167.

Kanth:2007:CSP

[Kan07] A. S. V. Ravi Kanth. Cubic spline polynomial for non-linear singular twopoint boundary value problems. Applied Mathematics and Computation, 189(2):2017–2022, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000215.

Kanas:2009:NPS

[Kan09a] Stanisława Kanas. Norm of pre-Schwarzian derivative for the class of k-uniformly convex and k-starlike functions. Applied Mathematics and Computation, 215(6):2275–2282, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007413.

Kang:2009:PSW

[Kan09b] Dongsheng Kang. Positive solutions to the weighted critical quasilinear problems. Applied Mathematics and Computation, 213(2):432–439, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002549.

Kasap:2008:GSF

[KAO08] Emin Kasap, F. Talay Akyildiz, and Keziban Orbay. A generalization of surfaces family with common spatial geodesic. Applied Mathematics and Computation, 201(1–2):781–789, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000465.

Khan:2006:SCB

[KAP06] Subuhi Khan, Babita Agrawal, and M. A. Pathan. Some connections between generalized Voigt functions with the different parameters. Applied Mathematics and Computation, 181(1):57–64, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600066X.

Khan:2007:ECE

[KAPM07] Subuhi Khan, Babita Agrawal, M. A. Pathan, and Fahim Mohammad. Evaluations of certain Euler type integrals. Applied Mathematics and Computation, 189(2):1993–2003, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000185.

Karaa:2005:ALS

[Kar05a] Samir Karaa. An accurate LOD scheme for two-dimensional parabolic problems. Applied Mathematics and Computation, 170(2):886–894, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000305.

Karaduman:2005:DMG

[Kar05b] Erdal Karaduman. On determinants of matrices with general Fibonacci numbers entries. Applied Mathematics and Computation, 167(1):670–676, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005223.

Karawia:2006:CAS

[Kar06] A. A. Karawia. A computational algorithm for solving periodic penta-diagonal linear systems. Applied Mathematics and Computation, 174(1):613–618, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004881.

Karawia:2007:TASb

[Kar07a] A. A. Karawia. Two algorithms for solving a general backward tridiagonal linear systems. Applied Mathematics and Computation, 194(2):534–539, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005061.

Karawia:2007:TASa

[Kar07b] A. A. Karawia. Two algorithms for solving comrade linear systems. Applied Mathematics and Computation, 189(1):291–297, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601602X.

Karakostas:2009:SLN

[Kar09a] George L. Karakostas. Solvability of the Φ-Laplacian with nonlocal boundary conditions. Applied Mathematics and Computation, 215(2):514– 523, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309005049.

Karoui:2009:WBG

[Kar09b] Abderrazek Karoui. Wavelet bases with a general integer dilation factor d ≥ 2 and better regularity properties. Applied Mathematics and Computation, 214(2):557–568, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003543.

Karpuz:2009:ABB

[Kar09c] Başak Karpuz. Asymptotic behaviour of bounded solutions of a class of higherorder neutral dynamic equations. Applied Mathematics and Computation, 215(6):2174–2183, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007267.

Kamal:2006:NSS

[KAS06a] M. Anwar Kamal, Muhammad Ashraf, and K. S. Syed. Numerical solution of steady viscous flow of a micropolar fluid driven by injection between two porous disks. Applied Mathematics and Computation, 179(1):1–10, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010015.

Kasap:2006:ERSb

[Kas06b] Emin Kasap. Erratum to: “Ruled surfaces with timelike rulings” [Appl. Math. Comput. 147 (2004) 241–253]. Applied Mathematics and Computation, 179(1):402–405, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010118.

See [AAABH04, KAK05, KAK06].

Kassem:2006:SSD

[Kas06c] Mohamed Abd El-Hady Kassem. Symmetric and self duality in vector optimization problem. Applied Mathematics and Computation, 183(2):1121–1126, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006461.

Kassem:2007:ISC

[Kas07] Mohamed Abd El-Hady Kassem. Interactive stability cutting-plane algorithm for multiobjective nonlinear programming problems. Applied Mathematics and Computation, 192(2):446–456, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003487.

Kassem:2009:HOS

[Kas09] Mohamed Abd El-Hady Kassem. Higher-order symmetric duality in vector optimization problem involving generalized coneinvex functions. Applied Mathematics and Computation, 209(2):405–409, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009788.

Katos:2005:RTB

[Kat05] Vasilios Katos. A randomness test for block ciphers. Applied Mathematics and Computation, 162(1):29– 35, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000530.

Kilic:2006:CGO

[KAT06] E. Kilic, B. Altunkaynak, and D. Tasci. On the computing of the generalized order-k Pell numbers in log time. Applied Mathematics and Computation, 181(1):511–515, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001202.

Katsikis:2007:CMP

[Kat07] Vasilios N. Katsikis. Computational methods in portfolio insurance. Applied Mathematics and Computation, 189(1):9–22, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015724.

Katsikis:2008:CML

[Kat08] V. N. Katsikis. Computational methods in latticesubspaces of C[a, b] with applications in portfolio insurance. Applied Mathematics and Computation, 200(1):204–219, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011009.

Katsikis:2009:MBR

[Kat09] Vasilios N. Katsikis. A Matlab-based rapid method for computing latticesubspaces and vector sublattices of R^{\}: Applications in portfolio insurance. Applied Mathematics and Computation, 215(3):961–972, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005852.

Kajani:2005:IMS

[KAV05] M. Tavassoli Kajani, B. Asady, and A. Hadi Vencheh. An iterative method for solving dual fuzzy nonlinear equations. Applied Mathematics and Computation, 167(1):316–323, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004941.

Kaya:2005:EUD

[Kay05a] Dogan Kaya. An experience using different synchronisation mechanisms on a shared memory multiprocessors. Applied Mathematics and Computation, 161(3):1027–1036, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000505.

Kaya:2005:IAG

[Kay05b] Doğan Kaya. An implementation of the ADM for generalized one-dimensional Klein–Gordon equation. Applied Mathematics and Computation, 166(2):426– 433, July 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304004461.

Kaya:2005:ERE

[Kay05c] Doğan Kaya. On experimental results and explicit exact solutions for the generalized Boussinesq type equation. Applied Mathematics and Computation, 165(2):303–311, June 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003832.

Kaya:2005:PARb

[Kay05d]  Doğan Kaya.  Parallel

algorithms for reduction of a symmetric matrix to tridiagonal form on a shared memory multiprocessor. Applied Mathematics and Computation, 169(2):1045–1062, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008355.

Kaya:2005:STE

[Kay05e] Doğan Kaya. The symmetric tridiagonal eigenproblem on a shared memory multiprocessor: Part II. Applied Mathematics and Computation, 163(1):213– 244, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030400222X.

Kazmi:2005:CSI

[KB05a] K. R. Kazmi and M. I. Bhat. Convergence and stability of iterative algorithms of generalized setvalued variational-like inclusions in Banach spaces. Applied Mathematics and Computation, 166(1):164– 180, July 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400428X.

Kirsche:2005:RAI

[KB05b] A. Kirsche and C. Böckmann.

Rational approximations for ill-conditioned equation systems. Applied Mathematics and Computation, 171(1):385–397, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001037.

Kucuk:2005:ACF

[KB05c] Ahmet Küçük and Mustafa Bayram. Automatic calculation of the fundamental group of an oriented surface of genus n with k boundary surfaces. Applied Mathematics and Computation, 162(1):1–6, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000414.

Kanth:2006:CSC

[KB06a] A. S. V. Ravi Kanth and Vishnu Bhattacharya. Cubic spline for a class of non-linear singular boundary value problems arising in physiology. Applied Mathematics and Computation, 174(1):768–774, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005163.

Kaslik:2006:CSS

[KB06b] E. Kaslik and St. Balint. Configurations of steady states for Hopfield-type neural networks. Applied Mathematics and Computation, 182(1):934–946, November 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306004097.

Kirsche:2006:PIM

[KB06c] A. Kirsche and C. Böckmann. Padé iteration method for regularization. Applied Mathematics and Computation, 180(2):648–663, September 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306000543.

Kumar:2007:OCL

[KB07] A. Vincent Antony Kumar and P. Balasubramaniam. Optimal control for linear singular system using genetic programming.

Applied Mathematics and Computation, 192(1):78– 89, September 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307002962.

Kammanee:2008:DPK

[KB08] A. Kammanee and C. Böckmann.

Determination of partially known Sturm–Liouville potentials. Applied Mathematics and Computation, 204(2):928–937, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005808.

Kammanee:2009:BVM

[KB09] A. Kammanee and C. Böckmann.

Boundary value method for inverse Sturm–Liouville problems. Applied Mathematics and Computation, 214(2):342–352, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003191.

Kaneko:2006:DGF

[KBH06a] Hideaki Kaneko, Kim S.

Bey, and Gene J. W. Hou. Discontinuous Galerkin finite element method for parabolic problems. Applied Mathematics and Computation, 182(1):388– 402, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306003468.

Kaneko:2006:DGM

[KBH06b] Hideaki Kaneko, Kim S. Bey, and Gene J. W. Hou. A discontinuous Galerkin Method for parabolic problems with modified hp-finite element approximation technique. Applied Mathematics and Computation, 182(2):1405–1417, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004589.

Kumaran:2009:MFP

[KBKV09] V. Kumaran, A. K. Banerjee, A. Vanav Kumar, and K. Vajravelu. MHD flow past a stretching permeable sheet. Applied Mathematics and Computation, 210(1):26–32, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007789.

Kaneko:2008:NEU

[KBLT08] Hideaki Kaneko, Kim S.

Bey, Yongwimon Lenbury, and Puntip Toghaw. Numerical experiments using hierarchical finite element method for nonlinear heat conduction in plates. Applied Mathematics and Computation, 201(1–2):414–430, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307012180.

Kohl:2008:PGO

[KBMZ08] Russell Kohl, Anjan Biswas, Daniela Milovic, and Essaid Zerrad. Perturbation of Gaussian optical solitons in dispersionmanaged fibers. Applied Mathematics and Computation, 199(1):250–258, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009927.

Kirmaci:2007:HTI

[KBÖP07] U. S. Kirmaci, M. Klaričić Bakula, M. E. Özdemir, and J. Pečarić. Hadamardtype inequalities for sconvex functions. Applied Mathematics and Computation, 193(1):26–35, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307003633.

Koudil:2007:UAB

[KBTS07] Mouloud Koudil, Karima Benatchba, Amina Tarabet, and El Batoul Sahraoui. Using artificial bees to solve partitioning and scheduling problems in codesign. Applied Mathematics and Computation, 186(2):1710–1722, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011064.

Kadilar:2005:NEU

[KC05a] Cem Kadilar and Hulya Cingi. A new estimator using two auxiliary variables. Applied Mathematics and Computation, 162(2):901–908, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001377.

Kwatny:2005:SCN

[KC05b] Harry G. Kwatny and Bor-Chin Chang. Symbolic computing of nonlinear observable and observer forms. Applied Mathematics and Computation, 171(2):1058–1080, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadilar:2006:REP

[KC06a] Cem Kadilar and Hulya Cingi. Ratio estimators for the population variance in simple and stratified random sampling. Applied Mathematics and Computation, 173(2):1047–1059, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004108.

Ke:2006:NSA

[KC06b] Jau-Chuan Ke and Yunn-Kuang Chu. Nonparametric and simulated analysis of intensity for a queueing system. Applied Mathematics and Computation, 183(2):1280–1291, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006758.

Kahya:2007:MSM

[KC07a] Emin Kahya and Jinhai Chen. A modified Secant method for unconstrained optimization. Applied Mathematics and Computation, 186(2):1000–1004, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010253.

Ke:2007:CAA

[KC07b] Jau-Chuan Ke and Yunn-Kuang Chu. Comparative analysis of availability for a redundant repairable system. Applied Mathematics and Computation, 188(1):332–338, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306013567.

Kumari:2007:SCS

[KC07c] Anchala Kumari and Soubhik Chakraborty. Software complexity: a statistical case study through insertion sort. Applied Mathematics and Computation, 190(1):40–50, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000264.

Kung:2009:EEA

[KC09] James J. Kung and Andrew P. Carverhill. An efficient ex-ante criterion for ranking investment strategies. Applied Mathematics and Computation, 210(1):258–268, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000034.

Kim:2008:TRG

[KCK08] Chihurn Kim, Geon Ho Choe, and Dong Han Kim. Tests of randomness by the gambler’s ruin algorithm. Applied Mathematics and Computation, 199(1):195–210, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009873.

See critical remarks [PJ10].

Ko:2007:OAB

[KCW07] Chia-Nan Ko, Ying-Pin Chang, and Chia-Ju Wu. An orthogonal-array-based particle swarm optimizer with nonlinear time-varying evolution. Applied Mathematics and Computation, 191(1):272–279, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002548.

Kudryashov:2009:TWS

[KD09] Nikolai A. Kudryashov and Maria V. Demina. Traveling wave solutions of the generalized nonlinear evolution equations. Applied Mathematics and Computation, 210(2):551–557, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008825.

Khalifa:2005:NSK

[KE05a] M. E. Khalifa and Mahmoud Elgamal. A numerical solution to Klein– Gordon equation with Dirichlet boundary condition. Applied Mathematics and Computation, 160(2):451–475, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303011731.

Kutluay:2005:FEA

[KE05b] S. Kutluay and A. Esen. Finite element approaches to the PTC thermistor problem. Applied Mathematics and Computation, 163(1):147–162, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002164.

Kutluay:2005:FES

[KE05c] S. Kutluay and A. Esen. Finite element solution of the thermistor problem with a ramp electrical conductivity. Applied Mathematics and Computation, 161(3):897–913, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000396.

Kutluay:2005:NST

[KE05d] S. Kutluay and A. Esen.

Numerical solutions of the thermistor problem by spline finite elements. Applied Mathematics and Computation, 162(1):475– 489, March 4, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304001109.

Kilic:2008:CAS

[KEM08] Emrah Kilic and Moawwad El-Mikkawy. A computational algorithm for special n th-order pentadiagonal Toeplitz determinants. Applied Mathematics and Computation, 199(2):820– 822, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010405.

Kaya:2005:NID

[KES05] Doğan Kaya and Salah M. El-Sayed. A numerical implementation of the decomposition method for the Lienard equation. Applied Mathematics and Computation, 171(2):1095–1103, December 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khosravifard:2009:ELA

[KES09] Mohammadali Khosravifard, Morteza Esmaeili, and Hossein Saidi. Extension of the Lasserre– Avrachenkov theorem on the integral of multilinear forms over simplices. Applied Mathematics and Computation, 212(1):94– 99, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001222.

Kouche:2008:PEP

[KeTL08] Mahiéddine Kouche, Nasser eddine Tatar, and Shengqiang Liu. Permanence and existence of a positive periodic solution to a periodic stagestructured system with infinite delay. Applied Mathematics and Computation, 202(2):620–638, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001513.

Keyanpour:2007:WDM

[KF07] M. Keyanpour and M. H.

Farahi. Wing drag minimization. Applied Mathematics and Computation, 186(1):685–692, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009805.

Kanan:2008:GBO

[KF08a] Hamidreza Rashidy Kanan and Karim Faez. GA-based optimal selection of PZMI features for face recognition. Applied Mathematics and Computation, 205(2):706–715, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003494.

Kanan:2008:IFS

[KF08b] Hamidreza Rashidy Kanan and Karim Faez. An improved feature selection method based on ant colony optimization (ACO) evaluated on face recognition system. Applied Mathematics and Computation, 205(2):716–725, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003500.

Kwon:2008:VAO

[KF08c] Changhyun Kwon and Terry L. Friesz. Valuation of American options by the gradient projection method. Applied Mathematics and Computation, 206(1):380–388, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006875.

Kano:2008:EDB

[KFM08] Hiroyuki Kano, Hiroyuki Fujioka, and Clyde F. Martin. Extrema detection of bivariate spline functions. Applied Mathematics and Computation, 200(1):58– 69, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010855.

Khorram:2006:LOF

[KG06] E. Khorram and A. Ghodousian. Linear objective function optimization with fuzzy relation equation constraints regarding max–av composition. Applied Mathematics and Computation, 173(2):872–886, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003991.

See note [Zim07].

Karsli:2008:SAP

[KG08] Harun Karsli and Vijay Gupta. Some approximation properties of q-Chlodowsky operators. Applied Mathematics and Computation, 195(1):220–229, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005322.

Kajani:2006:NSL

[KGB06] M. Tavassoli Kajani, M. Ghasemi, and E. Babolian. Numerical solution of linear integro-differential equation by using sine–cosine wavelets. Applied Mathematics and Computation, 180(2):569–574, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306000452.

Khorram:2006:SLO

[KGM06] Esmaile Khorram, Amin Ghodousian, and Ali Abbasi Molai. Solving linear optimization problems with max-star composition equation constraints. Applied Mathematics and Computation, 179(2):654–661, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010702.

Kao:2005:GBH

[KH05a] Shin-Shin Kao and Lih-Hsing Hsu. The globally bi-3* and hyper bi-3* connectedness of the spider web networks. Applied Mathematics and Computation, 170(1):597–610, November 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304009713.

Kao:2005:SWN

[KH05b] Shin-Shin Kao and Lih-Hsing Hsu. Spider web networks: a family of optimal, fault tolerant, Hamiltonian bipartite graphs. Applied Mathematics and Computation, 160(1):269–282, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303009020.

Keeling:2007:GMH

[KH07a] Stephen L. Keeling and Gundolf Haase. Geometric multigrid for high-order regularizations of early vision problems. Applied Mathematics and Computation, 184(2):536–556, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007296.

Kunze:2007:CCM

[KH07b] H. Kunze and K. Heidler.

The collage coding method and its application to an inverse problem for the Lorenz system. Applied Mathematics and Computation, 186(1):124–129, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009234.

Khan:2008:GRT

[KH08] Subuhi Khan and Nader Ali Makboul Hassan. Generating relations of Tricomi and Hermite–Tricomi functions using Lie algebra representation. Applied Mathematics and Computation, 202(1):86–101, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800057X.

Kuo:2009:SSC

[KH09] Nan-Hong Kuo and C. D. Hu. A study of the solutions of the combined sine–cosine-Gordon equation. Applied Mathematics and Computation, 215(3):1015–1019, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030900592X.

Khan:2007:ARC

[Kha07a] Rahmat Ali Khan. Approximations and rapid convergence of solutions of nonlinear three point boundary value problems. Applied Mathematics and Computation, 186(2):957–968, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010216.

Khan:2007:GAR

[Kha07b] Rahmat Ali Khan. Generalized approximations and rapid convergence of solutions of m-point boundary value problems. Applied Mathematics and Computation, 188(2):1878–1890, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015566.

Khan:2008:GAN

[Kha08a] Rahmat Ali Khan. Generalized approximations for nonlinear three-point boundary value problems. Applied Mathematics and Computation, 197(1):111–120, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700759X.

Khan:2008:PSF

[Kha08b] Rahmat Ali Khan. Positive solutions of fourpoint singular boundary value problems. Applied Mathematics and Computation, 201(1–2):762–773, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000441.

Khan:2009:GAM

[Kha09] Rahmat Ali Khan. Generalized approximation method for heat radiation equations. Applied Mathematics and Computation, 212(2):287–295, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900143X.

Kwon:2005:CLK

[KHKL05] Jeoung Ok Kwon, Jung Yeon Hwang, Changwook Kim, and Dong Hoon Lee. Cryptanalysis of Lee–Kim– Yoo password-based key agreement scheme. Applied Mathematics and Computation, 168(2):858–865, September 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304006447.

Ke:2006:PPA

[KHL06] Jau-Chuan Ke, Hsin-I Huang, and Chuen-Horng Lin. Parametric programming approach for batch arrival queues with vacation policies and fuzzy parameters. Applied Mathematics and Computation, 180(1):217–232, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305011161.

Khodabakhshi:2007:SEM

[Kho07]  M. Khodabakhshi.  A

super-efficiency model based on improved outputs in data envelopment analysis. Applied Mathematics and Computation, 184(2):695–703, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007454.

Konghua:2007:NIM

[KHZ07] Guo Konghua, Xiyan Hu, and Lei Zhang. A new iteration method for the matrix equation AX = B. Applied Mathematics and Computation, 187(2):1434–1441, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601280X.

Kaya:2005:CAA

[KI05a] Doğan Kaya and Ibrahim E.

Inan. A convergence analysis of the ADM and an application. Applied Mathematics and Computation, 161(3):1015–1025, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304000499.

Kaya:2005:NAD

[KI05b] Doğan Kaya and Ibrahim E. Inan. A numerical application of the decomposition method for the combined KdV–MKdV equation. Applied Mathematics and Computation, 168(2):915–926, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304006496.

Kitamura:2008:CCB

[KI08] Masahito Kitamura and Takehiro Inohara. A characterization of completeness of blockability relations with respect to unanimity. Applied Mathematics and Computation, 197(2):715–718, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008296.

Kianfar:2005:NMA

[Kia05] Farhad Kianfar. A numerical method to approximate optimal production and maintenance plan in a flexible manufacturing system. Applied Mathematics and Computation, 170(2):924–940, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000342.

Kilic:2008:EFI

[Kıl08a]  Emrah Kılıç.  Explicit

formula for the inverse of a tridiagonal matrix by backward continued fractions. Applied Mathematics and Computation, 197(1):345–357, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007825.

Kilic:2008:CDM

[Kil08b]  Emrah Kilic.  On a

constant-diagonals matrix. Applied Mathematics and Computation, 204(1):184–190, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004931.

Kim:2005:EOG

[Kim05a] Joocheol Kim. Estimation of optimality gap using stratified sampling. Applied Mathematics and Computation, 171(2):710– 720, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kim:2005:DIM

[Kim05b] Junseok Kim. A diffuseinterface model for axisymmetric immiscible twophase flow. Applied Mathematics and Computation, 160(2):589–606, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303011834.

Kim:2006:LTS

[Kim06a] Mi-Young Kim. Long-time stability of numerical solutions to Gurtin–MacCamy equations by method of characteristics. Applied Mathematics and Computation, 176(2):552–562, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008556.

Kim:2006:SDG

[Kim06b] Mi-Young Kim. Stabilized discontinuous Galerkin methods for scalar linear hyperbolic equations. Applied Mathematics and Computation, 182(2):1623– 1628, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306005091.

Kim:2007:LRN

[Kim07a] Taekyun Kim. Lebesgue– Radon–Nikodym theorem with respect to q-Volkenborn distribution on μ_{q}. Applied Mathematics and Computation, 187(1):266–271, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601160X.

Kim:2007:MPD

[Kim07b] Yong Chan Kim. Mapping properties of differential inequalities related to univalent functions. Applied Mathematics and Computation, 187(1):272–279, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011611.

Kim:2008:SPT

[Kim08] Gwang Hui Kim. On the stability of the pexiderized trigonometric functional equation. Applied Mathematics and Computation, 203(1):99–105, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002117.

Kim:2009:NGT

[Kim09]  Yong Chan Kim.  A

note on growth theorem of Bazilevič functions. Applied Mathematics and Computation, 208(2):542–546, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308009272.

Kincanon:2005:MCR

[Kin05] Eric Kincanon. A method to construct reflectionless potentials. Applied Mathematics and Computation, 165(3):565–569, June 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004023.

Kincanon:2006:EPG

[Kin06]  Eric Kincanon.  Er-

ror propagation in the Gel’Fand–Levitan equation. Applied Mathematics and Computation, 182(2):1639–1641, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600511X.

Kiss:2008:DRD

[Kis08] Krisztina Kiss. n-Dimensional ratio-dependent predator– prey systems with diffusion. Applied Mathematics and Computation, 205(1):325–335, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005912.

Kiymaz:2006:SAE

[Kıy06] Onur Kıymaz. A symbolic algorithm for exact power series solutions of n th order linear homogeneous differential equations with polynomial coefficients near an ordinary point. Applied Mathematics and Computation, 183(2):1052–1056, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006400.

Khosravi:2008:NLS

[KJ08] Alireza Khosravi and Aliakbar Jalali. A new LMI solution in the L_{1} optimal problem for wind turbine-induction generator unit. Applied Mathematics and Computation, 206(2):643–650, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003871.

Kulkarni:2009:RMT

[KJ09] D. B. Kulkarni and M. M. Joglekar. Residue minimization technique to analyze the efficiency of convective straight fins having temperature-dependent thermal conductivity. Applied Mathematics and Computation, 215(6):2184– 2191, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309007279.

Kim:2005:IYA

[KJY05] Kee-Won Kim, Jun-Cheol Jeon, and Kee-Young Yoo. An improvement on Yang et al.’s password authentication schemes. Applied Mathematics and Computation, 170(1):207–215, November 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030400918X.

Kang:2005:BSM

[KK05a] Tong Kang and Kwang Ik Kim. An E-based splitting method with a nonmatching grid for eddy current equations with discontinuous coefficients. Applied Mathematics and Computation, 170(1):462– 484, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304009622.

Katti:2005:EPA

[KK05b] C. P. Katti and Rama Kumari. Efficient parallel algorithm for quasi pentadiagonal systems on a hypercube. Applied Mathematics and Computation, 163(3):1215–1223, April 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002905.

Kumar:2005:DPS

[KK05c] B. V. Rathish Kumar and Bipin Kumar. On the development of parallel sparse CGM solver for CFD computations on Anu-cluster. Applied Mathematics and Computation, 160(3):739–761, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303011925.

Kumar:2005:PPH

[KK05d] B. V. Rathish Kumar and Bipin Kumar. Parallel preconditioners for heat transfer applications on ANU-cluster. Applied Mathematics and Computation, 163(3):1243–1263, April 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002929.

Kadalbajoo:2006:HRT

[KK06] M. K. Kadalbajoo and Ritesh Kumar. A high resolution total variation diminishing scheme for hyperbolic conservation law and related problems. Applied Mathematics and Computation, 175(2):1556–1573, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305007149.

Kadalbajoo:2007:GMF

[KK07a] Mohan K. Kadalbajoo and Devendra Kumar. Geometric mesh FDM for self-adjoint singular perturbation boundary value problems. Applied Mathematics and Computation, 190(2):1646–1656, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002093.

Kadalbajoo:2007:BSM

[KK07b] Mohan K. Kadalbajoo and Vivek Kumar. B-spline method for a class of singular two-point boundary value problems using optimal grid. Applied Mathematics and Computation, 188(2):1856–1869, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015542.

Kanaun:2007:NMS

[KK07c] S. Kanaun and S. Babaii Kochekseraii. A numerical method for the solution of 3D-integral equations of electro-static theory based on Gaussian approximating functions. Applied Mathematics and Computation, 184(2):754–768, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007557.

Kazmi:2007:APA

[KK07d] K. R. Kazmi and F. A.

Khan. Auxiliary problems and algorithm for a system of generalized variational-like inequality problems. Applied Mathematics and Computation, 187(2):789–796, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012124.

Khots:2007:OMM

[KK07e] Boris Khots and Dmitriy Khots. Observer’s mathematics — mathematics of relativity. Applied Mathematics and Computation, 187(1):228–238, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601157X.

Kadalbajoo:2008:FMB

[KK08a] M. K. Kadalbajoo and Devendra Kumar. Fitted mesh B-spline collocation method for singularly perturbed differential–difference equations with small delay. Applied Mathematics and Computation, 204(1):90–98, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004591.

Kadalbajoo:2008:NLS

[KK08b] Mohan K. Kadalbajoo and Devendra Kumar. A non-linear single step explicit scheme for nonlinear two-point singularly perturbed boundary value problems via initial value technique. Applied Mathematics and Computation, 202(2):738–746, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001720.

Kiss:2008:QBD

[KK08c] Krisztina Kiss and Sándor Kovács. Qualitative behavior of n-dimensional ratio-dependent predator– prey systems. Applied Mathematics and Computation, 199(2):535–546, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010296.

Kumar:2008:CHR

[KK08d] Ritesh Kumar and M. K.

Kadalbajoo. A class of high resolution shock capturing schemes for hyperbolic conservation laws. Applied Mathematics and Computation, 195(1):110–126, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005243.

Khan:2005:SPS

[KKA05] Arshad Khan, Islam Khan, and Tariq Aziz. A survey on parametric spline function approximation. Applied Mathematics and Computation, 171(2):983– 1003, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khan:2006:SSS

[KKA06] Arshad Khan, Islam Khan, and Tariq Aziz. Sextic spline solution of a singularly perturbed boundaryvalue problems. Applied Mathematics and Computation, 181(1):432–439, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001135.

Kazmi:2009:EIA

[KKA09a] K. R. Kazmi, Huzoor H. Khan, and Naeem Ahmad. Existence and iterative approximation of solutions of a system of general variational inclusions. Applied Mathematics and Computation, 215(1):110–117, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003890.

Khosravianarab:2009:SCA

[KKA09b] Sh. Khosravianarab, S. R.

Kulkarni, and O. P. Ahuja. Some classes of analytic functions involving differential subordinations. Applied Mathematics and Computation, 215(1):221–226, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004482.

Kamyad:2007:NAS

[KKF07] A. V. Kamyad, M. Keyanpour, and M. H. Farahi. A new approach for solving of optimal nonlinear control problems. Applied Mathematics and Computation, 187(2):1461–1471, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012847.

Kazmi:2009:SAS

[KKKM09] K. R. Kazmi, B. S. Komal, F. A. Khan, and Pankaj Mansotra. Sensitivity analysis for a system of parametric general quasi-variational-like inequality problems in uniformly smooth Banach spaces. Applied Mathematics and Computation, 215(2):716–726, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005529.

Koch:2006:ATD

[KKS06] Othmar Koch, Wolfgang Kreuzer, and Armin Scrinzi. Approximation of the time-dependent electronic Schrödinger equation by MCTDHF. Applied Mathematics and Computation, 173(2):960–976, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004054.

Kaebi:2005:AIP

[KKT05] Amer Kaebi, Asghar Kerayechian, and Faezeh Toutounian. Approximate inverse preconditioner by computing approximate solution of Sylvester equation. Applied Mathematics and Computation, 170(2):1067–1076, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000470.

Kaabi:2007:NVS

[KKT07] A. Kaabi, A. Kerayechian, and F. Toutounian. A new version of successive approximations method for solving Sylvester matrix equations. Applied Mathematics and Computation, 186(1):638–645, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009751.

Kebbiche:2007:EPI

[KKY07a] Z. Kebbiche, A. Keraghel, and A. Yassine. Extension of a projective interior point method for linearly constrained convex programming. Applied Mathematics and Computation, 193(2):553–559, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004067.

Kim:2007:GCC

[KKY07b] Changho Kim, Sang Dong Kim, and Jungho Yoon. Generalized Clenshaw– Curtis quadrature rule with application to a collocation least-squares method. Applied Mathematics and Computation, 190(1):781–789, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307001129.

Ke:2005:PSP

[KL05a] Hua Ke and Baoding Liu. Project scheduling problem with stochastic activity duration times. Applied Mathematics and Computation, 168(1):342–353, September 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304006083.

Kirane:2005:NGS

[KL05b] M. Kirane and Y. Laskri. Nonexistence of global solutions to a hyperbolic equation with a space– time fractional damping. Applied Mathematics and Computation, 167(2):1304–1310, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005685.

Ke:2006:FAQ

[KL06a] Jau-Chuan Ke and Chuen-Horng Lin. Fuzzy analysis of queueing systems with an unreliable server: a nonlinear programming approach. Applied Mathematics and Computation, 175(1):330– 346, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006119.

Ke:2006:MES

[KL06b] Jau-Chuan Ke and Chuen-Horng Lin. Maximum entropy solutions for batch arrival queue with an unreliable server and delaying vacations. Applied Mathematics and Computation, 183(2):1328–1340, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006801.

Kou:2007:FMS

[KL07a] Jisheng Kou and Yitian Li. A family of modified super-Halley methods with fourth-order convergence. Applied Mathematics and Computation, 189(1):366– 370, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016110.

Kou:2007:FNN

[KL07b] Jisheng Kou and Yitian Li. A family of new Newton-like methods. Applied Mathematics and Computation, 192(1):162–167, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307003050.

Kou:2007:IJM

[KL07c] Jisheng Kou and Yitian Li. An improvement of the Jarratt method. Applied Mathematics and Computation, 189(2):1816–1821, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017772.

Kou:2007:ICH

[KL07d] Jisheng Kou and Yitian Li. The improvements of Chebyshev–Halley methods with fifth-order convergence. Applied Mathematics and Computation, 188 (1):143–147, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0096300306013300.

Kou:2007:MCH

[KL07e] Jisheng Kou and Yitian Li. Modified Chebyshev– Halley methods with sixthorder convergence. Applied Mathematics and Computation, 188(1):681–685, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014068.

Kou:2007:MCM

[KL07f] Jisheng Kou and Yitian Li. Modified Chebyshev’s method free from second derivative for nonlinear equations. Applied Mathematics and Computation, 187(2):1027–1032, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012392.

Kou:2007:SVC

[KL07g] Jisheng Kou and Yitian Li. Some variants of Chebyshev–Halley methods with fifth-order convergence. Applied Mathematics and Computation, 189(1):49–54, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015773.

Kou:2007:ULS

[KL07h] Jisheng Kou and Yitian Li. A uniparametric LU-SGS method for systems of nonlinear equations. Applied Mathematics and Computation, 189(1):235–240, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015979.

Kudryashov:2008:ESE

[KL08] Nikolai A. Kudryashov and Nadejda B. Loguinova. Extended simplest equation method for nonlinear differential equations. Applied Mathematics and Computation, 205(1):396–402, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005997.

Klamka:2008:SCM

[Kla08] Jerzy Klamka. Stochastic controllability and minimum energy control of systems with multiple delays in control. Applied Mathematics and Computation, 206(2):704–715, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006413.

Kueng:2008:FTH

[KLL^{+}08] Tz-Liang Kueng, Cheng-Kuan Lin, Tyne Liang, Jimmy J. M. Tan, and Lih-Hsing Hsu. Fault-tolerant Hamiltonian connectedness of cycle composition networks. Applied Mathematics and Computation, 196(1):245–256, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006649.

Krejic:2007:NLM

[KLR07] Nataša Krejić, Zorana Lužanin, and Ivana Radeka. Newton-like method for nonlinear banded block diagonal system. Applied Mathematics and Computation, 189(2):1705–1711, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017620.

Krejic:2009:GNB

[KLS09] Nataša Krejić, Zorana Lužanin, and Irena Stojkovska. Gauss–Newtonbased BFGS method with filter for unconstrained minimization. Applied Mathematics and Computation, 211(2):354–362, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000824.

Kou:2006:ECN

[KLW06a] Jisheng Kou, Yitian Li, and Xiuhua Wang. Efficient continuation Newtonlike method for solving systems of non-linear equations. Applied Mathematics and Computation, 174(2):846–853, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005138.

Kou:2006:MNMb

[KLW06b] Jisheng Kou, Yitian Li, and Xiuhua Wang. A modification of Newton method with third-order convergence. Applied Mathematics and Computation, 181(2):1106–1111, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001949.

Kou:2006:MHM

[KLW06c] Jisheng Kou, Yitian Li, and Xiuhua Wang. Modified Halley’s method free from second derivative. Applied Mathematics and Computation, 183(1):704– 708, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306006047.

Kou:2006:FSD

[KLW06d] Jisheng Kou, Yitian Li, and Xiuhua Wang. On a family of second-derivative-free variants of Chebyshev’s method. Applied Mathematics and Computation, 181(2):982–987, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001792.

Kou:2006:MNMa

[KLW06e] Jisheng Kou, Yitian Li, and Xiuhua Wang. On modified Newton methods with cubic convergence. Applied Mathematics and Computation, 176(1):123– 127, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007770.

Kou:2007:CFO

[KLW07a] Jisheng Kou, Yitian Li, and Xiuhua Wang. A composite fourth-order iterative method for solving non-linear equations. Applied Mathematics and Computation, 184(2):471– 475, January 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306007223.

Kou:2007:FFOa

[KLW07b] Jisheng Kou, Yitian Li, and Xiuhua Wang. A family of fifth-order iterations composed of Newton and third-order methods. Applied Mathematics and Computation, 186(2):1258–1262, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010514.

Kou:2007:FFOb

[KLW07c] Jisheng Kou, Yitian Li, and Xiuhua Wang. A family of fourth-order methods for solving non-linear equations. Applied Mathematics and Computation, 188(1):1031–1036, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014512.

Kou:2007:FOI

[KLW07d] Jisheng Kou, Yitian Li, and Xiuhua Wang. Fourthorder iterative methods free from second derivative. Applied Mathematics and Computation, 184(2):880–885, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007764.

Kou:2007:VSH

[KLW07e] Jisheng Kou, Yitian Li, and Xiuhua Wang. A variant of super-Halley method with accelerated fourthorder convergence. Applied Mathematics and Computation, 186(1):535–539, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009647.

Kong:2006:NSI

[KLX06] Linghua Kong, Ruxun Liu, and Zhenli Xu. Numerical simulation of interaction between Schrödinger field and Klein–Gordon field by multisymplectic method. Applied Mathematics and Computation, 181(1):342–350, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001020.

Ko:2007:SMI

[KLY07a] Kwan Pyo Ko, Byung-Gook Lee, and Gang Joon Yoon. A study on the mask of interpolatory symmetric subdivision schemes. Applied Mathematics and Computation, 187(2):609–621, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011167.

Ko:2007:TPA

[KLY07b] Kwan Pyo Ko, Byung-Gook Lee, and Gang Joon Yoon. A ternary 4point approximating subdivision scheme. Applied Mathematics and Computation, 190(2):1563–1573, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001932.

Kong:2007:SSM

[KLZ07] Linghua Kong, Ruxun Liu, and Xiaohong Zheng. A survey on symplectic and multi-symplectic algorithms. Applied Mathematics and Computation, 186(1):670–684, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306009799.

Kheowan:2005:CSW

[KM05a] On-Uma Kheowan and Stefan C. Müller. Control of spiral waves in excitable media. Applied Mathematics and Computation, 164(2):373–390, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400325X.

Kiymaz:2005:NSCa

[KM05b] Onur Kıymaz and Şeref Mirasyedioğlu. A new symbolic computation for formal integration with exact power series. Applied Mathematics and Computation, 162(1):215–224, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000839.

Kiymaz:2005:NSCb

[KM05c] Onur Kıymaz and Şeref Mirasyedioğlu. A new symbolic computational approach to singular initial value problems in the second-order ordinary differential equations. Applied Mathematics and Computation, 171(2):1218–1225, December 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kumar:2005:TAS

[KM05d] B. V. Rathish Kumar and Mani Mehra. Timeaccurate solutions of Korteweg– de Vries equation using wavelet Galerkin method. Applied Mathematics and Computation, 162(1):447– 460, March 4, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304001080.

Kumar:2005:WBP

[KM05e] B. V. Rathish Kumar and Mani Mehra. Wavelet based preconditioners for sparse linear systems. Applied Mathematics and Computation, 171(1):203– 224, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305000901.

Kumar:2005:WMT

[KM05f] B. V. Rathish Kumar and Mani Mehra. Wavelet multilayer Taylor Galerkin schemes for hyperbolic and parabolic problems. Applied Mathematics and Computation, 166(2):312– 323, July 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304004394.

Kachiashvili:2006:POA

[KM06] K. J. Kachiashvili and D. I.

Melikdzhanian. Parameter optimization algorithms of difference calculation schemes for improving the solution accuracy of diffusion equations describing the pollutants transport in rivers. Applied Mathematics and Computation, 183(2):787–803, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006138.

Klimowicz:2007:MSD

[KM07] A. N. F. Klimowicz and M. D. Mihajlović. Modelling 3D semi-deformable tubes in real time. Applied Mathematics and Computation, 184(1):52–62, January 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600484X.

Kafini:2009:UDV

[KM09a] Mohammad Kafini and Salim A. Messaoudi. On the uniform decay in viscoelastic problems in R^{n}. Applied Mathematics and Computation, 215(3):1161–1169, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006298.

Khalique:2009:LGC

[KM09b] C. Masood Khalique and Ben Muatjetjeja. Lie group classification of the generalized Lane–Emden equation. Applied Mathematics and Computation, 210(2):405–410, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000083.

Karmitsa:2008:LMI

[KMA08] N. M. S. Karmitsa, M. M. Mäkelä, and M. M. Ali. Limited memory interior point bundle method for large inequality constrained nonsmooth minimization. Applied Mathematics and Computation, 198(1):382–400, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009009.

Kamyad:2005:BLA

[KMB05] Ali Vahidian Kamyad, Hamed Hashemi Mehne, and Akbar Hashemi Borzabadi. The best linear approximation for nonlinear systems. Applied Mathematics and Computation, 167(2):1041–1061, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400548X.

Knoll:2007:NSD

[KMK07] Peter Knoll, Siroos Mirzaei, and Horst Köhn. Numerical simulation of dynamic sand dunes. Applied Mathematics and Computation, 190(2):1747–1753, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002226.

Ke:2009:OMG

[KMN09] Hua Ke, Weimin Ma, and Yaodong Ni. Optimization models and a GA-based algorithm for stochastic time-cost tradeoff problem. Applied Mathematics and Computation, 215(1):308–313, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004615.

Karpuz:2009:OCC

[KMÖS09] B. Karpuz, J. V. Manojlović, Ö. Öcalan, and Y. Shoukaku. Oscillation criteria for a class of second-order neutral delay differential equations. Applied Mathematics and Computation, 210(2):303–312, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009983.

Khan:2009:ESS

[KMVL09] Hina Khan, Ram N. Mohapatra, K. Vajravelu, and S. J. Liao. The explicit series solution of SIR and SIS epidemic models. Applied Mathematics and Computation, 215(2):653–669, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005451.

Kulshreshtha:2005:PMF

[KNJ05] Kshitij Kulshreshtha, Neela Nataraj, and Michael Jung. A parallel mixed finite element implementation of fourth-order plate bending problems in distributed memory environments. Applied Mathematics and Computation, 163(1):253– 274, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002243.

Kano:2005:OCF

[KNM05] Hiroyuki Kano, Hiroaki Nakata, and Clyde F. Martin. Optimal curve fitting and smoothing using normalized uniform Bsplines: a tool for studying complex systems. Applied Mathematics and Computation, 169(1):96–128, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007556.

Kumar:2006:SRE

[KO06] Suresh Kumar and Mohamed Khaled Omar. Stochastic re-entrant line modeling for an environment stress testing in a semiconductor assembly industry. Applied Mathematics and Computation, 173(1):603–615, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305003796.

Kim:2008:DFC

[KO08] Jindae Kim and Changsoo Ok. Distributed feedback control algorithm for dynamic truck loading scheduling problem. Applied Mathematics and Computation, 199(1):275–284, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009952.

Koksoy:2006:MRD

[Kök06]  Onur Köksoy.  Mul-

tiresponse robust design: Mean square error (MSE) criterion. Applied Mathematics and Computation, 175(2):1716–1729, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305007332.

Koksoy:2008:NPS

[Kök08] Onur Köksoy. A nonlinear programming solution to robust multi-response quality problem. Applied Mathematics and Computation, 196(2):603–612, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700714X.

Koloskov:2006:CNS

[Kol06] G. Koloskov. Closed natural splines: an effective solution. Applied Mathematics and Computation, 177(1):111–117, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008933.

Koplinger:2006:DEH

[Köp06]  Jens Köplinger.  Dirac

equation on hyperbolic octonions. Applied Mathematics and Computation, 182(1):443–446, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003511.

Kara:2007:BSM

[KOP07a] Yakup Kara, Ugur Ozcan, and Ahmet Peker. Balancing and sequencing mixed-model just-in-time U-lines with multiple objectives. Applied Mathematics and Computation, 184(2):566–588, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007314.

Koplinger:2007:GEC

[Köp07b] Jens Köplinger. Gravity and electromagnetism on conic sedenions. Applied Mathematics and Computation, 188(1):948–953, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014421.

Koplinger:2007:HR

[Köp07c] Jens Köplinger. Hypernumbers and relativity. Applied Mathematics and Computation, 188(1):954– 969, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014433.

Koplinger:2007:SGC

[Köp07d] Jens Köplinger. Signature of gravity in conic sedenions. Applied Mathematics and Computation, 188 (1):942–947, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030601441X.

Koprowski:2022:NPN

[Kop22] Przemysław Koprowski. A note on the paper “A new approach for finding the determinant of matrices” by o. rezaifar and h. rezaee. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003198.

Karas:2008:GCS

[KOR08] Elizabeth W. Karas, Ana P.

Oening, and Ademir A. Ribeiro. Global convergence of slanting filter methods for nonlinear programming. Applied Mathematics and Computation, 200(2):486–500, July 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011290.

Kostreva:2005:MMC

[Kos05] Michael M. Kostreva. A mathematical model of a concurrent vehicle development system: estimating lead time. Applied Mathematics and Computation, 165(3):571–586, June 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004035.

Kose:2006:KDGb

[Kös06a] Ö. Köse. Kinematic differential geometry of a rigid body in spatial motion using dual vector calculus: Part I. Applied Mathematics and Computation, 183(1):17–29, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003419.

Kosmidou:2006:RCP

[Kos06b] O. I. Kosmidou. Robust control with pole shifting via performance index modification. Applied Mathematics and Computation, 182(1):596–606, November 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306003687.

Kosmidou:2007:GRE

[Kos07] O. I. Kosmidou. Generalized Riccati equations associated with guaranteed cost control: an overview of solutions and features. Applied Mathematics and Computation, 191(2):511–520, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002792.

Kryazhimskiy:2008:IHD

[KOS08] A. Kryazhimskiy, M. Obersteiner, and A. Smirnov. Infinite-horizon dynamic programming and application to management of economies effected by random natural hazards. Applied Mathematics and Computation, 204(2):609– 620, October 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308003093.

Kou:2007:FOV

[Kou07a]  Jisheng Kou.  Fourth-

order variants of Cauchy’s method for solving nonlinear equations. Applied Mathematics and Computation, 192(1):113–119, September 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307003001.

Kou:2007:IMN

[Kou07b] Jisheng Kou. The improvements of modified Newton’s method. Applied Mathematics and Computation, 189(1):602–609, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016341.

Kou:2007:CHM

[Kou07c] Jisheng Kou. On Chebyshev– Halley methods with sixthorder convergence for solving non-linear equations. Applied Mathematics and Computation, 190(1):126– 131, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000355.

Kou:2007:SDF

[Kou07d]  Jisheng Kou.  Second-

derivative-free variants of Cauchy’s method. Applied Mathematics and Computation, 190(1):339–344, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000549.

Kou:2007:SNS

[Kou07e] Jisheng Kou. Some new sixth-order methods for solving non-linear equations. Applied Mathematics and Computation, 189 (1):647–651, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306016389.

Kou:2007:TOM

[Kou07f] Jisheng Kou. A third-order modification of Newton method for systems of nonlinear equations. Applied Mathematics and Computation, 191(1):117–121, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001865.

Kovacs:2007:DHA

[Kov07]  Sándor Kovács.  Dy-

namics of an HIV/AIDS model — The effect of time delay. Applied Mathematics and Computation, 188(2):1597–1609, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601513X.

Kadalbajoo:2006:UCFa

[KP06a] Mohan K. Kadalbajoo and Kailash C. Patidar. ϵ-Uniformly convergent fitted mesh finite difference methods for general singular perturbation problems. Applied Mathematics and Computation, 179(1):248–266, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010349.

Kwon:2006:ERS

[KP06b] O. M. Kwon and Ju H. Park. Erratum to “Robust stabilization of uncertain systems with delays in control input: A matrix inequality approach” [Appl. Math. Comput. 172 (2) (2006) 1067– 1077]. Applied Mathematics and Computation, 182(2):1882, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004565.

See [KP06c].

Kwon:2006:RSU

[KP06c] O. M. Kwon and Ju H. Park. Robust stabilization of uncertain systems with delays in control input: a matrix inequality approach. Applied Mathematics and Computation, 172(2):1067–1077, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500278X.

See erratum [KP06b].

Kechagias:2007:CME

[KP07] P. S. Kechagias and Basil K. Papadopoulos. Computational method to evaluate fuzzy arithmetic operations. Applied Mathematics and Computation, 185(1):169–177, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008241.

Kwon:2008:ESU

[KP08a] O. M. Kwon and Ju H.

Park. Exponential stability for uncertain cellular neural networks with discrete and distributed timevarying delays. Applied Mathematics and Computation, 203(2):813–823, September 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308004086.

Kwon:2008:NDD

[KP08b] O. M. Kwon and Ju H.

Park. New delay-dependent robust stability criterion for uncertain neural networks with time-varying delays. Applied Mathematics and Computation, 205(1):417–427, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006073.

See comment [ZZW09b].

Kwon:2008:IDD

[KP08c] O. M. Kwon and Ju H. Park. On improved delaydependent stability criterion of certain neutral differential equations. Applied Mathematics and Computation, 199(1):385–391, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009733.

Kwon:2009:DRD

[KP09a] O. M. Kwon and Ju H. Park. Delay-range-dependent stabilization of uncertain dynamic systems with interval time-varying delays. Applied Mathematics and Computation, 208(1):58–68, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008473.

Kwon:2009:ESA

[KP09b] O. M. Kwon and Ju H.

Park. Exponential stability analysis for uncertain neural networks with interval time-varying delays. Applied Mathematics and Computation, 212(2):530–541, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001854.

Koutras:2007:SRR

[KPG07] Vasilis P. Koutras, Agapios N. Platis, and George A. Gravvanis. Software rejuvenation for resource optimization based on explicit approximate inverse preconditioning. Applied Mathematics and Computation, 189(1):163–177, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601589X.

Kwon:2008:DDR

[KPL08a] O. M. Kwon, Ju H.

Park, and S. M. Lee. On delay-dependent robust stability of uncertain neutral systems with interval time-varying delays. Applied Mathematics and Computation, 203(2):843–853, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800413X.

Kwon:2008:RSC

[KPL08b] O. M. Kwon, Ju H. Park, and S. M. Lee. On robust stability criterion for dynamic systems with time-varying delays and nonlinear perturbations. Applied Mathematics and Computation, 203(2):937–942, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004177.

Kwon:2008:SCU

[KPL08c] O. M. Kwon, Ju H. Park, and S. M. Lee. On stability criteria for uncertain delay-differential systems of neutral type with timevarying delays. Applied Mathematics and Computation, 197(2):864–873, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008569.

Kwon:2009:CPP

[KPL09a] O. M. Kwon, J. H. Park, and S. M. Lee. Corrigendum to published papers in Applied Mathematics and Computation. Applied Mathematics and Computation, 215(1):427–430, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003920.

Kwon:2009:ALF

[KPL09b] O. M. Kwon, Ju H. Park, and S. M. Lee. Augmented Lyapunov functional approach to stability of uncertain neutral systems with timevarying delays. Applied Mathematics and Computation, 207(1):202–212, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007807.

Kadalbajoo:2006:UCFb

[KPS06a] Mohan K. Kadalbajoo, Kailash C. Patidar, and Kapil K. Sharma. ϵ-Uniformly convergent fitted methods for the numerical solution of the problems arising from singularly perturbed general DDEs. Applied Mathematics and Computation, 182(1):119–139, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003195.

Karczmarek:2006:AJP

[KPS06b] Paweł Karczmarek, Dorota Pylak, and Michail A. Sheshko. Application of Jacobi polynomials to approximate solution of a singular integral equation with Cauchy kernel. Applied Mathematics and Computation, 181(1):694–707, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001494.

Kanth:2005:CSC

[KR05] A. S. V. Ravi Kanth and Y. N. Reddy. Cubic spline for a class of singular two-point boundary value problems. Applied Mathematics and Computation, 170(2):733–740, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000147.

Kadalbajoo:2007:HMN

[KR07a] Mohan K. Kadalbajoo and V. P. Ramesh. Hybrid method for numerical solution of singularly perturbed delay differential equations. Applied Mathematics and Computation, 187(2):797–814, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012136.

Kadalbajoo:2007:NMS

[KR07b] Mohan K. Kadalbajoo and V. P. Ramesh. Numerical methods on Shishkin mesh for singularly perturbed delay differential equations with a grid adaptation strategy. Applied Mathematics and Computation, 188(2):1816–1831, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015505.

Kim:2007:DSF

[KR07c] Kwangsoo Kim and Hong Seo Ryoo. Data separation via a finite number of discriminant functions: a global optimization approach. Applied Mathematics and Computation, 190(1):476– 489, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000793.

Kim:2007:NSD

[KR07d] Kwangsoo Kim and Hong Seo Ryoo. Nonlinear separation of data via Mixed 0–1 Integer and Linear Programming. Applied Mathematics and Computation, 193(1):183–196, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307003852.

Kannan:2008:FTP

[KR08] K. Kannan and K. R. Rajagopal. Flow through porous media due to high pressure gradients. Applied Mathematics and Computation, 199(2):748–759, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010673.

Kassem:2009:GST

[KR09] M. M. Kassem and A. S. Rashed. Group solution of a time dependent chemical convective process. Applied Mathematics and Computation, 215(5):1671–1684, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006675.

See comment [EM11].

Khalifa:2007:FDS

[KRA07] A. K. Khalifa, K. R.

Raslan, and H. M. Alzubaidi. A finite difference scheme for the MRLW and solitary wave interactions. Applied Mathematics and Computation, 189(1):346–354, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016080.

Karrakchou:2006:OCI

[KRG06] J. Karrakchou, M. Rachik, and S. Gourari. Optimal control and infectiology: Application to an HIV/AIDS model. Applied Mathematics and Computation, 177(2):807–818, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009616.

Kilbas:2007:ASS

[KRRGT07] A. A. Kilbas, M. Rivero, L. Rodríguez-Germá, and J. J. Trujillo. α-Analytic solutions of some linear fractional differential equations with variable coefficients. Applied Mathematics and Computation, 187(1):239–249, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011581.

Kim:2005:CLL

[KRY05] Kee-Won Kim, Eun-Kyung Ryu, and Kee-Young Yoo. Cryptanalysis of Lee–Lee authenticated key agreement scheme. Applied Mathematics and Computation, 163(1):193–198, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002206.

Khan:2005:IPD

[KS05a] T. Khan and A. Smirnova. 1D inverse problem in diffusion based optical tomography using iteratively regularized Gauss– Newton algorithm. Applied Mathematics and Computation, 161(1):149–170, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303012177.

Kumar:2005:DDN

[KS05b] B. V. Rathish Kumar and Shalini. Double diffusive natural convection in a doubly stratified wavy porous enclosure. Applied Mathematics and Computation, 171(1):180–202, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000895.

Kim:2006:HSC

[KS06a] Yong Chan Kim and H. M.

Srivastava. The Hardy space for a certain subclass of Bazilevič functions. Applied Mathematics and Computation, 183(2):1201–1207, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006643.

Kurt:2006:SDP

[KS06b] Nurcan Kurt and Mehmet Sezer. Solution of Dirichlet problem for a triangle region in terms of elliptic functions. Applied Mathematics and Computation, 182(1):73–81, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003146.

Kadalbajoo:2008:NMB

[KS08a] Mohan K. Kadalbajoo and Kapil K. Sharma. A numerical method based on finite difference for boundary value problems for singularly perturbed delay differential equations. Applied Mathematics and Computation, 197(2):692– 707, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008260.

Kaushik:2008:UCN

[KS08b] Aditya Kaushik and M. D. Sharma. A uniformly convergent numerical method on non-uniform mesh for singularly perturbed unsteady Burger–Huxley equation. Applied Mathematics and Computation, 195(2):688–706, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700608X.

Kavaklioglu:2008:FTR

[KS08c] Ömer Kavaklıoğlu and Baruch Schneider. On Floquet–Twersky representation for the diffraction of obliquely incident plane Hpolarized electromagnetic waves by an infinite grating of insulating dielectric circular cylinders. Applied Mathematics and Computation, 201(1–2):1–15, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011526.

Kavaklioglu:2008:ASF

[KS08d] Ömer Kavaklıoğlu and Baruch Schneider. On the asymptotic solution for the Fourier–Bessel multiple scattering coefficients of an infinite grating of insulating dielectric circular cylinders at oblique incidence. Applied Mathematics and Computation, 201(1–2):351–360, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030701212X.

Kim:2008:SSP

[KS08e] Yong Chan Kim and H. M.

Srivastava. Some subordination properties for spirallike functions. Applied Mathematics and Computation, 203(2):838–842, September 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308004116.

Kung:2008:ATS

[KS08f] Kuang Yuan Kung and H. M. Srivastava. Analytic transient solutions of a cylindrical heat equation with oscillating heat flux. Applied Mathematics and Computation, 195(2):745–753, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006121.

Kucuk:2009:NAO

[KS09a] Ismail Kucuk and Ibrahim Sadek. A numerical approach to an optimal boundary control of the viscous Burgers’ equation. Applied Mathematics and Computation, 210(1):126– 135, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308009442.

Kudryashov:2009:ESE

[KS09b] Nikolai A. Kudryashov and Dmitry I. Sinelshchikov. Exact solutions of equations for the Burgers hierarchy. Applied Mathematics and Computation, 215(3):1293–1300, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005761.

Kadalbajoo:2005:PUI

[KSA05] Mohan K. Kadalbajoo, K. K. Sharma, and A. Awasthi. A parameter-uniform implicit difference scheme for solving time-dependent Burgers’ equations. Applied Mathematics and Computation, 170(2):1365– 1393, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030500072X.

Kim:2008:FRU

[KSC^{+}08] Young-Gil Kim, Young-Jun Song, Un-Dong Chang, Dong-Woo Kim, Tae-Sung Yun, and Jae-Hyeong Ahn. Face recognition using a fusion method based on bidirectional 2DPCA. Applied Mathematics and Computation, 205(2):601–607, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800338X.

Kanwar:2006:MOC

[KSGM06] V. Kanwar, Sukhjit Singh, R. K. Guha, and Mamta. On method of osculating circle for solving nonlinear equations. Applied Mathematics and Computation, 176(1):379–382, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500812X.

Kose:2006:KDGa

[KSKK06] Ö. Köse, C. C. Sarıoğlu, B. Karabey, and İ. Karakılıç. Kinematic differential geometry of a rigid body in spatial motion using dual vector calculus: Part II. Applied Mathematics and Computation, 182(1):333–358, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003420.

Kanwar:2005:NFS

[KSM05] V. Kanwar, J. R. Sharma, and Mamta. A new family of Secant-like method with super-linear convergence. Applied Mathematics and Computation, 171(1):104–107, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000834.

Khattar:2009:NCA

[KSM09] Dinesh Khattar, Swarn Singh, and R. K. Mohanty. A new coupled approach high accuracy numerical method for the solution of 3D non-linear biharmonic equations. Applied Mathematics and Computation, 215(8):3036–3044, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008819.

Ko:2008:RAC

[KSP08] K. H. Ko, T. Sakkalis, and N. M. Patrikalakis. A reliable algorithm for computing the topological degree of a mapping in R ^{2}. Applied Mathematics and Computation, 196(2):666–678, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007205.

Kogan:2007:NIS

[KSS07] Tamara Kogan, Luba Sapir, and Amir Sapir. A nonstationary iterative second-order method for solving nonlinear equations. Applied Mathematics and Computation, 188 (1):75–82, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306013233.

Khalik:2007:PIP

[KSSA07] Mostafa A. Khalik, M. Sherif, S. Saraya, and F. Areed. Parameter identification problem: Real-coded GA approach. Applied Mathematics and Computation, 187(2):1495–1501, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012884.

Karimi:2006:BLS

[KT06] S. Karimi and F. Toutounian. The block least squares method for solving nonsymmetric linear systems with multiple righthand sides. Applied Mathematics and Computation, 177(2):852–862, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009665.

Kanwar:2007:MFN

[KT07a] V. Kanwar and S. K. Tomar. Modified families of Newton, Halley and Chebyshev methods. Applied Mathematics and Computation, 192(1):20– 26, September 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307002895.

Koulaei:2007:FSA

[KT07b] Mohammad H. Koulaei and Faezeh Toutounian. Factored sparse approximate inverse of block tridiagonal and block pentadiagonal matricies. Applied Mathematics and Computation, 184(2):223–234, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006990.

Kung:2007:RFB

[KT07c] Ching-Jing Kung and Hui-Chin Tang. A revised forward and backward heuristic for two-term multiple recursive random number generators. Applied Mathematics and Computation, 185(1):240–246, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008551.

Kohsaka:2008:SCN

[KT08a] Fumiaki Kohsaka and Wataru Takahashi. Strongly convergent net given by a fixed point theorem for firmly nonexpansive type mappings. Applied Mathematics and Computation, 202(2):760–765, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001768.

Krasovskii:2008:ACO

[KT08b] Andrey A. Krasovskii and Alexander M. Tarasyev. An algorithm for construction of optimal timing solutions in problems with a stochastic payoff function. Applied Mathematics and Computation, 204(2):632–643, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003111.

Kaabi:2006:PGM

[KTK06] A. Kaabi, F. Toutounian, and A. Kerayechian. Preconditioned Galerkin and minimal residual methods for solving Sylvester equations. Applied Mathematics and Computation, 181(2):1208–1214, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002402.

Komurcu:2008:EBB

[KTÖA08] Murat İhsan Kömürcü, Nedim Tutkun, İsmail Hakkı Özölçer, and Adem Akpınar. Estimation of the beach bar parameters using the genetic algorithms. Applied Mathematics and Computation, 195(1):49–60, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005164.

Kostreva:2007:MFF

[KTV07] Michael M. Kostreva, Latif Talybly, and Irina V. Viktorova. Modeling of fatigue fracture under stationary stochastic loading conditions. Applied Mathematics and Computation, 184(2):874–879, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007752.

Kellegoz:2008:CEG

[KTW08] Talip Kellegöz, Bilal Toklu, and John Wilson. Comparing efficiencies of genetic crossover operators for one machine total weighted tardiness problem. Applied Mathematics and Computation, 199(2):590–598, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030701034X.

Kucuk:2005:FGO

[Küç05a] Ahmet Küçük. The fundamental group of an oriented surface of genus n with k boundary surfaces. Applied Mathematics and Computation, 160(1):141–145, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303010968.

Kucukarslan:2005:ETB

[Küç05b] Semih Küçükarslan. An exact truncation boundary condition for incompressible– unbounded infinite fluid domains. Applied Mathematics and Computation, 163(1):61–69, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304002085.

Kudryashov:2009:SSS

[Kud09] Nikolai A. Kudryashov. Self-similar solutions of the Burgers hierarchy. Applied Mathematics and Computation, 215(5):1990–1993, November 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309007085.

Khattak:2008:CSN

[KuI08] A. J. Khattak and Siraj ul Islam. A comparative study of numerical solutions of a class of KdV equation. Applied Mathematics and Computation, 199(2):425–434, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010090.

Kumar:2003:SOS

[Kum03] Manoj Kumar. A second order spline finite difference method for singular two-point boundary value problems. Applied Mathematics and Computation, 142(2–3):283–290, October 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See erratum [Kum09c].

Kumar:2006:MMS

[Kum06] Avanish Kumar. A modified method for solving the unbalanced assignment problems. Applied Mathematics and Computation, 176(1):76– 82, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305007721.

Kumar:2007:HOM

[Kum07] Manoj Kumar. Higher order method for singular boundary-value problems by using spline function. Applied Mathematics and Computation, 192(1):175–179, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307003098.

Kumar:2008:NCP

[Kum08] R. Ravi Kumar. Nonlocal Cauchy problem for analytic resolvent integrodifferential equations in Banach spaces. Applied Mathematics and Computation, 204(1):352–362, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005134.

Kumar:2009:ASC

[Kum09a] Arun Kumar. An analytical solution for a coupled partial differential equation. Applied Mathematics and Computation, 212(1):245–250, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001404.

Kumar:2009:MHM

[Kum09b] Arun Kumar. A model for hyperparasitism in malaria with environmental fluctuations. Applied Mathematics and Computation, 215(1):206–211, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004391.

Kumar:2009:ESO

[Kum09c]  Manoj Kumar.  Erra-

tum to: “A second order spline finite difference method for singular twopoint boundary value problems” [Appl. Math. Comput. 142 (2003) 283–290]. Applied Mathematics and Computation, 209(2):430, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009624.

See [Kum03].

Kung:2008:MPA

[Kun08] James J. Kung. Multiperiod asset allocation by stochastic dynamic programming. Applied Mathematics and Computation, 199(1):341–348, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030701003X.

Kuo:2005:ADT

[Kuo05]  Bor-Lih Kuo.  Appli-

cation of the differential transformation method to the solutions of the free convection problem. Applied Mathematics and Computation, 165(1):63– 79, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003625.

Kutluay:2005:NSO

[Kut05]  S. Kutluay.  Numer-

ical schemes for onedimensional Stefan-like problems with a forcing term. Applied Mathematics and Computation, 168(2):1159–1168, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400671X.

Kajani:2005:SSK

[KV05a] M. Tavassoli Kajani and A. Hadi Vencheh. Solving second kind integral equations with Hybrid Chebyshev and Block– Pulse functions. Applied Mathematics and Computation, 163(1):71– 77, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002097.

Koleva:2005:BFD

[KV05b] Miglena N. Koleva and Lubin G. Vulkov. On the blow-up of finite difference solutions to the heat-diffusion equation with semilinear dynamical boundary conditions. Applied Mathematics and Computation, 161(1):69–91, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012104.

Kourdis:2006:SRD

[KV06] Panayotis D. Kourdis and Alexander F. Vakakis. Some results on the dynamics of the linear parametric oscillator with general time-varying frequency. Applied Mathematics and Computation, 183(2):1235–1248, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006692.

Korkut:2008:GFI

[KVŽŽ08] Luka Korkut, Domagoj Vlah, Darko Žubrinić, and Vesna Županović. Generalized Fresnel integrals and fractal properties of related spirals. Applied Mathematics and Computation, 206(1):236–244, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006723.

Kaya:2005:PARa

[KW05a] D. Kaya and K. Wright. Parallel algorithms for reduction of a general matrix to upper Hessenberg form on a shared memory multiprocessor. Applied Mathematics and Computation, 165(1):195–212, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030400373X.

Kaya:2005:PAL

[KW05b] Dogan Kaya and Ken Wright. Parallel algorithms for LU decomposition on a shared memory multiprocessor. Applied Mathematics and Computation, 163(1):179–191, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400219X.

Kou:2007:SOV

[KW07] Jisheng Kou and Xiuhua Wang. Sixth-order variants of Chebyshev–Halley methods for solving nonlinear equations. Applied Mathematics and Computation, 190(2):1839–1843, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002330.

Kang:2008:MPS

[KW08a] Ping Kang and Zhongli Wei. Multiple positive solutions of multi-point boundary value problems on the half-line. Applied Mathematics and Computation, 196(1):402–415, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006935.

Kang:2008:MSS

[KW08b] Ping Kang and Zhongli Wei. Multiple solutions of second-order three-point singular boundary value problems on the halfline. Applied Mathematics and Computation, 203(2):523–535, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002816.

Kang:2008:NEP

[KWdW08] Qi Kang, Lei Wang, and Qi di Wu. A novel ecological particle swarm optimization algorithm and its population dynamics analysis. Applied Mathematics and Computation, 205(1):61–72, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003184.

Kwon:2007:OTS

[Kwo07a] Hee-Dae Kwon. Optimal treatment strategies derived from a HIV model with drug-resistant mutants. Applied Mathematics and Computation, 188(2):1193–1204, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014639.

Kwong:2007:TDF

[Kwo07b] Harris Kwong. Two determinants with Fibonacci and Lucas entries. Applied Mathematics and Computation, 194(2):568–571, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004729.

Kuo:2007:SCM

[KWP07] Ming-Shin Kuo, Jer-Wei Wu, and Lie Pei. A soft computing method for selecting evaluation criteria of service quality. Applied Mathematics and Computation, 189(1):241–254, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015980.

Kartono:2005:ANO

[KWS05] Agus Kartono, Toto Winata, and Sukirno. Applications of non-orthogonal Laguerre function basis in helium atom. Applied Mathematics and Computation, 163(2):879–893, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030400267X.

Kang:2008:PSF

[KWX08] Ping Kang, Zhongli Wei, and Juanjuan Xu. Positive solutions to fourthorder singular boundary value problems with integral boundary conditions in abstract spaces. Applied Mathematics and Computation, 206(1):245–256, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006747.

Kong:2008:AAP

[KWZ08] Linghua Kong, Jinhuan Wang, and Sining Zheng. Asymptotic analysis to a parabolic equation with a weighted localized source. Applied Mathematics and Computation, 197(2):819– 827, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008429.

Koulouriotis:2008:RLE

[KX08] D. E. Koulouriotis and A. Xanthopoulos. Reinforcement learning and evolutionary algorithms for non-stationary multiarmed bandit problems.

Applied Mathematics and Computation, 196(2):913– 922, March 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307007448.

Kaya:2005:DMF

[KY05a] Doğan Kaya and Asif Yokus. A decomposition method for finding solitary and periodic solutions for a coupled higher-dimensional Burgers equations. Applied Mathematics and Computation, 164(3):857–864, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003522.

Kula:2005:SHS

[KY05b] L. Kula and Y. Yayli. On slant helix and its spherical indicatrix. Applied Mathematics and Computation, 169(1):600–607, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007970.

Koksoy:2006:MSE

[KY06] Onur Köksoy and Tankut Yalcinoz. Mean square error criteria to multiresponse process optimization by a new genetic algorithm. Applied Mathematics and Computation, 175(2):1657–1674, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007228.

Kadalbajoo:2008:BSC

[KY08] Mohan K. Kadalbajoo and Arjun Singh Yadaw. Bspline collocation method for a two-parameter singularly perturbed convection– diffusion boundary value problems. Applied Mathematics and Computation, 201(1–2):504–513, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030701226X.

Kasap:2005:NSC

[KYA05] Emin Kasap, Mustafa Yapici, and F. Talay Akyildiz. A numerical study for computation of geodesic curves. Applied Mathematics and Computation, 171(2):1206–1213, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadalbajoo:2008:CSS

[KYK08] M. K. Kadalbajoo, Arjun Singh Yadaw, and Devendra Kumar. Comparative study of singularly perturbed two-point BVPs via: Fitted-mesh finite difference method, Bspline collocation method and finite element method. Applied Mathematics and Computation, 204(2):713– 725, October 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308005456.

Kim:2005:MMN

[KYL05] Kyungsoo Kim, Dokkyun Yi, and Sungyun Lee. Mortar method for nonconforming finite elements. Applied Mathematics and Computation, 167(1):650– 669, August 5, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304005211.

Kilicman:2007:KOM

[KZ07] Adem Kilicman and Zeyad Abdel Aziz Al Zhour. Kronecker operational matrices for fractional calculus and some applications. Applied Mathematics and Computation, 187(1):250– 265, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011593.

Loghmani:2007:NSS

[LA07a] G. B. Loghmani and M. Ahmadinia. Numerical solution of sixth order boundary value problems with sixth degree Bspline functions. Applied Mathematics and Computation, 186(2):992–999, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010241.

Loghmani:2007:NSF

[LA07b] G. B. Loghmani and S. R. Alavizadeh. Numerical solution of fourthorder problems with separated boundary conditions. Applied Mathematics and Computation, 191(2):571–581, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002858.

Laksmikantham:2005:SDE

[Lak05] V. Laksmikantham. Set differential equations versus fuzzy differential equations. Applied Mathematics and Computation, 164(2):277–294, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400311X.

Li:2007:SOE

[LAL07] Yuan Li, Rong An, and Kaitai Li. Some optimal error estimates of biharmonic problem using conforming finite element. Applied Mathematics and Computation, 194(2):298–308, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004705.

Lal:2009:AFB

[Lal09] Shyam Lal. Approximation of functions belonging to the generalized Lipschitz Class by C^{1} · N_{p} summability method of Fourier series. Applied Mathematics and Computation, 209(2):346–350, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009703.

Landi:2007:FTL

[Lan07] G. Landi. A fast truncated Lagrange method for large-scale image restoration problems. Applied Mathematics and Computation, 186(2):1075–1082, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010332.

Lara:2005:NMS

[Lar05] Luis Lara. A numerical method for solving a system of nonautonomous linear ordinary differential equations. Applied Mathematics and Computation, 170(1):86–94, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008720.

Layeni:2008:RMA

[Lay08]  O. P. Layeni.  Re-

mark on modifications of Adomian decomposition method. Applied Mathematics and Computation, 197(1):167–171, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307007655.

Layeni:2009:NRA

[Lay09a] O. P. Layeni. A new rational auxiliary equation method and exact solutions of a generalized Zakharov system. Applied Mathematics and Computation, 215(8):2901–2907, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008595.

Layeni:2009:NES

[Lay09b]  O. P. Layeni.  Note

on exact solutions for a class of nonlinear diffusion equations. Applied Mathematics and Computation, 210(2):465–472, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000150.

Lazo:2008:GSN

[Laz08] Pedro Pablo Durand Lazo.

Global solutions for a nonlinear wave equation. Applied Mathematics and Computation, 200(2):596– 601, July 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011393.

Lin:2006:BSO

[LB06] Yiqin Lin and Liang Bao.

Block second-order Krylov subspace methods for large-scale quadratic eigenvalue problems. Applied Mathematics and Computation, 181(1):413–422, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001111.

Liang:2005:COP

[LBS05] Heng Liang, Fengshan Bai, and Leyuan Shi. Computing the optimal partition of variables in multi-homogeneous homotopy methods. Applied Mathematics and Computation, 163(2):825–840, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002632.

Liu:2008:CMD

[LBZY08] Yuan-Hua Liu, Jing-Xiu Bi, An-Ping Zeng, and Jing-Qi Yuan. A cybernetic model to describe the dynamics of myeloma cell cultivations. Applied Mathematics and Computation, 205(1):84–97, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003202.

Li:2005:IBM

[LC05a] Xiangxue Li and Kefei Chen. ID-based multiproxy signature, proxy multi-signature and multiproxy multi-signature schemes from bilinear pairings. Applied Mathematics and Computation, 169(1):437– 450, October 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304007842.

Lin:2005:SFG

[LC05b] Ming-Han Lin and Chin-Tai Chen. Study on the formation of Goertler vortices in natural convection flow over a rotating concave surface. Applied Mathematics and Computation, 169(2):778–796, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008100.

Liu:2005:NAM

[LC05c] Zhaoxia Liu and Qianshun Chang. Numerical analysis of the model of image processing with time-delay regularization. Applied Mathematics and Computation, 166(2):349–372, July 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004424.

Lu:2005:DVP

[LC05d] Rongxing Lu and Zhenfu Cao. Designated verifier proxy signature scheme with message recovery. Applied Mathematics and Computation, 169(2):1237– 1246, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304008483.

Lu:2005:NDA

[LC05e] Rongxing Lu and Zhenfu Cao. A new deniable authentication protocol from bilinear pairings. Applied Mathematics and Computation, 168(2):954–961, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304006538.

Li:2006:ETD

[LC06a] Zhong Li and Fengde Chen. Extinction in two dimensional nonautonomous Lotka–Volterra systems with the effect of toxic substances. Applied Mathematics and Computation, 182(1):684–690, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003778.

Lin:2006:NNU

[LC06b] Yiqin Lin and Yanhua Cao. A new nonlinear Uzawa algorithm for generalized saddle point problems. Applied Mathematics and Computation, 175(2):1432–1454, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006880.

Liu:2006:SPP

[LC06c] Chen-Chung Liu and Wen-Yuan Chen. Screw pitch precision measurement using simple linear regression and image analysis. Applied Mathematics and Computation, 178(2):390–404, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009847.

Luo:2006:BBA

[LC06d] Xiaochuan Luo and Feng Chu. A branch and bound algorithm of the single machine schedule with sequence dependent setup times for minimizing total tardiness. Applied Mathematics and Computation, 183(1):575–588, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005923.

Li:2007:GQR

[LC07] Hongwei Li and Kean Chen. Gauss quadrature rules for Cauchy principal value integrals with wavelets. Applied Mathematics and Computation, 186(1):357–364, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009453.

Lou:2008:EGA

[LC08a] Xuyang Lou and Baotong Cui. Existence and global attractivity of almost periodic solutions for neural field with time delay. Applied Mathematics and Computation, 200(1):465–472, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011472.

Lu:2008:SSS

[LC08b] Xueqin Lü and Minggen Cui. Solving a singular system of two nonlinear ODEs. Applied Mathematics and Computation, 198(2):534–543, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009198.

Li:2009:OSO

[LC09a] Chenshun Li and Shuming Chen. Oscillation of second-order functional differential equations with mixed nonlinearities and oscillatory potentials. Applied Mathematics and Computation, 210(2):504– 507, April 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309000290.

Lin:2009:APS

[LC09b] Xiao Lin and Fengde Chen. Almost periodic solution for a Volterra model with mutual interference and Beddington– DeAngelis functional response. Applied Mathematics and Computation, 214(2):548–556, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309003531.

Lee:2005:RCD

[LCC05a] Taih-Shyun Lee, Ye-Hwa Chen, and Jason C.-H. Chuang. Robust control design of fuzzy dynamical systems. Applied Mathematics and Computation, 164(2):555–572, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003376.

Liu:2005:NCM

[LCC05b] Zhong-Yun Liu, Han-Dong Cao, and Hui-Jiong Chen. A note on computing matrix–vector products with generalized centrosymmetric (centrohermitian) matrices. Applied Mathematics and Computation, 169(2):1332–1345, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008550.

Long:2005:DTC

[LCC05c] Yu Long, Zhenfu Cao, and Kefei Chen. A dynamic threshold commercial key escrow scheme based on conic. Applied Mathematics and Computation, 171(2):972–982, December 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:2006:CPS

[LCC06] Zhong-Yun Liu, Hui-Jiong Chen, and Han-Dong Cao. The computation of the principal square roots of centrosymmetric H-matrices. Applied Mathematics and Computation, 175(1):319–329, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006107.

Liu:2009:PGM

[LCH09] Qingbing Liu, Guoliang Chen, and Jian Huang. On the parallel GSAOR method for block diagonally dominant matrices. Applied Mathematics and Computation, 215(2):707–715, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005505.

Li:2008:MPSb

[LCJ08] Qiuyue Li, Fuzhong Cong, and Daqing Jiang. Multiplicity of positive solutions to second order Neumann boundary value problems with impulse actions. Applied Mathematics and Computation, 206(2):810–817, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007133.

Liu:2005:ESG

[LCL05] Xi-Qiang Liu, Han-Lin Chen, and Yong-Qiang Lü. Explicit solutions of the generalized KdV equations with higher order nonlinearity. Applied Mathematics and Computation, 171(1):315–319, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000974.

Li:2006:SSS

[LCL06] Tze Fen Li, Shui-Ching Chang, and Chung-Bow Lee. A simple statistical speech recognition of Mandarin monosyllables. Applied Mathematics and Computation, 177(2):644–651, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009458.

Loff:2007:CRI

[LCM07] Bruno Loff, José Félix Costa, and Jerzy Mycka. Computability on reals, infinite limits and differential equations. Applied Mathematics and Computation, 191(2):353–371, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002639.

Li:2008:GEB

[LCM08] Juan Li, Zejian Cui, and Chunlai Mu. Global existence and blow-up for degenerate and singular parabolic system with localized sources. Applied Mathematics and Computation, 199(1):292–300, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009976.

Cheng:2006:IMP

[lCS06] Xiao liang Cheng and Abdul Wasim Shaikh. Iterative methods for the Poisson equation in L-shape domain based on domain decomposition method. Applied Mathematics and Computation, 180(1):393– 400, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305011343.

Li:2008:DPP

[LCS08] Yongfeng Li, Jingan Cui, and Xinyu Song. Dynamics of a predator-prey system with pulses. Applied Mathematics and Computation, 204(1):269–280, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005043.

Lin:2009:CCI

[LCS09] Shy-Der Lin, Li-Fen Chang, and H. M. Srivastava. A certain class of incomplete elliptic integrals and associated definite integrals. Applied Mathematics and Computation, 215(3):1176–1184, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006328.

Cheng:2006:EEF

[lCX06] Xiao liang Cheng and Lian Xue. On the error estimate of finite difference method for the obstacle problem. Applied Mathematics and Computation, 183(1):416–422, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600573X.

Lu:2009:SEC

[LCX09] Min Lü, Guo-Liang Chen, and Xi-Rong Xu. On super edge-connectivity of product graphs. Applied Mathematics and Computation, 207(2):300–306, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007959.

Liu:2005:DPD

[LCZ05a] Bing Liu, Lansun Chen, and Yujuan Zhang. The dynamics of a prey-dependent consumption model concerning impulsive control strategy. Applied Mathematics and Computation, 169(1):305–320, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007763.

Lu:2005:TUS

[LCZ05b] Rong Xing Lu, Zhen Fu Cao, and Yuan Zhou.

Threshold undeniable signature scheme based on conic. Applied Mathematics and Computation, 162(1):165–177, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304000773.

Lu:2005:PBM

[LCZ05c] RongXing Lu, ZhenFu Cao, and Yuan Zhou. Proxy blind multi-signature scheme without a secure channel. Applied Mathematics and Computation, 164(1):179–187, May 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003042.

Lu:2005:RTP

[LCZ05d] RongXing Lu, ZhenFu Cao, and HaoJin Zhu. A robust (k, n) + 1 threshold proxy signature scheme based on factoring. Applied Mathematics and Computation, 166(1):35– 45, July 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004199.

Li:2008:CLV

[LCZT08] J. Li, Y. Chen, W. Zhang, and Y. Tian. Computation of Lyapunov values for two planar polynomial differential systems. Applied Mathematics and Computation, 204(1):240–248, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005006.

Li:2006:ICO

[LD06a] Xian-Fang Li and Xiang-Yang Duan. An interfaciallycracked orthotropic rectangular bi-material subjected to antiplane shear loading. Applied Mathematics and Computation, 174(2):1060–1079, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005333.

Lin:2006:CNU

[LD06b] Yiqin Lin and Yong Duan.

A corrected nonlinear Uzawa method for solving stabilized saddle point problems. Applied Mathematics and Computation, 183(2):1262–1269, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006722.

Li:2007:RFD

[LD07a] Changpin Li and Weihua Deng. Remarks on fractional derivatives. Applied Mathematics and Computation, 187(2):777–784, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012094.

Li:2007:MSB

[LD07b] Wanggen Li and Yongsheng Ding. A microfluidic systems-based DNA algorithm for solving special 0–1 integer programming problem. Applied Mathematics and Computation, 185(2):1160–1170, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600840X.

Li:2007:GAM

[LD07c] Yaotang Li and Pingfan Dai. Generalized AOR methods for linear complementarity problem. Applied Mathematics and Computation, 188(1):7–18, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030601304X.

Liu:2007:IPN

[LD07d] Hao Liu and Hua Dai. Inverse problems for nonsymmetric matrices with a submatrix constraint. Applied Mathematics and Computation, 189(2):1320–1330, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017188.

Liu:2008:ESS

[LD08] Changfu Liu and Zhengde Dai. Exact soliton solutions for the fifth-order Sawada–Kotera equation. Applied Mathematics and Computation, 206(1):272– 275, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308006085.

Li:2009:IBA

[LDC09] Gang Li, Changming Ding, and Ming Chen. Intertwined basins of attraction of dynamical systems. Applied Mathematics and Computation, 213(1):272– 274, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002288.

Lin:2008:UER

[LDL08] Xiaojie Lin, Zengji Du, and Wenbin Liu. Uniqueness and existence results for a third-order nonlinear multi-point boundary value problem. Applied Mathematics and Computation, 205(1):187–196, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004359.

Li:2006:INI

[LDZ06] Wencheng Li, Zichen Deng, and Suying Zhang. Improved Newton iteration for nonlinear matrix equations on quadratic Lie groups. Applied Mathematics and Computation, 183(1):447–457, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005777.

Lee:2005:IMI

[Lee05a] Wen-Chuan Lee. Inventory model involving controllable backorder rate and variable lead time demand with the mixtures of distribution. Applied Mathematics and Computation, 160(3):701–717, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303011895.

Lee:2005:ADS

[Lee05b] Zong-Yi Lee. Applying the double side method of weighted residual to the solution of shell deformation problem. Applied Mathematics and Computation, 167(1):622–634, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005193.

Lee:2005:HNM

[Lee05c] Zong-Yi Lee. Hybrid numerical method applied to 3-D multilayered hollow cylinder with periodic loading conditions. Applied Mathematics and Computation, 166(1):95– 117, July 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004242.

Leela:2005:UTC

[Lee05d] S. Leela. A unified topological concept of stability in terms of two measures. Applied Mathematics and Computation, 164(2):545–553, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003364.

Lee:2006:OBD

[Lee06a] Jeehyun Lee. An optimizationbased domain decomposition method for parabolic equations. Applied Mathematics and Computation, 175(2):1644–1656, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305007216.

Lee:2006:MBB

[Lee06b] Zong-Yi Lee. Method of bilaterally bounded to solution Blasius equation using particle swarm optimization. Applied Mathematics and Computation, 179(2):779–786, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030501088X.

Lei:2009:ROS

[Lei09] Caiming Lei. Remarks on oscillation of secondorder linear difference equation. Applied Mathematics and Computation, 215(8):2855–2857, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008261.

Lepik:2006:HWM

[Lep06] Ülo Lepik. Haar wavelet method for nonlinear integrodifferential equations. Applied Mathematics and Computation, 176(1):324– 333, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008076.

Lepik:2007:NSE

[Lep07]  Ülo Lepik.  Numeri-

cal solution of evolution equations by the Haar wavelet method. Applied Mathematics and Computation, 185(1):695–704, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009052.

Lepik:2008:SID

[Lep08] Ü. Lepik. Solving integral and differential equations by the aid of non-uniform Haar wavelets. Applied Mathematics and Computation, 198(1):326–332, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008922.

Lepik:2009:SFI

[Lep09] Ü. Lepik. Solving fractional integral equations by the Haar wavelet method. Applied Mathematics and Computation, 214(2):468–478, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003427.

Lesnic:2006:DMI

[Les06] D. Lesnic. The decomposition method for initial value problems. Applied Mathematics and Computation, 181(1):206–213, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000919.

Labidi:2005:UEB

[LeT05] Soraya Labidi and Nasser eddine Tatar. Unboundedness for the Euler– Bernoulli beam equation with a fractional boundary dissipation. Applied Mathematics and Computation, 161(3):697–706, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000219.

Li:2008:EST

[LF08] Lian-He Li and Tian-You Fan. Exact solutions of two semi-infinite collinear cracks in a strip of one dimensional hexagonal quasicrystal. Applied Mathematics and Computation, 196(1):1–5, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006212.

Liu:2008:GAS

[LFC08] Zhijun Liu, Meng Fan, and Lansun Chen. Globally asymptotic stability in two periodic delayed competitive systems. Applied Mathematics and Computation, 197(1):271–287, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007758.

Laporte:2006:DEM

[LFM06] Gilbert Laporte, Reza Zanjirani Farahani, and Elnaz Miandoabchi. Designing an efficient method for tandem AGV network design problem using tabu search. Applied Mathematics and Computation, 183(2):1410–1421, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006904.

Liu:2008:CSC

[LFW08] Hongjie Liu, Boqin Feng, and Guochang Wu. The compactly supported cardinal orthogonal vectorvalued wavelets with dilation factor α. Applied Mathematics and Computation, 205(1):309–316, November 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308005894.

Li:2006:NES

[LG06a] Huaying Li and Yucui Guo. New exact solutions to the Fitzhugh–Nagumo equation. Applied Mathematics and Computation, 180(2):524–528, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000142.

Liu:2006:APS

[LG06b] Mingming Liu and Yan Gao. An algorithm for portfolio selection in a frictional market. Applied Mathematics and Computation, 182(2):1629–1638, November 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306005108.

Liu:2006:NNA

[LG06c] Zhaoxia Liu and Boling Guo. New numerical algorithms for the nonlinear diffusion model of image denoising and segmentation. Applied Mathematics and Computation, 178(2):380–389, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009835.

Lin:2007:MRA

[LG07] Chi-Ming Lin and Mitsuo Gen. Multiobjective resource allocation problem by multistage decision-based hybrid genetic algorithm. Applied Mathematics and Computation, 187(2):574–583, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011131.

Luo:2008:OCF

[LG08] Zhixue Luo and Jinsheng Guo. Optimal control of free horizon problems for an age-dependent ndimensional food chain model. Applied Mathematics and Computation, 203(1):113–124, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002130.

Liu:2006:PVQ

[LGHS06] Xingping Liu, Tongxiang Gu, Xudeng Hang, and Zhiqiang Sheng. A parallel version of QMRCGSTAB method for large linear systems in distributed parallel environments. Applied Mathematics and Computation, 172(2):744–752, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304009087.

Lian:2006:SPSa

[LGJ06] Zhigang Lian, Xingsheng Gu, and Bin Jiao. A similar particle swarm optimization algorithm for permutation flowshop scheduling to minimize makespan. Applied Mathematics and Computation, 175(1):773– 785, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006442.

Lu:2007:OHP

[LGw07] Dianchen Lu, Jianjun Gu, and Xiaoming wang. Optimal harvesting problems for an age-dependent ndimensional food chain model with diffusion. Applied Mathematics and Computation, 184(2):659– 668, January 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306007405.

Liu:2007:OAN

[LGY07] M. Z. Liu, Jianfang Gao, and Z. W. Yang. Oscillation analysis of numerical solution in the θ-methods for equation x′(t) + ax(t) + a_{1}x([t − 1]) = 0. Applied Mathematics and Computation, 186(1):566–578, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009684.

Lehn:2005:TDA

[LH05] Frank Lehn and Eberhard P. Hofer. A timedelay approach for modeling chronic radiation injury dynamics. Applied Mathematics and Computation, 164(2):453–466, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003297.

Ling:2006:IAL

[LH06a] Xiaofeng Ling and Xiaozhe Hu. On the iterative algorithm for large sparse saddle point problems. Applied Mathematics and Computation, 178(2):372–379, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009823.

Liu:2006:NAC

[LH06b] Jianzhou Liu and Anqi He. A new algorithmic characterization of Hmatrices. Applied Mathematics and Computation, 183(1):603–609, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005959.

Lin:2007:DIO

[LH07a] Guojian Lin and Yiguang Hong. Delay induced oscillation in predator–prey system with Beddington– DeAngelis functional response. Applied Mathematics and Computation, 190(2):1296–1311, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001610.

Liu:2007:EBP

[LH07b] Fu-Ti Liu and Ting-Zhu Huang. An error bound for the preconditioned simultaneous displacement method. Applied Mathematics and Computation, 188(1):772–778, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014159.

Liu:2007:IIC

[LH07c] Jianzhou Liu and Anqi He.

An interleaved iterative criterion for H-matrices. Applied Mathematics and Computation, 186(1):727– 734, March 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306009933.

He:2008:BTWa

[lH08a] Tian lan He. Bifurcation of traveling wave solutions of (2 + 1) dimensional Konopelchenko– Dubrovsky equations. Applied Mathematics and Computation, 204(2):773– 783, October 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308005523.

Li:2008:DSM

[LH08b] Li Li and Bo Han. A dynamical system method for solving nonlinear illposed problems. Applied Mathematics and Computation, 197(1):399–406, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007886.

Liao:2008:HTE

[LH08c] Zhensong Liao and Yan Huang. Horseshoe and topological entropy estimate of a class of threedimensional cellular neural networks. Applied Mathematics and Computation, 197(1):382–388, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007862.

Lu:2008:NES

[LH08d] Dianchen Lu and Baojian Hong. New exact solutions for the (2 + 1)-dimensional generalized Broer–Kaup system. Applied Mathematics and Computation, 199(2):572– 580, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010326.

Liang:2006:HAC

[LHB06] Heng Liang, Songqi Huang, and Fengshan Bai. A hybrid algorithm for computing permanents of sparse matrices. Applied Mathematics and Computation, 172(2):708–716, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009051.

Liu:2007:FRP

[LHC07] Chao-Liang Liu, Gwoboa Horng, and Te-Yu Chen. Further refinement of pairing computation based on Miller’s algorithm. Applied Mathematics and Computation, 189(1):395–409, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016146.

Lu:2007:PTW

[LHF07] Junliang Lu, Tianlan He, and Dahe Feng. Persistence of travelling waves for a coupled nonlinear wave system. Applied Mathematics and Computation, 191(2):347–352, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002627.

See comment [Den08b].

Lin:2006:RNN

[LHGT06] Chi-Ming Lin, Jih-Jeng Huang, Mitsuo Gen, and Gwo-Hshiung Tzeng. Recurrent neural network for dynamic portfolio selection. Applied Mathematics and Computation, 175(2):1139–1146, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006661.

Lin:2005:IDV

[LHH05a] Chu-Hsing Lin, Ruei-Hau Hsu, and Lein Harn. Improved DSA variant for batch verification. Applied Mathematics and Computation, 169(1):75–81, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007532.

Lu:2005:NPS

[LHH05b] Eric Jui-Lin Lu, Min-Shiang Hwang, and Cheng-Jian Huang. A new proxy signature scheme with revocation. Applied Mathematics and Computation, 161(3):799–806, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000311.

Li:2006:INU

[LHL06] Hou-Biao Li, Ting-Zhu Huang, and Hong Li. An improvement on a new upper bound for moduli of eigenvalues of iterative matrices. Applied Mathematics and Computation, 173(2):977–984, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004066.

Lv:2008:NCI

[LHL08] Xiao-Guang Lv, Ting-Zhu Huang, and Jiang Le. A note on computing the inverse and the determinant of a pentadiagonal Toeplitz matrix. Applied Mathematics and Computation, 206(1):327–331, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006516.

Lott:2009:NSO

[LHSB09] Dawn A. Lott, Auris Henriquez, Benjamin J. M. Sturdevant, and Anjan Biswas. A numerical study of optical soliton-like structures resulting from the nonlinear Schrödinger’s equation with square-root law nonlinearity. Applied Mathematics and Computation, 207(2):319–326, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008138.

Lv:2007:PFMa

[LHW07] Yibing Lv, Tiesong Hu, and Zhongping Wan. A penalty function method for solving weak price control problem. Applied Mathematics and Computation, 186(2):1520–1525, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010782.

Lv:2007:PFMb

[LHWW07] Yibing Lv, Tiesong Hu, Guangmin Wang, and Zhongping Wan. A penalty function method based on Kuhn–Tucker condition for solving linear bilevel programming. Applied Mathematics and Computation, 188(1):808–813, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014196.

Lei:2005:NEP

[LHXL05] Aizhong Lei, DaiPing Hu, WeiGuo Xu, and Qian Shu Li. Noise enhanced the preponderant oscillation. Applied Mathematics and Computation, 171(2):824– 831, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lee:2005:NBS

[LHY05] Cheng-Chi Lee, Min-Shiang Hwang, and Wei-Pang Yang. A new blind signature based on the discrete logarithm problem for untraceability. Applied Mathematics and Computation, 164(3):837–841, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/S0096300304003509.

See comment [WW05c].

Li:2006:LRI

[LHZ06] Fan-Liang Li, Xi-Yan Hu, and Lei Zhang. Left and right inverse eigenpairs problem of skewcentrosymmetric matrices. Applied Mathematics and Computation, 177(1):105– 110, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008921.

Li:2009:LRI

[LHZ09] Fan-Liang Li, Xi-Yan Hu, and Lei Zhang. Left and right inverse eigenpairs problem of generalized centrosymmetric matrices and its optimal approximation problem. Applied Mathematics and Computation, 212(2):481–487, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001805.

Li:2005:AHP

[Li05a] Gong-Nong Li. Analysis for a homotopy path of complementarity problems based on μ-exceptional family. Applied Mathematics and Computation, 169(1):657–670, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400801X.

Li:2005:HOF

[Li05b] Jichun Li. High-order finite difference schemes for differential equations containing higher derivatives. Applied Mathematics and Computation, 171(2):1157– 1176, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2005:STS

[Li05c] Qinghong Li. On the stability of a two-step method for solving periodic IVPs. Applied Mathematics and Computation, 165(1):37– 44, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003601.

Li:2005:NFC

[Li05d] Steven Li. A new formula for computing implied volatility. Applied Mathematics and Computation, 170(1):611–625, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000032.

Li:2005:CES

[Li05e] Tseng-Kuei Li. Cycle embedding in star graphs with edge faults. Applied Mathematics and Computation, 167(2):891–900, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005375.

Li:2005:ESP

[Li05f] Xian-Fang Li. Erratum to: “Two perfectly-bonded dissimilar orthotropic strips with an interfacial crack normal to the boundaries” [Applied Mathematics and Computation 163 (2005) 961–975]. Applied Mathematics and Computation, 166(1):237–239, July 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002614.

See [Li05g].

Li:2005:TPB

[Li05g]  Xian-Fang Li.  Two

perfectly-bonded dissimilar orthotropic strips with an interfacial crack normal to the boundaries. Applied Mathematics and Computation, 163(2):961–975, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002735.

See erratum [Li05f].

Li:2006:DCC

[Li06a]  Gongsheng Li.  Data

compatibility and conditional stability for an inverse source problem in the heat equation. Applied Mathematics and Computation, 173(1):566–581, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003760.

Li:2006:ITK

[Li06b] Jian Li. Investigations on two kinds of two-level stabilized finite element methods for the stationary Navier–Stokes equations. Applied Mathematics and Computation, 182(2):1470–1481, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004747.

Li:2006:CP

[Li06c] Min Li. The charges on the paths. Applied Mathematics and Computation, 181(2):1163–1169, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306002013.

Li:2006:AAC

[Li06d]  Yong-Bin Li.  An al-

ternative algorithm for computing the pseudoremainder of multivariate polynomials. Applied Mathematics and Computation, 173(1):484–492, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305003681.

Li:2006:NAC

[Li06e]  Yong-Bin Li.  A new

approach for constructing subresultants. Applied Mathematics and Computation, 183(1):471–476, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005807.

Li:2007:PCP

[Li07a] Wei Li. Practical criteria for positive-definite matrix, M-matrix and Hurwitz matrix. Applied Mathematics and Computation, 185(1):397–401, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008733.

Li:2007:EAC

[Li07b] Yi Li. An effective algorithm of computing symbolic determinants with multivariate polynomial entries. Applied Mathematics and Computation, 192(2):382–388, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003402.

Li:2007:TDT

[Li07c]  Yiming Li.  A two-

dimensional thin-film transistor simulation using adaptive computing technique. Applied Mathematics and Computation, 184(1):73–85, January 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004863.

Li:2007:PPS

[Li07d] Yongkun Li. Positive periodic solutions of discrete Lotka–Volterra competition systems with state dependent and distributed delays. Applied Mathematics and Computation, 190(1):526–531, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000859.

Li:2007:EPS

[Li07e] Yongxiang Li. On the existence of positive solutions for the bending elastic beam equations. Applied Mathematics and Computation, 189(1):821–827, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016602.

Li:2008:KMF

[Li08a] Chong-Jun Li. A kind of multistep finite difference methods for arbitrary order linear boundary value problems. Applied Mathematics and Computation, 196(2):858–865, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307007394.

Li:2008:PBV

[Li08b] Jianli Li. Periodic boundary value problems for second-order impulsive integrodifferential equations. Applied Mathematics and Computation, 198(1):317– 325, April 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307008910.

Li:2008:PSN

[Li08c] Weiguo Li. Periodic solution to nonlinear differential equation at resonance. Applied Mathematics and Computation, 203(2):698–702, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003949.

Li:2008:MPSa

[Li08d]  Xiangfeng Li.  Multi-

ple positive solutions for some four-point boundary value problems with p-Laplacian. Applied Mathematics and Computation, 202(1):413–426, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000908.

Li:2008:OPS

[Li08e] Xiezhang Li. The optimal parameter of SOR- k method for p-cyclic matrices. Applied Mathematics and Computation, 197 (2):614–621, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307008181.

Li:2009:BSN

[Li09a] Fushan Li. Backward solutions to Neumann and Dirichlet problems of heatconduction equation. Applied Mathematics and Computation, 210(1):211– 214, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009958.

Li:2009:GHO

[Li09b]  Songxiao Li.  General-

ized Hilbert operator on the Dirichlet-type space. Applied Mathematics and Computation, 214(1):304– 309, August 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309003051.

Li:2009:EGE

[Li09c]  Xiaodi Li.  Existence

and global exponential stability of periodic solution for impulsive Cohen– Grossberg-type BAM neural networks with continuously distributed delays. Applied Mathematics and Computation, 215(1):292–307, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004603.

Li:2009:EHA

[Li09d] Yi Li. An effective hybrid algorithm for computing symbolic determinants. Applied Mathematics and Computation, 215(7):2495–2501, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007851.

Li:2009:SBE

[Li09e] Yiming Li. A simulationbased evolutionary approach to LNA circuit design optimization. Applied Mathematics and Computation, 209(1):57– 67, March 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004748.

Li:2009:CAU

[Li09f] Zhitao Li. Convergence analysis of an upwind mixed element method for advection diffusion problems. Applied Mathematics and Computation, 212(2):318–326, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001477.

Li:2009:NEK

[Li09g]  Zitian Li.  New exact

kink solutions, solitons and periodic form solutions for a combined KdV and Schwarzian KdV equation. Applied Mathematics and Computation, 215(8):2886–2890, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008297.

See comments [WDLM10].

Liao:2005:CBH

[Lia05] Shijun Liao. Comparison between the homotopy analysis method and homotopy perturbation method. Applied Mathematics and Computation, 169(2):1186–1194, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008446.

Liao:2008:IFO

[Lia08] Wenyuan Liao. An implicit fourth-order compact finite difference scheme for one-dimensional Burgers’ equation. Applied Mathematics and Computation, 206(2):755–764, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007042.

Liebrock:2006:PSW

[Lie06] Lorie M. Liebrock. Problem solving with well posedness analysis: Part

1. Applied Mathematics and Computation, 181 (2):1671–1682, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002980.

Lin:2005:MST

[Lin05a] Fuyong Lin. Multi-solution theory of signal processing based on finite element method. Applied Mathematics and Computation, 161(1):293–310, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012529.

Lin:2005:NMR

[Lin05b] Fuyong Lin. A new multiresolution theory based on orthogonal spline. Applied Mathematics and Computation, 168(2):1020–1030, September 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304006599.

Lin:2005:OBB

[Lin05c] Fuyong Lin. Orthogonal bases of 3-B-spline and its application in bending problem of plate and beam system. Applied Mathematics and Computation, 162(2):723–733, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001274.

Lin:2005:IRG

[Lin05d]  Yi-Qin Lin.  Implic-

itly restarted global FOM and GMRES for nonsymmetric matrix equations and Sylvester equations. Applied Mathematics and Computation, 167(2):1004–1025, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005466.

Ling:2005:MQI

[Lin05e]  Leevan Ling.  Multi-

variate quasi-interpolation schemes for dimensionsplitting multiquadric. Applied Mathematics and Computation, 161(1):195– 209, February 4, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300303012207.

Linss:2005:CDP

[Lin05f]  Torsten Linß.  On a

convection–diffusion problem with a weak layer. Applied Mathematics and Computation, 160(3):791– 795, January 27, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300303011950.

Lin:2006:MMP

[Lin06a] Chi-Ming Lin. Multicriteria– multistage planning for the optimal path selection using hybrid genetic algorithms. Applied Mathematics and Computation, 180(2):549–558, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306000439.

Lin:2006:MFC

[Lin06b]  Chi-Ming Lin.  Multi-

objective fuzzy competence set expansion problem by multistage decisionbased hybrid genetic algorithms. Applied Mathematics and Computation, 181(2):1402–1416, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306002591.

Lin:2006:TCM

[Lin06c]  Yi-Kuei Lin.  A two-

commodity multistate flow network with capacity weight varying with edges, nodes and types of commodity. Applied Mathematics and Computation, 183(1):142–151, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005479.

Lin:2006:MRM

[Lin06d] Yiqin Lin. Minimal residual methods augmented with eigenvectors for solving Sylvester equations and generalized Sylvester equations. Applied Mathematics and Computation, 181(1):487–499, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001184.

Lin:2007:PSV

[Lin07a]  Guojian Lin.  Periodic

solutions for van der Pol equation with time delay. Applied Mathematics and Computation, 187(2):1187–1198, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012550.

Lin:2007:EAS

[Lin07b] S.-S. Lin. An efficient algorithm for state estimation problems with coupling inequality constraints. Applied Mathematics and Computation, 194(2):492–501, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005000.

Lin:2007:ASG

[Lin07c] Shenggui Zhang Huiling Lin. Algorithms for symmetric groups of simplexes. Applied Mathematics and Computation, 188(2):1610–1634, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015141.

Lin:2007:EGT

[Lin07d] Shieh-Shing Lin. An efficient graph technique based dual-type algorithm for NMNF problems with large capacity constraints. Applied Mathematics and Computation, 190(1):309– 320, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000513.

Lin:2007:MBA

[Lin07e] Yi-Kuei Lin. MC-based Algorithm for a telecommunication network under node and budget constraints. Applied Mathematics and Computation, 190(2):1540–1550, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001907.

Lin:2007:SPT

[Lin07f]  Yi-Kuei Lin.  System

performance of a twocommodity stochastic-flow network with cost attributes in terms of minimal cuts. Applied Mathematics and Computation, 187(2):968–976, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012331.

Lin:2008:TIN

[Lin08] Ming-Han Lin. Thermal instability in natural convection flow over a rotating convex surface subject to external magnetic field. Applied Mathematics and Computation, 198(2):511– 525, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009174.

Lin:2009:TWN

[Lin09a]  Guojian Lin.  Travel-

ling waves in the Nicholson’s blowflies equation with spatio-temporal delay. Applied Mathematics and Computation, 209(2):314–326, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009673.

Lin:2009:DFR

[Lin09b] Hongwei Lin. On the derivative formula of a rational Bézier curve at a corner. Applied Mathematics and Computation, 210 (1):197–201, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300308009934.

Lisena:2006:CEP

[Lis06] Benedetta Lisena. Competitive exclusion in a periodic Lotka–Volterra system. Applied Mathematics and Computation, 177(2):761–768, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009574.

Liu:2005:DDN

[Liu05a] Jijun Liu. Determination of Dirichlet-to-Neumann map for a mixed boundary problem. Applied Mathematics and Computation, 161(3):843–864, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304000359.

Liu:2005:ULI

[Liu05b] JiJun Liu. On uniqueness and linearization of an inverse electromagnetic scattering problem. Applied Mathematics and Computation, 171(1):406–419, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001050.

Liu:2006:AFA

[Liu06a] Chun-Chu Liu. An analysis of the factors affecting change in an industry’s profit: The example of the Taiwan cement industry. Applied Mathematics and Computation, 183(1):695–703, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006035.

Liu:2006:NSN

[Liu06b] Kuo-Chi Liu. Numerical simulation for non-linear thermal wave. Applied Mathematics and Computation, 175(2):1385–1399, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006855.

Liu:2006:CMP

[Liu06c] Shiang-Tai Liu. Computational method for the profit bounds of inventory model with interval demand and unit cost. Applied Mathematics and Computation, 183(1):499–507, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005844.

Liu:2006:FTT

[Liu06d] Shiang-Tai Liu. Fuzzy total transportation cost measures for fuzzy solid transportation problem. Applied Mathematics and Computation, 174(2):927–941, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005229.

Liu:2006:GPA

[Liu06e] Shiang-Tai Liu. A geometric programming approach to profit maximization. Applied Mathematics and Computation, 182(2):1093–1097, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004243.

Liu:2006:HNM

[Liu06f]  X.-M. Liu.  Hybrid

numerical methods for exponential models of growth. Applied Mathematics and Computation, 179(2):772–778, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010878.

Liu:2006:FMS

[Liu06g] Xueming Liu. Fast methods for solving nonlinearly coupled equations describing fiber amplifiers. Applied Mathematics and Computation, 172(2):1018–1035, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002754.

Liu:2006:SQP

[Liu06h] Yuji Liu. A study on quasiperiodic boundary value problems for nonlinear higher order impulsive differential equations. Applied Mathematics and Computation, 183(2):842–857, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006187.

Liu:2007:EMK

[Liu07a] Chun-Chu Liu. An extended method for key factors in reducing CO 2 emissions. Applied Mathematics and Computation, 189(1):440–451, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016171.

Liu:2007:NLR

[Liu07b]  Huaning Liu.  A note

on local randomness in polynomial random number and random function generators. Applied Mathematics and Computation, 186(2):1360–1366, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010605.

Liu:2007:FPC

[Liu07c] Jin-Lin Liu. Further properties of a certain subclass of analytic and multivalent functions. Applied Mathematics and Computation, 187(1):290–294, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011635.

Liu:2007:CMM

[Liu07d] Shiang-Tai Liu. A computational method for the maximization of long-run and short-run profit. Applied Mathematics and Computation, 186(2):1104– 1112, March 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306010368.

Liu:2007:PMQ

[Liu07e] Shiang-Tai Liu. Profit maximization with quantity discount: an application of geometric programming. Applied Mathematics and Computation, 190(2):1723–1729, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002184.

Liu:2008:CRA

[Liu08a] Yan-Kui Liu. The convergent results about approximating fuzzy random minimum risk problems. Applied Mathematics and Computation, 205(2):608–621, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003391.

Liu:2008:PSM

[Liu08b] Yuji Liu. Positive solutions of mixed type multipoint non-homogeneous BVPs for p-Laplacian equations. Applied Mathematics and Computation, 206(2):796–805, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308007108.

Liu:2009:LEA

[Liu09a]  Guan-Ting Liu.  Lame

equation and asymptotic higher-order periodic solutions to nonlinear evolution equations. Applied Mathematics and Computation, 212(2):312–317, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001465.

Liu:2009:SPC

[Liu09b] Jin-Lin Liu. Some properties of certain meromorphically multivalent functions. Applied Mathematics and Computation, 210(1):136– 140, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009818.

Liu:2009:LAS

[Liu09c] Lingxia Liu. Local analytic solutions of a functional differential equation αz + βx′(z) = x(az + bx′(z)). Applied Mathematics and Computation, 215(2):644–652, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900544X.

Liu:2009:UGP

[Liu09d] Shiang-Tai Liu. Using geometric programming to profit maximization with interval coefficients and quantity discount. Applied Mathematics and Computation, 209(2):259–265, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009430.

Li:2007:SAE

[LJ07] Guoyuan Li and C. Rhett Jackson. Simple, accurate, and efficient revisions to MacCormack and Saulyev schemes: High Peclet numbers. Applied Mathematics and Computation, 186(1):610–622, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009726.

Lee:2008:TCM

[LJ08] Sunmi Lee and Eunok Jung. A two-chamber model of valveless pumping using the immersed boundary method. Applied Mathematics and Computation, 206(2):876–884, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308007145.

Lotfi:2007:AAE

[LJE07a] F. Hosseinzadeh Lotfi, G. R. Jahanshahloo, and M. Esmaeili. An alternative approach in the estimation of returns to scale under weight restrictions. Applied Mathematics and Computation, 189(1):719– 724, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016481.

Lotfi:2007:SAE

[LJE07b] F. Hosseinzadeh Lotfi, G. R. Jahanshahloo, and M. Esmaeili. Sensitivity analysis of efficient units in the presence of non-discretionary inputs. Applied Mathematics and Computation, 190(2):1185–1197, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700149X.

Lian:2006:SPSb

[LJG06] Zhigang Lian, Bin Jiao, and Xingsheng Gu. A similar particle swarm optimization algorithm for jobshop scheduling to minimize makespan. Applied Mathematics and Computation, 183(2):1008–1017, December 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306006369.

Li:2007:SEM

[LJK07] Shanling Li, G. R. Jahanshahloo, and M. Khodabakhshi. A superefficiency model for ranking efficient units in data envelopment analysis. Applied Mathematics and Computation, 184(2):638–648, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007387.

Liaw:2006:EOB

[LJL06] Horng-Twu Liaw, Wen-Shenq Juang, and Chi-Kai Lin. An electronic online bidding auction protocol with both security and efficiency. Applied Mathematics and Computation, 174(2):1487–1497, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005758.

Lou:2008:VAM

[LjZH08] Jian Lou, Li juan Zhang, and Zhen He. Viscosity approximation methods for asymptotically nonexpansive mappings. Applied Mathematics and Computation, 203(1):171–177, September 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300308002245.

Li:2008:OPG

[LK08] Jicheng Li and Xu Kong. Optimal parameters of GSOR-like methods for solving the augmented linear systems. Applied Mathematics and Computation, 204(1):150–161, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004864.

Lin:2005:DAC

[LKC05] Wenbin Lin, Narayan Kovvali, and Lawrence Carin. Direct algorithm for computation of derivatives of the Daubechies basis functions. Applied Mathematics and Computation, 170(2):1006–1013, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000421.

Leontitsis:2006:RSO

[LKP06] Alexandros Leontitsis, Dimitris Kontogiorgos, and Jenny Pagge. Repel the swarm to the optimum! Applied Mathematics and Computation, 173(1):265–272, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003425.

Liu:2007:EIA

[LKU07] Zeqing Liu, Shin Min Kang, and Jeong Sheok Ume. Existence and iterative approximations of nonoscillatory solutions of higher order nonlinear neutral delay differential equations. Applied Mathematics and Computation, 193(1):73–88, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700375X.

Liu:2009:EUM

[LKU09] Zeqing Liu, Shin Min Kang, and Jeong Sheok Ume. Existence of uncountably many bounded nonoscillatory solutions and their iterative approximations for second order nonlinear neutral delay difference equations. Applied Mathematics and Computation, 213(2):554–576, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002690.

Leung:2006:PCC

[LKW06] K. S. Leung, I. King, and Y. B. Wong. A probabilistic cooperative– competitive hierarchical model for global optimization. Applied Mathematics and Computation, 175(2):1092–1124, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006648.

Lee:2005:ENB

[LKY05a] Sung-Woon Lee, Hyun-Sung Kim, and Kee-Young Yoo. Efficient noncebased remote user authentication scheme using smart cards. Applied Mathematics and Computation, 167(1):355–361, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004977.

Lee:2005:EVB

[LKY05b] Sung-Woon Lee, Hyun-Sung Kim, and Kee-Young Yoo. Efficient verifier-based key agreement protocol for three parties without server’s public key. Applied Mathematics and Computation, 167(2):996–1003, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005454.

Lee:2005:IHA

[LKY05c] Sung-Woon Lee, Hyun-Sung Kim, and Kee-Young Yoo. Improvement of HWWM-authenticated key agreement protocol. Applied Mathematics and Computation, 162(3):1315– 1320, March 25, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304001894.

Lee:2005:ILL

[LKY05d] Sung-Woon Lee, Hyun-Sung Kim, and Kee-Young Yoo. Improvement of Lee and Lee’s authenticated key agreement scheme. Applied Mathematics and Computation, 162(3):1049–1053, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001481.

Lai:2005:FDS

[LL05a] Ming-Chih Lai and Hsi-Chi Liu. Fast direct solver for the biharmonic equation on a disk and its application to incompressible flows. Applied Mathematics and Computation, 164(3):679–695, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003169.

Lee:2005:CMA

[LL05b] Keon-Jik Lee and Byeong-Jik Lee. Cryptanalysis of the modified authenticated key agreement scheme. Applied Mathematics and Computation, 170(1):280–284, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009257.

Lee:2005:SSP

[LL05c] Narn-Yih Lee and Ming-Feng Lee. The security of a strong proxy signature scheme with proxy signer privacy protection. Applied Mathematics and Computation, 161(3):807–812, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000323.

Li:2005:RKM

[LL05d] D. Li and M. Z. Liu.

Runge–Kutta methods for the multi-pantograph delay equation. Applied Mathematics and Computation, 163(1):383–395, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002322.

Li:2005:AAS

[LL05e] Shuicai Li and Shi-Jun Liao. An analytic approach to solve multiple solutions of a strongly nonlinear problem. Applied Mathematics and Computation, 169(2):854–865, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008161.

Li:2005:PPS

[LL05f] Yongkun Li and Linghong Lu. Positive periodic solutions of discrete nspecies food-chain systems. Applied Mathematics and Computation, 167(1):324–344, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004953.

Ling:2005:CSS

[LL05g] Yi Ling and Fengshan Liu. The Choi–Saigo– Srivastava integral operator and a class of analytic functions. Applied Mathematics and Computation, 165(3):613–621, June 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004060.

Luo:2005:FCT

[LL05h] Jinwen Luo and Yijing Liu.

Fast computation of the target signal phase in noisy environments. Applied Mathematics and Computation, 166(1):224–232, July 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004989.

Liang:2006:CSL

[LL06a] Yew-Wen Liang and Der-Cherng Liaw. Common stabilizers for linear control systems in the presence of actuators outage. Applied Mathematics and Computation, 177(2):635–643, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009446.

Liang:2006:ESE

[LL06b] Yong Liang and Kwong-Sak Leung. Evolution strategies with exclusionbased selection operators and a Fourier series auxiliary function. Applied Mathematics and Computation, 174(2):1080–1109, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005345.

Liu:2006:FQA

[LL06c] Linzhong Liu and Yinzhen Li. The fuzzy quadratic assignment problem with penalty: New models and genetic algorithm. Applied Mathematics and Computation, 174(2):1229–1244, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005448.

Liu:2007:SCT

[LL07a] Yicheng Liu and Zhixiang Li. Strong convergence theorems in real reflexive Banach spaces for non-expansive mappings. Applied Mathematics and Computation, 193(2):438–445, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004171.

Llanas:2007:HIN

[LL07b] B. Llanas and S. Lantarón. Hermite interpolation by neural networks. Applied Mathematics and Computation, 191(2):429–439, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002706.

Lan:2008:CHB

[LL08a] Yaoyao Lan and Qing-Guo Li. Control of Hopf bifurcation in a simple plankton population model with a non-integer exponent of closure. Applied Mathematics and Computation, 200(1):220–230, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011010.

Li:2008:RAN

[LL08b] Pingkang Li and Kang Li.

A recursive algorithm for nonlinear model identification. Applied Mathematics and Computation, 205(2):511–516, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003299.

Li:2008:NTR

[LL08c] Shujun Li and Zhenhai Liu. A new trust region filter algorithm. Applied Mathematics and Computation, 204(1):485–489, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005365.

Li:2008:TWN

[LL08d] Xue-Shi Li and Guo Lin. Traveling wavefronts in nonlocal dispersal and cooperative Lotka– Volterra system with delays. Applied Mathematics and Computation, 204(2):738–744, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005481.

Li:2008:PFE

[LL08e] Yuan Li and Kaitai Li. Penalty finite element method for Stokes problem with nonlinear slip boundary conditions. Applied Mathematics and Computation, 204(1):216–226, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004979.

Liu:2008:LAS

[LL08f] Tongbo Liu and Hong Li. Local analytic solution of a second-order functional differential equation with a state derivative dependent delay. Applied Mathematics and Computation, 197(1):158–166, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007643.

Lv:2008:NSN

[LL08g] Xiao-Guang Lv and Jiang Le. A note on solving nearly penta-diagonal linear systems. Applied Mathematics and Computation, 204(2):707–712, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005432.

Liu:2009:GMF

[LL09] Yang Liu and Hong Li.

H^{1}-Galerkin mixed finite element methods for pseudo-hyperbolic equations. Applied Mathematics and Computation, 212(2):446–457, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001763.

See corrigendum [LL12].

Liu:2012:CGM

[LL12] Yang Liu and Hong Li.

Corrigendum to “H^{1}- Galerkin mixed finite element methods for pseudohyperbolic equations” [Appl. Math. Comput. 212 (2009) 446–457]. Applied Mathematics and Computation, 218(19):10008, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300312003268. See

[LL09].

Lin:2009:SCN

[LLAT09] R. Lin, F. Liu, V. Anh, and I. Turner. Stability and convergence of a new explicit finitedifference approximation for the variable-order nonlinear fractional diffusion equation. Applied Mathematics and Computation, 212(2):435–445, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001751.

Lu:2006:FBW

[LLC06a] Wei Lu, Hongtao Lu, and Fu-Lai Chung. Feature based watermarking using watermark template match. Applied Mathematics and Computation, 177(1):377–386, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009185.

Lu:2006:RDI

[LLC06b] Wei Lu, Hongtao Lu, and Fu-Lai Chung. Robust digital image watermarking based on subsampling. Applied Mathematics and Computation, 181(2):886–893, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001688.

Li:2009:SSD

[LLC09a] Shengguo Li, Housen Li, and Lizhi Cheng. Some second-derivative-free variants of Halley’s method for multiple roots. Applied Mathematics and Computation, 215(6):2192–2198, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007280.

Liu:2009:SDM

[LLC09b] Huan-Wen Liu, Li-Bin Liu, and Yanping Chen. A semi-discretization method based on quartic splines for solving one-space-dimensional hyperbolic equations. Applied Mathematics and Computation, 210(2):508–514, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000307.

Li:2008:NGR

[LLFM08] Gongsheng Li, Jinqing Liu, Xiaoping Fan, and Yu Ma. A new gradient regularization algorithm for source term inversion in 1D solute transportation with final observations. Applied Mathematics and Computation, 196(2):646–660, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007187.

Li:2006:NPA

[LLG06] L-L. Li, F. Li, and F. B. Gross. A new polynomial approximation for J_{ν} Bessel functions. Applied Mathematics and Computation, 183(2):1220–1225, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006679.

Li:2009:SBS

[LLG09] Xue-Zhi Li, Wen-Sheng Li, and Mini Ghosh. Stability and bifurcation of an SIR epidemic model with nonlinear incidence and treatment. Applied Mathematics and Computation, 210(1):141–150, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009831.

Liu:2009:LGC

[LLL09] Hanze Liu, Jibin Li, and Lei Liu. Lie group classifications and exact solutions for two variable-coefficient equations. Applied Mathematics and Computation, 215(8):2927–2935, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008649.

Lin:2006:MPS

[LLLW06] Xiaoyan Lin, Andong Liu, Yude Li, and Pengju Wu. A MATLAB programming for simulation of X-ray capillaries. Applied Mathematics and Computation, 172(1):188–197, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305001979.

Li:2007:GBN

[LLLX07] Shujuan Li, Yan Li, Yong Liu, and Yuefei Xu. A GA-based NN approach for makespan estimation. Applied Mathematics and Computation, 185(2):1003–1014, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008253.

Liu:2005:NSH

[LLW05] Kuo-Chi Liu, Chien-Nan Lin, and Jiun-Shen Wang. Numerical solutions for the hyperbolic heat conduction problems in a layered solid cylinder with radiation surface. Applied Mathematics and Computation, 164(3):805–820, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003479.

Liu:2007:PSS

[LLW07] Lishan Liu, Bingmei Liu, and Yonghong Wu. Positive solutions of singular boundary value problems for nonlinear differential systems. Applied Mathematics and Computation, 186(2):1163–1172, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010423.

Liu:2008:PSS

[LLW08] Bingmei Liu, Lishan Liu, and Yonghong Wu. Positive solutions for a singular second-order threepoint boundary value problem. Applied Mathematics and Computation, 196(2): 532–541, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307007059.

Liu:2009:PSS

[LLW09] Weiwei Liu, Lishan Liu, and Yonghong Wu. Positive solutions of a singular boundary value problem for systems of secondorder differential equations. Applied Mathematics and Computation, 208(2):511–519, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009247.

Lou:2007:EME

[LLWC07] Der-Chyuan Lou, Jui-Chang Lai, Chia-Long Wu, and Te-Jen Chang. An efficient Montgomery exponentiation algorithm by using signed-digit-recoding and folding techniques. Applied Mathematics and Computation, 185(1):31–44, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008010.

Li:2006:EHP

[lLWhL06] Ling lai Li, Ling Wang, and Li heng Liu. An effective hybrid PSOSA strategy for optimization and its application to parameter estimation. Applied Mathematics and Computation, 179(1):135–146, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010155.

Liang:2005:HSP

[LLX05] Jin Liang, James Liu, and Ti-Jun Xiao. Hyperbolic singular perturbations for integrodifferential equations. Applied Mathematics and Computation, 163(2):609–620, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002462.

Li:2007:ESS

[LLX07a] Chaolei Li, Jia Liu, and Shanhong Xia. English sentence stress detection system based on HMM framework. Applied Mathematics and Computation, 185(2):759–768, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007545.

Liang:2007:NSD

[LLX07b] Yong Liang, Kwong-Sak Leung, and Zong-Ben Xu. A novel splicing/ decomposable binary encoding and its operators for genetic and evolutionary algorithms. Applied Mathematics and Computation, 190(1):887–904, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001208.

Liu:2006:MFW

[LLY06] Linzhong Liu, Yinzhen Li, and Lixing Yang. The maximum fuzzy weighted matching models and hybrid genetic algorithm. Applied Mathematics and Computation, 181(1):662– 674, October 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306001469.

Liu:2007:SRK

[LLY07] M. Z. Liu, Hui Liang, and Z. W. Yang. Stability of Runge–Kutta methods in the numerical solution of linear impulsive differential equations. Applied Mathematics and Computation, 192(2):346–357, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003888.

Li:2006:EPS

[LLZ06] Xiaoping Li, Yuehua Liu, and Tiejun Zhou. The existence of positive solution and bounded oscillation for even-order neutral differential equations with unstable type. Applied Mathematics and Computation, 176(2):632–641, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008611.

Liu:2008:DDR

[LLZ08] Duyu Liu, Xinzhi Liu, and Shouming Zhong. Delaydependent robust stability and control synthesis for uncertain switched neutral systems with mixed delays. Applied Mathematics and Computation, 202(2):828–839, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001872.

Liu:2009:NND

[LLZ09] Bingchen Liu, Fengjie Li, and Sining Zheng. Note on a nonlinear diffusion system with convection. Applied Mathematics and Computation, 214(1):41–47, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002768.

Long:2007:FMS

[LM07a] Jun Long and Changfeng Ma. A filter method for solving nonlinear complementarity problems based on derivative-free line search. Applied Mathematics and Computation, 190 (1):271–286, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300307000471.

Lu:2007:OTS

[LM07b] Fang Lu and Fanwei Meng. Oscillation theorems for superlinear second-order damped differential equations. Applied Mathematics and Computation, 189 (1):796–804, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306016560.

Lee:2008:CBN

[LM08a] Yeunghak Lee and David Marshall. Curvature based normalized 3D component facial image recognition using fuzzy integral. Applied Mathematics and Computation, 205(2):815–823, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003603.

Li:2008:BSI

[LM08b] Lei Li and Jinwen Ma. A BYY scale-incremental EM algorithm for Gaussian mixture learning. Applied Mathematics and Computation, 205(2):832–840, November 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308003627.

Liu:2008:IOC

[LM08c] Haidong Liu and Fanwei Meng. Interval oscillation criteria for matrix differential systems with damping. Applied Mathematics and Computation, 201(1–2):305–312, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307012064.

Llibre:2009:BLC

[LM09] Jaume Llibre and Amar Makhlouf. Bifurcation of limit cycles from a 4dimensional center in 1: n resonance. Applied Mathematics and Computation, 215(1):140–146, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004123.

Liu:2008:NSS

[LMC08] Kaiyuan Liu, Xinzhu Meng, and Lansun Chen. A new stage structured predator–prey Gomportz model with time delay and impulsive perturbations on the prey. Applied Mathematics and Computation, 196(2):705–719, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007230.

Lu:2008:NAI

[LMD^{+}08] Yunfeng Lu, Jun Miao, Lijuan Duan, Yuanhua Qiao, and Ruixin Jia. A new approach to image segmentation based on simplified region growing PCNN. Applied Mathematics and Computation, 205(2):807–814, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003597.

Li:2006:PPS

[LMH06] Jian Li, Liquan Mei, and Yinnian He. A pressure-Poisson stabilized finite element method for the nonstationary Stokes equations to circumvent the inf–sup condition. Applied Mathematics and Computation, 182(1):24–35, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003080.

Li:2008:OPD

[LML08] Wenyan Li, Yongkui Man, and Guoliang Li. Optimal parameter design of input filters for general purpose inverter based on genetic algorithm. Applied Mathematics and Computation, 205(2):697–705, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003482.

Long:2007:NFM

[LMN07] Jun Long, Changfeng Ma, and Puyan Nie. A new filter method for solving nonlinear complementarity problems. Applied Mathematics and Computation, 185(1):705–718, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009064.

Laskari:2007:AEC

[LMSV07] E. C. Laskari, G. C.

Meletiou, Y. C. Stamatiou, and M. N. Vrahatis. Applying evolutionary computation methods for the cryptanalysis of Feistel ciphers. Applied Mathematics and Computation, 184(1):63–72, January 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004851.

Leelawattanachai:2009:MGA

[LMT^{+}09] Jeerapond Leelawattanachai, Charin Modchang, Wannapong Triampo, Darapond Triampo, and Yongwimon Lenbury. Modeling and genetic algorithm optimization of early events in signal transduction via dynamics of G-protein-coupled receptors: Internalization consideration. Applied Mathematics and Computation, 207(2):528–544, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008412.

Laskari:2009:ANI

[LMV09] E. C. Laskari, G. C. Meletiou, and M. N. Vrahatis. Aitken and Neville inverse interpolation methods for the Lucas logarithm problem. Applied Mathematics and Computation, 209(1):52– 56, March 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004736.

Liu:2007:SAR

[LMY07] M. Z. Liu, S. F. Ma, and Z. W. Yang. Stability analysis of Runge–Kutta methods for unbounded retarded differential equations with piecewise continuous arguments. Applied Mathematics and Computation, 191(1):57–66, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601705X.

Lee:2008:JTE

[LMZ08a] M. H. Lee, N. L. Manev, and Xiao-Dong Zhang. Jacket transform eigenvalue decomposition. Applied Mathematics and Computation, 198(2):858– 864, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009599.

Li:2008:BSS

[LmZ08b] Ping Li and Shou ming Zhong. BIBO stabilization for system with multiple mixed delays and nonlinear perturbations. Applied Mathematics and Computation, 196(1):207–213, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006601.

Li:2008:BST

[LmZ08c] Ping Li and Shou ming Zhong. BIBO stabilization of time-delayed system with nonlinear perturbation. Applied Mathematics and Computation, 195(1):264–269, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005371.

Li:2009:SNO

[LMZ09] Lianzhong Li, Fanwei Meng, and Zhaowen Zheng. Some new oscillation results for linear Hamiltonian systems. Applied Mathematics and Computation, 208(1):219–224, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008928.

Lu:2006:EPN

[LN06] Linzhang Lu and Michael K. Ng. Effects of a parameter on a nonsymmetric algebraic Riccati equation. Applied Mathematics and Computation, 172(2):753–761, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009099.

Liu:2007:SCA

[LN07] Jin-Lin Liu and Khalida Inayat Noor. On subordinations for certain analytic functions associated with Noor integral operator. Applied Mathematics and Computation, 187(2):1453–1460, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012823.

Liu:2008:GOI

[LN08a] Wenjun Liu and Quõc-Anh Ngô. A generalization of Ostrowski inequality on time scales for k points. Applied Mathematics and Computation, 203(2):754–760, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003998.

Lu:2008:QNT

[LN08b] Xiaoping Lu and Qin Ni. A quasi-Newton trust region method with a new conic model for the unconstrained optimization. Applied Mathematics and Computation, 204(1):373–384, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800516X.

Lamiri:2008:OHT

[LO08] I. Lamiri and A. Ouni. d-Orthogonality of Hermite type polynomials. Applied Mathematics and Computation, 202(1):24–43, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011563.

Lu:2005:PSS

[LOA05] Haishen Lü, Donal O’Regan, and Ravi P. Agarwal. Positive solutions for singular p-Laplacian equations with sign changing nonlinearities using inequality theory. Applied Mathematics and Computation, 165(3):587–597, June 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004047.

Lou:2007:BTW

[Lou07] Yijun Lou. Bifurcation of travelling wave solutions in generalized phifour equation. Applied Mathematics and Computation, 190(1):517–525, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000835.

Liao:2007:PSE

[LOZ07] Xinyuan Liao, Zigen Ouyang, and Shengfan Zhou. Permanence and stability of equilibrium for a twoprey one-predator discrete model. Applied Mathematics and Computation, 186(1):93–100, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009209.

Lizama:2006:SPV

[LP06a] Carlos Lizama and Humberto Prado. Singular perturbation for Volterra equations of convolution type. Applied Mathematics and Computation, 181(2):1624–1634, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002943.

Lizama:2006:SPI

[LP06b] Carlos Lizama and Humberto Prado. Singular perturbations of integrodifferential equations. Applied Mathematics and Computation, 175(2):1582– 1595, April 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305007162.

Landi:2007:FMM

[LP07a] G. Landi and E. Loli Piccolomini. Fast methods for MR spectroscopic imaging. Applied Mathematics and Computation, 191(2):389–396, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002664.

Lee:2007:OFM

[LP07b] S. M. Lee and Ju H. Park. Output feedback model predictive control for LPV systems using parameter-dependent Lyapunov function. Applied Mathematics and Computation, 190(1):671–676, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001026.

Lubuma:2007:NSM

[LP07c] Jean M.-S. Lubuma and Kailash C. Patidar. Nonstandard methods for singularly perturbed problems possessing oscillatory/layer solutions. Applied Mathematics and Computation, 187(2):1147–1160, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012525.

Lee:2008:RSD

[LP08a] S. M. Lee and Ju H.

Park. Robust stabilization of discrete-time nonlinear Lur’e systems with sector and slope restricted nonlinearities. Applied Mathematics and Computation, 200(1):429–436, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011484.

Liu:2008:CPM

[LP08b] Jin-Lin Liu and J. Patel. Certain properties of multivalent functions associated with an extended fractional differintegral operator. Applied Mathematics and Computation, 203(2):703–713, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003950.

Lozovanu:2008:CAM

[LP08c] Dmitrii Lozovanu and Stefan Pickl. A constructive algorithm for max– min paths problems on energy networks. Applied Mathematics and Computation, 204(2):602–608, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003081.

Lubuma:2009:TIS

[LP09] Jean M.-S. Lubuma and Kailash C. Patidar. Towards the implementation of the singular function method for singular perturbation problems. Applied Mathematics and Computation, 209(1):68– 74, March 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800475X.

Li:2009:HPT

[LPA^{+}09] Yaohang Li, Vladimir A. Protopopescu, Nikita Arnold, Xinyu Zhang, and Andrey Gorin. Hybrid parallel tempering and simulated annealing method. Applied Mathematics and Computation, 212(1):216– 228, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001362.

Lei:2007:MRA

[LpL07] Yuan Lei and An ping Liao.

A minimal residual algorithm for the inconsistent matrix equation AXB = C over symmetric matrices. Applied Mathematics and Computation, 188(1):499– 513, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013907.

Lee:2008:NSM

[LPP08] Moon Ho Lee, Subash Shree Pokhrel, and Dae Chul Park. A new sparse matrix analysis of discrete Fourier transform. Applied Mathematics and Computation, 205(1):482–486, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004001.

Liang:2008:PSG

[LPS08] Ruixi Liang, Jun Peng, and Jianhua Shen. Positive solutions to a generalized second order threepoint boundary value problem. Applied Mathematics and Computation, 196(2): 931–940, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307007461.

Lykina:2008:RAM

[LPW08] Valeriya Lykina, Sabine Pickenhain, and Marcus Wagner. On a resource allocation model with infinite horizon. Applied Mathematics and Computation, 204(2):595–601, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800307X.

Leonaite:2007:CAPa

[LR07a] Gintare Leonaite and Jonas Rimas. On computing of arbitrary positive integer powers for one type of odd order tridiagonal matrices with two zero rows and eigenvalues on imaginary axis-I. Applied Mathematics and Computation, 185(1):346–349, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008666.

Leonaite:2007:CAPb

[LR07b] Gintare Leonaite and Jonas Rimas. On computing of arbitrary positive integer powers for one type of odd order tridiagonal matrices with two zero rows. I. Applied Mathematics and Computation, 186(2):1537–1541, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010800.

Leyva-Ramos:2006:CDI

[LRMS06] J. Leyva-Ramos and J. A.

Morales-Saldaña. Computation of delay intervals for stability of timedelay systems. Applied Mathematics and Computation, 183(2):980–989, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006333.

Li:2005:DRD

[LS05] Wan-Tong Li and Hong-Rui Sun. Dynamics of a rational difference equation. Applied Mathematics and Computation, 163(2):577–591, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002449.

Lee:2006:ESC

[LS06a] S.-H. Lee and N.-Y. Su. Estimating the size of a closed population from a single mark–recapture regime by using a diffusion probability of marked individuals. Applied Mathematics and Computation, 182(1):403–411, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600347X.

Li:2006:MPS

[LS06b] Jianli Li and Jianhua Shen. Multiple positive solutions for a second-order threepoint boundary value problem. Applied Mathematics and Computation, 182(1):258–268, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003353.

Li:2006:PBV

[LS06c] Jianli Li and Jianhua Shen. Periodic boundary value problems for impulsive integro-differential equations of mixed type. Applied Mathematics and Computation, 183(2):890– 902, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306006230.

Liu:2006:NTA

[LS06d] Haihong Liu and Ning Su.

A note on a trivial assumption on kernel and model. Applied Mathematics and Computation, 180(1):374–375, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030501132X.

Liu:2006:RID

[LS06e] Hong Liu and Yongzhong Song. On a regularization of index 3 differential-algebraic equations. Applied Mathematics and Computation, 181(2):1369–1378, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002554.

laSen:2007:APC

[lS07a] M. De la Sen. About the positivity of a class of hybrid dynamic linear systems. Applied Mathematics and Computation, 189(1):852–868, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030601664X.

laSen:2007:SCL

[lS07b] M. De la Sen. Stability criteria for linear timeinvariant systems with point delays based on one-dimensional Routh– Hurwitz tests. Applied Mathematics and Computation, 187(2):1199–1207, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012562.

Liang:2007:PBV

[LS07c] Ruixi Liang and Jianhua Shen. Periodic boundary value problem for second-order impulsive functional differential equations. Applied Mathematics and Computation, 193(2):560–571, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004092.

Liu:2007:DTM

[LS07d] Hong Liu and Yongzhong Song. Differential transform method applied to high index differential– algebraic equations. Applied Mathematics and Computation, 184(2):748– 753, January 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306007521.

Liu:2007:ASC

[LS07e] Jian Liu and Jianguo Si.

Analytic solutions for a class of differential equation with delays depending on state. Applied Mathematics and Computation, 186(1):261–270, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009362.

laSen:2008:ACL

[lS08a] M. De la Sen. Adaptive control of linear singular time-invariant singleinput single-output systems with external point delay. Applied Mathematics and Computation, 203(1):319–332, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002567.

laSen:2008:SWO

[lS08b] M. De la Sen. Strong and weak oscillatory behavior in time-varying linear difference equations of arbitrary order in the presence of unmodeled dynamics and nominal parametrical errors. Applied Mathematics and Computation, 197(1):11–28, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007497.

Lee:2008:EPS

[LS08c] Sang-Hee Lee and Nan-Yao Su. Estimating the population size of subterranean termites by using a novel diffusionbased method. Applied Mathematics and Computation, 197(1):435–439, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700793X.

Li:2008:WCO

[LS08d] Songxiao Li and Stevo Stević. Weighted composition operators from Zygmund spaces into Bloch spaces. Applied Mathematics and Computation, 206(2):825–831, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007327.

Liang:2008:RBB

[LS08e] Y. S. Liang and W. Y. Su.

The relationship between the Box dimension of the Besicovitch functions and the orders of their fractional calculus. Applied Mathematics and Computation, 200(1):297–307, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011095.

Labbe:2009:EDP

[LS09a] Chantal Labbé and Kristina P.

Sendova. The expected discounted penalty function under a risk model with stochastic income. Applied Mathematics and Computation, 215(5):1852–1867, November 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300309006869.

Layton:2009:COA

[LS09b] William Layton and Iuliana Stanculescu. Chebychev optimized approximate deconvolution models of turbulence. Applied Mathematics and Computation, 208(1):106–118, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008527.

Li:2009:PCM

[LS09c] Min Li and Yuxiang Sun. Practical criteria for H-matrices. Applied Mathematics and Computation, 211(2):427–433, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000903.

Li:2009:CTO

[LS09d] Songxiao Li and Stevo Stević. Cesàro-type operators on some spaces of analytic functions on the unit ball. Applied Mathematics and Computation, 208(2):378–388, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009065.

Li:2009:ITOa

[LS09e] Songxiao Li and Stevo Stević. Integral-type operators from Bloch-type spaces to Zygmund-type spaces. Applied Mathematics and Computation, 215(2):464–473, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004871.

Li:2009:ITO

[LS09f] Songxiao Li and Stevo Stević. On an integraltype operator from iterated logarithmic Bloch spaces into Bloch-type spaces. Applied Mathematics and Computation, 215(8):3106–3115, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008972.

Li:2009:ITOb

[LS09g] Songxiao Li and Stevo Stević. On an integraltype operator from iterated logarithmic Bloch spaces into Bloch-type spaces. Applied Mathematics and Computation, 215(8):3106–3115, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309008972.

Li:2009:WHF

[LS09h] Songxiao Li and Stevo Stević. Weighted-Hardy functions with Hadamard gaps on the unit ball. Applied Mathematics and Computation, 212(1):229– 233, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001374.

laSen:2008:CTP

[lSAQ08] M. De la Sen and S. Alonso-Quesada. A Control Theory point of view on Beverton–Holt equation in population dynamics and some of its generalizations. Applied Mathematics and Computation, 199(2):464– 481, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010132.

Liang:2007:RPM

[LSBL07] Heng Liang, Linsong Shi, Fengshan Bai, and Xiaoyan Liu. Random path method with pivoting for computing permanents of matrices. Applied Mathematics and Computation, 185(1):59–71, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008034.

Shang:2007:FGM

[lSgPpJ07] You lin Shang, Ding guo Pu, and Ai ping Jiang. Finding global minimizer with one-parameter filled function on unconstrained global optimization. Applied Mathematics and Computation, 191(1):176–182, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700241X.

Li:2006:BTW

[LSL06] Hong Li, Shaorong Sun, and Jibin Li. Bifurcations of travelling wave solutions for the Boussinesq equations. Applied Mathematics and Computation, 175(1):61–71, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005916.

Long:2009:PBM

[LSN09] Guangqing Long, Mitali Madhumita Sahani, and Gnaneshwar Nelakanti. Polynomially based multiprojection methods for Fredholm integral equations of the second kind. Applied Mathematics and Computation, 215(1):147– 155, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300309004160.

Lopez:2009:MPT

[LST09] José L. López, Ester Pérez Sinusía, and Nico M. Temme. Multi-point Taylor approximations in onedimensional linear boundary value problems. Applied Mathematics and Computation, 207(2):519– 527, January 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300308008400.

Li:2006:MSP

[LSW06a] Qinghong Li, Yongzhong Song, and Yushun Wang. On multi-symplectic partitioned Runge–Kutta methods for Hamiltonian wave equations. Applied Mathematics and Computation, 177(1):36–43, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305008854.

Li:2006:MIB

[LSW06b] Qinghong Li, Yajuan Sun, and Yushun Wang. On multisymplectic integrators based on Runge– Kutta–Nyström methods for Hamiltonian wave equations. Applied Mathematics and Computation, 182(2):1056–1063, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004206.

Liao:2007:DMD

[LT07] Hung-Chang Liao and Ju-Hsin Tsai. Data mining for DNA viruses with breast cancer, fibroadenoma, and normal mammary tissue. Applied Mathematics and Computation, 188(1):989–1000, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014470.

Lagaris:2008:SRB

[LT08a] I. E. Lagaris and I. G.

Tsoulos. Stopping rules for box-constrained stochastic global optimization. Applied Mathematics and Computation, 197(2):622– 632, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008193.

Li:2008:NSM

[LT08b] Wei Li and Xiaoli Tian. Numerical solution method for general interval quadratic programming. Applied Mathematics and Computation, 202(2):589–595, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800146X.

Li:2009:IIK

[LT09] Chung-Wei Li and Gwo-Hshiung Tzeng. Identification of interrelationship of key customers’ needs based on structural model for services/capabilities provided by a Semiconductor-Intellectual-Property Mall. Applied Mathematics and Computation, 215(6):2001– 2010, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030900705X.

Liu:2007:GSP

[LTC07] Zhijun Liu, Ronghua Tan, and Lansun Chen. Global stability in a periodic delayed predator–prey system. Applied Mathematics and Computation, 186(1): 389–403, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306009489.

Liu:2008:SPS

[LTCC08] Zhijun Liu, Ronghua Tan, Yiping Chen, and Lansun Chen. On the stable periodic solutions of a delayed two-species model of facultative mutualism. Applied Mathematics and Computation, 196(1):105–117, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006492.

Li:2005:ISS

[LTH05] Li-Hua Li, Shiang-Feng Tzeng, and Min-Shiang Hwang. Improvement of signature scheme based on factoring and discrete logarithms. Applied Mathematics and Computation, 161(1):49–54, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303012086.

Liu:2006:CLS

[LTT06] Zhongyun Liu, Zhaolu Tian, and Yanxiang Tan. Computing the least-square solutions for centrohermitian matrix problems. Applied Mathematics and Computation, 174(1):566– 577, March 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305004856.

Lin:2007:ESC

[LTW07] Shy-Der Lin, Yen-Shung Tsai, and Pin-Yu Wang. Explicit solutions of a certain class of associated Legendre equations by means of fractional calculus. Applied Mathematics and Computation, 187(1):280– 289, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011623.

Liao:2009:ETW

[LTZ09] Xinyuan Liao, Xianhua Tang, and Shengfan Zhou. Existence of traveling wavefronts in a cooperative systems with discrete delays. Applied Mathematics and Computation, 215(5):1806–1812, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006808.

Lu:2007:NIA

[Lu07]  Junfeng Lu.  A note

on the iterative algorithm for large sparse saddle point problems. Applied Mathematics and Computation, 188(1):189–193, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013361.

Luo:2005:NNI

[Luo05a] Xing-Guo Luo. A note on the new iteration method for solving algebraic equations. Applied Mathematics and Computation, 171(2): 1177–1183, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Luo:2005:TSA

[Luo05b] Xing-Guo Luo. A two-step Adomian decomposition method. Applied Mathematics and Computation, 170(1):570–583, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304009695.

Luo:2006:OHC

[Luo06]  Zhixue Luo.  Optimal

harvesting control problem for an age-dependent competing system of n species. Applied Mathematics and Computation, 183(1):119–127, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005431.

Luo:2007:OHP

[Luo07]  Zhixue Luo.  Opti-

mal harvesting problem for an age-dependent ndimensional food chain diffusion model. Applied Mathematics and Computation, 186(2):1742–1752, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030601109X.

Lakshmikantham:2005:GQV

[LV05] V. Lakshmikantham and A. S. Vatsala. Generalized quasilinearization versus Newton’s method. Applied Mathematics and Computation, 164(2):523–530, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003340.

Li:2005:CTS

[LW05a] Qinghong Li and Xinyuan Wu. A class of twostep explicit methods for periodic IVPs. Applied Mathematics and Computation, 171(2):1239–1252, December 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:2005:MSS

[LW05b] Zhenqing Li and Weiming Wang. Mechanization for solving SPP by reducing order method. Applied Mathematics and Computation, 169(2):1028–1037, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400832X.

Lyuu:2005:CIH

[LW05c] Yuh-Dauh Lyuu and Ming-Luen Wu. Cryptanalysis of and improvement on the Hwang– Chen multi-proxy multisignature schemes. Applied Mathematics and Computation, 167(1):729–739, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304005260.

Li:2006:CAS

[LW06a] Zhenqing Li and Weiming Wang. Computer aided solving the highorder transition probability matrix of the finite Markov chain. Applied Mathematics and Computation, 172(1):267–285, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305002031.

See erratum [ESAH10].

Li:2006:DMG

[LW06b] Zhenqing Li and Weiming Wang. Distribution models of grazing animals around a grassland resource point. Applied Mathematics and Computation, 172(2):925–937, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030500247X.

Lin:2006:CUM

[LW06c] Yiqin Lin and Yimin Wei. Corrected Uzawa methods for solving large nonsymmetric saddle point problems. Applied Mathematics and Computation, 183(2):1108–1120, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600645X.

Liu:2006:SNS

[LW06d] Xinzhi Liu and Qing Wang.

Stability of nontrivial solution of delay differential equations with statedependent impulses. Applied Mathematics and Computation, 174(1):271– 288, March 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305004558.

Liu:2007:NSM

[LW07] Shiang-Tai Liu and Rong-Tsu Wang. A numerical solution method to interval quadratic programming. Applied Mathematics and Computation, 189(2):1274–1281, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017139.

Li:2009:BAE

[LW09a] Guihua Li and Wendi Wang. Bifurcation analysis of an epidemic model with nonlinear incidence. Applied Mathematics and Computation, 214(2):411– 423, August 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300309003269.

Li:2009:LOM

[LW09b] Xiang Li and Hau-San Wong. Logic optimality for multi-objective optimization. Applied Mathematics and Computation, 215(8):3045–3056, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309008820.

Liu:2009:NHB

[LW09c] M. Z. Liu and Qiubao Wang. Numerical Hopf bifurcation of linear multistep methods for a class of delay differential equations. Applied Mathematics and Computation, 208(2):462–474, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308009181.

Liu:2009:CIP

[LW09d] Zhenhai Liu and Baiyu Wang. Coefficient identification in parabolic equations. Applied Mathematics and Computation, 209(2):379–390, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009740.

Long:2009:NOC

[LW09e] Qi Long and Qi-Ru Wang.

New oscillation criteria of second-order nonlinear differential equations. Applied Mathematics and Computation, 212(2):357–365, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001672.

Li:2009:DDS

[LWG09] Wenhao Li, Shifeng Wu, and Siqing Gan. Delaydependent stability of symmetric schemes in Boundary Value Methods for DDEs. Applied Mathematics and Computation, 215(7):2445–2455, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007802.

Lee:2006:CAI

[LWH06a] Wen-Chuan Lee, Jong-Wuu Wu, and Jye-Wei Hsu. Computational algorithm for inventory model with a service level constraint, lead time demand with the mixture of distributions and controllable negative exponential backorder rate. Applied Mathematics and Computation, 175(2):1125–1138, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500665X.

Lee:2006:HNM

[LWH06b] Z. Y. Lee, C. C. Wang, and C. Y. Hu. Hybrid numerical method applied to multilayered hollow cylinder with vapor flow. Applied Mathematics and Computation, 179(1):11–29, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010027.

Lv:2005:PCA

[LWK05a] Jiqiang Lv, Xinmei Wang, and Kwangjo Kim. Practical convertible authenticated encryption schemes using self-certified public keys. Applied Mathematics and Computation, 169(2):1285–1297, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008513.

Lv:2005:SMS

[LWK05b] Jiqiang Lv, Xinmei Wang, and Kwangjo Kim. Security of a multisignature scheme for specified group of verifiers. Applied Mathematics and Computation, 166(1):58– 63, July 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004217.

Liu:2009:SIA

[LWKU09] Zeqing Liu, Lili Wang, Shin Min Kang, and Jeong Sheok Ume. Solvability and iterative algorithms for a higher order nonlinear neutral delay differential equation. Applied Mathematics and Computation, 215(7):2534–2543, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007917.

Liu:2006:PAP

[LWL06] Chun-Ho Liu, Chat-Ming Woo, and Dennis Y. C. Leung. Performance analysis of a parallel finite element solution to the direct numerical simulation of fluid turbulence on Linux PC clusters. Applied Mathematics and Computation, 172(2):731–743, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009075.

Lee:2007:CAP

[LWL07a] Wen-Chuan Lee, Jong-Wuu Wu, and Chia-Ling Lei. Computational algorithmic procedure for optimal inventory policy involving ordering cost reduction and back-order discounts when lead time demand is controllable. Applied Mathematics and Computation, 189(1):186– 200, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015918.

Liu:2007:IDC

[LWL07b] Shengping Liu, Licheng Wu, and Zhen Lu. Impact dynamics and control of a flexible dual-arm space robot capturing an object. Applied Mathematics and Computation, 185(2):1149–1159, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008393.

Lei:2005:CNS

[LWW05] Tiangang Lei, Yimin Wei, and Ching-Wah Woo. Condition numbers and structured perturbation of the W -weighted Drazin inverse. Applied Mathematics and Computation, 165 (1):185–194, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304003728.

Li:2009:PAN

[LWW09] Yongjun Li, Suyun Wang, and Hongqing Wu. Pullback attractors for nonautonomous reaction–diffusion equations in L ^{p}. Applied Mathematics and Computation, 207(2):373–379, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008230.

Li:2009:DQM

[LWWS09] Jianbing Li, Xuesong Wang, Tao Wang, and Chun Shen. Delaminating quadrature method for multi-dimensional highly oscillatory integrals. Applied Mathematics and Computation, 209(2):327– 338, March 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308009685.

Lin:2005:NIB

[LWZH05] Chih-Yin Lin, Tzong-Chen Wu, Fangguo Zhang, and Jing-Jang Hwang. New identity-based society oriented signature schemes from pairings on elliptic curves. Applied Mathematics and Computation, 160(1):245–260, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303008567.

Luo:2006:FOA

[LX06a] Xiaonan Luo and Qingzhen Xu. Fourth-order algorithm for solving 2D transient heat and moisture transfer simulation through fabric. Applied Mathematics and Computation, 182(2):1542–1555, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004954.

Luo:2006:NNI

[LX06b] Xiaonan Luo and Qingzhen Xu. A new numerical implementation on 2D heat and moisture transfer through fabric. Applied Mathematics and Computation, 174(2):1135–1150, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005370.

Li:2007:EOS

[LX07a] FengJun Li and ZongBen Xu. The essential order of simultaneous approximation for neural networks. Applied Mathematics and Computation, 194(1):120–127, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700450X.

Liu:2007:CRT

[LX07b] Jialan Liu and Aiguo Xiao.

Convergence results of twostep W -methods for twoparameter singular perturbation problems. Applied Mathematics and Computation, 189(1):669–681, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016420.

Lian:2009:HBA

[LX09a] Fuyun Lian and Yuantong Xu. Hopf bifurcation analysis of a predator– prey system with Holling type IV functional response and time delay. Applied Mathematics and Computation, 215(4):1484–1495, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900647X.

Liu:2009:OCF

[LX09b] Xiuxiang Liu and Zhiting Xu. Oscillation criteria for a forced mixed type Emden–Fowler equation with impulses. Applied Mathematics and Computation, 215(1):283–291, September 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309004597.

Li:2006:DTA

[LXH06] Wenxia Li, Dongmei Xiao, and Lin He. DNA ternary addition. Applied Mathematics and Computation, 182(2):977–986, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004127.

Li:2009:HPM

[LXJ09] Xicheng Li, Mingyu Xu, and Xiaoyun Jiang. Homotopy perturbation method to time-fractional diffusion equation with a moving boundary condition. Applied Mathematics and Computation, 208(2):434–439, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009132.

Luosheng:2009:CFN

[LXJL09] Wen Luosheng, Yang Xiaofan, Zhong Jiang, and Han Liang. Canonical form of a nonlinear monetary system. Applied Mathematics and Computation, 214(1):90–94, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002823.

Li:2006:CCE

[LXL06] Ruihong Li, Wei Xu, and Shuang Li. Chaos controlling of extended nonlinear Liénard system based on the Melnikov theory. Applied Mathematics and Computation, 178(2):405–414, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009859.

Ling-Xiao:2009:EMT

[LXML09] Li Ling-Xiao and Wang Ming-Liang. The (\frac{G′}{G})- expansion method and travelling wave solutions for a higher-order nonlinear Schrödinger equation. Applied Mathematics and Computation, 208(2):440–445, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009144.

Lin:2006:ACO

[LXSH06] Tony Lin, Shufang Xu, Qingyun Shi, and Pengwei Hao. An algebraic construction of orthonormal M-band wavelets with perfect reconstruction. Applied Mathematics and Computation, 172(2):717–730, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009063.

Lu:2007:PAB

[LXZ07] Quanyi Lü, Manyu Xiao, and Min Zhou. A parallel algorithm based on Galerkin theory for blocktridiagonal linear systems. Applied Mathematics and Computation, 187(2):1277–1285, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012665.

Liu:2009:MIS

[LXZ09] Yanjun Liu, Yongsong Xiao, and Xueliang Zhao. Multi-innovation stochastic gradient algorithm for multiple-input singleoutput systems using the auxiliary model. Applied Mathematics and Computation, 215(4):1477–1483, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006468.

Lee:2005:DIS

[LY05] Narn-Yih Lee and Shu-Ya Yang. The design of integrating subliminal channel with access control. Applied Mathematics and Computation, 171(1):573–580, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305001177.

Lai:2007:GADa

[LY07a] X. Lai and Yuan Yirang. Galerkin alternatingdirection methods for a kind of nonlinear hyperbolic equations on nonrectangular regions. Applied Mathematics and Computation, 187(2):1063–1075, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012434.

Lai:2007:GADb

[LY07b] X. Lai and Yirang Yuan. Galerkin alternating-direction method for a kind of second-order quasi-linear hyperbolic problems. Applied Mathematics and Computation, 189(2):1304– 1319, June 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306017176.

Li:2007:TPG

[LY07c] Junxiang Li and Bo Yu. Truncated partitioning group correction algorithms for large-scale sparse unconstrained optimization. Applied Mathematics and Computation, 190 (1):242–254, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307000458.

Liu:2007:ELS

[LY07d] Chunlian Liu and Zuodong Yang. Existence of large solutions for quasilinear elliptic problems with a gradient term. Applied Mathematics and Computation, 192(2):533–545, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003566.

Li:2008:MPP

[LY08a] Yao-Tang Li and Shunfeng Yang. A multiparameters preconditioned AOR iterative method for linear systems. Applied Mathematics and Computation, 206(1):465–473, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007017.

Yang:2008:GFN

[lY08b] Sheng liang Yang. On the k-generalized Fibonacci numbers and high-order linear recurrence relations. Applied Mathematics and Computation, 196(2):850– 857, March 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307007382.

Liu:2008:ELS

[LY08c] Chunlian Liu and Zuodong Yang. Existence of large solutions for a quasilinear elliptic problem via explosive sub-supersolutions. Applied Mathematics and Computation, 199(2):414– 424, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010089.

Luo:2008:CEF

[LY08d] Qiang Luo and Dongyun Yi. A co-evolving framework for robust particle swarm optimization. Applied Mathematics and Computation, 199(2):611– 622, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010387.

Li:2009:CND

[LY09a] Youfa Li and Shouzhi Yang. Construction of nonseparable dual Ω-wavelet frames in L^{2}(R^{s}). Applied Mathematics and Computation, 215(6):2082–2094, November 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309007176.

Lin:2009:ORS

[LY09b] Hwei-Jen Lin and Jih Pin Yeh. Optimal reduction of solutions for support vector machines. Applied Mathematics and Computation, 214(2):329–335, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003324.

Liu:2009:SCB

[LY09c] Gui Liu and Weidong Yu. Smart case-based indexing in worsted roving process: Combination of rough set and casebased reasoning. Applied Mathematics and Computation, 214(1):280–286, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003026.

Long:2009:MCM

[LY09d] Xiaohan Long and Yirang Yuan. Multistep characteristic method for incompressible flow in porous media. Applied Mathematics and Computation, 214(1):259–270, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003002.

Lee:2009:EHF

[LYCW09] Haw-Long Lee, Yu-Ching Yang, Win-Jin Chang, and Tser-Son Wu. Estimation of heat flux and thermal stresses in multilayer gun barrel with thermal contact resistance. Applied Mathematics and Computation, 209(2):211–221, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009375.

Li:2007:HMM

[LYGL07a] Shaoming Li, Qian Yin, Ping Guo, and Michael R. Lyu. A hierarchical mixture model for software reliability prediction. Applied Mathematics and Computation, 185(2):1120–1130, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008368.

Liu:2007:IFW

[LYGL07b] Shao-Hui Liu, Hong-Xun Yao, Wen Gao, and Yong-Liang Liu. An image fragile watermark scheme based on chaotic image pattern and pixelpairs. Applied Mathematics and Computation, 185(2):869–882, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008058.

Liu:2009:IMC

[LYH09] Xiaoji Liu, Yaoming Yu, and Chunmei Hu. The iterative methods for computing the generalized inverse A_{T,S}^{(2)} of the bounded linear operator between Banach spaces. Applied Mathematics and Computation, 214(2):391–410, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003257.

Li:2009:DAC

[LYJ09] Jinxian Li, Jurang Yan, and Xinchun Jia. Dynamical analysis of Cohen– Grossberg neural networks with finite and infinite delays. Applied Mathematics and Computation, 213(2):529–537, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002665.

Li:2007:PGE

[LYKC07] Kangshun Li, Xuezhi Yue, Lishan Kang, and Zhangxin Chen. A particle gradient evolutionary algorithm for solving multi-objective problems. Applied Mathematics and Computation, 187(2):1173–1186, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012549.

Lv:2007:SEM

[LYL07] W. J. Lv, Z. W. Yang, and M. Z. Liu. Stability of the Euler–Maclaurin methods for neutral differential equations with piecewise continuous arguments. Applied Mathematics and Computation, 186(2):1480–1487, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010733.

Lien:2008:RRH

[LYL^{+}08] Chang-Hua Lien, Ker-Wei Yu, Yen-Feng Lin, Yeong-Jay Chung, and Long-Yeu Chung. Robust reliable H_{∞} control for uncertain nonlinear systems via LMI approach. Applied Mathematics and Computation, 198(1):453–462, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008284.

Lin:2009:NFF

[LYM09] Youjiang Lin, Yongjian Yang, and M. Mammadov. A new filled function method for nonlinear equations. Applied Mathematics and Computation, 210(2):411–421, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000095.

Yan:2008:SRS

[lYpZqL08] Zhi lian Yan, Jian ping Zhou, and Xi qiang Liu. Symmetry reductions and similarity solutions of the (3 + 1)-dimensional breaking soliton equation. Applied Mathematics and Computation, 201(1–2):333–339, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030701209X.

Yan:2006:SSS

[lYqL06] Zhi lian Yan and Xi qiang Liu. Symmetry and similarity solutions of variable coefficients generalized Zakharov–Kuznetsov equation. Applied Mathematics and Computation, 180(1):288–294, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305011240.

Yan:2007:DSM

[lYqLW07] Zhi lian Yan, Xi qiang Liu, and Ling Wang. The direct symmetry method and its application in variable coefficients Schrödinger equation. Applied Mathematics and Computation, 187(2):701–707, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011258.

Liu:2009:DRW

[LYW09] Xiaoji Liu, Yaoming Yu, and Hongxing Wang. Determinantal representation of weighted generalized inverses. Applied Mathematics and Computation, 208(2):556–563, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009302.

Li:2008:FCF

[LYX08] Li Li, Fajun Yu, and Liu Xiaodong. Feedback control of fractional nonlinear differential algebraic systems with Hamiltonian function method. Applied Mathematics and Computation, 197(2):904–913, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008600.

Landi:2006:ASM

[LZ06a] G. Landi and F. Zama. The active-set method for nonnegative regularization of linear ill-posed problems. Applied Mathematics and Computation, 175(1):715– 729, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006405.

Li:2006:DAS

[LZ06b] Hui-Min Li and Ke-Cun Zhang. A decomposition algorithm for solving largescale quadratic programming problems. Applied Mathematics and Computation, 173(1):394–403, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003607.

Li:2006:PGC

[LZ06c] J. X. Li and H. W. Zhang. Partitioning group correction Cholesky techniques for large scale sparse unconstrained optimization. Applied Mathematics and Computation, 182(2):1010–1020, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004164.

Zhang:2006:RTR

[lZ06d] Ju liang Zhang. A robust trust region method for nonlinear optimization with inequality constraint. Applied Mathematics and Computation, 176(2):688–699, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008684.

Liu:2006:GPS

[LZ06e] LiYing Liu and XueSheng Zhang. Generalized pattern search methods for linearly equality constrained optimization problems. Applied Mathematics and Computation, 181(1):527–535, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001226.

Li:2007:CPG

[LZ07a] J. X. Li and H. W. Zhang.

On the convergence of partitioning group correction algorithms. Applied Mathematics and Computation, 186(1):365–378, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009465.

Li:2007:PAN

[LZ07b] Yongjun Li and Chengkui Zhong. Pullback attractors for the norm-to-weak continuous process and application to the nonautonomous reaction– diffusion equations. Applied Mathematics and Computation, 190(2):1020– 1029, July 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306015396.

Liu:2007:CSE

[LZ07c] Qian Liu and Yuqian Zhou. Construction of solutions with exactly k blow-up points for nonlinear Schrödinger equation with a harmonic potential. Applied Mathematics and Computation, 189(1):828– 835, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016614.

Lu:2007:AGO

[LZ07d] Xin-Wei Lu and Ke-Cun Zhang. An application of a global optimization parametrization method on macromolecular structure problem. Applied Mathematics and Computation, 193(1):127–134, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307003797.

Laszkiewicz:2008:NEA

[LZ08a] Beata Laszkiewicz and Krystyna Zieţak. Numerical experiments with algorithms for the ADI and Zolotarev coefficients. Applied Mathematics and Computation, 206(1):298– 312, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308006450.

Li:2008:ETW

[LZ08b] Jibin Li and Yi Zhang. Exact travelling wave solutions in a nonlinear elastic rod equation. Applied Mathematics and Computation, 202(2):504–510, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001252.

Li:2008:GMR

[LZ08c] Wenting Li and Hongqing Zhang. Generalized multiple Riccati equations rational expansion method with symbolic computation to construct exact complexiton solutions of nonlinear partial differential equations. Applied Mathematics and Computation, 197(1):288–296, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700776X.

Liang:2008:ECM

[LZ08d] Sihua Liang and Jihui Zhang. The existence of countably many positive solutions for onedimensional p-Laplacian with infinitely many singularities on the halfline. Applied Mathematics and Computation, 201(1–2):210–220, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011885.

Liang:2008:PST

[LZ08e] Sihua Liang and Jihui Zhang. Positive solutions of 2 n th-order ordinary differential equations with multi-point boundary conditions. Applied Mathematics and Computation, 197(1):262–270, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307007746.

Li:2009:NES

[LZ09a] Dong-Long Li and Jun-Xiao Zhao. New exact solutions to the (2 + 1)- dimensional Ito equation: Extended homoclinic test technique. Applied Mathematics and Computation, 215(5):1968–1974, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007073.

Li:2009:BSP

[LZ09b] Ping Li and Shou-Ming Zhong. BIBO stabilization of piecewise switched linear systems with delays and nonlinear perturbations. Applied Mathematics and Computation, 213(2):405–410, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002422.

Liu:2009:DTS

[LZ09c] Bing Liu and Lei Zhang.

Dynamics of a two-species Lotka–Volterra competition system in a polluted environment with pulse toxicant input. Applied Mathematics and Computation, 214(1):155–162, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002896.

Liu:2009:SGL

[LZ09d] Zuhan Liu and Ling Zhou.

The spinor Ginzburg– Landau model in dimension three. Applied Mathematics and Computation, 207(2):448–461, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008321.

Long:2009:CCR

[LZ09e] Han Long and Xinjian Zhang. Construction and calculation of reproducing kernel determined by various linear differential operators. Applied Mathematics and Computation, 215(2):759–766, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005591.

Liu:2007:SCD

[LZA^{+}07] F. Liu, P. Zhuang, V. Anh, I. Turner, and K. Burrage. Stability and convergence of the difference methods for the space– time fractional advection– diffusion equation. Applied Mathematics and Computation, 191(1):12–20, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012100.

Li:2005:WAE

[LZC05] Shiguang Li, Gang Zhou, and Dayue Chen. The WZ algorithm for the eigenvalue problem of complex matrix. Applied Mathematics and Computation, 165(2):347–353, June 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003868.

Liao:2007:PGS

[LZC07a] Xinyuan Liao, Shengfan Zhou, and Yuming Chen. On permanence and global stability in a general Gilpin–Ayala competition predator–prey discrete system. Applied Mathematics and Computation, 190(1):500–509, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000811.

Liao:2007:PDT

[LZC07b] Xinyuan Liao, Shengfan Zhou, and Yuming Chen. Permanence for a discrete time Lotka– Volterra type food-chain model with delays. Applied Mathematics and Computation, 186(1):279–285, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009386.

Liao:2006:ASD

[LZdC06] Xin-Yuan Liao, Hui-Yan Zhu, and Feng de Chen. On asymptotic stability of delay-difference systems. Applied Mathematics and Computation, 176(2):759–767, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008763.

Li:2007:ALB

[LZL^{+}07] Wei Li, Rong Zhang, Zhengkai Liu, Huijing Zhao, and Ryosuke Shibasaki. An approach of laser-based vehicle monitor. Applied Mathematics and Computation, 185(2):953–962, February 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306008125.

Liu:2007:CSR

[LZR07] Zhongyun Liu, Yulin Zhang, and Rui Ralha. Computing the square roots of matrices with central symmetry. Applied Mathematics and Computation, 186(1):715–726, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009921.

Liu:2007:EPS

[LZW07] Lishan Liu, Xinguang Zhang, and Yonghong Wu. On existence of positive solutions of a twopoint boundary value problem for a nonlinear singular semipositone system. Applied Mathematics and Computation, 192(1):223–232, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003256.

Luo:2009:RSA

[LZWK09] Mengzhuo Luo, Shouming Zhong, Rongjun Wang, and Wei Kang. Robust stability analysis for discretetime stochastic neural networks systems with timevarying delays. Applied Mathematics and Computation, 209(2):305–313, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009661.

Lin:2008:PPA

[LZX08] Ying Lin, Jun Zhang, and Jing Xiao. A pseudo parallel ant algorithm with an adaptive migration controller. Applied Mathematics and Computation, 205(2):677–687, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003469.

Zhao:2009:LIS

[lZXL09] Qiu lan Zhao, Xi-Xiang Xu, and Xin-Yue Li. The Liouville integrable systems associated with a new discrete 3 × 3 matrix spectral problem. Applied Mathematics and Computation, 215(7):2557–2564, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309007942.

Zhao:2005:SMS

[lZyT05] Chuan li Zhao and Heng yong Tang. Single machine scheduling problems with deteriorating jobs. Applied Mathematics and Computation, 161(3):865–874, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000372.

Zhang:2006:SLM

[lZZ06a] Ju liang Zhang and Xiangsun Zhang. A smoothing Levenberg–Marquardt method for NCP. Applied Mathematics and Computation, 178(2):212–228, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009689.

Liu:2006:ULB

[LZZ06b] Qian Liu, Yuqian Zhou, and Jian Zhang. Upper and lower bound of the blow-up rate for nonlinear Schrödinger equation with a harmonic potential. Applied Mathematics and Computation, 172(2):1121–1132, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500281X.

Liang:2009:SIP

[LZZC09] Xue-Zhang Liang, Jie-Lin Zhang, Ming Zhang, and Li-Hong Cui. Superposition interpolation process in C^{n}. Applied Mathematics and Computation, 215(1):227–234, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309004494.

Liu:2006:SCG

[LZZZ06] Qian Liu, Yuqian Zhou, Jian Zhang, and Weinian Zhang. Sharp condition of global existence for nonlinear Schrödinger equation with a harmonic potential. Applied Mathematics and Computation, 177(2):482–487, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009264.

Ma:2005:NSQ

[Ma05a] Changfeng Ma. A new smoothing quasi-Newton method for nonlinear complementarity problems. Applied Mathematics and Computation, 171(2):807– 823, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ma:2005:TAF

[Ma05b]  Jianwei Ma.  Towards

artifact-free characterization of surface topography using complex wavelets and total variation minimization. Applied Mathematics and Computation, 170(2):1014–1030, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000433.

Ma:2005:EPS

[Ma05c]  Ruyun Ma.  Existence

of positive solutions of a fourth-order boundary value problem. Applied Mathematics and Computation, 168(2):1219–1231, September 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304006964.

Maleknejad:2005:NSV

[MA05d] K. Maleknejad and N. Aghazadeh. Numerical solution of Volterra integral equations of the second kind with convolution kernel by using Taylor-series expansion method. Applied Mathematics and Computation, 161(3):915–922, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000402.

McCormack:2005:DED

[MA05e] Robert K. McCormack and Linda J. S. Allen. Disease emergence in deterministic and stochastic models for host and pathogen. Applied Mathematics and Computation, 168(2):1281–1305, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007027.

MirHassani:2005:EAC

[MA05f] S. A. MirHassani and M. R. Alirezaee. An efficient approach for computing non-Archimedean ϵ in DEA based on integrated models. Applied Mathematics and Computation, 166(2):449–456, July 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004497.

Mohamed:2005:SGS

[MA05g] A. S. Mohamed and H. A. Atia. Separation of the general second order elliptic differential operator with an operator potential in the weighted Hilbert spaces. Applied Mathematics and Computation, 162(1):155–163, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000761.

Ma:2006:MST

[Ma06a] Changfeng Ma. A method for solving time-dependent Maxwell’s equations. Applied Mathematics and Computation, 176(2):621– 631, May 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030500860X.

Ma:2006:SEC

[Ma06b] Changfeng Ma. A scheme for 3D eddy current model based on ungauged electromagnetic potentials. Applied Mathematics and Computation, 179(2):639–653, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010696.

Maleknejad:2006:NSFc

[MA06c] K. Maleknejad and A. Arzhang.

Numerical solution of the Fredholm singular integro-differential equation with Cauchy kernel by using Taylor-series expansion and Galerkin method. Applied Mathematics and Computation, 182(1):888–897, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004061.

Mat:2006:ENI

[MA06d] Mahmut D. Mat and Kemal Aldas. Experimental and numerical investigation of effect of particle size on particle distribution in particulate metal matrix composites. Applied Mathematics and Computation, 177(1):300–307, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009112.

Mohanty:2006:FNU

[MA06e] R. K. Mohanty and Urvashi Arora. A family of non-uniform mesh tension spline methods for singularly perturbed two-point singular boundary value problems with significant first derivatives. Applied Mathematics and Computation, 172(1):531–544, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002213.

Mohanty:2006:TIM

[MA06f] R. K. Mohanty and Urvashi Arora. A TAGE iterative method for the solution of non-linear singular two point boundary value problems using a sixth order discretization. Applied Mathematics and Computation, 180(2):538–548, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000427.

Ma:2007:FEM

[Ma07a] Jingtang Ma. Finite element methods for partial Volterra integro-differential equations on two-dimensional unbounded spatial domains. Applied Mathematics and Computation, 186(1):598–609, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306009714.

Ma:2007:PSB

[Ma07b] Ruyun Ma. Positive solutions for boundary value problems of functional differential equations. Applied Mathematics and Computation, 193(1):66–72, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307003748.

Ma:2007:SGN

[Ma07c] Zhengyi Ma. Symmetry group and non-propagating solitons in the (2 + 1)- dimensional sine-Gordon equation. Applied Mathematics and Computation, 194(1):67–73, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004444.

Malek:2007:NSL

[MA07d] A. Malek and M. Alipour. Numerical solution for linear and quadratic programming problems using a recurrent neural network. Applied Mathematics and Computation, 192(1):27– 39, September 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307002901.

Ma:2008:FEA

[Ma08a] Changfeng Ma. The finite element analysis of a decoupled T –ψ scheme for solving eddy-current problems. Applied Mathematics and Computation, 205(1):352–361, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005936.

Ma:2008:GCL

[Ma08b] Changfeng Ma. A globally convergent Levenberg– Marquardt method for the least l _{2}-norm solution of nonlinear inequalities. Applied Mathematics and Computation, 206(1):133–140, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006383.

Ma:2008:CSI

[Ma08c] Jianwei Ma. Compressed sensing by inverse scale space and curvelet thresholding. Applied Mathematics and Computation, 206(2):980–988, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007649.

Mabrouk:2008:LFD

[MA08d] Anouar Ben Mabrouk and Mekki Ayadi. A linearized finite-difference method for the solution of some mixed concave and convex non-linear problems. Applied Mathematics and Computation, 197(1):1–10, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307007485.

Mabrouk:2008:LTO

[MA08e] Anouar Ben Mabrouk and Mekki Ayadi. Lyapunov type operators for numerical solutions of PDEs. Applied Mathematics and Computation, 204(1):395–407, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005183.

Mostafa:2009:CSV

[MA09] A. O. Mostafa and M. K. Aouf. Certain subclasses of p-valent functions involving the Dziok– Srivastava operator. Applied Mathematics and Computation, 211(1):23– 32, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000630.

Ma:2011:CDM

[Ma11] N.-Y. Ma. A comment on “On the distribution of Ma and King”. Applied Mathematics and Computation, 218(1):202–203, September 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300311007296.

See [Nad07b].

Molavi-Arabshahi:2007:PTS

[MAD07] Seyedeh Mahboubeh Molavi-Arabshahi and Mehdi Dehghan. Preconditioned techniques for solving large sparse linear systems arising from the discretization of the elliptic partial differential equations. Applied Mathematics and Computation, 188(2):1371–1388, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014901.

Mabrouk:2008:WDA

[MAD08] A. Ben Mabrouk, N. Ben Abdallah, and Z. Dhifaoui. Wavelet decomposition and autoregressive model for time series prediction. Applied Mathematics and Computation, 199(1):334–340, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010028.

Mahmoudi:2005:WGM

[Mah05] Y. Mahmoudi. Wavelet Galerkin method for numerical solution of nonlinear integral equation. Applied Mathematics and Computation, 167(2):1119– 1129, August 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304005533.

Maheshwari:2009:FOI

[Mah09a] Amit Kumar Maheshwari. A fourth order iterative method for solving nonlinear equations. Applied Mathematics and Computation, 211(2):383–391, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000861.

Mahmoud:2009:GCS

[Mah09b] Magdi S. Mahmoud. Generalized H_{2} control of switched discrete-time systems with unknown delays. Applied Mathematics and Computation, 211(1):33– 44, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009338.

Maity:2005:NFB

[Mai05]  Damodar Maity.  A

novel far-boundary condition for the finite element analysis of infinite reservoir. Applied Mathematics and Computation, 170(2):1314–1328, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000664.

Momani:2005:NSSa

[MAK05] Shaher Momani and Kamel Al-Khaled. Numerical solutions for systems of fractional differential equations by the decomposition method. Applied Mathematics and Computation, 162(3):1351–1365, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001936.

Makinde:2006:LFL

[Mak06] O. D. Makinde. Laminar falling liquid film with variable viscosity along an inclined heated plate. Applied Mathematics and Computation, 175(1):80– 88, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030500593X.

Makinde:2007:ADA

[Mak07a] O. D. Makinde. Adomian decomposition approach to a SIR epidemic model with constant vaccination strategy. Applied Mathematics and Computation, 184(2):842–848, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007715.

Makinde:2007:SRD

[Mak07b] O. D. Makinde. Solving ratio-dependent predator– prey system with constant effort harvesting using Adomian decomposition method. Applied Mathematics and Computation, 186(1):17–22, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009118.

Makinde:2007:TSR

[Mak07c] Oluwole Daniel Makinde. Thermal stability of a reactive third grade fluid in a cylindrical pipe: an exploitation of Hermite–Padé approximation technique.

Applied Mathematics and Computation, 189(1):690– 697, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016456.

Makinde:2009:TSR

[Mak09] Oluwole Daniel Makinde.

On thermal stability of a reactive third-grade fluid in a channel with convective cooling the walls. Applied Mathematics and Computation, 213(1):170– 176, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900215X.

Maleknejad:2006:CFE

[Mal06] K. Maleknejad. A comparison of Fourier extrapolation methods for numerical solution of deconvolution. Applied Mathematics and Computation, 183(1):533–538, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005881.

Maleknejad:2006:NSFa

[MAM06] K. Maleknejad, N. Aghazadeh, and R. Mollapourasl. Numerical solution of Fredholm integral equation of the first kind with collocation method and estimation of error bound. Applied Mathematics and Computation, 179(1):352–359, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010428.

Mamode:2009:VIM

[Mam09]  Malik Mamode.  Vari-

ational iterative method and initial-value problems. Applied Mathematics and Computation, 215(1):276–282, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004585.

Mahdavi-Amiri:2006:DFN

[MAN06] N. Mahdavi-Amiri and S. H. Nasseri. Duality in fuzzy number linear programming by use of a certain linear ranking function. Applied Mathematics and Computation, 180(1):206–216, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030501115X.

Momani:2006:VIM

[MAO06] Shaher Momani, Salah Abuasad, and Zaid Odibat. Variational iteration method for solving nonlinear boundary value problems. Applied Mathematics and Computation, 183(2):1351–1358, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006825.

Mao:2009:NCT

[Mao09] Qing-Song Mao. A note on the Chebyshev-type inequality of Sugeno integrals. Applied Mathematics and Computation, 212 (1):275–279, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309001386.

Madan:2005:QFO

[MAR05a] Kailash C. Madan and Mohammad Al-Rawwash. On the M^{x}/G/1 queue with feedback and optional server vacations based on a single vacation policy. Applied Mathematics and Computation, 160(3):909–919, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303012025.

Marin:2005:MMN

[Mar05b] Liviu Marin. A meshless method for the numerical solution of the Cauchy problem associated with three-dimensional Helmholtz-type equations. Applied Mathematics and Computation, 165(2):355– 374, June 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030400387X.

Martins:2005:AAF

[Mar05c] João Luiz Martins. Approximate analytical formulas for Kirchhoff migration operator. Applied Mathematics and Computation, 163(3):1181–1195, April 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002887.

Maleknejad:2006:NSS

[MAR06] K. Maleknejad, N. Aghazadeh, and M. Rabbani. Numerical solution of second kind Fredholm integral equations system by using a Taylor-series expansion method. Applied Mathematics and Computation, 175(2):1229–1234, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006727.

Marchi:2007:MW

[Mar07] Stefano De Marchi. Mathematics and wine. Applied Mathematics and Computation, 192(1):180–190, September 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307003104.

Markolefas:2008:SGF

[Mar08] Stilianos Markolefas. Standard Galerkin formulation with high order Lagrange finite elements for option markets pricing. Applied Mathematics and Computation, 195(2):707–720, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006091.

Marik:2014:RPS

[Mar14] Robert Marík. Remarks on the paper by Sun and Meng, Appl. Math. Comput. 174 (2006). Applied Mathematics and Computation, 248(??):309–313, December 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300314013344. See

[SM06d].

Mahmoud:2008:PAS

[MAT08] Feras A. Mahmoud and Mohammad H. Al-Towaiq. Parallel algorithm for the solutions of PDEs in Linux clustered workstations. Applied Mathematics and Computation, 200(1):178–188, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010983.

Matsunaga:2009:SSS

[Mat09] Hideaki Matsunaga. Stability switches in a system of linear differential equations with diagonal delay. Applied Mathematics and Computation, 212(1):145– 152, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001271.

Mavrotas:2009:EIC

[Mav09]  George Mavrotas.  Ef-

fective implementation of the ϵ-constraint method in Multi-Objective Mathematical Programming problems. Applied Mathematics and Computation, 213(2):455–465, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002574.

Moghaddam:2009:ETW

[MAY09] M. Yaghobi Moghaddam, A. Asgari, and H. Yazdani. Exact travelling wave solutions for the generalized nonlinear Schrödinger (GNLS) equation with a source by Extended tanh– coth, sin–cos and Exp-Function methods. Applied Mathematics and Computation, 210(2):422–435, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000101.

Merdan:2005:EIL

[MB05a] Z. Merdan and M. Bayırlı.

The effect of the increase of linear dimensions on exponents obtained by finitesize scaling relations for the six-dimensional Ising model on the Creutz cellular automaton. Applied Mathematics and Computation, 167(1):212–224, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004862.

Morishita:2005:DCT

[MB05b] Helio Mitio Morishita and Yng Bolognani. Dynamics and control of a two-body floating system under realistic environmental loads. Applied Mathematics and Computation, 164(2):573–590, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003388.

Malek:2006:NSH

[MB06] A. Malek and R. Shekari Beidokhti. Numerical solution for high order differential equations using a hybrid neural network— Optimization method. Applied Mathematics and Computation, 183(1):260– 271, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306005583.

Mandal:2007:NSS

[MB07] B. N. Mandal and Subhra Bhattacharya. Numerical solution of some classes of integral equations using Bernstein polynomials. Applied Mathematics and Computation, 190(2):1707–1716, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002160.

Mukhopadhyay:2008:BAE

[MB08] B. Mukhopadhyay and R. Bhattacharyya. Bifurcation analysis of an ecological food-chain model with switching predator. Applied Mathematics and Computation, 201(1–2):260–271, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011939.

Massad:2008:OMA

[MBC08] Eduardo Massad, Marcelo Nascimento Burattini, and Francisco Antonio Bezerra Coutinho. An optimization model for antibiotic use. Applied Mathematics and Computation, 201(1–2):161–167, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011800.

Mansouri:2008:VFF

[MBDD08] R. Mansouri, M. Bettayeb, T. Djamah, and S. Djennoune. Vector Fitting fractional system identification using particle swarm optimization. Applied Mathematics and Computation, 206(2):510–520, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003731.

Maiti:2006:ARC

[MBM06] A. K. Maiti, A. K. Bhunia, and M. Maiti. An application of real-coded genetic algorithm (RCGA) for mixed integer nonlinear programming in twostorage multi-item inventory model with discount policy. Applied Mathematics and Computation, 183(2):903–915, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006242.

Mondal:2007:OTS

[MBM07] Biswajit Mondal, Asoke Kumar Bhunia, and Manoranjan Maiti. Optimal twospecies harvesting policy with price and size (biomass) of the fish population dependent catch rate function. Applied Mathematics and Computation, 187(2):600–608, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011155.

Mirevski:2007:RLF

[MBS07] S. P. Mirevski, L. Boyadjiev, and R. Scherer. On the Riemann–Liouville fractional calculus, g-Jacobi functions and F - Gauss functions. Applied Mathematics and Computation, 187(1):315–325, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000689.

McCartin:2006:ADA

[MC06] Brian J. McCartin and Matthew F. Causley. Angled derivative approximation of the hyperbolic heat conduction equations. Applied Mathematics and Computation, 182(2):1581–1607, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005066.

Mahdavi:2007:DLS

[MC07] Iraj Mahdavi and Namjae Cho. Dynamic link structure analysis using flow movement of customers in a cybermall. Applied Mathematics and Computation, 188(1):116–122, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306013270.

Meng:2008:DNS

[MC08] Xinzhu Meng and Lansun Chen. The dynamics of a new SIR epidemic model concerning pulse vaccination strategy. Applied Mathematics and Computation, 197(2):582– 597, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008041.

Ma:2009:DSE

[MC09] Wen-Xiu Ma and Min Chen. Direct search for exact solutions to the nonlinear Schrödinger equation. Applied Mathematics and Computation, 215(8):2835–2842, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008224.

Mahdavi:2008:NMH

[MCAF08] M. Mahdavi, M. Haghir Chehreghani, H. Abolhassani, and R. Forsati. Novel meta-heuristic algorithms for clustering web documents. Applied Mathematics and Computation, 201(1–2):441–451, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307012209.

Mahmood:2005:DMS

[MCAH05] Afrah Saad Mahmood, Luis Casasús, and Waleed Al-Hayani. The decomposition method for stiff systems of ordinary differential equations. Applied Mathematics and Computation, 167(2):964–975, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005429.

McCartin:2005:MAD

[McC05] Brian J. McCartin. The method of angled derivatives. Applied Mathematics and Computation, 170(1):440–461, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304009609.

McCartin:2006:EMA

[McC06] Brian J. McCartin. The exponential method of angled derivatives. Applied Mathematics and Computation, 177(2):665–685, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009483.

Meng:2007:TPD

[MCC07] Xinzhu Meng, Lansun Chen, and Huidong Cheng. Two profitless delays for the SEIRS epidemic disease model with nonlinear incidence and pulse vaccination. Applied Mathematics and Computation, 186(1):516–529, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009623.

Moura:2005:FLI

[MCGS05] André L. Moura, José R. Camacho, Sebastião C. Guimarães, Jr., and Carlos H. Salerno. A functional language to implement the divide-and-conquer Delaunay triangulation algorithm. Applied Mathematics and Computation, 168(1):178–191, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005971.

Chen:2008:FDM

[mCLB08] Chang ming Chen, F. Liu, and K. Burrage. Finite difference methods and a Fourier analysis for the fractional reaction– subdiffusion equation. Applied Mathematics and Computation, 198(2):754– 769, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009514.

Mandal:2007:EBA

[MCMA07] Prashanta Kumar Mandal, Santabrata Chakravarty, Arabinda Mandal, and Norsarahaida Amin. Effect of body acceleration on unsteady pulsatile flow of non-Newtonian fluid through a stenosed artery. Applied Mathematics and Computation, 189(1):766– 779, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016547.

Milovanovic:2008:EFF

[MCS08] Gradimir V. Milovanović, Aleksandar S. Cvetković, and Marija P. Stanić. Explicit formulas for fiveterm recurrence coefficients of orthogonal trigonometric polynomials of semiinteger degree. Applied Mathematics and Computation, 198(2):559–573, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009228.

Ma:2008:GSC

[MCW08] Changfeng Ma, Linjie Chen, and Desheng Wang. A globally and superlinearly convergent smoothing Broyden-like method for solving nonlinear complementarity problem. Applied Mathematics and Computation, 198(2):592– 604, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009253.

Mavrotas:2005:MCB

[MD05] G. Mavrotas and D. Diakoulaki. Multi-criteria branch and bound: a vector maximization algorithm for Mixed 0–1 Multiple Objective Linear Programming. Applied Mathematics and Computation, 171(1):53–71, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000809.

Maleknejad:2006:CMH

[MD06a] K. Maleknejad and H. Derili. The collocation method for Hammerstein equations by Daubechies wavelets. Applied Mathematics and Computation, 172(2):846–864, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002419.

Maleknejad:2006:NSIa

[MD06b] K. Maleknejad and H. Derili. Numerical solution of integral equations by using combination of Spline-collocation method and Lagrange interpolation. Applied Mathematics and Computation, 175(2):1235–1244, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006739.

Mersha:2006:LBP

[MD06c] Ayalew Getachew Mersha and Stephan Dempe. Linear bilevel programming with upper level constraints depending on the lower level solution. Applied Mathematics and Computation, 180(1):247–254, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305011197.

Maleknejad:2007:NSH

[MD07a] K. Maleknejad and H. Derili. Numerical solution of Hammerstein integral equations by using combination of spline-collocation method and Lagrange interpolation. Applied Mathematics and Computation, 190(2):1557–1562, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001920.

Mugisha:2007:DFM

[MD07b] J. Y. T. Mugisha and H. Ddumba. The dynamics of a fisheries model with feeding patterns and harvesting: Lates niloticus and Oreochromis niloticus in Lake Victoria. Applied Mathematics and Computation, 186(1):142–158, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009258.

Macias-Diaz:2008:BEM

[MD08a] Jorge Eduardo Macías-Díaz. On the bifurcation of energy in media governed by (2 + 1)-dimensional modified Klein–Gordon equations. Applied Mathematics and Computation, 206(1):221–235, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006711.

Mosic:2008:CNW

[MD08b] Dijana Mosić and Dragan S. Djordjević. Condition number of the W -weighted Drazin inverse. Applied Mathematics and Computation, 203(1):308–318, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002555.

Mu:2008:SPD

[MD08c] Lihua Mu and Hong Du.

The solution of a parabolic differential equation with non-local boundary conditions in the reproducing kernel space. Applied Mathematics and Computation, 202(2):708–714, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001604.

Meng:2009:SSO

[MD09a] Fanchao Meng and Zengji Du. Solvability of a secondorder multi-point boundary value problem at resonance. Applied Mathematics and Computation, 208(1):23–30, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008540.

Mohanty:2009:TOA

[MD09b] R. K. Mohanty and Deepika Dhall. Third order accurate variable mesh discretization and application of TAGE iterative method for the non-linear two-point boundary value problems with homogeneous functions in integral form. Applied Mathematics and Computation, 215(6):2024–2034, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007103.

Mosic:2009:CNR

[MD09c] Dijana Mosić and Dragan S. Djordjević. Condition number related to the outer inverse of a complex matrix. Applied Mathematics and Computation, 215(8):2826–2834, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008212.

Mei:2008:LPM

[MDZ08] Shu-Li Mei, Cheng-Jin Du, and Sen-Wen Zhang. Linearized perturbation method for stochastic analysis of a rill erosion model. Applied Mathematics and Computation, 200(1):289–296, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011083.

Muhammad:2005:CGM

[ME05] Abubakr Muhammad and Magnus Egerstedt. Connectivity graphs as models of local interactions. Applied Mathematics and Computation, 168(1):243– 269, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304006022.

Montegranario:2007:RAS

[ME07] Hebert Montegranario and Jairo Espinosa. A regularization approach for surface reconstruction from point clouds. Applied Mathematics and Computation, 188(1):583–595, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030601397X.

Mease:2005:MTS

[Mea05] K. D. Mease. Multiple time-scales in nonlinear flight mechanics: diagnosis and modeling. Applied Mathematics and Computation, 164(2):627–648, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030400342X.

Mousa:2008:GCE

[MED08] A. A. Mousa and I. M. El-Desoky. GENLS: Coevolutionary algorithm for nonlinear system of equations. Applied Mathematics and Computation, 197 (2):633–642, April 1, 2008.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030700820X.

Mehne:2006:NMS

[Meh06] H. H. Mehne. A numerical method for solving linear non-autonomous systems based on linear programming. Applied Mathematics and Computation, 178(2):287–294, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009744.

Mei:2006:MAS

[Mei06] Liquan Mei. Multigrid algorithms for solving multigroup neutron transport equations. Applied Mathematics and Computation, 179(2):473–483, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010532.

Melnik:2008:MCN

[Mel08] Roderick V. N. Melnik. Markov chain network training and conservation law approximations: Linking microscopic and macroscopic models for evolution. Applied Mathematics and Computation, 199(1):315–333, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010016.

Mena:2008:DTG

[Men08] J. B. Mena. Datum transformation by Gauss projection of the ellipsoid onto the sphere and bilinear function of complex variable: an application to ETRS89–ED50. Applied Mathematics and Computation, 195(1):190–195, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005309.

Mertz:2007:DW

[Mer07]  Joachim Mertz.  The

dragon war. Applied Mathematics and Computation, 186(1):907–914, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010137.

Mestrovic:2007:MDM

[Meš07] Mladen Meštrović. The modified decomposition method for eighth-order boundary value problems. Applied Mathematics and Computation, 188(2):1437–1444, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014962.

Medjden:2005:ABV

[MeT05] Mohamed Medjden and Nasser eddine Tatar. Asymptotic behavior for a viscoelastic problem with not necessarily decreasing kernel. Applied Mathematics and Computation, 167(2):1221–1235, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005636.

Monroy:2009:GMV

[MF09] Luisa Monroy and Francisco R. Fernández. A general model for voting systems with multiple alternatives. Applied Mathematics and Computation, 215(4):1537–1547, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006547.

Mahdavi:2007:IHS

[MFD07] M. Mahdavi, M. Fesanghary, and E. Damangir. An improved harmony search algorithm for solving optimization problems. Applied Mathematics and Computation, 188(2):1567–1579, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015098.

Mehne:2005:SND

[MFE05] H. H. Mehne, M. H. Farahi, and J. A. Esfahani. Slot nozzle design with specified pressure in a given subregion by using embedding method. Applied Mathematics and Computation, 168(2):1258–1272, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304007003.

Mavrotas:2009:SBO

[MFF09] George Mavrotas, José Rui Figueira, and Kostas Florios. Solving the biobjective multi-dimensional knapsack problem exploiting the concept of core. Applied Mathematics and Computation, 215(7):2502–2514, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007863.

Mirasyedioglu:2006:SAC

[MG06]  Şeref Mirasyedioğlu and

Tolga Güyer. A symbolic and algebraic computation based Lambda-Boolean reduction machine via PROLOG. Applied Mathematics and Computation, 176(1):65– 75, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500771X.

Ma:2009:SHB

[MG09] Junhai Ma and Qin Gao.

Stability and Hopf bifurcations in a business cycle model with delay. Applied Mathematics and Computation, 215(2):829–834, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005712.

Meng:2009:SSV

[MGG^{+}09] De-Xin Meng, Yi-Tian Gao, Xiao-Ling Gai, Lei Wang, Xin Yu, Zhi-Yuan Sun, Ming-Zhen Wang, and Xing Lü. Solitonic solutions for a variablecoefficient variant Boussinesq system in the long gravity waves. Applied Mathematics and Computation, 215(5):1744–1751, November 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309006742.

Meng:2008:CDC

[MGH08] Yimin Meng, Shangjiang Guo, and Lihong Huang. Convergence dynamics of Cohen–Grossberg neural networks with continuously distributed delays. Applied Mathematics and Computation, 202(1):188–199, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000738.

Mokhtarzadeh:2005:SIM

[MGMC05] M. R. Mokhtarzadeh, A. Golbabaee, R. Mokhtary, and N. G. Chegini. Suitable iterative methods for solving the linear system arising in the three fields domain decomposition method. Applied Mathematics and Computation, 170(2):741–751, November 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305000159.

Mancera:2006:SEHb

[MH06a] Paulo F. A. Mancera and R. Hunt. Some experiments with high order compact methods using a computer algebra software— Part II (non-uniform grid). Applied Mathematics and Computation, 180(1):233– 241, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305011173.

Mancera:2006:SEHa

[MH06b] Paulo F. A. Mancera and Roland Hunt. Some experiments with high order compact methods using a computer algebra software—Part I. Applied Mathematics and Computation, 174(2):775–794, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005096.

He:2006:NSE

[mH06c] Wen ming He. A new superconvergence and extrapolation for second order triangular element. Applied Mathematics and Computation, 174(1):300– 315, March 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305004571.

He:2007:AEO

[mH07a] Wen ming He. Asymptotic expansions of onedimensional second-order Helmholtz equation with rapidly oscillating coefficient. Applied Mathematics and Computation, 188(2):1921–1929, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015608.

He:2007:EEF

[mH07b]  Wen ming He.  Error

estimates of finite element method about elliptic problems with singular righthand side. Applied Mathematics and Computation, 188(1):824–832, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014214.

Merchant:2008:EVP

[MH08a] Bradley M. Merchant and Jeffrey J. Heys. Effects of variable permeability on aqueous humor outflow. Applied Mathematics and Computation, 196(1):371–380, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700690X.

Mrad:2008:OSD

[MH08b] Mehdi Mrad and Mohamed Haouari. Optimal solution of the discrete cost multicommodity network design problem. Applied Mathematics and Computation, 204(2):745–753, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005493.

Ma:2009:EMP

[MH09] Ruyun Ma and Xiaoling Han. Existence and multiplicity of positive solutions of a nonlinear eigenvalue problem with indefinite weight function. Applied Mathematics and Computation, 215(3):1077–1083, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006122.

Mostafa:2006:PPS

[MHA06] E. M. E. Mostafa, A. Hamdi, and A. Aboutahoun. A proximal-point SQP trust region method for solving some special class of nonlinear semi-definite programming problems. Applied Mathematics and Computation, 174(2):810– 832, March 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305005114.

Moitsheki:2009:CSH

[MHM09] R. J. Moitsheki, T. Hayat, and M. Y. Malik. Comparison of symmetry and homotopy solutions for nonlinear heat transfer systems. Applied Mathematics and Computation, 215(6):1995–2000, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006870.

He:2006:FEA

[mHsLzC06] Wen ming He, Chang sheng Lin, and Jun zhi Cui. Finite element approximations of Green function G_{x_{0}}^{ϵ} based on the method of multiscale asymptotic expansions. Applied Mathematics and Computation, 174(1): 199–222, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305004510.

He:2005:AEM

[mHzC05] Wen ming He and Jun zhi Cui. Asymptotic expansion method for the two point boundary value problem with small periodic structure. Applied Mathematics and Computation, 161(2):667–678, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000190.

Mahmoud:2008:RCN

[MI08] Magdi S. Mahmoud and Abdulla Ismail. Resilient control of non-linear discrete-time state-delay systems. Applied Mathematics and Computation, 206(2):561–569, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003792.

MirHassani:2006:CAE

[Mir06a] S. A. MirHassani. A computational approach to enhancing course timetabling with integer programming. Applied Mathematics and Computation, 175(1):814– 822, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006466.

MirHassani:2006:IPS

[Mir06b] S. A. MirHassani. Improving paper spread in examination timetables using integer programming. Applied Mathematics and Computation, 179(2):702–706, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010763.

MirHassani:2008:OPM

[Mir08] S. A. MirHassani. An operational planning model for petroleum products logistics under uncertainty. Applied Mathematics and Computation, 196(2):744– 751, March 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307007266.

Mishra:2005:OSA

[Mis05]  P. K. Mishra.  Opti-

mal systolic array algorithms for tensor product. Applied Mathematics and Computation, 168(1):496–518, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006204.

Mitra:2006:WFB

[Mit06] Sharmishtha Mitra. A wavelet filtering based analysis of macroeconomic indicators: the Indian evidence. Applied Mathematics and Computation, 175(2):1055–1079, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006624.

Mohanty:2005:CVM

[MJ05] R. K. Mohanty and Navnit Jha. A class of variable mesh spline in compression methods for singularly perturbed two point singular boundary value problems. Applied Mathematics and Computation, 168(1):704–716, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006332.

Ma:2007:LDL

[MJ07a] Changfeng Ma and Lihua Jiang. Lower-dimensional linear complementarity problem approaches to the solution of a biobstacle problem. Applied Mathematics and Computation, 190(2):1040–1049, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016390.

Ma:2007:SRL

[MJ07b] Changfeng Ma and Lihua Jiang. Some research on Levenberg–Marquardt method for the nonlinear equations. Applied Mathematics and Computation, 184(2):1032–1040, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007910.

Masjed-Jamei:2007:NTW

[MJ07c] Mohammad Masjed-Jamei. A new type of weighted quadrature rules and its relation with orthogonal polynomials. Applied Mathematics and Computation, 188(1):154–165, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013324.

Mehri:2007:LFM

[MJ07d] Bahman Mehri and Sadegh Jokar. Lebesgue function for multivariate interpolation by radial basis functions. Applied Mathematics and Computation, 187(1):306–314, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011659.

Mishra:2007:FPT

[MJ07e] Bimal Kumar Mishra and Navnit Jha. Fixed period of temporary immunity after run of antimalicious software on computer nodes. Applied Mathematics and Computation, 190(2):1207–1212, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001518.

Masjed-Jamei:2006:SSI

[MJD06] M. Masjed-Jamei and Mehdi Dehghan. On some statistical integral equations. Applied Mathematics and Computation, 179(1):87–91, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030501009X.

Masjed-Jamei:2007:PMQ

[MJD07] Mohammad Masjed-Jamei and Mehdi Dehghan. A probabilistic model for quadrature rules. Applied Mathematics and Computation, 187(2):1520–1526, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012926.

Masjed-Jamei:2005:NIG

[MJED05] M. Masjed-Jamei, M. R. Eslahchi, and M. Dehghan. On numerical improvement of Gauss– Radau quadrature rules. Applied Mathematics and Computation, 168(1):51– 64, September 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304005855.

Mahdavi:2007:DNM

[MJFAS07] Iraj Mahdavi, Babak Javadi, Kaveh Fallah-Alipour, and Jannes Slomp. Designing a new mathematical model for cellular manufacturing system based on cell utilization. Applied Mathematics and Computation, 190(1):662– 670, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001014.

Masjed-Jamei:2005:FKC

[MJHED05] M. Masjed-Jamei, S. M.

Hashemiparast, M. R. Eslahchi, and Mehdi Dehghan. The first kind Chebyshev–Lobatto quadrature rule and its numerical improvement. Applied Mathematics and Computation, 171(2):1104–1118, December 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Masjed-Jamei:2006:SKC

[MJHED06] M. Masjed-Jamei, S. M.

Hashemiparast, M. R. Eslahchi, and Mehdi Dehghan. The second kind Chebyshev quadrature rules of semi-open type and its numerical improvement. Applied Mathematics and Computation, 172(1):210–221, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001992.

Ma:2005:JSM

[MK05a] Changfeng Ma and Tong Kang. A Jacobian smoothing method for box constrained variational inequality problems. Applied Mathematics and Computation, 162(3):1397–1429, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001973.

Ma:2005:FDL

[MK05b] N.-Y. Ma and R. P. King. Fitting distributionlike data to exponential sums with genetic algorithms. Applied Mathematics and Computation, 169(1):82–95, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304007544.

Maleknejad:2005:NSN

[MK05c] K. Maleknejad and M. Karami. Numerical solution of nonlinear Fredholm integral equations by using multiwavelets in the Petrov– Galerkin method. Applied Mathematics and Computation, 168(1):102–110, September 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304005892.

Maleknejad:2005:UWM

[MK05d] K. Maleknejad and M. Karami.

Using the WPG method for solving integral equations of the second kind. Applied Mathematics and Computation, 166(1):123– 130, July 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004266.

Mohanty:2006:ATI

[MK06] R. K. Mohanty and Noopur Khosla. Application of TAGE iterative algorithms to an efficient third order arithmetic average variable mesh discretization for two-point nonlinear boundary value problems. Applied Mathematics and Computation, 172(1):148–162, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001943.

Mehra:2007:FWT

[MK07a] Mani Mehra and Vivek Kumar. Fast wavelet-Taylor Galerkin method for linear and non-linear wave problems. Applied Mathematics and Computation, 189(2):1292–1299, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017152.

Molai:2007:MAS

[MK07b] Ali Abbasi Molai and Esmaile Khorram. A modified algorithm for solving the proposed models by Ghodousian and Khorram and Khorram and Ghodousian. Applied Mathematics and Computation, 190(2):1161–1167, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001464.

Molabahrami:2008:NSH

[MK08a] A. Molabahrami and F. Khani. Numerical solutions of highly oscillatory integrals. Applied Mathematics and Computation, 198(2):657– 664, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009332.

Molai:2008:AMT

[MK08b] Ali Abbasi Molai and Esmaile Khorram. Another modification from two papers of Ghodousian and Khorram and Khorram et al. Applied Mathematics and Computation, 197(2):559–565, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008028.

See erratum [MK10].

Molai:2010:EAM

[MK10] Ali Abbasi Molai and Esmaile Khorram. Erratum to: “Another modification from two papers of Ghodousian and Khorram and Khorram et al.” [Appl. Math. Comput. 197 (2008) 559–565]. Applied Mathematics and Computation, 215(9):3488, January 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [MK08b].

Maleknejad:2005:NSH

[MKA05] K. Maleknejad, M. Karami, and N. Aghazadeh. Numerical solution of Hammerstein equations via an interpolation method. Applied Mathematics and Computation, 168(1):141– 145, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300304005934.

Mishra:2007:MMCb

[MKA07] Somna Mishra, V. K. Katiyar, and V. Arora. Mathematical modeling of chemotherapy strategies in vascular tumor growth using nanoparticles. Applied Mathematics and Computation, 189(2):1246–1254, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017097.

Martines:2009:CMR

[MKG09] Ian P. Martines, Hristo V. Kojouharov, and James P. Grover. A chemostat model of resource competition and allelopathy. Applied Mathematics and Computation, 215(2):573–582, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309005165.

Mamta:2005:CQC

[MKKS05a] Mamta, V. Kanwar, V. K. Kukreja, and Sukhjit Singh. On a class of quadratically convergent iteration formulae. Applied Mathematics and Computation, 166(3):633–637, July 26, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400462X.

Mamta:2005:STO

[MKKS05b] Mamta, V. Kanwar, V. K.

Kukreja, and Sukhjit Singh. On some thirdorder iterative methods for solving nonlinear equations. Applied Mathematics and Computation, 171(1):272–280, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000937.

Maleknejad:2006:UPG

[MKR06] K. Maleknejad, M. Karami, and M. Rabbani. Using the Petrov–Galerkin elements for solving Hammerstein integral equations. Applied Mathematics and Computation, 172(2):831–845, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002407.

Monovasilis:2009:FTF

[MKS09] Th. Monovasilis, Z. Kalogiratou, and T. E. Simos. A family of trigonometrically fitted partitioned Runge– Kutta symplectic methods. Applied Mathematics and Computation, 209(1):91– 96, March 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004773.

Mermri:2007:RAG

[MKSK07] El Bekkaye Mermri, Hideki Katagiri, Masatoshi Sakawa, and Kosuke Kato. Remarks on the application of genetic algorithm and tabu search method to nonlinear spanning tree problems. Applied Mathematics and Computation, 188(2):1071–1086, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007156.

Maleknejad:2005:NEM

[ML05] K. Maleknejad and T. Lotfi. Numerical expansion methods for solving integral equations by interpolation and Gauss quadrature rules. Applied Mathematics and Computation, 168(1):111–124, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005909.

Maleknejad:2006:EML

[ML06] K. Maleknejad and T. Lotfi.

Expansion method for linear integral equations by cardinal B-spline wavelet and Shannon wavelet as bases for obtain Galerkin system. Applied Mathematics and Computation, 175 (1):347–355, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305006120.

Malaquias:2007:ICA

[ML07] José Romildo Malaquias and Carlos Roberto Lopes. Implementing a computer algebra system in Haskell. Applied Mathematics and Computation, 192(1):120– 134, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307003013.

Ma:2008:FPT

[ML08a] Shiwei Ma and Kang Li. Fast parametric time– frequency modeling of nonstationary signals. Applied Mathematics and Computation, 205(1):170–177, November 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300308003275.

Mao:2008:PSI

[ML08b] Anmin Mao and Shixia Luan. Periodic solutions of an infinitedimensional Hamiltonian system. Applied Mathematics and Computation, 201(1–2):800–804, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011575.

Mao:2009:CPT

[ML09] Anmin Mao and Shixia Luan. Critical points theorems concerning strongly indefinite functionals and infinite many periodic solutions for a class of Hamiltonian systems. Applied Mathematics and Computation, 214(1):187–200, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002938.

Maleknejad:2007:NSFa

[MLM07] K. Maleknejad, T. Lotfi, and K. Mahdiani. Numerical solution of first kind Fredholm integral equations with wavelets-Galerkin method (WGM) and wavelets precondition. Applied Mathematics and Computation, 186(1):794– 800, March 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306010009.

Ma:2007:PEF

[MLP07] Meijie Ma, Guizhen Liu, and Xiangfeng Pan. Path embedding in faulty hypercubes. Applied Mathematics and Computation, 192(1):233–238, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003268.

Maleknejad:2007:NCM

[MLR07] K. Maleknejad, T. Lotfi, and Y. Rostami. Numerical computational method in solving Fredholm integral equations of the second kind by using Coifman wavelet. Applied Mathematics and Computation, 186(1):212–218, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009313.

Ma:2009:SAE

[MLW09] Yulan Ma, Bangqing Li, and Cong Wang. A series of abundant exact travelling wave solutions for a modified generalized Vakhnenko equation using auxiliary equation method. Applied Mathematics and Computation, 211(1):102– 107, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000435.

Maleknejad:2005:NSS

[MM05a] K. Maleknejad and F. Mirzaee. Numerical solution of stochastic linear heat conduction problem by using new algorithms. Applied Mathematics and Computation, 163(1):97–106, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002127.

Maleknejad:2005:URH

[MM05b] K. Maleknejad and F. Mirzaee. Using rationalized Haar wavelet for solving linear integral equations. Applied Mathematics and Computation, 160(2):579–587, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303011810.

Mena:2005:CIS

[MM05c] J. B. Mena and J. A.

Malpica. Color image segmentation based on three levels of texture statistical evaluation. Applied Mathematics and Computation, 161(1):1–17, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012062.

Morhac:2005:EAM

[MM05d] M. Morháč and V. Matoušek. Exact algorithm of multidimensional circulant deconvolution. Applied Mathematics and Computation, 164(1):155–166, May 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003029.

Makinde:2006:HPA

[MM06a] O. D. Makinde and P. Y.

Mhone. Hermite–Padé approximation approach to MHD Jeffery–Hamel flows. Applied Mathematics and Computation, 181(2):966–972, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001779.

Maleknejad:2006:NMN

[MM06b] K. Maleknejad and H. Mesgarani. Numerical method for a nonlinear boundary integral equation. Applied Mathematics and Computation, 182(2):1006–1009, November 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306004152.

M:2007:SDD

[MM07a] Eduardo Hernández M. and Mark A. McKibben. On state-dependent delay partial neutral functional– differential equations. Applied Mathematics and Computation, 186(1):294– 301, March 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306009404.

Mabrouk:2007:NSS

[MM07b] Anouar Ben Mabrouk and Mohamed Lakdar Ben Mohamed. Nodal solutions for some nonlinear elliptic equations. Applied Mathematics and Computation, 186(1):589–597, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009702.

Maiti:2007:DWS

[MM07c] Manas Kumar Maiti and Manoranjan Maiti. Determination of withdrawal schedule in single-species cultivation via genetic algorithm. Applied Mathematics and Computation, 188 (1):322–331, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306013555.

Makui:2007:BEL

[MM07d] Ahmad Makui and Alireza Madadi. The bullwhip effect and Lyapunov exponent. Applied Mathematics and Computation, 189(1):35– 40, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601575X.

Meng:2007:ORL

[MM07e] Fanwei Meng and Cuiqin Ma. Oscillation results for linear second order matrix differential systems with damping. Applied Mathematics and Computation, 187(2):844–855, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012161.

Moitsheki:2007:SIS

[MM07f] R. J. Moitsheki and O. D. Makinde. Some invariant solutions for a microwave heating model. Applied Mathematics and Computation, 191(2):308–313, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002585.

Ma:2008:ESC

[MM08a] Ruyun Ma and Huili Ma. Existence of sign-changing periodic solutions of second order difference equations. Applied Mathematics and Computation, 203(2):463–470, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004323.

Marcilio:2008:ALA

[MM08b] Débora Cintia Marcilio and Luiz Carlos Matioli. Augmented Lagrangian applied to convex quadratic problems. Applied Mathematics and Computation, 200(2):480– 485, July 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011289.

Mitchell:2008:ASM

[MM08c] S. L. Mitchell and T. G.

Myers. Approximate solution methods for onedimensional solidification from an incoming fluid. Applied Mathematics and Computation, 202(1):311– 326, August 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308001148.

Maleknejad:2007:NSV

[MMA07] K. Maleknejad, R. Mollapourasl, and M. Alizadeh. Numerical solution of Volterra type integral equation of the first kind with wavelet basis. Applied Mathematics and Computation, 194(2):400–405, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700481X.

Mahdavi:2008:DMF

[MMAHN08] Iraj Mahdavi, Nezam Mahdavi-Amiri, Armaghan Heidarzade, and Rahele Nourifar. Designing a model of fuzzy TOPSIS in multiple criteria decision making. Applied Mathematics and Computation, 206(2):607–617, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003846.

Mokhtarzadeh:2006:BCE

[MMC06] M. R. Mokhtarzadeh, R. Mokhtary, and N. G. Chegini. Bi-CG: an effective solver for three fields domain decomposition method in parallel environments. Applied Mathematics and Computation, 174(2):1196–1205, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005412.

Massad:2008:SFN

[MMC^{+}08] Eduardo Massad, Stefen Ma, Mark Chen, Cláudio José Struchiner, Nico Stollenwerk, and Maíra Aguiar. Scale-free network of a dengue epidemic. Applied Mathematics and Computation, 195(2):376–381, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005528.

Maiti:2006:TSI

[MMM06] A. K. Maiti, M. K. Maiti, and M. Maiti. Two storage inventory model with random planning horizon. Applied Mathematics and Computation, 183(2):1084–1097, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006436.

Maleknejad:2006:CNS

[MMN06a] K. Maleknejad, R. Mollapourasl, and K. Nouri. Convergence of numerical solution of the Fredholm integral equation of the first kind with degenerate kernel. Applied Mathematics and Computation, 181(2):1000–1007, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001834.

Momani:2006:DMSa

[MMN06b] Shaher Momani, Khaled Moadi, and Muhammad Aslam Noor. Decomposition method for solving a system of fourth-order obstacle boundary value problems. Applied Mathematics and Computation, 175(2):923–931, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006533.

Mabrouk:2007:FDA

[MMO07] Anouar Ben Mabrouk, Mohamed Lakdar Ben Mohamed, and Khaled Omrani. Finite difference approximate solutions for a mixed sub-superlinear equation. Applied Mathematics and Computation, 187(2):1007–1016, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012379.

Momoniat:2009:NIL

[MMR09] E. Momoniat, R. McIntyre, and R. Ravindran. Numerical inversion of a Laplace transform solution of a diffusion equation with a mixed derivative term. Applied Mathematics and Computation, 209(2):222–229, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009387.

Miret:2006:ACV

[MMS^{+}06] J. Miret, R. Moreno, D. Sadornil, J. Tena, and M. Valls. An algorithm to compute volcanoes of 2isogenies of elliptic curves over finite fields. Applied Mathematics and Computation, 176(2):739–750, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500874X.

Miret:2008:CHV

[MMS^{+}08] J. Miret, R. Moreno, D. Sadornil, J. Tena, and M. Valls. Computing the height of volcanoes of ℓ-isogenies of elliptic curves over finite fields. Applied Mathematics and Computation, 196(1):67–76, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006455.

Makinde:2007:SRE

[MMT07] O. D. Makinde, R. J.

Moitsheki, and B. A. Tau. Similarity reductions of equations for river pollution. Applied Mathematics and Computation, 188(2):1267–1273, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014731.

Ma:2009:CAD

[MMW09] Mingshu Ma, Wenjuan Ma, and Xiaofeng Wang. A compact alternate direct implicit difference method for solving parabolic equation of multi-dimension. Applied Mathematics and Computation, 212(2):281– 286, June 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300308008795.

Mehri:2005:SCR

[MN05] B. Mehri and M. A. Niksirat. Some computable results on the existence of periodic solutions for singular non-autonomous third order systems. Applied Mathematics and Computation, 163(1):51– 60, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002073.

Momani:2006:NMF

[MN06] Shaher Momani and Muhammad Aslam Noor. Numerical methods for fourthorder fractional integrodifferential equations. Applied Mathematics and Computation, 182(1):754– 760, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306003845.

Momani:2007:NCM

[MN07] Shaher Momani and Muhammad Aslam Noor. Numerical comparison of methods for solving a special fourth-order boundary value problem. Applied Mathematics and Computation, 191(1):218–224, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002482.

Mahapatra:2009:ASM

[MNG09] T. Ray Mahapatra, S. K.

Nandy, and A. S. Gupta. Analytical solution of magnetohydrodynamic stagnationpoint flow of a power-law fluid towards a stretching surface. Applied Mathematics and Computation, 215(5):1696–1710, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006699.

Mitra:2007:SDR

[MNM07] Sharmishtha Mitra, Basab Nandi, and Amit Mitra. Study of dynamic relationships between financial and real sectors of economies with wavelets. Applied Mathematics and Computation, 188(1):83– 95, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306013245.

Modarres:2005:FLR

[MNN05] M. Modarres, E. Nasrabadi, and M. M. Nasrabadi. Fuzzy linear regression models with least square errors. Applied Mathematics and Computation, 163(2):977–989, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002747.

Maleknejad:2007:DCL

[MNY07] K. Maleknejad, K. Nouri, and M. Yousefi. Discussion on convergence of Legendre polynomial for numerical solution of integral equations. Applied Mathematics and Computation, 193(2):335–339, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700402X.

Malek:2005:NSC

[MO05] A. Malek and H. G.

Oskoei. Numerical solutions for constrained quadratic problems using high-performance neural networks. Applied Mathematics and Computation, 169(1):451–471, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007854.

Momani:2006:AST

[MO06] Shaher Momani and Zaid Odibat. Analytical solution of a time-fractional Navier–Stokes equation by Adomian decomposition method. Applied Mathematics and Computation, 177(2):488–494, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009276.

Mestrovic:2007:ARE

[MO07] Mladen Meštrović and Eva Ocvirk. An application of Romberg extrapolation on quadrature method for solving linear Volterra integral equations of the second kind. Applied Mathematics and Computation, 194(2):389–393, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004791.

Moazeni:2006:FSP

[Moa06] Somayeh Moazeni. Fuzzy shortest path problem with finite fuzzy quantities. Applied Mathematics and Computation, 183(1):160–169, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005492.

Mohanty:2005:OST

[Moh05a]  R. K. Mohanty.  An

operator splitting technique for an unconditionally stable difference method for a linear three space dimensional hyperbolic equation with variable coefficients. Applied Mathematics and Computation, 162(2):549–557, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001146.

Mohanty:2005:USF

[Moh05b] R. K. Mohanty. An unconditionally stable finite difference formula for a linear second order one space dimensional hyperbolic equation with variable coefficients. Applied Mathematics and Computation, 165(1):229–236, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003777.

Mohanty:2006:CNU

[Moh06] R. K. Mohanty. A class of non-uniform mesh three point arithmetic average discretization for y′ = f(x, y, y′) and the estimates of y′. Applied Mathematics and Computation, 183(1):477–485, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005819.

Mohanty:2007:IHA

[Moh07a] R. K. Mohanty. An implicit high accuracy variable mesh scheme for 1-D non-linear singular parabolic partial differential equations. Applied Mathematics and Computation, 186(1):219–229, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009325.

Mohanty:2007:SBA

[Moh07b]  R. K. Mohanty.  The

smart-BLAGE algorithm for singularly perturbed 2D elliptic partial differential equations. Applied Mathematics and Computation, 190(1):321–331, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000525.

Mohanty:2007:SIE

[Moh07c] R. K. Mohanty. Stability interval for explicit difference schemes for multidimensional second-order hyperbolic equations with significant first-order space derivative terms. Applied Mathematics and Computation, 190(2):1683–1690, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002135.

Mohanty:2009:VMC

[Moh09] R. K. Mohanty. A variable mesh C-SPLAGE method of accuracy O(k^{2}h_{ℓ}^{−1} + kh_{ℓ} + h_{ℓ}^{3}) for 1D nonlinear parabolic equations. Applied Mathematics and Computation, 213(1):79– 91, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002045.

Momani:2005:AASb

[Mom05a] Shaher Momani. Analytic and approximate solutions of the space- and time-fractional telegraph equations. Applied Mathematics and Computation, 170(2):1126–1134, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000536.

Momani:2005:AASa

[Mom05b] Shaher Momani. Analytical approximate solution for fractional heatlike and wave-like equations with variable coefficients using the decomposition method. Applied Mathematics and Computation, 165(2):459–472, June 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003959.

Momani:2005:NSSb

[Mom05c] Shaher Momani. A numerical scheme for the solution of Sivashinsky equation. Applied Mathematics and Computation, 168(2):1273–1280, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304007015.

Momani:2006:NSS

[Mom06] Shaher Momani. A numerical scheme for the solution of multi-order fractional differential equations. Applied Mathematics and Computation, 182(1):761–770, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003857.

Mondal:2007:DSS

[Mon07] Biswajit Mondal. Discussion on single species fish population with time dependent growth rate. Applied Mathematics and Computation, 194(2):514– 519, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307005036.

Morillas:2005:DAS

[Mor05] Patricia Mariela Morillas. Dykstra’s algorithm with strategies for projecting onto certain polyhedral cones. Applied Mathematics and Computation, 167(1):635–649, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400520X.

Mordukhovich:2008:SMD

[Mor08a] Boris Mordukhovich. Suboptimal minimax design of constrained parabolic systems with mixed boundary control. Applied Mathematics and Computation, 204(2):580–588, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003056.

Morhac:2008:EFA

[Mor08b] Miroslav Morháč. Errorfree algorithms to solve special and general discrete systems of linear equations. Applied Mathematics and Computation, 202(1):1–23, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010272.

Morigi:2009:FSP

[Mor09] S. Morigi. Feature-sensitive parameterization of polygonal meshes. Applied Mathematics and Computation, 215(4):1561–1572, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006596.

Moudafi:2008:DTM

[Mou08] Abdellatif Moudafi. On the difference of two maximal monotone operators: Regularization and algorithmic approaches. Applied Mathematics and Computation, 202(2):446–452, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308000702.

Massoudi:2008:MSG

[MP08a] Mehrdad Massoudi and Tran X. Phuoc. On the motion of a second grade fluid due to longitudinal and torsional oscillations of a cylinder: A numerical study. Applied Mathematics and Computation, 203(2):471–481, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004335.

Maurer:2008:DOM

[MP08b] Helmut Maurer and Hans Josef Pesch. Direct optimization methods for solving a complex state-constrained optimal control problem in microeconomics. Applied Mathematics and Computation, 204(2):568–579, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003044.

Munyakazi:2008:REF

[MP08c] Justin B. Munyakazi and Kailash C. Patidar. On Richardson extrapolation for fitted operator finite difference methods. Applied Mathematics and Computation, 201(1–2): 465–480, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307012222.

Majak:2009:WFB

[MPEL09] Jüri Majak, Meelis Pohlak, Martin Eerme, and Toomas Lepikult. Weak formulation based Haar wavelet method for solving differential equations. Applied Mathematics and Computation, 211(2):488–494, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001076.

Mainardi:2007:SAF

[MPG07] Francesco Mainardi, Gianni Pagnini, and Rudolf Gorenflo. Some aspects of fractional diffusion equations of single and distributed order. Applied Mathematics and Computation, 187(1):295–305, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011647.

Maiti:2008:ETD

[MPS08] A. Maiti, A. K. Pal, and G. P. Samanta. Effect of time-delay on a food chain model. Applied Mathematics and Computation, 200(1):189–203, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010995.

Maurer:2008:NAD

[MPT08] Helmut Maurer, Tapio Palokangas, and Alexander Tarasyev. New approaches in dynamic optimization to assessment of economic and environmental systems: Introduction. Applied Mathematics and Computation, 204(2):517–518, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003032.

Manyu:2007:PIM

[MQ07] Xiao Manyu and Lu Quanyi.

A parallel iterative method for solving periodical blocktridiagonal linear equations. Applied Mathematics and Computation, 184(2):599–607, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007338.

Mo:2005:MSC

[MQW05] Jiangtao Mo, Liqun Qi, and Zengxin Wei. A manufacturing supply chain optimization model for distilling process. Applied Mathematics and Computation, 171(1):464–485, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001098.

McGough:2005:PBR

[MR05a] Jeff S. McGough and Kyle L. Riley. A priori bounds for reaction– diffusion systems arising in chemical and biological dynamics. Applied Mathematics and Computation, 163(1):1–16, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303011044.

Moustafa:2005:EMT

[MR05b] H. M. Moustafa and S. G. Ramadan. Errors of misclassification and their probability distributions when the parent populations are Gompertz. Applied Mathematics and Computation, 163(1):423– 442, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002346.

Mahapatra:2006:FMO

[MR06a] G. S. Mahapatra and T. K. Roy. Fuzzy multi-objective mathematical programming on reliability optimization model. Applied Mathematics and Computation, 174(1):643–659, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004911.

Maleknejad:2006:NSFb

[MR06b] K. Maleknejad and M. Rabbani. Numerical solution for the Fredholm integral equation of the second kind with Toeplitz kernels by using preconditioners. Applied Mathematics and Computation, 180(1):128–133, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305011033.

Mandal:2006:MII

[MR06c] Nirmal Kumar Mandal and Tapan Kumar Roy. Multiitem imperfect production lot size model with hybrid number cost parameters. Applied Mathematics and Computation, 182(2):1219–1230, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600436X.

Malkowsky:2007:MDT

[MR07a] Eberhard Malkowsky and Vladimir Rakočević. On matrix domains of triangles. Applied Mathematics and Computation, 189(2):1146–1163, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016997.

Matucci:2007:NSS

[MŘ07b] Serena Matucci and Pavel Řehák. Nonoscillatory solutions of a second-order nonlinear discrete system. Applied Mathematics and Computation, 190(1):833– 845, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001154.

Mir:2007:TOF

[MR07c] N. A. Mir and Naila Rafiq. Third-order and fourthorder iterative methods for finding multiple and distinct zeros of nonlinear equations. Applied Mathematics and Computation, 190(1):432–440, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000732.

Mittal:2009:EDP

[MR09] Reena Mittal and U. S. Rana. Effect of dust particles on a layer of micropolar ferromagnetic fluid heated from below saturating a porous medium. Applied Mathematics and Computation, 215(7):2591–2607, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007978.

Matbuly:2009:NFF

[MRN09] M. S. Matbuly, Ola Ragb, and M. Nassar. Natural frequencies of a functionally graded cracked beam using the differential quadrature method. Applied Mathematics and Computation, 215(6):2307–2316, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007553.

Maksay:2006:CMS

[MS06a] Stefan Maksay and Diana M. Stoica. Considerations on modeling some distribution laws. Applied Mathematics and Computation, 175(1):238–246, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500603X.

Mohanty:2006:NFO

[MS06b] R. K. Mohanty and Swarn Singh. A new fourth order discretization for singularly perturbed two dimensional non-linear elliptic boundary value problems. Applied Mathematics and Computation, 175(2):1400–1414, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006867.

Momani:2006:DMSb

[MS06c] Shaher Momani and Nabil Shawagfeh. Decomposition method for solving fractional Riccati differential equations. Applied Mathematics and Computation, 182(2):1083–1092, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004231.

Majid:2007:IFP

[MS07a] Zanariah Abdul Majid and Mohamed Bin Suleiman. Implementation of fourpoint fully implicit block method for solving ordinary differential equations. Applied Mathematics and Computation, 184(2):514–522, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007272.

Maleknejad:2007:NSFb

[MS07b] K. Maleknejad and S. Sohrabi.

Numerical solution of Fredholm integral equations of the first kind by using Legendre wavelets. Applied Mathematics and Computation, 186(1):836–843, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601006X.

Marosevic:2007:AFI

[MS07c] Tomislav Marošević and Rudolf Scitovski. An application of a few inequalities among sequences in electoral systems. Applied Mathematics and Computation, 194(2):480–485, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004985.

Mishra:2007:MMCa

[MS07d] Bimal Kumar Mishra and Dinesh Saini. Mathematical models on computer viruses. Applied Mathematics and Computation, 187(2):929–936, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012252.

Mishra:2007:SEM

[MS07e] Bimal Kumar Mishra and Dinesh Kumar Saini. SEIRS epidemic model with delay for transmission of malicious objects in computer network. Applied Mathematics and Computation, 188(2):1476–1482, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015001.

Maleknejad:2008:NSS

[MS08a] K. Maleknejad and A. Salimi Shamloo. Numerical solution of singular Volterra integral equations system of convolution type by using operational matrices. Applied Mathematics and Computation, 195(2):500–505, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005796.

Malinowska:2008:CCD

[MS08b] Iwona Malinowska and Dominik Szynal. On characterization of certain distributions of k th lower (upper) record values. Applied Mathematics and Computation, 202(1):338–347, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001173.

Mittal:2008:SSF

[MS08c] M. L. Mittal and Uaday Singh. T · C_{1} summability of a sequence of Fourier coefficients. Applied Mathematics and Computation, 204(2):702–706, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005420.

Ma:2009:UEL

[MS09a] Yajun Ma and Jitao Sun.

Uniform eventual Lipschitz stability of impulsive systems on time scales. Applied Mathematics and Computation, 211(1):246– 250, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000472.

Misra:2009:BVF

[MS09b] J. C. Misra and G. C.

Shit. Biomagnetic viscoelastic fluid flow over a stretching sheet. Applied Mathematics and Computation, 210(2):350–361, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308009636.

Martinez-Serrano:2009:DIB

[MSCG09] M. F. Martínez-Serrano and N. Castro-González. On the Drazin inverse of block matrices and generalized Schur complement. Applied Mathematics and Computation, 215(7):2733–2740, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030900811X.

Mazrooei-Sebdani:2006:DNL

[MSD06a] Reza Mazrooei-Sebdani and Mehdi Dehghan. Dynamics of a non-linear difference equation. Applied Mathematics and Computation, 178(2):250–261, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009719.

Mazrooei-Sebdani:2006:GSI

[MSD06b] Reza Mazrooei-Sebdani and Mehdi Dehghan. Global stability of y_{n+1} = (p + qy_{n} + ry_{n−k})/(1 + y_{n}). Applied Mathematics and Computation, 182(1):621– 630, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306003729.

Mazrooei-Sebdani:2006:SCR

[MSD06c] Reza Mazrooei-Sebdani and Mehdi Dehghan. The study of a class of rational difference equations. Applied Mathematics and Computation, 179(1):98–107, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030501012X.

Miao:2006:PPT

[MSD06d] Weihua Miao, Wenyu Sun, and Chuangying Dang. A practical penalty trust-region method for equality-constrained optimization problems. Applied Mathematics and Computation, 174(1):578–587, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004868.

Merabet:2008:AAD

[MSH08] N. Merabet, H. Siyyam, and M. H. Hamdan. Analytical approach to the Darcy–Lapwood–Brinkman equation. Applied Mathematics and Computation, 196(2):679–685, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307007217.

Messaoudi:2007:GEA

[MSHeT07] Salim A. Messaoudi, Belkacem Said-Houari, and Nasser eddine Tatar. Global existence and asymptotic behavior for a fractional differential equation. Applied Mathematics and Computation, 188(2):1955–1962, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015633.

Maleknejad:2005:NSI

[MSK05] K. Maleknejad, M. Shahrezaee, and H. Khatami. Numerical solution of integral equations system of the second kind by Block– Pulse functions. Applied Mathematics and Computation, 166(1):15– 24, July 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004175.

Maeda:2007:QNA

[MSM07a] Michiharu Maeda, Masaya Suenaga, and Hiromi Miyajima. Qubit neuron according to quantum circuit for XOR problem. Applied Mathematics and Computation, 185(2):1015–1025, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008265.

Mittal:2007:SNS

[MSM07b] Madan Lal Mittal, Uaday Singh, and Vishnu N. Mishra. On the strong Nörlund summability of conjugate Fourier series. Applied Mathematics and Computation, 187(1):326– 331, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011672.

Mohsen:2008:NSE

[MSM08] A. Mohsen, L. F. Sedeek, and S. A. Mohamed. New smoother to enhance multigrid-based methods for Bratu problem. Applied Mathematics and Computation, 204(1):325–339, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005110.

Maleknejad:2007:NSN

[MSR07] K. Maleknejad, S. Sohrabi, and Y. Rostami. Numerical solution of nonlinear Volterra integral equations of the second kind by using Chebyshev polynomials. Applied Mathematics and Computation, 188(1):123– 128, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306013282.

Modani:2008:NSE

[MSR08] Manish Modani, Maithili Sharan, and S. Chandra Sekhara Rao. Numerical solution of elliptic partial differential equations on parallel systems. Applied Mathematics and Computation, 195(1):162–182, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005280.

Manojlovic:2006:OCS

[MSTY06] J. Manojlović, Y. Shoukaku, T. Tanigawa, and N. Yoshida. Oscillation criteria for second order differential equations with positive and negative coefficients. Applied Mathematics and Computation, 181(2):853–863, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001652.

Ma:2008:QCS

[MT08] Changfeng Ma and Jia Tang. The quadratic convergence of a smoothing Levenberg–Marquardt method for nonlinear complementarity problem. Applied Mathematics and Computation, 197(2):566– 581, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700803X.

Mustafa:2009:ALS

[MT09] Octavian G. Mustafa and Cemil Tunç. Asymptotically linear solutions of differential equations via Lyapunov functions. Applied Mathematics and Computation, 215(8):3076–3081, December 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300309008881.

Ma:2007:GCL

[MTC07] Changfeng Ma, Jia Tang, and Xiaohong Chen. A globally convergent Levenberg– Marquardt method for solving nonlinear complementarity problem. Applied Mathematics and Computation, 192(2):370– 381, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307003396.

Manolova:2005:RTP

[MTD05] Petia Manolova, Charles Lai Tong, and Christophe Deissenberg. Real taxation and production in a monetary economy with spatially differentiated agents. Applied Mathematics and Computation, 164(2):591–603, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030400339X.

Ma:2007:PSN

[MTL07] Manjun Ma, Hengsheng Tang, and Wei Luo. Periodic solutions for nonlinear second-order difference equations. Applied Mathematics and Computation, 184(2):685–694, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007442.

Mu:2008:SPS

[MTL08] Chunlai Mu, Ya Tian, and Yuhuan Li. Support properties of solution to degenerate equation with variable coefficient and absorption. Applied Mathematics and Computation, 198(2):824–832, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009551.

Mamedov:2005:CGE

[MTM05] B. A. Mamedov, R. Tapramaz, and Z. Merdan. Calculation of generalized elliptic type integrals using the binomial expansion theorem. Applied Mathematics and Computation, 168(1):333–341, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006071.

Tang:2007:NVL

[mTxWyL07] Chun ming Tang, Zeng xin Wei, and Guo yin Li. A new version of the Liu–Storey conjugate gradient method. Applied Mathematics and Computation, 189(1):302–313, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016043.

Mugdadi:2005:LST

[Mug05]  A. R. Mugdadi.  The

least squares type estimation of the parameters in the power hazard function. Applied Mathematics and Computation, 169(2):737–748, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008070.

Mukaidani:2007:ENP

[Muk07a] Hiroaki Mukaidani. Efficient numerical procedures for solving closedloop Stackelberg strategies with small singular perturbation parameter. Applied Mathematics and Computation, 188(2):1173–1183, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014615.

Mukaidani:2007:NMS

[Muk07b] Hiroaki Mukaidani. Newton’s method for solving cross-coupled signindefinite algebraic Riccati equations for weakly coupled large-scale systems. Applied Mathematics and Computation, 188(1):103– 115, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013269.

Mukaidani:2007:NAF

[Muk07c]  Hiroaki Mukaidani.  A

numerical algorithm for finding solution of signindefinite algebraic Riccati equations for general multiparameter singularly perturbed systems. Applied Mathematics and Computation, 189(1):255–270, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015992.

Mukaidani:2008:NCH

[Muk08a] Hiroaki Mukaidani. Numerical computation for H_{∞} output feedback control for strongly coupled large-scale systems. Applied Mathematics and Computation, 197(1):212– 227, March 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307007710.

Mukaidani:2008:NCC

[Muk08b] Hiroaki Mukaidani. Numerical computation of cross-coupled algebraic Riccati equations related to H_{∞}-constrained LQG control problem. Applied Mathematics and Computation, 199(2):663–676, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010545.

Murty:2006:ETM

[Mur06] Y. Narasimha Murty. Effect of throughflow and magnetic field on Marangoni convection in micro-polar fluids. Applied Mathematics and Computation, 173(2):1288–1299, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305004315.

Mustafa:2007:OSC

[Mus07] Octavian G. Mustafa. On oscillatory solutions of certain first order ordinary differential equations. Applied Mathematics and Computation, 190(2):1290–1295, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001609.

Mustafa:2008:NOI

[Mus08] Octavian G. Mustafa.

A note on oscillatory integration. Applied Mathematics and Computation, 199(2):637–643, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030701051X.

Magosso:2009:WBE

[MUZG09] Elisa Magosso, Mauro Ursino, Anna Zaniboni, and Elena Gardella. A wavelet-based energetic approach for the analysis of biomedical signals: Application to the electroencephalogram and electrooculogram. Applied Mathematics and Computation, 207(1):42–62, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000180.

Melton:2008:IGQ

[MV08] Tanya G. Melton and A. S.

Vatsala. Improved generalized quasilinearization method and rapid convergence for reaction diffusion equations. Applied Mathematics and Computation, 203(2):563–572, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800283X.

Martinez:2009:BCD

[MV09a] Antonio Martinez and Jose C. Valverde. Bifurcations of continuous dynamical systems under extreme and new transversality conditions. Applied Mathematics and Computation, 210(1):33– 38, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007856.

Mitchell:2009:FDM

[MV09b] S. L. Mitchell and M. Vynnycky. Finite-difference methods with increased accuracy and correct initialization for one-dimensional Stefan problems. Applied Mathematics and Computation, 215(4):1609–1621, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006638.

Ma:2007:MTD

[MW07] Changfeng Ma and Desheng Wang. A method for time-dependent Maxwell’s equations based on explicitmagnetic-field scheme. Applied Mathematics and Computation, 190(2):1273– 1283, July 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307001580.

Ma:2009:KBA

[MW09] Limin Ma and Zongmin Wu. Kernel based approximation in Sobolev spaces with radial basis functions. Applied Mathematics and Computation, 215(6):2229–2237, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007358.

Ma:2009:NEC

[MWQ09] Hong-Cai Ma, Yan Wang, and Zhen-Yun Qin. New exact complex traveling wave solutions for (2 + 1)-dimensional BKP equation. Applied Mathematics and Computation, 208(2):564–568, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308009326.

Miele:2005:FLV

[MWW05] A. Miele, T. Wang, and P. N. Williams. On feasibility of launch vehicle designs. Applied Mathematics and Computation, 164(2):295–312, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003121.

Meng:2007:KTO

[MX07a] Fanwei Meng and Run Xu. Kamenev-type oscillation criteria for even order neutral differential equations with deviating arguments. Applied Mathematics and Computation, 190(2):1402–1408, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001713.

Meng:2007:OCC

[MX07b] Fanwei Meng and Run Xu. Oscillation criteria for certain even order quasi-linear neutral differential equations with deviating arguments. Applied Mathematics and Computation, 190(1):458–464, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700077X.

Ma:2008:HIN

[MX08a] Lin-Lin Ma and Xi-Xiang Xu. A hierarchy of integrable nonlinear lattice equations and its two integrable coupling systems. Applied Mathematics and Computation, 203(1):7–13, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001951.

Ma:2008:GDS

[MX08b] Zhanyou Ma and Qingzhen Xu. General decrementing service M / G / 1 queue with multiple adaptive vacations. Applied Mathematics and Computation, 204(1):478–484, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005353.

Meng:2007:PDN

[MXC07] Xinzhu Meng, Weijian Xu, and Lansun Chen. Profitless delays for a nonautonomous Lotka– Volterra predator–prey almost periodic system with dispersion. Applied Mathematics and Computation, 188(1):365–378, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013609.

Malek:2005:PDS

[MY05] A. Malek and A. Yari. Primal–dual solution for the linear programming problems using neural networks. Applied Mathematics and Computation, 167(1):198–211, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004849.

Maleknejad:2006:CMS

[MY06a] K. Maleknejad and M. R. Fadaei Yami. A computational method for system of Volterra–Fredholm integral equations. Applied Mathematics and Computation, 183(1):589–595, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005935.

Maleknejad:2006:NSIb

[MY06b] K. Maleknejad and M. Yousefi.

Numerical solution of the integral equation of the second kind by using wavelet bases of Hermite cubic splines. Applied Mathematics and Computation, 183(1):134–141, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005467.

Miao:2007:EPS

[MY07a] Qing Miao and Zuodong Yang. Entire positive solution to the system of quasilinear elliptic equations. Applied Mathematics and Computation, 192(2):496–506, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003530.

Min:2007:PFD

[MY07b] Tian Min and Danping Yang. Parallel finite difference schemes for heat equation based upon overlapping domain decomposition. Applied Mathematics and Computation, 186(2):1276–1292, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010538.

Miao:2008:EMP

[MY08] Qing Miao and Zuodong Yang. On the existence of multiple positive entire solutions for a quasilinear elliptic systems. Applied Mathematics and Computation, 198(1):12– 23, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008624.

Mycka:2006:ACB

[Myc06]  Jerzy Mycka.  Analog

computation beyond the Turing limit. Applied Mathematics and Computation, 178(1):103–117, July 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008386.

Ma:2008:NET

[MYG08] Hong-Cai Ma, Yao-Dong Yu, and Dong-Jie Ge. New exact traveling wave solutions for the modified form of Degasperis–Procesi equation. Applied Mathematics and Computation, 203(2):792–798, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004050.

Maleknejad:2007:CMI

[MYN07] K. Maleknejad, M. Yousefi, and K. Nouri. Computational methods for integrals involving functions and Daubechies wavelets. Applied Mathematics and Computation, 189(2):1828–1840, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017796.

Mahdy:2005:MAC

[MYSA05] Y. B. Mahdy, A. Younes, M. A. Soliman, and M. H. Abdellha. A modified algorithm for computing the upper-bound reliabilty of computer networks. Applied Mathematics and Computation, 160(2):523–538, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303011792.

Mir:2007:SQB

[MZ07] Nazir Ahmad Mir and Tooba Zaman. Some quadrature based threestep iterative methods for non-linear equations. Applied Mathematics and Computation, 193(2):366– 373, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307004079.

Mei:2009:SRE

[MZ09] Jianqin Mei and Hongqing Zhang. Symmetry reductions and explicit solutions of a (3 + 1)- dimensional PDE. Applied Mathematics and Computation, 211(2):347–353, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000800.

Mu:2007:NBB

[MZL07] Xuewen Mu, Yaling Zhang, and Sanyang Liu. A new branch and bound method with pretreatment for the binary quadratic programming. Applied Mathematics and Computation, 192(1):252–259, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003293.

See erratum [Pak11].

Mo:2005:NTR

[MZW05] Jiangtao Mo, Kecun Zhang, and Zengxin Wei. A nonmonotone trust region method for unconstrained optimization. Applied Mathematics and Computation, 171(1):371–384, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001025.

Ma:2008:DDS

[MZW08] Shuping Ma, Chenghui Zhang, and Zhen Wu. Delay-dependent stability and H_{∞} control for uncertain discrete switched singular systems with timedelay. Applied Mathematics and Computation, 206(1):413–424, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006917.

Zhang:2005:NAS

[mZzH05] Qi min Zhang and Chong zhao Han. Numerical analysis for stochastic age-dependent population equations. Applied Mathematics and Computation, 169(1):278–294, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030400774X.

Noor:2006:FOC

[NA06a] Muhammad Aslam Noor and Faizan Ahmad. Fourthorder convergent iterative method for nonlinear equation. Applied Mathematics and Computation, 182(2):1149–1153, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004292.

Noor:2006:NCI

[NA06b] Muhammad Aslam Noor and Faizan Ahmad. Numerical comparison of iterative methods for solving nonlinear equations. Applied Mathematics and Computation, 180(1):167– 172, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305011094.

Nguetchue:2008:CAS

[NA08] S. N. Neossi Nguetchue and S. Abelman. A computational algorithm for solving nearly penta-diagonal linear systems. Applied Mathematics and Computation, 203(2):629–634, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002932.

Nabli:2009:OSA

[Nab09] Hédi Nabli. An overview on the simplex algorithm. Applied Mathematics and Computation, 210(2):479–489, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000198.

Nacevska:2009:IMC

[Nač09] Biljana Načevska. Iterative methods for computing generalized inverses and splittings of operators. Applied Mathematics and Computation, 208(1):186–188, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008849.

Nadarajah:2005:PRB

[Nad05] Saralees Nadarajah. Products, and ratios for a bivariate gamma distribution. Applied Mathematics and Computation, 171(1):581–595, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305001189.

Nadarajah:2006:FIE

[Nad06a] Saralees Nadarajah. Fisher information for the elliptically symmetric Pearson distributions. Applied Mathematics and Computation, 178(2):195–206, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009677.

Nadarajah:2006:IMB

[Nad06b] Saralees Nadarajah. Information matrix for the bivariate Gumbel distribution. Applied Mathematics and Computation, 172(1):394–405, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305002109.

Nadarajah:2007:LE

[Nad07a] Saralees Nadarajah. Letter to the Editor. Applied Mathematics and Computation, 190(1):987, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000252.

Nadarajah:2007:DMK

[Nad07b] Saralees Nadarajah. On the distribution of Ma and King. Applied Mathematics and Computation, 189 (1):732–733, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030601650X. See

comment [Ma11].

Noor:2006:TSIa

[NAJ06] Muhammad Aslam Noor, Faizan Ahmad, and Shumaila Javeed. Two-step iterative methods for nonlinear equations. Applied Mathematics and Computation, 181(2):1068–1075, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001901.

Nakamori:2007:CTR

[Nak07] Seiichi Nakamori. Chandrasekhartype recursive Wiener estimation technique in linear discrete-time stochastic systems. Applied Mathematics and Computation, 188(2):1656–1665, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015189.

Nakamori:2008:SRA

[Nak08]  S. Nakamori.  Square-

root algorithms of RLS Wiener filter and fixedpoint smoother in linear discrete stochastic systems. Applied Mathematics and Computation, 203(1):186–193, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002269.

Nilam:2007:MEC

[NAM^{+}07] Nilam, Murray E. Alexander, Ruchi Mathur, Seyed M. Moghadas, and Pappur N. Shivakumar. Modelling the effect of CSII on the control of glucose concentration in type 1 diabetes. Applied Mathematics and Computation, 187(2):1476–1483, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012860.

Nam:2008:ICE

[Nam08] Phan T. Nam. An improved criterion for exponential stability of linear systems with multiple time delays. Applied Mathematics and Computation, 202(2):870–876, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001914.

Nam:2009:ESC

[Nam09] Phan T. Nam. Exponential stability criterion for timedelay systems with nonlinear uncertainties. Applied Mathematics and Computation, 214(2):374–380, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003233.

Narayaninsamy:2007:FIP

[Nar07a] T. Narayaninsamy. Fractional iterates for piecewise differentiable maps. Applied Mathematics and Computation, 192(1):274– 279, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307003037.

Narayaninsamy:2007:MCI

[Nar07b]  T. Narayaninsamy.  A

method to construct irregular fractal curves. Applied Mathematics and Computation, 192(1):260–273, September 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307002925.

Nataraj:2009:MFE

[Nat09] Neela Nataraj. A mixed finite element method for fourth order eigenvalue problems. Applied Mathematics and Computation, 213(1):60–72, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030900201X.

Noor:2009:UCC

[NAUH09] Khalida Inayat Noor, Mohammad Arif, and Wasim Ul-Haq. On kuniformly close-to-convex functions of complex order. Applied Mathematics and Computation, 215(2):629–635, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309005414.

Navarro:2008:CPS

[Nav08a] Juan F. Navarro. Computation of periodic solutions in perturbed secondorder ODEs. Applied Mathematics and Computation, 202(1):171–177, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000714.

Navarro:2008:IPL

[Nav08b] Juan F. Navarro. On the implementation of the Poincaré–Lindstedt technique. Applied Mathematics and Computation, 195(1):183–189, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005292.

Noor:2007:IAG

[NB07] Muhammad Aslam Noor and Abdellah Bnouhachem. On an iterative algorithm for general variational inequalities. Applied Mathematics and Computation, 185(1):155–168, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008228.

Noor:2009:AFR

[NB09] Khalida Inayat Noor and Syed Zakar Hussain Bukhari. On analytic functions related with generalized Robertson functions. Applied Mathematics and Computation, 215(8):2965– 2970, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309008686.

Nguyen-Ba:2009:OSS

[NBHYV09] Truong Nguyen-Ba, Han Hao, Hemza Yagoub, and Rémi Vaillancourt. Onestep 5-stage Hermite– Birkhoff–Taylor ODE solver of order 12. Applied Mathematics and Computation, 211(2):313–328, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000770.

Nikkhah-Bahrami:2009:EIM

[NBO09] M. Nikkhah-Bahrami and R. Oftadeh. An effective iterative method for computing real and complex roots of systems of nonlinear equations. Applied Mathematics and Computation, 215(5):1813–1820, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900681X.

Nakamori:2006:LSO

[NCÁHC^{+}06] S. Nakamori, R. Caballero-Águila, A. Hermoso-Carazo, J. Jiménez-López, and J. Linares-Pérez. Leastsquares νth-order polynomial estimation of signals from observations affected by non-independent uncertainty. Applied Mathematics and Computation, 176(2):642–653, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008623.

Nakamori:2005:RES

[NCÁHCLP05a] S. Nakamori, R. Caballero-Águila, A. Hermoso-Carazo, and J. Linares-Pérez. Recursive estimators of signals from measurements with stochastic delays using covariance information. Applied Mathematics and Computation, 162(1):65– 79, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000566.

Nakamori:2005:DRW

[NCÁHCLP05b] Seiichi Nakamori, R. Caballero-Águila, A. Hermoso-Carazo, and J. Linares-Pérez. Design of recursive Wiener fixed-point smoothers based on innovations approach in linear discrete-time stochastic systems. Applied Mathematics and Computation, 165(3):731–747, June 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400414X.

Navarro:2007:GAD

[NCGdSM07] Hélio Aparecido Navarro, Luben Cabezas-Gómez, Renato César da Silva, and Arlindo Neto Montagnoli. A generalized alternatingdirection implicit scheme for incompressible magnetohydrodynamic viscous flows at low magnetic Reynolds number. Applied Mathematics and Computation, 189(2):1601–1613, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017450.

Nemeti:2006:RCT

[ND06] István Németi and Gyula Dávid. Relativistic computers and the Turing barrier. Applied Mathematics and Computation, 178 (1):118–142, July 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300305008398.

Natesan:2007:RCM

[ND07] Srinivasan Natesan and Briti Sundar Deb. A Robust computational method for singularly perturbed coupled system of reaction– diffusion boundary-value problems. Applied Mathematics and Computation, 188(1):353–364, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306013592.

Nasri:2005:SSN

[NDD05] Mostafa Nasri, Mehdi Dehghan, and Majid Jaberi Douraki. Study of a system of non-linear difference equations arising in a deterministic model for HIV infection. Applied Mathematics and Computation, 171(2):1306–1330, December 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nedzhibov:2005:AIP

[Ned05] Gyurhan H. Nedzhibov.

An acceleration of iterative processes for solving nonlinear equations. Applied Mathematics and Computation, 168(1):320– 332, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030400606X.

Neher:2007:NSA

[Neh07a] Markus Neher. A note on a sum associated with the generalized hypergeometric function. Applied Mathematics and Computation, 187(2):1527–1534, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012938.

Neher:2007:VIL

[Neh07b]  Markus Neher.  Veri-

fied integration of linear n th order ODEs using large steps. Applied Mathematics and Computation, 186(1):879–890, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010101.

Neta:2008:NTO

[Net08a] B. Neta. New third order nonlinear solvers for multiple roots. Applied Mathematics and Computation, 202(1):162–170, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000696.

Neta:2008:PMN

[Net08b] B. Neta. On Popovski’s method for nonlinear equations. Applied Mathematics and Computation, 201(1–2):710–715, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000374.

Najafi:2006:MMR

[NG06a] H. Saberi Najafi and H. Ghazvini. A modification on minimum restarting method in the Arnoldi algorithm for computing the eigenvalues of a nonsymmetric matrix. Applied Mathematics and Computation, 181(2):1455–1461, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002748.

Najafi:2006:WRM

[NG06b] H. Saberi Najafi and H. Ghazvini. Weighted restarting method in the weighted Arnoldi algorithm for computing the eigenvalues of a nonsymmetric matrix. Applied Mathematics and Computation, 175(2):1276–1287, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006776.

Noor:2007:MHIb

[NG07] Muhammad Aslam Noor and Vijay Gupta. Modified Householder iterative method free from second derivatives for nonlinear equations. Applied Mathematics and Computation, 190(2):1701–1706, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002159.

Nikas:2009:BZI

[NG09] I. A. Nikas and T. N. Grapsa. Bounding the zeros of an interval equation. Applied Mathematics and Computation, 213(2):466–478, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002586.

Nakamori:2006:DFI

[NGLHCLP06] S. Nakamori, M. J. García-Ligero, A. Hermoso-Carazo, and J. Linares-Pérez. Derivation of fixed-interval smoothing algorithm using covariance information in distributed parameter systems. Applied Mathematics and Computation, 176(2):662–672, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008659.

Ngo:2009:SMV

[Ngô09] Quô´ c-Anh Ngô. Some mean value theorems for integrals on time scales. Applied Mathematics and Computation, 213(2):322– 328, July 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309002343.

Noor:2007:SRI

[NH07a] Muhammad Aslam Noor and Zhenyu Huang. Some resolvent iterative methods for variational inclusions and nonexpansive mappings. Applied Mathematics and Computation, 194(1):267–275, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004663.

Noor:2007:TSM

[NH07b] Muhammad Aslam Noor and Zhenyu Huang. Threestep methods for nonexpansive mappings and variational inequalities. Applied Mathematics and Computation, 187(2):680– 685, April 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306011234.

Noor:2007:WHE

[NH07c] Muhammad Aslam Noor and Zhenyu Huang. Wiener– Hopf equation technique for variational inequalities and nonexpansive mappings. Applied Mathematics and Computation, 191(2):504–510, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002780.

Nakamori:2007:DFL

[NHCLP07] Seiichi Nakamori, Aurora Hermoso-Carazo, and Josefa Linares-Pérez. Design of fixed-lag smoother using covariance information based on innovations approach in linear discrete-time stochastic systems. Applied Mathematics and Computation, 193(1):162–174, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003839.

Nine:2006:EWF

[NHKZ06] B. Nine, O. Haif-Khaif, and A. Zerarka. The eigenenergies of the wave function through the nonvariational Galerkin–Bspline approach. Applied Mathematics and Computation, 178(2):486–492, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009951.

Nishimoto:2007:FCP

[Nis07] Katsuyuki Nishimoto. N-Fractional calculus of products of some power functions and some doubly infinite sums. Applied Mathematics and Computation, 187(1):340–349, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011696.

Niu:2008:PSO

[NJG08] Qun Niu, Bin Jiao, and Xingsheng Gu. Particle swarm optimization combined with genetic operators for job shop scheduling problem with fuzzy processing time. Applied Mathematics and Computation, 205(1):148–158, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003251.

Nakhi:2005:GMD

[NK05a] Y. Ben Nakhi and S. L. Kalla. On a generalized mixture distribution. Applied Mathematics and Computation, 169(2):943–952, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008240.

Ng:2005:IRC

[NK05b] Michael K. Ng and Wilson C. Kwan. Image restoration by cosine transform-based iterative regularization. Applied Mathematics and Computation, 160(2):499–515, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303011767.

Nadarajah:2007:SBB

[NK07a] Saralees Nadarajah and Samuel Kotz. Some bivariate Bessel distributions. Applied Mathematics and Computation, 187(1):332– 339, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011684.

Nistazakis:2007:FBA

[NK07b] Hector E. Nistazakis and Athanassios Katsis. A flexible Bayesian algorithm for sample size calculations in misclassified data. Applied Mathematics and Computation, 184(1):86–92, January 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004875.

Najafi:2005:CQC

[NKE05] H. Saberi Najafi, S. Kordrostami, and M. Esmaeilzadeh. Computational Quartic c ^{3}-spline and Quintic c ^{2}-spline algorithms for solving secondorder initial value problems. Applied Mathematics and Computation, 167(1):237–251, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004886.

Noor:2007:NMH

[NKH07] Muhammad Aslam Noor, Waseem Asghar Khan, and Akhtar Hussain. A new modified Halley method without second derivatives for nonlinear equation. Applied Mathematics and Computation, 189(2):1268–1273, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017127.

Navarro:2008:WGM

[NKR^{+}08] Hélio A. Navarro, Magda K. Kaibara, José B. Rubert, Arlindo N. Montagnoli, Luben Cabezas-Gómez, and Renato C. da Silva. Wavelet-Galerkin method for one-dimensional elastoplasticity and damage problems: Constitutive modeling and computational aspects. Applied Mathematics and Computation, 198(2):904–915, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009289.

Ngoc:2005:SSC

[NL05] Pham Huu Anh Ngoc and Byung Soo Lee. Some sufficient conditions for exponential stability of linear neutral functional differential equations. Applied Mathematics and Computation, 170(1):515–530, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009658.

Niakas:2008:ANS

[NLD08] N. Niakas, V. C. Loukopoulos, and C. Douskos. An accurate numerical solver on second order PDEs with variable coefficients in three dimensions. Applied Mathematics and Computation, 204(1):50–57, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004530.

Najafi:2008:NRM

[NM08a] H. Saberi Najafi and H. Moosaei. A new restarting method in the harmonic projection algorithm for computing the eigenvalues of a nonsymmetric matrix. Applied Mathematics and Computation, 198(1):143–149, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008752.

Noor:2008:AFG

[NM08b] Khalida Inayat Noor and Ali Muhammad. On analytic functions with generalized bounded Mocanu variation. Applied Mathematics and Computation, 196(2):802–811, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307007321.

See comments [WZ13].

Navarro:2009:UFD

[NM09] Juan F. Navarro and M. C. Martínez. Universality of fractal dimension on time-independent Hamiltonian systems. Applied Mathematics and Computation, 214(2):462–467, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309003415.

Noor:2007:VIT

[NMD07] Muhammad Aslam Noor and Syed Tauseef Mohyud-Din. Variational iteration technique for solving higher order boundary value problems. Applied Mathematics and Computation, 189(2):1929–1942, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000124.

Naz:2008:CDA

[NMM08] R. Naz, F. M. Mahomed, and D. P. Mason. Comparison of different approaches to conservation laws for some partial differential equations in fluid mechanics. Applied Mathematics and Computation, 205(1):212–230, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004918.

Noor:2009:SCV

[NMM09] Khalida Inayat Noor, Saima Mustafa, and Bushra Malik. On some classes of p-valent functions involving Carlson–Shaffer operator. Applied Mathematics and Computation, 214(2):336–341, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003087.

Niamsup:2008:IES

[NMP08] P. Niamsup, K. Mukdasai, and V. N. Phat. Improved exponential stability for time-varying systems with nonlinear delayed perturbations. Applied Mathematics and Computation, 204(1):490–495, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005389.

Naby:2006:HFG

[NMS06] Abd El Hakeem Abd El Naby, A. E. M. El Misiery, and Ibrahim El Shamy. Hydromagnetic flow of generalized Newtonian fluid through a uniform tube with peristalsis. Applied Mathematics and Computation, 173(2):856–871, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030500398X.

Ntaganda:2007:MBP

[NMS07] Jean Marie Ntaganda, Benjamin Mampassi, and Djibril Seck. Modelling blood partial pressures of the human cardiovascular/ respiratory system. Applied Mathematics and Computation, 187(2):1100–1108, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601246X.

Najafi:2006:GAR

[NN06a] Amir Abbas Najafi and Seyed Taghi A. Niaki. A genetic algorithm for resource investment problem with discounted cash flows. Applied Mathematics and Computation, 183(2):1057–1070, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006412.

Noor:2006:SIS

[NN06b] Muhammad Aslam Noor and Khalida Inayat Noor. Some iterative schemes for nonlinear equations. Applied Mathematics and Computation, 183(2):774– 779, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306006114.

Noor:2006:TSIb

[NN06c] Muhammad Aslam Noor and Khalida Inayat Noor. Three-step iterative methods for nonlinear equations. Applied Mathematics and Computation, 183(1):322–327, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005649.

Noor:2007:IMF

[NN07a] Khalida Inayat Noor and Muhammad Aslam Noor. Iterative methods with fourth-order convergence for nonlinear equations. Applied Mathematics and Computation, 189(1):221– 227, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015943.

Noor:2007:PCH

[NN07b] Khalida Inayat Noor and Muhammad Aslam Noor. Predictor–corrector Halley method for nonlinear equations. Applied Mathematics and Computation, 188(2):1587–1591, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015116.

Noor:2007:FOI

[NN07c] Muhammad Aslam Noor and Khalida Inayat Noor.

Fifth-order iterative methods for solving nonlinear equations. Applied Mathematics and Computation, 188(1):406–410, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013750.

Noor:2007:IIM

[NN07d] Muhammad Aslam Noor and Khalida Inayat Noor. Improved iterative methods for solving nonlinear equations. Applied Mathematics and Computation, 184(2):270–275, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007028.

Noor:2007:MIM

[NN07e] Muhammad Aslam Noor and Khalida Inayat Noor. Modified iterative methods with cubic convergence for solving nonlinear equations. Applied Mathematics and Computation, 184(2):322–325, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007090.

Neuman:2008:MTS

[NN08] Yair Neuman and Ophir Nave. A mathematical theory of sign-mediated concept formation. Applied Mathematics and Computation, 201(1–2):72–81, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011721.

Noor:2006:IMN

[NNKA06] Muhammad Aslam Noor, Khalida Inayat Noor, Waseem Asghar Khan, and Faizan Ahmad. On iterative methods for nonlinear equations. Applied Mathematics and Computation, 183(1):128–133, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005455.

Noor:2007:MHIa

[NNM07] Khalida Inayat Noor, Muhammad Aslam Noor, and Shaher Momani. Modified Householder iterative method for nonlinear equations. Applied Mathematics and Computation, 190(2):1534–1539, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001890.

Noor:2006:IMC

[NNMDS06] Muhammad Aslam Noor, Khalida Inayat Noor, Syed Tauseef Mohyud-Din, and Asim Shabbir. An iterative method with cubic convergence for nonlinear equations. Applied Mathematics and Computation, 183(2):1249–1255, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006709.

Nasrabadi:2005:FLR

[NNN05] Mohammad Mehdi Nasrabadi, Ebrahim Nasrabadi, and Ali Reza Nasrabady. Fuzzy linear regression analysis: a multi-objective programming approach. Applied Mathematics and Computation, 163(1):245–251, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002231.

Noor:2007:TOI

[NNuH07] Muhammad Aslam Noor, Khalida Inayat Noor, and Mahmood ul Hassan. Third-order iterative methods free from second derivatives for nonlinear equation. Applied Mathematics and Computation, 190(2):1551–1556, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001919.

Nishiwaki:2007:CCA

[NO07] Junichi Nishiwaki and Shigeyoshi Owa. Certain classes of analytic functions concerned with uniformly starlike and convex functions. Applied Mathematics and Computation, 187(1):350–355, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011830.

Nardo:2009:CCB

[NO09] E. Di Nardo and I. Oliva. On the computation of classical, Boolean and free cumulants. Applied Mathematics and Computation, 208(2):347–354, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008801.

Noor:2006:IOD

[Noo06a] Khalida Inayat Noor. Integral operators defined by convolution with hypergeometric functions. Applied Mathematics and Computation, 182(2):1872–1881, November 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306005364.

Noor:2006:PCM

[Noo06b] Muhammad Aslam Noor.

Predictor–corrector methods for multi-valued hemiequilibrium problems. Applied Mathematics and Computation, 181(1):721–731, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001512.

Noor:2007:SAC

[Noo07a] Khalida Inayat Noor. On some applications of certain integral operators. Applied Mathematics and Computation, 188(1):814– 823, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014202.

Noor:2007:NCI

[Noo07b] Muhammad Aslam Noor.

New classes of iterative methods for nonlinear equations. Applied Mathematics and Computation, 191(1):128–131, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002354.

See note [Raf07a].

Noor:2007:NFI

[Noo07c] Muhammad Aslam Noor. New family of iterative methods for nonlinear equations. Applied Mathematics and Computation, 190(1):553–558, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000884.

See note [Raf07b].

Noor:2007:NIS

[Noo07d] Muhammad Aslam Noor. New iterative schemes for nonlinear equations. Applied Mathematics and Computation, 187(2):937– 943, April 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306012264.

Noor:2007:SIM

[Noo07e] Muhammad Aslam Noor. Some iterative methods free from second derivatives for nonlinear equations. Applied Mathematics and Computation, 192(1):101–106, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002986.

See note [Raf07a].

Noor:2008:CAF

[Noo08a] Khalida Inayat Noor. On certain analytic functions related with strongly closeto-convex functions. Applied Mathematics and Computation, 197(1):149– 157, March 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307007631.

Noor:2008:DNC

[Noo08b] Muhammad Aslam Noor.

Differentiable non-convex functions and general variational inequalities. Applied Mathematics and Computation, 199(2):623– 630, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010417.

Nedzhibov:2005:FIM

[NP05] Gyurhan H. Nedzhibov and Milko G. Petkov. On a family of iterative methods for simultaneous extraction of all roots of algebraic polynomial. Applied Mathematics and Computation, 162(1):427–433, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001055.

Nabiyev:2008:TSP

[NP08] Vasif V. Nabiyev and Hüseyin Pehlivan. Tournament scoring problem. Applied Mathematics and Computation, 199(1):211– 222, May 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307009885.

Nian:2009:NSA

[NPGW09] Xiaohong Nian, Huan Pan, Weihua Gui, and Haibo Wang. New stability analysis for linear neutral system via state matrix decomposition. Applied Mathematics and Computation, 215(5):1830–1837, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006845.

Natalini:2009:HPS

[NPR09] Pierpaolo Natalini, Roberto Patrizi, and Paolo E. Ricci. Heat problems for a starlike shaped plate. Applied Mathematics and Computation, 215(2):495–502, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004901.

Najafi:2007:FIV

[NR07a] H. Saberi Najafi and A. Refahi. FOM-inverse vector iteration method for computing a few smallest (largest) eigenvalues of pair ( A B ). Applied Mathematics and Computation, 188(1):641–647, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014020.

Najafi:2007:NRM

[NR07b] H. Saberi Najafi and A. Refahi. A new restarting method in the Lanczos algorithm for generalized eigenvalue problem. Applied Mathematics and Computation, 184(2):421–428, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600717X.

Naimzada:2008:CDM

[NR08a] Ahmad K. Naimzada and Giorgio Ricchiuti. Complex dynamics in a monopoly with a rule of thumb. Applied Mathematics and Computation, 203(2):921– 925, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300308002270.

Naimzada:2008:HFI

[NR08b] Ahmad K. Naimzada and Giorgio Ricchiuti. Heterogeneous fundamentalists and imitative processes. Applied Mathematics and Computation, 199(1):171–180, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009848.

Navickas:2009:HFO

[NR09] Zenonas Navickas and Minvydas Ragulskis. How far one can go with the Expfunction method? Applied Mathematics and Computation, 211(2):522–530, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001118.

Najafi:2007:WFI

[NRA07] H. Saberi Najafi, A. Refahi, and M. Akbari. Weighted FOM-inverse vector iteration method for computing a few smallest (largest) eigenvalues of pair ( A B ). Applied Mathematics and Computation, 192(1):239–246, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700327X.

Nielsen:2007:OMA

[NRLT07] Bjørn Fredrik Nielsen, Tomas Syrstad Ruud, Glenn Terje Lines, and Aslak Tveito. Optimal monodomain approximations of the bidomain equations. Applied Mathematics and Computation, 184(2):276–290, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600703X.

Najafi:2005:SAM

[NS05] H. Saberi Najafi and M. Shams. Shifting algorithms with Maple and implicit shift in the QR algorithm. Applied Mathematics and Computation, 161(3):947–962, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400044X.

Najafi:2006:CAC

[NS06a] H. Saberi Najafi and M. Shams Solary. Computational algorithms for computing the inverse of a square matrix, quasiinverse of a non-square matrix and block matrices. Applied Mathematics and Computation, 183(1):539–550, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005893.

Nilsrakoo:2006:NTS

[NS06b] Weerayuth Nilsrakoo and Satit Saejung. A new three-step fixed point iteration scheme for asymptotically nonexpansive mappings. Applied Mathematics and Computation, 181(2):1026–1034, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600186X.

Nammanee:2007:MNI

[NS07a] Kamonrat Nammanee and Suthep Suantai. The modified Noor iterations with errors for non-Lipschitzian mappings in Banach spaces. Applied Mathematics and Computation, 187(2):669– 679, April 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306011222.

Nilsrakoo:2007:RCT

[NS07b] Weerayuth Nilsrakoo and Satit Saejung. A reconsideration on convergence of three-step iterations for asymptotically nonexpansive mappings. Applied Mathematics and Computation, 190(2):1472–1478, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001804.

Najafi:2008:NAC

[NS08] H. Saberi Najafi and M. Shams Solary. A new algorithm for computing the inertia of eigenproblems (Ax = λx) and (Ax = λBx). Applied Mathematics and Computation, 201(1–2):138–148, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011782.

Nilsrakoo:2009:CTC

[NS09] Weerayuth Nilsrakoo and Satit Saejung. Convergence theorems for a countable family of Lipschitzian mappings. Applied Mathematics and Computation, 214(2):595–606, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309003592.

Naresh:2006:MEI

[NSS06a] Ram Naresh, Shyam Sundar, and J. B. Shukla. Modeling the effect of an intermediate toxic product formed by uptake of a toxicant on plant biomass. Applied Mathematics and Computation, 182(1):151–160, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003213.

Naresh:2006:MRP

[NSS06b] Ram Naresh, Shyam Sundar, and J. B. Shukla. Modeling the removal of primary and secondary pollutants from the atmosphere of a city by rain. Applied Mathematics and Computation, 179(1):282–295, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010362.

Naresh:2006:MSA

[NTO06] Ram Naresh, Agraj Tripathi, and Sandip Omar. Modelling the spread of AIDS epidemic with vertical transmission. Applied Mathematics and Computation, 178(2):262–272, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009720.

Noguera:2008:AEG

[NTZ08] José-Vicente Noguera, Leandro Tortosa, and Antonio Zamora. Analysis and efficiency of the GNG3D algorithm for mesh simplification. Applied Mathematics and Computation, 197(1):29–40, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007503.

Naseri:2007:EVK

[NV07] Rasool Naseri and Azizollah Valinejad. An extended variant of Karmarkar’s interior point algorithm. Applied Mathematics and Computation, 184(2):737–742, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007508.

Nwogugu:2006:DMR

[Nwo06a] Michael Nwogugu. Decisionmaking, risk and corporate governance: New dynamic models/algorithms and optimization for bankruptcy decisions. Applied Mathematics and Computation, 179(1):386–401, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010465.

Nwogugu:2006:ESO

[Nwo06b] Michael Nwogugu. Employeestock-options, production/ service functions and game theory. Applied Mathematics and Computation, 181(1):552–562, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001275.

Nwogugu:2006:FCC

[Nwo06c] Michael Nwogugu. A further critique of cumulative prospect theory and related approaches. Applied Mathematics and Computation, 179(2):451–465, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010490.

Nwogugu:2006:FCG

[Nwo06d] Michael Nwogugu. Further critique of GARCH/ ARMA/VAR/EVT Stochastic-Volatility models and related approaches. Applied Mathematics and Computation, 182(2):1735–1748, November 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306001810.

Nwogugu:2006:RMW

[Nwo06e] Michael Nwogugu. Regret minimization, willingnessto-accept-losses and framing. Applied Mathematics and Computation, 179(2):440–450, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010489.

Nwogugu:2006:SSU

[Nwo06f] Michael Nwogugu. Site selection in the US retailing industry. Applied Mathematics and Computation, 182(2):1725–1734, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001251.

Nwogugu:2006:VRM

[Nwo06g] Michael Nwogugu. Volatility, risk modeling and utility. Applied Mathematics and Computation, 182(2):1749–1754, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001822.

Nwogugu:2007:DMR

[Nwo07a] Michael Nwogugu. Decisionmaking, risk and corporate governance: a critique of methodological issues in bankruptcy/recovery prediction models. Applied Mathematics and Computation, 185(1):178–196, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008332.

Nwogugu:2007:IDR

[Nwo07b]  Michael Nwogugu.  Is-

sues in disintermediation in the real estate brokerage sector. Applied Mathematics and Computation, 186(2):1054–1064, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010319.

Nwogugu:2007:SGT

[Nwo07c] Michael Nwogugu. Some game theory and financial contracting issues in corporate transactions. Applied Mathematics and Computation, 186(2):1018–1030, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010289.

Nwogugu:2007:SID

[Nwo07d] Michael Nwogugu. Some issues in disintermediation and securitization. Applied Mathematics and Computation, 186(2):1031–1039, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010290.

Xing:2006:IGA

[nXwCpC06] Li ning Xing, Ying wu Chen, and Huai ping Cai. An intelligent genetic algorithm designed for global optimization of multiminima functions. Applied Mathematics and Computation, 178(2):355–371, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009811.

Noor:2007:TSI

[NY07] Muhammad Aslam Noor and Yonghong Yao. Threestep iterations for variational inequalities and nonexpansive mappings. Applied Mathematics and Computation, 190(2):1312– 1321, July 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307001622.

Najafi:2008:NCG

[NZ08] H. Saberi Najafi and H. Zareamoghaddam. A new computational GM-RES method. Applied Mathematics and Computation, 199(2):527–534, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010284.

Niu:2007:MMS

[NZHW07] Ben Niu, Yunlong Zhu, Xiaoxian He, and Henry Wu. MCPSO: a multi-swarm cooperative particle swarm optimizer. Applied Mathematics and Computation, 185(2):1050–1062, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008290.

Ozis:2005:SAA

[ÖA05] Turgut Öziş and Yaşar Aslan. The semi-approximate approach for solving Burgers’ equation with high Reynolds number. Applied Mathematics and Computation, 163(1):131–145, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002152.

Omrani:2008:NCF

[OAAK08] Khaled Omrani, Faycal Abidi, Talha Achouri, and Noomen Khiari. A new conservative finite difference scheme for the Rosenau equation. Applied Mathematics and Computation, 201(1–2):35–43, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030701154X.

Omar:2009:CMM

[OAHR09] M. A. Omar, A. Aboul-Hassan, and S. I. Rabia. The composite Milstein methods for the numerical solution of Stratonovich stochastic differential equations. Applied Mathematics and Computation, 215(2):727–745, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309005530.

Osman:2005:EGAa

[OASM05a] M. S. Osman, M. A. Abo-Sinna, and A. A. Mousa. An effective genetic algorithm approach to multiobjective resource allocation problems (MORAPs). Applied Mathematics and Computation, 163(2):755–768, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002589.

Osman:2005:EGAb

[OASM05b] M. S. Osman, M. A. Abo-Sinna, and A. A. Mousa.

An effective genetic algorithm approach to multiobjective routing problems (MORPs). Applied Mathematics and Computation, 163(2):769–781, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002590.

Osman:2005:CGA

[OASM05c] M. S. Osman, Mahmoud A. Abo-Sinna, and A. A. Mousa. A combined genetic algorithmfuzzy logic controller (GA– FLC) in nonlinear programming. Applied Mathematics and Computation, 170(2):821–840, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000263.

Osman:2006:ICI

[OASM06] M. S. Osman, Mahmoud A.

Abo-Sinna, and A. A. Mousa. IT-CEMOP: an iterative co-evolutionary algorithm for multiobjective optimization problem with nonlinear constraints. Applied Mathematics and Computation, 183(1):373–389, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005698.

Oliveira:2009:PAE

[OB09] Jáuber C. Oliveira and Fermín S. Viloche Bazán. Poisson approach for evaluating numerical methods for the two-dimensional wave equation constrained to absorbing boundary conditions. Applied Mathematics and Computation, 209(2):273–284, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308009478.

Ozturk:2006:DEI

[OBO06] I. Ozturk, F. Bozkurt, and S. Ozen. On the difference equation y_{n+1} = (α +β exp(y_{n}))/(γ +y_{n−1}). Applied Mathematics and Computation, 181(2):1387– 1393, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306002578.

Ocana:2007:APC

[Oca07] Francisco A. Ocaña. An approximation problem in computing electoral volatility. Applied Mathematics and Computation, 192(2):299–310, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700330X.

Occorsio:2009:ELI

[Occ09] D. Occorsio. Extended Lagrange interpolation in weighted uniform norm. Applied Mathematics and Computation, 211(1):10– 22, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000447.

Odibat:2006:AFI

[Odi06a] Zaid Odibat. Approximations of fractional integrals and Caputo fractional derivatives. Applied Mathematics and Computation, 178(2):527–533, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010003.

Odibat:2006:RDM

[Odi06b] Zaid M. Odibat. Rectangular decomposition method for fractional diffusionwave equations. Applied Mathematics and Computation, 179(1):92–97, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010106.

Odibat:2006:RMR

[Odi06c] Zaid M. Odibat. A reliable modification of the rectangular decomposition method. Applied Mathematics and Computation, 183(2):1226–1234, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006680.

Odibat:2007:NMH

[Odi07] Zaid M. Odibat. A new modification of the homotopy perturbation method for linear and nonlinear operators. Applied Mathematics and Computation, 189(1):746–753, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016523.

Odibat:2008:CSC

[Odi08] Zaid M. Odibat. Compact structures in a class of nonlinearly dispersive equations with time-fractional derivatives. Applied Mathematics and Computation, 205(1):273–280, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004906.

Odibat:2009:CEB

[Odi09] Zaid M. Odibat. Computing eigenelements of boundary value problems with fractional derivatives. Applied Mathematics and Computation, 215(8):3017–3028, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008741.

Ozis:2005:NSB

[ÖEK05] T. Özis, A. Esen, and S. Kutluay. Numerical solution of Burgers’ equation by quadratic B-spline finite elements. Applied Mathematics and Computation, 165(1):237–249, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003789.

Oh:2006:EAS

[OF06] Jiyeon Oh and Donald A.

French. Error analysis of a specialized numerical method for mathematical models from neuroscience. Applied Mathematics and Computation, 172(1):491–507, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305002183.

Ou:2007:SMC

[OHJ07] Hsia-Hung Ou, Min-Shiang Hwang, and Jinn-Ke Jan. A simple mobile communication billing system among charged parties. Applied Mathematics and Computation, 192(2):487–495, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307003529.

Ohmi:2008:SBP

[Ohm08]  Kazuo Ohmi.  SOM-

based particle matching algorithm for 3D particle tracking velocimetry. Applied Mathematics and Computation, 205(2):890–898, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003688.

Ong:2005:MIA

[OHT05a] Chorng-Shyong Ong, Jih-Jeng Huang, and Gwo-Hshiung Tzeng. Model identification of ARIMA family using genetic algorithms. Applied Mathematics and Computation, 164(3):885–912, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003546.

Ong:2005:NHM

[OHT05b] Chorng-Shyong Ong, Jih-Jeng Huang, and Gwo-Hshiung Tzeng. A novel hybrid model for portfolio selection. Applied Mathematics and Computation, 169(2):1195–1210, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008458.

Ohtsubo:2008:SSP

[Oht08] Yoshio Ohtsubo. Stochastic shortest path problems with associative accumulative criteria. Applied Mathematics and Computation, 198(1):198–208, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700879X.

Onyango:2009:IRB

[OIL09] T. T. M. Onyango, D. B. Ingham, and D. Lesnic. Inverse reconstruction of boundary condition coefficients in one-dimensional transient heat conduction. Applied Mathematics and Computation, 207(2):569–575, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800845X.

Othman:2007:AIB

[OISM07] Khairil Iskandar Othman, Zarina Bibi Ibrahim, Mohamed Suleiman, and Zanariah Majid. Automatic intervalwise block partitioning using Adams type method and backward differentiation formula for solving ODEs. Applied Mathematics and Computation, 188(2):1642–1646, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015165.

Ozer:2005:ANM

[ÖK05] S. Özer and S. Kutluay. An analytical– numerical method for solving the Korteweg–de Vries equation. Applied Mathematics and Computation, 164(3):789–797, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003455.

Oh:2005:HLM

[OKEL05] Sang-Keon Oh, Min-Soeng Kim, Tae-Dok Eom, and Ju-Jang Lee. Heterogeneous local model networks for time series prediction. Applied Mathematics and Computation, 168(1):164–177, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400596X.

Olagunju:2006:LSS

[Ola06] David O. Olagunju. Local similarity solutions for boundary layer flow of a FENE-P fluid. Applied Mathematics and Computation, 173(1):593–602, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003784.

Odibat:2006:ASB

[OM06] Zaid M. Odibat and Shaher Momani. Approximate solutions for boundary value problems of time-fractional wave equation. Applied Mathematics and Computation, 181(1):767–774, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001573.

Omran:2008:GBH

[OM08] Mahamed G. H. Omran and Mehrdad Mahdavi. Global-best harmony search. Applied Mathematics and Computation, 198(2):643–656, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009320.

Omarah:2006:TSQ

[Oma06] Al-Seedy Ragab Omarah. A transient solution of queues with variable channel considering balking concept when S = 2 and σ = 1. Applied Mathematics and Computation, 174(1):337–344, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305004601.

Odibat:2008:GDT

[OME08] Zaid Odibat, Shaher Momani, and Vedat Suat Erturk. Generalized differential transform method: Application to differential equations of fractional order. Applied Mathematics and Computation, 197(2):467–477, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007916.

Ouyang:2009:CPS

[OML09] Yao Ouyang, Radko Mesiar, and Jun Li. On the comonotonic- * -property for Sugeno integral. Applied Mathematics and Computation, 211(2):450– 458, May 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309000939.

Omrani:2006:CFD

[Omr06] Khaled Omrani. The convergence of fully discrete Galerkin approximations for the Benjamin–Bona– Mahony (BBM) equation. Applied Mathematics and Computation, 180(2):614–621, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306000506.

Omrani:2007:NME

[Omr07a] Khaled Omrani. Numerical methods and error analysis for the nonlinear Sivashinsky equation. Applied Mathematics and Computation, 189(1):949– 962, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016742.

Ostadi:2007:NLP

[OMR07b] B. Ostadi, D. Moazzami, and K. Rezaie. A nonlinear programming model for optimization of the electrical energy consumption in typical factory. Applied Mathematics and Computation, 187(2):944–950, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012288.

Olguin:2007:BSS

[OMST07] M. C. Olguín, M. A. Medina, M. C. Sanziel, and D. A. Tarzia. Behavior of the solution of a Stefan problem by changing thermal coefficients of the substance. Applied Mathematics and Computation, 190(1):765–780, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001117.

Olamaei:2008:APS

[ONG08] J. Olamaei, T. Niknam, and G. Gharehpetian. Application of particle swarm optimization for distribution feeder reconfiguration considering distributed generators. Applied Mathematics and Computation, 201(1–2):575–586, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000027.

Onyejekwe:2005:EGE

[Ony05a] Okey Oseloka Onyejekwe.

An efficient Green element algorithm for radial flow. Applied Mathematics and Computation, 165(3):635–645, June 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004084.

Onyejekwe:2005:SGE

[Ony05b] Okey Oseloka Onyejekwe.

A simplified Green element method solution for a the biharmonic equation and its application to Stokes flows. Applied Mathematics and Computation, 169(2):1405–1418, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008616.

Ozis:2006:SST

[ÖÖ06] T. Öziş and S. Özer. A simple similarity-transformation-iterative scheme applied to Korteweg–de Vries equation. Applied Mathematics and Computation, 173(1):19–32, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305002997.

Ozen:2007:RSI

[OOB07] S. Ozen, I. Ozturk, and F. Bozkurt. On the recursive sequence y_{n+1} = (α + y_{n−1})/(β + y_{n}) + y_{n−1}/y_{n}. Applied Mathematics and Computation, 188(1):180– 188, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030601335X.

Obradovic:2009:SCC

[OP09] M. Obradović and S. Ponnusamy. Starlike and close-to-convex functions defined by differential inequalities. Applied Mathematics and Computation, 215(7):2642–2648, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008017.

ORegan:2008:SIE

[O’R08] Donal O’Regan. Singular integral equations arising in draining and coating flows. Applied Mathematics and Computation, 205(1):438–441, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006048.

ORegan:2009:FPI

[O’R09] Donal O’Regan. The fixed point index for compact absorbing contractive admissible maps. Applied Mathematics and Computation, 215(5):1975–1983, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007644.

Ord:2006:MFH

[Ord06a] Toby Ord. The many forms of hypercomputation. Applied Mathematics and Computation, 178(1):143– 153, July 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008404.

Ordokhani:2006:SNV

[Ord06b]  Y. Ordokhani.  Solu-

tion of nonlinear Volterra– Fredholm–Hammerstein integral equations via rationalized Haar functions. Applied Mathematics and Computation, 180(2):436– 443, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306000051.

Overby:2009:RMH

[ORG^{+}09] Darryl R. Overby, Jeffrey W. Ruberti, Haiyan Gong, Thomas F. Freddo, and Mark Johnson. Response to “Merchant and Heys, Effects of Variable Permeability on Aqueous Humor Outflow, Applied Mathematics and Computation 196 (2008) 371– 380”. Applied Mathematics and Computation, 211 (1):251–252, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300308009612.

Ozer:2006:FSG

[ÖŞ06] Saadet Özer and Erdoğan S.

Şuhubi. Flow of a second grade fluid through a cylindrical permeable tube. Applied Mathematics and Computation, 179(2):672–687, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010726.

Odibat:2007:GTF

[OS07] Zaid M. Odibat and Nabil T. Shawagfeh. Generalized Taylor’s formula. Applied Mathematics and Computation, 186(1):286– 293, March 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306009398.

ORegan:2008:NCT

[OS08] Donal O’Regan and Reza Saadati. Nonlinear contraction theorems in probabilistic spaces. Applied Mathematics and Computation, 195(1):86–93, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005206.

Osilike:2009:EAC

[OS09] M. O. Osilike and Y. Shehu. Explicit averaging cyclic algorithm for common fixed points of asymptotically strictly pseudocontractive maps. Applied Mathematics and Computation, 213(2):548–553, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002689.

Ozdemir:2009:ITL

[ÖSÖG09] Halim Özdemir, Murat Sarduvan, Ahmet Yaşar Özban, and Nesrin Güler. On idempotency and tripotency of linear combinations of two commuting tripotent matrices. Applied Mathematics and Computation, 207(1):197–201, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308007790.

Ossipov:2005:HOS

[Oss05a] P. Ossipov. Heuristic optimization of sequence of customer orders. Applied Mathematics and Computation, 162(3):1303–1313, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001870.

Ossipov:2005:SHA

[Oss05b] P. Ossipov. Simple heuristic algorithm for profile reduction of arbitrary sparse matrix. Applied Mathematics and Computation, 168(2):848–857, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006435.

Opps:2009:RFA

[OSS09] Sheldon B. Opps, Nasser Saad, and H. M. Srivastava. Recursion formulas for Appell’s hypergeometric function F _{2} with some applications to radiation field problems. Applied Mathematics and Computation, 207(2):545–558, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008424.

Ostrovska:2008:BPC

[Ost08]  Sofiya Ostrovska.  q-

Bernstein polynomials of the Cauchy kernel. Applied Mathematics and Computation, 198(1):261–270, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008867.

Owa:2007:STF

[OSY07] Shigeyoshi Owa, Tadayuki Sekine, and Rikuo Yamakawa. On Sakaguchi type functions. Applied Mathematics and Computation, 187(1):356–361, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011842.

Omar:2006:MSE

[OT06] Mohamed K. Omar and Siew C. Teo. Minimizing the sum of earliness/ tardiness in identical parallel machines schedule with incompatible job families: an improved MIP approach. Applied Mathematics and Computation, 181(2):1008–1017, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001846.

Ocal:2005:RGF

[ÖTA05] Ahmet Ali Öcal, Naim Tuglu, and Ercan Altinişik. On the representation of k-generalized Fibonacci and Lucas numbers. Applied Mathematics and Computation, 170(1):584–596, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009701.

Ouaf:2005:EST

[Oua05] Mahmoud E. M. Ouaf.

Exact solution of thermal radiation on MHD flow over a stretching porous sheet. Applied Mathematics and Computation, 170(2):1117–1125, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000524.

Oujja:2006:NMC

[Ouj06] Rachid Oujja. A new method for cavitations approximation in some general lubrication devices. Applied Mathematics and Computation, 181(2):1645–1656, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002967.

Oh:2005:OFE

[OY05] Seungtaik Oh and Jin-Whan Yim. Optimal finite element mesh for elliptic equation of divergence form. Applied Mathematics and Computation, 162(2):969–989, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001432.

ORegan:2007:LWN

[OZ07] Donal O’Regan and Mirosława Zima. Leggett–Williams norm-type fixed point theorems for multivalued mappings. Applied Mathematics and Computation, 187(2):1238–1249, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012604.

Ozalp:2005:CMH

[Öza05] Nuri Özalp. A computational method for hypersurfaces in terms of gamma function. Applied Mathematics and Computation, 161(3):721–732, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304000244.

Ozban:2006:NMA

[Özb06] Ahmet Yaşar Özban. New methods for approximating square roots. Applied Mathematics and Computation, 175(1):532–540, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006272.

Ozban:2007:SRD

[Özb07a] Ahmet Yaşar Özban. On the system of rational difference equations x_{n} = a/y_{n−3}, y_{n} = by_{n−3}/(x_{n−q}y_{n−q}). Applied Mathematics and Computation, 188(1):833– 837, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014226.

Ozbilge:2007:ACF

[Ozb07b] Ebru Ozbilge. An analysis of conservative finite difference schemes for differential equations with discontinuous coefficients. Applied Mathematics and Computation, 191(1):183–192, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002445.

Ozdemir:2005:VIA

[Ozd05] Müjgan Sagir Ozdemir. Validity and inconsistency in the analytic hierarchy process. Applied Mathematics and Computation, 161(3):707–720, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304000232.

Ozkan:2007:SSR

[Özk07] Öznur Özkan. Some subordination results on the classes starlike and convex functions of complex order. Applied Mathematics and Computation, 187(1):362– 368, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011854.

Oztop:2005:NSF

[Özt05] Hakan F. Öztop. Numerical study of flow and heat transfer in curvilinear ducts: applications of elliptic grid generation. Applied Mathematics and Computation, 168(2):1449–1460, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007428.

Piqueira:2009:MEM

[PA09] José Roberto C. Piqueira and Vanessa O. Araujo. A modified epidemiological model for computer viruses. Applied Mathematics and Computation, 213(2):355–360, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002379.

Pakdaman:2011:ENB

[Pak11] Morteza Pakdaman. Erratum to “A new branch and bound method with pretreatment for the binary quadratic programming” [Appl. Math. Comput. 192 (2007) 252– 259]. Applied Mathematics and Computation, 217 (13):6441, March 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [MZL07].

Palubeckis:2007:ITS

[Pal07] Gintaras Palubeckis. Iterated tabu search for the maximum diversity problem. Applied Mathematics and Computation, 189 (1):371–383, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306016122.

Palokangas:2008:EPE

[Pal08] Tapio Palokangas. Emission policy in an economic union with Poisson technological change. Applied Mathematics and Computation, 204(2):589–594, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003068.

Palubeckis:2009:NBP

[Pal09] Gintaras Palubeckis. A new bounding procedure and an improved exact algorithm for the Max-2-SAT problem. Applied Mathematics and Computation, 215(3):1106–1117, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006183.

Pamuk:2005:ALN

[Pam05a] Serdal Pamuk. An application for linear and nonlinear heat equations by Adomian’s decomposition method. Applied Mathematics and Computation, 163(1):89–96, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002115.

Pamuk:2005:DMC

[Pam05b] Serdal Pamuk. The decomposition method for continuous population models for single and interacting species. Applied Mathematics and Computation, 163(1):79–88, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002103.

Pamuk:2006:SSP

[Pam06] Nevin Pamuk. Series solution for porous medium equation with a source term by Adomian’s decomposition method. Applied Mathematics and Computation, 178(2):480–485, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500994X.

Pantokratoras:2006:CLB

[Pan06] A. Pantokratoras. Comment on “Laminar boundary layer flow over a horizontal permeable flat plate”. Applied Mathematics and Computation, 182(1):1–2, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002827.

See [AK05i].

Pan:2008:PDB

[Pan08a]  Ping-Qi Pan.  A pri-

mal deficient-basis simplex algorithm for linear programming. Applied Mathematics and Computation, 196(2):898–912, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007436.

Pantokratoras:2008:CFH

[Pan08b] Asterios Pantokratoras. Comment on “Flow and heat transfer in a micropolar fluid past a stretching surface embedded in a non-Darcian porous medium with a uniform free stream” by Emad M. Abo-Eldahab and Mohamed A. El Aziz [Appl. Math. Comput. 162 (2005) 881–899]. Applied Mathematics and Computation, 202(2):895–896, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/S0096300308000581.

See [AEA05a].

Pantokratoras:2008:CTD

[Pan08c] Asterios Pantokratoras.

Comment on “Thermaldiffusion and diffusionthermo effects on mixed free-forced convection and mass transfer boundary layer flow for non-Newtonian fluid with temperature dependent viscosity”, by N. T. Eldabe, A. G. El-Saka and Ashraf Fouad [Applied Mathematics and Computation 152 (2004) 867– 883]. Applied Mathematics and Computation, 196(2): 959–961, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307006534. See

[EESF04].

Pantokratoras:2012:CCF

[Pan12] Asterios Pantokratoras.

Comment on “Chebyshev finite difference method for MHD flow of a micropolar fluid past a stretching sheet with heat transfer”, by N.T. Eldabe, E.F. Elshehawey, Elsayed M.E. Elbarbary and Nasser S. Elgazery [Applied Mathematics and Computation, 160, 2005, 437–450]. Applied Mathematics and Computation, 218(9):5827–5828, January 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300311013555.

See [EEEE05].

Pao:2007:MIM

[Pao07] C. V. Pao. Monotone iterative methods for numerical solutions of nonlinear integro-elliptic boundary problems. Applied Mathematics and Computation, 186(2):1624–1642, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030601099X.

Park:2005:LOA

[Par05a] Ju H. Park. LMI optimization approach to asymptotic stability of certain neutral delay differential equation with time-varying coefficients. Applied Mathematics and Computation, 160(2):355–361, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303011494.

Park:2005:NRS

[Par05b] Ju H. Park. Novel robust stability criterion for a class of neutral systems with mixed delays and nonlinear perturbations. Applied Mathematics and Computation, 161(2):413–421, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012955.

Park:2005:DDO

[Par05c]  Ju H. Park.  On de-

sign of dynamic output feedback controller for GCS of large-scale systems with delays in interconnections: LMI optimization approach. Applied Mathematics and Computation, 161(2):423–432, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012967.

Park:2006:FRA

[Par06a] Ju H. Park. Further result on asymptotic stability criterion of cellular neural networks with time-varying discrete and distributed delays. Applied Mathematics and Computation, 182(2):1661–1666, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005145.

Park:2006:GES

[Par06b] Ju H. Park. Global exponential stability of cellular neural networks with variable delays. Applied Mathematics and Computation, 183(2):1214–1219, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006667.

Park:2006:NSA

[Par06c] Ju H. Park. A new stability analysis of delayed cellular neural networks. Applied Mathematics and Computation, 181(1):200–205, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306000907.

Park:2007:FNG

[Par07] Ju H. Park. Further note on global exponential stability of uncertain cellular neural networks with variable delays. Applied Mathematics and Computation, 188(1):850–854, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601424X.

Park:2008:PMV

[Par08]  Dong-Chul Park.  Pre-

diction of MPEG video traffic over ATM networks using dynamic bilinear recurrent neural network. Applied Mathematics and Computation, 205(2):648–657, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003433.

Parkes:2009:NTW

[Par09a] E. J. Parkes. A note on travelling-wave solutions to Lax’s seventh-order KdV equation. Applied Mathematics and Computation, 215(2):864–865, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005621.

Parlakci:2009:RDD

[Par09b] M. N. Alpaslan Parlakçı.

Robust delay-dependent guaranteed cost controller design for uncertain neutral systems. Applied Mathematics and Computation, 215(8):2936–2949, December 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300309008650.

Patidar:2005:DIP

[Pat05a] Kailash C. Patidar. Dispersion induced by the pollution for the wave equation. Applied Mathematics and Computation, 160(2):329–341, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030301138X.

Patidar:2005:HOF

[Pat05b] Kailash C. Patidar. High order fitted operator numerical method for selfadjoint singular perturbation problems. Applied Mathematics and Computation, 171(1):547–566, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305001141.

Patidar:2007:HOP

[Pat07] Kailash C. Patidar. High order parameter uniform numerical method for singular perturbation problems. Applied Mathematics and Computation, 188 (1):720–733, May 1, 2007.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030601410X.

Padma:2007:AHP

[PB07a] T. Padma and P. Balasubramanie. Analytic hierarchy process to assess occupational risk for shoulder and neck pain. Applied Mathematics and Computation, 193(2):321–324, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004006.

Pakdemirli:2007:GRF

[PB07b] Mehmet Pakdemirli and Hakan Boyacı. Generation of root finding algorithms via perturbation theory and some formulas. Applied Mathematics and Computation, 184(2):783–788, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007570.

Pang:2009:ULS

[PB09] Yuanyuan Pang and Zhanbing Bai. Upper and lower solution method for a fourth-order four-point boundary value problem on time scales. Applied Mathematics and Computation, 215(6):2243–2247, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007371.

Pal:2007:OPI

[PBG07] P. Pal, A. K. Bhunia, and S. K. Goyal. On optimal partially integrated production and marketing policy with variable demand under flexibility and reliability considerations via Genetic Algorithm. Applied Mathematics and Computation, 188(1):525– 537, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013920.

Papakostas:2009:PCU

[PBKM09] G. A. Papakostas, Y. S. Boutalis, D. A. Karras, and B. G. Mertzios. Pattern classification by using improved wavelet Compressed Zernike Moments. Applied Mathematics and Computation, 212(1):162– 176, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001313.

Papakostas:2008:NSF

[PBPF08] G. A. Papakostas, Y. S.

Boutalis, C. N. Papaodysseus, and D. K. Fragoulis. Numerical stability of fast computation algorithms of Zernike moments. Applied Mathematics and Computation, 195(1):326–345, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005565.

Park:2005:NDS

[PC05] Je Hong Park and Seongtaek Chee. A note on digital signature scheme using a self-pairing map. Applied Mathematics and Computation, 169(1):472–475, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007866.

Pan:2007:TUO

[PC07a] Shaohua Pan and Jein-Shan Chen. Two unconstrained optimization approaches for the Euclidean κ-centrum location problem. Applied Mathematics and Computation, 189(2):1368–1383, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017231.

Patel:2007:CSC

[PC07b] J. Patel and N. E. Cho. On certain sufficient conditions for close-to-convexity. Applied Mathematics and Computation, 187(1):369– 378, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011866.

Pei:2008:QMS

[PC08] Minghe Pei and Sung Kag Chang. A quasilinearization method for secondorder four-point boundary value problems. Applied Mathematics and Computation, 202(1):54–66, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308000568.

Pavaloiu:2009:SHM

[PC09] I. Păvăloiu and E. Cătinaş. On a Steffensen–Hermite method of order three. Applied Mathematics and Computation, 215(7):2663– 2672, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309008042.

Posadas-Castillo:2008:SCN

[PCCHLG08] C. Posadas-Castillo, C. Cruz-Hernández, and R. M. López-Gutiérrez. Synchronization of chaotic neural networks with delay in irregular networks. Applied Mathematics and Computation, 205(1):487–496, November 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030800581X.

Podisuk:2007:SSF

[PCS07] Maitree Podisuk, Ungsana Chundang, and Wannaporn Sanprasert. Single step formulas and multistep formulas of the integration method for solving the initial value problem of ordinary differential equation. Applied Mathematics and Computation, 190(2):1438–1444, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001762.

Phongthanapanich:2009:CFV

[PD09]  Sutthisak  Phongthana-

panich and Pramote Dechaumphai.

Combined finite volume element method for singularly perturbed reaction– diffusion problems. Applied Mathematics and Computation, 209(2):177–185, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008126.

Pushpam:2006:NSH

[PDA06] A. Emimal Kanaga Pushpam, D. Paul Dhayabaran, and E. C. Henry Amirtharaj. Numerical solution of higher order systems of IVPs using generalized STWS technique. Applied Mathematics and Computation, 180(1):200–205, September 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305011148.

Pelaez:2007:AOO

[PDGR07] J. I. Peláez, J. M. Doña, and J. A. Gómez-Ruiz. Analysis of OWA operators in decision making for modelling the majority concept. Applied Mathematics and Computation, 186(2):1263–1275, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010526.

Pang:2006:GFP

[PDW06] Changci Pang, Wei Dong, and Zhongli Wei. Green’s function and positive solutions of nth order mpoint boundary value problem. Applied Mathematics and Computation, 182(2):1231–1239, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004383.

Pamuk:2007:MLN

[PE07a] Serdal Pamuk and Arzu Erdem. The method of lines for the numerical solution of a mathematical model for capillary formation: The role of endothelial cells in the capillary. Applied Mathematics and Computation, 186(1):831–835, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010058.

Potgieter:2007:GAS

[PE07b] G. Potgieter and A. P. Engelbrecht. Genetic algorithms for the structural optimisation of learned polynomial expressions.

Applied Mathematics and Computation, 186(2):1441– 1466, March 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306010708.

Petela:2009:ACK

[Pe¸ 09] Joanna Peţela. Average conditions for Kolmogorov systems. Applied Mathematics and Computation, 215(2):481–494, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309004895.

Peng:2005:IEP

[Pen05]  Zhen-Yun Peng.  The

inverse eigenvalue problem for Hermitian antireflexive matrices and its approximation. Applied Mathematics and Computation, 162(3):1377–1389, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030400195X.

Peng:2006:GAS

[Pen06]  Yuehui Peng.  Global

asymptotic stability for nonlinear difference equations. Applied Mathematics and Computation, 182(1):67–72, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003134.

Peng:2008:ETW

[Pen08] Yan-Ze Peng. Exact travelling wave solutions for the Zakharov–Kuznetsov equation. Applied Mathematics and Computation, 199 (2):397–405, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307008594.

Peng:2009:ATA

[Pen09] Xingxuan Peng. Adaptive thinning algorithm of nonsingular triangulation. Applied Mathematics and Computation, 215(6):2162–2173, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007255.

Petkovic:2007:CBS

[Pet07] Miodrag S. Petković. Comments on the Basto– Semiao–Calheiros root finding method. Applied Mathematics and Computation, 184(2):143–148, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006850.

See [BSC06].

Petkovic:2008:CES

[Pet08a]  Ivan Petković.  Com-

putational efficiency of some combined methods for polynomial equations. Applied Mathematics and Computation, 204(2):949–956, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005857.

Petkovic:2008:HER

[Pet08b] Miodrag S. Petković. A highly efficient root-solver of very fast convergence. Applied Mathematics and Computation, 205(1):298– 302, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300308005821.

Pourakbar:2007:JEL

[PFA07] Morteza Pourakbar, Reza Zanjirani Farahani, and Nasrin Asgari. A joint economic lot-size model for an integrated supply network using genetic algorithm. Applied Mathematics and Computation, 189(1):583– 596, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016328.

Pang:2008:MPS

[PFG08] Huihui Pang, Hanying Feng, and Weigao Ge. Multiple positive solutions of quasi-linear boundary value problems for finite difference equations. Applied Mathematics and Computation, 197(1):451–456, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007114.

Peng:2009:FVM

[PFL09] Yehui Peng, Heying Feng, and Qiyong Li. A filtervariable-metric method for nonsmooth convex constrained optimization. Applied Mathematics and Computation, 208(1):119– 128, February 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308008564.

Parida:2006:IRF

[PG06] P. K. Parida and D. K. Gupta. An improved regula-falsi method for enclosing simple zeros of nonlinear equations. Applied Mathematics and Computation, 177(2):769–776, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009586.

Polat:2007:CEE

[PG07a] Kemal Polat and Salih Güneş. Classification of epileptiform EEG using a hybrid system based on decision tree classifier and fast Fourier transform. Applied Mathematics and Computation, 187(2):1017–1026, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012380.

Polat:2007:DEA

[PG07b] Kemal Polat and Salih Güneş. Detection of ECG Arrhythmia using a differential expert system approach based on principal component analysis and least square support vector machine. Applied Mathematics and Computation, 186(1):898–906, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010125.

Polat:2007:NAE

[PG07c] Kemal Polat and Salih Güneş. A novel approach to estimation of E. coli promoter gene sequences: Combining feature selection and least square support vector machine (FS LSSVM). Applied Mathematics and Computation, 190(2):1574–1582, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001944.

Polat:2007:PHD

[PG07d] Kemal Polat and Salih Güneş. Prediction of hepatitis disease based on principal component analysis and artificial immune recognition system. Applied Mathematics and Computation, 189(2):1282–1291, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017140.

Parhi:2008:SOM

[PG08a] S. K. Parhi and D. K. Gupta. A sixth order method for nonlinear equations. Applied Mathematics and Computation, 203(1): 50–55, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300308001926.

Parida:2008:IMF

[PG08b] P. K. Parida and D. K. Gupta. An improved method for finding multiple roots and it’s multiplicity of nonlinear equations in R. Applied Mathematics and Computation, 202(2):498–503, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308001240.

Polat:2008:NDR

[PG08c] Kemal Polat and Salih Güneş. A novel data reduction method: Distance based data reduction and its application to classification of epileptiform EEG signals. Applied Mathematics and Computation, 200(1):10– 27, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307012106.

Park:2008:PHE

[PGdSS08] Ho-Hyun Park, Alexandre Grings, Marcus Vinicius dos Santos, and Alexsandro Santos Soares. Parallel hybrid evolutionary computation: Automatic tuning of parameters for parallel gene expression programming. Applied Mathematics and Computation, 201(1–2):108–120, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011757.

Phohomsiri:2006:APR

[PH06] Phailaung Phohomsiri and Byungrin Han. An alternative proof for the recursive formulae for computing the Moore–Penrose M-inverse of a matrix. Applied Mathematics and Computation, 174(1):81– 97, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004431.

Patricio:2009:SAR

[PH09] P. Patrı´ cio and R. E.

Hartwig. Some additive results on Drazin inverses. Applied Mathematics and Computation, 215(2):530–538, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309005074.

Phat:2006:GSL

[Pha06] Vu N. Phat. Global stabilization for linear continuous time-varying systems. Applied Mathematics and Computation, 175(2):1730–1743, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305007344.

Pan:2006:FRC

[PHL06] Ping-Qi Pan, Jian-Feng Hu, and Chen Li. Feasible region contraction interior point algorithm. Applied Mathematics and Computation, 182(2):1361–1368, November 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306004504.

Petkovic:2009:SMN

[PHP09] M. S. Petković, – D. Herceg, and I. Petković. On a simultaneous method of Newton–Weierstrass’ type for finding all zeros of a polynomial. Applied Mathematics and Computation, 215(7):2456–2463, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007814.

Peng:2005:IMS

[PHZ05] Ya-Xin Peng, Xi-Yan Hu, and Lei Zhang. An iteration method for the symmetric solutions and the optimal approximation solution of the matrix equation AXB = C. Applied Mathematics and Computation, 160(3):763–777, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303011937.

Peng:2006:KIE

[PHZ06a] Juan Peng, Xi-Yan Hu, and Lei Zhang. A kind of inverse eigenvalue problems of Jacobi matrix. Applied Mathematics and Computation, 175(2):1543–1555, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007137.

Peng:2006:IMS

[PHZ06b] Ya-Xin Peng, Xi-Yan Hu, and Lei Zhang. An iterative method for symmetric solutions and optimal approximation solution of the system of matrix equations A _{1} XB _{1} = C _{1}, A _{2} XB 2 = C _{2}. Applied Mathematics and Computation, 183(2):1127–1137, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006473.

Peng:2008:SPP

[PHZ08] Juan Peng, Xi-Yan Hu, and Lei Zhang. The (M, N)-symmetric Procrustes problem. Applied Mathematics and Computation, 198(1):24– 34, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008636.

Pinasco:2006:CEL

[Pin06] Juan Pablo Pinasco. Comparison of eigenvalues for the p-Laplacian with integral inequalities. Applied Mathematics and Computation, 182(2):1399–1404, November 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306004577.

Pinheiro:2007:MWT

[Pin07] M. R. Pinheiro. Medium worlds theories I. Applied Mathematics and Computation, 188(2):1061–1070, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006734.

Plesser:2010:RSI

[PJ10] Hans Ekkehard Plesser and Anders Grønvik Jahnsen. Re-seeding invalidates tests of random number generators. Applied Mathematics and Computation, 217 (1):339–346, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [KCK08].

Park:2008:STD

[PJWL08] Ju H. Park, D. H. Ji, S. C. Won, and S. M. Lee. H_{∞} synchronization of time-delayed chaotic systems. Applied Mathematics and Computation, 204(1):170–177, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800489X.

Park:2005:SLA

[PK05a] Je Hong Park and Bo Gyeong Kang. Security of Lu et al.’s proxy-protected signature scheme. Applied Mathematics and Computation, 169(2):1010–1015, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008306.

Park:2005:GCC

[PK05b] Ju H. Park and O. Kwon. On guaranteed cost control of neutral systems by retarded integral state feedback. Applied Mathematics and Computation, 165(2):393–404, June 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003893.

Park:2005:NSC

[PK05c] Ju H. Park and O. Kwon. On new stability criterion for delay-differential systems of neutral type. Applied Mathematics and Computation, 162(2):627– 637, March 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304001195.

Park:2005:RSN

[PK05d] Ju H. Park and O. Kwon.

On robust stabilization for neutral delay-differential systems with parametric uncertainties and its application. Applied Mathematics and Computation, 162(3):1167–1182, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001766.

Park:2008:EUU

[PK08a] Jong Yeoul Park and Jum Ran Kang. Existence, uniqueness and uniform decay for the nonlinear degenerate equation with memory condition at the boundary. Applied Mathematics and Computation, 202(2):481–488, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800043X.

Park:2008:DSE

[PK08b] Ju H. Park and O. M.

Kwon. Design of state estimator for neural networks of neutral-type. Applied Mathematics and Computation, 202(1):360–369, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308001197.

Park:2008:IDD

[PK08c] Ju H. Park and O. M. Kwon. On improved delaydependent criterion for global stability of bidirectional associative memory neural networks with timevarying delays. Applied Mathematics and Computation, 199(2):435–446, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010107.

Peiravi:2008:FAC

[PK08d] Ali Peiravi and Hossein Tolooei Kheibari. A fast algorithm for connectivity graph approximation using modified Manhattan distance in dynamic networks. Applied Mathematics and Computation, 201(1–2):319–332, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307012088.

Park:2009:FRS

[PK09] Ju H. Park and O. M. Kwon. Further results on state estimation for neural networks of neutraltype with time-varying delay. Applied Mathematics and Computation, 208(1):69–75, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008485.

Park:2005:CZA

[PKH05] Je Hong Park, Bo Gyeong Kang, and Jae Woo Han. Cryptanalysis of Zhou et al.’s proxy-protected signature schemes. Applied Mathematics and Computation, 169(1):192–197, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007696.

Park:2008:LOAb

[PKL08a] Ju H. Park, O. M. Kwon, and S. M. Lee. LMI optimization approach on stability for delayed neural networks of neutraltype. Applied Mathematics and Computation, 196(1):236–244, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006637.

Park:2008:LOAa

[PKL08b] Ju H. Park, O. M. Kwon, and S. M. Lee. LMI optimization approach to stabilization of Genesio–Tesi chaotic system via dynamic controller. Applied Mathematics and Computation, 196(1):200–206, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006595.

Park:2008:SEN

[PKL08c] Ju H. Park, O. M. Kwon, and S. M. Lee. State estimation for neural networks of neutral-type with interval time-varying delays. Applied Mathematics and Computation, 203(1):217–223, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002324.

Polat:2006:ANSa

[PKT06a] Necat Polat, Doğan Kaya, and H. Ilhan Tutalar. A analytic and numerical solution to a modified Kawahara equation and a convergence analysis of the method. Applied Mathematics and Computation, 179(2):466–472, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010520.

Polat:2006:ANSb

[PKT06b] Necat Polat, Doğan Kaya, and H. Ilhan Tutalar. A analytic and numerical solution to a modified Kawahara equation and a convergence analysis of the method. Applied Mathematics and Computation, 181(1):193–199, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000890.

Pan:2005:EAE

[PL05] Shaohua Pan and Xingsi Li. An efficient algorithm for the Euclidean r-centrum location problem. Applied Mathematics and Computation, 167(1):716–728, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005259.

Pan:2006:EAS

[PL06a] Shaohua Pan and Xingsi Li. An efficient algorithm for the smallest enclosing ball problem in high dimensions. Applied Mathematics and Computation, 172(1):49–61, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305001852.

Peng:2006:DFF

[PL06b] Yehui Peng and Zhenhai Liu. A derivativefree filter algorithm for nonlinear complementarity problem. Applied Mathematics and Computation, 182(1):846–853, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004024.

Peng:2006:VMM

[PL06c] Yehui Peng and Zhenhai Liu. A variable metric method for nonsmooth convex constrained optimization. Applied Mathematics and Computation, 183(2):961–971, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006308.

Pon:2005:MDS

[PLJ05a] Shun-Fu Pon, Erl-Huei Lu, and Albert B. Jeng.

Meta-He digital signatures based on factoring and discrete logarithms. Applied Mathematics and Computation, 165(1):171–176, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003704.

Pon:2005:OPK

[PLJ05b] Shun-Fu Pon, Erl-Huei Lu, and Albert B. Jeng. One private-key for all DL-based cryptosystems. Applied Mathematics and Computation, 170(1):666– 672, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305000081.

Pang:2008:CSI

[PLL08] W. K. Pang, Ronghua Li, and Min Liu. Convergence of the semi-implicit Euler method for stochastic age-dependent population equations. Applied Mathematics and Computation, 195(2):466–474, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005772.

Liu:2006:HAC

[pLT06] Ya ping Liu and Lü Tao. High accuracy combination algorithm and a posteriori error estimation for solving the first kind Abel integral equations. Applied Mathematics and Computation, 178(2):441–451, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009896.

Peng:2007:DBA

[PLTW07] Jian-Xun Peng, Kang Li, Steve Thompson, and Peter A. Wieringa. Distributionbased adaptive bounding genetic algorithm for continuous optimisation problems. Applied Mathematics and Computation, 185(2):1063–1077, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008307.

Piccolomini:2005:BSP

[PLZ05] E. Loli Piccolomini, G. Landi, and F. Zama. A B-spline parametric model for high resolution dynamic magnetic resonance imaging. Applied Mathematics and Computation, 164(1):133– 148, May 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003005.

Pan:2007:IDK

[PLZ07] Haiwei Pan, Jianzhong Li, and Wei Zhang. Incorporating domain knowledge into medical image clustering. Applied Mathematics and Computation, 185(2):844–856, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007661.

Petkovic:2005:IHL

[PM05] M. S. Petković and D. Milošević.

Improved Halley-like methods for the inclusion of polynomial zeros. Applied Mathematics and Computation, 169(1):417–436, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007830.

Pal:2009:ITD

[PM09] Dulal Pal and Hiranmoy Mondal. Influence of temperature-dependent viscosity and thermal radiation on MHD forced convection over a nonisothermal wedge. Applied Mathematics and Computation, 212(1):194–208, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001349.

Petkovic:2008:HOE

[PMP08] M. S. Petković, D. M. Milošević, and L. D. Petković. High order Euler-like method for the inclusion of polynomial zeros. Applied Mathematics and Computation, 196(2):762–773, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700728X.

Pour-Mahmoud:2005:NSS

[PMRAS05a] J. Pour-Mahmoud, M. Y. Rahimi-Ardabili, and S. Shahmorad. Numerical solution of the system of Fredholm integro-differential equations by the Tau method. Applied Mathematics and Computation, 168(1):465–478, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006186.

Pour-Mahmoud:2005:NSV

[PMRAS05b] J. Pour-Mahmoud, M. Y.

Rahimi-Ardabili, and S. Shahmorad. Numerical solution of Volterra integrodifferential equations by the Tau method with the Chebyshev and Legendre bases. Applied Mathematics and Computation, 170(1):314–338, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009282.

Pathak:2009:RDF

[PMS09] Sweta Pathak, Alakes Maiti, and G. P. Samanta. Rich dynamics of a food chain model with Hassell– Varley type functional responses. Applied Mathematics and Computation, 208(2):303–317, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009168.

Pasandideh:2006:MRS

[PN06] Seyed Hamid Reza Pasandideh and Seyed Taghi Akhavan Niaki. Multi-response simulation optimization using genetic algorithm within desirability function framework. Applied Mathematics and Computation, 175(1):366–382, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006144.

Pasandideh:2008:GAA

[PN08] Seyed Hamid Reza Pasandideh and Seyed Taghi Akhavan Niaki. A genetic algorithm approach to optimize a multi-products EPQ model with discrete delivery orders and constrained space. Applied Mathematics and Computation, 195(2):506–514, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005802.

Pan:2006:NPS

[PNB06] Jian-Yu Pan, Michael K. Ng, and Zhong-Zhi Bai. New preconditioners for saddle point problems. Applied Mathematics and Computation, 172(2):762– 771, January 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304009105.

Plienpanich:2005:CSP

[PNL05] Tidarut Plienpanich, Piyapong Niamsup, and Yongwimon Lenbury. Controllability and stability of the perturbed Chen chaotic dynamical system. Applied Mathematics and Computation, 171(2):927–947, December 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Podisuk:2006:OFR

[Pod06] Maitree Podisuk. Open formula of Runge–Kutta method for solving autonomous ordinary differential equation. Applied Mathematics and Computation, 181(1):536–542, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001238.

Pogany:2007:CGK

[Pog07a] Tibor K. Pogány. Convergence of generalized Kapteyn expansion. Applied Mathematics and Computation, 190(2):1844– 1847, July 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307001683.

Pogany:2007:WTD

[Pog07b] Tibor K. Pogány. Whittakertype derivative sampling reconstruction of stochastic L^{α}(Ω)-processes. Applied Mathematics and Computation, 187(1):384–394, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030601188X.

Pourdarvish:2006:RSI

[Pou06] Ahmad Pourdarvish. A reliable symbolic implementation of algorithm for calculating Adomian polynomials. Applied Mathematics and Computation, 172(1):545–550, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305002225.

Petkovic:2007:NSR

[PP07a] Ljiljana D. Petković and Miodrag S. Petković. A note on some recent methods for solving nonlinear equations. Applied Mathematics and Computation, 185(1):368–374, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008691.

Petkovic:2007:CZF

[PP07b] M. S. Petković and L. D.

Petković. Construction of zero-finding methods by Weierstrass functions. Applied Mathematics and Computation, 184(2):351–359, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007120.

Petkovic:2007:OPS

[PP07c] Miodrag S. Petković and Ljiljana D. Petković. A one parameter square root family of two-step root-finders. Applied Mathematics and Computation, 188(1):339– 344, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013579.

Plubtieng:2008:NIM

[PP08] Somyot Plubtieng and Rattanaporn Punpaeng. A new iterative method for equilibrium problems and fixed point problems of nonexpansive mappings and monotone mappings. Applied Mathematics and Computation, 197(2):548– 558, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008016.

Palligkinis:2009:RKM

[PPF09] S. Ch. Palligkinis, G. Papageorgiou, and I. Th. Famelis. Runge–Kutta methods for fuzzy differential equations. Applied Mathematics and Computation, 209(1):97–105, March 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004785.

Park:2008:NSC

[PPKL08] Ju H. Park, C. H. Park, O. M. Kwon, and S. M. Lee. A new stability criterion for bidirectional associative memory neural networks of neutral-type. Applied Mathematics and Computation, 199(2):716– 722, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010582.

Pandey:2009:SGS

[PPS09] Manoj Pandey, B. D. Pandey, and V. D. Sharma. Symmetry groups and similarity solutions for the system of equations for a viscous compressible fluid. Applied Mathematics and Computation, 215(2):681–685, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005475.

Pecaric:2008:EMW

[PPV08] J. Pečarić, M. Ribičić Penava, and A. Vukelić. Euler’s method for weighted integral formulae. Applied Mathematics and Computation, 206(1):445–456, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006966.

Park:2008:MCE

[PQJ08] Jeong-Soo Park, Guoqi Q.

Qian, and Yuna Jun. Monte Carlo EM algorithm in logistic linear models involving non-ignorable missing data. Applied Mathematics and Computation, 197(1):440–450, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007941.

Petkovic:2005:NFO

[PR05] M. S. Petković and L. Rančić.

A new fourth-order family of simultaneous methods for finding polynomial zeros. Applied Mathematics and Computation, 164(1):227–248, May 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030400308X.

Prasad:2006:FFP

[PR06] Sharat C. Prasad and K. R. Rajagopal. Flow of a fluid with pressure dependent viscosity due to a boundary that is being stretched. Applied Mathematics and Computation, 173(1):50–68, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305003115.

Pinter:2007:ZSB

[PR07] Á. Pintér and Cs. Rakaczki. On the zeros of shifted Bernoulli polynomials. Applied Mathematics and Computation, 187(1):379– 383, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011878.

Parida:2009:RNI

[PR09a] Nigam Chandra Parida and Soumyendu Raha. Regularized numerical integration of multibody dynamics with the generalized α method. Applied Mathematics and Computation, 215(3):1224–1243, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006377.

Pogany:2009:EQT

[PR09b] Tibor K. Pogány and Arjun K. Rathie. Extension of a quadratic transformation due to Exton. Applied Mathematics and Computation, 215(1):423–426, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004500.

Prabha:2007:IER

[Pra07] Hem Prabha. H_{J} integrals and evaluation of rectangular source integral. Applied Mathematics and Computation, 184(2):680–684, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007430.

Prastaro:2008:ECPa

[Prá08a] Agostino Prástaro. On the extended crystal PDE’s stability. I: The n - d’Alembert extended crystal PDE’s. Applied Mathematics and Computation, 204(1):63–69, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004554.

Prastaro:2008:ECPb

[Prá08b] Agostino Prástaro. On the extended crystal PDE’s stability. II: Entropyregular-solutions in MHD-PDE’s. Applied Mathematics and Computation, 204(1):82–89, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800458X.

Pastor:2009:GVP

[PRA^{+}09] G. Pastor, M. Romera, G. Alvarez, D. Arroyo, A. B. Orue, V. Fernandez, and F. Montoya. A general view of pseudoharmonics and pseudoantiharmonics to calculate external arguments of Douady and Hubbard. Applied Mathematics and Computation, 213(2):484–497, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002604.

Prentice:2006:CNI

[Pre06a] J. S. C. Prentice. Comments on numerically improved Gauss–Legendre quadrature. Applied Mathematics and Computation, 174(2):1454–1459, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005631.

See erratum [BJE05, Pre06b].

Prentice:2006:ECN

[Pre06b] J. S. C. Prentice. Erratum to “Comments on numerically improved Gauss–Legendre quadrature” [Appl. Math. Comput. 174 (2) (2006) 1454]. Applied Mathematics and Computation, 183(1):709, December 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306004711. See

[Pre06a].

Parmananda:2005:CCU

[PRGH05] P. Parmananda, M. Rivera, B. J. Green, and J. L. Hudson. Controlling complexity using forcing: simulations and experiments. Applied Mathematics and Computation, 164(2):467–491, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003303.

Pang:2006:ESN

[PRM06] W. K. Pang, Li Ronghua, and Liu Ming. Exponential stability of numerical solutions to stochastic age-dependent population equations. Applied Mathematics and Computation, 183(1):152–159, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005480.

Papi:2006:OAL

[PS06a] Marco Papi and Simone Sbaraglia. Optimal asset– liability management with constraints: A dynamic programming approach. Applied Mathematics and Computation, 173(1):306– 349, February 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305003541.

Patidar:2006:UCN

[PS06b] Kailash C. Patidar and Kapil K. Sharma. ϵ-Uniformly convergent nonstandard finite difference methods for singularly perturbed differential difference equations with small delay. Applied Mathematics and Computation, 175(1):864–890, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006508.

Piejko:2006:DSO

[PS06c] Krzysztof Piejko and Janusz Sokół. On the Dziok– Srivastava operator under multivalent analytic functions. Applied Mathematics and Computation, 177(2):839–843, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009641.

Padhi:2008:MPS

[PS08] Seshadev Padhi and Shilpee Srivastava. Multiple periodic solutions for a nonlinear first order functional differential equations with applications to population dynamics. Applied Mathematics and Computation, 203(1):1–6, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308001781.

Polak:2008:ANN

[PSBM08] Sebastian Polak, Agnieszka Skowron, Jerzy Brandys, and Aleksander Mendyk. Artificial neural networks based modeling for pharmacoeconomics application. Applied Mathematics and Computation, 203(2):482–492, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010818.

Prax:2007:EHO

[PSD07] C. Prax, H. Sadat, and E. Dabboura. Evaluation of high order versions of the diffuse approximate meshless method. Applied Mathematics and Computation, 186(2):1040–1053, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010307.

Psihoyios:2007:STD

[Psi07] G. Psihoyios. Solving time dependent PDEs via an improved modified extended BDF scheme. Applied Mathematics and Computation, 184(1):104–115, January 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004899.

Psihoyios:2009:EAS

[Psi09] G. Psihoyios. Explicit advanced step-point (EAS) methods and the EAS2 multistep scheme for the solution of non-stiff initial value problems. Applied Mathematics and Computation, 209(1):106–124, March 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004797.

Platkowski:2005:EPM

[PSP05] Tadeusz Płatkowski and Joanna Stachowska-Pietka. ESSs in n-player mixed games. Applied Mathematics and Computation, 167(1):592–606, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005168.

Peer:2006:CAL

[PSRN06] S. K. Peer, Dinesh K.

Sharma, K. Ravindranath, and M. M. Naidu. Computation of arc length in the presence of barriers in networks. Applied Mathematics and Computation, 175(1):541–556, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006284.

Peer:2007:FSP

[PSS07] S. K. Peer, Dinesh K. Sharma, and Rakesh K. Sharma. Finding shortest path in the presence of barriers: An alternate approach. Applied Mathematics and Computation, 185(1):333–345, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008642.

Pogany:2006:SFM

[PST06] Tibor K. Pogány, H. M. Srivastava, and Živorad Tomovski. Some families of Mathieu a-series and alternating Mathieu a-series. Applied Mathematics and Computation, 173(1):69–108, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305003127.

Pang:2007:TPD

[PSW07] Li-Ping Pang, Jie Shen, and Wei Wang. Two parallel distribution algorithms for convex constrained minimization problems. Applied Mathematics and Computation, 186(2):1762– 1771, March 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030601112X.

Paramanathan:2008:ACF

[PU08] P. Paramanathan and R. Uthayakumar. An algorithm for computing the fractal dimension of waveforms. Applied Mathematics and Computation, 195(2):598–603, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005899.

Pudmenzky:2005:TEM

[Pud05]  Alexander  Pudmenzky.

Teleonomic entropy: measuring the phase-space of end-directed systems. Applied Mathematics and Computation, 162(2):695– 705, March 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304001250.

Pavlidis:2007:ECS

[PVM07] N. G. Pavlidis, M. N. Vrahatis, and P. Mossay. Existence and computation of short-run equilibria in economic geography. Applied Mathematics and Computation, 184(1):93–103, January 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004887.

Pereira:2006:FAM

[PVN06a] Fábio Henrique Pereira, Sérgio Luís Lopes Verardi, and Silvio Ikuyo Nabeta. A fast algebraic multigrid preconditioned conjugate gradient solver. Applied Mathematics and Computation, 179(1):344–351, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010416.

Pereira:2006:WBA

[PVN06b] Fábio Henrique Pereira, Sérgio Luís Lopes Verardi, and Sílvio Ikuyo Nabeta. A wavelet-based algebraic multigrid preconditioner for sparse linear systems. Applied Mathematics and Computation, 182(2):1098–1107, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004255.

Perez:2008:HTB

[PVS^{+}08] J. Pérez, F. R. Villatoro, J. L. G. Santander, M. A. Borovsky, Yu. L. Ratis, E. A. Izzheurov, and P. Fernández de Córdoba. Heat transfer between a gas and an ultralow thermal conductivity porous structure. Applied Mathematics and Computation, 204(2):687–693, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005390.

Pandey:2009:FDS

[PVV09] K. Pandey, Lajja Verma, and Amit K. Verma. On a finite difference scheme for Burgers’ equation. Applied Mathematics and Computation, 215(6):2206–2214, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007310.

Ponnusamy:2009:RVC

[PVY09] S. Ponnusamy, A. Vasudevarao, and H. Yanagihara. Region of variability for close-to-convex functions-II. Applied Mathematics and Computation, 215(3):901–915, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005773.

Pang:2005:NMS

[PW05] Liao-Jun Pang and Yu-Min Wang. A new ( t n ) multi-secret sharing scheme based on Shamir’s secret sharing. Applied Mathematics and Computation, 167(2):840–848, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005338.

Pao:2008:NST

[PW08a] C. V. Pao and Yuan-Ming Wang. Numerical solutions of a three-competition Lotka–Volterra system. Applied Mathematics and Computation, 204(1):423– 440, October 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308005213.

Peng:2008:UBO

[PW08b] Xiao-Fei Peng and WenLi.

On unsymmetric block overrelaxation-type methods for saddle point problems. Applied Mathematics and Computation, 203(2):660–671, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800297X.

Pan:2009:TWF

[PW09] Shuxia Pan and Shimin Wan. Traveling wave fronts in a delayed population model of Daphnia magna. Applied Mathematics and Computation, 215(3):1118–1123, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006201.

Pan:2006:PSO

[PWL06] Hui Pan, Ling Wang, and Bo Liu. Particle swarm optimization for function optimization in noisy environment. Applied Mathematics and Computation, 181(2):908–919, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001718.

Peng:2008:SPH

[PWY08] Rui Peng, Mingxin Wang, and Guoying Yang. Stationary patterns of the Holling–Tanner prey–predator model with diffusion and cross-diffusion. Applied Mathematics and Computation, 196(2):570–577, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307007102.

Pakdemirli:2007:ERP

[PY07] M. Pakdemirli and H. A. Yurtsever. Estimating roots of polynomials using perturbation theory. Applied Mathematics and Computation, 188(2):2025–2028, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015712.

Peng:2008:BDP

[PY08a] Congming Peng and Zuodong Yang. Blow-up for a degenerate parabolic equation with a nonlocal source. Applied Mathematics and Computation, 201(1–2):250–259, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011927.

Petkovic:2008:SLF

[PY08b] Miodrag S. Petković and Beong In Yun. Sigmoidlike functions and root finding methods. Applied Mathematics and Computation, 204(2):784–793, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005614.

Qadir:2007:GSF

[Qad07] Asghar Qadir. The generalization of special functions. Applied Mathematics and Computation, 187(1):395– 402, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011891.

Qesmi:2006:CMN

[QBH06a] R. Qesmi, M. Ait Babram, and M. L. Hbid. Center manifolds and normal forms for a class of retarded functional differential equations with parameter associated with Fold-Hopf singularity. Applied Mathematics and Computation, 181(1):220–246, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000932.

Qesmi:2006:MPC

[QBH06b] R. Qesmi, M. Ait Babram, and M. L. Hbid. A Maple program for computing a terms of a center manifold, and element of bifurcations for a class of retarded functional differential equations with Hopf singularity. Applied Mathematics and Computation, 175(2):932–968, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006545.

Qiu:2005:IOS

[QC05] Weidong Qiu and Kefei Chen. Identity oriented signature scheme based on quadratic residues. Applied Mathematics and Computation, 168(1):235–242, September 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304006010.

Qualls:2007:POD

[QC07] C. Qualls and D. A. Cicci. Preliminary orbit determination of a tethered satellite. Applied Mathematics and Computation, 188(1):462–471, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013865.

Qiu:2008:ESC

[QC08] Jianlong Qiu and Jinde Cao. Exponential stability of a competitive Lotka– Volterra system with delays. Applied Mathematics and Computation, 201(1–2):819–829, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011617.

Qian:2005:CLT

[QCB05a] Haifeng Qian, ZhenFu Cao, and Haiyong Bao. Cryptanalysis of Li–Tzeng– Hwang’s improved signature schemes based on factoring and discrete logarithms. Applied Mathematics and Computation, 166(3):501–505, July 26, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004473.

Qian:2005:SPL

[QCB05b] Haifeng Qian, Zhenfu Cao, and Haiyong Bao. Security of Pon–Lu–Jeng’s Meta-He digital signature schemes. Applied Mathematics and Computation, 170(1):724–730, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000135.

Qi:2009:APM

[QCDS09] Feng Qi, Pietro Cerone, Sever S. Dragomir, and H. M. Srivastava. Alternative proofs for monotonic and logarithmically convex properties of oneparameter mean values. Applied Mathematics and Computation, 208(1):129– 133, February 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308008709.

Qi:2008:GVC

[QCN08] Feng Qi, Jian Cao, and Da-Wei Niu. A generalization of van der Corput’s inequality. Applied Mathematics and Computation, 203(2):770–777, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004025.

Qin:2009:CAI

[QCS09] Xiaolong Qin, Yeol Je Cho, and Meijuan Shang. Convergence analysis of implicit iterative algorithms for asymptotically nonexpansive mappings. Applied Mathematics and Computation, 210(2):542–550, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000368.

Qian:2009:NRS

[QCSF09] Wei Qian, Shen Cong, Youxian Sun, and Shumin Fei. Novel robust stability criteria for uncertain systems with time-varying delay. Applied Mathematics and Computation, 215(2):866–872, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005657.

Qi:2008:APS

[QD08] Wang Qi and Bingxiang Dai. Almost periodic solution for n-species Lotka– Volterra competitive system with delay and feedback controls. Applied Mathematics and Computation, 200(1):133–146, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010946.

Du:2008:GCM

[qDyC08] Shou qiang Du and Yuan yuan Chen. Global convergence of a modified spectral FR conjugate gradient method. Applied Mathematics and Computation, 202(2):766–770, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800177X.

Qian:2006:MMH

[QFF06] Zhi Qian, Chu-Li Fu, and Xiao-Li Feng. A modified method for high order numerical derivatives. Applied Mathematics and Computation, 182(2):1191–1200, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004334.

Qian:2007:MMB

[QFS07] Zhi Qian, Chu-Li Fu, and Rui Shi. A modified method for a backward heat conduction problem. Applied Mathematics and Computation, 185(1):564–573, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008915.

Qian:2006:MMN

[QFX06] Zhi Qian, Chu-Li Fu, and Xiang-Tuan Xiong. A modified method for a non-standard inverse heat conduction problem. Applied Mathematics and Computation, 180(2):453–468, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000087.

Qian:2006:FTM

[QFXW06] Zhi Qian, Chu-Li Fu, Xiang-Tuan Xiong, and Ting Wei. Fourier truncation method for high order numerical derivatives. Applied Mathematics and Computation, 181(2):940–948, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001743.

Qi:2008:WGI

[QG08] Feng Qi and Bai-Ni Guo.

Wendel’s and Gautschi’s inequalities: Refinements, extensions, and a class of logarithmically completely monotonic functions. Applied Mathematics and Computation, 205(1):281–290, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005377.

Qian:2008:NWM

[Qia08]  Ailin Qian.  A new

wavelet method for solving an ill-posed problem. Applied Mathematics and Computation, 203(2):635–640, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002944.

Qiang:2009:ASA

[Qia09]  Jing Qiang.  Additive

Schwarz algorithms for parabolic problems. Applied Mathematics and Computation, 208(2):531– 541, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308009260.

Qiu:2005:CND

[Qiu05] Weidong Qiu. Converting normal DLP-based signatures into blind. Applied Mathematics and Computation, 170(1):657–665, November 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030500007X.

Jin:2006:PM

[qJ06]  Xiao qing Jin.  M-

preconditioner for Mmatrices. Applied Mathematics and Computation, 172(2):701–707, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400904X.

Qi:2006:STD

[QL06] Mingnan Qi and Sanyang Liu. Selecting two different defective coins. Applied Mathematics and Computation, 180(2):559–568, September 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306000440.

Qian:2008:ADE

[Ql08] Weiyi Qian and Ajun li. Adaptive differential evolution algorithm for multiobjective optimization problems. Applied Mathematics and Computation, 201(1–2):431–440, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307012192.

Qi:2008:DFI

[QLG08] Feng Qi, Qiu-Ming Luo, and Bai-Ni Guo. Darboux’s formula with integral remainder of functions with two independent variables. Applied Mathematics and Computation, 199(2):691– 703, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010569.

Liu:2005:ESM

[qLJ05] Xi qiang Liu and Song Jiang. Exact solutions of multi-component nonlinear Schrödinger and Klein– Gordon equations. Applied Mathematics and Computation, 160(3):875–880, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303012001.

Qikui:2005:EAA

[QM05] Du Qikui and Tang Minxia.

Exact and approximate artificial boundary conditions for the hyperbolic problems in unbounded domains. Applied Mathematics and Computation, 169(1):544–562, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007933.

Ma:2009:IMS

[qMpXjH09] Hui qiang Ma, Jiu ping Xu, and Nan jing Huang. An iterative method for a system of linear complementarity problems with perturbations and interval data. Applied Mathematics and Computation, 215(1):175–184, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004263.

Mei:2005:NSL

[qMqZ05] Jian qin Mei and Hong qing Zhang. New solitonlike and periodic-like solutions for the KdV equation. Applied Mathematics and Computation, 169(1):589–599, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304007969.

Qin:2008:GWH

[QN08] Xiaolong Qin and Muhammad Aslam Noor. General Wiener–Hopf equation technique for nonexpansive mappings and general variational inequalities in Hilbert spaces. Applied Mathematics and Computation, 201(1–2):716–722, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308000386.

Qin:2005:TGS

[qQcM05] Xin qiang Qin and Yi chen Ma. Two-grid scheme for characteristics finiteelement solution of nonlinear convection diffusion problems. Applied Mathematics and Computation, 165(2):419–431, June 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003911.

Qatanani:2006:ANI

[QS06] Naji Qatanani and Monika Schulz. Analytical and numerical investigation of the Fredholm integral equation for the heat radiation problem. Applied Mathematics and Computation, 175(1):149–170, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005989.

Qin:2009:AAS

[QS09] Xiao-Ying Qin and Yan-Ping Sun. Approximate analytic solutions for a twodimensional mathematical model of a packedbed electrode using the Adomian decomposition method. Applied Mathematics and Computation, 215(1):270–275, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004573.

Qi:2008:CRD

[QSS08] Jinpeng Qi, Shihuang Shao, and Yizhen Shen. Cellular responding DNA damage: an improved modeling of P53 gene regulatory networks under ion radiation (IR). Applied Mathematics and Computation, 205(1):73– 83, November 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300308003196.

Qin:2008:SCG

[QSZ08] Xiaolong Qin, Meijuan Shang, and Haiyun Zhou. Strong convergence of a general iterative method for variational inequality problems and fixed point problems in Hilbert spaces. Applied Mathematics and Computation, 200(1):242–253, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011034.

Quang:2007:IMS

[Qua07]  A. Dang Quang.  It-

erative method for solving a nonlinear boundary value problem. Applied Mathematics and Computation, 190(2):1675–1682, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002123.

Quintanilla:2009:SBS

[Qui09]  R. Quintanilla.  Spa-

tial behaviour of solutions of the three-phase-lag heat equation. Applied Mathematics and Computation, 213(1):153–162, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002124.

Qin:2005:DSW

[QW05] Mei Qin and Guanglin Wang. Displacement structure of W -weighted Drazin inverse A_{d,W} and its perturbation. Applied Mathematics and Computation, 162(1):403–419, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001031.

Qu:2006:OPQ

[QW06] Chaochun Qu and Ping Wang. Optimization in production quota problem with convex cost function. Applied Mathematics and Computation, 177(2):652–658, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500946X.

Qin:2008:TTR

[QW08a] H. H. Qin and D. W.

Wen. Tikhonov type regularization method for the Cauchy problem of the modified Helmholtz equation. Applied Mathematics and Computation, 203(2):617–628, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002920.

Qiu:2008:LSS

[QW08b] Yuyang Qiu and Anding Wang. Least squares solutions to the equations AX = B, XC = D with some constraints. Applied Mathematics and Computation, 204(2):872–880, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005729.

Qian:2009:CMC

[QW09] Ailin Qian and Yujiang Wu. A computational method for a Cauchy problem of Laplace’s equation. Applied Mathematics and Computation, 207(2):478–485, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008345.

Quan:2008:MPN

[QWX08] Jin-Juan Quan, Xian-Bin Wen, and Xue-Quan Xu. Multiscale probabilistic neural network method for SAR image segmentation. Applied Mathematics and Computation, 205(2):578–583, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003366.

Xu:2005:PPP

[qXbL05] Gui qiong Xu and Zhi bin Li. PDEPtest: a package for the Painlevé test of nonlinear partial differential equations. Applied Mathematics and Computation, 169(2):1364–1379, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008574.

Qian:2008:PSO

[QYYL08] Weiyi Qian, Yingjie Yang, Ningning Yang, and Chun Li. Particle swarm optimization for SNP haplotype reconstruction problem. Applied Mathematics and Computation, 196(1):266–272, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006674.

Qi:2009:PEA

[QZ09] Yongfeng Qi and Jiashu Zhang. (2D)^{2} PCALDA: an efficient approach for face recognition. Applied Mathematics and Computation, 213(1):1–7, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002021.

Qu:2007:GOA

[QZJ07a] Shao-Jian Qu, Ke-Cun Zhang, and Ying Ji. A global optimization algorithm using parametric linearization relaxation. Applied Mathematics and Computation, 186(1):763– 771, March 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306009969.

Qu:2007:NGO

[QZJ07b] Shao-Jian Qu, Ke-Cun Zhang, and Ying Ji. A new global optimization algorithm for signomial geometric programming via Lagrangian relaxation. Applied Mathematics and Computation, 184(2):886–894, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007776.

Qiu:2007:MEI

[QZL07a] Yuyang Qiu, Zhenyue Zhang, and Junfeng Lu. The matrix equations AX = B, XC = D with P X = sXP constraint. Applied Mathematics and Computation, 189(2):1428–1434, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017292.

Qiu:2007:MIS

[QZL07b] Yuyang Qiu, Zhenyue Zhang, and Junfeng Lu. Matrix iterative solutions to the least squares problem of BXA^{T} = F with some linear constraints. Applied Mathematics and Computation, 185(1):284–300, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008605.

Qu:2007:EAG

[QZZ07] Shao-Jian Qu, Ke-Cun Zhang, and Jia-Kun Zhao. An efficient algorithm for globally minimizing sum of quadratic ratios problem with nonconvex quadratic constraints. Applied Mathematics and Computation, 189(2):1624–1636, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017474.

Ren:2009:CMN

[RA09] Hongmin Ren and Ioannis K. Argyros. On convergence of the modified Newton’s method under Hölder continuous Fréchet derivative. Applied Mathematics and Computation, 213(2):440–448, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002550.

Rababah:2005:DRT

[Rab05] Abedallah Rababah. Distances with rational triangular Bézier surfaces. Applied Mathematics and Computation, 160(2):379– 386, January 14, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300303011639.

Radwan:2005:HSG

[Rad05a] Ahmed E. Radwan. Hydromagnetic stability of gravitational streaming coaxial cylinders with double perturbed interfaces. Applied Mathematics and Computation, 160(1):213–244, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303007574.

Radwan:2005:ICA

[Rad05b] Ahmed E. Radwan. Instability of a compressible annular magnetized fluid. Applied Mathematics and Computation, 160(1):283–301, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303009317.

Radu:2009:SAG

[Rad09] Cristina Radu. On statistical approximation of a general class of positive linear operators extended in q-calculus. Applied Mathematics and Computation, 215(6):2317–2325, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007565.

Radwan:2005:PCA

[RAES05] Mohamad Radwan, M. K.

Ahmed, and F. A. Abd El-Salam. The problem of close approach of two small masses in the field of massive oblate primary. Applied Mathematics and Computation, 168(1):389–397, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006113.

Rafiq:2006:MNI

[Raf06]  Arif Rafiq.  Modified

Noor iterations for nonlinear equations in Banach spaces. Applied Mathematics and Computation, 182(1):589–595, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003675.

Rafiq:2007:NNC

[Raf07a] Arif Rafiq. A note on “New classes of iterative methods for nonlinear equations” and “Some iterative methods free from second derivatives for nonlinear equations”. Applied Mathematics and Computation, 193(2):572–576, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004328.

See [Noo07b, Noo07e].

Rafiq:2007:NNF

[Raf07b] Arif Rafiq. A note on “New family of iterative methods for nonlinear equations”. Applied Mathematics and Computation, 192(1):280–282, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003062.

See [Noo07c].

Rafiq:2008:NFM

[Raf08]  Arif Rafiq.  A note

on “A family of methods for solving nonlinear equations”. Applied Mathematics and Computation, 195(2):819–821, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005607.

See [UEL07].

Rezaie:2007:UEM

[RAG^{+}07] K. Rezaie, M. S. Amalnik, A. Gereie, B. Ostadi, and M. Shakhseniaee. Using extended Monte Carlo simulation method for the improvement of risk management: Consideration of relationships between uncertainties. Applied Mathematics and Computation, 190(2):1492–1501, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001841.

Rueda:2005:RMM

[RAGA^{+}05] M. Rueda, A. Arcos, S. González-Aguilera, M. D. Martínez-Miranda, Y. Román, and S. Martínez-Puertas. Ratio methods to the mean estimation with known quantiles. Applied Mathematics and Computation, 170(2):1031–1044, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000457.

Rajan:2008:PPC

[Raj08] M. P. Rajan. A posteriori parameter choice with an efficient discretization scheme for solving illposed problems. Applied Mathematics and Computation, 204(2):891–904, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005766.

Ramos:2005:EMOa

[Ram05a]  J. I. Ramos.  Equiva-

lence of C^{0} and C^{1} methods for ODE’s. Applied Mathematics and Computation, 167(2):1331–1338, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005703.

Ramos:2005:EMOb

[Ram05b]  J. I. Ramos.  Expo-

nential methods for onedimensional reaction–diffusion equations. Applied Mathematics and Computation, 170(1):380–398, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009567.

Ramos:2005:EFMa

[Ram05c] J. I. Ramos. An exponentiallyfitted method for singularly perturbed, onedimensional, parabolic problems. Applied Mathematics and Computation, 161(2):513–523, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000086.

Ramos:2005:EFMb

[Ram05d] J. I. Ramos. An exponentiallyfitted method for singularlyperturbed ordinary differential equations with turning points and parabolic problems. Applied Mathematics and Computation, 165(3):549–564, June 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004011.

Ramos:2005:EFMc

[Ram05e] J. I. Ramos. Exponentiallyfitted methods on layeradapted meshes. Applied Mathematics and Computation, 167(2):1311–1330, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005697.

Ramos:2005:LTSa

[Ram05f] J. I. Ramos. Linearization techniques for singular initial-value problems of ordinary differential equations. Applied Mathematics and Computation, 161(2):525–542, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000098.

Ramos:2005:LTSb

[Ram05g] J. I. Ramos. Linearization techniques for singularlyperturbed initial-value problems of ordinary differential equations. Applied Mathematics and Computation, 163(3):1143–1163, April 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002863.

Ramos:2005:PAM

[Ram05h] J. I. Ramos. A piecewiseanalytical method for singularly perturbed parabolic problems. Applied Mathematics and Computation, 161(2):501–512, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000074.

Ramos:2005:SLA

[Ram05i] J. I. Ramos. A smooth locally-analytical technique for singularly perturbed two-point boundary-value problems. Applied Mathematics and Computation, 163(3):1123–1142, April 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002851.

Ramachandran:2006:TAE

[Ram06a] M. P. Ramachandran. Two-axis Euler rotation determination. Applied Mathematics and Computation, 175(2):980–983, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006569.

Ramos:2006:DCF

[Ram06b]  J. I. Ramos.  Damp-

ing characteristics of finite difference methods for one-dimensional reaction– diffusion equations. Applied Mathematics and Computation, 182(1):607– 609, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306003699.

Ramos:2006:EFD

[Ram06c] J. I. Ramos. Explicit finite difference methods for the EW and RLW equations. Applied Mathematics and Computation, 179(2):622–638, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010684.

Ramos:2006:EMS

[Ram06d]  J. I. Ramos.  Expo-

nential methods for singularly perturbed ordinary differential–difference equations. Applied Mathematics and Computation, 182(2):1528–1541, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004942.

Ramos:2006:INI

[Ram06e] J. I. Ramos. Iterative and non-iterative, full and approximate factorization methods for multidimensional reaction–diffusion equations. Applied Mathematics and Computation, 174(2):1586–1608, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005825.

Ramos:2006:LIA

[Ram06f] J. I. Ramos. Linearlyimplicit, approximate factorization, exponential methods for multi-dimensional reaction–diffusion equations. Applied Mathematics and Computation, 174(2):1609–1633, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005837.

Ramos:2006:PLM

[Ram06g] J. I. Ramos. Piecewiselinearized methods for initial-value problems with oscillating solutions. Applied Mathematics and Computation, 181(1):123– 146, October 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030600083X.

Ramos:2007:NSE

[Ram07a] Higinio Ramos. A nonstandard explicit integration scheme for initialvalue problems. Applied Mathematics and Computation, 189(1):710–718, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030601647X.

Ramos:2007:AAC

[Ram07b]  J. I. Ramos.  Asymp-

totic analysis of channel flows with slip lengths that depend on the pressure. Applied Mathematics and Computation, 188(2):1310–1318, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014858.

Ramos:2007:FVM

[Ram07c]  J. I. Ramos.  A fi-

nite volume method for one-dimensional reaction– diffusion problems. Applied Mathematics and Computation, 188(1):739–748, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014123.

Ramos:2007:NMNa

[Ram07d] J. I. Ramos. Numerical methods for nonlinear second-order hyperbolic partial differential equations. I. Time-linearized finite difference methods for 1-D problems. Applied Mathematics and Computation, 190(1):722–756, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001099.

Ramos:2007:NMNb

[Ram07e] J. I. Ramos. Numerical methods for nonlinear second-order hyperbolic partial differential equations. II. Rothe’s techniques for 1-D problems. Applied Mathematics and Computation, 190(1):804– 832, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001142.

Ramos:2007:LPT

[Ram07f] J. I. Ramos. On Linstedt– Poincaré techniques for the quintic Duffing equation. Applied Mathematics and Computation, 193(2):303–310, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307003980.

Ramos:2007:PQT

[Ram07g] J. I. Ramos. Piecewisequasilinearization techniques for singularly perturbed Volterra integrodifferential equations. Applied Mathematics and Computation, 188(2):1221– 1233, May 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306014664.

Ramos:2007:PIM

[Ram07h] J. I. Ramos. Propagation and interaction of moving fronts in polymer crystallization. Applied Mathematics and Computation, 189(1):780–795, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016559.

Ramos:2008:LCN

[Ram08a] J. I. Ramos. Limit cycles of non-smooth oscillators. Applied Mathematics and Computation, 199(2):738– 747, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010612.

Ramos:2008:LGA

[Ram08b] J. I. Ramos. Linearized Galerkin and artificial parameter techniques for the determination of periodic solutions of nonlinear oscillators. Applied Mathematics and Computation, 196(2):483–493, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006996.

Ramos:2008:NID

[Ram08c] J. I. Ramos. A noniterative derivative-free method for nonlinear ordinary differential equations. Applied Mathematics and Computation, 203(2):672–678, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002981.

Ramos:2008:PLM

[Ram08d] J. I. Ramos. On the Picard–Lindelof method for nonlinear second-order differential equations. Applied Mathematics and Computation, 203(1):238– 242, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308002440.

Ramos:2008:VIM

[Ram08e] J. I. Ramos. On the variational iteration method and other iterative techniques for nonlinear differential equations. Applied Mathematics and Computation, 199(1):39– 69, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009691.

Ramos:2008:PHM

[Ram08f] J. I. Ramos. Piecewise homotopy methods for nonlinear ordinary differential equations. Applied Mathematics and Computation, 198(1):92–116, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008697.

Ramos:2009:INI

[Ram09a] J. I. Ramos. Iterative and non-iterative methods for non-linear Volterra integro-differential equations. Applied Mathematics and Computation, 214(1):287–296, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003038.

Ramos:2009:PIM

[Ram09b] J. I. Ramos. Picard’s iterative method for nonlinear advection–reaction– diffusion equations. Applied Mathematics and Computation, 215(4):1526– 1536, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309006535.

Rababah:2007:WDF

[RAN07] Abedallah Rababah and Mohammad Al-Natour. The weighted dual functionals for the univariate Bernstein basis. Applied Mathematics and Computation, 186(2):1581–1590, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010940.

Rababah:2008:WDF

[RAN08a] Abedallah Rababah and Mohammad Al-Natour. Weighted dual functions for Bernstein basis satisfying boundary constraints. Applied Mathematics and Computation, 199(2):456– 463, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010120.

Rang:2008:PCI

[Ran08b] J. Rang. Pressure corrected implicit θ-schemes for the incompressible Navier– Stokes equations. Applied Mathematics and Computation, 201(1–2):747–761, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308000428.

Ranjbar:2008:SRC

[Ran08c] Mohammad Ranjbar. Solving the resource-constrained project scheduling problem using filter-and-fan approach. Applied Mathematics and Computation, 201(1–2):313–318, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307012076.

Rao:2005:NMO

[Rao05] Prasada Rao. Numerical modeling of open channel flows using a multiple grid ENO scheme. Applied Mathematics and Computation, 161(2):599– 610, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304000141.

Rao:2006:SMM

[Rao06a] P. S. Rama Chandra Rao. Special multistep methods based on numerical differentiation for solving the initial value problem. Applied Mathematics and Computation, 181(1):500–510, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001196.

Rao:2006:NMO

[Rao06b] Prasada Rao. Numerical modeling of open channel flows with moving fronts using a variable boundary formulation. Applied Mathematics and Computation, 182(1):369–382, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003444.

Rasulov:2005:FDM

[RAP05] Mahir Rasulov, Zafer Aslan, and Ozkan Pakdil. Finite differences method for shallow water equations in a class of discontinuous functions. Applied Mathematics and Computation, 160(2):343–353, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030301141X.

Raslan:2005:CMU

[Ras05a]  K. R. Raslan.  Col-

location method using quartic B-spline for the equal width (EW) equation. Applied Mathematics and Computation, 168(2):795–805, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400640X.

Raslan:2005:CMR

[Ras05b] K. R. Raslan. A computational method for the regularized long wave (RLW) equation. Applied Mathematics and Computation, 167(2):1101–1118, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005521.

Rasulov:2005:NSO

[Ras05c] Mahir Rasulov. Numerical solution of one dimensional filtration of three phase compressible fluid through porous medium in a class of discontinuous functions. Applied Mathematics and Computation, 167(2):1249–1266, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400565X.

Rahimi-Ardabili:2007:INS

[RAS07] M. Y. Rahimi-Ardabili and S. Shahmorad. Iterative numerical solution of nonlinear integro-differential equations by the Tau method. Applied Mathematics and Computation, 193(2):514–522, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004353.

Roghanian:2008:IGP

[RAS08] E. Roghanian, M. B. Aryanezhad, and S. J. Sadjadi. Integrating goal programming, Kuhn–Tucker conditions, and penalty function approaches to solve linear bi-level programming problems. Applied Mathematics and Computation, 195(2):585– 590, February 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307005875.

Rawashdeh:2006:NSF

[Raw06a] E. A. Rawashdeh. Numerical solution of fractional integro-differential equations by collocation method. Applied Mathematics and Computation, 176(1):1–6, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305007551.

Rawashdeh:2006:NSS

[Raw06b] E. A. Rawashdeh. Numerical solution of semidifferential equations by collocation method. Applied Mathematics and Computation, 174(2):869–876, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005151.

Ray:2006:NSC

[Ray06] S. Saha Ray. A numerical solution of the coupled sine-Gordon equation using the modified decomposition method. Applied Mathematics and Computation, 175(2):1046–1054, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006612.

Ray:2008:NAA

[Ray08] Santanu Saha Ray. A new approach for the application of Adomian decomposition method for the solution of fractional space diffusion equation with insulated ends. Applied Mathematics and Computation, 202(2):544–549, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001392.

Rafiq:2007:MEV

[RAZ07] Arif Rafiq, Muhammad Awais, and Fiza Zafar. Modified efficient variant of super-Halley method. Applied Mathematics and Computation, 189(2):2004–2010, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000197.

Ray:2005:ASB

[RB05a] S. Saha Ray and R. K. Bera. Analytical solution of the Bagley Torvik equation by Adomian decomposition method. Applied Mathematics and Computation, 168(1):398–410, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006125.

Ray:2005:ASN

[RB05b] S. Saha Ray and R. K. Bera. An approximate solution of a nonlinear fractional differential equation by Adomian decomposition method. Applied Mathematics and Computation, 167(1):561–571, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005132.

Ray:2006:ASF

[RB06] S. Saha Ray and R. K.

Bera. Analytical solution of a fractional diffusion equation by Adomian decomposition method. Applied Mathematics and Computation, 174(1):329–336, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004595.

Raina:2007:SPC

[RB07] R. K. Raina and D. K.

Bansal. Some properties of a certain class of rational functions. Applied Mathematics and Computation, 187(1):403–407, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011908.

Rakkiyappan:2008:DDA

[RB08a] R. Rakkiyappan and P. Balasubramaniam. Delaydependent asymptotic stability for stochastic delayed recurrent neural networks with time varying delays. Applied Mathematics and Computation, 198(2):526– 533, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009186.

Rakkiyappan:2008:LCG

[RB08b] R. Rakkiyappan and P. Balasubramaniam. LMI conditions for global asymptotic stability results for neutral-type neural networks with distributed time delays. Applied Mathematics and Computation, 204(1):317–324, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005109.

Rathinasamy:2008:MSS

[RB08c] A. Rathinasamy and K. Balachandran. Mean square stability of semi-implicit Euler method for linear stochastic differential equations with multiple delays and Markovian switching. Applied Mathematics and Computation, 206(2):968–979, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308007698.

Rafikov:2008:MMC

[RBvB08] M. Rafikov, J. M. Balthazar, and H. F. von Bremen. Mathematical modeling and control of population systems: Applications in biological pest control. Applied Mathematics and Computation, 200(2):557– 573, July 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011356.

Ren:2006:NSCa

[RC06] Fengli Ren and Jinde Cao.

Novel α-stability criterion of linear systems with multiple time delays. Applied Mathematics and Computation, 181(1):282–290, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600097X.

Ray:2006:AAS

[RCB06] S. Saha Ray, K. S. Chaudhuri, and R. K. Bera. Analytical approximate solution of nonlinear dynamic system containing fractional derivative by modified decomposition method. Applied Mathematics and Computation, 182(1):544–552, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003626.

Ray:2008:AMD

[RCB08] S. Saha Ray, K. S. Chaudhuri, and R. K. Bera. Application of modified decomposition method for the analytical solution of space fractional diffusion equation. Applied Mathematics and Computation, 196(1):294–302, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006716.

Rafei:2007:VIM

[RDG07] M. Rafei, H. Daniali, and D. D. Ganji. Variational iteration method for solving the epidemic model and the prey and predator problem. Applied Mathematics and Computation, 186(2):1701–1709, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011052.

Rafei:2007:SPP

[RDGP07] M. Rafei, H. Daniali, D. D. Ganji, and H. Pashaei. Solution of the prey and predator problem by homotopy perturbation method. Applied Mathematics and Computation, 188(2):1419–1425, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014937.

Rebholz:2006:MDF

[Reb06] Leo G. Rebholz. A multiscale V –P discretization for flow problems. Applied Mathematics and Computation, 177(1):24– 35, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008842.

Ramadan:2005:IPD

[REDES05] Mohamed A. Ramadan, Talaat S. El-Danaf, and Naglaa M. El-Shazly. Iterative positive definite solutions of the two nonlinear matrix equations X ± A^{T} X^{−2}A = I. Applied Mathematics and Computation, 164(1):189–200, May 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003054.

Reese:2007:QSR

[Ree07] Andrea Reese. A Quasishrinking rectangle algorithm for complex zeros of a function. Applied Mathematics and Computation, 185(1):96–114, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008149.

Remesikova:2007:NST

[Rem07]  M. Remešíková.  Nu-

merical solution of twodimensional convection– diffusion–adsorption problems using an operator splitting scheme. Applied Mathematics and Computation, 184(1):116–130, January 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004905.

Rigozo:2005:CSB

[REN^{+}05] Nivaor Rodolfo Rigozo, Ezequiel Echer, Daniel Jean Roger Nordemann, Luís Eduardo Antunes Vieira, and Heloisa Helena de Faria. Comparative study between four classical spectral analysis methods. Applied Mathematics and Computation, 168(1):411–430, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006150.

Ren:2006:NSCc

[Ren06] Hongmin Ren. New sufficient convergence conditions of the Secant method for nondifferentiable operators. Applied Mathematics and Computation, 182(2):1255–1259, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004401.

Ren:2007:ICT

[Ren07] Hongmin Ren. An improved convergence theorem for a class of Secantlike methods. Applied Mathematics and Computation, 189(1):472–481, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016201.

Ren:2008:NPD

[Ren08a] Hongmin Ren. A note on a paper by D. K. R. Babajee and M. Z. Dauhoo. Applied Mathematics and Computation, 201(1–2):830–833, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011952.

See [BD06a].

Ren:2008:SDF

[Ren08b] Hongmin Ren. A secondderivative-free modified Secant-like method with order 2.732 … for unconstrained optimization. Applied Mathematics and Computation, 202(2):688– 692, August 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308001574.

Ramadan:2006:MEI

[RES06] Mohamed A. Ramadan and Naglaa M. El-Shazly. On the matrix equation X + A^{2} + ^{2^{n}} \sqrt{X^{−1}A} = I. Applied Mathematics and Computation, 173 (2):992–1013, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500408X.

Ramadan:2007:MEA

[RES07] Mohamed A. Ramadan and Ehab A. El-Sayed. On the matrix equation XH = HX and the associated controllability problem. Applied Mathematics and Computation, 186(1):844–859, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010071.

Reyes:2005:LNG

[Rey05a] Pedro M. Reyes. Logistics networks: a game theory application for solving the transshipment problem. Applied Mathematics and Computation, 168(2):1419–1431, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007404.

Reyes:2005:MET

[Rey05b] Pedro M. Reyes. A mathematical example of the twoechelon inventory model with asymmetric market information. Applied Mathematics and Computation, 162(1):257–264, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000864.

Roman-Flores:2007:FIM

[RFFFCC07] H. Román-Flores, A. Flores-Franulic, and Y. Chalco-Cano. The fuzzy integral for monotone functions. Applied Mathematics and Computation, 185(1):492–498, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008848.

Roman-Flores:2008:CTI

[RFFFCC08a] H. Román-Flores, A. Flores-Franulič, and Y. Chalco-Cano. A convolution type inequality for fuzzy integrals. Applied Mathematics and Computation, 195(1):94–99, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700522X.

Roman-Flores:2008:HTI

[RFFFCC08b] H. Román-Flores, A. Flores-Franulič, and Y. Chalco-Cano. A Hardy-type inequality for fuzzy integrals. Applied Mathematics and Computation, 204(1):178–183, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800492X.

Rashidinia:2005:CNS

[RG05] J. Rashidinia and A. Golbabaee. Convergence of numerical solution of a fourth-order boundary value problem. Applied Mathematics and Computation, 171(2):1296–1305, December 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rueda:2005:IMI

[RGA05] M. M. Rueda, S. González, and A. Arcos. Indirect methods of imputation of missing data based on available units. Applied Mathematics and Computation, 164(1):249–261, May 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003091.

Rueda:2006:GCE

[RGA06] M. M. Rueda, S. González, and A. Arcos. A general class of estimators with auxiliary information based on available units. Applied Mathematics and Computation, 175(1):131– 148, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005977.

Reyes:2007:NMV

[RGAFVA07] J. A. Reyes, F. García-Alonso, J. M. Ferrándiz, and J. Vigo-Aguiar. Numeric multistep variable methods for perturbed linear system integration.

Applied Mathematics and Computation, 190(1):63– 79, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000318.

Rafei:2007:SEM

[RGD07] M. Rafei, D. D. Ganji, and H. Daniali. Solution of the epidemic model by homotopy perturbation method. Applied Mathematics and Computation, 187(2):1056–1062, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012422.

Rashidinia:2007:SASa

[RGM07] J. Rashidinia, M. Ghasemi, and Z. Mahmoodi. Spline approach to the solution of a singularly-perturbed boundary-value problems. Applied Mathematics and Computation, 189(1):72– 78, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015815.

Ran:2006:IBT

[RH06] Rui-Sheng Ran and Ting-Zhu Huang. The inverses of block tridiagonal matrices. Applied Mathematics and Computation, 179(1):243–247, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010337.

Rostami:2008:MLA

[RH08] Habib Rostami and Jafar Habibi. Minimum linear arrangement of Chord graphs. Applied Mathematics and Computation, 203(1):358–367, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002592.

Rafiq:2007:NIM

[RHAA07] Arif Rafiq, Sifat Hussain, Farooq Ahmad, and Muhammad Awais. New iterative methods. Applied Mathematics and Computation, 189(2):1260–1267, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017115.

Ronghua:2006:ESN

[RHQ06] Li Ronghua, Meng Hongbing, and Chang Qin. Exponential stability of numerical solutions to SDDEs with Markovian switching. Applied Mathematics and Computation, 174(2):1302–1313, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005503.

Ronghua:2006:CNS

[RHY06] Li Ronghua, Meng Hongbing, and Dai Yonghong. Convergence of numerical solutions to stochastic delay differential equations with jumps. Applied Mathematics and Computation, 172(1):584–602, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002250.

Rigatos:2008:CMA

[Rig08] Gerasimos G. Rigatos. Coordinated motion of autonomous vehicles with the use of a distributed gradient algorithm. Applied Mathematics and Computation, 199(2):494–503, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010156.

Rimas:2005:CAPb

[Rim05a] Jonas Rimas. On computing of arbitrary positive integer powers for one type of even order tridiagonal matrices with zero row–I. Applied Mathematics and Computation, 164(1):149– 154, May 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003133.

Rimas:2005:CAPd

[Rim05b] Jonas Rimas. On computing of arbitrary positive integer powers for one type of odd order symmetric circulant matrices– I. Applied Mathematics and Computation, 165(1):137– 141, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003674.

Rimas:2005:CAPf

[Rim05c] Jonas Rimas. On computing of arbitrary positive integer powers for one type of odd order symmetric circulant matrices. II. Applied Mathematics and Computation, 169(2):1016–1027, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008318.

Rimas:2005:CAPg

[Rim05d] Jonas Rimas. On computing of arbitrary positive integer powers for one type of odd order tridiagonal matrices with zero row. I. Applied Mathematics and Computation, 169(2):1390–1394, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008598.

Rimas:2005:CAPe

[Rim05e] Jonas Rimas. On computing of arbitrary positive integer powers for one type of symmetric tridiagonal matrices of even order. Applied Mathematics and Computation, 168(2):783–787, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304006381.

Rimas:2005:CAPh

[Rim05f] Jonas Rimas. On computing of arbitrary positive integer powers for one type of symmetric tridiagonal matrices of odd order. I. Applied Mathematics and Computation, 171(2):1214– 1217, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rimas:2005:CAPa

[Rim05g] Jonas Rimas. On computing of arbitrary positive integer powers for one type of tridiagonal matrices. Applied Mathematics and Computation, 161(3):1037–1040, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000517.

Rimas:2005:CAPc

[Rim05h] Jonas Rimas. On computing of arbitrary positive integer powers for one type of tridiagonal matrices of even order. Applied Mathematics and Computation, 164(3):829–835, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003492.

Rimas:2006:CAPf

[Rim06a] Jonas Rimas. On computing of arbitrary positive integer powers for one type of even order skewsymmetric tridiagonal matrices with eigenvalues on imaginary axis-. I. Applied Mathematics and Computation, 174(2):997–1000, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005266.

Rimas:2006:CAPg

[Rim06b] Jonas Rimas. On computing of arbitrary positive integer powers for one type of even order skewsymmetric tridiagonal matrices with eigenvalues on imaginary axis. II. Applied Mathematics and Computation, 181(2):1120–1125, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001962.

Rimas:2006:CAPa

[Rim06c] Jonas Rimas. On computing of arbitrary positive integer powers for one type of even order symmetric circulant matrices. I. Applied Mathematics and Computation, 172(1):86–90, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305001888.

Rimas:2006:CAPd

[Rim06d] Jonas Rimas. On computing of arbitrary positive integer powers for one type of even order symmetric circulant matrices. II. Applied Mathematics and Computation, 174(1):511–523, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004820.

Rimas:2006:CAPh

[Rim06e] Jonas Rimas. On computing of arbitrary positive integer powers for one type of odd order skew-symmetric tridiagonal matrices with eigenvalues on imaginary axis. I. Applied Mathematics and Computation, 183(2):1378–1380, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006849.

Rimas:2006:CAPc

[Rim06f] Jonas Rimas. On computing of arbitrary positive integer powers for one type of odd order tridiagonal matrices with zero row. II. Applied Mathematics and Computation, 174(1):490–499, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305004807.

Rimas:2006:CAPb

[Rim06g] Jonas Rimas. On computing of arbitrary positive integer powers for one type of symmetric tridiagonal matrices of even order. II. Applied Mathematics and Computation, 172(1):245–251, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305002018.

Rimas:2006:CAPe

[Rim06h] Jonas Rimas. On computing of arbitrary positive integer powers for one type of symmetric tridiagonal matrices of odd order. II. Applied Mathematics and Computation, 174(1):676–683, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305004947.

Rimas:2007:CAPf

[Rim07a] Jonas Rimas. On computing of arbitrary positive integer powers for one type of even order tridiagonal matrices with eigenvalues on imaginary axis. I. Applied Mathematics and Computation, 189(2):1916–1920, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000100.

Rimas:2007:CAPg

[Rim07b] Jonas Rimas. On computing of arbitrary positive integer powers for one type of even order tridiagonal matrices with eigenvalues on imaginary axis. II. Applied Mathematics and Computation, 190(2):1466–1471, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001798.

Rimas:2007:CAPa

[Rim07c] Jonas Rimas. On computing of arbitrary positive integer powers for one type of odd order skewsymmetric tridiagonal matrices with eigenvalues on imaginary axis-II. Applied Mathematics and Computation, 186(1):872–878, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010095.

Rimas:2007:CAPd

[Rim07d] Jonas Rimas. On computing of arbitrary positive integer powers for tridiagonal matrices with elements −1, 0, 0, …, 0, 1 in principal and 1, 1, 1, …, 1 in neighbouring diagonals. I. Applied Mathematics and Computation, 188(1):634– 637, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014007.

Rimas:2007:CAPh

[Rim07e] Jonas Rimas. On computing of arbitrary positive integer powers for tridiagonal matrices with elements −1, 0, 0, …, 0 in principal and 1, 1, 1, …, 1 in neighbouring diagonals. I. Applied Mathematics and Computation, 193(1):253–256, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003943.

Rimas:2007:CAPe

[Rim07f] Jonas Rimas. On computing of arbitrary positive integer powers for tridiagonal matrices with elements −1, 0, 0, …, 0, 1 in principal and 1, 1, 1, …, 1 in neighbouring diagonals. II. Applied Mathematics and Computation, 188(2):2020–2024, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015700.

Rimas:2007:CAPb

[Rim07g] Jonas Rimas. On computing of arbitrary positive integer powers for tridiagonal matrices with elements 1, 0, 0, …, 0, 1 in principal and 1, 1, 1, …, 1 in neighbouring diagonals. I. Applied Mathematics and Computation, 186(2):1254–1257, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010502.

Rimas:2007:CAPc

[Rim07h] Jonas Rimas. On computing of arbitrary positive integer powers for tridiagonal matrices with elements 1, 0, 0, …, 0, 1 in principal and 1, 1, 1, …, 1 in neighbouring diagonals. II. Applied Mathematics and Computation, 187(2):1472–1475, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012859.

Rimas:2008:CAIa

[Rim08a] Jonas Rimas. On computing of arbitrary integer powers for one type of symmetric antitridiagonal matrices of even order. Applied Mathematics and Computation, 204(1):288–298, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005067.

Rimas:2008:CAIb

[Rim08b] Jonas Rimas. On computing of arbitrary integer powers of even order anti-tridiagonal matrices with zeros in main skew diagonal and elements 1, 1, 1, …, 1; −1, −1, −1, …, −1 in neighbouring diagonals. Applied Mathematics and Computation, 204(2):754–763, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800550X.

Rimas:2008:CAPa

[Rim08c] Jonas Rimas. On computing of arbitrary positive integer powers for one type of symmetric antitridiagonal matrices of odd order. Applied Mathematics and Computation, 203(2):573–581, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002853.

Rimas:2008:CAPb

[Rim08d] Jonas Rimas. On computing of arbitrary positive integer powers for one type of symmetric pentadiagonal matrices of even order. Applied Mathematics and Computation, 203(2):582–591, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002877.

Rimas:2008:CAPc

[Rim08e] Jonas Rimas. On computing of arbitrary positive integer powers for one type of symmetric pentadiagonal matrices of odd order. Applied Mathematics and Computation, 204(1):120–129, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004669.

Rimas:2009:CAPb

[Rim09a] Jonas Rimas. On computing of arbitrary positive integer powers for one type of even order symmetric antipentadiagonal matrices. Applied Mathematics and Computation, 211(1):54– 74, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900037X.

Rimas:2009:CAPa

[Rim09b] Jonas Rimas. On computing of arbitrary positive integer powers of odd order anti-tridiagonal matrices with zeros in main skew diagonal and elements 1, 1, 1, …, 1; −1, −1, −1, …, −1 in neighbouring diagonals. Applied Mathematics and Computation, 210(1):64– 71, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008394.

Ryu:2005:STJ

[RIY05] Eun-Kyung Ryu, Jae-Yuel Im, and Kee-Young Yoo. Security of Tseng–Jan’s conference key distribution system. Applied Mathematics and Computation, 167(2):833–839, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005326.

Rashidinia:2007:AGS

[RJ07] J. Rashidinia and E. Jafarzadeh. Accelerated generalized successive overrelaxation method for least squares problems. Applied Mathematics and Computation, 186(1):175–183, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009271.

Rashidinia:2007:SASb

[RJF07] J. Rashidinia, R. Jalilian, and K. Farajeyan. Spline approximate solution of fifth-order boundary-value problem. Applied Mathematics and Computation, 192(1):107–112, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002998.

Rashidinia:2007:SMS

[RJK07] J. Rashidinia, R. Jalilian, and V. Kazemi. Spline methods for the solutions of hyperbolic equations. Applied Mathematics and Computation, 190(1):882– 886, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001191.

Rashidinia:2007:CPC

[RJM07a] J. Rashidinia, R. Jalilian, and R. Mohammadi. Comment on the paper “A class of methods based on non-polynomial spline functions for the solution of a special fourth-order boundary-value problems with engineering applications”. Applied Mathematics and Computation, 186(2):1572–1580, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010848.

See [uITA06].

Rashidinia:2007:NPS

[RJM07b] J. Rashidinia, R. Jalilian, and R. Mohammadi. Non-polynomial spline methods for the solution of a system of obstacle problems. Applied Mathematics and Computation, 188(2):1984–1990, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015670.

Rashidinia:2007:SSM

[RJMG07] J. Rashidinia, R. Jalilian, R. Mohammadi, and M. Ghasemi. Sextic spline method for the solution of a system of obstacle problems. Applied Mathematics and Computation, 190(2):1669–1674, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002111.

Rahman:2007:SAP

[RJW07] M. Rahman, Z. Jackiewicz, and B. D. Welfert. Stochastic approximations of perturbed Fredholm Volterra integro-differential equation arising in mathematical neurosciences. Applied Mathematics and Computation, 186(2):1173–1182, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010435.

Ranjbar:2007:SDT

[RK07a] Mohammad R. Ranjbar and Fereydoon Kianfar. Solving the discrete time/ resource trade-off problem in project scheduling with genetic algorithms. Applied Mathematics and Computation, 191(2):451–456, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002731.

Rao:2007:OBS

[RK07b] S. Chandra Sekhara Rao and Mukesh Kumar. Optimal B-spline collocation method for self-adjoint singularly perturbed boundary value problems. Applied Mathematics and Computation, 188(1):749– 761, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014135.

Ramesh:2008:UMU

[RK08] V. P. Ramesh and M. K. Kadalbajoo. Upwind and midpoint upwind difference methods for timedependent differential difference equations with layer behavior. Applied Mathematics and Computation, 202(2):453–471, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011502.

Rady:2009:NPW

[RKOK09] A. S. Abdel Rady, A. H. Khater, E. S. Osman, and Mohammed Khalfallah. New periodic wave and soliton solutions for system of coupled Korteweg– de Vries equations. Applied Mathematics and Computation, 207(2):406–414, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800828X.

Ranjbar:2008:SRA

[RKS08] Mohammad Ranjbar, Fereydoon Kianfar, and Shahram Shadrokh. Solving the resource availability cost problem in project scheduling by path relinking and genetic algorithm. Applied Mathematics and Computation, 196(2):879–888, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007412.

Rocha:2007:AMR

[RLdLSF07] Marcus P. C. Rocha, Lourenildo W. B. Leite, Mauro de L. Santos, and Valcir J.da C. Farias. Attenuation of multiple in reflection seismic data using Kalman–Bucy filter. Applied Mathematics and Computation, 189(1):805– 815, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016584.

Rababah:2006:SMF

[RLY06] Abedallah Rababah, Byung-Gook Lee, and Jaechil Yoo. A simple matrix form for degree reduction of Bézier curves using Chebyshev–Bernstein basis transformations. Applied Mathematics and Computation, 181(1):310–318, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306000993.

Ramadan:2007:PNS

[RLZ07] M. A. Ramadan, I. F. Lashien, and W. K. Zahra. Polynomial and nonpolynomial spline approaches to the numerical solution of second order boundary value problems. Applied Mathematics and Computation, 184(2):476–484, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007235.

Ramadan:2008:HOA

[RLZ08] M. A. Ramadan, I. F. Lashien, and W. K. Zahra. High order accuracy nonpolynomial spline solutions for 2μth order two point boundary value problems. Applied Mathematics and Computation, 204(2):920–927, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005791.

Rao:2005:MDA

[RM05] Prasada Rao and Miguel A.

Medina, Jr. A multiple domain algorithm for modeling one-dimensional transient contaminant transport flows. Applied Mathematics and Computation, 167(1):1–15, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004692.

Rao:2006:MDA

[RM06] Prasada Rao and Miguel A.

Medina, Jr. A multiple domain algorithm for modeling two dimensional contaminant transport flows. Applied Mathematics and Computation, 174(1):117–133, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004455.

Ren:2008:ASE

[RM08] Junbai Ren and Xuan Ma. Asymptotic smoothing effect of solutions of two-dimensional micropolar fluid flows. Applied Mathematics and Computation, 204(1):363–367, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005146.

Rabbani:2007:NCS

[RMA07a] M. Rabbani, K. Maleknejad, and N. Aghazadeh. Numerical computational solution of the Volterra integral equations system of the second kind by using an expansion method. Applied Mathematics and Computation, 187(2):1143–1146, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012513.

Richard:2007:CGM

[RMA07b] M. J. Richard, J. J. McPhee, and R. J. Anderson. Computerized generation of motion equations using variational graphtheoretic methods. Applied Mathematics and Computation, 192(1):135–156, September 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307003025.

Rabbani:2007:CPM

[RMAM07] M. Rabbani, K. Maleknejad, N. Aghazadeh, and R. Mollapourasl. Computational projection methods for solving Fredholm integral equation. Applied Mathematics and Computation, 191(1):140–143, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002378.

Rashidinia:2007:PSM

[RMG07a] J. Rashidinia, Z. Mahmoodi, and M. Ghasemi. Parametric spline method for a class of singular two-point boundary value problems. Applied Mathematics and Computation, 188(1):58–63, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013191.

Rashidinia:2007:CSS

[RMG07b] J. Rashidinia, R. Mohammadi, and M. Ghasemi. Cubic spline solution of singularly perturbed boundary value problems with significant first derivatives. Applied Mathematics and Computation, 190(2):1762–1766, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700224X.

Rashidinia:2007:NSN

[RMJ07] J. Rashidinia, R. Mohammadi, and R. Jalilian. The numerical solution of non-linear singular boundary value problems arising in physiology. Applied Mathematics and Computation, 185(1):360–367, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600868X.

Rashidinia:2007:CCS

[RMJG07] J. Rashidinia, R. Mohammadi, R. Jalilian, and M. Ghasemi. Convergence of cubic-spline approach to the solution of a system of boundary-value problems. Applied Mathematics and Computation, 192(2):319–331, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003323.

Rajkovic:2005:NKM

[RMS05] Predrag M. Rajković, Sladjana D. Marinković, and Miomir S. Stanković. On q-Newton–Kantorovich method for solving systems of equations. Applied Mathematics and Computation, 168(2):1432–1448, September 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304007416.

Ronghua:2009:CNS

[RMWk09] Li Ronghua, Liu Min, and Pang Wan-kai. Convergence of numerical solutions to stochastic pantograph equations with Markovian switching. Applied Mathematics and Computation, 215(1):414– 422, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309004846.

Rajaram:2009:ATC

[RN09] R. Rajaram and M. Najafi. Analytical treatment and convergence of the Adomian decomposition method for a system of coupled damped wave equations. Applied Mathematics and Computation, 212(1):72– 81, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001192.

Rolo-Naranjo:2009:PPP

[RNRRMRCH09] Alberto Rolo-Naranjo, Rocio Rebollido-Rios, Kenia Melchor-Rodriguez, and Edelsys Codorniu-Hernández. Pseudophase portrait applied to pattern recognition in flavonoid–protein interactions. Applied Mathematics and Computation, 215(1):156–167, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900424X.

Rathod:2007:NIS

[RNV07a] H. T. Rathod, K. V. Nagaraja, and B. Venkatesudu. Numerical integration of some functions over an arbitrary linear tetrahedra in Euclidean threedimensional space. Applied Mathematics and Computation, 191(2):397–409, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002676.

Rathod:2007:ATS

[RNV07b] H. T. Rathod, K. V. Nagaraja, and B. Venkatesudu. On the application of two symmetric Gauss Legendre quadrature rules for composite numerical integration over a triangular surface. Applied Mathematics and Computation, 190(1):21– 39, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000240.

Rathod:2007:SGL

[RNV07c] H. T. Rathod, K. V. Nagaraja, and B. Venkatesudu. Symmetric Gauss Legendre quadrature formulas for composite numerical integration over a triangular surface. Applied Mathematics and Computation, 188(1):865–876, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014263.

Rezaie:2007:MMO

[RO07] K. Rezaie and B. Ostadi.

A mathematical model for optimal and phased implementation of flexible manufacturing systems. Applied Mathematics and Computation, 184(2):729–736, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007491.

Rady:2009:MSS

[ROK09] A. S. Abdel Rady, E. S. Osman, and Mohammed Khalfallah. Multi soliton solution for the system of Coupled Korteweg– de Vries equations. Applied Mathematics and Computation, 210(1):177–181, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308009892.

Roman:2009:SFA

[Rom09a] Adam Roman. Synchronizing finite automata with short reset words. Applied Mathematics and Computation, 209(1):125–136, March 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004803.

Romeiras:2009:ETW

[Rom09b] Filipe J. Romeiras. Exact travelling wave solutions of the generalized Bretherton equation. Applied Mathematics and Computation, 215(5):1791–1805, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900678X.

Rost:2007:GAC

[Rös07] Gergely Röst. On the global attractivity controversy for a delay model of hematopoiesis. Applied Mathematics and Computation, 190(1):846–850, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001166.

Rehman:2006:FGG

[RR06] Tahir Rehman and Carlos Romero. Formulating generalised ‘goal games’ against nature: An illustration from decisionmaking under uncertainty in agriculture. Applied Mathematics and Computation, 175(1):486–496, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006235.

Rezaifar:2007:NAF

[RR07] Omid Rezaifar and Hossein Rezaee. A new approach for finding the determinant of matrices. Applied Mathematics and Computation, 188(2):1445–1454, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014974.

See note [Kop22] invalidating claims of linear convergence.

Ruan:2006:EAC

[RS06] Ning Ruan and Xiaoling Sun. An exact algorithm for cost minimization in series reliability systems with multiple component choices. Applied Mathematics and Computation, 181(1):732–741, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001536.

Rao:2008:SLS

[RS08a] S. Chandra Sekhara Rao and Sarita. A symmetric linear system solver. Applied Mathematics and Computation, 203(1):368– 379, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308002609.

Repsys:2008:MAS

[RS08b] S. Repsys and V. Skakauskas.

Modelling of the age– sex–structured population dynamics with maternal care. Applied Mathematics and Computation, 198(1):163–183, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008776.

Roghanian:2007:PBL

[RSA07] E. Roghanian, S. J. Sadjadi, and M. B. Aryanezhad. A probabilistic bi-level linear multi-objective programming problem to supply chain planning. Applied Mathematics and Computation, 188(1):786–800, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014172.

Romera:2007:DRS

[RSD07] Miguel Romera, Michael Small, and Marius-F. Danca. Deterministic and random synthesis of discrete chaos. Applied Mathematics and Computation, 192(1):283–297, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003074.

Rashed:2008:EOT

[RSDC08] G. I. Rashed, H. I. Shaheen, X. Z. Duan, and S. J. Cheng. Evolutionary optimization techniques for optimal location and parameter setting of TCSC under single line contingency. Applied Mathematics and Computation, 205(1):133–147, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800324X.

Ramachandran:2007:UCU

[RSSS07] C. Ramachandran, T. N. Shanmugam, H. M. Srivastava, and A. Swaminathan. A unified class of k-uniformly convex functions defined by the Dziok– Srivastava linear operator. Applied Mathematics and Computation, 190(2):1627–1636, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700207X.

Rossler:2008:MLS

[RSZ08] Andreas Rößler, Mohammed Seaïd, and Mostafa Zahri. Method of lines for stochastic boundary-value problems with additive noise. Applied Mathematics and Computation, 199(1):301– 314, May 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307009988.

Rus:2008:NMB

[RV08] Francisco Rus and Francisco R. Villatoro. Numerical methods based on modified equations for nonlinear evolution equations with compactons. Applied Mathematics and Computation, 204(1):416–422, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005201.

Rus:2009:REC

[RV09] Francisco Rus and Francisco R. Villatoro. A repository of equations with cosine/sine compactons. Applied Mathematics and Computation, 215(5):1838–1851, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006857.

Rathod:2007:ATG

[RVN07] H. T. Rathod, B. Venkatesudu, and K. V. Nagaraja.

On the application of two Gauss–Legendre quadrature rules for composite numerical integration over a tetrahedral region. Applied Mathematics and Computation, 189(1):131–162, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015888.

Rathod:2007:GLG

[RVNI07] H. T. Rathod, B. Venkatesudu, K. V. Nagaraja, and Md. Shafiqul Islam. Gauss Legendre–Gauss Jacobi quadrature rules over a tetrahedral region. Applied Mathematics and Computation, 190(1):186– 194, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000392.

Ren:2007:CBM

[RW07] Hongmin Ren and Qingbiao Wu. Convergence ball of a modified secant method with convergence order 1.839…. Applied Mathematics and Computation, 188(1):281–285, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013518.

Ren:2008:CMS

[RW08] Hongmin Ren and Qingbiao Wu. A class of modified Secant methods for unconstrained optimization. Applied Mathematics and Computation, 206(2):716–720, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006681.

Ren:2009:CBE

[RW09] Hongmin Ren and Qingbiao Wu. Convergence ball and error analysis of a family of iterative methods with cubic convergence. Applied Mathematics and Computation, 209(2):369–378, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009739.

Ren:2009:CTS

[RWB09] Hongmin Ren, Qingbiao Wu, and Weihong Bi. A class of two-step Steffensen type methods with fourthorder convergence. Applied Mathematics and Computation, 209(2):206–210, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308009363.

Ren:2009:EUS

[RX09a] Yong Ren and Ningmao Xia. Existence, uniqueness and stability of the solutions to neutral stochastic functional differential equations with infinite delay. Applied Mathematics and Computation, 210(1):72–79, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008436.

Ren:2009:NEU

[RX09b] Yong Ren and Ningmao Xia. A note on the existence and uniqueness of the solution to neutral stochastic functional differential equations with infinite delay. Applied Mathematics and Computation, 214(2):457–461, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309003312.

Rui:2009:INI

[RX09c] Shao-Ping Rui and Cheng-Xian Xu. Inexact noninterior continuation method for solving large-scale monotone SDCP. Applied Mathematics and Computation, 215(7):2521–2527, December 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309007887.

Ryu:2005:SEU

[RY05a] Eun-Kyung Ryu and Kee-Young Yoo. On the security of efficient user identification scheme. Applied Mathematics and Computation, 171(2):1201–1205, December 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ryu:2005:OCA

[RY05b] Sang-Uk Ryu and Atsushi Yagi. Optimal control for an adsorbateinduced phase transition model. Applied Mathematics and Computation, 171(1):420–432, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001062.

Ronghua:2006:CSN

[RY06] Li Ronghua and Hou Yingmin. Convergence and stability of numerical solutions to SDDEs with Markovian switching. Applied Mathematics and Computation, 175(2):1080–1091, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006636.

Raimundo:2007:MET

[RYE07] Silvia Martorano Raimundo, Hyun Mo Yang, and Alejandro B. Engel. Modelling the effects of temporary immune protection and vaccination against infectious diseases. Applied Mathematics and Computation, 189(2):1723–1736, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017644.

Ren:2006:NSCb

[RYW06] Hongmin Ren, Shijun Yang, and Qingbiao Wu. A new semilocal convergence theorem for the Secant method under Hölder continuous divided differences. Applied Mathematics and Computation, 182(1):41–48, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003110.

Ye:2007:SCD

[rYzS07] Chao rong Ye and Zhi zhong Sun. On the stability and convergence of a difference scheme for an one-dimensional parabolic inverse problem. Applied Mathematics and Computation, 188(1):214–225, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013403.

Rashidinia:2005:NSL

[RZ05] J. Rashidinia and M. Zarebnia. Numerical solution of linear integral equations by using Sinc–collocation method. Applied Mathematics and Computation, 168(2):806–822, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006411.

Rashidinia:2007:NAN

[RZ07a] J. Rashidinia and M. Zarebnia. New approach for numerical solution of Hammerstein integral equations. Applied Mathematics and Computation, 185(1):147–154, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008216.

Rashidinia:2007:NSI

[RZ07b] J. Rashidinia and M. Zarebnia. The numerical solution of integro-differential equation by means of the Sinc method. Applied Mathematics and Computation, 188(2):1124–1130, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030601455X.

Ronghua:2007:CNS

[RZ07c] Li Ronghua and Chang Zhaoguang. Convergence of numerical solution to stochastic delay differential equation with Poisson jump and Markovian switching. Applied Mathematics and Computation, 184(2):451–463, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600720X.

Rahal:2008:NEP

[RZ08] Mohamed Rahal and Abdelkader Ziadi. A new extension of Piyavskii’s method to Hölder functions of several variables. Applied Mathematics and Computation, 197(2):478– 488, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007953.

Ren:2007:NLB

[RZX07] Yong-Hong Ren, Li-Wei Zhang, and Xian-Tao Xiao. A nonlinear Lagrangian based on Fischer-Burmeister NCP function. Applied Mathematics and Computation, 188(2):1344–1363, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014883.

Saleh:2005:RDE

[SA05a] M. Saleh and M. Aloqeili. On the rational difference equation y_{n+1} = A + y_{n−k}/y_{n}. Applied Mathematics and Computation, 171(2):862–869, December 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Syam:2005:NSS

[SA05b] Muhammed I. Syam and Basem S. Attili. Numerical solution of singularly perturbed fifth order two point boundary value problem. Applied Mathematics and Computation, 170(2):1085–1094, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000494.

Saldanha:2006:SMM

[SA06a] Godfrey Saldanha and Sujatha D. Achar. Symmetric multistep methods with zero phase-lag for periodic initial value problems of second order differential equations. Applied Mathematics and Computation, 175(1):401–412, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006168.

Saleh:2006:DE

[SA06b] M. Saleh and M. Aloqeili. On the difference equation y_{n+1} = A + y_{n}/yn − k with A < 0. Applied Mathematics and Computation, 176(1):359– 363, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008106.

Saleh:2006:RDE

[SA06c] M. Saleh and M. Aloqeili. On the rational difference equation y_{n+1} = A + y_{n}/y_{n−k}. Applied Mathematics and Computation, 177(1):189–193, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009008.

Selim:2006:PSD

[SA06d] M. M. Selim and M. K.

Ahmed. Plane strain deformation of an initially stressed orthotropic elastic medium. Applied Mathematics and Computation, 175(1):221–237, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006028.

Selim:2006:PAS

[SA06e] M. M. Selim and M. K.

Ahmed. Propagation and attenuation of seismic body waves in dissipative medium under initial and couple stresses. Applied Mathematics and Computation, 182(2):1064–1074, November 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306004218.

Siddiqi:2006:SFO

[SA06f] Shahid S. Siddiqi and Ghazala Akram. Solution of fifth order boundary value problems using nonpolynomial spline technique. Applied Mathematics and Computation, 175(2):1574–1581, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305007150.

Siddiqi:2006:STO

[SA06g] Shahid S. Siddiqi and Ghazala Akram. Solutions of twelfth order boundary value problems using thirteen degree spline. Applied Mathematics and Computation, 182(2):1443–1453, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004723.

Syam:2006:WRM

[SA06h] Muhammed I. Syam and Basem S. Attili. Weighted residual method for obtaining positive solutions of two-point nonlinear boundary value problems. Applied Mathematics and Computation, 176(2):775–784, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008799.

Siddiqi:2007:NSSb

[SA07a] Shahid S. Siddiqi and Ghazala Akram. Numerical solution of a system of fourth order boundary value problems using cubic non-polynomial spline method. Applied Mathematics and Computation, 190(1):652–661, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001002.

Siddiqi:2007:SOB

[SA07b] Shahid S. Siddiqi and Ghazala Akram. Solution of 10th-order boundary value problems using nonpolynomial spline technique. Applied Mathematics and Computation, 190 (1):641–651, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307000987.

Siddiqi:2007:SSFb

[SA07c] Shahid S. Siddiqi and Ghazala Akram. Solution of the system of fourth order boundary value problems using cubic spline. Applied Mathematics and Computation, 187(2):1219–1227, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012586.

Siddiqi:2007:SSFa

[SA07d] Shahid S. Siddiqi and Ghazala Akram. Solution of the system of fourthorder boundary value problems using non-polynomial spline technique. Applied Mathematics and Computation, 185(1):128–135, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008174.

Siddiqi:2007:STO

[SA07e] Shahid S. Siddiqi and Ghazala Akram. Solutions of tenth-order boundary value problems using eleventh degree spline. Applied Mathematics and Computation, 185(1):115–127, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008162.

Saeed:2008:IMQ

[SA08a] Rostam K. Saeed and Kawa M. Aziz. An iterative method with quartic convergence for solving nonlinear equations. Applied Mathematics and Computation, 202(2):435–440, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001458.

Seliaman:2008:OID

[SA08b] M. E. Seliaman and Ab Rahman Ahmad. Optimizing inventory decisions in a multi-stage supply chain under stochastic demands. Applied Mathematics and Computation, 206(2):538–542, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003767.

Sen:2008:BDR

[SA08c] S. K. Sen and Ravi P. Agarwal. Best k-digit rational approximation of irrational numbers: Pre-computer versus computer era. Applied Mathematics and Computation, 199(2):770– 786, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030701065X.

Shamilov:2008:FMM

[SA08d] Aladdin Shamilov and Senay Asma. Finite mixtures of MaxEnt distributions. Applied Mathematics and Computation, 206(2):530–537, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003755.

Siddiqi:2008:SOB

[SA08e] Shahid S. Siddiqi and Ghazala Akram. Solutions of 12th order boundary value problems using non-polynomial spline technique. Applied Mathematics and Computation, 199(2):559–571, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010314.

Samli:2009:NRG

[SA09] Ruya Samli and Sabri Arik. New results for global stability of a class of neutraltype neural systems with time delays. Applied Mathematics and Computation, 210(2):564–570, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000332.

Saad:2005:IGP

[Saa05] Omar M. Saad. An iterative goal programming approach for solving fuzzy multiobjective integer linear programming problems. Applied Mathematics and Computation, 170(1):216–225, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009191.

Saleh:2006:DHO

[SAB06] M. Saleh and S. Abu-Baha. Dynamics of a higher order rational difference equation. Applied Mathematics and Computation, 181(1):84–102, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000701.

Shahpar:2008:DPF

[SAB08] Amir H. Shahpar, Hedayat Z. Aashtiani, and Abbas Babazadeh. Dynamic penalty function method for the side constrained traffic assignment problem. Applied Mathematics and Computation, 206(1):332–345, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800653X.

Sezer:2006:TPS

[SAD06] Mehmet Sezer and Ayşegül Akyüz-Daşcıoğlu. Taylor polynomial solutions of general linear differential– difference equations with variable coefficients. Applied Mathematics and Computation, 174(2):1526– 1538, March 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305005771.

Sadegheih:2007:SOD

[Sad07] A. Sadegheih. Sequence optimization and design of allocation using GA and SA. Applied Mathematics and Computation, 186(2):1723–1730, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011076.

Saddeek:2009:GIP

[Sad09] A. M. Saddeek. Generalized iterative process and associated regularization for J-Pseudomonotone mixed variational inequalities. Applied Mathematics and Computation, 213(1): 8–17, July 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309002069.

Saen:2005:DNM

[Sae05] Reza Farzipoor Saen. Developing a nondiscretionary model of slacks-based measure in data envelopment analysis. Applied Mathematics and Computation, 169(2):1440–1447, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008641.

Saen:2006:ART

[Sae06a] Reza Farzipoor Saen. An algorithm for ranking technology suppliers in the presence of nondiscretionary factors. Applied Mathematics and Computation, 181(2):1616–1623, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306002931.

Saen:2006:DMSa

[Sae06b] Reza Farzipoor Saen. A decision model for selecting slightly non-homogeneous technologies. Applied Mathematics and Computation, 177(1):149–158, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008969.

Saen:2006:DMSb

[Sae06c] Reza Farzipoor Saen. A decision model for selecting technology suppliers in the presence of nondiscretionary factors. Applied Mathematics and Computation, 181(2):1609–1615, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600292X.

Saen:2006:DMT

[Sae06d] Reza Farzipoor Saen. A decision model for technology selection in the existence of both cardinal and ordinal data. Applied Mathematics and Computation, 181(2):1600–1608, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002918.

Saen:2006:TRP

[Sae06e] Reza Farzipoor Saen. Technologies ranking in the presence of both cardinal and ordinal data. Applied Mathematics and Computation, 176(2):476–487, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008258.

Saen:2007:NMA

[Sae07] Reza Farzipoor Saen. A new mathematical approach for suppliers selection: Accounting for non-homogeneity is important. Applied Mathematics and Computation, 185(1):84–95, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008137.

Siddiqi:2008:QSS

[SAE08] Shahid S. Siddiqi, Ghazala Akram, and Arfa Elahi. Quartic spline solution of linear fifth order boundary value problems. Applied Mathematics and Computation, 196(1):214–220, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006613.

Saitoh:2007:GPS

[Sai07] Hitoshi Saitoh. Geometric properties of solutions of a class of ordinary linear differential equations. Applied Mathematics and Computation, 187(1):408– 416, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030601191X.

Saker:2005:OCD

[Sak05] S. H. Saker. Oscillation of continuous and discrete diffusive delay Nicholson’s blowflies models. Applied Mathematics and Computation, 167(1):179–197, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004837.

Saka:2006:FEM

[Sak06] Bülent Saka. A finite element method for equal width equation. Applied Mathematics and Computation, 175(1):730–747, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006417.

Saremi:2008:WSI

[SAK08] Hamed Qahri Saremi, Babak Abedin, and Amirhosein Meimand Kermani. Website structure improvement: Quadratic assignment problem approach and ant colony metaheuristic technique. Applied Mathematics and Computation, 195(1):285– 298, January 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307005504.

Salman:2005:NNA

[Sal05] Raied Salman. Neural networks of adaptive inverse control systems. Applied Mathematics and Computation, 163(2):931–939, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002711.

Salama:2006:NMB

[Sal06a] A. A. Salama. Numerical methods based on extended one-step methods for solving optimal control problems. Applied Mathematics and Computation, 183(1):243–250, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600556X.

Salkuyeh:2006:CTA

[Sal06b] Davod Khojasteh Salkuyeh. CG-type algorithms to solve symmetric matrix equations. Applied Mathematics and Computation, 172(2):985–999, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305002730.

Salkuyeh:2006:CNT

[Sal06c] Davod Khojasteh Salkuyeh. Comments on “A note on a three-term recurrence for a tridiagonal matrix”. Applied Mathematics and Computation, 176(2):442–444, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008209.

See [EM03].

Salkuyeh:2006:NACa

[Sal06d] Davod Khojasteh Salkuyeh.

A new approach to compute sparse approximate inverse factors of a matrix. Applied Mathematics and Computation, 174(2):1110–1121, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005357.

Salkuyeh:2006:FDA

[Sal06e] Davod Khojasteh Salkuyeh.

On the finite difference approximation to the convection– diffusion equation. Applied Mathematics and Computation, 179(1):79–86, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010088.

Salman:2006:GET

[Sal06f]  M. Salman.  Gener-

alized elliptic-type integrals and their representations. Applied Mathematics and Computation, 181(2):1249–1256, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600244X.

Salahi:2007:FTM

[Sal07a] Maziar Salahi. A finite termination Mehrotra-type predictor–corrector algorithm. Applied Mathematics and Computation, 190(2):1740–1746, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002214.

Saldanha:2007:SCH

[Sal07b] Godfrey Saldanha. Single cell high order difference schemes for Poisson’s equation in three variables. Applied Mathematics and Computation, 186(1):548– 557, March 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306009660.

Saleh:2007:EMS

[Sal07c] Mohamed Saleh. Estimating market shares in each market segment using the information entropy concept. Applied Mathematics and Computation, 190(2):1735–1739, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002202.

Salkuyeh:2007:ENS

[Sal07d] Davod Khojasteh Salkuyeh.

Erratum to: “A numerical solution technique for a one-dimensional inverse nonlinear parabolic problem” [Appl. Math. Comput. 184 (2) (2007) 308– 315]. Applied Mathematics and Computation, 189(2):2023, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017681.

See [SFPE07].

Salas:2008:SST

[Sal08a] Alvaro Salas. Some solutions for a type of generalized Sawada–Kotera equation. Applied Mathematics and Computation, 196(2): 812–817, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307007345.

Salas:2008:ESG

[Sal08b]  Alvaro H. Salas.  Ex-

act solutions for the general fifth KdV equation by the exp function method. Applied Mathematics and Computation, 205(1):291–297, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005444.

Salem:2008:ECS

[Sal08c] Hussein A. H. Salem. On the existence of continuous solutions for a singular system of nonlinear fractional differential equations. Applied Mathematics and Computation, 198(1):445–452, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009137.

Salido:2008:NBC

[Sal08d] Miguel A. Salido. A nonbinary constraint ordering heuristic for constraint satisfaction problems. Applied Mathematics and Computation, 198(1):280–295, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008880.

Salahi:2009:SN

[Sal09] Maziar Salahi. A short note on min_{x∈R^{n}} \frac{||Ax−b||^{2}}{1+||x||^{2}} . Applied Mathematics and Computation, 212(1):270– 272, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000848.

Sampaio:2006:IPP

[Sam06] Jorge H. B. Sampaio, Jr. An iterative procedure for perpendicular offsets linear least squares fitting with extension to multiple linear regression. Applied Mathematics and Computation, 176(1):91– 98, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007745.

Siddiqi:2007:NSSa

[SAM07] Shahid S. Siddiqi, Ghazala Akram, and Salman A. Malik. Nonpolynomial sextic spline method for the solution along with convergence of linear special case fifth-order two-point boundary value problems. Applied Mathematics and Computation, 190(1):532– 541, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000860.

Siddiqi:2007:QSS

[SAN07] Shahid S. Siddiqi, Ghazala Akram, and Saima Nazeer. Quintic spline solution of linear sixth-order boundary value problems.

Applied Mathematics and Computation, 189(1):887– 892, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016675.

Sabzpoushan:2007:SLR

[SANN07] S. H. Sabzpoushan, A. Ayatollahi, D. Noble, and P. J. Noble. A survey on linear relations between main electrical components during action potential in ventricular cell. Applied Mathematics and Computation, 188(1):148–153, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013312.

Sunil:2005:EMF

[SAS05] Sunil, Anupama, and R. C. Sharma. The effect of magnetic field dependent viscosity on thermosolutal convection in ferromagnetic fluid. Applied Mathematics and Computation, 163(3):1197–1214, April 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002899.

Samawi:2007:AMI

[SAS07] Hani M. Samawi and Mohammad F. Al-Saleh. On the approximation of multiple integrals using multivariate ranked simulated sampling. Applied Mathematics and Computation, 188(1):345–352, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013580.

Savas:2007:SND

[Sav07] Ekrem Savaş. On some new double sequence spaces defined by a modulus. Applied Mathematics and Computation, 187(1):417– 424, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011921.

Sharaf:2005:GSD

[SB05a] Amr A. Sharaf and H. O.

Bakodah. A good spatial discretisation in the method of lines. Applied Mathematics and Computation, 171(2):1253–1263, December 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Soufyane:2005:SLN

[SB05b] A. Soufyane and M. Boulmalf. Solution of linear and nonlinear parabolic equations by the decomposition method. Applied Mathematics and Computation, 162(2):687–693, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001249.

Salido:2006:DCG

[SB06a] Miguel A. Salido and Federico Barber. Distributed CSPs by graph partitioning. Applied Mathematics and Computation, 183(1):491–498, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005832.

Sangwine:2006:QSV

[SB06b] Stephen J. Sangwine and Nicolas Le Bihan. Quaternion singular value decomposition based on bidiagonalization to a real or complex matrix using quaternion Householder transformations. Applied Mathematics and Computation, 182(1):727–738, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003808.

Saha:2008:DAD

[SB08a] Tapan Saha and Malay Bandyopadhyay. Dynamical analysis of a delayed ratio-dependent prey–predator model within fluctuating environment. Applied Mathematics and Computation, 196(1):458– 478, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030700700X.

Sun:2008:IGM

[SB08b] Yin Sun and Árpád Baricz. Inequalities for the generalized Marcum Qfunction. Applied Mathematics and Computation, 203(1):134–141, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002154.

Sassaman:2009:TNT

[SB09] Ryan Sassaman and Anjan Biswas. Topological and non-topological solitons of the generalized Klein–Gordon equations. Applied Mathematics and Computation, 215(1):212–220, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309004470.

Singh:2009:WMS

[SBA09] Bani Singh, Anuj Bhardwaj, and Rashid Ali. A wavelet method for solving singular integral equation of MHD. Applied Mathematics and Computation, 214(1):271–279, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309003014.

Saad:2008:NAR

[SBS08] A. S. Saad, M. Banaszkiewicz, and G. Sitarski. A new algorithm of the recurrent power series method for the nongravitational motion of comets. Applied Mathematics and Computation, 197 (2):874–879, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307008570.

Shao:2005:NEV

[SC05a] Jun Shao and Zhenfu Cao. A new efficient ( t n ) verifiable multi-secret sharing (VMSS) based on YCH scheme. Applied Mathematics and Computation, 168(1):135–140, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005922.

Shi:2005:RCA

[SC05b] Yuying Shi and Qianshun Chang. Remark on convergence of algebraic multigrid in the form of matrix decomposition. Applied Mathematics and Computation, 170(2):1170–1184, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000561.

Stephanides:2005:GAK

[SC05c] George Stephanides and Nicolae Constantinescu. The GN-authenticated key agreement. Applied Mathematics and Computation, 170(1):531–544, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400966X.

Sun:2005:IPK

[SC05d] Da-Zhi Sun and Zhen-Fu Cao. Improved public key authentication scheme for non-repudiation. Applied Mathematics and Computation, 168(2):927–932, September 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304006502.

Shi:2006:NTD

[SC06a] Yuying Shi and Qianshun Chang. New time dependent model for image restoration. Applied Mathematics and Computation, 179(1):121–134, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010143.

Song:2006:SCT

[SC06b] Yisheng Song and Rudong Chen. Strong convergence theorems on an iterative method for a family of finite nonexpansive mappings. Applied Mathematics and Computation, 180(1):275–287, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305011239.

Sun:2006:EAM

[SC06c] Jiping Sun and Lingfei Cheng. Equivalent analysis method of electromagnetic waves propagation in arched tunnel. Applied Mathematics and Computation, 178(2):332–337, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009781.

Sheng:2007:FIM

[SC07a] Xingping Sheng and Guoliang Chen. A finite iterative method for solving a pair of linear matrix equations (AXB, CXD) = (E, F). Applied Mathematics and Computation, 189(2):1350–1358, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017218.

Song:2007:VAM

[SC07b] Yisheng Song and Rudong Chen. Viscosity approximative methods to Cesàro means for non-expansive mappings. Applied Mathematics and Computation, 186(2):1120–1128, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010381.

Su:2007:MSF

[SC07c] Ke Su and Jianren Che. A modified SQP-filter method and its global convergence. Applied Mathematics and Computation, 194(1):92–101, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700447X.

Sun:2007:EPPa

[SC07d] Shulin Sun and Lansun Chen. Existence of positive periodic solution of an impulsive delay Logistic model. Applied Mathematics and Computation, 184(2):617–623, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007351.

Sun:2007:CLV

[SC07e] Tingkai Sun and Songcan Chen. Class label versus sample labelbased CCA. Applied Mathematics and Computation, 185(1):272–283, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008599.

Sundararajan:2007:NSA

[SC07f] Kiran Kumar Sundararajan and Soubhik Chakraborty. A new sorting algorithm. Applied Mathematics and Computation, 188(1):1037– 1041, May 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306014524.

Shen:2008:DFE

[SC08a] Yaliang Shen and Nanbin Cao. The double F - expansion approach and novel nonlinear wave solutions of soliton equation in (2 + 1)-dimension. Applied Mathematics and Computation, 198(2):683– 690, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009356.

Sheng:2008:SGI

[SC08b] Xingping Sheng and Guoliang Chen. Some generalized inverses of partition matrix and quotient identity of generalized Schur complement. Applied Mathematics and Computation, 196(1):174–184, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700656X.

Shi:2008:SSL

[SC08c] Ruiqing Shi and Lansun Chen. Staged-structured Lotka–Volterra predator– prey models for pest management. Applied Mathematics and Computation, 203(1):258–265, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002476.

Sheng:2009:OPI

[SC09] Xingping Sheng and Guoliang Chen. An oblique projection iterative method to compute Drazin inverse and group inverse. Applied Mathematics and Computation, 211(2):417–421, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000885.

Sunil:2007:DDC

[SCB07] Sunil, Prakash Chand, and Pavan Kumar Bharti.

Double-diffusive convection in a micropolar ferromagnetic fluid. Applied Mathematics and Computation, 189(2):1648–1661, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017498.

Schmidt:2005:MPL

[Sch05] Andreas U. Schmidt. A model for prejudiced learning in noisy environments. Applied Mathematics and Computation, 168(1):354–379, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006095.

Schnurr:2008:CSE

[Sch08] Marco Schnurr. Computing slope enclosures by exploiting a unique point of inflection. Applied Mathematics and Computation, 204(1):249–256, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005018.

Stanimirovic:2008:SMS

[SCI08] Predrag S. Stanimirović and Dragana S. Cvetković-Ilić. Successive matrix squaring algorithm for computing outer inverses. Applied Mathematics and Computation, 203(1):19– 29, September 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308002348.

Shen:2005:DCN

[SCL^{+}05a] Judong Shen, Shing I. Chang, E. Stanley Lee, Youping Deng, and Susan J. Brown. Determination of cluster number in clustering microarray data. Applied Mathematics and Computation, 169(2):1172–1185, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008434.

Su:2005:IBT

[SCL05b] Pin-Chang Su, Henry Ker-Chang Chang, and Erl-Huei Lu. ID-based threshold digital signature schemes on the elliptic curve discrete logarithm problem. Applied Mathematics and Computation, 164(3):757–772, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003431.

Chang:2009:CTC

[sCLC09] Shih sen Chang, H. W. Joseph Lee, and Chi Kin Chan. Convergence theorem of common fixed points for Lipschitzian pseudocontraction semigroups in Banach spaces. Applied Mathematics and Computation, 212(1):60–65, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001179.

Shen:2009:SSQ

[SCM09] Peiping Shen, Yongqiang Chen, and Yuan Ma. Solving sum of quadratic ratios fractional programs via monotonic function. Applied Mathematics and Computation, 212(1):234– 244, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001398.

Sun:2009:PSS

[SCO09] Yan Sun, Yeol Je Cho, and Donal O’Regan. Positive solutions for singular second order Neumann boundary value problems via a cone fixed point theorem. Applied Mathematics and Computation, 210(1):80–86, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008734.

Shi:2006:DMP

[SCQ06] Hongli Shi, Yuanli Cai, and Zulian Qiu. On design of multiwavelet prefilters. Applied Mathematics and Computation, 172(2):1175–1187, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305002845.

Scott:2008:PPS

[SCR08a] Luis P. B. Scott, Jorge Chahine, and José R. Ruggiero. Predicting peptides structure with solvation potential and rotamer library dependent of the backbone. Applied Mathematics and Computation, 199(1):155–161, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009824.

Scott:2008:UGAa

[SCR08b] Luis P. B. Scott, Jorge Chahine, and José R. Ruggiero. Use of genetic algorithms and solvation potential to study peptides structure. Applied Mathematics and Computation, 195(2):515–522, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005814.

Scott:2008:UGAb

[SCR08c] Luis P. B. Scott, Jorge Chahine, and José R. Ruggiero. Using genetic algorithm to design protein sequence. Applied Mathematics and Computation, 200(1):1–9, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009757.

Shen:2005:NCB

[SCS05a] Victor R. L. Shen, Tzer-Shyong Chen, and Kai-Quan Shai. A novel cryptosystem based on grey system theory and genetic algorithm. Applied Mathematics and Computation, 170(2):1290–1302, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000640.

Sun:2005:CCL

[SCS05b] Da-Zhi Sun, Zhen-Fu Cao, and Yu Sun. Comment: cryptanalysis of Lee–Hwang–Li’s key authentication scheme. Applied Mathematics and Computation, 164(3):675– 678, May 25, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304003078. See

[ZK05].

Sun:2005:RNK

[SCS05c] Da-Zhi Sun, Zhen-Fu Cao, and Yu Sun. Remarks on a new key authentication scheme based on discrete logarithms. Applied Mathematics and Computation, 167(1):572–575, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304005144.

Sun:2006:HCM

[SCS06] Da-Zhi Sun, Zhen-Fu Cao, and Yu Sun. How to compute modular exponentiation with large operators based on the rightto-left binary algorithm. Applied Mathematics and Computation, 176(1):280– 292, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007927.

Shi:2007:CFP

[SCX07] Yuying Shi, Qianshun Chang, and Jing Xu. Convergence of fixed point iteration for deblurring and denoising problem. Applied Mathematics and Computation, 189(2):1178–1185, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017012.

Subramani:2006:CVC

[SD06] K. Subramani and D. Desovski. On contrasting vertex contraction with relaxation-based approaches for negative cost cycle detection. Applied Mathematics and Computation, 173(1):273–305, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003437.

Saka:2009:QBS

[SD09] Bülent Saka and İdris DaḠ.

Quartic B-spline Galerkin approach to the numerical solution of the KdVB equation. Applied Mathematics and Computation, 215(2):746–758, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309005554.

Shidfar:2005:NSS

[SDG05] A. Shidfar, M. Djalalvand, and M. Garshasbi. A numerical scheme for solving special class of nonlinear diffusion–convection equation. Applied Mathematics and Computation, 167(2):1080–1089, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304005508.

Smink:2009:FAC

[SdHT09] R. W. Smink, B. P. de Hon, and A. G. Tijhuis. Fast and accurate computation of Bessel functions with (large) complex order and argument. Applied Mathematics and Computation, 207(2):442–447, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800831X.

Soleimani-damaneh:2006:CTP

[SdJA06] M. Soleimani-damaneh, G. R. Jahanshahloo, and S. Abbasbandy. Computational and theoretical pitfalls in some current performance measurement techniques; and a new approach. Applied Mathematics and Computation, 181(2):1199–1207, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002049.

Shi:2009:AEF

[SDL09] Yeqiong Shi, Zhengde Dai, and Donglong Li. Application of Exp-function method for 2D cubic– quintic Ginzburg–Landau equation. Applied Mathematics and Computation, 210(1):269–275, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309000046.

Shen:2008:RSA

[SDM08] Peiping Shen, Yunpeng Duan, and Yuan Ma. A robust solution approach for nonconvex quadratic programs with additional multiplicative constraints. Applied Mathematics and Computation, 201(1–2):514–526, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307012271.

Scalassara:2009:REM

[SDM^{+}09] Paulo Rogério Scalassara, María Eugenia Dajer, Carlos Dias Maciel, Rodrigo Capobianco Guido, and José Carlos Pereira. Relative entropy measures applied to healthy and pathological voice characterization. Applied Mathematics and Computation, 207(1):95–108, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308000234.

Shidfar:2007:SNA

[SDR07] A. Shidfar, J. Damirchi, and P. Reihani. An stable numerical algorithm for identifying the solution of an inverse problem. Applied Mathematics and Computation, 190(1):231– 236, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000434.

Sunil:2005:EDP

[SDS05] Sunil, Divya, and R. C. Sharma. Effect of dust particles on a ferromagnetic fluid heated and soluted from below saturating a porous medium. Applied Mathematics and Computation, 169(2):832–853, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400815X.

Sermutlu:2007:NQR

[SE07] E. Sermutlu and H. T.

Eyyuboğlu. A new quadrature routine for improper and oscillatory integrals. Applied Mathematics and Computation, 189(1):452– 461, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016183.

Seaid:2006:SNM

[Sea06] Mohammed Seaïd. Stable numerical methods for conservation laws with discontinuous flux function. Applied Mathematics and Computation, 175(1):383– 400, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006156.

Seleim:2005:EFT

[Sel05] Adel A. S. Abo Seleim.

Electromagnetic field of transient signals above a dielectric layer. Applied Mathematics and Computation, 160(1):261–267, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303008646.

Selim:2006:OEM

[Sel06] M. M. Selim. Orthotropic elastic medium under the effect of initial and couple stresses. Applied Mathematics and Computation, 181(1):185–192, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306000889.

Selim:2007:SDI

[Sel07]  M. M. Selim.  Static

deformation of an irregular initially stressed medium. Applied Mathematics and Computation, 188(2):1274–1284, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014834.

Sendroiu:2007:STC

[Şen07] Elena Şendroiu. Sheaf tools for computation. Applied Mathematics and Computation, 184(1):131–141, January 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004917.

Sermutlu:2005:CNC

[Ser05] Emre Sermutlu. Comparison of Newton–Cotes and Gaussian methods of quadrature. Applied Mathematics and Computation, 171(2):1048–1057, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Serrano:2009:AOA

[Ser09] Sergio Serrano. Asymptotic ordination of artificial satellite Hamiltonian. Applied Mathematics and Computation, 215(3):1124–1134, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006225.

Srivastava:2009:INP

[SEŞ09] H. M. Srivastava, Sevtap Sümer Eker, and Bilal Şeker. Inclusion and neighborhood properties for certain classes of multivalently analytic functions of complex order associated with the convolution structure.

Applied Mathematics and Computation, 212(1):66– 71, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001180.

Sever:2007:LBS

[Sev07] Ali Sever. A linearizationbased solution to an inverse problem in financial markets. Applied Mathematics and Computation, 188(2):1666–1670, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015207.

Seyidmamedov:2005:CME

[Sey05] Zahir Muradoglu Seyidmamedov. Computational modeling of elastoplastic behaviour of homogeneous materials under indentation. Applied Mathematics and Computation, 163(3):1225–1241, April 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002917.

Saadatmandi:2007:CFD

[SF07] Abbas Saadatmandi and Jalal Askari Farsangi. Chebyshev finite difference method for a nonlinear system of secondorder boundary value problems. Applied Mathematics and Computation, 192(2):586–591, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003220.

Si:2008:SDS

[SFA08] Zhiyong Si, Xinlong Feng, and Abdurishit Abduwali. The semi-discrete streamline diffusion finite element method for timedependented convection– diffusion problems. Applied Mathematics and Computation, 202(2):771–779, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001793.

Song:2008:RHC

[SFFY08] Yang Song, Jian Fan, Minrui Fei, and Taicheng Yang. Robust H_{∞} control of discrete switched system with time delay. Applied Mathematics and Computation, 205(1):159–169, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003263.

Srivastava:2007:DRC

[SFG07] H. M. Srivastava, Zoltán Finta, and Vijay Gupta. Direct results for a certain family of summation– integral type operators. Applied Mathematics and Computation, 190(1):449– 457, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000768.

Siyyam:2009:FOA

[SFH09] H. I. Siyyam, R. A. Ford, and M. H. Hamdan. Firstorder accurate finite difference schemes for boundary vorticity approximations in curvilinear geometries. Applied Mathematics and Computation, 215(6):2378–2387, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309007619.

Shidfar:2007:NST

[SFPE07] A. Shidfar, M. Fakhraie, R. Pourgholi, and M. Ebrahimi. A numerical solution technique for a one-dimensional inverse nonlinear parabolic problem. Applied Mathematics and Computation, 184(2):308–315, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007077.

See erratum [Sal07d].

Sezer:2005:NPA

[SG05a] Mehmet Sezer and Mustafa Gülsu. A new polynomial approach for solving difference and Fredholm integrodifference equations with mixed argument. Applied Mathematics and Computation, 171(1):332–344, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000998.

Shakhno:2005:IAO

[SG05b] S. M. Shakhno and O. P.

Gnatyshyn. On an iterative algorithm of order 1.839 … for solving the nonlinear least squares problems. Applied Mathematics and Computation, 161(1):253–264, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012475.

Shidfar:2005:NSC

[SG05c] A. Shidfar and M. Garshasbi. Numerical study of in-cylinder pressure in an internal combustion engine. Applied Mathematics and Computation, 165(1):163–170, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003698.

Sotiropoulos:2005:OCB

[SG05d] D. G. Sotiropoulos and T. N. Grapsa. Optimal centers in branch-and-prune algorithms for univariate global optimization. Applied Mathematics and Computation, 169(1):247–277, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007738.

Sezer:2006:ASC

[SG06] Mehmet Sezer and Mustafa Gülsu. Approximate solution of complex differential equations for a rectangular domain with Taylor collocation method. Applied Mathematics and Computation, 177(2):844–851, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009653.

Sezer:2007:PSM

[SG07a] Mehmet Sezer and Mustafa Gülsu. Polynomial solution of the most general linear Fredholm–Volterra integrodifferential-difference equations by means of Taylor collocation method. Applied Mathematics and Computation, 185(1):646– 657, February 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030600899X.

Sharma:2007:FMO

[SG07b] J. R. Sharma and R. K.

Guha. A family of modified Ostrowski methods with accelerated sixth order convergence. Applied Mathematics and Computation, 190(1):111–115, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000331.

Sun:2007:SIP

[SG07c] Bo Sun and WeiGao Ge. Successive iteration and positive pseudosymmetric solutions for a three-point second-order p-Laplacian boundary value problems. Applied Mathematics and Computation, 188(2):1772–1779, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030601544X.

Salas:2008:CES

[SG08a] Alvaro H. Salas and Cesar A. Gómez. Computing exact solutions for some fifth KdV equations with forcing term. Applied Mathematics and Computation, 204(1):257–260, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800502X.

Sharomi:2008:CSD

[SG08b] O. Sharomi and A. B. Gumel. Curtailing smoking dynamics: a mathematical modeling approach. Applied Mathematics and Computation, 195(2):475–499, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005784.

Shen:2008:DBA

[SG08c] Peiping Shen and Minna Gu. A duality-bounds algorithm for non-convex quadratic programs with additional multiplicative constraints. Applied Mathematics and Computation, 198(1):1–11, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008065.

Sari:2009:SOC

[SG09a] Murat Sari and Gürhan Gürarslan. A sixth-order compact finite difference scheme to the numerical solutions of Burgers’ equation. Applied Mathematics and Computation, 208(2):475–483, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009193.

Shen:2009:CDM

[SG09b] Jindong Shen and Chuanqing Gu. Computing the determinants of matrix Padé approximation. Applied Mathematics and Computation, 214(2):433–441, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003282.

Srinivas:2009:PTN

[SG09c] S. Srinivas and R. Gayathri. Peristaltic transport of a Newtonian fluid in a vertical asymmetric channel with heat transfer and porous medium. Applied Mathematics and Computation, 215(1):185– 196, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300309004275.

Sharma:2006:FGP

[SGA06] Dinesh K. Sharma, Debasis Ghosh, and Julius A. Alade. A fuzzy goal programming approach for regional rural development planning. Applied Mathematics and Computation, 176(1):141–149, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305007800.

Singh:2005:MSA

[SGB05] Sanjeet Singh, Pankaj Gupta, and Davinder Bhatia. Multiparametric sensitivity analysis of the constraint matrix in linear-plus-linear fractional programming problem. Applied Mathematics and Computation, 170(2):1243–1260, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000603.

Salas:2009:NSM

[SGC09] Alvaro H. Salas, Cesar A.

Gómez S., and Jairo E. Castillo H. New solutions for the modified generalized Degasperis–Procesi equation. Applied Mathematics and Computation, 215(7):2608–2615, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900798X.

Spadotto:2009:CNS

[SGG^{+}09] André Augusto Spadotto, Ana Rita Gatto, Rodrigo Capobianco Guido, Arlindo Neto Montagnoli, Paula Cristina Cola, José Carlos Pereira, and Arthur Oscar Schelp. Classification of normal swallowing and oropharyngeal dysphagia using wavelet. Applied Mathematics and Computation, 207(1):75– 82, January 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300308000209.

Shukla:2007:SMS

[SGP07] D. Shukla, Surendra Gadewar, and R. K. Pathak. A stochastic model for spacedivision switches in computer networks. Applied Mathematics and Computation, 184(2):235–269, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007004.

Sharma:2009:SVH

[SGS09] J. R. Sharma, R. K. Guha, and Rajni Sharma. Some variants of Hansen– Patrick method with third and fourth order convergence. Applied Mathematics and Computation, 214(1):171–177, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002914.

Sezer:2006:TCM

[SGT06] Mehmet Sezer, Mustafa Gülsu, and Bekir Tanay. A Taylor collocation method for the numerical solution of complex differential equations with mixed conditions in elliptic domains. Applied Mathematics and Computation, 182(1):498–508, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003572.

Shamilov:2008:NCR

[SGUK08] Aladdin Shamilov, Cigdem Giriftinoglu, Ilhan Usta, and Yeliz Mert Kantar. A new concept of relative suitability of moment function sets. Applied Mathematics and Computation, 206(2):521–529, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003743.

Shang:2005:EHO

[SH05a] Desheng Shang and Maoan Han. The existence of homoclinic orbits to saddle-focus. Applied Mathematics and Computation, 163(2):621–631, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002474.

Soliman:2005:CSR

[SH05b] A. A. Soliman and M. H.

Hussien. Collocation solution for RLW equation with septic spline. Applied Mathematics and Computation, 161(2):623–636, February 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304000165.

Shahraki:2006:CHO

[SH06a] M. Mokarram Shahraki and S. Mohammad Hosseini. Comparison of a higher order method and the simple upwind and non-monotone methods for singularly perturbed boundary value problems. Applied Mathematics and Computation, 182(1):460–473, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003535.

Shouzhi:2006:HOB

[SH06b] Yang Shouzhi and Wang Hongyong. High-order balanced multiwavelets with dilation factor a. Applied Mathematics and Computation, 181(1):362–369, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001056.

Syam:2006:EMS

[SH06c] Muhammed I. Syam and Abdelrahem Hamdan. An efficient method for solving Bratu equations. Applied Mathematics and Computation, 176(2):704–713, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008702.

Sajadifar:2007:OST

[SH07a] S. M. Sajadifar and R. Haji.

Optimal solution for a two-level inventory system with information exchange leading to a more computationally efficient search. Applied Mathematics and Computation, 189(2):1341–1349, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017206.

Sajid:2007:NSS

[SH07b] M. Sajid and T. Hayat.

Non-similar series solution for boundary layer flow of a third-order fluid over a stretching sheet. Applied Mathematics and Computation, 189(2):1576–1585, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017437.

Sankar:2007:NLM

[SH07c] D. S. Sankar and K. Hemalatha. Non-linear mathematical models for blood flow through tapered tubes. Applied Mathematics and Computation, 188(1):567– 582, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306013968.

Shang:2007:NSF

[SH07d] Xufeng Shang and Danfu Han. Numerical solution of Fredholm integral equations of the first kind by using linear Legendre multi-wavelets. Applied Mathematics and Computation, 191(2):440–444, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002718.

Shiraishi:2007:DSC

[SH07e] Fumihide Shiraishi and Yuji Hatoh. Dynamic sensitivities in chaotic dynamical systems. Applied Mathematics and Computation, 186(2):1347–1359, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010599.

Sugisaka:2007:DPC

[SH07f] M. Sugisaka and D. Hazry.

Development of a proportional control method for a mobile robot. Applied Mathematics and Computation, 186(1):74– 82, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009180.

Sajid:2008:WCA

[SH08a] M. Sajid and T. Hayat.

Wire coating analysis by withdrawal from a bath of Sisko fluid. Applied Mathematics and Computation, 199(1):13–22, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009666.

Shen:2008:NCR

[SH08b] Shu-Qian Shen and Ting-Zhu Huang. New comparison results for parallel multisplitting iterative methods. Applied Mathematics and Computation, 206(2):738–747, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006991.

Shan:2009:FDS

[SH09a] Li Shan and Yanren Hou. A fully discrete stabilized finite element method for the time-dependent Navier–Stokes equations. Applied Mathematics and Computation, 215(1):85– 99, September 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309003865.

Sun:2009:DEU

[SH09b] Shinn Sun and Shing-Cheng Hu. Discriminating efficient units using MAJ FDH. Applied Mathematics and Computation, 215(8):3116–3123, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008984.

Shahmorad:2005:NSG

[Sha05a] S. Shahmorad. Numerical solution of the general form linear Fredholm–Volterra integro-differential equations by the Tau method with an error estimation. Applied Mathematics and Computation, 167(2):1418–1429, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005776.

Shao:2005:IEP

[Sha05b] Zuhua Shao. Improvement of efficient proxy signature schemes using self-certified public keys. Applied Mathematics and Computation, 168(1):222–234, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006009.

Shao:2005:NKA

[Sha05c]  Zuhua Shao.  A new

key authentication scheme for cryptosystems based on discrete logarithms. Applied Mathematics and Computation, 167(1):143– 152, August 5, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304004801.

Shao:2005:SMD

[Sha05d]  Zuhua Shao.  Security

of meta-He digital signature scheme based on factoring and discrete logarithms. Applied Mathematics and Computation, 170(2):976–984, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030500038X.

Sharma:2005:CTO

[Sha05e] J. R. Sharma. A composite third order Newton– Steffensen method for solving nonlinear equations. Applied Mathematics and Computation, 169(1):242–246, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304007726.

Shang:2006:NES

[Sha06a]  Yadong Shang.  New

exact special solutions with solitary patterns for Boussinesq-like B ( m n ) equations with fully nonlinear dispersion. Applied Mathematics and Computation, 173(2):1137–1148, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305004194.

Shang:2006:NSW

[Sha06b] Yadong Shang. New solitary wave solutions with compact support for the KdV-like K ( m n ) equations with fully nonlinear dispersion. Applied Mathematics and Computation, 173(2):1124–1136, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004182.

Sharaf:2006:CCD

[Sha06c] M. A. Sharaf. Computations of the cosmic distance equation. Applied Mathematics and Computation, 174(2):1269–1278, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005473.

Shang:2007:BTL

[Sha07a] Yadong Shang. Bäcklund transformation, Lax pairs and explicit exact solutions for the shallow water waves equation. Applied Mathematics and Computation, 187(2):1286–1297, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012677.

Shanmugam:2007:HFG

[Sha07b] T. N. Shanmugam. Hypergeometric functions in the geometric function theory. Applied Mathematics and Computation, 187(1):433– 444, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011945.

Sharma:2007:FTO

[Sha07c] J. R. Sharma. A family of third-order methods to solve nonlinear equations by quadratic curves approximation. Applied Mathematics and Computation, 184(2):210–215, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006965.

Sharma:2007:OPF

[Sha07d] J. R. Sharma. A oneparameter family of secondorder iteration methods. Applied Mathematics and Computation, 186(2):1402– 1406, March 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306010666.

Shahraki:2008:CEA

[Sha08a] M. S. Shiri Shahraki. A computationally efficient algorithm for determination of efficient max-ordering optimal solutions in multiple objective programming. Applied Mathematics and Computation, 196(2):720–723, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007242.

Shampine:2008:DAN

[Sha08b] L. F. Shampine. Dissipative approximations to neutral DDEs. Applied Mathematics and Computation, 203(2):641–648, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002956.

Shampine:2008:M

[Sha08c] L. F. Shampine. Matlab. Applied Mathematics and Computation, 202(1):266–274, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000982.

Shang:2008:DNI

[Sha08d] Li Shang. Denoising natural images based on a modified sparse coding algorithm. Applied Mathematics and Computation, 205(2):883–889, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003676.

Shampine:2009:SLM

[Sha09] L. F. Shampine. Stability of the leapfrog/ midpoint method. Applied Mathematics and Computation, 208(1):293–298, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008758.

Sun:2007:EPPb

[SHD07] Yan Sun, Maoan Han, and Lokenath Debnath. Existence of positive periodic solutions for a class of functional differential equations. Applied Mathematics and Computation, 190(1):699–704, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001063.

Shen:2005:LMG

[She05a] Peiping Shen. Linearization method of global optimization for generalized geometric programming. Applied Mathematics and Computation, 162(1):353– 370, March 4, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304001006.

Shen:2005:IAN

[She05b] Zhizhang Shen. The impact of the apex node on routing inside a pyramid structure. Applied Mathematics and Computation, 169(1):157–178, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007672.

Shen:2005:SGC

[She05c] Zhizhang Shen. A study of a generalization of a card problem. Applied Mathematics and Computation, 166(2):385–410, July 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004448.

Shen:2006:BPB

[She06] Zhizhang Shen. A bypassing path based routing algorithm for the pyramid structures. Applied Mathematics and Computation, 181(2):1523–1543, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306002840.

Shen:2007:PGA

[She07a] Chunxia Shen. Permanence and global attractivity of the food-chain system with Holling IV type functional response. Applied Mathematics and Computation, 194(1):179–185, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004560.

Shen:2007:PPS

[She07b]  Chunxia Shen.  Posi-

tive periodic solution of a kind of nonlinear foodchain system. Applied Mathematics and Computation, 194(1):234–242, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004626.

Shen:2007:GBS

[She07c] Zhizhang Shen. A generalized broadcasting schema for the mesh structures. Applied Mathematics and Computation, 186(2):1293– 1310, March 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030601054X.

Shen:2008:SWS

[She08]  Jianwei Shen.  Shock

wave solutions of the compound Burgers–Korteweg– de Vries equation. Applied Mathematics and Computation, 196(2):842–849, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007370.

Song:2007:SBK

[SHG07] Lingyu Song, Yinnian He, and Zhihao Ge. Stability and bifurcation for a kind of nonlinear delayed differential equations. Applied Mathematics and Computation, 190(1):677– 685, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001038.

Shim:2005:LPG

[Shi05a]  Kyungah Shim.  Off-

line password-guessing attacks on the generalized key agreement and password authentication protocol. Applied Mathematics and Computation, 169(1):511–515, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304007908.

Shirazi:2005:FPA

[Shi05b] Kourosh H. Shirazi. Five points of accuracy synthesis of motion generator linkages using Maple-V. Applied Mathematics and Computation, 169(2):1142–1171, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008422.

Shirazi:2005:SLF

[Shi05c] Kourosh H. Shirazi. Synthesis of linkages with four points of accuracy using Maple-V. Applied Mathematics and Computation, 164(3):731–755, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003248.

Shi:2009:CPC

[Shi09a] Weichen Shi. Collinear periodic cracks and/or rigid line inclusions of antiplane sliding mode in one-dimensional hexagonal quasicrystal. Applied Mathematics and Computation, 215(3):1062–1067, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005992.

Shi:2009:NMS

[Shi09b]  Yong-Guo Shi.  Non-

monotonic solutions and continuously differentiable solutions of conjugacy equations. Applied Mathematics and Computation, 215(7):2399–2404, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007632.

Shaojian:2006:GOA

[SHK06] Qu Shaojian, Yin Hongyou, and Zhang Kecun. A global optimization algorithm using linear relaxation. Applied Mathematics and Computation, 178(2):510–518, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009987.

Sun:2007:CEA

[SHSZ07] Da-Zhi Sun, Jin-Peng Huai, Ji-Zhou Sun, and Jia-Wan Zhang. Computational efficiency analysis of Wu et al.’s fast modular multi-exponentiation algorithm. Applied Mathematics and Computation, 190(2):1848–1854, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002299.

Shtern:2007:SVW

[Sht07]  A. I. Shtern.  Stabil-

ity of the van der Waerden theorem on the continuity of homomorphisms of compact semisimple Lie groups. Applied Mathematics and Computation, 187(1):455–465, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011969.

Shubov:2005:BTV

[Shu05] Marianna A. Shubov. Bending-torsion vibration model with two-end energy dissipation. Applied Mathematics and Computation, 164(2):351–372, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003236.

Shyur:2006:CEU

[Shy06] Huan-Jyh Shyur. COTS evaluation using modified TOPSIS and ANP. Applied Mathematics and Computation, 177(1):251–259, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009070.

Shang:2008:EHF

[SHY08] Yadong Shang, Yong Huang, and Wenjun Yuan. The extended hyperbolic functions method and new exact solutions to the Zakharov equations. Applied Mathematics and Computation, 200(1):110–122, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010922.

Soh:2006:SAS

[SI06] C. Wafo Soh and O. J.

Ilegbusi. Semi-analytical solution of a constrained fourth-order integro-differential equation of steady flow– structure interaction in a model collapsible tube. Applied Mathematics and Computation, 174(2):1564– 1575, March 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305005801.

Sidi:2005:NIS

[Sid05] Avram Sidi. Numerical integration over smooth surfaces in R^{3} via class S_{⇕} variable transformations. Part I: Smooth integrands. Applied Mathematics and Computation, 171(1):646–674, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305001220.

Sidi:2006:NIS

[Sid06] Avram Sidi. Numerical integration over smooth surfaces in R^{∋} via class S_{⇕} variable transformations. Part II: Singular integrands. Applied Mathematics and Computation, 181(1):291–309, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306000981.

Siddique:2009:SCN

[Sid09] Mohammad Siddique. Smoothing of Crank–Nicolson scheme for the two-dimensional diffusion with an integral condition. Applied Mathematics and Computation, 214(2):512–522, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003476.

Silver:2005:ATDb

[Sil05a]  G. L. Silver.  Analy-

sis of three-dimensional grids: cubes and cubic coefficients. Applied Mathematics and Computation, 166(1):196–203, July 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004308.

See erratum [Sil06h].

Silver:2005:ATDd

[Sil05b] G. L. Silver. Analysis of three-dimensional grids: Cubic equations for ninepoint prismatic arrays. Applied Mathematics and Computation, 170(2):752– 760, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Silver:2005:ATDa

[Sil05c]  G. L. Silver.  Analy-

sis of three-dimensional grids: five- and nine-point cubes. Applied Mathematics and Computation, 160(1):133–140, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303010956.

Silver:2005:ATDc

[Sil05d]  G. L. Silver.  Analy-

sis of three-dimensional grids: four- and fivepoint cubes. Applied Mathematics and Computation, 166(2):457–463, July 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004503.

Silver:2006:ATDd

[Sil06a] G. L. Silver. Analysis of three-dimensional grids: Alternative methods for five- and nine-point cubes. Applied Mathematics and Computation, 176(1):37– 43, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030500768X.

Silver:2006:ATDf

[Sil06b]  G. L. Silver.  Analy-

sis of three-dimensional grids: Alternative polynomial equations for the nine-point prismatic array. Applied Mathematics and Computation, 182(2):1201–1207, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004346.

Silver:2006:ATDb

[Sil06c] G. L. Silver. Analysis of three-dimensional grids: Five-point cubes and curvature coefficients. Applied Mathematics and Computation, 174(2):1033–1038, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005308.

Silver:2006:ATDc

[Sil06d]  G. L. Silver.  Analy-

sis of three-dimensional grids: Interpolation of data in prismatic array and the estimation of a missing datum. Applied Mathematics and Computation, 174(2):1398–1407, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005588.

Silver:2006:ATDg

[Sil06e] G. L. Silver. Analysis of three-dimensional grids: Sine-and-cosine equation for the nine-point prismatic array. Applied Mathematics and Computation, 183(1):617–621, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005972.

Silver:2006:ATDe

[Sil06f] G. L. Silver. Analysis of three-dimensional grids: The cube and the octahedron. Applied Mathematics and Computation, 179(2):553–558, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030501060X.

Silver:2006:ATDa

[Sil06g]  G. L. Silver.  Analy-

sis of three-dimensional grids: The four-point cube. Applied Mathematics and Computation, 172(2):963–970, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030500250X.

Silver:2006:EAT

[Sil06h]  G. L. Silver.  Erra-

tum to: “Analysis of three-dimensional grids: Cubes and cubic coefficients” [Applied Mathematics and Computation 166 (2005) 196–203]. Applied Mathematics and Computation, 174(2):1659, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005655.

See [Sil06c].

Silver:2007:ATDb

[Sil07a] G. L. Silver. Analysis of three-dimensional grids: Interpolating the ninepoint prism. The shifting operator as an analytical instrument. Applied Mathematics and Computation, 186(2):1694–1700, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011040.

Silver:2007:ATDa

[Sil07b]  G. L. Silver.  Analy-

sis of three-dimensional grids: The estimation of two missing data. Applied Mathematics and Computation, 184(2):743–747, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600751X.

Silver:2007:CPG

[Sil07c] G. L. Silver. Center points and grid equations from sub-grid equations. Applied Mathematics and Computation, 189(2):1467–1475, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017334.

Silver:2007:OCP

[Sil07d] G. L. Silver. Operational center points. Applied Mathematics and Computation, 188(1):1042–1048, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014536.

Silver:2007:OD

[Sil07e] G. L. Silver. Operational discretization. Applied Mathematics and Computation, 188(1):440–445, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306013828.

Silver:2007:OMC

[Sil07f]  G. L. Silver.  Opera-

tional measures of central tendency. Applied Mathematics and Computation, 186(2):1379–1384, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010630.

Simsek:2007:THZ

[Sim07]  Yilmaz Simsek.  On

twisted q-Hurwitz zeta function and q-two-variable L-function. Applied Mathematics and Computation, 187(1):466–473, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306011970.

Simic:2009:BPG

[Sim09a] Slavko Simic. Best possible global bounds for Jensen’s inequality. Applied Mathematics and Computation, 215(6):2224–2228, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007346.

Simos:2009:HOC

[Sim09b]  T. E. Simos.  High

order closed Newton– Cotes trigonometricallyfitted formulae for the numerical solution of the Schrödinger equation. Applied Mathematics and Computation, 209(1):137– 151, March 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300308004815.

Simos:2009:P

[Sim09c]  T. E. Simos.  Pref-

ace. Applied Mathematics and Computation, 209(1): 1, March 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004670.

Singh:2005:TGT

[Sin05] Baljeet Singh. On the theory of generalized thermoelasticity for piezoelectric materials. Applied Mathematics and Computation, 171(1):398–405, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305001049.

Singh:2007:WPG

[Sin07a]  Baljeet Singh.  Wave

propagation in a generalized thermoelastic material with voids. Applied Mathematics and Computation, 189(1):698–709, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016468.

Singh:2007:RMF

[Sin07b]  S. R. Singh.  A ro-

bust method of forecasting based on fuzzy time series. Applied Mathematics and Computation, 188(1):472– 484, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013877.

Singh:2007:SMF

[Sin07c]  S. R. Singh.  A sim-

ple method of forecasting based on fuzzy time series. Applied Mathematics and Computation, 186(1): 330–339, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030600943X.

Singh:2008:EHI

[Sin08a] Baljeet Singh. Effect of hydrostatic initial stresses on waves in a thermoelastic solid half-space. Applied Mathematics and Computation, 198(2):494–505, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009150.

Singh:2008:IGR

[Sin08b] Vimal Singh. Improved global robust stability of interval delayed neural networks via split interval: Generalizations. Applied Mathematics and Computation, 206(1):290–297, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006346.

Singh:2009:MCG

[Sin09]  Vimal Singh.  Modi-

fied criteria for global robust stability of interval delayed neural networks. Applied Mathematics and Computation, 215(8):3124–3133, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309009060.

Sitarz:2006:HMM

[Sit06]  Sebastian Sitarz.  Hy-

brid methods in multicriteria dynamic programming. Applied Mathematics and Computation, 180(1): 38–45, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305010957.

Stevic:2009:ITO

[SiU09] Stevo Stević and Sei ichiro Ueki. Integral-type operators acting between weighted-type spaces on the unit ball. Applied Mathematics and Computation, 215(7):2464–2471, December 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309007826.

Sivakumar:2007:GIT

[Siv07] K. C. Sivakumar. Generalized inverses of tridiagonal operators. Applied Mathematics and Computation, 189(2):1300–1303, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017164.

Sun:2005:RKM

[SJ05] Wei Sun and Yao-Lin Jiang. Runge–Kutta methods of dynamic iteration for index-2 differential– algebraic equations. Applied Mathematics and Computation, 169(2):1346– 1363, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304008562.

Shao:2006:FMP

[SJ06a] Zhen Shao and Xiaoyu Ji. Fuzzy multi-product constraint newsboy problem. Applied Mathematics and Computation, 180(1):7–15, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010921.

Shen:2006:LMC

[SJ06b] Peiping Shen and Hongwei Jiao. Linearization method for a class of multiplicative programming with exponent. Applied Mathematics and Computation, 183(1):328–336, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005650.

Shen:2006:NRB

[SJ06c] Peiping Shen and Hongwei Jiao. A new rectangle branch-and-pruning approach for generalized geometric programming. Applied Mathematics and Computation, 183(2):1027– 1038, December 15, 2006.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306006382.

Sun:2006:NDD

[SJ06d] Wei Sun and Yao-Lin Jiang. A note on discrete dynamic iterations of stiff index-2 DAE’s. Applied Mathematics and Computation, 172(2):1201–1211, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002870.

Saeidpourazar:2008:NRMa

[SJ08a] Reza Saeidpourazar and Nader Jalili. Nano-robotic manipulation using a RRP nanomanipulator: Part A

— Mathematical modeling and development of a robust adaptive driving mechanism. Applied Mathematics and Computation, 206(2):618–627, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003858.

Saeidpourazar:2008:NRMb

[SJ08b] Reza Saeidpourazar and Nader Jalili. Nano-robotic manipulation using a RRP nanomanipulator: Part B

— Robust control of manipulator’s tip using fused visual servoing and force sensor feedbacks. Applied Mathematics and Computation, 206(2):628–642, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800386X.

Shi:2009:EIV

[SJC09] Ruiqing Shi, Xiaowu Jiang, and Lansun Chen. The effect of impulsive vaccination on an SIR epidemic model. Applied Mathematics and Computation, 212(2):305–311, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001453.

Sjoberg:2007:DRP

[Sjö07] A. Sjöberg. Double reduction of PDEs from the association of symmetries with conservation laws with applications. Applied Mathematics and Computation, 184(2):608–616, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600734X.

Song:2009:ASI

[SJW09] Xinyu Song, Yu Jiang, and Huiming Wei. Analysis of a saturation incidence SVEIRS epidemic model with pulse and two time delays. Applied Mathematics and Computation, 214(2):381–390, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003245.

Shidfar:2005:NSI

[SK05] A. Shidfar and G. R.

Karamali. Numerical solution of inverse heat conduction problem with nonstationary measurements. Applied Mathematics and Computation, 168(1):540–548, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006228.

Saxena:2006:UMD

[SK06a] R. K. Saxena and S. L.

Kalla. On a unified mixture distribution. Applied Mathematics and Computation, 182(1):325–332, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003407.

Sharma:2006:SDO

[SK06b] K. K. Sharma and Aditya Kaushik. A solution of the discrepancy occurs due to using the fitted mesh approach rather than to the fitted operator for solving singularly perturbed differential equations. Applied Mathematics and Computation, 181(1):756–766, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001561.

Seckiner:2007:SAA

[SK07a] Serap Ulusam Seçkiner and Mustafa Kurt. A simulated annealing approach to the solution of job rotation scheduling problems. Applied Mathematics and Computation, 188(1):31– 45, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306013166.

Singh:2007:WRV

[SK07b] Baljeet Singh and Rajneesh Kumar. Wave reflection at viscoelastic–micropolar elastic interface. Applied Mathematics and Computation, 185(1):421–431, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008770.

Saxena:2008:SCF

[SK08a] R. K. Saxena and S. L.

Kalla. On the solutions of certain fractional kinetic equations. Applied Mathematics and Computation, 199(2):504–511, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010168.

Schneider:2008:PBF

[SK08b] Baruch Schneider and Ömer Kavaklıoḡlu. Poincaré– Bertrand formula on a piecewise Liapunov curve in two-dimensional. Applied Mathematics and Computation, 202(2):814– 819, August 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308001859.

Seckiner:2008:ACO

[SK08c] Serap Ulusam Seçkiner and Mustafa Kurt. Ant colony optimization for the job rotation scheduling problem. Applied Mathematics and Computation, 201(1–2):149–160, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011794.

Srinivas:2009:IHM

[SK09a] S. Srinivas and M. Kothandapani. The influence of heat and mass transfer on MHD peristaltic flow through a porous space with compliant walls. Applied Mathematics and Computation, 213(1):197–208, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002185.

Susanto:2009:NMB

[SK09b] H. Susanto and N. Karjanto. Newton’s method’s basins of attraction revisited. Applied Mathematics and Computation, 215(3):1084–1090, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006134.

Saidi:2006:ESS

[SKHZ06] H. Saidi, N. Khelil, S. Hassouni, and A. Zerarka. Energy spectra of the Schrödinger equation and the differential quadrature method: Improvement of the solution using particle swarm optimization. Applied Mathematics and Computation, 182(1):559–566, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600364X.

Singh:2006:RGT

[SKS06] Baljeet Singh, Ajay Kumar, and Jagdish Singh. Reflection of generalized thermoelastic waves from a solid half-space under hydrostatic initial stress. Applied Mathematics and Computation, 177(1):170– 177, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008982.

Scitovski:2006:MMC

[SKSJ06] Rudolf Scitovski, Gordana Kralik, Kristian Sabo, and Tatjana Jelen. A mathematical model of controlling the growth of tissue in pigs. Applied Mathematics and Computation, 181(2):1126–1138, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001974.

Salkuyeh:2006:PAA

[SKT06] Davod Khojasteh Salkuyeh, Saeed Karimi, and Faezeh Toutounian. A parallel algorithm to approximate inverse factors of a matrix via sparse–sparse iterations. Applied Mathematics and Computation, 181(1):782–792, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001597.

Shim:2005:SFA

[SL05a] Kyungah Shim and Young-Ran Lee. Security flaws in authentication and key establishment protocols for mobile communications. Applied Mathematics and Computation, 169(1):62–74, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304007520.

Su:2005:GAS

[SL05b] You-Hui Su and Wan-Tong Li. Global asymptotic stability of a secondorder nonlinear difference equation. Applied Mathematics and Computation, 168(2):981–989, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006563.

Song:2006:GSP

[SL06a] Xinyu Song and Yongfeng Li. Global stability and periodic solution of a model for HTLV-I infection and ATL progression. Applied Mathematics and Computation, 180(1):401–410, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305011355.

Su:2006:SCT

[SL06b] Yongfu Su and Suhong Li. Strong convergence theorems on two iterative method for non-expansive mappings. Applied Mathematics and Computation, 181(1):332–341, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001019.

Sun:2006:MPS

[SL06c] Hong-Rui Sun and Wan-Tong Li. Multiple positive solutions for p-Laplacian m-point boundary value problems on time scales. Applied Mathematics and Computation, 182(1):478–491, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003559.

Shen:2007:PSK

[SL07a] Jianhua Shen and Ruixi Liang. Periodic solutions for a kind of second order neutral functional differential equations. Applied Mathematics and Computation, 190(2):1394–1401, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001701.

Shih:2007:NIE

[SL07b] Yintzer Shih and Thomas R. Lucas. New implementation of Elliptic Systems Method for time dependent diffusion tomography with back reflected and transmitted boundary data. Applied Mathematics and Computation, 188(1):64– 74, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601321X.

Shen:2009:ABS

[SL09a] Jianhua Shen and Yanjun Liu. Asymptotic behavior of solutions for nonlinear delay differential equation with impulses. Applied Mathematics and Computation, 213(2):449–454, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002562.

Shi:2009:CSA

[SL09b] Dongyang Shi and Hui Liang. Convergence and superconvergence analysis of a new quadratic Hermite-type triangular element on anisotropic meshes. Applied Mathematics and Computation, 212(1):257–269, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309001428.

Slavov:2006:STR

[Sla06] Zdravko Dimitrov Slavov.

Scalarization techniques or relationship between a social welfare function and a Pareto optimality concept. Applied Mathematics and Computation, 172(1):464–471, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002158.

Su:2005:KPK

[SLC05] Pin-Chang Su, Erl-Huei Lu, and Henry Ker-Chang Chang. A knapsack public-key cryptosystem based on elliptic curve discrete logarithm. Applied Mathematics and Computation, 168(1):40– 46, September 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304005831.

Shin:2007:OCP

[SLC07] Yong Hyun Shin, Byung Hwa Lim, and U Jin Choi. Optimal consumption and portfolio selection problem with downside consumption constraints. Applied Mathematics and Computation, 188(2):1801–1811, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015487.

Shi:2009:SUT

[SLC09] Zhen-Xia Shi, Wan-Tong Li, and Cui-Ping Cheng.

Stability and uniqueness of traveling wavefronts in a two-dimensional lattice differential equation with delay. Applied Mathematics and Computation, 208(2):484–494, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009211.

Song:2008:PSN

[SLL08] Xinyu Song, Senlin Li, and An Li. Practical stability of nonlinear differential equation with initial time difference. Applied Mathematics and Computation, 203(1):157–162, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002208.

See:2007:IRU

[SLLL07] K. W. See, K. S. Loke, P. A. Lee, and K. F. Loe. Image reconstruction using various discrete orthogonal polynomials in comparison with DCT. Applied Mathematics and Computation, 193(2):346–359, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004043.

Slodicka:2007:EMM

[Slo07] Marián Slodička. Estimation of the Michaelis– Menten parameters for reductive dechlorination of chlorinated solvents. Applied Mathematics and Computation, 189(1):105– 115, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015852.

Sinkala:2008:OSG

[SLO08] W. Sinkala, P. G. L. Leach, and J. G. O’Hara. An optimal system and groupinvariant solutions of the Cox–Ingersoll–Ross pricing equation. Applied Mathematics and Computation, 201(1–2):95–107, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011745.

Su:2008:PSS

[SLS08] You-Hui Su, Wan-Tong Li, and Hong-Rui Sun. Positive solutions of singular p-Laplacian dynamic equations with sign changing nonlinearity. Applied Mathematics and Computation, 200(1):352–368, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011241.

Song:2008:CSO

[SLW08] Guangxing Song, Wenjuan Li, and Guotao Wang. A class of second order impulsive integro-differential equations on unbounded domain in a Banach space. Applied Mathematics and Computation, 203(2):733–744, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003974.

Shi:2006:FMC

[SLWX06] Xiaohong Shi, LuoLiang, Yan Wan, and Jin Xu. A finding maximal clique algorithm for predicting loop of protein structure. Applied Mathematics and Computation, 180(2):676–682, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000567.

Shen:2005:BTW

[SLX05a] Jianwei Shen, Jibin Li, and Wei Xu. Bifurcations of travelling wave solutions in a model of the hydrogenbonded systems. Applied Mathematics and Computation, 171(1):242–271, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000925.

Sugisaka:2005:DAB

[SLX^{+}05b] M. Sugisaka, A. Loukianov, F. Xiongfeng, T. Kubik, and K. B. Kubik. Development of an artificial brain for LifeRobot. Applied Mathematics and Computation, 164(2):507–521, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003339.

Shi:2005:ETB

[SLZ05a] Chenggen Shi, Jie Lu, and Guangquan Zhang. An extended Kth-best approach for linear bilevel programming. Applied Mathematics and Computation, 164(3):843–855, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003510.

Shi:2005:EKT

[SLZ05b] Chenggen Shi, Jie Lu, and Guangquan Zhang. An extended Kuhn–Tucker approach for linear bilevel programming. Applied Mathematics and Computation, 162(1):51–63, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000554.

Shi:2006:EBB

[SLZZ06] Chenggen Shi, Jie Lu, Guangquan Zhang, and Hong Zhou. An extended branch and bound algorithm for linear bilevel programming. Applied Mathematics and Computation, 180(2):529–537, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000415.

Saati:2005:RWF

[SM05a] S. Saati and A. Memariani. Reducing weight flexibility in fuzzy DEA. Applied Mathematics and Computation, 161(2):611– 622, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300304000153.

Shadman:2005:NHH

[SM05b] Dariush Shadman and Bahman Mehri. A nonhomogeneous Hill’s equation. Applied Mathematics and Computation, 167(1):68–75, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004734.

Sjoberg:2005:ANL

[SM05c] A. Sjöberg and F. M. Mahomed. The association of non-local symmetries with conservation laws: applications to the heat and Burger’s equations. Applied Mathematics and Computation, 168(2):1098–1108, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304006666.

Sun:2005:ORM

[SM05d] Yuan Gong Sun and Fan Wei Meng. Oscillation results for matrix differential systems with damping. Applied Mathematics and Computation, 170(1):545–555, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009671.

Samanta:2006:EBT

[SM06a] Bablu Samanta and Sanat Kumar Majumder. Entropy based transportation problem use geometric programming method. Applied Mathematics and Computation, 179(2):431–439, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010519.

Shavandi:2006:FQL

[SM06b] Hassan Shavandi and Hashem Mahlooji. A fuzzy queuing location model with a genetic algorithm for congested systems. Applied Mathematics and Computation, 181(1):440–456, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001147.

Shi:2006:MIF

[SM06c] Guoliang Shi and Xianrui Meng. Monotone iterative for fourth-order p-Laplacian boundary value problems with impulsive effects. Applied Mathematics and Computation, 181(2):1243–1248, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306002438.

Sun:2006:NPD

[SM06d] Yuan Gong Sun and Fan Wei Meng. Note on the paper of Džurina and Stavroulakis. Applied Mathematics and Computation, 174(2):1634–1641, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005849.

See remarks [?].

Sun:2006:ORM

[SM06e] Yuan Gong Sun and Fan Wei Meng. Oscillation results for matrix differential systems with damping. Applied Mathematics and Computation, 175(1):212– 220, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006016.

Siddiqui:2007:WOF

[SM07a] Azad A. Siddiqui and M. T. Mustafa. Wavelet optimized finite difference method with nonstatic re-gridding. Applied Mathematics and Computation, 186(1):203–211, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009301.

Siddiqui:2007:SMC

[SM07b] S. U. Siddiqui and Shailesh Mishra. A study of modified Casson’s fluid in modelled normal and stenotic capillary-tissue diffusion phenomena. Applied Mathematics and Computation, 189(2):1048–1057, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601681X.

Sun:2007:OHO

[SM07c] Y. G. Sun and A. B.

Mingarelli. Oscillation of higher-order forced nonlinear differential equations. Applied Mathematics and Computation, 190(1):905– 911, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700121X.

Sun:2007:LAS

[SM07d] Yuan Gong Sun and Fan Wei Meng. LMI approach to stability for a competitive Lotka–Volterra system with time-varying delays. Applied Mathematics and Computation, 194(2):291–297, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004687.

Sugie:2008:NTH

[SM08a] Jitsuro Sugie and Kouhei Matsumura. A nonoscillation theorem for halflinear differential equations with periodic coefficients. Applied Mathematics and Computation, 199(2):447– 455, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010119.

Sun:2008:ICO

[SM08b] Yuan Gong Sun and Fan Wei Meng. Interval criteria for oscillation of second-order differential equations with mixed nonlinearities. Applied Mathematics and Computation, 198(1):375–381, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008983.

Sun:2008:NOS

[SM08c] Yuan Gong Sun and Fan Wei Meng. Nonoscillation and oscillation of second order half-linear difference equations. Applied Mathematics and Computation, 197(1):121–127, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007606.

Sunil:2008:NSA

[SM08d] Sunil and Amit Mahajan. A nonlinear stability analysis for rotating magnetized ferrofluid heated from below. Applied Mathematics and Computation, 204(1):299–310, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005079.

Sunil:2008:TMF

[SM08e] Sunil and Amit Mahajan. Thermoconvective magnetized ferrofluid with internal angular momentum saturating a porous medium: a nonlinear stability analysis. Applied Mathematics and Computation, 205(1):403–416, November 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300308006024.

Karmakar:2008:MEA

[SM08f] Kakali Karmakar (Sur) and Sanat Kumar Majumder. Maximum entropy approach in a traffic stream. Applied Mathematics and Computation, 195(1):61–65, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005176.

Srisuphab:2009:GKA

[SM09a] Ananta Srisuphab and Jarernsri L. Mitrpanont. Gaussian kernel approximation algorithm for feedforward neural network design. Applied Mathematics and Computation, 215(7):2686–2693, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309008066.

Sun:2009:STD

[SM09b] Tongjun Sun and Keying Ma. A space–time discontinuous Galerkin method for linear convection-dominated Sobolev equations. Applied Mathematics and Computation, 210(2):490–503, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000289.

Sun:2009:OSO

[SM09c] Yuan Gong Sun and Fan Wei Meng. Oscillation of second-order delay differential equations with mixed nonlinearities. Applied Mathematics and Computation, 207(1):135–139, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007728.

Szidarovszky:2009:SDC

[SM09d] Ferenc Szidarovszky and Akio Matsumoto. Spatial duopoly with cost subsidy. Applied Mathematics and Computation, 215(5):1659–1670, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006663.

Salahi:2006:PTS

[SMA06] Maziar Salahi and Nezam Mahdavi-Amiri. Polynomial time second order Mehrotra-type predictor– corrector algorithms. Applied Mathematics and Computation, 183(1):646– 658, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030600600X.

Sadeghi:2008:CTS

[SMA08] M. Sha Sadeghi, H. R. Momeni, and R. Amirifar. H_{∞} and L_{1} control of a teleoperation system via LMIs. Applied Mathematics and Computation, 206(2):669–677, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003901.

Sharma:2008:AHP

[SMB08] Mithun J. Sharma, Ilkyeong Moon, and Hyerim Bae. Analytic hierarchy process to assess and optimize distribution network. Applied Mathematics and Computation, 202(1):256–265, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000957.

Sarkar:2007:TLC

[SMBB07] R. R. Sarkar, B. Mukhopadhyay, R. Bhattacharyya, and Sandip Banerjee. Time lags can control algal bloom in two harmful phytoplankton–zooplankton system. Applied Mathematics and Computation, 186(1):445–459, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306009532.

Shukla:2008:MAA

[SMC08] J. B. Shukla, A. K.

Misra, and Peeyush Chandra. Modeling and analysis of the algal bloom in a lake caused by discharge of nutrients. Applied Mathematics and Computation, 196(2):782–790, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007308.

Smirlis:2006:DEA

[SMD06] Yannis G. Smirlis, Elias K.

Maragos, and Dimitris K. Despotis. Data envelopment analysis with missing values: an interval DEA approach. Applied Mathematics and Computation, 177(1):1–10, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008817.

Shen:2007:BAT

[SMG07] Jianwei Shen, Baojun Miao, and Lingling Guo. Bifurcation analysis of travelling wave solutions in the nonlinear Klein– Gordon model with anharmonic coupling. Applied Mathematics and Computation, 188(2):1975–1983, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015669.

Sun:2008:SIP

[SMG08] Bo Sun, Chunmei Miao, and WeiGao Ge. Successive iteration and positive symmetric solutions for some Sturm–Liouvillelike four-point p-Laplacian boundary value problems. Applied Mathematics and Computation, 201(1–2):481–488, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307012246.

Siyyam:2007:SNS

[SMH07] H. Siyyam, N. Merabet, and M. H. Hamdan. Standard numerical schemes for coupled parallel flow over porous layers. Applied Mathematics and Computation, 194(1):38–45, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004407.

Sun:2006:NCS

[SMHQ06] Jian-Qiang Sun, Zhong-Qi Ma, Wei Hua, and Meng-Zhao Qin. New conservation schemes for the nonlinear Schrödinger equation. Applied Mathematics and Computation, 177(1):446– 451, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009239.

Smith:2006:CTM

[Smi06a] Warren D. Smith. Church’s thesis meets the N-body problem. Applied Mathematics and Computation, 178(1):154–183, July 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305008490.

Smith:2006:TCR

[Smi06b] Warren D. Smith. Three counterexamples refuting Kieu’s plan for “quantum adiabatic hypercomputation”; and some uncomputable quantum mechanical tasks. Applied Mathematics and Computation, 178(1):184–193, July 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008507.

Smietanski:2007:NCP

[Śmi07] Marek J. Śmietański. On a new class parametrized Newton-like method for semismooth equations. Applied Mathematics and Computation, 193(2):430– 437, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030700416X.

Saen:2005:ECC

[SML05a] R. Farzipoor Saen, A. Memariani, and F. Hosseinzadeh Lotfi. The effect of correlation coefficient among multiple input vectors on the efficiency mean in data envelopment analysis. Applied Mathematics and Computation, 162(2):503–521, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001067.

Saen:2005:DRE

[SML05b] Reza Farzipoor Saen, Azizollah Memariani, and Farhad Hosseinzadeh Lotfi. Determining relative efficiency of slightly nonhomogeneous decision making units by data envelopment analysis: a case study in IROST. Applied Mathematics and Computation, 165(2):313–328, June 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003844.

Sadhukhan:2008:DAS

[SMM08] D. Sadhukhan, B. Mondal, and M. Maiti. Discrete age-structured population model with age dependent harvesting and its stability analysis. Applied Mathematics and Computation, 201(1–2):631–639, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800009X.

Sun:2006:SER

[SMQ06] Jian-Qiang Sun, Zhong-Qi Ma, and Meng-Zhao Qin. Simulation of envelope Rossby solitons in a pair of cubic Schrödinger equations. Applied Mathematics and Computation, 183(2):946–952, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600628X.

Sagara:2008:NSS

[SMY08] Muneomi Sagara, Hiroaki Mukaidani, and Toru Yamamoto. Numerical solution of stochastic Nash games with state-dependent noise for weakly coupled largescale systems. Applied Mathematics and Computation, 197(2):844–857, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008545.

Saberi-Nadjafi:2007:GBB

[SNH07a] Jafar Saberi-Nadjafi and Mahdi Heidari. A generalized block-by-block method for solving linear Volterra integral equations. Applied Mathematics and Computation, 188(2):1969–1974, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015657.

Saberi-Nadjafi:2007:QMV

[SNH07b] Jafar Saberi-Nadjafi and Mahdi Heidari. A quadrature method with variable step for solving linear Volterra integral equations of the second kind. Applied Mathematics and Computation, 188(1):549– 554, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013944.

Saberi-Nadjafi:2007:SLI

[SNH07c] Jafar Saberi-Nadjafi and Mahdi Heidari. Solving linear integral equations of the second kind with repeated modified trapezoid quadrature method. Applied Mathematics and Computation, 189(1):980– 985, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016778.

Srivastava:2006:SGL

[SNK06] H. M. Srivastava, Saralees Nadarajah, and Samuel Kotz. Some generalizations of the Laplace distribution. Applied Mathematics and Computation, 182(1):223–231, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003316.

Sarduvan:2008:LCT

[SÖ08] Murat Sarduvan and Halim Özdemir. On linear combinations of two tripotent, idempotent, and involutive matrices. Applied Mathematics and Computation, 200(1):401–406, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011447.

Sofo:2009:IFS

[Sof09]  Anthony Sofo.  Inte-

gral forms of sums associated with harmonic numbers. Applied Mathematics and Computation, 207(2):365–372, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008229.

Sogabe:2007:TTR

[Sog07]  Tomohiro Sogabe.  On

a two-term recurrence for the determinant of a general matrix. Applied Mathematics and Computation, 187(2):785–788, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012112.

Sogabe:2008:FNA

[Sog08a] Tomohiro Sogabe. A fast numerical algorithm for the determinant of a pentadiagonal matrix. Applied Mathematics and Computation, 196(2):835–841, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007369.

See note [Sog08c].

Sogabe:2008:NASb

[Sog08b] Tomohiro Sogabe. New algorithms for solving periodic tridiagonal and periodic pentadiagonal linear systems. Applied Mathematics and Computation, 202(2):850–856, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001896.

Sogabe:2008:NFN

[Sog08c]  Tomohiro Sogabe.  A

note on “A fast numerical algorithm for the determinant of a pentadiagonal matrix”. Applied Mathematics and Computation, 201(1–2):561–564, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307012337.

See [Sog08a].

Sogabe:2008:NASa

[Sog08d] Tomohiro Sogabe. Numerical algorithms for solving comrade linear systems based on tridiagonal solvers. Applied Mathematics and Computation, 198(1):117–122, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008715.

Sohn:2005:SPC

[Soh05] Ayoung Sohn. Stokes problem of the cornered domain in the plane. Applied Mathematics and Computation, 161(1):93–99, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303012116.

Soh:2006:NPS

[Soh06] Célestin Wafo Soh. Nonperturbative semi-analytical source-type solutions of thin-film equation. Applied Mathematics and Computation, 174(2):1576–1585, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005813.

Sokol:2007:SCC

[Sok07a] Janusz Sokół. On sufficient condition to be in a certain subclass of starlike functions defined by subordination. Applied Mathematics and Computation, 190(1):237–241, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000446.

Sokol:2007:SHP

[Sok07b] Janusz Sokół. Starlikeness of Hadamard product of certain analytic functions. Applied Mathematics and Computation, 190(2):1157–1160, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700135X.

Sokol:2008:CAM

[Sok08a] Janusz Sokół. On a class of analytic multivalent functions. Applied Mathematics and Computation, 203(1):210–216, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002312.

Sokol:2008:FDS

[Sok08b] Janusz Sokół. On functions with the derivative satisfying a geometric condition. Applied Mathematics and Computation, 204(1):116–119, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004657.

Sokol:2008:SAD

[Sok08c] Janusz Sokół. On some applications of the Dziok– Srivastava operator. Applied Mathematics and Computation, 201(1–2): 774–780, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300308000453.

Sokol:2008:SLT

[Sok08d] Janusz Sokół. Starlikeness of the Libera transform of functions with bounded turning. Applied Mathematics and Computation, 203(1):273–276, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800249X.

Sokol:2009:RPC

[Sok09]  Janusz Sokół.  Radius

problems in the class SL^{∗}. Applied Mathematics and Computation, 214(2):569–573, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003555.

Solak:2005:NCM

[Sol05a] Süleyman Solak. On the norms of circulant matrices with the Fibonacci and Lucas numbers. Applied Mathematics and Computation, 160(1):125–132, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303010920.

Soliman:2005:CPL

[Sol05b] A. A. Soliman. On cone perturbing Liapunov function for impulsive differential systems. Applied Mathematics and Computation, 163(3):1069–1079, April 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002802.

Soliman:2005:LSC

[Sol05c] A. A. Soliman. On Lipschitz stability for comparison systems of differential equations via limiting equation. Applied Mathematics and Computation, 163(3):1061–1067, April 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002796.

Soliman:2005:PSP

[Sol05d]  A. A. Soliman.  On

practical stability of perturbed differential systems. Applied Mathematics and Computation, 163(3):1055–1060, April 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002784.

Solak:2007:ENC

[Sol07] Süleyman Solak. Erratum to “On the norms of circulant matrices with the Fibonacci and Lucas numbers” [Appl. Math. Comput. 160 (2005) 125– 132]. Applied Mathematics and Computation, 190(2):1855–1856, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002421.

See [UNS07].

Song:2005:PMI

[Son05] Yongzhong Song. On parallel multisplitting iterative methods for singular linear systems. Applied Mathematics and Computation, 162(2):585–604, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001171.

Song:2007:ISM

[Son07] Yisheng Song. Iterative selection methods for the common fixed point problems in a Banach space. Applied Mathematics and Computation, 193(1):7–17, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700361X.

Song:2008:NSC

[Son08a] Yisheng Song. A new sufficient condition for the strong convergence of Halpern type iterations. Applied Mathematics and Computation, 198(2):721– 728, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009484.

Song:2008:SCA

[Son08b]  Yisheng Song.  Strong

convergence of averaged iteration for asymptotically non-expansive mappings. Applied Mathematics and Computation, 204(2):854–861, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005705.

Soriano:2006:ECZ

[Sor06]  J. M. Soriano.  Exis-

tence and computation of zeros of perturbed mappings. Applied Mathematics and Computation, 173(1):457–467, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305003668.

Srivastava:2007:AFT

[SOS07] H. M. Srivastava, Shigeyoshi Owa, and Tadayuki Sekine. Analytic function theory, fractional calculus and their applications. Applied Mathematics and Computation, 187(1):1–2, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011295.

Sekine:2007:IMC

[SOT07] Tadayuki Sekine, Shigeyoshi Owa, and Kazuyuki Tsurumi. Integral means of certain analytic functions for fractional calculus. Applied Mathematics and Computation, 187(1):425– 432, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011933.

Shidfar:2005:AFD

[SP05] A. Shidfar and R. Pourgholi.

Application of finite difference method to analysis an ill-posed problem. Applied Mathematics and Computation, 168(2):1400–1408, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007374.

Shidfar:2006:NAS

[SP06a] A. Shidfar and R. Pourgholi.

Numerical approximation of solution of an inverse heat conduction problem based on Legendre polynomials. Applied Mathematics and Computation, 175(2):1366–1374, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006831.

Song:2006:SBA

[SP06b] Yongli Song and Yahong Peng. Stability and bifurcation analysis on a Logistic model with discrete and distributed delays. Applied Mathematics and Computation, 181(2):1745–1757, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003067.

Sonia:2006:BTM

[SP06c] Sonia and M. C. Puri. Bilevel time minimizing assignment problem. Applied Mathematics and Computation, 183(2):990–999, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006345.

Sonia:2006:CCA

[SP06d] Sonia and M. C. Puri. On critical capacity of arcs in a directed network. Applied Mathematics and Computation, 181(2):1568–1576, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002876.

Stanimirovic:2006:CGI

[SP06e] Predrag S. Stanimirović and Marko D. Petković. Computing generalized inverse of polynomial matrices by interpolation. Applied Mathematics and Computation, 172(1):508– 523, January 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305002195.

Simos:2007:P

[SP07] T. E. Simos and G. Psihoyios. Preface. Applied Mathematics and Computation, 184(1):1, January 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004929.

Sharma:2008:VSA

[SP08] Shambhu N. Sharma and H. Parthasarathy. Volterra series arising from the discrete Schrödinger wave equation in L^{2}[a, b] space. Applied Mathematics and Computation, 196(2):563– 569, March 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307007084.

Shi:2009:LOC

[SP09] Dongyang Shi and Lifang Pei. Low order Crouzeix– Raviart type nonconforming finite element methods for the 3D timedependent Maxwell’s equations. Applied Mathematics and Computation, 211(1): 1–9, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000381.

Senthilkumaran:2008:NEE

[SPV08] M. Senthilkumaran, D. Pandiaraja, and B. Mayil Vaganan. New exact explicit solutions of the generalized KdV equations. Applied Mathematics and Computation, 202(2):693–699, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308001586.

Shan:2007:ALA

[SPX07] Feng Shan, Li-Ping Pang, and Zun-Quan Xia. An approximate U-Lagrangian and algorithm to UV decomposition. Applied Mathematics and Computation, 184(2):924–930, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600782X.

Su:2007:GIA

[SQ07] Yongfu Su and Xiaolong Qin. General iteration algorithm and convergence rate optimal model for common fixed points of nonexpansive mappings. Applied Mathematics and Computation, 186(1):271–278, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009374.

Shang:2011:CPO

[SQ11] Nina Shang and Huizeng Qin. Comments on the paper: “Oscillation of secondorder nonlinear ODE with damping” [Applied Mathematics and Computation 199 (2008) 644– 652]. Applied Mathematics and Computation, 218(6):2979–2980, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300311010125.

See [HM08b].

Sun:2008:EIP

[SQG08] Bo Sun, Ying Qu, and WeiGao Ge. Existence and iteration of positive solutions for a multipoint one-dimensional p-Laplacian boundary value problem. Applied Mathematics and Computation, 197(1):389–398, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307007874.

Shang:2008:AEE

[SQHY08] Yadong Shang, Jinghong Qin, Yong Huang, and Wenjun Yuan. Abundant exact and explicit solitary wave and periodic wave solutions to the Sharma–Tasso–Olver equation. Applied Mathematics and Computation, 202(2):532–538, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308001379.

Shanthi:2006:ANM

[SR06] V. Shanthi and N. Ramanujam. Asymptotic numerical method for boundary value problems for singularly perturbed fourthorder ordinary differential equations with a weak interior layer. Applied Mathematics and Computation, 172(1):252–266, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500202X.

Shi:2009:NSO

[SRH09] Dongyang Shi, Jincheng Ren, and Xiaobin Hao. A new second order nonconforming mixed finite element scheme for the stationary Stokes and Navier– Stokes equations. Applied Mathematics and Computation, 207(2):462–477, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008333.

Srivastava:2007:STF

[Sri07] Bhaskar Srivastava. Some theta-function identities. Applied Mathematics and Computation, 187(1):474– 478, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011982.

Srivastava:2009:CSI

[Sri09] Rekha Srivastava. Classes of series identities and associated hypergeometric reduction formulas. Applied Mathematics and Computation, 215(1):118–124, September 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300309003907.

Santos:2008:EWE

[SRMRGRL08] Lucelina Batista Santos, Marko Rojas-Medar, Gabriel Ruiz-Garzón, and Antonio Rufián-Lizana. Existence of weakly efficient solutions in nonsmooth vector optimization. Applied Mathematics and Computation, 200(2):547– 556, July 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011344.

Saito:2005:NFD

[SS05a] Norikazu Saito and Takashi Suzuki. Notes on finite difference schemes to a parabolic-elliptic system modelling chemotaxis. Applied Mathematics and Computation, 171(1):72–90, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000810.

Shi:2005:CDM

[SS05b] Zhen-Jun Shi and Jie Shen. Convergence of descent method without line search. Applied Mathematics and Computation, 167(1):94–107, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400476X.

Shi:2005:NDA

[SS05c] Zhen-Jun Shi and Jie Shen. A new descent algorithm with curve search rule. Applied Mathematics and Computation, 161(3):753–768, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400027X.

Shi:2005:NSM

[SS05d] Zhen-Jun Shi and Jie Shen.

A new super-memory gradient method with curve search rule. Applied Mathematics and Computation, 170(1):1–16, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008677.

Singh:2005:PMN

[SS05e] Sumeet Singh and M. P. Singh. Prediction of missed number from pin code of ATM card with back propagation method of artificial neural network. Applied Mathematics and Computation, 171(1):152–162, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000871.

Sun:2005:EPP

[SS05f] Y. G. Sun and S. H. Saker. Existence of positive periodic solutions of nonlinear discrete model exhibiting the Allee effect. Applied Mathematics and Computation, 168(2):1086–1097, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304006654.

Sun:2005:OSO

[SS05g] Y. G. Sun and S. H. Saker. Oscillation for secondorder nonlinear neutral delay difference equations. Applied Mathematics and Computation, 163(2):909–918, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002693.

Sun:2005:OAB

[SS05h] Y. G. Sun and S. H. Saker.

Oscillatory and asymptotic behavior of positive periodic solutions of nonlinear discrete model exhibiting the Allee effect. Applied Mathematics and Computation, 168(2):1205–1218, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400685X.

Santos:2006:PAG

[SS06a] P. S. M. Santos and S. Scheimberg. A projection algorithm for general variational inequalities with perturbed constraint sets. Applied Mathematics and Computation, 181(1):649–661, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001457.

Shi:2006:CPM

[SS06b] Zhen-Jun Shi and Jie Shen. Convergence of PRP method with new nonmonotone line search. Applied Mathematics and Computation, 181(1):423– 431, October 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306001123.

Shi:2006:NCS

[SS06c] Zhen-Jun Shi and Jie Shen. A new class of supermemory gradient methods. Applied Mathematics and Computation, 183(2):748–760, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006096.

Shi:2006:SSE

[SS06d] Zhen-Jun Shi and Jie Shen. On step-size estimation of line search methods. Applied Mathematics and Computation, 173(1):360–371, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003565.

ShuJin:2006:PAI

[SS06e] Li ShuJin and Li ShengHong. Pricing American interest rate option on zerocoupon bond numerically. Applied Mathematics and Computation, 175(1):834– 850, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500648X.

Sun:2006:FOH

[SS06f] Y. G. Sun and S. H.

Saker. Forced oscillation of higher-order nonlinear differential equations. Applied Mathematics and Computation, 173(2):1219–1226, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500425X.

Sun:2006:PPS

[SS06g] Y. G. Sun and S. H. Saker.

Positive periodic solutions of discrete three-level foodchain model of Holling type II. Applied Mathematics and Computation, 180(1):353–365, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305011306.

Sang:2007:SSC

[SS07a] Yanbin Sang and Hua Su. Several sufficient conditions of solvability for a nonlinear higher order three-point boundary value problem on time scales. Applied Mathematics and Computation, 190(1):566– 575, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000902.

Shanmugam:2007:DST

[SS07b] T. N. Shanmugam and A. Singaravelu. Differential sandwich theorems for some subclasses of analytic functions associated with linear operator. Applied Mathematics and Computation, 187(1):445–454, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011957.

Siyyam:2007:FSP

[SS07c] Hani I. Siyyam and Mohamad A. Samarah. Finding special points using matrix-free predictor– corrector methods. Applied Mathematics and Computation, 185(1):554–563, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008903.

Stankewitz:2007:SJS

[SS07d] Rich Stankewitz and Hiroki Sumi. Structure of Julia sets of polynomial semigroups with bounded finite postcritical set. Applied Mathematics and Computation, 187(1):479–488, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011994.

Sun:2007:FOH

[SS07e] Y. G. Sun and S. H.

Saker. Forced oscillation of higher order nonlinear difference equations. Applied Mathematics and Computation, 187(2):868–872, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012185.

Sekhar:2008:SST

[SS08a] T. Raja Sekhar and V. D.

Sharma. Similarity solutions for three dimensional Euler equations using Lie group analysis. Applied Mathematics and Computation, 196(1):147–157, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006546.

Sharma:2008:HPD

[SS08b] Kapil K. Sharma and Paramjeet Singh. Hyperbolic partial differentialdifference equation in the mathematical modeling of neuronal firing and its numerical solution. Applied Mathematics and Computation, 201(1–2):229–238, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011903.

Sierra:2008:GTC

[SS08c] Cesar Augusto Gomez Sierra and Alvaro H. Salas. The generalized tanh-coth method to special types of the fifth-order KdV equation. Applied Mathematics and Computation, 203(2):873–880, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004165.

Soheili:2008:MMM

[SS08d] Ali R. Soheili and Soheil Salahshour. Moving mesh method with local time step refinement for blow-up problems. Applied Mathematics and Computation, 195(1):76–85, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700519X.

Sacks:2009:CMS

[SS09a] Paul Sacks and Jaemin Shin. Computational methods for some inverse scattering problems. Applied Mathematics and Computation, 207(1):111–123, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007315.

Sanches:2009:PTD

[SS09b] Carlos Alberto Alonso Sanches and Nei Yoshihiro Soma. A polynomialtime DNA computing solution for the Bin-Packing Problem. Applied Mathematics and Computation, 215(6):2055–2062, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007139.

Sang:2009:SSC

[SS09c] Yanbin Sang and Hua Su. Several sufficient conditions of solvability for a nonlinear higher-order three-point boundary value problem with all derivatives. Applied Mathematics and Computation, 215(3):1154–1160, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006286.

Shen:2009:NVS

[SS09d] Shoufeng Shen and Junquang Song. Nonlinear variable separation solutions of some mKdV-type equations and the unified construction method. Applied Mathematics and Computation, 212(1):251– 256, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001416.

Samidurai:2009:GAS

[SSA09] R. Samidurai, R. Sakthivel, and S. M. Anthoni. Global asymptotic stability of BAM neural networks with mixed delays and impulses. Applied Mathematics and Computation, 212(1):113–119, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001246.

Safaei:2006:EGA

[SSB06] N. Safaei, S. J. Sadjadi, and M. Babakhani. An efficient genetic algorithm for determining the optimal price discrimination. Applied Mathematics and Computation, 181(2):1693–1702, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003006.

Saleem:2006:ESS

[SSEA06] M. Saleem, A. H. Sadiyal, Prasetyo Edi, and P. R. Arora. Evolutionarily stable strategies for defensive switching. Applied Mathematics and Computation, 177(2):697–713, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009501.

Shelokar:2007:PSA

[SSJK07] P. S. Shelokar, Patrick Siarry, V. K. Jayaraman, and B. D. Kulkarni. Particle swarm and ant colony algorithms hybridized for improved continuous optimization. Applied Mathematics and Computation, 188(1):129– 142, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013294.

Sunil:2006:EMF

[SSK06] Sunil, Anupama Sharma, and Pardeep Kumar. Effect of magnetic field dependent viscosity and rotation on ferroconvection in the presence of dust particles. Applied Mathematics and Computation, 182(1):82–88, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003158.

Stanimirovic:2009:CMR

[SSMI09] Stefan Stanimirović, Predrag Stanimirović, Marko Miladinović, and Aleksandar Ilić. Catalan matrix and related combinatorial identities. Applied Mathematics and Computation, 215(2):796–805, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005670.

Srivastava:2005:UPG

[SSR05] H. M. Srivastava, R. K.

Saxena, and Chena Ram. A unified presentation of the Gamma-type functions occurring in diffraction theory and associated probability distributions. Applied Mathematics and Computation, 162(2):931–947, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001407.

Sunil:2006:ERF

[SSS06] Sunil, Anu Sharma, and R. G. Shandil. Effect of rotation on a ferromagnetic fluid heated and soluted from below in the presence dust particles. Applied Mathematics and Computation, 177(2):614–628, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009422.

Sivaloganathan:2005:VAM

[SSTD05a] S. Sivaloganathan, M. Stastna, G. Tenti, and J. M. Drake. A viscoelastic approach to the modelling of hydrocephalus. Applied Mathematics and Computation, 163(3):1097–1107, April 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002826.

Sivaloganathan:2005:VMB

[SSTD05b] S. Sivaloganathan, M. Stastna, G. Tenti, and J. M. Drake. A viscoelastic model of the brain parenchyma with pulsatile ventricular pressure. Applied Mathematics and Computation, 165(3):687–698, June 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004114.

Saucez:2009:MIM

[SSW09] P. Saucez, L. Some, and A. Vande Wouwer. Matlab implementation of a moving grid method based on the equidistribution principle. Applied Mathematics and Computation, 215(5):1821–1829, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006821.

Salkuyeh:2005:BVA

[ST05a] D. Khojasteh Salkuyeh and F. Toutounian. A block version algorithm to approximate inverse factors. Applied Mathematics and Computation, 162(3):1499–1509, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002036.

Simakhina:2005:CSR

[ST05b] Svetlana V. Simakhina and Charles Tier. Computing the stability regions of Hill’s equation. Applied Mathematics and Computation, 162(2):639–660, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001201.

Stoyanov:2005:MCS

[ST05c] Jordan Stoyanov and Leonid Tolmatz. Method for constructing Stieltjes classes for M-indeterminate probability distributions. Applied Mathematics and Computation, 165(3):669– 685, June 27, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304004102.

Salkuyeh:2006:NAS

[ST06a] Davod Khojasteh Salkuyeh and Faezeh Toutounian.

New approaches for solving large Sylvester equations. Applied Mathematics and Computation, 173(1):9–18, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002985.

Salkuyeh:2006:NACb

[ST06b] Davod Khojasteh Salkuyeh and Faezeh Toutounian. Numerical accuracy of a certain class of iterative methods for solving linear system. Applied Mathematics and Computation, 176(2):727–738, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008738.

Singh:2006:RTT

[ST06c] Jaswant Singh and S. K. Tomar. Reflection and transmission of transverse waves at a plane interface between two different porous elastic solid half-spaces. Applied Mathematics and Computation, 176(1):364–378, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008118.

Shampine:2007:MAS

[ST07] L. F. Shampine and S. Thompson. Moving averages of solutions of ODEs. Applied Mathematics and Computation, 193(1):175–182, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003840.

Sophocleous:2008:DIQ

[ST08] C. Sophocleous and R. Tracinà.

Differential invariants for quasi-linear and semilinear wave-type equations. Applied Mathematics and Computation, 202(1):216–228, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000854.

Srivastava:2009:FCI

[ST09] H. M. Srivastava and Živorad Tomovski. Fractional calculus with an integral operator containing a generalized Mittag-Leffler function in the kernel. Applied Mathematics and Computation, 211 (1):198–210, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030900071X.

Stamova:2006:LMB

[Sta06a] Ivanka M. Stamova. Lyapunov method for boundedness of solutions of nonlinear impulsive functional differential equations. Applied Mathematics and Computation, 177(2):714–719, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009513.

Stanislavsky:2006:NRF

[Sta06b] Aleksander A. Stanislavsky. Nonlinear reaction with fractional dynamics. Applied Mathematics and Computation, 174(2):1122– 1134, March 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305005369.

Stannett:2006:CH

[Sta06c]  Mike Stannett.  The

case for hypercomputation. Applied Mathematics and Computation, 178(1):8–24, July 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008301.

Stevic:2008:NOL

[Ste08] Stevo Stević. On a new operator from the logarithmic Bloch space to the Bloch-type space on the unit ball. Applied Mathematics and Computation, 206(1):313–320, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006474.

Stevic:2009:BTF

[Ste09a] Stevo Stević. Bloch-type functions with Hadamard gaps. Applied Mathematics and Computation, 208(2):416–422, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009090.

Stevic:2009:GSD

[Ste09b] Stevo Stević. Global stability of a difference equation with maximum. Applied Mathematics and Computation, 210(2):525–529, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000320.

Stevic:2009:NEN

[Ste09c] Stevo Stević. Norm and essential norm of composition followed by differentiation from α-Bloch spaces to H_{μ}^{∞}. Applied Mathematics and Computation, 207(1):225–229, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308007820.

Stevic:2009:NWC

[Ste09d]  Stevo Stević.  Norm

of weighted composition operators from α-Bloch spaces to weighted-type spaces. Applied Mathematics and Computation, 215(2):818–820, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309005694.

Stevic:2009:NSO

[Ste09e]  Stevo Stević.  Norms

of some operators on the Bergman and the Hardy space in the unit polydisk and the unit ball. Applied Mathematics and Computation, 215(6):2199–2205, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007309.

Stevic:2009:NBT

[Ste09f] Stevo Stević. On new Bloch-type spaces. Applied Mathematics and Computation, 215(2):841–849, September 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309005748.

Stevic:2009:WCO

[Ste09g] Stevo Stević. Weighted composition operators between Fock-type spaces in C^{N}. Applied Mathematics and Computation, 215(7):2750–2760, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008145.

Stevic:2009:WCOa

[Ste09h] Stevo Stević. Weighted composition operators from weighted Bergman spaces to weighted-type spaces on the unit ball. Applied Mathematics and Computation, 212(2):499–504, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001829.

Stevic:2009:WDC

[Ste09i] Stevo Stević. Weighted differentiation composition operators from mixednorm spaces to weightedtype spaces. Applied Mathematics and Computation, 211(1):222–233, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000745.

Shampine:2005:NNS

[STKB05] L. F. Shampine, S. Thompson, J. A. Kierzenka, and G. D. Byrne. Non-negative solutions of ODEs. Applied Mathematics and Computation, 170(1):556–569, November 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304009683.

Surla:2009:RLR

[STU09] Katarina Surla, Ljiljana Teofanov, and Zorica Uzelac. A robust layerresolving spline collocation method for a convection– diffusion problem. Applied Mathematics and Computation, 208(1):76–89, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008497.

Stanimirovic:2009:EFA

[STV09] Predrag S. Stanimirović, Milan B. Tasić, and Ky M. Vu. Extensions of Faddeev’s algorithms to polynomial matrices. Applied Mathematics and Computation, 214(1):246–258, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002999.

Su:2007:EDG

[Su07a] Ching-Liang Su. Edge distance and gray level extraction and orientation invariant transform for Chinese seal recognition. Applied Mathematics and Computation, 193(2):325–334, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004018.

Su:2007:FRU

[Su07b]  Ching-Liang Su.  Face

recognition by using characteristic matrix and ring rotation invariant transform. Applied Mathematics and Computation, 189 (1):116–125, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306015864.

Su:2007:FEF

[Su07c] Ching-Liang Su. Finger extraction, finger image automatic registration, and finger identification by image phase matching. Applied Mathematics and Computation, 188(1):912–923, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014299.

Su:2007:FSE

[Su07d] Ching-Liang Su. Finger shape expansion and recognition by wavelet transform. Applied Mathematics and Computation, 190(2):1583–1592, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002032.

Su:2008:FEE

[Su08a]  Ching-Liang Su.  Fin-

ger edge extraction and finger shape comparisons for person’s identification by using index, middle, and ring fingers. Applied Mathematics and Computation, 197(2):643–653, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008211.

Su:2008:LMD

[Su08b]  Ching-Liang Su.  Lo-

cating MRI discrepancy by orientation invariant method. Applied Mathematics and Computation, 195(2):412–419, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005620.

Su:2008:PSO

[Su08c]  Hua Su.  Positive so-

lutions for n-order mpoint p-Laplacian operator singular boundary value problems. Applied Mathematics and Computation, 199(1):122–132, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009782.

Su:2009:DFF

[Su09] Ninghu Su. N-dimensional fractional Fokker–Planck equation and its solutions for anomalous radial two-phase flow in porous media. Applied Mathematics and Computation, 213(2):506–515, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900263X.

Subasi:2007:RCK

[Sub07] Murat Subaşı. Reply to the Comment of E. Kahya on “An improvement on Fibonacci search method in optimization theory” by Murat Subaşı, Necmettin Yildirim, Bünyamin Yildiz, [An improvement on Fibonacci search method in optimization theory, Applied Mathematics and Computation 147 (2004) 893–901]. Applied Mathematics and Computation, 188(1):1059–1060, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013464.

See [SYY04].

Sumi:2007:RDP

[Sum07] Hiroki Sumi. Random dynamics of polynomials and devil’s-staircase-like functions in the complex plane. Applied Mathematics and Computation, 187(1):489– 500, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012008.

Sun:2005:SKG

[Sun05a] Qiang Sun. Solving the Klein–Gordon equation by means of the homotopy analysis method. Applied Mathematics and Computation, 169(1):355–365, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007799.

Sun:2005:ONO

[Sun05b] Yuan Gong Sun. Oscillation and non-oscillation for second-order linear difference equations. Applied Mathematics and Computation, 170(2):1095–1103, November 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305000500.

Sun:2006:AED

[Sun06a] Jiachang Sun. Approximate eigen-decomposition preconditioners for solving numerical PDE problems. Applied Mathematics and Computation, 172(2):772–787, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304009117.

Sun:2006:OSO

[Sun06b] Yuan Gong Sun. Oscillation of second order functional differential equations with damping. Applied Mathematics and Computation, 178(2):519–526, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009999.

Sun:2007:NCL

[Sun07a]  Da-Zhi Sun.  A note

on Chang–Lai’s modular square algorithm based on the generalized Chinese remainder theorem. Applied Mathematics and Computation, 188(1):411–416, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306013762.

Sun:2007:RPP

[Sun07b]  Yuan Gong Sun.  Re-

mark on the paper of Park. Applied Mathematics and Computation, 187(2):1250–1255, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012616.

Sun:2008:CVP

[Sun08] Yajuan Sun. A class of volume-preserving numerical algorithms. Applied Mathematics and Computation, 206(2):841–852, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007492.

Sommariva:2006:MCG

[SV06] Alvise Sommariva and Marco Vianello. Meshless cubature by Green’s formula. Applied Mathematics and Computation, 183(2):1098–1107, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006448.

Suvorova:2007:NHT

[SV07] J. V. Suvorova and I. V.

Viktorova. The nonlinear hereditary-type stress– strain relations for metals with temperature effects. Applied Mathematics and Computation, 189(1):844– 851, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016638.

Siddiqui:2009:MMP

[SVMG09] S. U. Siddiqui, N. K. Verma, Shailesh Mishra, and R. S. Gupta. Mathematical modelling of pulsatile flow of Casson’s fluid in arterial stenosis. Applied Mathematics and Computation, 210(1):1–10, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006704.

Shen:2005:NPT

[SW05a] Peiping Shen and Yanjun Wang. A new pruning test for finding all global minimizers of nonsmooth functions. Applied Mathematics and Computation, 168(2):739–755, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304006356.

Shim:2005:WIB

[SW05b] Kyungah Shim and Sungsik Woo. Weakness in IDbased one round authenticated tripartite multiplekey agreement protocol with pairings. Applied Mathematics and Computation, 166(3):523–530, July 26, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004552.

Shen:2006:GOS

[SW06a] Pei-Ping Shen and Chun-Feng Wang. Global optimization for sum of linear ratios problem with coefficients. Applied Mathematics and Computation, 176(1):219–229, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007861.

Sun:2006:NPA

[SW06b] Huan Sun and Yimin Wei.

A note on the PageRank algorithm. Applied Mathematics and Computation, 179(2):799–806, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030501091X.

Sun:2006:NAR

[SW06c] Huijun Sun and Jianjun Wu. A new approach for ranking fuzzy numbers based on fuzzy simulation analysis method. Applied Mathematics and Computation, 174(1):755–767, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005084.

Shi:2009:LOA

[SW09a] Dong-Yang Shi and Xiao-Ling Wang. A low order anisotropic nonconforming characteristic finite element method for a convection-dominated transport problem. Applied Mathematics and Computation, 213(2):411–418, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002513.

Shi:2009:AAO

[SW09b] Xue Rong Shi and Zuo Lei Wang. Adaptive addedorder anti-synchronization of chaotic systems with fully unknown parameters. Applied Mathematics and Computation, 215(5):1711–1717, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006717.

Stevic:2009:DCO

[SW09c] Stevo Stević and Elke Wolf. Differences of composition operators between weighted-type spaces of holomorphic functions on the unit ball of C^{N}. Applied Mathematics and Computation, 215(5):1752– 1760, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309006754.

Sun:2009:DSB

[SW09d] Zhi-Zhong Sun and Xiao-Nan Wu. A difference scheme for Burgers equation in an unbounded domain. Applied Mathematics and Computation, 209(2):285–304, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009648.

Shi:2005:CCM

[SWC05] Jinlin Shi, Aizhi Weng, and Jiyang Chen. The calculation for characteristic multiplier of Hill’s equation in case q(t) < 0. Applied Mathematics and Computation, 167(2):1130–1149, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304005545.

Shengwei:2007:NAW

[SWH07] Yao Shengwei, Zengxin Wei, and Hai Huang. A note about WYL’s conjugate gradient method and its applications. Applied Mathematics and Computation, 191(2):381–388, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002652.

Saha:2006:GPM

[SWK^{+}06] Nilanjan Saha, Layne T. Watson, Karen Kafadar, Alexey Onufriev, Naren Ramakrishnan, Cecilia Vasquez-Robinet, and Jonathan Watkinson. A general probabilistic model of the PCR process. Applied Mathematics and Computation, 182(1):232–243, November 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306003328.

Wang:2007:SIE

[sWMX07] La sheng Wang, Changlin Mei, and Hong Xue. The semi-implicit Euler method for stochastic differential delay equation with jumps. Applied Mathematics and Computation, 192(2):567–578, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003608.

Su:2006:EPS

[SWX06a] Hua Su, Zhongli Wei, and Fuyi Xu. The existence of positive solutions for nonlinear singular boundary value system with p-Laplacian. Applied Mathematics and Computation, 181(2):826–836, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001639.

Sun:2006:DDR

[SWX06b] Yuan Gong Sun, Long Wang, and Guangming Xie. Delay-dependent robust stability and stabilization for discrete-time switched systems with mode-dependent time-varying delays. Applied Mathematics and Computation, 180(2):428–435, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600004X.

Wang:2007:CNS

[sWX07a] La sheng Wang and Hong Xue. Convergence of numerical solutions to stochastic differential delay equations with Poisson jump and Markovian switching. Applied Mathematics and Computation, 188(2):1161–1172, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014603.

Sun:2007:DDR

[SWX07b] Yuan Gong Sun, Long Wang, and Guangming Xie. Delay-dependent robust stability and H_{∞} control for uncertain discretetime switched systems with mode-dependent time delays. Applied Mathematics and Computation, 187(2):1228–1237, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012598.

Su:2007:PSO

[SWZL07] Hua Su, Zhongli Wei, Xiaoyan Zhang, and Jian Liu. Positive solutions of n-order and m-order multipoint singular boundary value system. Applied Mathematics and Computation, 188(2):1234–1243, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014676.

Shen:2006:SNS

[SX06] Jianwei Shen and Wei Xu. Smooth and nonsmooth travelling wave solutions of generalized Degasperis–Procesi equation. Applied Mathematics and Computation, 182(2):1418–1429, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004693.

Shen:2005:BMT

[SXJ05] Jianwei Shen, Wei Xu, and Yanfei Jin. Bifurcation method and traveling wave solution to Whitham– Broer–Kaup equation. Applied Mathematics and Computation, 171(2):677– 702, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shengguo:2009:NFO

[SXL09a] Li Shengguo, Liao Xiangke, and Cheng Lizhi. A new fourth-order iterative method for finding multiple roots of nonlinear equations. Applied Mathematics and Computation, 215(3):1288–1292, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006444.

Song:2009:LGB

[SXL09b] Zigen Song, Jian Xu, and Qunhong Li. Local and global bifurcations in an SIRS epidemic model. Applied Mathematics and Computation, 214(2):534–547, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900352X.

Shen:2005:TWS

[SXX05] Jianwei Shen, Wei Xu, and Yong Xu. Travelling wave solutions in the generalized Hirota– Satsuma coupled KdV system. Applied Mathematics and Computation, 161(2):365–383, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012918.

Sarikaya:2007:WSC

[SY07a] Mehmet Zeki Sarikaya and Hüseyin Yildirim. On the weak solutions of the compound Bessel ultrahyperbolic equation. Applied Mathematics and Computation, 189(1):910– 917, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016705.

Shi:2007:CMV

[SY07b] Hang Shi and Yu Yang.

A computational model of visual attention based on saliency maps. Applied Mathematics and Computation, 188(2):1671–1677, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015219.

Shouzhi:2007:TDR

[SY07c] Yang Shouzhi and Li Youfa.

Two-direction refinable functions and two-direction wavelets with dilation factor m. Applied Mathematics and Computation, 188(2):1908–1920, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015591.

Song:2008:EUM

[SY08a] Huailing Song and Yirang Yuan. The expanded upwind-mixed multi-step method for the miscible displacement problem in three dimensions. Applied Mathematics and Computation, 195(1):100–109, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005231.

Sun:2008:PEE

[SY08b] Tongjun Sun and Danping Yang. A priori error estimates for interior penalty discontinuous Galerkin method applied to nonlinear Sobolev equations. Applied Mathematics and Computation, 200(1):147–159, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010958.

Sarikaya:2009:CBD

[SY09] Mehmet Zeki Sarikaya and Hüseyin Yildirim. On the B-convolutions of the Bessel diamond kernel of Riesz. Applied Mathematics and Computation, 208(1):18–22, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008539.

Syam:2005:ADM

[Sya05a] Muhammed I. Syam. Adomian decomposition method for approximating the solution of the Korteweg–de Vries equation. Applied Mathematics and Computation, 162(3):1465–1473, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002000.

Syam:2005:NOE

[Sya05b] Muhammed I. Syam. Nonlinear optimization exclusion tests for finding all solutions of nonlinear equations. Applied Mathematics and Computation, 170(2):1104–1116, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000512.

Sheng:2007:USP

[SYH07] Zhiqiang Sheng, Guangwei Yuan, and Xudeng Hang. Unconditional stability of parallel difference schemes with second order accuracy for parabolic equation. Applied Mathematics and Computation, 184(2):1015–1031, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007909.

Sun:2007:BOP

[SYL07a] Chao Sun, Jing-Yang Yang, and Sheng-Hong Li. On barrier option pricing in binomial market with transaction costs. Applied Mathematics and Computation, 189(2):1505–1516, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017383.

Sun:2007:ROP

[SYL07b] Chao Sun, Jing-Yang Yang, and Sheng-Hong Li. On reset option pricing in binomial market with both fixed and proportional transaction costs. Applied Mathematics and Computation, 193(1):143–153, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307003815.

Samimi:2006:AOC

[SYT06] Ahmad Jafari Samimi, Sohrabali Yousefi, and Amir Mansor Tehranchian. An application of optimal control theory in economics: optimal fiscal and monetary policies in Iran. Applied Mathematics and Computation, 172(1):198–209, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001980.

Srivastava:2008:SSM

[SYX08] H. M. Srivastava, Ding-Gong Yang, and N-Eng Xu. Some subclasses of meromorphically multivalent functions associated with a linear operator. Applied Mathematics and Computation, 195(1):11–23, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005127.

Subasi:2004:IFS

[SYY04] Murat Subasi, Necmettin Yildirim, and Bünyamin Yildiz. An improvement on Fibonacci search method in optimization theory. Applied Mathematics and Computation, 147(3):893– 901, January 16, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See comment [Kah06a] and reply [Sub07].

Shidfar:2005:NTB

[SZ05] Abdollah Shidfar and Ali Zakeri. A numerical technique for backward inverse heat conduction problems in one-dimensional space. Applied Mathematics and Computation, 171(2):1016– 1024, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shu:2006:SCG

[SZ06a] Ji Shu and Jian Zhang. Sharp criterion of global existence for a class of nonlinear Schrödinger equation with critical exponent. Applied Mathematics and Computation, 182(2):1482–1487, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004759.

Song:2006:NES

[SZ06b] Lina Song and Hongqing Zhang. New exact solutions of the Broer–Kaup equations in (2 + 1)- dimensional spaces. Applied Mathematics and Computation, 180(2):664– 675, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306000555.

Shahzad:2007:SCI

[SZ07a] Naseer Shahzad and Habtu Zegeye. Strong convergence of an implicit iteration process for a finite family of generalized asymptotically quasi-nonexpansive maps. Applied Mathematics and Computation, 189(2):1058–1065, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016821.

Song:2007:NCS

[SZ07b] Lina Song and Hongqing Zhang. New complexiton solutions of the nonlinear evolution equations using a generalized rational expansion method with symbolic computation. Applied Mathematics and Computation, 190(1):974–986, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001269.

Song:2007:NES

[SZ07c] Lina Song and Hongqing Zhang. New exact solutions for the Konopelchenko– Dubrovsky equation using an extended Riccati equation rational expansion method and symbolic computation. Applied Mathematics and Computation, 187(2):1373–1388, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012756.

Song:2007:NKV

[SZ07d] Lina Song and Hongqing Zhang. A new Korteweg– de Vries equation-based sub-equation method and its application to the (2 + 1)-dimensional Korteweg– de Vries equation. Applied Mathematics and Computation, 187(2):1368–1372, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012744.

Song:2007:NVC

[SZ07e] Lina Song and Hongqing Zhang. A new variable coefficient Korteweg– de Vries equation-based sub-equation method and its application to the (3 + 1)-dimensional potential-YTSF equation. Applied Mathematics and Computation, 189(1):560–566, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016298.

Song:2008:AEH

[SZ08a] Lina Song and Hongqing Zhang. Application of the extended homotopy perturbation method to a kind of nonlinear evolution equations. Applied Mathematics and Computation, 197(1):87–95, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007576.

Su:2008:SHO

[SZ08b] Hua Su and Mingchuan Zhang. Solutions for higher-order dynamic equations on time scales. Applied Mathematics and Computation, 200(1):413– 428, June 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307011460.

Sun:2008:NNI

[SZ08c] Ye-Peng Sun and Hong-Qing Zhao. New nonisospectral integrable couplings of the AKNS system. Applied Mathematics and Computation, 203(1):163–170, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002221.

Shen:2005:PCM

[SZF05] Yuming Shen, Jinxi Zhao, and Hongjun Fan. Properties and computations of matrix pseudospectra. Applied Mathematics and Computation, 161(2):385–393, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030301292X.

Shi:2005:DLB

[SZL05] Chenggen Shi, Guangquan Zhang, and Jie Lu. On the definition of linear bilevel programming solution. Applied Mathematics and Computation, 160(1):169–176, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303011147.

Shi:2007:TBA

[SZL^{+}07] Chenggen Shi, Hong Zhou, Jie Lu, Guangquan Zhang, and Zhongwei Zhang. The Kth-best approach for linear bilevel multifollower programming with partial shared variables among followers. Applied Mathematics and Computation, 188(2):1686–1698, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015232.

Sun:2005:WIW

[SZS05] Da-Zhi Sun, Ji-Dong Zhong, and Yu Sun. Weakness and improvement on Wang–Li–Tie’s userfriendly remote authentication scheme. Applied Mathematics and Computation, 170(2):1185–1193, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000573.

Song:2006:MTW

[SZXL06] Y. Q. Song, Y. C. Zhang, H. Y. Xu, and B. H. Lu. Magneto-thermoviscoelastic wave propagation at the interface between two micropolar viscoelastic media. Applied Mathematics and Computation, 176(2):785–802, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008805.

Tiryaki:2005:FSS

[TA05a] Fatma Tiryaki and Mehmet Ahlatcioglu. Fuzzy stock selection using a new fuzzy ranking and weighting algorithm. Applied Mathematics and Computation, 170(1):144–157, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008781.

Torun:2005:ESP

[TA05b] Gülten Torun and İbrahim Ethem Anar. The electromagnetic scattering problem: the case of TE polarized waves. Applied Mathematics and Computation, 169(1):339–354, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304007787.

Torun:2006:PFF

[TA06] Gülten Torun and Ülkü (Dinlemez) Ateş. Properties of the far field operator in the inverse conductive scattering problem. Applied Mathematics and Computation, 175(2):1503–1514, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006934.

Tadi:2008:IMS

[Tad08] M. Tadi. An iterative method for the solution of ill-posed parabolic systems. Applied Mathematics and Computation, 201(1–2):843–851, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307012313.

Takahashi:2005:AMP

[Tak05] Daisuke Takahashi. An algorithm for multipleprecision floating-point multiplication. Applied Mathematics and Computation, 166(2):291–298, July 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004370.

Takano:2007:LTS

[Tak07a]  Katsuo Takano.  On

Laplace transforms of some probability densities. Applied Mathematics and Computation, 187(1):501– 506, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012021.

Tsitsas:2007:RAI

[TAK07b] Nikolaos L. Tsitsas, Emmanouil G. Alivizatos, and Grigorios H. Kalogeropoulos. A recursive algorithm for the inversion of matrices with circulant blocks. Applied Mathematics and Computation, 188(1):877– 894, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014275.

Tang:2005:UCW

[Tan05a] Lei Tang. Upwind and central WENO schemes. Applied Mathematics and Computation, 166(2):434– 448, July 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304004485.

Tang:2005:NSP

[Tan05b]  Yanbin Tang.  Numer-

ical simulations of periodic travelling waves to a generalized Ginzburg– Landau equation. Applied Mathematics and Computation, 165(1):155–161, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003807.

Tanaka:2006:EAI

[Tan06a]  Yasuhito Tanaka.  On

the equivalence of the Arrow impossibility theorem and the Brouwer fixed point theorem. Applied Mathematics and Computation, 172(2):1303–1314, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002936.

Tanaka:2006:TEA

[Tan06b]  Yasuhito Tanaka.  On

the topological equivalence of the Arrow impossibility theorem and Amartya Sen’s liberal paradox. Applied Mathematics and Computation, 181(2):1490–1498, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002785.

Tanaka:2006:TAA

[Tan06c]  Yasuhito Tanaka.  A

topological approach to the Arrow impossibility theorem when individual preferences are weak orders. Applied Mathematics and Computation, 174(2):961–981, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005242.

Tanaka:2006:TPE

[Tan06d] Yasuhito Tanaka. A topological proof of Eliaz’s unified theorem of social choice theory. Applied Mathematics and Computation, 176(1):83– 90, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305007733.

Tanaka:2007:EHGa

[Tan07a] Yasuhito Tanaka. Equivalence of the HEX game theorem and the Arrow impossibility theorem. Applied Mathematics and Computation, 186(1):509–515, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009611.

Tanaka:2007:GST

[Tan07b] Yasuhito Tanaka. The Gibbard–Satterthwaite theorem of social choice theory in an infinite society and LPO (limited principle of omniscience). Applied Mathematics and Computation, 193(2):475–481, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004304.

Tanaka:2007:EHGb

[Tan07c] Yasuhito Tanaka. On the equivalence of the HEX game theorem and the Duggan–Schwartz theorem for strategy-proof social choice correspondences. Applied Mathematics and Computation, 188(1):303– 313, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013531.

Tanaka:2007:TTC

[Tan07d] Yasuhito Tanaka. Type two computability of social choice functions and the Gibbard–Satterthwaite theorem in an infinite society. Applied Mathematics and Computation, 192(1):168–174, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003086.

Tanaka:2008:CBS

[Tan08a]  Yasuhito Tanaka.  On

the computability of binary social choice rules in an infinite society and the halting problem. Applied Mathematics and Computation, 197(2):598–603, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008053.

Tanaka:2008:UUE

[Tan08b] Yasuhito Tanaka. Undecidability of Uzawa equivalence theorem and LLPO (Lesser limited principle of omniscience). Applied Mathematics and Computation, 201(1–2):378–383, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307012155.

Toloo:2008:MOP

[TARm08] Mehdi Toloo, Nazila Aghayi, and Mohsen Rostamymalkhalifeh. Measuring overall profit efficiency with interval data. Applied Mathematics and Computation, 201(1–2):640–649, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000106.

Tawhid:2006:SSM

[Taw06]  M. A. Tawhid.  On

strictly semi-monotone (semi-monotone) properties in nonsmooth functions. Applied Mathematics and Computation, 175(2):1609–1618, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007186.

Tawhid:2007:MIL

[Taw07] M. A. Tawhid. On minimizing the implicit Lagrangian for nonlinear complementarity problems under H-differentiability. Applied Mathematics and Computation, 185(1):484– 491, February 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306008836.

Tascan:2009:ASD

[TB09a] Filiz Taşcan and Ahmet Bekir. Analytic solutions of the (2 + 1)-dimensional nonlinear evolution equations using the sine–cosine method. Applied Mathematics and Computation, 215(8):3134–3139, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030900825X.

Tascan:2009:TWS

[TB09b] Filiz Taşcan and Ahmet Bekir. Travelling wave solutions of the Cahn–Allen equation by using first integral method. Applied Mathematics and Computation, 207(1):279–282, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800787X.

Tsaur:2005:EAS

[TC05] Woei-Jiunn Tsaur and Chih-Ho Chou. Efficient algorithms for speeding up the computations of elliptic curve cryptosystems. Applied Mathematics and Computation, 168(2):1045–1064, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304006629.

Tlelo-Cuautle:2007:DCC

[TCDV07] E. Tlelo-Cuautle and M. A. Duarte-Villaseñor. Designing Chua’s circuit from the behavioral to the transistor level of abstraction. Applied Mathematics and Computation, 184(2):715–720, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007478.

Tlelo-Cuautle:2007:SCC

[TCMP07] E. Tlelo-Cuautle and Jesús M.

Muñoz-Pacheco. Simulation of Chua’s circuit by automatic control of step-size. Applied Mathematics and Computation, 190(2):1526–1533, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001889.

Tlelo-Cuautle:2007:FSS

[TCMPMC07] E. Tlelo-Cuautle, J. M.

Muñoz-Pacheco, and J. Martínez-Carballido. Frequency scaling simulation of Chua’s circuit by automatic determination and control of step-size. Applied Mathematics and Computation, 194(2):486–491, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004997.

Tang:2008:NLA

[TCW08] Jianliang Tang, Wen-Sheng Chen, and Jie Wang. A novel linear algorithm for P5P problem. Applied Mathematics and Computation, 205(2):628–634, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800341X.

Tunga:2005:FHD

[TD05] M. Alper Tunga and Metin Demiralp. A factorized high dimensional model representation on the nodes of a finite hyperprismatic regular grid. Applied Mathematics and Computation, 164(3):865–883, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003534.

Tavakoli:2006:NSG

[TD06] Rohallah Tavakoli and Parviz Davami. New stable group explicit finite difference method for solution of diffusion equation. Applied Mathematics and Computation, 181(2):1379–1386, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002566.

Tavakoli:2007:PSG

[TD07a] Rohallah Tavakoli and Parviz Davami. 2D parallel and stable group explicit finite difference method for solution of diffusion equation. Applied Mathematics and Computation, 188(2):1184–1192, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014627.

Tavakoli:2007:NPG

[TD07b] Rohallah Tavakoli and Parviz Davami. A new parallel Gauss–Seidel method based on alternating group explicit method and domain decomposition method. Applied Mathematics and Computation, 188(1):713– 719, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014093.

Tian:2009:IMP

[TD09] Xia Tian and Hua Dai. Inverse mode problems for the finite element model of a vibrating rod. Applied Mathematics and Computation, 214(2):479–486, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309003439.

Tremblay:2007:UFD

[TF07] R. Tremblay and B.-J. Fugère. The use of fractional derivatives to expand analytical functions in terms of quadratic functions with applications to special functions. Applied Mathematics and Computation, 187(1):507–529, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012033.

Tahmasbi:2008:NSL

[TF08] Ali Tahmasbi and Omid S. Fard. Numerical solution of linear Volterra integral equations system of the second kind. Applied Mathematics and Computation, 201(1–2):547–552, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307012295.

Tian:2007:NDM

[TFX07] Hongjiong Tian, Liqiang Fan, and Jiaxiang Xiang. Numerical dissipativity of multistep methods for delay differential equations. Applied Mathematics and Computation, 188(1):934– 941, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014408.

Tao:2006:ASM

[TG06a] Youshan Tao and Qian Guo. Analysis and simulation of a model for intracellular drug accumulation in tumors. Applied Mathematics and Computation, 176(2):577–593, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500857X.

Taskin:2006:EAR

[TG06b] Alev Taskin and Ali Fuat Guneri. Economic analysis of risky projects by ANNs. Applied Mathematics and Computation, 175(1):171– 181, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005990.

Tian:2007:IMC

[TG07] Zhaolu Tian and Chuanqing Gu. The iterative methods for centrosymmetric matrices. Applied Mathematics and Computation, 187(2):902–911, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012227.

Tian:2008:ASC

[TG08a] Hongjiong Tian and Ni Guo. Asymptotic stability, contractivity and dissipativity of one-leg θ-method for non-autonomous delay functional differential equations. Applied Mathematics and Computation, 203(1):333–342, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002579.

Tian:2008:NAL

[TG08b] Zhaolu Tian and Chuanqing Gu. A numerical algorithm for Lyapunov equations. Applied Mathematics and Computation, 202(1):44–53, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308000040.

Taheri:2007:HRA

[TGSNB07] S. H. Taheri, H. Ghazvini, J. Saberi-Nadjafi, and J. Biazar. A hybrid of the restarted Arnoldi and electromagnetism metaheuristic methods for calculating eigenvalues and eigenvectors of a nonsymmetric matrix. Applied Mathematics and Computation, 191(1):79–88, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000847.

Tang:2007:BTW

[TH07a] Shengqiang Tang and Wentao Huang. Bifurcations of travelling wave solutions for the generalized double sinh-Gordon equation. Applied Mathematics and Computation, 189(2):1774–1781, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601770X.

Tavazoei:2007:CDO

[TH07b] Mohammad Saleh Tavazoei and Mohammad Haeri. Comparison of different one-dimensional maps as chaotic search pattern in chaos optimization algorithms. Applied Mathematics and Computation, 187(2):1076–1085, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012446.

Tang:2008:BTW

[TH08] Shengqiang Tang and Wentao Huang. Bifurcations of travelling wave solutions for the K(n, −n, 2n) equations. Applied Mathematics and Computation, 203(1): 39–49, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300308000611.

Thakur:2007:WSC

[Tha07] Balwant Singh Thakur. Weak and strong convergence of composite implicit iteration process. Applied Mathematics and Computation, 190(1):965–973, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001257.

Tsai:2007:IIA

[THLC07] Jinn-Tsong Tsai, Wen-Hsien Ho, Tung-Kuan Liu, and Jyh-Horng Chou. Improved immune algorithm for global numerical optimization and jobshop scheduling problems. Applied Mathematics and Computation, 194(2):406–424, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004821.

Thukral:2006:IDF

[Thu06]  R. Thukral.  Introduc-

tion to the determinantal formulae for the Levintype algorithms. Applied Mathematics and Computation, 181(1):516–526, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001214.

Thukral:2007:DOT

[Thu07a]  R. Thukral.  Develop-

ment of the Overholt transformation for accelerating the convergence of sequences. Applied Mathematics and Computation, 189(2):1459–1466, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017322.

Thukral:2007:FDN

[Thu07b] R. Thukral. Further development of the new algorithms for accelerating the convergence of functionaltype sequences. Applied Mathematics and Computation, 186(1):749–762, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009957.

Thukral:2007:NMA

[Thu07c]  R. Thukral.  A new

method for accelerating convergence of alternating series. Applied Mathematics and Computation, 187(2):1502–1510, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012896.

Thuraisingham:2007:EPP

[Thu07d] R. A. Thuraisingham. Enhancing Poincaré plot information via sampling rates. Applied Mathematics and Computation, 186(2):1374–1378, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010629.

Thukral:2008:INT

[Thu08]  R. Thukral.  Introduc-

tion to a Newton-type method for solving nonlinear equations. Applied Mathematics and Computation, 195(2):663–668, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005942.

Tian:2009:ACM

[Tia09] Yongge Tian. Approximation and continuity of Moore–Penrose inverses of orthogonal row block matrices. Applied Mathematics and Computation, 210 (1):18–25, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308007686.

Tien:2005:RPM

[Tie05] Tzu-Li Tien. A research on the prediction of machining accuracy by the deterministic grey dynamic model DGDM(1, 1, 1). Applied Mathematics and Computation, 161(3):923–945, February 25, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304000426.

Tirmizi:2008:NPS

[TiHuI08] Ikram A. Tirmizi, Fazal i Haq, and Siraj ul Islam. Non-polynomial spline solution of singularly perturbed boundaryvalue problems. Applied Mathematics and Computation, 196(1):6–16, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006224.

Tian:2008:MPS

[TJG08] Yu Tian, Daqing Jiang, and Weigao Ge. Multiple positive solutions of periodic boundary value problems for second order impulsive differential equations. Applied Mathematics and Computation, 200(1):123–132, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010934.

Toutounian:2006:GLS

[TK06a] F. Toutounian and S. Karimi. Global least squares method (Gl-LSQR) for solving general linear systems with several right-hand sides. Applied Mathematics and Computation, 178(2):452–460, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009902.

Tsai:2006:CTL

[TK06b] Chin-Lang Tsai and Wei-Wen Kao. Constrained total least-square solution for GPS compass attitude determination. Applied Mathematics and Computation, 183(1):106–118, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600542X.

Tsai:2006:ICG

[TK06c] Chin-Lang Tsai and Wei-Wen Kao. Ill-conditioned GPS compass attitude determination using regularization approach. Applied Mathematics and Computation, 183(1):184–198, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005510.

Tavakoli:2007:IRC

[TK07] Ali Tavakoli and Asghar Kerayechian. Improvement of the rate of convergence estimates for multigrid algorithm. Applied Mathematics and Computation, 187(2):922–928, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012240.

Tabari:2008:NML

[TKA^{+}08] Mojtaba Tabari, Amin Kaboli, M. B. Aryanezhad, Kamran Shahanaghi, and Ali Siadat. A new method for location selection: a hybrid analysis. Applied Mathematics and Computation, 206(2):598–606, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003834.

Trinh:2008:FBM

[TKJ08] Hoang-Hon Trinh, Dae-Nyeon Kim, and Kang-Hyun Jo. Facet-based multiple building analysis for robot intelligence. Applied Mathematics and Computation, 205(2):537–549, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003329.

Tang:2006:BTW

[TL06] Shengqiang Tang and Ming Li. Bifurcations of travelling wave solutions in a class of generalized KdV equation. Applied Mathematics and Computation, 177(2):589–596, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009392.

Tang:2007:NMD

[TL07] Bao-Qing Tang and Xian-Fang Li. A new method for determining the solution of Riccati differential equations. Applied Mathematics and Computation, 194(2):431–440, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004924.

Tang:2008:SCV

[TL08] B.-Q. Tang and X.-F. Li.

Solution of a class of Volterra integral equations with singular and weakly singular kernels. Applied Mathematics and Computation, 199(2):406–413, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010077.

Tao:2007:HAM

[TLB07] Yuhui Tao, Heng Liang, and Fengshan Bai. A hybrid algorithm for multihomogeneous Bézout number. Applied Mathematics and Computation, 189(2):1755–1764, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030601767X.

Tao:2005:LON

[TLC05] Qing Tao, Xin Liu, and Xunxue Cui. A linear optimization neural network for associative memory. Applied Mathematics and Computation, 171(2): 1119–1128, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tsai:2005:CAE

[TLH05] Chwei-Shyong Tsai, Shu-Chen Lin, and Min-Shiang Hwang. Cryptanalysis of an authenticated encryption scheme using selfcertified public keys. Applied Mathematics and Computation, 166(1):118– 122, July 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004254.

Li:2008:PEN

[tLlWpL08] Yao tang Li, Shi liang Wu, and Da ping Liu. Positive eigenvector of nonlinear perturbations of nonsymmetric M-matrix and Newton iterative solution. Applied Mathematics and Computation, 200(1):308–320, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011101.

Tang:2009:OSS

[TLM09] Jia Tang, Sanyang Liu, and Changfeng Ma. Onestep smoothing Newton method for solving the mixed complementarity problem with a P_{0} function. Applied Mathematics and Computation, 215(6):2326–2336, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007577.

Li:2007:IAM

[tLxLlW07] Yao tang Li, Cui xia Li, and Shi liang Wu. Improving AOR method for consistent linear systems. Applied Mathematics and Computation, 186(1):379– 388, March 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300306009477.

Tasic:2005:ESV

[TM05] Bratislav Tasić and Robert M. M. Mattheij. Explicitly solving vectorial ODEs. Applied Mathematics and Computation, 164(3):913–933, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003558.

Tanriverdi:2007:GEC

[TM07] Tanfer Tanriverdi and John Bryce Mcleod. Generalization of the eigenvalues by contour integrals. Applied Mathematics and Computation, 189(2):1765–1773, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017693.

Tian:2008:ISS

[TM08] Ya Tian and Chunlai Mu. Instantaneous shrinking of the solution to a nonlinear degenerate equation with variable coefficient and convection. Applied Mathematics and Computation, 196(1):137–146, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006522.

Tavakkoli-Moghaddam:2005:SDC

[TMASA05] R. Tavakkoli-Moghaddam, M. B. Aryanezhad, N. Safaei, and A. Azaron. Solving a dynamic cell formation problem using metaheuristics. Applied Mathematics and Computation, 170(2):761–780, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000238.

Tang:2006:ECS

[TmChL06] Qiong Tang, Chuan miao Chen, and Luo hua Liu. Energy conservation and symplectic properties of continuous finite element methods for Hamiltonian systems. Applied Mathematics and Computation, 181(2):1357–1368, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002542.

Tavakkoli-Moghaddam:2007:DFL

[TMJJS07] R. Tavakkoli-Moghaddam, N. Javadian, B. Javadi, and N. Safaei. Design of a facility layout problem in cellular manufacturing systems with stochastic demands. Applied Mathematics and Computation, 184(2):721–728, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600748X.

Tavakkoli-Moghaddam:2005:HMS

[TMJV^{+}05] R. Tavakkoli-Moghaddam, F. Jolai, F. Vaziri, P. K. Ahmed, and A. Azaron. A hybrid method for solving stochastic job shop scheduling problems. Applied Mathematics and Computation, 170(1):185–206, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009178.

Tumwiine:2007:MMD

[TML07] J. Tumwiine, J. Y. T. Mugisha, and L. S. Luboobi. A mathematical model for the dynamics of malaria in a human host and mosquito vector with temporary immunity. Applied Mathematics and Computation, 189(2):1953–1965, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000148.

Tavakkoli-Moghaddam:2005:OSS

[TMMVA05] R. Tavakkoli-Moghaddam, G. Moslehi, M. Vasei, and A. Azaron. Optimal scheduling for a single machine to minimize the sum of maximum earliness and tardiness considering idle insert. Applied Mathematics and Computation, 167(2):1430–1450, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005788.

Tavakkoli-Moghaddam:2006:BBA

[TMMVA06] R. Tavakkoli-Moghaddam, G. Moslehi, M. Vasei, and A. Azaron. A branch-and-bound algorithm for a single machine sequencing to minimize the sum of maximum earliness and tardiness with idle insert. Applied Mathematics and Computation, 174(1):388–408, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004662.

Tavakkoli-Moghaddam:2006:MCS

[TMRV06] R. Tavakkoli-Moghaddam and A. R. Rahimi-Vahed.

Multi-criteria sequencing problem for a mixed-model assembly line in a JIT production system. Applied Mathematics and Computation, 181(2):1471–1481, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002761.

Tabatabaei:2008:NSN

[TMS08] S. Amir Hossein A. E. Tabatabaei, Alaeddin Malek, and Elham Shakour. Numerical solution for the nonlinear eigenvalue problems in bifurcation points. Applied Mathematics and Computation, 195(2):397– 401, February 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307005590.

Tavakkoli-Moghaddam:2006:HSA

[TMSG06] R. Tavakkoli-Moghaddam, N. Safaei, and Y. Gholipour. A hybrid simulated annealing for capacitated vehicle routing problems with the independent route length. Applied Mathematics and Computation, 176(2):445–454, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008210.

Tavakkoli-Moghaddam:2006:MAV

[TMSZ06] R. Tavakkoli-Moghaddam, A. R. Saremi, and M. S. Ziaee. A memetic algorithm for a vehicle routing problem with backhauls. Applied Mathematics and Computation, 181(2):1049–1060, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001883.

Truong:2009:OIS

[TNL09] Le Xuan Truong, Le Thi Phuong Ngoc, and Nguyen Thanh Long. The n-order iterative schemes for a nonlinear Kirchhoff–Carrier wave equation associated with the mixed inhomogeneous conditions. Applied Mathematics and Computation, 215(5):1908–1925, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006936.

Tripathi:2007:MES

[TNS07] Agraj Tripathi, Ram Naresh, and Dileep Sharma. Modeling the effect of screening of unaware infectives on the spread of HIV infection. Applied Mathematics and Computation, 184(2):1053–1068, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007946.

Tohumoglu:2005:APT

[Toh05] Gülay Tohumoğlu. Analysis of periodically timevarying discrete-time systems in spectral domain. Applied Mathematics and Computation, 162(3):1151–1165, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001754.

Toksari:2006:ACO

[Tok06] M. Duran Toksari. Ant colony optimization for finding the global minimum. Applied Mathematics and Computation, 176 (1):308–316, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305007952.

Touma:2009:CUF

[Tou09]  R. Touma.  Central

unstaggered finite volume schemes for hyperbolic systems: Applications to unsteady shallow water equations. Applied Mathematics and Computation, 213(1):47– 59, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002008.

Tian:2009:PAP

[TQC09] Baodan Tian, Yanhong Qiu, and Ning Chen. Periodic and almost periodic solution for a nonautonomous epidemic predatorprey system with timedelay. Applied Mathematics and Computation, 215(2):779–790, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005633.

Taniar:2008:ERA

[TRLD08] David Taniar, Wenny Rahayu, Vincent Lee, and Olena Daly. Exception rules in association rule mining. Applied Mathematics and Computation, 205(2):735–750, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003524.

Tomar:2005:TLW

[TS05] S. K. Tomar and Jaswant Singh. Transmission of longitudinal waves through a plane interface between two dissimilar porous elastic solid half-spaces. Applied Mathematics and Computation, 169(1):671– 688, October 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304008021.

Tharwat:2006:CNA

[TS06] A. A. Tharwat and M. Saleh. A comprehensive numerical algorithm for solving service points location problems. Applied Mathematics and Computation, 176(1):44–57, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305007691.

Tasic:2008:SRC

[TS08a] Milan B. Tasić and Predrag S. Stanimirović. Symbolic and recursive computation of different types of generalized inverses. Applied Mathematics and Computation, 199(1):349– 367, May 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307010041.

Tfaili:2008:NCA

[TS08b] Walid Tfaili and Patrick Siarry. A new charged ant colony algorithm for continuous dynamic optimization. Applied Mathematics and Computation, 197 (2):604–613, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030700817X.

Trojnar-Spelina:2008:CAH

[TS08c] Lucyna Trojnar-Spelina.

On certain applications of the Hadamard product. Applied Mathematics and Computation, 199(2):653– 662, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010533.

Tavares:2009:KPE

[TS09] C. C. S. Tavares and M. L.

Santos. On the Kirchhoff plates equations with thermal effects and memory boundary conditions. Applied Mathematics and Computation, 213(1):25– 38, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002240.

Tsaur:2005:SSS

[Tsa05] Woei-Jiunn Tsaur. Several security schemes constructed using ECC-based self-certified public key cryptosystems. Applied Mathematics and Computation, 168(1):447–464, September 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304006174.

Tabatabaei:2007:SIM

[TSD07] Amir Hossein A. E. Tabatabaei, Elham Shakour, and Mehdi Dehghan. Some implicit methods for the numerical solution of Burgers’ equation. Applied Mathematics and Computation, 191(2):560–570, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002846.

Tsekouras:2005:FVQ

[Tse05] George E. Tsekouras. A fuzzy vector quantization approach to image compression. Applied Mathematics and Computation, 167(1):539–560, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005120.

Tsoulos:2008:MRC

[Tso08] Ioannis G. Tsoulos. Modifications of real code genetic algorithm for global optimization. Applied Mathematics and Computation, 203(2):598–607, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002907.

Tsoulos:2009:SCO

[Tso09] Ioannis G. Tsoulos. Solving constrained optimization problems using a novel genetic algorithm. Applied Mathematics and Computation, 208(1):273–283, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008990.

Tasic:2007:SCW

[TSP07] Milan B. Tasić, Predrag S. Stanimirović, and Marko D. Petković. Symbolic computation of weighted Moore–Penrose inverse using partitioning method. Applied Mathematics and Computation, 189(1):615– 640, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016365.

Taneja:2009:FBD

[TSS09] Reema Taneja, Kennon C. Shelton, and Ajit Sadana. Fractal binding and dissociation kinetics of prion proteins on biosensor surfaces. Applied Mathematics and Computation, 207(1):5–22, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308000167.

Tareghian:2006:ARM

[TT06a] Hamed R. Tareghian and Hassan Taheri. An application of randomized minimum cut to the project time/cost tradeoff problem. Applied Mathematics and Computation, 173(2):1200–1207, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305004236.

Tareghian:2006:DTC

[TT06b] Hamed R. Tareghian and Seyyed Hassan Taheri. On the discrete time, cost and quality trade-off problem. Applied Mathematics and Computation, 181(2):1305–1312, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002499.

Tareghian:2007:SPD

[TT07] Hamed R. Tareghian and Seyyed Hassan Taheri. A solution procedure for the discrete time, cost and quality tradeoff problem using electromagnetic scatter search. Applied Mathematics and Computation, 190(2):1136–1145, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001336.

Tuan:2009:NRM

[TT09] Nguyen Huy Tuan and Dang Duc Trong. A new regularized method for two dimensional nonhomogeneous backward heat problem. Applied Mathematics and Computation, 215(3):873–880, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/S0096300309005827.

See corrigendum [Tua13].

Tseng:2008:ACR

[TTC08] Din-Chang Tseng, Hsiao-Ting Tseng, and Chun-Liang Chien. Automatic cloud removal from multi-temporal SPOT images. Applied Mathematics and Computation, 205(2):584–600, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003378.

Tsai:2007:OEP

[TTG07] Jason S. H. Tsai, Tzong J. Tsai, and Shu-Mei Guo. Observer with evolutionaryprogramming-based optimal digital tracker for analog neutral multiple timedelay systems. Applied Mathematics and Computation, 188(1):1001–1015, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014482.

Teng:2008:PHP

[TTH08] Yuan-Hsiang Teng, Jimmy J. M. Tan, and Lih-Hsing Hsu. Panpositionable Hamiltonicity and panconnectivity of the arrangement graphs. Applied Mathematics and Computation, 198(1):414–432, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009113.

Thompson:2009:UCD

[TTZ09] B. C. Thompson, J. V.

Tucker, and J. I. Zucker. Unifying computers and dynamical systems using the theory of synchronous concurrent algorithms. Applied Mathematics and Computation, 215(4):1386–1403, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004184.

Tuan:2013:CSN

[Tua13] Nguyen Huy Tuan. Corrigendum to: “A new regularized method for two dimensional nonhomogeneous backward heat problem” [Appl. Math. Comput. 215 (2009) 873– 880]. Applied Mathematics and Computation, 219(9):4931–4932, January 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/S0096300312010594.

See [TT09].

Triki:2009:SOM

[TW09] Houria Triki and Abdul-Majid Wazwaz. Sub-ODE method and soliton solutions for the variablecoefficient mKdV equation. Applied Mathematics and Computation, 214(2):370–373, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309003221.

Tsaur:2005:EUA

[TWL05] Woei-Jiunn Tsaur, Chia-Chun Wu, and Wei-Bin Lee. An enhanced user authentication scheme for multi-server Internet services. Applied Mathematics and Computation, 170(1):258–266, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304009233.

Tan:2007:TDM

[TWZ07] Junyu Tan, Yong Wu, and Linhua Zhang. Twodimensional multiple nonhomogeneous harmonic equation and its boundary integral equations. Applied Mathematics and Computation, 187(2):893–901, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012215.

Tan:2008:HSM

[TX08] Xiaobin Tan and Hongsheng Xi. Hidden semi-Markov model for anomaly detection. Applied Mathematics and Computation, 205(2):562–567, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003342.

Tang:2009:TWSa

[TXW09] Shengqiang Tang, Yixin Xiao, and Zhaojuan Wang. Travelling wave solutions for a class of nonlinear fourth order variant of a generalized Camassa– Holm equation. Applied Mathematics and Computation, 210(1):39–47, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008096.

Tsai:2007:ALS

[TY07] Chiung-Chiou Tsai and Suh-Yuh Yang. Analysis of the L^{2} least-squares finite element method for a velocity–vorticity problem arising in incompressible inviscid rotational flows. Applied Mathematics and Computation, 185(1):604–613, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008952.

Tsai:2006:MFS

[TYCL06] C. C. Tsai, D. L. Young, J. H. Chiang, and D. C. Lo. The method of fundamental solutions for solving options pricing models. Applied Mathematics and Computation, 181(1):390–401, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001093.

Tzekis:2007:NAS

[Tze07] P. A. Tzekis. A new algorithm for the solution of a polynomial matrix Diophantine equation. Applied Mathematics and Computation, 193(2):395–407, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004134.

Tang:2009:TWS

[TZH09a] Shengqiang Tang, Jianxian Zheng, and Wentao Huang. Travelling wave solutions for a class of generalized KdV equation. Applied Mathematics and Computation, 215(7):2768–2774, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008170.

Tang:2009:TWSb

[TZH09b] Shengqiang Tang, Jianxian Zheng, and Wentao Huang. Travelling wave solutions for a class of generalized KdV equation. Applied Mathematics and Computation, 215(7):2768–2774, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008170.

Tzoumas:2008:SSC

[Tzo08] Michael G. Tzoumas. On sign symmetric circulant matrices. Applied Mathematics and Computation, 195(2):604–617, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005905.

Tang:2009:IFM

[TZW09] Shengqiang Tang, Jie Zheng, and Zhaojuan Wang. The integral factor method for solving a class of generalized KdV equation. Applied Mathematics and Computation, 211(1):185–189, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000691.

Tian:2007:DDA

[TZX07] Junkang Tian, Shouming Zhong, and Lianglin Xiong. Delay-dependent absolute stability of Lurie control systems with multiple time-delays. Applied Mathematics and Computation, 188(1):379–384, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306013610.

Upadhyay:2008:CEE

[UBKC08] R. K. Upadhyay, N. Bairagi, K. Kundu, and J. Chattopadhyay. Chaos in eco-epidemiological problem of the Salton Sea and its possible control. Applied Mathematics and Computation, 196(1):392–401, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006923.

Udwadia:2005:P

[Udw05a] F. E. Udwadia. Preface. Applied Mathematics and Computation, 164(2):275–276, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003108.

Udwadia:2005:BCQ

[Udw05b] Firdaus E. Udwadia. Boundary control, quiet boundaries, super-stability and super-instability. Applied Mathematics and Computation, 164(2):327–349, May 16, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003212.

Udwadia:2005:CM

[Udw05c] Firdaus E. Udwadia. On constrained motion. Applied Mathematics and Computation, 164(2):313– 320, May 16, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304003157.

Ujevic:2007:GCM

[UE07] Nenad Ujević and Goran Erceg. A generalization of the corrected midpointtrapezoid rule and error bounds. Applied Mathematics and Computation, 184(2):216–222, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006977.

Ujevic:2007:FMS

[UEL07] Nenad Ujević, Goran Erceg, and Ivan Lekić. A family of methods for solving nonlinear equations. Applied Mathematics and Computation, 192(2):311–318, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307003311.

See note [Raf08].

Uddin:2009:MFN

[UHuI09] Marjan Uddin, Sirajul Haq, and Siraj ul Islam. A mesh-free numerical method for solution of the family of Kuramoto– Sivashinsky equations. Applied Mathematics and Computation, 212(2):458– 469, June 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309001775.

Siraj-ul-Islam:2006:NMB

[uIK06] Siraj ul Islam and Muhammad Azam Khan. A numerical method based on polynomial sextic spline functions for the solution of special fifth-order boundary-value problems. Applied Mathematics and Computation, 181(1):356– 361, October 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306001044.

Siraj-ul-Islam:2005:NPS

[uIKTT05] Siraj ul Islam, Muhammad Azam Khan, Ikram A. Tirmizi, and E. H. Twizell. Non polynomial spline approach to the solution of a system of third-order boundary-value problems. Applied Mathematics and Computation, 168(1):152–163, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304005958.

Siraj-ul-Islam:2006:QNP

[uINTK06] Siraj ul Islam, Muhammad Aslam Noor, Ikram A. Tirmizi, and Muhammad Azam Khan. Quadratic non-polynomial spline approach to the solution of a system of second-order boundary-value problems. Applied Mathematics and Computation, 179(1):153– 160, August 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305010179.

Siraj-ul-Islam:2006:NSA

[uIT06] Siraj ul Islam and Ikram A. Tirmizi. Nonpolynomial spline approach to the solution of a system of second-order boundaryvalue problems. Applied Mathematics and Computation, 173(2):1208–1218, February 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305004248.

Siraj-ul-Islam:2006:CMB

[uITA06] Siraj ul Islam, Ikram A. Tirmizi, and Saadat Ashraf. A class of methods based on non-polynomial spline functions for the solution of a special fourth-order boundary-value problems with engineering applications. Applied Mathematics and Computation, 174(2):1169–1180, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005394.

See comment [RJM07a].

Siraj-ul-Islam:2008:NPS

[uITiHK08] Siraj ul Islam, Ikram A.

Tirmizi, Fazal i Haq, and M. Azam Khan. Nonpolynomial splines approach to the solution of sixth-order boundaryvalue problems. Applied Mathematics and Computation, 195(1):270–284, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005474.

Ujevic:2006:MSN

[Uje06a] Nenad Ujević. A method for solving nonlinear equations. Applied Mathematics and Computation, 174(2):1416–1426, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005606.

Ujevic:2006:NIM

[Uje06b] Nenad Ujević. A new iterative method for solving linear systems. Applied Mathematics and Computation, 179(2):725–730, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010805.

Ujevic:2009:NPC

[Uje09] Nenad Ujević. Notes on a paper considering nonlinear equations. Applied Mathematics and Computation, 212(1):273–274, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001295.

Uslu:2007:ENC

[UNS07] K. Uslu, A. Nalli, and M. Sen. Erratum to “On the norms of circulant matrices with the Fibonacci and Lucas numbers” [Appl. Math. Comput. 160 (1) (2005) 125–132]. Applied Mathematics and Computation, 190(1):997–998, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000306.

See [Sol07].

Udwadia:2007:GIM

[UP07a] Firdaus E. Udwadia and Phailaung Phohomsiri. Generalized LM-inverse of a matrix augmented by a column vector. Applied Mathematics and Computation, 190(2):999–1006, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012914.

Uthayakumar:2007:ACF

[UP07b] R. Uthayakumar and P. Paramanathan. An algorithm for computing fractal dimension of rectifiable irregular graphs. Applied Mathematics and Computation, 190(1):305–308, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000501.

Uthayakumar:2007:FDI

[UP07c] R. Uthayakumar and P. Paramanathan. Fractal dimension of irregular digitalized curves by divider method. Applied Mathematics and Computation, 189(1):68– 71, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015803.

Upadhyay:2009:DEM

[Upa09] Ranjit Kumar Upadhyay. Dynamics of an ecological model living on the edge of chaos. Applied Mathematics and Computation, 210(2):455–464, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000149.

Utkin:2006:RPP

[Utk06]  Lev V. Utkin.  Rank-

ing procedures by pairwise comparison using random sets and the imprecise Dirichlet model. Applied Mathematics and Computation, 183(1):394–408, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005716.

Vajargah:2007:DSA

[Vaj07a] Behrouz Fathi Vajargah. Different stochastic algorithms to obtain matrix inversion. Applied Mathematics and Computation, 189(2):1841–1846, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017802.

Vajargah:2007:NAO

[Vaj07b] Behrouz Fathi Vajargah.

New advantages to obtain accurate matrix inversion. Applied Mathematics and Computation, 189(2):1798–1804, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017747.

Vajargah:2007:NAM

[Vaj07c] Behrouz Fathi Vajargah.

A new algorithm with maximal rate convergence to obtain inverse matrix. Applied Mathematics and Computation, 191(1):280–286, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700255X.

Vajargah:2007:WOM

[Vaj07d] Behrouz Fathi Vajargah.

A way to obtain Monte Carlo matrix inversion with minimal error. Applied Mathematics and Computation, 191(1):225–233, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002494.

Vajargah:2008:SIM

[Vaj08] Behrouz Fathi Vajargah. Stochastic inverse matrix computation with minimum variance of errors. Applied Mathematics and Computation, 199(1):181–185, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700985X.

Valli:2007:NSM

[Val07] Pramada Valli. Numerical and statistical modelling for stabilisation of landslides. Applied Mathematics and Computation, 186(2):1083–1093, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010344.

Vigo-Aguiar:2007:EFB

[VAMV07] J. Vigo-Aguiar and J. Martín-Vaquero. Exponential fitting BDF algorithms and their properties. Applied Mathematics and Computation, 190(1):80–110, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700032X.

VandeVyver:2005:MER

[Van05] Hans Van de Vyver. Modified explicit Runge–Kutta methods for the numerical solution of the Schrödinger equation. Applied Mathematics and Computation, 171(2):1025–1036, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

VandeVyver:2007:AEH

[Van07] Hans Van de Vyver. An adapted explicit hybrid method of Numerov-type for the numerical integration of perturbed oscillators. Applied Mathematics and Computation, 186(2):1385–1394, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010642.

VanGorder:2009:CCI

[Van09] Robert A. Van Gorder.

Computation of certain infinite series of the form ∑ f(n)n^{k} for arbitrary real-valued k. Applied Mathematics and Computation, 215(3):1209–1216, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006353.

Vajravelu:2006:FFN

[VC06] K. Vajravelu and J. R. Cannon. Fluid flow over a nonlinearly stretching sheet. Applied Mathematics and Computation, 181(1):609–618, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001317.

Velez-Cerrada:2009:EBP

[VCRCG09] J. Y. Vélez-Cerrada, J. Robles, and N. Castro-González. Error bounds for the perturbation of the Drazin inverse under some geometrical conditions. Applied Mathematics and Computation, 215(6):2154–2161, November 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300309007243.

Velupillai:2006:AFC

[Vel06] K. Vela Velupillai. Algorithmic foundations of computable general equilibrium theory. Applied Mathematics and Computation, 179(1):360–369, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030501043X.

Velupillai:2009:UUE

[Vel09] K. Vela Velupillai. Uncomputability and undecidability in economic theory. Applied Mathematics and Computation, 215(4):1404–1416, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004172.

Verma:2007:PPA

[Ver07] Ram U. Verma. Proximal point algorithms and generalized nonlinear variational problems. Applied Mathematics and Computation, 187(1):535–543, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012057.

Vieru:2008:FGO

[VFF08a] D. Vieru, Corina Fetecau, and C. Fetecau. Flow of a generalized Oldroyd-B fluid due to a constantly accelerating plate. Applied Mathematics and Computation, 201(1–2):834–842, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307012234.

Vieru:2008:FVF

[VFF08b] D. Vieru, Corina Fetecau, and C. Fetecau. Flow of a viscoelastic fluid with the fractional Maxwell model between two side walls perpendicular to a plate. Applied Mathematics and Computation, 200(1):459–464, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011265.

Valverde:2007:CPP

[VG07] Jose C. Valverde and Laura Gardini. On the changes of periodicities in a piecewise linear rotation model. Applied Mathematics and Computation, 194(2):381–388, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700478X.

Varon:2007:ODT

[VGMV^{+}07] R. Varón, M. García-Moreno, E. Valero, M. E.

Fuentes, M. L. L. Amo, E. Arias, E. Arribas, and F. García-Sevilla. Optimized derivation of transfer functions and a software giving it. Application to biological systems. Applied Mathematics and Computation, 184(2):823–841, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007703.

Vieru:2008:FPS

[VHFF08] D. Vieru, T. Hayat, Corina Fetecau, and C. Fetecau. On the first problem of Stokes for Burgers’ fluids. II: The cases γ = λ^{2}/4 and γ > λ^{2}/4. Applied Mathematics and Computation, 197(1):76–86, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007552.

Viazminsky:2008:NSC

[Via08] C. P. Viazminsky. Necessary and sufficient conditions for a function to be separable. Applied Mathematics and Computation, 204(2):658–670, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308000945.

Vicent:2009:URS

[Vic09] Luis Edgar Vicent. Unitary rotation of squarepixellated images. Applied Mathematics and Computation, 211(1):111–117, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000460.

Vinh:2009:ESE

[Vin09]  Pham Chi Vinh.  Ex-

plicit secular equations of Rayleigh waves in elastic media under the influence of gravity and initial stress. Applied Mathematics and Computation, 215(1):395–404, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309004822.

Viswanathan:2008:IHW

[Vis08] V. Viswanathan. Information hiding in wave files through frequency domain. Applied Mathematics and Computation, 201(1–2):121–127, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011769.

Vitanov:2009:NPW

[VJD09] Nikolay K. Vitanov, Ivan P.

Jordanov, and Zlatinka I. Dimitrova. On nonlinear population waves. Applied Mathematics and Computation, 215(8):2950–2964, December 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309008662.

Velazquez:2005:MOR

[VKVHN05] Ismael Velazquez, David Kaplan, Jorge X. Velasco-Hernandez, and Sergio A. Navarrete. Multistability in an open recruitment food web model. Applied Mathematics and Computation, 163(1):275–294, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002255.

Voglis:2009:TIM

[VL09] C. Voglis and I. E. Lagaris. Towards “Ideal Multistart”. A stochastic approach for locating the minima of a continuous function inside a bounded domain. Applied Mathematics and Computation, 213(1):216–229, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002215.

Vencheh:2005:UFE

[VMK05] A. Hadi Vencheh, R. Kazemi Matin, and M. Tavassoli Kajani. Undesirable factors in efficiency measurement. Applied Mathematics and Computation, 163(2):547–552, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002413.

Vogel:2007:FBF

[Vog07] Judith A. Vogel. Flexible BiCG and flexible Bi-CGSTAB for nonsymmetric linear systems. Applied Mathematics and Computation, 188(1):226–233, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306013427.

Viswanath:2008:MTL

[VR08] Asha Viswanath and D. Roy. A multistep transversal linearization (MTL) method in non-linear dynamics through a Magnus characterization. Applied Mathematics and Computation, 198(2):799–823, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700954X.

Verma:2006:GTD

[VRG06] Rajeev Verma, Niranjan Roy, and Ranjan Ganguli. Gas turbine diagnostics using a soft computing approach. Applied Mathematics and Computation, 172(2):1342–1363, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002961.

Verde-Star:2007:SMC

[VS07] Luis Verde-Star. Symmetric multivariate Chebyshev polynomials. Applied Mathematics and Computation, 187(1):530–534, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012045.

Vajravelu:2005:PPH

[VSB05] K. Vajravelu, S. Sreenadh, and V. Ramesh Babu. Peristaltic pumping of a Herschel–Bulkley fluid in a channel. Applied Mathematics and Computation, 169(1):726–735, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008069.

Vulkov:2007:BSQ

[Vul07] Lubin G. Vulkov. Blow up for some quasilinear equations with dynamical boundary conditions of parabolic type. Applied Mathematics and Computation, 191(1):89–99, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000999.

Vajargah:2006:PMC

[VV06] Behrouz Fathi Vajargah and Kianoush Fathi Vajargah. Parallel Monte Carlo computations for solving SLAE with minimum communications. Applied Mathematics and Computation, 183(1):1–9, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306002906.

Vynnycky:2007:MTP

[Vyn07] M. Vynnycky. On the modelling of two-phase flow in the cathode gas diffusion layer of a polymer electrolyte fuel cell. Applied Mathematics and Computation, 189(2):1560–1575, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017425.

Wang:2004:NAS

[Wan04] Lei Wang. A new algorithm for solving classical Blasius equation. Applied Mathematics and Computation, 157(1):1–9, September 27, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See comments [Has06b].

Wang:2005:NSSb

[Wan05a] Hanquan Wang. Numerical studies on the split-step finite difference method for nonlinear Schrödinger equations. Applied Mathematics and Computation, 170(1):17–35, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008689.

Wang:2005:CDA

[Wan05b]  Lianwen Wang.  Con-

vergence of discrete approximations to optimization problems of neutral functional–differential inclusions. Applied Mathematics and Computation, 165(2):375–391, June 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003881.

Wang:2005:SCG

[Wan05c] Lin Wang. Stability of Cohen–Grossberg neural networks with distributed delays. Applied Mathematics and Computation, 160(1):93–110, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303010300.

Wang:2005:HGAa

[Wan05d]  Ling Wang.  A hy-

brid genetic algorithm– neural network strategy for simulation optimization. Applied Mathematics and Computation, 170(2):1329–1343, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000676.

Wang:2005:NSSa

[Wan05e] Quan-Fang Wang. Numerical solution for series sine-Gordon equations using variational method and finite element approximation. Applied Mathematics and Computation, 168(1):567–599, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006253.

Wang:2005:SCR

[Wan05f] Shiuh-Jeng Wang. Steganography of capacity required using modulo operator for embedding secret image. Applied Mathematics and Computation, 164(1):99– 116, May 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400298X.

Wang:2005:ASD

[Wan05g] Weiming Wang. An algorithm for solving DAEs with mechanization. Applied Mathematics and Computation, 167(2):1350– 1372, August 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304005727.

Wang:2005:ASN

[Wan05h] Weiming Wang. An algorithm for solving nonlinear singular perturbation problems with mechanization. Applied Mathematics and Computation, 169(2):995–1009, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400829X.

Wang:2006:UEE

[Wan06a]  Bingxing Wang.  Un-

biased estimations for the exponential distribution based on stepstress accelerated lifetesting data. Applied Mathematics and Computation, 173(2):1227–1237, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004261.

Wang:2006:URC

[Wan06b] Chi-Chang Wang. Use residual correction method to calculate the upper and lower solutions of initial value problem. Applied Mathematics and Computation, 181(1):29–39, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000634.

Wang:2006:NSH

[Wan06c] He-Sheng Wang. Numerical solutions of Hamiltonian DAE’s and applications. Applied Mathematics and Computation, 180(1):97–110, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030501101X.

Wang:2006:SMS

[Wan06d] Ji-Bo Wang. Single machine scheduling with common due date and controllable processing times. Applied Mathematics and Computation, 174(2):1245– 1254, March 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030500545X.

Wang:2006:AMA

[Wan06e] Qi Wang. An application of the modified Adomian decomposition method for (N + 1)-dimensional sine-Gordon field. Applied Mathematics and Computation, 181(1):147–152, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000841.

Wang:2006:NSF

[Wan06f]  Qi Wang.  Numerical

solutions for fractional KdV–Burgers equation by Adomian decomposition method. Applied Mathematics and Computation, 182(2):1048–1055, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600419X.

Wang:2006:ASH

[Wan06g] Weiming Wang. An algorithm for solving the highorder nonlinear Volterra– Fredholm integro-differential equation with mechanization. Applied Mathematics and Computation, 172(1):1–23, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305001815.

Wang:2006:MASc

[Wan06h] Weiming Wang. Mechanical algorithm for solving the second kind of Volterra integral equation. Applied Mathematics and Computation, 173(2):1149–1162, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004200.

Wang:2006:MASb

[Wan06i] Weiming Wang. A mechanical algorithm for solving the Volterra integral equation. Applied Mathematics and Computation, 172(2):1323–1341, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500295X.

Wang:2006:NMA

[Wan06j] Weiming Wang. A new mechanical algorithm for solving the second kind of Fredholm integral equation. Applied Mathematics and Computation, 172(2):946–962, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002493.

Wang:2006:NSDa

[Wan06k] Xing Tao Wang. Numerical solution of delay systems containing inverse time by hybrid functions. Applied Mathematics and Computation, 173(1):535–546, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305003735.

Wang:2006:LGP

[Wan06l] Ying-Ming Wang. On lexicographic goal programming method for generating weights from inconsistent interval comparison matrices. Applied Mathematics and Computation, 173(2):985–991, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305004078.

Wang:2007:NMF

[Wan07a]  Jian Wang.  A note

on multisymplectic Fourier pseudospectral discretization for the nonlinear Schrödinger equation. Applied Mathematics and Computation, 191(1):31– 41, August 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306013038.

Wang:2007:ANM

[Wan07b] Jin Wang. On the accuracy of a numerical method for linear interfacial motion. Applied Mathematics and Computation, 188(2):1733–1741, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015402.

Wang:2007:ECC

[Wan07c] Li Wang. Equivalent conditions for the convergence of extrapolated method for linear systems. Applied Mathematics and Computation, 184(2):917–923, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007818.

Wang:2007:HPM

[Wan07d]  Qi Wang.  Homotopy

perturbation method for fractional KdV equation. Applied Mathematics and Computation, 190(2):1795– 1802, July 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307002287.

Wang:2007:GRG

[Wan07e] Wu-Sheng Wang. A generalized retarded Gronwalllike inequality in two variables and applications to BVP. Applied Mathematics and Computation, 191(1):144–154, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700238X.

Wang:2007:NSO

[Wan07f] Xing Tao Wang. Numerical solutions of optimal control for time delay systems by hybrid of blockpulse functions and Legendre polynomials. Applied Mathematics and Computation, 184(2):849–856, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007727.

Wang:2008:ACC

[Wan08a]  Chunpeng Wang.  Ap-

proximate controllability of a class of degenerate systems. Applied Mathematics and Computation, 203(1):447–456, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002786.

Wang:2008:ENF

[Wan08b]  Ji-Bo Wang.  Erratum

to: A note on flow shop scheduling problems with a learning effect on noidle dominant machines [Appl. Math. Comput. 184 (2007) 945–949]. Applied Mathematics and Computation, 202(2):897–898, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010788.

See [CSY07].

Wang:2008:MSG

[Wan08c]  Jian Wang.  On mul-

tisymplecticity of Sinc– Gauss–Legendre collocation discretizations for some Hamiltonian PDEs. Applied Mathematics and Computation, 199(1):105– 121, May 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307009770.

Wang:2008:TIC

[Wan08d] Quan-Fang Wang. Theoretical issue of controlling nucleus in Klein– Gordon–Schrödinger dynamics with perturbation in control field. Applied Mathematics and Computation, 206(1):276–289, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006097.

Wang:2008:NST

[Wan08e] Xing Tao Wang. Numerical solution of time-varying systems with a stretch by general Legendre wavelets. Applied Mathematics and Computation, 198(2):613– 620, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009277.

Wang:2008:GAS

[Wan08f] Yuan-Ming Wang. Global asymptotic stability of 3-species Lotka–Volterra models with diffusion and time delays. Applied Mathematics and Computation, 195(1):34–48, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005152.

Wangkeeree:2008:VAM

[Wan08g] Rabian Wangkeeree. Viscosity approximative methods to Cesàro mean iterations for nonexpansive nonself-mappings in Banach spaces. Applied Mathematics and Computation, 201(1–2):239–249, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011915.

Wang:2009:EAD

[Wan09a] Jian Wang. Error analysis of discrete conservation laws for Hamiltonian PDEs under the central box discretizations. Applied Mathematics and Computation, 211(1):155– 166, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000654.

Wang:2009:CIP

[Wan09b] Quan-Fang Wang. Computation issue of pointwise controls for diffusion Hopfield neural network system. Applied Mathematics and Computation, 207(1):148–163, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007741.

Wang:2009:SSC

[Wan09c] Zhanping Wang. Stability of solution to a class of investment system. Applied Mathematics and Computation, 207(2):340–345, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008163.

Wazwaz:2005:ADMb

[Waz05a] Abdul-Majid Wazwaz. Adomian decomposition method for a reliable treatment of the Bratu-type equations. Applied Mathematics and Computation, 166(3):652–663, July 26, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004643.

Wazwaz:2005:ADMa

[Waz05b] Abdul-Majid Wazwaz. Adomian decomposition method for a reliable treatment of the Emden–Fowler equation. Applied Mathematics and Computation, 161(2):543–560, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304000104.

Wazwaz:2005:AST

[Waz05c] Abdul-Majid Wazwaz. Analytical solution for the time-dependent Emden– Fowler type of equations by Adomian decomposition method. Applied Mathematics and Computation, 166(3):638–651, July 26, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004631.

Wazwaz:2005:CHK

[Waz05d] Abdul-Majid Wazwaz. The Camassa–Holm–KP equations with compact and noncompact travelling wave solutions. Applied Mathematics and Computation, 170(1):347–360, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009300.

Wazwaz:2005:CNF

[Waz05e] Abdul-Majid Wazwaz. A class of nonlinear fourth order variant of a generalized Camassa–Holm equation with compact and noncompact solutions. Applied Mathematics and Computation, 165(2):485–501, June 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003972.

Wazwaz:2005:CNP

[Waz05f] Abdul-Majid Wazwaz. Compact and noncompact physical structures for the ZK– BBM equation. Applied Mathematics and Computation, 169(1):713–725, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008057.

Wazwaz:2005:CSP

[Waz05g] Abdul-Majid Wazwaz. Compactons, solitons and periodic solutions for variants of the KdV and the KP equations. Applied Mathematics and Computation, 161(2):561–575, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304000116.

Wazwaz:2005:ESC

[Waz05h] Abdul-Majid Wazwaz. Exact solutions of compact and noncompact structures for the KP–BBM equation. Applied Mathematics and Computation, 169(1):700–712, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008045.

Wazwaz:2005:GBT

[Waz05i] Abdul-Majid Wazwaz. Generalized Boussinesq type of equations with compactons, solitons and periodic solutions. Applied Mathematics and Computation, 167(2):1162–1178, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005582.

Wazwaz:2005:NCN

[Waz05j] Abdul-Majid Wazwaz. New compact and noncompact solutions for two variants of a modified Camassa– Holm equation. Applied Mathematics and Computation, 163(3):1165–1179, April 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002875.

Wazwaz:2005:NDS

[Waz05k] Abdul-Majid Wazwaz. Nonlinear dispersive special type of the Zakharov– Kuznetsov equation ZK(n, n) with compact and noncompact structures. Applied Mathematics and Computation, 161(2):577–590, February 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304000128.

Wazwaz:2005:RTP

[Waz05l] Abdul-Majid Wazwaz. A reliable treatment of the physical structure for the nonlinear equation K(m, n). Applied Mathematics and Computation, 163(3):1081–1095, April 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002814.

Wazwaz:2005:TSC

[Waz05m] Abdul-Majid Wazwaz. The tanh and the sine–cosine methods for compact and noncompact solutions of the nonlinear Klein–Gordon equation. Applied Mathematics and Computation, 167(2):1179–1195, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005600.

Wazwaz:2005:TME

[Waz05n] Abdul-Majid Wazwaz. The tanh method: exact solutions of the sine-Gordon and the sinh-Gordon equations. Applied Mathematics and Computation, 167(2):1196–1210, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005612.

Wazwaz:2005:MRT

[Waz05o] Abdul-Majid Wazwaz. The tanh method for a reliable treatment of the K(n, n) and the KP (n, n) equations and its variants. Applied Mathematics and Computation, 170(1):361–379, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009312.

Wazwaz:2005:TMG

[Waz05p] Abdul-Majid Wazwaz. The tanh method for generalized forms of nonlinear heat conduction and Burgers– Fisher equations. Applied Mathematics and Computation, 169(1):321–338, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007775.

Wazwaz:2005:TWS

[Waz05q] Abdul-Majid Wazwaz. Travelling wave solutions of generalized forms of Burgers, Burgers–KdV and Burgers–Huxley equations. Applied Mathematics and Computation, 169(1):639– 656, October 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304008008.

Wazwaz:2006:ASS

[Waz06a] Abdul-Majid Wazwaz. Abundant solitons solutions for several forms of the fifth-order KdV equation by using the tanh method. Applied Mathematics and Computation, 182(1):283–300, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003377.

Wazwaz:2006:ASF

[Waz06b] Abdul-Majid Wazwaz. Analytic study of the fifth order integrable nonlinear evolution equations by using the tanh method. Applied Mathematics and Computation, 174(1):289– 299, March 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030500456X.

Wazwaz:2006:CSW

[Waz06c] Abdul-Majid Wazwaz. Compactons and solitary wave solutions for the Boussinesq wave equation and its generalized form. Applied Mathematics and Computation, 182(1):529–535, November 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306003602.

Wazwaz:2006:CSB

[Waz06d] Abdul-Majid Wazwaz. A comparison study between the modified decomposition method and the traditional methods for solving nonlinear integral equations. Applied Mathematics and Computation, 181(2):1703–1712, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003018.

Wazwaz:2006:ETWbb

[Waz06e] Abdul-Majid Wazwaz. Exact travelling wave solutions to seventh-order and ninth-order KdV-like equations. Applied Mathematics and Computation, 182(1):771–780, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003869.

Wazwaz:2006:ETWa

[Waz06f] Abdul-Majid Wazwaz. Explicit travelling wave solutions of variants of the K(n, n) and the ZK(n, n) equations with compact and noncompact structures. Applied Mathematics and Computation, 173(1):213–230, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003218.

Wazwaz:2006:KSS

[Waz06g] Abdul-Majid Wazwaz. Kinks and solitons solutions for the generalized KdV equation with two power nonlinearities. Applied Mathematics and Computation, 183(2):1181–1189, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600662X.

Wazwaz:2006:MDMb

[Waz06h] Abdul-Majid Wazwaz. The modified decomposition method and Padé approximants for a boundary layer equation in unbounded domain. Applied Mathematics and Computation, 177(2):737–744, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009549.

Wazwaz:2006:MDMa

[Waz06i] Abdul-Majid Wazwaz. The modified decomposition method for analytic treatment of differential equations. Applied Mathematics and Computation, 173(1):165–176, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003188.

Wazwaz:2006:NSW

[Waz06j] Abdul-Majid Wazwaz. New solitary wave solutions to the Kuramoto–Sivashinsky and the Kawahara equations. Applied Mathematics and Computation, 182(2):1642–1650, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005121.

Wazwaz:2006:PAA

[Waz06k] Abdul-Majid Wazwaz. Padé approximants and Adomian decomposition method for solving the Flierl– Petviashivili equation and its variants. Applied Mathematics and Computation, 182(2):1812–1818, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005303.

Wazwaz:2006:PKC

[Waz06l] Abdul-Majid Wazwaz. Peakons, kinks, compactons and solitary patterns solutions for a family of Camassa–Holm equations by using new hyperbolic schemes. Applied Mathematics and Computation, 182(1):412–424, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003481.

Wazwaz:2006:SCT

[Waz06m] Abdul-Majid Wazwaz. The sine–cosine and the tanh methods: Reliable tools for analytic treatment of nonlinear dispersive equations. Applied Mathematics and Computation, 173(1):150–164, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305003176.

Wazwaz:2006:SWS

[Waz06n] Abdul-Majid Wazwaz. Solitary wave solutions and periodic solutions for higherorder nonlinear evolution equations. Applied Mathematics and Computation, 181(2):1683–1692, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002992.

Wazwaz:2006:TWSa

[Waz06o] Abdul-Majid Wazwaz. Travelling wave solutions for combined and double combined sine–cosine-Gordon equations by the variable separated ODE method. Applied Mathematics and Computation, 177(2):755– 760, June 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305009562.

Wazwaz:2006:TWSb

[Waz06p] Abdul-Majid Wazwaz. Travelling wave solutions for the MKdV-sine-Gordon and the MKdV-sinh-Gordon equations by using a variable separated ODE method. Applied Mathematics and Computation, 181(2):1713–1719, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600302X.

Wazwaz:2006:VSO

[Waz06q] Abdul-Majid Wazwaz. The variable separated ODE and the tanh methods for solving the combined and the double combined sinh–cosh-Gordon equations. Applied Mathematics and Computation, 177(2):745–754, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009550.

Wazwaz:2007:ETMb

[Waz07a] Abdul-Majid Wazwaz. The extended tanh method for abundant solitary wave solutions of nonlinear wave equations. Applied Mathematics and Computation, 187(2):1131–1142, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012501.

Wazwaz:2007:ETMa

[Waz07b] Abdul-Majid Wazwaz. The extended tanh method for new solitons solutions for many forms of the fifth-order KdV equations. Applied Mathematics and Computation, 184(2):1002–1014, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007892.

Wazwaz:2007:MFS

[Waz07c] Abdul-Majid Wazwaz. Multiplefront solutions for the Burgers equation and the coupled Burgers equations. Applied Mathematics and Computation, 190(2):1198–1206, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001506.

Wazwaz:2007:MSSb

[Waz07d] Abdul-Majid Wazwaz. Multiplesoliton solutions for the Boussinesq equation. Applied Mathematics and Computation, 192(2):479– 486, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307003517.

Wazwaz:2007:MSSa

[Waz07e] Abdul-Majid Wazwaz. Multiplesoliton solutions for the KP equation by Hirota’s bilinear method and by the tanh–coth method. Applied Mathematics and Computation, 190(1):633–640, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000975.

Wazwaz:2007:NSW

[Waz07f] Abdul-Majid Wazwaz. New solitary wave solutions to the modified forms of Degasperis–Procesi and Camassa–Holm equations. Applied Mathematics and Computation, 186(1):130– 141, March 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306009246.

Wazwaz:2007:NSK

[Waz07g] Abdul-Majid Wazwaz. New solitons and kinks solutions to the Sharma–Tasso– Olver equation. Applied Mathematics and Computation, 188(2):1205–1213, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014640.

Wazwaz:2007:MNC

[Waz07h] Abdul-Majid Wazwaz. The tanh–coth method for new compactons and solitons solutions for the K(n, n) and the K(n + 1, n + 1) equations. Applied Mathematics and Computation, 188(2):1930–1940, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601561X.

Wazwaz:2007:TCM

[Waz07i] Abdul-Majid Wazwaz. The tanh–coth method for solitons and kink solutions for nonlinear parabolic equations. Applied Mathematics and Computation, 188(2):1467–1475, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014998.

Wazwaz:2007:VIMb

[Waz07j] Abdul-Majid Wazwaz. The variational iteration method for a reliable treatment of the linear and the nonlinear Goursat problem. Applied Mathematics and Computation, 193(2):455–462, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004286.

Wazwaz:2007:VIMa

[Waz07k] Abdul-Majid Wazwaz. The variational iteration method for solving two forms of Blasius equation on a halfinfinite domain. Applied Mathematics and Computation, 188(1):485–491, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013889.

Wazwaz:2008:ASB

[Waz08a] Abdul-Majid Wazwaz. Analytic study on Burgers, Fisher, Huxley equations and combined forms of these equations. Applied Mathematics and Computation, 195(2):754–761, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006133.

Wazwaz:2008:EMS

[Waz08b] Abdul-Majid Wazwaz. Erratum to: “Multiplesoliton solutions for the ninth-order KdV equation and sixth-order Boussinesq equation” [Appl. Math. Comput. 203 (2008) 277– 283]. Applied Mathematics and Computation, 206(2):1005, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006759.

See [Waz08o].

Wazwaz:2008:HBM

[Waz08c] Abdul-Majid Wazwaz. The Hirota’s bilinear method and the tanh–coth method for multiple-soliton solutions of the Sawada– Kotera–Kadomtsev–Petviashvili equation. Applied Mathematics and Computation, 200(1):160–166, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030701096X.

Wazwaz:2008:HDMa

[Waz08d] Abdul-Majid Wazwaz. The Hirota’s direct method and the tanh–coth method for multiple-soliton solutions of the Sawada– Kotera–Ito seventh-order equation. Applied Mathematics and Computation, 199(1):133–138, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009794.

Wazwaz:2008:HDMb

[Waz08e] Abdul-Majid Wazwaz. The Hirota’s direct method for multiple-soliton solutions for three model equations of shallow water waves. Applied Mathematics and Computation, 201(1–2):489–503, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307012258.

Wazwaz:2008:IKE

[Waz08f] Abdul-Majid Wazwaz. The integrable KdV6 equations: Multiple soliton solutions and multiple singular soliton solutions. Applied Mathematics and Computation, 204(2):963–972, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308005870.

Wazwaz:2008:MFS

[Waz08g] Abdul-Majid Wazwaz. Multiplefront solutions for the Burgers–Kadomtsev–Petviashvili equation. Applied Mathematics and Computation, 200(1):437–443, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011496.

Wazwaz:2008:MKS

[Waz08h] Abdul-Majid Wazwaz. Multiple kink solutions and multiple singular kink solutions for the (2 + 1)- dimensional Burgers equations. Applied Mathematics and Computation, 204(2):817–823, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005663.

Wazwaz:2008:MSSh

[Waz08i] Abdul-Majid Wazwaz. Multiple soliton solutions and multiple singular soliton solutions for the (3 + 1)- dimensional Burgers equations. Applied Mathematics and Computation, 204(2):942–948, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005845.

Wazwaz:2008:MSSg

[Waz08j] Abdul-Majid Wazwaz. Multiplesoliton solutions for the Calogero–Bogoyavlenskii– Schiff, Jimbo–Miwa and YTSF equations. Applied Mathematics and Computation, 203(2):592–597, September 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300308002890.

Wazwaz:2008:MSSa

[Waz08k] Abdul-Majid Wazwaz. Multiplesoliton solutions for the fifth order Caudrey–Dodd– Gibbon (CDG) equation. Applied Mathematics and Computation, 197(2):719– 724, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008302.

Wazwaz:2008:MSSe

[Waz08l] Abdul-Majid Wazwaz. Multiplesoliton solutions for the generalized (1+1)-dimensional and the generalized (2 + 1)-dimensional Ito equations. Applied Mathematics and Computation, 202(2):840–849, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308001884.

Wazwaz:2008:MSSc

[Waz08m] Abdul-Majid Wazwaz. Multiplesoliton solutions for the Lax–Kadomtsev–Petviashvili (Lax–KP) equation. Applied Mathematics and Computation, 201(1–2): 168–174, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307011812.

Wazwaz:2008:MSSb

[Waz08n] Abdul-Majid Wazwaz. Multiplesoliton solutions for the Lax seventh-order equation. Applied Mathematics and Computation, 198 (2):877–881, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307009629.

Wazwaz:2008:MSSf

[Waz08o] Abdul-Majid Wazwaz. Multiplesoliton solutions for the ninth-order KdV equation and sixth-order Boussinesq equation. Applied Mathematics and Computation, 203(1):277–283, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002518.

See [Waz08b].

Wazwaz:2008:MSSd

[Waz08p] Abdul-Majid Wazwaz. Multiplesoliton solutions of two extended model equations for shallow water waves. Applied Mathematics and Computation, 201(1–2):790–799, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000477.

Wazwaz:2008:SSC

[Waz08q] Abdul-Majid Wazwaz. Nsoliton solutions for the combined KdV–CDG equation and the KdV–Lax equation. Applied Mathematics and Computation, 203(1):402–407, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002646.

Wazwaz:2008:NSD

[Waz08r] Abdul-Majid Wazwaz. New solutions of distinct physical structures to highdimensional nonlinear evolution equations. Applied Mathematics and Computation, 196(1):363–370, February 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307006893.

Wazwaz:2008:RSS

[Waz08s] Abdul-Majid Wazwaz. Regular soliton solutions and singular soliton solutions for the modified Kadomtsev– Petviashvili equations. Applied Mathematics and Computation, 204(1):227– 232, October 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308004980.

Wazwaz:2008:SMS

[Waz08t] Abdul-Majid Wazwaz. Single and multiple-soliton solutions for the (2 + 1)- dimensional KdV equation. Applied Mathematics and Computation, 204(1):20–26, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004347.

Wazwaz:2008:SWS

[Waz08u] Abdul-Majid Wazwaz. Solitary wave solutions of the generalized shallow water wave (GSWW) equation by Hirota’s method, tanh– coth method and Expfunction method. Applied Mathematics and Computation, 202(1):275–286, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000994.

Wazwaz:2008:SSS

[Waz08v] Abdul-Majid Wazwaz. Solitons and singular solitons for the Gardner– KP equation. Applied Mathematics and Computation, 204(1):162–169, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004888.

Wazwaz:2008:TCS

[Waz08w] Abdul-Majid Wazwaz. The tanh–coth and the sine– cosine methods for kinks, solitons, and periodic solutions for the Pochhammer– Chree equations. Applied Mathematics and Computation, 195(1):24–33, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005139.

Wazwaz:2009:DNE

[Waz09a] Abdul-Majid Wazwaz. A (3 + 1)-dimensional nonlinear evolution equation with multiple soliton solutions and multiple singular soliton solutions. Applied Mathematics and Computation, 215(4):1548–1552, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006559.

Wazwaz:2009:CHT

[Waz09b] Abdul-Majid Wazwaz. The Cole–Hopf transformation and multiple soliton solutions for the integrable sixth-order Drinfeld–Sokolov– Satsuma–Hirota equation. Applied Mathematics and Computation, 207(1):248– 255, January 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308007947.

Wazwaz:2009:FDI

[Waz09c] Abdul-Majid Wazwaz. Four (2 + 1)-dimensional integrable extensions of the KdV equation: Multiplesoliton and multiple singular soliton solutions. Applied Mathematics and Computation, 215(4):1463– 1476, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309006341.

Wazwaz:2009:MSS

[Waz09d] Abdul-Majid Wazwaz. Multiplesoliton solutions and multiplesingular soliton solutions for two higher-dimensional shallow water wave equations. Applied Mathematics and Computation, 211(2):495–501, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001088.

Wazwaz:2009:TST

[Waz09e] Abdul-Majid Wazwaz. Two systems of two-component integrable equations: Multiple soliton solutions and multiple singular soliton solutions. Applied Mathematics and Computation, 207(2):397–405, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008266.

Wazwaz:2009:VIM

[Waz09f] Abdul-Majid Wazwaz. The variational iteration method for analytic treatment for linear and nonlinear ODEs. Applied Mathematics and Computation, 212(1):120– 134, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001258.

Wang:2006:CAT

[WB06] Jing Wang and Zhong-Zhi Bai. Convergence analysis of two-stage waveform relaxation method for the initial value problems. Applied Mathematics and Computation, 172(2):797–808, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009130.

Wang:2009:PDI

[WB09] G. Q. Wang and Y. Q. Bai.

A primal-dual interiorpoint algorithm for secondorder cone optimization with full Nesterov–Todd step. Applied Mathematics and Computation, 215(3):1047–1061, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005980.

Wang:2005:NSD

[WC05a] QuanFang Wang and DaiZhan Cheng. Numerical solution of damped nonlinear Klein–Gordon equations using variational method and finite element approach. Applied Mathematics and Computation, 162(1):381–401, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400102X.

Wu:2005:CAC

[WC05b] Tzong-Mou Wu and Cha’o-Kuang Chen. Computeraided curvature analyses of planar four-bar linkage mechanism. Applied Mathematics and Computation, 168(2):1175–1188, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304006733.

Wu:2005:MMS

[WC05c] Tzong-Mou Wu and Chao-Kuang Chen. Mathematical model and its simulation of exactly mechanism synthesis with adjustable link. Applied Mathematics and Computation, 160(2):309–316, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030300866X.

Wu:2005:PMU

[WC05d] Wann-Yih Wu and Shuo-Pei Chen. A prediction method using the grey model GMC(1, n) combined with the grey relational analysis: a case study on Internet access population forecast. Applied Mathematics and Computation, 169(1):198–217, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007702.

Wang:2006:CBA

[WC06a] Kuo-Hsiung Wang and Liang-Wei Chiu. Cost benefit analysis of availability systems with warm standby units and imperfect coverage. Applied Mathematics and Computation, 172(2):1239–1256, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002900.

Wang:2006:GLA

[WC06b] Qi Wang and Yong Chen.

Generalized Q–S (lag, anticipated and complete) synchronization in modified Chua’s circuit and Hindmarsh–Rose systems. Applied Mathematics and Computation, 181(1):48– 56, October 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306000658.

Wang:2006:RFS

[WC06c] Xingyuan Wang and Peijun Chang. Research on fractal structure of generalized M–J sets utilized Lyapunov exponents and periodic scanning techniques. Applied Mathematics and Computation, 175(2):1007–1025, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006582.

Wang:2006:EMG

[WC06d] Ying-Ming Wang and Kwai-Sang Chin. An eigenvector method for generating normalized interval and fuzzy weights. Applied Mathematics and Computation, 181(2):1257–1275, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002463.

Wang:2007:BVP

[WC07a] Haihua Wang and Haibo Chen. Boundary value problem for second-order impulsive functional differential equations. Applied Mathematics and Computation, 191(2):582–591, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700286X.

Wang:2007:PSN

[WC07b] Haihua Wang and Haibo Chen. Positive solutions of a nonlinear secondorder n-point boundary value problem. Applied Mathematics and Computation, 186(2):1129–1136, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010393.

Wang:2008:ASD

[WC08a] Fei Wang and Xiao-Liang Cheng. An algorithm for solving the double obstacle problems. Applied Mathematics and Computation, 201(1–2):221–228, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011897.

Wu:2008:DSS

[WC08b] Hongling Wu and Caisheng Chen. Decaying solution for singular p-Laplacian equation. Applied Mathematics and Computation, 204(2):833–838, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005687.

Wang:2009:IEM

[WC09a] Hai-Jun Wang and De-Xin Cao. Interval expansion method for nonlinear equation in several variables. Applied Mathematics and Computation, 212(1):153– 161, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001283.

Wang:2009:CAA

[WC09b] Kuo-Hsiung Wang and Yu-Ju Chen. Comparative analysis of availability between three systems with general repair times, reboot delay and switching failures. Applied Mathematics and Computation, 215(1):384–394, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309004810.

Wu:2009:ETP

[WC09c] Xian Wu and Jilin Chen. Existence theorems of periodic solutions for a class of damped vibration problems. Applied Mathematics and Computation, 207(1):230–235, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007844.

Wang:2008:NMS

[WCH08a] Wenyan Wang, Minggen Cui, and Bo Han. A new method for solving a class of singular twopoint boundary value problems. Applied Mathematics and Computation, 206(2):721–727, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006905.

Wu:2008:RSC

[WCH08b] Lai-Yun Wu, Lap-Loi Chung, and Hsu-Hui Huang. Radial spline collocation method for static analysis of beams. Applied Mathematics and Computation, 201(1–2):184–199, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011848.

Wang:2008:SSC

[WCL08] Qing-Wen Wang, Hai-Xia Chang, and Chun-Yan Lin. P -(skew)symmetric common solutions to a pair of quaternion matrix equations. Applied Mathematics and Computation, 195(2):721–732, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006108.

Wang:2005:NET

[WCLZ05] Qi Wang, Yong Chen, Biao Li, and Hongqing Zhang. New exact travelling wave solutions for the shallow long wave approximate equations. Applied Mathematics and Computation, 160(1):77–88, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303009573.

Wang:2008:CSS

[WCN08] Qing-Wen Wang, Hai-Xia Chang, and Qun Ning. The common solution to six quaternion matrix equations with applications. Applied Mathematics and Computation, 198(1):209–226, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008806.

Wang:2008:DTM

[WCO08] Fulong Wang, Yuqing Chen, and Donal O’Regan.

Degree theory for (S_{+}) mappings in non-reflexive Banach spaces. Applied Mathematics and Computation, 202(1):229–232, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000866.

Wang:2009:GDC

[WCO09] Fulong Wang, Yuqing Chen, and Donal O’Regan. Generalized degree for classes of finite dimensional upper hemi-continuous mappings in separable Banach spaces. Applied Mathematics and Computation, 211(1):173–176, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000678.

Wang:2007:IAS

[WCW07a] Minghui Wang, Xuehan Cheng, and Musheng Wei. Iterative algorithms for solving the matrix equation AXB + CX^{T} D = E. Applied Mathematics and Computation, 187(2):622–629, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011179.

Wu:2007:DZS

[WCW07b] Hong-Wu Wu, Sui Sun Cheng, and Qi-Ru Wang. Distribution of zeros of solutions of functional differential equations. Applied Mathematics and Computation, 193(1):154–161, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307003827.

Wei:2005:NCP

[WCX05] Yimin Wei, Yanhua Cao, and Hua Xiang. A note on the componentwise perturbation bounds of matrix inverse and linear systems. Applied Mathematics and Computation, 169(2):1221–1236, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008471.

Wang:2006:GFE

[WCX06] Zhihua Wang, Xiaofeng Chen, and Bing Xu. Generalization of functional equation for the square root spiral. Applied Mathematics and Computation, 182(2):1355–1360, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004498.

Wei:2008:CMB

[WCX08] Zhou Wei, Zhou Changbin, and Chang Xiaolin. A contact model based on the SQP algorithm and engineering application. Applied Mathematics and Computation, 198(2):916– 924, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009587.

Wang:2006:GAM

[WCZ06] Qi Wang, Yong Chen, and Hongqing Zhang. Generalized algebraic method and new exact traveling wave solutions for (2 + 1)- dimensional dispersive long wave equation. Applied Mathematics and Computation, 181(1):247–255, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000944.

Wang:2009:LBC

[WCZ09] Zhaoxia Wang, Xingwu Chen, and Weinian Zhang. Local bifurcations of critical periods in a generalized 2D LV system. Applied Mathematics and Computation, 214(1):17–25, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002732.

Wang:2006:CRS

[WD06] Zhengsheng Wang and Hua Dai. On the construction of a real symmetric fivediagonal matrix from its three eigenpairs. Applied Mathematics and Computation, 175(1):597–608, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006338.

Wang:2008:EAD

[WD08] Yuping Wang and Chuangyin Dang. An evolutionary algorithm for dynamic multi-objective optimization. Applied Mathematics and Computation, 205(1):6–18, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003135.

Wang:2006:CRA

[WDK06] Kuo-Hsiung Wang, Wen-Li Dong, and Jyh-Bin Ke. Comparison of reliability and the availability between four systems with warm standby components and standby switching failures. Applied Mathematics and Computation, 183(2):1310–1322, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006783.

Wang:2007:CHE

[WDL07] Lingyun Wang, Feng Ding, and Peter X. Liu. Convergence of HLS estimation algorithms for multivariable ARX-like systems. Applied Mathematics and Computation, 190(2):1081–1093, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001282.

Wang:2010:CNE

[WDLM10] Chuanjian Wang, Zhengde Dai, Songqing Lin, and Gui Mu. Comments on “New exact kink solutions, solitons and periodic form solutions for a combined KdV and Schwarzian KdV equation”. Applied Mathematics and Computation, 216(2):361–363, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [Li09g].

Wei:2006:DIS

[WDN06] Yimin Wei, Huaian Diao, and Michael K. Ng. On Drazin inverse of singular Toeplitz matrix. Applied Mathematics and Computation, 172(2):809–817, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304009142.

White:2007:MEU

[WdRS07] S. Hoya White, A. Martín del Rey, and G. Rodríguez Sánchez. Modeling epidemics using cellular automata. Applied Mathematics and Computation, 186(1):193–202, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306009295.

Wu:2006:SME

[WDY06] Ai-Guo Wu, Guang-Ren Duan, and Hai-Hua Yu. On solutions of the matrix equations XF − AX = C and XF − AX̄ = C. Applied Mathematics and Computation, 183(2):932–941, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006266.

Wei:2006:NSC

[Wei06a] Zhongli Wei. A necessary and sufficient condition for the existence of positive solutions of singular superlinear m-point boundary value problems. Applied Mathematics and Computation, 179(1):67–78, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010076.

Wei:2006:PST

[Wei06b]  Zhongli Wei.  Positive

solutions for 2 n thorder singular sub-linear m-point boundary value problems. Applied Mathematics and Computation, 182(2):1280–1295, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004437.

Wei:2009:CNS

[Wei09]  Mao Wei.  Conver-

gence of numerical solutions for variable delay differential equations driven by Poisson random jump measure. Applied Mathematics and Computation, 212(2):409–417, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001738.

Welfert:2008:QFB

[Wel08] B. D. Welfert. On quadrature formulae based on derivative collocation. Applied Mathematics and Computation, 204(2):647– 657, October 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308000635.

Wenhui:2005:ASS

[Wen05]  Zhou Wenhui.  Analy-

sis of a single-server retrial queue with FCFS orbit and Bernoulli vacation. Applied Mathematics and Computation, 161(2):353–364, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012906.

Weyburne:2006:MDF

[Wey06] David W. Weyburne. A mathematical description of the fluid boundary layer. Applied Mathematics and Computation, 175(2):1675–1684, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500723X.

See erratum [Wey08].

Weyburne:2008:EMD

[Wey08] David W. Weyburne. Erratum to “A mathematical description of the fluid boundary layer” [Appl. Math. Comput. 175 (2) (2006) 1675–1684]. Applied Mathematics and Computation, 197(1):466, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007096.

See [Wey06].

Wang:2007:GDA

[WF07a] Ying-Ming Wang and Zhi-Ping Fan. Group decision analysis based on fuzzy preference relations: Logarithmic and geometric least squares methods. Applied Mathematics and Computation, 194(1):108–119, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004493.

Wu:2007:NES

[WF07b] Hongxia Wu and Tianyou Fan. New explicit solutions of the nonlinear wave equations with damping term. Applied Mathematics and Computation, 191(2):457–465, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002743.

Wu:2007:NDE

[WF07c] Hongxia Wu and Tianyou Fan. Nonlinear dispersive ZK(n, n) equations: New compacton solutions and solitary pattern solutions. Applied Mathematics and Computation, 187(2):1308–1314, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012690.

Wu:2007:WIR

[WF07d] Xinyuan Wu and Yonglei Fang. Wilkinson’s iterative refinement of solution with automatic step-size control for linear system of equations. Applied Mathematics and Computation, 193(2):506–513, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700433X.

Wu:2009:CFS

[WFHD09] Ai-Guo Wu, Gang Feng, Junqiang Hu, and Guang-Ren Duan. Closed-form solutions to the nonhomogeneous Yakubovichconjugate matrix equation. Applied Mathematics and Computation, 214(2):442–450, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003294.

Wang:2006:MBP

[WFL06] Lin-Lin Wang, Yong-Hong Fan, and Wan-Tong Li. Multiple bifurcations in a predator–prey system with monotonic functional response. Applied Mathematics and Computation, 172(2):1103–1120, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002808.

Wang:2008:OFO

[WFS08] Rong Wang, Hui Feng, and Raymond J. Spiteri. Observations on the fifth-order WENO method with non-uniform meshes. Applied Mathematics and Computation, 196(1):433–447, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006972.

Wang:2007:VME

[WFX07] Gang Wang, Enmin Feng, and Zhilong Xiu. Vector measure for explicit nonlinear impulsive system of glycerol bioconversion in fed-batch cultures and its parameter identification. Applied Mathematics and Computation, 188(2):1151–1160, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014597.

Wang:2005:CNR

[WG05] Guorong Wang and Chao Gu. Condition number related with W -weighted Drazin inverse and singular linear systems. Applied Mathematics and Computation, 162(1):435–446, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001079.

Gao:2007:EBS

[wG07a] Pei wang Gao. An efficient bound-and-stopped algorithm for integer linear programs on the objective function hyperplane. Applied Mathematics and Computation, 185(1):301–311, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008617.

Wei:2007:EUL

[WG07b] Yingjie Wei and Wenjie Gao. Existence and uniqueness of local solutions to a class of quasilinear degenerate parabolic systems. Applied Mathematics and Computation, 190(2):1250–1257, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001555.

Walczak:2008:NCB

[WG08] Zbigniew Walczak and Vijay Gupta. A note on the convergence of Baskakov type operators. Applied Mathematics and Computation, 202(1):370–375, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308001215.

Weng:2005:PSM

[WGCH05] Shiyou Weng, Haiyin Gao, Renzhao Chen, and Xuezhang Hou. Penalty shifting method on calculation of optimal linewise control of dam-detouring osimotic system for concrete dams with heterogemous. Applied Mathematics and Computation, 169(2):1129–1141, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008410.

Wang:2006:ESS

[WGF^{+}06] Jianfeng Wang, Gao Ge, Yubo Fan, Lilin Chen, Shengxue Liu, Yali Jin, and Jinguo Yu. The estimation of sample size in multi-stage sampling and its application in medical survey. Applied Mathematics and Computation, 178(2):239–249, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009707.

Wang:2008:ABS

[WGL08] Guolian Wang, Boling Guo, and Yangrong Li. The asymptotic behavior of the stochastic Ginzburg– Landau equation with additive noise. Applied Mathematics and Computation, 198(2):849–857, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009575.

Wang:2006:RCH

[WH06a] Zhu Wang and Bing Hu.

Research of combined hybrid method applied in the Reissner–Mindlin plate model. Applied Mathematics and Computation, 182(1):49–66, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003122.

Wu:2006:FIM

[WH06b] Peng Wu and Danfu Han.

A family of iterative methods with higher-order convergence. Applied Mathematics and Computation, 182(1):474–477, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003547.

Wang:2008:UGA

[WH08a] Chao-Hung Wang and Li-Chang Hsu. Using genetic algorithms grey theory to forecast high technology industrial output. Applied Mathematics and Computation, 195(1):256–263, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700536X.

Wang:2008:SMB

[WH08b] Chi-Chang Wang and Hai-Ping Hu. Study on monotonicity of boundary value problems of differential equations. Applied Mathematics and Computation, 202(1):383–394, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308001434.

Wu:2009:ECP

[WH09] Shi-Liang Wu and Ting-Zhu Huang. Erratum to: “Convergence of the preconditioned AOR method for irreducible L-matrices” [Appl. Math. Comput. 201 (2008) 56–64]. Applied Mathematics and Computation, 212(2):551–552, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001726.

See [YK08].

Wang:2008:UDE

[WHC08] Xue-Song Wang, Ming-Lin Hao, and Yu-Hu Cheng. On the use of differential evolution for forward kinematics of parallel manipulators. Applied Mathematics and Computation, 205(2):760–769, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003548.

Wang:2008:CPL

[WHD^{+}08] Xiao-Feng Wang, De-Shuang Huang, Ji-Xiang Du, Huan Xu, and Laurent Heutte. Classification of plant leaf images with complicated background. Applied Mathematics and Computation, 205(2):916–926, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003718.

Wen:2005:URE

[WHH05] Hsiang-An Wen, Sheng-Yu Hwang, and Tzonelih Hwang. On the unlinkability of randomizationenhanced Chaum’s blind signature scheme. Applied Mathematics and Computation, 164(3):799–803, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003583.

Wang:2007:EGB

[WHH07] Yi Wang, Jiankun Hu, and Fengling Han. Enhanced gradient-based algorithm for the estimation of fingerprint orientation fields. Applied Mathematics and Computation, 185(2):823–833, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007648.

White:2006:RTD

[Whi06]  Luther White.  A re-

duced three-dimensional dynamic structural model for structural health assessment. Applied Mathematics and Computation, 182(1):572–588, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003663.

Wang:2006:TGN

[WHL06] Cheng Wang, Ziping Huang, and Likang Li. Twogrid nonconforming finite element method for second order elliptic problems. Applied Mathematics and Computation, 177(1):211– 219, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009021.

Wu:2008:NWR

[WHL08] Shulin Wu, Chengming Huang, and Yong Liu. Newton waveform relaxation method for solving algebraic nonlinear equations. Applied Mathematics and Computation, 201(1–2):553–560, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307012301.

Wang:2006:NSDb

[WiN06] Quan-Fang Wang and Shin ichi Nakagiri. Numerical solutions of distributed diffusion Hopfield neural network equations. Applied Mathematics and Computation, 174(2):1343–1362, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005540.

Wang:2005:GMS

[WJ05] Xiao-Yun Wang and Yao-Lin Jiang. A general method for solving singular perturbed impulsive differential equations with twopoint boundary conditions. Applied Mathematics and Computation, 171(2):775– 806, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2006:FEA

[WJ06] Xiang Wang and Zhou Jituan. A fast eigenvalue algorithm for Pascal matrices. Applied Mathematics and Computation, 183(2):711–716, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006059.

Wang:2008:VSC

[WJZG08] Youqing Wang, Canghua Jiang, Donghua Zhou, and Furong Gao. Variable structure control for a class of nonlinear systems with mismatched uncertainties. Applied Mathematics and Computation, 200(1):387–400, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011435.

Wang:2008:VJM

[WKL08] Xiuhua Wang, Jisheng Kou, and Yitian Li. A variant of Jarratt method with sixth-order convergence. Applied Mathematics and Computation, 204(1):14–19, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004128.

Wang:2005:DMG

[WL05a] Weiming Wang and Zhenqing Li. The distribution models of grazing animals between two grassland resource points. Applied Mathematics and Computation, 169(2):1395–1404, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008604.

Wang:2005:CMR

[WL05b] Weiming Wang and Xinze Lian. Computations of multi-resultant with mechanization. Applied Mathematics and Computation, 170(1):237–257, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304009221.

Wang:2005:NAS

[WL05c] Weiming Wang and Xinze Lian. A new algorithm for symbolic integral with application. Applied Mathematics and Computation, 162(2):949–968, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001419.

Wang:2005:NAI

[WL05d] Weiming Wang and Changsheng Lin. A new algorithm for integral of trigonometric functions with mechanization. Applied Mathematics and Computation, 164(1):71– 82, May 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002966.

Wang:2005:DGO

[WL05e] Yanjun Wang and Zhian Liang. A deterministic global optimization algorithm for generalized geometric programming. Applied Mathematics and Computation, 168(1):722–737, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001839.

Wen:2005:NSL

[WL05f] Liping Wen and Shoufu Li. Nonlinear stability of linear multistep methods for stiff delay differential equations in Banach spaces. Applied Mathematics and Computation, 168(2):1031–1044, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006617.

White:2005:MBD

[WL05g] Luther W. White and Yiqi Luo. Model-based data assessment for terrestrial carbon processes: implications for sampling strategy in FACE experiments. Applied Mathematics and Computation, 167(1):419–434, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005028.

Wang:2006:STTa

[WL06a] Peiguang Wang and Hairong Lian. Stability in terms of two measures of impulsive integro-differential equations via variation of the Lyapunov method. Applied Mathematics and Computation, 177(1):387–395, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009197.

Wang:2006:MASa

[WL06b] Weiming Wang and Zhenqing Li. A mechanical algorithm for solving ordinary differential equation. Applied Mathematics and Computation, 172(1):568–583, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305002249.

Wang:2006:FAS

[WL06c] Xiang Wang and Linzhang Lu. A fast algorithm for solving linear systems of the Pascal type. Applied Mathematics and Computation, 175(1):441–451, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006193.

Wang:2006:GJS

[WL06d] Xingyuan Wang and Chao Luo. Generalized Julia sets from a non-analytic complex mapping. Applied Mathematics and Computation, 181(1):113–122, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000828.

Wang:2006:DEA

[WL06e] Ying-Ming Wang and Ying Luo. DEA efficiency assessment using ideal and anti-ideal decision making units. Applied Mathematics and Computation, 173(2):902–915, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004017.

Wei:2006:HON

[WL06f] T. Wei and M. Li. High order numerical derivatives for one-dimensional scattered noisy data. Applied Mathematics and Computation, 175(2):1744–1759, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305007368.

Wu:2006:CAL

[WL06g] Jong-Wuu Wu and Wen-Chuan Lee. Computational algorithm of least absolute deviation method for determining number of outliers under normality. Applied Mathematics and Computation, 175(1):609– 617, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030500634X.

Wang:2007:EEN

[WL07a] Deng-Shan Wang and Hong-Bo Li. Elliptic equation’s new solutions and their applications to two nonlinear partial differential equations. Applied Mathematics and Computation, 188(1):762–771, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014147.

Wang:2007:ESI

[WL07b] Qing Wang and Xinzhi Liu. Exponential stability of impulsive cellular neural networks with time delay via Lyapunov functionals. Applied Mathematics and Computation, 194(1):186–198, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004572.

Wang:2007:OLM

[WL07c] Wan-Sheng Wang and Shou-Fu Li. On the oneleg θ-methods for solving nonlinear neutral functional differential equations. Applied Mathematics and Computation, 193(1):285–301, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003979.

Wang:2007:DCG

[WL07d] Xiao Wang and Zhixiang Li. Dynamics for a class of general hematopoiesis model with periodic coefficients. Applied Mathematics and Computation, 186(1):460– 468, March 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306009544. See

erratum [WL08a].

Wang:2007:GDB

[WL07e] Xiao Wang and Zhixiang Li. Globally dynamical behaviors for a class of nonlinear functional difference equation with almost periodic coefficients. Applied Mathematics and Computation, 190(2):1116–1124, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001312.

Wang:2007:NIS

[WL07f] Xin Wang and Zhiping Li. A numerical iterative scheme for computing finite order rank-one convex envelopes. Applied Mathematics and Computation, 185(1):19–30, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008009.

Wang:2008:EDC

[WL08a] Xiao Wang and Zhixiang Li. Erratum to: “Dynamics for a class of general hematopoiesis model with periodic coefficients” [Appl. Math. Comput. 186 (2007) 460–468]. Applied Mathematics and Computation, 200(1):473–476, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010636.

See [WL07d].

Whang:2008:ABC

[WL08b] Sungim Whang and Sunmi Lee. Absorbing boundary conditions for the stationary forced KdV equation. Applied Mathematics and Computation, 202(2):511–519, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001264.

White:2008:MIN

[WL08c] Luther White and Yiqi Luo. Modeling and inversion of net ecological exchange data using an Ito stochastic differential equation approach. Applied Mathematics and Computation, 196(2):686–704, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007229.

Wang:2009:GES

[WL09a] Lingna Wang and Yiping Lin. Global exponential stability for shunting inhibitory CNNs with delays. Applied Mathematics and Computation, 214(1):297–303, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030900304X.

Wang:2009:TNF

[WL09b] Xia Wang and Liping Liu. Two new families of sixthorder methods for solving non-linear equations. Applied Mathematics and Computation, 213(1):73– 78, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002033.

Wang:2009:NSI

[WL09c] Xing Tao Wang and Yuan Min Li. Numerical solutions of integrodifferential systems by hybrid of general block-pulse functions and the second Chebyshev polynomials. Applied Mathematics and Computation, 209(2):266–272, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308009454.

Wang:2009:NDP

[WLA09] Fanglei Wang, Wanjun Li, and Yukun An. Nonnegative doubly periodic solutions for nonlinear telegraph system with twinparameters. Applied Mathematics and Computation, 214(1):310–317, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003063.

Wu:2005:PIF

[WLC05] Jong-Wuu Wu, Wen-Chuan Lee, and Sheau-Chiann Chen. Prediction intervals of future observation from one-parameter exponential distribution based on multiply type II censored samples. Applied Mathematics and Computation, 167(2):741–806, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005272.

Wu:2006:CCW

[WLC06] Jong-Wuu Wu, Wen-Chuan Lee, and Sheau-Chiann Chen. Computational comparison for weighted moments estimators and BLUE of the scale parameter of a Pareto distribution with known shape parameter under type II multiply censored sample. Applied Mathematics and Computation, 181(2):1462–1470, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600275X.

Wu:2007:FBM

[WLC07a] Chia-Long Wu, Der-Chyuan Lou, and Te-Jen Chang. Fast binary multiplication by performing dot counting and complement recoding. Applied Mathematics and Computation, 191(1):132–139, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002366.

Wu:2007:CCPa

[WLC07b] Jong-Wuu Wu, Wen-Chuan Lee, and Sheau-Chiann Chen. Computational comparison of prediction future lifetime of electronic components with Pareto distribution based on multiply type II censored samples. Applied Mathematics and Computation, 184(2):374–406, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007144.

Wu:2007:CCPb

[WLC07c] Jong-Wuu Wu, Wen-Chuan Lee, and Sheau-Chiann Chen. Computational comparison of prediction intervals of future observation for twoparameter exponential distribution. Applied Mathematics and Computation, 184(2):1084–1117, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007995.

Wu:2007:CCPc

[WLC07d] Jong-Wuu Wu, Wen-Chuan Lee, and Sheau-Chiann Chen. Computational comparison of the prediction intervals of future observation for three-parameter Pareto distribution with known shape parameter. Applied Mathematics and Computation, 190(1):150–178, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000379.

Wong:2006:FES

[WLCL06] K. W. Wong, Edward C. W. Lee, L. M. Cheng, and Xiaofeng Liao. Fast elliptic scalar multiplication using new doublebase chain and point halving. Applied Mathematics and Computation, 183(2):1000–1007, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006357.

Wang:2009:EAA

[WLH09] Chi-Chang Wang, David. T. W. Lin, and Hai-Ping Hu. Error analysis of approximate solutions of non-linear Burgers’ equation. Applied Mathematics and Computation, 212(2):387–395, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001702.

Wen:2005:TRB

[WLHH05] Hsiang-An Wen, Kuo-Chang Lee, Sheng-Yu Hwang, and Tzonelih Hwang. On the traceability on RSA-based partially signature with low computation. Applied Mathematics and Computation, 162(1):421–425, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001043.

Wang:2008:ARC

[WLK08] Chi-Chang Wang, Zong-Yi Lee, and Yiyo Kuo. Application of residual correction method in calculating upper and lower approximate solutions of fifth-order boundary-value problems. Applied Mathematics and Computation, 199(2):677–690, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307010557.

Wu:2007:CTA

[WLL07a] Jong-Wuu Wu, Hsiu-Mei Lee, and Chia-Ling Lei. Computational testing algorithmic procedure of assessment for lifetime performance index of products with two-parameter exponential distribution. Applied Mathematics and Computation, 190(1):116– 125, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000343.

Wu:2007:MFD

[WLL07b] Yan Wu, Ming Li, and Guisheng Liao. Multiple features data fusion method in color texture analysis. Applied Mathematics and Computation, 185(2):784–797, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007612.

Wu:2009:GCA

[WLL09] H. X. Wu, H. Z. Luo, and S. L. Li. The global convergence of augmented Lagrangian methods based on NCP function in constrained nonconvex optimization. Applied Mathematics and Computation, 207(1):124–134, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308007716.

Wu:2007:FMM

[WLLC07] Chia-Long Wu, Der-Chyuan Lou, Jui-Chang Lai, and Te-Jen Chang. Fast modular multi-exponentiation using modified complex arithmetic. Applied Mathematics and Computation, 186(2):1065–1074, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010320.

Wang:2007:AAC

[WLLL07] Feng Wang, Yuanxiang Li, Li Li, and Kangshun Li. Automated analog circuit design using two-layer genetic programming. Applied Mathematics and Computation, 185(2):1087–1097, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008320.

Wang:2007:TRN

[WLP07] Xinyu Wang, Junwei Lei, and Changpeng Pan. Trigonometric RBF neural robust controller design for a class of nonlinear system with linear input unmodeled dynamics. Applied Mathematics and Computation, 185(2):989–1002, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600823X.

Wei:2006:NNC

[WLQ06a] Zengxin Wei, Guoyin Li, and Liqun Qi. New nonlinear conjugate gradient formulas for largescale unconstrained optimization problems. Applied Mathematics and Computation, 179(2):407–430, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010507.

Wei:2006:NQN

[WLQ06b] Zengxin Wei, Guoyin Li, and Liqun Qi. New quasi-Newton methods for unconstrained optimization problems. Applied Mathematics and Computation, 175(2):1156–1188, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006685.

Wang:2008:ESJ

[WLS08] Pin-Yu Wang, Shy-Der Lin, and H. M. Srivastava. Explicit solutions of Jacobi and Gauss differential equations by means of operators of fractional calculus. Applied Mathematics and Computation, 199(2):760–769, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307010661.

Wen:2009:NES

[WLS09] Zhenshu Wen, Zhengrong Liu, and Ming Song. New exact solutions for the classical Drinfel’d–Sokolov– Wilson equation. Applied Mathematics and Computation, 215(6):2349–2358, November 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309007590.

Wang:2005:ORM

[WLT05a] Ling Wang, Ling-Lai Li, and Fang Tang. Optimal reduction of models using a hybrid searching strategy. Applied Mathematics and Computation, 168(2):1357–1369, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007349.

Wang:2005:SAI

[WLT05b] Yingjie Wang, J. H. Li, and L. Tie. Security analysis and improvement of a user-friendly remote authentication protocol. Applied Mathematics and Computation, 168(1):47– 50, September 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304005843.

Wang:2007:SFC

[WLT07a] Pin-Yu Wang, Shy-Der Lin, and Shih-Tong Tu.

A survey of fractionalcalculus approaches to the solutions of the Bessel differential equation of general order. Applied Mathematics and Computation, 187(1):544–555, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012069.

Wu:2007:CAP

[WLT07b] Jong-Wuu Wu, Wen-Chuan Lee, and Hui-Yin Tsai. Computational algorithmic procedure of optimal inventory policy involving a negative exponential crashing cost and variable lead time demand. Applied Mathematics and Computation, 184(2):798–808, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007594.

Wang:2008:PSS

[WLW08] Ying Wang, Lishan Liu, and Yonghong Wu. Positive solutions of singular boundary value problems on the half-line. Applied Mathematics and Computation, 197(2):789–796, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008387.

Wang:2009:KTO

[WLW09] Peiguang Wang, Ping Li, and Yonghong Wu. k th Order convergence for semilinear elliptic boundary value problems. Applied Mathematics and Computation, 214(1):4–9, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002197.

White:2005:CSI

[WLX05] L. W. White, Y. Luo, and T. Xu. Carbon sequestration: inversion of FACE data and prediction. Applied Mathematics and Computation, 163(2):783–800, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002607.

Wu:2006:CCM

[WLY06] Jong-Wuu Wu, Wen-Chuan Lee, and Chi-Chin Yang. Computational comparison for ML estimator of quadratic functions of the Bernoulli parameter in IS and FSS methods.

Applied Mathematics and Computation, 176(1):317– 323, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007964.

Wei:2008:SCN

[WLY08] Zengxin Wei, Liying Liu, and Shengwei Yao. The superlinear convergence of a new quasi-Newton-SQP method for constrained optimization. Applied Mathematics and Computation, 196(2):791–801, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700731X.

Wang:2005:SCF

[WLZ05] Deng Shan Wang, Yi Fang Liu, and Hong-Qing Zhang. Symbolic computation and families of Jacobi elliptic function solutions of the (2 + 1)-dimensional Nizhnik–Novikov–Veselov equation. Applied Mathematics and Computation, 168(2):823–847, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006423.

Wu:2006:NAA

[WLZ06] Fengjie Wu, Jie Lu, and Guangquan Zhang. A new approximate algorithm for solving multiple objective linear programming problems with fuzzy parameters. Applied Mathematics and Computation, 174(1):524–544, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004832.

Wang:2008:NSA

[WLZ08] B. Wang, X. Liu, and S. Zhong. New stability analysis for uncertain neutral system with timevarying delay. Applied Mathematics and Computation, 197(1):457–465, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007564.

Wang:2008:NSR

[WLZN08] Xiang Wang, Linzhang Lu, Chuanxi Zhu, and Qiang Niu. A note on the spectral for real Hankel matrices. Applied Mathematics and Computation, 206(2):994–996, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308007005.

Wang:2007:BSS

[WM07] Ying Wang and Chunlai Mu. Blow-up and scattering of solution for a generalized Boussinesq equation. Applied Mathematics and Computation, 188(2):1131–1141, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014573.

Wang:2009:FSM

[WM09] Yanfei Wang and Shiqian Ma. A fast subspace method for image deblurring. Applied Mathematics and Computation, 215(6):2359–2377, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309007607.

Wu:2007:FFM

[WMBY07] Z. Y. Wu, M. Mammadov, F. S. Bai, and Y. J. Yang. A filled function method for nonlinear equations. Applied Mathematics and Computation, 189(2):1196–1204, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017036.

Wu:2008:SCM

[WMC08] Fuke Wu, Xuerong Mao, and Kan Chen. Strong convergence of Monte Carlo simulations of the meanreverting square root process with jump. Applied Mathematics and Computation, 206(1):494–505, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007121.

Wang:2005:DMF

[WML05] Jianzhou Wang, Zhixin Ma, and Lian Li. Detection, mining and forecasting of impact load in power load forecasting. Applied Mathematics and Computation, 168(1):29– 39, September 1, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030400582X.

Wang:2007:IAM

[WML07] Jizhong Wang, Fanwei Meng, and Sanyang Liu. Integral average method for oscillation of second order partial differential equations with delays. Applied Mathematics and Computation, 187(2):815–823, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012148.

Wang:2007:BSP

[WMX07] Yulan Wang, Chunlai Mu, and Zhaoyin Xiang. Blowup of solutions to a porous medium equation with nonlocal boundary condition. Applied Mathematics and Computation, 192(2):579–585, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307003645.

Wang:2008:NMM

[WN08a] Haijun Wang and Qin Ni. A new method of moving asymptotes for large-scale unconstrained optimization. Applied Mathematics and Computation, 203(1): 62–71, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030800194X.

Withers:2008:FEN

[WN08b] Christopher S. Withers and Saralees Nadarajah. Fredholm equations for nonsymmetric kernels with applications to iterated integral operators. Applied Mathematics and Computation, 204(1):499–507, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004840.

Withers:2008:MBP

[WN08c] Christopher S. Withers and Saralees Nadarajah. Multivariate Bell polynomials and their applications to powers and fractionary iterates of vector power series and to partial derivatives of composite vector functions. Applied Mathematics and Computation, 206(2):997–1004, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007157.

Woods:2009:OC

[WN09a] Damien Woods and Thomas J.

Naughton. Optical computing. Applied Mathematics and Computation, 215(4):1417–1430, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309004238.

Wozny:2009:MSS

[WN09b] Paweł Woźny and Rafał Nowak. Method of summation of some slowly convergent series. Applied Mathematics and Computation, 215(4):1622–1645, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030900664X.

Wang:2006:MTC

[WNCL06] Ji-Bo Wang, C. T. Daniel Ng, T. C. E. Cheng, and Li-Li Liu. Minimizing total completion time in a two-machine flow shop with deteriorating jobs. Applied Mathematics and Computation, 180(1):185–193, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305011124.

Woodbury:2005:SCE

[Woo05] Greg Woodbury. Sufficient conditions for equilibrium in dynamical systems. Applied Mathematics and Computation, 160(1):89–92, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303009937.

Wei:2005:PSS

[WP05] Zhongli Wei and Changci Pang. Positive solutions of some singular m-point boundary value problems at non-resonance. Applied Mathematics and Computation, 171(1):433–449, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001074.

Wu:2006:TWS

[WP06] Liping Wu and Chunping Pang. Travelling wave solutions for the generalized coupled Hirota–Satsuma KdV system. Applied Mathematics and Computation, 175(1):101–118, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005953.

Wang:2009:SOL

[WQL09] Wansheng Wang, Tingting Qin, and Shoufu Li. Stability of one-leg θ-methods for nonlinear neutral differential equations with proportional delay. Applied Mathematics and Computation, 213(1):177–183, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002161.

Wu:2006:CBM

[WR06] Qingbiao Wu and Hongmin Ren. Convergence ball of a modified secant method for finding zero of derivatives. Applied Mathematics and Computation, 174(1):24– 33, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030500439X.

Wang:2007:CSE

[WR07a] Zhengping Wang and Lizhi Ruan. On a class of semilinear elliptic problems with singular nonlinearities. Applied Mathematics and Computation, 193(1):89–105, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307003761.

Wu:2007:NSN

[WR07b] Qingbiao Wu and Hongmin Ren. A note on some new iterative methods with third-order convergence. Applied Mathematics and Computation, 188(2):1790–1793, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015463.

Wu:2007:CBE

[WRB07] Qingbiao Wu, Hongmin Ren, and Weihong Bi. Convergence ball and error analysis of Müller’s method. Applied Mathematics and Computation, 184(2):464–470, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007211.

Wu:2006:ESP

[WRL06] Chen Wu, Jiong Ruan, and Wei Lin. On the existence and stability of the periodic solution in the Cohen–Grossberg neural network with time delay and high-order terms. Applied Mathematics and Computation, 177(1):194– 210, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500901X.

Wang:2005:UPI

[WS05a] Jinting Wang and Raj Srinivasan. Unreliable production-inventory systems with hyper-exponential renewal demand processes. Applied Mathematics and Computation, 166(2):475– 488, July 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304004527.

White:2005:OWF

[WS05b] Luther White and Alan Shapiro. Optimization of wind field retrieval procedures. Applied Mathematics and Computation, 171(1):25–52, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000780.

Wu:2006:NFD

[WS06a] Qingbiao Wu and Airong Sheng. A note on finite difference method to analysis an ill-posed problem. Applied Mathematics and Computation, 182(2):1040–1047, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004188.

Wu:2006:QNB

[WS06b] Ting Wu and Linping Sun.

A quasi-Newton based pattern search algorithm for unconstrained optimization. Applied Mathematics and Computation, 183(1):685–694, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006023.

Wang:2007:PSD

[WS07a] Changzhong Wang and Jinlin Shi. Periodic solution for a delay multispecies Logarithmic population model with feedback control. Applied Mathematics and Computation, 193(1):257–265, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003955.

Wang:2007:PAP

[WS07b] Changzhong Wang and Jinlin Shi. Positive almost periodic solutions of a class of Lotka–Volterra type competitive system with delays and feedback controls. Applied Mathematics and Computation, 193(1):240–252, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307003931.

Wang:2007:HAS

[WS07c] Haixing Wang and Jinsheng Shen. Heuristic approaches for solving transit vehicle scheduling problem with route and fueling time constraints. Applied Mathematics and Computation, 190(2):1237–1249, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001543.

Wang:2007:PSM

[WS07d] Weibing Wang and Jianhua Shen. Positive solutions to a multi-point boundary value problem with delay. Applied Mathematics and Computation, 188(1):96– 102, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306013257.

Wang:2007:GSP

[WS07e] Xia Wang and Xinyu Song. Global stability and periodic solution of a model for HIV infection of CD4 ^{+} T cells. Applied Mathematics and Computation, 189(2):1331–1340, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601719X.

White:2007:ORS

[WS07f] Luther White and Alan Shapiro. Optimization of radar scanning strategies using an ensemble relative error criterion. Applied Mathematics and Computation, 188(1):693–712, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014081.

White:2007:RNS

[WS07g] Luther White and Alan Shapiro. Radar network scanning coordination based on ensemble transform Kalman filtering variance optimization. Applied Mathematics and Computation, 188(2):1285–1309, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014846.

Witula:2007:CDI

[WS07h] Roman Wituła and Damian Słota. On computing the determinants and inverses of some special type of tridiagonal and constantdiagonals matrices. Applied Mathematics and Computation, 189(1):514–527, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016250.

Wu:2007:SIG

[WS07i] Shan-He Wu and H. M. Srivastava. Some improvements and generalizations of Steffensen’s integral inequality. Applied Mathematics and Computation, 192(2):422–428, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003451.

Wang:2008:GOA

[WS08a] Chun-Feng Wang and Pei-Ping Shen. A global optimization algorithm for linear fractional programming. Applied Mathematics and Computation, 204(1):281–287, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005055.

Weng:2008:OSO

[WS08b] Aizhi Weng and Jitao Sun. Oscillation of second order delay differential equations. Applied Mathematics and Computation, 198(2):930–935, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009654.

Witula:2008:PFD

[WS08c] Roman Wituła and Damian Słota. Partial fractions decompositions of some rational functions. Applied Mathematics and Computation, 197(1):328–336, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007801.

Witula:2008:SPP

[WS08d] Roman Wituła and Damian Słota. Some phenomenon of the powers of certain tridiagonal and asymmetric matrices. Applied Mathematics and Computation, 202(1):348–359, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001185.

Wu:2008:FRJ

[WS08e] Shan-He Wu and H. M. Srivastava. A further refinement of a Jordan type inequality and its application. Applied Mathematics and Computation, 197(2):914–923, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008612.

Wang:2009:CBL

[WS09a] Zuo-Lei Wang and Xue-Rong Shi. Chaotic bursting lag synchronization of Hindmarsh–Rose system via a single controller. Applied Mathematics and Computation, 215(3):1091–1097, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006158.

Wei:2009:VAD

[WS09b] Han Wei and Zhang Suying. A variational approach to (m + 1)-dimensional n-order field. Applied Mathematics and Computation, 210(1):169–176, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009879.

Weng:2009:ISS

[WS09c] Aizhi Weng and Jitao Sun. Impulsive stabilization of second-order nonlinear delay differential systems. Applied Mathematics and Computation, 214(1):95–101, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002847.

Wang:2006:SOB

[WSB06] Bin Wang, Yangfeng Su, and Zhaojun Bai. The second-order biorthogonalization procedure and its application to quadratic eigenvalue problems. Applied Mathematics and Computation, 172(2):788– 796, January 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304009129.

Wauer:2007:DML

[WSCL07] J. Wauer, D. Schwarzer, G. Q. Cai, and Y. K. Lin. Dynamical models of love with time-varying fluctuations. Applied Mathematics and Computation, 188(2):1535–1548, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015062.

Wang:2006:PFS

[WSJW06] Ji-Bo Wang, Feng Shan, Bo Jiang, and Li-Yan Wang. Permutation flow shop scheduling with dominant machines to minimize discounted total weighted completion time. Applied Mathematics and Computation, 182(1):947–954, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004103.

Wang:2007:NER

[WSKZ07] Dan Wang, Weiwei Sun, Cuicui Kong, and Hongqing Zhang. New extended rational expansion method and exact solutions of Boussinesq equation and Jimbo–Miwa equations. Applied Mathematics and Computation, 189(1):878– 886, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016663.

Wang:2007:ERS

[WSL07] Qing-Wen Wang, Guang-Jing Song, and Chun-Yan Lin. Extreme ranks of the solution to a consistent system of linear quaternion matrix equations with an application. Applied Mathematics and Computation, 189(2):1517–1532, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017395.

Wang:2008:RER

[WSL08] Qing-Wen Wang, Guang-Jing Song, and Chun-Yan Lin. Rank equalities related to the generalized inverse A_{T,S}^{(2)} with applications. Applied Mathematics and Computation, 205(1):370–382, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800595X.

Wang:2009:STP

[WSL09] Yong-Gang Wang, Hui-Fang Song, and Dan Li. Solving two-point boundary value problems using combined homotopy perturbation method and Green’s function method. Applied Mathematics and Computation, 212(2):366– 376, June 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309001684.

Wan:2007:GMN

[WSLZ07] Ying Wan, Lina Song, LiYin, and Hongqing Zhang. Generalized method and new exact wave solutions for (2 + 1)- dimensional Broer–Kaup– Kupershmidt system. Applied Mathematics and Computation, 187(2):644– 657, April 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306011209.

Wu:2009:NMR

[WSXL09] Zhu-Wu Wu, Xue-Feng Song, Ying-Ying Xu, and Kun Liu. A note on a minimax rule for portfolio selection and equilibrium price system. Applied Mathematics and Computation, 208(1):49–57, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008461.

Wang:2007:BEM

[WsZ07a] Han Wang and Hui sheng Zhang. Boundary element method for simulating the coupled motion of a fluid and a threedimensional body. Applied Mathematics and Computation, 190(2):1328–1343, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001646.

Wang:2007:NCS

[WSZ07b] Qi Wang, Lina Song, and Hongqing Zhang. A new coupled sub-equations expansion method and novel complexiton solutions of (2 + 1)-dimensional Burgers equation. Applied Mathematics and Computation, 186(1):632–637, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600974X.

Wang:2007:FFM

[WSZ07c] Weixiang Wang, Youlin Shang, and Liansheng Zhang. A filled function method with one parameter for box constrained global optimization. Applied Mathematics and Computation, 194(1):54–66, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004432.

Wang:2008:NSMa

[WT08] Jin Wang and Jianjun Tian. Numerical study for a model of tumor virotherapy. Applied Mathematics and Computation, 196(1):448–457, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006984.

Wang:2009:SRG

[WT09] Guangbin Wang and Fuping Tan. Some results on a generalized alternating iterative method. Applied Mathematics and Computation, 210(1):100–106, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008953.

Wijaya:2007:PSM

[WTG07] Sastra Wijaya, Syn Kiat Tan, and Sheng-Uei Guan. Permutation and sampling with maximum length CA or pseudorandom number generation. Applied Mathematics and Computation, 185(1):312–321, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008629.

Wang:2009:PVS

[WTS09] Xia Wang, Youde Tao, and Xinyu Song. Pulse vaccination on SEIR epidemic model with nonlinear incidence rate. Applied Mathematics and Computation, 210(2):398–404, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000071.

Wang:2005:HGAb

[WTW05] Ling Wang, Fang Tang, and Hao Wu. Hybrid genetic algorithm based on quantum computing for numerical optimization and parameter estimation. Applied Mathematics and Computation, 171(2):1141– 1156, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:2008:MIM

[WTW08] Peiguang Wang, Shuhuan Tian, and Yonghong Wu. Monotone iterative method for first-order functional difference equations with nonlinear boundary value conditions. Applied Mathematics and Computation, 203(1):266–272, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002488.

Wang:2006:STTb

[WTZ06] Peiguang Wang, Shuhuan Tian, and Ping Zhao. Stability in terms of two measures for difference equations. Applied Mathematics and Computation, 182(2):1309–1315, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004450.

Wu:2005:MFA

[Wu05a]  Tzong-Mou Wu.  A

modified formula of ancient Chinese algorithm by the homotopy continuation technique. Applied Mathematics and Computation, 165(1):31– 35, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003595.

Wu:2005:OOM

[Wu05b]  Tzong-Mou Wu.  One-

to-one mapping matrix. Applied Mathematics and Computation, 169(2):963– 970, October 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304008264.

Wu:2005:SCN

[Wu05c] Tzong-Mou Wu. A study of convergence on the Newton-homotopy continuation method. Applied Mathematics and Computation, 168(2):1169–1174, September 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304006721.

Wu:2005:IMM

[Wu05d] Xinyuan Wu. Improved Muller method and Bisection method with global and asymptotic superlinear convergence of both point and interval for solving nonlinear equations. Applied Mathematics and Computation, 166(2):299–311, July 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004382.

Wu:2005:RAE

[Wu05e] Xinyuan Wu. Recursive algorithm without extra function evaluations for the Jacobian matrix of Viéta’s polynomial system with applications. Applied Mathematics and Computation, 162(2):539–548, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001134.

Wu:2006:NDE

[Wu06a]  Desheng Wu.  A note

on DEA efficiency assessment using ideal point: an improvement of Wang and Luo’s model. Applied Mathematics and Computation, 183(2):819–830, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006163.

Wu:2006:NFS

[Wu06b] Tzong-Mou Wu. A new formula of solving nonlinear equations by Adomian and homotopy methods. Applied Mathematics and Computation, 172(2):903–907, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002456.

Wu:2006:SNE

[Wu06c] Tzong-Mou Wu. Solving the nonlinear equations by the Newton-homotopy continuation method with adjustable auxiliary homotopy function. Applied Mathematics and Computation, 173(1):383–388, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003589.

Wu:2006:EPC

[Wu06d]  Xinyuan Wu.  An ef-

fective predictor–corrector process for large scale linear system of equations. Applied Mathematics and Computation, 180(1):160–166, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305011082.

Wu:2006:NPC

[Wu06e] Xinyuan Wu. Note on predictor–corrector process for ill-conditioned linear system of equations. Applied Mathematics and Computation, 183(1):596–602, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005947.

Wu:2006:ROC

[Wu06f] Xinyuan Wu. The recursions of order conditions for some Runge–Kuttalike methods. Applied Mathematics and Computation, 179(1):147–152, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010167.

Wu:2007:UFS

[Wu07a]  Hsien-Chung Wu.  Us-

ing fuzzy sets theory and Black–Scholes formula to generate pricing boundaries of European options. Applied Mathematics and Computation, 185(1):136–146, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008198.

Wu:2007:TCC

[Wu07b] Min Wu. Two conditions concerning Newton’s method. Applied Mathematics and Computation, 194(1):102–107, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004481.

Wu:2007:ESR

[Wu07c] Shujin Wu. The Euler scheme for random impulsive differential equations. Applied Mathematics and Computation, 191(1):164–175, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002408.

Wu:2007:EAP

[Wu07d] Xinyuan Wu. Error analysis for the predictor– corrector process relating to ill-conditioned linear system of equations. Applied Mathematics and Computation, 186(1):530– 534, March 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306009635.

Wu:2007:NLM

[Wu07e] Xinyuan Wu. Newtonlike method with some remarks. Applied Mathematics and Computation, 188 (1):433–439, May 1, 2007.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306013816.

Wu:2007:NCA

[Wu07f] Xinyuan Wu. A note on computational algorithm for the inverse of a square matrix. Applied Mathematics and Computation, 187(2):962–964, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012318.

Wu:2007:NIN

[Wu07g] Xinyuan Wu. Note on the improvement of Newton’s method for system of nonlinear equations. Applied Mathematics and Computation, 189(2):1476–1479, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017358.

Wu:2008:LBN

[Wu08a]  Honglin Wu.  On the

lower bounds for the norms of Cauchy–Toeplitz and Cauchy–Hankel matrices. Applied Mathematics and Computation, 205(1):391–395, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005973.

Wu:2008:NSL

[Wu08b] Min Wu. A new semilocal convergence theorem for the inexact Newton methods. Applied Mathematics and Computation, 200(1):80–86, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010880.

Wu:2008:NSG

[Wu08c] Shanhe Wu. A new sharpened and generalized version of Hölder’s inequality and its applications. Applied Mathematics and Computation, 197(2):708– 714, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008272.

Wu:2008:SUO

[Wu08d] Ting Wu. Solving unconstrained optimization problem with a filter-based nonmonotone pattern search algorithm. Applied Mathematics and Computation, 203(1):380–386, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002610.

Wu:2009:WOM

[Wu09] J. L. Wu. A wavelet operational method for solving fractional partial differential equations numerically. Applied Mathematics and Computation, 214(1):31–40, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002756.

Wang:2005:HOM

[WW05a] Yushun Wang and Bin Wang. High-order multisymplectic schemes for the nonlinear Klein–Gordon equation. Applied Mathematics and Computation, 166(3):608–632, July 26, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004618.

Wei:2005:UPA

[WW05b] Fengying Wei and Ke Wang. Uniform persistence of asymptotically periodic multispecies competition predator–prey systems with Holling. III. type functional response. Applied Mathematics and Computation, 170(2):994–998, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000408.

Wu:2005:CNB

[WW05c] Ting Wu and Jin-Rong Wang. Comment: a new blind signature based on the discrete logarithm problem for untraceability. Applied Mathematics and Computation, 170(2):999–1005, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500041X.

See [LHY05].

Wu:2005:IRS

[WW05d] Weixing Wu and Yongxiang Wang. Investment with restricted stock and the value of information. Applied Mathematics and Computation, 163(2):811–824, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002620.

Wang:2006:SMM

[WW06] Yen-Zen Wang and Tzong-Mou Wu. Some mathematical models for deciding the number of examination questions on computerized adaptive tests. Applied Mathematics and Computation, 180(1):178– 184, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305011112.

Wang:2008:COQ

[WW08a] Weiping Wang and Tianming Wang. Commentary on an open question. Applied Mathematics and Computation, 196(1):353–355, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700687X.

Wang:2008:NTW

[WW08b] Yang Wang and Long Wei. New travelling wave solutions to some nonlinear equations via a combined method. Applied Mathematics and Computation, 204(2):726–732, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005468.

Wisniowska-Wajnryb:2009:SEB

[WW09]  Agnieszka  Wiśniowska-

Wajnryb. Some extremal bounds for subclasses of univalent functions. Applied Mathematics and Computation, 215(7):2634–2641, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008005.

Wu:2006:MEE

[WWC06] Chin-Chuan Wu, Shu-Fei Wu, and Huang-Yu Chan. MLE and the estimated expected test time for the two-parameter Gompertz distribution under progressive censoring with binomial removals. Applied Mathematics and Computation, 181(2):1657–1670, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002979.

Wu:2009:PAB

[WWC09] Jingli Wu, Jianxin Wang, and Jian’er Chen. A practical algorithm based on particle swarm optimization for haplotype reconstruction. Applied Mathematics and Computation, 208(2):363–372, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308009028.

Wang:2009:EUP

[WWD09] Qi Wang, Yue Wang, and Binxiang Dai. Existence and uniqueness of positive periodic solutions for a neutral Logarithmic population model. Applied Mathematics and Computation, 213(1):137–147, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002100.

Wu:2008:IAT

[WWHL08] Qing’e Wu, Tuo Wang, Yongxuan Huang, and Jisheng Li. Improvement of algorithm on the track recognition. Applied Mathematics and Computation, 205(2):635–647, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003421.

Wang:2006:ERQ

[WWL06] Qing-Wen Wang, Zhong-Cheng Wu, and Chun-Yan Lin. Extremal ranks of a quaternion matrix expression subject to consistent systems of quaternion matrix equations with applications. Applied Mathematics and Computation, 182(2):1755–1764, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005224.

Wang:2008:NSE

[WWL08a] Wansheng Wang, Liping Wen, and Shoufu Li. Nonlinear stability of explicit and diagonally implicit Runge–Kutta methods for neutral delay differential equations in Banach space. Applied Mathematics and Computation, 199(2):787– 803, June 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010648.

Wang:2008:NSMb

[WWL08b] Wansheng Wang, Liping Wen, and Shoufu Li. Nonlinear stability of θ-methods for neutral differential equations in Banach space. Applied Mathematics and Computation, 198(2):742–753, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009502.

White:2006:EPC

[WWLX06] Luther White, Frances White, Yiqi Luo, and Tao Xu. Estimation of parameters in carbon sequestration models from net ecosystem exchange data. Applied Mathematics and Computation, 181(2):864–879, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001664.

Wang:2005:MEA

[WWP05] Kuo-Hsiung Wang, Tsung-Yin Wang, and Wen-Lea Pearn. Maximum entropy analysis to the N policy M/G/1 queueing system with server breakdowns and general startup times. Applied Mathematics and Computation, 165(1):45– 61, June 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003613.

Wu:2005:OPE

[WWT05] Jong-Wuu Wu, Chin-Chuan Wu, and Mei-Huei Tsai. Optimal parameter estimation of the twoparameter bathtub-shaped lifetime distribution based on a type II right censored sample. Applied Mathematics and Computation, 167(2):807–819, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005284.

Wei:2003:CND

[WWW03] Yimin Wei, Guanglin Wang, and Dingkun Wang. Condition number of Drazin inverse and their condition numbers of singular linear systems. Applied Mathematics and Computation, 146(2–3):455–467, December 31, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See comment [CX08b].

Wang:2007:GCA

[WWWL07] Guangmin Wang, Xianjia Wang, Zhongping Wan, and Yibing Lv. A globally convergent algorithm for a class of bilevel nonlinear programming problem. Applied Mathematics and Computation, 188(1):166– 172, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306013336.

Wang:2007:RLI

[WWXX07] Zhongshi Wang, Wei Wu, Xinhe Xu, and Dingyu Xue. Recognition and location of the internal corners of planar checkerboard calibration pattern image. Applied Mathematics and Computation, 185(2):894–906, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008071.

Wen:2008:DMC

[WWY08] Liping Wen, Wansheng Wang, and Yuexin Yu. Dissipativity of θ-methods for a class of nonlinear neutral delay differential equations. Applied Mathematics and Computation, 202(2):780–786, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800180X.

Wen:2009:DRK

[WWY09] Liping Wen, Suxia Wang, and Yuexin Yu. Dissipativity of Runge–Kutta methods for neutral delay integro-differential equations. Applied Mathematics and Computation, 215(2):583–590, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900530X.

Wen:2008:NSA

[WWYL08] Liping Wen, Wansheng Wang, Yuexin Yu, and Shoufu Li. Nonlinear stability and asymptotic stability of implicit Euler method for stiff Volterra functional differential equations in Banach spaces. Applied Mathematics and Computation, 198(2):582–591, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009241.

Wang:2005:SKR

[WX05] Guorong Wang and Zhaoliang Xu. Solving a kind of restricted matrix equations and Cramer rule. Applied Mathematics and Computation, 162(1):329–338, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000980.

Wu:2006:SSM

[WX06] Xinyuan Wu and Jianlin Xia. Some substantial modifications and improvements for derivativefree iterative methods and derivative-free transformation for multiple zeros. Applied Mathematics and Computation, 181(2):1585–1599, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600289X.

Wang:2008:GSG

[WX08a] Shubin Wang and Hongxia Xue. Global solution for a generalized Boussinesq equation. Applied Mathematics and Computation, 204(1):130–136, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004827.

Wu:2008:NSO

[WX08b] Xianping Wu and Qingzhen Xu. A new second order optimality conditions for the extremal problem under inclusion constraints. Applied Mathematics and Computation, 199(1):99–104, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009745.

Wang:2009:UAC

[WX09] Haiyong Wang and Shuhuang Xiang. Uniform approximations to Cauchy principal value integrals of oscillatory functions. Applied Mathematics and Computation, 215(5):1886–1894, November 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309006894.

Wang:2006:DPS

[WXLH06] Zhaocai Wang, Dongmei Xiao, Wenxia Li, and Lin He. A DNA procedure for solving the shortest path problem. Applied Mathematics and Computation, 183(1):79–84, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600539X.

Wang:2009:RBI

[WXW09] Lingyun Wang, Li Xie, and Xianfang Wang. The residual based interactive stochastic gradient algorithms for controlled moving average models. Applied Mathematics and Computation, 211(2):442–449, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000927.

Wang:2005:HOS

[WY05a] Shiuh-Jeng Wang and Cheng-Hsing Yang. Hierarchyoriented searching algorithms using alternative duplicate codewords for vector quantization mechanism. Applied Mathematics and Computation, 162(2):559–576, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001158.

Wang:2005:DSI

[WY05b] Shiying Wang and Aiming Yang. DNA solution of integer linear programming. Applied Mathematics and Computation, 170(1):626–632, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000044.

Wu:2005:SIA

[WY05c] Jong-Wuu Wu and Huei-Yi Yu. Statistical inference about the shape parameter of the Burr type XII distribution under the failurecensored sampling plan. Applied Mathematics and Computation, 163(1):443– 482, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002358.

Wu:2005:CTR

[WY05d] Lin-Chuan Wu and Yi-Shiung Yeh. Comment on traceability on RSA-based partially signature with low computation. Applied Mathematics and Computation, 170(2):1344–1348, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500069X.

Wang:2006:NSS

[WY06] Chia Hung Wang and Ru Yang. A numerical study for slip flow heat transfer. Applied Mathematics and Computation, 173(2):1246–1264, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004285.

Wang:2007:JSS

[WY07] Xingyuan Wang and Xuejing Yu. Julia sets for the standard Newton’s method, Halley’s method, and Schröder’s method. Applied Mathematics and Computation, 189(2):1186–1195, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017024.

Wu:2008:EBB

[WY08] Mingzhu Wu and Zuodong Yang. Existence of boundary blow-up solutions for a class of quasilinear elliptic systems with critical case. Applied Mathematics and Computation, 198(2):574– 581, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700923X.

Wang:2009:CAQ

[WY09] Kuo-Hsiung Wang and Dong-Yuh Yang. Controlling arrivals for a queueing system with an unreliable server: Newton-Quasi method. Applied Mathematics and Computation, 213(1):92–101, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002057.

Wang:2007:WGC

[WYD07] Zizhong J. Wang, Huiqing H.

Yang, and Jiu Ding. Web graphics for the computation of invariant measures. Applied Mathematics and Computation, 187(2):1442–1452, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012811.

Wei:2006:CPS

[WYL06a] Zengxin Wei, Shengwei Yao, and Liying Liu. The convergence properties of some new conjugate gradient methods. Applied Mathematics and Computation, 183(2):1341–1350, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006813.

Wen:2006:SED

[WYL06b] Liping Wen, Yuexin Yu, and Shoufu Li. Stability of explicit and diagonal implicit Runge– Kutta methods for nonlinear Volterra functional differential equations in Banach spaces. Applied Mathematics and Computation, 183(1):68–78, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005388.

Wen:2006:TPC

[WYL06c] Liping Wen, Yuexin Yu, and Shoufu Li. The test problem class of Volterra functional differential equations in Banach space. Applied Mathematics and Computation, 179(1):30–38, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010039.

Wu:2006:EAC

[WYL06d] Desheng (Dash) Wu, Zijiang Yang, and Liang Liang. Efficiency analysis of cross-region bank branches using fuzzy data envelopment analysis. Applied Mathematics and Computation, 181(1):271– 281, October 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306000968.

Wang:2008:ERL

[WYL08a] Qing-Wen Wang, Shao-Wen Yu, and Chun-Yan Lin. Extreme ranks of a linear quaternion matrix expression subject to triple quaternion matrix equations with applications. Applied Mathematics and Computation, 195(2):733–744, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700611X.

Wang:2008:REC

[WYL08b] Shiying Wang, Jun Yuan, and Aixia Liu. k-Restricted edge connectivity for some interconnection networks. Applied Mathematics and Computation, 201(1–2):587–596, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000039.

Wenbin:2005:MFE

[WYQ05] Chen Wenbin, Shen Yifan, and Xia Qing. A mortar finite element approximation for the linear Poisson– Boltzmann equation. Applied Mathematics and Computation, 164(1):11– 23, May 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002930.

Wang:2005:IWG

[WYX05] Ying-Ming Wang, Jian-Bo Yang, and Dong-Ling Xu. Interval weight generation approaches based on consistency test and interval comparison matrices. Applied Mathematics and Computation, 167(1):252–273, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004898.

Wang:2008:FME

[WYZL08] Fei Wang, Zejian Yuan, Nanning Zheng, and Yuehu Liu. A fast, memoryefficient and parallelizable arc/circle segmentation algorithm. Applied Mathematics and Computation, 205(2):841–848, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003639.

Wei:2005:NHG

[WZ05] Lingyun Wei and Mei Zhao. A niche hybrid genetic algorithm for global optimization of continuous multimodal functions. Applied Mathematics and Computation, 160(3):649–661, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303011858.

Wang:2006:TAA

[WZ06a] Fusheng Wang and Kecun Zhang. The theoretical analysis and algorithm on a class of optimal curve fitting problems. Applied Mathematics and Computation, 183(1):622–633, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005984.

Wang:2006:ABS

[WZ06b] Hailing Wang and Shouming Zhong. Asymptotic behavior of solutions in nonautonomous predator– prey patchy system with beddington-type functional response. Applied Mathematics and Computation, 172(1):122–140, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500192X.

Wang:2006:IFL

[WZ06c] Ke Wang and Bing Zheng. Inconsistent fuzzy linear systems. Applied Mathematics and Computation, 181(2):973–981, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001780.

Wang:2006:SSO

[WZ06d] Ke Wang and Bing Zheng. Symmetric successive overrelaxation methods for fuzzy linear systems. Applied Mathematics and Computation, 175(2):891– 901, April 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030500651X.

Wang:2006:SOU

[WZ06e] Ling Wang and Liang Zhang. Stochastic optimization using simulated annealing with hypothesis test. Applied Mathematics and Computation, 174(2):1329–1342, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005539.

Wang:2006:NFF

[WZ06f] Xiaoli Wang and Guobiao Zhou. A new filled function for unconstrained global optimization. Applied Mathematics and Computation, 174(1):419–429, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004686.

Wu:2006:BEP

[WZ06g] Hongping Wu and Guofen Zhang. Bayesian estimation of the parameters in n non-independent and non-identical series system. Applied Mathematics and Computation, 174(1):223– 235, March 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305004522.

Wu:2006:CTF

[WZ06h] Qingbiao Wu and Yueqing Zhao. The convergence theorem for a family deformed Chebyshev method in Banach space. Applied Mathematics and Computation, 182(2):1369–1376, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004516.

Wu:2006:NKC

[WZ06i] Qingbiao Wu and Yueqing Zhao. A Newton– Kantorovich convergence theorem for the inversefree Jarratt method in Banach space. Applied Mathematics and Computation, 179(1):39–46, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010040.

Wu:2006:TOC

[WZ06j] Qingbiao Wu and Yueqing Zhao. Third-order convergence theorem by using majorizing function for a modified Newton method in Banach space. Applied Mathematics and Computation, 175(2):1515–1524, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006946.

Wu:2006:COP

[WZ06k] Zhen Wu and Liyan Zhang. The corporate optimal portfolio and consumption choice problem in the real project with borrowing rate higher than deposit rate. Applied Mathematics and Computation, 175(2):1596–1608, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007174.

Wang:2007:NSC

[WZ07a] B. Wang and S. M. Zhong.

On novel stability criterion of neutral system with time-varying delay. Applied Mathematics and Computation, 189(2):1124–1130, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016870.

Wang:2007:GOI

[WZ07b] Yong-Jun Wang and Jiang-She Zhang. Global optimization by an improved differential evolutionary algorithm. Applied Mathematics and Computation, 188(1):669–680, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014056.

Wang:2007:MNK

[WZ07c] Zhen Wang and Hong-Qing Zhang. Many new kinds exact solutions to (2 + 1)-dimensional Burgers equation and Klein– Gordon equation used a new method with symbolic computation. Applied Mathematics and Computation, 186(1):693–704, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009817.

Wu:2007:NKT

[WZ07d] Qingbiao Wu and Yueqing Zhao. Newton– Kantorovich type convergence theorem for a family of new deformed Chebyshev method. Applied Mathematics and Computation, 192(2):405–412, September 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307003438.

Wan:2008:ESS

[WZ08a] Li Wan and Qinghua Zhou. Exponential stability of stochastic reaction– diffusion Cohen–Grossberg neural networks with delays. Applied Mathematics and Computation, 206(2):818–824, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007303.

Wang:2008:GAP

[WZ08b] Kai Wang and Yanling Zhu. Global attractivity of positive periodic solution for a Volterra model. Applied Mathematics and Computation, 203(2):493–501, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002063.

Wang:2008:CCO

[WZ08c] Shanshan Wang and Luming Zhang. A class of conservative orthogonal spline collocation schemes for solving coupled Klein– Gordon–Schrödinger equations. Applied Mathematics and Computation, 203(2):799–812, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004062.

Wei:2008:SOE

[WZ08d] Zhengyuan Wei and Xinsheng Zhang. Second order exponential differential operator and generalized Hermite polynomials. Applied Mathematics and Computation, 206(2):781–787, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308007078.

Wei:2008:PSS

[WZ08e] Zhongli Wei and Mingchuan Zhang. Positive solutions of singular sub-linear boundary value problems for fourth-order and secondorder differential equation systems. Applied Mathematics and Computation, 197(1):135–148, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700762X.

Wu:2008:NAM

[WZ08f] Huizhuo Wu and KeCun Zhang. A new accelerating method for global non-convex quadratic optimization with non-convex quadratic constraints. Applied Mathematics and Computation, 197(2):810– 818, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008417.

Wang:2009:FTS

[WZ09a] Ping Wang and Na Zhao. Finite termination of a smoothing-type algorithm for the monotone affine variational inequality problem. Applied Mathematics and Computation, 211 (1):177–184, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030900068X.

Wang:2009:RAI

[WZ09b] Ping Wang and Kewang Zheng. Regularization approach for inverting the exponential radon transforms (ERT). Applied Mathematics and Computation, 207(2):514–518, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008382.

Wang:2009:ASO

[WZ09c] Yunjuan Wang and Detong Zhu. An affine scaling optimal path method with interior backtracking curvilinear technique for linear constrained optimization. Applied Mathematics and Computation, 207(1):178–196, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007777.

Wang:2013:CSA

[WZ13] Zhi-Gang Wang and Zhi-Hua Zhang. Comments on “On analytic functions with generalized bounded Mocanu variation” by K. I. Noor and A. Muhammad [Appl. Math. Comput. 196 (2008) 802–811]. Applied Mathematics and Computation, 219(11):5754–5756, February 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300312012465.

See [NM08b].

Wang:2007:CSS

[WZC07] Tingchun Wang, Luming Zhang, and Fangqi Chen. Conservative schemes for the symmetric Regularized Long Wave equations. Applied Mathematics and Computation, 190(2):1063–1080, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001270.

Wang:2008:NAM

[WZC08] Tingchun Wang, Luming Zhang, and Fangqi Chen. Numerical analysis of a multi-symplectic scheme for a strongly coupled Schrödinger system. Applied Mathematics and Computation, 203(1):413–431, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800266X.

Wang:2009:AFP

[WZDgH09] Chao Wang, Jinghao Zhu, Boško Damjanović, and Liang gen Hu. Approximating fixed points of a pair of contractive type mappings in generalized convex metric spaces. Applied Mathematics and Computation, 215(4):1522–1525, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006523.

Wu:2008:CMK

[WZF08] Hongxia Wu, Yunbo Zeng, and Tianyou Fan. Complexitons of the modified KdV equation by Darboux transformation. Applied Mathematics and Computation, 196(2):501–510, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007023.

Wang:2008:ASC

[WZL08a] B. Wang, S. Zhong, and X. Liu. Asymptotical stability criterion on neural networks with multiple time-varying delays. Applied Mathematics and Computation, 195(2):809–818, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006200.

Wang:2008:AET

[WZL08b] Mingliang Wang, Jinliang Zhang, and Xiangzheng Li. Application of the (\frac{G′}{G})-expansion to travelling wave solutions of the Broer–Kaup and the approximate long water wave equations. Applied Mathematics and Computation, 206(1):321–326, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006486.

Wu:2006:OCS

[WZM06] Hong-Wu Wu, Rong-Kun Zhuang, and Ronald M. Mathsen. Oscillation criteria for second-order nonlinear neutral variable delay dynamic equations. Applied Mathematics and Computation, 178(2):321–331, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500977X.

Wang:2007:ESB

[WZS07] Guangwa Wang, Mingru Zhou, and Li Sun. Existence of solutions of boundary value problem for 3rd order nonlinear system. Applied Mathematics and Computation, 189(2):1131–1138, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016973.

Wang:2008:DGS

[WZSL08] Xingyuan Wang, Xu Zhang, Yuanyuan Sun, and Fanping Li. Dynamics of the generalized M set on escape-line diagram. Applied Mathematics and Computation, 206(1):474–484, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007029.

Wang:2006:NGA

[WZSZ06] Jing Wang, Xiaoling Zhang, Lina Song, and Hongqing Zhang. A new general algebraic method with symbolic computation and its application to two nonlinear differential equations with nonlinear terms of any order. Applied Mathematics and Computation, 182(2):1330–1340, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004474.

Wang:2006:FPA

[WZT06] Fusheng Wang, Kecun Zhang, and Xiaolong Tan. A fractional programming algorithm based on conic quasi-Newton trust region method for unconstrained minimization. Applied Mathematics and Computation, 181(2):1061–1067, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001895.

Wang:2008:VTR

[WZWW08] Fusheng Wang, Kecun Zhang, Chuanlong Wang, and Li Wang. A variant of trust-region methods for unconstrained optimization. Applied Mathematics and Computation, 203(1):297–307, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002543.

Wenrui:2007:DZS

[WZY07] Shan Wenrui, Niu Zhilei, and Guo Yanping. The distribution of zeros of solutions for a class of neutral delay differential equations. Applied Mathematics and Computation, 186(2):1137–1150, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601040X.

Wang:2006:GMF

[WZZ06a] Jie Wang, Weijun Zhong, and Jun Zhang. A general meshsize fourth-order compact difference discretization scheme for 3D Poisson equation. Applied Mathematics and Computation, 183(2):804–812, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600614X.

Wang:2006:HOC

[WZZ06b] Jie Wang, Weijun Zhong, and Jun Zhang. High order compact computation and nonuniform grids for streamfunction vorticity equations. Applied Mathematics and Computation, 179(1):108–120, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010131.

Xingyuan:2006:JSS

[XB06] Wang Xingyuan and Liu Bo. Julia sets of the Schröder iteration functions of a class of oneparameter polynomials with high degree. Applied Mathematics and Computation, 178(2):461–473, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009926.

Xu:2006:NNA

[XBZ06] H. Xu, R. Barron, and C. Zhang. A new numerical algorithm based on transformed equations and its applications to very low Re fluid flows. Applied Mathematics and Computation, 181(1):619–633, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001433.

Xu:2005:ETR

[XC05] Qiu-Liang Xu and Tzer-Shyong Chen. An efficient threshold RSA digital signature scheme. Applied Mathematics and Computation, 166(1):25– 34, July 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004187.

Xingyuan:2006:RCP

[XC06a] Wang Xingyuan and Luo Chao. Researches on chaos phenomenon of EEG dynamics model. Applied Mathematics and Computation, 183(1):30–41, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003717.

Xiong:2006:SRF

[XC06b] Zhiguang Xiong and Chuanmiao Chen. Superconvergence of rectangular finite element with interpolated coefficients for semilinear elliptic problem. Applied Mathematics and Computation, 181(2):1577–1584, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002888.

Xiong:2007:STQ

[XC07] Zhiguang Xiong and Chuanmiao Chen. Superconvergence of triangular quadratic finite element with interpolated coefficients for semilinear parabolic equation. Applied Mathematics and Computation, 184(2):901–907, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600779X.

Xu:2005:GSS

[XCD05a] Rui Xu, M. A. J. Chaplain, and F. A. Davidson. Global stability of a stage-structured predatorprey model with prey dispersal. Applied Mathematics and Computation, 171(1):293–314, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000950.

Xu:2005:PSP

[XCD05b] Rui Xu, M. A. J. Chaplain, and F. A. Davidson. Periodic solutions for a predator–prey model with Holling-type functional response and time delays. Applied Mathematics and Computation, 161(2):637–654, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000177.

Xu:2006:GCR

[XCD06] Rui Xu, M. A. J. Chaplain, and F. A. Davidson. Global convergence of a reaction– diffusion predator–prey model with stage structure for the predator. Applied Mathematics and Computation, 176(1):388–401, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008155.

Xie:2005:CSW

[XCFX05] Wen-Xian Xie, Li Cai, Jian-Hu Feng, and Wei Xu. Computations of shallow water equations with high-order central-upwind schemes on triangular meshes. Applied Mathematics and Computation, 170(1):296–313, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009270.

Xu:2005:PSD

[XCH05a] Rui Xu, Lansun Chen, and Feilong Hao. Periodic solutions of a discrete time Lotka–Volterra type food-chain model with delays. Applied Mathematics and Computation, 171(1):91–103, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000822.

Xu:2005:PSS

[XCH05b] Rui Xu, Lansun Chen, and Feilong Hao. Periodic solutions of an n-species Lotka–Volterra type foodchain model with time delays. Applied Mathematics and Computation, 171(1):511–530, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001128.

Xingyuan:2009:NDR

[XCJ09] Wang Xingyuan, Luo Chao, and Meng Juan. Nonlinear dynamic research on EEG signals in HAI experiment. Applied Mathematics and Computation, 207(1):63–74, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000192.

Xu:2009:CBD

[XCL09] Junyan Xu, Lijuan Chen, and Zhong Li. Convergence behavior of delayed bidirectional associative memory cellular neural networks with asymptotically periodic coefficients. Applied Mathematics and Computation, 215(3):928–935, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005797.

Xiang:2007:CCD

[XCM07] Zhaoyin Xiang, Qiong Chen, and Chunlai Mu. Critical curves for degenerate parabolic equations coupled via non-linear boundary flux. Applied Mathematics and Computation, 189(1):549–559, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016286.

Xing:2007:MGA

[XCS07] Li-Ning Xing, Ying-Wu Chen, and Xue-Shi Shen. Multiprogramming genetic algorithm for optimization problems with permutation property. Applied Mathematics and Computation, 185(1):473–483, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008824.

Xie:2008:DSN

[XCW08] Xiaoping Xie, Zhitong Cao, and Xuchu Weng. Detecting spatiotemporal nonlinear dynamics in resting state of human brain based on fMRI datasets. Applied Mathematics and Computation, 205(1):19–25, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003147.

Xia:2007:PCA

[XD07] Fu Quan Xia and Xie Ping Ding. Predictor–corrector algorithms for solving generalized mixed implicit quasi-equilibrium problems. Applied Mathematics and Computation, 188 (1):173–179, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306013348.

Xiong:2006:CDR

[XF06] Xiang-Tuan Xiong and Chu-Li Fu. Central difference regularization method for the Cauchy problem of the Laplace’s equation. Applied Mathematics and Computation, 181(1):675–684, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001470.

Xiong:2008:SRM

[XF08a] Xiang-Tuan Xiong and Chu-Li Fu. A spectral regularization method for solving surface heat flux on a general sideways parabolic. Applied Mathematics and Computation, 197(1):358–365, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007837.

Xue:2008:SCN

[XF08b] Zhiqun Xue and Ruiqin Fan. Some comments on Noor’s iterations in Banach spaces. Applied Mathematics and Computation, 206(1):12–15, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006504.

Xiong:2006:CDM

[XFL06] Xiang-Tuan Xiong, Chu-Li Fu, and Hong-Fang Li. Central difference method of a non-standard inverse heat conduction problem for determining surface heat flux from interior observations. Applied Mathematics and Computation, 173(2):1265–1287, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305004303.

Xiong:2006:TNM

[XFQ06] Xiang-Tuan Xiong, Chu-Li Fu, and Zhi Qian. Two numerical methods for solving a backward heat conduction problem. Applied Mathematics and Computation, 179(1):370–377, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010441.

Xiao-Fang:2009:SDD

[XFTY09] Wang Xiao-Fang and Fan Tian-You. Study on the dynamics of the double cantilever-beam specimen of decagonal Al–Ni– Co quasicrystals. Applied Mathematics and Computation, 211(2):336–346, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000794.

Xiang:2008:GQR

[XG08] Shuhuang Xiang and Weihua Gui. On generalized quadrature rules for fast oscillatory integrals. Applied Mathematics and Computation, 197(1):60–75, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007540.

Xu:2007:SEE

[XGTS07] Wei Xu, Liang Gao, Yaning Tang, and Jianwei Shen. A series of explicit and exact travelling wave solutions of the B ( m n ) equations. Applied Mathematics and Computation, 185(1):748–754, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009106.

Gu:2007:IBC

[xGyZtZ^{+}07] Tong xiang Gu, Xian yu Zuo, Li tao Zhang, Wan qin Zhang, and Zhi qiang Sheng. An improved bi-conjugate residual algorithm suitable for distributed parallel computing. Applied Mathematics and Computation, 186(2):1243–1253, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010496.

Gao:2008:FRU

[xGZZ08] Quan xue Gao, Lei Zhang, and David Zhang. Face recognition using FLDA with single training image per person. Applied Mathematics and Computation, 205(2):726–734, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003512.

Xu:2006:SIO

[XHH06] Zhenli Xu, Jingsong He, and Houde Han. Semiimplicit operator splitting Padé method for higherorder nonlinear Schrödinger equations. Applied Mathematics and Computation, 179(2):596–605, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010659.

Xie:2007:IMM

[XHL07] Qingming Xie, Anqi He, and Jianzhou Liu. On the iterative method for H-matrices. Applied Mathematics and Computation, 189(1):41–48, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015761.

Xu:2005:SSW

[XHLL05] WeiGuo Xu, DaiPing Hu, AiZhong Lei, and Qin Shu Li. Single spiral wave induced by noise. Applied Mathematics and Computation, 171(2):703–709, December 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xu:2006:EIS

[XHSL06] WeiGuo Xu, DaiPing Hu, HuiZhang Shen, and Mengyu Li. Explicit internal signal stochastic resonance in a chemical oscillator. Applied Mathematics and Computation, 172(2):1188–1194, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002857.

Xu:2009:RBA

[XHSL09] Wangtu Xu, Shiwei He, Rui Song, and Juan Li. Reliability based assignment in stochastic-flow freight network. Applied Mathematics and Computation, 211(1):85– 94, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000411.

Xu:2007:EPH

[XHX07] Min Xu, Xiao-Dong Hu, and Jun-Ming Xu. Edgepancyclicity and Hamiltonian laceability of the balanced hypercubes. Applied Mathematics and Computation, 189(2):1393–1401, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017255.

Xu:2006:EBS

[XHZ06] Min Xu, Xiao-Dong Hu, and Qiang Zhu. Edgebipancyclicity of star graphs under edge-fault tolerant. Applied Mathematics and Computation, 183(2):972–979, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006321.

Xiang:2006:LTE

[Xia06] Xinmin Xiang. The large time error estimation of Chebyshev pseudospectral approximation for a class of semilinear parabolic equations. Applied Mathematics and Computation, 177(2):720–728, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009525.

Xiaojian:2008:MCH

[Xia08] Zhou Xiaojian. Modified Chebyshev–Halley methods free from second derivative. Applied Mathematics and Computation, 203(2):824–827, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004098.

Xie:2005:ITA

[Xie05] Qi Xie. Improvement of Tzeng et al.’s nonrepudiable threshold proxy signature scheme with known signers. Applied Mathematics and Computation, 168(2):776–782, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030400637X.

Xingyuan:2006:GSC

[Xin06]  Wang Xingyuan.  The

generalized M sets from a class complex mapping system. Applied Mathematics and Computation, 175(2):1484–1494, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006910.

Xu:2005:NST

[XL05] Qingzhen Xu and Xiaonan Luo. Numerical solution of a two-dimensional simulation on heat and mass transfer through cloth. Applied Mathematics and Computation, 171(2):843– 852, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xiang:2006:FCC

[XL06a] Wang Xiang and Lu Linzhang. Factorizations of confluent Cauchy–Vandermonde matrices. Applied Mathematics and Computation, 182(2):1667–1672, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005157.

Xu:2006:DTC

[XL06b] Qingzhen Xu and Xiaonan Luo. Dynamic thermal comfort numerical simulation model on 3D garment CAD. Applied Mathematics and Computation, 182(1):106–118, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306003183.

Xu:2007:RAC

[XL07] M. H. Xu and G. F. Luan.

A rapid algorithm for a class of linear complementarity problems. Applied Mathematics and Computation, 188(2):1647–1655, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015177.

Xie:2008:APS

[XL08a] Yi Xie and Xiangao Li.

Almost periodic solutions of single population model with hereditary effects. Applied Mathematics and Computation, 203(2):690–697, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003937.

Xie:2008:ETS

[XL08b] Yi Xie and Haishen Lü. An existence theorem for singular boundary value problems with sign changing nonlinearities. Applied Mathematics and Computation, 197(1):412–425, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307007904.

Xu:2009:CRG

[XL09a] Bin Xu and Xi-Qiang Liu. Classification, reduction, group invariant solutions and conservation laws of the Gardner-KP equation. Applied Mathematics and Computation, 215(3):1244–1250, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006389.

Xu:2009:HRS

[XL09b] Weiwei Xu and Wen Li. The Hermitian reflexive solutions to the matrix inverse problem AX = B. Applied Mathematics and Computation, 211(1):142– 147, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000629.

Xu:2009:SCH

[XL09c] Xiubin Xu and Yonghui Ling. Semilocal convergence for Halley’s method under weak Lipschitz condition. Applied Mathematics and Computation, 215(8):3057–3067, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008844.

Xu:2007:IDS

[XLH^{+}07] M. H. Xu, Y. Q. Liu, Q. L. Huang, Y. X. Zhang, and G. F. Luan. An improved Dijkstra’s shortest path algorithm for sparse network. Applied Mathematics and Computation, 185(1):247–254, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008563.

Xu:2009:MAS

[XLHK09] Haiyan Xu, Kevin W.

Li, Keith W. Hipel, and D. Marc Kilgour. A matrix approach to status quo analysis in the graph model for conflict resolution. Applied Mathematics and Computation, 212(2):470–480, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001799.

Xu:2009:MBA

[XLKH09] Haiyan Xu, Kevin W. Li, D. Marc Kilgour, and Keith W. Hipel. A matrixbased approach to searching colored paths in a weighted colored multidigraph. Applied Mathematics and Computation, 215(1):353–366, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004779.

Xu:2006:GES

[XLL06] Bingji Xu, Xinzhi Liu, and Xiaoxin Liao. Global exponential stability of high order Hopfield type neural networks. Applied Mathematics and Computation, 174(1):98–116, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004443.

Xu:2006:HAN

[XLSL06] M. H. Xu, William H. K.

Lam, H. Shao, and G. F. Luan. A heuristic algorithm for network equilibration. Applied Mathematics and Computation, 174(1):430–446, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305004777.

Xiang:2007:IPS

[XLwW07] Tao Xiang, Xiaofeng Liao, and Kwok wo Wong. An improved particle swarm optimization algorithm combined with piecewise linear chaotic map. Applied Mathematics and Computation, 190(2):1637–1645, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002081.

Li:2006:CPG

[xLyW06] Mei xia Li and Chang yu Wang. Convergence property of gradienttype methods with nonmonotone line search in the presence of perturbations. Applied Mathematics and Computation, 174(2):854–868, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030500514X.

Xu:2006:SNO

[XM06] Run Xu and Fanwei Meng. Some new oscillation criteria for second order quasi-linear neutral delay differential equations. Applied Mathematics and Computation, 182(1):797–803, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003973.

Xu:2007:NKT

[XM07a] Run Xu and Fanwei Meng. New Kamenev-type oscillation criteria for second order neutral nonlinear differential equations. Applied Mathematics and Computation, 188(2):1364–1370, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306014895.

Xu:2007:OCS

[XM07b] Run Xu and Fanwei Meng. Oscillation criteria for second order quasi-linear neutral delay differential equations. Applied Mathematics and Computation, 192(1):216–222, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003244.

Xu:2008:DTG

[XM08] Qingzhen Xu and Zhanyou Ma. Discrete time Geo / G / 1 queue with Bernoulli gated service simulation system. Applied Mathematics and Computation, 204(1):37–44, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004517.

Xu:2009:ASW

[XMGW09] Xiao-Ge Xu, Xiang-Hua Meng, Yi-Tian Gao, and Xiao-Yong Wen. Analytic N-solitary-wave solution of a variable-coefficient Gardner equation from fluid dynamics and plasma physics. Applied Mathematics and Computation, 210(2):313–320, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007832.

Xenophontos:2007:NSF

[XO07] C. Xenophontos and L. Oberbroeckling. A numerical study on the finite element solution of singularly perturbed systems of reaction– diffusion problems. Applied Mathematics and Computation, 187(2):1351–1367, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012732.

Xu:2007:NAD

[XPS07] Xiubin Xu, Yunlan Pan, and Robert D. Small. Nonlinear approximations to the derivative of delta function. Applied Mathematics and Computation, 185(1):583–593, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008939.

Xingyuan:2006:GMJ

[XQ06a] Wang Xingyuan and Shi Qijiang. The generalized Mandelbort–Julia sets from a class of complex exponential map. Applied Mathematics and Computation, 181(2):816–825, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001627.

Xining:2006:PAB

[XQ06b] Cui Xining and Lü Quanyi.

A parallel algorithm for band linear systems. Applied Mathematics and Computation, 181(1):40– 47, October 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306000646.

Xingyuan:2009:GMS

[XRZ09] Wang Xingyuan, Jia Ruihong, and Zhang Zhenfeng. The generalized Mandelbrot set perturbed by composing noise of additive and multiplicative. Applied Mathematics and Computation, 210(1):107– 118, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009119.

Xu:2006:RII

[XS06] Run Xu and Yuan Gong Sun. On retarded integral inequalities in two independent variables and their applications. Applied Mathematics and Computation, 182(2):1260–1266, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004413.

Xu:2009:SCP

[XS09] Degang Xu and Zhifang Su. Synchronization criterions and pinning control of general complex networks with time delay. Applied Mathematics and Computation, 215(4):1593–1608, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006626.

Xu:2008:PPO

[XSG08] Zhaoliang Xu, Jie Sun, and Chao Gu. Perturbation for a pair of oblique projectors AA_{MN}^{†} and BB_{MN}^{†}. Applied Mathematics and Computation, 203(1):432–446, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002762.

Xu:2007:ACM

[XSKC07] Guangjun Xu, Erfang Shan, Liying Kang, and T. C. E. Cheng. The algorithmic complexity of the minus clique-transversal problem. Applied Mathematics and Computation, 189(2):1410–1418, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017279.

Maolong:2008:IQB

[XSX08] Maolong Xi, Jun Sun, and Wenbo Xu. An improved quantum-behaved particle swarm optimization algorithm with weighted mean best position. Applied Mathematics and Computation, 205(2):751–759, November 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308003536.

Xiao:2006:EAP

[XT06] Aiguo Xiao and Yifa Tang. Equilibrium attractive properties of a class of multistep Runge–Kutta methods. Applied Mathematics and Computation, 173(2):1068–1081, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004121.

Xue:2008:MPS

[XT08] Yan-Fang Xue and Chun-Lei Tang. Multiple periodic solutions for superquadratic second-order discrete Hamiltonian systems. Applied Mathematics and Computation, 196(2): 494–500, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307007011.

Xue:2008:CBI

[XTCC08] Yun Xue, Chong Sze Tong, Ying Chen, and Wen-Sheng Chen. Clusteringbased initialization for nonnegative matrix factorization. Applied Mathematics and Computation, 205(2):525–536, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003317.

Xu:2006:RDP

[Xu06]  Rui Xu.  A reaction–

diffusion predator–prey model with stage structure and nonlocal delay. Applied Mathematics and Computation, 175(2):984–1006, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006570.

Xu:2007:PSM

[Xu07] Fuyi Xu. Positive solutions for multipoint boundary value problems with onedimensional p-Laplacian operator. Applied Mathematics and Computation, 194(2):366–380, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004778.

Xu:2008:MPS

[Xu08a] Fuyi Xu. Multiple positive solutions for nonlinear singular m-point boundary value problem. Applied Mathematics and Computation, 204(1):450–460, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004360.

Xu:2008:NOS

[Xu08b] Zhiting Xu. Nonlinear oscillation of second order differential equations with state-dependent delays. Applied Mathematics and Computation, 206(1):198–204, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006553.

Xu:2009:NKT

[Xu09] Yancong Xu. New Kamenevtype theorems for superlinear matrix differential equations. Applied Mathematics and Computation, 209(2):410–414, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800982X.

Xingyuan:2006:JSN

[XW06] Wang Xingyuan and Liu Wei. The Julia set of Newton’s method for multiple root. Applied Mathematics and Computation, 172(1):101–110, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001906.

Xiao:2007:NSL

[XW07a] Yunhai Xiao and Zengxin Wei. A new subspace limited memory BFGS algorithm for large-scale bound constrained optimization. Applied Mathematics and Computation, 185(1):350–359, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008678.

Xu:2007:PDD

[XW07b] Changyong Xu and Meijuan Wang. Permanence for a delayed discrete three-level food-chain model with Beddington– DeAngelis functional response. Applied Mathematics and Computation, 187(2):1109–1119, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012471.

Xu:2008:COS

[XW08] Qingzhen Xu and Xianping Wu. The characterizations of optimal solution set in programming problem under inclusion constrains. Applied Mathematics and Computation, 198(1):296–304, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008892.

Xiang:2009:RRC

[XW09] Zhengrong Xiang and Ronghao Wang. Robust L_{∞} reliable control for uncertain nonlinear switched systems with time delay. Applied Mathematics and Computation, 210(1):202– 210, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308009946.

Xu:2008:MST

[XWSL08] Feng Xu, Ting Wen, Keith Seffen, and Tianjian Lu. Modeling of skin thermal pain: a preliminary study. Applied Mathematics and Computation, 205(1):37–46, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308003160.

Xiang:2007:NPS

[XwWL07] Tao Xiang, Kwok wo Wong, and Xiaofeng Liao. A novel particle swarm optimizer with time-delay. Applied Mathematics and Computation, 186(1):789–793, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009994.

Xiao:2006:SSK

[XWZ06] Jin-You Xiao, Li-Hua Wen, and Duo Zhang. Solving second kind Fredholm integral equations by periodic wavelet Galerkin method. Applied Mathematics and Computation, 175(1):508– 518, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006259.

Xingyuan:2007:VGF

[XX07] Wang Xingyuan and Yu Xuejing. Visualizing generalized 3x + 1 function dynamics based on fractal. Applied Mathematics and Computation, 188(1):234– 243, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013439.

Xie:2005:PID

[XXC05] W. X. Xie, W. Xu, and L. Cai. Path integration of the Duffing–Rayleigh oscillator subject to harmonic and stochastic excitations. Applied Mathematics and Computation, 171(2):870– 884, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Xie:2006:SDR

[XXC06] W. X. Xie, W. Xu, and L. Cai. Study of the Duffing–Rayleigh oscillator subject to harmonic and stochastic excitations by path integration. Applied Mathematics and Computation, 172(2):1212–1224, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305002882.

Xie:2008:NMP

[XXC08] Wen-Xian Xie, Wei Xu, and Li Cai. Numerical meshfree path integration method for non-linear dynamic systems. Applied Mathematics and Computation, 197(1):426–434, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307007928.

Xue:2006:MVP

[XXF06] Hong-Gang Xue, Cheng-Xian Xu, and Zong-Xian Feng. Mean– variance portfolio optimal problem under concave transaction cost. Applied Mathematics and Computation, 174(1):1–12, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305004376.

Xiao:2009:AEF

[XXL09] Guang-Can Xiao, Da-Quang Xian, and Xi-Qiang Liu. Application of Exp-function method to Dullin–Gottwald–Holm equation. Applied Mathematics and Computation, 210(2):536–541, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000356.

Xiao:2009:ACS

[XXTL09] Yuzhu Xiao, Wei Xu, Sufang Tang, and Xiuchun Li. Adaptive complete synchronization of the noise-perturbed two bi-directionally coupled chaotic systems with timedelay and unknown parametric mismatch. Applied Mathematics and Computation, 213(2):538–547, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002677.

Xu:2006:MPF

[XXX06] Fengmin Xu, Chengxian Xu, and Honggang Xue. A multiple penalty function method for solving Max-Bisection problems. Applied Mathematics and Computation, 173(2):757–766, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305003905.

Xie:2007:CRF

[XY07] Changzhen Xie and Shouzhi Yang. Construction of refinable function vector via GTST. Applied Mathematics and Computation, 194(2):425–430, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004912.

Xu:2011:CBF

[XY11] Dehua Xu and Yunqiang Yin. Comments on “A bicriteria flowshop scheduling problem with setup times”. Applied Mathematics and Computation, 217(17):7361–7364, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [EG06b].

Xu:2010:CNM

[XYL10] Dehua Xu, Yunqiang Yin, and Hongxing Li. Comments on “A note on minimizing maximum lateness in an m-machine scheduling problem with a learning effect”. Applied Mathematics and Computation, 217 (2):939–943, September 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [Ere09b].

Xiao:2009:HHB

[XYZ09] Jing Xiao, YuPing Yan, and Jun Zhang. HP-BILc: a histogram-based EDA for continuous optimization. Applied Mathematics and Computation, 215(3):973–982, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005864.

Xie:2005:SCC

[XZ05] Zheng Xie and Hong-Qing Zhang. Symbolic computation and construction of soliton-like solutions for a (2 + 1)- dimensional breaking soliton equation. Applied Mathematics and Computation, 162(1):283–291, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000888.

Xia:2007:REF

[XZ07a] Yan Xia and Guofen Zhang. Reliability equivalence factors in Gamma distribution. Applied Mathematics and Computation, 187(2):567–573, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008204.

Xiong:2007:EGA

[XZ07b] Xinsheng Xiong and Zhengqiu Zhang. Existence and global attractivity of a periodic solution for a prey–predator model with sex-structure. Applied Mathematics and Computation, 190(2):1213–1224, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700152X.

Xu:2007:ESS

[XZ07c] Rui Xu and ZhienMa. The effect of stage-structure on the permanence of a predator–prey system with time delay. Applied Mathematics and Computation, 189(2):1164–1177, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306017000.

Xing:2008:NOC

[XZ08a] Lihong Xing and Zhaowen Zheng. New oscillation criteria for forced second order half-linear differential equations with damping.

Applied Mathematics and Computation, 198(2):481– 486, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008995.

Xu:2008:ASL

[XZ08b] Y. Xu and J. J. Zhao.

Asymptotic stability of linear non-autonomous difference equations with fixed delay. Applied Mathematics and Computation, 198(2):787–798, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009538.

Xiao:2009:CPM

[XZ09a] Aiguo Xiao and Yongxiang Zhao. Convergence of parallel multistep hybrid methods for singular perturbation problems. Applied Mathematics and Computation, 215(6):2139–2148, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900722X.

Xu:2009:MPS

[XZ09b] Yumei Xu and Haijun Zhang. Multiple positive solutions of a boundary value problem for a class of 2n th-order m-point singular integrodifferential equations in Banach spaces. Applied Mathematics and Computation, 214(2):607–617, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003609.

Xu:2005:NNA

[XZB05] H. Xu, C. Zhang, and R. Barron. A new numerical approach to solve an elliptic equation. Applied Mathematics and Computation, 171(1):1–24, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000779.

Xu:2006:NSC

[XZB06] H. Xu, C. Zhang, and R. Barron. Numerical simulation of cavity flows based on transformed equations. Applied Mathematics and Computation, 176(2):506–515, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008271.

Xiang:2009:IAA

[XZM^{+}09] Xiuqiao Xiang, Jianzhong Zhou, Li Mo, Zhimeng Luo, and Chaoshun Li. Improved algorithm about NSFOT. Applied Mathematics and Computation, 215(3):881–888, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002616.

Xu:2008:SFO

[XZY08] Yu-Guang Xu, Xin-Wei Zhou, and Li Yao. Solving the fifth order Caudrey– Dodd–Gibbon (CDG) equation using the exp-function method. Applied Mathematics and Computation, 206(1):70–73, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006292.

Yamac:2005:EAS

[Yam05a] K. Yamaç. Erratum to “An algorithm for stability of discrete-time interval matrices”. Applied Mathematics and Computation, 168(2):1469, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/S0096300304006836.

See [YB03].

Yamac:2005:ESD

[Yam05b] K. Yamac. Erratum to “On stability of discrete-time interval matrices” [Applied Mathematics and Computation 152 (1) (2004) 163– 167]. Applied Mathematics and Computation, 170 (1):731–732, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [YB04].

Yamamoto:2008:ECL

[Yam08] Yusaku Yamamoto. Equality conditions for lower bounds on the smallest singular value of a bidiagonal matrix. Applied Mathematics and Computation, 200(1):254–260, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011046.

Yan:2005:AJE

[Yan05a] Zhenya Yan. Approximate Jacobi elliptic function solutions of the modified KdV equation via the decomposition method. Applied Mathematics and Computation, 166(3):571–583, July 26, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004588.

Yan:2005:CHN

[Yan05b]  Zhenya Yan.  Control-

ling hyperchaos in the new hyperchaotic Chen system. Applied Mathematics and Computation, 168(2):1239–1250, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006988.

Yan:2005:NDP

[Yan05c]  Zhenya Yan.  Numeri-

cal doubly periodic solution of the KdV equation with the initial condition via the decomposition method. Applied Mathematics and Computation, 168(2):1065–1078, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006630.

Yang:2005:GAS

[Yan05d] Xiaofan Yang. On the global asymptotic stability of the difference equation x_{n} = (x_{n−1}x_{n−2} + x_{n−3} + a)/(x_{n−1} + x_{n−2}x_{n−3} + a). Applied Mathematics and Computation, 171(2):857– 861, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2006:BTD

[Yan06a] Shouzhi Yang. Biorthogonal two-direction refinable function and two-direction wavelet. Applied Mathematics and Computation, 182(2):1717–1724, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005212.

Yang:2006:SMN

[Yan06b] Sung-Dae Yang. Shooting methods for numerical solutions of exact controllability problems constrained by linear and semilinear wave equations with local distributed controls. Applied Mathematics and Computation, 177(1):128– 148, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008957.

Yang:2006:UPP

[Yan06c] Xitao Yang. Uniform persistence for a predator– prey system with delays. Applied Mathematics and Computation, 173(1):523–534, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003723.

Yang:2006:EEP

[Yan06d] Zuodong Yang. Existence of explosive positive solutions of quasilinear elliptic equations. Applied Mathematics and Computation, 177(2):581–588, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009380.

Yan:2007:CSM

[Yan07a]  Tao Yan.  A class

of smoothing methods for mathematical programs with complementarity constraints. Applied Mathematics and Computation, 186(1):1–9, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007016.

Yan:2007:SHB

[Yan07b]  Xiang-Ping Yan.  Sta-

bility and Hopf bifurcation for a delayed prey– predator system with diffusion effects. Applied Mathematics and Computation, 192(2):552–566, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307003591.

Yang:2007:STA

[Yan07c] Jianhua Yang. Schwarz type algorithms for parabolic problems. Applied Mathematics and Computation, 188(1):206–213, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306013397.

Yang:2007:FER

[Yan07d] Sheldon Yang. A further expansion of the Riemann ζ-function. Applied Mathematics and Computation, 187(1):556–558, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012070.

Yang:2007:BSM

[Yan07e]  Zhi-Ming Yang.  The

block smoothing method for estimating the spectral radius of a nonnegative matrix. Applied Mathematics and Computation, 185(1):266–271, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008587.

Yan:2008:MKE

[Yan08a] Zhenya Yan. The modified KdV equation with variable coefficients: Exact uni/bi-variable travelling wave-like solutions. Applied Mathematics and Computation, 203(1):106– 112, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308002129.

Yang:2008:NRF

[Yan08b] G. C. Yang. New results of Falkner–Skan equation arising in boundary layer theory. Applied Mathematics and Computation, 202(1):406–412, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000891.

Yang:2008:PMN

[Yan08c] Lu Yang. A perturbation method for numerical differentiation. Applied Mathematics and Computation, 199(1):368–374, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010053.

Yang:2008:UAN

[Yan08d] Lu Yang. Uniform attractor for non-autonomous hyperbolic equation with critical exponent. Applied Mathematics and Computation, 203(2):895–902, September 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030800430X.

Yang:2008:ASO

[Yan08e] Min Yang. Analysis of second order finite volume element methods for pseudo-parabolic equations in three spatial dimensions. Applied Mathematics and Computation, 196(1):94–104, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006480.

Yang:2009:PES

[Yan09a] Ming-Chien Yang. Path embedding in star graphs. Applied Mathematics and Computation, 207(2):283– 291, January 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308007881.

Yang:2009:NTS

[Yan09b] Q. W. Yang. A numerical technique for structural damage detection. Applied Mathematics and Computation, 215(7):2775–2780, December 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030900767X.

Yang:2009:ITO

[Yan09c]  Weifeng Yang.  On

an integral-type operator between Bloch-type spaces. Applied Mathematics and Computation, 215(3):954–960, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005839.

Yang:2009:HBT

[Yan09d] Yu Yang. Hopf bifurcation in a two-competitor, one-prey system with time delay. Applied Mathematics and Computation, 214(1):228–235, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002975.

Yao:2008:SST

[Yao08a] Baoheng Yao. A series solution to the Thomas–Fermi equation. Applied Mathematics and Computation, 203(1):396–401, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002634.

Yao:2008:MIM

[Yao08b] Qingliu Yao. Monotonically iterative method of nonlinear cantilever beam equations. Applied Mathematics and Computation, 205(1):432–437, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006036.

Yalcin:2009:FVA

[YAO09] Hasan Serter Yalcin, Aytac Arikoglu, and Ibrahim Ozkol. Free vibration analysis of circular plates by differential transformation method. Applied Mathematics and Computation, 212(2):377–386, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001696.

Yaxuan:2007:RFS

[Yax07] Yu Yaxuan. Rational formal solutions of hybrid lattice equation. Applied Mathematics and Computation, 186(1):474–482, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009568.

Yamac:2003:ASD

[YB03] Kerem Yamaç and Durmu Bozkurt. An algorithm for stability of discrete-time interval matrices. Applied Mathematics and Computation, 139(1):121–131, July 1, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yamac:2004:SDT

[YB04] Kerem Yamaç and Durmuş Bozkurt. On stability of discrete-time interval matrices. Applied Mathematics and Computation, 152(1):163–167, April 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See erratum [Yam05b].

Yurusoy:2005:NMA

[YB05] Muhammet Yürüsoy and Hüseyin Bayrakçeken. Numerical method in the analysis of slider bearing with Power-law fluid. Applied Mathematics and Computation, 162(1):491–501, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001110.

Yousefi:2006:NSF

[YB06] S. Yousefi and A. Banifatemi. Numerical solution of Fredholm integral equations by using CAS wavelets. Applied Mathematics and Computation, 183(1):458–463, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005789.

Yusufoglu:2007:TWS

[YB07] E. Yusufoğlu and A. Bekir. A travelling wave solution to the Ostrovsky equation. Applied Mathematics and Computation, 186(1):256– 260, March 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306009350.

Yebdri:2005:IMF

[YBA05] Mustapha Yebdri, Sidi Mohammed Bouguima, and Ovide Arino. An iterative method for functional differential equations. Applied Mathematics and Computation, 161(1):265–269, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012487.

Ye:2005:FSM

[YC05a] Xingde Ye and Xiaoliang Cheng. The Fourier spectral method for the Cahn– Hilliard equation. Applied Mathematics and Computation, 171(1):345–357, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001001.

Yu:2005:ETG

[YC05b] Yuan-Lung Yu and Tzer-Shyong Chen. An efficient threshold group signature scheme. Applied Mathematics and Computation, 167(1):362–371, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004990.

Yang:2006:NAC

[YC06a] Lihong Yang and Minggen Cui. New algorithm for a class of nonlinear integro-differential equations in the reproducing kernel space. Applied Mathematics and Computation, 174(2):942–960, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005230.

Yang:2006:USP

[YC06b] Yu Yang and Wencheng Chen. Uniformly strong persistence of a nonlinear asymptotically periodic multispecies competition predator–prey system with general functional response. Applied Mathematics and Computation, 183(1):423–426, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005741.

Yang:2006:SIE

[YC06c] Yu-Ching Yang and Win-Jin Chang. Simultaneous inverse estimation for boundary heat and moisture fluxes of a double-layer annular cylinder with interface resistance. Applied Mathematics and Computation, 176(2):594–608, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008581.

Yong:2006:DRN

[YC06d] Xuelin Yong and Yufu Chen. Direct reductions of a nonlinear partial differential system. Applied Mathematics and Computation, 183(2):942–945, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006278.

Yao:2007:NAC

[YC07a] Huanmin Yao and Minggen Cui. A new algorithm for a class of singular boundary value problems. Applied Mathematics and Computation, 186(2):1183–1191, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010447.

Yong:2007:TAF

[YC07b] Xuelin Yong and Yufu Chen. On two approaches of finding exact solutions to nonlinear evolution equations. Applied Mathematics and Computation, 194(1):74–84, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004456.

Yang:2008:FNN

[YC08a] Yongqing Yang and Jinde Cao. A feedback neural network for solving convex constraint optimization problems. Applied Mathematics and Computation, 201(1–2):340–350, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307012118.

Yi:2008:EBA

[YC08b] Fahuai Yi and Yingshan Chen. Exercise boundary of American-style Asian option. Applied Mathematics and Computation, 204(1):70–81, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004566.

Yuan:2008:NVE

[YC08c] Hua Yuan and Guoqing Chen. Network virusepidemic model with the point-to-group information propagation. Applied Mathematics and Computation, 206(1):357–367, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006735.

Yao:2009:NAS

[YC09a] Baoheng Yao and Jianping Chen. A new analytical solution branch for the Blasius equation with a shrinking sheet. Applied Mathematics and Computation, 215(3):1146–1153, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006274.

Yao:2009:SSF

[YC09b] Baoheng Yao and Jianping Chen. Series solution to the Falkner– Skan equation with stretching boundary. Applied Mathematics and Computation, 208(1):156–164, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800876X.

Yeh:2009:SDC

[YC09c] Wei-Chang Yeh and H. W.

Corley. A simple direct cosine simplex algorithm. Applied Mathematics and Computation, 214(1):178–186, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002926.

Yeh:2008:RDT

[YCC08] Ken Yeh, Chen-Yuan Chen, and Cheng-Wu Chen. Robustness design of time-delay fuzzy systems using fuzzy Lyapunov method. Applied Mathematics and Computation, 205(2):568–577, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003354.

Ye:2007:SOC

[YCL07] Yuquan Ye, Chi Kin Chan, and B. P. K. Leung. Some optimality conditions of quasilinear elliptic obstacle optimal control problems. Applied Mathematics and Computation, 188(2):1757–1771, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015438.

Yang:2009:APT

[YCL09a] Pei Yang, Yong Chen, and Zhi-Bin Li. ADM-padé technique for the nonlinear lattice equations. Applied Mathematics and Computation, 210(2):362–375, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000022.

Ye:2009:ERO

[YCL09b] Yuquan Ye, Chi Kin Chan, and H. W. J. Lee. The existence results for obstacle optimal control problems. Applied Mathematics and Computation, 214(2):451–456, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003300.

Yang:2005:IMM

[YCW05] Jen-Ho Yang, Chin-Chen Chang, and Chih-Hung Wang. An iterative modular multiplication algorithm in RNS. Applied Mathematics and Computation, 171(1):637–645, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001219.

Yuelin:2005:NTL

[YCYL05] Gao Yuelin, Xu Chengxian, Wang Yanjun, and Zhang Lianshen. A new two-level linear relaxed bound method for geometric programming problems. Applied Mathematics and Computation, 164(1):117– 131, May 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002991.

Yang:2006:EPS

[YCZ06] Yu Yang, Wencheng Chen, and Cuimei Zhang. Existence of periodic solutions of a delayed predator–prey system with general functional response. Applied Mathematics and Computation, 181(2):1076–1083, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001913.

Yang:2007:IFA

[YDL07] Yongjian Yang, Xuewu Du, and Mingming Li. A integral filter algorithm for unconstrained global optimization. Applied Mathematics and Computation, 184(2):173–180, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600693X.

Ye:2006:ISI

[Ye06] Qiang Ye. An iterated shift-and-invert Arnoldi algorithm for quadratic matrix eigenvalue problems. Applied Mathematics and Computation, 172(2):818–827, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009154.

Yazdi:2007:ESC

[YE07] Hadi Sadoghi Yazdi and Sohrab Effati. Eigenvalue spread criteria in the particle swarm optimization algorithm for solving of constraint parametric problems. Applied Mathematics and Computation, 192(1): 40–50, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307002913.

Youssef:2007:PIA

[YEA07] I. K. Youssef and H. A. El-Arabawy. Picard iteration algorithm combined with Gauss–Seidel technique for initial value problems. Applied Mathematics and Computation, 190(1):345– 355, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000550.

Yang:2005:NAD

[YEM05] Xiaofan Yang, David J. Evans, and Graham M. Megson. On a nonautonomous difference equation. Applied Mathematics and Computation, 168(1):380–388, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304006101.

Yang:2006:TRD

[YEM06] Xiaofan Yang, David J. Evans, and Graham M. Megson. On two rational difference equations. Applied Mathematics and Computation, 176(2):422–430, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008180.

Yousefpoor:2008:LSA

[YEN08] P. Yousefpoor, M. S. Esfahani, and H. Nojumi. Looking for systematic approach to select chaos tests.

Applied Mathematics and Computation, 198(1):73– 91, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008685.

Yazdi:2007:PCS

[YES07] Hadi Sadoghi Yazdi, Sohrab Effati, and Zahra Saberi. The probabilistic constraints in the support vector machine. Applied Mathematics and Computation, 194(2):467–479, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004961.

Yang:2008:AEV

[YF08] Hyun Mo Yang and Cláudia Pio Ferreira. Assessing the effects of vector control on dengue transmission. Applied Mathematics and Computation, 198(1):401–413, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009101.

Ye:2009:EEP

[YFLX09] Jianxiong Ye, Enmin Feng, Hansheng Lian, and Zhilong Xiu. Existence of equilibrium points and stability of the nonlinear dynamical system in microbial continuous cultures. Applied Mathematics and Computation, 207(2):307–318, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008102.

Yang:2005:AAB

[YGL05] Wu-Chuan Yang, D. J. Guan, and Chi-Sung Laih. Algorithm of asynchronous binary signed-digit recoding on fast multiexponentiation. Applied Mathematics and Computation, 167(1):108–117, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004771.

Yu:2009:GCP

[YGW09] Gaohang Yu, Lutai Guan, and Zengxin Wei. Globally convergent Polak– Ribière–Polyak conjugate gradient methods under a modified Wolfe line search. Applied Mathematics and Computation, 215(8):3082–3090, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008923.

Yuan:2009:OMV

[YGY09] Li Yuan, Hong Guo, and Zhaohua Yin. On optimal message vector length for block single parallel partition algorithm in a three-dimensional ADI solver. Applied Mathematics and Computation, 215(7):2565–2577, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007954.

Yang:2009:CCC

[YH09] Shouzhi Yang and Yongdong Huang. Construction of a class of compactly supported symmetric and balanced refinable function vector by GTST. Applied Mathematics and Computation, 207(1):83–89, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000210.

Yuan:2008:ESM

[YHL08a] Quan Yuan, Zhiqing He, and Huinan Leng. An evolution strategy method for computing eigenvalue bounds of interval matrices. Applied Mathematics and Computation, 196(1):257–265, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006650.

See erratum [YHL09b].

Yuan:2008:HGA

[YHL08b] Quan Yuan, Zhiqing He, and Huinan Leng. A hybrid genetic algorithm for a class of global optimization problems with box constraints. Applied Mathematics and Computation, 197(2):924–929, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008703.

See erratum [YHL09a].

Yuan:2008:ICC

[YHL08c] Quan Yuan, Zhiqing He, and Huinan Leng. An improvement for Chebyshev collocation method in solving certain Sturm– Liouville problems. Applied Mathematics and Computation, 195(2):440–447, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005644.

Yuan:2009:EHG

[YHL09a] Quan Yuan, Zhiqing He, and Huinan Leng. Erratum to: “A hybrid genetic algorithm for a class of global optimization problems with box constraints” [Appl. Math. Comput. 197 (2008) 924–929]. Applied Mathematics and Computation, 215(4):1646, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006456.

See [YHL08b].

Yuan:2009:EES

[YHL09b] Quan Yuan, Zhiqing He, and Huinan Leng. Erratum to: “An evolution strategy method for computing eigenvalue bounds of interval matrices” [Appl. Math. Comput. 196 (2008) 257–265]. Applied Mathematics and Computation, 215(3):1307, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006432.

See [YHL08a].

Yan:2009:MPM

[YHS09] Xihong Yan, Deren Han, and Wenyu Sun. A modified projection method with a new direction for solving variational inequalities. Applied Mathematics and Computation, 211(1):118–129, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000484.

Yang:2005:IBN

[YHTR05] Fenghong Yang, Miao He, Yun Tang, and Ming Rao. Inexact block Newton methods for solving nonlinear equations. Applied Mathematics and Computation, 162(3):1207–1218, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030400181X.

Yu:2006:CHI

[YHW06] Qian Yu, Chongchao Huang, and Xianjia Wang. A combined homotopy interior point method for the linear complementarity problem. Applied Mathematics and Computation, 179(2):696–701, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010751.

Yi:2005:GCS

[Yi05] Fahuai Yi. Global classical solution of quasi-stationary Stefan free boundary problem. Applied Mathematics and Computation, 160(3):797–817, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303011962.

Yigit:2008:AAN

[Yig08]  Faruk Yigit.  Approxi-

mate analytical and numerical solutions for a twodimensional Stefan problem. Applied Mathematics and Computation, 202(2):857–869, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001902.

Yilmaz:2006:CWR

[Yil06]  L. Yilmaz.  Chaos in

the water resources system. Applied Mathematics and Computation, 183(2):761–773, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006102.

Yin:2007:MMC

[Yin07] Peng-Yeng Yin. Multilevel minimum cross entropy threshold selection based on particle swarm optimization. Applied Mathematics and Computation, 184(2):503–513, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007260.

Yin:2008:CAP

[Yin08a] Qingxiang Yin. On computing of arbitrary positive powers for antitridiagonal matrices of even order. Applied Mathematics and Computation, 203(1):252–257, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002464.

Yingxin:2008:ENS

[Yin08b]  Guo Yingxin.  Erra-

tum to “Nontrivial solution of third-order nonlinear eigenvalue problems” [Appl. Math. Comput. 176 (2006) 714–721]. Applied Mathematics and Computation, 200(1):477, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307010363.

See [ZLW06c].

Yin:2009:PBS

[Yin09] Jun-Feng Yin. Preconditioner based on the Sherman–Morrison formula for regularized least squares problems. Applied Mathematics and Computation, 215(8):3007–3016, December 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030900873X.

Yang:2007:SAP

[YJ07] Geng Yang and Peng Jiang.

SSOR and ASSOR preconditioners for Block– Broyden method. Applied Mathematics and Computation, 188(1):194–205, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013385.

Yang:2009:MCM

[YJ09] Suxiang Yang and Ziwen Jiang. Mixed covolume method for parabolic problems on triangular grids. Applied Mathematics and Computation, 215(3):1251–1265, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006407.

Yuan:2008:EUP

[YJO08] Chengjun Yuan, Daqing Jiang, and Donal O’Regan. Existence and uniqueness of positive solutions for fourth-order nonlinear singular continuous and discrete boundary value problems. Applied Mathematics and Computation, 203(1):194–201, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002282.

Yeh:2006:AFV

[YJW06] Yen-Liang Yeh, Ming-Jyi Jang, and Cheng Chi Wang. Analyzing the free vibrations of a plate using finite difference and differential transformation method. Applied Mathematics and Computation, 178(2):493–501, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009963.

Yildiz:2003:FSM

[YK03] Bünyamin Yildiz and Erdal Karaduman. On Fibonacci search method with k-Lucas numbers. Applied Mathematics and Computation, 143(2–3):523–531, November 10, 2003. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See comment [Kah05a].

Yuchi:2006:EAU

[YK06a] Ming Yuchi and Jong-Hwan Kim. Evolutionary algorithm using feasibilitybased grouping for numerical constrained optimization problems. Applied Mathematics and Computation, 175(2):1298–1319, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500679X.

Yun:2006:PRM

[YK06b] Jae Heon Yun and Sang Wook Kim. Parallel relaxed multisplitting methods for a symmetric positive definite matrix. Applied Mathematics and Computation, 176(1):150–165, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007812.

Yun:2008:CPA

[YK08] Jae Heon Yun and Sang Wook Kim. Convergence of the preconditioned AOR method for irreducible Lmatrices. Applied Mathematics and Computation, 201(1–2):56–64, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011599.

See erratum [WH09].

Yuchi:2007:PEI

[YKJ07] Ming Yuchi, Jong-Hwan Kim, and Jun Jo. A population ecology inspired parent selection strategy for numerical constrained optimization problems. Applied Mathematics and Computation, 190(1):292–304, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000495.

Yi:2009:CFP

[YKKY09] Dokkyun Yi, Do-Hyung Kim, Eunyoun Kim, and Sung-Dae Yang. Convergence of a fixed point iteration method for the OSV model. Applied Mathematics and Computation, 215(5):1780–1790, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006778.

Yang:2008:ACA

[YKL08] Xiaojing Yang, Yong-In Kim, and Kueiming Lo. Approximation for constant e and its applications. Applied Mathematics and Computation, 206(1):50–55, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006279.

Yun:2006:AMF

[YKR06] Beong In Yun, Hyun Chul Kim, and Kyung Soo Rim. An averaging method for the Fourier approximation to discontinuous functions. Applied Mathematics and Computation, 183(1):272–284, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005595.

Yan:2006:HBG

[YL06a] Xiang-Ping Yan and Wan-Tong Li. Hopf bifurcation and global periodic solutions in a delayed predator–prey system. Applied Mathematics and Computation, 177(1):427– 445, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009227.

Yao:2006:CMP

[YL06b] Ruo-Xia Yao and Zhi-Bin Li. CONSLAW: a Maple package to construct the conservation laws for nonlinear evolution equations. Applied Mathematics and Computation, 173(1):616–635, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305003802.

Yi:2006:FOP

[YL06c] Dokkyun Yi and Sungyun Lee. Fourth-order partial differential equations for image enhancement. Applied Mathematics and Computation, 175(1):430– 440, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006181.

Yuan:2007:SPC

[YL07] Shao-Mou Yuan and Zai-Ming Liu. Some properties of α-convex and αquasiconvex functions with respect to n-symmetric points. Applied Mathematics and Computation, 188(2):1142–1150, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014585.

Yang:2008:CMM

[YL08a] Ding-Gong Yang and Jin-Lin Liu. A class of meromorphically multivalent functions defined by means of a linear operator. Applied Mathematics and Computation, 204(2):862–871, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005717.

Yang:2008:WUD

[YL08b] Hu Yang and Hanyu Li. Weighted UDV ^{∗}- decomposition and weighted spectral decomposition for rectangular matrices and their applications. Applied Mathematics and Computation, 198(1):150–162, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008764.

Yang:2009:SWN

[YL09a] Minbo Yang and Baorong Li. Solitary waves for nonhomogeneous Schrödinger– Maxwell system. Applied Mathematics and Computation, 215(1):66– 70, September 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030900383X.

Yu:2009:ICD

[YL09b] Fajun Yu and Li Li. An integrable couplings of Dirac soliton hierarchy with selfconsistent sources. Applied Mathematics and Computation, 207(1):171–177, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308007765.

Yu:2009:NML

[YL09c] Fajun Yu and Li Li. A new matrix Lie algebra, the multicomponent Yang hierarchy and its super-integrable coupling system. Applied Mathematics and Computation, 207(2):380–387, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008242.

Yang:2005:SHO

[YLB05] Xiaofan Yang, Yaoxin Liu, and Sen Bai. On the system of high order rational difference equations x_{n} = a/y_{n−p}, y_{n} = by_{n−p}/(x_{n−q}y_{n−q}). Applied Mathematics and Computation, 171(2):853–856, December 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2006:NSS

[YLB06] Xiaofan Yang, Yaoxin Liu, and Sen Bai. A numerical solution of second-order linear partial differential equations by differential transform. Applied Mathematics and Computation, 173(2):792–802, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003930.

Yang:2009:HCT

[YLC09] Xiao-Song Yang, Qingdu Li, and Shijie Cheng. Horseshoe chaos and topological entropy estimate in a simple power system. Applied Mathematics and Computation, 211(2):467–473, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001040.

Lan:2009:GMO

[yLcCsL09] Heng you Lan, Le cai Cai, and Zi shan Liu. General A-monotone operators and perturbed iterations for nonlinear setvalued relaxed cocoercive operator inclusion problems. Applied Mathematics and Computation, 215(4):1583–1592, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006614.

Yin:2009:CHT

[YLH09] Jingxue Yin, Yinghua Li, and Rui Huang. The Cahn–Hilliard type equations with periodic potentials and sources. Applied Mathematics and Computation, 211(1):211–221, May 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000721.

Yao:2009:SCI

[YLK09] Yonghong Yao, Yeong-Cheng Liou, and Shin Min Kang. Strong convergence of an iterative algorithm on an infinite countable family of nonexpansive mappings. Applied Mathematics and Computation, 208(1):211–218, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008916.

Yang:2005:GEP

[YLL^{+}05] Xiaofan Yang, Chuandong Li, Xiaofeng Liao, David J. Evans, and Graham M. Megson. Global exponential periodicity of a class of bidirectional associative memory networks with finite distributed delays. Applied Mathematics and Computation, 171(1):108–121, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000846.

Yang:2006:DEM

[YLL06a] Xiaofan Yang, Xiaofeng Liao, and Chuandong Li. On a difference equation with maximum. Applied Mathematics and Computation, 181(1):1–5, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306000609.

Yang:2006:FTH

[YLL06b] Xiaofan Yang, Jun Luo, and Shuangqing Li. Faulttolerant Hamiltonicity in a class of faulty meshes. Applied Mathematics and Computation, 182(2):1696– 1708, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306005182.

You:2007:NAM

[YLL07] Junyong You, Guizhong Liu, and Hongliang Li. A novel attention model and its application in video analysis. Applied Mathematics and Computation, 185(2):963–975, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008150.

Ling:2006:ABC

[yLLjP06] Liu yi Ling, Liang Liang, and Xu jin Pu. ATC based coordination of distributed production planning and supplier selection. Applied Mathematics and Computation, 182(2):1154–1168, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004309.

Yang:2008:BCS

[YLLL08] Xiaofan Yang, Qing Lu, Chuandong Li, and Xiaofeng Liao. Biological computation of the solution to the quadratic assignment problem. Applied Mathematics and Computation, 200(1):369–377, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011253.

Yen:2007:ICM

[YLM07] Sung-Ming Yen, Wei-Chih Lien, and SangJae Moon. Inefficiency of commonmultiplicand multiplication and exponentiation algorithms by performing binary complements. Applied Mathematics and Computation, 189(1):285–290, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016018.

Yang:2005:SRK

[YLS05] Zhanwen Yang, Mingzhu Liu, and Minghui Song. Stability of Runge–Kutta methods in the numerical solution of equation u′(t) = au(t) + a_{0}u([t]) + a_{1}u([t − 1]). Applied Mathematics and Computation, 162(1):37–50, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000542.

Yang:2008:MLS

[YLX08] Hu Yang, Deqiang Liu, and Jianwen Xu. Matrix left symmetry factor and its applications in generalized inverses A_{T,S}^{(2,4)}. Applied Mathematics and Computation, 197(2):836– 843, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008533.

Yao:2007:IAA

[YLY07] Yonghong Yao, Yeong-Cheng Liou, and Jen-Chih Yao. An iterative algorithm for approximating convex minimization problem. Applied Mathematics and Computation, 188(1):648– 656, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014032.

Yan:2006:GAS

[YLZ06] Xing-Xue Yan, Wan-Tong Li, and Zhu Zhao. Global asymptotic stability for a higher order nonlinear rational difference equations. Applied Mathematics and Computation, 182(2):1819–1831, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005315.

Yan:2006:ADD

[YM06] Wen-Jing Yan and Yi-Chen Ma. The application of domain derivative for heat conduction with mixed condition in shape reconstruction. Applied Mathematics and Computation, 181(2):894–902, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600169X.

Yan:2007:FEM

[YM07a] Wen-Jing Yan and Yi-Chen Ma. Finite element methods for the viscous incompressible fluid. Applied Mathematics and Computation, 185(1):547–553, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008897.

Yashtini:2007:SCV

[YM07b] M. Yashtini and A. Malek. Solving complementarity and variational inequalities problems using neural networks. Applied Mathematics and Computation, 190(1):216–230, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000422.

Yashtini:2008:DTN

[YM08a] M. Yashtini and A. Malek. A discrete-time neural network for solving nonlinear convex problems with hybrid constraints. Applied Mathematics and Computation, 195(2):576–584, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005863.

Yuan:2008:P

[YM08b] Jin Yun Yuan and To Fu Ma. Preface. Applied Mathematics and Computation, 200(2):479, July 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011277.

Yang:2006:LBS

[YMCL06] Xiaofan Yang, Graham M.

Megson, Jianqiu Cao, and Jun Luo. A lower bound on the size of kneighborhood in generalized cubes. Applied Mathematics and Computation, 179(1):47–54, August 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010052.

Yan:2007:MPS

[YML07] Wei Yan, Rong Miao, and Shurong Li. Multi-period semi-variance portfolio selection: Model and numerical solution. Applied Mathematics and Computation, 194(1):128–134, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004511.

Matsushita:2008:AFP

[yMT08] Shin ya Matsushita and Wataru Takahashi. Approximating fixed points of nonexpansive mappings in a Banach space by metric projections. Applied Mathematics and Computation, 196(1):422–425, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307006959.

Yingzhen:2005:RES

[YMY05] Lin Yingzhen, Cui Minggen, and Zheng Yi. Representation of the exact solution for infinite system of linear equations. Applied Mathematics and Computation, 168(1):636–650, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304006290.

Nie:2005:FMS

[yN05]  Pu yan Nie.  A fil-

ter method for solving nonlinear complementarity problems. Applied Mathematics and Computation, 167(1):677–694, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304005235.

Nie:2006:CLA

[yN06] Pu yan Nie. CDT like approaches for the system of nonlinear equations. Applied Mathematics and Computation, 172(2):892–902, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002444.

Yao:2007:CTS

[YN07] Yonghong Yao and Muhammad Aslam Noor. Convergence of three-step iterations for asymptotically nonexpansive mappings. Applied Mathematics and Computation, 187(2):883–892, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012203.

Yao:2008:SCT

[YNCL08] Yonghong Yao, Muhammad Aslam Noor, Rudong Chen, and Yeong-Cheng Liou. Strong convergence of three-step relaxed hybrid steepestdescent methods for variational inequalities. Applied Mathematics and Computation, 201(1–2):175–183, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011836.

Nie:2006:TRF

[yNfM06] Pu yan Nie and Chang feng Ma. A trust region filter method for general non-linear programming. Applied Mathematics and Computation, 172(2):1000–1017, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002742.

Nie:2005:DFF

[yNyF05] Pu yan Nie and Jin yan Fan. A derivative-free filter method for solving nonlinear complementarity problems. Applied Mathematics and Computation, 161(3):787–797, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400030X.

Yoon:2005:IAA

[Yoo05] Jungho Yoon. Improved accuracy of L_{p}-approximation to derivatives by radial basis function interpolation. Applied Mathematics and Computation, 161(1):109–119, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030301213X.

Youness:2005:LSA

[You05] Ebrahim A. Youness. Level set algorithm for solving convex multiplicative programming problems. Applied Mathematics and Computation, 167(2):1412–1417, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005764.

See comment [gZW08].

Yousefi:2006:LWM

[You06a] Sohrab Ali Yousefi. Legendre wavelets method for solving differential equations of Lane–Emden type. Applied Mathematics and Computation, 181(2):1417–1422, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002608.

Yousefi:2006:NSA

[You06b] Sohrab Ali Yousefi. Numerical solution of Abel’s integral equation by using Legendre wavelets. Applied Mathematics and Computation, 175(1):574–580, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006302.

Youssef:2006:GMT

[You06c] H. M. Youssef. Generalized magneto-thermoelasticity in a conducting medium with variable material properties. Applied Mathematics and Computation, 173(2):822–833, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305003954.

Yousuf:2008:CHO

[You08] M. Yousuf. On the class of high order time stepping schemes based on Padé approximations for the numerical solution of Burgers’ equation. Applied Mathematics and Computation, 205(1):442–453, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800605X.

Yousuf:2009:ESM

[You09] M. Yousuf. Efficient Lstable method for parabolic problems with application to pricing American options under stochastic volatility. Applied Mathematics and Computation, 213(1):121–136, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002094.

Peng:2005:IML

[yP05] Zhen yun Peng. An iterative method for the least squares symmetric solution of the linear matrix equation AXB = C. Applied Mathematics and Computation, 170(1):711–723, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000123.

Yanjun:2005:BBA

[YPZ05] Wang Yanjun, Shen Peiping, and Liang Zhian. A branch-and-bound algorithm to globally solve the sum of several linear ratios. Applied Mathematics and Computation, 168(1):89–101, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005880.

Yigui:2009:NTR

[YQ09] Ou Yigui and Zhou Qian.

A nonmonotonic trust region algorithm for a class of semi-infinite minimax programming. Applied Mathematics and Computation, 215(2):474–480, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004883.

Yuan:2007:SSL

[YQYL07] Zejian Yuan, Yanyun Qu, Chao Yang, and Yuehu Liu. A sequential subspace learning method and its application to dynamic texture analysis. Applied Mathematics and Computation, 185(2):834–843, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600765X.

Yoon:2005:CFI

[YRY05a] Eun-Jun Yoon, Eun-Kyung Ryu, and Kee-Young Yoo. Cryptanalysis and further improvement of Peinado’s improved LHL-key authentication scheme. Applied Mathematics and Computation, 168(2):788–794, September 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300304006393.

Yoon:2005:FPL

[YRY05b] Eun-Jun Yoon, Eun-Kyung Ryu, and Kee-Young Yoo. Fixing problems in Lin et al.’s OSPA protocol. Applied Mathematics and Computation, 166(1):46– 57, July 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004205.

Yoon:2005:ICJ

[YRY05c] Eun-Jun Yoon, Eun-Kyung Ryu, and Kee-Young Yoo. Improvement of Chien– Jan’s authenticated multiplekey agreement protocol without using conventional one-way function. Applied Mathematics and Computation, 167(1):711–715, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005302.

Yoon:2005:IFA

[YRY05d] Eun-Jun Yoon, Eun-Kyung Ryu, and Kee-Young Yoo. Improvement of Fan et al.’s deniable authentication protocol based on Diffie–Hellman algorithm. Applied Mathematics and Computation, 167(1):274–280, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004904.

Yalcinbas:2006:TCM

[YS06a] Salih Yalçinbaş and Mehmet Sezer. A Taylor collocation method for the approximate solution of general linear Fredholm–Volterra integro-difference equations with mixed argument. Applied Mathematics and Computation, 175(1):675– 690, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006387.

Yang:2006:NFF

[YS06b] Yongjian Yang and Youlin Shang. A new filled function method for unconstrained global optimization. Applied Mathematics and Computation, 173(1):501–512, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030500370X.

Sun:2006:CTI

[yS06c] Li ying Sun. A comparison theorem of the improving Gauss–Seidel method for H-matrix and its comparison matrix. Applied Mathematics and Computation, 183(1):390–393, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005704.

Yu:2006:LBM

[YS06d] Xiaomei Yu and Baochang Shi. A lattice Boltzmann model for reaction dynamical systems with time delay. Applied Mathematics and Computation, 181(2):958–965, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001767.

Yang:2007:NBV

[YS07] Xuxin Yang and Jianhua Shen. Nonlinear boundary value problems for first order impulsive functional differential equations. Applied Mathematics and Computation, 189(2):1943–1952, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000136.

Yang:2005:RS

[YSC^{+}05] Xiaofan Yang, Weifeng Su, Bill Chen, Graham M. Megson, and David J. Evans. On the recursive sequence x_{n} = (ax_{n−1} + bx_{n−2})/(c + dx_{n−1}x_{n−2}). Applied Mathematics and Computation, 162(3):1485–1497, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002024.

Yepez:2005:SAG

[YSCC05] E. Yépez, J. R. Soto, J. J. Castro, and A. Calles. A simple algorithm for the group theoretical classification of quantum states II. The case of molecular electronic states. Applied Mathematics and Computation, 164(3):719–729, May 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304003224.

Yago:2006:WPD

[YSdS^{+}06] J. M. Yago, F. García Sevilla, C. Garrido del Solo, R. G. Duggleby, and R. Varón. A Windows program for the derivation of steady-state equations in enzyme systems. Applied Mathematics and Computation, 181(2):837–852, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001640.

Yang:2005:NES

[YSL05] Shu Min Yang, Dong Guo Shao, and Yang Jie Luo. A novel evolution strategy for multiobjective optimization problem. Applied Mathematics and Computation, 170(2):850–873, November 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305000287.

Yalcinbas:2009:LPS

[YSS09] Salih Yalçinbaş, Mehmet Sezer, and Hüseyin Hilmi Sorkun. Legendre polynomial solutions of highorder linear Fredholm integro-differential equations. Applied Mathematics and Computation, 210(2):334–349, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009466.

Yamamura:2009:LNN

[YST09] Kiyotaka Yamamura, Koki Suda, and Naoya Tamura. LP narrowing: a new strategy for finding all solutions of nonlinear equations. Applied Mathematics and Computation, 215(1):405–413, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004834.

Yang:2008:CGE

[YSXL08] Xinan Yang, Xiao Sun, Jianming Xie, and Zuhong Lu. Comparability of gene expression in human blood, immune and carcinoma cells. Applied Mathematics and Computation, 205(1):178–184, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003287.

Yang:2005:CAO

[YSZG05] S. J. Yang, B. Shi, D. C. Zhang, and M. J. Gai. Counterexamples to the article “Oscillation and global attractivity in hematopoiesis model with delay time”. Applied Mathematics and Computation, 168(2):973–980, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006551.

Yu:2006:SCM

[YT06] Rachung Yu and Gwo-Hshiung Tzeng. A soft computing method for multi-criteria decision making with dependence and feedback. Applied Mathematics and Computation, 180(1):63–75, September 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305010982.

Yang:2006:ODP

[YTLH06] Xiaofan Yang, Yuan Yan Tang, Qing Lu, and Zhenyu He. Optimal doublecast path in hexagonal honeycomb mesh. Applied Mathematics and Computation, 182(2):1267–1279, November 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306004425.

Yang:2005:SEA

[YTWY05] Chou-Chen Yang, Yuan-Liang Tang, Ren-Chiun Wang, and Hung-Wen Yang. A secure and efficient authentication protocol for anonymous channel in wireless communications. Applied Mathematics and Computation, 169(2):1431–1439, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400863X.

Yu:2005:LAR

[Yu05]  Li Yu.  An LMI ap-

proach to reliable guaranteed cost control of discrete-time systems with actuator failure. Applied Mathematics and Computation, 162(3):1325–1331, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001912.

Yu:2007:CPC

[Yu07] Ming-Miin Yu. The capacity productivity change and the variable input productivity change: a new decomposition of the Malmquist productivity index. Applied Mathematics and Computation, 185(1):375–381, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008708.

Yu:2008:WAF

[Yu08a] Dansheng Yu. Weighted approximation of functions with singularities by combinations of Bernstein operators. Applied Mathematics and Computation, 206(2):906–918, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007522.

Yu:2008:CDN

[Yu08b] Fajun Yu. Constructing (2+1)-dimensional nonlinear soliton equations with a generalized zero curvature equation. Applied Mathematics and Computation, 201(1–2):739–746, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000416.

Yu:2008:NIH

[Yu08c] Fajun Yu. A new integrable hierarchy and constrained flows, its expanding integrable system. Applied Mathematics and Computation, 203(2):864–872, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004153.

Yu:2009:NII

[Yu09]  Fajun Yu.  A non-

isospectral integrable couplings of Volterra lattice hierarchy with self-consistent sources. Applied Mathematics and Computation, 215(3):1217–1223, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006365.

Yuan:2005:CPM

[Yua05] Dongjin Yuan. On the convergence of parallel multisplitting asynchronous GAOR method for Hmatrix. Applied Mathematics and Computation, 160(2):477–485, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303011743.

Yuan:2008:OSC

[Yua08] Yubo Yuan. Optimal solutions to a class of nonconvex minimization problems with linear inequality constraints. Applied Mathematics and Computation, 203(1):142–152, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002166.

Yuan:2009:SIE

[Yua09] Yongxin Yuan. A symmetric inverse eigenvalue problem in structural dynamic model updating. Applied Mathematics and Computation, 213(2):516–521, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002641.

Yucel:2008:HAM

[Yüc08]  Uḡur Yücel.  Homo-

topy analysis method for the sine-Gordon equation with initial conditions. Applied Mathematics and Computation, 203(1):387–395, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002622.

Yueting:2007:IMS

[Yue07] Yang Yueting. The iterative method for solving nonlinear matrix equation X^{s} + A^{∗}X^{−t}A = Q. Applied Mathematics and Computation, 188(1):46– 53, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013178.

Yujian:2007:ASE

[Yuj07]  Li Yujian.  An ana-

lytic solution for estimating two-dimensional hidden Markov models. Applied Mathematics and Computation, 185(2):810–822, February 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306007636.

Yun:2007:CTS

[Yun07a] Jae Heon Yun. Convergence of two-stage and multisplitting methods using various types of splittings. Applied Mathematics and Computation, 189(1):87– 100, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015839.

Yun:2007:NIM

[Yun07b] Jae Heon Yun. A note on the improving modified Gauss–Seidel (IMGS) method. Applied Mathematics and Computation, 184(2):674–679, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007429.

Yun:2007:NMS

[Yun07c] Jae Heon Yun. A note on the modified SOR method for Z-matrices. Applied Mathematics and Computation, 194(2):572–576, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004973.

Yun:2008:NIM

[Yun08a] Beong In Yun. A noniterative method for solving non-linear equations. Applied Mathematics and Computation, 198(2):691– 699, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009368.

Yun:2008:NTS

[Yun08b]  Jae Heon Yun.  A

note on three-step iterative method for nonlinear equations. Applied Mathematics and Computation, 202(1):401–405, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000878.

Yun:2009:CTS

[Yun09] Jae Heon Yun. Convergence of two-stage multisplitting method using AOR or SSOR multisplittings. Applied Mathematics and Computation, 215(7):2694–2700, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309008078.

Yurekli:2007:IFI

[Yür07]  O. Yürekli.  Identities

on fractional integrals and various integral transforms. Applied Mathematics and Computation, 187(1):559– 566, April 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012082.

Yadavalli:2006:STP

[YVU06] V. S. S. Yadavalli, C. De W. Van Schoor, and S. Udayabaskaran. A substitutable two-product inventory system with joint-ordering policy and common demand. Applied Mathematics and Computation, 172(2):1257–1271, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002912.

Yang:2005:CIA

[YW05] Chou-Chen Yang and Ren-Chiun Wang. Cryptanalysis of improved authenticated multiple-key agreement protocol without using conventional oneway function. Applied Mathematics and Computation, 162(1):211–214, March 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304000815.

Yin:2006:ACO

[YW06a] Peng-Yeng Yin and Jing-Yu Wang. Ant colony optimization for the nonlinear resource allocation problem. Applied Mathematics and Computation, 174(2):1438–1453, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500562X.

Yin:2006:PSO

[YW06b] Peng-Yeng Yin and Jing-Yu Wang. A particle swarm optimization approach to the nonlinear resource allocation problem. Applied Mathematics and Computation, 183(1):232–242, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005558.

Yang:2008:PSO

[YW08] Jingbao Yang and Zhongli Wei. Positive solutions of nth order m-point boundary value problem. Applied Mathematics and Computation, 202(2):715–720, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308001616.

Yu:2009:DCI

[YW09] Yaoming Yu and Guorong Wang. DFT calculation for the 2-inverse of a polynomial matrix with prescribed image and kernel. Applied Mathematics and Computation, 215(7):2741–2749, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309008133.

Yang:2005:IYS

[YWC05] Chou-Chen Yang, Ren-Chiun Wang, and Ting-Yi Chang. An improvement of the Yang-Shieh password authentication schemes. Applied Mathematics and Computation, 162(3):1391–1396, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001961.

Yeh:2007:UFD

[YWJ07] Yen-Liang Yeh, Cheng Chi Wang, and Ming-Jyi Jang. Using finite difference and differential transformation method to analyze of large deflections of orthotropic rectangular plate problem. Applied Mathematics and Computation, 190(2):1146–1156, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001348.

Wang:2007:APG

[yWjCnG07] Xing yuan Wang, Pei jun Chang, and Ni ni Gu. Additive perturbed generalized Mandelbrot–Julia sets. Applied Mathematics and Computation, 189(1):754– 765, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016535.

Wang:2006:GCP

[yWjL06] Chang yu Wang and Shu jun Lian. Global convergence properties of the two new dependent Fletcher– Reeves conjugate gradient methods. Applied Mathematics and Computation, 181(2):920–931, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600172X.

Yu:2007:NSR

[YWL07a] Yuexin Yu, Liping Wen, and Shoufu Li. Nonlinear stability of Runge– Kutta methods for neutral delay integro-differential equations. Applied Mathematics and Computation, 191(2):543–549, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002822.

Yu:2007:SAG

[YWL07b] Yuexin Yu, Liping Wen, and Shoufu Li. Stability analysis of general linear methods for nonlinear neutral delay differential equations. Applied Mathematics and Computation, 187(2):1389–1398, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012768.

Yang:2009:ESP

[YWL09] Jingbao Yang, Zhongli Wei, and Ke Liu. Existence of symmetric positive solutions for a class of Sturm– Liouville-like boundary value problems. Applied Mathematics and Computation, 214(2):424–432, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003270.

Wang:2006:FIS

[yWyCqD06] Chang yu Wang, Yuan yuan Chen, and Shou qiang Du. Further insight into the Shamanskii modification of Newton method. Applied Mathematics and Computation, 180(1):46– 52, September 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305010969.

Yi:2007:BNN

[YWZW07] Jianqiang Yi, Qian Wang, Dongbin Zhao, and John T. Wen. BP neural network prediction-based variableperiod sampling approach for networked control systems. Applied Mathematics and Computation, 185(2):976–988, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008186.

Yang:2006:IES

[YX06a] Zhichun Yang and Daoyi Xu. Impulsive effects on stability of Cohen– Grossberg neural networks with variable delays. Applied Mathematics and Computation, 177(1):63– 78, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008878.

Yao:2006:CDA

[YX06b] Hong-Xing Yao and Feng Xu. Complex dynamics analysis for a duopoly advertising model with nonlinear cost. Applied Mathematics and Computation, 180(1):134–145, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305011045.

You:2008:MCD

[YX08] Fucai You and Tiecheng Xia. The multi-component dispersive long wave equation hierarchy, its integrable couplings and their Hamiltonian structures.

Applied Mathematics and Computation, 201(1–2):44– 55, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307011551.

Ye:2009:OCS

[YX09] Luhong Ye and Zhiting Xu. Oscillation criteria for second order quasilinear neutral delay differential equations. Applied Mathematics and Computation, 207(2):388–396, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008254.

See remarks [HLSS10].

Xiao:2006:GQR

[yXhWZ06] Jin you Xiao, Li hua Wen, and Duo Zhang. Gauss quadrature rules for partial support and logarithmic singular integrals with compactly supported wavelets. Applied Mathematics and Computation, 179(2):572–580, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010635.

Yang:2005:RSN

[YXSX05] Xiaoli Yang, Wei Xu, Zhongkui Sun, and Yong Xu. Responses of strongly non-linear oscillator parametrically excited by random narrow-band noise. Applied Mathematics and Computation, 171(2):885– 899, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2007:EPB

[YXW07] Zuodong Yang, Bing Xu, and Mingzhu Wu. Existence of positive boundary blow-up solutions for quasilinear elliptic equations via sub and supersolutions. Applied Mathematics and Computation, 188(1):492– 498, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306013890.

Yang:2005:ASM

[YY05a] Jianhua Yang and Danping Yang. Additive Schwarz methods for parabolic problems. Applied Mathematics and Computation, 163(1):17–28, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304002048.

Yoon:2005:CZX

[YY05b] Eun-Jun Yoon and Kee-Young Yoo. Cryptanalysis of Zhang–Xiao’s multisignature scheme for specified group of verifiers. Applied Mathematics and Computation, 170(1):226–229, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304009208.

Yoon:2005:SWL

[YY05c] Eun-Jun Yoon and Kee-Young Yoo. On the security of Wu–Lin’s robust key authentication scheme. Applied Mathematics and Computation, 169(1):1–7, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007386.

Yoon:2005:WSY

[YY05d] Eun-Jun Yoon and Kee-Young Yoo. Weakness and solution of Yang et al.’s protected password changing scheme. Applied Mathematics and Computation, 169(1):295–304, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304007751.

Yao:2006:FRB

[YY06a] Chih-Chia Yao and Pao-Ta Yu. Fuzzy regression based on asymmetric support vector machines. Applied Mathematics and Computation, 182(1):175–193, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306003249.

Yin:2006:SNR

[YY06b] Honghui Yin and Zuodong Yang. Some new results on the existence of bounded positive entire solutions for quasilinear elliptic equations. Applied Mathematics and Computation, 177(2):606–613, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009410.

Yuan:2006:MCD

[YY06c] Jin Yun Yuan and Plamen Y. Yalamov. A method for constructing diagonally dominant preconditioners based on Jacobi rotations. Applied Mathematics and Computation, 174(1):74–80, March 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500442X.

Yao:2007:MIM

[YY07a] Yonghong Yao and Jen-Chih Yao. On modified iterative method for nonexpansive mappings and monotone mappings. Applied Mathematics and Computation, 186(2):1551–1558, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010824.

Yin:2007:NRE

[YY07b] Honghui Yin and Zuodong Yang. New results on the existence of bounded positive entire solutions for quasilinear elliptic systems. Applied Mathematics and Computation, 190(1):441– 448, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000744.

Yin:2007:VMM

[YY07c] Xianggang Yin and Weidong Yu. The virtual manufacturing model of the worsted yarn based on artificial neural networks and grey theory. Applied Mathematics and Computation, 185(1):322–332, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008630.

Yang:2008:DE

[YY08a] Yi Yang and Xiaofan Yang. On the difference equation x_{n+1} = \frac{px_{n−s}+x_{n−t}}{qx_{n−s}+x_{n−t}}. Applied Mathematics and Computation, 203(2):903–907, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308001811.

Yau:2008:CSD

[YY08b] Her-Terng Yau and Jun-Juh Yan. Chaos synchronization of different chaotic systems subjected to input nonlinearity. Applied Mathematics and Computation, 197(2):775–788, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008375.

Yuan:2008:ELS

[YY08c] Junli Yuan and Zuodong Yang. Existence of large solutions for a class of quasilinear elliptic equations. Applied Mathematics and Computation, 201(1–2):852–858, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307012349.

Yuan:2009:CAV

[YY09] Quan Yuan and Xiao-Song Yang. Computer-assisted verification of chaos in the model of nuclear spin generator. Applied Mathematics and Computation, 213(1):148–152, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002112.

Yang:2007:DEI

[YYL07] Xiaofan Yang, Yi Yang, and Jun Luo. On the difference equation x_{n} = (p+x_{n−1})/(qx_{n−1} +x_{n−1}). Applied Mathematics and Computation, 189(1):918– 926, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016717.

Yuan:2009:SND

[YYMZ09] Zhiling Yuan, Xingwei Yuan, Fanwei Meng, and Hongxia Zhang. Some new delay integral inequalities and their applications. Applied Mathematics and Computation, 208(1):231–237, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008941.

Yin:2007:MOT

[YYWW07] Peng-Yeng Yin, Shiuh-Sheng Yu, Pei-Pei Wang, and Yi-Te Wang. Multiobjective task allocation in distributed computing systems by hybrid particle swarm optimization. Applied Mathematics and Computation, 184(2):407–420, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007168.

Yang:2007:MPS

[YYYM07] Xueming Yang, Jinsha Yuan, Jiangye Yuan, and Huina Mao. A modified particle swarm optimizer with dynamic adaptation. Applied Mathematics and Computation, 189(2):1205–1213, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017048.

Yin:2006:OFV

[YZ06] Youbing Yin and Ke-Qin Zhu. Oscillating flow of a viscoelastic fluid in a pipe with the fractional Maxwell model. Applied Mathematics and Computation, 173(1):231–242, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003395.

Yang:2007:NIW

[YZ07a] Shouyuan Yang and Xingwei Zhou. A note on irregular wavelet frames. Applied Mathematics and Computation, 187(2):1410–1416, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012781.

Yang:2007:PSC

[YZ07b] Zhihui Yang and Zhihu Zhou. Periodic solutions of a class of neutral differential models with feedback control. Applied Mathematics and Computation, 189(1):996–1009, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016791.

Yu:2007:CMF

[YZ07c] Fajun Yu and Hongqing Zhang. Constructing a multi-function equation hierarchy and its integrable coupling system with MAPLE. Applied Mathematics and Computation, 187(2):1208–1218, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012574.

Yu:2007:NFO

[YZ07d] Fajun Yu and Hongqing Zhang. A new fractional order soliton equation hierarchy and its integrable coupling system. Applied Mathematics and Computation, 194(1):259–266, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307004651.

Zhang:2007:DME

[yZ07e] Zhen yu Zhang. A development of modal expansion method for eigensensitivity analysis of a defective matrix. Applied Mathematics and Computation, 188(2):1995–2019, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015694.

Yu:2008:HSI

[YZ08] Fajun Yu and Hongqing Zhang. Hamiltonian structure of the integrable couplings for the multicomponent Dirac hierarchy. Applied Mathematics and Computation, 197(2):828– 835, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008521.

Yang:2009:ESS

[YZ09a] Yang Yang and Jihui Zhang. Existence of solutions for some discrete boundary value problems with a parameter. Applied Mathematics and Computation, 211(2):293–302, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000733.

Yang:2009:IMM

[YZ09b] Yang Yang and Jihui Zhang. Infinitely many mountain pass solutions on a kind of fourthorder Neumann boundary value problem. Applied Mathematics and Computation, 213(1):262–271, July 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002276.

Yu:2008:DAP

[YZD08] Shiwei Yu, Kejun Zhu, and Fengqin Diao. A dynamic all parameters adaptive BP neural networks model and its application on oil reservoir prediction. Applied Mathematics and Computation, 195(1):66–75, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005188.

Ye:2007:NAW

[YZDW07] Dengpan Ye, Changfu Zou, Yuewei Dai, and Zhiquan Wang. A new adaptive watermarking for real-time MPEG videos. Applied Mathematics and Computation, 185(2):907–918, February 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306008083.

Yu:2008:SES

[YZFL08] Jianjiang Yu, Kanjian Zhang, Shumin Fei, and Tao Li. Simplified exponential stability analysis for recurrent neural networks with discrete and distributed time-varying delays. Applied Mathematics and Computation, 205(1):465–474, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006115.

Yu:2008:PEE

[YZLD08] Li Yu, Jiabo Zhang, Yuwu Liao, and Jie Ding. Parameter estimation error bounds for Hammerstein nonlinear finite impulsive response models. Applied Mathematics and Computation, 202(2):472–480, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000143.

Zhang:2006:CEE

[yZsZ06] Zhen yu Zhang and Hui sheng Zhang. Calculation of eigenvalue and eigenvector derivatives of a defective matrix. Applied Mathematics and Computation, 176(1):7–26, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305007666.

Yang:2007:GSA

[YZW07a] Fengjian Yang, Chaolong Zhang, and Dongqing Wu. Global stability analysis of impulsive BAM type Cohen– Grossberg neural networks with delays. Applied Mathematics and Computation, 186(1):932–940, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306010162.

Yu:2007:DNC

[YZW07b] Gaohang Yu, Yanlin Zhao, and Zengxin Wei. A descent nonlinear conjugate gradient method for large-scale unconstrained optimization. Applied Mathematics and Computation, 187(2):636–643, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306011192.

Yu:2007:SGC

[YZW07c] Zhensheng Yu, Weiguo Zhang, and Baofeng Wu. Strong global convergence of an adaptive nonmonotone memory gradient method. Applied Mathematics and Computation, 185(1):681–688, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009039.

Yang:2008:LSC

[YZY08] Changcai Yang, Sheng Zheng, and Jin Ye. Level set contour extraction method based on support value filter. Applied Mathematics and Computation, 205(2):688–696, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003470.

Ye:2008:FAI

[YZZP08] Yalan Ye, Zhi-Lin Zhang, Jiazhi Zeng, and Lei Peng. A fast and adaptive ICA algorithm with its application to fetal electrocardiogram extraction. Applied Mathematics and Computation, 205(2):799–806, November 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003585.

Zahran:2009:EWS

[Zah09] Yousef Hashem Zahran. An efficient WENO scheme for solving hyperbolic conservation laws. Applied Mathematics and Computation, 212(1):37–50, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001155.

Zhang:2005:ITA

[ZAX05] Zhang Zhang, Shunsuke Araki, and Guozhen Xiao. Improvement of Tseng et al.’s authenticated encryption scheme with message linkages. Applied Mathematics and Computation, 162(3):1475–1483, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002012.

Zhang:2006:IEI

[ZBH06] Wen Dong Zhang, Yan Ping Bai, and Hong Ping Hu. The incorporation of an efficient initialization method and parameter adaptation using selforganizing maps to solve the TSP. Applied Mathematics and Computation, 172(1):603–623, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002262.

Zhang:2005:AVC

[ZC05a] Ming-Heng Zhang and Qian-Sheng Cheng. An approach to VaR for capital markets with Gaussian mixture. Applied Mathematics and Computation, 168(2):1079–1085, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006642.

Zhang:2005:OFS

[ZC05b] Xianfu Zhang and Zhaolin Cheng. Output feedback stabilization of nonlinear systems with delays in the input. Applied Mathematics and Computation, 167(2):1026–1040, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005478.

Zhang:2006:VCA

[ZC06a] Wei Zhang and Li-Qun Chen. Vibration control of an axially moving string system: Wave cancellation method. Applied Mathematics and Computation, 175(1):851–863, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305006491.

Zhang:2006:CDI

[ZC06b] Xiangyun Zhang and Guoliang Chen. The computation of Drazin inverse and its application in Markov chains. Applied Mathematics and Computation, 183(1):292–300, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005613.

Zhao:2006:CFD

[ZC06c] Jichao Zhao and Robert M. Corless. Compact finite difference method for integrodifferential equations. Applied Mathematics and Computation, 177(1):271– 288, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009094.

Zheng:2006:GTL

[ZC06d] Zhaowen Zheng and Shaozhu Chen. GKN theory for linear Hamiltonian systems. Applied Mathematics and Computation, 182(2):1514–1527, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004930.

Zhao:2008:MMF

[ZC08a] Shengjie Zhao and Yufu Chen. Mixed moving finite element method. Applied Mathematics and Computation, 196(1):381–391, February 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307006911.

Zhao:2008:MST

[ZC08b] Yulin Zhao and Haibo Chen. Multiplicity of solutions to two-point boundary value problems for second-order impulsive differential equations. Applied Mathematics and Computation, 206(2):925–931, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007650.

Zhang:2009:IMS

[ZC09] Jiawei Zhang and Binghe Chen. An iterative method for solving the Falkner– Skan equation. Applied Mathematics and Computation, 210(1):215–222, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030800996X.

Zhou:2007:NST

[ZCFX07] Jun Zhou, Li Cai, Jian-Hu Feng, and Wen-Xian Xie. Numerical simulation for two-phase flows using hybrid scheme. Applied Mathematics and Computation, 186(2):980–991, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030601023X.

Zhou:2007:EAD

[ZCH07] Zhong Zhou, Anthony Chen, and Deren Han. An extended alternating direction method for variational inequality problems with linear equality and inequality constraints. Applied Mathematics and Computation, 184(2):769–782, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007569.

Zhou:2005:PSP

[ZCL05] Yuan Zhou, ZhenFu Cao, and RongXing Lu. Provably secure proxy-protected signature schemes based on factoring. Applied Mathematics and Computation, 164(1):83–98, May 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002978.

Zhang:2007:SLO

[ZCMQ07] Qing-Hua Zhang, Shuiming Chen, Jian Ma, and Yuanyuan Qu. Solutions of linear ordinary differential equations with non-singular varying coefficients by using the corrected Fourier series. Applied Mathematics and Computation, 187(2):765–776, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011829.

Zhu:2005:NAC

[ZCW05] Yonggui Zhu, Qianshun Chang, and Shengchang Wu. A new algorithm for calculating Adomian polynomials. Applied Mathematics and Computation, 169(1):402–416, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304007829.

Zhang:2006:IIM

[ZCW06] Xian Zhang, Jianfeng Cai, and Yimin Wei. Interval iterative methods for computing Moore–Penrose inverse. Applied Mathematics and Computation, 183(1):522–532, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030600587X.

Zhang:2005:NFR

[ZCZ05] Daoqiang Zhang, Songcan Chen, and Zhi-Hua Zhou. A new face recognition method based on SVD perturbation for single example image per person. Applied Mathematics and Computation, 163(2):895–907, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002681.

Zhang:2007:PEP

[ZCZ07] Hong Zhang, Lansun Chen, and Rongping Zhu. Permanence and extinction of a periodic predator–prey delay system with functional response and stage structure for prey. Applied Mathematics and Computation, 184(2):931–944, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007831.

Zhou:2008:NHR

[ZCZ08] Jun Zhou, Li Cai, and Feng-Qi Zhou. New highresolution scheme for threedimensional nonlinear hyperbolic conservation laws. Applied Mathematics and Computation, 198(2):770– 786, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009526.

Zhu:2007:GBQ

[ZCZX07] Zhengyu Zhu, Xinghuan Chen, Qingsheng Zhu, and Qihong Xie. A GA-based query optimization method for web information retrieval. Applied Mathematics and Computation, 185(2):919–930, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008095.

Zhang:2006:GPS

[ZD06a] Su-Ying Zhang and Zi-Chen Deng. Group preserving schemes for nonlinear dynamic system based on RKMK methods. Applied Mathematics and Computation, 175(1):497–507, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006247.

Zhao:2006:DAS

[ZD06b] Hongyong Zhao and Nan Ding. Dynamic analysis of stochastic Cohen– Grossberg neural networks with time delays. Applied Mathematics and Computation, 183(1):464–470, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005790.

Zhang:2007:PRK

[ZDL07] Suying Zhang, Zichen Deng, and Wencheng Li. A precise Runge–Kutta integration and its application for solving nonlinear dynamical systems. Applied Mathematics and Computation, 184(2):496–502, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007259.

Zhang:2008:CNT

[ZDY08] Li-Hua Zhang, Li-Hua Dong, and Li-Mei Yan. Construction of non-travelling wave solutions for the generalized variable-coefficient Gardner equation. Applied Mathematics and Computation, 203(2):784–791, September 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308004049.

Zeng:2009:DNA

[Zen09]  Zhijun Zeng.  Dynam-

ics of a non-autonomous ratio-dependent food chain model. Applied Mathematics and Computation, 215(3):1274–1287, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006420.

Zhang:2005:CSS

[ZF05] Zhenfeng Zhang and Dengguo Feng. Cryptanalysis of some signature schemes with message recovery. Applied Mathematics and Computation, 170(1):103–114, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008744.

Zhou:2009:PON

[ZFL^{+}09] Jianzhong Zhou, Rengcun Fang, Yuhong Li, Yongchuan Zhang, and Bing Peng. Parameter optimization of nonlinear grey Bernoulli model using particle swarm optimization. Applied Mathematics and Computation, 207(2):292–299, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308007911.

Zamirian:2007:AMS

[ZFN07] M. Zamirian, M. H. Farahi, and A. R. Nazemi. An applicable method for solving the shortest path problems. Applied Mathematics and Computation, 190(2):1479–1486, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001816.

Zhou:2006:SAM

[ZFT06] P. Zhou, L. W. Fan, and H. W. Tang. On stability analysis of multiple objective dynamic input–output model. Applied Mathematics and Computation, 177(1):79– 84, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500888X.

Zhang:2007:OHO

[ZFY07] Chaolong Zhang, Weizhen Feng, and Fengjian Yang. Oscillations of higher order nonlinear functional differential equations with impulses. Applied Mathematics and Computation, 190 (1):370–381, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307000586.

Zhang:2008:OSO

[ZFYH08] Chaolong Zhang, Weizhen Feng, Jianfu Yang, and Mugen Huang. Oscillations of second order impulses nonlinear FDE with forcing term. Applied Mathematics and Computation, 198(1):271–279, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008879.

Zhang:2008:MSC

[ZG08a] Haomin Zhang and Siqing Gan. Mean square convergence of one-step methods for neutral stochastic differential delay equations. Applied Mathematics and Computation, 204(2):884–890, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005754.

Ziani:2008:ALM

[ZG08b] M. Ziani and F. Guyomarc’h. An autoadaptative limited memory Broyden’s method to solve systems of nonlinear equations. Applied Mathematics and Computation, 205(1):202–211, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004839.

Zayed:2009:SAE

[ZG09] E. M. E. Zayed and Khaled A. Gepreel. Some applications of the (\frac{G′}{G})- expansion method to nonlinear partial differential equations. Applied Mathematics and Computation, 212(1):1–13, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001301.

Zandieh:2006:IAA

[ZGH06] M. Zandieh, S. M. T. Fatemi Ghomi, and S. M. Moattar Husseini. An immune algorithm approach to hybrid flow shops scheduling with sequence-dependent setup times. Applied Mathematics and Computation, 180(1):111–127, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305011021.

Zhao:2007:EPB

[ZGH07] Xinchao Zhao, Xiao-Shan Gao, and Ze-Chun Hu. Evolutionary programming based on non-uniform mutation. Applied Mathematics and Computation, 192(1):1–11, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008654.

Zare:2006:DHA

[ZGK06] H. Khademi Zare, S. M.

T. Fatemi Ghomi, and B. Karimi. Developing a heuristic algorithm for order production planning using network models under uncertainty conditions. Applied Mathematics and Computation, 182(2):1208–1218, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004358.

Zheng:2008:FCM

[ZGL08] Sining Zheng, Haojie Guo, and Jing Liu. A food chain model for two resources in un-stirred chemostat. Applied Mathematics and Computation, 206(1):389–402, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308006887.

Zhou:2005:UDF

[ZH05] Xiaolin Zhou and Yinnian He. Using divergence free wavelets for the numerical solution of the 2-D stationary Navier– Stokes equations. Applied Mathematics and Computation, 163(2):593–607, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002450.

Zhang:2007:CNS

[ZH07a] Qimin Zhang and Chongzhao Han. Convergence of numerical solutions to stochastic age-structured population system with diffusion. Applied Mathematics and Computation, 186(2):1234–1242, March 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306010484.

Zhu:2007:SBT

[ZH07b] Huiyan Zhu and Lihong Huang. Stability and bifurcation in a tri-neuron network model with discrete and distributed delays. Applied Mathematics and Computation, 188(2):1742–1756, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015414.

Zhao:2008:IGP

[ZH08a] Xi-Le Zhao and Ting-Zhu Huang. On the inverse of a general pentadiagonal matrix. Applied Mathematics and Computation, 202(2):639–646, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001525.

Zhou:2008:GEP

[ZH08b] Liqun Zhou and Guangda Hu. Global exponential periodicity and stability of cellular neural networks with variable and distributed delays. Applied Mathematics and Computation, 195(2):402–411, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307005619.

Zhou:2008:CWR

[ZH08c] Sheng Zhou and Ting-Zhu Huang. Convergence of waveform relaxation methods for Hermitian positive definite linear systems. Applied Mathematics and Computation, 203(2):943–952, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004311.

Zou:2008:AOA

[ZH08d] Yun-Zhi Zou and Nan-Jing Huang. H(·, ·)-accretive operator with an application for solving variational inclusions in Banach spaces. Applied Mathematics and Computation, 204(2):809–816, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800564X.

Zhao:2009:NIM

[ZH09a] Jing Zhao and Songnian He. A new iterative method for equilibrium problems and fixed point problems of infinitely nonexpansive mappings and monotone mappings. Applied Mathematics and Computation, 215(2):670–680, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005463.

Zou:2009:NSV

[ZH09b] Yun-Zhi Zou and Nan-Jing Huang. A new system of variational inclusions involving H(·, ·)-accretive operator in Banach spaces. Applied Mathematics and Computation, 212(1):135– 144, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900126X.

Zhang:2005:CCD

[Zha05a]  Luming Zhang.  Con-

vergence of a conservative difference scheme for a class of Klein–Gordon– Schrödinger equations in one space dimension. Applied Mathematics and Computation, 163(1):343– 355, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304002292.

Zhang:2005:FDS

[Zha05b] Luming Zhang. A finite difference scheme for generalized regularized long-wave equation. Applied Mathematics and Computation, 168(2):962–972, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030400654X.

Zhang:2005:FEN

[Zha05c] Zhiyue Zhang. The finite element numerical analysis for a class nonlinear evolution equations. Applied Mathematics and Computation, 166(3):489–500, July 26, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304004539.

Zhao:2005:AMC

[Zha05d] Caidi Zhao. Algorithm and mechanization to Cauchy problem of parabolic equation. Applied Mathematics and Computation, 169(2):1038–1044, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304008331.

Zhao:2005:AFM

[Zha05e] Caidi Zhao. Algorithmization formula and mechanization to Cauchy problem of high dimensional wave equation. Applied Mathematics and Computation, 167(2):957–963, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005417.

Zhang:2006:CMP

[Zha06a] Yunxin Zhang. Convergence of meshless Petrov– Galerkin method using radial basis functions. Applied Mathematics and Computation, 183(1):307– 321, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306005637.

Zhang:2006:SPDb

[Zha06b] Yunxin Zhang. Solve partial differential equations by two or more radial basis functions. Applied Mathematics and Computation, 181(1):793–801, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306001615.

Zhao:2006:AMC

[Zha06c] Caidi Zhao. Algorithm and mechanization to Cauchy problem of wave equation with linear damping and source terms. Applied Mathematics and Computation, 173(1):1–8, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305002973.

Zhang:2007:SCC

[Zha07a]  Aihua Zhang.  A se-

cret to create a complete market from an incomplete market. Applied Mathematics and Computation, 191(1):253–262, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307002524.

Zhang:2007:NTR

[Zha07b] Liping Zhang. A new trust region algorithm for nonsmooth convex minimization. Applied Mathematics and Computation, 193(1):135–142, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003803.

Zhang:2007:GAE

[Zha07c] Sheng Zhang. A generalized auxiliary equation method and its application to the (2 + 1)- dimensional KdV equations. Applied Mathematics and Computation, 188(1): 1–6, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013026.

Zhang:2007:GNA

[Zha07d] Sheng Zhang. A generalized new auxiliary equation method and its application to the (2 + 1)-dimensional breaking soliton equations. Applied Mathematics and Computation, 190(1):510– 516, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307000823.

Zhang:2007:PMS

[Zha07e] Wei-Guo Zhang. Possibilistic mean–standard deviation models to portfolio selection for bounded assets. Applied Mathematics and Computation, 189(2):1614–1623, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017462.

Zhang:2007:RMF

[Zha07f] Yunxin Zhang. Reconstruct multiscale functions using different RBFs in different subdomains. Applied Mathematics and Computation, 189(1):893–901, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016687.

Zhang:2007:SPD

[Zha07g] Yunxin Zhang. Solving partial differential equations by meshless methods using radial basis functions. Applied Mathematics and Computation, 185(1):614–627, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008964.

Zhao:2007:HAC

[Zha07h] Jichao Zhao. Highly accurate compact mixed methods for two point boundary value problems. Applied Mathematics and Computation, 188(2):1402–1418, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014925.

Zhang:2008:SSC

[Zha08a]  Liping Zhang.  Solv-

ability of semidefinite complementarity problems. Applied Mathematics and Computation, 196(1):86–93, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006479.

Zhang:2008:EFMa

[Zha08b] Sheng Zhang. Exp-function method exactly solving a KdV equation with forcing term. Applied Mathematics and Computation, 197(1):128–134, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007618.

Zhang:2008:EFMb

[Zha08c]  Sheng Zhang.  Exp-

function method for constructing explicit and exact solutions of a lattice equation. Applied Mathematics and Computation, 199(1):242–249, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009915.

Zhang:2008:FIE

[Zha08d] Sheng Zhang. A further improved extended Fan sub-equation method for (2+1)-dimensional breaking soliton equations. Applied Mathematics and Computation, 199(1):259– 267, May 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307009939.

Zhang:2008:EPS

[Zha08e]  Xingqiu Zhang.  Exis-

tence of positive solutions for multi-point boundary value problems on infinite intervals in Banach spaces. Applied Mathematics and Computation, 206(2):932–941, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007637.

Zhang:2008:EHO

[Zha08f] Xinguang Zhang. Eigenvalue of higher-order semipositone multi-point boundary value problems with derivatives. Applied Mathematics and Computation, 201(1–2):361–370, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307012131.

Zhang:2008:SRL

[Zha08g]  Xiuyun Zhang.  Some

results of linear fractional order time-delay system. Applied Mathematics and Computation, 197(1):407–411, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007898.

Zhang:2008:AFC

[Zha08h] Yang Zhang. AD–FDSD for convection–diffusion problems. Applied Mathematics and Computation, 206(1):257–271, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308001203.

Zhang:2009:NCA

[Zha09a] Guoqing Zhang. A note on “A Continuous Approach to Nonlinear Integer Programming”. Applied Mathematics and Computation, 215(6):2388–2389, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309007334.

Zhang:2009:IWY

[Zha09b] Li Zhang. An improved Wei–Yao–Liu nonlinear conjugate gradient method for optimization computation. Applied Mathematics and Computation, 215(6):2269–2274, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309007401.

Zhang:2009:TWS

[Zha09c] Li-Hua Zhang. Travelling wave solutions for the generalized Zakharov– Kuznetsov equation with higher-order nonlinear terms. Applied Mathematics and Computation, 208(1):144– 155, February 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308008722.

Zhang:2009:OGM

[Zha09d] Naimin Zhang. An online gradient method with momentum for two-layer feedforward neural networks. Applied Mathematics and Computation, 212(2):488–498, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309001817.

Zhang:2009:FDM

[Zha09e] Yang Zhang. A finite difference method for fractional partial differential equation. Applied Mathematics and Computation, 215(2):524–529, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005062.

Zhao:2009:SNF

[Zha09f] Hong Zhao. Some new fractals behaviors in (2 + 1)-dimensional integrable system. Applied Mathematics and Computation, 215(5):1653–1658, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006651.

Zhao:2009:ESC

[Zha09g] Zengqin Zhao. Exact solutions of a class of secondorder nonlocal boundary value problems and applications. Applied Mathematics and Computation, 215(5):1926–1936, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309006948.

Zhao:2009:EFP

[Zha09h]  Zengqin Zhao.  Exis-

tence of fixed points for some convex operators and applications to multipoint boundary value problems. Applied Mathematics and Computation, 215(8):2971–2977, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309008698.

Zheng:2008:CBC

[Zhe08] Z. C. Zheng. A consistent boundary condition for vorticity–streamfunction simulation of wall-bounded vortex flow. Applied Mathematics and Computation, 206(1):205–213, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006656.

Zhang:2008:CRM

[ZHGG08] Li-Tao Zhang, Ting-Zhu Huang, Tong-Xiang Gu, and Xin-Lan Guo. Convergence of relaxed multisplitting USAOR methods for H-matrices linear systems. Applied Mathematics and Computation, 202(1):121–132, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000659.

Zhi:2009:SRL

[Zhi09] Hongyan Zhi. Symmetry reductions of the Lax pair for the 2 + 1-dimensional Konopelchenko–Dubrovsky equation. Applied Mathematics and Computation, 210(2):530–535, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000344.

Zhang:2007:GTP

[ZHL07] Yong Zhang, Ting-Zhu Huang, and Xing-Ping Liu. Gauss type preconditioning techniques for linear systems. Applied Mathematics and Computation, 188(1):612–633, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013993.

Zhang:2007:CPB

[ZHLG07] Yong Zhang, Ting-Zhu Huang, Xing-Ping Liu, and Tong-Xiang Gu. A class of preconditioners based on the (I +S(α))-type preconditioning matrices for solving linear systems. Applied Mathematics and Computation, 189(2):1737–1748, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017656.

Zhou:2005:PAD

[Zho05] Wen-Hui Zhou. Performance analysis of discretetime queue GI / G / 1 with negative arrivals. Applied Mathematics and Computation, 170(2):1349–1355, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000706.

Zhou:2006:EJS

[Zho06]  Xinlong Zhou.  Esti-

mates for the joint spectral radius. Applied Mathematics and Computation, 172(1):332–348, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002067.

Zhong:2007:EAT

[Zho07] Sheng Zhong. An efficient algorithm for topology discovery of a blackbox communication network. Applied Mathematics and Computation, 186(1): 506–508, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030600960X.

Zhong:2008:EAA

[Zho08a] Sheng Zhong. An efficient approximation algorithm for counting n-cycles in a graph. Applied Mathematics and Computation, 196(1):479–482, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006662.

Zhou:2008:EPS

[Zho08b] Xiaoliang Zhou. Eventually positive solutions of higher order nonlinear neutral differential equations. Applied Mathematics and Computation, 201(1–2):859–863, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000118.

Zhou:2009:DAL

[Zho09a]  Guangming Zhou.  A

descent algorithm without line search for unconstrained optimization. Applied Mathematics and Computation, 215(7):2528– 2533, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309007899.

Zhou:2009:NSH

[Zho09b] Xiaojian Zhou. A note on “Some higher-order modifications of Newton’s method for solving nonlinear equations”. Applied Mathematics and Computation, 208(1):14–17, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008278.

Zhou:2009:ENZ

[Zho09c] Xiaoliang Zhou. The estimate for number of zeros of solutions of second order functional differential equations. Applied Mathematics and Computation, 215(3):1266–1273, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006419.

Zhu:2005:ASP

[Zhu05a] Detong Zhu. An affine scaling projective reduced Hessian algorithm for minimum optimization with nonlinear equality and linear inequality constraints. Applied Mathematics and Computation, 166(1):131– 163, July 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004278.

Zhu:2005:NBI

[Zhu05b] Detong Zhu. Nonmonotone backtracking inexact quasi-Newton algorithms for solving smooth nonlinear equations. Applied Mathematics and Computation, 161(3):875–895, February 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304000384.

Zhu:2005:CCD

[Zhu05c] Ling Zhu. On the convergent condition of Durand– Kerner method in parallel circular iteration of multi-step. Applied Mathematics and Computation, 169(1):179–191, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304007684.

Zhu:2005:CCN

[Zhu05d] Ling Zhu. On the convergent condition of Newtonlike method in parallel circular iteration for simultaneously finding all multiple zeros of a polynomial, II. Applied Mathematics and Computation, 168(1):677–685, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304006319.

Zhu:2005:CFF

[Zhu05e] Wenxing Zhu. A class of filled functions for box constrained continuous global optimization. Applied Mathematics and Computation, 169(1):129–145, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007568.

Zhu:2005:FIP

[Zhu05f]  Zhibin Zhu.  A feasi-

ble interior-point algorithm for nonconvex nonlinear programming. Applied Mathematics and Computation, 163(2):745–753, April 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002577.

Zhu:2005:GSC

[Zhu05g]  Zhibin Zhu.  A glob-

ally and superlinearly convergent feasible QP-free method for nonlinear programming. Applied Mathematics and Computation, 168(1):519–539, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304006216.

Zhu:2006:ASI

[Zhu06a] Detong Zhu. An affine scaling interior trust-region method for LC ^{1} minimization subject to bounds on variables. Applied Mathematics and Computation, 172(2):1272–1302, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305002924.

Zhu:2006:USO

[Zhu06b] Wenxing Zhu. Unsolvability of some optimization problems. Applied Mathematics and Computation, 174(2):921–926, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305005217.

Zhu:2006:SFS

[Zhu06c] Zhibin Zhu. A simple feasible SQP algorithm for inequality constrained optimization. Applied Mathematics and Computation, 182(2):987–998, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306004139.

Zhu:2007:ASR

[Zhu07] Detong Zhu. An affine scaling reduced preconditional conjugate gradient path method for linear constrained optimization. Applied Mathematics and Computation, 184(2):181–198, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006941.

Zhu:2008:GES

[Zhu08]  Wei Zhu.  Global ex-

ponential stability of impulsive reaction–diffusion equation with variable delays. Applied Mathematics and Computation, 205(1):362–369, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005948.

Zhu:2009:AST

[Zhu09a]  Congxu Zhu.  Adap-

tive synchronization of two novel different hyperchaotic systems with partly uncertain parameters. Applied Mathematics and Computation, 215(2):557–561, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005116.

Zhu:2009:ITO

[Zhu09b] Xiangling Zhu. Integraltype operators from iterated logarithmic Bloch spaces to Zygmund-type spaces. Applied Mathematics and Computation, 215(3):1170–1175, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006304.

Huang:2007:ECE

[zHWH07] Fu zhuo Huang, Ling Wang, and Qie He. An effective co-evolutionary differential evolution for constrained optimization. Applied Mathematics and Computation, 186(1):340– 356, March 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306009441.

Zhao:2006:RNL

[ZHX06] Narisa Zhao, Tang Huanwen, and Luo Xiaona. Research on non-linear input– output model based on production function theory and a new method to update IO coefficients matrix. Applied Mathematics and Computation, 181(1):478–486, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001172.

Zhang:2007:IKR

[ZHZ07a] Jiang-She Zhang, Xing-Fang Huang, and Cheng-Hu Zhou. An improved kernel regression method based on Taylor expansion. Applied Mathematics and Computation, 193(2):419–429, November 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004158.

Zou:2007:NSM

[ZHZ07b] Yongkui Zou, Qingwan Hu, and Ran Zhang. On numerical studies of multipoint boundary value problem and its fold bifurcation. Applied Mathematics and Computation, 185(1):527–537, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008873.

Zhao:2008:LRP

[ZHZ08] Lijun Zhao, Xiyan Hu, and Lei Zhang. Linear restriction problem of Hermitian reflexive matrices and its approximation. Applied Mathematics and Computation, 200(1):341–351, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011228.

Ziegler:2009:PRC

[Zie09] Martin Ziegler. Physicallyrelativized Church–Turing Hypotheses: Physical foundations of computing and complexity theory of computational physics. Applied Mathematics and Computation, 215(4):1431–1447, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309004226.

Zimmermann:2007:NPK

[Zim07] Karel Zimmermann. A note on a paper by E. Khorram and A. Ghodousian. Applied Mathematics and Computation, 188 (1):244–245, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306013440. See

[KG06].

Zu:2008:ETM

[ZJO08] Li Zu, Daqing Jiang, and Donal O’Regan. Existence theory for multiple solutions to semipositone Dirichlet boundary value problems with singular dependent nonlinearities for second-order impulsive differential equations. Applied Mathematics and Computation, 195(1):240–255, January 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005346.

Zhang:2005:CLH

[ZK05] Fangguo Zhang and Kwangjo Kim. Cryptanalysis of Lee– Hwang–Li’s key authentication scheme. Applied Mathematics and Computation, 161(1):101–107, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012128.

See comment [SCS05b].

Zhang:2005:EHM

[ZL05a] Yunong Zhang and W. E.

Leithead. Exploiting Hessian matrix and trustregion algorithm in hyperparameters estimation of Gaussian process. Applied Mathematics and Computation, 171(2):1264–1281, December 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhao:2005:MCE

[ZL05b] Ze-Mao Zhao and Feng-Yu Liu. Method of constructing elliptic curve authenticated encryption scheme. Applied Mathematics and Computation, 168(1):146–151, September 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304005946.

Zhang:2006:CPS

[ZL06a] Qi Zhang and Yirong Liu. A cubic polynomial system with seven limit cycles at infinity. Applied Mathematics and Computation, 177(1):319–329, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009136.

Zheng:2006:FVR

[ZL06b] Yongshuang Zheng and Baoding Liu. Fuzzy vehicle routing model with credibility measure and its hybrid intelligent algorithm. Applied Mathematics and Computation, 176(2):673–683, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008660.

Zhou:2006:DTQ

[ZL06c] Wen-Hui Zhou and Yong-Jun Li. A discrete-time queueing model of EMD policy in high-speed networks. Applied Mathematics and Computation, 181(1):543–551, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600124X.

Zhang:2007:PSFb

[ZL07a] Xinguang Zhang and Lishan Liu. Positive solutions of fourth-order multi-point boundary value problems with bending term. Applied Mathematics and Computation, 194(2):321–332, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004730.

Zhou:2007:BTW

[ZL07b] Yuqian Zhou and Qian Liu.

Bifurcation of travelling wave solutions for a (2+1)- dimensional nonlinear dispersive long wave equation. Applied Mathematics and Computation, 189(1):970– 979, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306016766.

Zhu:2007:SUL

[ZL07c] Shihui Zhu and Xiaoguang Li. Sharp upper and lower bounds on the blow-up rate for nonlinear Schrödinger equation with potential. Applied Mathematics and Computation, 190(2):1267–1272, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001579.

Zhang:2008:PCN

[ZL08a] Liang Zhang and Hong-Xu Li. Periodicity on a class of neutral impulsive delay system. Applied Mathematics and Computation, 203(1):178–185, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002257.

Zhang:2008:NST

[ZL08b] Xinguang Zhang and Lishan Liu. Nontrivial solution of third-order nonlinear eigenvalue problems (II). Applied Mathematics and Computation, 204(1):508–512, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308005092.

Zheng:2008:NSB

[ZL08c] Z. C. Zheng and W. Li.

Numerical stabilities and boundary conditions in time-domain Eulerian simulations of acoustic wave propagations with and without background flow. Applied Mathematics and Computation, 202(1):146– 161, August 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308000684.

Zhou:2008:GDL

[ZL08d] Lujun Zhou and Yaqiong Li. A generalized dynamic IS-LM model with delayed time in investment processes. Applied Mathematics and Computation, 196(2):774–781, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007291.

Zhou:2008:ENZ

[ZL08e] Xin Zhou and Cuiping Li.

Estimate of the number of zeros of Abelian integrals for a kind of quartic Hamiltonians with two centers. Applied Mathematics and Computation, 204(1):202–209, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004955.

Zhou:2009:MTD

[ZL09a] Wuneng Zhou and Minghao Li. Mixed time-delays dependent exponential stability for uncertain stochastic high-order neural networks. Applied Mathematics and Computation, 215(2):503–513, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309004913.

Zhou:2009:SID

[ZL09b] Yongfang Zhou and Yingzhen Lin. Solving integrodifferential equations with Cauchy kernel. Applied Mathematics and Computation, 215(7):2438–2444, December 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630030900770X.

Zhang:2006:BEC

[ZLC06] Qi Zhang, Yirong Liu, and Haibo Chen. Bifurcation at the equator for a class of quintic polynomial differential system. Applied Mathematics and Computation, 181(1):747–755, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030600155X.

Zhou:2007:ECM

[ZLC07] Yongfang Zhou, Yinzhen Lin, and Minggen Cui. An efficient computational method for second order boundary value problems of nonlinear differential equations. Applied Mathematics and Computation, 194(2):354–365, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004766.

Zhang:2006:SLE

[ZLCZ06] Jian Jun Zhang, Wen Bin Liu, Tai Yong Chen, and Hui Xing Zhang. Solvability of p-Laplace equations subject to threepoint boundary value problems. Applied Mathematics and Computation, 179(2):688–695, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305010738.

Zhou:2009:SFM

[ZLDW09] Bin Zhou, Zhao-Yan Li, Guang-Ren Duan, and Yong Wang. Solutions to a family of matrix equations by using the Kronecker matrix polynomials. Applied Mathematics and Computation, 212(2):327–336, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001489.

Zhang:2009:EME

[ZLH^{+}09] Junling Zhang, Changyong Liang, Yongqing Huang, Jian Wu, and Shanlin Yang. An effective multiagent evolutionary algorithm integrating a novel roulette inversion operator for engineering optimization. Applied Mathematics and Computation, 211(2):392–416, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309000873.

Zhang:2006:WMM

[ZLL06a] Xinming Zhang, Ke’an Liu, and Jiaqi Liu. The wavelet multiscale method for inversion of porosity in the fluid-saturated porous media. Applied Mathematics and Computation, 180(2):419–427, September 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000026.

Zhou:2006:EGE

[ZLL06b] Tiejun Zhou, Yirong Liu, and Yuehua Liu. Existence and global exponential stability of periodic solution for discretetime BAM neural networks. Applied Mathematics and Computation, 182(2):1341–1354, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306004486.

Li:2008:EFM

[zLmLL08] Ya zhou Li, Kai ming Li, and Chang Lin. Expfunction method for solving the generalized-Zakharov equations. Applied Mathematics and Computation, 205(1):197–201, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308004384.

Zhang:2005:CMA

[ZLMW05] Juan Zhang, Jie Lou, Zhien Ma, and Jianhong Wu. A compartmental model for the analysis of SARS transmission patterns and outbreak control measures in China. Applied Mathematics and Computation, 162(2):909–924, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304001389.

Zhong:2007:GIP

[ZLR07] Wenyong Zhong, Wei Lin, and Jiong Ruan. The generalized invariance principle for dynamic equations on time scales. Applied Mathematics and Computation, 184(2):557–565, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306007302.

Zhang:2009:DBN

[ZLT09] Tailei Zhang, Junli Liu, and Zhidong Teng. Dynamic behavior for a nonautonomous SIRS epidemic model with distributed delays. Applied Mathematics and Computation, 214(2):624–631, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003622.

Zhang:2006:RIA

[ZLW06a] Bing-Quan Zhang, Xing-Guo Luo, and Qing-Biao Wu. The restrictions and improvement of the Adomian decomposition method. Applied Mathematics and Computation, 177(1):99–104, June 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500891X.

Zhang:2006:AEP

[ZLW06b] Wei-Guo Zhang, Wen-An Liu, and Ying-Luo Wang. On admissible efficient portfolio selection: Models and algorithms. Applied Mathematics and Computation, 176(1):208– 218, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500785X.

Zhang:2006:NST

[ZLW06c] Xinguang Zhang, Lishan Liu, and Congxin Wu. Nontrivial solution of third-order nonlinear eigenvalue problems. Applied Mathematics and Computation, 176(2):714–721, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008714.

See erratum [Yin08b].

Zhang:2007:EPSa

[ZLW07] Xinguang Zhang, Lishan Liu, and Yonghong Wu. Existence of positive solutions for second-order semipositone differential equations on the halfline. Applied Mathematics and Computation, 185(1):628–635, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008976.

Zhang:2008:UPS

[ZLW08] Haijun Zhang, Lishan Liu, and Yonghong Wu. A unique positive solution for n th-order nonlinear impulsive singular integrodifferential equations on unbounded domains in Banach spaces. Applied Mathematics and Computation, 203(2):649–659, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002968.

Zhang:2009:RBR

[ZLW09a] Jian Zhang, Xiaoguang Li, and Yonghong Wu. Remarks on the blowup rate for critical nonlinear Schrödinger equation with harmonic potential. Applied Mathematics and Computation, 208(2):389–396, February 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009077.

Zhang:2009:GSP

[ZLW09b] Qi Zhang, Yirong Liu, and Gui Weihua. Generalized singular point quantity and integrability of degenerate resonant singular point. Applied Mathematics and Computation, 215(4):1507–1512, October 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900650X.

Zhang:2008:HBS

[ZLY08] Jia-Fang Zhang, Wan-Tong Li, and Xiang-Ping Yan. Hopf bifurcation and stability of periodic solutions in a delayed ecoepidemiological system. Applied Mathematics and Computation, 198(2):865– 876, May 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009617.

Zhou:2008:NSL

[ZLYL08] Jiquan Zhou, Xiangli Liu, Xiaozhong Yang, and Yifang Liu. Novel solitonlike and multi-solitary wave solutions of (3 + 1)- dimensional Burgers equation. Applied Mathematics and Computation, 204(1):461–467, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800533X.

Zeng:2006:PSD

[ZLZ06] Zhi-Bin Zeng, Yan Li, and Wenxing Zhu. Partner selection with a due date constraint in virtual enterprises. Applied Mathematics and Computation, 175(2):1353–1365, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500682X.

Zhang:2007:PSFa

[ZLZ07] Xinguang Zhang, Lishan Liu, and Huichao Zou. Positive solutions of fourth-order singular three point eigenvalue problems. Applied Mathematics and Computation, 189(2):1359–1367, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601722X.

Zhou:2008:BTW

[ZLZ08] Yuqian Zhou, Qian Liu, and Weinian Zhang. Bifurcation of travelling wave solutions for the (2 + 1)-dimensional Broer– Kau–Kupershmidt equation. Applied Mathematics and Computation, 204(1):210–215, October 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308004967.

Zhang:2006:RJO

[ZLZG06] J. Zhang, X. Q. Li, X. T.

Zhang, and J. F. Gao. Rationality judge and optimization design of lottery scheme. Applied Mathematics and Computation, 183(2):780–786, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306006126.

Zhou:2006:ESD

[ZLZZ06] Yuqian Zhou, Qian Liu, Jian Zhang, and Weinian Zhang. Exact solution for (2 + 1)-dimension nonlinear dispersive long wave equation. Applied Mathematics and Computation, 177(2):495–499, June 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305009288.

Zhang:2007:SAC

[ZLZZ07] Yuping Zhang, Xinzhi Liu, Hong Zhu, and Shouming Zhong. Stability analysis and control synthesis for a class of switched neutral systems. Applied Mathematics and Computation, 190(2):1258–1266, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307001567.

Zaman:2006:IDI

[ZM06a] F. D. Zaman and Khalid Masood. Investigation of the direct and inverse solutions for torsional waves propagating in a cylinder. Applied Mathematics and Computation, 173(2):1238–1245, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305004273.

Zhao:2006:OSO

[ZM06b] Xueqin Zhao and Fanwei Meng. Oscillation of second-order nonlinear ODE with damping. Applied Mathematics and Computation, 182(2):1861–1871, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005352.

See comment [ÇT07a].

Zhi:2007:QBN

[ZM07] Yuanhong Zhi and Chunlai Mu. The quenching behavior of a nonlinear parabolic equation with nonlinear boundary outflux. Applied Mathematics and Computation, 184(2):624–630, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007363.

Zhang:2008:IIT

[ZM08a] Hongxia Zhang and Fanwei Meng. Integral inequalities in two independent variables for retarded Volterra equations. Applied Mathematics and Computation, 199(1):90– 98, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307009721.

Zhang:2008:CII

[ZM08b] Hongxia Zhang and Fanwei Meng. On certain integral inequalities in two independent variables for retarded equations. Applied Mathematics and Computation, 203(2):608–616, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308002919.

Zhao:2008:OCS

[ZM08c] Junhua Zhao and Fanwei Meng. Oscillation criteria for second-order neutral equations with distributed deviating argument. Applied Mathematics and Computation, 206(1):485–493, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308006929.

Zhi:2008:NSQ

[ZM08d] Yuanhong Zhi and Chunlai Mu. Non-simultaneous quenching in a semilinear parabolic system with weak singularities of logarithmic type. Applied Mathematics and Computation, 196(1):17–23, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006236.

Zhang:2009:ASS

[ZM09] Pingping Zhang and Lufang Mi. Analytic solutions of a second order iterative functional differential equation. Applied Mathematics and Computation, 210(2):277–283, April 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800920X.

Zaman:2006:ISM

[ZMM06] F. D. Zaman, Khalid Masood, and Z. Muhiameed. Inverse scattering in multilayer inverse problem in the presence of damping. Applied Mathematics and Computation, 176(2):455–461, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305008222.

Zunic:2009:LMC

[ŽMO09] Joviša Žunić and Carlos Martinez-Ortiz. Linearity measure for curve segments. Applied Mathematics and Computation, 215(8):3098–3105, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309008960.

Meng:2008:DID

[zMsCxL08] Xin zhu Meng, Lan sun Chen, and Qing xue Li. The dynamics of an impulsive delay predator– prey model with variable coefficients. Applied Mathematics and Computation, 198(1):361–374, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307008971.

Zorzano:2005:NID

[ZMV05] M.-P. Zorzano, A. M. Mancho, and L. Vázquez. Numerical integration of the discrete-ordinate radiative transfer equation in strongly nonhomogeneous media. Applied Mathematics and Computation, 164(1):263– 274, May 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304003947.

Zhang:2005:NTA

[ZMW05] Xinzhen Zhang, Fengming Ma, and Yiju Wang. A Newton-type algorithm for generalized linear complementarity problem over a polyhedral cone. Applied Mathematics and Computation, 169(1):388–401, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304007817.

Zhi:2008:QPN

[ZMY08] Yuanhong Zhi, Chunlai Mu, and Daming Yuan. The quenching phenomenon of a nonlocal semilinear heat equation with a weak singularity. Applied Mathematics and Computation, 201(1–2):701–709, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308000362.

Zhang:2005:AEP

[ZN05] Wei-Guo Zhang and Zan-Kan Nie. On admissible efficient portfolio selection policy. Applied Mathematics and Computation, 169(1):608–623, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304007982.

Zhang:2009:NIK

[ZnCxX09] Yi Zhang, Xiao na Cai, and Hong xian Xu. A note on “The integrable KdV6 equation: Multiple soliton solutions and multiple singular soliton solutions”. Applied Mathematics and Computation, 214(1):1–3, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002835.

Zhang:2005:AFE

[ZO05] Shangyou Zhang and David O. Olagunju. Axisymmetric finite element solution of non-isothermal parallelplate flow. Applied Mathematics and Computation, 171(2):1081–1094, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zotos:2007:INS

[Zot07a] Kostas Zotos. Improving numerical software. Applied Mathematics and Computation, 192(1):247– 251, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307003281.

Zotos:2007:OOD

[Zot07b]  Kostas Zotos.  Object-

oriented design principles in mathematics. Applied Mathematics and Computation, 188(2):1430–1436, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014950.

Zotos:2007:OOP

[Zot07c]  Kostas Zotos.  Object-

oriented programming in Mathematics. Applied Mathematics and Computation, 188(2):1562–1566, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306015086.

Zotos:2007:PCM

[Zot07d] Kostas Zotos. Performance comparison of Maple and Mathematica. Applied Mathematics and Computation, 188(2):1426–1429, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014949.

Zotos:2008:CAS

[Zot08]  Kostas Zotos.  Com-

puter Algebra Systems — New strategies and techniques. Applied Mathematics and Computation, 198(1):123–127, April 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307008727.

Zhang:2008:MSE

[ZOZZ08] Lin Zhang, Jie Ouyang, Xiaohua Zhang, and Wenbin Zhang. On a multiscale element-free Galerkin method for the Stokes problem. Applied Mathematics and Computation, 203(2):745–753, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003986.

Zhang:2006:HOI

[ZPH06] Xin Zhang, Hong Peng, and Guiwu Hu. A high order iteration formula for the simultaneous inclusion of polynomial zeros. Applied Mathematics and Computation, 179(2):545–552, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010593.

Zahibo:2006:ANS

[ZPT^{+}06] Narcisse Zahibo, Efim Pelinovsky, Tatiana Talipova, Andrey Kozelkov, and Andrey Kurkin. Analytical and numerical study of nonlinear effects at tsunami modeling. Applied Mathematics and Computation, 174(2):795–809, March 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305005102.

Peng:2006:NAS

[zPY06] Yan ze Peng and Emmanuel Yomba. New applications of the singular manifold method to the (2 + 1)-dimensional Burgers equations. Applied Mathematics and Computation, 183(1):61–67, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005376.

Zheng:2006:NSB

[ZQ06] Sining Zheng and Lan Qiao. Non-simultaneous blow-up in a reactiondiffusion system. Applied Mathematics and Computation, 180(1):309–317, September 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305011264.

Zheng:2008:NSV

[ZQ08] Sining Zheng and Lan Qiao. Non-simultaneous versus simultaneous blowup for coupled heat equations with positive–negative sources. Applied Mathematics and Computation, 202(1):141–145, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308000672.

Zhang:2009:INS

[ZQJ09] Chengjian Zhang, Tingting Qin, and Jie Jin. An improvement of the numerical stability results for nonlinear neutral delayintegro-differential equations. Applied Mathematics and Computation, 215(2):548–556, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005104.

Zecevic:2005:CLS

[ZŠ05] A. I. Zečević and D. D.

Šiljak. Control of largescale systems in a multiprocessor environment. Applied Mathematics and Computation, 164(2):531– 543, May 16, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300304003352.

Zheng:2006:QRD

[ZS06] Sining Zheng and Han Su. A quasilinear reactiondiffusion system coupled via nonlocal sources. Applied Mathematics and Computation, 180(1):295– 308, September 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305011252.

Zegeye:2007:VAMb

[ZS07a] Habtu Zegeye and Naseer Shahzad. Viscosity approximation methods for a common fixed point of finite family of nonexpansive mappings. Applied Mathematics and Computation, 191(1):155–163, August 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002391.

Zhang:2007:EES

[ZS07b] Lijuan Zhang and Ligeng Si. Existence and exponential stability of almost periodic solution for BAM neural networks with variable coefficients and delays. Applied Mathematics and Computation, 194(1):215–223, December 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307004602.

Zhang:2007:ECM

[ZS07c] Lisha Zhang and Zhengxing Sun. An experimental comparison of machine learning for adaptive sketch recognition. Applied Mathematics and Computation, 185(2):1138–1148, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008381.

Ziaee:2007:MBI

[ZS07d] Mohsen Ziaee and S. J.

Sadjadi. Mixed binary integer programming formulations for the flow shop scheduling problems. A case study: ISD projects scheduling. Applied Mathematics and Computation, 185(1):218–228, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008538.

Zhang:2008:ESO

[ZS08a] Chao Zhang and Yuming Shi. Eigenvalues of secondorder symmetric equations on time scales with periodic and antiperiodic boundary conditions. Applied Mathematics and Computation, 203(1):284–296, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002531.

Zhao:2008:SSC

[ZS08b] Huailin Zhao and Masanori Sugisaka. Simulation study of CMAC control for the robot joint actuated by McKibben muscles. Applied Mathematics and Computation, 203(1):457–462, September 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308002841.

Zhang:2009:MPS

[ZS09a] Xiaoping Zhang and Yongping Sun. Monotone positive solutions of secondorder multi-point boundary value problems. Applied Mathematics and Computation, 207(2):493– 500, January 15, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308008369.

Zheng:2009:AEF

[ZS09b] Zhi Zheng and Wen-Rui Shan. Application of Exp-function method to the Whitham–Broer–Kaup shallow water model using symbolic computation. Applied Mathematics and Computation, 215(6):2390–2396, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007620.

Zhang:2009:GEC

[ZSCW09] Yunong Zhang, Yanyan Shi, Ke Chen, and Chaoli Wang. Global exponential convergence and stability of gradient-based neural network for online matrix inversion. Applied Mathematics and Computation, 215(3):1301–1306, October 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006213.

Zerarka:2006:EBS

[ZSHB06] A. Zerarka, H. Saidi, S. Hassouni, and N. Bensalah. Evaluation of the bound states of a quantum system via the differential quadrature method: Extended to coupled differential equations. Applied Mathematics and Computation, 182(1):665–671, November 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306003754.

Zhou:2005:PSF

[ZSL05] Jin Zhou, Shu Sun, and Zengrong Liu. Periodic solutions of forced Liénardtype equations. Applied Mathematics and Computation, 161(2):655–666, February 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300304000189.

Zegeye:2007:VAMa

[ZSM07] Habtu Zegeye, Naseer Shahzad, and Tefera Mekonen. Viscosity approximation methods for pseudocontractive mappings in Banach spaces. Applied Mathematics and Computation, 185(1):538–546, February 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306008885.

Zhang:2005:AHA

[ZSNS05] Fangguo Zhang, Reihaneh Safavi-Naini, and Willy Susilo. Attack on Han et al.’s ID-based confirmer (undeniable) signature at ACM-EC’03. Applied Mathematics and Computation, 170(2):1166–1169, November 15, 2005. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630030500055X.

Zhou:2008:ANP

[ZSS08a] Xueyong Zhou, Xiangyun Shi, and Xinyu Song. Analysis of nonautonomous predator-prey model with nonlinear diffusion and time delay. Applied Mathematics and Computation, 196(1):129–136, February 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307006510.

Zhou:2008:ASH

[ZSS08b] Xueyong Zhou, Xinyu Song, and Xiangyun Shi. Analysis of stability and Hopf bifurcation for an HIV infection model with time delay. Applied Mathematics and Computation, 199(1):23–38, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009678.

Zhou:2009:DBV

[ZSZS09] Xueyong Zhou, Xiangyun Shi, Zhonghua Zhang, and Xinyu Song. Dynamical behavior of a virus dynamics model with CTL immune response. Applied Mathematics and Computation, 213(2):329–347, July 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002355.

Zhang:2005:SPD

[ZT05] Yun-Xin Zhang and Yong-Ji Tan. Solving partial differential equations by BKM combined with DDM. Applied Mathematics and Computation, 171(2):1004–1015, December 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:2006:SPDa

[ZT06] Yun-Xin Zhang and Yong-Ji Tan. Solve partial differential equations by meshless subdomains method combined with RBFs. Applied Mathematics and Computation, 174(1):700– 709, March 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300305004960.

Zhang:2007:PCP

[ZT07a] Long Zhang and Zhidong Teng. Permanence for a class of periodic time-dependent competitive system with delays and dispersal in a patchyenvironment. Applied Mathematics and Computation, 188(1):855–864, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014251.

Zhang:2007:GWT

[ZT07b] Xinhong Zhang and Dengke Tong. A generalized Weber transform and its inverse formula. Applied Mathematics and Computation, 193(1):116–126, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003785.

Zhang:2009:WDF

[ZTWL09] Li Zhang, Jieqing Tan, Hongyi Wu, and Zhi Liu. The weighted dual functions for Wang–Bézier type generalized Ball bases and their applications. Applied Mathematics and Computation, 215(1):22– 36, September 1, 2009. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300309003506.

Zuo:2009:SPS

[Zuo09] Jin-Ming Zuo. Solitons and periodic solutions for the Rosenau–KdV and Rosenau–Kawahara equations. Applied Mathematics and Computation, 215(2):835–840, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309005736.

Zhang:2005:ISS

[ZW05a] Jianhong Zhang and Yumin Wang. An improved signature scheme without using one-way Hash functions. Applied Mathematics and Computation, 170(2):905–908, November 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305000329.

Zhang:2005:SCA

[ZW05b] Jianhong Zhang and Yumin Wang. On the security of a convertible authenticated encryption. Applied Mathematics and Computation, 169(2):1063–1069, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008367.

Zhang:2005:SMA

[ZW05c] Jianhong Zhang and Yumin Wang. On the security of Miyaji et al. group signature scheme. Applied Mathematics and Computation, 167(2):1373–1382, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005739.

Zhang:2005:TPS

[ZW05d] Jianhong Zhang and Yumin Wang. Two provably secure k-out-of-n oblivious transfer schemes. Applied Mathematics and Computation, 169(2):1211–1220, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400846X.

Zhang:2005:NSE

[ZW05e] Naimin Zhang and Yimin Wei. A note on solving EP inconsistent linear systems. Applied Mathematics and Computation, 169(1):8–15, October 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400743X.

Zhao:2005:OVU

[ZW05f] Shan Zhao and G. W.

Wei. Option valuation by using discrete singular convolution. Applied Mathematics and Computation, 167(1):383–418, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304005016.

Zheng:2005:LAC

[ZW05g] Bing Zheng and Guorong Wang. Leverrier’s algorithm and Cayley–Hamilton theorem for 2-D system. Applied Mathematics and Computation, 160(3):725–738, January 27, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300303011913.

Zheng:2005:SSO

[ZW05h] Bing Zheng and Ke Wang. Symmetric successive overrelaxation methods for solving the rank deficient linear least squares problem. Applied Mathematics and Computation, 169(2):1305–1323, October 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300304008537.

Zhou:2005:DDM

[ZW05i] Shuzi Zhou and Dehua Wang. Domain decomposition method for a parabolic variational inequality. Applied Mathematics and Computation, 166(1):213– 223, July 6, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304004321.

Zhuang:2005:SCT

[ZW05j] Rong-Kun Zhuang and Hong-Wu Wu. Sturm comparison theorem of solution for second order nonlinear differential equations. Applied Mathematics and Computation, 162(3):1227–1235, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001833.

Zhao:2006:SBD

[ZW06a] Hongyong Zhao and Lei Wang. Stability and bifurcation for discrete-time Cohen–Grossberg neural network. Applied Mathematics and Computation, 179(2):787–798, August 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305010908.

Zheng:2006:CSS

[ZW06b] Bing Zheng and Ke Wang.

Comments on “Symmetric successive overrelaxation methods for rank deficient linear systems”:

[Applied Mathematics and Computation 173 (2006) 404–420]. Applied Mathematics and Computation, 182(2):1709–1713, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005194.

See [DKA06b].

Zheng:2006:GFL

[ZW06c] Bing Zheng and Ke Wang. General fuzzy linear systems. Applied Mathematics and Computation, 181(2):1276–1286, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002475.

Zheng:2006:AOM

[ZW06d] Bing Zheng and Ke Wang. On accelerate overrelaxation methods for rank deficient linear systems. Applied Mathematics and Computation, 173(2):951– 959, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300305004042.

Zhou:2006:ARD

[ZW06e] Jieyong Zhou and Yimin Wei. The analysis of restart DGMRES for solving singular linear systems. Applied Mathematics and Computation, 176(1):293– 301, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305007939.

Zhou:2006:TSA

[ZW06f] Jieyong Zhou and Yimin Wei. A two-step algorithm for solving singular linear systems with index one. Applied Mathematics and Computation, 175(1):472– 485, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006223.

Zhang:2007:MWS

[ZW07a] Guosheng Zhang and Yifu Wang. Minimal wave speed in a diffusive epidemic model with temporal delay. Applied Mathematics and Computation, 188(1):275– 280, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013506.

Zhang:2007:EPSb

[ZW07b] Mingchuan Zhang and Zhongli Wei. Existence of positive solutions for fourth-order mpoint boundary value problem with variable parameters. Applied Mathematics and Computation, 190(2):1417–1431, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001737.

Zhou:2007:BIM

[ZW07c] Jinhua Zhou and Guorong Wang. Block idempotent matrices and generalized Schur complement. Applied Mathematics and Computation, 188(1):246– 256, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013452.

Zhang:2008:POU

[ZW08a] Kai Zhang and Song Wang. Pricing options under jump diffusion processes with fitted finite volume method. Applied Mathematics and Computation, 201(1–2):398–413, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307012179.

Zhao:2008:CST

[ZW08b] Na Zhao and Wei Wu. Convergence of a smoothingtype algorithm for the monotone affine variational inequality problem. Applied Mathematics and Computation, 202(2):820–827, August 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308001860.

Zhao:2008:NKT

[ZW08c] Yueqing Zhao and Qingbiao Wu. Newton– Kantorovich theorem for a family of modified Halley’s method under Hölder continuity conditions in Banach space. Applied Mathematics and Computation, 202(1):243–251, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000921.

Zhou:2008:ESS

[ZW08d] Qinghua Zhou and Li Wan.

Exponential stability of stochastic delayed Hopfield neural networks. Applied Mathematics and Computation, 199(1):84– 89, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030700971X.

Zeng:2009:GER

[ZW09a] Xin Zeng and Deng-Shan Wang. A generalized extended rational expansion method and its application to (1 + 1)- dimensional dispersive long wave equation. Applied Mathematics and Computation, 212(2):296–304, June 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001441.

Zhan:2009:EOC

[ZW09b] Zaidong Zhan and W. Wei. On existence of optimal control governed by a class of the first-order linear dynamic systems on time scales. Applied Mathematics and Computation, 215(6):2070–2081, November 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309007164.

Zhu:2009:NSB

[ZW09c] Chun-Gang Zhu and Ren-Hong Wang. Numerical solution of Burgers’ equation by cubic B-spline quasiinterpolation. Applied Mathematics and Computation, 208(1):260–272, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008989.

Zhou:2006:SCT

[ZWC06] Haiyun Zhou, Li Wei, and Yeol Je Cho. Strong convergence theorems on an iterative method for a family of finite nonexpansive mappings in reflexive Banach spaces. Applied Mathematics and Computation, 173(1):196–212, February 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305003206.

Zhu:2009:NSF

[ZWC09] Shengfeng Zhu, Qingbiao Wu, and Xiaoliang Cheng. Numerical solution of the Falkner–Skan equation based on quasilinearization. Applied Mathematics and Computation, 215(7):2472–2485, December 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309007838.

Zhao:2007:DGG

[ZWD07] Xuan Zhao, Shengjin Wang, and Xiaoqing Ding. A dynamic generating graphical model for pointsets matching. Applied Mathematics and Computation, 185(2):1109–1119, February 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306008356.

Zhang:2009:ISN

[ZWG09] Liping Zhang, Soon-Yi Wu, and Tingran Gao. Improved smoothing Newton methods for P_{0} nonlinear complementarity problems. Applied Mathematics and Computation, 215(1):324–332, September 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030900472X.

Zhang:2006:ETS

[ZWL06] Bing-Quan Zhang, Qing-Biao Wu, and Xing-Guo Luo. Experimentation with two-step Adomian decomposition method to solve evolution models. Applied Mathematics and Computation, 175(2):1495–1502, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006922.

Zhou:2008:ACS

[ZWLX08] Xiaobing Zhou, Yue Wu, Yi. Li, and Hongquan Xue. Adaptive control and synchronization of a novel hyperchaotic system with uncertain parameters. Applied Mathematics and Computation, 203(1):80– 85, September 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308002087.

Zhu:2008:IMM

[ZWM08] Rongbo Zhu, Jiangqing Wang, and Maode Ma. Intelligent MAC model for traffic scheduling in IEEE 802.11e wireless LANs. Applied Mathematics and Computation, 205(1):109–122, November 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003226.

Zhang:2008:GVD

[ZWS08] Zhi-Hua Zhang, Yu-Dong Wu, and H. M. Srivastava. Generalized Vandermonde determinants and mean values. Applied Mathematics and Computation, 202(1):300–310, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308001136.

Zhang:2009:DBW

[ZWT09a] Li Zhang, Hongyi Wu, and Jieqing Tan. Dual bases for Wang–Bézier basis and their applications. Applied Mathematics and Computation, 214(1):218–227, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309002963.

Zhang:2009:DBF

[ZWT09b] Li Zhang, Hongyi Wu, and Jieqing Tan. Dual basis functions for the NSpower basis and their applications. Applied Mathematics and Computation, 207(2):434–441, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308008308.

Zhang:2009:GEM

[ZWT09c] Sheng Zhang, Wei Wang, and Jing-Lin Tong. A generalized (\frac{G′}{G})-expansion method and its application to the (2 + 1)- dimensional Broer–Kaup equations. Applied Mathematics and Computation, 209(2):399–404, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009776.

Zhang:2005:GES

[ZWX05] Qiang Zhang, Xiaopeng Wei, and Jin Xu. On global exponential stability of delayed cellular neural networks with timevarying delays. Applied Mathematics and Computation, 162(2):679–686, March 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304001237.

Zhao:2008:BTW

[ZWX08] Xiaoshan Zhao, Mingchun Wang, and Wei Xu. Bifurcations of traveling wave solutions for a class of the nonlinear equations. Applied Mathematics and Computation, 197(1):228–242, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307007722.

Wei:2007:MPO

[zWxY07] Shu zhi Wei and Zhong xing Ye. Multi-period optimization portfolio with bankruptcy control in stochastic market. Applied Mathematics and Computation, 186(1):414–425, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306009507.

Zhao:2007:DTD

[ZWY07] Zhengrong Zhao, Wei Wang, and Bin Yang. Delay and its time-derivative dependent robust stability of neutral control system. Applied Mathematics and Computation, 187(2):1326–1332, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306012719.

Zhao:2005:OHP

[ZWZ05] Chun Zhao, Miansen Wang, and Ping Zhao. Optimal harvesting problems for age-dependent interacting species with diffusion. Applied Mathematics and Computation, 163(1):117– 129, April 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304002140.

Zheng:2009:DNI

[ZWZ09] Kelong Zheng, Yu Wu, and Shouming Zhong. Discrete nonlinear integral inequalities in two variables and their applications. Applied Mathematics and Computation, 207(1):140–147, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030800773X.

Zhu:2006:AFT

[ZWZL06] Xuegui Zhu, Yi Wang, Jianming Zan, and Chuanbing Li. Application of fractal theory in generation and refinement of finite element mesh. Applied Mathematics and Computation, 175(2):1039–1045, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305006600.

Zhou:2007:CAD

[ZWZZ07] Limei Zhou, Yue Wu, Liwei Zhang, and Guang Zhang. Convergence analysis of a differential equation approach for solving nonlinear programming problems. Applied Mathematics and Computation, 184(2):789–797, January 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306007582.

Zheng:2009:SLM

[ZWZZ09] Quan Zheng, Jing Wang, Peng Zhao, and Li Zhang. A Steffensen-like method and its higher-order variants. Applied Mathematics and Computation, 214(1):10–16, August 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309002720.

Zhang:2006:FIE

[ZX06a] Sheng Zhang and TieCheng Xia. Further improved extended Fan sub-equation method and new exact solutions of the (2 + 1)-dimensional Broer– Kaup–Kupershmidt equations. Applied Mathematics and Computation, 182(2):1651–1660, November 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306005133.

Zhang:2006:GEM

[ZX06b] Sheng Zhang and TieCheng Xia. A generalized F - expansion method and new exact solutions of Konopelchenko–Dubrovsky equations. Applied Mathematics and Computation, 183(2):1190–1200, December 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306006631.

Zhang:2006:SCN

[ZX06c] Sheng Zhang and Tiecheng Xia. Symbolic computation and new families of exact non-travelling wave solutions to (3 + 1)- dimensional Kadomstev– Petviashvili equation. Applied Mathematics and Computation, 181(1):319– 331, October 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306001007.

Zhang:2007:GEM

[ZX07a] Sheng Zhang and Tiecheng Xia. A generalized F - expansion method with symbolic computation exactly solving Broer–Kaup equations. Applied Mathematics and Computation, 189(1):836–843, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300306016626.

Zhao:2007:TWS

[ZX07b] Xiaoshan Zhao and Wei Xu. Travelling wave solutions for a class of the generalized Benjamin– Bona–Mahoney equations. Applied Mathematics and Computation, 192(2):507– 519, September 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300307003542.

Zhou:2008:TLC

[ZX08] Hongxian Zhou and Wei Xu. Thirteen limit cycles for a class of cubic Hamiltonian system with higher-order perturbed terms. Applied Mathematics and Computation, 204(2):905–913, October 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308001446.

Zhao:2005:SRM

[ZXDL05] J. J. Zhao, Y. Xu, S. Y. Dong, and M. Z. Liu. Stability of the Rosenbrock methods for the neutral delay differential-algebraic equations. Applied Mathematics and Computation, 168(2):1128–1144, September 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030400668X.

Zhang:2007:SHB

[ZXF07] Ying Zhang, Wei Xu, and Tong Fang. Stochastic Hopf bifurcation and chaos of stochastic Bonhoeffer– van der Pol system via Chebyshev polynomial approximation. Applied Mathematics and Computation, 190(2):1225–1236, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307001531.

Zhao:2005:SAN

[ZXL05] J. J. Zhao, Y. Xu, and M. Z. Liu. Stability analysis of numerical methods for linear neutral Volterra delay-integro-differential system. Applied Mathematics and Computation, 167(2):1062–1079, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005491.

Zhu:2006:EFT

[ZXL06] Qiang Zhu, Jun-Ming Xu, and Min Lv. Edge fault tolerance analysis of a class of interconnection networks. Applied Mathematics and Computation, 172(1):111–121, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300305001918.

Zheng:2009:PDA

[ZXL09] Xuedong Zheng, Jin Xu, and Wu Li. Parallel DNA arithmetic operation based on n-moduli set. Applied Mathematics and Computation, 212(1):177–184, June 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300309001325.

Zhao:2006:BTW

[ZXLS06] Xiaoshan Zhao, Wei Xu, Shuang Li, and Jianwei Shen. Bifurcations of traveling wave solutions for a class of the generalized Benjamin–Bona– Mahony equation. Applied Mathematics and Computation, 175(2):1760–1774, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030500737X.

Zhou:2007:NLC

[ZXLZ07] Hongxian Zhou, Wei Xu, Shuang Li, and Ying Zhang. On the number of limit cycles of a cubic polynomials Hamiltonian system under quintic perturbation. Applied Mathematics and Computation, 190(1):490–499, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030700080X.

Zhao:2006:SCR

[ZXWL06] J. J. Zhao, Y. Xu, H. X.

Wang, and M. Z. Liu. Stability of a class of Runge– Kutta methods for a family of pantograph equations of neutral type. Applied Mathematics and Computation, 181(2):1170–1181, October 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306002025.

Zhu:2006:GES

[ZXY06] Wei Zhu, Daoyi Xu, and Zhichun Yang. Global exponential stability of impulsive delay difference equation. Applied Mathematics and Computation, 181(1):65–72, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306000671.

Zhou:2007:DFM

[ZXZL07] Hongxian Zhou, Wei Xu, Xiaoshan Zhao, and Shuang Li. Detection function method and its application to a class of quintic Hamiltonian systems with quintic perturbations. Applied Mathematics and Computation, 191(2):490–503, August 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307002779.

Zhou:2007:BLC

[ZXZZ07] Hongxian Zhou, Wei Xu, Xiaoshan Zhao, and Bingchang Zhou. Bifurcations of limit cycles for a perturbed quintic Hamiltonian system with four infinite singular points. Applied Mathematics and Computation, 187(2):686–700, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306011246.

Zhou:2007:PSH

[ZY07] Zhihu Zhou and Zhihui Yang. Periodic solutions in higher-dimensional Lotka– Volterra neutral competition systems with statedependent delays. Applied Mathematics and Computation, 189(1):986–995, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030601678X.

Zhang:2008:ESS

[ZY08a] Dongmei Zhang and Li Yu.

Equivalence of some stability criteria for linear timedelay systems. Applied Mathematics and Computation, 202(1):395–400, August 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030701185X.

Zhang:2008:GEB

[ZY08b] Rui Zhang and Zuodong Yang. Global existence and blow-up solutions and blow-up estimates for a non-local quasilinear degenerate parabolic system. Applied Mathematics and Computation, 200(1):267–282, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030701106X.

Zhao:2008:WCC

[ZY08c] Jinling Zhao and Qingzhi Yang. Weak co-coercivity and its applications in several algorithms for solving variational inequalities. Applied Mathematics and Computation, 201(1–2):200–209, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011861.

Zhou:2008:ATW

[ZY08d] Ming-Chun Zhou and Yan-Fang Yang. An application of traveling wave analysis in economic growth model. Applied Mathematics and Computation, 200(1):261–266, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307011058.

Zhao:2009:ESE

[ZY09] Cunlu Zhao and Chun Yang. Exact solutions for electro-osmotic flow of viscoelastic fluids in rectangular micro-channels. Applied Mathematics and Computation, 211(2):502–509, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630030900109X.

Zhao:2009:BNP

[ZYC09] Zhong Zhao, Li Yang, and Lansun Chen. Bifurcation of nontrivial periodic solutions for a biochemical model with impulsive perturbations. Applied Mathematics and Computation, 215(8):2806–2814, December 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309006316.

Zheng:2008:GIN

[ZYCI08] Bing Zheng, Lijuan Ye, and Dragana S. Cvetkovic-Ilic. Generalized inverses of a normal matrix. Applied Mathematics and Computation, 206(2):788–795, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308007091.

Zhu:2007:SCA

[ZYD07] Kejun Zhu, Shiwei Yu, and Fengqin Diao. Soft computing applications to estimate the quantitative contribution of education on economic growth. Applied Mathematics and Computation, 187(2):1038–1055, April 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306012410.

Zhang:2009:NLB

[ZYD09] Jianying Zhang, Guangwu Yan, and Yinfeng Dong. A new lattice Boltzmann model for the Laplace equation. Applied Mathematics and Computation, 215(2):539–547, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005098.

Zhang:2008:IBT

[ZYH08] Tong Zhang, Hanchun Yang, and Yinnian He. Interactions between two rarefaction waves for the pressure-gradient equations in the gas dynamics. Applied Mathematics and Computation, 199(1):231–241, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009903.

Zhang:2006:SRD

[ZYME06] Yu Zhang, Xiaofan Yang, Graham M. Megson, and David J. Evans. On the system of rational difference equations x_{n} = A + 1/y_{n−p}, y_{n} = A + y_{n−1}/(x_{n−r}y_{n−s}). Applied Mathematics and Computation, 176(2):403–408, May 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305008143.

Zhang:2009:NSC

[ZYT09] Yijun Zhang, Dong Yue, and Engang Tian. New stability criteria of neural networks with interval timevarying delay: a piecewise delay method. Applied Mathematics and Computation, 208(1):249–259, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308008977.

Zhang:2008:PSS

[ZYW08] Mingchuan Zhang, Yanmin Yin, and Zhongli Wei. Positive solution of singular higher-order mpoint boundary value problem with nonlinearity that changes sign. Applied Mathematics and Computation, 201(1–2):678–687, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300308000349.

Zhu:2006:HOP

[ZYZ06] Shaohong Zhu, Zhiling Yu, and Jennifer Zhao. A highorder parallel finite difference algorithm. Applied Mathematics and Computation, 183(1):365–372, December 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306005686.

Zhou:2007:ESP

[ZYZ07] Jie Zhou, Zuodong Yang, and Jianqing Zhao. Existence of singular positive solutions for a class quasilinear elliptic equations. Applied Mathematics and Computation, 190(1):423– 431, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300307000720.

Zhang:2005:PAN

[ZZ05a] Guang Zhang and Lijie Zhang. Periodicity and attractivity of a nonlinear higher order difference equation. Applied Mathematics and Computation, 161(2):395–401, February 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012931.

Zhang:2005:TEO

[ZZ05b] Jun Zhang and Jennifer J.

Zhao. Truncation error and oscillation property of the combined compact difference scheme. Applied Mathematics and Computation, 161(1):241–251, February 4, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303012438.

Zhen:2006:SCM

[ZZ06a] Wang Zhen and Hongqing Zhang. A symbolic computational method for constructing exact solutions to difference-differential equations. Applied Mathematics and Computation, 178(2):431–440, July 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305009884.

Zhu:2006:SCS

[ZZ06b] Z. B. Zhu and K. C.

Zhang. A superlinearly convergent SQP algorithm for mathematical programs with linear complementarity constraints. Applied Mathematics and Computation, 172(1):222–244, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002006.

Zhang:2007:NGR

[ZZ07a] Xiao-Ling Zhang and Hong-Qing Zhang. A new generalized Riccati equation rational expansion method to a class of nonlinear evolution equations with nonlinear terms of any order. Applied Mathematics and Computation, 186(1):705–714, March 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306009829.

Zhao:2007:EAS

[ZZ07b] Dianli Zhao and Lei Zhang. Exponential asymptotic stability of nonlinear Volterra equations with random impulses. Applied Mathematics and Computation, 193(1):18–25, October 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307003621.

Zheng:2007:ATP

[ZZ07c] Baodong Zheng and Yang Zhang. On additive transformations preserving a multiplicative matrix function. Applied Mathematics and Computation, 189(2):1402–1409, June 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306017267.

Zhi:2007:ACT

[ZZ07d] Hongyan Zhi and Hongqing Zhang. Applications of the combined tanh function method with symmetry method to the nonlinear evolution equations. Applied Mathematics and Computation, 188(1):385– 393, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306013622.

Zhou:2007:RSH

[ZZ07e] Shuzi Zhou and Zhanyong Zou. A relaxation scheme for Hamilton– Jacobi–Bellman equations. Applied Mathematics and Computation, 186(1):806– 813, March 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306010022.

Zeng:2008:MPP

[ZZ08a] Zhijun Zeng and Zhongcheng Zhou. Multiple positive periodic solutions for a class of state-dependent delay functional differential equations with feedback control. Applied Mathematics and Computation, 197(1):306–316, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007783.

Zhou:2008:OAP

[ZZ08b] Xiaoliang Zhou and Weinian Zhang. Oscillatory and asymptotic properties of higher order nonlinear neutral difference equations. Applied Mathematics and Computation, 203(2):679–689, September 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308003925.

Zhou:2009:GPI

[ZZ09] Yuan-Yuan Zhou and Guo-Feng Zhang. A generalization of parameterized inexact Uzawa method for generalized saddle point problems. Applied Mathematics and Computation, 215(2):599–607, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309005360.

Zhang:2008:ATW

[ZZB08] Jian Zhang, Jiao Zhang, and Liling Bo. Abundant travelling wave solutions for KdV–Sawada– Kotera equation with symbolic computation. Applied Mathematics and Computation, 203(1):233–237, September 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300308002439.

Zhao:2006:CCF

[ZZC06] Jichao Zhao, Tie Zhang, and Robert M. Corless. Convergence of the compact finite difference method for second-order elliptic equations. Applied Mathematics and Computation, 182(2):1454–1469, November 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300306004735.

Zhang:2007:SND

[ZZDL07] Suying Zhang, Minzhen Zhang, Zichen Deng, and Wencheng Li. Solution of nonlinear dynamic differential equations based on numerical Laplace transform inversion. Applied Mathematics and Computation, 189(1):79–86, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300306015827.

Zang:2006:BLC

[ZZH06] Hong Zang, Tonghua Zhang, and Maoan Han. Bifurcations of limit cycles in a cubic system with cubic perturbations. Applied Mathematics and Computation, 176(1):341– 358, May 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630030500809X.

Zhang:2009:GRA

[ZZH09] Jian-Chen Zhang, Shu-Zi Zhou, and Xi-Yan Hu. The (P Q) generalized reflexive and anti-reflexive solutions of the matrix equation AX = B. Applied Mathematics and Computation, 209(2):254–258, March 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308009429.

Zhao:2005:TLF

[ZZKY05] Jennifer J. Zhao, Jun Zhang, Ning Kang, and Fuqian Yang. A two level finite difference scheme for one dimensional Pennes’ bioheat equation. Applied Mathematics and Computation, 171(1):320–331, December 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305000986.

Zhang:2005:IPC

[ZZL05] Shan-Qing Zhang, Gui-Xu Zhang, and Zhi-Bin Li. ICS: a package for computing involutive characteristic set of algebraic partial differential equation systems. Applied Mathematics and Computation, 160(2):551–577, January 14, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300303011822.

Zhao:2008:UAT

[ZZL08] Caidi Zhao, Shengfan Zhou, and Yongsheng Li. Uniform attractor for a two-dimensional nonautonomous incompressible non-Newtonian fluid. Applied Mathematics and Computation, 201(1–2):688–700, July 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300308000350.

Zhang:2007:HPS

[ZZLL07] Jing-Ru Zhang, Jun Zhang, Tat-Ming Lok, and Michael R. Lyu. A hybrid particle swarm optimization–backpropagation algorithm for feedforward neural network training. Applied Mathematics and Computation, 185(2):1026–1037, February 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306008277.

Zhang:2008:GSS

[ZzLZ08] Fengpan Zhang, Zi zhen Li, and Feng Zhang. Global stability of an SIR epidemic model with constant infectious period. Applied Mathematics and Computation, 199(1):285–291, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307009964.

Zhao:2006:TWS

[ZZTJ06] Xiaoshan Zhao, Hongxian Zhou, Yaning Tang, and Huabing Jia. Travelling wave solutions for modified Zakharov–Kuznetsov equation. Applied Mathematics and Computation, 181(1):634–648, October 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306001445.

Zhang:2008:ESU

[ZZW08] Chun-Xia Zhang, Jiang-She Zhang, and Guan-Wei Wang. An empirical study of using Rotation Forest to improve regressors. Applied Mathematics and Computation, 195(2):618–629, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005917.

Zhang:2009:BLC

[ZZW09a] Qi Zhang, Zhao Zhengquan, and Gui Weihua. Bifurcation of limit cycles at the equator. Applied Mathematics and Computation, 211(2):422–426, May 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309000897.

Zheng:2009:CES

[ZZW09b] Cheng-De Zheng, Huaguang Zhang, and Zhanshan Wang. Comments on “Exponential stability analysis for uncertain neural networks with interval time varying delays” and “New delay-dependent robust stability criterion for uncertain neural networks with time-varying delays”. Applied Mathematics and Computation, 214(2):648, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003488.

See [KP08b].

Zhao:2005:MMC

[ZZX05] Haibo Zhao, Chuguang Zheng, and Minghou Xu. Multi-Monte Carlo method for particle coagulation: description and validation. Applied Mathematics and Computation, 167(2):1383–1399, August 15, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300304005740.

Zhou:2008:CTI

[ZZYT08] Jiali Zhou, Shuyou Zhang, Guoping Yang, and Jianrong Tan. A convergence theorem for the inexact Newton methods based on Hölder continuous Fréchet derivative. Applied Mathematics and Computation, 197(1):206–211, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307007709.

Zhang:2008:DDM

[ZZYY08] Kai Zhang, Ran Zhang, Yunguang Yin, and Shi Yu. Domain decomposition methods for linear and semilinear elliptic stochastic partial differential equations. Applied Mathematics and Computation, 195(2):630–640, February 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300307005929.

Zhao:2006:NRE

[ZZYZ06a] Xueqin Zhao, Hongyan Zhi, Yaxuan Yu, and Hongqing Zhang. A new Riccati equation expansion method with symbolic computation to construct new travelling wave solution of nonlinear differential equations. Applied Mathematics and Computation, 172(1):24–39, January 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305001827.

Zhi:2006:MHF

[ZZYZ06b] Hongyan Zhi, Xueqin Zhao, Zhongyuan Yang, and Hongqing Zhang. A modified hyperbolic function method and its application to the Toda lattice. Applied Mathematics and Computation, 172(2):938–945, January 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300305002481.

Zheng:2007:GRE

[ZZZ07] Ying Zheng, Yuanyuan Zhang, and Hongqing Zhang. Generalized Riccati equation rational expansion method and its application. Applied Mathematics and Computation, 189(1):490–499, June 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306016225.

Zhang:2009:FPT

[ZZZ09] Guowei Zhang, Tongshan Zhang, and Tie Zhang. Fixed point theorems of Rothe and Altman types about convex-power condensing operator and application. Applied Mathematics and Computation, 214(2):618–623, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300309003610.